A Complete Bibliography of ACM Transactions on Graphics

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

10 May 2024 Version 1.157

Title word cross-reference

2 [AWL^{+}19, BKLP16, BHR13, BSW02, BGF^{+}23, BSM^{+}07, DBB^{+}17, EPD09, GIZ09, HGRT04, Hil87, HDK07, JSKJ12, KFCO^{+}07, LT09, LPL^{+}17, LHVT17a, LHVT17b, Mai92, MC21, MU22, NG18, RMD12, SLV^{+}13, Shn92, XCS^{+}14, YPL^{+}23]. 2.5 [RID10]. 3 [AJS20, AKZ^{+}17, AL13, ALX^{+}14, AXZ^{+}15, AZB09, AAR05, AVB^{+}23, AS21, AIH^{+}08, ARS14, BVF^{+}17a, BIP01, BLC^{+}22, BP07, BSS^{+}11, BSK^{+}16, BBN^{+}12, BVG11, BGK^{+}13, BWSS12, BVS16, Bly06, BSM^{+}07, BB22, BR07, BAU15, BATU18, CCA^{+}12, CB04, CWLZ13, CAD^{+}21, CMZP14, CK10, CKGK11, CGF09, CSPF12, CZS^{+}13, CLD^{+}13, CZL^{+}15b, CKIW15, CLF^{+}18, CPY^{+}22, CPW^{+}23, CSL^{+}22, CGP^{+}21, CRCM23, DNZ^{+}17b, DNZ^{+}17a, DS15, DTP15, DLSCS08, DSAF^{+}13, DIP^{+}18, DHL14, DDP02, DSC^{+}20, ESCK16, EBGB14, EDF^{+}16, EST^{+}20, EPD09, ESZ^{+}17, ERP^{+}19, EM96, FZBR16, FLJK21, FFBB21, FJL^{+}16, FH10, FRS^{+}12, FSL^{+}15, FBS^{+}23, FMK^{+}03, GDAB^{+}17a, GDAB^{+}17b, GZW^{+}16, GZC^{+}16, GIZ09, GM05, GF08, GGS03, GTDS10, GKHH12, GWN^{+}03, GWB05, GHL^{+}20, GFD^{+}12, GRT13, GZC15, HGRT04, HGY17, HASK17, HK18a, HNH19, HLP^{+}22, HLHR09, HLZ10, HZP^{+}22, HDK07]. 3 [HMC11, HLV^{+}17a, HLV^{+}17b, HHL^{+}24, HTWB11, HCTW11, HMT^{+}15, HDGN17, HMM^{+}21, HZC^{+}22, Hud92, HOM15, IBP15, ICG17, JTRS12, JBM^{+}17, JLF^{+}09, JBX^{+}20, JCG^{+}21, JHR22, JHS^{+}23, JZH07, KMM^{+}02, KHS10, KH06, KSH^{+}14, KWS^{+}23, KDM^{+}16, KDR^{+}16, KDW^{+}17, KKLD23, KC23, KSES14, KMYG12, KLM^{+}12, KRD^{+}12, KLM^{+}13, KLKL13, KNK^{+}22, KTL^{+}04, KDMW17, KL22, KSS^{+}15, KS04b, KYC^{+}17, LMS13, LHW^{+}10, LRAT08, LHKR10, LCXS09, LOMI11, LRA^{+}07, LACS08, LSH^{+}10, LVG^{+}13, LYF^{+}20, LLHF21, LJZ^{+}23, LHM^{+}18, LCOZ^{+}11, LYC18, LOW18, LFZ18, LGJA09, LWCT14, LHLF15, LGB^{+}21, LHH^{+}23, LKG^{+}03b, LFL09, LvBK^{+}10, LSZ^{+}14, LBRM12, MLZ^{+}16, MPF^{+}18, MHS^{+}19a, MLYZ19, MWH^{+}13, MPI^{+}18, MSHS06, MPN^{+}02, MP04, MRA^{+}22, MAN^{+}16, MTN^{+}15, MSS^{+}17, MPE^{+}23, MGP10, MGP06, MYWI15, MLS^{+}18, NLGK18, NAH^{+}18, NISA07, NRDR05, NPLX22, NZIS13, NIR^{+}21, NPC^{+}22, OHB^{+}11, OLGM11, ONOI04, PQF^{+}23, Par17, PGP^{+}19]. 3 [PMW^{+}08, PK05, PXW18, PZ17, PRM14, PS04, PAR21, PDF^{+}22, PAAG21, PFB^{+}20, PSG^{+}06, PWLSH13, RSL16, RPC^{+}21, RSI^{+}08, RDI10, RB23, RHHL02, RMBB^{+}13, SS14, SHM^{+}18, SCH^{+}14, SLV^{+}13, SSGS11, SKSK09, SBR^{+}15, SSW^{+}23, SHL^{+}17, SHZ^{+}20, SAA^{+}21, SF07, SGXT20, SGX^{+}21, SSS^{+}08, SARW^{+}15, SSS06, SDW^{+}16, SVB^{+}12, SQRH^{+}16, SRB^{+}19, SWS^{+}22, SSK^{+}17, MBU22, TD16, TDM11, TMB18, TS08, TTZ^{+}20, TFK^{+}03, TMB14, UZB^{+}23, UTB^{+}19, VVC^{+}15, VSHJ12, WBF^{+}17a, WBF^{+}17b, WAO^{+}09, WWY^{+}13, WGW^{+}13, WSXC16, WLG^{+}17, WSLT18, WSH^{+}18, WLX^{+}18, WKHA18, WMB^{+}20, Wan23, WLLS22, WXLY17, WLHR11, WDB^{+}07, WSW^{+}12, WZQ^{+}18, WWL^{+}19, WZL^{+}20, WZ22, XSL^{+}22, XLF^{+}11, XIAP^{+}17, XZP^{+}23, XZT^{+}09, XZZ^{+}11, XZCOC12, XCF^{+}13, XCS^{+}14, XSZ^{+}16, XNZ^{+}22, YI17, YSL^{+}14, YSC^{+}16, YLJ18, YSW^{+}23, YWS^{+}11, YLC^{+}20, YZL^{+}22, YML^{+}23, YKC^{+}16, YZX21, YZX^{+}18, YAB^{+}22, YSHWSH16, ZLP^{+}15, ZAC^{+}17, ZYX^{+}21, ZCM22, ZQL^{+}23]. 3 [ZTNW23, ZWK14, ZPYX23, ZSW^{+}10, ZBW^{+}20, ZPW^{+}23, ZSMS14, ZK14, ZZCJ13, ZPKG02]. 360^{∘} [Kop16, LLZ18]. 4 [Che13, DKD^{+}16, HTCH15, IGP^{+}17, LHG^{+}09, LBB^{+}17b, LHZ^{+}21, MHS^{+}19a, MPDW03, NKS^{+}23, PS04, PMPHB17, RAWV08, TDL^{+}18, YMRD15, Zho18]. 5 [BSS^{+}13, OHX^{+}14]. 6 [HMT^{+}15]. 8 [BMBRD24]. ^{2} [LZ04]. ^{∘} [JMY^{+}07, BYLR20, CLJ^{+}20, KC19, TWLT19]. ^{TM} [SMG^{+}05]. C^{2} [MP09c, Pet89, SW05, Yuk20]. C_{0} [Jam20]. d [EPM^{+}14]. Δ [LWF^{+}23, YBAF22]. ℓ_{1} [ASGCO10]. f [LWO19]. G^{1} [LFS16, Sar00]. G^{k} [PU06]. γ [CXW^{+}05]. K [FLHCO10, TS12, Tsa15, EPM^{+}14, LWO19, MSDL17, YSW^{+}17]. L_{0} [HS13, XLXJ11]. L_{1} [BHY15, HWCO^{+}13, PMA^{+}14, HJS^{+}14, WYL^{+}14]. L_{p} [LL10]. λ [MWM23]. N [RVLL08, RS14b, Ten20, BSEH18]. p [GA20]. Q [LH17a, LH17b]. r [DS92]. R^{3} [Sar00]. s [SR00]. SO(3, R) [CGM11]. T [MPKZ10]. v [XH18].

*Cages [GPCP13].

-analysis [WYL^{+}14]. -axis [ERP^{+}19]. -Based [HJS^{+}14]. -bit [BMBRD24]. -Clustered [Tsa15, TS12]. -curves [YSW^{+}17]. -D [BR07, Hil87, Mai92, Shn92]. -dimensional [EPM^{+}14, Ten20]. -direction [PS04]. -DOF [HMT^{+}15]. -grooves [XH18]. -Learning [LH17a, LH17b]. -medial [HWCO^{+}13]. -meshes [MPKZ10]. -nearest [MSDL17]. -power [SR00]. -Prox [LWF^{+}23]. -set [FLHCO10]. -Sets [DS92]. -Sparse [ASGCO10]. -subdivision [MWM23]. -Sweep [CZS^{+}13]. -Symmetry [BSEH18, RS14b, RVLL08]. -ton [CXW^{+}05]. -Willmore [GA20].

2 [MKRH11]. 2-manifolds [Man86]. 2PAC [TFD^{+}18].

360 [JMY^{+}07, WPL^{+}21]. 3D [WW82]. 3QNet [HZC^{+}22].

4 [BAM13]. 4-points [AMCO08].

5DOF [WPGM16].

6D [FRSL08].

77 [VCA^{+}22].

A-Patches [BCX95]. AA [AHD15]. aberration [CLS^{+}17, WLM^{+}15]. Aberrations [CFP^{+}21, HLBR12, HWBR14, POAR12]. ABF [SLMB05]. absolute [KS04a]. absorbent [CT05]. Absorption [BBS14a]. Abstract [KK91, YXFH21, YL10]. Abstracted [XZP^{+}23, LMLH07]. Abstraction [ACP^{+}01, MZL^{+}09, BSM^{+}13, DS02, FLB17, LYC18, LFZ18, NSX^{+}11, WOG06, YC21, YK12]. Abstractions [JGMR23]. Accelerate [MHNT15]. Accelerated [BSSJ23, HYS23, JRSS21, KGL16, ZCT22, BDT99, CW17, KB12, LYK^{+}21, NPP^{+}11, PVL^{+}05]. Accelerating [BJ10a, BKKL15, LNLB16, RV89, LVS^{+}16, Wan15, YPB16]. acceleration [CZJ12, JLBM05, MA06, PDZ^{+}18, MA07]. accelerations [KLF^{+}19]. Accelerometer [SH08, TZK^{+}11]. Accelerometer-based [SH08]. Access [VSW^{+}23, KCYW13, LSK^{+}06, NH08]. Accommodation [KPM^{+}17, KBBD17, CLS^{+}17, KNL^{+}22, MWH^{+}13]. Accommodation-invariant [KPM^{+}17]. Accommodative [KNL^{+}22]. account [CLC96]. accuracy [CKH18, LDS02, SHD^{+}18]. Accurate [BSSJ23, BOFN18, GM09, GGHS03, HCH22, KWS^{+}23, LBHH23, MSHS06, SBN15, SSR20, Ste20, VJ19, WZC12, WZYR19, XLC^{+}23, XNZ^{+}22, ZBGB19, BBB07, BHK14, CGP^{+}21, Dee05, DDP99, HHM19, JBP06, LBB17a, LD14, LKYU12, MLT17, MG03, SXH^{+}21, VMTF09, XSTN14, YTJR15]. Achieving [JLF^{+}09]. achromat [Fre16]. achromatic [SDP^{+}18]. Acknowledgment [Ano10]. Acknowledgments [Hod02a]. ACM [Kro82, Spe03]. Acorn [MLL^{+}21]. Acoustic [LFZ15, LLMZ16, PFP^{+}22, ACSM12, JBP06, JLWM22, LZJ16, OHR14, WJ19]. acoustic-potential [OHR14]. Acquiring [Bou18, KMYG12, NGD^{+}06, TDG18, TFG^{+}13, DWd^{+}08, OEE^{+}18]. Acquisition [Did18, HED05, HHA^{+}10, LCC^{+}22, TG17b, BGK16, BJTK18, BTFN^{+}08, DJ18b, GHP^{+}08, GGHS03, GLL^{+}04, GRB^{+}18, GTR^{+}06, GLT^{+}21, HLZ10, HCTW11, HJM^{+}22, LLW^{+}08, MP04, NLW^{+}16, NLGK18, NJR15, PCK^{+}08, RHHL02, SWTC14, TG17a, XSZ^{+}16, XNY^{+}16, ZCD^{+}16, ZRL^{+}09]. Across [JNK^{+}23, MGS^{+}21]. acting [DYP03]. Action [ANBH23, ACCO05, MLZ^{+}16, DWT^{+}02, GCR13, SCH^{+}14]. Action-driven [MLZ^{+}16]. actions [ACOYL08, BDG15, YM16]. activations [SNF05]. Active [CHP07, DPD22, FLP14, MNBN07, PM18, RV89, WAvK^{+}12, YSW^{+}23, PZM13, SWR^{+}21, YKC^{+}16]. activities [KLF^{+}19]. Activity [FSL^{+}15, LY23, SFC^{+}23, FCW^{+}17]. activity-associated [FCW^{+}17]. Activity-centric [FSL^{+}15]. Actor [LXZ^{+}19, LHR^{+}21]. actors [CTMS03, LHR^{+}21, WSVT13]. Actuated [JHS^{+}23, KMM17b, Ano03, GMB17, KMM17c, LPLL19, STC^{+}13, YKZ^{+}22]. actuation [Ano03, HPC21, JWDL19]. actuators [WHDK12]. Acuity [MGDA^{+}15]. Adaptation [SP05, XXA^{+}23, YNK^{+}22, DE05, GPM^{+}22, HKT10, VMGM15]. adaptations [HGRT04]. adapted [BBR^{+}21, Sze06]. AdaptiBrush [RAR^{+}21]. adapting [PSK^{+}12, YCBvdP08]. Adaptive [BSSJ23, BMW^{+}09, BO04, BF08, BDW13, CGG^{+}04, CBK20, CJM21, DWX^{+}21, EC96, FCW^{+}17, FBLS07, GO12, HYS23, HWRH13, Hil87, HWH^{+}16, HWZ^{+}20, JLS^{+}03, KD13a, KTS^{+}14, KYS^{+}15, LPC22, MAKWL22, MCY14, MIGYM15, MMMG16, MNV^{+}21, NSO12, ODR09, PNdJO14, RGL05, RKZ11, RKZ12, Sah18, SMR^{+}22, SHFH11, SW18, WFP12, WSLT18, WSND^{+}23, WYW23, WK21, ZTD^{+}23, AGL^{+}17, ANHD17, ATW13, AB20, BAM13, BLDA11, BFK^{+}16, CYFW14, CTH^{+}14, DJ18b, EB14, FFB^{+}09, GTJS17, GWAB19, GKS02, HJW^{+}08, HJ11a, HTC^{+}14, HGF14, KJM10, KSP13, LHKR10, LWC12, LDN^{+}18, MDK08, MLL^{+}21, MB12, NNC^{+}20, NPO13, NLMD12, PO08, PBvdP16, PLC^{+}21, RAR^{+}21, RWL^{+}22, SABS14, THKM13, VdFG99, VKK18, VKJ19, WPC^{+}14, WFDH18, WLT22, cWP10, XCW^{+}20, YI17, YW13, YIC^{+}10, YSC^{+}18, ZSKS18, ZD20]. Adaptively [APKG07]. adaptivity [TMDK15, WHK17, ZLB16a]. AdaptNet [XXA^{+}23]. add [MRK^{+}13, GHZ^{+}20]. add-on [MRK^{+}13]. Adding [DKD^{+}17a, Hud92, DKD^{+}17b, SKC^{+}14]. Additive [LJM^{+}16, AHB18, BR21b, MDL16, MSDL17, MHSL18, PZM^{+}15, WWW22, YIO^{+}15, ZJ18]. address [KWB^{+}13]. adhesion [AAT13]. adjacency [SNCH08]. Adjoint [PM95, TOG22, VK16, MTPS04]. Adjoint-driven [VK16]. Adjunct [Bea88]. Adjustable [SB93]. Adjusting [WXZ^{+}23, DIO^{+}12]. Adjustment [WZHL23, YZW^{+}16, ZZL^{+}21, BSK^{+}16, LFUS06, Wan18a]. ADMM [BN21]. ADOP [RFS22]. Advanced [CFS^{+}18, SM15, Zha18, FLG15]. advection [BNTS07, ZNT18, ZBG15a]. advection-projection [ZBG15a]. advection-reflection [ZNT18].

Adversarial [SSII18a, GDG^{+}17, PMA^{+}21, PGH^{+}22, WAH^{+}10, ZZB^{+}18]. Aerial [HNH19, Liu18, LCX^{+}21, LLH^{+}22, OKH^{+}16, SMGH18, NOP^{+}18, ZYX^{+}21, ZXH^{+}20]. Aerobatics [WPL18]. aerodynamic [DYN03, UB18]. Aerodynamics [XIM18, MUB15]. Aerophones [AR15]. Aesthetic [ZLH^{+}21]. Aesthetic-guided [ZLH^{+}21]. aesthetically [GSH18]. Aether [ALLD17]. affect [HCOB10]. Affine [BT19, LKL^{+}22, SR09, GAB20, HSS98, JSS^{+}15, RV11]. Affinities [PABE^{+}21]. affinity [IMF^{+}21, XLJ^{+}09]. affinity-based [XLJ^{+}09]. Affordable [IRN^{+}22]. after [HXM^{+}13]. against [Wan14]. age [APCO21]. Agent [SGD21, BDM^{+}20, LSCC20]. agglomeration [GJTP17]. Aggregate [NGCL09, CHPR07, GvdBL^{+}12]. aggregates [RBF08]. aggregation [SHM22, TJM15, ZZW^{+}22b]. AgileGAN [SLL^{+}21b]. Aging [BW22, ZCS^{+}22]. Agnostic [HFW^{+}19, SACO22, AWL^{+}19]. Ahead [Fol91, ZK22, ZMN^{+}19]. AI [BDM^{+}20]. Aided [BG89b, Gol84, Gol85a, SMPZ15]. AIM [Ano10]. Air [AS21, MCKM15, CFW13, LSCS14, SPGI13]. Airbrush [SMPZ15]. AIREAL [SPGI13]. airplanes [UKSI14]. Airy [MIB15]. Alain [Fiu00]. Albedo [GXY^{+}17a, DDTP15, GXY^{+}17b, HR13]. algebra [Gol85b, KW03, LKJC21, MDP^{+}04]. Algebraic [ACC90, BI92, BIW93, Bli82, CCOST05, GM84, GG07, MD94, PBS20, PP93, AB89, BWSK12, CZXL23, Gol02, LT09, LB06, SHM22, War89]. Algebras [Duf17a, Duf17b]. Algorithm [CG89, Day90, EPO91, HA92, KM97, LWF^{+}23, LMR83, LM97, Mey91, MPB17a, PF89, Sah18, Sai89, SG82, SO92, WFS^{+}21, WS85, Zyd88, AAM03, BCRK^{+}10, BSFG09, CS00, EKA84, EPD09, GI04, GMP09, GD04, LZF10, MMT18, MBF04, MPB17b, RSH05b, SYBF06, SOA11, SSBD03, XW09, XCM^{+}14, YXH14]. Algorithms [Bak94, CMS95, CLS85, DGHM93, Dun83, EM90, Jan91, JDH^{+}22, Kla91b, Kro82, MGA^{+}22, MD94, MST89, RV89, VN85, EKA84, GYGS22, HDN^{+}16, KW03, LJGH11, RKAP^{+}12, Spr82, WDB^{+}08]. alias [SOA11]. alias-free [SOA11]. Aliasing [Tur82, WCZ^{+}22, BAM13, HSD13]. Aliasing-Aware [WCZ^{+}22]. Aligned [MLS^{+}18, SOG^{+}22, ZVC^{+}20, CPMS14, GDC15, HTWB11, JTPSH15, MWR12, MWRD13, MYRD14, MPKZ10, MPZ14, PLS^{+}15, STJ^{+}17, TPP^{+}11, XCOJ^{+}09]. ALIGNet [HFW^{+}19]. Alignment [HFW^{+}19, HXM^{+}18, LHJ^{+}14, LCC^{+}22, SDA^{+}23, ARS14, BR07, CWTW17, FZZ^{+}20, HZM^{+}08, SPKS16, SRL^{+}15, WGP^{+}10, WSH19, XLY09, YSW^{+}20, ZLE14, BZL^{+}17]. alive [CMT^{+}12, HLYK08, LBB02]. All-frequency [NRH03, TS06, WTL05, WTL06b, WRG^{+}09, ADM^{+}08, NRH04, XCM^{+}14]. All-hex [FXBH16, LLX^{+}12, LZS^{+}21]. all-pairs [AP08]. along [WSH19]. Alpha [EM94, MNB23, PRLH^{+}22]. alternative [HGRT04, LD06]. Always [DLP^{+}23]. Ambient [GAF^{+}10, ZRSM18]. Ambiguity [LTH^{+}23]. Ambiguity-Aware [LTH^{+}23]. amendment [MB21, TBTS08]. Ames [STXJ15]. AMFS [CTH^{+}14]. amodal [YZL^{+}22]. among [SGG^{+}06, WWOH08]. amorphous [ZYL^{+}20]. Amortized [YNS^{+}09]. Amortizing [WWWZ23]. AMP [PMA^{+}21]. Amplification [PGP^{+}19]. analog [HSHF10]. analogies [WWH06]. analogue [SR97]. analogy [LYY^{+}17]. Analysis [ASHW23, BBS14a, CM83, DTPC23, DKD^{+}17a, EC93, HNO^{+}23, KP92, Kla91b, LLZM10, LTDD16, LDW97, Mai92, MOR^{+}18, ÖG12, SPV^{+}16, VFK^{+}14, WBCPS19, WMP^{+}06, Wu92, YKGA17a, YZX^{+}18, ZCT22, ZXTZ15, BHR13, BBB^{+}14, BWWM10, CCOST05, CZXL23, DHS^{+}05, DKD^{+}17b, ETH^{+}09, EHDR11, FKY08, FV96, GF08, HXZW20, HSTP11, HRV97, HvKW^{+}16, HSS^{+}13, HWK15, HHA^{+}10, JSKJ12, KSHG18, hKPS03, KCGF14, LSD^{+}16, LHG^{+}09, LLH04, MC12, OK10, OHX^{+}14, Par17, PSC^{+}15, PCHF18, RMB07, ST14, SJ22b, SJ17, SK13, TOS^{+}03, WAvK^{+}12, WGW^{+}13, WYL^{+}14, WLG^{+}17, WW11, XHS^{+}15, YKGA17b, ZTS09, ZN06, ZXJ^{+}13, ZPZ13, vKXZ^{+}13]. analysis/synthesis [TOS^{+}03]. Analytic [Cas91, NL13, SMR^{+}22, SKSK09, SDK19, WR18, WYW23, ZWRY21, BLPW14, HW12, SRNN05, SJR18, WAK20]. analytical [GBAM11]. analytically [GHZ^{+}20]. analytics [SHK^{+}14]. analyze [GSMCO09]. analyze-and-edit [GSMCO09]. Analyzing [Che13, SHH99, HWG14, KGFF14]. anatomical [KIL^{+}16]. Anatomically [LJL23, CCGB22, DZS08, SZK15, WBGB16]. anatomically-based [SZK15]. anatomically-constrained [CCGB22, WBGB16]. Anatomy [AHLG^{+}13, WMB21, ZWHB22, ABL^{+}21]. anchor [BHB^{+}11]. Anchors [XHWW22]. Anderson [PDZ^{+}18]. angle [CAA09, PRP^{+}15, SLMB05, SLL19, TAV^{+}10]. angle-based [PRP^{+}15]. angles [CKMR^{+}21, LS07]. Angular [DLW^{+}22, KZP^{+}13]. animal [WP09a, XWL^{+}08]. animals [WPP14]. Animatable [BZH^{+}23, ZQL^{+}23, BLS^{+}21, FFBB21, SGdA^{+}10, XPB^{+}21]. Animated [FZLM11, JK23, NPC^{+}22, TGBE16, VKJ^{+}17, ZJY^{+}22, BCC17, CS09, HRvdP04, LCR^{+}02, MBB12, MA06, NSB13, OHR14, SN17, SS17, SDO^{+}04, TLJP18, WIK^{+}06, WG09]. Animating [BDWR12, CJ11, CGZ^{+}05, CTTL15, EB08, FOA03, FOK05, GPH^{+}18, KA08, MWTK13, SRH^{+}15, SvTSH14, SJM17, SZL^{+}23, XXK^{+}06, XWL^{+}08, YL08, ZB05, BAAR12, BWHT07, BBS^{+}13, CMT04, CLQW08, GBO04, HHV^{+}21, LZJ16, PH06, PTG12, PNDN12, SB12, TMB14, WCF07, WCL^{+}20]. Animation [AJS20, BC14, CZB23, EMF02, EHSN20, EAPL06, HSX^{+}22, HTCH15, JW15, MMG06, RPC^{+}10, SDN18, TBW^{+}12, ZWL22, ZZC^{+}23, AWL^{+}20, AHSS04, ASK^{+}05, BKLP16, BP07, BdSP09, BJS^{+}08, BCK^{+}13, BWP13, BFA02, CTFP05, CWLZ13, CHZ14, CWW^{+}16, CAD^{+}21, CH05, CB05, CÖS19, Cor18, DRvdP15, DYP03, DBB^{+}17, DSC^{+}20, Erl07, EGP02, FL04, FYK10, GSZ^{+}18, GB13, GMP^{+}16, GRGC15, HYL12, HZP^{+}22, HDK07, IKKP17, IWZL09, JTCW07, JGGN15, KIL^{+}16, KAL^{+}17, KSKL14, KPMP^{+}17, KGP^{+}16, KFCO06, KCD09, LJ14, LLL18, LYYB13, LWL17, LMM^{+}22, LXC^{+}15, LCC21, MCC09, MCW^{+}21, MCP^{+}09, NZC^{+}18, NSCL08, NKAS08, NFJ02, OBH02, OLSL16, PKA^{+}05, PB02, RP03, RP07, SHW19, SSK^{+}11, SY05, SKSY08, SKM10, SZZK21, SKP08, TKY^{+}17, TLP07, TCL21, VBMP08, WP06, WAH^{+}10, WDAC06, WHRO10, WSXC16, WQLJ18, WSS^{+}19, WBLP11, WSL13, WFL^{+}19, YL10, YRPF09]. animation [YCZ11, YGM97, ZSCS04, ZM13, ZXL^{+}18, ZHS^{+}20, ZPBC19, ZMCF05, ZBBB18, dSDP09]. Animation-Ready [CZB23]. Animations [PM18, YBMN^{+}23, DLKS18, FJS^{+}17, GSKJ03, GJ22, HOKP16, JT05, JFA^{+}15, KG06, LP02, LMY^{+}13, ODGK03, cWP03, XWSY15, YKH04]. animator [ELFS16, ZXL^{+}18]. animator-centric [ELFS16, ZXL^{+}18]. animatronic [HPC21]. Anime [LXZ^{+}23]. AniMesh [JGGN15]. AnisoMPM [WCL^{+}20]. Anisotropic [ACSD^{+}03, BX03, BSTY15, CZM^{+}23, FLSG14, GFL^{+}22, GZD08, JGT17, KDI19, LWSF10, LLR^{+}15, LDS^{+}22, McC99, Ste20, XSD^{+}13, CK11, FZZ^{+}20, JAM^{+}10, NSO12, PPTSH14, PLMR17, PTC^{+}10, PH15a, SJ17, TOII08, WZT^{+}08b, WNEH22, WCL^{+}20, XLZ^{+}10, YT13, ZJ18, ZHLB10, ZWDR16, ZGW^{+}13]. anisotropy [BLdG^{+}16, KFR04]. annealed [YYW^{+}12a]. annealing [DH96]. annotated [BUSB13, LCL06]. annotation [YKC^{+}16]. annotations [AFO03, GIZ09, TFK^{+}03]. Anti [Tur82, BAM13]. Anti-Aliasing [Tur82, BAM13]. Antialiased [Kla91a, DHI^{+}13]. Antialiasing [BYRN17a, BYRN17b, YSLH11, CS00, GT96]. antiradiance [DSDD07]. Antithetic [ZDDZ21]. Any [GRH^{+}12, GZ05, MYWI15]. Aperture [PC82, RKB^{+}23, BCN08, GSMD07, GWGB10, LFDF07, LCV^{+}04, LLW^{+}08, VRA^{+}07]. Apparent [DER^{+}10, IM10, JDA07]. Appeal [WZC^{+}20]. Appear [ZSAF21]. Appearance [BSK23, CBKM15, CZM^{+}23, DLW^{+}22, DBP^{+}15, DCP^{+}14b, DWMG15, FR22, HXM^{+}18, KSZ^{+}15, LH06a, SPSH14, TBTA^{+}24, TWZ20, VADWG15, VPB^{+}18, WTL^{+}06a, WJHY23, XBS^{+}22, AYL^{+}12, AP08, ATDP11, BBP21, BUSB13, BLS^{+}21, CLL^{+}21, DCP14a, GGN18, GXZ^{+}13, GRB^{+}18, GTR^{+}06, GLZ^{+}21, JFA^{+}15, JSB^{+}10, KWN^{+}17, KRK11, KBC^{+}13, KFB10, LMS^{+}19, LEN09, LDPT13, LKG^{+}03a, LDPT17, LSSS18, LXJ^{+}22, MKZ^{+}21, MWAM05, MDLW15, MGZJ20, NIR^{+}21, ODAO15, PL07, PLMR17, RTD^{+}21, RPK^{+}12, SBdDJ13, SGM^{+}16, SLS^{+}16, TDG18, WM14, WZYR19, XMR^{+}11, ZJMB11, ZJMB12, ZCB^{+}22, ZZW^{+}22b]. appearance-driven [PL07]. Appearance-from-motion [DCP^{+}14b]. Appearance-mimicking [SPSH14]. Appearance-preserving [TWZ20]. Appearance-space [LH06a, AP08, ATDP11]. AppGen [DTPG11]. AppIm [DCP14a]. Application [BBC^{+}23, BLDA11, CA00, DRC^{+}15, RO85, RO87, AG05, BA83]. Applications [APH^{+}14, BIP01, BBG24, BF01, OF01, SR00, YSHWSH16, AČMS10, BZL^{+}15, CH89, DRE^{+}12, DEM96, Fat09a, GKHH12, Gue07, HSGL11, HGS23, JSKJ12, Jia21, KDR^{+}16, KKW21, LWA^{+}12, LL10, MMCK14, MASS15, SMG^{+}20, SCW^{+}21, XLC^{+}16, YGL^{+}14]. applied [BLR^{+}11, SABS14, YBAF22]. AppProp [AP08]. Appreciation [Fiu00]. Approach [AOCBC15, Bar86, Cas91, DPVA23, DKD^{+}17a, EM96, FH97, GM84, JWD^{+}23, MC92, MHCT23, MGDA^{+}15, PPV95, SLGS01, Shn92, SHS^{+}18, US24, WKMH^{+}23, YSC^{+}23, BLR^{+}11, CWW13b, CDM^{+}02, DWT^{+}02, DK09, DIO^{+}12, DKD^{+}17b, DSC^{+}20, FLB16, GSMCO09, GD04, HZW12, HLHZ08, HZG^{+}12, HWJ^{+}15, JHC^{+}21, KBS15, KZ11, LdPS84, MM06, Mor11, MMTD07, NXS12, NØ13, OPOD10, RPE^{+}05, Sha03, SXZ^{+}12, SHS^{+}17, SFWG04, TB21, TZCT20, TKY^{+}17, TWGT10, VBCG10, VdFG99, VMTF09, WFA^{+}05, WWZ^{+}09, Wan15, WWB^{+}19, WGH20, WMW15, Wym05, XRLF15, ZCLJ20, ZCW^{+}17, ZRL^{+}09]. Approaches [Mil87, CWZ^{+}21a, FH04b]. Approximate [DYYT17, HLZCO14, LW15, McI83, NFD07, TLJP18, VFK^{+}14, WLLS22, AFO05, KCŽO08, MS04, MGP06, MCK13, RFS22, SSK^{+}05b, TL04, Wym05]. approximated [KDH22]. approximately [CZM^{+}10]. Approximating [Hub96, LS08, LSNC09, GI04, LYLL08, SOS04]. Approximation [BIW93, LFZ15, TGBE16, Tsa15, BO04, CB17, CZXL23, CPWAP08, CH89, CSAD04, FD17, IRHSH20, MCSA15, NRH03, PZM13, SSR20, TGB13, TS06, TS12, WWS^{+}05, WYY^{+}14, WDB^{+}08, YLJ18, ZYWK08]. Approximations [CJM21, DLTW90, Tau94, ZLZ^{+}23, BODO18, HW16, KFB10, ZFO^{+}22]. AppWand [PL07]. AppWarp [ATDP11]. ARAP [LCK22]. Arbitrarily [HA92, KG06]. Arbitrary [CSZZ20, EPO91, JPL22, LDW97, Sar00, Sei93, WB23, XWX^{+}22, ZTD^{+}23, AGK^{+}22, AFC^{+}10, BVG11, BW13, FDBH22, GD02, GLD^{+}19, GJ22, GH98, GHZ18, HF06, POC05, Sta03, TZL^{+}02, WZ14, WPGM16, YZ04, ZZV^{+}03, ZJ12, ZWL^{+}18]. arc [BPK^{+}11]. Architectural [JWT^{+}23, XZP^{+}23, CKX^{+}08, DAB15, EKS^{+}10, KW11, LHL10, NSX^{+}11, NHAH03, PKM^{+}11, SSS^{+}08]. Architecture [CFS^{+}18, FHL^{+}18, HSV^{+}22, Lev84, NKK^{+}14, RYW^{+}22, SRX^{+}23, Wes88, YIO^{+}15, AMS03, ASF^{+}13, CTM13, DN02, DHW^{+}11, JTC09, KKSS18, LCOZ^{+}11, LWW08, PLW^{+}07, SM15, SCS^{+}08, WFH^{+}07, WWSR03]. Architecture-scale [YIO^{+}15]. Architectures [HMLB16, ZZC^{+}22, LSA05, LSH^{+}10]. Arcimboldo [HZZ11]. Arcimboldo-like [HZZ11]. arclength [KSH^{+}16]. Arcs [FH93]. Area [NMLH14, XBLZ23, BLD20, HJ11b, MAC22, NMLH11, WR18]. Arithmetic [FBCA23, KLN91, CLSA20, FV96, HSS98]. Arnold [GIF^{+}18, KCSG18]. arrangement [WLY20, YYT^{+}11]. ArrangementNet [HZD^{+}23]. Arrangements [HZD^{+}23, JTMW20, LHVT17a, CLSA20, FRS^{+}12, LHVT17b, MMBM15, SMZ^{+}14, TNWK22, ZGZJ16]. Array [CBYJ23, CSL^{+}23, HSHF10, VLD^{+}13]. Arrays [GPHSH19, JMA06, LKK^{+}16, SMH^{+}11, WJV^{+}05]. Art [PAR21, SLWL23, WCFL22, ZGXF23, AVR^{+}22, HHC18, KYYL08, KL11, LHE^{+}07, LEN09, LCL^{+}17, LZF^{+}19, MP09b, MSSG^{+}21, XZW10, YXFH21, YNS19, APW23]. Art-directable [PAR21]. Artemis [LXJ^{+}22]. arterial [LLZM10]. Articulated [ACP02, AFP^{+}95, TGTL11, TTT^{+}17, VBMP08, ZB94, BBJP12, CCA^{+}12, CZ11, CBL^{+}16, JL11b, LKB22a, LXJ^{+}22, RGL05, TK14, TOK14, WWB^{+}19, XZK^{+}20, YHL^{+}18, ZRLK07]. articulation [DSF22, JPG^{+}14, JMD^{+}07, KS12]. articulations [LAH^{+}21]. artifacts [ARNL05, CHM^{+}12, GRBN09]. artificial [PTSG09]. Artist [BKLP16, BSM88, LFL^{+}23, SSK^{+}11, LRS18, SPJT10]. Artist-directed [BKLP16]. artist-intended [LRS18]. Artistic [BST09, CAA10, NJS^{+}11, RRS13]. artists [SLD17]. arts [SZZK21]. As-Continuous-As-Possible [ZXZL23]. As-locally-uniform-as-possible [AVR^{+}22]. As-Rigid-As-Possible [NI22, NI24, IMH05]. Ascent [CM21]. ASCII [XZW10]. ASE [PGH^{+}22]. assemblable [ACA^{+}19]. assembled [DFZ^{+}17]. assemblies [BDCDA11, JMM09, JHC^{+}21, KTS^{+}14, MYY^{+}10, TZZ21, WSP18, WSP21]. Assembling [DPW^{+}14, GSKJ03]. Assembly [WKMH^{+}23, APH^{+}03, CCA^{+}12, CKGK11, DYY16, FSY^{+}15, FL16, LTT^{+}20, SLR^{+}16, SFCH12, YNW16]. assembly-based [CKGK11]. Assessing [Erl18, SK13]. Assessment [NDD^{+}23, AMMS08, AČMS10]. Asset [LCC^{+}22, ZZZ^{+}22, LKZ^{+}20, LSH^{+}22]. assets [LS02]. assistance [LFTC13]. assisted [BPD09, BPB13, ILB15, PB02, SARW^{+}15, YIO^{+}15]. associated [FCW^{+}17]. asymmetric [CLQW08, VRM^{+}18]. asymptotic [CZXZ14, Jia21]. Asynchronous [GLX^{+}22, HVS^{+}09, AVGT12, BAM13]. athletic [YYVY21]. Atlanta [SSJ^{+}11]. Atlantis [SSJ^{+}11]. Atlas [LFY^{+}19, LPRM02, LVS18, MVD^{+}18, PKC^{+}17, PKCH18]. atlases [CH02, KOWD21]. Atmosphere [Kla87]. atmospheres [WVBR^{+}21]. atmospheric [KMM^{+}17a]. atomic [Bel18]. attack [MLD^{+}08]. Attend [CAV^{+}23]. Attend-and-Excite [CAV^{+}23]. Attention [ASK^{+}22, CAV^{+}23, KKW23, YPG01, CLC14, PCLC16, VPHB^{+}21, XSZ^{+}16, YNL^{+}21]. Attention-aware [KKW23]. Attention-Based [CAV^{+}23]. attention-directing [CLC14]. attention-driven [XSZ^{+}16]. Attentional [ZPW^{+}23]. attenuation [NSJ14, WLHR11, WVBR^{+}21]. attenuation-based [WLHR11]. Attraction [BVF17b, AVF17]. attractiveness [HRZ^{+}13, LCODL08]. Attractors [TFD^{+}18]. Attribute [AZMW21, ZDT^{+}23, LYY^{+}17, TYS09]. Attribute-Aware [ZDT^{+}23]. attribute-based [TYS09]. Attribute-conditioned [AZMW21]. Attributes [KAEE20, LRT^{+}14, OLAH14]. attribution [Ano10]. audeosynth [LYGC15]. Audio [ANBH23, KAL^{+}17, DZS08, EML^{+}18, JMD^{+}17, LLZ18, LLM21, LXC^{+}15, SSKS17, TGD04, YCL^{+}20, ZXL^{+}18]. Audio-Driven [ANBH23, KAL^{+}17, ZXL^{+}18]. audio-visual [EML^{+}18]. Augmentation [SSII18a, JSP17]. Augmented [LNZ^{+}23, SH23, SMPZ15, SSJ^{+}14, YCP16, ALK^{+}17, BP12, CGP^{+}21, GMW16, HK18b, JBM^{+}17, KJS^{+}19, KKW20, LJM^{+}16, LDPT17, LLHY22, MGDB05, MLR^{+}14, MGK17]. Augmenting [BBG^{+}13, RPC^{+}10]. auroral [BWRB05]. Authentic [CSK^{+}22]. Author [Ano85a, Ano90b, Ano92a, Ano93, Ano94, Ano95, Ano96]. Authoring [BBS^{+}13, CGG^{+}17, ENCC^{+}21, PGCG23, PRMG16, SPF^{+}23, AGP^{+}20, CDM^{+}02, GDG^{+}17, KGG^{+}20, MCS15, PTSO15, ZB13]. Authors [Ano82, Ano83, Ano84, Ano86, Ano87, Ano88, Ano89, Ano90c]. AutoCollage [RBHB06]. Autocomplete [PXW18, XCW14, XWSY15]. AutoConnect [KSS^{+}15]. Autocuts [PTH^{+}17]. autodiff [YBAF22]. autoencoder [CKS^{+}17, KCW^{+}18, SYS^{+}21, XLLW20, YI17]. Autoencoders [LPX^{+}19, LXC^{+}17, SHM^{+}18, SMK22]. autofocus [ZMN^{+}19]. AutoHair [CSW^{+}16]. automata [CLM^{+}13, Ols84]. AutoMate [JHC^{+}21]. Automated [ASN^{+}20, Cas91, FZBR16, HK12, KG04, LACS08, LJH13b, SaLY^{+}08, DHL14, NMD^{+}17, POT17]. Automatic [AB89, APS^{+}14, AFP^{+}95, BP07, BPK05, CCL12, CLJ^{+}20, CYW^{+}16, CLW^{+}14, FNO89, GYQ^{+}18, GASP08, GKTT13, HMAM09, HEH05, KSH^{+}14, KAB^{+}10, LHM09, LdPS84, LSZ^{+}22, LYO^{+}10, NAH^{+}18, SWTC14, SNF05, VAZH^{+}09, WYY^{+}14, YZW^{+}16, BJD^{+}12, CSW^{+}16, CXY^{+}15, DK09, DIP^{+}18, HFTF15, HZG09, ISSI16, JBK^{+}12, JTRS12, JHC^{+}21, KC19, LPRM02, LRFN04, LSH^{+}10, LHM^{+}18, LZT^{+}19, LKvK^{+}14, MPBC16, Pel05, PHBC21, RKKS^{+}07, RCOL09, Sha03, WLY20, XLY09, XSTN14, YYT^{+}11, YYTC12, BZL^{+}17, MYH^{+}10]. Automatically [LNLB16, MSQ^{+}18, MAS^{+}16, BKD^{+}08, DIO^{+}12, RMBB^{+}13]. Automating [LLN^{+}14, Mac86, SG91]. automultiscopic [DSAF^{+}13, DDD^{+}14, EDF^{+}16]. Autonomous [CXW^{+}23a, LGHL23, XZY^{+}17, DE05, LXS^{+}18]. Autoregressive [WSML23, LSH^{+}22, VPHB^{+}21]. Autoscanning [GLX^{+}22, XHS^{+}15, WSL^{+}14]. autostereoscopic [MP04, SMG^{+}05]. auxetic [CPSP21, KCD^{+}16]. auxetics [KLPCP18]. auxiliary [YNL^{+}21]. Avatar [DWS^{+}23, HSW^{+}17, ZWS^{+}24, IBP15, XPB^{+}21]. AvatarCLIP [HZP^{+}22]. AvatarReX [ZZZ^{+}23]. Avatars [BZH^{+}23, MPE^{+}23, XBS^{+}22, ZZZ^{+}23, BWS^{+}21, BBG^{+}13, CWW^{+}16, CSK^{+}22, HZP^{+}22, LCR^{+}02, NSX^{+}18, SQRH^{+}16]. AvatarStudio [MPE^{+}23]. AverageExplorer [ZLE14]. Averages [BF01, PBDSH13]. avoidance [KOOP11]. avoiding [Fat09a]. Aware [AGL^{+}22, BGF^{+}23, CBS^{+}22, CKT^{+}23, LTH^{+}23, MJJG18, PBM^{+}22, SFC^{+}23, SWS^{+}22, TB22, TZS^{+}18, WZK^{+}23, WCZ^{+}22, XZP^{+}23, YBMN^{+}23, ZLC^{+}22, ZDT^{+}23, ALL^{+}20, AMG^{+}19, AFTCO07, AS07, BWS^{+}21, BWKS11, BN21, CAO09, CAD19, CPD07, CLMK17, DAD^{+}18, DLSCS08, DRE^{+}12, DWX^{+}21, EMU15, ESZ^{+}17, FFL10, FSGF16, GO11, GYGS22, GLT^{+}21, HPSZ11, HPG^{+}22, HK18b, HWG^{+}13, KE18, KH10, KRK11, KP18, KKW23, LSD^{+}16, LVKS21, LLZ18, LHLY21, LQGY24, LYC18, LWCT14, LWH15, LFJG17, LXS^{+}18, LGG^{+}07, LSC^{+}12, LLR13, MLPP09, NID20, OHHD18, PQW^{+}08, PHK11, PGZ^{+}19, PLR^{+}16, PLKD18, RvBB^{+}03, RNd^{+}07, RAWV08, RGH^{+}22, RVAL09, SLS^{+}07, SNW21, SRB^{+}19, SYM^{+}24, TB21, TSL^{+}16, TFK^{+}03, TAKW^{+}19, VPB^{+}22, WFS^{+}09, WLLS22, WLP16, WWL^{+}19, XWY^{+}09, YWS^{+}11, ZAC^{+}17, ZJMB12, ZQCL19, ZQPM12, ZHS^{+}20]. awareness [SGX^{+}21, XXL^{+}21]. axes [YSC^{+}16, YLJ18]. Axial [PVY90, TAV^{+}10]. Axial-cones [TAV^{+}10]. Axis [CCW93, LLF^{+}20, LWS^{+}15, MWR12, MWRD13, MLS^{+}18, WWWG22, BO04, DWW^{+}18, ERP^{+}19, FZZ^{+}20, MYRD14, MGP10]. Axis-Aligned [MLS^{+}18, MWR12, MWRD13, MYRD14]. Azimuthal [KM17].

B [JNK^{+}23, BS88, BS90, CCL^{+}22, CG89, FW12, GLP^{+}22, Pra89, RLU95, WPL06, YJY23]. B-rep [JNK^{+}23, GLP^{+}22]. B-Spline [BS88, BS90, CCL^{+}22, RLU95, YJY23, CG89, WPL06]. B-Splines [Pra89, FW12]. Back [Fol91]. background [PEL^{+}21, ZWZ^{+}16, ZYQ^{+}14]. backlighting [WLHR12]. backpropagation [VSJ21]. Backward [MEMS06, TJ08]. badly [SBHH16]. BakedAvatar [DWS^{+}23]. Baking [DWS^{+}23, KPWG24]. balance [BBPD12, HMP^{+}08, MZS09, TZCT20]. balancing [PWLSH13]. Ball [Sai89]. ballistic [RP03, SP05]. Band [BBPA15, HC04, LHZ^{+}18]. Band-Sifting [BBPA15]. bandelets [PM05]. Bar [Ols92]. Barbershop [ZAFW21]. barcodes [MLYZ19]. Barrier [HCLK24, CLL^{+}22]. barriers [LHKR10]. Barycentric [DSSS23, BPC16, BLTD16, ZDL^{+}14]. Bas [SKC^{+}14, WDB^{+}07]. Bas-relief [SKC^{+}14, WDB^{+}07]. Base [War92, GDC15, LVS^{+}13]. base-complex [GDC15, LVS^{+}13]. baseball [TAH^{+}04]. Based [AASP17b, ASK^{+}22, AFP^{+}95, BD86, BVF17b, BBPA15, CPAB22, CAV^{+}23, DFM88, Erl18, FHXW22, GNHM15, HSX^{+}22, HWZ^{+}14, HC86, HZE^{+}19, HSV^{+}22, HCW^{+}23, HWP^{+}23, HJS^{+}14, JLF^{+}23, KM97, LYC^{+}22, LXY^{+}23, LZCX19, LVY16, LH17a, LPC22, LSSW19, Mer23, MMHP23, MCY14, NI22, NI24, NBHSB22, ST16, SMH^{+}23, SLGS01, SS00, TB22, US24, WHHY20, WWWZ23, Wes88, XRW^{+}22, XXA^{+}23, YPL^{+}23, YFFA21, YSCL22, YYL^{+}19, YXK^{+}22, YIC^{+}14, ZZT^{+}21, ZZZ^{+}22, ZXZL23, ZZLH23, AHSS04, AvdP16, AASP17a, AVF17, APCO21, AVB08, ATW^{+}17, AG05, ASF^{+}13, AAM03, BLDL21, BBPP10, BP08, BDD11, BC02, BLAE22, BBG^{+}13, BCK^{+}23, BME21, BSHK04, BKR17, BSPP13, BJD^{+}12, BNB13, BD02b, BLdG^{+}16, CWW^{+}16, CH07, CFL^{+}15, CKGK11, CDSHD13, CNX^{+}08, CKIW15, CWL12, CBL^{+}16, CGZ08, CT17, CTM13, CBvdP09, CWW13b, DBG14, DSB^{+}12, DJ17, DS15, DCP14a, DYN03, DKNY08]. based [DDTP15, DFL^{+}15, DCOY03, DYY16, DBB^{+}17, DLKS18, ERB^{+}12, EC96, EVC^{+}15, EHSN20, Erl07, FCA09, FJL^{+}16, FJS^{+}17, FH10, FRS^{+}12, FH04b, FTZ^{+}19, FKN17, GHBCO21, GAA^{+}23, GLC^{+}23, GPCP13, GZ05, GvdPvdS13, GPD^{+}18, GGG^{+}13, GB13, GLA^{+}19, GBFP11, GZ08, GJ22, DJ18a, GMP09, GBC^{+}13, GMHP04, GDG^{+}17, GJZ21, GBK05, GS04, HMS05, HR05, HW16, HGY17, HLW^{+}18, HTG14, HCL^{+}18, HTER04, HRDB16, HPP^{+}18, HLR^{+}17, HMG03, HHC^{+}19, HZW^{+}13, HLG^{+}22, HESL11, IKKP17, IWZL09, JLS^{+}03, JYL09, JL11a, JZW^{+}15, JRPW20, JMD^{+}17, JLWM22, JWJ^{+}14, JTSB16, JZvdP^{+}08, KIL^{+}16, KSB^{+}13, KWR16, KJM10, KCKK12, KRFB06, KTY09, hKPS03, KLM^{+}13, KO11, KWB^{+}15, KNC^{+}08, KLS^{+}13, KEBK05, LWA^{+}12, LK02, LdPS84, Lee05, LLB24, LAD08, LKG^{+}03a, LWP10, LWC12, LWL17, LPL^{+}18, LCL^{+}23, LLX^{+}01, LWO19, LHP05, LYvdP^{+}10, LCL^{+}17, LH17b, LCT19, LSZ^{+}22, LCL^{+}22, LYFD12, LFB^{+}13]. based [MM13, MHM^{+}09, MS05, MTGG11, MPN^{+}02, MSM11, MLH^{+}09, MRA^{+}13, MWRD13, MBT^{+}15, MS04, MWH^{+}09, MGZJ20, MdLH10, MRC05, MHTG05, MZWV07, NSACO05, NKAS08, NDD^{+}23, NFD07, NFJ02, NIR^{+}21, ODGK03, OPOD10, ÖG12, ÖG15, PGK^{+}22, PRP^{+}15, PIC^{+}21, PSN20, PKG06, PAK^{+}19, PALvdP18, PMA^{+}21, PTV^{+}17, PLC^{+}21, PHS^{+}18, QHY^{+}16, QTZ^{+}06, RCL21, RYL13, RDL^{+}15, RMBCO23, RCPO21, ROA^{+}13, SML^{+}12, SS14, SZK15, SGH^{+}22, SS19, SDKN18, SNM^{+}13, SHHD17, SJJ12, SKY^{+}12, SAN23, Sha03, SMZ^{+}14, SACO04, SLMB05, SLWL23, SZT^{+}08, SH08, SSY^{+}04, SKG^{+}12, SaLY^{+}08, SKM10, SV19, SWR^{+}21, SKB^{+}14, SGdA^{+}10, SLW22, SSD09b, SZGP05, Sun06, TK05, TB21, TPSHSH13, TZW^{+}07, TEG18, TWL^{+}18, TYS09, TTD22, TD16, TDM11, TCG^{+}14, TWGT10, TZZ21, TOS^{+}03, VRC^{+}13, VT04, VBK05, VBFG12, VBBF16, VSHJ12, WPC^{+}14, WRDF13, Wam16]. based [WPL06, WZT^{+}08b, WYZG09, WWZ^{+}09, WHRO10, WLSL10, WYX11, WFP12, WHDK12, WHY^{+}13, WMZ^{+}13, Wan15, WZB17, WLG^{+}17, WZK^{+}17, WSLT18, WQLJ18, Wan21, WFS^{+}21, WLZ^{+}21, WZHL23, Wan23, WTS^{+}23, WBLP11, WP10, Wes21, WLHR11, WMP^{+}06, WGH22, WLH^{+}13, WDR11, WZN^{+}14, WBG^{+}16, WWYW21, WWW22, XLY09, XWM^{+}20, XFT^{+}08, XFZ^{+}09, XYH^{+}21, XXK^{+}06, XGC07, XLJ^{+}09, XZW10, XLS^{+}11, XFAT12, XCF^{+}13, XUC^{+}14, XB17, XSHR18, XWZ^{+}21, YI17, YTS^{+}11, YCL^{+}15, YRPF09, YZ04, YZX^{+}04, YT13, ZG04, ZWGS02, ZMT05, ZHLB10, ZM11, ZCW^{+}17, ZZMC13, ZZL^{+}21, ZJ12, ZGW^{+}13, ZIH^{+}11, ZAFW21, ZCX^{+}22, ZDI^{+}15, ZPKG02, dGWH^{+}15, dLMH10, vW02, vFTS06, WCL^{+}23]. baseline [XZZ^{+}21]. bases [DCD15, HTC^{+}14, LDF14, WST09]. Basis [CXW^{+}23b, ASK^{+}12, Coh87, HRV97, SR97, SR00, SSC10, Sze06, TS06, ZM11]. basketball [LH18]. Batch [FHL^{+}18]. Batch-Shading [FHL^{+}18]. Bayesian [DTB06, VKK18]. BD [JP04]. BD-tree [JP04]. Be [BMBRD24, FKI^{+}14, ISSI16, SZC^{+}22]. beadwork [IIM12]. Beady [IIM12]. beam [PKLI^{+}19]. Beams [BSR^{+}23, JWT^{+}23, BJ17, JNSJ11, JNT^{+}11, KGH^{+}14]. beat [DA18, hKPS03]. Beating [CH14]. beautification [Zit13]. before [HXM^{+}13]. before-and-after [HXM^{+}13]. Behave [ZSAF21]. behavior [BBO^{+}10, LP10, SHP04, WT08]. behavior-specific [SHP04]. behavioral [VABW09]. behaviors [JWW^{+}20, MTP12, SKL07, WGH20]. belief [HRL15]. believing [EMO10]. Bellman [dSDP09]. below [WAK20]. Beltrami [NH22]. Benchmark [WFS^{+}21, BLN^{+}13, CGF09, SMGH18, YVG20]. benchmarking [KPZK17]. Bend [XKCB18]. Bend-it [XKCB18]. BendFields [IBB15]. Bending [FHXW22, WYW23]. BendSketch [LPL^{+}17]. bent [GGP^{+}20]. Bernstein [Pat87, Pat85, TTWM14]. Bernstein-Bézier [Pat87, Pat85]. Bespoke [WKMH^{+}23]. Best [McI83, ALS^{+}18]. best-buddies [ALS^{+}18]. Beta [BB83, Joe90a, Joe90b, TB87, Joe89, NCVMO05]. Beta-connection [NCVMO05]. Beta-Spline [Joe90a]. Beta-Splines [Joe90b, TB87, BB83, Joe89]. better [AFSR03, Jam20, ZAE^{+}14]. Between [MPB17a, BDG15, BWS10, CMT04, CFW13, CNR08, ESBC19, GKJ^{+}05, MPB17b, MRF06, OBCS^{+}12, TMY^{+}11, WM14, YM16]. Betweening [QZZ22, HYNP20]. Beyond [BJ17, Csé19, GJZ21, Hac18, KCD^{+}16, RSM^{+}23, TCT23, WKF^{+}21, ZB14]. Bézier [BC14, DeR88, DKA23, Gal99, GPSZ11, KC23, LJG14, LD89, Pat85, Pat87, War92]. Bézier-Splines [DKA23]. Bi [HWB23, LDPT13, MP09c, SLSS03, FW12, IDN12, WDR11, WDR13]. Bi-3 [MP09c]. Bi-directed [HWB23]. bi-Laplacians [FW12]. Bi-scale [LDPT13, SLSS03, IDN12, WDR11, WDR13]. Bias [BB83, SK13]. Biased [GIGM22, MBGJ22]. Bicubic [Fol87, KP07, LM91, LS08]. bicycle [TGLT14]. BiDi [HLHR09]. Bidirectional [NID20, RLU95, WKB12, BNTS07, CRS^{+}16, FCGH08, GYGS22, HP03, HHA^{+}10, KBD07, QZG^{+}19, SOHK16, SLW22, TZL^{+}02, YTS^{+}11, YHCOZ18]. BiggerPicture [WLL^{+}14]. BigSUR [KFWM17]. Biharmonic [IKCM13, LRF10, FW12, JBPS11]. bijections [APL14]. Bijective [CSZ16, HC23, JSZP20, JZH^{+}21, SS15, JSP17]. Bilateral [CGW^{+}13, CAWH16, CLKL14, FDCO03, CPD07, DD02b, GCB^{+}17, KCLU07, Wei06]. Bilinear [ASK^{+}12]. Billboard [DDSD03]. Binary [Kou16]. Binaural [LLM21]. binding [LZT^{+}19]. Binocular [AKG^{+}23, YZWH12, CAD^{+}21, HXFW20, VWB^{+}12]. bio [IZE^{+}21]. bio-inspired [IZE^{+}21]. biological [Sun06]. Biologically [BW22, JWDL19, WHDK12]. biologically-based [WHDK12]. Biomechanical [SSB^{+}15, SLST14, LT06, LST09, NZC^{+}18]. Biomechanically [KWS^{+}23]. biomechanics [WZB17]. biomimetic [NZC^{+}18]. biped [CLS03, CBvdP10, LKL10, LLK^{+}15, SKL07, VSHJ12, YLvdP07]. bipedal [GvdPvdS13, cWP10]. bird [cWP03]. Birefractive [BGK16]. birefringency [WW08]. Bisector [EK98, ZWK14]. bispectral [HHA^{+}10]. Bistable [CPSP21]. bit [BMBRD24]. Bitmap [PMKB23, BB22, GS82, Pik83]. black [LYC18, TYY^{+}19]. black-and-white [LYC18]. black-box [TYY^{+}19]. blackboard [SBLD15]. blackboard-style [SBLD15]. blend [GBC^{+}13, LD13, LAH^{+}21]. Blended [AFL23, KLF11, ZBK18]. Blending [Fil89, NPC^{+}22, RWTT14, Roc89, VCA^{+}22, War89, XLY^{+}22b, ALX^{+}14, ATW^{+}17, HPP^{+}18, KCŽO08, NSS^{+}19]. blendshape [SILN11]. blendshapes [SLS^{+}12]. Blind [BTS^{+}15, YSQS08]. blink [LSL^{+}18]. blink-induced [LSL^{+}18]. Blister [HR05]. Block [MLS^{+}18, YNW16]. Blocking [SLS^{+}16]. Blocks [LW15, CLF^{+}18, LCL06]. Blockwise [KIM^{+}19]. Blossoming [DGHM93]. Blue [ARW22, Fat11, HSD13, JZW^{+}15, MEA^{+}18, QCHC17b, dGBOD12, APC^{+}16, AW20, CGW^{+}13, GWN^{+}03, KTBV16, KCODL06, LWSF10, ODJ04, QCHC17a, SLS^{+}16, SZG^{+}13, Wei10]. blue-c [GWN^{+}03]. Blue-Noise [MEA^{+}18, Fat11, AW20, SZG^{+}13]. Blur [SLL^{+}21a, VMCS15, AXR09, BHR13, BSS^{+}13, ETH^{+}09, HCOB10, HQL^{+}10, LES10, LSR18, WKF^{+}21]. Blur-Invariant [SLL^{+}21a]. blurred [YSQS07]. blurred/noisy [YSQS07]. Boards [FBS^{+}23]. Bodies [BC14, CMT04, CFW13, CPMK21, DBB^{+}17, GBF03, HRZ^{+}13, IGLF06, JTSB16, KEP05, LHLK10, PMS12, RGL05, RTB17, SZK15, WMW15, YKZ^{+}22, ZFL^{+}10]. Body [JPL22, KNK^{+}22, LXY^{+}23, PQF^{+}23, SQRH^{+}16, YZH^{+}23, ACP02, ACP03, BWS^{+}21, CZJ12, CBK20, EMO10, FLS^{+}21, FTP16, GHZ^{+}20, HHC^{+}19, HFG^{+}18, KIL^{+}16, KE18, KP11b, LKL^{+}22, LJ14, LST09, LTK09, LVKS21, LYWG13, LHZ^{+}21, MTP^{+}18, MEM^{+}19, MTA^{+}20, PRMG16, PYA^{+}24, PSE03, SPS^{+}11, TB21, TTL12, Ten20, TBV12, TJ08, VSK^{+}17, WY16, WSJP17, WZC12, WP12, WWW22, ZSZ^{+}14, ZJ10, ZZZ^{+}23, ZBG15b]. body-aware [LVKS21]. Body-mounted [YZH^{+}23, SPS^{+}11]. bodybuilding [SZK15]. BodyFormer [PQF^{+}23]. Bokode [MWH^{+}09]. bone [MK16]. Bones [JS11, LD12, LZQ^{+}22]. Bookmarks [Ols92]. Books [XZM^{+}18]. boolean [AD03, HR05, Man86, RNP^{+}22]. Booleans [CPAL22, TNWK22]. BoolSurf [RNP^{+}22]. Boom [TFK^{+}03]. Boosting [DMB^{+}14]. bootstrapping [DWT^{+}10]. Botanical [WZB17, WLX^{+}18, IOOI05, LKK^{+}21, PSK^{+}12, PJH^{+}17]. Bounce [RSM^{+}23, WSJP17, MDKD16, WJF^{+}22]. Boundaries [BGI^{+}18, BHW16, KGB^{+}09, LFB^{+}13, LCBK19, SS15, TBBC^{+}22, WZHB09, WZ14]. Boundary [ASGS23, BSSJ23, CPAB22, CDY23, DS92, DZCJ21, HTWB11, KC23, LL23, MSCG23, RS98, RV89, SC18a, SMGC23, SV93, SVB17a, SVB17b, SGWJ18, SJWG20, SCJ^{+}23, CCS^{+}21, DF88, HW15, HW16, HDS^{+}18, IKCM13, PTSG09, SKM10, SS17, WAK20, YLB^{+}22, ZLB16a]. Boundary-Conforming [KC23]. Boundary-Dominated [LL23]. Boundary-Respecting [CPAB22]. Boundary-sampled [DZCJ21]. Bounded [CW15, CCW16, CLW16, JBPS11, Lip12, LYP^{+}14, AD03, AL13, BDT99, CWKBC13, FOL^{+}21, KABL15, LW16, LFY^{+}19, PMHD19, ZG02]. bounded-error [BDT99]. Bounding [CB17, CGM11, SHH99, VAZH^{+}09, WBS07]. Bounds [CCK92, LAKL11]. Box [HHX^{+}18, LVS18, CBYJ23, CGM11, JBLL18, SRL^{+}15, TYY^{+}19]. Boxelization [ZSMS14]. boxes [SHH99, ZSMS14]. braided [HML^{+}14]. Branching [GJB^{+}20]. BRDF [BAOR06, BAERD08, CDP^{+}14, EBJ^{+}06, HDMR21, LK02, LRR04, LWL^{+}23b, LKYU12, NJR15, Pet21, RGB16, TUGM22, XNY^{+}16]. BRDF-based [LK02]. BRDFs [BSN16, BLPW14, LGX^{+}13, SZC^{+}07, SJR18, XCM^{+}14, ZZW^{+}22a]. Break [STXJ15]. Breaking [SLM^{+}23]. Breathing [TMB14]. bridge [MRF06]. Bridging [DHL14]. Bright [JGC^{+}15]. Brightness [DGH16, WZC^{+}20]. Bringing [AECOKC17]. bristle [CKIW15]. Brittle [FCK22, HW15, HW16]. Broadcast [ZYM^{+}23]. Brook [BFH^{+}04]. browsing [GJ22, KCSC10, TJ07]. Brush [PF89, CTW09, HTER04, RAR^{+}21]. brushes [DJ17]. Brushstroke [SLF22]. Brute [GIF^{+}18]. Brute-Force [GIF^{+}18]. BSDF [GLH^{+}23]. BSDFs [GHZ18, HHdD16, RBSM19, WJF^{+}22]. BSGP [HZG08]. BSP [GMP09]. BSP-based [GMP09]. BSSRDF [DLR^{+}09, YSJR17]. bubble [BDWR12, KySK10, PCK^{+}19]. Bubbles [HLYK08, XAW^{+}23, DBWG15, GAB20, HIK^{+}20, KLL^{+}07, LZJ16, WFS22]. Bubbling [CPPK07]. buddies [ALS^{+}18]. budget [HHGH13, WYM^{+}16]. Buffer [FF88, BBO91, JLBM05, LCD06]. buffers [CM14]. Build [LZCX19, LSZ^{+}14]. Build-to-last [LSZ^{+}14]. Building [BD86, LW15, LBW^{+}23, MG03, ZJMB11, BYMW13, CLF^{+}18, KGFF14, KGS^{+}18, LCL06, MSK10, YNW16, MRF06]. Buildings [FW16, SW14, MWH^{+}06, WOD09, WSW^{+}12]. bulk [GJZ21, HZG08]. bulk-synchronous [HZG08]. Bundle [WXZ^{+}23]. Bundle-Adjusting [WXZ^{+}23]. Bundled [LYTS13]. BundleFusion [DNZ^{+}17b, DNZ^{+}17a]. bunnies [SBHH16]. bunny [WKHA18]. burr [XLF^{+}11]. Burst [HSG^{+}16, LYT^{+}14]. bursts [LEPM22]. Bush [GM84]. Bush-Trajectory [GM84]. Butterfly [CLT^{+}22, DLG90]. Buzo [DSZ17]. BVH [DFM13, KOY^{+}11]. BxDF [YJB^{+}14]. By-example [DLL^{+}15, LHL10, RRS13].

C [OCNG21, OGN^{+}23, GWN^{+}03, BSR^{+}23, MGAK03]. C-like [MGAK03]. C-Shells [BSR^{+}23]. C1x6 [KKB^{+}11]. Cache [MBK^{+}10, YLPM05, WS99]. Cache-oblivious [MBK^{+}10, YLPM05]. Caching [MJJG18, MSCG23, WTS^{+}23, JDZJ08, MA07, MRNK21, MHC^{+}16, PFHA10, SJJ12, SSM15]. CAD [GLP^{+}22, JHC^{+}21, JNK^{+}23, LPBM20, LPBM22, LGHL23, SXZ^{+}17, WPL^{+}21]. cage [GPCP13, JZvdP^{+}08]. cage-based [JZvdP^{+}08]. cages [BC18, SVJ15, TMB18]. Calculating [MC92]. Calculations [SWZ96]. calculus [ZJ18, dGDMD16]. Calibrated [RPK^{+}12, MKRH11, MYC^{+}22]. Call [Ano85b, Ano92b, Ols88]. calligrams [ZCR^{+}16]. Cam [CSL^{+}22, CSSL21]. cam-follower [CSSL21]. Cam-Linkage [CSL^{+}22]. Camera [GXY^{+}17a, JCW^{+}21, JGN16, PHM^{+}23, PC82, SCCB22, SZD^{+}20, TMM^{+}21, WLS^{+}23, CZL^{+}15a, FKI^{+}14, FSH^{+}06, GSH18, GRBN09, GXY^{+}17b, HST^{+}14, HGG^{+}11, HOM15, JWW^{+}20, JMA06, JRT^{+}15, LSC^{+}22, LKK^{+}16, LD21, LFDF07, LC15, LYTS13, MRK^{+}13, MSS^{+}17, MWH^{+}09, MDB^{+}19, OHB^{+}11, PCPW20, PRAV09, RTF^{+}04, RAWV08, SMG^{+}20, SXZ^{+}12, SLL19, SHHW16, VLD^{+}13, VCA^{+}22, WGJ^{+}18, WSXC16, WZC12, WLM^{+}15, WJV^{+}05, WSVT13, XYH^{+}18, YPL21, ZWW^{+}18, ZZZX21, ZNI^{+}14]. camera-in-the-loop [PCPW20]. Cameras [CKH18, CSL^{+}23, DPW15, LR15, YLC^{+}20, APS^{+}14, CWL12, HSG^{+}16, KWB^{+}13, KWR16, LHG^{+}09, RRC^{+}16, RH16, RZK11, SPS^{+}11, TAV^{+}10, VRA^{+}07, WFDH18, WZN^{+}14, ZSZ^{+}14, ZK14]. Camouflage [CHM^{+}10]. CamP [PHM^{+}23]. Can [BMBRD24, BDM^{+}20, SC20, SZC^{+}22]. Candid [FAC11]. Canonical [VMW18, FKY08]. canvas [SSGS11]. Canvases [BCV^{+}15, YBMN^{+}23]. CAP [SMPZ15, DHB17]. Capacity [BSD09, XLC^{+}16]. Capacity-constrained [BSD09]. Capture [BBO^{+}09, CPY^{+}22, DXG^{+}23, FJA^{+}14, GPHSH19, HXZ^{+}19, HTCH15, LSM23, PBS04, SBSH18, SGPT23, XCZ^{+}18, YZH^{+}23, AWL13, AWL15, Ari06, AIH^{+}08, BGKS17, BBB^{+}10a, BHB^{+}11, BBN^{+}14, BBGB16, BBA^{+}07, BPS^{+}08, BHPS10, CBZB15, CWZ^{+}21b, CLS03, DAD^{+}18, DWT^{+}10, DKD^{+}16, DDF^{+}17, FKI^{+}14, GFT^{+}11, GITH14, GSH^{+}20, Hol18, HMLL14, HCTW11, ITM^{+}14, JCRA11, KCW^{+}18, KP06, KN06, LMB14, LLR13, MBPY^{+}18, MCE^{+}17, MPH^{+}20, MRC05, NZV^{+}11, PRMG16, PMPHB17, PB02, RNd^{+}07, RRC^{+}16, SMP03, SLH^{+}20, SGXT20, SGX^{+}21, SPS^{+}11, SNF05, TFK^{+}03, VWB^{+}12, VPB^{+}18, VAV^{+}07, VPB^{+}09b, VSHJ12, WMZ^{+}13, WWY^{+}15, WZK^{+}17, WZC12, WZC^{+}22, WSVT13, WBGB16, XWW^{+}14, ZSCS04, ZN06, ZSZ^{+}14, ZMCF05, ZBGB19, dAST^{+}08]. captured [BBPP10, Leh07, YZL^{+}22]. Capturing [AHM^{+}15, ASN^{+}20, CPMK21, CZB23, EBGB14, HML^{+}14, JMM09, KUDC07, PH06, PNDN12, WCF07, Zho18, BDCDA11, BLCD02, DBDB11, LRAT08, RTB17, TMB14, VWJ^{+}13]. Cardinality [MS13]. Cardinality-constrained [MS13]. caricature [CLY18, HGY17, JJJ^{+}21]. CariGANs [CLY18]. CARL [LSCC20]. Carlo [AW20, ALLD17, BVM^{+}17, BAGL19, CKS^{+}17, CGMS22, CHY21, DMB^{+}14, GLA^{+}19, GHZ18, HET^{+}14, HRV^{+}18, IMF^{+}21, JM12, KBS15, LADL18, McC99, OKH^{+}17, PSC^{+}15, RAMN12, RLSÖ^{+}22, RMGH15, SGH^{+}22, SSJC22, SMGC23, SHHD17, SD12, SWZ96, SJ17, YNL^{+}21, ZSGJ21, ZDDZ21, ZZXY21]. Carpentry [ZWZ^{+}22]. Carry [MTA^{+}20]. cartography [TBW^{+}12]. cartography-intrinsic [TBW^{+}12]. Cartoon [BCV^{+}15, ZWL22, BOD^{+}13, DLKS18, RID10, WDAC06]. cartoons [BLCD02, WWH06]. carve [MAYZ^{+}20, ZZX^{+}18]. carving [AS07, DZPZ09, FHM^{+}21, RSA08, SSZCO10].

Cascaded [HLR^{+}14, PCI^{+}21, WLT16].

cascading [SZT^{+}07]. case

[McK87, PRZ17, SZB18, ZPZ13]. Cases [EM90]. Casteljau [Pra89]. casting [AMB^{+}21, KGB^{+}09]. Casual [AECO15, HASK17, BYLR20, DSC^{+}20, HWV^{+}18, TT09, ZMN^{+}19]. casually [BBPP10]. CAT [HGRT04]. Catacaustics [KLR^{+}22]. catadioptric [KN06, TAV^{+}10]. catadioptrical [NYY04]. catalog [BUSB13]. catalogue [DFL^{+}15]. cataracts [PPZ^{+}11]. Catch [MTA^{+}20]. catching [MLH^{+}09]. Categorical [ZZLH23]. Catmull [DB88, LFS16, LJG14, LS08, MRF06, NLMD12]. Catmull-Rom [DB88]. CATRA [PPZ^{+}11]. Cauchy [LCK22]. causal [RCLM19]. causality [HMO12]. caustic [MMT18, STTP14]. Caustics [YIC^{+}14, GSLM^{+}08]. CD [WFL^{+}19]. CD-MPM [WFL^{+}19]. cel [LMY^{+}13]. Cell [BC23, WCZ^{+}22, AA06, CMSA20, CM11, FGG^{+}17, JSS^{+}15, QLDJ22]. Cell-Controllable [WCZ^{+}22]. cellular [HSF07]. Center [TFD^{+}18]. centered [GB08a]. centers [LH16]. Centimeter [BWC^{+}23]. Centimeter-wave [BWC^{+}23]. Centric [GWBN24, ELFS16, FSL^{+}15, KCGF14, RCCO22, ZXL^{+}18]. Centroidal [XLC^{+}16, KLV20, LWL^{+}09, LXY^{+}16, LL10]. CFL [WLF^{+}20]. CFL-Rate [WLF^{+}20]. Cg [MGAK03]. Chain [JM12, YYL^{+}19, GLP^{+}22, OKH^{+}17, RCLM19]. Chain-Based [YYL^{+}19]. chaining [XYH^{+}18]. Chainmail [TCT23]. Chains [FHG^{+}23, Gol84, Gol85a]. challenging [DKD^{+}16]. chameleon [TFK^{+}03]. chandeliers [PCK^{+}19]. Change [CM21, GSP^{+}23, BW13, SSJ^{+}14, SXH^{+}21, ZPBK17]. Changes [TD23, DFW20, HRvdP04, KBC^{+}13, WM14, WTGT10, WRS^{+}12]. changing [MBF04, PH15a]. channel [HLR^{+}17, WYL^{+}20]. Character [ANL^{+}23, BCV^{+}15, BVF17b, Cor18, DSF22, EHSN20, HDK07, HTCH15, LZCV20, MSK^{+}23, WAH^{+}10, WZC^{+}20, XXA^{+}23, ZZLH23, AWL^{+}19, AVF17, BB22, CKP^{+}21, DYP03, GCR13, GRGC15, HYL12, HKT10, HSK16, HKS17, IWZL09, JPG^{+}14, JMD^{+}07, KS12, KHKL09, LLP09, LWB^{+}10, LLL18, LMLL21, LWH^{+}12, LWS02, LP02, MZS^{+}11, MMG06, MG03, PALvdP18, PMA^{+}21, RP03, RP07, RTK^{+}15, SH08, SKSY08, SZKZ20, TBvdP04, TLP07, VGB^{+}14, WLO^{+}14, WGH22, YL10, dSDP09]. character-agnostic [AWL^{+}19]. Characterization [CSBC^{+}17a, CO19, CSBC^{+}17b, RZK11, SMCT18, SD89]. characterizations [CI97]. characterizes [ZCL18]. Characterizing [FSH11b]. Characters [LVY16, LH17a, PAR21, YSCL22, BBJP12, BP07, BBS^{+}13, BVS16, BDI^{+}02, CBL^{+}16, CBvdP09, CTN^{+}13, DE05, EAPL06, FBH21, HLX^{+}21, HXK^{+}19, JL11a, JL11b, JSMH12, JHS12, KP11b, KLF^{+}19, LYWG13, LH17b, LZT^{+}19, MP07, MLPP09, MPP11, PGH^{+}22, STC^{+}13, SGdA^{+}10, SDO^{+}04, SKC^{+}14, TWH^{+}22, TCG^{+}14, WGH20, WGH21, XLS^{+}11, XKCB18, XZK^{+}20, YL08]. Charades [ANL^{+}23]. Charcoal [BSM88]. CHARMS [GKS02]. chart [BHMK^{+}18, GP09]. Charted [Pan17]. Charter [Fol94, Fol95b]. Chebyshev [Wan15]. Chemomechanical [HIK^{+}20]. Chen [YXH14, XW09]. chi [LLZ^{+}20]. chi-squared [LLZ^{+}20]. Chief [Bea91]. Children [SZL^{+}23]. Chimeras [LLL22]. Chinese [XXK^{+}06]. choices [HFF16]. Cholesky [CSHD21, HLSO12, HA18, HSH20, LLKC21]. Chopper [LBRM12]. chopsticks [YYL22]. Chordal [CLJL20]. Choreography [LDD^{+}23, CTL^{+}21]. choreography-oriented [CTL^{+}21]. ChoreoMaster [CTL^{+}21]. ChromaBlur [CLS^{+}17]. chromatic [CLS^{+}17, GKJ^{+}05]. Chromium [HHN^{+}02]. CIELAB [HRV97].

Cinema [EDF^{+}16]. Cinemagraphs [MSL^{+}23]. cinematic [HPB06, PTG02]. Cinematographic [GLC^{+}18]. Cinematography [ASN^{+}20, JWW^{+}20, NMD^{+}17, PVL^{+}05]. Circle [PF89, KSS06]. Circle-Brush [PF89]. Circles [McI83, MST89, SHWP09, Bak94]. Circuit [FBS^{+}23]. Circular [BPK^{+}11]. Circularly [GCP^{+}10]. circulation [DBWG15, ETK^{+}07]. circulation-preserving [DBWG15, ETK^{+}07]. City [LWL17, XFZ^{+}09]. City-scale [LWL17]. Clark [LFS16, LJG14, LS08, MRF06, NLMD12]. Class [Ree83, SGSS22, Yuk20, PKLI^{+}19, SKB^{+}21, Wei10]. Classes [CQS^{+}23, LZZ^{+}21, SS10b]. Classification [Jan91, JTMW20, CSHH21, ISSI16, Man86, ST14, TTWM14]. classification-driven [ST14]. classifications [WYXJ21]. classifiers [BWSS09]. classify [NXS12]. clean [NHS^{+}13]. cleanup [SSISI16, YVG20]. Clearance [Kal14]. Clebsch [CKPS17, XWWZ22, YXZ^{+}21]. climate [PMG^{+}22]. climate-response [PMG^{+}22]. climbing [NRH17]. clip [AVR^{+}22, LHE^{+}07, LEN09, LLLL21, AZL23, BAC^{+}23, GPM^{+}22, Mir98]. clip-art [AVR^{+}22]. CLIP-guided [BAC^{+}23, GPM^{+}22]. CLIPasso [VPB^{+}22]. Clipless [LAKL11]. clipmaps [LH04]. Clipped [BXH^{+}18]. Clipping [ABE^{+}20, LB84, Mai92, GH98]. cloaking [SBB^{+}22]. Clone [MLD^{+}08]. cloning [BKS^{+}12, FHL^{+}09, LSC^{+}12, SLS^{+}12]. Close [CPS15, SYZ^{+}23, FKI^{+}14]. close-range [FKI^{+}14]. Close-to-conformal [CPS15]. Closed [LM91, PFX^{+}22, BWSS12, FXBH16, JSW05, YZL^{+}22, vW09]. closed-form [FXBH16]. Closed-loop [PFX^{+}22]. Closest [KTT13, KC21]. closing [SKSJ20]. Closure [LWH15]. Closure-aware [LWH15]. Cloth [BME22, HSX^{+}22, LKL23, LDW^{+}23, WWYW21, ZDF^{+}22, AJM12, BWK03, BFA02, CFW13, CK02, CLMMO14, FYK10, GHF^{+}07, IM12, JGT17, KJM08, KJM10, KGBS11, KKN^{+}13, LWS^{+}18, LDN^{+}18, LTT^{+}20, MTB^{+}13, NSO12, OKRC10, RPC^{+}10, SBdDJ13, SNW20, TJM15, TWL^{+}18, VMTF09, WOR11, Wan21, WPLS18, WCF07, WWW22, ZLB16b, LTT^{+}20, TWL^{+}18]. ClothCap [PMPHB17]. ClothCombo [LKL23]. Clothed [KNK^{+}22, TWR^{+}23]. Clothed-Human [KNK^{+}22]. Clothes [LKL23]. Clothing [CPY^{+}22, IH03, XBS^{+}22, BRB^{+}19, HTC^{+}14, PMPHB17, WHRO10, XPB^{+}21, XUC^{+}14, YKJM12, dASTH10]. clothoids [CBD13]. Cloud [HZC^{+}22, MSQ^{+}18, MHGCO21, WTS^{+}23, Che13, DKNY08, FSP^{+}22, GSC^{+}15, HMP^{+}20, HWCO^{+}13, TZCO09]. Cloud-native [WTS^{+}23]. Clouds [HLP^{+}22, LXSW23, WSL^{+}19, WZX^{+}23, Wan23, XDW^{+}23, BDS^{+}18, DDSD03, DIO^{+}12, GAF^{+}10, HRV^{+}18, HLZ^{+}09, KMM^{+}17a, KL22, LGB^{+}21, LYO^{+}10, MHZ^{+}21b, WPL06, WNEH22, XTZ^{+}21, YC21, YHZ^{+}14]. cluster [WWLC21]. Clustered [SHHS03, Tsa15, TWZ22, TS06, TS12]. Clustering [CLSS97, KT03, SvKK^{+}11]. clusters [HHN^{+}02, VLV^{+}21]. cluttered [NXS12]. CNC [BBR^{+}21]. CNN [CT17, LSQ^{+}15, WLG^{+}17, WSLT18, WSL^{+}19]. CNN-based [CT17]. CNNs [EKD^{+}17]. Co [AGL^{+}22, HLV^{+}17a, HLV^{+}17b, YZX^{+}18, YK12, YK14, ZWZ^{+}22, ZZL^{+}23, vKXZ^{+}13, BAS14, HvKW^{+}16, KKB^{+}11, ML22, SvKK^{+}11, WAvK^{+}12, XCF^{+}13]. Co-abstraction [YK12]. Co-Analysis [YZX^{+}18, HvKW^{+}16, WAvK^{+}12]. Co-constrained [YK14]. Co-Decomposition [ZZL^{+}23]. co-dimensional [ML22]. Co-hierarchical [vKXZ^{+}13]. co-located [KKB^{+}11]. Co-Locating [HLV^{+}17a, HLV^{+}17b]. Co-Optimization [ZWZ^{+}22]. co-placement [XCF^{+}13]. co-representation [BAS14]. co-retrieval [XCF^{+}13]. co-segmentation [SvKK^{+}11]. Co-Speech [AGL^{+}22]. Coaching [HL14]. Coarse [WYXJ21, EB14, JZH^{+}21, LZF10, RPC^{+}10, SDW^{+}16]. Coarse-to-fine [WYXJ21, SDW^{+}16]. Coarsely [CCK^{+}21]. coarsening [CLMK17, CBW^{+}18, CLJL20, FCA09, GAB20, KMOD09, LJO19, TREO16]. coaxial [HLZ10]. cocktail [EML^{+}18]. Code [HTS^{+}22, GKK^{+}21, HBD^{+}14]. Coded [GWGB10, KWB^{+}13, RAT06, SZD^{+}20, CZN10, LFDF07, VRA^{+}07]. Codes [GWBN24, CCLM13, Kan15]. Codimensional [LKJ21, WJL^{+}20, ZQC^{+}14, ZLQF15]. Coding [GVNB18, LCD^{+}19, CRG^{+}20, ORK12, PK05, RS14a, RS18]. coefficient [ZF03]. coefficients [SSJC22, WR18]. CofiFab [SDW^{+}16]. cognitive [MCS15, SSRB^{+}17]. coherence [HZ82, WFS^{+}09]. Coherent [GLHL11, KDMF03, KP11a, LBP^{+}12, RKB^{+}23, WKMH^{+}23, YCZ11, ASC^{+}14, HTG14, HKAK16, LLV^{+}12, RSI^{+}08, WIK^{+}06, WHSL11, XFCT18]. Cohomology [YNW^{+}23]. cold [GGP^{+}20]. Collaborating [JNK^{+}23]. collaboration [KKB^{+}11]. Collaborative [GLX^{+}22, CSTP16, DXZ^{+}19, SSTP15, TGY^{+}09]. collage [HZZ11, KSH^{+}16]. Collapsing [BC23]. collection [HZG^{+}12, SW85]. Collections [SSB^{+}17a, FAR07, FvKBCO16, HSGL13, HSG13, HSS^{+}13, HWG14, HWK15, HLW^{+}19, HOM15, KLM^{+}12, KLM^{+}13, KOC^{+}22, LBP^{+}12, OLGM11, PRFS18, SSB^{+}17b, SSS^{+}08, SSS06, SHM^{+}14, TKKT12, WSH^{+}16, XMZ^{+}14, YKC^{+}16, YK12, YK14, ZLE14]. Collision [LLF^{+}20, VJ19, WFS^{+}21, WCL^{+}23, WG09, BJ10b, BEB12, GKJ^{+}05, Hub96, JP04, KOOP11, LHLK10, Mir98, MASS15, MCKM15, SPO10, TMY^{+}11, TTWM14, TWL^{+}18, Wan14, WLLS22, WLH^{+}13, ZRLK07, ZJ12]. collision-aware [WLLS22]. Collision-free [WG09]. Collision-Ready [LLF^{+}20]. Collisions [OD01, WWYW21, BFA02, HVTG08, KTS^{+}14, MZS^{+}11, MTM16, VMT06]. Color [AAPS16, AASP17b, BVF^{+}17a, BAU15, CKT^{+}23, COSG^{+}06, DCT^{+}22, Fat14, GW90, HLC^{+}19, HDC07, KP92, LR90, LR91, MSM^{+}23, NRS15, OAH11, PCB23, PH15a, PH15b, SCB87, SFB92, SCB88, Sto92, WZHL23, WC90, WC91, Wu92, Xia97, ZZL^{+}21, ZK14, AAPS17, AASP17a, AHB18, BCN08, BSPP13, BATU18, CGZ08, DK99, GOTG05, HSGL13, HCE03, HRV97, ISSI16, JSB^{+}10, KYS^{+}15, KWK09, KJDL09, KRK11, KC21, KL12, LRFH13, LDS02, MHP^{+}19, SMH^{+}11, SMHW16, SLS^{+}16, SLD17, TOS^{+}03, WP09b, WYW^{+}10, WYX11, WAM02, ZRL^{+}09]. color-by-numbers [LRFH13]. color-difference [HRV97]. color-filtered [BCN08]. Color-Lines [Fat14]. Color-Perception-Guided [DCT^{+}22]. color-to-gray [KJDL09]. Color2Gray [GOTG05]. Colored [SMB^{+}19, BGB^{+}05, DLKS18, LD06]. ColorfulCurves [CKT^{+}23]. colorimetric [LDS02]. Colorization [HZL22, LLW04, XHWW22, CZG^{+}11, HCL^{+}18, ISSI16, LWQ^{+}08, QWH06, ZZI^{+}17, ZLW^{+}18]. colorizations [LRFH13]. colors [YKH10]. column [HPB07, HWH^{+}16]. CoLux [Par17]. combination [Ale02, dSDP09]. combinations [HR05]. Combined [OKH^{+}16]. Combiner [GIGM22, BHHM20]. Combining [BWG03, DKH^{+}10, JASR99, PS04, CWZ^{+}21a, CGG^{+}17, DSB^{+}12, EB14, HP17, NRDR05]. Combustion [NBHSB22, PJH^{+}17]. comfort [DMHG13, KBBD17]. comics [KL12]. commands [LPBM22]. Comments [Pav90, WP90]. Commodity [YLC^{+}20, CM14, GM05, HDGN17]. Commonsense [JWD^{+}23]. Communication [ANL^{+}23, GWBN24, Hil86, JGN16]. Compact [BKGK17, JBY^{+}19, LLP09, SKOA14, ZZW^{+}22a, ACSM12, GLLR11, HNB^{+}06, KCYW13, MC12, PvBM^{+}06, ZCR^{+}16]. companding [LSA05]. Comparative [HRV97, RGSS10]. compare [KC21]. Comparison [BBB^{+}93, LC96, SHG^{+}22, SCB87, EKA84, HXZW20, KKN^{+}14, MGT^{+}03]. compatibility [LHLF15, OAH11]. Compatible [SG01, Tak22, WLY^{+}16, KS04b, MEMS06]. compensate [POAR12]. compensated [ZRL^{+}08]. compensating [WM14]. compensation [BHW13, SRB^{+}19]. competitive [BDM^{+}20, WGH21]. compilable [LKJC21]. compilation [LS02]. Compiler [YXK^{+}22, HLY^{+}21, JSRV22, MAN^{+}16]. Compilers [LNI^{+}23]. compiling [HBD^{+}14]. Complement [CZY17b, PAK^{+}19, CZY17a, LMAS16]. Complement-based [PAK^{+}19]. Complementary [BZC^{+}23, ZBLJ20]. Complementme [SSK^{+}17]. Complete [JTMW20, RHJD18]. Completion [CPW21, DLP^{+}23, ASK^{+}05, DCOY03, FMLW14, HTG14, HE07, HYG^{+}13, HKAK14, HKAK16, HWH^{+}16, ISSI17, KKDK12, LLV^{+}12, SACO04, SYJS05, SKAG15]. Complex [BYRN17a, CKSV23, DBP^{+}15, HJS^{+}14, SW14, VADWG15, BYRN17b, BAOR06, CAC^{+}02, CJAMJ05, DDP02, EHDR11, EMF02, FGBP11, GGN18, GDC15, GM05, GLY^{+}03, GLP^{+}22, JBP06, JSP17, KH06, KTS^{+}14, KBT17, KSSCO08, LZJ16, LDS03, LRA^{+}07, LTT^{+}20, LP02, LVS^{+}13, MZD05, MB12, MTP12, PMS12, PKZ04, RSM^{+}10a, RBF08, SS14, SILN11, SWF^{+}21, TGD04, WSP21, WM03, YMR^{+}13, ZBX^{+}21]. Complexes [BC23, PBCF93, AA06, DRvdP14, GD02, ZQC^{+}14]. ComplexGen [GLP^{+}22]. complexity [CI84, ME05]. Compliant [DTPC23, ZAB21, MZB^{+}17, TZCT20]. component [KCKK12, SSK^{+}17, YWS^{+}11]. component-based [KCKK12]. Components [WLZ^{+}21, DYY16, HFH^{+}17, NKGR06, NVW^{+}13, SHHS03, SFWG04, WZF^{+}18]. Components-based [WLZ^{+}21]. composable [FH11]. compose [Rit18]. composed [YZL^{+}22]. Composing [DeR88]. Composite [MPP11, XSZK23, AMG^{+}19, CSSL21, SPSH^{+}17, WMZ^{+}13, ZKBT17]. composites [XADR12]. Compositing [Duf17a, KSH^{+}14, Aga07, BSS^{+}11, BPB13, CGC^{+}03, DWT^{+}02, Duf17b, HLR^{+}17, RGF^{+}20, SGW06, YTBK11, ZAFW21]. Composition [DGHM93, GXSD23, LM97, BGKS17, CLC14, GB08b, HGCO^{+}12, LYvdPG12, ZJ18, ZXC^{+}18]. Compound [TMM^{+}21]. comprehensible [BF08]. Comprehensive [LST09, JdJM14, JNSJ11]. Compressed [MHU19, SLM^{+}17a, NNSM07, SLM^{+}17b, WYL^{+}14]. Compressible [CCL^{+}22, LBW^{+}23, GHB^{+}20]. Compressing [LSA05]. Compression [Ari06, BIP01, HZC^{+}22, MHU19, MM22, SILN11, SWWW15, VSW^{+}23, AFSR03, BCG05, FLW02, GD02, IG03, LAJJ14, LD13, LVGO21, MEMS06, MCHAM06, Nah20, PM05, RAI06, TDL^{+}18, TR98, WCSC22, YGM97]. Compressive [ITM^{+}14, MWH^{+}13, MWBR13, PML^{+}09, HWRH13, HWR14, LLWD14, WLHR12]. Computation [JCY23, PM95, PVY90, VKW^{+}23, VMKK00, WJZL08, DZCJ22, FBC18, FHM^{+}21, GSCO12, GS85, GJZ21, HZ82, ILSS06, JTL^{+}12, LK02, LFH15, LWL^{+}09, MIB15, PSBM07, QHY^{+}16, RGK^{+}08, She13, SGG^{+}06, TLK09, TK14, WCSC22, XLC^{+}16]. computation-efficient [WCSC22]. Computational [AHB18, BGKS17, BAD10, BM07, BLT^{+}15, CWSB22, CTN^{+}13, DSZ^{+}16, FGN84, FSY^{+}15, GJG16, GGP^{+}20, GA20, HGG^{+}11, IWHH20, JMZ^{+}22, KGL^{+}22, LDTA17, LZF^{+}19, LXG^{+}22, MZL^{+}17, MLB16, MDKD16, OKH^{+}16, PIC^{+}21, PKPP21, PYB^{+}16, PRM14, POT17, PDF^{+}22, PCB23, RRMG10, SZK15, SPG^{+}16, SHHW16, STC^{+}13, SWT^{+}17, SZ15, TKG^{+}23, TCT23, TCG^{+}14, VRP^{+}23, WHG84, WCFL22, XZM^{+}18, YCC17, ZYZZ15, ZFS^{+}19, ZAB21, ZBJ^{+}23, ZGXF23, AJD^{+}10, AMG^{+}18, BPK^{+}13, DYN03, DKNY08, FV96, Fre16, HRH^{+}13, HWBR14, HPK^{+}17, JWI^{+}21, KCD^{+}16, KPM^{+}17, KSS^{+}15, KS11, LHG^{+}09, LLMZ16, MDZ^{+}21, MPI^{+}18, MZB^{+}17, OHR14, STTP14, WFDH18, XKF^{+}18, XDF^{+}19, XRLF15, ZHPY21]. computationally [KTY09]. computations [WJF^{+}22]. compute [LMAS16]. Computed [SSW^{+}23, Bae18, IYYI14]. Computer [BG89b, CT82, Coo86, Gol84, Gol85a, Hil86, KP92, MSK10, MRC^{+}86, Pav90, SMPZ15, SLGS01, WP90, Ano03, AČMS10, Gol02, HCW15, ILB15, KFS13, PVL^{+}05, RLR^{+}21, SHL^{+}17, TL04, WQLJ18, WQF^{+}21, YGM97, ZAJ^{+}15]. Computer-Aided [BG89b, Gol84]. computer-assisted [ILB15]. computer-controlled [Ano03]. Computer-generated [MSK10, WQF^{+}21, ZAJ^{+}15]. Computing [ACP^{+}01, BHK14, CCW93, DLSCS08, DEM96, FOL^{+}21, FCJ07, FLG15, FL16, GOMP98, HBLM11, LWS^{+}15, LFO^{+}22, LPS^{+}13, PYW14, PV06, SS19, WC21b, WWWG22, XWX^{+}22, YYL^{+}19, YXW^{+}23, ZWL^{+}18, BFH^{+}04, CWW13b, OK10, PNH^{+}14, SCS^{+}08, WGS23, YPB16]. concatenated [KDH22]. concatenative [AJM12]. concavity [WLLS22]. Concept [IBB15, LB84, VVCOS23, GHL^{+}20, SBSS12, SLZ^{+}13]. Concurrency [Hil86]. concurrent [BSL12]. condensation [TMDK15]. Conditional [AZMW21, ALY^{+}21, GDG^{+}17, WGH22]. Conditioned [LXZ^{+}23, ZWS^{+}24, AZMW21, CSHD21]. Conditions [BS88, SMGC23, SGWJ18, BBPD12, KO11, MKRH11, MAF^{+}09]. Conduction [BBC^{+}23]. Cone [Lev23, LFZ^{+}23, SSZCO10, LSVT15, SSC18, WSP21]. cone-joints [WSP21]. Cones [CSZZ20, NCB23, FOL^{+}21, LFO^{+}22, TAV^{+}10, Van06]. conferencing [KPB^{+}12]. configurable [Pel05]. configuring [RvBB^{+}03]. conflation [WJ19]. confocal [LCV^{+}04]. Conformal [GA20, LFZ^{+}23, SSP08, VMW15, CCS^{+}21, CPS15, CPS13, FOL^{+}21, GSC21b, KSS06, LPRM02, LFO^{+}22, SSC18, WG10]. Conformation [BGFAO17]. Conforming [Ale20, KC23, ACA^{+}19, HGCO^{+}12]. congruent [AMCO08]. Conic [Pav83, PK83, Pot91]. conical [LPW^{+}06]. conics [Far89]. Conjoining [NSX^{+}11]. conjugate [BFGS03, LXW^{+}11]. Connect [LKvK^{+}14]. Connect-The-Dots [LKvK^{+}14]. Connected [ZGH^{+}16, ICG17]. Connecting [SJR18, GITH14]. Connection [LTDD16, BWS10, GKDS12, NCVMO05]. connections [PVG19, SLW22, TH19]. Connectivity [PZKW11, GLLR11, YLL^{+}22]. connectors [KSS^{+}15, LOMI11]. conquer [Mor11]. conservation [KUJH21]. conservative [ANZS18, QZG^{+}19]. conserving [CKMR^{+}21, ISF07]. Considerations [VW94, VW95]. Consistency [RO94, SAA^{+}21, BTS^{+}15, HSGL13, LWA^{+}12]. Consistent [ACBCO17, DNZ^{+}17b, LYO^{+}23, QLH^{+}22, RSM10b, XDW^{+}23, ZCT^{+}21, ASL^{+}17, CRA11, DNZ^{+}17a, DDTP15, ENCC^{+}21, HZG^{+}12, ISSI17, KOWD21, LLJ22, LCK^{+}14, MBGJ22, SLL^{+}21b]. consistently [LWC^{+}11]. consolidating [LRS18]. Consolidation [HLZ^{+}09, LABS23, MHGCO21, WHG^{+}15, ZSW^{+}10]. Constant [DLW^{+}22, MU22, WHHY20, PCL^{+}12, VSJ21]. Constant-Cost [DLW^{+}22]. Constrained [BR94, DPVA23, KUJH21, MVH^{+}17, SW18, SCD^{+}21, WLJ^{+}22, BSD09, CCGB22, CBYvdP08, DKZ^{+}21, KSG03, LFO^{+}22, LZC^{+}18, MS13, MZ13, SJLP11, TBTS08, TNGF15, WBGB16, YYPM11, YK14, ZJL14, ZHCJ15]. Constraining [SWW^{+}20, YCP16]. Constraint [BCK^{+}23, BD86, CH07, GAB20, Sha03, BML^{+}14, HK12, JASR99, KHD14, SAZK06, WG09]. Constraint-Based [BD86, BCK^{+}23, CH07, Sha03]. constraint-solving [JASR99]. Constraints [FH97, Gol84, KF93, RHW94, SW14, TQ94, UZB^{+}23, AFC^{+}10, BGFAO17, HSG^{+}19, HZ82, IOOI05, JTCW07, KOOP11, ML22, SvTSH14, XLC^{+}16, YL08, YYW^{+}12a]. ConstructAide [KGFF14]. Constructing [LFXH17, MHS^{+}19a, KSG03]. Construction [AFH20, FG90, HJS^{+}14, LMAH^{+}18, LFZ^{+}23, MMT23, SH07, SB95, WLY^{+}16, BO04, BLTD16, CGG^{+}04, DS15, DKP11, DFM13, FZLM11, IIM12, KGFF14, LXFH15, LVS^{+}13, WWT^{+}06, WG09, WPL^{+}21, XK07, YZ04, ZM11, ZHWG08, vTSSH13]. Constructions [DB88]. Constructive [CCK92, DZCJ21, FH97, JASR99, LDF14]. Constructor [VKJ^{+}17]. Consumer [CKH18, LWCT14, WZN^{+}14, ZK14]. Contact [Erl18, IRWP23, KL17a, LFP21, LDW^{+}23, MHNT15, MLPP09, PAK^{+}19, RCCO22, TB22, TFD^{+}18, XLYJ23, AVGT12, AFC^{+}10, BLT^{+}15, BFA02, CKMR^{+}21, DJBDDT13, GHZ^{+}20, GHF^{+}18, HVS^{+}09, JTL^{+}12, JGT17, JLF^{+}09, KJM10, KL17b, KSJP08, KP03, LYK^{+}21, LLJ^{+}11, LDN^{+}18, LFS^{+}20, LKJ21, LYvdP^{+}10, LCBD^{+}18, LJBBD20, MZS^{+}11, MTP12, MWTK13, PRWH^{+}18, RCPO21, RLR^{+}21, RLZ^{+}21, SZKZ20, TB20, TB21, TOK14, TZZ21, VBG^{+}13, YL12, ZJ11, LFL^{+}23]. Contact-Aware [TB22, MLPP09, TB21]. contact-based [TZZ21]. Contact-centric [RCCO22]. contact-invariant [MTP12, MWTK13]. contact-rich [LYvdP^{+}10]. contact-space [JTL^{+}12]. Contacting [FSKP23]. contacts [BBG21, Dav20, JL11a]. Content [AVB^{+}23, KSP13, LHKR10, LGJA09, THKM13, ZQCL19, AFR^{+}07, AS07, BLDA11, BDM^{+}21, CAA09, HDGN17, MRC05, WWOH08, XLZ^{+}10]. Content-adaptive [KSP13, LHKR10, THKM13, BLDA11]. Content-aware [ZQCL19, AS07]. content-based [MRC05]. Content-preserving [LGJA09, CAA09]. Contention [HC86]. ConTesse [LBHH23]. Context [FH10, HTG14, LGG^{+}07, SACO04, HZvK^{+}15, KP18, LMS13, LSD^{+}16, LPBM20, PKM^{+}11, WLP16, YCL^{+}20, MGT^{+}03]. Context-aware [LGG^{+}07, KP18, LSD^{+}16, WLP16]. Context-based [FH10, HTG14, SACO04]. contexts [MGS^{+}21]. contextual [CLW^{+}14, XMZ^{+}14]. Contingent [KAW20, ATM^{+}17, KKW20, MSM^{+}17]. continua [NØ13]. Continuation [YCBvdP08, SAJ21]. Continuity [BS88, DB88, FSRS22, GP09, SYSP14, Far89, HH10, HHP^{+}21, HB89, Pot91]. Continuous [AZMW21, KP03, LWH^{+}12, MM08, PP93, PMA^{+}14, RPWO18, SMP03, Sei93, SHD^{+}14, TMOT12, TSLP14, TWY^{+}20, WFS^{+}21, WCL^{+}23, YIC^{+}14, ZRLK07, ZYX^{+}21, ZLW^{+}16, ZXZL23, BGSF10, BEB12, DTP15, Kou16, LVGO21, Lev06, OLGM11, PRJ^{+}13, SMGH18, SXZ^{+}20, TMY^{+}11, TTWM14, TBC^{+}16, TLP07, TFG^{+}13, Wan14, WHK17, WLH^{+}13]. continuously [TDMS16, ZIT^{+}18]. Continuum [TCP06, YSB^{+}15, CLC^{+}20, DBD16, MSW^{+}09, WFL^{+}19, YSC^{+}18]. contoning [BVF^{+}17a]. Contour [DLTW90, Zyd88, PV06, VMT06]. contouring [BGOS06, CTFZ22, JLSW02]. Contours [EPO91, LBHH23, MSS92, MNB23, BHK14, DFRS03, SPO10]. contraction [ATC^{+}08]. contraptions [RCLM19]. Contrast [MC92, TD23, DRE^{+}12, HSHF10, MAC22, STTP14, THG99, TAKW^{+}19]. Contrastive [LDD^{+}23, ZTD^{+}23, CHY21, SAN23]. Contributing [BDD11]. Control [BB83, BSM88, BVF17b, CKT^{+}23, CJM21, DLG90, EHSN20, GFK^{+}23, Hil87, Lev23, LXY^{+}23, LHJ^{+}14, LVY16, LH17a, NRN^{+}23, PM17b, RYPZ23, SLST14, WGH21, XXA^{+}23, XSZK23, ZZLH23, AVF17, BP08, BdSP09, CH05, CWC11, CLL^{+}21, CSSL21, CKP^{+}21, CÖS19, CBvdP09, CBvdP10, DZS08, DKNY08, HYL12, HRL15, HGG^{+}11, HSvTP12, HKS17, HHC^{+}19, HZM^{+}08, IWZL09, ITM^{+}14, JL11b, JCW^{+}21, JWL^{+}13, KLL^{+}07, KCD09, LCR^{+}02, LT06, LKL10, LES10, LPKL14, LYP^{+}18, LLL18, LPLL19, LMLL21, LWH^{+}12, LC15, LYvdP^{+}10, LYvdPG12, LYWG13, LH17b, LHR^{+}21, MTP^{+}18, MZS09, MTPS04, MB21, MLPP09, MPP11, MRKN20, NZC^{+}18, OHB^{+}11, PM17a, PMA^{+}21, PFX^{+}22, PSE03, RSH^{+}05a, RTK^{+}15, RCOL09, RJN16, SSB^{+}15, SBR^{+}15, SJJ12, SGM^{+}16, SH08, SMD^{+}15, TER^{+}20, TMPS03, TLP07, TJ07, VSHJ12, WMZ^{+}13, WWH04, WPKL17, WPL18, WGH20, cWP10, XYJ13, YL10, YLvdP07, YHZ^{+}14, ZSKS18, ZZMC13, dSDP09]. Controllable [LDD^{+}23, SY05, SG01, WG10, WCZ^{+}22, XCLT14, YJLL22, ZSAF21, ZPW^{+}23, HAB20, JYQ^{+}22, LH05, LSCC20, MDLW15, Pot91, TWH^{+}22, TiABI07]. Controlled [CCW93, MZ13, PMLB22, AHD15, Ano03, ESCK16, FZZ^{+}20, FSH11a, HSD13, HGS23, HZCJ17, LHZ^{+}21]. Controlled-distortion [MZ13]. Controller [AFP^{+}95, Gla90, SCCB22, BG84, XDF^{+}19]. Controller-Based [AFP^{+}95]. Controllers [YSCL22, CHP07, LLP09, LKTK10, LLKP11, LZCV20, MTA^{+}20, MK16, WFH09, WFH10, WHDK12, WGH22, dLMH10]. Controlling [JL11a, KABL14, KH17a, RMGH15, KH17b]. controls [CTS^{+}21]. ControlVAE [YSCL22]. Convection [BBC^{+}23]. conventional [LFDF07]. Conventions [FSRS22]. Convergence [SJ17]. Conversational [SGD21, SDO^{+}04]. conversations [EMO10]. Conversion [RWW90, SV93, DIP^{+}18, KDW^{+}17, XLLW20]. Converting [LOMI11, EPD09]. Convex [Day90, DA21, MPB17a, TM84, BDD11, BLTD16, FLSG14, HZ82, MDK^{+}16, MPB17b, MCK13, TLJP18, WLLS22]. Convexity [VFK^{+}14, AA09, LW16]. conveying [DFRS03]. Convolution [FFL11, HLG^{+}22, HMM^{+}21, HRV^{+}18, LLDD09, NFA^{+}15, PSNB13, PO18]. Convolutional [GZC15, HKC^{+}18, MGA^{+}17, SFD^{+}22, SdGP^{+}15, TSLP14, AML18, BVM^{+}17, BB15, KHL19, LDPT17, SED16, SSISI16, WLG^{+}17, WSCR18]. Convolutions [NDS^{+}23]. Coons [KOY^{+}11]. cooperation [EAPL06]. Coordinate [Tur82, MLL^{+}21, PEVBC21]. Coordinated [LKM^{+}23]. Coordinates [CZ23, DSSS23, FHL^{+}09, BPC16, BLTD16, GSC21a, HF06, JMD^{+}07, JSW05, LJH13a, LSLCO05, PBH15, TMB18, YL08, ZDL^{+}14, LLCO08]. coordination [YLNP12]. Cope [EM90]. copresence [MWHL21]. copy [LvBK^{+}10]. core [CGG^{+}04, IG03, NNSM07, SCS^{+}08, SBZ09, WWS^{+}05, WHY^{+}13]. CoreCavity [NAI^{+}18]. cores [YLJ18]. Corner [Ros20]. Corner-operated [Ros20]. corners [LD06]. corotational [HLSO12, TREO16]. corrected [WKR99]. Correcting [HLBR12, HWBR14, KLF^{+}19, RMD12, WFDH18]. Correction [CFP^{+}21, KPB^{+}12, MHM^{+}17]. corrections [RCPO21]. Corrective [GZW^{+}16, SP09]. correctives [LYYB13]. correlated [BHHM20, GCH^{+}19, JAG18]. Correlation [GNHM15, WZK^{+}23, CHWH17, FKY08, ÖG12]. Correlation-Aware [WZK^{+}23]. Correlation-Based [GNHM15]. Correlations [ABGL21, SCO17b, SCO17a]. Correspondence [ASGS23, HPP^{+}22, Sah18, XLY^{+}22b, ALS^{+}18, AXZ^{+}15, BSFG09, HSGL11, LF09, SPKS16, ZYL^{+}17]. Correspondences [HLC^{+}19, HKC^{+}18, KLM^{+}12, LMS13, RPWO18, TMRL14, TBC^{+}16]. Corrigenda [Bak94, LR91, RO87, WC91]. Corrigendum [Ano90a, BK87, Pat87, RY93, VW95]. cosines [HDHN16]. Cost [DLW^{+}22, WWY^{+}13, CSHH21, LDS02, MCE^{+}17]. Cost-effective [WWY^{+}13]. COTS [Ros20]. Coulomb [BDCDA11, DBDB11]. couple [CZ17]. Coupled [BBN^{+}12, CMZP14, DAB15, XAW^{+}23, FQL^{+}20, XHS^{+}15]. Coupling [BBC^{+}23, GPB^{+}19, GSLF05, LXY^{+}23, LD23, TB22, XLYJ23, ANZS18, AIA^{+}12, BBB07, DFW20, HLW^{+}12, HFG^{+}18, IGLF06, LCD^{+}20a, LLDL21, NGL10, RMSG^{+}08, TB20, TB21, TLK16, YMR^{+}13, YSC^{+}18]. couture [UKIG11]. covariance [BSS^{+}13]. covariances [KEE13]. Covariant [LTDD16]. Covector [NWRC22]. Covering [HCW^{+}23]. covers [MGA^{+}17]. CPF [PEVBC21]. CPPM [LLZ^{+}20]. CPU [WWB^{+}14, WQS^{+}20]. CPUs [BSL^{+}16, FSP^{+}22]. crack [FFB^{+}09]. crack-free [FFB^{+}09]. cracking [PNdJO14]. crafting [ILB15]. create [BDM09, BL15]. created [HRE^{+}08]. Creating [KLY^{+}14, KLF^{+}19, LCK^{+}14, SHOW02, War92, XYH^{+}18, FNvD82, SDO^{+}04, XLS^{+}11, ZHG^{+}16]. Creation [BLC^{+}22, GHCG17, NI24, QLH^{+}22, ALX^{+}14, HDGN17, IBP15, JKT^{+}15, LZ04, LFB^{+}13, NKAS08, GRG04]. creativity [CK10]. creatures [GvdPvdS13, GPD^{+}18, MTN^{+}15, TGTL11, WPL18]. CRFs [ST16]. critical [Hub96, LMLL21]. crop [WLSL10]. crop-and-warp [WLSL10]. cropping [ZLH^{+}21]. Cross [KS04b, LYC^{+}22, ZVC^{+}20, ALS^{+}18, ACBCO17, BVG11, FBC18, HTWB11, HZCJ17, NCVMO05, PPTSH14, SBSS12, SZC^{+}22, SMGE11, ZHCJ15]. cross-domain [ALS^{+}18, SMGE11]. Cross-Editing [LYC^{+}22]. cross-frame [HTWB11]. Cross-parameterization [KS04b]. cross-section [SBSS12]. cross-sections [BVG11, HZCJ17, ZHCJ15]. crossbreed [PSN20]. crossing [AG05]. crossing-based [AG05]. CrossLink [HOM15]. CrossShade [SBSS12]. CrossY [AG05]. Crowd [CPV^{+}23, FYY^{+}16, KSHG18, DHOO05, GvdBL^{+}12, HXZW20, HOKP16, KSNG17, KSKL14, MLD^{+}08, NGCL09, OPOD10, WLP16]. Crowd-driven [FYY^{+}16]. crowds [JCP^{+}10, KSNG17, KSSI17, KOOP11, MLH^{+}09, TCP06, CPV^{+}23]. crowdshaping [SQRH^{+}16]. crowdsourced [OLAH14]. crowdsourcing [LFTC13, ZAE^{+}14]. CRT [MC92]. Crumpling [CLG^{+}16, SRH^{+}15, NPO13]. Crystals [Ste20, WW08]. cSculpt [CSTP16]. CSG [DIP^{+}18, Jan91, RV89, SV93]. CT [ZJMB11]. CT2Hair [SSW^{+}23]. cubature [AKJ08]. cubes [CZ21, LEQ^{+}07]. Cubic [BCX95, BHN98, Hob91, Kla91a, Kla91b, LJH13a, PP93, vW84, GI04, Joe89, LJG14, SD89]. cubic-order [GI04]. Cubics [Kla94]. cuboid [LZS^{+}21, SMZ^{+}14, YC21, ZCC^{+}12]. Cues [WF96, HCW15, NAB^{+}15]. culling [AHAM15, BJ10b, HAM07, HMAM09, LHLK10, TMY^{+}11, WLH^{+}13, ZRLK07, ZJ12]. Cumulative [Ano90b]. cumuliform [DKNY08]. Cup [WCFL22]. Cups [BCK^{+}23]. cured [ZBK18]. Curl [BHN07, CPAB22]. Curl-Flow [CPAB22]. Curl-noise [BHN07]. Cursor [Hud92, JX96]. Curvature [BS90, DKA23, Far89, IBB15, MWM23, BBR^{+}21, CPS13, GMB17, KNS^{+}09, Lev06, PCL^{+}12, Pot91, WPL06, YSW^{+}17]. curvature-based [WPL06]. Curve [LHJ^{+}14, Pat85, Pav83, Sai89, TZCO09, ULP^{+}15, VN85, BGAM12, Gal99, GSV^{+}17, Gos00, HSG^{+}19, IKCM13, KYC^{+}17, LRS18, LB05, PLS^{+}15, SXD^{+}12, TWY^{+}20, XCS^{+}14, YHZ^{+}14, ZCT16, ZM11, ZZCJ13, Pat87]. Curve-Drawing [VN85]. curve-driven [YHZ^{+}14]. Curved [BSR^{+}23, FAB^{+}18, KC23, KFC^{+}08, KMM17b, SYSP14, SJWG20, ERP^{+}19, KMM17c, KLPCP18, PSB^{+}08, RPC^{+}21, TCL21]. Curved-Knot [SYSP14]. CurveFusion [LCC^{+}18]. Curves [ACC90, AS21, BSSJ23, Che92, EK98, FG90, Hob90, Hob91, Joe90a, Kla91a, MD94, Mil87, Pet89, Rap91, Sei93, Tau94, YSC21, AB89, BWSS12, DSF22, DJBDT10, GMP09, HB21, HB89, JCW09a, JCW09b, KST08, NISA07, OBW^{+}08, PZ08, SS14, SBSS12, SSJ^{+}20, SD89, STZ14, WPL06, XSTN14, YSW^{+}17, ZS00]. CurveUps [GMB17]. curvilinear [XLY09]. CurviSlicer [ERP^{+}19]. custom [SBK^{+}18, WPMR09]. custom-ink [SBK^{+}18]. customizable [NQC^{+}21, SSM15]. Customization [RO94, JFH^{+}15]. Customizing [MGDA^{+}15]. Cut [BMBZ02, CMSA20, CPWAP08, KWL^{+}21, LSS05, PTH^{+}17, ZCLJ20]. Cut-and-paste [BMBZ02]. cut-cell [CMSA20]. cutaway [LRA^{+}07]. cutaways [BF08]. cutout [BWSS09, BJS^{+}08, FZL^{+}15, WBC^{+}05, ZQPM12]. cuts [BLA12, GF08, KT03, KSE^{+}03, LKK^{+}18, LVS^{+}13, RKB04, TDM^{+}14, WHY20]. cutter [LVS18]. Cutting [YCP16, FDBH22, KMB^{+}09, KBT17, LLKC21, SC18b]. cycles [ZZCJ13]. Cyclic [ACXG09, CZM^{+}23, HAK^{+}22]. cylinder [ZYH^{+}15]. Cylinders [BK85, AMZ99, BK87].

D [BIP01, Bou18, GIZ09, SLV^{+}13, AJS20, AKZ^{+}17, AWL^{+}19, AL13, ALX^{+}14, AXZ^{+}15, AZB09, AAR05, AVB^{+}23, AS21, AIH^{+}08, ARS14, BVF^{+}17a, BKLP16, BHR13, BLC^{+}22, BP07, BSS^{+}11, BSK^{+}16, BSW02, BBN^{+}12, BSS^{+}13, BGF^{+}23, BVG11, BGK^{+}13, BWSS12, BVS16, Bly06, BSM^{+}07, BB22, BR07, BAU15, BATU18, CCA^{+}12, CB04, CWLZ13, CKH18, CAD^{+}21, CMZP14, CK10, CKGK11, CGF09, CSPF12, Che13, CLD^{+}13, CLW^{+}14, CZL^{+}15b, CKIW15, CLF^{+}18, CPY^{+}22, CPW^{+}23, CSL^{+}22, CGP^{+}21, CRCM23, DNZ^{+}17b, DNZ^{+}17a, DS15, DLSCS08, DSAF^{+}13, DKD^{+}16, DIP^{+}18, DHL14, DDP02, DBB^{+}17, DSC^{+}20, ESCK16, EBGB14, EDF^{+}16, EST^{+}20, EPD09, ESZ^{+}17, EM96, FZBR16, FLJK21, FFBB21, FJL^{+}16, FH10, FRS^{+}12, FSL^{+}15, FBS^{+}23, FMK^{+}03, GDAB^{+}17a, GDAB^{+}17b, GZW^{+}16, GZC^{+}16, GIZ09, GM05, GF08, GGS03, GTDS10, GKHH12, GWN^{+}03, GWB05, GHL^{+}20, GFD^{+}12, GRT13]. D [GZC15, GXY^{+}17a, HGRT04, HGY17, HASK17, HK18a, HNH19, Hil87, HLP^{+}22, HLHR09, HLZ10, HZP^{+}22, HDK07, HMC11, HLV^{+}17a, HLV^{+}17b, HHL^{+}24, HTWB11, HCTW11, HTCH15, HMT^{+}15, HDGN17, HMM^{+}21, HZC^{+}22, Hud92, HOM15, IBP15, IGP^{+}17, ICG17, JTRS12, JBM^{+}17, JSKJ12, JLF^{+}09, JBX^{+}20, JCG^{+}21, JHR22, JHS^{+}23, JZH07, KMM^{+}02, KHS10, KH06, KSH^{+}14, KWS^{+}23, KDM^{+}16, KDR^{+}16, KKLD23, KC23, KSES14, KMYG12, KLM^{+}12, KRD^{+}12, KLM^{+}13, KLKL13, KNK^{+}22, KTL^{+}04, KDMW17, KFCO^{+}07, KL22, KSS^{+}15, KS04b, KYC^{+}17, LMS13, LHW^{+}10, LRAT08, LHKR10, LCXS09, LOMI11, LHG^{+}09, LRA^{+}07, LACS08, LT09, LSH^{+}10, LVG^{+}13, LPL^{+}17, LBB^{+}17b, LYF^{+}20, LLHF21, LJZ^{+}23, LHM^{+}18, LCOZ^{+}11, LYC18, LOW18, LFZ18, LGJA09, LWCT14, LHLF15, LHZ^{+}21, LGB^{+}21, LHH^{+}23, LKG^{+}03b, LFL09, LvBK^{+}10, LHVT17a, LHVT17b, LSZ^{+}14, LBRM12, MLZ^{+}16, MPF^{+}18, MHS^{+}19a, MLYZ19, Mai92, MWH^{+}13, MPI^{+}18, MC21, MSHS06, MPDW03, MPN^{+}02, MP04]. D [MRA^{+}22, MAN^{+}16, MTN^{+}15, MSS^{+}17, MPE^{+}23, MGP10, MGP06, MYWI15, MU22, MLS^{+}18, NG18, NLGK18, NAH^{+}18, NKS^{+}23, NISA07, NRDR05, NPLX22, NZIS13, NIR^{+}21, NPC^{+}22, OHX^{+}14, OHB^{+}11, OLGM11, ONOI04, PQF^{+}23, Par17, PGP^{+}19, PMW^{+}08, PK05, PXW18, PZ17, PRM14, PAR21, PDF^{+}22, PAAG21, PFB^{+}20, PSG^{+}06, PMPHB17, PWLSH13, RAWV08, RSL16, RPC^{+}21, RSI^{+}08, RID10, RDI10, RMD12, RB23, RHHL02, RMBB^{+}13, SS14, SHM^{+}18, SCH^{+}14, SLV^{+}13, SSGS11, SKSK09, SBR^{+}15, SSW^{+}23, SHL^{+}17, SHZ^{+}20, SAA^{+}21, SF07, SGXT20, SGX^{+}21, Shn92, SSS^{+}08, SARW^{+}15, SSS06, SDW^{+}16, SVB^{+}12, SQRH^{+}16, SRB^{+}19, SWS^{+}22, SSK^{+}17, MBU22, TDL^{+}18, TZY^{+}23, TD16, TDM11, TMB18, TS08, TTZ^{+}20, TFK^{+}03, TMB14, UZB^{+}23, UTB^{+}19, VVC^{+}15, VSHJ12, WBF^{+}17a, WBF^{+}17b, WAO^{+}09, WWY^{+}13, WGW^{+}13, WSXC16, WLG^{+}17, WSLT18, WSH^{+}18, WLX^{+}18, WKHA18, WMB^{+}20, Wan23, WLLS22]. D [WXLY17, WLHR11, WDB^{+}07, WSW^{+}12, WZQ^{+}18, WWL^{+}19, WZL^{+}20, WZ22, XSL^{+}22, XLF^{+}11, XIAP^{+}17, XZP^{+}23, XLJ^{+}09, XZT^{+}09, XZZ^{+}11, XZCOC12, XCF^{+}13, XCS^{+}14, XSZ^{+}16, XNZ^{+}22, YI17, YSL^{+}14, YMRD15, YSC^{+}16, YLJ18, YSW^{+}23, YWS^{+}11, YLC^{+}20, YZL^{+}22, YML^{+}23, YPL^{+}23, YKC^{+}16, YZX21, YZX^{+}18, YAB^{+}22, YSHWSH16, ZLP^{+}15, ZAC^{+}17, ZWW^{+}18, ZYX^{+}21, ZCM22, ZQL^{+}23, ZTNW23, ZWK14, ZPYX23, ZSW^{+}10, ZBW^{+}20, ZPW^{+}23, ZSMS14, ZK14, Zho18, ZZCJ13, ZNI^{+}14, ZPKG02]. D-Aware [SWS^{+}22, LWCT14, YWS^{+}11]. D-modeling [TS08]. D-printable [KSS^{+}15, LBRM12, MTN^{+}15]. D-Printed [JHS^{+}23]. D-printing [CCA^{+}12]. D-shape [WWL^{+}19]. D-to- [GIZ09]. DAG [MB21].

DAGs [KSA13]. damage [WFL^{+}19, WCL^{+}20]. damping [XB17]. dance [CTL^{+}21, VPHB^{+}21]. Dapper [CZL^{+}15b]. Dappled [VRA^{+}07]. Dark [JGC^{+}15, KF09, WZMM22]. Darkroom [HBD^{+}14]. DART [MGDB05]. Darts [MEA^{+}18, EPM^{+}14]. Data [CK10, CLSM15, CT17, Fol87, GLL^{+}16, HFL14, JHS12, JWL^{+}13, KNS^{+}09, KGG^{+}20, KPMP^{+}17, LJS^{+}15, LKL10, Lev90, LCODL08, LCX16, MTP^{+}15, NRS15, PH08, PYA^{+}24, RFW^{+}23, RO85, SRX^{+}23, SPDF13, SMGE11, SKAG15, Tsa15, WYW^{+}10, WOR11, XNZ^{+}22, AA09, ACP02, BCG05, BKR^{+}05, Che13, CLW^{+}14, CWZ^{+}21b, CLS03, CLZ^{+}22, DH06, FKY08, FCGH08, Hol18, HDK07, JLBM05, JLSW02, KHS03, KG04, LBJK09, LCR^{+}02, LCL06, LSK^{+}06, LGZ^{+}13, LGF04, MPO21, MUB15, MPBM03, MRC05, RPE^{+}05, SNF05, SSII18b, SKL07, SWR^{+}21, SJR18, TZK^{+}11, WAO^{+}09, WWS^{+}05, WLL^{+}14, WS21, WSL13, ZCW^{+}17, ZLE14, JTCW07, RO87]. Data-Driven [GLL^{+}16, NRS15, Tsa15, CK10, CLSM15, CT17, HFL14, JHS12, JWL^{+}13, KNS^{+}09, KGG^{+}20, KPMP^{+}17, LJS^{+}15, LKL10, LCODL08, LCX16, MTP^{+}15, PH08, RFW^{+}23, SPDF13, SMGE11, SKAG15, WYW^{+}10, WOR11, MUB15, MPBM03, RPE^{+}05, SSII18b, SJR18, WLL^{+}14, WSL13, ZCW^{+}17, JTCW07]. Data-Oriented [SRX^{+}23]. Database [GF82, MXZ^{+}23, HMLL15, SBHH16, XLS^{+}11]. databases [Ari06, MPF^{+}18]. dataflow [HZG09]. Dataset [LXZ^{+}23, NDD^{+}23, WGY^{+}18, WTD^{+}22, JHC^{+}21, WPL^{+}21, XZZ^{+}21]. datasets [BZL^{+}15, IZE^{+}21, KGB^{+}09, OAH11]. day [SPDF13, WM14]. DCT [MYC^{+}22]. DCT-net [MYC^{+}22]. de-animating [BAAR12]. dead [KHS03]. DeadWood [PGG^{+}24]. Deblurring [SLL^{+}21a, CL09, CWL12, JKZS10, LSC^{+}22, LWC^{+}13, RAT06, SJA08, WHB^{+}12, YSQS07]. Debugging [HZG09, DNB^{+}05]. Decaf [SGPT23]. decal [SGW06]. Decay [PGG^{+}24]. decimation [DTB06]. decision [DPF03]. Declarations [GF82]. Declarative [JDH^{+}22]. DecoBrush [LBW^{+}14]. decodable [KPM16]. Decomposable [Zyd88]. decompose [CZL^{+}15b, MAYZ^{+}20, Rit18, ZZX^{+}18]. decompose-and-carve [MAYZ^{+}20]. decompose-and-pack [CZL^{+}15b]. decompose-and-spiral-carve [ZZX^{+}18]. Decomposed [LGL^{+}19]. Decomposing [TDSG15, TLG17a, TLG17b]. Decomposition [BBPA15, CA24, CXW^{+}23b, DLW^{+}22, JHR22, LW15, MLS^{+}18, SBN15, TM84, VVCOS23, ZWL22, ZZL^{+}23, AMB^{+}21, AFO05, Bel18, BHY15, CRA11, CLJL20, DKT^{+}23, GLDZ15, GNS^{+}12, GKJ^{+}05, HXZW20, HLZCO14, KT03, Kou16, KHLN17, LD12, LZS^{+}21, LGZ^{+}13, LGB^{+}21, NAI^{+}18, PK05, SSD09b, TEG18, TLJP18, TLHD03, WLLS22, XXK^{+}06, ZZWC12, ZCB^{+}22, ZYH^{+}15]. decompositions [FFLS08, MSM^{+}17, MCK13]. deconvolution [KWB^{+}13, YSQS08]. Decorative [FPSG22, YKGA17a, LBW^{+}14, YKGA17b, ZHHZ20]. decorator [CXY^{+}15]. Decorrelating [SLK^{+}24]. decors [CML^{+}17]. Decoupled [RKLC^{+}11, CTM13]. Decoupling [RKAP^{+}12, SHD^{+}18, WYL^{+}14, LFJG17]. dedicated [RLR^{+}21]. Deducing [LYLL08]. Deep [ACOH^{+}18, BHHM20, BSK23, BLS^{+}21, CXW^{+}23a, CPV^{+}23, CK20, CPW21, CM14, Duf17a, Duf17b, EKM17, GLD^{+}19, GCPD16, GCB^{+}17, GZC15, HWH^{+}18, HCL^{+}18, HPP^{+}18, HKA^{+}18, HWZ^{+}20, KR17, KMM^{+}17a, KHL19, KP18, KGT^{+}18, KNC^{+}08, LLW17, LH17a, LSSS18, MHP^{+}19, MPH^{+}20, NZC^{+}18, NDD^{+}23, SMR^{+}22, SCO17a, SCO17b, SBK^{+}18, SYM^{+}24, WSCR18, WHG^{+}15, WSS18, XBS^{+}22, XSHR18, XBS^{+}19, YZW^{+}16, YHL^{+}18, ZYM^{+}20, ZCM22, ALL^{+}20, BODO18, CLL^{+}21, CYT^{+}18, DAD^{+}18, EKD^{+}17, GWY^{+}21, HLX^{+}21, HGY17, HLW^{+}18, HSK16, LT20, LGA^{+}18, LYY^{+}17, LOW18, LH17b, LH18, LZT^{+}19, LCL^{+}22, MTP^{+}18, MRA^{+}22, PBvdP16, PBYV17, PALvdP18, PHS^{+}18, SJ22b, SSR20, TKY^{+}17, WSLT18, WLY20, WNEH22, XYH^{+}21, YSW^{+}20, ZZI^{+}17, PAAG21]. DeepFaceEditing [CLL^{+}21]. DeepFaceVideoEditing [LCL^{+}22]. DeepFocus [XKF^{+}18]. DeepFormableTag [YMJ^{+}21]. DeepJoin [LBB22]. DeepLens [WSZ^{+}18]. DeepLoco [PBYV17]. DeepMag [CM21]. DeepMimic [PALvdP18]. DeepPhase [SMK22]. DeepSketch2Face [HGY17]. DeepToF [MHM^{+}17]. DeepVecFont [WL21]. DEF [MRA^{+}22]. Defending [Wan14]. Deferred [GCD^{+}20, TZN19, CTM13]. deficiencies [SMHW16]. Defined [Kaj83, vW84]. Defining [AK04, HLV^{+}17a, HLV^{+}17b]. Definition [vOV96]. Defocus [MMP^{+}05, VMCS15, BSS^{+}13, HQL^{+}10, ZN06, ZMN^{+}19]. defocused [MLR^{+}14]. Deformable [BdSP09, BC14, CSAP21, CMT^{+}12, MEM^{+}19, PM18, VJ19, YSC^{+}23, BJ05, BSG12, CFW13, DSP06, DLL^{+}18, FGBP11, GKJ^{+}05, GSLF05, HSvTP12, HNB^{+}06, HTYW22, IM10, ISF07, JF03, JP04, KUJH21, KJ09, KP11b, KS21, MCC09, MB12, NKJF09, PYW14, RMSG^{+}08, SvTSH14, STC^{+}13, SLS05, SGG^{+}06, TTZ^{+}20, WBS07, WMW15, WWW22, XWY^{+}09, YMJ^{+}21, YLX^{+}15, ZBYX19, vTSSH13]. deformables [KBT17]. Deformation [AXZ^{+}15, BSB16, CO19, GLL^{+}16, GPHSH19, JS11, JWJ^{+}14, LLF^{+}20, NI24, SGPT23, SP04, SJA^{+}20, WWY^{+}15, ZYL^{+}17, ACP02, BODO18, BVGP09, BZ11, BCWG09, BME21, BBO^{+}10, BSB17, BWKS11, BJD^{+}12, BN21, CW17, CSvRV18, DTPC23, FH07, FLJK21, FKY08, FYK10, GB08a, GYQ^{+}18, GPCP13, HSL^{+}06, JBPS11, JP02, Jam20, Jia21, JTSB16, LFS^{+}20, LCH^{+}21, MJC^{+}08, NFA^{+}15, NVW^{+}13, POB09, PH06, PH08, RS98, RTD^{+}10, RJ07, RCCO22, SMP03, SMW06, SYBF06, SZT^{+}07, SNW21, SSP07, VBG^{+}13, WJBK15, WG10, WY04, WBGB16, YK14, YCHK15, ZHS^{+}05, ZPBK17]. Deformation-driven [AXZ^{+}15, ZYL^{+}17, MJC^{+}08]. Deformations [BR94, NI22, AKJ08, CGC^{+}02, CPSS10, CPMK21, CPS15, HZ13, JZvdP^{+}08, KG05, LKF12, MZL^{+}17, MJBF02, MHTG05, TMDK15, VMW15, Wam16, ZJ12, vFTS06]. deformer [BBG21]. deformers [KS12, PMS12]. Deforming [WTGT09, KG06, SSJ^{+}20, SSW^{+}13, TMY^{+}11, XZY^{+}07, ZIT^{+}18, ZIT^{+}19]. DeformSyncNet [SJA^{+}20]. Degenerate [EM90, FNO89]. degenerations [GPSZ11]. Degree [Sei93, SJ94, CADS09, CLS85, PU06]. degree-raising [CLS85]. Degrees [IWC22]. Dehazing [Fat14, Fat08]. Delaunay [Ale20, BSTY15, DPVA23, FAB^{+}18, ILSS06, KLN91, LXFH15, LFXH17, TWAD09, WWX^{+}22, YLH18]. Delay [AMN03]. Delayed [RLLG^{+}20]. delta [LL19, LVGO21]. DeltaConv [WNEH22]. Demarcating [KST08]. demonstration [GAL^{+}09]. Demosaicking [MGA^{+}22, GCPD16]. Denoise [ANBH23]. Denoiser [SFD^{+}22]. Denoising [SFD^{+}22, VRM^{+}18, BVM^{+}17, CKS^{+}17, FDCO03, GCPD16, GLA^{+}19, HS13, Hol18, IMF^{+}21, LYT^{+}14, WLT16, YNL^{+}21, ZZXY21]. Dense [CZM^{+}23, CRCM23, HLC^{+}19, SB95, ZK13, BNB13, CKS18, DXZ^{+}19, HSGL11, KRF^{+}18, LD13, NGCL09, OCH^{+}16, SWW^{+}20, XIAP^{+}17, ZZZX21]. dense-weight [LD13]. Densely [YSHWSH16]. densification [HK18b].

Density [YR23, DLC^{+}15, DJBJ19, Fat11, GHV^{+}18, HJJ10, WHSG97]. Dental [ZEF^{+}22]. departures [WDW^{+}15]. Dependency [GF82]. Dependent [PNTK23, YSB^{+}15, KKW21, WWT^{+}03]. depict [CSD^{+}09]. Depicting [GSLM^{+}08, LMPB^{+}13, RBD06]. Depiction [PGG^{+}24, TDR^{+}12, VPB^{+}09a]. Depixelizing [KL11]. Deployable [BSR^{+}23, LPC^{+}23, PKLI^{+}19, PM21]. deployment [KLPCP18, RKP^{+}22]. Depolarized [NKKJ23]. Depth [CDSHD13, CSN^{+}12, HMI23, Jan91, LES09, LKE18, PBM^{+}22, VKB^{+}18, WSP^{+}23, ZZT^{+}21, AHAM15, BGK16, BCN08, BHR13, BBO91, CSHH21, CZL^{+}15a, CZN10, FKI^{+}14, FG11, GWM^{+}08, HLHR09, HK18b, JTL^{+}12, KHKR11, KKW20, LSR18, LFDF07, LHG^{+}09, LCD06, McC00, MDB^{+}19, PZM13, RTF^{+}04, STXJ15, SDP^{+}18, SSD^{+}09a, SHM^{+}14, TK14, WGJ^{+}18, WSZ^{+}18, WZC12, WM03, WZMM22, WZN^{+}14, XSZ^{+}16, ZSZ^{+}14, ZCT^{+}21, ZK14]. Depth-of-field [LES09, KHKR11, LSR18, WGJ^{+}18]. Depth-Order [Jan91]. Depth-presorted [CSN^{+}12]. depth-sensing [HLHR09]. Depths [Che92]. Deringing [WWH06]. derivation [WKR99]. Derivative [LTDD16, LC96]. Derivatives [AOCBC15, XLY^{+}22a, OKRC10]. derive [Spr82]. Descent [LLJ^{+}23, WY16, YLYW18]. descreening [KP18]. describing [RBvB^{+}04]. Description [dFP95]. descriptive [GSV^{+}17]. Descriptor [MOR^{+}18, GMW16, HZvK^{+}15, KSH^{+}16, SvKK^{+}11]. descriptor-space [SvKK^{+}11]. Descriptors [HKC^{+}18, CT17, TD16]. DESIA [WSP18]. Design [AKG^{+}23, BI92, BTSB23, BG89b, BWSS12, BBO^{+}10, BR94, BSBC12, BAC^{+}18, Cas91, FSDH07, GDAB^{+}17a, Gol84, Gol85a, HB23, JMB^{+}20, JTSW17, JMZ^{+}22, LCT23, LTDD16, LHVT17a, Mac86, MDH^{+}23, MWC^{+}23, MSK^{+}23, NPP22, PMLB22, PPV95, PTC^{+}15, RHW94, RFW^{+}23, RSSH24, SSL^{+}14, SW14, SG91, TBWP16, TMM^{+}21, US24, VHWP12, XZM^{+}18, XLCB15, YSC^{+}23, YKGA17a, ZAB21, ZBJ^{+}23, ZWZ^{+}22, ZGXF23, AMG^{+}18, AMG^{+}19, ASB22, AHB18, ACBCO17, BB15, BCC17, BLT^{+}15, CK14b, CZXZ14, CLSM15, CLMK17, CWSB22, CPWAP08, CTN^{+}13, DLC^{+}15, DSZ^{+}16, DYYT15, FYY^{+}16, GDAB^{+}17b, GSFD^{+}14, GJG16, GGP^{+}20, GPD^{+}18, GSV^{+}17, HB21, IIM12, JWI^{+}21, KP09, KP10, KGL^{+}22, KCD^{+}16, KSS^{+}15, KSSI17, KAMJ05, LSD^{+}16, LWS^{+}18, LKB22a, LXW^{+}11, LZF^{+}19, LKvK^{+}14, LHVT17b, LCBD^{+}18, MZL^{+}17, MDZ^{+}21, MGDB05, MPBC16, MPI^{+}18, MDLW15, MSS^{+}19, MZD05, MTN^{+}15, MZB^{+}17, MSL^{+}11]. design [MMT18, MLB16, MIWI16, MI07, PZ07, PRK^{+}17, PCS23a, PIC^{+}21, PTG02, PKPP21, PYB^{+}16, POT17, PTV^{+}17, RVLL08, RRS13, SXZ^{+}17, SWC^{+}18, STTP14, STC^{+}13, SCGT15, SWT^{+}17, SZ15, SWF^{+}21, TGY^{+}09, TCG^{+}14, UBW99, UMK17, UIM12, UKSI14, UPSW16, UB18, VABW09, VGDA^{+}12, VBFG12, WJBK15, WCPM18, WLM^{+}15, WPL^{+}21, WDR11, WDR13, WZL^{+}20, XSZB15, XB17, XKCB18, XDF^{+}19, YWVW13, YXFH21, YKGA17b, YCC17, ZKBT17, ZMT06, ZFS^{+}19, ZHPY21, ZXKL^{+}20, ZQCL19]. Design-driven [BWSS12]. Designing [APH^{+}03, CLM^{+}13, HPC21, MHCT23, PBSH13, PPW18, RCLM19, STK^{+}14, TAN^{+}21, ZCT16, Coh87, JRT^{+}15, NISA07, ONOI04, TZCT20, TTZ^{+}20, WSP18]. Designs [WZHL23, ZZL^{+}21, CKX^{+}08, DFL^{+}15, LYH^{+}15, MGS^{+}21, PKM^{+}11, PCLC16, ZCL18]. desired [BBO^{+}10, MZL^{+}17, ZKBT17]. desktop [LRFN04]. destination [KAB^{+}10]. Destruction [SLM^{+}23]. Detail [FH07, HYS23, HK10a, MSW^{+}09, SK16, WZK^{+}23, ALY^{+}21, CH04, CHPR07, ECBK14, FFLS08, FAR07, FKY^{+}10, HFTF15, KGS^{+}18, LKG^{+}03a, MLR^{+}22, NSACO05, PSNB13, PKZ04, RBD06, WWA^{+}16, YKJM12, ZNT18, ZZW^{+}22b]. Detail-Preserving [SK16, HK10a, ALY^{+}21, NSACO05, WWA^{+}16, ZNT18]. Detailed [ANL^{+}23, BBK^{+}15, EB14, KNK^{+}22, LJL23, AFO05, CPMK21, DKH^{+}10, FFBB21, GVWT13, GMP^{+}06, KMB^{+}09, YL12]. details [Bae18, BWDL21, BC18, Bou18, Cor18, Did18, Gup18, Hac18, Iza18, JCW09b, Kal18, Kau18, Kim18, Lau18, Lee18, Li18, Lip18, Liu18, Mit18, Pan18, Rit18, Ter18, Wan18b, Xu18, Zha18, Zho18, Zhu18a, Zhu18b]. Detecting [YLL^{+}22, HHV^{+}21]. Detection [HLP^{+}22, LYF^{+}20, RV89, WFS^{+}21, BEB12, CMZP14, DAB15, GKJ^{+}05, Hub96, JP04, MSHS06, Mir98, MGP06, RTF^{+}04, SPO10, TTWM14, Wan14, WLH^{+}13, XZJ^{+}12, YNS19, ZRLK07]. Detections [NY94]. Determination [EM96, JTMW20, SNF05]. deterministic [GGY18]. deterministic-stochastic [GGY18]. Dev2PQ [VVHSH22]. Developability [SGC18, ZFO^{+}22]. Developability-driven [ZFO^{+}22]. Developable [IRWP23, JHR^{+}15, RHSH18a, SRH^{+}15, TBWP16, VVHSH22, ZLZ^{+}23, EB08, IRHSH20, KCD^{+}16, LPW^{+}06]. Development [WW82, HFH^{+}17]. developmental [PNH^{+}14]. Deviation [HWB23, WDW^{+}15]. Device [GFMS95, GMP^{+}16, JPG^{+}14]. Device-Directed [GFMS95]. Devices [LMR83, NKK^{+}14, RvE93, DWS^{+}20, HHGH13, XBZN19]. Dexterous [LYP^{+}18, HMT^{+}15]. Dextrous [Liu09]. DFlow [DTP15]. DHFSlicer [YAV^{+}20]. Diagram [WWWG22]. Diagrams [LFXH17, XWX^{+}22, AGL^{+}17, GS85, IOOI05, LACS08, MHSL18, SGG^{+}06, XLC^{+}16, dGWH^{+}15]. DiagSplit [FFB^{+}09].

Dialogue [Gre86, LDTA17].

dialogue-driven [LDTA17]. Dichoptic [WC21a]. dictionaries [GZB^{+}13, MWBR13]. Dictionary [BLDL21, CXW^{+}23b, XZZ^{+}14]. Dictionary-based [BLDL21]. diff [BDG15]. DiffAqua [MDZ^{+}21]. DiffCloth [LDW^{+}23]. diffeomorphisms [WGS23]. difference [CHWH17, GWAB19, HRV97, MDC^{+}21]. Differences [VMKK00, WGY^{+}18, BDG15, Nah20, ROA^{+}13]. Differencing [Kla91b, Kla94, Rap91]. different [SPDF13]. Differentiable [BWL^{+}23, DWM^{+}22, GFK^{+}23, LWF^{+}23, LADL18, LGA^{+}18, LLGRK20, LDW^{+}23, LXY^{+}23, LWL^{+}20, MDH^{+}23, PBM^{+}22, RAM^{+}21, SLWL23, SQSL22, TMM^{+}21, VSJ22, WVY^{+}22, WA23, WCRZ21, WLL23, XLY^{+}22a, YKC^{+}21, YZN^{+}22, ZRJ23, ZWRY21, BLD20, DWS^{+}20, GHZ^{+}20, JSRV22, LHK^{+}20, MDZ^{+}21, RGF^{+}20, RFS22, SYS^{+}21, SLH^{+}20, TYY^{+}19, YLB^{+}22, ZDDZ21, ZYZ21]. Differential [HMI23, MNB23, WW11, XBLZ23, LXY^{+}16, NDMKJ22, SSR20, WC21b, YLH18, ZSGJ21, ZD20]. differential-informed [ZD20]. differentiate [SWF^{+}21]. differentiating [BMM^{+}21, VSJ21]. Differentiation [NDS^{+}23, YZN^{+}22, Gue07]. Differently [XSL^{+}22]. DifferSketching [XSL^{+}22]. DiffFR [LXY^{+}23]. difficult [JM12]. Diffing [RSP23, DP13]. DiffPD [DWM^{+}22]. diffracted [JBY^{+}19]. Diffraction [TG17b, CHB^{+}12, HP17, SMM14, SMC21, TG17a, YJB^{+}14]. diffractive [Fre16, SBB^{+}22, WVJH17]. diffuse [DDTP15, MWRD13, SMM14, SJR18]. diffuse-specular [SJR18]. Diffusion [ANBH23, AZL23, AFL23, BSSJ23, BAU15, CAV^{+}23, CZN10, FFL10, HHL^{+}24, JLF^{+}23, Knu87, LDD^{+}23, LTH^{+}23, MSL^{+}24, OBW^{+}08, SXD^{+}12, TMRL14, UZB^{+}23, VKM^{+}23, ZDT^{+}23, ZTNW23, ZPW^{+}23, AW20, BX03, CAO09, DI11, DFW20, DJ05, IKCM13, JCW09a, JCW09b, McC99, STZ14, TSNI10, WZT^{+}08a, XSTN14, XSH^{+}20, ZF03]. DiffusionNet [SACO22]. Digital [ASHW23, BLC^{+}22, GRS93, KWS^{+}23, Knu87, KL12, MBS^{+}11, PSA^{+}04, PGG^{+}24, RFW^{+}23, RSSH24, SFB92, SCB88, WDB^{+}07, ZSSJL20, ADA^{+}04, CXY^{+}15, CZX^{+}16, GB08b, ITM^{+}14, PLKD18, RSSF02, RMD12, RC22, Sha03, SJ21]. Digitization [GARP^{+}23, HSW^{+}17]. Dihedral [NI24, PRP^{+}15, LS07]. dilution [NGD^{+}06]. Dimension [PBCF93, GZ05]. Dimension-Independent [PBCF93]. Dimensional [CKH18, Day90, EM94, Gla90, KM97, MEA^{+}18, OF01, AGDL09, BBO91, BJ17, Boi84, CH05, COSL98, EPM^{+}14, GO12, IGLF06, JSMH12, LWH^{+}12, ML22, MSRB07, MdLH10, PSH^{+}21, RSSH24, SHP04, Ten20, UB18, WWS^{+}05, ZWL^{+}18]. dimensionally [GMP09]. Dimensions [WF96]. DINUS [MFR^{+}10]. Dip [AKZ^{+}17]. Dipole [FHK14, FD17, MHZ^{+}21b]. Direct [HPB06, Jac86, KTB07, LL19, LVGO21, Lev21, SB95, SF09, SWZ96, ZHX^{+}07, BSK^{+}16, MIB15, NKGR06, PFX^{+}22, SILN11, TCL21, VKK18, Bly06]. Direct-Manipulation [Jac86]. Direct-to-indirect [HPB06]. Directable [HG09, BMWG07, PAR21]. Directed [GFMS95, BKLP16, HWB23, MTP^{+}18, PRMG16, RYPZ23, SRL^{+}15]. Directing [GLC^{+}18, PCLC16, CLC14]. Direction [RS14b, TWLT19, BMSG09, GI04, KCPS13, LXW^{+}11, PS04, RVLL08, RVAL09, ZLP^{+}15]. Directional [BV22, CRG^{+}20, CV20, FHK14, Pag98, EDR11, KWN^{+}17, MUB15, PO18, RS18, SV19, WLHR12]. directly [KMM^{+}02]. Dirty [DSJA^{+}21, GRBN09]. DISCO [GLL^{+}04, JGN16]. discontinuities [BMM^{+}21]. Discontinuity [BGF^{+}23, ZQPM12, HFG^{+}18]. Discontinuity-Aware [BGF^{+}23, ZQPM12]. Discontinuous [HK05, CBW^{+}18, EB14, YBAF22]. Discovering [JGMR23, NRH17, PMW^{+}08, RGF^{+}20, YYVY21, BLPW14, LWC^{+}11, WL16]. Discovery [HGM14, MTP12, JCG^{+}21]. discrepancy [APC^{+}16, DEM96]. Discrete [AFH20, AW11, AHL17a, AOCBC15, BUAG12, BWR^{+}08, BAV^{+}10, CQS^{+}23, ESBC19, FW12, FGC23, GSC21b, JHY^{+}14, KSS06, LTDD16, MWT11, Mal89, MOR^{+}18, MGP10, RHSH18a, TCT23, Tau94, TLHD03, WYW23, AHL17b, ABA02, CCS^{+}21, CPS11, DBWG15, HPC21, LCCS18, LZZ^{+}21, LZH^{+}17, QHY^{+}16, RHSH18b, SGW06, SS10b, SRGB14, SGG^{+}06, VBCG10, XW09, YWH13, YXH14, YSC^{+}18]. Discretization [SACO22, XSH^{+}20]. discriminative [ARS14]. Discriminators [RMP^{+}23]. Disentangled [GWLG23, JCFG23, SWS^{+}22, XHWW22, YML^{+}23, CLL^{+}21]. disentanglement [NBLCO20]. Disentangling [HAK^{+}22, KPACO22, TBTA^{+}24]. disk [BWWM10, EDP^{+}11, EBJ^{+}06, GM09, Wei08, YW13, DH06]. Disney [BAC^{+}18]. Disparities [AKG^{+}23]. disparity [DRE^{+}11, DRE^{+}12, FKN17, KDM^{+}16, LHW^{+}10]. dispersed [KySK10]. Dispersion [CMT^{+}16, CT05]. Dispersive [JW23]. Displaced [BR21b, CHZ14]. Displacement [BvdPPH11, Roc89, DHI^{+}13, HFG^{+}18, MJC^{+}08, NFA^{+}15, NL13, WWT^{+}03, WZYR19, YHCOZ18]. displacement-mapped [WZYR19]. Display [DVC09, DCT^{+}22, Jan91, JGN16, LMR83, MDK08, PRM14, RO85, RO87, SBSH18, WK95, Zyd88, AWGB04, ALK^{+}17, BNK10, BSW02, BGB^{+}05, DER^{+}10, Did18, DD02b, EDF^{+}16, FH04b, FRS19, GZL14, GWN^{+}03, HWBR14, JBM^{+}17, JBLL18, JMY^{+}07, KYS^{+}15, KJS^{+}19, Kou16, KKB^{+}11, LWH^{+}11, LCTS05, LTO^{+}15, MWH^{+}13, MP04, NBB04, PMOR10, SMG^{+}05, SHS^{+}04, SST^{+}83, TFK^{+}03, THG99, YJB^{+}14, ZN06, ZJY^{+}21]. Display-Camera [JGN16]. displaying [SDIN18]. Displays [CBYJ23, Dun83, MMHP23, PFP^{+}22, QCOS23, VN85, AFR^{+}07, BF12, CB04, CTS^{+}20, CKS18, CGP^{+}21, DSAF^{+}13, DDD^{+}14, FRSL08, GWB05, HWRH13, HLR^{+}14, HLBR12, HWBR14, HCW15, HPK^{+}17, KNL^{+}22, KPM^{+}17, KBBD17, LHKR10, LL13, LJM^{+}16, MLR^{+}14, MGK17, MS05, MFL17, MWHL21, MSM^{+}17, NAB^{+}15, POAR12, SLV^{+}13, SHK^{+}17, TDMS16, WLHR11, WLHR12, XKF^{+}18]. dissections [DYYT17]. dissipation [FGW^{+}21]. dissipative [BOFN18]. Distance [HCW^{+}23, LBB22, MPB17a, PP94, XWX^{+}22, ZXS^{+}23, BR21b, COSL98, CWW13b, HCOB10, KvKSHCO15, LRF10, LCDF10, ML22, MWH^{+}09, MPB17b, TLK09, TTT^{+}17, VSJ22, WPL06, WDB^{+}08, Xia97, ZDI^{+}15]. distances [AWGB04, SRGB14, SdGP^{+}15]. Distinctive [LYF^{+}20, SF07, LRFN04]. distinctiveness [HRZ^{+}13]. Distortion [LYP^{+}14, SLL19, SDK19, SJWG20, AL13, APL14, CWKBC13, CW15, CCW16, CLW16, FOL^{+}21, KLS03, KABL15, LW16, Lip12, MZ13, PTH^{+}17, SdS02, TBTS08, ZBK18]. Distortion-free [SLL19]. Distortions [WTD^{+}22, VRC^{+}13, WFDH18]. Distributed [KSH10, LN84, QRL^{+}23]. Distributing [MSQ^{+}18]. Distribution [YMRD15, HDCD15, HHA^{+}10, LD05, LAC^{+}11, MYRD14]. Distributional [PP94]. Distributions [SMR^{+}22, BSD09, DHB17, OFCD02, ÖG12, XH18, YHMR16]. Disturbance [CHTK24, PGG^{+}24]. dithering [MBU22]. diverse [HSC^{+}22, WLO^{+}14, WGH20, XZCOC12, YYL22, YYVY21]. divide [Mor11]. divide-and-conquer [Mor11]. division [ABJN85]. DLayout [PAAG21]. Dlite [HDGN17]. Do [AFR^{+}07, XSL^{+}22, CGL^{+}08, CSD^{+}09, EHA12, JMB^{+}14, WKHA18].

DOC [GFK^{+}23]. dockers [BWKS11].

Document [LQGY24, JLS^{+}03].

Documents [XZZ18, FNvD82]. DOF

[HMT^{+}15, SHX^{+}22]. Domain

[AVF17, BVF17b, DMZ^{+}17, GAA^{+}23, GO11, LLN^{+}14, SHD^{+}14, ALS^{+}18, Aga07, AWL13, ALLD17, BPE17, BZCC10, BDT^{+}08, FLW02, FN20, GPM^{+}22, GNS^{+}12, GHV^{+}18, HSRG07, HHL^{+}24, HSL^{+}06, KH08, KSH10, KMA^{+}15, KHL19, KLS^{+}13, LKL^{+}13, Lév03, MRK^{+}14, MKD^{+}16, MP08, MYC^{+}22, PKCH18, SMGE11, WJ19, WW11, XZY^{+}07, YWVW13, ZLC^{+}13]. domain-calibrated [MYC^{+}22]. domains [FDBH22, HZCJ17, MC21, NRC21, SdGP^{+}15, TPP^{+}11, WMW15]. dome [HW12]. Dominant [SRUL16, GJTP17, RLZ^{+}21, SPGT18, SRUL17]. Dominated [LL23]. dominates [EMO10]. doodles [TBvdP04]. Doppler [HHHW15, KJGP23, WKR99]. Dot [Knu87]. Dots [LKvK^{+}14]. Double [DBWG15, HCW^{+}23, RY92, SR09, YAV^{+}20, MFR^{+}10]. double- [SR09]. Double-Step [RY92]. Downsampling [ZWDR16]. downscaling [GO17, KSP13, ÖG15, WWA^{+}16]. DR [WLL23]. DR-Occluder [WLL23]. DR.JIT [JSRV22]. Drag [JSTS06]. Drag-and-drop [JSTS06]. dragon [WPKL17]. Drape [FHXW22, GRH^{+}12]. Draping [LKL23]. draw [CGL^{+}08]. Drawing [AS21, Bli82, DH96, Kla91a, SLF22, VN85, AG05, FLB16, FTP03, Gal99, GTDS10, JDA07, KMM^{+}02, KNS^{+}09, KLKL13, LZC11, LFTC13, LBW^{+}14, PLKD18, PNCB21, SKSK09, Spr82]. Drawings [BCV^{+}15, BS19, OCNG21, OGN^{+}23, SZL^{+}23, BVS16, BKR^{+}05, CSD^{+}09, FZLM11, LMLH07, LPBM22, LRS18, NSX^{+}11, NHS^{+}13, RRS19, VA88, WQF^{+}21]. Drawn [YBMN^{+}23, JSMH12, SBHH16, SKC^{+}14, XWSY15]. DreamFace [ZQL^{+}23]. dress [CYT^{+}18]. Dressing [XBS^{+}22, CTTL15, CYT^{+}18, GRH^{+}12].

DressUp [YYTC12]. dribbling [HHC^{+}19, LH18]. Driven [ANBH23, CWL22, GLL^{+}16, JSSH15, JWD^{+}23, MPE^{+}23, NRS15, Tsa15, VKM^{+}23, YSW^{+}23, ZZZ^{+}22, ZXS^{+}22, Aca07, AXZ^{+}15, AJM12, BSK^{+}16, BAC^{+}23, BDM09, BWSS12, CTFP05, CGC^{+}02, CK10, CLSM15, CTL^{+}21, CZXL23, CT17, DPF03, FL04, FKY08, FYY^{+}16, GHBCO21, HZP^{+}22, HDS^{+}18, HZW^{+}13, HYG^{+}13, HFL14, JWW^{+}20, JYQ^{+}22, JHS12, JWL^{+}13, KNS^{+}09, KGG^{+}20, KAL^{+}17, KYS^{+}15, KP11b, KPMP^{+}17, LJS^{+}15, LS02, LDTA17, LKL10, LTK09, LCODL08, LYGC15, LT00, LYWG13, LXC^{+}15, LABS23, LCX16, MJC^{+}08, MLZ^{+}16, MPF^{+}18, MTP^{+}15, MUB15, MPBM03, MCW^{+}21, NHS^{+}13, PH08, PSF09, PL07, PNA^{+}21, PNCB21, RPE^{+}05, RFW^{+}23, ST14, SPDF13, SMGE11, SSII18b, SJR18, SKAG15, VK16, WYW^{+}10, WOR11, WLL^{+}14, WSL13, WSL^{+}14, XZZ^{+}11, XSZ^{+}16, YKZ^{+}22, YHZ^{+}14, ZCW^{+}17, ZFO^{+}22, ZXL^{+}18, ZYL^{+}17, JTCW07]. Driving [BWS^{+}21, FJA^{+}14]. Driving-signal [BWS^{+}21]. Drone [LLH^{+}22, NMD^{+}17]. Drones [ASN^{+}20, GLC^{+}18]. drop [JSTS06]. drops [BNK10, WMT05]. Drucker [KGP^{+}16]. Dry [LDW^{+}23, LJBBD20]. DS [DML17]. DSCarver [ZZX^{+}18]. DSG [YML^{+}23]. DSG-Net [YML^{+}23]. DShape2VecSet [ZTNW23]. DSL [BSL^{+}16]. DTV [KDW^{+}17, SLV^{+}13]. Dual [CBK12, CK14b, GARP^{+}23, JLSW02, Lév03, LFXH17, LPC22, SCG^{+}05, WLT22, ZYWK08, CTFZ22, HPK^{+}17, KCŽO08, LSC^{+}22, LAKL11, LHKR10, ORK12, WSM11, WL21]. dual-frame [HPK^{+}17]. dual-layer [LHKR10]. dual-modality [WL21]. Dual-scale [GARP^{+}23, WSM11]. dual-space [LAKL11]. ductile [OBH02]. due [GRBN09]. during [AKG^{+}23, DYT05, HRvdP04, MBF04]. dust [OHR14]. Dyadic [ASHW23, KBZ15, AW21]. Dyna [PMRMB15]. Dynamic [ASP07, AMMS08, BLDL21, BMBRD24, BAM14, BSM^{+}07, CWW^{+}13a, CLX^{+}22, CM10, DGH16, DJ18a, HL14, IBP15, JKH^{+}22, Kal14, KC21, KH17a, LCTS05, LLL22, LKZ^{+}20, MWLT13, MLL^{+}22, MPE^{+}23, PBvdP15, PAR21, SLR^{+}16, TQ94, VPB^{+}09b, WSL^{+}19, WSP^{+}23, WRK^{+}10, Wu92, WS17a, XWW^{+}14, YPG01, ZWCM21, ZCM22, ZIH^{+}11, ZMCF05, ADM^{+}08, BBB^{+}14, BI08, CHZ14, CWW^{+}16, CCWL18, CGC^{+}02, CH07, CZ11, DJBDT10, DJBDDT13, DHW^{+}11, DD02b, FLW02, GVWT13, GRB^{+}18, HLX^{+}21, HSG^{+}16, HKAK16, JP02, JF03, JSB^{+}10, KSB^{+}13, KR17, KNS^{+}09, KUWS03, KYYL08, KFCO06, KLF^{+}19, KH17b, LWH^{+}11, LEPM22, LSA05, LLV^{+}12, LTT^{+}20, LP02, LYvdPG12, LNWB03, LSS^{+}19, MRK^{+}13, MKMS04, MEMS06, MP04, MWI18, MLPP09, MK16, MCK13, MCHAM06, Mus13, NSX^{+}18, NHAH03, PBH15, PBYV17, PMRMB15, RSM^{+}10a, RWS^{+}06, RAI06, SMC21, SHS^{+}04, SKY^{+}12, SHX^{+}22, SZT^{+}08, SCT^{+}15, SKS02]. dynamic [SKK^{+}12, SKB^{+}14, SJLP11, SM06, SZC^{+}07, SZS^{+}08, SPW^{+}18, TAHL07, TPWG02, Van06, VBK05, WBS07, WRG^{+}09, WLHR11, WFL^{+}19, WS17b, YPB16, YL08, YPL21, ZWZ^{+}16, ZJY^{+}21, ZHL^{+}05]. Dynamical [LCCS18]. dynamically [KJS^{+}19, RH16, SSJ^{+}20]. dynamically-foveated [KJS^{+}19]. Dynamics [BZC^{+}23, CLMK17, DWM^{+}22, MEM^{+}19, MHNT15, BKLP16, BWRB05, BAC^{+}06, BML^{+}14, BEH18, CLL^{+}22, DBDB11, DYN03, DKNY08, Erl07, FLS^{+}21, FTP16, GHZ^{+}20, GvdBL^{+}12, HMP^{+}20, ISN^{+}20, KEP05, KUJH21, KPH18, KLV20, LKL^{+}22, LMY^{+}22, LT08, LLR^{+}15, LFS^{+}20, LLKC21, LCX16, LJBBD20, NGCL09, RGL05, RCPO21, Ten20, TNGF15, TJ07, Wan15, WWB^{+}19, WP12, WST09, XB16, YPL21, ZSZ^{+}14, ZBLJ20, ZPBK17]. Dynamics-aware [CLMK17]. dynamism [LJH13b]. DyRT [JP02].

EARS [RGH^{+}22]. Earth [SRGB14]. easier [FGW^{+}21]. Easily [LZCX19]. Easy [Pet95, RKAP^{+}12, SFG^{+}13]. EasyFont [LZCX19]. Ebb [BSL^{+}16]. eccentricity [KKW21, MAC22]. eccentricity-dependent [KKW21]. Ecoclimates [PMG^{+}22]. ecosystem [CGG^{+}17]. ecosystems [KGG^{+}20, MHS^{+}19b]. Edge [FFLS08, Fat09a, FCA09, HLP^{+}22, HWG^{+}13, KRK11, SGM12, SSD09b, WWT^{+}06, BHY15, CPD07, FFL10, Fat07, GO11, HHF^{+}19, KTY09, LADL18, LSVT15, PHK11, RTF^{+}04, WSM11, WCSC22]. Edge-avoiding [Fat09a]. Edge-aware [HWG^{+}13, KRK11, CPD07, FFL10, GO11, PHK11]. Edge-based [FCA09, KTY09]. edge-cone [LSVT15]. Edge-guided [SGM12]. Edge-preserving [FFLS08, SSD09b, BHY15]. edgebreaker [AFSR03]. edges [BWG03, LD06, Nai98, SNCH08, SC20, WXLY17]. Edit [GJWW14, AP08, CZZT12, GSMCO09, JMB^{+}14, KvKSHCO15, XLJ^{+}09, LFL^{+}23]. Editable [ZLY^{+}21, CZS^{+}13, EPD09, LD21]. Editing [BL18, BAC^{+}23, BBPA15, CKT^{+}23, GLC^{+}23, HRS^{+}23, JSSH15, JDH^{+}22, JZH07, KG06, LYC^{+}22, LZKW10, MLL^{+}22, MPE^{+}23, PABE^{+}21, RMBCO23, SDN18, SSSH17, SWS^{+}22, VKM^{+}23, YFFA21, ZMW^{+}23, AYL^{+}12, APS^{+}14, AFTCO07, BCT15, BPK^{+}13, BSG12, BSFG09, BC02, BSK^{+}16, BAOR06, BAERD08, BSHK04, BMBZ02, BWSK12, BST^{+}14, Bou18, BD02b, CLL^{+}21, CZM^{+}10, CBL^{+}16, CSRP10, DTP15, DCP14a, DDTP15, DKT^{+}23, FH04a, FH07, FFL10, FTD21, FTZ^{+}19, GZ08, GCSS06, HR13, HPG^{+}22, HSK16, HXM^{+}13, HZW^{+}13, IDN12, JCW09a, JGGN15, KOWD21, KBD07, KRFB06, KN02, KHKL09, KLLT08, LRT^{+}14, LBAD^{+}06, LDTA17, LHdG^{+}14, LLGRK20, LWW08, LTJ18, LSH^{+}22, LCL^{+}22, LKG^{+}03b, LSS^{+}17, MBWB02, NSACO05, PHT^{+}13, PL07, Pel10, PZKW11, PGB03, PHS^{+}18, RTD^{+}21, RAKRF08, ROTS09, SSTP15, SFLM04, SSRB^{+}17, STPP09, SSJ^{+}11, TPSHSH13, UKIG11, WYXJ21, XZY^{+}07, XMR^{+}11, XYJ13, YZX^{+}04, YCHK15, ZWZ^{+}16, ZPKG02]. Editor [GW90, Tan83, Bea91, Ber82a, Ber82b, Fol86a, Fol86b, Fol86c, FGN84, Fuc82, Pha18]. Editor-in-Chief [Bea91]. Editorial [Bea91, Fol91, Fol92, Fol95a, Gla95, Gla97, Har03a, Har03b, Har04, Har05, Hod00, Hod02b, Hod03]. Editors [BG89b, BG89a, BG90, FR87]. edits [HLR^{+}17, IAF09]. Effect [JHS^{+}23, Kla87, WTS^{+}23, DK99, HOKP16, MBB12, SCW^{+}21, ZAJ^{+}15]. Effect-based [WTS^{+}23]. effective [APH^{+}03, BSW02, WWY^{+}13]. Effects [BYRN17a, KFB10, TG17b, YMRD15, ZSSJL20, ZCS^{+}22, BYRN17b, CLC96, CFW13, GGN18, HAK^{+}22, KKN^{+}13, LES10, LAC^{+}11, MYRD14, PH15a, RAWV08, SSBG10, SKC^{+}14, TG17a, WKR99]. Efficiency [GYGS22, EKA84, LFY^{+}19, RGH^{+}22, Wan18a]. Efficiency-aware [GYGS22, RGH^{+}22]. Efficient [AJS20, Aga07, AONA22, Bel18, BFK^{+}16, BEB12, CCS^{+}21, CCL^{+}22, DII23, Dun83, EDP^{+}11, FP03, FBCA23, GLL^{+}16, GHF^{+}07, GH98, Gue07, GZS^{+}22, HH16, IGLF06, IH20, JTMW20, KJM10, KCW^{+}18, KLN91, KM97, KFS13, LLF^{+}20, LRR04, Lev90, LHZ16, LMLD22, LFZ^{+}23, LXFH15, MZS^{+}11, MWM08, MK16, MRC05, MPG^{+}16, NMLH14, PZM13, PM17a, PM17b, QZG^{+}19, RSV^{+}23, SNCH08, SS00, SBN15, TEG18, TBC^{+}16, VJ19, WAO^{+}09, WSND^{+}23, WLS^{+}23, XLJ^{+}09, YLB^{+}22, YPG01, YZN^{+}22, YXK^{+}22, YSHWSH16, AKJ08, BZL^{+}15, BGFAO17, BSS^{+}13, CBCG02, CGG^{+}04, DHI^{+}13, DJ18b, EDR11, FV96, GSC^{+}15, GWAB19, GAB20, HGRT04, HDN^{+}16, HZ11, HJ11b, IZT^{+}07, KV05, Kan15, KHD14, KTY09, KSSI17, LSK^{+}06, LSR18, LVS18, LC15, LSS^{+}21, LKYU12, MDK^{+}16, MG03, NSF12, PCHF18, RZW^{+}21, RGK^{+}08, SPO10, She13, SOA11, SSBD03, SdGP^{+}15, SFWG04, TNWK22, VAZH^{+}09]. efficient [WWB^{+}14, WWZ^{+}09, WWB^{+}19, WHY20, WCSC22, WSS18, YHMR16, YSJR17, YJR17, YLJ18, ZM11, ZHRB13, ZXH^{+}20, ZSTB10, ZZCJ13, vTSSH13, NMLH11]. Efficiently [ACP^{+}01, CSAP21, NRDR05, XWX^{+}22, CJAMJ05]. EgoCap [RRC^{+}16]. Egocentric [EMT^{+}20, JMK^{+}22, RRC^{+}16]. EgoLocate [YZH^{+}23]. eigenfluids [CSK18]. eigenfunctions [DLF12]. Eigenmode [LAJJ14]. Eigenmodes [BZC^{+}23]. Eigenproblems [NH22]. Eigensystem [WYW23]. Eigensystems [SDK19]. Eikonal [IZT^{+}07]. Elastic [LFP21, PZM^{+}15, SPV^{+}16, TB22, TLZ^{+}24, VRP^{+}23, XLC^{+}23, BWR^{+}08, BBG21, CPSS10, CZXZ14, GBFP11, HB21, KMOD09, LHdG^{+}14, LCBD^{+}18, MKB^{+}10, MTGG11, MAKWL22, PMS12, PM21, PLR^{+}16, RKP^{+}22, RLR^{+}21, SJM17, WOR11, WY16, YLYW18, ZSTB10]. Elastic-Solid [TB22]. Elastic-viscoplastic [TLZ^{+}24]. elastica [CK14b]. Elastically [VJ19]. Elasticity [KS12, CS09, DJ17, KDI19, LHZ^{+}21, MZS^{+}11, NKJF09, SBR^{+}15, SWW^{+}20]. Elasticity-inspired [KS12]. Elastics [LLF^{+}20, LYK^{+}21]. elastodynamic [MSW14, MLT17]. Elastodynamics [FSKP23, DJ18a, HLSO12, LSNP13, LGL^{+}19]. elastomeric [JCRA11]. ElastoMonolith [TB22]. elastoplastic [FLGJ19, GTJS17, JWJ^{+}14, WRK^{+}10]. elastoplasticity [JGT17, KGP^{+}16]. elastostatic [JP03]. Electromyography [ZLC^{+}22]. electrostatics [WSSK13]. Element [IRWP23, LHJ^{+}14, LHVT17a, SDG^{+}19, SHG^{+}22, SVB17a, BWHT07, HW16, ISF07, KDI19, LdPS84, LHVT17b, MWT11, MWLT13, SVB17b, TCL21]. Elements [BC14, FPSG22, HLV^{+}17a, LKS15, SHG^{+}22, ARS14, BBB10b, CLC14, CLSM15, CZM^{+}10, EB08, HW15, HLV^{+}17b, IKCM13, JMB^{+}20, KTY09, KBT17, LJM^{+}16, LLK^{+}20, SCGT15, XFAT12]. Eliminating [Xia21]. Elimination [And82, RV89, LVS18]. Ellipses [FH93, McI92]. Ellipsoidal [PVG19]. Ellipsoids [JTMW20]. ellipsometry [HJM^{+}22]. Elliptic [SHG^{+}22]. Elliptical [FH93, KM17]. Embedded [BC23, RK13, SSP07, ALLD17, HCE03, Jam20, LLK^{+}20, NKJF09]. Embedding [JYW^{+}23, PZWW23, XZZ18, JWJ^{+}14, LCDF10, SJZP19, TER^{+}20, ZWL^{+}18]. Embeddings [AGL^{+}22, FBCA23, AL15, AL16, AKL17, CWK^{+}20, LWH^{+}12, LSQ^{+}15, PGH^{+}22]. EMBER [TNWK22]. Embodied [RTB17]. Embree [WWB^{+}14]. Emerging [MCL^{+}09]. Emotion [WZC^{+}20, KAL^{+}17]. Empirical [CMS95, DLR^{+}09, ZBBB18]. Emptying [ZCC16]. EMS [LJZ^{+}23]. Emulating [TDMS16]. Enabling [NFL12]. enclosed [GOMP98]. Encoded [LWP^{+}23, LLWD14, Tar16]. Encoder [GAA^{+}23, TAN^{+}21]. Encoder-based [GAA^{+}23]. Encoding [Van06, HZG^{+}12, LDS03, MKMS04, MESK22]. End [DSJA^{+}21, LSM23, SDP^{+}18, SZD^{+}20, SWF^{+}21, TMM^{+}21, ISSI16, KAL^{+}17, YMJ^{+}21]. End-to-end [DSJA^{+}21, LSM23, SDP^{+}18, SZD^{+}20, SWF^{+}21, TMM^{+}21, ISSI16, KAL^{+}17, YMJ^{+}21]. Endless [HHV^{+}21]. ends [VSK^{+}17]. energetic [BB12]. Energetically [LLJ22]. Energies [BSEH18, ERT14, SDK19]. Energy [CTE05, HP04, HCLK24, LCK22, MCP^{+}09, SJWG20, ZJ12, DLL^{+}18, HGMRT20, Kan15, KUJH21, LWL^{+}09, NSCL08, SSBD03, WCSC22, YCR^{+}15, YTL18]. energy- [WCSC22]. Energy-based [ZJ12]. energy-efficient [Kan15]. Energy-minimizing [HP04]. energy-momentum [KUJH21]. Energy-preserving [MCP^{+}09]. Enforcing [WZX^{+}23]. Engine [MMHP23, SLF22, DNB^{+}05, FMK^{+}03, NPP^{+}11, PBD^{+}10, PVL^{+}05]. Enhanced [CLJ^{+}20, Hud94, Ols92, DFL^{+}15, KK87, VRA^{+}07, VPB^{+}09a]. Enhancement [JLF^{+}23, BM05, BBB^{+}14, BF12, DER^{+}10, ED04, FAR07, GSC^{+}15, GCB^{+}17, HSGL11, JMAK10, KNC^{+}08, LCODL08, LCD06, RSI^{+}08, SGM12, TTD22, WYW^{+}10, WYX11]. Enhancing [UZB^{+}23, MBPY^{+}18]. Enlighten [WZC^{+}20]. enriching [LSGV18]. Enrichment [KMB^{+}09]. Ensemble [BRSMD22, ZZXY21, JTCW07]. ensembles [Xia21]. Entropic [SPKS16]. enveloping [WPP07]. Environment [Ols86, PM18, ARBJ03, JKH^{+}22, LF02, NOP^{+}18, RH02, RZL^{+}10, WPL^{+}21, XMR^{+}11]. environment-independent [NOP^{+}18]. environmentally [CSvRV18]. Environments [CSS96, YPG01, GLY^{+}03, GB08b, KMYG12, KKB^{+}11, LCL06, LNWB03, MIWB02, NHAH03, SCH^{+}14, SMM14, SSC10, SKS02, SBK11, TGD04, WFH10, WS99, WM03]. envyLight} [Pel10]. Epipolar [ABW^{+}17, GF12]. epsilon [DD02a, ITM^{+}14]. Equation [BSSJ23, ABW14, CK11, WZT^{+}08a]. equational [JASR99]. Equations [PM95, AZB09, CI97]. equilibrated [FLGJ19]. Equilibrium [SPV^{+}16, dGAOD13]. Equipped [XWD^{+}22]. equitable [VCA^{+}22]. Equivalence [CQS^{+}23, CCS^{+}21, GSC21b, LZZ^{+}21, RFWB07, SS10b, SSP08]. Equivalent [FM84, MRA^{+}13]. equivariant [PO18]. erasure [LFJG17]. Erosion [CJP^{+}23, SPF^{+}23, YSC^{+}16, CGG^{+}17]. Errata [NMLH14, Spe03]. Error [AAR05, BAU15, CGMS22, LWS^{+}15, LGC^{+}23, WBF^{+}17a, AW20, BDT99, BHW13, CAO09, HJJ10, PSF09, RKZ11, SJJ12, SLWF14, TGB13, WBF^{+}17b, YRPF09, ZG02, ZF03]. error-bounded [ZG02]. error-driven [PSF09]. Error-resilient [AAR05]. error-tolerant [SLWF14]. errors [PMOR10, RP03, Wan14, Xia21]. Escher [NI24, OCNG21, OGN^{+}23]. Escher-like [NI24]. Escherization [NI22]. estimates [BHHM20]. Estimating [Che92, SHM^{+}14, WSM11, ZS00, BB22, CDP^{+}14, HLHZ08, NSJ14, PMOR10]. Estimation [FHXW22, HMI23, LYO^{+}23, SLL^{+}21a, SSBL^{+}22, ZWL22, ZWTP23, ZK22, DJBJ19, GLD^{+}19, GWP^{+}19, GHV^{+}18, HJJ10, HMP^{+}08, JNSJ11, LZHJ20, MRA^{+}22, MSS^{+}17, MTB^{+}13, NOP^{+}18, WHSG97, Xia21, YLB^{+}22, YZX21]. estimator [KDPN21]. estimators [MBGJ22, PCS^{+}20, SOHK16, ZSGJ21]. ETC2 [Nah20]. Étendue [CBYJ23]. ETER [XLC^{+}23]. euclidean [KDH22, ZWL^{+}18]. Eulerian [CCL^{+}22, CM11, DWK^{+}22, FLLP13, HK10a, KDW^{+}17, LLJ^{+}11, LFZ18, MSL^{+}23, MSQ^{+}18, MMTD07, NØ13, QLY^{+}23, SBRBO20, TLK16, WPLS18, WRS^{+}12]. Eulerian-on-Lagrangian [FLLP13, SBRBO20, WPLS18]. Evaluating [HRZ^{+}13, ODGK03, RP07, WF96, CHM^{+}12, CJAMJ05, KP09, KP10, LWC^{+}13, WQF^{+}21]. Evaluation [LCTS05, LC96, MAF^{+}09, MRC^{+}86, RV89, AFR^{+}07, GRG04, ML22, UHT17, WB08]. Event [AECO15, LZHJ20, SSRB^{+}17]. Eventfulness [SDA^{+}23]. events [VBK05]. everyday [VAV^{+}07]. Evolution [BAC^{+}18, CKSV23, MOR^{+}18, LXY^{+}16, MLZ^{+}16, XZCOC12, YLH18].

Evolutionary [ZLZ^{+}23]. evolving [BHLW12, IYAH17, ISN^{+}20, PV06, PKC^{+}17]. Exact [CSL^{+}22, FBCA23, Kla94, RvE93, BDCDA11, BEB12, FV96, QHY^{+}16, SSK^{+}05b, TTWM14, TNWK22]. Exaggerated [RBD06]. examination [WC21a]. Example [BSPP13, DFM88, DBB^{+}17, FJS^{+}17, FRS^{+}12, JWW^{+}20, JTSB16, LWP10, LCL^{+}23, MTGG11, Mer23, RYL13, SDKN18, ST16, SZT^{+}08, WYZG09, WHRO10, WYX11, WHHY20, WZ22, XB17, AVB08, BCK^{+}13, DLL^{+}15, DLKS18, EVC^{+}15, FJL^{+}16, FKS^{+}04, GLLD12, GDG^{+}17, GJWW15, JST^{+}19, JMAK10, KEBK05, LHL10, LYFD12, LBW^{+}14, LFB^{+}13, PCSS06, PALvdP18, RRS13, SSL^{+}14, VSLD13, Wam16, WZT^{+}08b, WPKL17, XUC^{+}14]. Example-Based [Mer23, ST16, WHHY20, BSPP13, DBB^{+}17, FJS^{+}17, FRS^{+}12, JTSB16, LWP10, LCL^{+}23, MTGG11, SDKN18, SZT^{+}08, WYZG09, WHRO10, WYX11, XB17, AVB08, DLKS18, EVC^{+}15, FJL^{+}16, GDG^{+}17, KEBK05, LYFD12, LFB^{+}13, Wam16, WZT^{+}08b, XUC^{+}14]. Example-driven [JWW^{+}20]. Example-guided [RYL13, PALvdP18, WPKL17]. Examples [CPV^{+}23, Gol85a, AF02, FF11, HMLL14, LVGO21, LBDF13, MG03, RTK^{+}15]. excess [WHDS04]. exchange [ZLB16a]. Excite [CAV^{+}23]. exemplar [HCL^{+}18]. exemplar-based [HCL^{+}18]. Exemplars [DBP^{+}15, KFCO^{+}07]. exhaustive [KKN^{+}13]. EXIM [LHH^{+}23]. existing [EKA84]. expanded [JBLL18]. Expanding [LM97]. Expansion [BVF17b, CBYJ23, NCB23, AVF17, DSAF^{+}13, ZZB^{+}18]. Expansions [BXH^{+}18]. Expediting [YLX^{+}15]. Experience [AFP^{+}95, JGC^{+}15]. experiences [MGDB05, SPGI13]. Experimental [BBB^{+}93, MRC^{+}86, SCB87, AJD^{+}10, FNvD82, KKN^{+}14]. Experiments [GHCC88]. Explicit [LHH^{+}23, RBSM19, WWX^{+}22, WYL^{+}20].

Explicit-Implicit [LHH^{+}23]. exploded [LACS08]. Exploiting [PKH^{+}17a, PKH^{+}17b, YRPF09]. Exploration [AZMW21, DPD22, MM22, OLGM11, VRP^{+}23, VVCOS23, BBPP10, BBP21, DFL^{+}15, HFF16, JM12, LZ04, LCX^{+}21, MGDB05, MVH^{+}17, ROA^{+}13, SXZ^{+}17, SWC^{+}18, UIM12, YYPM11, ZLE14]. explorative [YXFH21]. Exploratory [OLAH14, TGY^{+}09, PCS23a]. Exploring [KSSGS11, KLM^{+}12, PBJW14, BYMW13, GBLM16, HWG14, MGS^{+}21, SSS06, TKKT12, YRPF09]. explosions [FOA03, SRF05, YY17]. Exponential [CSAP21, MSW14, BRM^{+}18, SGW06, VJK21]. exponentiation [RWS^{+}06]. Exposing [KOF13, KOF14, OF12]. Exposure [HHX^{+}18, TKG^{+}23, ARNL05, EKD^{+}17, KBC^{+}13, MAF^{+}09, RAT06, TAH^{+}04]. exposures [BM05]. Expression [HTS^{+}22, SGD21, YWS^{+}11, CHZ14, LBB^{+}17b, SLS^{+}12, TZN^{+}15]. expressions [BB14, BBGO11, Gol85b, LCXS09]. Expressive [CTFP05, ZZZ^{+}23, CB05, DBB^{+}17, ELFS16, GCR13, KHS03]. Extended [BN90, MRF06, ANZS18, CMSA20, CZN10, KWK09, SDP^{+}18, KBT17]. Extending [HGF14, RT90]. Extensible [SRX^{+}23, HFF18]. Extension [DS92, AML18, BB17, HPJ12, LHG^{+}09, PSF09, XLC^{+}16, ZLC^{+}13]. extensions [NM16]. Exterior [SW14, dGDMD16]. Exteriors [FW16]. Extracting [BCN08, CZS^{+}13, HZG^{+}12, NGH04, TOS^{+}03]. Extraction [ASK^{+}22, HCW^{+}23, JYW^{+}23, SMH^{+}23, ULP^{+}15, ATC^{+}08, EBCK13, KG04, LLW17, LSA^{+}16, LBK16, RKB04, SAN23, TZCO09, XYXJ12, ZTS09]. Extraordinary [MWM23, CADS09]. Extrapolation [LLK^{+}19, Lév03, WLL^{+}14, ZM13].

Extrema [DKA23, SSD09b]. extreme [DDSD03, ZPBK17]. Extrinsic [CSBC^{+}17a, WBCPS19, CSBC^{+}17b]. Extrusion [HSST10, ZXZL23]. Extrusion-Based [ZXZL23]. extrusions [KW11]. Eye [AKG^{+}23, CBYJ23, DLP^{+}23, MLH^{+}09, ALK^{+}17, BBGB16, CTS^{+}20, CLS^{+}17, Dee05, HCW15, JBM^{+}17, JBLL18, JLF^{+}09, KPM^{+}17, LSL^{+}18, LL13, MGK17, SHL^{+}17, SRL^{+}15, TDM^{+}14, WSXC16]. Eye-box [CBYJ23, JBLL18]. Eye-catching [MLH^{+}09]. Eyebrow [LJZ^{+}23]. Eyecatch [YLNP12]. Eyeglasses [HWBR14, MLR^{+}14]. Eyeglasses-free [HWBR14]. eyelash [XZZ^{+}21]. EyelashNet [XZZ^{+}21]. eyelids [BBK^{+}15, WXLY17]. EyeNeRF [LMM^{+}22]. EyeOpener [SSSH17]. Eyepiece [CBYJ23]. Eyes [LBB02, NN04, SSSH17, BBN^{+}14, LMM^{+}22]. eyeSelfie [SRL^{+}15]. EZ [SLWF14]. EZ-sketching [SLWF14].

Fab [SSM15]. Fabric [GHCG17, FBGZ18, KWN^{+}17, ZFS^{+}19, ZJMB11, ZJMB12]. Fabricable [CML^{+}17, LFZ18]. fabricatable [LOMI11]. Fabricated [IWHH20]. Fabricating [BBJP12, DWP^{+}10, LGX^{+}13, PRJ^{+}13, SDIN18, WPMR09, CLM^{+}13, HBLM11, WW13]. Fabrication [CZM^{+}23, PMLB22, PCB23, SMB^{+}19, TISM16, US24, ZWZ^{+}22, ZXZL23, BBO^{+}10, CZX^{+}16, CLMK17, CLF^{+}18, EBGB14, HZH^{+}16, JMB^{+}20, JWI^{+}21, KCD^{+}16, LDPT13, LSD^{+}16, LMAH^{+}18, LZZ^{+}21, MZL^{+}17, Mit18, NAI^{+}18, PZM^{+}15, PTC^{+}15, POT17, PLKD18, PWLSH13, RMD12, SSL^{+}14, SSM15, SDW^{+}16, VWRKM13, XKCB18, ZKBT17, ZGH^{+}16]. Fabrics [DGVG^{+}23, KSZ^{+}15, MGZJ20, SSBL^{+}22]. Fabulously [Bae18]. facade [BSW13, FMLW14, WYD^{+}14, XFT^{+}08]. facades [CMZP14, MZWV07, SHFH11, ZXJ^{+}13, GGP^{+}20]. Face [AJS20, BKD^{+}08, EST^{+}20, GZC^{+}16, JCFG23, LSC^{+}22, LCXS09, LCC^{+}22, NBLCO20, QLH^{+}22, SGPT23, VBPP05, ZCS^{+}22, BLDA11, BKS^{+}12, CCWL18, CWZ^{+}21a, CLL^{+}21, DSJ^{+}11, FFBB21, GVWT13, GFT^{+}11, HGY17, IKKP17, KS21, LCL^{+}22, LSSS18, PHS^{+}18, SSR20, TDM11, WBGB16, WSS18, YWS^{+}11, YNS19]. face-rig [KS21]. Faces [KCMP23, Li18, WTD^{+}22, ZQL^{+}23, BLS^{+}21, BKD^{+}08, KHS03, WMP^{+}06, ZAJ^{+}15, ZSCS04]. Faceshop [PHS^{+}18]. Facets [RB23]. FaceVR [TZS^{+}18]. Facial [BBB^{+}14, FJA^{+}14, GZX^{+}22, GLC^{+}23, LTO^{+}15, MJC^{+}08, TZS^{+}18, WZC^{+}22, ZZZ^{+}22, ZWS^{+}24, BZL^{+}17, BBB^{+}10a, BHB^{+}11, BBN^{+}12, BB14, BBA^{+}07, BWP13, BHPS10, CTFP05, CWLZ13, CHZ14, CBZB15, CWW^{+}16, CAD^{+}21, CCGB22, FJS^{+}17, GSZ^{+}18, GHP^{+}08, GMP^{+}06, GRG04, GRB^{+}18, HCTW11, JSB^{+}10, KAL^{+}17, LCXS09, LCODL08, LWP10, LYYB13, LBB^{+}17b, LKZ^{+}20, LXC^{+}15, MHP^{+}19, MPK09, MCW^{+}21, OLSL16, PTMD07, SSK^{+}11, SWTC14, SNF05, TZN^{+}15, VWB^{+}12, WSS^{+}19, WBLP11, WYXJ21, XCLT14, YSN^{+}18, ZBGB19]. Facility [LBW^{+}23]. Factor [BSN16, HA18, LRFH13, YBY^{+}13]. Factored [MYRD14, SMPR07, HCW15, KYS^{+}15, LRR04, LCDF10, PvBM^{+}06]. Factoring [WWOH08]. factorization [HPK^{+}17, LHKR10, LK02, LSCO03, NSF12, ZSD^{+}21]. factorizations [HA18]. FactorMatte [GXSD23]. factors [HLSO12]. Fair [NGH04]. fairing [CPS13]. Fairy [OKH^{+}16]. Falling [HYL12]. families [CI97, Wim14]. Family [PP93, LLLL21, LKvK^{+}14, NCVMO05]. Far [GM05, YJR17]. fashion [Bae18]. Fast [Ada21, AFH20, AYL^{+}12, AFO05, APH^{+}14, BODO18, BSSJ23, BWL^{+}23, BDS^{+}18, BZC^{+}23, BZH^{+}23, BDT^{+}08, CGM11, CMMK15, CLSA20, CPWAP08, CL09, DE05, DDP99, DD02b, GAA^{+}23, GDAB^{+}17a, GDAB^{+}17b, GWBN24, HW16, HLP^{+}22, HK18b, HCLK24, JBK^{+}12, KEP05, KWN^{+}17, KP11b, KLV20, LCD^{+}19, LCD^{+}20a, LFH15, LBOK13, LYT^{+}14, LLDL21, MGA^{+}22, ML22, Mai92, MSM^{+}17, Nah20, NSCL08, NKGR06, ODJ04, QHY^{+}16, QJ21, RWW90, SNB07, SMC21, SS10a, SLJT08, SGG^{+}06, STZ14, SSK^{+}05b, FFWL^{+}22, TTWM14, VKJ^{+}17, WPC^{+}14, Wam16, WS21, WCL^{+}23, Wei06, WT08, WWYW21, YMRD15, YCR^{+}15, ZWRY21, ZXS^{+}23, AGDL09, BBB07, BML^{+}14, DLL^{+}18, DFM13, DH06, FDBH22, FHM^{+}21, GS04, LS07, LKL^{+}22, LWO19, LWL^{+}09, Mir98, OK10, PFHA10, PKHK15, PMA^{+}14, RJ07, SHM22, SLMB05, SYBF06, STP12, TTT^{+}17, ZB14, ZZZX21, ZYWK08, TMY^{+}11]. Fast-Multipole-Accelerated [BSSJ23]. Faster [MPB17a, WV92, LAKL11, MPB17b]. Fastest [DLP^{+}23]. FastLSM [RJ07]. fauna [ENCC^{+}21]. FD [NNC^{+}20]. FDM [FZZ^{+}20]. Feasibility [KL17a, KL17b, LW16]. feasible [RH16]. feathers [CXGS02, WG09]. Feature [CMS95, FKY^{+}10, KIM^{+}19, Lee05, LHJ^{+}14, LYP^{+}14, LCC^{+}22, MPKZ10, NLMD12, TBTA^{+}24, WWWG22, WY04, XCOJ^{+}09, XWD^{+}22, ZWGS02, ZMT05, ZVC^{+}20, dLMH10, CWK^{+}20, CT17, HGCO^{+}12, JJJ^{+}21, JDD03, LFB^{+}13, PZ08, PNA^{+}21, TFBW^{+}10, Wes21, XLY09, YNL^{+}21]. Feature-adaptive [NLMD12]. Feature-Aligned [ZVC^{+}20, MPKZ10, XCOJ^{+}09]. Feature-based [Lee05, ZWGS02, ZMT05, dLMH10]. feature-conforming [HGCO^{+}12]. Feature-Line [XWD^{+}22, PNA^{+}21]. Feature-preserving [FKY^{+}10, JDD03]. Features [HWZ^{+}14, PGP^{+}19, DCB^{+}22, FCOS05, GCO06, IMF^{+}21, MRA^{+}22, RSH^{+}05a, WYL^{+}14, WT08, WTGT10].

Feedback [DKNY08, BP12, DK99, YL10]. Femto [VWJ^{+}13]. Femto-photography [VWJ^{+}13]. Femtosecond [OKH^{+}16]. Femtoseconds [OKH^{+}16]. FEPR [DLL^{+}18]. Fermat [ZGH^{+}16]. ferrofluids [HHM19, HM20]. fiber [BDCDA11, JMM09, XWM^{+}20]. FiberMesh [NISA07]. Fibers [KM17, PRM14, MJC^{+}03]. Fibonacci [KISS15]. Fidelity [BLC^{+}22, FLB16, IRG^{+}23, SSW^{+}23, CBZB15, HCTW11, LGA^{+}21, ODGK03, OLSL16, RFWB07, SWTC14, WZC^{+}22, WSS18, XCLT14, YSN^{+}18, ZZC^{+}23, ZWS^{+}24]. fiducial [YMJ^{+}21]. Field [CPY^{+}22, CKSV23, CPMS14, DPW15, HGCO^{+}12, HWZ^{+}20, KKLD23, KQG^{+}23, LBB22, LR15, LTDD16, MHU19, MLS^{+}18, NDS^{+}23, PP94, Pag98, PBM^{+}22, RS14b, STJ^{+}17, SHD^{+}14, SOG^{+}22, VMCS15, WSND^{+}23, XDW^{+}23, ZWS^{+}24, ACBCO17, BHR13, BGL20, CZ17, CRG^{+}20, CBCG02, CNX^{+}08, COSL98, CNR08, CZN10, FBC18, FRSL08, GJTP17, HWR14, HTWB11, HWBR14, HCW15, JTPSH15, JBM^{+}17, JFH^{+}15, JMY^{+}07, KWR16, KHKR11, LHKR10, LWH^{+}11, LL13, LES09, LJM^{+}16, LAC^{+}11, LALD12, LSR18, LHG^{+}09, LNA^{+}06, LLX^{+}12, LK20, LLW^{+}08, LXW^{+}11, MWGZ09, MLR^{+}14, MRK^{+}13, MDC^{+}21, MWBR13, MWHL21, MHP^{+}19, MSOC^{+}19, MPZ14, OHR14, OEE^{+}18, PZ07, PRK^{+}17, RS14a, RVLL08, RVAL09, RSL16, SHL^{+}17, SSY^{+}04, SDP^{+}18, SSD^{+}09a, SHK^{+}17, TAV^{+}10, TPP^{+}11, TLHD03, TLZ^{+}24, WGJ^{+}18, WZK^{+}17, WSZ^{+}18, WLM^{+}15, WLHR11, WLHR12, XNY^{+}16, YJR17, YAV^{+}20, YZX^{+}04, ZWGS02, ZMT06, ZBW^{+}20, vFTS06]. Field-Aligned [SOG^{+}22, CPMS14, STJ^{+}17, JTPSH15, MPZ14]. Field-guided [HGCO^{+}12, CZ17, GJTP17]. field-of-view [MDC^{+}21]. Fields [AOCBC15, BGF^{+}23, BS19, BSB16, BV22, BSEH18, CXW^{+}23b, CO19, CV20, DXG^{+}23, DWS^{+}23, GLC^{+}23, HCW^{+}23, IBB15, IRG^{+}23, JCFG23, LB23, MHU19, OKH^{+}16, PBS20, PLPZ12, PHM^{+}23, RYW^{+}22, RSV^{+}23, RMP^{+}23, SVB17a, TCS^{+}23, WXZ^{+}23, YSHWSH16, ZVC^{+}20, ZTNW23, AGK^{+}22, BSB17, BR21b, CBCG02, CLZ^{+}22, DVPSH15, EHDR11, FSDH07, FBLS07, GRT13, GCH^{+}19, HLHR09, JMB^{+}14, KHH^{+}11, KZP^{+}13, KCPS13, LRAT08, LWH^{+}11, LWB^{+}10, LZC^{+}18, MPDW03, MHP^{+}19, NSB13, PPTSH14, PSH^{+}21, PEVBC21, SZC^{+}22, SVB17b, SV19, TTT^{+}17, VRA^{+}07, WWT^{+}06, XZY^{+}17, ZMSS18, ZHL^{+}05, BMSR20]. Figure [GM84, SZL^{+}23, AHM^{+}15]. Figures [AFP^{+}95, ZB94, HPC21, WFY^{+}10]. Filament [PGK^{+}22, SMB^{+}19, WP10, FZZ^{+}20]. Filament-based [WP10]. Filaments [IWC22]. filigrees [CZX^{+}16]. fill [ZCLJ20]. Filling [Dun83, LMR83, Shn92, TOII08, XLLW20]. Film [ZWTP23, DWK^{+}22, HIK^{+}20, ISN^{+}20, WDK^{+}21]. Filming [SCCB22]. films [DBWG15, IYAH17, TL04, VRBC18]. Filter [MU22, SMH^{+}11, TK05, WDAC06, WFL^{+}15]. Filtered [SGSS22, BCN08]. Filtering [LD11, NMLH14, YMRD15, ZJNZ23, AGDL09, BZCC10, CLKL14, DHC^{+}21, DSAF^{+}13, DD02b, EHDR11, EDR11, GGN18, GO12, HSRG07, KBS15, MS13, MWR12, MWRD13, MYRD14, Nai98, NMLH11, NM16, RKZ12, SD12, Wei06]. Filters [APH^{+}14, YZN^{+}22, Ada21, BJ10a, KS10, LLMZ16, PHK11]. final [GD04, REG^{+}09]. Find [CGM91, Day90, SC20]. Finding [SGSS08, VPR19, CZM^{+}10, TSG^{+}14, ZZCJ13]. Fine [CCK^{+}21, HSG13, VKM^{+}23, KvKSHCO15, SDW^{+}16, WZF^{+}18, WYXJ21]. Fine-grained [HSG13, KvKSHCO15, WZF^{+}18]. finger [GWB05, JHS12]. Fingertips [VVC^{+}15]. finished [MWAM05]. Finishing [TZJ^{+}22].

Finite [BC14, SDG^{+}19, SHG^{+}22, BWHT07, CLSM15, GWAB19, ISF07, KTY09, KBT17, LdPS84, LLK^{+}20]. Finite-Element [SDG^{+}19, LdPS84]. Fire [HBP^{+}21, CJ11, HG09, NFJ02]. First [ASN^{+}20, KCS14, SC18a, RMB07]. first-order [RMB07]. First-Person [ASN^{+}20, KCS14]. fish [IZE^{+}21]. Fisher [ST14]. fishes [SHU^{+}16]. fisheye [RRC^{+}16]. Fit [XZCOC12]. fitted [WVBR^{+}21]. Fitting [CG89, CS09, FB95, Pav83, WPL06, ZLB16b, FCOS05, Gos00, LWC^{+}11, VLV^{+}21, OBS04, YAB^{+}22]. Five [Ano90b, CCW93]. Five-Axis [CCW93]. Five-Year [Ano90b]. fixed [WZ14, YAV^{+}20]. fixed-height [YAV^{+}20]. flakes [PLMR17]. flames [HSF07, LF02]. flapping [JWL^{+}13, WPKL17]. flare [HESL11, BZH^{+}23]. Flash [ED04, SLKS06, ARNL05, HDMR21, HJM^{+}22, KF09, MDKD16, NLGK18, PSA^{+}04, RTF^{+}04]. flash-exposure [ARNL05]. flat [EPM^{+}14, GMB17, MPI^{+}18]. flatland [AR15]. Flattening [FW22, LQGY24, SC18a, MZ12, SLMB05, SSC18]. Flesh [SDK18]. Flexible [DTPC23, GLL^{+}16, GvdPvdS13, SMH^{+}23, ZBJ^{+}23, DML17, GAB20, HDD^{+}16, HST^{+}14, KLV20, MPBC16, MPI^{+}18, NQC^{+}21, OBCS^{+}12, PTC^{+}15, STP12, WWB^{+}19]. FlexISP [HST^{+}14]. FlexiStickers [TT09]. FlexMaps [MPI^{+}18]. FlexMolds [MPBC16]. flicker [KKW21]. Flight [BWC^{+}23, CLT^{+}22, GNHM15, GVNB18, HMI23, KZSR16, KJGP23, ABW^{+}17, CHWH17, HHHW15, JWL^{+}13, KWB^{+}13, MHM^{+}17, NZV^{+}11, SHHW16, UKSI14, WPKL17, cWP03]. flipping [SC20]. Floating [FG14, CLSA20]. floating-point [CLSA20]. floorplans [SWL^{+}22]. flora [ENCC^{+}21]. Floral [IOOI05]. Flow [BSHK04, CCL^{+}22, CPAB22, CHTK24, DYZ^{+}23, GA20, HWB23, LD23, LL23, PLS^{+}15, SS14, SDN18, VBBF16, WSL13, XRW^{+}22, BWHT07, BLR^{+}11, BHN07, CWW13b, CPS13, DWS^{+}20, GGT17, HIK^{+}20, IYAH17, KySK10, LAD08, LZF10, LPL^{+}18, LCD^{+}20a, LLDL21, PCLC16, PNCB21, RXL21, SHM22, SAL^{+}08, TB21, UB18, WJL^{+}20, XCW^{+}20, XFCT18, XTZ^{+}21, XWWZ22, YWS^{+}11, ZQC^{+}14, vW02]. Flow-based [BSHK04]. Flow-complex-based [SS14]. Flow-guided [VBBF16, LPL^{+}18]. Flower [IYYI14, IOOI05, ROC^{+}21]. flowing [NGL10]. FlowRep [GSV^{+}17]. Flows [AZMW21, HWZ^{+}14, LWP^{+}23, QLY^{+}23, Sta03, YSB^{+}15, Aca07, AIH^{+}08, ABO16, CT17, GPH^{+}18, HAB20, LMLD22, NFD07, TWGT10, VBFG12]. Fluid [CZY17b, CHTK24, DLF12, DYZ^{+}23, GPB^{+}19, HH16, JWL^{+}24, KFCO06, LXY^{+}23, LD23, LNZ^{+}23, LWF^{+}22, MTP^{+}18, MSQ^{+}18, MTPS04, ODAO15, PGG^{+}23, RLY^{+}14, RYPZ23, RLSÖ^{+}22, SDN18, XLYJ23, YR23, YNW^{+}23, ZIH^{+}11, ANZS18, AIA^{+}12, ABO16, BGOS06, BGFAO17, BBB07, BHW16, BB12, BHN07, BBB10b, CMT04, CZY17a, CLSK21, DYN03, DKNY08, FHM^{+}21, GPH^{+}18, GNS^{+}12, GAB20, HLW^{+}12, JFA^{+}15, KTJG08, Kim10, KD13b, LJS^{+}15, LAD08, LMH^{+}15, LLDL21, MYH^{+}10, MCP^{+}09, PTC^{+}10, RXL21, RLZ^{+}21, SKM10, TB20, TBBC^{+}22, TLK16, WDK^{+}21, WST09, WTGT10, XFCT18, XIAP^{+}17, YCR^{+}15, YXZ^{+}21, ZNT18, ZM13, ZBG15a, ZLB16a, ZB05, dGWH^{+}15]. Fluid-directed [RYPZ23]. Fluid-Rigid [GPB^{+}19, LXY^{+}23]. Fluid-Solid [LD23, LLDL21]. Fluidic [LDS^{+}22, DWS^{+}20]. Fluids [CPAB22, APKG07, AAT13, Ang17, CKPS17, CTS^{+}21, ETK^{+}07, FQL^{+}20, GBO04, GKHH12, GITH14, HK05, KAGS20, LMAS16, MM13, NWRC22, NSS^{+}19, PICT15, PTG12, QZG^{+}19, RMSG^{+}08, TLP06, YJL^{+}16, YT13, ZB14, ZJ09, ZLQF15]. Fluorescence [LCD^{+}19]. Fluorescent [HFI^{+}08]. fluttered [RAT06]. flux [ZHRB13]. Fluxed [SS17]. Fly [DNZ^{+}17b, DNZ^{+}17a, LYYB13, RTS^{+}07, VSLD13, XDF^{+}19, XNZ^{+}22]. Flycon [NOP^{+}18]. flying [WPL18]. Foam [YSB^{+}15, BDWR12, DWK^{+}22, KLL^{+}07, WFS22]. foams [DBWG15, IYAH17, MDL16, MSDL17]. Focal [MFL17, AWGB04, CKS18, PMOR10, XMZ^{+}14, ZJY^{+}21]. Focus [DPW15, MWH^{+}13, WSP^{+}23, HCW15, KHKR11, LES10, NAB^{+}15, MGT^{+}03]. Focused [OKH^{+}16]. Foids [IZE^{+}21]. Foldabilizing [LHAZ15]. Foldable [FBS^{+}23]. Folded [KMM17b, KMM17c]. Folding [NPO13, KFC^{+}08, NKS^{+}23, ZSMS14]. Foldover [GKK^{+}21]. Foldover-free [GKK^{+}21]. Folds [JHR^{+}15, LSGV18]. Foldsketch [LSGV18]. Foley [LJ14]. foliage [BNB13]. Foliations [HC23, CSZ16, VZF^{+}19]. follower [CSSL21]. font [OLAH14]. FontCode [XZZ18]. Fonts [BLAE22, LZCX19, CK14a, WL21]. Fool [YRPF09]. footage [APS^{+}14]. Force [GIF^{+}18, ZLC^{+}22, RP09, UMK17]. Force-Aware [ZLC^{+}22]. force-sensing [RP09]. forces [BP08, BOFN18, TMOT12]. foreground [RKB04]. foremost [STZ^{+}16]. Forestry [NPA^{+}22]. forests [LJS^{+}15]. Form [TSG^{+}14, BBG12, FXBH16, GSV^{+}17, HR05, KH06, KG08, MKZ^{+}21, Nas87, UPSW16, WP09a]. Form-finding [TSG^{+}14]. Formal [DFM88]. formation [DFW20, DKNY08]. Formed [JTMW20, SY22, UKSI14]. Forming [CJP^{+}23]. Forms [Sei93, KPWP17, SSM15, WC21b]. Formula [LXSW23, LSSW19, HRV97]. formulas [LKF12]. Formulation [KM97, HGMRT20, KTY09, MRA^{+}13, MGJ19, PICT15, SXH^{+}21]. formulations [LDS02]. formwork [ZFS^{+}19]. Forward [Kla91b, Kla94, PR06, Rap91, FD17, GITH14, ZD20]. foundation [LBDA21, MCS15]. foundations [Gol02, LKB^{+}22b]. Four [CCW93, ZCW^{+}17]. Four- [CCW93]. four-view [ZCW^{+}17]. Fourier [AMZ99, Les20, Mal93, Ng05, SHD^{+}14, SSD^{+}09a, SK13, WPC^{+}14, XSH^{+}20]. Fournier [Fiu00]. Foveated [GFD^{+}12, KJS^{+}19, KKW23, KKW21, PSK^{+}16, TTD22, TAKW^{+}19, WKF^{+}21]. foveation [SHK^{+}17]. FovVideoVDP [MDC^{+}21]. Fractal [OCNG21, VR94]. Fractional [GCH^{+}19, OKRC10]. Fracture [FCK22, SLM^{+}23, BDW13, CYFW14, HW15, HW16, MCK13, OBH02, WFL^{+}19, ZJ10, ZBG15b]. fractured [HFG^{+}06]. Fracturing [CCL^{+}22, PKA^{+}05]. Fragment [DCOY03, FBH^{+}10]. Fragment-based [DCOY03]. Fragments [LH17a, BTFN^{+}08, LH17b, MP07, TFBW^{+}10]. Frame [CK20, FF88, GBFP11, HZ82, JFH^{+}15, LB23, PBS20, PPTSH14, SOG^{+}22, WKMH^{+}23, Wes88, BDM^{+}21, CDP^{+}14, HB89, HTWB11, HZH^{+}16, HPK^{+}17, JKT^{+}15, LCORL07, RSL16, SFG^{+}13, TDMS16, WHSL11, WWY^{+}13, WGDE^{+}19]. Frame-based [GBFP11]. Frame-to-frame [HZ82]. FrameFab [HZH^{+}16]. Frames [CC23, LLK^{+}19, ZVC^{+}20, BHB^{+}11, CC19, WJZL08, YGM97]. Framework [GRS93, HHX^{+}18, HPP^{+}22, HZL22, KK91, LR15, MHU19, NCB23, PCB23, SGD21, TLZ^{+}24, ZTD^{+}23, ZZC^{+}22, AZB09, BGKS17, BT19, BAGL19, BBB07, BLDA11, BZCC10, BRM^{+}18, BK04, DFL^{+}15, GM05, GWAB19, GKS02, GMG^{+}20, HJJ10, HST^{+}14, HK10a, HMG03, HSK16, HMC11, HHN^{+}02, JAM^{+}10, JdJM14, JMM^{+}14, JAG18, JSP17, KKN^{+}14, KS98, Leh07, LSD^{+}22, MMG06, MJBF02, MSSG^{+}21, NIR^{+}21, PTSO15, RH04, RLR^{+}21, SAN23, SHM22, SY21a, WWB^{+}14, WSP18, YCL^{+}17, YKC^{+}16]. Frankencamera [AJD^{+}10]. FrankenGAN [KGS^{+}18]. Free [ASGS23, BWC^{+}23, CTMS03, CTFH22, CRCM23, HWZ^{+}14, HWP^{+}23, KG08, LSM23, MKZ^{+}21, NGL10, PMGD21, SMGC23, AZB09, BBG12, CMMK15, CCS^{+}15, Csé19, DWW^{+}18, FLS^{+}21, FFB^{+}09, FL16, FKN17, GCD^{+}20, GKK^{+}21, GSV^{+}17, GKTT13, GHZ18, HR05, HPP^{+}18, HTYW22, HWBR14, KH06, LKL^{+}22, LMY^{+}22, LD21, LFS^{+}20, LCOLTE07, LHR^{+}21, LCBK19, MMT18, Nas87, SSJC22, SLL19, SOA11, SS15, SKM10, SPGI13, TBV12, UKSI14, UPSW16, Wan18a, WJF^{+}22, WG09, XWWZ22, XRLF15, YCR^{+}15, YZL^{+}22, ZLY^{+}21, ZYQ^{+}14]. free-flight [UKSI14]. Free-flowing [NGL10]. Free-form [KG08, MKZ^{+}21, BBG12, GSV^{+}17, HR05, KH06, Nas87, UPSW16]. free-formed [UKSI14]. Free-space [BWC^{+}23]. free-surface [XWWZ22]. free-view [LD21, LHR^{+}21]. Free-viewpoint [CTMS03, PMGD21, CCS^{+}15, GCD^{+}20, HPP^{+}18, YZL^{+}22, ZLY^{+}21]. Free2CAD [LPBM22]. Freedom [IWC22]. Freeform [DGH16, FSH11a, JRPW20, PSB^{+}08, BK04, EKS^{+}10, EC96, JMB^{+}20, KOY^{+}11, LPL^{+}17, LPL^{+}18, NISA07, PLW^{+}07, RKP^{+}22, TISM16]. freehand [GHL^{+}20, HFL14, LZC11, LPBM22, WQF^{+}21]. Freely [TTZ^{+}20]. FreeStyleGAN [LD21]. Frenet [HB89]. Frequency [BBS14a, ETH^{+}09, EHDR11, FN20, HSRG07, HMI23, RH02, AWL13, ADM^{+}08, BDT^{+}08, CTH^{+}14, DHS^{+}05, LHG^{+}09, MAC22, NKGR06, NRH03, NRH04, OHX^{+}14, SKS02, SXZ^{+}20, TS06, WTL05, WTL06b, WRG^{+}09, XCM^{+}14]. Frequency-domain [FN20, BDT^{+}08]. fresco [BTFN^{+}08, TFBW^{+}10]. Friction [MHNT15, BDCDA11, BFA02, CFW13, DBDB11, LCBD^{+}18, MTB^{+}13]. Frictional [LFP21, LDW^{+}23, Dav20, DJBDDT13, GHZ^{+}20, GHF^{+}18, JGT17, KEP05, KSJP08, LDN^{+}18, LJBBD20, RLR^{+}21]. FrictionalMonolith [TB21]. friendly [SPJT10, SSK^{+}11]. From-Region [VKW^{+}23, LSCO03]. frothing [CPPK07]. Full [CK20, PYA^{+}24, WLZ^{+}21, WZQ^{+}18, YXW^{+}23, ZZZ^{+}23, BWS^{+}21, Fre16, HHC^{+}19, HW12, KE18, PRMG16, TMDK15, WZC12, ZSZ^{+}14]. Full-body [PYA^{+}24, ZZZ^{+}23, BWS^{+}21, HHC^{+}19, KE18, PRMG16, WZC12, ZSZ^{+}14]. Full-frame [CK20]. Full-Wave [YXW^{+}23]. Fully [YI17, CSW^{+}16, HK10a, LHM^{+}18, SSISI16]. fully-Eulerian [HK10a]. Fun [Mit18]. Function [GRS^{+}17a, LBB22, LKM^{+}23, US24, XWC^{+}16, ATW15, GXZ^{+}13, GRS^{+}17b, HvKW^{+}16, JP03, LD05, MAC22, Rus19, VSJ22]. Function-Based [US24]. Functional [CSBC^{+}17a, CSBC^{+}17b, CO19, DGHM93, DWS^{+}20, HWG14, OBCS^{+}12, ACBCO17, CI97, FSL^{+}15, FD17, PYB^{+}16, RPWO18]. Functionality [LKWS16, ZAC^{+}17, HZvK^{+}15, HYZ^{+}18, LMS13]. Functionality-aware [ZAC^{+}17]. functioned [HKS17]. Functions [GVNB18, NID20, SWWW15, BX03, BHSH^{+}22, CTW^{+}04, CBW^{+}18, CJAMJ05, DLC^{+}15, DZCJ22, FLSG14, GJWW14, HHA^{+}10, KBD07, MSS^{+}12, MIB15, NGH04, PSF09, RWG^{+}13, TZL^{+}02, TS06, VRM^{+}18, YYW12b, ZM11, ZDI^{+}15]. Fundamental [SHW19, DJ17, DJ18a]. Fundamentals [GGS03]. Furniture [YKGA17a, FSY^{+}15, LOMI11, LHAZ15, LHLF15, MSL^{+}11, SLR^{+}16, SFJ^{+}17, UIM12, YKGA17b, YYT^{+}11]. Further [AFP^{+}95]. Fused [SMB^{+}19]. Fusing [OKH^{+}17, BML^{+}14]. Fusion [FG11, WLS^{+}23, DMB^{+}14, KKW21, LSC^{+}22, LK20, LOW18, MSOC^{+}19, DKD^{+}16, TZY^{+}23, WPL^{+}21, XNZ^{+}22]. Future [EST^{+}20, CTH^{+}14]. Fuzzy [Ree83, KT03, KLM^{+}12].

Gabor [GLLD12, LLDD09, LD11].

GADGET [FH04b]. gait [WP09a]. Galaxy [HC23]. Galerkin [EB14, HCH22, SSW^{+}13]. galleries [XZCOC12]. gallery [WPL^{+}21]. Game [MSL^{+}24]. games [KGBS11, SHK^{+}14, WAH^{+}10]. Gaming [AKG^{+}23]. Gamut [SCB88]. gamuts [MGS^{+}21]. GAN [GWLG23, LLHF21, WBZ22, XFCT18, ZAFW21]. GAN-based [ZAFW21]. GANeRF [RMP^{+}23]. GANimator [LAZ^{+}22]. GANs [GSZ^{+}18, KGS^{+}18]. Gap [YW13, DHL14, HYG^{+}13]. gaps [ABO16]. GARM [LNZ^{+}23]. GARM-LS [LNZ^{+}23]. Garment [CZL^{+}15a, LXL^{+}23, RKS^{+}14, YPA^{+}18, BSK^{+}16, BME21, BPS^{+}08, BSBC12, PDF^{+}22, SMD^{+}15, UKIG11, WCPM18, WSH19]. GarmentCode [KSH23]. Garments [ZCM22, BGK^{+}13, KWL^{+}21, KL22, LSGV18, LHZ^{+}21, RC22, ZWCM21]. gas [AIH^{+}08]. gases [FOK05]. gated [PVG19, WCRZ21]. gathering [QSH^{+}15, REG^{+}09, SZLG10]. gauge [YXZ^{+}21]. Gauss [FTP16, HCLK24, LXSW23, LSSW19, SY21b, ZCT22]. Gaussian [AGDL09, ARW22, BJ10a, DKA23, GCH^{+}19, IAF09, KKLD23, KWN^{+}17, LLR^{+}15, PBW19, ZFWW18]. Gaussian-product [PBW19]. Gaussians [XSD^{+}13]. Gaze [JSSH15, KAW20, KPB^{+}12, TZS^{+}18, ATM^{+}17, BMSG09, KKW20, MSM^{+}17, PSK^{+}16, PRMG16, WSXC16, WKHA18]. Gaze-Aware [TZS^{+}18]. Gaze-Contingent [KAW20, ATM^{+}17, KKW20, MSM^{+}17]. Gaze-Driven [JSSH15]. gaze-tracked [PSK^{+}16]. GazeStereo [KDM^{+}16]. GCN [SFD^{+}22]. GCN-Denoiser [SFD^{+}22]. gems [GS04]. gemstones [GS04]. genBRDF [BLPW14]. General [CPW^{+}23, FH93, GUPZ20, HPP^{+}22, KK91, Lev84, LXW^{+}11, MSSG^{+}21, ZPYX23, AW11, GS85, GMG^{+}20, HTYW22, MMT18, NH08, PBD^{+}10, RAR^{+}21, SJ22b, STXJ15, TLK09, WSP18, ZHWW12, ZZCJ13]. General-Purpose [Lev84]. Generalization [Bli82, GNHM15, LD89]. Generalized [BHW16, BK85, BK87, CBvdP10, FHM^{+}21, Lew87, LKB^{+}22b, Pet89, PM21, PLC^{+}21, Sai89, SPGT18, SM06, ZYH^{+}15, AMZ99, CDP^{+}14, GTJS17, JKSH13, TK14, TKY^{+}17]. Generalizing [IAF09, JW23, RTK^{+}15, WPP14]. Generate [WZ22, JBX^{+}20, SWL^{+}22]. Generated [AZMW21, BS88, BS90, KPACO22, RBSM19, MSK10, OHR14, TL04, WQF^{+}21, YGM97, ZAJ^{+}15]. Generating [BYMW13, GAL^{+}09, HA92, LY23, RH16, WLO^{+}14, WLL23, ZSSJL20, IZE^{+}21, KSH^{+}16, LDS^{+}11, MPK09, PGML^{+}19, NCVMO05]. Generation [CBYJ23, CWL22, CSL^{+}22, GLC^{+}23, HHL^{+}24, LYC^{+}22, LHH^{+}23, PCS^{+}23b, PC82, SFC^{+}23, VW94, VLA15, WV92, XZP^{+}23, YML^{+}23, YIC^{+}14, ZQL^{+}23, ZPW^{+}23, Zyd88, AF02, BDK^{+}16, CLL^{+}21, CSHD03, DK09, DH06, FH04b, GJTP17, GGG^{+}13, GLY^{+}03, GASP08, GLP^{+}22, HPG^{+}22, HZP^{+}22, JBP06, JJJ^{+}21, JFH^{+}15, JYQ^{+}22, KAB^{+}10, LHM09, LdPS84, LPRM02, LACS08, LKZ^{+}20, LLM21, LLHF21, LKvK^{+}14, MCC09, RSL16, RCOL09, SP16, TPSHSH13, TS08, TWAD09, VPHB^{+}21, WMC11, YMJ^{+}21, YCL^{+}20, Zhu18b, VW95]. Generative [BTSB23, HDMR21, JCFG23, LPX^{+}19, LCL^{+}23, NAH^{+}22, YSCL22, ZYM^{+}20, ZTNW23, ZCP^{+}23, BSP^{+}19, BHMK^{+}18, GHBCO21, GWY^{+}21, GDG^{+}17, GHS^{+}22, GSH^{+}20, HYZ^{+}18, LXC^{+}17, MC12, TTR^{+}17, WSH^{+}18, WWL^{+}19, ZQCL19]. Generator [CLX^{+}22, LLB24, QLH^{+}22, PGML^{+}19]. Generators [YSC^{+}23, GPM^{+}22, PV06]. Generic [CRCM23, GGT17, SY21a, LSK^{+}06]. Genetic [Sah18, SAMWL11, BLPW14]. gentle [BP08]. Genus [CSZZ20]. Geo [WTD^{+}22]. Geo-Metric [WTD^{+}22]. Geodesic [AFH20, CSRP10, LFXH17, NPP22, PZWW23, PHD^{+}10, RHSH18a, LXY^{+}16, PM21, PO18, QHY^{+}16, RHSH18b, SC20, VZF^{+}19, XW09, YWH13]. Geodesics [CWW13b, SSK^{+}05b, YXH14]. GeoLatent [YSC^{+}23]. Geometric [ACP^{+}01, BG89b, Boi84, BR94, BBGO11, CCK92, DB88, EM90, FH97, Gol84, Gol85a, KCŽO08, KMP07, LPW^{+}06, Mil87, NN90, PPV95, POK23, SPSH^{+}17, TWBO03, TR98, TQ94, WYW23, YSC^{+}23, BLTD16, CPSS10, DLX^{+}21, GCO06, GP08, Gol02, GJWW14, HPSZ11, HB89, HZvK^{+}15, HFG^{+}06, IYAH17, JASR99, KOY^{+}11, KGL16, LdPS84, LKG^{+}03a, LZ14, LJGH11, LJO19, MRA^{+}22, MJBF02, PCK^{+}08, PKZ04, PM05, SAZK06, SdGP^{+}15, SD89, THW^{+}14, WFL^{+}15, WS21, WBZ22, WNEH22, YNS19, ZHW^{+}06]. geometrical [VABW09]. Geometrically [Sei93, BEB12, JBP06, RvBB^{+}03]. geometries [WDW^{+}15]. Geometry [BBR^{+}21, CCK92, CSBC^{+}17a, FGN84, GGH02, GXY^{+}17a, HZC^{+}22, LMS13, LWL^{+}23b, LH04, OHHD18, PK05, PLW^{+}07, RVAL09, SRH^{+}15, SGWJ18, SRB^{+}19, TLG17a, UZB^{+}23, WBCPS19, WLJ^{+}22, WC90, WA23, XZP^{+}23, YML^{+}23, ZSSJL20, ZRJ23, Zhu18b, dGMMD14, AMD02, AAM03, ABO16, BBB^{+}10a, BW13, BBA^{+}07, Bou18, BBB10b, CLSM15, CLL^{+}21, CK11, CSBC^{+}17b, DLSCS08, DHOO05, FKY^{+}10, FV96, FMR20, GVWT13, GSC21a, GF12, GMP^{+}06, GXY^{+}17b, HDA17, HLZ10, KV05, Kal18, KS04a, LAGP09, LCOLTE07, ML22, MZPS21, MGP10, MGP06, Mit18, MMTD07, NRDR05, NJJ21, PBS04, PKKG03, PMW^{+}08, PDZ^{+}18, PGZ^{+}19, RMBB^{+}13, SR00, SSM15, SS21, SNW21, TLG17b, TEG18, WYZG09, WGP^{+}10, YSN^{+}18, YHZ^{+}14, ZGZJ16, dGDMD16, WC91]. Geometry-Aware [XZP^{+}23, OHHD18, RVAL09, SRB^{+}19, DLSCS08, PGZ^{+}19]. geometry-based [AAM03].

Geometry-Constrained [WLJ^{+}22]. Geometry-guided [PK05]. geometry/impostor [DHOO05]. Geopostors [DHOO05]. Geostatistical [MK05]. geoTangle [NPP22]. Gestalt [NSX^{+}11]. Gesticulator [AGL^{+}22]. gestural [GWB05]. Gesture [AGL^{+}22, AZL23, LKTK10, NKAS08, PQF^{+}23, BVS16, LGK^{+}16, SN17, TFK^{+}03, YCL^{+}20, BVS16]. GestureDiffuCLIP [AZL23]. gestures [RTK^{+}15]. gesturing [JHS12]. Get [Xu18]. Getting [McI92]. Ghost [SB12, FKN17, GKTT13]. ghost-free [FKN17, GKTT13]. ghosting [SLV^{+}13]. GHz [VCA^{+}22]. gigantic [CGG^{+}04, IG03]. Gigapixel [HLSH18, KUDC07]. Gigascale [QRL^{+}23]. Gimbal [SCCB22]. GIPC [HCLK24]. Girth [XWC^{+}16]. GJK [MPB17a, MPB17b]. GKS [DFM88]. Glacial [CJP^{+}23]. glaciers [AGP^{+}20]. Glare [RAWV08, TAHL07]. glass [GGP^{+}20, WGL^{+}18]. glasses [FKN17, SLV^{+}13]. Glift [LSK^{+}06]. glints [YHJ^{+}14]. Glinty [DLW^{+}22]. Global [BYRN17a, BR07, CZ11, CSS96, CLSS97, CC23, DPD22, GZS^{+}22, MZ12, PTSZ11, RWG^{+}13, VMKK00, WHSG97, WSH^{+}18, XHWW22, AFO05, BYRN17b, BAERD08, BLDA11, BMW^{+}09, BCW17, CBK15, CNR08, DSDD07, DKH^{+}10, DKZ^{+}21, DDP99, FLB16, GD04, ISSI16, JSKJ12, KJDL09, KFB10, LALD12, LWC^{+}11, LXY^{+}16, MA06, MZ13, MPZ14, NKGR06, OHX^{+}14, RLL^{+}06, RLP^{+}20, SZC^{+}22, SL17, SFWG04, SKC^{+}14, TL04, TMRL14, TPWG02, VAZH^{+}09, WWZ^{+}09, WS99, YNW16, YSJR17, ZCLJ20]. Global-to-local [WSH^{+}18]. Globally [DNZ^{+}17b, FW22, ISSI17, KLS03, KCPS13, XDW^{+}23, ZLWH16, DNZ^{+}17a, FLJK21, HPC21]. GlobFit [LWC^{+}11]. Gloss [BOD^{+}13, PFB^{+}20, TDR^{+}12, WAKB09]. Glossy [CSS96, CLSS97, RLP^{+}20, DKH^{+}10, HKWB09, IDN12, LKYU12, SM06, WTL06b, WSM11]. glove [GWP^{+}19, WP09b]. Glyph [XZZ18, LN22]. Goal [YIC^{+}14]. Goal-Based [YIC^{+}14]. GoLD [BGB^{+}05]. Goniochromatic [RB23]. Good [SLM^{+}23, BYMW13, LS07, PL14, YLJ18]. google [BBGO11]. GPU [ASA^{+}09, BFGS03, BFK^{+}16, CKIW15, CW17, DKHS14, GWW^{+}18, GBK05, HR05, HG09, HZG08, HZG09, JRSS21, JCW09a, KKSS18, KB12, KPM16, KW03, LMY^{+}22, LSK^{+}06, LHZ16, LTT^{+}20, LB06, MPO21, NMLH11, NMLH14, NLMD12, RSLL18, SF09, SJP05, SKB^{+}14, TWL^{+}18, WWZ^{+}09, WHY^{+}13, WY16, Wan21, WCL^{+}23, WWYW21, WWW22]. GPU-accelerated [CW17, KB12]. GPU-based [WCL^{+}23, CKIW15, GBK05, HR05, TWL^{+}18, WWZ^{+}09, WHY^{+}13, Wan21, WWYW21, WWW22]. GPU-decodable [KPM16]. GPU-efficient [NMLH11, NMLH14]. GPUs [BSL^{+}16, BFH^{+}04, CM14, FBH^{+}10, KGB^{+}09, SS10a, SKK^{+}12, ZHX^{+}07, ZHR^{+}09]. GrabCut [RKB04]. Gradient [BPE17, CM21, FLW02, GHV^{+}18, KMA^{+}15, LKL^{+}13, LNZ^{+}23, PKCH18, SMH^{+}23, XZY^{+}07, Aga07, ARNL05, BZCC10, DKT^{+}23, GFT^{+}11, GBC^{+}13, HSL^{+}06, KH08, KSH10, KHL19, KLS^{+}13, LHM09, MRK^{+}14, MKD^{+}16, MP08, MHP^{+}19, NIR^{+}21, SLWS07, XLXJ11, YZX^{+}04]. Gradient-Augmented [LNZ^{+}23]. Gradient-Based [SMH^{+}23, GBC^{+}13, NIR^{+}21]. Gradient-domain [BPE17, GHV^{+}18, KMA^{+}15, LKL^{+}13, PKCH18, Aga07, BZCC10, KH08, KSH10, KHL19, MRK^{+}14, MKD^{+}16, MP08]. gradients [BFGS03, BBG12, CKMR^{+}21, FLB17, MHM^{+}09]. GradientShop [BZCC10]. grading [BSPP13]. Grain [ZWTP23]. grained [HSG13, KvKSHCO15, WZF^{+}18]. grains [YSC^{+}18, LPX^{+}19]. grammar [LCK^{+}14, SP16, ZXKL^{+}20]. Grammars [Mer23, DLC^{+}15, LWW08]. GRAMPS [SFB^{+}09]. granular [DBD16, MPH^{+}15, MPG^{+}16, NGL10, TB21, YSC^{+}18]. Graph [FH97, KL17a, MWC^{+}23, Mer23, PZWW23, SFD^{+}22, WSL^{+}19, ACXG09, FSH11b, GSRN21, KL17b, KSE^{+}03, LZT^{+}19, LVS^{+}13, PRAV09, RKB04, WLL^{+}14, WLW^{+}19, WLT22, YWH13, ZXKL^{+}20, ZHS^{+}05]. Graph-constructive [FH97]. graph-cuts [LVS^{+}13]. Graphcut [KSE^{+}03]. Graphic [Cas91, WZHL23, ZZL^{+}21, ZCL18, ZQCL19]. Graphical [Bar86, HC86, Mac86, OBH02, PK83, Res87, SG91, UTB^{+}19, FNvD82, LZH^{+}17]. Graphics [AMS03, CM83, CT82, Coo86, DMZ^{+}17, GF82, GS04, LMR83, LN84, Lev84, MRC^{+}86, OKH^{+}16, Pik83, Wes88, WW82, ZMW^{+}23, AMN03, AHAM15, AAM03, AČMS10, BKKL15, BDM09, BFH^{+}04, CHM^{+}12, CTH^{+}14, DRvdP14, DRvdP15, DN02, DNB^{+}05, EPD09, FSH11a, FH11, GLdFN14, GM05, Gol02, Gue07, GFD^{+}12, HGF14, HMG03, HCW15, JP02, KKSS18, KTL^{+}04, KKW21, KFS13, LWA^{+}12, LLGRK20, LHLK10, LB05, MGAK03, MESK22, MCHAM06, NH08, OHR14, PTSO15, PBMH02, RKLC^{+}11, RLR^{+}21, SFLM04, SHL^{+}17, SFB^{+}09, WP06, ZHWG08, Pav90, WP90, Bea88]. Graphs [AFH20, HWB23, HTCH15, LVY16, RSP23, BDK^{+}16, DH96, DHC^{+}21, FCW^{+}17, KGP02, LHLY21, LRFH13, LCK^{+}14, MC12, PSBM07, PKC^{+}16, RP07, RCLM19, SH07, SPGT18, She13, SLH^{+}20, SWL^{+}22, YBY^{+}13, JTCW07]. grasping [Liu09, SHX^{+}22, ZZMC13]. GRASS [LXC^{+}17]. Gray [DSZ17, KJDL09]. grayscale [XLW18]. greedy [RKZ11]. Green [AAPS16, LLCO08, AAPS17, JP03, LCK22]. Green-Screen [AAPS16, AAPS17].

Gregory [LSNC09]. GREIL [CPV^{+}23]. GREIL-Crowds [CPV^{+}23]. greyscale [WAM02]. Grid [And82, CPAB22, LPC22, SSJC22, SMGC23, BCE^{+}13, CPD07, CMSA20, CM11, CMMK15, JLS^{+}03, LZF10, LHZ^{+}18, PLC^{+}21, SABS14, SdS02, WIK^{+}06, XCW^{+}20, YY17, ZLC^{+}13]. Grid-Based [CPAB22, JLS^{+}03, PLC^{+}21]. Grid-Free [SMGC23, SSJC22, CMMK15]. GRIDiron [MCS15]. Grids [McI83, AGL^{+}17, EB14, LH04, NG18, PM21, ZG02]. Gridshells [BSR^{+}23]. gripper [SHX^{+}22]. gripper-object [SHX^{+}22]. grippers [KGL^{+}22]. gripping [YYL22]. GroomGen [ZCP^{+}23]. grooves [XH18]. Group [KLLT08, LDD^{+}23, BT19, CGM11, KCD09, LZH^{+}17, WZF^{+}18]. Grouping [Gos00]. Growing [SB95, Che13]. growth [DFW20]. GST [KPM16]. gTangle [SP16]. Guaranteed [FLJK21, MC21, YJY23, MGT^{+}03, SJ22b, VSK^{+}17]. Guaranteed-quality [MC21]. guarantees [GMP09]. Guarding [KC23]. Guest [Tan83, BG89b, BG89a, BG90, Fol86a, Fol86b, Fol86c, FGN84, FR87, Fuc82, Pha18]. Guidance [CLJ^{+}20, CAV^{+}23, SFC^{+}23, ZQL^{+}23, HXZW20, HKAK14, LZC11, LXS^{+}18, WFL^{+}15]. Guide [NB11]. Guided [CZL^{+}14, DCT^{+}22, LWL23a, LVY16, LHH^{+}23, MSL^{+}23, SZW^{+}23, UIM12, ULP^{+}15, WWZ^{+}06, WFS22, WSML23, ZCM22, ALY^{+}21, BAC^{+}23, BLR^{+}11, BBPD12, CZ17, CAWH16, FJL^{+}16, FHM^{+}21, GPM^{+}22, GJTP17, GSZ^{+}18, HGCO^{+}12, KGS^{+}18, KLF12, LPL^{+}18, LLHF21, LLHY22, MSL^{+}24, MTA^{+}20, MTM16, MGC^{+}19, PQF^{+}23, PCSS06, PK05, PALvdP18, RHJD18, RYL13, SGM12, SDK21, SHK^{+}17, VBBF16, WLZ^{+}09, WPKL17, WSL^{+}14, XK07, XZY^{+}17, YSW^{+}20, YXFH21, YNL^{+}21, YCZ11, ZZI^{+}17, ZLH^{+}21, ZXS^{+}12]. guidelines [MSL^{+}11, MSOC^{+}19]. Guiding [FHG^{+}23, HZE^{+}19, ZXS^{+}22, FN20, SY05, ZXS^{+}21].

Haar [LF08]. HACK [ZZC^{+}23]. Hair [BW22, HWP^{+}23, KM17, PCK^{+}08, SSW^{+}23, XWH^{+}23, YSK09, ZCP^{+}23, BBN^{+}12, BAC^{+}06, CWW^{+}12, CWW^{+}13a, CLS^{+}15, CSW^{+}16, DBDB11, DJBDDT13, EBGB14, FMB^{+}17, HZW12, HMLL14, HMLL15, JMM09, JGT17, KN02, LHM^{+}18, LLR13, MJC^{+}03, MSW^{+}09, MM06, MWM08, PBS04, RZL^{+}10, SPJT10, SHM^{+}18, SLF08, WYZG09, WOQS05, WZC^{+}22, XYH^{+}21, XMR^{+}11, XWW^{+}14, ZCW^{+}17, ZWW^{+}18, ZYWK08, ZRL^{+}09]. hairs [CZZ14]. hairstyle [HMLL15]. hairstyles [HML^{+}14, PCK^{+}08]. Haisor [SYM^{+}24]. halfspaces [DZCJ21]. Halftone [CCLM13, KP18, PH15a]. halftoned [KL12]. Halftones [Knu87]. Halftoning [GRS93, PQW^{+}08]. Halfway [LLN^{+}14, EBJ^{+}06]. halide [AMA^{+}19, LGA^{+}18, MAS^{+}16]. Hallucination [KAEE20, GWM^{+}08, SPDF13]. Hamiltonian [LLR^{+}15]. Han [YXH14, XW09]. Hand [ANL^{+}23, LFL^{+}23, SGPT23, WMB19, YBMN^{+}23, ZZT^{+}21, ZWHB22, CWL12, DLKS18, GWP^{+}19, HLW^{+}18, IBP15, JSMH12, LZQ^{+}22, SWW^{+}20, SKP08, SKC^{+}14, TBC^{+}16, TPT16, TTR^{+}17, WP09b, WMZ^{+}13, WMB^{+}20, XWSY15, YL12, ZBYX19, ZYSK21, ZYQ^{+}14]. hand-colored [DLKS18]. Hand-Drawn [YBMN^{+}23, JSMH12, SKC^{+}14, XWSY15]. hand-held [CWL12, IBP15, ZYQ^{+}14]. Hand-Object [LFL^{+}23, ZZT^{+}21, ZBYX19, ZYSK21]. hand-tracking [WP09b]. handed [LKG^{+}03b]. Handheld [WGDE^{+}19, HWV^{+}18]. Handle [AFTCO07, DLSCS08, RCPO21, She13]. Handle-aware [AFTCO07]. handle-based [RCPO21]. handles [YK14, YCHK15]. Handling [FG90, WCL^{+}23, MCKM15, TWL^{+}18, WAK20]. Hands [TSLP14, DYY16, MDB^{+}19, RTB17, SSB^{+}15, SDO^{+}04, TTT^{+}17]. hands-on [DYY16]. Handwriting [GWLG23, HAB16, LZCX19, Zit13]. haptic [LSCS14, OL03]. Hardware [HYS23, NKK^{+}14, VKJ^{+}17, AMN03, AMS03, AHAM15, AAM03, BKKL15, BFH^{+}04, CTS^{+}20, CBCG02, DFM13, FH11, HBD^{+}14, HDD^{+}16, HMG03, JP02, LB05, LSNC09, MGAK03, MCHAM06, NPP^{+}11, NL13, PVL^{+}05, PBMH02, WFH^{+}07, ZHWG08, JLBM05]. Hardware-Accelerated [HYS23, PVL^{+}05]. hardware-in-the-loop [CTS^{+}20]. harmful [SLS^{+}16]. Harmonic [Ale19, BCW17, CAJ09, ESBC19, FW22, JMD^{+}07, WSSK13, ZJ09, BCWG09, CW15, CCW16, LW16, NSF12, RWS^{+}06, TZCT20, TFG^{+}13, WR18, WS21, WGS23]. Harmonics [BXH^{+}18, MWM08]. harmonization [COSG^{+}06, SJMP10]. hash [MESK22]. hashing [ASA^{+}09, GLHL11, LH06b, NZIS13]. Hatching [PK22, KNBH12]. Hausdorff [TLK09]. HAvatar [ZWS^{+}24]. HDR [AFR^{+}07, ASC^{+}14, CWL22, DGH16, DTPG12, EKD^{+}17, GKTT13, LYO^{+}23, MKRH11, SKY^{+}12, TKTS11]. HDR-VDP-2 [MKRH11]. Head [BLC^{+}22, DWS^{+}23, MLL^{+}22, MPE^{+}23, ZZC^{+}23, ZWS^{+}24, FTZ^{+}19, FRS19, Iza18, KBBD17, LTO^{+}15, LCC21, RTD^{+}21, SED16, ZHS^{+}20]. head-mounted [FRS19, KBBD17, LTO^{+}15]. Headon [TZT^{+}18]. Heads [KQG^{+}23, LT06, YFFA21]. Headset [AKG^{+}23]. headshot [SPB^{+}14]. HEART [LKJC21]. Heat [SSC19b, CWW13b, HHP^{+}21, VBCG10]. Height [MLS^{+}18, PP94, Pag98, NSB13, YAV^{+}20]. Height-Field [MLS^{+}18, YAV^{+}20]. held [CWL12, IBP15, ZYQ^{+}14]. helices [BAC^{+}06]. Helmholtz [YCR^{+}15]. helper [MK16]. HelpingHand [LYFD12]. hemoglobin [TOS^{+}03]. here [CLC14]. Hermite [AA09, BI92, JLSW02, Pet89, SY21b]. Hessian [BLdG^{+}16, LLR^{+}15, SJJ12, WZX^{+}23]. Hessian-based [BLdG^{+}16, SJJ12]. Heterodyne [HMI23]. heterodyned [VRA^{+}07]. heterogeneous [BBO^{+}09, DWd^{+}08, HLW^{+}19, KHLN17, LMAS16, MPG^{+}16, PvBM^{+}06, STPP09, WZT^{+}08a, XWCH15, XMZ^{+}14]. heuristic [XGC07]. heuristic-based [XGC07]. Hex [PCS^{+}23b, FXBH16, GJTP17, GPW^{+}17, LLX^{+}12, LZS^{+}21, LSVT15]. hex-dominant [GJTP17]. Hex-Mesh [PCS^{+}23b, LSVT15]. hexagonal [PEVBC21]. Hexahedral [BC23, GDC15, SHG^{+}22, SRUL16, SRUL17, BBC22, LZC^{+}18, LBK16, PLC^{+}21]. Hexahedral-Dominant [SRUL16, SRUL17]. hexahedrizations [VPR19]. hexahedron [PVR18]. HexEx [LBK16]. Hexmeshing [LPC22]. Hidden [And82, IWC22, KL23, SO92, HZ82, KK87, McK87]. hidden-surface [McK87]. Hidden-Volume [KL23]. Hiding [FKN17, PH15b]. Hierarchical [AGL^{+}22, FB95, HNB^{+}06, KT03, KLM24, NH22, SCA02, TH19, WLF^{+}20, XSTN14, XLY^{+}22b, YHB05, ZCP^{+}23, ZXS^{+}21, dFP95, AW20, BCRK^{+}10, DF88, DDP99, JB02, LZT^{+}08, ODJ04, PBYV17, SPO10, Sze06, VdFG99, YWVW13, YGH^{+}17, vKXZ^{+}13]. hierarchies [BSW02, WBS07]. hierarchy [YY17]. HiGAN [GWLG23]. High [AAPS16, BMBRD24, BLC^{+}22, BGAM12, BBB^{+}10a, BHB^{+}11, BBN^{+}14, BV22, BHPS10, CKH18, CLS^{+}15, CJN^{+}17, CCS^{+}15, DGH16, FJA^{+}14, GHCC88, GBAM11, GLT^{+}23, HW15, HRH^{+}13, IRG^{+}23, KSA13, KUWS03, KKN^{+}22, KLM24, LEPM22, LWP^{+}23, MHZ^{+}21a, MEA^{+}18, MCHAM06, Mus13, OLSL16, RAI06, SMM14, STTP14, SHS^{+}04, SRX^{+}23, SJA08, SYS^{+}21, SSW^{+}23, SWS^{+}22, TWR^{+}23, TREO16, Tsa15, US24, Van82, WHB^{+}12, WSP^{+}23, WJV^{+}05, YSN^{+}18, YR23, YJLL22, ZZC^{+}23, ZRB14, ZWS^{+}24, ZZC^{+}22, ZCP^{+}23, ZKU^{+}04, AGL^{+}17, AGDL09, AAPS17, AYL^{+}12, BWDL21, BWG03, BTFN^{+}08, CS00, CBZB15, CADS09, CWZ^{+}21a, CCOST05, CTW09, CWSB22, DD02b, ESCK16, FLW02, GLD^{+}19, GO12, GT96, HSG^{+}16, HFH^{+}17, HBD^{+}14, HG09, HSHF10, HCTW11, JZH^{+}21, KSB^{+}13, KR17, KKSS18, KZP^{+}13, KLF^{+}19, LRT^{+}14, LHK^{+}20, LGA^{+}21, LGX^{+}13, LSA05, LCX^{+}21, LSH^{+}22, MRK^{+}13, MKMS04, MEMS06, MHP^{+}19, NKGR06, NB11]. high [SHX^{+}22, SWTC14, SFWG04, SXZ^{+}20, TAHL07, TAH^{+}04, THG99, Van06, VLD^{+}13, WAC07, WL21, WLHR11, WZC^{+}22, WSS18, XCLT14, YHJ^{+}14, ZSCS04, ZHRB13, ZJ11, ZWL^{+}18, ZJY^{+}21, ZSTB10, LCTS05]. High-accuracy [CKH18]. High-contrast [STTP14]. high-degree [CADS09]. High-Dimensional [MEA^{+}18, AGDL09, GO12, ZWL^{+}18]. high-DOF [SHX^{+}22]. high-dynamic-range [DD02b, ZJY^{+}21]. High-Fidelity [BLC^{+}22, IRG^{+}23, SSW^{+}23, OLSL16, YSN^{+}18, ZZC^{+}23, ZWS^{+}24, CBZB15, HCTW11, LGA^{+}21, SWTC14, WSS18, XCLT14]. high-frequency [SXZ^{+}20]. High-Level [Van82, CWSB22, HBD^{+}14, LRT^{+}14]. High-Order [BV22, LWP^{+}23, SMM14, SYS^{+}21, ZRB14, JZH^{+}21]. high-pass [CCOST05]. High-Performance [MHZ^{+}21a, SRX^{+}23, Tsa15, ZZC^{+}22, KKSS18, LHK^{+}20]. High-Precision [US24]. High-Quality [AAPS16, KKN^{+}22, ZCP^{+}23, BGAM12, BBB^{+}10a, BHB^{+}11, BBN^{+}14, CLS^{+}15, CJN^{+}17, CCS^{+}15, GBAM11, HRH^{+}13, SJA08, WHB^{+}12, ZKU^{+}04, AAPS17, BWG03, CS00, CWZ^{+}21a, LCX^{+}21, MHP^{+}19, WL21, ZJ11].

High-Resolution [FJA^{+}14, SWS^{+}22, TWR^{+}23, YJLL22, HW15, KLM24, Mus13, TREO16, AYL^{+}12, BWDL21, GLD^{+}19, HG09, YHJ^{+}14, ZHRB13, ZSTB10]. high-speed [TAH^{+}04]. high-volume [BTFN^{+}08]. Higher [BIW93, BSEH18, LLK^{+}20, BJ17, Csé19, MC21, PSH^{+}21]. higher-dimensional [BJ17, PSH^{+}21]. Higher-Order [BIW93, BSEH18, LLK^{+}20, MC21]. Highlight [GLT^{+}21, TDR^{+}12, RRMG10]. Highlight-aware [GLT^{+}21]. Highlighted [KHKR11]. highlighting [BDG15]. Highlights [PNTK23]. Highly [ATW13, ZB94, HRE^{+}08, IDN12, LYvdPG12, SJLP11]. hinting [Sha03]. histogram [BPC16, DMB^{+}14, KS10]. histories [SSTP15]. history [HXM^{+}13]. HLBVH [VKJ^{+}17]. HMDs [OLSL16]. Hodge [MMdGD11]. Hodge-optimized [MMdGD11]. HodgeNet [SS21]. Hogel [CTFH22]. Hogel-Free [CTFH22]. holes [BW13]. holey [BW13]. holodeck [WS99]. Holographic [CBYJ23, JBLL18, MGK17, OKH^{+}16, CTS^{+}20, CGP^{+}21, JMB^{+}20, KNL^{+}22, LJM^{+}16, SHL^{+}17, TDG18]. Holography [CTFH22, CBS^{+}22, KSL^{+}23, NKKJ23, CGP^{+}21, KNL^{+}22, PDSH17, PCPW20, RRMG10, WCSC22]. holonomy [BCW17, SZC^{+}22]. home [KDW^{+}17, KPB^{+}12, YYT^{+}11]. Homogeneous [Kan15, FAW19, HJ11b, KSSCO08, TWL^{+}05]. Homogenized [SNW20]. Homomorphic [LK02]. Hookean [SDK18]. HOT [MMdGD11]. HRBF [XNZ^{+}22]. HRBF-Fusion [XNZ^{+}22]. HSV [SCB87]. huge [BGB^{+}05, GM05]. Hull [Day90]. hulls [MPN^{+}02]. Human [AVB^{+}23, BLC^{+}22, CPY^{+}22, DXG^{+}23, DKD^{+}17a, GRG04, GWBN24, HL14, HXZ^{+}19, Hil86, KNK^{+}22, KQG^{+}23, KH17a, LJL23, LY23, LWL23a, LXZ^{+}19, MLL^{+}22, MPE^{+}23, SAA^{+}21, SYZ^{+}23, SLST14, SZL^{+}23, SYM^{+}24, TSLP14, WLZ^{+}21, XCZ^{+}18, YPL21, ZJY^{+}22, AHM^{+}15, ACP03, ACOYL08, CTMS03, CPMK21, CTTL15, CYT^{+}18, Dee05, DWd^{+}08, DK99, DKD^{+}17b, FKI^{+}14, FP03, GSCO12, HRZ^{+}13, HPP05, HKA^{+}18, JWDL19, JYQ^{+}22, KE18, KWK09, KCGF14, KPMP^{+}17, KLF^{+}19, KH17b, LCR^{+}02, LPLL19, LMM^{+}22, LHR^{+}21, LCX16, MJC^{+}03, MSS^{+}17, MCC09, MWTK13, NOP^{+}18, NZC^{+}18, PRWH^{+}18, PH06, PMRMB15, RPE^{+}05, RSH^{+}05a, SHP04, SZK15, SGX^{+}21, SKL07, SGdA^{+}10, SDO^{+}04, TZT^{+}18, TMB14, Van06, VPB^{+}18, WC10, WMC11, WMP^{+}06, WPL^{+}21, WL16, XWCH15, XPB^{+}21, XLS^{+}11, YKH04, YZX21, YIO^{+}15, YM16, ZZMC13, ZFL^{+}10, dSAP08]. human-assisted [YIO^{+}15]. Human-aware [SYM^{+}24]. Human-Centric [GWBN24, KCGF14]. Human-Computer [Hil86]. Human-Made [AVB^{+}23]. Human-Scene [LY23]. humanoid [NRH17]. humanoids [HRL15, LPKL14]. HumanRF [IRG^{+}23]. Humans [IRG^{+}23, KWS^{+}23, TWR^{+}23, EHA12, JTST10, KE18, MBB12]. Hybrid [BSSJ23, CSAP21, EC93, HWP^{+}23, HTCH15, Kla94, LHH^{+}23, MSQ^{+}18, NN95, OTS06, QLY^{+}23, Rap91, SQSL22, VR94, WLS^{+}23, YSC^{+}18, ZYM^{+}20, ZZL^{+}23, DBDB11, FOK05, LMM^{+}22, PVL^{+}05, PPW18, SWL11, WZK^{+}17, XDF^{+}19, ZXS^{+}21]. hybridization [FBT^{+}18]. hybrids [RHDG10]. hydrodynamics [WDK^{+}21, WAK20]. hydrographic [ZYZZ15]. hydrology [GGG^{+}13]. Hydrophilic [LWF^{+}22]. Hydrophobic [LWF^{+}22]. Hyper [BEH18, KCS14]. hyper-lapse [KCS14]. Hyper-reduced [BEH18]. Hyperbolic [AL16, IYAH17]. Hyperelastic [LBK17a, WB23, LBK17b]. Hyperelasticity [LLJ^{+}23]. Hyperion [BAC^{+}18]. hyperlapse [JKT^{+}15]. HyperNeRF [PSH^{+}21]. Hyperparameter [TYY^{+}19]. Hyperspectral [CBKM15, SS19, BKGK17, CJN^{+}17, JBY^{+}19, KRD^{+}12, LLWD14]. Hypersurfaces [MHS^{+}19a].

I-cloth [TWL^{+}18]. ICARUS [RYW^{+}22]. ICON [HZvK^{+}15, WSML23]. Icons [HH90, BL15, LRFN04]. IconShop [WSML23]. ICP [Rus19]. ICTree [PHBC21]. IDE [SWS^{+}22]. IDE- [SWS^{+}22]. ideal [WDW^{+}15]. identification [XSZ^{+}16, ZWZ^{+}16]. identifying [DFL^{+}15]. identity [NBLCO20, YCL^{+}20]. IGT [GP08]. II [Gol85a]. IK [Wam16]. iLamps [RvBB^{+}03]. ill [CSHD21]. ill-conditioned [CSHD21]. Illicit [Gol85b]. illuminant [BDM09, LEN09]. illuminating [YY17]. Illumination [BYRN17a, CRA11, CSS96, CLSS97, DPD22, GZS^{+}22, LTH^{+}23, PM95, RLU95, VADWG15, VMKK00, AFO05, AS02, BYRN17b, BAERD08, BBG^{+}13, CGZ08, CNR08, DSDD07, DKH^{+}10, Fat09b, FJL^{+}16, GSY^{+}17, GCP^{+}10, GFT^{+}11, GD04, JSKJ12, Kan15, KKN^{+}14, KFB10, LALD12, MHP^{+}19, MA06, NKGR06, Pel10, RWG^{+}13, RGK^{+}08, RLP^{+}20, SCW^{+}21, SL17, SFWG04, SKC^{+}14, TL04, TPWG02, TFG^{+}13, VAZH^{+}09, VKK18, WHSG97, WFA^{+}05, WWZ^{+}09, WS99, WGT^{+}05, XDPT16, YSJR17, ZSD^{+}21]. illumination-guided [FJL^{+}16]. illumination-invariant [CGZ08]. illuminators [RNd^{+}07]. illusion [STXJ15]. illusory [CLQW08]. Illustrating [MYY^{+}10]. Illustration [ZIH^{+}11, ASP07, ACCO05, GAGH14, KNBH12, KYYL08, KST08, LEQ^{+}07, ONOI04]. illustrations [GRG04, LRA^{+}07]. IM [Kim18]. IM-material [Kim18]. IM6D [HMT^{+}15]. Image [AASP17b, ARMCO23, BIP01, BAC^{+}23, BLR^{+}11, BBPA15, BNB13, CA24, CAA10, CLC96, CAV^{+}23, CLX^{+}22, CZL^{+}14, DSB^{+}12, DCD15, DBP^{+}15, DSJA^{+}21, DKT^{+}23, DCB^{+}22, FYW^{+}18, Fat07, FF11, GAA^{+}23, GGY18, GHCG17, GLT^{+}23, HM92, HKAK14, HXM^{+}18, HRS^{+}23, IVH^{+}23, JLF^{+}23, JKZS10, KRFB06, KLS^{+}13, LKG^{+}03a, LFDF07, LW15, LLN^{+}14, LT00, LCL^{+}17, LZH^{+}20, LXL^{+}23, LNLB16, LCD06, MPN^{+}02, MZWV07, PNTK23, PC82, QTZ^{+}06, RDL^{+}15, RO85, RO87, RJN16, SMW06, SKG^{+}12, SYJS05, SLWS07, TZW^{+}07, TCS^{+}23, TOS^{+}03, VKM^{+}23, VBK05, WSP^{+}23, XHWW22, XFT^{+}08, XFZ^{+}09, XK07, XLXJ11, YTS^{+}11, YPA^{+}18, YSQS07, ZK22, Zhu18a, ZCX^{+}22, vW02, AASP17a, ALY^{+}21, AS07, AMMS08, BGKS17, BSFG09, BC02, BSP^{+}19, BMSR20, BZCC10, BHY15, BKR17, BPB13, BA83, CHM^{+}12, CWW^{+}16, CSW^{+}16, CKS^{+}17, CDSHD13, CPD07, CTW09, CCT^{+}09, CZM^{+}10, CHY21, CGZ08, CSHD03, CSRP10, DMIF15, DAD^{+}18]. image [DCP14a, DZPZ09, DTPG11, DCOY03, EKD^{+}17, FH07, FHL^{+}09, FFL10, FAR07, Fat08, FCA09, FLB17, FRS19, GPM^{+}22, GSY^{+}17, GO11, GO17, GCB^{+}17, GRBN09, GMW16, GLT^{+}21, HSGL11, HSB^{+}12, HRDB16, HPP^{+}18, HBD^{+}14, HDD^{+}16, HWRH13, HST^{+}14, HDN^{+}16, HLR^{+}17, HMG03, HXM^{+}13, HZW^{+}13, HSW^{+}17, HYG^{+}13, HWK15, HOM15, ISSI16, ISSI17, JCW09a, JTC09, JYQ^{+}22, KEE13, KP02, KKDK12, KSP13, Kou16, KOC^{+}22, KSE^{+}03, LHM09, LWA^{+}12, LEPM22, LDF14, LSQ^{+}15, LGA^{+}18, LXR^{+}18, LK20, LQGY24, LYY^{+}17, LTJ18, LFB^{+}13, LSS^{+}17, LSC^{+}12, MHM^{+}09, MAS^{+}16, NFD07, PHL^{+}09, PHK11, PSN20, PGB03, PSA^{+}04, PTSZ11, PAAG21, PHS^{+}18, RKAP^{+}12, RFWB07, RPK^{+}12, RHDG10, RGSS10, SFLM04, SLJT08, SJA08, SLS^{+}16, SMGE11, SSY^{+}04, SDP^{+}18, SHM^{+}14, SLWF14, SSD09b, SBT^{+}19, SJMP10, TFX^{+}08, TYS09, TER^{+}20, TZN19, TVLF20, TS08, TAN^{+}21, TYY^{+}19, VRC^{+}13, VT04]. image [VBFG12, VBBF16, WWH06, WTSL08, WYW^{+}10, WYX11, WFP12, WHB^{+}12, WLL^{+}14, WSZ^{+}18, WYL^{+}20, WWA^{+}16, WSS^{+}19, WLHR11, WSTS08, Wym05, XLY09, Xia97, XKF^{+}18, XYH^{+}21, XSTN14, XYJ13, XSHR18, XWZ^{+}21, XADR12, YSN^{+}18, YSQS08, YJHS12, ZZXZ09, ZN06, ZCW^{+}17, ZZI^{+}17, ZCC^{+}12, ZLH^{+}21, ZAFW21]. Image-Based [BBPA15, BNB13, KRFB06, KLS^{+}13, LKG^{+}03a, LCL^{+}17, MPN^{+}02, MZWV07, QTZ^{+}06, SKG^{+}12, TZW^{+}07, TOS^{+}03, VBK05, XFT^{+}08, XFZ^{+}09, YTS^{+}11, BKR17, CWW^{+}16, CDSHD13, DCP14a, HRDB16, HPP^{+}18, HLR^{+}17, HMG03, LWA^{+}12, NFD07, SSY^{+}04, VRC^{+}13, VT04, VBFG12, VBBF16, WFP12, XSHR18, XWZ^{+}21, ZCW^{+}17]. Image-driven [LT00]. Image-guided [BLR^{+}11, XK07]. image-noise [CTW09]. Image-space [DCD15, RJN16, Wym05]. image/video [SLJT08]. Imagery [MRC^{+}86, MGDA^{+}15, HH10, KH10, KCSC10, NAB^{+}15, SSJ^{+}11]. Images [AZMW21, DRC^{+}15, KPACO22, LR90, LJZ^{+}23, LWL^{+}23b, PABE^{+}21, RMBCO23, SB95, SS19, SPG^{+}23, SCB88, TLG17a, WS17a, YNK^{+}22, ZLW^{+}16, AM10, BBS14b, Bou18, BPD09, CAA09, CWW^{+}13a, CWC11, CLQW08, CZG^{+}11, CHM^{+}10, DSB^{+}12, DER^{+}10, DTPG12, DD02b, FKY^{+}10, FFBB21, GLD^{+}19, GGH02, GSLM^{+}08, HHV^{+}21, HCS13, HDMR21, HCE03, HC04, HDC07, HZZ11, IKCM13, JMAK10, KE18, KH08, KSH10, KP18, KUDC07, LBP^{+}12, LSA05, LSQ^{+}15, LKK^{+}21, LYT^{+}14, LSS^{+}19, MCL^{+}09, MPK09, MNBN07, NFL12, ODAO15, OTS06, OBW^{+}08, ÖG15, PBS04, RSSF02, SDIN18, STZ^{+}16, STXJ15, SHZ^{+}20, TLG17b, TEG18, TD16, TAH^{+}04, THG99, TT09, WWOH08, WSH^{+}16, WAM02, WS17b, WYXJ21, XLX^{+}16, XBS^{+}19, ZCC^{+}12, ZFL^{+}10, ZTF^{+}18, LR91]. ImageSpirit [CZL^{+}14]. Imageworks [KCSG18]. Imaging [ABGL21, BWC^{+}23, BMBRD24, CFP^{+}21, DMZ^{+}17, GNHM15, GVNB18, HOZ^{+}19, HMI23, JYW^{+}23, KZSR16, LCD^{+}19, RSM^{+}23, ZWHB22, ABW^{+}17, BGK16, BKGK17, CHWH17, CSHH21, CWK^{+}20, Fre16, GKHH12, HSG^{+}16, HRH^{+}13, HHGH13, HHHW15, IGP^{+}17, ITM^{+}14, JBY^{+}19, KR17, Kan15, KRD^{+}12, KN06, Lee18, LCV^{+}04, LLWD14, LOW18, LWO19, MRK^{+}13, MHM^{+}17, NZV^{+}11, Par17, PKHK15, PH15b, RTF^{+}04, RRFG17, SHHW16, SDP^{+}18, SRL^{+}15, TAHL07, WZK^{+}17, WMB19, WJV^{+}05, WW13, XIAP^{+}17, ZJMB11]. Imagining [SMZ^{+}14]. iMapper [MGC^{+}19]. Imitation [GWLG23]. immersed [GAB20]. Immersion [LB18, HFI^{+}08]. immersive [GWN^{+}03, HCW15, LNWB03]. impact [KLF^{+}19, SKV^{+}12, SN17, VSK^{+}17, WSJP17]. impactful [KLF^{+}19]. Imperceptible [KOOP11, SMG^{+}20, LSL^{+}18, MWH^{+}09]. Imperfect [RGK^{+}08, SPGT18]. Implementation [Day90, Mai92, KW03]. Implemented [LS00]. Implications [AKG^{+}23]. Implicit [BIW93, BGI^{+}18, BRB^{+}19, CSAP21, DSDD07, IWC22, JWL^{+}24, JCFG23, KSNG17, LHH^{+}23, PGP^{+}19, Roc89, TZY^{+}23, Tau94, VBG^{+}13, WLJ^{+}22, WSP^{+}23, WZX^{+}23, YKZ^{+}22, ZEF^{+}22, ATW^{+}17, BMSR20, CKMR^{+}21, CH89, DBD16, DZCJ22, FLGJ19, GHB^{+}20, GMP09, GBC^{+}13, HPG^{+}22, HCJ19, HGMRT20, LT09, LDN^{+}18, LTT^{+}20, MASS15, NPLX22, OBS04, PICT15, PV06, SSGS11, SJ22b, SOS04, SS11, TO02, WG09, YYW12b]. Implicitization [Hob91]. Implicitizing [SG17]. Implicits [JK23, XNZ^{+}22, OBA^{+}03]. Importance [CSS96, FHG^{+}23, MMR^{+}19, SLGS01, WZC^{+}20, ARBJ03, CJAMJ05, GKH^{+}13, GYGS22, KVG^{+}19, LRR04, LKB^{+}22b, ODJ04, Pet21]. Importance-Based [SLGS01]. Important [ANL^{+}23]. imposed [Fat07]. impossible [WFY^{+}10]. impostor [DHOO05]. Improper [ACC90]. Improve [MGDA^{+}15, VMKK00]. Improved [LR90, LR91, MRK^{+}14, RSA08, WHG84, XAW^{+}23, CZJ12, SMG^{+}20, TDR^{+}12, WSJP17]. Improvements [DKHS14]. improves [CLS^{+}17]. Improving [DDD^{+}14, MPB17a, MPB17b, Per02, WLM^{+}15, XW09, ZF03, KNL^{+}22, SVB^{+}12, WZMM22, XADR12]. IMUs [KLF^{+}19]. In-Betweening [QZZ22, HYNP20]. in-the-wild [FFBB21]. In-Timestep [FSKP23]. in-volume [HJ11b]. inaccessible [YSL^{+}14]. Incident [HWZ^{+}20, MPDW03]. Include [RT90]. Including [PGG^{+}24]. Inclusion [WFS^{+}21]. Inclusion-based [WFS^{+}21]. incomplete [TZCO09, YHZ^{+}14]. Incompressible [BGI^{+}18, CPAB22, XTZ^{+}21, AIA^{+}12, LZF10, SHM22, SAL^{+}08, SP09, WHK17, XCW^{+}20, dGWH^{+}15]. inconsistent [DSB^{+}12, KOF13, OF12]. Incorporating [LNWB03]. Incremental [Fie85, LFS^{+}20, RY92, LYK^{+}21, LKJ21, MZ12, TWL^{+}18, WM03, WSS18]. Incrementor [Res87]. Independent [PBCF93, AMMS08, BHHM20, BBG21, EML^{+}18, LB05, NOP^{+}18, SXD^{+}12, YM16]. Index [Ano85a, Ano90b, Ano92a, Ano93, Ano94, Ano95, Ano96]. Indexing [ZWK14]. indirect [HPB06, LALD12, MWRD13, RGK^{+}08]. individualization [YI17]. Indoor [HZD^{+}23, LYO^{+}23, PMGD21, RFW^{+}23, SYM^{+}24, WLJ^{+}22, ZXTZ15, CLW^{+}14, CXY^{+}15, FCW^{+}17, GSY^{+}17, KMYG12, MLZ^{+}16, MDKD16, NXS12, PAAG21, SXZ^{+}12, WSCR18, WLW^{+}19, WXZ^{+}22, XZY^{+}17, XWZ^{+}21, ZCC16, LPX^{+}19]. induced [CSvRV18, FXBH16, KBC^{+}13, LSL^{+}18]. induction [YHL^{+}18]. Inelastic [QLY^{+}23]. inelasticity [LLJ22]. inertia [BWBSH14]. Inertial [WVY^{+}22, HKA^{+}18, JKZS10, YZX21]. Inexact [YLYW18]. inextensible [GHF^{+}07]. Inference [KSH^{+}14, MHP^{+}19, YSN^{+}18]. Inferring [GJB^{+}20, KF93, SCH^{+}14]. Infill [CZM^{+}23]. Infinite [WHK17, NRC21, NM16, SPW^{+}18, VSLD13]. inflatable [STK^{+}14]. inflatables [PIC^{+}21]. inflorescences [IOOI05, OCH^{+}16]. influence [DCB^{+}22, VLD07]. Information [Ano82, Ano83, Ano84, Ano86, Ano87, Ano88, Ano89, Ano90c, JYW^{+}23, Mac86, WK95, WF96, XZZ18, CLW^{+}14, TOS^{+}03, WW13]. Informative [HXZW20]. informed [CLZ^{+}22, ZD20]. InfoTypography [LN22]. Infrared [BBC^{+}23, JYW^{+}23]. InfraStructs [WW13]. Inhomogeneous [Ste20, KMOD09, YIC^{+}10]. Initialization [HWP^{+}23, HTYW22]. Injective [AL13, FW22, RPPSH17a, CW17, FLJK21, FLG15, LCH^{+}21, RPPSH17b, WZ14]. injectivity [DKZ^{+}21]. Ink [SKC^{+}14, AAMSB20, CT05, KNBH12, PCK^{+}19, PFX^{+}22, SBK^{+}18]. Ink-and-ray [SKC^{+}14]. inking [SSII18b]. Inkjet [NKS^{+}23]. inpainting [XXL^{+}21]. input [GMP^{+}16, IBP15, JPG^{+}14, NM16, RP09]. inputs [WFH10]. Insertion [Joe90a, MFR^{+}10, CAR^{+}09, JMD^{+}17]. inside-out [HRDB16]. inside-outside [JKSH13]. Insitu [PKM^{+}11]. Inspiration [VVCOS23]. Inspired [BW22, HL14, OCNG21, OGN^{+}23, YPA^{+}18, CYFW14, DZS08, IZE^{+}21, KGBS11, KS12, WTGT10, XZZ^{+}11]. inspiring [XZCOC12]. Instant [HK18a, JTPSH15, MESK22, NG18, PSNB13, WWSR03, FHL^{+}09]. Instantaneous [HMI23]. instantiating [WLW^{+}19]. instantiation [SSBD03]. instructions [APH^{+}03, SLR^{+}16]. instrument [UPSW16]. instruments [AR15]. Integer [BCE^{+}13, FBC18, GSC21a, Kla91b, Kla94, LSM23, McI83, PK83, AAMSB20, ASB22, BZK09, FV96, LFO^{+}22]. integer-constrained [LFO^{+}22]. Integer-grid [BCE^{+}13]. Integer-only [FBC18]. Integrable [DVPSH15]. Integral [BSSJ23, FD17, MGJ19, SM06]. Integrals [SBN15, XBLZ23, LZHJ20, NRH04, SR09, YLB^{+}22]. Integrated [BDI^{+}02]. Integrating [BXH^{+}18]. Integration [OF01, Özt16, WLF^{+}20, AKJ08, BJ05, DNZ^{+}17a, FGW^{+}21, HZ13, LLJ22, PSC^{+}15, SGH^{+}22, SK13]. Integrator [CSAP21, KSNG17, LGL^{+}19, MLT17]. Integrators [DLK18, BOFN18, KCD09, LTT^{+}20, MSW14, MCP^{+}09]. Intelligently [LNLB16]. intended [LRS18, YLL^{+}22]. Intensity [ABGL21, ME05]. Inter [LKL23, SAPH04, MCSK^{+}17, YSQS08]. Inter-Cloth [LKL23]. inter-scale [YSQS08]. Inter-surface [SAPH04, MCSK^{+}17]. interacting [LSSF06, MDB^{+}19, RBvB^{+}04, TTT^{+}17]. Interaction [Hil86, HZvK^{+}15, KP06, LKL23, LWF^{+}22, Ols86, PKH^{+}17a, SB93, SKSY08, YSW^{+}23, ZLC^{+}22, ZWK14, CB04, FKI^{+}14, GWB05, HGRT04, HLHR09, HMT^{+}15, MWH^{+}09, MGC^{+}19, PLR^{+}16, PKH^{+}17b, RLZ^{+}21, SCH^{+}16, SHX^{+}22, SY21b, TREO16]. interaction-aware [PLR^{+}16]. Interaction-Driven [YSW^{+}23]. interaction-guided [MGC^{+}19]. InteractionFusion [ZBYX19]. Interactions [LFL^{+}23, LY23, PM18, SGPT23, SYZ^{+}23, ZZT^{+}21, BDWR12, CWSO13, FMB^{+}17, FBGZ18, HMO12, HvKW^{+}16, KPH18, WMB^{+}20, WLO^{+}14, ZBYX19]. Interactive [ABL^{+}21, AD03, ADA^{+}04, AAPS16, AAPS17, AVB08, AMD02, ACSM12, AF02, BAS14, BIP01, BSG12, BBO91, BCC17, BST^{+}14, BR94, CRS^{+}16, CGC^{+}02, CLJ^{+}20, CKS^{+}17, CEW^{+}08, CAR^{+}09, CPAL22, CK11, DLC^{+}15, GWP^{+}19, GLY^{+}03, GKJ^{+}05, GDG^{+}17, HR13, HSTP11, HSvTP12, HLP^{+}22, IMF^{+}21, IDN12, KC19, KBD07, KW11, KN02, KSKL14, LWS^{+}18, LCR^{+}02, LLL18, LRA^{+}07, LFZ15, LZS^{+}21, LLKC21, LLHY22, LWW08, LFUS06, MTN^{+}15, MSL^{+}11, MM22, MCC09, NPP22, NGDA^{+}16, Ols88, PHT^{+}13, PKZ04, PJH^{+}17, RHW94, RRS13, RZL^{+}10, ROTS09, RTD^{+}10, Ros94, SCCB22, SM17a, SM17b, SGW06, SPF^{+}23, SXZ^{+}17, SWC^{+}18, SWL11, SLS^{+}07, SLF22, SSS^{+}08, SCGT15, SSJ^{+}11, SZC^{+}07, SZS^{+}08, SWS^{+}22, SDA^{+}23, TLK09, TK14, TBWP16, TDM11, TQ94, TPWG02, US24, VVC^{+}15, VABW09, WBC^{+}05, WSTS08, WS17a, WS17b, XMR^{+}11, XLCB15, XLX^{+}16, YMRD15, YCYW20, YKGA17a, YKGA17b]. Interactive [ZB13, ZCC^{+}12, dSAP08, AR15, BCT15, BWG03, BBPP10, BAERD08, BDI^{+}02, BGB^{+}05, CK14b, CZZ14, CRG^{+}20, CTW09, CÖS19, DSDD07, DKP11, DE05, DTPG11, DPF03, EVC^{+}15, FNvD82, GM05, HSH20, HZW^{+}13, HHN^{+}02, IIM12, IOOI05, JYL09, JP03, JF03, JX96, JMY^{+}07, JRT^{+}15, KTL^{+}04, KYC^{+}17, LWB^{+}10, LMLL21, LACS08, LTT^{+}20, MTP^{+}15, MWR12, MWRD13, MCS15, MI07, NAI^{+}18, NSZ^{+}10, NHAH03, OHB^{+}11, PMOR10, PPZ^{+}11, PCS23a, PTG02, PSK^{+}12, RKKS^{+}07, RMD04, RLP^{+}20, RKB04, SMM14, SXZ^{+}12, SHL^{+}17, SLD17, SSY^{+}04, SSII18b, SPGI13, TWL^{+}18, TBC^{+}16, UBW99, UKIG11, UKSI14, UPSW16, UB18, VGB^{+}14, WTL05, WAC07, WWZ^{+}09, WSZ^{+}14, WS99, WTBS07b, WDR11, WZL^{+}20, Wym05, YJL^{+}16, YMR^{+}13, YHZ^{+}14, ZG04, ZHR^{+}09, ZLE14, ZPKG02, vdHDT^{+}07, LCXS09]. Interactively [ESCK16, SRH^{+}15, YCP16]. interception [YLNP12]. Interchange [KP92]. Interchangeable [DYY16]. intercluster [Xia97]. interest [ZK13]. Interface [BL18, Fol86a, Fol86b, Fol86c, HC86, Hud94, RvE93, RO94, ZLC^{+}22, BJS^{+}08, DK99, FQL^{+}20, FH04b, GCR13, HK10a, IWZL09, KP09, KP10, MB12, NSACO05, Ols84, PTG02, Pel10, TBvdP04]. Interfaces [Bar86, BD86, Jac86, SG91, ZZC^{+}22, Ano03, LRFN04, SH08]. Interference [HPSZ11, RV89, KWB^{+}13, MHM^{+}17]. Interference-aware [HPSZ11]. interferometry [GLDZ15]. Interior [LLJ^{+}23, MSL^{+}11]. Interior-point [LLJ^{+}23]. Interiors [YSW^{+}23]. interleaved [JGGN15]. Interleaving [TWAD09]. Interlinked [GPB^{+}19]. Interlocking [CRCM23, OGN^{+}23, TCT23, CWSB22, FSY^{+}15, SCGT15, SFCO12, SFJ^{+}17, WSP18]. Interlocking-Free [CRCM23]. internal [MTB^{+}13, ONOI04]. Internet [CCT^{+}09, CZG^{+}11, HZZ11, MBGS15, STZ^{+}16]. interplay [CMT04]. interpolant [Jam20]. interpolants [BDT99]. interpolate [TO02]. interpolated [SH07]. Interpolating [DKA23, FG90, SOS04, Yuk20, LYLL08, RP09]. Interpolation [BI92, BIW93, BF01, CPAB22, CK20, DLG90, Fie85, Fol87, JW15, Pet89, RY92, SDN18, VTSSH15, WX91, BT19, BMSR20, BvdPPH11, BDM^{+}21, CWKBC13, CCW16, Csé19, FZL^{+}15, GTJS17, GAF^{+}10, LVKS21, MDZ^{+}21, MHM^{+}09, Mal89, MK05, PR97a, RSM10b, SV19, VW97, VBK05, WS21, WG10, YSW^{+}17, ZPBK17, ZKU^{+}04]. Interpolations [Thu17a, Thu17b]. Interpolatory [AA09, DM13, ZM11]. interpretation [CKX^{+}08]. Interpreting [SLZ^{+}13]. interreflections [CRA11, DDTP15, XCM^{+}14]. Intersecting [CCW93, CDY23, KS95, MD94, LB18]. Intersection [ACC90, CGM91, FLS^{+}21, KM97, MST89, Mil87, NY94, YJY23, CZXL23, LKL^{+}22, LFS^{+}20, NPP^{+}11, SHH99, VMT06, WFP12, Bak94]. intersection-and [LFS^{+}20]. Intersection-free [FLS^{+}21, LKL^{+}22]. Intersections [FNO89, MD94, SJ94]. interval [CZXL23]. intervals [ZS00]. interview [BLA12]. intra [YSQS08]. intra-scale [YSQS08]. intricate [TBBC^{+}22]. Intrinsic [BBS14b, CA24, CSBC^{+}17a, DRC^{+}15, LWQ^{+}08, LFXH17, LGC^{+}23, WP06, XWC^{+}16, YGL^{+}14, AGK^{+}22, BHY15, BST^{+}14, BPD09, CSBC^{+}17b, ED04, GSC21a, KLF11, LBP^{+}12, LZBCJ21, MZRT16, ROA^{+}13, SSC19a, Tak22, TBW^{+}12, XZT^{+}09, XZJ^{+}12]. Introduction [BG89b, BG89a, BG90, Ber82a, Ber82b, Fol86a, Fol86b, Fol86c, FGN84, FR87, Fuc82, Pha18, Ros94, Tan83]. Intuitive [BL18, LFL^{+}23, LC15, RZW^{+}21, BK04, GCR13, SGM^{+}16]. Invariant [NY94, SLL^{+}21a, ASGS23, BHR13, BBGO11, CGZ08, KPM^{+}17, LSC^{+}08, LSLCO05, MTP12, MWTK13, PR97a]. Invariants [LCK22]. invasive [NHAH03]. Inverse [BJNJ18, DSP06, DJBDDT13, GDAB^{+}17a, GZB^{+}13, GJB^{+}20, HMLB16, HXM^{+}13, HHD^{+}22, LJ14, LBAD^{+}06, LCBD^{+}18, LTH^{+}23, MDH^{+}23, NRN^{+}23, PMLB22, VGDA^{+}12, WWWZ23, WHZ^{+}08, WDR13, WYD^{+}14, ZB94, ASB22, BBP21, BWS10, CZXZ14, DJBDT10, DIO^{+}12, GLD^{+}19, GDAB^{+}17b, GP08, GITH14, GMHP04, KE18, LP10, LHP05, LCX16, MB21, NJJ21, NDMKJ22, PIC^{+}21, SZT^{+}07, SZGP05, WPP14]. Inverse-Foley [LJ14]. InverseCSG [DIP^{+}18]. Inversion [BAC^{+}23, HHL^{+}24, JK23, FL16, KDI19, LFS^{+}20, SLL^{+}21b]. inversion-consistent [SLL^{+}21b]. inversion-free [FL16, LFS^{+}20]. inversion-safety [KDI19]. Inverted [KH17a, KH17b, SKB^{+}21]. Invertible [AXR09, XLW18]. investigating [MBB12]. Investigation [BS90]. iOrthoPredictor [YSW^{+}20]. IPC [HCLK24, LYK^{+}21]. iPSR [HWW^{+}22]. IQ [FQL^{+}20]. IQ-MPM [FQL^{+}20]. iridal [POB09]. iridescence [BB17, WVJH17]. iridescences [Sun06]. irradiance [AFO05, SJJ12]. Irregular [YPL^{+}23, JLBM05, LZKW10, LCOZ^{+}11, ZXJ^{+}13]. irregularity [WLM^{+}15]. irregularly [Gos00]. Islamic [KS04a]. Islands [HA92]. iso [VGB^{+}14]. iso-surface [VGB^{+}14]. isocurve [EC96]. isocurve-based [EC96]. isolines [AFTCO07]. Isometric [Sah18]. isometries [JWI^{+}21]. isometry [TMRL14]. Isosurface [LS07, SMH^{+}23, VW94, VW95, WV92]. isosurfaces [LDS03, WHDS04]. Isothetic [PVY90]. Isotopic [MCSA15]. Isotropic [BSN16, CZB23, LCK22, MHS^{+}19a, SDK19, SBN15, TWAD09, WOR10]. Issue [BG89b, Fol86a, Fol86b, Fol86c, FGN84, Pha18, Ros94, Sto92]. iterated [RKB04]. Iteration [NH22]. Iterative [CK20, HL14, HWW^{+}22, LKE18, LZH^{+}20, YFFA21, DBDB11, DHC^{+}21, JTL^{+}12, JDD03, Wan15]. IV [AB89]. iWIRES [GSMCO09].

Jacobian [AGK^{+}22]. Jagged [Nai98]. JALI [ELFS16]. James [GIGM22]. jaw [ZBBB18, ZBGB19]. jewelry [ILB15]. Jigsaw [KP02]. jitter [TBV12]. jitter-free [TBV12]. Joinery [YKGA17a, CPMS14, YKGA17b]. Joint [DSAF^{+}13, GKH^{+}13, HKG11, JWD^{+}23, KPWG24, KCLU07, LBB22, LSQ^{+}15, TWLT19, WCSC22, XWY^{+}09, ZHPY21, GCPD16, HWK15, HOM15, ISSI16, JWDL19, KAL^{+}17, LKK^{+}18, VLV^{+}21, TBC^{+}16, YC21]. Joint-aware [XWY^{+}09]. joints [LT08, SZ15, WSP21]. Judder [CAD19]. Jump [BJNJ18, ZG04, YYW^{+}12a, AGB^{+}16]. JumpCut [FZL^{+}15]. jumping [YYVY21]. junctions [KPP17]. just [JSRV22, SC20]. just-in-time [JSRV22]. Juxtaform [PCS23a].

K-D [XLJ^{+}09]. K-Surfaces [DKA23]. kaleidoscope [HP03]. Kaleidoscopic [AGS21]. KD [AGDL09, ZHWG08]. KD-tree [ZHWG08]. KD-trees [AGDL09].

Kelvin [NRC21]. kelvinlets [DJ17, DJ18a]. Kernel [BVM^{+}17, CPW21, SBN15, WDT^{+}09, CLC^{+}20, Fat11, FKY08, GLA^{+}19, GLZ^{+}21, LSR18, LDF14, SJP05, VRM^{+}18, WWB^{+}14]. Kernel-predicting [BVM^{+}17]. kernel-splatting [GLA^{+}19]. kernels [ASL^{+}17, CMT^{+}16, FSH11b, VBCG10, YT13]. Key [MA07]. Keyframe [AHSS04, TMPS03]. Keyframe-based [AHSS04]. keyframing [JCW^{+}21]. Keying [AAPS16, AAPS17]. Keypoint [PNCB21]. Keypoint-driven [PNCB21]. Kinematics [HMLB16, MSK^{+}23, ZB94, BCT15, DSP06, GMHP04, SZT^{+}07, SZGP05, ZSZ^{+}14]. Kinetic [BL20, LD23, LWP^{+}23, LL23, LMLD22, XKCB18]. Kirchhoff [BJ05, HB23, KTY09, POT17, WB23]. Kirchhoff-plateau [POT17]. kirigami [JRPW20]. KleinPAT [WJ19]. Knit [JMZ^{+}22, KWL^{+}21, JGT17]. KnitKit [NQC^{+}21]. Knittable [WSY19]. knitted [KJM08, YKJM12]. Knitting [LNI^{+}23, LHZ^{+}21, NAH^{+}18, MAN^{+}16, NQC^{+}21, NWYM19]. Knot [Joe90a, SYSP14, Joe89]. Knots [VRP^{+}23, LIY^{+}22]. Knowledge [JWD^{+}23, XGC07, MYWI15]. Knowledge-Driven [JWD^{+}23]. Kontrol [Ols86]. KRISM [SS19]. Krylov [SS19].

L [GJB^{+}20, LLB24]. L-Systems [GJB^{+}20, LLB24]. LAB [SCB87]. Laban [DKD^{+}17a, DKD^{+}17b]. Label [CMS95, LSA^{+}16, RMBB^{+}13, WZF^{+}18]. label-map [LSA^{+}16]. Labeling [GZC15, ST16, VVC^{+}15, HSG13, HFL14, KHS10, YGH^{+}17]. labelled [HZCJ17]. labels [HLW^{+}18]. laboratory [ZJ18]. laden [GPH^{+}18]. Lagrangian [BGOS06, BvdPPH11, CCL^{+}22, CWSO13, DWK^{+}22, FLLP13, HGMRT20, KDW^{+}17, KAGS20, PTG12, QLY^{+}23, SBRBO20, SSJ^{+}20, WPLS18, XLYJ23, YCZ11].

Lagrangian/Eulerian [CCL^{+}22, QLY^{+}23]. Laminar [SOG^{+}22]. lamps [RBvB^{+}04]. Lampshades [ZLW^{+}16]. landing [ATM^{+}17, HYL12]. landmark [YNS19]. landscape [BLDA11]. Landscapes [PKH^{+}17a, CGG^{+}17, ENCC^{+}21, PKH^{+}17b]. Language [DGVG^{+}23, DMZ^{+}17, Jac86, KKRK^{+}16, MPF^{+}18, SFC^{+}23, Van82, WZHL23, ALLD17, GS82, HFF18, LTK09, MGAK03]. Language-based [WZHL23]. Language-driven [MPF^{+}18]. Languages [BK16, YPB16]. Laplace [NH22]. Laplacian [APH^{+}14, CSK18, DLF12, JCW09a, KFS13, LSR18, PHK11, ZHS^{+}05]. LaplacianFusion [KNK^{+}22]. Laplacians [AW11, FW12]. Lapped [TOII08]. lapse [BM07, HAK^{+}22, KCS14, LEN09, MBGS15, SMPR07, TDSG15]. Large [DTPC23, GNS^{+}12, KABL15, LWF^{+}23, LZCX19, MHZ^{+}21a, NI22, NDD^{+}23, NJJ21, SM17a, SHG^{+}22, SPF^{+}23, SJLP11, WFS^{+}21, WXZ^{+}23, XLC^{+}23, ZHS^{+}05, BZ11, BWHT07, BZL^{+}15, CB04, DFZ^{+}17, EDF^{+}16, FAW19, GB13, HSG13, HWG14, HHM19, HGMRT20, IGLF06, JP03, KGG^{+}20, KH08, KFWM17, KLM^{+}13, KSKL14, KPZK17, KG04, LGL^{+}19, LFS^{+}20, LCX^{+}21, MRA^{+}13, OAH11, PRFS18, PGH^{+}22, RNGF03, SMM14, SM17b, SWL11, SHM22, SDW^{+}16, SZLG10, WFDH18, WJV^{+}05, YMR^{+}13]. Large-deformation [DTPC23, BZ11, LFS^{+}20]. Large-Scale [LWF^{+}23, LZCX19, MHZ^{+}21a, SHG^{+}22, WXZ^{+}23, XLC^{+}23, GNS^{+}12, KABL15, NDD^{+}23, SPF^{+}23, SJLP11, WFS^{+}21, DFZ^{+}17, FAW19, GB13, HHM19, JP03, KGG^{+}20, KFWM17, KSKL14, KPZK17, LCX^{+}21, PRFS18, PGH^{+}22, SWL11, SHM22, WFDH18]. large-step [LGL^{+}19]. Larrabee [SCS^{+}08]. Laser [OKH^{+}16, XGC07]. laser-scanned [XGC07]. last [LSZ^{+}14]. latency [DCB^{+}22, FSP^{+}22]. Latent [AFL23, LLB24, RMBCO23, SLF22, YSC^{+}23, ZCP^{+}23, NBLCO20, WYXJ21]. Latent-based [RMBCO23]. Latents [AZL23]. lateral [SMG^{+}20]. lattice [ANHD17, FHM^{+}21, PMS12, RJ07]. lattice-guided [FHM^{+}21]. Lattices [Ros20]. Laughing [DZS08]. laughter [DZS08]. layer [DKT^{+}23, IM10, LHKR10, LWH^{+}11, LD13, PLW^{+}07, SBK^{+}18, XPB^{+}21, ZJ18]. LayerCode [MLYZ19]. Layered [DYP03, GLH^{+}23, JYW^{+}23, KOWD21, LKL23, LCD^{+}20b, RCOL09, VMCS15, WJHY23, WLHR11, ZMSS18, ZXJ^{+}13, ZZC^{+}22, BNK10, BBP21, Bel18, BRB^{+}19, BDW13, DS15, DJ05, DWd^{+}08, FLB17, GHP^{+}08, GHZ18, JdJM14, LVKS21, RCL21, ZLY^{+}21, ZGH^{+}16, ZKU^{+}04]. layering [MP09a, SZZK21]. Layers [TLG17a, HLR^{+}14, PTSG09, Pik83, SMH^{+}11, TDSG15, TLG17b, ZLB16a]. Layout [LQGY24, ULP^{+}15, XLY^{+}22b, AVB08, BSW13, CCL12, FYY^{+}16, JLS^{+}03, MSL^{+}11, PAAG21, YWVW13]. Layout-aware [LQGY24]. layouts [BYMW13, CBK12, CK14b, FMLW14, KS21, MSK10, PYW14, RRS13, WYD^{+}14, YLPM05, ZQCL19]. Lazy [LSTS04, XFAT12]. LazyFluids [JFA^{+}15]. LCD [HLHR09]. LCDs [LWH^{+}11]. LDR [AFR^{+}07]. Ldr2Hdr [RTS^{+}07]. leaf [RFL^{+}05]. Learned [BMBRD24, CTS^{+}20, CWK^{+}20, GLH^{+}23, HKPP20, JHR22, LABS23, QLH^{+}22, SZD^{+}20, ZK22, CCWL18, DCB^{+}22, KLV20, VKJ19, XKF^{+}18, ZZI^{+}17]. Learning [AWL^{+}19, AMA^{+}19, BLDL21, BB15, BZH^{+}23, CK14a, CXW^{+}23a, CRB23, CPV^{+}23, CXW^{+}23b, CTS^{+}21, CYT^{+}18, FKI^{+}14, FFBB21, FHXW22, FR22, FTP03, GSY^{+}17, GTB15, GJWW15, HFW^{+}19, HvKW^{+}16, HLV^{+}17c, HKC^{+}18, HPP^{+}22, HZD^{+}23, HWZ^{+}20, KWR16, KHS10, KNBH12, Kal18, KLM^{+}13, LBB22, LP10, LLLL21, LMLL21, LLL22, LBB^{+}17b, LXR^{+}18, LKK^{+}21, LAH^{+}21, LY23, LZCX19, LHP05, LVY16, LH17a, LH17b, LH18, LLH^{+}22, LZH^{+}17, NDD^{+}23, PZWW23, PABE^{+}21, RSH^{+}05a, RYPZ23, Rit18, RCPO21, SACO22, SHX^{+}22, SLF22, SSISI16, SWR^{+}21, SYM^{+}24, TGLT14, UB18, VVC^{+}15, WZF^{+}18, WCPM18, WSL^{+}19, WWLC21, WZ22, XDF^{+}19, XSZK23, XGZ^{+}23, YYL22, YML^{+}23, YPL^{+}23, YSCL22, YNK^{+}22, YGH^{+}17, YTL18, ZYL^{+}20, ZZC^{+}23, ZYM^{+}23, ZTD^{+}23, ZPYX23, AGK^{+}22, BDI^{+}02, CHY21, CGP^{+}21, CHP07, FYW^{+}18, FBH21, GCB^{+}17, HGY17, HRV^{+}18, HSK16, HXC^{+}20, HKA^{+}18, ISSI16, JWW^{+}20, JLWM22, JBX^{+}20, JHC^{+}21, KBS15, KAL^{+}17, LLL18]. learning [LWL17, LGA^{+}18, LLGRK20, LLM21, LSS^{+}19, LKS15, LSCC20, MTP^{+}18, NZC^{+}18, PZM13, PEL^{+}21, PBvdP15, PBvdP16, PBYV17, PALvdP18, PKM^{+}18, PFX^{+}22, RCCO22, SBHH16, SCH^{+}16, SAN23, SHZ^{+}20, SS21, SLL^{+}21b, SZKZ20, SMK22, SWL^{+}22, SSKS17, TKY^{+}17, VKS^{+}14, WZK^{+}17, WLY20, WL21, WLT22, WNEH22, WPL18, WSS18, XZZ^{+}14, YC21, ZTF^{+}18]. Learning-Based [FHXW22, LZCX19, KWR16, NDD^{+}23, JLWM22, WZK^{+}17]. Least [BIW93, DMZ^{+}17, LPRM02, LZH^{+}20, MHZ^{+}21a, CLC^{+}20, FCOS05, HFG^{+}18, SMW06, WJL^{+}20]. Least-Squares [BIW93, MHZ^{+}21a, FCOS05]. leaves [WWD^{+}05]. lecture [SBLD15]. legacy [KHFH11, RTS^{+}07]. Legged [GFK^{+}23]. Legible [ZCR^{+}16]. LEGO [LYH^{+}15]. Legolization [LYH^{+}15]. LEGO(R) [ZGXF23]. legs [GPD^{+}18]. length [HRvdP04]. Lens [ANL^{+}23, CBYJ23, PC82, HESL11, LES10, PHN^{+}12, SWF^{+}21]. lenses [GRBN09, HRH^{+}13, RAWV08]. Lensing [DGH16]. Lenslet [LR15]. lenticular [THKM13]. leopard [DFW20]. less [FGW^{+}21, RRC^{+}16]. Let [ISSI16]. letter [LN22]. Level [Aca07, CH14, ECBK14, HYS23, HCW^{+}23, LNZ^{+}23, MMHP23, MBWB02, Van82, YCL^{+}15, BHY15, CKIW15, CWSB22, CLMMO14, DE05, FKY^{+}10, FPBCO20, HFTF15, HBD^{+}14, HNB^{+}06, KJM08, Kim10, KCSC10, LRT^{+}14, LWS^{+}18, Lee18, LWS02, MLR^{+}22, MASS15, NZWC20, NNSM07, OBA^{+}03, RSH05b, SNW20, SSBL^{+}22, SLWF14, YKJM12, ZZW^{+}22b]. Level-of-Detail [HYS23, ECBK14, FKY^{+}10, HFTF15, MLR^{+}22, ZZW^{+}22b]. Level-Set [HCW^{+}23, LNZ^{+}23, NZWC20]. Level-set-based [YCL^{+}15]. Levels [WZK^{+}23, KWK09, Wan21]. Leveraging [HCTW11, JCFG23, RMP^{+}23, ZSZ^{+}14, LSL^{+}18]. levitated [OHR14]. Levitation [PFP^{+}22]. Lexical [SG91]. library [HZW^{+}13]. library-driven [HZW^{+}13]. LiDAR [LGZ^{+}13]. lie [BT19, Duf17b, Duf17a, KCD09]. life [AECOKC17, TMB14]. Lifetime [LCD^{+}19]. Lifted [APL14]. Lifting [GHL^{+}20]. Light [BRSMD22, BBS14a, BSB16, BJNJ18, CBCG02, CNR08, DPW15, DKHS14, DII23, DJ05, FBS^{+}23, GKDS12, GZS^{+}22, GWBN24, HSHF10, HMP^{+}08, JLF^{+}23, Kla87, LNA^{+}06, LLR^{+}15, LR15, MJC^{+}03, MMT18, MHU19, NID20, OF01, PRM14, RLLG^{+}20, SHD^{+}14, SXZ^{+}20, VMCS15, VPB^{+}09a, WZK^{+}17, YNK^{+}22, YSHWSH16, ZFT^{+}21, ZHM^{+}23, AGS21, BH21, BHR13, BMSR20, BDM09, BSB17, BJ17, CDP^{+}14, DHS^{+}05, EHDR11, FAR07, Fat09b, GTHD03, GLDZ15, GGHS03, HPJ12, HKD14, Hac18, HPB07, HKWB09, HSG^{+}16, HDHN16, HDC07, HLHR09, HWR14, HWBR14, HCW15, IZT^{+}07, JBM^{+}17, JMB^{+}14, JMY^{+}07, KWR16, KHD14, KHKR11, KHH^{+}11, KZP^{+}13, KBC^{+}13, KO11, KGH^{+}14, LHKR10, LWH^{+}11, LL13, LJM^{+}16, Leh07, LZT^{+}08, LAC^{+}11, LALD12, LKL^{+}13, LK20, LLW^{+}08, MSRB07, MLR^{+}14, MRK^{+}13, MRK^{+}14, MWBR13, MPDW03, MWHL21, MSOC^{+}19, MGJ19, MCT15, OK10, ORK12]. light [OHX^{+}14, OHHD18, OEE^{+}18, POB09, Pan17, PML^{+}09, QSH^{+}15, RHJD18, SNM^{+}13, SHL^{+}17, SLS^{+}16, SSY^{+}04, SOHK16, SY21a, SY21b, SHK^{+}17, TAV^{+}10, VRA^{+}07, VWJ^{+}13, VSJ21, VKS^{+}14, VK16, WDT^{+}09, WHY20, WLM^{+}15, WLHR11, WLHR12, ZSGJ21, ZWGS02, ZBW^{+}20]. light- [BMSR20]. light-driven [BDM09]. light-field [MRK^{+}13]. light-matter [SY21b]. lightcuts [WABG06, WKB12, WFA^{+}05]. Lighting [HZW12, JCFG23, LYO^{+}23, NBB04, PBMF07, RFW^{+}23, SW14, SWZ96, SHS^{+}18, SS00, YY17, ZSSJL20, BAOR06, BBPD12, BPB13, CPWAP08, DWT^{+}02, DCP^{+}14b, GGN18, GCD^{+}20, KP09, KAMJ05, LK02, LYL^{+}16, MWRD13, NRH03, NJS^{+}11, RKKS^{+}07, RMB07, RNd^{+}07, RZL^{+}10, SHS^{+}17, SKS02, VWB^{+}12, WSM11, XMR^{+}11]. Lightness [CKT^{+}23]. Lights [OKH^{+}16, DKH^{+}10, HKWB09, HWJ^{+}15, KWN^{+}17, NNDJ12, OP11, Pet21, WHY^{+}13, WR18, WWLC21]. LightSlice [OP11]. lightspeed [RKKS^{+}07]. Lightweight [BBGB16, HLP^{+}22, UMK17, VWB^{+}12]. Like [ZSAF21, DSG^{+}12, HZZ11, KLY^{+}14, NI24, MGAK03]. Lillicon [BL15]. limbs [MWTK13]. limit [TSL^{+}16, OCNG21]. Limited [DBP^{+}15]. limiting [WOR10]. Limits [BAU15, WP06]. Linde [DSZ17]. Line [And82, BS19, BKR^{+}05, KYYL08, LMLH07, LB84, RWW90, RSM^{+}23, SZLG10, SZG^{+}13, VA88, XWD^{+}22, BGAM12, CSHH21, CWK^{+}20, CSD^{+}09, FLB16, FZLM11, GTDS10, GCR13, GRT13, HOZ^{+}19, IH20, JDA07, KNS^{+}09, KLKL13, KSSI17, LWO19, MSSG^{+}21, NHS^{+}13, PSBM07, PNA^{+}21, PNCB21, Spr82, VKS^{+}14, Wes21]. Line-art [KYYL08]. line-drawing [Spr82]. Linear [Ale02, BSB16, DPW15, DMZ^{+}17, DLTW90, DHI^{+}13, Fie85, GTHD03, HGM14, KW03, LS00, LSLCO05, MHZ^{+}21a, Mey91, NON85, OF01, RY92, SZW^{+}23, WJBK15, WS85, dSDP09, BBO91, BBO^{+}09, BSB17, CDP^{+}14, CS09, DCP14a, DZCJ22, DKT^{+}23, FLB17, HSB^{+}12, HDA17, HKG11, LKJC21, LMR^{+}15, MMG06, MIGYM15, MHR^{+}16, NRH03, PLR^{+}16, SdS02, TDM11, VSJ21, WHSG97, WB08]. linearization [KJM10]. linearly [HDHN16]. Lines [Bak94, CH14, Fat14, MST89, YZX^{+}18, CGL^{+}08, FTP03, GKK^{+}21, KK87, LLW17, OBS04]. LineUp [YYL^{+}19]. Linkage [CSL^{+}22, BCT15, TCG^{+}14]. linkage-based [TCG^{+}14]. LinkEdit [BCT15]. linking [QJ21]. lip [ELFS16, SSKS17]. lips [GZW^{+}16]. Liquid [BHW13, Ste20, TB22, Thu17a, ATW13, ATW15, AB20, BDWR12, CWSO13, FMB^{+}17, FBGZ18, KTT13, MBT^{+}15, NB11, PHT^{+}13, Thu17b, UHT17, WLZ^{+}09, YCYW20]. Liquid-Crystals [Ste20]. liquid-fabric [FBGZ18]. liquid-hair [FMB^{+}17]. liquids [AGL^{+}17, CMSA20, CWSO13, CPPK07, DHB^{+}16, GB13, HGMRT20, KySK09, LBB17a, LSSF06, MYH^{+}10, NNC^{+}20, RWTT14, SXH^{+}21]. List [TOP03]. Listen [ANBH23, EML^{+}18]. listeners [CRG^{+}20]. lists [CSN^{+}12]. Live [DXG^{+}23, LCC21, MZRT16, DWT^{+}02, KDMW17]. live-action [DWT^{+}02]. live-streaming [KDMW17]. LiveCap [HXZ^{+}19]. lization [MPK09]. Lloyd [BSD09]. lobes [LPC^{+}11]. Local [ABGL21, APH^{+}14, BR21a, BB83, BBS14a, CCGB22, GSV^{+}14, HKC^{+}18, JHR22, Kal14, Les20, MP09a, MSOC^{+}19, MCY14, PHK11, Pet89, SLS05, SZKZ20, WGY^{+}18, ZDL^{+}14, ASC^{+}14, CDSHD13, CH89, Coh87, DKH^{+}10, DMIF15, FF11, FLSG14, GGY18, HZ13, ISSI16, KS10, KAMJ05, LFUS06, MHR^{+}16, RKZ12, SCF^{+}04, SL17, SSD09b, TMRL14, TNWK22, VMGM15, WHSG97, WSH^{+}18, WRK^{+}10, WBGB16, YSW^{+}17, ZSW^{+}10]. locality [SNB07]. Localization [YZH^{+}23]. Localized [HDA17, HC23, WLZ^{+}21, BWSS09, NVW^{+}13, PHT^{+}13]. Locally [BSB16, HYS23, LB23, Pot91, RPPSH17a, SW18, Sze06, TiABI07, WZ14, ZPW^{+}23, AVR^{+}22, BSB17, CW17, FLG15, ISSI17, LCH^{+}21, MSRB07, RPPSH17b, YYW^{+}12a]. Locally-Adaptive [HYS23]. located [KKB^{+}11]. Locating [HLV^{+}17a, HLV^{+}17b]. location [EKA84, UMK17]. Locking [FAER21]. Locking-Proof [FAER21]. Locomotion [CKJ^{+}11, FSRS22, KL17a, LPKL14, AvdP16, CLS03, GvdPvdS13, KL17b, KLV20, LWB^{+}10, LLK^{+}15, LLKP11, LSCC20, MdLH10, PBvdP16, PBYV17, TTL12, WP09a, WPP14, WHDK12, cWP10, YLvdP07, YTL18, dSAP08, dLMH10]. LOD [VLA15, WWH04]. Logarithmic [LGQ^{+}08]. Lohmann [QCOS23]. Loki [LSD^{+}22]. Long [SCCB22, TKG^{+}23, AAC^{+}06]. Look [CLC14, ZK22, BPD06, DSG^{+}12, Lau18, WKHA18, ZMN^{+}19]. Look-Ahead [ZK22, ZMN^{+}19]. Looking [EML^{+}18, Fol91, RPC^{+}10]. LookinGood [MBPY^{+}18]. LookOut [SCCB22]. Loop [Lev23, CTS^{+}20, HGG^{+}11, PCPW20, PFX^{+}22]. looping [LJH13b]. Loops [HLSH18, CBK12, DLSCS08, HHV^{+}21, LFH15, She13]. loopy [QJ21]. Lorenz [FCJ07]. Lorenz-Mie [FCJ07]. Loss [ZK22, VRM^{+}18]. Lossless [YGM97, GD02, PK05]. loud [DZS08]. Love [WB23]. Low [APC^{+}16, CSHH21, CPW^{+}23, HHGH13, JLF^{+}23, Lee18, MCE^{+}17, MMHP23, ME05, WS17a, APL14, CH05, CLW^{+}14, FSP^{+}22, GKH^{+}13, HSG^{+}16, KLS03, KO11, LHKR10, LWH^{+}12, MSRB07, MdLH10, MK16, PU06, SHP04, SKS02, WZMM22, WS17b, YTL18]. Low-budget [HHGH13]. Low-complexity [ME05]. Low-cost [CSHH21, MCE^{+}17]. low-dimensional [CH05, LWH^{+}12, MdLH10, SHP04]. Low-discrepancy [APC^{+}16]. Low-Dynamic [WS17a, WS17b]. low-energy [YTL18]. low-frequency [SKS02]. low-latency [FSP^{+}22].

Low-Level [MMHP23, Lee18]. Low-Light [JLF^{+}23, HSG^{+}16, KO11]. low-order [GKH^{+}13]. Low-Poly [CPW^{+}23]. low-quality [CLW^{+}14]. low-rank [LHKR10, MK16]. Lower [KM97, SJ94, MWTK13]. Lower-Dimensional [KM97]. Lpics [PVL^{+}05]. LR [GLLR11]. LS [LNZ^{+}23]. LuisaRender [ZZC^{+}22]. Luma [Nah20]. Luminaires [VADWG15, ZBX^{+}21]. Luminance [CAD19, MC92, TAKW^{+}19, DRE^{+}12, KWK09, MKRH11, MAC22, SCT^{+}15, WZMM22]. Luminance-aware [CAD19]. Luminance-contrast-aware [TAKW^{+}19, DRE^{+}12].

M. [OCNG21, OGN^{+}23]. Machinability [CCW93]. Machine [LNI^{+}23, NAH^{+}18, PMKB23, KBS15, KWL^{+}21, MAN^{+}16, NQC^{+}21, NWYM19, SARW^{+}15]. machine-knit [KWL^{+}21]. Machine-Made [PMKB23]. Machines [CCW93]. machining [BBR^{+}21]. macro [JCG^{+}21]. macros [BLDA11]. Made [AVB^{+}23, Pet95, PMKB23, FCODS08, LMS13, MZL^{+}09, MMBM15, SFG^{+}13, SSJ^{+}11, TSG^{+}14]. Magic [CXY^{+}15, PHN^{+}12]. magnetic [HMT^{+}15, KPH18, NZWC20, PLMR17, WMB19]. Magnetization [KPH18]. magnetized [SNZ^{+}21]. magnetoelastic [CNZ^{+}22]. Magnets [TGPS08]. Magnification [CM21, LTF^{+}05, WDW^{+}15, WRS^{+}12, ZTF^{+}18]. maintain [HK12]. Maintained [vOV96]. Maintaining [RO94]. majorization [SPSH^{+}17]. Make [PWLSH13, WCZ^{+}22, YYT^{+}11]. MakeltTalk [ZHS^{+}20]. makes [DSG^{+}12]. Making [MS04, XLF^{+}11, PDF^{+}22, BSW02]. man [FCODS08, LMS13, MZL^{+}09, MMBM15]. man-made [FCODS08, LMS13, MZL^{+}09, MMBM15]. management [BPD06, LDS02, Ols84]. Manga [QWH06, CCL12, CLC14, LLW17, QPWH08, XXL^{+}21, XLLW20]. Manifold [CZZT12, DS92, DWT^{+}10, FHG^{+}23, JM12, LXY^{+}16, CK14a, CHY21, LD21, MASS15, RRS19, YZ04]. manifold-based [YZ04]. Manifolds [NRS15, WLY^{+}16, CBK12, GO12, HP04, LVS^{+}16, Man86, ÖAG10, SMK22, WTL^{+}06a]. ManipNet [ZYSK21]. manipulate [ZYL^{+}20]. Manipulating [KAEE20, Res87]. Manipulation [AASP17b, HHL^{+}24, HRS^{+}23, Jac86, KOF14, vOV96, AASP17a, BSL12, BSP^{+}19, BLDA11, CAA10, CWW^{+}12, CWW^{+}13a, DCD15, FFLS08, FSGF16, GSMCO09, GAL^{+}09, GS82, GS85, IH03, IMH05, IM10, KOF13, KSES14, KLF12, KSKL14, KS21, LYP^{+}18, LLZM10, LLHF21, LCORL07, LLH04, Liu09, OF12, SNM^{+}13, SILN11, SMG^{+}20, SSP07, TAN^{+}21, VBBF16, WMZ^{+}13, XWY^{+}09, YKH04, YZX^{+}04, YJHS12, ZYSK21, ZCC^{+}12, ZHX^{+}07]. manipulations [BLDA11, KDM^{+}16, YL12]. Manual [PK22]. Manufacturing [ZZL^{+}23, AHB18, BR21b, MAYZ^{+}20, MDL16, MSDL17, MHSL18, YIO^{+}15, ZZX^{+}18]. Manuka [FHL^{+}18]. Many [SRX^{+}23, TJ07, GJ22, HPB07, HKWB09, HWJ^{+}15, JLF^{+}09, LPKL14, OP11, SCS^{+}08, WHY^{+}13, WWLC21]. many-core [SCS^{+}08]. many-light [HPB07, HKWB09]. many-lights [HWJ^{+}15, OP11, WHY^{+}13]. many-muscle [LPKL14]. Many-World [SRX^{+}23]. Many-worlds [TJ07, GJ22]. manycore [KGB^{+}09]. Map [LNZ^{+}23, ROA^{+}13, ASP07, HSRG07, HWG14, JJJ^{+}21, LSA^{+}16, NFA^{+}15, RH02, ZG04, ZK14]. Map-based [ROA^{+}13, ZG04]. Mapped [KH17a, KH17b, WZYR19, YHJ^{+}14]. Mapping [GFL^{+}22, HC23, Lip18, NCB23, SW18, SCB88, SWK16, TB87, WC21a, YZH^{+}23, ASC^{+}14, BKR17, CS00, CBCG02, DHI^{+}13, EMU15, EKM17, GP09, HOJ08, HJ09, HSST10, KD13a, KISS15, KJDL09, KO11, KZ11, LHW^{+}10, LCTS05, LW16, LLZ^{+}20, Lip12, MCSK^{+}17, MDK08, MAF^{+}09, MWI18, MGC^{+}19, MM06, NL13, NBLCO20, PSNB13, POC05, PTH^{+}17, QZG^{+}19, RTS^{+}07, SAPH04, SHHD17, SdS02, SCT^{+}15, SCA02, SXD^{+}12, TT09, WWT^{+}03, YZWH12, ZMT05]. Mappings [BJNJ18, DFYL19, GKK^{+}24, RPPSH17a, AL13, APL14, APL15, AGK^{+}22, CW15, DFZ^{+}17, FLG15, FL16, KSS06, KABL15, PL14, RPPSH17b]. Maps [ASGS23, DYZ^{+}23, ESBC19, HC23, HHD^{+}22, HJS^{+}14, RLU95, Ros20, Shn92, THCM04, ARBJ03, BCWG09, BCE^{+}13, CSZ16, CZ17, CKPS17, DK09, FFL10, Fat09b, FG11, GKK^{+}21, GASP08, HSB^{+}12, HZG^{+}12, HLW^{+}19, JSP17, KLF11, KAB^{+}10, KSG03, LSO07, LPRM02, LGQ^{+}08, MJC^{+}08, McC00, NG18, OBCS^{+}12, PRP^{+}15, PBFJ05, RPWO18, RMOW20, RGK^{+}08, RCOL09, SCH^{+}14, SGW06, SCH03, SD02, Tar16, TWBO03, WSJP17, WGS23, WDB^{+}08, WG10, vW09]. MapTree [RMOW20]. marching [CZ21, KDPN21, ZRL^{+}08]. marionettes [ZPBC19]. markdown [LKJC21]. Marker [HPP^{+}22, HLW^{+}18, RRC^{+}16, SNF05]. marker-based [HLW^{+}18]. marker-less [RRC^{+}16]. Markerless [BPS^{+}08, MWI18, ZBGB19]. markers [HMT^{+}15, LMB14, RNd^{+}07, YMJ^{+}21]. Markov [CNX^{+}08, Gol84, Gol85a, JM12, OKH^{+}17]. Markovian [GRS93]. martial [SZZK21]. mash [DSC^{+}20]. mask [FZL^{+}15, VRA^{+}07]. Masked [AHAM15, MSL^{+}24]. masking [LCD06, RSI^{+}08]. masks [ZAFW21]. Masonry [CQS^{+}23, PBSH13, WOD09, WSW^{+}12, dGAOD13]. Mass [SHS^{+}18, TBV12, TFD^{+}18, BvdPPH11, KGS^{+}18, LBOK13, SLF08, SHS^{+}17]. mass-spring [LBOK13]. masses [AMS03]. Massive [JRSS21, PFHA10, SSJ^{+}11]. Massively [GLdFN14, KS95, WQS^{+}20]. Massively-parallel [GLdFN14]. Mastering [SSII18a, PBC^{+}22]. masters [BLCD02]. MAT [LWS^{+}15]. MatBuilder [PBC^{+}22]. match [PDSH17, SLH^{+}20]. matched [LSO07]. Matching [BBB^{+}93, BBB10b, CRB23, JNK^{+}23, KSZ^{+}15, LYC^{+}22, LCL^{+}23, LYP^{+}14, MOR^{+}18, SH23, BTFN^{+}08, DML17, GCO06, HKPP20, HFG^{+}06, KFR04, MHTG05, RJ07, ST04, STZ^{+}16, SMGE11, TFBW^{+}10, TCL21, WLL^{+}14, WY04]. Matchmaker [KSG03]. Material [BBPA15, FCK22, GARP^{+}23, JTRS12, KPWP17, LL11, LSM23, LCT23, QRL^{+}23, VSW^{+}23, XLCB15, YSB^{+}15, BBP21, CSHH21, CRA11, CLSM15, CXY^{+}15, CKMR^{+}21, CPWAP08, DBD16, DTPG11, DJ18b, FLGJ19, FQL^{+}20, FGW^{+}21, GTJS17, GWW^{+}18, GZB^{+}13, GHF^{+}18, HFG^{+}18, HSC^{+}22, KP10, KRFB06, Kim18, KFB10, LMS^{+}19, LBAD^{+}06, LHdG^{+}14, MIWI16, MWI18, RXL21, SGM^{+}16, SCW^{+}21, SLH^{+}20, SARW^{+}15, SSC^{+}13, SNZ^{+}21, VLD07, VWRKM13, WQS^{+}20, WFL^{+}19, WDR11, WDR13, XSZB15, YCL^{+}17, ZAJ^{+}15, ZFWW18]. Material-minimizing [KPWP17]. MaterialGAN [GSH^{+}20]. Materialistic [SPG^{+}23]. Materials [BAU15, CZB23, GLH^{+}23, HM92, IRN^{+}22, LBK17a, RT90, SPG^{+}23, TCT23, WB23, AWL15, ATDP11, Bel18, BBO^{+}10, DI11, DBD16, DJ05, GTJS17, GHS^{+}22, GMG^{+}20, HFM^{+}10, HR13, IDN12, JAM^{+}10, JdJM14, JAG18, JB02, KMOD09, KCD^{+}16, KMX^{+}21, LKL^{+}22, LBK17b, LMPB^{+}13, MTGG11, MPH^{+}15, NGL10, PRJ^{+}13, PRFS18, PL07, RCL21, RGB16, SMCT18, SSJ^{+}14, SNZ^{+}21, Ter18, TWL^{+}05, WTL^{+}06a, WZT^{+}08a, WZB17, YAV^{+}20, ZYL^{+}20]. MatFormer [GHS^{+}22]. mathematical [LZ04, LBDA21]. MathPad [LZ04]. mating [JHC^{+}21]. Matrices [Gol85a, YCP16, KFS13, WWS^{+}05]. Matrix [HPB07, MU22, BFGS03, HWJ^{+}15, HWH^{+}16, LTT^{+}20, OP11, RCL21]. matte [BCN08]. matter [APCO21, SY21b].

Matting [GXSD23, YTBK11, CAC^{+}02, CGC^{+}03, JMA06, LL11, MMP^{+}05, SJTS04, SLKS06, WAC07, WTBS07a, XZZ^{+}21]. maximal [EDP^{+}11, YW13]. maximization [ZXJ^{+}13]. maximum [ME05, Xia97, YSW^{+}17]. maze [XK07]. MCMC [SLK^{+}24, YYW^{+}12a, ZD20]. MDE [LXY^{+}16]. Me [WZC^{+}20, MBB12, YRPF09]. Mean [HF06, JSW05, TMB18, LJH13a, PCL^{+}12]. Meandering [PGCG23]. meaningful [CTS^{+}21]. means [ABJN85, RKZ12, Zit13]. measure [GAGH14, GvdBL^{+}12, LMS^{+}19]. Measured [DWMG15, MXZ^{+}23, ZZW^{+}22a, ATDP11, PL07, STPP09, SJR18]. Measurement [DDTP15, BBO91, JKZS10, WOR11, WMP^{+}06, YTJR15]. Measurement-based [DDTP15, WMP^{+}06]. Measurements [IRN^{+}22, CHM^{+}12, HKA^{+}18]. measures [MIWB02]. Measuring [HP03, MWAM05, KRD^{+}12, PRWH^{+}18, PPZ^{+}11]. Mechanical [SMCT18, CLM^{+}13, CTN^{+}13, KLY^{+}14, MSS^{+}19, MYY^{+}10, TZCT20, XBZN19, ZKBT17, ZXS^{+}12]. Mechanics [SNW21, AVGT12, HVS^{+}09, WCL^{+}20]. Mechanics-aware [SNW21]. Mechanism [ASK^{+}22, XLX^{+}16]. Mechanisms [CSL^{+}22, CSSL21, HFF18, MZB^{+}17, RKP^{+}22, ZAC^{+}17]. Media [Ste20, BAGL19, BGL20, BRM^{+}18, Fat09b, FCJ07, GCH^{+}19, HED05, HWH^{+}16, JDZJ08, LBDF13, MPG^{+}16, NGD^{+}06, NNDJ12, NSJ14, RSA09, WZHB09, YIC^{+}10, YSC^{+}18, ZYZ21, ZWDR16]. Medial [LLF^{+}20, LYK^{+}21, LWS^{+}15, WWWG22, BO04, HWCO^{+}13, YSC^{+}16, YLJ18]. medial-axis [BO04]. Median [MU22, Ada21, Wei06]. Medical [ZWHB22]. Megapixel [WFDH18]. melanin [TOS^{+}03]. melding [DSB^{+}12]. memex [JTRS12]. Memory [JHS^{+}23, RSV^{+}23, BAM13, VSJ21]. Memory-Efficient [RSV^{+}23]. Menu [Ols86]. MERF [RSV^{+}23]. merge [WTGT09]. MergeTree [VKJ^{+}17]. Merging [RSP23, CBK20, DP13, FBH^{+}10, GKDS12]. Mesh [ASGS23, ACP^{+}01, BYG96, BZH^{+}23, CPAL22, Erl18, GZC15, HS13, HWB23, HLG^{+}22, JTCW07, JDH^{+}22, LVJ05, MMT23, NDD^{+}23, PCS^{+}23b, SK16, SFD^{+}22, SMH^{+}23, SZT^{+}07, SLMR14, SZGP05, TGBE16, ULP^{+}15, WLT16, XWX^{+}22, YZX^{+}04, YXK^{+}22, YKH10, ZJY^{+}22, ZHW^{+}06, ZGZJ16, ZXS^{+}23, ACXG09, ATC^{+}08, ACBCO17, BAS14, BCG05, CGF09, CLSA20, CPMS14, DBG14, DSSC08, DKP11, DTP15, EBCK13, FDCO03, GDC15, GJTP17, GPCP13, GSFD^{+}14, GF08, HSH20, HSL^{+}06, JT05, Jia21, JDD03, KHS10, KT03, KG05, KBZ15, LT20, LHM09, LDS^{+}16, LD14, LdPS84, LXW^{+}11, LFJG17, LSVT15, LBK16, LWL^{+}20, MPO21, MBF04, NSACO05, NGH04, PK05, SNCH08, SHD^{+}18, SYBF06, TPP^{+}11, TWGT10, TWAD09, TNWK22, VMW15, VBMP08, WZHB09, Wam16, XZY^{+}07, YLH18, YLPM05, ZZWC12, ZJ12, ZHS^{+}05]. Mesh-Based [Erl18, YXK^{+}22, SZGP05, DBG14, TWGT10]. Meshable [LB23]. MeshCNN [HHF^{+}19]. Meshed [CH02, CCK^{+}21, Wil92]. Meshes [BSTY15, CDY23, ERT14, IRWP23, KCMP23, LS00, MWM23, NAH^{+}18, Sar00, TGBE16, WSY19, YCP16, ZDF^{+}23, AGK^{+}22, AW11, ATW13, AFSR03, BBJP12, BC18, CSPF12, CS09, CWSO13, DM13, DP13, EB14, EPD09, FOK05, FKY^{+}10, FTD21, FSK04, GWY^{+}21, GGS03, GLLR11, HV04, HA18, IG03, JTPSH15, JZH^{+}21, JSW05, KFCO06, LS07, LZKW10, LSLCO05, Lip12, LPW^{+}06, LXFH15, LXY^{+}16, MS04, MCKM15, MPKZ10, OBS04, PRP^{+}15, PZKW11, PPW18, PTC^{+}15, PKC^{+}16, PKC^{+}17, RKP^{+}22, SPGT18, SC20, SBZ09, SS21, SSP08, SSW^{+}13, SGC18, SP04, SLWS07, SSK^{+}05b, SKC^{+}14, TPSHSH13, TMY^{+}11, TSG^{+}14, TLJP18, TPT16, VMW17, WLLS22, WM03, WTGT09, WPGM16, YYPM11, YSK09, YKJM12, ZFO^{+}22, ZBG15b, TGB13]. MeshFlow [DKP11]. MeshGit [DP13]. MeshHisto [SSTP15]. Meshing [ABE^{+}20, BC23, CPW^{+}23, KC23, Pan18, SRUL16, ACSYD05, BCE^{+}13, BBC22, CBK12, DA21, ECBK14, FXBH16, FLSG14, GPW^{+}17, HZG^{+}18, LLX^{+}12, LZS^{+}21, LZC^{+}18, LCBK19, MC21, PLC^{+}21, PEVBC21, SRUL17, WGF^{+}18, ZGW^{+}13]. Meshless [MHTG05, PKA^{+}05, RSLL18, FGBP11, HLW^{+}12, LZT^{+}08]. MeshTaichi [YXK^{+}22]. MeshWalker [LT20]. Meso [RB23]. Meso-Facets [RB23]. Mesoscale [QLY^{+}23, HBP^{+}21]. Meta [FR22, GLH^{+}23, Wim14]. Meta-Learned [GLH^{+}23]. Meta-Learning [FR22]. Meta-representation [Wim14]. Metal [BTSB23, DWMG15, PH15a]. MetaLayer [GLH^{+}23]. metallic [HCE03, PH15b]. metallophone [BLT^{+}15]. Metamaterial [LCT23, MWC^{+}23, MSS^{+}19]. Metamaterials [MWC^{+}23]. metamers [WKF^{+}21]. metamodel [LWL17]. metamodel-based [LWL17]. Metamolds [AMG^{+}18]. metamorphosis [COSL98, PSN20]. Metaphor [SB93]. Metappearance [FR22]. MetaSilicone [ZKBT17]. Metasurface [NKKJ23]. Method [BSSJ23, CCL^{+}22, CC23, FG90, LR90, LR91, LNZ^{+}23, LXZ^{+}23, LB84, Mai92, MHNT15, NH22, PGG^{+}23, PK83, QRL^{+}23, RLSÖ^{+}22, Roc89, RT90, Sar00, SMGC23, SDG^{+}19, SHG^{+}22, SSC19b, SZL^{+}23, SCJ^{+}23, WWYW21, XLYJ23, YSB^{+}15, ANZS18, BSD09, BGOS06, BWHT07, BBP21, CZXZ14, CMSA20, CLC^{+}20, CKMR^{+}21, CLL^{+}22, DBD16, DWK^{+}22, DTB06, FLGJ19, FQL^{+}20, FGW^{+}21, FTP16, FGG^{+}17, Gal99, GTJS17, GBO04, GHF^{+}18, HZ11, HFG^{+}18, JSS^{+}15, JZW^{+}15, KLL^{+}07, LXY^{+}16, MHM^{+}09, MTPS04, NZWC20, QLDJ22, SRF05, SMGH18, SSC^{+}13, SS17, SNZ^{+}21, TCL21, UBW99, WDT^{+}09, WQS^{+}20, XZZ^{+}21, XWWZ22, ZHLB10, ZB14, ZSTB10, ZZCJ13]. Methodology [Erl18]. Methods [CCK92, Erl18, LBK17a, LC96, MEM^{+}19, NN95, PP94, SPV^{+}16, WHG84, GWW^{+}18, JP03, LBK17b, Nas87, NNSM07, THG99, UHT17, WY16, WFL^{+}19, YLYW18, YCBvdP08]. Metric [CZ23, KH10, NDD^{+}23, WTD^{+}22, CKPS18, DMHG13, FCGH08, JFH^{+}15, LWC^{+}13, MKRH11, SPKS16, ZWL^{+}18]. Metric-aware [KH10]. Metrics [LGC^{+}23, WGY^{+}18, CHM^{+}12, CLW16, MSS^{+}19, PHBC21, RP03, TGB13]. Metropolis [BRSMD22, BJNJ18, GRS^{+}17a, GRS^{+}17b, HKD14, LKL^{+}13, LLR^{+}15, MRK^{+}14, OHHD18, Pan17, RLLG^{+}20, TLL^{+}11]. metropolised [SOHK16]. Micro [GSCO12, KSZ^{+}15, MMT23, REG^{+}09, KWN^{+}17, Par17, ZJMB11]. Micro-Appearance [KSZ^{+}15, KWN^{+}17]. Micro-Mesh [MMT23]. micro-motion [Par17]. Micro-rendering [REG^{+}09]. microcylinder [SBdDJ13]. microdisparity [TDR^{+}12]. Microfacet [BSN16, RBSM19, SHHD17, BB17, HHdD16, HP17, JHY^{+}14, WZT^{+}08b, WJF^{+}22]. Microfacet-based [SHHD17]. Microflake [WJHY23, HDCD15]. Microgeometry [DWMG15, JCRA11, RBSM19, WPMR09, YHW^{+}18]. micrography [MBS^{+}11]. micromirror [HSHF10]. Micron [GLDZ15]. Micron-scale [GLDZ15]. microphone [DRW^{+}14]. microphysical [GJZ21]. Micropolygon [HQL^{+}10, FFB^{+}09, HZ11]. microscale [NLW^{+}16]. microscopy [LNA^{+}06]. Microstructure [WHHY20, NFA^{+}15, YHMR16]. Microstructures [SBR^{+}15, ZBJ^{+}23, ZSCM17b, PRZ17, TTZ^{+}20, ZSCM17a]. Mid [AS21, FYY^{+}16, LSCS14, ZF03].

Mid-Air [AS21, LSCS14]. mid-scale [FYY^{+}16]. mid-tone [ZF03]. MIDAS [MWI18]. Mie [FCJ07, GJZ21]. migration [LWO19]. MIKE [Ols86]. millimeter [LGK^{+}16]. milling [YAV^{+}20]. million [LHLK10]. millions [HE07]. mimicking [SPSH14]. Min [HWB23]. Min-Deviation-Flow [HWB23]. Mind [HYG^{+}13]. Minimal [MHS^{+}19a, XNY^{+}16, NJR15, WC21b]. Minimization [LWS^{+}15, HS13, RKZ11, VMT06, WPL06, XLXJ11]. minimize [SdS02]. minimizers [LZ14]. minimizing [HP04, HXK^{+}19, KPWP17, MCSK^{+}17, WJZL08, Xia97]. mining [MBGS15]. Minkowski [BDD11]. MIP [CS00, GFL^{+}22]. MIPNet [GFL^{+}22]. MIPS [FLG15, TZY^{+}23]. MIPS-Fusion [TZY^{+}23]. Mirror [WCFL22, ZAE^{+}14, WGL^{+}18]. Mirrors [RSM^{+}23]. missing [ZBG15a]. Mix [PDSH17]. Mix-and-match [PDSH17]. Mixed [AAMSB20, ASB22, BSS^{+}11, BZK09, HPK^{+}17, LSM23, MM22, Wan18b, BBPD12]. Mixed-Integer [LSM23, BZK09]. Mixed-order [BSS^{+}11]. Mixed-primary [HPK^{+}17]. mixer [HHGH13, SLD17]. Mixing [NSS^{+}19, GKHH12, SJ21]. Mixture [LDS^{+}22, LSS^{+}21, RLY^{+}14, GPH^{+}18, HMP^{+}08, HGS23, VKS^{+}14]. mixtures [PRJ^{+}13, TGK^{+}17]. Mobile [NKK^{+}14, TKG^{+}23, WLS^{+}23, AMS03, HSG^{+}16, LSC^{+}22, WGJ^{+}18, XBZN19]. mobility [HLV^{+}17c]. Möbius [LF09, VMW15, VMW18]. mocap [CLM^{+}13, CWZ^{+}21b]. MoCap-solver [CWZ^{+}21b]. MOCCA [WSP21]. Modal [HZL22, JL11b, LFZ15, BDT^{+}08, DCD15, HSTP11, JLWM22, LAJJ14, RYL13, SGD21, ZJ11]. Modal-space [JL11b]. modality [WL21]. Mode [GLX^{+}22, ZSKS18, WJ19]. Mode-adaptive [ZSKS18]. Model [AZL23, BSN16, BW22, CAD19, CLT^{+}22, CT82, DK99, EHSN20, FW16, FHK14, GHCG17, GLH^{+}23, Hud94, LDS^{+}22, LMH^{+}15, PGG^{+}23, PC82, RLY^{+}14, Sar00, TUGM22, TLP06, TD23, TGZ18, VKM^{+}23, WLZ^{+}21, WJHY23, WBG^{+}16, XWH^{+}23, XLCB15, YSCL22, ZEF^{+}22, ZZC^{+}23, ZWS^{+}24, ZCP^{+}23, APCO21, ARS14, BBGB16, BWSK12, CAJ09, CH07, CZZ14, CZ11, CPSS10, CLD^{+}13, CHB^{+}12, DI11, DF88, DDSD03, Dee05, DRE^{+}11, DRE^{+}12, DWd^{+}08, DLR^{+}09, DCB^{+}22, ELFS16, EML^{+}18, Fat11, FMB^{+}17, FBGZ18, FFBB21, FD17, GHBCO21, GWM^{+}08, GMP^{+}06, GHS^{+}22, GSH^{+}20, HHdD16, HP17, HW12, HOM15, ISN^{+}20, JSB^{+}10, KCKK12, KDR^{+}16, KJ09, KNC^{+}08, KKW21, LWS02, LBB^{+}17b, LZQ^{+}22, LHM^{+}18, LMR^{+}15, MAC22, MPBM03, MM08, MC12, MGZJ20, NSS^{+}19, PLR^{+}16, PMRMB15, RGB16, RHHL02, SBdDJ13, SLF08, SCW^{+}21, SYS^{+}21, SFB^{+}09, SRNN05, TOII08, TTR^{+}17, TS12]. model [UKSI14, Van06, VMGM15, VKJ19, VJK21, WSH^{+}16, WSJP17, WSH^{+}18, WMP^{+}06, WVBR^{+}21, WBGB16, XWM^{+}20, XZZ^{+}11, XYJ13, YSJR17, YJR17, YSW^{+}20, YXFH21, YCL^{+}15, YL10]. Model-Based [YSCL22, WBG^{+}16, KNC^{+}08]. model-driven [XZZ^{+}11]. model-guided [YSW^{+}20, YXFH21]. Model-reduced [LMH^{+}15]. Modeled [DLP^{+}23]. Modeling [AMZ99, BCX95, BCV^{+}15, BR94, BSEH18, CXGS02, CFW13, CBKM15, FKS^{+}04, GLL^{+}16, GJB^{+}20, HM92, HHD^{+}22, HXM^{+}18, HZD^{+}23, Iza18, KWK09, KDH22, Kla87, LWF^{+}23, LFL^{+}23, LBJK09, LKL23, LDS^{+}11, LDPT17, LJZ^{+}23, LZZ^{+}21, Mer23, MTB^{+}13, NI22, NY94, OCH^{+}16, PBCF93, RHSH18a, Ree83, ROC^{+}21, RFL^{+}05, SSW^{+}23, TCT23, TDM^{+}14, TWL^{+}05, TB87, WZT^{+}08b, WZT^{+}08a, WMB21, WOQS05, WFY^{+}10, XPB^{+}21, ZWW^{+}18, ZYM^{+}20, ZJY^{+}22, AAL16, AZB09, AGP^{+}20, ASF^{+}13, BAS14, BB17, BHMK^{+}18, BBO^{+}09, BWS10, BJD^{+}12, BK04, BWP13, BRB^{+}19, CWW^{+}12, CLS^{+}15, CSW^{+}16, CK10, CKGK11, CEW^{+}08, CNX^{+}08, CLW^{+}14, CZL^{+}15a, DP13, DJBDDT13, DZS08, DA21, DTPG11, DZCJ21, DSC^{+}20, EBJ^{+}06, FSL^{+}15, GHP^{+}08, GIZ09, GRB^{+}18, GKTT13, GTR^{+}06, GCH^{+}19, HGY17, HPSZ11, HSTP11, HMG03, HMLL15, IKKP17, IOOI05, IYYI14, JTC09, JGGN15, KBD07, KW11, KMP07, KN02, KYC^{+}17]. modeling [KCYW13, LF02, LRAT08, LCXS09, Lee05, LT06, LST09, LT09, LPL^{+}17, LPL^{+}18, LPBM20, LPW^{+}06, MHS^{+}19b, MWAM05, MPH^{+}15, MWH^{+}06, MZWV07, NKAS08, NFD07, NFJ02, OBH02, ODAO15, PMG^{+}22, PPZ^{+}11, PCL^{+}12, PH08, PKKG03, PKZ04, PLKD18, QTZ^{+}06, RS98, RZW^{+}21, RMGH15, RDI10, RC22, RTB17, SZK15, SSTP15, SM15, SXZ^{+}12, SLR^{+}16, SSY^{+}04, SSS^{+}08, SSK^{+}17, TAV^{+}10, TSNI10, TGY^{+}09, TLL^{+}11, TZW^{+}07, TFX^{+}08, TS08, TPT16, TMB14, UKIG11, VBG^{+}13, VABW09, VBK05, VPB^{+}18, WTL^{+}06a, WLZ^{+}09, WOR11, WWY^{+}15, WMB19, WYL^{+}20, WSP21, WC10, WOD09, cWP03, WYD^{+}14, WWL^{+}19, XFT^{+}08, XFZ^{+}09, XGC07, XZZ^{+}11, XLX^{+}16, YTJR15, YCYW20, YKJM12, ZSCS04, ZCW^{+}17, ZQCL19, ZXS^{+}12]. modelless [MWI18]. Modelling [TO02, ABL^{+}21, DYY16, HDMR21, LPC^{+}11, vdHDT^{+}07]. Models [ANBH23, CAV^{+}23, CPW^{+}23, EST^{+}20, GAA^{+}23, GZX^{+}22, GDAB^{+}17a, Gre86, JLF^{+}23, KSZ^{+}15, KH17a, MSL^{+}24, NON85, NPC^{+}22, PM18, Roc89, SCB87, TLZ^{+}24, UZB^{+}23, VR94, VJ19, WLX^{+}18, XLC^{+}23, XZP^{+}23, ZDT^{+}23, ZTNW23, ZXZL23, ASK^{+}12, AAR05, BJ05, BLS^{+}21, BPK05, BGB^{+}05, CCA^{+}12, CGP^{+}21, CGG^{+}04, CDM^{+}02, gDGPR02, DS15, DAB15, DSP06, DLSCS08, DIP^{+}18, ESCK16, FGBP11, FH10, FMK^{+}03, GDAB^{+}17b, GGG^{+}13, GBFP11, GM05, GAB20, GKJ^{+}05, HBLM11, HMC11, ISF07, JHY^{+}14, JP04, Ju04, JZH07, KIL^{+}16, KMM^{+}02, KGFF14, KGS^{+}18, KSES14, KWN^{+}17, KOY^{+}11, KLM^{+}12, KS04b, KSSCO08, KH17b, LAJJ14, LOMI11, LdPS84, LRA^{+}07, LSH^{+}10, LHLF15, LSSS18, LKYU12, LBRM12, MCC09, NKJF09, NGDA^{+}16, NCVMO05, ONOI04, PHL^{+}09, POB09, PBSH13, PDF^{+}22, PNDN12, PSK^{+}12, PNH^{+}14, PJH^{+}17, PHBC21, RID10, SXZ^{+}17, SLF^{+}11, SILN11, SHOW02, SSBD03, SSBL^{+}22, SWR^{+}21]. models [SGG^{+}06, TLK09, TK14, TDM11, TREO16, TCL21, VGDA^{+}12, VBPP05, VKS^{+}14, WOR11, WMC11, WLH^{+}13, XLF^{+}11, XWY^{+}09, XCF^{+}13, ZRLK07, ZLP^{+}15, ZJMB11, ZLB16b]. Models-Past [EST^{+}20]. Modes [DTPC23, SLM^{+}23]. Modified [Lev06, LSSW19]. modifying [DMIF15]. Modular [BWL^{+}23, LHK^{+}20, LAM^{+}11, WST09, ZHRB13, FH11, GMP^{+}16, HFH^{+}17, JPG^{+}14, LLMZ16, XBZN19]. modulation [JJJ^{+}21, ZF03]. MoGlow [HAB20]. moiré [HC04, CH14]. molding [MPBC16]. molds [AMG^{+}18, AMG^{+}19, NAI^{+}18]. Moment [LWP^{+}23, BWBSH14, PKHK15]. Moment-Encoded [LWP^{+}23]. moments [GOMP98, PMHD19]. Momentum [KH17a, KH17b, MZS09, CKMR^{+}21, KUJH21]. momentum-conserving [CKMR^{+}21]. Momentum-Mapped [KH17a, KH17b]. Monitor [LR90, LR91]. Monocular [GZX^{+}22, GZC^{+}16, HXZ^{+}19, RKS^{+}14, SAA^{+}21, SGPT23, XCZ^{+}18, GVWT13, GZW^{+}16, MGC^{+}19, SWTC14, SGXT20, SGX^{+}21, WMB^{+}20, WC10, WBGB16, YPL21]. Monolith [TB20]. Monolithic [TB22, TB20, TB21, VLD^{+}13]. Mononizing [HXFW20]. MonoPerfCap [XCZ^{+}18]. monotone [LVS^{+}13]. Monster [DSC^{+}20]. montage [CCT^{+}09, LYGC15]. Monte [JM12, AW20, ALLD17, BVM^{+}17, BAGL19, CKS^{+}17, CGMS22, CHY21, DMB^{+}14, GLA^{+}19, GHZ18, HET^{+}14, HRV^{+}18, IMF^{+}21, KBS15, LADL18, McC99, OKH^{+}17, PSC^{+}15, RAMN12, RLSÖ^{+}22, RMGH15, SGH^{+}22, SSJC22, SMGC23, SHHD17, SD12, SWZ96, SJ17, YNL^{+}21, ZSGJ21, ZDDZ21, ZZXY21]. Mood [CB05]. Morfit [YHZ^{+}14]. Morph [JHS^{+}23]. Morphable [EST^{+}20, JCP^{+}10, ZEF^{+}22, MZD05]. Morphing [LLN^{+}14, SG01, AMZ99, ZWGS02]. morphogenesis [PNDN12]. morphologies [HRE^{+}08]. morphology [CB14]. morphs [RV11]. morse [FBT^{+}18, NGH04]. morse-parameterization [FBT^{+}18]. mosaics [BA83, KP02, RAKRF08]. MoSh [LMB14]. motifs [ACOH^{+}18]. Motion [ANBH23, AJM12, AFO03, ACOYL08, AFP^{+}95, CKP^{+}21, DKD^{+}17a, GSP^{+}23, GXY^{+}17a, HTCH15, IRG^{+}23, JTCW07, JPL22, KDR^{+}16, KGP02, LCL06, LWB^{+}10, LLL22, LSC^{+}08, LWS02, LCL^{+}23, LWL23a, LTF^{+}05, MWGZ09, MC12, PYA^{+}24, PSE03, PKC^{+}16, PB02, QZZ22, SAA^{+}21, SPS^{+}11, SLL^{+}21a, TZK^{+}11, TBvdP04, WFS^{+}09, WLSL10, WF96, XSZK23, YZH^{+}23, ZCM22, ZMW^{+}23, ZXS^{+}12, AJS20, AWL^{+}19, ALL^{+}20, AWL^{+}20, AXR09, AF02, Ari06, ACOH^{+}18, BHR13, BBR^{+}21, BSS^{+}13, BBA^{+}07, BLCD02, CMZP14, CH07, CWZ^{+}21b, CSSL21, CLQW08, CL09, CLS03, CBL^{+}16, CGZ^{+}05, CYT^{+}18, CHP07, DWW^{+}18, DCP^{+}14b, DMHG13, DKD^{+}17b, ETH^{+}09, EMO10, FP03, FBH21, GSH18, GPD^{+}18, GSKJ03, GXY^{+}17b, HYL12, HET^{+}14, HRvdP04, HYNP20, HRE^{+}08, HAB20, HKT10, HSK16, Hol18, HKPP20, HHC^{+}19, HDK07, HXK^{+}19, HQL^{+}10, HPP05, HCTW11, HMT^{+}15, IAF09, JYL09, JWDL19, JHS12, KA08, hKPS03, KHKL09, KG08, KLLT08]. motion [LBJK09, LCR^{+}02, LLLL21, LSR18, LAGP09, LHdG^{+}14, LAZ^{+}22, LZCV20, LP02, LHP05, LYvdP^{+}10, LWC^{+}13, LHZ^{+}21, LMB14, LXJ^{+}22, LCX16, MP07, MCC09, MYWI15, MK05, MRC05, PHT^{+}13, Par17, PH06, PCSS06, PRMG16, PMA^{+}21, PMRMB15, RAT06, RNd^{+}07, RP03, RP07, RPE^{+}05, RSH^{+}05a, RRC^{+}16, SHP04, SH07, SHU^{+}16, SSBG10, SMG^{+}20, SJA08, SGXT20, SGX^{+}21, SNF05, SKL07, SZKZ20, SMK22, SP05, TK05, TWH^{+}22, TBW^{+}12, TAH^{+}04, TGPS08, VKB^{+}18, VAV^{+}07, VSHJ12, WRDF13, WAO^{+}09, WB08, WMZ^{+}13, WC10, WMC11, WZC12, WLP16, WL16, XWCH15, XWL^{+}08, YM16, ZSKS18, ZZZX21, ZZMC13, ZMCF05, BZL^{+}17]. Motion-aware [WFS^{+}09]. Motion-based [WLSL10]. motion-beat [hKPS03]. Motion-driven [AJM12]. Motion-guided [ZXS^{+}12]. Motion-invariant [LSC^{+}08]. Motion2fusion [DDF^{+}17]. MotioNet [SAA^{+}21]. Motions [ANL^{+}23, GFK^{+}23, KH17a, WLZ^{+}21, DJ18a, HRZ^{+}13, HOKP16, KG04, KH17b, LJ14, LYvdPG12, PCSS06, RV11, TZCT20]. motivated [MKMS04]. motor [LLLL21]. motorcycle [SPGT18]. mountainous [BST09]. Mounted [YLC^{+}20, FRS19, KBBD17, LTO^{+}15, SPS^{+}11, YZH^{+}23]. mouse [HGRT04]. move [Lau18, WLY20]. Movement [DLP^{+}23, DKD^{+}17a, DKD^{+}17b]. Movements [AKG^{+}23, NRH17, SZKZ20, SZZK21, YPL21]. mover [SRGB14, WLY20]. moves [XYH^{+}18]. Movie [CFS^{+}18, FHL^{+}18, SSRB^{+}17]. MovieReshape [JTST10]. Moving [CC23, JX96, JCY23, MHM^{+}09, MLR^{+}22, SG17, CLC^{+}20, CPMK21, CC19, DWK^{+}22, DER^{+}10, FCOS05, HFG^{+}18, LDS^{+}11, LCORL07, SMW06, WJL^{+}20, ZCT^{+}21]. moving-least-squares [WJL^{+}20]. MoXi [CT05]. MPEG [MEMS06]. MPI [LK20]. MPM [FQL^{+}20, SSJ^{+}14, SXH^{+}21, TLZ^{+}24, WLF^{+}20, WFL^{+}19]. Multi [Ang17, BHMK^{+}18, BBA^{+}07, CQD^{+}18, DXZ^{+}19, GSMD07, GWB05, GLX^{+}22, HNH19, HHC18, HZL22, KL17a, KL17b, Kim10, KHH^{+}11, KL23, KQG^{+}23, KIM^{+}19, LKL23, LSA^{+}16, MEM^{+}19, MPH^{+}15, OBA^{+}03, PGZ^{+}19, PMGD21, PO18, RGB16, RYPZ23, RSH05b, RSA09, SGSS22, SM17a, SKB^{+}21, SGD21, SOG^{+}22, SJMP10, TGK^{+}17, TZY^{+}23, TFBW^{+}10, TFD^{+}18, WOR10, Wei10, WTS^{+}23, XZJ^{+}12, YR23, AAC^{+}06, ASL^{+}17, BNK10, BDW13, CTH^{+}14, DWW^{+}18, DE05, DJ05, FZBR16, FZZ^{+}20, FFLS08, FAR07, FMB^{+}17, FBGZ18, GPCP13, GHZ^{+}20, GP09, HSB^{+}12, HGF14, HDD^{+}16, HLR^{+}17, HZCJ17, JKH^{+}22, KHKL09, Kou16, KMX^{+}21, LWH^{+}11, LLL18, LTT^{+}20, LLM21, LTJ18, LMR^{+}15, MHS^{+}19b, NMD^{+}17, NOP^{+}18, NAB^{+}15, ODAO15, Par17, PLW^{+}07, RTF^{+}04, RP09, SM17b, SBK^{+}18, SHHW16, SKSY08, SCT^{+}15, SARW^{+}15, SZKZ20, TAH^{+}04, VSLD13, VBCG10, VWRKM13, VBMP08, VPB^{+}09b, WWS^{+}05]. multi [WQS^{+}20, WLO^{+}14, WGDE^{+}19, XLS^{+}11, XLX^{+}16, YCL^{+}17, ZJY^{+}21, dAST^{+}08]. Multi-aperture [GSMD07]. multi-axis [DWW^{+}18, FZZ^{+}20]. Multi-body [MEM^{+}19, GHZ^{+}20]. multi-cage-based [GPCP13]. multi-camera [SHHW16]. multi-channel [HLR^{+}17]. multi-character [KHKL09, SKSY08, WLO^{+}14]. Multi-chart [BHMK^{+}18, GP09]. Multi-Class [SGSS22, SKB^{+}21, Wei10]. Multi-Contact [KL17a, TFD^{+}18, KL17b, SZKZ20]. multi-CPU [WQS^{+}20]. multi-dimensional [WWS^{+}05]. Multi-directional [PO18]. multi-exposure [TAH^{+}04]. Multi-feature [TFBW^{+}10]. Multi-finger [GWB05]. multi-flash [RTF^{+}04]. Multi-Fluid [YR23]. multi-focal [ZJY^{+}21]. multi-frame [WGDE^{+}19]. multi-frequency [CTH^{+}14]. multi-GPU [LTT^{+}20]. Multi-Implicit-Submaps [TZY^{+}23]. multi-labelled [HZCJ17]. Multi-Laminar [SOG^{+}22]. multi-layer [LWH^{+}11, PLW^{+}07, SBK^{+}18].

Multi-Layered [LKL23, BNK10, BDW13, DJ05]. Multi-level [OBA^{+}03, RSH05b, DE05]. multi-light [FAR07]. multi-material [SARW^{+}15, VWRKM13, YCL^{+}17]. Multi-Modal [HZL22, SGD21]. multi-object [FZBR16, Par17]. multi-objective [LLL18]. Multi-operator [RSA09]. Multi-Order [KIM^{+}19]. multi-person [LMR^{+}15]. Multi-perspective [KHH^{+}11]. Multi-phase [Kim10, YCL^{+}17]. multi-plane [NAB^{+}15]. multi-projection [SCT^{+}15]. multi-rate [HGF14, HDD^{+}16]. Multi-resolution [WOR10, HSB^{+}12, KMX^{+}21, VBCG10]. Multi-Robot [GLX^{+}22, DXZ^{+}19]. Multi-scale [Ang17, BBA^{+}07, CQD^{+}18, LSA^{+}16, MPH^{+}15, RGB16, SJMP10, XZJ^{+}12, ASL^{+}17, FFLS08, FMB^{+}17, FBGZ18, MHS^{+}19b, VSLD13]. Multi-Source [SM17a, SM17b]. Multi-species [TGK^{+}17]. Multi-task [RYPZ23, LLM21]. multi-touch [RP09]. multi-user [JKH^{+}22]. Multi-View [HNH19, KL23, HHC18, KQG^{+}23, PGZ^{+}19, PMGD21, Kou16, LTJ18, NMD^{+}17, NOP^{+}18, ODAO15, VBMP08, VPB^{+}09b, XLS^{+}11, XLX^{+}16, dAST^{+}08]. Multi-viewer [WTS^{+}23]. multi-viewpoint [AAC^{+}06]. Multibody [MHNT15, PAK^{+}19, CLL^{+}22, Erl07, KSJP08, LT08, TJ07]. multicopter [DSZ^{+}16]. Multidimensional [HJW^{+}08, HH90, RO85, RO87, WABG06, GM09]. MultiFab [SARW^{+}15]. MultiFLIP [BB12]. Multifocal [QCOS23, CKS18, MSM^{+}17]. Multigrid [KS11, BFGS03, KH08, LZBCJ21, SHM22, SYBF06, SBZ09, TJM15, ZSTB10]. multilayer [HBLM11, HLBR12, WLHR12, YJB^{+}14]. multilegged [KLV20]. multilevel [GPCP13, KS11, WWW22]. multilinear [TS12, VT04, VBPP05]. Multimaterial [DBG14]. Multimodal [MMHP23, PABE^{+}21, VPHB^{+}21, WCPM18]. multipath [KWB^{+}13, MHM^{+}17]. Multiphase [YJL^{+}16, CLC^{+}20, YYW12b]. multiphysics [LSD^{+}22]. multiplane [ZTF^{+}18]. Multiple [EPO91, HHdD16, HC86, Joe89, KF93, LSSF06, NID20, RLY^{+}14, SYZ^{+}23, XH18, AWGB04, APS^{+}14, FG11, GYGS22, KGB^{+}09, KVG^{+}19, LJJ^{+}18, MYRD14, MM06, MWM08, PBS04, RMOW20, RXL21, SDIN18, WTL05, WSZ^{+}14, WYL^{+}20, WJF^{+}22, WOQS05, WSVT13, WMW15, YCR^{+}15, ZYWK08]. multiple-bounce [WJF^{+}22]. Multiple-Fluid [RLY^{+}14, RXL21, YCR^{+}15]. Multiple-knot [Joe89]. Multiple-scattering [HHdD16]. Multiplexed [HKD14, LLW^{+}08, NZV^{+}11, RNd^{+}07, WGT^{+}05]. Multiplexing [NKKJ23]. Multipole [BSSJ23, STZ14]. Multiprocessor [GHCC88]. Multiresolution [JP03, LDW97, VR94, dFP95, BMBZ02, BA83, CGG^{+}04, DHW^{+}11, GM05, KN02, KS98, Lee05, MESK22]. Multiscale [CSHD21, FAR07, HRRG08, WYZG11, HH10, HMC11, PKG06, SSD09b, TWGT10, TLHD03]. multisensory [EMO10]. Multisided [War92, LD89]. Multisource [KSL^{+}23]. multispectral [LYL^{+}16, MRK^{+}13, SBK^{+}18]. Multistable [VRP^{+}23, ZAB21]. Multithreaded [HMLB16]. Multiuser [DFYL19]. multivalued [MASS15]. Multivariate [CGM91]. Multiview [DRC^{+}15, GFT^{+}11, HKC^{+}18, KN06, LWL^{+}23b, KDW^{+}17, LES09, WSS^{+}19]. Multiway [Tsa15]. muscle [GvdPvdS13, LPKL14, LPLL19, PH08, SNF05]. muscle-actuated [LPLL19]. muscle-based [GvdPvdS13]. muscles [LYP^{+}18, LZQ^{+}22]. musculoskeletal [ABL^{+}21, FLP14]. Musculotendon [SKP08].

Musculotendons [WVY^{+}22]. mush [LL19, LVGO21]. music [CTL^{+}21, LYGC15]. music-driven [CTL^{+}21, LYGC15]. Mutations [SLK^{+}24, LLR^{+}15]. My [HAB16]. MyStyle [NAH^{+}22].

NADA [GPM^{+}22]. Naive [Mor11]. Nanophotonic [CSL^{+}23]. nanostructural [AHB18]. narration [JMD^{+}17]. narratives [CM10]. Narrow [LHZ^{+}18, ABO16]. Narrow-band [LHZ^{+}18]. native [WTS^{+}23]. Natural [JMA06, KAEE20, SJ94, SGWJ18, WTBS07a, ZLC^{+}22, BAC^{+}06, KHD14, Pel10, RPE^{+}05, ZMSS18]. natural-constraint [KHD14]. naturalistic [NB11]. Nature [PGG^{+}24]. Nautilus [LSS^{+}17]. navigate [RMBB^{+}13]. Navigating [LS00, SSC19a]. Navigation [DWX^{+}21, CDSHD13, KC19, SAZK06]. NC [HA92]. NDFs [RBSM19]. Near [ALK^{+}17, BHR13, CBYJ23, HGM14, JYW^{+}23, KKN^{+}13, LL13, LLH04, SHL^{+}17, TLP07, BGL20, CAJ09, CTS^{+}20, HCW15, JBM^{+}17, JBLL18, KPM^{+}17, MGK17, XNY^{+}16, YJR17, YL08]. Near-exhaustive [KKN^{+}13]. Near-eye [ALK^{+}17, LL13, SHL^{+}17, CTS^{+}20, HCW15, JBM^{+}17, JBLL18, KPM^{+}17, MGK17]. near-field [BGL20, XNY^{+}16]. Near-infrared [JYW^{+}23]. Near-invariant [BHR13]. Near-optimal [TLP07]. Near-Regular [HGM14, LLH04]. near-rigid [CAJ09]. near-unactuated [YL08]. nearest [MSDL17]. NeAT [RWL^{+}22]. Neck [ZZC^{+}23, LT06]. need [KRF^{+}18]. needed [Nai98]. needle [CAR^{+}09]. Neighborhood [FFWL^{+}22, Man86]. Neighboring [LLK^{+}19]. NeLT [ZHM^{+}23]. Neo [SDK18]. Neo-Hookean [SDK18]. NeRF [GZX^{+}22]. NeRFactor [ZSD^{+}21]. NeRFFaceLighting [JCFG23]. NeRO [LWL^{+}23b]. NeROIC [KOC^{+}22]. NeRSemble [KQG^{+}23]. Nested [SVJ15, LH04]. Nesterov [MHNT15]. net [YHCOZ18, GWY^{+}21, YHCOZ18, HXC^{+}20, YML^{+}23, MYC^{+}22]. NETRA [PMOR10]. Nets [BC18, IRWP23, LXG^{+}22, RHSH18a, WF96, AW21, RHSH18b]. Network [HLP^{+}22, HZC^{+}22, DAD^{+}18, GLA^{+}19, GLT^{+}21, GLZ^{+}21, HHF^{+}19, LK20, PYB^{+}16, PGZ^{+}19, TVLF20, WWL^{+}19, Xia21]. Networks [GZC15, HLG^{+}22, HKC^{+}18, KLM24, LCT23, PZWW23, SFD^{+}22, TSLP14, YZW^{+}16, ALL^{+}20, Ada21, AML18, BVM^{+}17, BB15, DZCJ22, FvKBCO16, GWY^{+}21, GSV^{+}17, GDG^{+}17, HKS17, HYZ^{+}18, HWG14, KMM^{+}17a, LDPT17, LZT^{+}19, MGA^{+}17, MLL^{+}21, PLS^{+}15, PO18, RDL^{+}15, SED16, SSISI16, WLG^{+}17, WLW^{+}19, WLT22, XCS^{+}14, ZCT16, ZSKS18, ZZCJ13]. Neumann [SMGC23]. NeuMIP [KMX^{+}21]. Neural [ALS^{+}18, AGK^{+}22, ARMCO23, AGL^{+}22, AONA22, BWC^{+}23, BGF^{+}23, BME22, BDM^{+}21, CZ21, CTFZ22, CXW^{+}23b, CGP^{+}21, CPW21, DYZ^{+}23, DPD22, DWS^{+}23, GLC^{+}23, GFL^{+}22, GSRN21, GIGM22, GZC15, HCZ21, HLP^{+}22, HHL^{+}24, HCH22, IRG^{+}23, JCFG23, KLM24, KLR^{+}22, LCT23, LXZ^{+}19, LHR^{+}21, LWL^{+}23b, LSS^{+}19, MLL^{+}22, MHGCO21, MMR^{+}19, MRKN20, NDS^{+}23, PZWW23, PSN20, PHM^{+}23, PCPW20, PMGD21, RYW^{+}22, RKB^{+}23, RMP^{+}23, SGX^{+}21, SLF22, SZZK21, TZY^{+}23, TZJ^{+}22, VSW^{+}23, WLJ^{+}22, WSND^{+}23, WSP^{+}23, WZX^{+}23, WZK^{+}23, WXZ^{+}23, XGZ^{+}23, YZW^{+}16, YZL^{+}22, YFFA21, ZFT^{+}21, ZZZ^{+}22, ZTNW23, ZJY^{+}22, ZWS^{+}24, ZBW^{+}20, ZZW^{+}22a, ZHM^{+}23, ZBX^{+}21, ZXS^{+}22, ZZLH23, AAL16, ASB22, AML18, BB15, BMSR20, BME21, CAD^{+}21, CWZ^{+}21a, CWZ^{+}21b, CLZ^{+}22, GCD^{+}20, HKS17, HYZ^{+}18, IMF^{+}21, KMM^{+}17a, KOWD21, KAGS20, KOC^{+}22, KMX^{+}21, LDPT17, LPL^{+}18, LAH^{+}21, LAZ^{+}22, LGB^{+}21, LSS^{+}21, LCD^{+}20b, LXJ^{+}22, MKZ^{+}21, MGA^{+}17, MLL^{+}21, MBPY^{+}18, MPH^{+}20]. neural [MTA^{+}20, MRNK21, MESK22, NPLX22, PSH^{+}21, PO18, RDL^{+}15, RWL^{+}22, SED16, SJ22b, TZN19, WLG^{+}17, WCSC22, WXZ^{+}22, Xia21, XZK^{+}20, YKZ^{+}22, ZSKS18, ZWCM21, ZLY^{+}21, ZYSK21, ZSD^{+}21, ZXS^{+}21]. Neural-Singular-Hessian [WZX^{+}23]. NeuralMarker [HPP^{+}22]. NeuralRoom [WLJ^{+}22]. NeuralSound [JLWM22]. NeuralTailor [KL22]. NeuralVDB [KLM24]. neuromuscular [LT06]. NeuroSkinning [LZT^{+}19]. never [DFW20]. newton [LBK17b, BDCDA11, CLL^{+}22, HCLK24, LBK17a, MEM^{+}19, ZCT22, ZBK18]. Newtonian [HNO^{+}23, ZLQF15]. next [LZHJ20]. nicely [DH96]. Night [DII23, WM14]. NIMBLE [LZQ^{+}22]. no [LWC^{+}13, PSA^{+}04, SLV^{+}13]. no-flash [PSA^{+}04]. no-reference [LWC^{+}13]. nodal [Dav20]. Node [RSP23, MBF04]. NodeGit [RSP23]. Noise [ARW22, CTW09, MEA^{+}18, QCHC17b, TTD22, YLC^{+}20, APC^{+}16, AW20, BHN07, CZJ12, CGW^{+}13, CD05, EMU15, Fat11, GLLD12, GSV^{+}14, GZD08, GKTT13, HSD13, JZW^{+}15, KTBV16, KBS15, KP11a, KCODL06, LLDD09, LD11, LWSF10, McC99, ODJ04, Per02, QCHC17a, SD12, SZG^{+}13, TEZ^{+}19, WYL^{+}14, Wei10, ZHWW12, dGBOD12]. noise-aware [EMU15]. Noise-based [TTD22]. Noise-Resilient [YLC^{+}20]. noisy [YSQS07]. Non [BSN16, DMZ^{+}17, HSGL11, HNO^{+}23, HOZ^{+}19, IH20, JDD03, JHR^{+}15, KSSCO08, MEM^{+}19, MHZ^{+}21a, MASS15, MHR^{+}16, NHAH03, RTF^{+}04, RSM^{+}23, SSW^{+}13, TUGM22, ZSW^{+}10, ZZB^{+}18, AIH^{+}08, BBO^{+}09, BRM^{+}18, BHW16, BR07, CCA^{+}12, CSHH21, CADS09, CWK^{+}20, DMIF15, FZL^{+}15, FLGJ19, FQL^{+}20, FAW19, HSB^{+}12, HRV^{+}18, KP11a, KDH22, LBAD^{+}06, LFS16, LZQ^{+}22, LWO19, MKZ^{+}21, NRH03, PPTSH14, PLR^{+}16, RCOL09, RKZ12, VJK21, WW11, XSH^{+}20, YSQS08, ZLQF15, ZNI^{+}14]. non-assembly [CCA^{+}12]. non-blind [YSQS08]. non-equilibrated [FLGJ19]. non-euclidean [KDH22]. non-exponential [BRM^{+}18, VJK21]. Non-homogeneous [KSSCO08, FAW19]. Non-invasive [NHAH03]. Non-iterative [JDD03]. Non-Line-of-Sight [RSM^{+}23, HOZ^{+}19, IH20, CSHH21, CWK^{+}20, LWO19]. Non-Linear [DMZ^{+}17, MHZ^{+}21a, MHR^{+}16, BBO^{+}09, HSB^{+}12, NRH03, PLR^{+}16]. Non-local [ZSW^{+}10, DMIF15, RKZ12]. Non-manifold [MASS15]. Non-Newtonian [HNO^{+}23, ZLQF15]. non-orthogonal [PPTSH14]. Non-Parametric [BSN16, TUGM22, LBAD^{+}06]. Non-photorealistic [RTF^{+}04, KP11a]. Non-Planar [JHR^{+}15, MKZ^{+}21]. Non-polynomial [SSW^{+}13]. non-reflecting [BHW16]. Non-rigid [HSGL11, BR07, LZQ^{+}22, ZNI^{+}14]. Non-smooth [MEM^{+}19]. Non-stationary [ZZB^{+}18, AIH^{+}08, RCOL09]. non-sticky [FQL^{+}20]. non-successive [FZL^{+}15]. non-uniform [CADS09, LFS16, WW11]. non-uniformly [HRV^{+}18]. nonconforming [EB08]. Nonconstant [FG90]. Nonconvex [GBF03, BDD11]. nondiffuse [WS99]. nondissipative [SSK05a]. nonhomogeneous [GMP09]. Nonlinear [CWC11, DTPC23, FMR20, HMG03, LHW^{+}10, LCT23, VTSSH15, XSZB15, ZB94, CAJ09, CPWAP08, CQD^{+}18, FQL^{+}20, KJDL09, LHP05, MLPP09, PMS12, SYS^{+}21, TZCT20, TOG22, VMTF09]. nonlinearity [KTS^{+}14]. nonlinearly [SNZ^{+}21]. nonminimal [ABJN85]. Nonparametric [Hob90]. nonphotorealistic [HTER04]. Nonplanar [Mil87]. nonreflective [SKM10]. Nonrigid [SK16, WAO^{+}09]. Nonsingular [BHN98]. Nonsmooth [SRH^{+}15, BDCDA11]. Nonuniform [BSB16, BSB17, MFR^{+}10]. norm [TK14]. Normal [GFL^{+}22, LBB22, XDW^{+}23, FSK04, HSRG07, RSM10b, SHHD17, TWBO03, VW97, WFL^{+}15, WLT16, WTBS07b, WSTS08, YHJ^{+}14, YHMR16]. normal-mapped [YHJ^{+}14]. Normal-to-Anisotropic-Roughness [GFL^{+}22]. normalising [HAB20]. Normalizing [AZMW21]. Normals [HOZ^{+}19, LXSW23, BJTK18, HLHZ08, NLW^{+}16, NRDR05]. Notebook [Ols88]. notes [SBLD15]. Novel [HSV^{+}22, KLR^{+}22, WBF^{+}17a, GI04, LZF10, MPK09, WBF^{+}17b, XSH^{+}20, YWH13]. Novel-View [KLR^{+}22]. novice [KP09, KP10]. nowcasting [HHP^{+}21]. NPR [KMM^{+}02]. null [MGJ19]. null-scattering [MGJ19]. Number [RvE93, XDW^{+}23, GLD^{+}19]. Numbers [FGC23, BDS^{+}18, JKSH13, LRFH13, QJ21, RAD12]. Numeric [EC93]. Numerical [CBW^{+}18, KMOD09, OF01, CZXZ14, CLMK17, Jia21, KW03, SAJ21, XSH^{+}20]. Numerically [CCW93, Hob91]. NURB [LC96]. NURBS [CADS09, GBK05, MRF06, SF09, SFL^{+}08, TQ94, XLC^{+}23]. NURCCs [SZBN03]. Nyström [WDT^{+}09].

O [ASF^{+}13, WLG^{+}17, WSLT18]. O-CNN [WLG^{+}17, WSLT18]. O-snap [ASF^{+}13]. Obama [SSKS17]. Object [ABJN85, BC02, Bar86, JWD^{+}23, KSH^{+}14, LFL^{+}23, LWL23a, LXS^{+}18, LXG^{+}22, PKH^{+}17a, SB93, YSW^{+}23, YYW12b, YSHWSH16, ZZT^{+}21, ZHM^{+}23, BWSS09, BSL12, BdSP09, DF88, FZBR16, FRS^{+}12, FCW^{+}17, HYZ^{+}18, HWV^{+}18, KSES14, KPH18, LD05, LSS05, MYWI15, Par17, PKH^{+}17b, SHX^{+}22, SHH99, TK14, VPB^{+}22, XZZ^{+}11, XHS^{+}15, XSZ^{+}16, YLNP12, YHL^{+}18, ZBYX19, ZYSK21, ZQPM12]. Object-aware [LXS^{+}18]. Object-based [BC02]. Object-Oriented [Bar86, SB93, ZHM^{+}23]. Object-space [YYW12b]. Object-Wrapping [LXG^{+}22]. objective [GGT17, LLL18, Rus19]. objectives [WHDK12]. Objects [CSAP21, CRCM23, Kaj83, KK91, LWL^{+}23b, MPB17a, NKS^{+}23, RHW94, Ree83, RYPZ23, XSL^{+}22, vW84, ALY08, BWBSH14, BBO91, BVG11, CZS^{+}13, CMT^{+}12, CNR08, Dav20, DCD15, DLL^{+}18, EHA12, FCODS08, GLL^{+}04, GOMP98, GMB17, HSvTP12, HK10b, HvKW^{+}16, HFG^{+}06, IM10, IZT^{+}07, ICG17, JTRS12, JP03, KHFH11, KUJH21, KRD^{+}12, KLY^{+}14, KOC^{+}22, LKB22a, LNWB03, LSZ^{+}14, LWL^{+}20, MZL^{+}17, MPI^{+}18, MPB17b, NLGK18, OHR14, PLR^{+}16, SvTSH14, SY05, SSM15, SOA11, SDW^{+}16, SVB^{+}12, SBK11, SM06, SZS^{+}08, TISM16, VA88, WTL05, WTL06b, WWY^{+}13, WWY^{+}15, WKHA18, WW13, WZQ^{+}18, YZL^{+}22, YTBK11, ZIT^{+}18, ZIT^{+}19, ZBYX19, ZCT^{+}21, ZSMS14, vTSSH13]. oblivious [MBK^{+}10, YLPM05]. Obscuring [HRvdP04]. observations [SCH^{+}16]. obstacles [ABO16]. obstruction [XRLF15]. obstruction-free [XRLF15]. obstructions [SBB^{+}22]. Occluded [KZSR16, WCF07]. Occluder [WLL23]. Occluders [HOZ^{+}19, WLL23, EHDR11, GRBN09, LRAT08]. Occluding [LBHH23]. Occlusion [MJJG18, EDR11, HK18b, KE18, PFHA10]. Occlusion-Aware [MJJG18, HK18b, KE18]. Occupancy [DXG^{+}23, LBB22]. ocean [HQT^{+}21, DKD^{+}17a, DKD^{+}17b]. Octahedral [SVB17a, ZVC^{+}20, LZC^{+}18, SVB17b]. OctFormer [Wan23]. Octree [BD02a, FFWL^{+}22, Wan23, AB20, GWAB19, LGF04, PK05, VA88, WLG^{+}17, WLT22]. Octree-based [Wan23, WLG^{+}17]. octree-represented [VA88]. Octrees [BN90, WV92, ABJN85]. Ocular [KAW20]. off [MHM^{+}17]. off-the-shelf [MHM^{+}17]. offline [LCX^{+}21]. offs [LDS02, SWC^{+}18]. offset [HLR^{+}14, MAB^{+}15, PRLH^{+}22]. offsets [Far89]. Offsite [ZXH^{+}20]. omni [MUB15]. omni-directional [MUB15]. OmniAD [MUB15]. omnidirectional [JMK^{+}22]. OmniPhotos [BYLR20]. Omnistereoscopic [SBSH18]. On-line [VKS^{+}14, PSBM07]. On-Sensor [CSL^{+}23]. On-set [WSVT13]. on-surface [RTD^{+}10]. On-the-Fly [DNZ^{+}17b, VSLD13, XNZ^{+}22, DNZ^{+}17a, LYYB13, RTS^{+}07]. once [HA18]. One [OF01, JLF^{+}09, RFS22]. One-Dimensional [OF01]. one-pixel [RFS22]. one-to-many [JLF^{+}09]. Online [BVG11, BWP13, HET^{+}14, HRL15, HLW^{+}18, HMM^{+}21, LGHL23, TZY^{+}23, TTR^{+}17, VKW^{+}23, ZXTZ15, ZPYX23, CKP^{+}21, KJ09, KOC^{+}22, LCX^{+}21, RMBB^{+}13, STJ^{+}17, VKK18, YGH^{+}17, ZZZX21]. Only [APCO21, DHB^{+}16, FBC18, HM20, LZF10]. onto [GFK^{+}23, YAB^{+}22]. Opacity [GRT13, MPN^{+}02]. opaque [SOA11]. open [MRA^{+}13, YYW^{+}12a]. OpenFab [VWRKM13]. Opening [SKSJ20]. OpenMPD [MMHP23]. OpenSurfaces [BUSB13]. OpenSVBRDF [MXZ^{+}23]. operated [Ros20]. Operation [BN90, JCG^{+}21]. Operations [JHR22, RNP^{+}22, YXK^{+}22, AD03, HSB^{+}12, IM10, KH08, LZKW10, Man86]. Operator [AOCBC15, BDK^{+}16, LKG^{+}03b, RSA09]. Operators [EC93, MNB23, ACSM12, AML18, Bel18, KW03, LCTS05, LJO19, MBWB02, WNEH22]. Opponent [SCB87]. Opt [DMZ^{+}17]. OptCuts [LKK^{+}18]. Optical [CFP^{+}21, GARP^{+}23, OK10, PRM14, SS19, CWZ^{+}21b, HLW^{+}18, Hol18, HLZ10, HLBR12, JMB^{+}20, LJM^{+}16, MLYZ19, SGM12, WVJH17]. Optically [Ste20, SZD^{+}20]. Optics [TMM^{+}21, XWH^{+}23, Fre16, GJZ21, IGP^{+}17, LGX^{+}13, NYY04, SDP^{+}18, SSY22, WFDH18, XWM^{+}20, YHW^{+}18]. Optimal [AHL17a, AHL17b, BLdG^{+}16, GFK^{+}23, GVNB18, KVG^{+}19, LM97, LPC22, MWM23, MSK^{+}23, NAB^{+}15, PFP^{+}22, SGSS22, SW18, SSC18, SV19, WP09a, XLY^{+}22a, XLY^{+}22b, YL10, BdSP09, BPC16, BC19, FAB^{+}18, GKK^{+}24, HPC21, JKH^{+}22, JKT^{+}15, KCPS13, LCCS18, LDS02, MGS^{+}21, McK87, MSM^{+}17, NJR15, SH07, SdGP^{+}15, TLP07, WPP14, WJ19, XSHR18, ZLWH16, dGBOD12]. optimality [BCG05]. Optimisation [POK23]. optimised [DFM13]. Optimization [ASF^{+}13, CZ23, CKT^{+}23, CGMS22, DMZ^{+}17, HNH19, HTS^{+}22, HCLK24, JYL09, KPWG24, LWF^{+}23, LCD^{+}19, LDS^{+}22, LSM23, MHCT23, SMH^{+}23, SYM^{+}24, TB22, TMM^{+}21, WLF^{+}20, WKMH^{+}23, ZWZ^{+}22, ZSCM17b, BZCC10, BKR17, BOFN18, CH07, CGM11, CNZ^{+}22, CÖS19, DWS^{+}20, FH04b, GWW^{+}18, GPD^{+}18, GRT13, HFF16, HDN^{+}16, HMG03, HZG^{+}12, JTSW17, KSNG17, KGL16, KSSI17, KEBK05, LDK^{+}18, LZ14, LLW04, LWC12, LHdG^{+}14, LLMZ16, LWL17, LKK^{+}18, LGL^{+}19, LLJ22, LHP05, LXY^{+}16, LH18, LHZ^{+}18, LSVT15, MDLW15, MTP12, MWTK13, MAB^{+}15, MHR^{+}16, NIR^{+}21, PL07, PDZ^{+}18, PTH^{+}17, RKAP^{+}12, SXZ^{+}17, SZB18, SZT^{+}07, SPSH^{+}17, SCT^{+}15, SaLY^{+}08, SDP^{+}18, SHOW02, SMGH18, SLWF14, TB21, TBC^{+}16, TWAD09, TYY^{+}19, TWZ20, UKSI14, WHSL11, WSW^{+}12, XWW^{+}14, YLYW18, YCL^{+}15, YYT^{+}11, YYTC12, ZCLJ20, ZZZX21, ZK14, ZSCM17a, ZBK18]. Optimization-Based [TB22, ASF^{+}13, JYL09, FH04b, GPD^{+}18, TB21]. Optimize [RMP^{+}23, AMA^{+}19]. Optimized [DZPZ09, WTSL08, WK21, LH16, LKK^{+}16, MWBR13, MMdGD11, OHB^{+}11, SLWS07, WLSL10, XUC^{+}14, ZXKL^{+}20]. Optimizing [AW21, AKJ08, CAA09, DKZ^{+}21, GSH18, HSGL13, HH10, KS21, KKW20, Ter18, TZZ21, WFH09, WFH10, WHDK12, BWBSH14, LHKR10, LYH^{+}15, TDM^{+}14, WSP21]. OptiTrap [PFP^{+}22]. OptiX [PBD^{+}10]. Orbifold [AL15, AL16, AKL17].

Order [ASGS23, BIW93, BV22, BSEH18, EC93, Jan91, KIM^{+}19, LWP^{+}23, MJJG18, BSS^{+}11, GKH^{+}13, GI04, JZH^{+}21, LLJ^{+}23, LLK^{+}20, MC21, MAB^{+}15, RMB07, SMM14, SYS^{+}21, SXH^{+}21, ZRB14]. Order-invariant [ASGS23]. Ordered [BSW02, RMGH15]. Ordering [AECO15, Wil92, AW20]. Ordinal [CA24]. organization [HSS^{+}13]. Organizing [XMZ^{+}14, PHL^{+}09]. Orientable [CZM^{+}23, TTZ^{+}20]. Orientation [XDW^{+}23, FCODS08, HZM^{+}08, LSL^{+}18, RPWO18]. orientation-preserving [RPWO18]. Oriented [Bar86, SB93, SRX^{+}23, ZHM^{+}23, CGM11, CTL^{+}21, FvKBCO16, MC11, QHY^{+}16]. Orienting [MHZ^{+}21b]. Origami [NKS^{+}23]. ornamental [ZCT16]. ortho [TS08]. ortho-image [TS08]. orthogonal [PPTSH14, RHSH18b, LF08]. Orthotropic [MSDL17, WYW23]. OSCAM [OHB^{+}11]. oscillatory [KA08]. Oslo [Mey91]. OT [PK05]. Out-of-core [IG03, NNSM07, SBZ09, WWS^{+}05, CGG^{+}04, WHY^{+}13]. outdoor [LRT^{+}14, SPDF13]. Outdoors [DRC^{+}15]. outfit [YYTC12]. Output [PK83, SO92, JP04, JBP06]. Output-Sensitive [SO92, JP04, JBP06]. outward [ZLH^{+}21]. OverCoat [SSGS11]. overcomplete [MWBR13]. overdraw [SNB07]. overlap [KA08]. overlapping [Pik83]. overlay [SdS02]. overview [ACOYL08]. Owen [APW23]. Own [WCZ^{+}22].

P [LTT^{+}20]. P-cloth [LTT^{+}20]. P2M [ZXS^{+}23]. P2P [YHCOZ18]. P2P-NET [YHCOZ18]. pack [CZL^{+}15b, HXC^{+}20]. Packed [FPSG22]. packets [JW17]. Packing [CRCM23, SHWP09, XGZ^{+}23, YPL^{+}23, ZPYX23, LVS18, LFY^{+}19, RRS13, YCL^{+}15]. pad [RP09]. paGAN [NSX^{+}18]. paged [AGL^{+}17, SABS14]. Paint [LSS09, PBMF07]. painterly [BBS^{+}13, BOD^{+}13, ZZXZ09]. Painting [ARS14, CH04, gDGPR02, LFB^{+}13, SED16, SMPZ15, ZSSJL20, CKIW15, LBDF13, MP08, SSGS11, SBK^{+}18, SJ21, XCW14]. Painting-to- [ARS14]. paintings [BSS^{+}11, BTFN^{+}08, TDSG15, XXK^{+}06]. pair [BDD11, HXM^{+}13, ÖG12]. pairs [AP08, PSA^{+}04, YHL^{+}18, YSQS07]. pairwise [AMCO08]. Palette [CFL^{+}15, CKT^{+}23, MVH^{+}17, SLD17, TEG18]. Palette-Aware [CKT^{+}23]. Palette-based [CFL^{+}15, TEG18]. palette-space [MVH^{+}17]. palettes [DLX^{+}21, KC21]. panchromatic [WHB^{+}12]. Panel [XIM18]. Paneling [EKS^{+}10]. Panels [JWT^{+}23, PSB^{+}08]. PanoMan [WLZ^{+}21]. Panorama [CWL22, HLSH18, STP12, ZCB^{+}22]. Panoramas [YLC^{+}20, AAC^{+}06]. Panoramic [AZP^{+}05, DK09, HCS13, LKK^{+}16, OEE^{+}18, PAAG21]. PantaRay [PFHA10]. Paparazzi [LTJ18]. Paper [SRH^{+}15, CT05, LBDA21, LSH^{+}10]. papercraft [MS04]. Papers [Ano85b, Ano92b, Spe03]. Paradigm [BBB^{+}93]. paradigms [KP09, KP10]. paradise [HBP^{+}21]. Parallax [KAW20, KDR^{+}16, LHKR10]. Parallel [BWWM10, CG89, CZY17b, HMLB16, KS95, LH05, NM16, WDB^{+}08, Wei08, AVGT12, ASA^{+}09, CZY17a, FFB^{+}09, GLdFN14, GLHL11, REG^{+}09, SS10a, TBV12, WQS^{+}20, YXH14]. Parallelepipeds [PVY90]. Parameter [FHXW22, FG90, JW15, Pag98, Pat85, Pat87, ZWTP23, MMT18, YLYW18, ZS00]. parameter-free [MMT18]. Parameterization [LCOLTE07, MLL^{+}22, AB89, ACP03, BN21, DKZ^{+}21, DHB17, DJ18b, FBT^{+}18, GDC15, GGS03, HSH20, KG04, KS04b, LKK^{+}18, LYvdPG12, VLV^{+}21, PKC^{+}17, RLL^{+}06, SZC^{+}22, SS15, TBTS08, WSSK13, ZMT05].

Parameterization-free [LCOLTE07]. Parameterizations [LFZ^{+}23, FOL^{+}21, KLS03, LFO^{+}22, LYNF18]. parameterized [BWSK12, LLKP11]. Parameterizing [HSH10, Gos00]. Parameters [DB88, Res87, DIO^{+}12, GJZ21, LN22, SD12, ZWDR16]. Parametric [BSN16, Fil89, JCY23, KSH23, LL23, MD94, MIB15, QLH^{+}22, RS14a, RS18, SSB^{+}17a, SLM^{+}17a, WA23, ZEF^{+}22, ZZC^{+}23, ZFL^{+}10, BMM^{+}21, BBGB16, HB89, LBAD^{+}06, MB21, RS98, SSB^{+}17b, SLM^{+}17b, SLD17, SD89, TUGM22, VKS^{+}14, WDB^{+}08]. Parametrization [CSZZ20, CC23, Lev21, Lev23, LCBK19, BCW17, BBC22, CBK15, CLW16, MZ12, MZ13, MPZ14, PTSZ11, PH03, TPP^{+}11, WZ14]. Parametrizations [BHN98, PU06]. Parametrizing [LXSW23]. Pareto [LDS02, MGS^{+}21]. Pareto-optimal [LDS02]. Paris [DSG^{+}12]. Parquetry [IWHH20]. Parsing [BGK^{+}13, CZL^{+}14, FNO89, LXZ^{+}23, LPBM22, ZZXZ09]. Parsing-Conditioned [LXZ^{+}23]. Part [HKC^{+}18, JPL22, BJD^{+}12, HPG^{+}22, HLV^{+}17c, KLM^{+}13, SFCH12, XLZ^{+}10, YHL^{+}18, Gol84]. part-based [BJD^{+}12, KLM^{+}13]. Partial [HFW^{+}19, HOZ^{+}19, Lev23, MGP06, XZT^{+}09, BWS10, BC19, GCO06, XZJ^{+}12]. Partial-Shape [HFW^{+}19]. Participating [Fat09b, FCJ07, HED05, HWH^{+}16, JDZJ08, NGD^{+}06, NNDJ12, NSJ14, YIC^{+}10, ZYZ21]. Particle [MMHP23, Ree83, ZGW^{+}13, APKG07, CLC^{+}20, DWK^{+}22, FOA03, FGG^{+}17, GPH^{+}18, HRL15, JSS^{+}15, LAD08, MMCK14, MBT^{+}15, NFD07, QLDJ22, RXL21, SRF05, SG11, TBBC^{+}22, WDK^{+}21, WAK20, XIAP^{+}17, YCL^{+}17, YT13, ZLB16a]. Particle-Based [MMHP23, ZGW^{+}13, LAD08, MBT^{+}15, YT13]. particle-in-cell [FGG^{+}17, JSS^{+}15, QLDJ22]. particle-in-polyhedron [TBBC^{+}22]. particle-laden [GPH^{+}18]. particles [MC11, PTC^{+}10, WJL^{+}20, YHK07, dGWH^{+}15]. partition [ACA^{+}19, OBA^{+}03]. partitioned [ANZS18]. partitioning [JKH^{+}22, LBRM12, SHFH11, YCL^{+}15]. parts [LOMI11, LBRM12, YSL^{+}14]. party [EML^{+}18]. pass [CCOST05]. Passive [BCK^{+}23, BHB^{+}11, BHPS10, CB04, DRW^{+}14, FRSL08, HMT^{+}15, KGL^{+}22]. Past [EST^{+}20]. paste [BMBZ02, LSS05, LvBK^{+}10]. pasting [JSTS06]. Patch [BKR17, KSB^{+}13, LLX^{+}01, XLY09, BZL^{+}15, CWL12, DSB^{+}12, FPBCO20, HZW^{+}13, SKY^{+}12, WSLT18]. Patch-based [BKR17, KSB^{+}13, LLX^{+}01, XLY09, CWL12, DSB^{+}12, HZW^{+}13, SKY^{+}12, WSLT18]. patch-level [FPBCO20]. Patches [BCX95, GPSZ11, LCL06, LS08, LSNC09, SKSY08]. Patching [Pet01]. PatchMatch [BSFG09]. PatchNet [HZW^{+}13]. PatchTable [BZL^{+}15]. PATEX [GBLM16]. Path [BBC^{+}23, BYRN17a, CA00, CDY23, CSL^{+}22, CFS^{+}18, DHC^{+}21, FHG^{+}23, FHL^{+}18, GIF^{+}18, HZE^{+}19, JRSS21, KIM^{+}19, LLH^{+}22, NID20, PCS^{+}20, SNM^{+}13, VSJ21, WHY20, XBLZ23, YZN^{+}22, ZYZ21, ZXS^{+}22, ZD20, BPE17, BYRN17b, CRS^{+}16, CHY21, CTE05, FZBR16, FSP^{+}22, HJ11a, HPJ12, HR13, KHD14, KMA^{+}15, KB12, LHZ16, LCX^{+}21, MHM^{+}09, MKD^{+}16, MGJ19, MRNK21, PVG19, SHHD17, SMGH18, SLW22, YLB^{+}22, ZYX^{+}21, ZXH^{+}20, ZXS^{+}21]. path-based [MHM^{+}09]. Path-Space [BBC^{+}23, BYRN17a, YZN^{+}22, SNM^{+}13, ZYZ21, BYRN17b, YLB^{+}22]. path-traced [HR13]. Path-Tracing [CFS^{+}18, KIM^{+}19]. pathfinding [SMC21]. Paths [HA92, NID20, SH23, KGH^{+}14, LYTS13, RHJD18, SC20, SGSS08, VSJ21]. Pattern [BWKS11, HSX^{+}22, LXL^{+}23, YCZ11, BSK^{+}16, DFW20, GBLM16, KL22, LRFH13, POB09, PDF^{+}22, PH15b, RGF^{+}20, SCA02, SNW21, Wan18a, WSH19, YWVW13]. Pattern-aware [BWKS11]. Pattern-Based [HSX^{+}22]. Pattern-guided [YCZ11]. Patterns [FPSG22, HRS^{+}23, KSH23, MDH^{+}23, NPP22, Ros20, WWSP23, AHD15, BGK^{+}13, BSM^{+}07, CLQW08, DEM96, DLL^{+}15, HHV^{+}21, HCE03, HSF07, JTV^{+}15, KS04a, KWL^{+}21, KSS06, KRD^{+}12, KCPS15, LWS^{+}18, LBW^{+}14, LZH^{+}17, MV21, PPW18, PHD^{+}10, ROC^{+}21, RFL^{+}05, SP16, VMW17, YBY^{+}13, ZJL14]. Patternshop [HRS^{+}23]. PAVEL [FPSG22]. PBNS [BME21]. PCBend [FBS^{+}23]. PCEDNet [HLP^{+}22]. PCH [YXH14]. PCU [HAM07]. PDE [UBW99]. PDEs [SSJC22, SMGC23, SHG^{+}22]. PDF [HSB^{+}12]. pearlescent [GMG^{+}20]. peeling [LZF^{+}19]. Pen [And83, KNBH12]. pen-and-ink [KNBH12]. Penalty [GA20, TMOT12]. Pendulum [KH17a, KH17b]. Penetration [LMY^{+}22, JTL^{+}12, PZM13, TK14]. Penetration-free [LMY^{+}22]. Penner [CZ23]. People [XSL^{+}22, ASK^{+}05, CGL^{+}08, JMB^{+}14, Lau18, LCD^{+}20b, WKHA18]. per-frame [WHSL11]. per-pixel [BM05]. per-triangle [SOA11]. perceived [HCOB10, YLL^{+}22]. perceiving [HMO12]. Perception [CAD19, DSJA^{+}21, DWX^{+}21, DCT^{+}22, HDS^{+}18, LN22, LABS23, MKMS04, OD01, PLKD18, RBF08, VRC^{+}13, BOD^{+}13, CGZ08, KWK09, KKW20, MBB12, SCW^{+}21, VLD07, WZMM22, ZAJ^{+}15, MLD^{+}08]. Perception-aware [DWX^{+}21, PLKD18]. perception-based [CGZ08]. Perception-driven [HDS^{+}18, LABS23]. Perception-motivated [MKMS04]. perceptions [SN17]. Perceptual [CGMS22, DKD^{+}17a, FRS19, HOKP16, MS05, RP03, SLF^{+}11, SFWG04, TGD04, TD23, TGZ18, UHT17, WTD^{+}22, ZLP^{+}15, DRE^{+}11, DCB^{+}22, DKD^{+}17b, GSCO12, KKW21, LKS15, PLR^{+}16, PHBC21, SMD^{+}15, WAKB09, YI17]. perceptual-based [YI17]. Perceptually [DPF03, HTER04, KO11, ÖG15, SFLM04, SHK^{+}17, GWM^{+}08, KYS^{+}15, KNL^{+}22]. Perceptually-driven [DPF03, KYS^{+}15]. Perceptually-guided [SHK^{+}17]. Perceptually-supported [SFLM04]. Perfect [LH06b, CZ17]. Perforated [ZLW^{+}16]. Perforation [LNLB16]. PERFORM [DKD^{+}17a, DKD^{+}17b]. Performance [CM83, CH05, DXG^{+}23, FJA^{+}14, HXZ^{+}19, HTCH15, IWZL09, MWM23, MHZ^{+}21a, SRX^{+}23, Tsa15, VMKK00, WGT^{+}05, XCZ^{+}18, ZJY^{+}22, ZZC^{+}22, dAST^{+}08, BHB^{+}11, BBB^{+}14, BHPS10, CBZB15, CCGB22, DKD^{+}16, DDF^{+}17, DK99, HFH^{+}17, HCTW11, KKSS18, LHK^{+}20, LTO^{+}15, MJC^{+}08, MBPY^{+}18, MPH^{+}20, PTMD07, SN17, SDO^{+}04, VWB^{+}12, VLD^{+}13, WBLP11, WJV^{+}05, WGP^{+}10, WZC^{+}22, WSVT13, XCLT14, ZBGB19]. Performance-based [IWZL09, WBLP11]. performances [SWTC14, TDL^{+}18, XLS^{+}11, Zho18]. performative [BJS^{+}08]. performed [SP05]. Performing [NN90, WGH21]. Peridynamics [YR23]. Periodic [RLL^{+}06, HHV^{+}21, LWS^{+}18, SMK22, TZCT20]. peripheral [WWH04]. Periphery [TD23]. Permission [ZG02]. Person [ASN^{+}20, KCS14, LMR^{+}15, GRH^{+}12]. Personal [JMAK10]. personalities [ZCL18]. Personality [DKD^{+}17a, SGD21, DKD^{+}17b, SN17]. Personalization [ARMCO23, GAA^{+}23, ZDT^{+}23, TTR^{+}17]. Personalized [GZX^{+}22, GZC^{+}16, NAH^{+}22, WMB21, KIL^{+}16]. Perspective [CPW21, FSGF16, LSC^{+}12, SD02, UZB^{+}23, CAA10, GB08a, HJ11b, KHH^{+}11, LGQ^{+}08, SBK11, VRC^{+}13]. Perspective-aware [FSGF16, LSC^{+}12]. Perturbation [CA00, XZZ18]. pets [LXJ^{+}22]. PH [PEVBC21]. PH-CPF [PEVBC21]. Phace [IKKP17]. Phase [HKS17, LD23, TLZ^{+}24, WRDF13, BB12, CTS^{+}20, FKN17, GSV^{+}14, GXZ^{+}13, Kim10, LMLD22, SMK22, SSJ^{+}14, SXH^{+}21, WCSC22, YCL^{+}17]. Phase-based [WRDF13, FKN17]. phase-change [SSJ^{+}14]. Phase-field [TLZ^{+}24]. Phase-functioned [HKS17]. phases [SZKZ20]. Phasor [GNHM15, TEZ^{+}19]. Phenomena [LL23, BWRB05, BLR^{+}11, HMS05, RNGF03]. phone [CSK^{+}22, WGJ^{+}18]. Phones [WLS^{+}23, AMS03, LSC^{+}22, SLL19]. Phong [BA08, Jam20, VW97]. Photo [HHX^{+}18, HSC^{+}22, KOF14, LHE^{+}07, SSS06, TZJ^{+}22, WZHL23, XZZ^{+}11, YZW^{+}16, ZZL^{+}21, BSP^{+}19, BLDA11, CLY18, CLS^{+}15, CFL^{+}15, CYW^{+}16, CZS^{+}13, GSZ^{+}18, GSC^{+}15, GAL^{+}09, HSGL13, HEH05, JMAK10, KOF13, KNC^{+}08, LBP^{+}12, OF12, SPDF13, SSS^{+}08]. Photo-Finishing [TZJ^{+}22]. Photo-inspired [XZZ^{+}11]. photo-to-caricature [CLY18]. Photo-to-shape [HSC^{+}22]. Photo2clipart [FLB17]. PhotoApp [RTD^{+}21]. photobios [KSSGS11]. photobooth [PCK^{+}08]. photogrammetric [TT09]. photograph [FH04a, FSH^{+}06, KSES14, KNC^{+}08, LDPT17]. Photographic [RSSF02, BPD06, BPB13]. Photographing [AAC^{+}06]. photographs [BKD^{+}08, DS02, DIO^{+}12, GCD^{+}20, HE07, KHFH11, KGFF14, RMD04, RTS^{+}07]. Photography [Les20, TKG^{+}23, AJD^{+}10, ARNL05, BPK^{+}13, BYLR20, CZN10, ED04, GSMD07, HSG^{+}16, HASK17, HK18a, HJM^{+}22, ITM^{+}14, KHKR11, KF09, KS11, LSC^{+}08, LLW^{+}08, MKZ^{+}21, MWBR13, MPN^{+}02, MCE^{+}17, NLGK18, Ng05, PSA^{+}04, RAT06, RAWV08, SCG^{+}05, VRA^{+}07, VWJ^{+}13, XRLF15]. photometric [HLHZ08, MS05, PCK^{+}08, VPB^{+}09b, WGP^{+}10, XBS^{+}19, ZRL^{+}09]. photomontage [ADA^{+}04]. Photon [DJBJ19, GRS^{+}17a, ZXS^{+}22, BJ17, Dee05, GRS^{+}17b, GHV^{+}18, HOJ08, HJ09, HJJ10, HJ11a, JNSJ11, JNT^{+}11, KD13a, KZ11, LLZ^{+}20, LOW18, MM06, QSH^{+}15, SJ13, ZXS^{+}21]. Photon-Driven [ZXS^{+}22]. photonic [HHGH13]. Photons [MSM^{+}23]. Photorealistic [GN06, POB09, Tsa15, XBS^{+}22, KP11a, LMM^{+}22, LCC21, PRFS18, RTD^{+}21, RTF^{+}04]. Photos [AECO15, FSGF16, MBGS15, SGSS08]. photosensing [RBvB^{+}04, RNd^{+}07]. PhotoShape [PRFS18]. PhysCap [SGXT20]. Physical [BSL^{+}16, BKS^{+}12, CSvRV18, HFM^{+}10, KKRK^{+}16, RLR^{+}21, SY21b, YYL^{+}19, BBG^{+}13, LBDF13, MIWI16, PKM^{+}18, SGX^{+}21, SY21a, SSY22, SWK16, WW13]. physical-optics [SSY22]. Physically [HMS05, HESL11, LCT19, NFJ02, SML^{+}12, WLZ^{+}09, WMC11, Wes21, WDR11, XBS^{+}22, YTJR15, ZZZ^{+}22, ZYM^{+}23, ZPYX23, BP08, BME21, FP03, FBH21, GS04, LSGV18, MWRD13, MPP11, ODGK03, PGH^{+}22, RYL13, SHP04, SNM^{+}13, SGXT20, SH08, TK05, UIM12, WC10, WGH20, WGH21]. Physically-accurate [YTJR15]. Physically-Based [ZZZ^{+}22, HMS05, HESL11, LCT19, SML^{+}12, Wes21, WDR11, GS04, MWRD13, SNM^{+}13, TK05]. Physics [BSK^{+}16, BVF17b, CYFW14, CLZ^{+}22, DLK18, EHSN20, GB13, HHC^{+}19, KGBS11, LVY16, LH17a, NBHSB22, WWWZ23, WTGT10, WGH22, XXA^{+}23, YPA^{+}18, YSCL22, ZZLH23, AVF17, CBvdP09, GJ22, HMT^{+}12, IKKP17, JL11a, KIL^{+}16, KPMP^{+}17, LHP05, LH17b, LSZ^{+}22, MMCK14, MTM16, MdLH10, PDZ^{+}18, PALvdP18, PMA^{+}21, SWR^{+}21, YKZ^{+}22, YRPF09, ZZMC13]. Physics-Based [BVF17b, LVY16, LH17a, NBHSB22, WWWZ23, XXA^{+}23, YSCL22, ZZLH23, EHSN20, GB13, HHC^{+}19, WGH22, AVF17, CBvdP09, GJ22, IKKP17, JL11a, KIL^{+}16, LHP05, LH17b, LSZ^{+}22, MdLH10, PALvdP18, PMA^{+}21, SWR^{+}21, YRPF09, ZZMC13]. Physics-driven [BSK^{+}16, YKZ^{+}22]. physics-guided [MTM16]. Physics-Inspired [YPA^{+}18, CYFW14, KGBS11, WTGT10]. Physiological [MIWB02]. PIC [NNC^{+}20]. PiCam [VLD^{+}13]. picker [DK99]. Pictures [KCSG18, Van82, CGZ^{+}05, HDK07]. PIE [TER^{+}20]. piece [AMB^{+}21, NAI^{+}18]. pieceable [LBDA21]. Piecewise [CJM21, DLTW90, LM91, YAB^{+}22, ZLZ^{+}23, DZCJ22, Far89, GOMP98, LT09, LB06, ZFO^{+}22]. Piecewise-polynomial [CJM21]. Piecewise-smooth [YAB^{+}22]. pigment [PRJ^{+}13, SJ21]. pigmentation [DFW20, ROC^{+}21]. Pigmented [HM92]. PiGraphs [SCH^{+}16]. Piko [PTSO15]. pile [HK12]. Piles [HK10b]. Pinlight [MLR^{+}14]. Pipeline [HHD^{+}22, SBSH18, TMM^{+}21, BKKL15, DNB^{+}05, HGF14, KKSS18, MDZ^{+}21, VWRKM13]. pipelined [LTT^{+}20]. Pipelines [LNLB16, HBD^{+}14, MAS^{+}16, PTSO15, RKLC^{+}11, RKAP^{+}12, SFB^{+}09]. Pitching [TAH^{+}04]. Pivotal [RMBCO23]. Pixel [SLL^{+}21a, XLC^{+}23, YZN^{+}22, BHHM20, BM05, HLR^{+}14, KL11, RFS22, SGM12, SCT^{+}15, SaLY^{+}08]. Pixel-Accurate [XLC^{+}23]. Pixelization [WCZ^{+}22, HWH^{+}18]. Pixelor [BDM^{+}20]. Pixels [DSJA^{+}21, IWHH20, AW20, WHB^{+}12]. Pixie [OHR14]. Placement [CMS95, HK12, XCF^{+}13]. placements [GJWW15]. placing [BLA12]. plain [ACXG09]. plain-weaving [ACXG09]. Plan [HNH19]. Planar [CWKBC13, EPO91, JWT^{+}23, JHR^{+}15, SG01, VVHSH22, WX91, ZAB21, ZBJ^{+}23, ZPBK17, vW84, ASP07, FDBH22, GMP09, HF06, HKAK14, KSH10, LXW^{+}11, MKZ^{+}21, MSM11, MLB16, NCVMO05, PEVBC21, PSG^{+}06, PL14]. planar-reflective [PSG^{+}06]. planar-rod [MLB16]. Plane [BS88, Pag98, CW15, HB21, JX96, LKF12, NAB^{+}15]. Planes [JCY23, SG17, MMBM15]. PlanIT [WLW^{+}19]. planner [SHU^{+}16]. Planning [CLS03, LLH^{+}22, WKMH^{+}23, BBR^{+}21, EAPL06, FZBR16, LLKP11, LYvdPG12, LCX^{+}21, MdLH10, NMD^{+}17, SMGH18, WLW^{+}19, WLY20, ZYX^{+}21, ZXH^{+}20]. Plans [ZWZ^{+}22, MCSK^{+}17]. plant [MHS^{+}19b, QTZ^{+}06, SSBD03, WWD^{+}05]. plants [Che13, ZB13]. plasma [PGK^{+}22]. Plastic [PSK^{+}12, WMB21, JTSB16, MCS15]. Plastics [QLY^{+}23]. plate [FSH11a]. plateau [POT17]. plates [BDW13, GMB17]. platform [AJD^{+}10, SARW^{+}15]. platforms [GM05, LMAS16]. plausible [CDSHD13, DCD15, MHM^{+}09, SGXT20]. playback [KC19]. player [SHK^{+}14, WAH^{+}10, WGH21]. Players [ZSAF21]. Playful [SLD17]. pleasing [GSH18]. plethysmography [VCA^{+}22]. Plotting [And83]. plush [MI07]. Plushie [MI07]. plushies [BCC17]. ply [MGZJ20]. ply-based [MGZJ20]. PML [SKM10]. PML-based [SKM10]. pneumatic [MZL^{+}17]. Pocket [RWS^{+}11]. Pockets [HA92]. Point [AA06, AML18, CB14, CMS95, Erl18, HLP^{+}22, HZC^{+}22, HRS^{+}23, Jan91, KLR^{+}22, LXSW23, MDK^{+}16, MHGCO21, NON85, Özt16, PKG06, QRL^{+}23, RHW94, TFD^{+}18, WX91, WSL^{+}19, WZX^{+}23, Wan23, WS85, XDW^{+}23, YSB^{+}15, ZHWW12, ZXS^{+}23, AHD15, ANHD17, AA09, AK04, ASGCO10, BSD09, Che13, CKMR^{+}21, CLSA20, DVS03, DBD16, EKA84, FLGJ19, FQL^{+}20, Fat11, FGW^{+}21, FCOAS03, GTJS17, GWW^{+}18, GAF^{+}10, GG07, GHF^{+}18, HRV^{+}18, HFG^{+}18, HLZ^{+}09, HWG^{+}13, HWCO^{+}13, HCJ19, JWJ^{+}14, KTB07, KTT13, KL22, LLJ^{+}23, LdPS84, LGB^{+}21, LYO^{+}10, MLR^{+}14, MHZ^{+}21b, ÖG12, PKKG03, RFS22, SSC^{+}13, SNZ^{+}21, TZCO09, WPL06, WQS^{+}20, WNEH22, WFL^{+}19, YC21, YHZ^{+}14, YHCOZ18, ZPKG02, MA07]. Point-based [PKG06, JWJ^{+}14, LdPS84, ZPKG02]. Point-Feature [CMS95]. point-location [EKA84]. Point-sampled [AA06, PKKG03]. point-set [AK04]. Point-Visible [WS85]. Points [Day90, FCK22, War92, AMCO08, BWG03, BJ17, CADS09, CSPF12, Gos00, HWW^{+}22, JNSJ11, KGH^{+}14, STZ14, WHG^{+}15, XMZ^{+}14, ZK13]. Pointshop [ZPKG02]. Pointwise [CPAB22]. Poisson [BWWM10, CK11, DH06, EDP^{+}11, GM09, HWW^{+}22, JCW09a, KH13, PGB03, SJ22a, SJTS04, Wei08, WSL^{+}14, YW13, YZX^{+}04, YIC^{+}14]. Poisson-Based [YIC^{+}14, YZX^{+}04]. Poisson-disk [DH06, EDP^{+}11, GM09, YW13]. Poisson-guided [WSL^{+}14]. Poking [CZB23]. Polar [Sei93, KP07, MP09c, SV19]. Polarimetric [BH21, BJTK18, HJM^{+}22]. Polarization [LWH^{+}11, NKKJ23, RRFG17, MRK^{+}13]. Polarization-Multiplexing [NKKJ23]. polarized [GCP^{+}10, GFT^{+}11]. policies [CBvdP09]. Policy [Kro82, XXA^{+}23]. Pollution [DII23]. Poly [CPW^{+}23, SDG^{+}19]. Poly-Spline [SDG^{+}19]. Polycube [HJS^{+}14, FXBH16, LVS^{+}13, THCM04]. PolyCube-Maps [THCM04]. PolyCut [LVS^{+}13]. PolyDepth [JTL^{+}12]. polydisperse [MPG^{+}16]. Polygon [BYG96, Dun83, Mai92, SG82, WS85, BPK05, IG03, SOS04]. Polygon-Filling [Dun83]. Polygonal [XWD^{+}22, ACXG09, AW11, ACSD^{+}03, BF08, CGG^{+}04, DP13, HDHN16, Ju04, Pet21, PNDN12, POC05, TLK09, VMW17, WR18]. polygonal-light [HDHN16]. Polygons [CCW93, FM84, TM84, BHSH^{+}22, GH98, HF06, SW85]. Polyhedra [Pet95, Wil92, BDD11, BHSH^{+}22, Hub96, PR97b]. Polyhedral [JTV^{+}15, MHSL18, Nas87, DA21, GJTP17, GSC21b, KGB^{+}09, Mir98, PKD^{+}19, TSG^{+}14]. polyhedron [TBBC^{+}22]. Polylines [RS14b]. Polynomial [PGG^{+}23, SB95, BAERD08, CJM21, FGG^{+}17, GOMP98, MJC^{+}08, MMMG16, SR97, SR00, SSW^{+}13]. Polynomials [Kla91b, LM97]. polyomino [LFL09]. polyominoes [Ost07]. PolyStokes [PGG^{+}23]. polytopes [BLTD16, KDH22]. Polyvector [BS19, DVPSH15, PNCB21]. Pop [SSY^{+}04, XZM^{+}18, HEH05, LJGH11]. Pop-up [SSY^{+}04, XZM^{+}18, HEH05]. pop-ups [LJGH11]. PopStage [LYC^{+}22]. populated [LHZ^{+}18]. Popup [LSH^{+}10]. Porous [LAD08, RXL21, TGK^{+}17]. portable [HJM^{+}22]. portal [GWN^{+}03]. Portrait [CLX^{+}22, SHS^{+}17, SHS^{+}18, SWS^{+}22, TCS^{+}23, YJLL22, YNK^{+}22, BSM^{+}13, CWW^{+}12, CLS^{+}15, FAC11, FSGF16, LD21, MYC^{+}22, SBT^{+}19, TER^{+}20, TZT^{+}18, WYL^{+}20, WYXJ21]. portraits [AECOKC17, KS16, KGT^{+}18, LVG^{+}13, LCC21, MDKD16, PEL^{+}21, RTD^{+}21, SED16, SPB^{+}14, SLL19, SLL^{+}21b, YNS19, ZAE^{+}14]. Pose [ALY^{+}21, EM96, TSLP14, XB16, AZB09, ACCO05, BME21, BB22, GWP^{+}19, HKA^{+}18, HOM15, KAL^{+}17, Liu09, LHR^{+}21, MSS^{+}17, MDB^{+}19, NOP^{+}18, TBC^{+}16, YZX21]. pose-free [AZB09]. pose-guided [ALY^{+}21]. Pose-space [XB16]. poser [HKA^{+}18, LCXS09]. poses [ZBYX19]. posing [BVS16, GCR13]. Position [GHZ18, MM13, PTV^{+}17, RMD12, WJF^{+}22, XRW^{+}22, YHMR16, ATM^{+}17, LSL^{+}18, Wan15]. Position-Based [XRW^{+}22, PTV^{+}17, Wan15]. Position-correcting [RMD12]. Position-free [GHZ18, WJF^{+}22]. Position-normal [YHMR16]. Positioning [Bae82, ZB94]. positions [NRDR05]. Possible [NI22, NI24, ZXZL23, AVR^{+}22, IMH05, ZCD^{+}16]. Post [HHX^{+}18, PTMD07, BGKS17, ITM^{+}14]. post-capture [BGKS17, ITM^{+}14].

Post-Processing [HHX^{+}18]. Post-production [PTMD07]. Posterior [LTH^{+}23]. Postprocessing [CFP^{+}21]. potential [CS00, LYK^{+}21, LFS^{+}20, LKJ21, OHR14]. Potentially [KL23, VKW^{+}23]. Power [AGL^{+}17, BLTD16, DCT^{+}22, FF88, QLY^{+}23, WWWG22, dGWH^{+}15, MMT18, PEVBC21, QLDJ22, SR97, SR00, WYM^{+}16, XLC^{+}16]. PPPM [ZB14]. Practical [AWL13, CLT^{+}22, DPVA23, EDR11, GHP^{+}08, GRB^{+}18, LWA^{+}12, LYL^{+}16, LJJ^{+}18, LSVT15, MC92, NLGK18, RSL16, RZK11, SJJ12, SJ21, SCJ^{+}23, TG17a, TG17b, VAV^{+}07, XWH^{+}23, ZZW^{+}22b, AB20, BB17, CAJ09, EKA84, FTP16, JSB^{+}10, KySK10, MSOC^{+}19, MGZJ20, SBdDJ13, SSY22, SRNN05, TWAD09, XCM^{+}14, YJR17, ZG02, ZRL^{+}09]. Practice [ABGL21]. Prager [KGP^{+}16]. Prakash [RNd^{+}07]. pre [HMAM09, YZL^{+}22]. pre-captured [YZL^{+}22]. pre-tessellation [HMAM09]. precise [NRDR05, TBC^{+}16]. Precision [SFB92, US24, TVLF20, Wan18a]. precomputation [KKN^{+}13, WJ19, YLX^{+}15]. Precomputed [CZJ12, JBP06, KAMJ05, RSM^{+}10a, SKS02, XIM18, ZHL^{+}05, BAERD08, Leh07, RS14a, RS18, SL17, SKOA14, SHHS03, SLS05, TS06, ZJ10]. Precomputing [JF03]. Preconditioner [CZY17b, CZY17a, WWW22]. preconditioners [KS11]. Preconditioning [PHM^{+}23, CSHD21, KFS13, Sze06]. predict [GSY^{+}17, HLV^{+}17c, SHZ^{+}20]. predictable [RAR^{+}21]. Predicting [BWDL21, DWMG15, WGY^{+}18, BVM^{+}17, BAC^{+}06, KMM^{+}17a]. Prediction [SSII18a, WBF^{+}17a, ATM^{+}17, GLZ^{+}21, KKDK12, LPL^{+}18, VRM^{+}18, WBF^{+}17b, WLP16, YSW^{+}20]. predictions [MKRH11, MIGYM15]. Predictive [EHSN20, HYZ^{+}18, SP09, KSHG18, ZJMB12]. Predictive-corrective [SP09].

Predictor [VMKK00, MDC^{+}21].

Predictors [KL17a, KL17b]. preference [SLF^{+}11, ZLP^{+}15]. Prefiltering [BSK23, DLW^{+}22, WZK^{+}23, GT96, WZYR19]. prescribed [SZC^{+}22]. prescriptive [MSOC^{+}19]. Presence [RO94, MIWB02, SSC10]. Present [EST^{+}20]. Presentation [MMHP23, NAB^{+}15]. Presentations [Cas91, Mac86]. presenting [FNvD82]. Preservation [WWWG22, LCORL07]. Preserving [AVB^{+}23, ABO16, NKJF09, SK16, WX91, ALY^{+}21, BHY15, BSBC12, CAA09, CZZT12, DBWG15, DHB17, ETK^{+}07, FH07, FFLS08, FKY^{+}10, GOTG05, HK10a, HKT10, JDD03, KEE13, LHM09, LCOZ^{+}11, LGJA09, LKWS16, MSW^{+}09, MCP^{+}09, NSACO05, OL03, QPWH08, RPWO18, SLS^{+}16, SSD09b, TWZ20, WWA^{+}16, WZYR19, ZNT18]. presorted [CSN^{+}12]. Pressure [BGI^{+}18, GPB^{+}19, LBB17a, TB20, ZSZ^{+}14]. pressure-viscosity-contact [TB20]. preview [RKKS^{+}07]. Primal [ORK12]. Primal-dual [ORK12]. primaries [SMH^{+}11]. Primary [CJM21, HPK^{+}17]. Primary-space [CJM21]. Primitive [PK22, FTD21]. Primitives [GS85, JGMR23, LHK^{+}20, LWC^{+}11, LSS^{+}21, MESK22, SNCH08]. Principal [Wu92, GI04, SHHS03, TISM16, XSZB15]. Print [NKS^{+}23, OCNG21, OGN^{+}23, NIR^{+}21, UPSW16]. print-wind [UPSW16]. printable [KSS^{+}15, LBRM12, MTN^{+}15, SVB^{+}12, YCL^{+}15]. Printed [AJS20, JHS^{+}23, PRM14, ZLW^{+}16, LSZ^{+}14, MLYZ19, TDG18]. printer [LDS02, WPGM16]. printers [ERP^{+}19]. Printing [BAU15, DTPG12, LR90, LR91, MSS^{+}12, MAG^{+}09, NKS^{+}23, PLMR17, RB23, SCB88, UTB^{+}19, WPGM16, BVF^{+}17a, BATU18, CCA^{+}12, CZL^{+}15b, DWW^{+}18, DHL14, ESZ^{+}17, ICG17, PFB^{+}20, SBR^{+}15, SBK^{+}18, SARW^{+}15, SRB^{+}19, MBU22, WWY^{+}13, ZYZZ15, ZLP^{+}15, ZBW^{+}20]. Printone [UPSW16]. prints [CLD^{+}13, PH15a, THKM13, TTZ^{+}20]. Prior [CPW21, JCFG23, NAH^{+}22, BSP^{+}19, CCWL18, CJN^{+}17, MYWI15, WLW^{+}19]. Priors [VR94, ZZLH23, ISSI16, LCXS09, PMA^{+}21, SKAG15, WSCR18, ZZI^{+}17, ZXC^{+}18]. prism [BKGK17]. Proactive [YSL^{+}14, XHS^{+}15]. Probabilistic [CKGK11, FW16, LRFH13, RHW94, CLS03, DCB^{+}22, HAB20, KCKK12, KZ11, LCK^{+}14, NKAS08, SLW22, VPHB^{+}21, WLP16]. Probability [DLP^{+}23, DLC^{+}15]. Probability-Modeled [DLP^{+}23]. probable [DTB06]. probe [BBO91, ORK12, RLP^{+}20]. Probes [GZS^{+}22, SL17]. probing [BH21, OHX^{+}14]. problem [DIO^{+}12, HPB07, LW16, OP11, XW09, YWH13]. Problems [FM84, Gol84, OF01, SCJ^{+}23, CSHD21, DML17, GITH14, MSW14, MLT17, PKHK15, SPKS16]. Procedural [BSW13, GDAB^{+}17a, GJB^{+}20, HHD^{+}22, LLDD09, LIY^{+}22, LSM23, MWC^{+}23, MDL16, Mer23, MWH^{+}06, NPA^{+}22, SW14, TEZ^{+}19, WOD09, BDK^{+}16, BWS10, BHN07, CH02, CEW^{+}08, CDM^{+}02, EVC^{+}15, GDAB^{+}17b, GGG^{+}13, GSV^{+}14, GHS^{+}22, GSLM^{+}08, HSS98, KW11, LD05, LWW08, MZWV07, NSCL08, NGDA^{+}16, RMGH15, SP16, SM15, SLH^{+}20, TLL^{+}11, VGDA^{+}12, WYD^{+}14, ZLB16b]. Procedurally [Kaj83]. procedures [MCS15]. Process [MOR^{+}18]. Processes [Özt16, ZZW^{+}22a, IAF09]. Processing [BBG24, DSJA^{+}21, HHX^{+}18, PCS^{+}23b, SGWJ18, TMM^{+}21, XWC^{+}16, dGMMD14, CPD07, CKPS17, CGZ08, CK11, FLJK21, FMR20, GO11, GSC21a, HBD^{+}14, HDD^{+}16, HST^{+}14, HDA17, HHN^{+}02, KSH10, KH10, KG08, KWB^{+}15, LGA^{+}18, LHLK10, LTJ18, MZPS21, MASS15, MAS^{+}16, MMTD07, OEE^{+}18, PHK11, PKCH18, RKAP^{+}12, RH04, RVAL09, SR00, SDP^{+}18, SLMR14, STP12, TWBO03, TYY^{+}19, WRDF13, WFL^{+}15, WSS05, YW13, Zhu18a, dGDMD16]. Processor [KS95]. processors [CTH^{+}14]. Product [SG17, BB15, NRH04, PBW19, SM06, SR09]. Production [FHL^{+}18, GIF^{+}18, Pha18, ZZZ^{+}22, ZCS^{+}22, LF02, LSD^{+}22, LZT^{+}19, MCW^{+}21, PTMD07, SSBL^{+}22, TKTS11]. Production-Ready [ZCS^{+}22]. products [CJAMJ05]. Professional [ZSAF21, LVS^{+}16]. profile [DSF22]. profiles [KWB^{+}13]. Program [NN90, SFC^{+}23, Spr82]. Programmable [GTDS10, LLW^{+}08, LHVT17a, LHVT17b, SSBG10, HAM07, HGG^{+}11, HMG03, KLPCP18, LB05, NJS^{+}11, PTSO15, PBMH02, VAZH^{+}09, VWRKM13, WSS05]. programmatic [WPL^{+}21]. Programming [BWL^{+}23, BK16, GF82, HGM14, KSH23, PPV95, Wu92, ZB94, BLPW14, HZG08, HKG11, KABL14, LGA^{+}18, MGAK03, NWYM19, SAMWL11, SFB^{+}09]. Programs [JGMR23, AMA^{+}19, HZG09, JBX^{+}20, JCG^{+}21, RMGH15, YBAF22]. Progressive [DKHS14, FCOAS03, GD02, HOJ08, HLC^{+}19, JNT^{+}11, KZ11, LDS03, LYNF18, NCB23, SJZP19, SJ13, VMKK00, YSQS08, ZDF^{+}22, ZQL^{+}23, ZDF^{+}23, HJ09, HJJ10, KD13a, LLKC21, LJH13b, LLZ^{+}20, PK05]. progressively [ZZV^{+}03]. progressively-variant [ZZV^{+}03]. Project [LGA^{+}21, Ano10, ZIT^{+}19]. Projected [And82, YZX^{+}18]. Projection [DGH16, ZN06, ARNL05, DLL^{+}18, GWGB10, HWR14, HSHF10, JBM^{+}17, JTL^{+}12, JSZP20, KYS^{+}15, LZF10, LCOLTE07, MS05, MWI18, ME05, PMA^{+}14, SCT^{+}15, SSW^{+}13, ZBG15a]. projection-based [MS05]. projections [AYL^{+}12, BML^{+}14, CAA09, KSJP08, MWBR13, MHR^{+}16, PBC^{+}22, SBK11]. Projective [BML^{+}14, DWM^{+}22, LJBBD20, Pat85, WGW^{+}13, ZRJ23, ZLW^{+}16, BEH18, KUJH21, LMY^{+}22, LLKC21, Wan15, Pat87]. Projectively [NY94]. projector [BBG^{+}13]. projector-based [BBG^{+}13]. projectors [RvBB^{+}03, RBvB^{+}04, SGM12]. prolongation [LZBCJ21]. Prompt [ZDT^{+}23]. Promptable [MSL^{+}24]. Proof [FAER21]. Propagation [SM17a, AP08, ACSM12, CRS^{+}16, CRG^{+}20, CZZT12, CGP^{+}21, Erl07, Fat09b, GJWW14, HRL15, Liu18, MRA^{+}13, MHZ^{+}21b, QHY^{+}16, RSM^{+}10a, RS14a, RS18, SMM14, SM17b, SMC21, SYJS05, VWJ^{+}13, XLJ^{+}09, YMR^{+}13, ZRSM18]. properties [AHD15, FCJ07, NGD^{+}06, ODJ04, SZG^{+}13, WSM11, ZKBT17]. prosody [LTK09]. prosody-driven [LTK09]. ProSpect [ZDT^{+}23]. Protected [KTL^{+}04]. prototype [AWGB04]. prototypes [KLY^{+}14, YPB16]. Provably [PL14, DML17, YLJ18]. Prox [LWF^{+}23]. proxies [CB17, JSMH12, TYY^{+}19, ZCC^{+}12]. Proximal [LWF^{+}23, HDN^{+}16, HDN^{+}16]. proximity [SGG^{+}06]. Proxy [HXM^{+}18, LSM23, XLYJ23, KGL16, MSM11]. Proxy-Free [LSM23]. Proxy-to-Image [HXM^{+}18]. pruning [TMRL14]. psychophysical [AFR^{+}07, GRG04]. psychophysically [FCGH08]. psychophysics [SSC10]. Pteromys [UKSI14]. Pupil [CBS^{+}22, JBM^{+}17, POB09]. Pupil-Aware [CBS^{+}22]. pupil-tracked [JBM^{+}17]. PuppetMaster [ZPBC19]. Puppetry [SLGS01, BJS^{+}08, SZT^{+}07]. pure [WHY20]. purification [LSQ^{+}15]. Purpose [Lev84, PBD^{+}10]. Push [HMO12, LLK^{+}15]. Push-recovery [LLK^{+}15]. Pushdown [Ols84]. Pushing [BAU15]. PushPull [LWM14]. Putting [BW13]. Puzzle [JPL22, LFL09]. puzzles [CWSB22, LKvK^{+}14, SFCO12, SZ15, XLF^{+}11]. Pyramid [KSH^{+}16, ZJNZ23, PHK11]. pyramidal [CLF^{+}18, HLZCO14]. pyramids [FFL11].

Q [FTD21, LWS^{+}15]. Q-MAT [LWS^{+}15]. Q-zip [FTD21]. Qasistatics [ZDF^{+}23]. QEx [EBCK13]. QR [CCLM13]. Quad [HSV^{+}22, IRWP23, ULP^{+}15, BCE^{+}13, CBK12, CK14b, EBCK13, ECBK14, ESCK16, FBH^{+}10, FTD21, JRPW20, LCBK19, PPW18, PNA^{+}21, SW05, SPGT18, TPSHSH13, TPP^{+}11, TMB18]. Quad-Based [HSV^{+}22, JRPW20]. quad-dominant [SPGT18]. quad-fragment [FBH^{+}10]. quad-remeshing [PNA^{+}21]. quadrangulated [SZC^{+}22]. Quadrangulation [FBT^{+}18, LHJ^{+}14, ACBCO17, BWSS12, BZK09, DBG^{+}06, HZM^{+}08, MTP^{+}15, ZHLB10]. quadrangulations [PBJW14, VPR19]. Quadratic [BC14, ERT14, LWS^{+}15, BHSH^{+}22, KGL16]. Quadrature [GT96, FQL^{+}20]. Quadric [CGM91, FNO89, GZ05, Mil87, TGB13]. Quadric-based [GZ05]. Quadric-Surface [FNO89]. Quadrics [SJ94]. Quadrilateral [DSSC08, MWM23, VVHSH22, DM13, LXW^{+}11, PZKW11]. quadrotor [GSH18, JRT^{+}15, RH16, XYH^{+}18]. quadruped [LSCC20, ZSKS18]. quadrupeds [CKJ^{+}11]. QuadStream [HSV^{+}22]. Quadtrees [LS00, Aga07, ABJN85, BFK^{+}16, SW85]. Qualitative [HSS^{+}13]. Quality [AAPS16, KKN^{+}22, KKDK12, NDD^{+}23, WSL^{+}14, ZCP^{+}23, AAPS17, AMMS08, AČMS10, BWG03, BGAM12, BBB^{+}10a, BHB^{+}11, BBN^{+}14, CHM^{+}12, CBK12, CS00, CLS^{+}15, CWZ^{+}21a, CTW09, CLW^{+}14, CJN^{+}17, CCS^{+}15, Csé19, DDD^{+}14, GBAM11, GT96, HRH^{+}13, LWC^{+}13, LCX^{+}21, MC21, MKRH11, MHP^{+}19, SHD^{+}18, SJA08, SFWG04, WAC07, WHB^{+}12, WL21, ZJ11, ZF03, ZKU^{+}04]. Quality-driven [WSL^{+}14]. QuanTaichi [HLY^{+}21]. quantifying [RPE^{+}05]. Quantitative [CM83, TGZ18]. Quantization [HWB23, HZC^{+}22, Wu92, BBC22, CCOST05, HRV97, LSZ^{+}22, LCBK19, Xia97]. Quantized [CBK15, DI11, HLY^{+}21]. quantized-diffusion [DI11]. quantum [BSW02]. quartet [HSS^{+}13]. Quartic [Joe90b, Pet89]. Quasi [LBK17a, LBK17b, TWL^{+}05, WS21, WGS23, ZBK18]. quasi-harmonic [WS21, WGS23]. quasi-homogeneous [TWL^{+}05]. Quasi-newton [LBK17b, LBK17a, ZBK18]. quasiconformal [LKF12]. quasistatic [SWR^{+}21]. Quasistatics [ZDF^{+}22]. Quaternion [HFK94, KCŽO08]. Queries [MPB17a, BZL^{+}15, HJ11b, MPB17b, SJ22b]. Querying [ZXS^{+}23]. Quest [GKK^{+}24]. QuickETC2 [Nah20]. quilting [ZHW^{+}06]. quilts [LBDA21].

R2E2 [FSP^{+}22]. Rack [TE82]. radar [LGK^{+}16, VCA^{+}22]. Radial [WLH^{+}13, KN06, TS06]. Radiance [BDT99, DXG^{+}23, GLC^{+}23, HWZ^{+}20, IRG^{+}23, JDZJ08, JCFG23, KKLD23, KQG^{+}23, MJJG18, PHM^{+}23, RYW^{+}22, RSV^{+}23, RMP^{+}23, TCS^{+}23, WSND^{+}23, WXZ^{+}23, ZWS^{+}24, HW12, JNSJ11, KMM^{+}17a, KAMJ05, MRNK21, PSH^{+}21, RWG^{+}13, SL17, SKS02, SLSS03, SHHS03, SLS05, TS06, WKR99, WVBR^{+}21, LAM^{+}11]. radiance-predicting [KMM^{+}17a]. radiant [SSBD03]. Radiative [BBC^{+}23, ABW14, BRM^{+}18, JAM^{+}10, JAG18, PCS^{+}20, ZRB14]. Radiosity [ACP^{+}01, NN95, RT90, DDP99, HCZ21]. RAID [GMW16]. rain [GN06, LCT19]. Rainbow [XIAP^{+}17]. rainbows [SML^{+}12]. raising [CLS85]. Random [HZE^{+}19, NH08, PM95, VSW^{+}23, AMA^{+}19, CNX^{+}08, GSV^{+}14, HAK^{+}22, KCYW13, LT20, LSK^{+}06, SD12, ZZZX21]. Random-Access [VSW^{+}23, NH08, KCYW13, LSK^{+}06]. random-phase [GSV^{+}14]. Randomized [GF08, BSFG09]. Range [BMBRD24, SB95, WSP^{+}23, WS17a, ACP02, ACP03, AMMS08, BI08, CZ11, DD02b, FKI^{+}14, FLW02, HSG^{+}16, HFI^{+}08, KSB^{+}13, KR17, KUWS03, LEPM22, LSA05, MRK^{+}13, MKMS04, MEMS06, MCHAM06, PMOR10, PTSZ11, RAI06, SHS^{+}04, SJ22b, TAHL07, Van06, WLHR11, WS17b, ZJY^{+}21, BZL^{+}17, LCTS05]. Rank [SW18, LHKR10, MK16]. Rank-Constrained [SW18]. ranking [WLO^{+}14]. Rapid [KLPCP18, LKB22a, LCC^{+}22, RvE93, WWA^{+}16, HFTF15, HFF16, JB02, MGDB05, vdHDT^{+}07]. Rapidly [Fol87, TMRL14, ZIT^{+}19]. RAPter [MMBM15]. Raster [Dun83, Lev84, McI92, VN85, WW82]. Rasterization [Hob90, AMS03, FRS19, LAKL11, LHZ16, LLGRK20, PR06]. Rasterizing [Tau94]. Rate [WLF^{+}20, HGF14, HDD^{+}16, KLF^{+}19]. rates [TDMS16]. Ratio [YR23, NSJ14]. Rational [BHN98, Che92, EK98, HB89, JCY23, KLN91, SG17, War92, AB89, BCW17, CADS09, CZXL23, Gal99, Joe89, ZS00]. raw [LEPM22, LRS18]. Ray [BK85, GHCC88, HYS23, KGB^{+}09, Kaj83, Lev90, LSCO03, NKK^{+}14, PP94, PBMH02, RLU95, SLM^{+}17a, VKJ^{+}17, WIK^{+}06, WBS07, WHG84, vW84, BDT99, BAM14, DMB^{+}14, DHW^{+}11, EDR11, HJW^{+}08, HQL^{+}10, HZ11, IYYI14, KDPN21, LAA^{+}05, LADL18, MBK^{+}10, Mor11, MHC^{+}16, NPP^{+}11, NNDJ12, PFHA10, PBD^{+}10, RAWV08, RSH05b, SLM^{+}17b, SWF^{+}21, SKC^{+}14, TOG22, WWB^{+}14, WS99, WSS05, YMR^{+}13, ZRL^{+}08, BK87]. ray-marching [KDPN21]. ray-traced [EDR11, PFHA10]. Ray-Tracing [NKK^{+}14, Mor11]. RayCore [NKK^{+}14]. Razor [DHW^{+}11]. RBF [NNC^{+}20]. RBF-FD [NNC^{+}20]. Re [GXSD23, JSSH15, Pav90, WC21a, WP90, ZCS^{+}22, BHW16, DNZ^{+}17a, GDC15, GPW^{+}17, KD13b, MBPY^{+}18, NKAS08]. Re-Aging [ZCS^{+}22]. Re-Composition [GXSD23]. re-creation [NKAS08]. Re-Editing [JSSH15]. Re-examination [WC21a]. re-integration [DNZ^{+}17a]. re-meshing [GPW^{+}17]. re-parameterization [GDC15]. re-rendering [MBPY^{+}18]. re-simulation [BHW16, KD13b]. reaching [SHX^{+}22]. reaching-and-grasping [SHX^{+}22]. reaction [DFW20, DCB^{+}22, RCLM19]. Ready [CZB23, LLF^{+}20, ZCS^{+}22, ZB13]. Real [ASA^{+}09, ADM^{+}08, BHN98, BJ05, BP08, BZ11, BAC^{+}23, BAOR06, CBZB15, CWW^{+}16, CKH18, CAD^{+}21, CPD07, CM11, CHTK24, DNZ^{+}17b, DLK18, DYN03, DFYL19, DWS^{+}23, EMU15, FKY08, GXY^{+}17a, GXY^{+}17b, GZS^{+}22, HXZ^{+}19, HLX^{+}21, HV04, HRE^{+}08, HDHN16, JTL^{+}12, JKT^{+}15, KKLD23, KSZ^{+}15, KKN^{+}22, KIM^{+}19, LH16, LES10, LTK09, LLX^{+}01, LCH^{+}21, LFTC13, LHLK10, LBK17a, LZH^{+}20, LB06, MP08, MDB^{+}19, MNV^{+}21, MCK13, MRNK21, NMD^{+}17, NZIS13, PZ08, PO08, POC05, PYA^{+}24, RSV^{+}23, RWS^{+}06, RMBCO23, RHHL02, SBSH18, SCT^{+}15, SL17, SSII18b, TDL^{+}18, TWH^{+}22, TZN^{+}15, TZS^{+}18, TSLP14, TCS^{+}23, VRBC18, VTSSH15, WWD^{+}05, WPP07, WP09b, WYM^{+}16, WXLY17, WOG06, WZN^{+}14, XLC^{+}23, YNK^{+}22, YZH^{+}23, ZXTZ15, ZZI^{+}17, ZZT^{+}21, ZHHZ20, ZZZ^{+}23, ZHWG08, ZRL^{+}08, ZNI^{+}14, ALY08, BK04, CWLZ13, CHZ14, CCWL18, CH02, CBI13, CT05, CHP07, CNR08, DNZ^{+}17a, DvGNK99]. real [DLL^{+}18, DHOO05, DFZ^{+}17, DKD^{+}16, DDF^{+}17, FYK10, GO12, GCB^{+}17, GB08b, HFF18, HMO12, HSW^{+}17, HKA^{+}18, HESL11, JBPS11, JP02, KNS^{+}09, KUJH21, KCODL06, KRF^{+}18, KAMJ05, LZC11, LXC^{+}15, LBK17b, LCX^{+}21, LNWB03, LCC21, MMCK14, MHM^{+}17, MBPY^{+}18, MP04, MBB12, MSS^{+}17, NSX^{+}18, NOP^{+}18, PRWH^{+}18, PCK^{+}08, RSM^{+}10a, RTK^{+}15, RJ07, SZT^{+}08, SGXT20, SKS02, SRNN05, TZT^{+}18, TPT16, TLP06, TS12, VBG^{+}13, WKF^{+}21, WAO^{+}09, WJBK15, WSJP17, WMB^{+}20, XUC^{+}14, YZX21, ZBYX19, dASTH10]. Real-Time [BJ05, CHTK24, DNZ^{+}17b, DLK18, DWS^{+}23, GXY^{+}17a, GZS^{+}22, HXZ^{+}19, KKLD23, KIM^{+}19, LBK17a, MNV^{+}21, TZS^{+}18, TSLP14, TCS^{+}23, VTSSH15, XLC^{+}23, ZXTZ15, ZZT^{+}21, ASA^{+}09, ADM^{+}08, BP08, BZ11, BAOR06, CBZB15, CWW^{+}16, CKH18, CAD^{+}21, CPD07, CM11, DYN03, EMU15, FKY08, GXY^{+}17b, HLX^{+}21, HV04, HRE^{+}08, HDHN16, JTL^{+}12, JKT^{+}15, LH16, LES10, LTK09, LLX^{+}01, LCH^{+}21, LFTC13, LHLK10, LZH^{+}20, LB06, MP08, MDB^{+}19, MRNK21, NMD^{+}17, NZIS13, PZ08, PO08, POC05, PYA^{+}24, RSV^{+}23, RWS^{+}06, RHHL02, SCT^{+}15, SL17, SSII18b, TDL^{+}18, TWH^{+}22, TZN^{+}15, VRBC18, WWD^{+}05, WPP07, WP09b, WYM^{+}16, WXLY17, WOG06, WZN^{+}14, YZH^{+}23, ZZI^{+}17, ZHHZ20, ZZZ^{+}23, ZHWG08, ZRL^{+}08, ZNI^{+}14, BK04, CWLZ13, CHZ14, CCWL18, CH02, CBI13, CT05, CHP07, DNZ^{+}17a, DLL^{+}18, DHOO05, DKD^{+}16, DDF^{+}17, FYK10, GO12, GCB^{+}17, HFF18, HSW^{+}17, HESL11, JBPS11, KNS^{+}09, KUJH21]. real-time [KCODL06, KRF^{+}18, KAMJ05, LZC11, LXC^{+}15, LBK17b, LCX^{+}21, LCC21, MMCK14, MHM^{+}17, MBPY^{+}18, MP04, MSS^{+}17, NSX^{+}18, NOP^{+}18, RSM^{+}10a, RTK^{+}15, RJ07, SKS02, TZT^{+}18, TPT16, TLP06, TS12, VBG^{+}13, WKF^{+}21, WAO^{+}09, WJBK15, WSJP17, WMB^{+}20, XUC^{+}14, YZX21, ZBYX19, dASTH10]. Real-World [SBSH18, ALY08, DvGNK99]. RealBrush [LBDF13]. realism [CLS^{+}17, XADR12]. Realistic [CLT^{+}22, HM92, SLST14, SBK11, WW08, cWP03, CXGS02, CPWAP08, DFW20, DPF03, HRZ^{+}13, JWDL19, KNL^{+}22, RPC^{+}10, SHP04, SQRH^{+}16, WC10, WVBR^{+}21, WFS22, ZLB16b, CKX^{+}08]. realistic-looking [RPC^{+}10]. Reality [AS21, DFYL19, DWX^{+}21, DCT^{+}22, FSRS22, JWD^{+}23, KAW20, MNV^{+}21, TZS^{+}18, ALK^{+}17, AGB^{+}16, BP12, CKX^{+}08, CGP^{+}21, Did18, HK18b, JBM^{+}17, KJS^{+}19, KDMW17, KKW20, LSL^{+}18, LJM^{+}16, LHLY21, LLHY22, MGDB05, MLR^{+}14, MGK17, OEE^{+}18, PSK^{+}16, RRS19, SMG^{+}05, SSRB^{+}17, SMG^{+}20, SWK16, SPW^{+}18, Wan18b, ZJY^{+}21]. Realizable [ZPYX23]. realization [LJM^{+}16]. reallocation [HSHF10]. really [WKHA18]. Realtime [LYYB13, SLM^{+}23, WSXC16, WBLP11, XWCH15, BWP13, SJP05, WAC07, WZC12, WSS05, ZZMC13]. Reanimating [KHS03]. Reasoning [JWD^{+}23, CKGK11]. Reassembling [HFG^{+}06, BTFN^{+}08]. Reassembly [LLL22]. rebuilding [MMBM15]. recipes [GSC^{+}15]. Reciprocal [SFG^{+}13, CRG^{+}20, SRL^{+}15]. recognition [CWK^{+}20, LMS13, YMJ^{+}21, ZYL^{+}17]. recoloring [CRA11, CFL^{+}15, DLX^{+}21, TEG18]. recommendation [CKP^{+}21]. Recomposition [LGHL23]. Reconciling [SPV^{+}16]. Reconfigurable [SFJ^{+}17, MRK^{+}13]. reconfigurables [GJG16]. reconnection [WP10]. reconstruct [HKA^{+}18, LXR^{+}18, LKK^{+}21]. Reconstructability [LLH^{+}22]. reconstructed [RMBB^{+}13]. Reconstructing [GZX^{+}22, GVWT13, KIL^{+}16, LALD12, SYZ^{+}23, WGL^{+}18, XWD^{+}22, YT13, KL22, LCC^{+}18]. Reconstruction [BL20, CKH18, CXW^{+}23a, CHTK24, DNZ^{+}17b, GZC^{+}16, GXY^{+}17a, GLX^{+}22, HNH19, HSV^{+}22, HCH22, HWZ^{+}20, IH20, KNK^{+}22, KKN^{+}22, KQG^{+}23, KIM^{+}19, LXSW23, LWL^{+}23b, LXL^{+}23, LSSW19, PYA^{+}24, PMKB23, RKB^{+}23, SMR^{+}22, SG82, SJ22a, SHD^{+}14, SAA^{+}21, TWR^{+}23, TZY^{+}23, VMCS15, WBF^{+}17a, WLJ^{+}22, WWX^{+}22, XNZ^{+}22, YSW^{+}23, YLC^{+}20, YZN^{+}22, YSHWSH16, ZXTZ15, ZZT^{+}21, ZXS^{+}22, AKZ^{+}17, ACP03, ASL^{+}17, ASGCO10, BBN^{+}12, BLN^{+}13, BVG11, BBK^{+}15, BNB13, CKS^{+}17, CZ11, CBI13, CHY21, CJN^{+}17, CLZ^{+}22, DNZ^{+}17a, DTB06, DXZ^{+}19, ETH^{+}09, EKD^{+}17, FG14, GZW^{+}16, GKTT13, GD04, GXY^{+}17b, GLP^{+}22, HJW^{+}08, HZW12, HWW^{+}22, HWV^{+}18, HLZ^{+}09, HWK15, HZCJ17, JMK^{+}22, KHS03, KH13, KFWM17, KHL19, KZP^{+}13, KPZK17, KL12, LDK^{+}18, LAC^{+}11, LSR18, LKG^{+}03a, LAGP09, LLZM10, LWC12, LGZ^{+}13, LCOLTE07, LCL^{+}17, LXS^{+}18, LCX^{+}21, LGB^{+}21, LYO^{+}10, LWL^{+}20, McC00, MTM16, MDB^{+}19, MHC^{+}16, NSZ^{+}10, NZIS13]. reconstruction [PZ17, PAAG21, PMA^{+}14, RMD04, RZW^{+}21, RKZ11, SS14, STJ^{+}17, SKY^{+}12, SSZCO10, SLS^{+}07, SAL^{+}08, SL17, SMGH18, SSR20, SGdA^{+}10, TZK^{+}11, TBW^{+}12, TVLF20, WBF^{+}17b, WAO^{+}09, WSL13, WSTS08, WBG^{+}16, WZQ^{+}18, XZZ^{+}14, XHS^{+}15, XZY^{+}17, YHZ^{+}14, ZBYX19, ZYX^{+}21, ZZZX21, ZK13, ZK14, ZXH^{+}20, ZXS^{+}21, ZNI^{+}14, ZHCJ15]. reconstructive [MCS15]. recording [HZG09]. recordings [SDO^{+}04]. recover [KWB^{+}13]. Recovering [XDPT16, DCP^{+}14b, HXM^{+}13, LSS^{+}17, RMOW20]. Recovery [GLT^{+}23, TSLP14, YPA^{+}18, DRW^{+}14, HWJ^{+}15, LLK^{+}15, SFCH12, YSLH11]. Rectangles [Bae82]. Rectangling [HCS13]. rectification [GGY18, LSVT15]. Rectifying [WWSP23]. Recurrent [LCC^{+}22, SLL^{+}21a, CKS^{+}17]. Recursive [KCODL06, LPX^{+}19, NMLH14, NRN^{+}23, Sai89, SFCO12, XLY^{+}22b, ISD04, LXC^{+}17, NMLH11, NM16, ZXC^{+}18]. Recursively [BS88, BS90]. Redefining [GXSD23, UTB^{+}19]. Redirected [DFYL19, JKH^{+}22]. redirection [LSL^{+}18, SPW^{+}18]. redistribution [CTE05]. RedMax [WWB^{+}19]. Reduce [HC86]. Reduced [FW22, MAB^{+}15, PM18, BdSP09, BEH18, CZZ14, DSP06, GAB20, JP04, LMH^{+}15, SNB07, TVLF20, YLX^{+}15, vTSSH13]. Reduced-order [MAB^{+}15]. reduced-precision [TVLF20]. reducer [CLD^{+}13]. reducer-tuner [CLD^{+}13]. Reducing [And83, FBH^{+}10, ABA02, RAWV08, TBTS08]. Reduction [DCT^{+}22, PGG^{+}23, XLCB15, FCGH08, KJ09, LHdG^{+}14, McC99, SYS^{+}21, TLP06]. Redundancy [RV89]. Reeb [PSBM07, She13]. Reenactment [LXZ^{+}19, TZS^{+}18, TZN^{+}15, TZT^{+}18]. Reference [ASK^{+}22, LNZ^{+}23, YXW^{+}23, LWC^{+}13, SAN23]. reference-based [SAN23]. Reference-Map [LNZ^{+}23]. Refined [Pet01]. Refinement [CCK92, WK21, CYFW14, LVS18, LFY^{+}19, PLC^{+}21, SCF^{+}04, TWAD09, WZN^{+}14, ZDI^{+}15]. refining [SHK^{+}14]. Reflectance [AAL16, CDP^{+}14, CT82, CHB^{+}12, DvGNK99, MXZ^{+}23, TG17b, YXW^{+}23, ZK22, BDM09, DTPG12, DCP^{+}14b, DWd^{+}08, DHI^{+}13, FBLS07, FRSL08, GHP^{+}08, GZL14, GSH^{+}20, Gup18, HP03, HLZ10, HP17, HHA^{+}10, KCW^{+}18, LXR^{+}18, MSS^{+}12, MPBM03, MAG^{+}09, MHP^{+}19, NZV^{+}11, NLW^{+}16, NJR15, PTMD07, TG17a, TFG^{+}13, VLD07, WZT^{+}08b, WRG^{+}09, WYL^{+}20, WGT^{+}05, WMP^{+}06, WPMR09, XDPT16, XNY^{+}16, YSN^{+}18, YTJR15, YJR17, ZJ18, ZSD^{+}21]. Reflecting [RT90, BHW16]. Reflection [SLWL23, XWH^{+}23, HP17, IM12, RH04, ROTS09, ZNT18]. reflectional [SHZ^{+}20, XZT^{+}09]. Reflections [KLR^{+}22, SKV^{+}12, OF12, SMM14, SKG^{+}12, XWZ^{+}21]. Reflective [LWL^{+}23b, TB87, PSG^{+}06]. reflectometry [GTHD03, GCP^{+}10, RWS^{+}11, RRFG17]. reflectors [SDIN18]. reflex [POB09]. refocusing [MNBN07, VRA^{+}07]. refraction [LWL^{+}20, Wym05]. refraction-tracing [LWL^{+}20]. Refractive [ABW14, TB87, IZT^{+}07, PMOR10, PHN^{+}12, PCS^{+}20, SZS^{+}08, WZHB09, YTBK11]. refurnishing [ZCC16]. Region [JHR22, SB95, VKW^{+}23, ZZL^{+}21, KEE13, LSCO03, TDM11, WW13, YKC^{+}16]. Region-based [ZZL^{+}21, TDM11]. Regional [STZ^{+}16, Kim10, LSS^{+}17]. Regions [LMR83, LYF^{+}20, MWM23, VKW^{+}23, GAB20, SF07]. Registration [HSX^{+}22, AMCO08, CZ11, HGCO^{+}12, HYG^{+}13, MDK^{+}16, YNW16]. Regression [KIM^{+}19, MCY14, SGH^{+}22, SWWW15, APCO21, BPC16, CWLZ13, CHZ14, LJS^{+}15, RWG^{+}13, VKK18, WPP14, WPP07, WLT16]. Regression-based [SGH^{+}22]. Regular [HGM14, SYSP14, VMW17, ANHD17, LLH04, LPS^{+}13, LH04, MMBM15, vW09]. regularities [THW^{+}14]. regularity [PMW^{+}08]. regularization [XCS^{+}14]. Regularized [DJ17, IBB15]. Regularizing [XDW^{+}23]. regulated [WPL18]. rehabilitation [KDI19]. Reinforced [FZZ^{+}20]. Reinforcement [CXW^{+}23a, CPV^{+}23, GUPZ20, HWZ^{+}20, RYPZ23, SYM^{+}24, XGZ^{+}23, CYT^{+}18, HXC^{+}20, LP10, LH18, LSCC20, MTP^{+}18, PBvdP15, PBvdP16, PBYV17, PALvdP18, PKM^{+}18, PFX^{+}22, WLY20, XDF^{+}19]. Reintegration [DNZ^{+}17b]. Rejection [RLLG^{+}20]. Relating [THW^{+}14]. relation [FCW^{+}17, GMW16, WLW^{+}19]. relation-augmented [GMW16]. Relational [Mac86, DNB^{+}05]. Relations [KK91, vOV96, LWC^{+}11, ZRB14]. Relationship [ZHG^{+}16, GJWW14, HKT10]. relationships [FSH11b]. relative [XYXJ12]. relativistic [CLC96]. relativity [WKR99]. relaxation [DML17, KOOP11, MDK^{+}16, SJ13]. Release [PMLB22]. Reliable [MPB17a, PNA^{+}21, BCE^{+}13, MPB17b, Wam16]. Relief [ZTS09, PRZ17, POC05, SVB^{+}12, SKC^{+}14, WDB^{+}07]. Reliefs [AM10]. Relight [YNK^{+}22, PEL^{+}21]. Relightable [BZH^{+}23, BLS^{+}21, MPH^{+}20]. Relighting [DRC^{+}15, KE18, MPDW03, PMGD21, WS17a, ZFT^{+}21, ED04, GCD^{+}20, HPB06, LMM^{+}22, NRH04, NN04, PEL^{+}21, PTMD07, PVL^{+}05, PGZ^{+}19, RDL^{+}15, SZC^{+}07, SZS^{+}08, SBT^{+}19, SXZ^{+}20, WTL05, WTL06b, WYL^{+}20, WGT^{+}05, WS17b, XSHR18, ZCC16]. Relying [CH14]. remapping [GO17]. Remarks [Las90]. Remeshing [FSKP23, VVHSH22, AMD02, ACSD^{+}03, ESCK16, KS04b, NSO12, PNA^{+}21, PH03, WRK^{+}10]. remote [KTL^{+}04]. Removal [SO92, GOTG05, McK87]. Remove [GTB15]. Removing [ARNL05, FSH^{+}06, GRBN09, WHDS04]. Render [MBB12]. renderable [LSS^{+}19]. RenderAnts [ZHR^{+}09]. Rendered [OKH^{+}16, BDM^{+}21]. Renderer [BAC^{+}18]. renderers [PGML^{+}19, Sun06]. Rendering [BYG96, BGL20, BBC^{+}23, CWZ^{+}21a, CGMS22, CFS^{+}18, FH93, GFMS95, Gup18, IH20, JCW09b, JMY^{+}07, KHFH11, KKLD23, KJGP23, KL23, KAW20, KKW23, LXZ^{+}19, LSCS14, LC96, LTH^{+}23, Mal93, MCY14, MNV^{+}21, NRN^{+}23, Pha18, PBM^{+}22, RYW^{+}22, Rap91, SM17a, SLWL23, Ste20, SY22, Sun06, TG17b, Tsa15, TB87, VADWG15, WHHY20, WSND^{+}23, WWWZ23, WTS^{+}23, WA23, WLL23, WXZ^{+}23, XLY^{+}22a, XLC^{+}23, YHJ^{+}14, YMRD15, YHW^{+}18, YPG01, YZN^{+}22, ZWTP23, ZRJ23, ZJY^{+}22, ZZC^{+}22, ALLD17, ATM^{+}17, BWG03, BBPP10, BLD20, BAGL19, BGAM12, BKKL15, Bel18, BOD^{+}13, BFK^{+}16, BST09, BF08, CBCG02, CXGS02, CLS^{+}17, DI11, gDGPR02, DMB^{+}14, DAD^{+}18, Did18, DYN03, DIO^{+}12, DHOO05, DWd^{+}08, DPF03, DJ18b, ETH^{+}09, EC96, EMF02, FFB^{+}09, GLD^{+}19, GN06, GZB^{+}13, GM05, GGHS03, GTDS10, GSRN21, GBAM11, GYGS22, GTR^{+}06, GCH^{+}19, GS04]. rendering [HR05, HV04, HKWB09, HRDB16, HPP^{+}18, HMC11, HSW^{+}17, HGS23, HESL11, HHN^{+}02, HWJ^{+}15, HWH^{+}16, IZT^{+}07, JAM^{+}10, JM12, JdJM14, JSRV22, JMM^{+}14, JB02, KV05, KMM^{+}17a, KE18, KP11a, KHL19, KWN^{+}17, KB12, KDH22, KTL^{+}04, KLS^{+}13, KKW20, KCYW13, KOC^{+}22, KHLN17, LHK^{+}20, LS02, LES09, LAC^{+}11, LD21, LHZ16, LSSS18, LSS^{+}21, LB05, LB06, LH04, LKYU12, LCD^{+}20b, MBPY^{+}18, MYRD14, MPH^{+}20, MPH^{+}15, MBGJ22, MIGYM15, MMMG16, MPG^{+}16, NH08, NJJ21, NLMD12, NDMKJ22, NNDJ12, OL03, OKH^{+}17, ODR09, OEE^{+}18, PZ08, PSK^{+}16, PVG19, PMHD19, RH02, RCL21, RTF^{+}04, RGB16, RMD04, RZL^{+}10, REG^{+}09, RKZ12, RJN16, RFS22, SBdDJ13, SM17b, SD12, SHL^{+}17, SSY^{+}04, SKG^{+}12, SKS02, SSY22, SFWG04, SRNN05, SM06, SR09, TAV^{+}10, TTD22, TZN19, TG17a, TWL^{+}05, TS12, TGD04, TAKW^{+}19, VRC^{+}13, VT04, VSJ22, WKF^{+}21]. rendering [WWD^{+}05, WZT^{+}08a, WRG^{+}09, WHY^{+}13, WYM^{+}16, WHY20, WWLC21, WS99, WW08, WVJH17, Wes21, WFY^{+}10, WZYR19, WCRZ21, WXZ^{+}22, XMR^{+}11, XCM^{+}14, XWZ^{+}21, YTJR15, YHMR16, YSJR17, YLB^{+}22, YZL^{+}22, YKC^{+}21, YIC^{+}10, ZZXZ09, ZDDZ21, ZYZ21, ZHRB13, ZWDR16, ZLB16b, ZHHZ20, ZRL^{+}08, ZHR^{+}09, ZBX^{+}21]. rendering-aware [DAD^{+}18]. Renderings [GIGM22, BVM^{+}17, FJL^{+}16, ZZXY21]. RenderMan [CFS^{+}18]. Reodesics [SH23]. reordering [MBK^{+}10, SNB07]. rep [GLP^{+}22, JNK^{+}23]. Repair [LBB22, Ju04]. Reparameterization [XBLZ23]. Repeated [NDS^{+}23, CZM^{+}10, CLQW08, WWOH08, ZHRB13]. repetition [KMYG12]. repetitions [XCW14]. RepFinder [CZM^{+}10]. Rephotography [WBF^{+}17a, BAD10, LZY^{+}21, WBF^{+}17b]. Replacement [RKS^{+}14, DSJ^{+}11, JMD^{+}17, PEL^{+}21, TSL^{+}16, ZYQ^{+}14]. replacing [BKD^{+}08]. replay [VSJ21]. repositories [YGH^{+}17]. represent [PMHD19]. Representation [ARMCO23, BN90, DK99, GLL^{+}16, JCFG23, KLM24, LHH^{+}23, SLM^{+}17a, WSP^{+}23, WZX^{+}23, ZTNW23, ZZW^{+}22a, ABA02, ABJN85, BAS14, BAERD08, Boi84, CBCG02, DF88, DZCJ21, FKY^{+}10, GLLR11, HNB^{+}06, HZW^{+}13, KV05, KHD14, KCYW13, LRR04, LBAD^{+}06, LKK^{+}16, LZT^{+}08, LMM^{+}22, MLL^{+}21, MASS15, MWI18, OBW^{+}08, OBCS^{+}12, PSH^{+}21, PKG06, PvBM^{+}06, RS98, RAKRF08, SPSH18, SLM^{+}17b, SHX^{+}22, STPP09, STZ14, WSLT18, Wim14, YKZ^{+}22, ZLY^{+}21, ZYSK21, ZBX^{+}21, ZKU^{+}04]. Representations [DS92, GWLG23, PBS20, WLY^{+}16, ZYM^{+}20, ZCP^{+}23, MGP10, NPLX22, VJK21, WLT22]. represented [VA88]. Representing [BDK^{+}16]. reproducible [LSGV18]. Reproducing [HCE03, ZJY^{+}21, CLC^{+}20, DTPG12, LDF14]. Reproduction [FR22, SFB92, AAMSB20, DWT^{+}02, ESZ^{+}17, HFM^{+}10, LYL^{+}16, PFB^{+}20, RSSF02, RPK^{+}12, SBK^{+}18]. reprojection [RLP^{+}20, SaLY^{+}08, YTS^{+}11]. reprojection-based [SaLY^{+}08]. Repulsion [WWYW21]. Repulsive [YSC21, YBSC21]. Requirements [SFB92]. reradiation [HHA^{+}10]. resampled [LKB^{+}22b]. Resampling [NID20, HWG^{+}13]. Resampling-aware [NID20]. rescreening [KP18]. ReShader [PNTK23]. Reshaping [AVB^{+}23, AVR^{+}22, JTST10, ZFL^{+}10]. Residential [FW16, RFW^{+}23, LGZ^{+}13, MSK10]. Residual [NSJ14]. Resilient [YLC^{+}20, AAR05]. Resizing [WWF^{+}10, AS07, DZPZ09, KSSCO08, WTSL08, WFS^{+}09, WHSL11]. Resolution [BF12, FJA^{+}14, GLT^{+}23, LSO07, LB05, QRL^{+}23, SWS^{+}22, TWR^{+}23, YJLL22, AGL^{+}17, AYL^{+}12, AFC^{+}10, AB20, BWDL21, BHPS10, DER^{+}10, ESCK16, GLD^{+}19, GGY18, HSB^{+}12, HW15, HG09, KSA13, KZP^{+}13, KLM24, KMX^{+}21, LEPM22, LGX^{+}13, LFJG17, Mus13, NB11, SGM12, SDP^{+}18, SXD^{+}12, SZD^{+}20, SXZ^{+}20, TREO16, VBCG10, VSK^{+}17, WOR10, WAK20, WGDE^{+}19, XFCT18, YHJ^{+}14, ZSCS04, ZHRB13, ZSTB10]. Resolution-matched [LSO07]. resolutions [LSH^{+}22]. resolved [AIH^{+}08]. Resolving [VMT06, ZLB16a]. resonance [UPSW16, WMB19]. Respecting [CPAB22]. Response [VJ19, JP02, KNL^{+}22, PMG^{+}22, PNH^{+}14, TDM^{+}14, ZMCF05]. responses [LMLL21]. Responsive [MP07, CK02, YL08]. ReSTIR [LKB^{+}22b, SLK^{+}24, WWWZ23]. restitution [WSJP17]. restoration [ALY08, BPK05]. Restoring [ZBG15a]. Restricted [WWWG22, WWX^{+}22, CM11, LLX^{+}12]. Restructuring [DWX^{+}21]. Resultant [SG17]. Resultants [CGM91]. Retargeting [GFK^{+}23, YFFA21, AWL^{+}19, ALL^{+}20, ATDP11, CCGB22, HRE^{+}08, HXK^{+}19, KLHG09, LCOZ^{+}11, PMPHB17, RSA08, RSA09, RGSS10, SSTP15, SSK^{+}11, TK05, WLSL10, ZAC^{+}17, BZL^{+}17]. Rethinking [FSRS22]. Retiling [AFSR03]. retiming [LCD^{+}20b]. retina [JBM^{+}17]. Retinal [JBM^{+}17]. Retrieval [JWD^{+}23, SK16, SSB^{+}17a, SSB^{+}17b, BBGO11, CTS^{+}20, ERB^{+}12, MRC05, WCSC22, XCF^{+}13]. retrieve [SBHH16]. Returning [BSM88]. Reusable [JZvdP^{+}08, MTA^{+}20, PGH^{+}22]. Reuse [MNV^{+}21, HZ11]. reusing [BPE17, HA18]. reveal [CHM^{+}12]. Revealing [DMIF15, WDW^{+}15, WRS^{+}12]. REVEL [BP12]. reverse [EKM17, MAF^{+}09, RTS^{+}07]. Reversible [BJNJ18, ESBC19, LMAH^{+}18, YYW^{+}12a]. reviewers [TOP03]. revision [CWC11]. revisited [GS04, VW97]. Revisiting [FGW^{+}21, HGS23]. Revolution [MHS^{+}19a, HSST10]. Rewriting [WBZ22]. Reyes [PO08, ZHR^{+}09]. Reyes-style [PO08]. RFEPS [XWD^{+}22]. RFIG [RBvB^{+}04]. RGB [BIP01, BLC^{+}22, Bou18, CKH18, CLW^{+}14, GXY^{+}17a, MSS^{+}17, SCB87, SHZ^{+}20, Sun06, TLG17a, TLG17b, TZY^{+}23, WSXC16, WMB^{+}20, XNZ^{+}22, YLC^{+}20, ZWW^{+}18, ZSSJL20, ZNI^{+}14]. RGB-based [Sun06]. RGB-D [BLC^{+}22, CKH18, CLW^{+}14, GXY^{+}17a, SHZ^{+}20, TZY^{+}23, XNZ^{+}22, YLC^{+}20, ZWW^{+}18, ZNI^{+}14]. RGB-space [TLG17b, TLG17a, ZSSJL20]. RGB2Hands [WMB^{+}20]. RGBA [UTB^{+}19]. RGBD [GXY^{+}17b, LCC^{+}18, SXZ^{+}12]. RGBXY [TEG18, XLY^{+}22a]. RGBXY-space [TEG18]. RGYB [WC91, WC90]. Rhizomorph [LKM^{+}23]. Rhythm [AGL^{+}22, DA18]. Rhythm-Aware [AGL^{+}22]. Rhythmic [AGL^{+}22, hKPS03]. Rhythmic-motion [hKPS03]. ribbon [RAR^{+}21]. ribbons [RPC^{+}21]. rich [LYvdP^{+}10]. Rich360 [LKK^{+}16]. richly [BUSB13]. Richness [QPWH08]. Richness-preserving [QPWH08]. Ridge [OBS04]. Ridge-valley [OBS04]. ridges [JDA07]. Riemannian [ZWL^{+}18]. Rig [GMP^{+}16, HMT^{+}12, PAR21, KS21, LKZ^{+}20, SSR20, ZBBB18]. Rig-space [HMT^{+}12]. Rigel [HDD^{+}16]. rigging [BP07, BJD^{+}12, LD14, LWP10, XUC^{+}14, XZK^{+}20]. Right [McI92]. Rigid [BB14, CMT04, GPB^{+}19, LXY^{+}23, NI22, NI24, ZJ10, AIA^{+}12, AMB^{+}21, BR07, CAJ09, CZJ12, CBK20, DBB^{+}17, FLS^{+}21, GBF03, GSLF05, HSGL11, HFG^{+}18, IMH05, JTSB16, KEP05, LJ14, LD12, LZQ^{+}22, MTP^{+}18, NAI^{+}18, PSE03, RMSG^{+}08, TB20, TB21, TK14, Ten20, TBV12, TJ08, VSK^{+}17, WSJP17, WP12, ZBG15b, ZNI^{+}14]. Rigid-body [ZJ10, CZJ12, LJ14, PSE03, TB21, WSJP17]. rigid-fluid [AIA^{+}12, TB20]. rigid-rigid [TB20]. rigidification [MAKWL22]. rigidity [CCWL18]. RigMesh [BJD^{+}12]. RigNet [XZK^{+}20]. Rigs [GZC^{+}16]. Ring [AECO15]. Ring-Ordering [AECO15]. RingIt [AECO15]. rings [PCK^{+}19, WPS14]. Rivers [PGCG23]. RLE [HNB^{+}06]. roadmaps [CLS03]. RoboGrammar [ZXKL^{+}20]. Robot [GLX^{+}22, YLC^{+}20, DXZ^{+}19, ZXKL^{+}20]. Robot-Mounted [YLC^{+}20]. Robotic [MHCT23, MSK^{+}23, GPD^{+}18, HXK^{+}19, HZH^{+}16, MTN^{+}15, ZPBC19]. Robots [GFK^{+}23]. Robust [BFA02, CRB23, CPW^{+}23, CPAL22, CBvdP09, CPS13, DPVA23, DZCJ22, DD02a, FH93, FCOS05, GJTP17, GPW^{+}17, HJ11a, HVTG08, HYNP20, HWZ^{+}14, Hol18, HCW^{+}23, HMLL14, JKSH13, Ju04, Kal14, KJDL09, KBT17, LDK^{+}18, LD14, LAGP09, LPL^{+}18, MdLH10, MPZ14, PCL^{+}12, PSBM07, RS14b, SBRBO20, SKY^{+}12, SOHK16, TZY^{+}23, VGB^{+}14, XZZ^{+}14, ZWZ^{+}16, ZZMC13, AMCO08, BWSS09, BRB^{+}19, CCS^{+}21, CLSA20, CWTW17, DJBJ19, DA21, EBCK13, FDBH22, HPJ12, HSG^{+}19, KSNG17, LBK16, Mir98, MCKM15, RJ07, SHHD17, SLMB05, TNWK22, VCA^{+}22, YLJ18]. robustly [DBDB11, TMRL14]. rod [KTS^{+}14, MLB16, PTC^{+}15]. Rodent [PGML^{+}19]. Rods [HB23, BWR^{+}08, MKB^{+}10, SBRBO20, SJM17]. role [GXZ^{+}13]. Rolling [JGN16, WFL^{+}15]. Rom [DB88]. roof [RZW^{+}21]. room [STXJ15, YZL^{+}22]. Rooms [HC86]. Roots [LKM^{+}23]. Rope [LXG^{+}22]. ROSEFusion [ZZZX21]. Rotation [HFK94, Hil87, ACXG09, BN21, CGM11, JBY^{+}19, LH16, LSLCO05, NSF12, PBH15, WJZL08, Xia21]. rotation-aware [BN21]. rotation-invariant [LSLCO05]. rotation-strain [PBH15]. Rotational [PZ07, SHZ^{+}20, WPP07]. rotations [PR97a]. Roto [LVS^{+}16]. rotoscoping [AHSS04, LVS^{+}16]. Rough [IBB15, SY22, LJJ^{+}18, SSISI16, SSII18b, YVG20]. Roughness [GFL^{+}22, TGZ18]. roulette [RGH^{+}22, TH19, VK16]. Round [Pra89]. Route [DLP^{+}23]. Routing [PRM14]. row [HPB07]. row-column [HPB07]. RPU [WSS05]. rubber [FLGJ19]. Rule [Wan18a]. Rule-free [Wan18a]. rules [NSX^{+}11, WBZ22]. run [GSKJ03]. run-time [GSKJ03]. runner [LYvdPG12]. Russian [RGH^{+}22, TH19, VK16]. RXMesh [MPO21].

Saccade [ATM^{+}17, DCB^{+}22]. saccadic [SPW^{+}18]. Saddle [YWH13]. Safe [WWYW21]. Safer [BC23]. safety [KDI19]. Sag [HWP^{+}23, HTYW22]. Sag-Free [HWP^{+}23, HTYW22]. SAGE [DN02]. SAGNet [WWL^{+}19]. SAH [DFM13]. SAH-optimised [DFM13]. SAILOR [DXG^{+}23]. Saint [KTY09]. salience [GOTG05]. salience-preserving [GOTG05]. saliency [LDS^{+}16, LVJ05, MLH^{+}09, SLMR14]. Salient [GCO06]. Sample [GLA^{+}19, DH06, WLM^{+}15]. Sample-based [GLA^{+}19]. Sampled [HWZ^{+}14, YSHWSH16, APKG07, AA06, BGAM12, DZCJ21, HRV^{+}18, MWR12, PKKG03]. sampler [ANHD17]. Samplers [SLK^{+}24]. Samples [LNLB16, WWWZ23, BJ17, XSHR18, ZXS^{+}21]. Sampling [Coo86, FHG^{+}23, HSS98, HWZ^{+}20, KWB^{+}15, LLX^{+}01, LYvdP^{+}10, MHGCO21, MEA^{+}18, MMR^{+}19, Ost07, Pav90, QCHC17b, Sah18, SGSS22, SMR^{+}22, WP90, ZRJ23, ARBJ03, ARNL05, APC^{+}16, AW20, ALLD17, BLD20, BMW^{+}09, BWWM10, CGW^{+}13, CJAMJ05, CTM13, EDP^{+}11, Fat11, FBLS07, GM09, GKH^{+}13, GYGS22, HJW^{+}08, HPB07, HSD13, HGS23, HWJ^{+}15, HWH^{+}16, JZW^{+}15, KTBV16, KVG^{+}19, LRR04, LDF14, LWSF10, LWC12, LADL18, LKB^{+}22b, MRK^{+}14, MSOC^{+}19, NJR15, ODJ04, OP11, ÖAG10, PCI^{+}21, PBC^{+}22, Pet21, QCHC17a, RKLC^{+}11, RAMN12, RAWV08, RHJD18, RKZ11, SJ17, SK13, SZG^{+}13, VKK18, WPC^{+}14, Wei08, Wei10, WW11, WWZ^{+}06, XNY^{+}16, YW13, YL12, YIC^{+}10, ZDDZ21, ZHWW12, ZD20, EPM^{+}14]. sampling-and-recovery [HWJ^{+}15]. Sampling-based [LYvdP^{+}10]. sand [KGP^{+}16, TGK^{+}17, ZB05]. SANM [Jia21]. sans [DBWG15]. Sassafras [Hil86]. Saucer [WCFL22]. sauces [NSS^{+}19]. scaffoldings [DHL14]. scaffolds [SKSK09]. Scalable [CBI13, CZY17b, CSK18, CRCM23, GGN18, HRDB16, LPLL19, PTC^{+}10, RPPSH17a, RPPSH17b, SGSS22, TZY^{+}23, WHSL11, WXZ^{+}22, WXZ^{+}23, XWZ^{+}21, AFTCO07, BDT^{+}08, CZY17a, Dav20, DML17, FZBR16, LCD^{+}20a, LMAS16, MP04, MGT^{+}03, REG^{+}09, WFA^{+}05, WQS^{+}20, WGH20, YKC^{+}16]. scalar [PSF09]. Scale [GKK^{+}24, LWF^{+}23, LZCX19, LYC18, MHZ^{+}21a, SHG^{+}22, WXZ^{+}23, XLC^{+}23, ZSCM17b, Ang17, ASL^{+}17, BPD06, BL15, BBA^{+}07, CQD^{+}18, DFZ^{+}17, EDF^{+}16, FFLS08, FMB^{+}17, FBGZ18, FYY^{+}16, FSP^{+}22, FAW19, FG14, GARP^{+}23, GB13, GLDZ15, GNS^{+}12, HP17, HHM19, IDN12, JP03, KGG^{+}20, KFWM17, KSKL14, KPZK17, KABL15, LDPT13, LWL17, LCX^{+}21, LSA^{+}16, MHS^{+}19b, MPH^{+}15, MGP10, NDD^{+}23, NZIS13, PRFS18, PGH^{+}22, PCHF18, RNGF03, RGB16, SPF^{+}23, SWL11, SHM22, SLSS03, SG11, SJLP11, SJMP10, VSLD13, WTSL08, WSM11, WFDH18, WFS^{+}21, WDR11, WDR13, XZJ^{+}12, YIO^{+}15, YSQS08, ZSCM17a]. scale-and-stretch [WTSL08]. Scale-aware [LYC18]. Scale-optimal [GKK^{+}24]. scales [FG11, XLZ^{+}10]. scaling [DZPZ09]. Scan [RWW90, ACP02, CSK^{+}22, LKZ^{+}20, ZSW^{+}10]. Scan-Conversion [RWW90]. ScanBot [CXW^{+}23a]. ScaNeRF [WXZ^{+}23]. scanline [LHZ16]. scanned [XGC07]. Scanner [PCHF18, HLZ10, WAO^{+}09]. Scanning [LGHL23, CDP^{+}14, FZBR16, HWV^{+}18, HCTW11, HDGN17, HFI^{+}08, MKZ^{+}21, YSL^{+}14]. Scans [FJA^{+}14, ACP03, BR07, CZ11, LBB^{+}17b, YNW16]. SCAPE [ASK^{+}05]. Scattering [BBS14a, ESZ^{+}17, FHK14, KM17, BAGL19, BGL20, BCRK^{+}10, DWP^{+}10, FD17, FCJ07, GKH^{+}13, GJZ21, HFM^{+}10, HHdD16, KMM^{+}17a, LJJ^{+}18, MJC^{+}03, MGJ19, MM06, MWM08, NZV^{+}11, NGD^{+}06, PvBM^{+}06, STPP09, SRB^{+}19, SRNN05, SZLG10, VKJ19, WZHB09, WTL05, XWM^{+}20, XH18, ZWDR16, ZYWK08]. Scattering-aware [ESZ^{+}17]. Scenarios [TFD^{+}18]. Scene [DWX^{+}21, GLX^{+}22, HE07, HSV^{+}22, HZD^{+}23, KSH^{+}14, KZP^{+}13, KKN^{+}22, LLZ18, LY23, LGHL23, RO85, RO87, SFC^{+}23, SYM^{+}24, WLY20, WLJ^{+}22, WXZ^{+}23, YBMN^{+}23, ZXTZ15, ZYM^{+}20, BHY15, CZM^{+}10, DXZ^{+}19, FSL^{+}15, GSRN21, HXZW20, JMK^{+}22, KWB^{+}15, KPZK17, KN06, LHLY21, LLHY22, LCK^{+}14, LXS^{+}18, LSH^{+}22, MLZ^{+}16, MPF^{+}18, MLL^{+}21, MGC^{+}19, NXS12, NKGR06, RSI^{+}08, SMZ^{+}14, STZ^{+}16, SMGH18, VJK21, WSCR18, WXZ^{+}22, XMZ^{+}14, XHS^{+}15, XWZ^{+}21, YTS^{+}11, YZL^{+}22, ZN06, ZYX^{+}21, ZHG^{+}16, ZK13, ZXH^{+}20, vdHDT^{+}07]. Scene-Aware [SFC^{+}23, YBMN^{+}23, LLZ18, LHLY21]. scene-level [BHY15]. scene-space [KWB^{+}15]. SceneGrok [SCH^{+}14]. Scenes [DPD22, DRC^{+}15, JGC^{+}15, JRSS21, KAEE20, LPX^{+}19, RSV^{+}23, SM17a, VLA15, YLC^{+}20, ZWK14, AAC^{+}06, AZB09, ADM^{+}08, BSM^{+}07, BF08, CLW^{+}14, CXY^{+}15, CAC^{+}02, DKD^{+}16, FSH11b, FSP^{+}22, FCW^{+}17, GTDS10, HKWB09, JM12, JF03, KR17, KNS^{+}09, LRT^{+}14, LDTA17, LGZ^{+}13, LCX^{+}21, MPF^{+}18, MP04, MRA^{+}13, MMBM15, NPLX22, NNDJ12, PFHA10, RSM^{+}10a, RWS^{+}06, SM17b, SKY^{+}12, SXZ^{+}12, SKG^{+}12, SZLG10, TPWG02, WIK^{+}06, WBS07, WLW^{+}19, WDB^{+}07, WGL^{+}18, XZY^{+}17, YMR^{+}13, ZSW^{+}10, ZHL^{+}05]. Schedule [LH17a, LH17b]. schedules [RKAP^{+}12]. Scheduling [LNI^{+}23, MHZ^{+}21a, BDK^{+}16, MAS^{+}16, SKK^{+}12, SKB^{+}14]. Schelling [CSPF12]. Schematic [GCSS06]. Scheme [DLG90, LCD^{+}19, MWM23, DM13, FGW^{+}21, PR97b, VB06, ZM11]. Schemes [LPC22, CADS09, LYLL08, WWT^{+}06]. Schrödinger [CKP^{+}16]. Schur [CZY17a, CZY17b, LMAS16, PAK^{+}19]. Schur-complement [LMAS16]. Schwarz [WWW22]. scissors [WAC07]. Scope [Fol94, Fol95b]. SCORES [ZXC^{+}18]. Scrambling [APW23]. Scratch [SLWL23, WVJH17]. Scratch-based [SLWL23]. scratched [RGB16]. Screen [AW20, AAPS16, PCB23, AAPS17, HLHR09]. Screen-space [AW20]. Screened [KH13, CK11]. screening [QPWH08]. screens [ALK^{+}17]. screentone [XLLW20]. scribble [XFAT12]. scribble-based [XFAT12]. Scroll [Ols92]. Sculpting [RAD12, Ros94, TQ94, CSTP16, DJ17, JX96, PXW18]. SCULPTOR [QLH^{+}22]. SDF [ZPW^{+}23]. Seam [AS07, DZPZ09, FHM^{+}21, LFJG17, RSA08, STP12]. seam-aware [LFJG17]. Seamless [APL15, CSZZ20, Lev21, Lev23, SMHW16, XXL^{+}21, FPBCO20, KDM^{+}16, LFH15, LSC^{+}12, MGA^{+}17, PMPHB17, LFJG17]. seamlessness [MS05]. seams [RC22, WSH19]. Search [FFWL^{+}22, AMA^{+}19, FH10, FMK^{+}03, HPC21, KSSI17, NXS12, SH07, TYS09, WLLS22]. search-classify [NXS12]. searches [EPM^{+}14]. Searching [MGA^{+}22]. Searchlight [WKR99]. Second [EC93, LLJ^{+}23, MJJG18, SXH^{+}21].

Second-Order [EC93, MJJG18, LLJ^{+}23, SXH^{+}21]. secondary [DJ18a, KKN^{+}13]. SecondSkin [DS15]. section [SBSS12]. Sections [PK83, BVG11, HZCJ17, MSM11, NCVMO05, ZHCJ15]. sediment [GPH^{+}18]. see [ALK^{+}17]. see-through [ALK^{+}17]. Seeing [EMO10, MSM^{+}23, SBB^{+}22]. segment [SZG^{+}13, XTZ^{+}21]. Segmentation [AASP17b, BLAE22, HMM^{+}21, ST16, VFK^{+}14, YSHWSH16, AASP17a, AOP^{+}18, ACA^{+}19, CGF09, DAB15, HKG11, HFL14, JKSH13, KHS10, SSRB^{+}17, SvKK^{+}11, WGW^{+}13, YC21, YGH^{+}17, ZAFW21]. Segments [KPACO22, Gal99]. Seidel [FTP16]. Selected [KP92]. Selecting [SPG^{+}23, TMRL14, Xia21]. selection [AAMSB20, ACCO05, FAC11, JKT^{+}15, LSS09, OLAH14, XFAT12]. Selective [RHJD18, ZZL^{+}21, MLH^{+}09, XCS^{+}14]. Selectively [BAAR12]. Self [BMBRD24, BD02b, CDY23, CLQW08, JHS^{+}23, MHS^{+}19a, MHGCO21, NKS^{+}23, OCNG21, PHL^{+}09, SHK^{+}14, WWYW21, ZWL22, BJ10b, DPW^{+}14, FF11, LVG^{+}13, LDPT17, LB18, LPS^{+}13, MIB15, MASS15, PSK^{+}12, RvBB^{+}03, RBvB^{+}04, SPO10, SRL^{+}15, TOK14, VHWP12, WPL18, WLH^{+}13, Xia21, YNL^{+}21, YY17, ZJ12]. self-adapting [PSK^{+}12]. Self-animating [CLQW08]. self-attention [YNL^{+}21]. self-augmented [LDPT17]. self-collision [BJ10b, MASS15, SPO10, WLH^{+}13, ZJ12]. self-configuring [RvBB^{+}03]. self-contact [TOK14]. self-describing [RBvB^{+}04]. self-examples [FF11]. Self-folding [NKS^{+}23]. self-illuminating [YY17]. Self-Intersecting [CDY23, LB18]. Self-organizing [PHL^{+}09]. self-portraits [LVG^{+}13]. Self-refining [SHK^{+}14]. self-regulated [WPL18]. Self-Sampling [MHGCO21]. self-selecting [Xia21]. Self-Shaping [JHS^{+}23]. Self-Similar [OCNG21]. Self-similarity [BD02b]. Self-supervised [BMBRD24, ZWL22]. Self-Supporting [MHS^{+}19a, DPW^{+}14, LPS^{+}13, MIB15, VHWP12]. Selfies [BLC^{+}22]. Semantic [AOP^{+}18, BVGP09, BSP^{+}19, CAV^{+}23, CZG^{+}11, GZX^{+}22, HLC^{+}19, HWV^{+}18, HMM^{+}21, IVH^{+}23, LGZ^{+}13, YCHK15, CLW^{+}14, HXM^{+}13, LMS13, LSH^{+}22, MC12, SXZ^{+}12, TD16, TER^{+}20, TSL^{+}16, WXLY17]. semantic-aware [TSL^{+}16]. semantic-based [TD16]. Semantically [ZCB^{+}22, VPB^{+}22]. semantically-aware [VPB^{+}22]. SemanticPaint [VVC^{+}15]. Semantics [LNI^{+}23, PQF^{+}23, LLHY22, XXL^{+}21]. Semantics-guided [PQF^{+}23]. Semi [CSAP21, MCW^{+}21, SAN23, WAK20, YZX^{+}18, BGOS06, DBD16, GBAM11, HDS^{+}18, HSG13, Wan15]. Semi-analytic [WAK20]. semi-analytical [GBAM11]. Semi-Implicit [CSAP21, DBD16]. semi-iterative [Wan15]. semi-Lagrangian [BGOS06]. semi-structured [HDS^{+}18]. Semi-Supervised [YZX^{+}18, MCW^{+}21, SAN23, HSG13]. semidefinite [KABL14]. Sensation [OL03]. Sensing [MHU19, PRM14, XGZ^{+}23, CSHH21, GWP^{+}19, HLHR09, LTO^{+}15, LGK^{+}16, MYWI15, PML^{+}09, RP09, VCA^{+}22, WYL^{+}14]. Sensitive [SO92, UKIG11, JP04, JBP06, NBB04]. Sensitivity [XUC^{+}14, YPG01, ZCT22, MAC22, RP03]. Sensitivity-optimized [XUC^{+}14]. Sensor [CSL^{+}23, GPHSH19, JCRA11, LOW18]. sensorimotor [NZC^{+}18]. Sensors [JGN16, KZSR16, YZH^{+}23, CHWH17, JKZS10, YZX21, ZSZ^{+}14]. separable [Ada21]. separate [XPB^{+}21]. Separating [CCW93]. Separation [SV93, CTW09, EML^{+}18, FGW^{+}21, NKGR06, SJR18, XLZ^{+}10]. Separators [BR21a]. Sequence [GW90, LAZ^{+}22, WL16]. Sequences [ASHW23, RKS^{+}14, CLM^{+}13, CKS^{+}17, DKP11, HAK^{+}22, LEN09, LD14, LCC^{+}18, TS08, WC10, WPL^{+}21, XZY^{+}07]. Sequential [DVS03, KSSI17, HET^{+}14, LPBM20, RMGH15]. series [CYW^{+}16]. Session [Bae18, BC18, Bou18, Cor18, Did18, Gup18, Hac18, Iza18, Kal18, Kau18, Kim18, Lau18, Lee18, Li18, Lip18, Liu18, Mit18, Pan18, Rit18, Ter18, Wan18b, Xu18, Zha18, Zho18, Zhu18a, Zhu18b]. Set [Day90, HCW^{+}23, JK23, LNZ^{+}23, PVY90, SZB18, Aca07, AA09, AK04, ASGCO10, FCOAS03, FLHCO10, GG07, HNB^{+}06, HWG^{+}13, HCJ19, MBWB02, NZWC20, NNSM07, SvKK^{+}11, WAvK^{+}12, WSVT13, XZCOC12, YCL^{+}15, ZM11]. Set-in-stone [SZB18]. Sets [DS92, VKW^{+}23, AHD15, AMCO08, KTB07, Kim10, KG04, MASS15, PTSZ11]. sew [KWL^{+}21]. Sewing [KSH23, LXL^{+}23, BGK^{+}13, KL22, Wan18a]. SFV [PKM^{+}18]. SGGX [HDCD15]. SGN [ZCT22]. SH [NSF12]. shade [LBAD^{+}06, LMPB^{+}13]. shaded [OBW^{+}08]. Shader [BWL^{+}23, HFH^{+}17, LS02, MDP^{+}04, HFTF15, HFF16, Pel05, SAMWL11, SaLY^{+}08, WYY^{+}14]. Shader-driven [LS02]. shaders [FH11, HSS98, VAZH^{+}09, YBAF22]. Shading [CA24, FHL^{+}18, GZ08, KOF14, MVD^{+}18, MNV^{+}21, NON85, PAR21, RV89, ZDI^{+}15, AB08, BSM^{+}07, CDP^{+}14, CTM13, CTH^{+}14, CM14, FBH^{+}10, HGF14, HFF18, HDHN16, HZ11, LMLH07, RMB07, RBD06, SPJT10, SBSS12, TiABI07, VBFG12, WZN^{+}14]. Shading-based [GZ08, ZDI^{+}15, WZN^{+}14]. Shadow [CGC^{+}03, McC00, MP09b, SCH03, WZC^{+}20, WL16, AAM03, BCRK^{+}10, EHDR11, GLY^{+}03, LAA^{+}05, LSO07, LGQ^{+}08, PTG02, RGK^{+}08, SOA11, SD02, WTBS07a, ZHL^{+}05]. ShadowDraw [LZC11]. Shadows [GTB15, Hud92, KOF14, ADM^{+}08, KOF13, MWR12, NRH03, PSNB13, RMB07, RWS^{+}06, SKOA14]. shake [FSH^{+}06]. Shallow [JW23, WSZ^{+}18]. Shape [BBB^{+}93, BL20, BBGO11, CRB23, CPY^{+}22, CKPS18, CPW21, DB88, GSP^{+}23, HFW^{+}19, HHL^{+}24, HKC^{+}18, IRHSH20, JS11, JHR22, JHS^{+}23, KFR04, LBB22, LHH^{+}23, MOR^{+}18, NI22, OFCD02, PMLB22, PKKG03, SK16, Sah18, SPSH18, SSB^{+}17a, MBU22, VFK^{+}14, VR94, VTSSH15, WLX^{+}18, WBCPS19, XWC^{+}16, YYPM11, YML^{+}23, YSC^{+}23, YPL^{+}23, YZX^{+}18, ZTNW23, ZPW^{+}23, AKZ^{+}17, ALX^{+}14, AXZ^{+}15, ASK^{+}05, AFTCO07, BAS14, BBB^{+}14, Boi84, BWKS11, BWSK12, BJD^{+}12, BHSH^{+}22, CB17, CWLZ13, CI84, CWKBC13, CZXZ14, CW17, CBW^{+}18, CCW16, CSAD04, CSD^{+}09, DLC^{+}15, DFRS03, DYT05, ERB^{+}12, FH07, FAR07, FvKBCO16, GCO06, GSMCO09, GYQ^{+}18, GJWW15, HK12, HLZCO14, HSC^{+}22, HKG11, HGCO^{+}12, HZG^{+}12, HSG13, HWG14, HWK15, HLW^{+}19, HOM15, HJM^{+}22, IMH05, JBX^{+}20, JCG^{+}21, KCKK12, KMP07, KCGF14, KvKSHCO15, KST08, LVS^{+}16, LXC^{+}17, LBB^{+}17b, LMAH^{+}18, LXR^{+}18, LLHF21, LCORL07, LFZ18, LFJG17]. shape [LMB14, LKS15, LKWS16, MDZ^{+}21, MDLW15, MSM11, MDB^{+}19, MHTG05, MAB^{+}15, MHR^{+}16, PCS23a, PRFS18, PMRMB15, RHSH18b, RKP^{+}22, RJ07, RCOL09, RBD06, ROA^{+}13, SS14, SSB^{+}17b, SCW^{+}21, SHM^{+}14, SSP07, SKAG15, SJA^{+}20, TBW^{+}12, TGB13, TCL21, TMB14, TFG^{+}13, VLD07, VBBF16, VKJ19, VPB^{+}09b, WAO^{+}09, WGW^{+}13, WJBK15, WLG^{+}17, WZF^{+}18, WCPM18, WLT22, WG10, Wim14, WWL^{+}19, XDPT16, XCOJ^{+}09, XZCOC12, XFAT12, YC21, YKC^{+}16, YGH^{+}17, YHCOZ18, YK12, YK14, YCHK15, ZAJ^{+}15, ZSD^{+}21, ZYL^{+}17, ZXC^{+}18, vKXZ^{+}13, vFTS06, Ano10]. shape-adaptive [VKJ19]. shape-complexity [CI84]. Shape-Matching [BBB^{+}93]. shape-proxy [MSM11]. Shape2Pose [KCGF14]. Shape2Vec [TD16]. ShapeAssembly [JBX^{+}20]. ShapeCoder [JGMR23]. Shaped [EPO91, HA92, MSS^{+}19]. ShapeMOD [JCG^{+}21]. ShapePalettes [WTBS07b]. Shapes [BBG24, CH14, EM94, FBS^{+}23, HLV^{+}17a, HJS^{+}14, LYF^{+}20, MLS^{+}18, PMKB23, WZ22, ZPYX23, ACP03, GSV^{+}17, HR05, HPG^{+}22, HLV^{+}17b, HSS^{+}13, HZH^{+}16, KH06, KLM^{+}13, KSH^{+}16, LMS13, LLV^{+}12, LSQ^{+}15, LAH^{+}21, LYC18, LKG^{+}03b, LSCS14, MLYZ19, MSHS06, MRA^{+}22, MZL^{+}09, MB21, NB11, OLGM11, OBCS^{+}12, PSG^{+}06, PWLSH13, SHZ^{+}20, SvKK^{+}11, TD16, THW^{+}14, UIM12, WAvK^{+}12, WSLT18, WSH^{+}18, XZT^{+}09, YSC^{+}16, ZAC^{+}17]. Shapes-Theory [BBG24]. Shaping [JHS^{+}23, CLC96, GMB17, MPI^{+}18]. Shared [FSRS22, BAM13, KKB^{+}11, WCPM18]. sharing [SGM12, SSTP15, SMHW16]. sharp [ASGCO10, FCOS05, MRA^{+}22]. Shear [YSB^{+}15, NSS^{+}19]. Shear-Dependent [YSB^{+}15]. Sheared [YMRD15, ETH^{+}09, EHDR11]. SHED [KvKSHCO15]. shedding [WP10]. Sheet [BTSB23, SMCT18]. sheets [BUAG12, DBWG15, NPO13, PTG12, PNdJO14]. shelf [MHM^{+}17]. Shell [CTW^{+}04, GUPZ20, PBFJ05, ZDF^{+}23, CSvRV18, CQD^{+}18, JSZP20, LCBD^{+}18, NAI^{+}18, CQS^{+}23]. Shells [BSR^{+}23, CCK^{+}21, JHS^{+}23, MM22, WSND^{+}23, WB23, BMWG07, CAJ09, CLF^{+}18, CNZ^{+}22, GSLF05, GHF^{+}18, KMB^{+}09, MPBC16, MPI^{+}18, MKB^{+}10, RK13, RMSG^{+}08, PKLI^{+}19]. Shield [LRAT08]. shiftable [SMH^{+}11]. Shifting [POK23]. Shining [KHKR11]. Ships [HQT^{+}21]. Shock [CCL^{+}22, Erl07]. shooting [HHC^{+}19]. Shoots [LKM^{+}23]. Shortest [CDY23, DLP^{+}23]. Shot [CWL22, AWL15, BGK16, BKGK17, BBB^{+}10a, HZP^{+}22, XNY^{+}16]. Shots [ASN^{+}20, JRT^{+}15, LWCT14]. shoulder [HOKP16]. SHRED [JHR22]. Shrink [JHS^{+}23]. Shutter [JGN16, RAT06]. side [XFZ^{+}09]. SIERE [CSAP21]. Sifting [BBPA15]. SIGGRAPH [Spe03]. Sight [RSM^{+}23, CSHH21, CWK^{+}20, HOZ^{+}19, IH20, LWO19]. sign [TTWM14]. signal [BWS^{+}21, RH04, RTD^{+}10, WYY^{+}14]. signal-processing [RH04]. signals [CH05, PMHD19]. signatures [ACOH^{+}18, SZC^{+}22]. Signed [LBB22, BR21b, VSJ22, ZDI^{+}15]. silhouette [RSH^{+}05a, SCH03]. Silhouettes [JHR^{+}15, KDMF03, RDI10, VBMP08, WL16]. silicone [AMG^{+}18, ZKBT17]. Silly [FLGJ19]. silviculture [MHS^{+}19b]. SIMBICON [YLvdP07]. Similar [OCNG21, SPG^{+}23, BDG15, Ros20]. Similarity [CZ17, HNO^{+}23, LLN^{+}14, BB15, BD02b, DAB15, GCO06, GAGH14, GvdBL^{+}12, KvKSHCO15, LMS^{+}19, LKS15, SMGE11, ZRB14]. Simit [KKRK^{+}16]. Simple [BR94, Dav20, FM84, LR90, LR91, LKF12, MD94, SO92, TPP^{+}11, TM84, CPSS10, Gal99, GKS02, HRH^{+}13, LP02, SSJ^{+}11, TSG^{+}14, VMTF09, YLvdP07, YZ04]. simplest [PR97b]. simplex [FL16]. simplexes [DeR88]. Simplicial [JSP17, PBCF93, CSZ16, ETK^{+}07, FLSG14, GD02, MZD05, MB12, ZQC^{+}14, dGAOD13]. Simplicity [EM90, FLB16, PSBM07]. Simplification [LGC^{+}23, ABA02, CHPR07, DSSC08, DDSD03, GPW^{+}17, GZ05, LT00, LWH15, LXFH15, OL03, Pel05, SCF^{+}04, SAMWL11, WYY^{+}14, YLH18, ZG02, ZCLJ20]. simplify [SSISI16]. Simplifying [WM03]. Simulated [XBS^{+}22, ZYM^{+}23, CKJ^{+}11, DH96, FBH21, HRL15, HMLL14, MPP11, PGH^{+}22, SH08, WGH20, WGH21, YCBvdP08]. Simulating [BWRB05, CSAP21, CWSO13, FCK22, JGC^{+}15, KJM08, LDHM16, LL23, LGF04, MM06, PGCG23, SSC10, SKL07, TOK14, WM14, ZBG15b, FLGJ19, FMB^{+}17, FBGZ18, GTJS17, HMP^{+}20, SSJ^{+}20, SSBD03, SXH^{+}21, YLNP12]. Simulation [AGP^{+}20, BCK^{+}23, BSL^{+}16, BK16, BME22, CPV^{+}23, CFP^{+}21, CLT^{+}22, CNZ^{+}22, CZY17b, DKHS14, DYZ^{+}23, EM90, GDAB^{+}17a, HWZ^{+}14, HH16, HWP^{+}23, JWL^{+}24, KLL^{+}07, KKRK^{+}16, LJL23, LDW^{+}23, LD23, LNZ^{+}23, LWP^{+}23, LYWG13, LBK17a, MSL^{+}24, NBHSB22, PGG^{+}23, PMS12, RLY^{+}14, RLSÖ^{+}22, SPF^{+}23, SRX^{+}23, SLST14, SDK18, SS00, SQSL22, WVY^{+}22, XIM18, YR23, ZDF^{+}22, ZWHB22, AR15, BGOS06, BGFAO17, BME21, BHW16, BML^{+}14, BB12, BBB10b, BDW13, CMT^{+}16, CXW^{+}05, CKIW15, CSvRV18, CLC^{+}20, CAR^{+}09, CM11, CZY17a, CLMMO14, CQD^{+}18, CBK20, CGG^{+}17, CLSK21, DBD16, DLF12, DWK^{+}22, DLL^{+}18, FLLP13, GDAB^{+}17b, GKDS12, GHB^{+}20, GNS^{+}12, GHF^{+}07, GITH14, GKS02, GHZ18, HMS05, HPJ12, HBP^{+}21, HTC^{+}14, HW15, HW16, HXZW20, HG09, HHM19, HIK^{+}20, IGLF06, IZE^{+}21, JP02, JP03, JWJ^{+}14, KHD14, KSNG17, Kau18, KGBS11, KUJH21, KTJG08, KJ09]. simulation [KySK10, KP11b, KD13b, KGH^{+}14, KP03, LKL^{+}22, LST09, LPLL19, LSD^{+}22, LLJ^{+}11, LDN^{+}18, LCD^{+}20a, LTT^{+}20, LMLD22, LBOK13, LMH^{+}15, LBK17b, LCT19, LSZ^{+}22, LLK^{+}20, LLDL21, MKB^{+}10, MSW^{+}09, MBF04, MYH^{+}10, MC11, NGCL09, NSO12, NZWC20, NB11, NØ13, OPOD10, OKRC10, PBH15, PDZ^{+}18, PTC^{+}10, QSH^{+}15, RSM^{+}10a, RNGF03, RK13, SSB^{+}15, SML^{+}12, SHD^{+}18, SLF08, SABS14, SWL11, SHM22, SMD^{+}15, SOHK16, SG11, SSBL^{+}22, SSC^{+}13, SKP08, SJLP11, TGK^{+}17, TJM15, TWL^{+}18, TBBC^{+}22, TBV12, TJ08, UHT17, UPSW16, VMTF09, VKS^{+}14, VK16, WY16, WMB19, Wan21, WPLS18, WRK^{+}10, WLP16, WFS22, WMW15, WZL^{+}20, WWW22, XCW^{+}20, XTZ^{+}21, XWWZ22, YJL^{+}16, YLX^{+}15, YCR^{+}15, ZNT18, ZB13, ZSTB10, dSAP08]. simulation-ready [ZB13]. Simulations [AONA22, JW23, MSQ^{+}18, FFWL^{+}22, Thu17a, ATW13, ATW15, BP08, BSG12, HTYW22, HLY^{+}21, ISF07, Kim10, LJS^{+}15, LAD08, MBT^{+}15, NRC21, PSE03, RPC^{+}10, SDK21, Thu17b, TMPS03, YCL^{+}17, YSC^{+}18]. Simulator [YXW^{+}23, AB20]. simulators [RLR^{+}21]. Simultaneous [BJTK18, NLW^{+}16, HVTG08, ISSI16, PTH^{+}17, SKV^{+}12, TFK^{+}03, VSK^{+}17]. Single [BMBRD24, BBC^{+}23, CWW^{+}12, DAD^{+}18, Fat08, GHCG17, GXY^{+}17a, GLT^{+}23, HMLL15, HWK15, LJZ^{+}23, LQGY24, LOW18, LXL^{+}23, NZV^{+}11, PNTK23, SYSP14, SBT^{+}19, TFX^{+}08, TCS^{+}23, VKM^{+}23, WZHB09, WYL^{+}20, WS17a, WZ22, YPA^{+}18, ZZT^{+}21, ZK22, BGK16, BKGK17, BSW13, BCRK^{+}10, BBB^{+}10a, CLS^{+}15, CSW^{+}16, CZS^{+}13, DMIF15, DTPG11, DSC^{+}20, EKD^{+}17, FSH^{+}06, GSY^{+}17, GSZ^{+}18, GXY^{+}17b, GLT^{+}21, GSLM^{+}08, HSW^{+}17, HLV^{+}17c, JTC09, KSES14, KYC^{+}17, LLLL21, LAGP09, LDPT17, LXR^{+}18, LKZ^{+}20, LK20, LKK^{+}21, LAZ^{+}22, MSS^{+}17, MDB^{+}19, PSB^{+}08, SJA08, STXJ15, SHZ^{+}20, SPDF13, SRNN05, SZLG10, WGJ^{+}18, WTL05, WSXC16, WSZ^{+}18, WZC12, WSTS08, WS17b, ZCB^{+}22]. single-camera [WGJ^{+}18]. Single-Image [TCS^{+}23, ZK22, DAD^{+}18, LQGY24, GLT^{+}21]. single-photo [GSZ^{+}18]. Single-photon [LOW18]. single-shot [BGK16, BKGK17, BBB^{+}10a]. Single-View [LJZ^{+}23, YPA^{+}18, CWW^{+}12, HMLL15, HWK15, DSC^{+}20, LAGP09, SHZ^{+}20]. Singular [WZX^{+}23, KABL14]. singularities [SSC18]. Singularity [JCY23, LFZ^{+}23, LZC^{+}18, FTD21, LLX^{+}12]. Singularity-constrained [LZC^{+}18]. singularity-restricted [LLX^{+}12]. sites [KGFF14]. six [KKB^{+}11, YZX21]. six-user [KKB^{+}11]. Size [LHJ^{+}14, HCOB10]. Sizing [Bae82]. Skaterbots [GPD^{+}18]. Skeletal [HTCH15, JS11, LD14, LH16, LAH^{+}21, LYO^{+}10, WLH^{+}13]. Skeletal-Surface [HTCH15]. Skeleton [ALL^{+}20, ATC^{+}08, KWS^{+}23, QLH^{+}22, SAA^{+}21, ULP^{+}15, BAS14, CGC^{+}02, HWCO^{+}13, KP11b, LYWG13, TZCO09]. Skeleton-aware [ALL^{+}20]. Skeleton-Consistent [QLH^{+}22]. skeleton-driven [CGC^{+}02, KP11b, LYWG13]. skeleton-mesh [BAS14]. Skeletonization [BR21a]. Sketch [ATW^{+}17, ASK^{+}22, CNX^{+}08, ERB^{+}12, GLC^{+}23, LABS23, ST14, ST16, TPSHSH13, XSL^{+}22, ZGXF23, ZIH^{+}11, BDM^{+}20, BB22, CBL^{+}16, DS15, EHA12, FPBCO20, LPL^{+}18, LWH15, LCL^{+}22, NSACO05, PHS^{+}18, SAN23, SSISI16, SSII18b, XYH^{+}21, XCF^{+}13, YVG20, YLL^{+}22, ZLW^{+}18]. Sketch-based [ATW^{+}17, CNX^{+}08, ERB^{+}12, GLC^{+}23, TPSHSH13, ZIH^{+}11, CBL^{+}16, DS15, LPL^{+}18, LCL^{+}22, NSACO05, PHS^{+}18, XYH^{+}21, XCF^{+}13]. Sketch2CAD [LPBM20]. Sketch2Photo [CCT^{+}09]. Sketch2Pose [BB22]. Sketch2Scene [XCF^{+}13]. Sketches [IBB15, MNB23, PMKB23, GHL^{+}20, HFL14, KH06, LZ04, LRS18, SBSS12, SLZ^{+}13, TD16, XCS^{+}14, YCYW20, YAB^{+}22]. SketchFaceNeRF [GLC^{+}23]. SketchHairSalon [XYH^{+}21]. SketchiMo [CBL^{+}16]. Sketching [BSM88, CKX^{+}08, JHR^{+}15, KG05, SSII18a, BSM^{+}13, BDM^{+}20, GRGC15, HGY17, JZH07, KWL^{+}21, LPL^{+}17, LPBM20, MSSG^{+}21, NGDA^{+}16, PKM^{+}11, PSE03, SLWF14, TBvdP04, VPB^{+}22, WTBS07b]. SketchPatch [FPBCO20]. sketchy [SBHH16]. skill [PGH^{+}22]. Skills [HL14, ZYM^{+}23, ZPYX23, CBYvdP08, CKJ^{+}11, LLLL21, LH18, PBvdP15, PBvdP16, PBYV17, PALvdP18, PKM^{+}18, YCBvdP08]. Skin [CBKM15, KWS^{+}23, NFA^{+}15, PCB23, BBN^{+}12, DWd^{+}08, LSNP13, LZT^{+}19, PH06, PH08, SMP03, TOS^{+}03, VBG^{+}13, WWY^{+}13, WMP^{+}06]. skin-frame [WWY^{+}13]. Skin-Screen [PCB23]. SkinMixer [NPC^{+}22]. skinned [BBJP12, FKY^{+}10, LMR^{+}15]. Skinning [BL18, BZC^{+}23, JT05, LJG14, JBK^{+}12, JZvdP^{+}08, KCŽO08, LD12, LD13, LH16, LL19, LVGO21, MZS^{+}11, MK16, SZT^{+}08, VBG^{+}13, VGB^{+}14]. Skins [MHCT23, MG03]. Skipping [KJ09, LNLB16]. Skippy [KYC^{+}17]. skull [KHS03]. Sky [DII23, TSL^{+}16, HW12, TYS09]. sky-dome [HW12]. skydome [KKN^{+}14]. SkyFinder [TYS09]. Slang [HFF18]. SLANG.D [BWL^{+}23]. slender [RLR^{+}21]. Slice [Les20, CWTW17, Ng05, OP11]. Sliced [SGSS22, BC19]. Slices [MSM11]. Slicing [AHL17a, AHL17b, ERP^{+}19, YAV^{+}20]. slide [KCSC10]. sliding [BWKS11]. slightly [ERP^{+}19]. Slippage [AVB^{+}23, ZYQ^{+}14]. Slippage-free [ZYQ^{+}14]. Slippage-Preserving [AVB^{+}23]. Slope [LZHJ20]. Slope-space [LZHJ20]. Small [DFM88, VPR19]. Smart [RO94, XFAT12, ZCC^{+}12]. SmartBoxes [NSZ^{+}10]. smartphone [VKB^{+}18]. SMASH [MTM16]. Smith [HHdD16, WJF^{+}22]. Smocking [RSSH24]. Smoke [BLDL21, PM17b, RNGF03, Thu17a, WPS14, CKP^{+}16, CT17, CLZ^{+}22, FL04, FN20, GSLF05, LGF04, PM17a, SDK21, SRF05, SABS14, SY05, Thu17b, TMPS03, WP10, YCZ11, ZRL^{+}08]. Smooth [DFZ^{+}17, DFYL19, LFP21, LD12, LM91, PR97a, Pet01, RHW94, RLU95, BHK14, HTWB11, KLS03, KP03, MEM^{+}19, ML22, Mal89, OBW^{+}08, WP06, WWT^{+}06, YAB^{+}22, ZWL^{+}18]. smooth-shaded [OBW^{+}08]. Smoothed [ERT14, KS10, TJM15, WDK^{+}21, WAK20]. Smoothing [LZH^{+}20, Pet95, SdS02, SGWJ18, BHY15, FYW^{+}18, JDD03, KEE13, PR97b, XLXJ11]. Smoothness [SJWG20, LWL^{+}09, PKD^{+}19, YZ04]. SmoothSketch [KH06]. SMPL [LMR^{+}15]. snakes [LLZM10]. Snap [GSKJ03, ASF^{+}13]. Snap-together [GSKJ03]. SnapCut [BWSS09]. snapping [ASF^{+}13, LSTS04]. Snapshot [CHWH17, HLV^{+}17c, JBY^{+}19]. Snapshots [KF93, SCH^{+}16]. SNeRF [NPLX22]. Snippets [LY23]. snow [GHB^{+}20, SSC^{+}13]. soap [DBWG15, HIK^{+}20, ISN^{+}20]. Sobol’ [PCI^{+}21]. soccer [HHC^{+}19]. social [APS^{+}14, MWHL21]. SofGAN [CLX^{+}22]. Soft [AASP17b, GPHSH19, GTB15, LAA^{+}05, PZ17, TTL12, WAC07, AASP17a, AOP^{+}18, AAM03, BBO^{+}09, FTP16, GWP^{+}19, JL11a, KPMP^{+}17, LYWG13, MZL^{+}17, MDZ^{+}21, MWR12, MA07, PRWH^{+}18, RWS^{+}06, WWY^{+}15, YKZ^{+}22]. Softshell [SKK^{+}12]. Software [Fol86a, Fol86b, Fol86c, Mai92, WW82, KKSS18]. SOHO [LF08]. solar [KKN^{+}14]. Soli [LGK^{+}16]. Solid [BN90, CCK92, KFCO^{+}07, LD23, LWF^{+}22, MC11, NY94, RYPZ23, Roc89, RLZ^{+}21, TB22, XLYJ23, ANZS18, ABA02, BBB07, CH02, CS09, CWSO13, CDM^{+}02, DF88, DA21, DZCJ21, HLW^{+}12, JDR04, KRD^{+}12, LD11, LIY^{+}22, LLJ^{+}11, LDHM16, LLDL21, NGL10, RS98, SS10a, TOII08, TBBC^{+}22, TLK16, WZYG10, ZGZJ16]. Solid-Fluid [LWF^{+}22, XLYJ23, RLZ^{+}21, BBB07, HLW^{+}12, TLK16]. solid-liquid [CWSO13]. Solids [CCL^{+}22, KS95, LFP21, AD03, FLGJ19, FQL^{+}20, FGBP11, Lee05, LB18, MKB^{+}10, MAKWL22, PKA^{+}05, RMSG^{+}08, YJL^{+}16, ZSTB10]. Solution [SAZK06, BRB^{+}19, YWH13]. Solutions [GM84, OF01, DJ17, DJ18a, HDA17, RMOW20, SHW19]. Solver [HCH22, LL23, PM17b, QLY^{+}23, TB22, XIM18, ZXS^{+}23, ATW15, BDCDA11, BBG12, CWZ^{+}21b, DBDB11, GHB^{+}20, JCW09a, Jia21, LBB17a, LDN^{+}18, LMAS16, NNC^{+}20, PM17a, SBZ09, TB20, XSH^{+}20, ZNT18, dGWH^{+}15]. Solvers [GPB^{+}19, MHZ^{+}21a, XWX^{+}22, BFGS03, ZBG15a]. Solving [FH97, PKHK15, SHG^{+}22, Hol18, JASR99]. Some [CI97, GM84]. Sonar [RKB^{+}23]. Sony [KCSG18]. sort [CTM13, KC21]. sort-based [CTM13]. sorting [Ada21]. Sound [LFZ15, SM17a, XAW^{+}23, ACSM12, CRS^{+}16, CAJ09, CJ11, CZJ12, CRG^{+}20, CLG^{+}16, CQD^{+}18, DRW^{+}14, DYN03, DLL^{+}15, JBP06, JLWM22, LAJJ14, LJ14, LCT19, MRA^{+}13, MYH^{+}10, RSM^{+}10a, RS14a, RS18, RYL13, SMM14, SM17b, SMC21, SJM17, WQLJ18, WOD09, YI17, YMR^{+}13, ZCT16, ZRSM18, ZJ10, ZJ11, ZHHZ20]. soundbanks [ZJ10]. Sounding [MYH^{+}10]. sounds [AJM12, BLT^{+}15, BDT^{+}08]. soundscapes [ZHHZ20]. soup [SOS04]. soups [BDS^{+}18]. Source [CM21, SM17a, GTHD03, GGHS03, MRA^{+}13, SM17b]. Source-Specific [CM21]. Sources [NON85, OF01, CRG^{+}20, CDP^{+}14, JBP06, MLR^{+}14, RSM^{+}10a]. SP [LLHF21]. SP-GAN [LLHF21]. Space [ARMCO23, BYG96, BBC^{+}23, BYRN17a, EK98, GRGC15, HB23, HC86, LLKP11, LHdG^{+}14, Pet89, SAL^{+}08, Shn92, SLF22, TLG17a, TBTA^{+}24, WLX^{+}18, YSC^{+}23, YZN^{+}22, ZIT^{+}18, AB89, AW20, ACP03, AP08, ATDP11, BWC^{+}23, BWDL21, BSW02, BYRN17b, BKCO16, BCWG09, BBB^{+}14, BME21, CBD13, CLW16, CGZ08, CÖS19, CJM21, DCD15, DHC^{+}21, HPJ12, HB21, HMT^{+}12, JL11b, JTL^{+}12, JKH^{+}22, JTSW17, KHD14, KSHG18, KMP07, KWB^{+}15, LAKL11, LH06a, LSCO03, LC15, LKG^{+}03b, LZHJ20, MVH^{+}17, MMG06, MHC^{+}16, NBLCO20, RHSH18b, RH02, RMOW20, RJN16, SNM^{+}13, SXZ^{+}17, SGM^{+}16, SvKK^{+}11, SMD^{+}15, SAZK06, SY21b, SZLG10, TEG18, TMDK15, WCPM18, WAKB09, WYXJ21, Wym05, XB16, YLB^{+}22, YYPM11, YYW12b, ZSSJL20, ZYZ21, TLG17b]. Space-Filling [Shn92]. Space-Time [ARMCO23, GRGC15, LLKP11, LHdG^{+}14, SAL^{+}08, ZIT^{+}18]. space-warp [LKG^{+}03b]. spaced [Gos00]. Spaces [FSRS22, KP92, RFW^{+}23, DCP14a, HRV97, KDH22, Lip12, OKH^{+}17, SHP04, SJA^{+}20, TGY^{+}09, VABW09, ZCC16, dASTH10]. Spacetime [PM17b, SLS^{+}12, ZSCS04, HSvTP12, PM17a, SvTSH14, SAJ21, XWW^{+}14]. SPAD [CSHH21, SZD^{+}20]. SPAGHETTI [HPG^{+}22]. Spark [FH11]. Sparse [ASGCO10, BFGS03, CBYJ23, CKT^{+}23, FGBP11, HSB^{+}12, HSX^{+}22, HSH20, HJM^{+}22, KLM24, NVW^{+}13, NSF12, PYA^{+}24, QRL^{+}23, TUGM22, WLY^{+}16, WLZ^{+}21, YZH^{+}23, ZCT22, ZCD^{+}16, AGL^{+}17, ALS^{+}18, BBN^{+}12, CLZ^{+}22, FOL^{+}21, HLSO12, HDA17, HKA^{+}18, KWB^{+}13, KSA13, LLDD09, LD13, LFO^{+}22, LMB14, Mus13, ODAO15, RTK^{+}15, SvTSH14, SABS14, SNF05, SL17, TZK^{+}11, TKKT12, TS12, XYJ13, XSHR18, XBS^{+}19, dAST^{+}08]. Sparse-as-possible [ZCD^{+}16]. Sparse-View [HSX^{+}22]. Sparsely [HWZ^{+}14, LHZ^{+}18]. SparsePoser [PYA^{+}24]. Sparsity [HTS^{+}22, SHD^{+}14]. Sparsity-Specific [HTS^{+}22]. Spatial [BSB16, CSSL21, GRS^{+}17a, HKT10, KPACO22, LLWD14, BSB17, CKMR^{+}21, DLX^{+}21, DH06, GB08a, GAB20, GRS^{+}17b, LBJK09, LH06b, LKG^{+}03a, LGX^{+}13, WLW^{+}19, YI17, ZYSK21]. Spatial-spectral [LLWD14]. Spatial-temporal [CSSL21, DLX^{+}21]. Spatially [MXZ^{+}23, WK21, BJ10a, BATU18, DWP^{+}10, DTPG12, DCP^{+}14b, GWN^{+}03, GCH^{+}19, HMP^{+}08, JAG18, LXR^{+}18, MAG^{+}09, PFB^{+}20, SSJC22, TDG18, TFK^{+}03, WRG^{+}09, XDPT16]. spatially-aware [TFK^{+}03]. spatially-correlated [GCH^{+}19, JAG18]. Spatially-Varying [MXZ^{+}23, DWP^{+}10, DTPG12, LXR^{+}18, MAG^{+}09, WRG^{+}09, XDPT16]. Spatio [DLW^{+}22, LYC^{+}22, ZM13, BH21, BBK^{+}15, GBAM11, KZP^{+}13, KKW21, MAC22, VBK05]. Spatio-Angular [DLW^{+}22, KZP^{+}13]. Spatio-Temporal [LYC^{+}22, ZM13, BH21, BBK^{+}15, GBAM11, KKW21, MAC22, VBK05]. Spatiotemporal [PKC^{+}17, YPG01, ASK^{+}12, HLR^{+}14]. Spatiotemporally [LYO^{+}23]. SPCBPT [SLW22]. speaker [EML^{+}18, NKAS08, YCL^{+}20, ZHS^{+}20]. speaker-aware [ZHS^{+}20]. speaker-independent [EML^{+}18]. Speaking [SDO^{+}04]. Spec2Fab [CLD^{+}13]. Special [BG89b, Fol86a, Fol86b, Fol86c, FGN84, Pha18, Ros94, Sto92, WKR99]. Specialized [RYW^{+}22]. species [TGK^{+}17]. Specific [CM21, DMZ^{+}17, HTS^{+}22, ALLD17, SHP04]. Specification [DFM88, GM84, Hud94, Jac86, JDH^{+}22, RvE93]. specifications [CLD^{+}13, PYB^{+}16]. specified [HFM^{+}10, LZZ^{+}21, WPC^{+}14]. Specifying [Van82]. Speckle [ABGL21, SY22, BAGL19, BGL20, Par17]. spectra [BDM09, SJ17, WPC^{+}14]. Spectral [BBG24, CRB23, CRCM23, DBG^{+}06, FHL^{+}18, GO17, HZM^{+}08, IRN^{+}22, KBC^{+}13, KHLN17, LHJ^{+}14, LJO19, ÖAG10, POK23, WBCPS19, YM16, AHD15, AAMSB20, BCG05, CLJL20, CJN^{+}17, CLSK21, FMR20, HW12, KYS^{+}15, LLWD14, PMHD19, RZK11, SvKK^{+}11, SS21, SLMR14]. spectroscopy [KRD^{+}12]. Spectrum [ZDT^{+}23, BWWM10, Fre16, ZHWW12]. Specular [CA00, FHG^{+}23, IM12, JM12, KYYL08, LZHJ20, SJR18, WHY20, XH18, YHJ^{+}14, YHMR16, YHW^{+}18]. Specularly [RT90]. Speculative [AVGT12]. Speech [AGL^{+}22, YCL^{+}20, CTFP05, CB05, EML^{+}18, EGP02, LCC21, OLSL16, TKY^{+}17, ZXL^{+}18]. speech-driven [CTFP05]. Speed [GHCC88, KRF^{+}18, PSBM07, TAH^{+}04]. Spelunking [SJ22b]. SPGrid [SABS14]. SPH [AIA^{+}12, AAT13, BGI^{+}18, GPB^{+}19, GHB^{+}20, HWZ^{+}14, JWL^{+}24, JZW^{+}15, LXY^{+}23, PICT15, RLY^{+}14, SB12, SP09, FFWL^{+}22, WHK17, WK21, YJL^{+}16]. SPH-Based [HWZ^{+}14, LXY^{+}23, JZW^{+}15]. Sphere [HH16, TGBE16, TPT16, BO04, LF08, LLHF21, VPR19, TGB13]. sphere-guided [LLHF21]. Sphere-Meshes [TGBE16, TPT16, TGB13]. sphere-tree [BO04]. Spheres [MSCG23, Hub96, SHWP09]. Spherical [AKL17, BXH^{+}18, BF01, CCW93, KISS15, KCMP23, PH03, SBN15, DHB17, GCP^{+}10, GFT^{+}11, GGS03, HKWB09, HIK^{+}20, KSH10, KH10, KWN^{+}17, LKK^{+}16, MWM08, PAAG21, RWS^{+}06, SHL^{+}17, TAV^{+}10, TGB13, TS06, TFG^{+}13, WR18, XSD^{+}13]. Spin [BWBSH14, CPS11]. Spin-it [BWBSH14]. spinnable [BWBSH14]. SpinVR [KDMW17]. Spiral [CLSK21, OGN^{+}23, ZZX^{+}18]. Spiral-spectral [CLSK21]. spirals [ZGH^{+}16]. splash [YCYW20]. splashes [HQT^{+}21]. splashing [GB13]. Splatting [KKLD23, GLA^{+}19, LSR18, WFH^{+}07]. Spline [BS88, BS90, BL18, CCL^{+}22, Fol87, Joe90a, KPP17, Kla91a, LT08, RLU95, SDG^{+}19, Sei93, SYSP14, YJY23, vOV96, BA83, CG89, PU06, SCF^{+}04, WPL06, ZCX^{+}22, GBK05]. Splines [BBB^{+}93, BF01, DB88, DKA23, FB95, Joe90b, Las90, PP93, Pav83, Pra89, TB87, Yuk20, vOV96, vW84, BB83, CZ17, CLS85, Coh87, FW12, FSH11a, HP04, Joe89, KA08, LT09, LJG14, Pot91, SZBN03, YHB05].

Split [QCOS23, WTGT09].

Split-Lohmann [QCOS23]. Splitting [XLYJ23, RGH^{+}22, TBV12, VK16, YWVW13]. Spoke [MEA^{+}18]. Spoke-Darts [MEA^{+}18]. SpongeCake [WJHY23]. sports [WGH21]. SPOT [BC19]. spots [DFW20]. spray [IGP^{+}17, NØ13]. spray-on [IGP^{+}17]. Spreadsheet [Hud94]. spring [LBOK13, SLF08]. Sprite [ZWL22, ZWL22]. Sprite-from-Sprite [ZWL22]. Sprites [WCZ^{+}22, ZSAF21]. square [CLC^{+}20, OCNG21]. squared [LLZ^{+}20, WPL06]. Squares [BIW93, DMZ^{+}17, LZH^{+}20, MHZ^{+}21a, FCOS05, HFG^{+}18, LPRM02, MZPS21, SMW06, WJL^{+}20]. St. [BJ05]. Stability [YKGA17a, LLK^{+}15, SMZ^{+}14, YKGA17b]. stability-based [SMZ^{+}14]. Stabilization [CK20, TWLT19, BB14, FL11, GF12, Kop16, LGJA09, LGW^{+}11, LYTS13]. Stabilized [CCWL18, WMB19]. Stabilizing [DLK18]. Stable [CK02, DJBDT10, ETK^{+}07, Hob91, HCLK24, SH23, SDK18, SSK05a, TNGF15, WYW23, dASTH10, FTP16, LKL^{+}22, MLB16]. Stack [WSP^{+}23]. Stackabilization [LAZ^{+}12]. stacking [GBF03]. stacks [CKS18]. Stage [LYC^{+}22, QZZ22, YNK^{+}22, ALY08, HTYW22, SXZ^{+}20, ZLW^{+}18]. Staggered [HLW^{+}12, KSJP08, XCW^{+}20]. staggered-tilted [XCW^{+}20]. staging [ZCB^{+}22]. Stair [KTBV16]. stand [PWLSH13]. Standard [UTB^{+}19, RFWB07]. Star [MSS^{+}19, SPO10, KS04a]. Star-contours [SPO10]. Star-shaped [MSS^{+}19]. starline [LGA^{+}21]. Stars [SMGC23]. state [OKH^{+}17]. Static [FV96, SPV^{+}16, FKY^{+}10, HLV^{+}17c]. stationary [AWL15, AIH^{+}08, RCOL09, ZZB^{+}18, MFR^{+}10]. Statistical [KV05, MA06, SY22, Bel18, CH07, GGY18, GMP^{+}06, GvdBL^{+}12, LWS02, LWL17, LCT19, WMC11]. Statistics [AKG^{+}23, BAGL19, BGL20, Fat07]. Steady [RV11, DHL14]. Steerable [AS02]. steering [CAR^{+}09, OPOD10]. steganography [PHN^{+}12]. Stein [GIGM22]. Steklov [WBCPS19]. stelaCSF [MAC22]. Stencil [LLJ^{+}23]. Step [RY92, APH^{+}03, LGL^{+}19, WSM11]. step-by-step [APH^{+}03]. step-edge [WSM11]. Stepping [WLF^{+}20]. steps [KJ09, NJJ21, TJ08]. Stereo [HNH19, PMGD21, WF96, ZTF^{+}18, AWGB04, BGK16, FKN17, HGG^{+}11, KDR^{+}16, KDW^{+}17, KKW20, MCE^{+}17, VPB^{+}09b, WZMM22, WSVT13, ZJY^{+}21]. stereo-to-multiview [KDW^{+}17]. Stereological [JDR04]. stereoscope [HCW15]. Stereoscopic [DLP^{+}23, KLKL13, LvBK^{+}10, DMHG13, KKB^{+}11, LHW^{+}10, LSC^{+}12, NFL12, OHB^{+}11, TDM^{+}14]. Stereoscopizing [LMY^{+}13]. stereoscopy [KHH^{+}11]. Sticky [vOV96, FQL^{+}20]. Stiff [CSAP21, PAK^{+}19, LKL^{+}22, MSW14]. stiffly [MLT17]. Stiffness [FHXW22, WYW23, VMTF09]. still [HHV^{+}21, XWL^{+}08]. stills [OEE^{+}18]. stippling [DSZ17, SKB^{+}21]. Stitch [WGF^{+}18, WSY19, YKJM12]. Stochastic [Coo86, CHPR07, GKHH12, HJ09, LSD^{+}16, Lew87, Özt16, SJ22a, VR94, CGZ^{+}05, GGY18, JHY^{+}14, LAKL11, SK13, YIC^{+}10, Pav90, WP90]. stochastically [RMGH15]. stochastically-ordered [RMGH15]. stock [KSES14]. Stokes [DWS^{+}20, LBB17a]. stone [SZB18]. Stop [AJS20]. Stop-motion [AJS20]. Storage [WHHY20]. Store [Wes88]. Storing [SW85]. Stormscapes [HMP^{+}20]. storyboarding [GCSS06]. storytelling [LLHY22]. straight [MSW^{+}09]. strain [PBH15, WOR10]. Strains [WMB21]. Strand [HWP^{+}23]. Strand-Based [HWP^{+}23]. strands [SJLP11]. strategies [SK13, WGH21, YYVY21]. stratified [ZD20]. streaks [GN06]. Stream [SDK21, ZZC^{+}22, ATW15, BAM14, BFH^{+}04, GLT^{+}21, HZG09, HHN^{+}02]. Stream-guided [SDK21]. stream-processing [HHN^{+}02]. streamable [CCS^{+}15]. Streaming [HSV^{+}22, ILSS06, KH08, KDMW17, KLHG09, MVD^{+}18, SBZ09, TDL^{+}18]. streams [AMN03]. Street [KCSC10, CEW^{+}08, XFZ^{+}09]. street-side [XFZ^{+}09]. strength [FZZ^{+}20, LSZ^{+}14, SVB^{+}12, ZLB16a]. Stress [MHS^{+}19a, SVB^{+}12, MIB15, PRZ17, PNH^{+}14]. stressful [MIWB02]. stretch [GWP^{+}19, WTSL08]. stretch-sensing [GWP^{+}19]. Stretchable [GPHSH19, JS11]. stretches [XSZB15]. Stretching [KySK09]. Strict [LZ14]. String [KMM17b, KMM17c]. Strip [VVHSH22, WWSP23, CK14b, LPC^{+}23, MS04]. strip-based [MS04]. Stripe [KCPS15, MDH^{+}23]. StripMaker [LABS23]. strips [CK14b, TISM16]. Stroke [BLAE22, LYFD12, VLV^{+}21, XXK^{+}06]. Stroke-based [BLAE22, XXK^{+}06]. StrokeAggregator [LRS18]. strokes [HTER04, KMM^{+}02]. StrokeStrip [VLV^{+}21]. StrokeStyles [BLAE22]. Strong [GPB^{+}19, TB20]. strongly [FQL^{+}20]. Structural [LF02, LLN^{+}14, WSW^{+}12, ALX^{+}14, BSFG09, FSH11b, IOOI05, LSD^{+}16, LLW17, PMW^{+}08, SVB^{+}12, SKAG15, ZPZ13]. structurally [DLL^{+}15, WOD09, ZCT16]. structurally-sound [WOD09, ZCT16]. Structure [CAO09, FMLW14, FvKBCO16, HGM14, KEE13, LCOZ^{+}11, LLR13, MDLW15, PQW^{+}08, SFCH12, TBTA^{+}24, XZW10, XYXJ12, YML^{+}23, ZXTZ15, ZJMB12, CMZP14, DH06, GPW^{+}17, HYG^{+}13, HKAK14, JAM^{+}10, JBX^{+}20, LDHM16, LGF04, MPO21, NGH04, RGF^{+}20, SABS14, SYJS05, UMK17, WVJH17, WWL^{+}19, WYXJ21, ZLC^{+}13, YCZ11].

Structure-aware [CAO09, LLR13, PQW^{+}08, ZJMB12, WWL^{+}19]. Structure-based [XZW10]. structure-driven [HYG^{+}13]. structure-from-motion [CMZP14]. Structure-oriented [FvKBCO16]. Structure-preserving [KEE13, LCOZ^{+}11]. Structured [ARBJ03, GIZ09, Kau18, LN84, MDH^{+}23, SSII18a, AGS21, HDS^{+}18, KFWM17, LKK^{+}16, LBW^{+}14, MCT15, RGB16, RHDG10, SMCT18]. Structures [BTSB23, CQS^{+}23, DTPC23, GUPZ20, GJB^{+}20, JYW^{+}23, JWT^{+}23, PMLB22, SOG^{+}22, WKMH^{+}23, ZAB21, vOV96, BPK^{+}11, Boi84, CPSP21, DPW^{+}14, HSC^{+}22, JTSW17, JLBM05, KPWP17, KL22, LSK^{+}06, LXC^{+}17, LCC^{+}18, LPC^{+}23, LYO^{+}10, MLB16, PKLI^{+}19, PLW^{+}07, QJ21, RLR^{+}21, SZB18, STK^{+}14, SHOW02, SFG^{+}13, Ter18, WWY^{+}13, YCC17, ZHRB13, dGAOD13, vKXZ^{+}13]. Study [CMS95, LJGH11, RGSS10]. stuffing [LS07]. stunts [TGLT14]. Style [AONA22, BSM^{+}13, GMHP04, HPP05, HLV^{+}17a, IWHH20, JPL22, LZCX19, LHLF15, SPB^{+}14, SLF22, XLZ^{+}10, YJLL22, ZTD^{+}23, AWL^{+}20, APCO21, ALY^{+}21, CWZ^{+}21a, FTP03, GHBCO21, GAGH14, HLV^{+}17b, JCW^{+}21, KGS^{+}18, KAGS20, LJGH11, LHP05, LKS15, LKWS16, MBB12, NKAS08, PO08, SDKN18, SED16, SBLD15, WPP14, WYX11, XWCH15, XLLW20, YNS19, YM16]. Style-based [GMHP04, APCO21, GHBCO21]. Style-content [XLZ^{+}10]. Style-Defining [HLV^{+}17a, HLV^{+}17b]. style-synchonized [KGS^{+}18]. StyleCariGAN [JJJ^{+}21]. StyleFlow [AZMW21]. StyleFusion [KPACO22]. StyleGAN [AZMW21, ALY^{+}21, BAC^{+}23, GPM^{+}22, JJJ^{+}21, KPACO22, TAN^{+}21]. StyleGAN-Generated [AZMW21, KPACO22]. StyleGAN-NADA [GPM^{+}22]. Styles [YZX^{+}18, LP10, SHU^{+}16, YYL22]. Styling [CLX^{+}22]. Stylised [PAR21]. stylistic [CCL12]. StyLit [FJL^{+}16]. Stylization [BLAE22, DS02, FJL^{+}16, FPBCO20, GLZ^{+}21, LYFD12, MYC^{+}22, ZAJ^{+}15]. stylize [ZAJ^{+}15]. stylized [FJS^{+}17, KDMF03, LMPB^{+}13, NPLX22, PMA^{+}21, RTF^{+}04, TiABI07, Wam16, dSAP08]. Stylizing [BCK^{+}13, JST^{+}19, EBGB14, GLZ^{+}21, SLL^{+}21b]. Sub [NID20, CMSA20, HA18]. sub-grid [CMSA20]. sub-meshes [HA18]. Sub-Paths [NID20]. subband [LSA05]. Subdivision [AB08, Che92, CV20, DLG90, Gol85a, HLG^{+}22, Kla94, Lew87, LBHH23, Rap91, dGDMD16, BFK^{+}16, CADS09, DM13, HSH10, ISD04, KP07, KS98, KBZ15, Lev06, LYLL08, LJG14, LS08, LSNC09, MWM23, MRF06, MFR^{+}10, MP09c, Nas87, NLMD12, PO08, PR97b, PS04, PBW19, SW05, SJP05, VB06, VMW18, WP06, WWT^{+}06, ZHX^{+}07]. Subdivision-based [HLG^{+}22]. subdivisions [GS85, PVR18]. SubEdit [STPP09]. Subjective [ASN^{+}20, SY22]. Submaps [TZY^{+}23]. submillimeter [Wan21]. Submissions [Ols88]. Subspace [BJ10b, FW22, HTC^{+}14, HZ13, HSL^{+}06, KD13b, LGW^{+}11, MA07, NH22, PBH15, SS19, TMDK15, AKJ08, BJ05, MHR^{+}16, RCPO21, SSR20, SLW22, TOK14, WJBK15, WMW15, XB16]. Subspace-based [SS19, SLW22]. substance [NZWC20]. substrate [PH15b]. substructure [ZXC^{+}18]. Substructuring [PAK^{+}19, BZ11]. Subsurface [FHK14, DWP^{+}10, HFM^{+}10, PvBM^{+}06, STPP09, VKJ19]. Subtle [BMSG09, WRS^{+}12]. subtractive [MAYZ^{+}20, ZJ18, ZZX^{+}18]. successive [FZL^{+}15]. Suction [BCK^{+}23]. suggesting [LRFH13]. suggestion [CXY^{+}15]. suggestions [CK10, JTRS12, SSK^{+}17]. Suggestive [DFRS03]. Sum [MZPS21, BDD11]. Sum-of-squares [MZPS21]. summarization [DTP15, PCS23a, WWF^{+}10]. summation [ZB14]. Summed [NMLH14, NMLH11]. Summed-Area [NMLH14, NMLH11]. Super [BAC^{+}06, CBD13, NYY04, SZD^{+}20, GGY18, LEPM22, SDP^{+}18, SXZ^{+}20, WGDE^{+}19, XFCT18]. Super-helices [BAC^{+}06]. Super-resolution [SZD^{+}20, GGY18, LEPM22, SDP^{+}18, SXZ^{+}20, WGDE^{+}19, XFCT18]. supercompressed [KPM16]. superimposed [AYL^{+}12]. Superimposing [BI08]. superresolution [HLR^{+}14]. supersampling [DVC09, DEM96, YNS^{+}09]. SuperTrack [FBH21]. Supervised [YZX^{+}18, BMBRD24, CHY21, FBH21, HSG13, MCW^{+}21, SAN23, SSK^{+}17, ZWL22, ZCB^{+}22]. Supervoxel [HMM^{+}21]. Support [DWW^{+}18, AFR^{+}07, CK10, ISD04]. Support-free [DWW^{+}18]. supported [SFLM04]. Supporting [Hil86, JWT^{+}23, MHS^{+}19a, TLZ^{+}24, DPW^{+}14, LPS^{+}13, MIB15, VHWP12]. suppression [LSL^{+}18]. Supra [WWH04]. Supra-threshold [WWH04]. SURE [LWC12]. SURE-based [LWC12]. Surface [BI92, Bli82, CG89, CC23, DHB^{+}16, DNZ^{+}17b, DLG90, EC93, EK98, FNO89, FG90, FB95, GLL^{+}16, HWZ^{+}14, HOZ^{+}19, HH16, HTCH15, HM20, HCH22, JW23, JWL^{+}24, KM97, LXSW23, LZBCJ21, LGC^{+}23, LSSW19, LC96, MBT^{+}15, Mil87, PM05, SJ22a, SO92, SYSP14, TG17b, VBFG12, WWX^{+}22, WJHY23, XRW^{+}22, XWD^{+}22, YJY23, YXW^{+}23, YIC^{+}14, ZWK14, ZXZL23, ZZL^{+}23, ZXS^{+}23, Zyd88, dFP95, AMCO08, APL14, APL15, AAT13, AB20, ABA02, ACA^{+}19, ASL^{+}17, BUSB13, BHMK^{+}18, BHK14, BLN^{+}13, BHW13, BBB10b, CBCG02, CSPF12, CBI13, CMSA20, CPSP21, CKMR^{+}21, CZXL23, CMMK15, DBG14, DNZ^{+}17a, DTB06, DBG^{+}06, DCP^{+}14b, DZCJ22, EB14, FG14, GZ08, GWM^{+}08, GTR^{+}06, HTG14, HSTP11, HLZ10, HWW^{+}22, HNB^{+}06, HLZ^{+}09, HZ82, HGMRT20, JCW09b, JSMF^{+}18, KH13, KG06, LDK^{+}18, LDPT17, LKK^{+}18, LPL^{+}18, LF09, LTJ18, MCSK^{+}17, MFL17, McK87, MASS15, MBWB02]. surface [NGH04, OBS04, PIC^{+}21, PO08, PKG06, RAM^{+}21, RTD^{+}10, RLZ^{+}21, STJ^{+}17, SAPH04, SS10a, SSZCO10, SACO04, SLS^{+}07, SAL^{+}08, SC18b, SCGT15, SWW^{+}20, SKM10, SS11, TWBO03, TWGT10, TG17a, TCL21, VGB^{+}14, VPB^{+}09a, VMT06, WZT^{+}08b, WLZ^{+}09, WYY^{+}14, WJL^{+}20, WVJH17, WFH^{+}07, WPMR09, XDPT16, XZZ^{+}14, XWWZ22, YHZ^{+}14, YAB^{+}22, ZJ18, ZMT05, ZM11, ZGW^{+}13, ZQC^{+}14, ZBG15b, ZHCJ15, ZPKG02]. surface-based [PIC^{+}21]. Surface-only [DHB^{+}16, HM20]. surface-surface [CZXL23]. surface-tension-dominant [RLZ^{+}21]. Surface/Surface [YJY23]. Surface2Volume [ACA^{+}19]. SurfaceBrush [RRS19]. Surfaces [And82, AS21, AOCBC15, BIW93, BHN98, BS88, BS90, BSTY15, Che92, CGM91, DKA23, DWMG15, ESBC19, FGC23, Fil89, JCY23, Joe90a, JHR^{+}15, KPP17, KMM17b, LM91, LBHH23, LDW97, LC96, MHS^{+}19a, MDH^{+}23, MSS92, NPP22, RHSH18a, Rap91, RS14b, RNP^{+}22, Roc89, SB95, Sar00, SLM^{+}17a, SACO22, SG17, SJWG20, SY22, TBWP16, VVHSH22, WLJ^{+}22, War92, XWX^{+}22, AB89, ACXG09, AA09, AK04, ASGCO10, BX03, BW13, BMBZ02, BHLW12, BWWM10, BFK^{+}16, CI97, CS09, CPS11, DvGNK99, DJBJ19, EKS^{+}10, EC96, EB08, EMF02, FCOAS03, FLHCO10, GSC21b, GOMP98, GG07, GBK05, HSH10, HCJ19, KNBH12, KMM17c, KYYL08, KTT13, KCPS15, KLPCP18, KP03, LCCS18, LJJ^{+}18, Lev06, LFS16, LPL^{+}17, LB18, LPW^{+}06, LPS^{+}13, LJG14, LD89, LB06, LS08, LSNC09, LKYU12, MGA^{+}17, MV21, MLR^{+}22, MIB15, MRF06, MFR^{+}10, MAB^{+}15]. surfaces [Nas87, NISA07, NLMD12, PZ07, PCL^{+}12, PLPZ12, PBDSH13, PSF09, PKD^{+}19, PKPP21, POT17, PV06, POC05, PSB^{+}08, PU06, PBW19, RRS19, SHWP09, SF09, SPSH14, SLM^{+}17b, SKSJ20, SJ22b, SOS04, SF07, SS10b, SSJ^{+}20, SCD^{+}21, SRGB14, Sta03, TSNI10, TDG18, TZL^{+}02, TO02, VBCG10, VdFG99, VHWP12, WMT05, WSM11, WC21b, War89, WDB^{+}08, WG09, WGL^{+}18, WZYR19, YHJ^{+}14, YZ04, YT13, YBSC21, ZMSS18, ZZV^{+}03, ZMT06, ZS00, ZHX^{+}07, vW09]. SurfaceVoronoi [XWX^{+}22]. surfacing [PLS^{+}15]. surfel [AD03]. surfel-bounded [AD03]. surgery [MCS15, TR98]. surgical [CAR^{+}09]. surroundings [VAV^{+}07]. Survey [DKHS14, Gre86, PCS^{+}23b, GB08a]. suspended [FOA03]. SV [RGB16]. SV-BRDF [RGB16]. SVBRDF [AWL13, AWL15, BJTK18, DAD^{+}18, DWT^{+}10, GLD^{+}19, GSH^{+}20, GLT^{+}21, GLT^{+}23, HHD^{+}22, HJM^{+}22, NLGK18, Zha18, ZCD^{+}16, ZK22]. SVG [YWH13]. SWAGAN [GHBCO21]. swapping [BKD^{+}08]. Sweep [CZS^{+}13]. Sweeping [vW84]. Swept [SAJ21]. swimmers [MDZ^{+}21]. Swimming [SLST14, SHU^{+}16, TGTL11]. swings [CB05]. SwingWrapper [AFSR03]. Switchable [SMH^{+}11]. Switching [GLX^{+}22]. Symbolic [EC93, BCT15, Gue07, Jia21]. Symmetric [ASGS23, CC19, JTC09, vW09, GWAB19, LF08, PLPZ12, Rus19, SR97, YTL18]. symmetries [MSHS06, SHZ^{+}20, THW^{+}14]. Symmetrization [MGP07]. Symmetry [BSEH18, KLF12, LCDF10, RS14b, BWS10, CMZP14, LSS^{+}17, MGP06, PZ07, PSG^{+}06, RVLL08, WWF^{+}10, XZT^{+}09, XZJ^{+}12, ZXJ^{+}13]. Symmetry-guided [KLF12]. symmetry-summarization [WWF^{+}10].

SymmetryNet [SHZ^{+}20]. sync [SSKS17]. synchonized [KGS^{+}18]. Synchronization [Hil86, ELFS16, WSZ^{+}14]. Synchronized [KHKL09, SJA^{+}20]. synchronizing [HLW^{+}19, LJ14]. synchronous [HLZ10, HZG08]. Synergizing [DXG^{+}23]. synopsis [ACCO05]. Syntactic [SG91]. Synthesis [ASHW23, ANBH23, AGL^{+}22, AFP^{+}95, BSL12, CZX^{+}16, CBYvdP08, DBP^{+}15, DWS^{+}23, HM92, JWDL19, KLR^{+}22, LW15, LCL^{+}23, LWL23a, LLX^{+}01, LP02, MSL^{+}23, PNTK23, PQF^{+}23, RSV^{+}23, RO85, RO87, SCO17b, SOG^{+}22, SWS^{+}22, TZL^{+}02, TCS^{+}23, WB08, WSML23, XAW^{+}23, YL12, YBY^{+}13, ZZV^{+}03, ZYM^{+}20, ZFT^{+}21, AAL16, ALY^{+}21, AVB08, AJM12, AFO03, BSHK04, BDT^{+}08, BNB13, CDSHD13, CTL^{+}21, CWL12, CT17, CLG^{+}16, CWTW17, DSB^{+}12, DLL^{+}15, DLKS18, EVC^{+}15, FP03, FH04a, FJS^{+}17, FPBCO20, FRS^{+}12, FSL^{+}15, FRS19, FAW19, FCW^{+}17, GGY18, GPD^{+}18, GMP^{+}06, HET^{+}14, HRRG08, HWRH13, HAB20, HSK16, JYL09, JLWM22, JBX^{+}20, JHS12, KWR16, KCKK12, KGS^{+}18, hKPS03, KLF12, KFCO^{+}07, KP06, KSE^{+}03, KEBK05, LES09, LH05, LH06a, LHL10, LSR18, LDF14, LTK09, LWS02, LMM^{+}22, LAZ^{+}22, LHR^{+}21, LSA^{+}16, LXJ^{+}22, MJC^{+}08, MWGZ09, MPF^{+}18, MM08, MSOC^{+}19, MC12, MYH^{+}10]. synthesis [NSCL08, ÖG12, PHL^{+}09, PCSS06, PZ17, PB02, RYL13, RZW^{+}21, RCOL09, SHM^{+}18, SCO17a, TZN19, TOS^{+}03, WZT^{+}08b, WYZG09, WHRO10, WSCR18, WQLJ18, WHZ^{+}08, WLHR11, WLHR12, WY04, XKF^{+}18, XYH^{+}21, XUC^{+}14, XBS^{+}19, YYTC12, ZG04, ZYSK21, ZJMB12, ZHW^{+}06, ZJL14, ZZB^{+}18, ZTF^{+}18, ZFWW18]. Synthesizing [LK20, LHLY21, NSB13, RHDG10, SHP04, SSKS17, YKH04, YYW^{+}12a, CYT^{+}18, NRH17, SZZK21, SWL^{+}22, WL21].

Synthetic [LCV^{+}04, MHS^{+}19b, PTSG09, PC82, RKB^{+}23, WGJ^{+}18, YNK^{+}22, ZMN^{+}19, BDI^{+}02, CNR08, IZE^{+}21, KHFH11, OPOD10]. Synthetic-to-Real [YNK^{+}22]. synthetic-vision [OPOD10]. System [AJS20, CM83, EHSN20, GARP^{+}23, GF82, LZCX19, SG86, Bly06, BTFN^{+}08, CSTP16, DHOO05, FNvD82, GPCP13, HGY17, HFTF15, HFF16, HGG^{+}11, HWR14, HMT^{+}15, JLF^{+}09, KLHG09, LZ04, LGA^{+}21, MGAK03, MP04, MIWI16, MI07, NQC^{+}21, NJS^{+}11, OEE^{+}18, RKKS^{+}07, RXL21, SPJT10, SSY^{+}04, TL04, TKTS11, WZK^{+}17, WS99, YCL^{+}17, ZPKG02]. Systematic [CZB23, GJZ21]. Systematically [BMM^{+}21]. Systems [FH97, GJB^{+}20, JNK^{+}23, LN84, PAK^{+}19, Ree83, WW82, ZIH^{+}11, ACXG09, FLP14, GHZ^{+}20, HFF18, HDA17, HPC21, KSJP08, LLB24, LTT^{+}20, LBOK13, SSB^{+}15, SHS^{+}04, SHHW16, SAZK06, TZCT20].

T [CZ17, GBK05, HWB23, KPP17, KBZ15, SZBN03, SCF^{+}04]. T&I [NPP^{+}11]. T-junctions [KPP17]. T-Mesh [HWB23, KBZ15]. T-NURCCs [SZBN03]. T-Spline [GBK05, SCF^{+}04]. T-splines [CZ17, SZBN03]. Tables [NMLH14, NMLH11]. tabletop [Ano03]. Tactile [LDS^{+}16, TGZ18, BP12, SPGI13, TWZ20]. tags [MWH^{+}09, RBvB^{+}04]. Tailored [DWX^{+}21, POAR12]. Takes [SCCB22]. taking [CLC96]. talk [SQRH^{+}16]. Talking [YFFA21, FTZ^{+}19, LCC21, ZHS^{+}20]. talking-head [FTZ^{+}19, LCC21, ZHS^{+}20]. Talking-Heads [YFFA21]. tall [CM11]. Tangent [BS88, CPW21, CÖS19, PP93, FSDH07, VB06]. Tangent-space [CÖS19]. tangents [HLHZ08]. Tangible [JPG^{+}14, Ano03, GMP^{+}16]. Tangle [NPP22, SP16]. Tanks [KPZK17]. TAP [HXC^{+}20]. TAP-Net [HXC^{+}20]. tapestries [BGSF10]. Target [FL04, GRS^{+}17a, GRS^{+}17b]. Target-driven [FL04]. Task [AvdP16, Cas91, XSZK23, CBvdP09, LLM21, RYPZ23, SKB^{+}14]. Task-Analytic [Cas91]. Task-based [AvdP16, CBvdP09, SKB^{+}14]. Tasks [GXSD23, BSL12, GSCO12, MTA^{+}20, YKH04]. Tattoos [PCB23]. Tau [Las90]. Tau-Splines [Las90]. Taylor [ZRLK07]. TCB [ZCX^{+}22]. TCB-spline-based [ZCX^{+}22]. tearing [LLKC21, PNdJO14]. Technique [EM90, Ree83, Res87, JM12, JB02, KSHG18]. Techniques [And83, HL14, Jan91, Kaj83, Ols88, RO85, RO87, SWZ96, UBW99, CB04, IGLF06, JDR04, JASR99]. technology [BP12]. teeth [VPB^{+}18, WBG^{+}16, YSW^{+}20]. tele [HYG^{+}13]. tele-registration [HYG^{+}13]. teleconferencing [JLF^{+}09]. Telepointer [RO94]. Telepointers [RO94]. teleport [LHLY21]. telepresence [GWN^{+}03, LGA^{+}21]. telescoping [YCC17]. Templates [JMZ^{+}22, JZvdP^{+}08, KLM^{+}13, PYW14, ZHG^{+}16]. temples [KPZK17]. tempoGAN [XFCT18]. Temporal [AECO15, JK23, LYC^{+}22, LAC^{+}11, MKD^{+}16, OHX^{+}14, TD23, WKMH^{+}23, WGP^{+}10, BH21, BGSF10, BBK^{+}15, BTS^{+}15, CSSL21, DLX^{+}21, GBAM11, KKW21, LWA^{+}12, LBJK09, MAC22, VBK05, WFS^{+}09, ZRLK07, ZM13]. Temporally [ASC^{+}14, HKAK16, LLV^{+}12, MNV^{+}21, XFCT18]. tendinous [SSB^{+}15]. Tennis [ZSAF21, ZYM^{+}23]. tensegrity [PTV^{+}17]. tensile [VMTF09]. Tension [BB83, DLG90, JWL^{+}24, MM22, XRW^{+}22, AAT13, CMSA20, CKMR^{+}21, GMB17, RLZ^{+}21, SZB18, TWGT10, WJL^{+}20, ZQC^{+}14]. tension-actuated [GMB17]. Tension-Compression [MM22]. tensioned [Coh87]. Tensor [DLW^{+}22, HLW^{+}19, PRK^{+}17, SG17, Tsa15, WLHR12, TS06, TS12, WWS^{+}05, XZY^{+}17]. TensorTextures [VT04]. terabyte [FSP^{+}22]. terabyte-scale [FSP^{+}22]. terahertz [WW13]. Terrain [GGG^{+}13, LYvdPG12, PGP^{+}19, PBvdP16, SPF^{+}23, cWP10, BST09, CGG^{+}17, GDG^{+}17, LH04, PBvdP15, ZXKL^{+}20]. Terrain-adaptive [PBvdP16, cWP10]. terrain-optimized [ZXKL^{+}20]. Terrains [CJP^{+}23]. tesselation [VdFG99]. Tessellated [NKS^{+}23]. Tessellation [XLC^{+}23, FFB^{+}09, GBK05, HMAM09, LWL^{+}09, LSNC09, NL13, ZS00, BA08, LL10]. tessellations [BLdG^{+}16, LXY^{+}16, ZMSS18]. Testbed [WW82]. Tester [FHXW22]. Testing [LBW^{+}23]. Tetrahedra [FAER21, PVR18]. Tetrahedral [HZG^{+}18, KC23, SHG^{+}22, ACSYD05, ATW13, JZH^{+}21, KTY09, LS07, PRP^{+}15]. Tetrahedrization [DPVA23]. tetrahedron [TWAD09]. tetrapuzzles [CGG^{+}04]. Text [ARMCO23, BAC^{+}23, CAV^{+}23, CWL22, FTZ^{+}19, GAA^{+}23, HAB16, LHH^{+}23, MSL^{+}23, MSL^{+}24, MPE^{+}23, VKM^{+}23, WSML23, XZZ18, YFFA21, ZQL^{+}23, HZP^{+}22, JYQ^{+}22, JMD^{+}17, RMBB^{+}13, SFLM04, YCL^{+}20]. Text-Based [YFFA21, FTZ^{+}19, JMD^{+}17]. Text-Driven [CWL22, MPE^{+}23, VKM^{+}23, BAC^{+}23, HZP^{+}22, JYQ^{+}22]. Text-Guided [LHH^{+}23, MSL^{+}23, WSML23, MSL^{+}24]. Text-to-Image [ARMCO23, CAV^{+}23, GAA^{+}23]. Text2Human [JYQ^{+}22]. Text2Light [CWL22]. textiles [NQC^{+}21]. Textual [PABE^{+}21]. Texture [CS00, DYT05, KPWG24, KEBK05, LLX^{+}01, LPC^{+}11, LHVT17a, MZD05, MHC^{+}16, SCO17b, SS00, SWWW15, TBTS08, TB87, WK95, XZP^{+}23, ZJNZ23, AAL16, BKCO16, BKR17, BNTS07, BD02b, CTW^{+}04, CLKL14, CSHD03, DvGNK99, ESZ^{+}17, FH04a, FCGH08, HP03, HRRG08, KBD07, KLF12, KFCO^{+}07, KSG03, LH05, LH06a, LPRM02, LWS02, LLH04, LDHM16, LSA^{+}16, LHVT17b, LFB^{+}13, MWGZ09, MS13, MCHAM06, Nah20, PKCH18, RAI06, SCO17a, SdS02, SXD^{+}12, TZL^{+}02, TOS^{+}03, TT09, WSH^{+}16, WHZ^{+}08, WY04, XYXJ12, ZG04, ZMT05, ZHW^{+}06, ZZB^{+}18]. Texture-Based [SS00]. Texture-lobes [LPC^{+}11]. Textured [KKN^{+}22, NDD^{+}23, BGB^{+}05, GWY^{+}21, PKC^{+}16, WM03]. TextureMe [KKN^{+}22]. TextureMontage [ZWT^{+}05]. Textures [VSW^{+}23, AZP^{+}05, AS02, BD02a, CGZ^{+}05, gDGPR02, DYN03, FAW19, GP08, GP09, HDMR21, JDR04, JP02, KMB^{+}09, KPM16, KSE^{+}03, LHL10, LGG^{+}07, MWT11, MWLT13, MZD05, MPH^{+}20, NSX^{+}18, ONOI04, PZM^{+}15, PZ08, RCOL09, SXD^{+}12, TOII08, TZN19, TWY^{+}20, TWZ22, WZYG10, ZZV^{+}03, ZHHZ20]. Textureshop [FH04a]. texturing [CH02, GSV^{+}14, LIY^{+}22, PB02, VSLD13, XCOJ^{+}09]. Thallo [MHZ^{+}21a]. theatre [WL16]. their [Fat09a]. theme [WYW^{+}10]. theories [LJGH11]. Theory [ABGL21, APH^{+}14, BBG24, CA00, HZE^{+}19, JSKJ12, BB17, DPF03, FCJ07, GJZ21, JNSJ11, LDF14, MSRB07, RAMN12]. Theran [BTFN^{+}08]. There [PVR18, ISSI16]. thermal [HZW12]. thermoforming [SPG^{+}16]. thickness [ISN^{+}20, YSC^{+}16]. Thin [CSL^{+}23, CCK^{+}21, HWZ^{+}14, LSNP13, WDK^{+}21, ASL^{+}17, ABO16, BMWG07, BDW13, CAJ09, CSvRV18, CNZ^{+}22, CQD^{+}18, Dav20, DWK^{+}22, FSH11a, GRBN09, GSLF05, GHF^{+}18, HLHR09, LCC^{+}18, PNdJO14, RK13, VRBC18, VLD^{+}13, WT08, WTGT10]. Thin-film [WDK^{+}21]. thin-plate [FSH11a]. thin-shell [CQD^{+}18]. things [Iza18]. think [BDM^{+}20]. thinning [NSS^{+}19]. Third [RSM^{+}23]. thousands [FSP^{+}22]. threads [BAV^{+}10]. Three [CKH18, CCW93, CGM91, COSL98, Day90, EM94, Gre86, JSMH12, RSSH24, SG17, WF96, BBO91, Boi84, IGLF06, SLWF14, UB18]. Three- [CCW93]. Three-Dimensional [CKH18, Day90, EM94, COSL98, JSMH12, RSSH24, BBO91, Boi84, UB18]. three-level [SLWF14]. threshold [WWH04, ZF03]. tight [DML17]. TightCap [CPY^{+}22]. Tightness [CPY^{+}22]. tilable [FLHCO10]. tile [CML^{+}17, WPC^{+}14]. tile-based [WPC^{+}14]. tiled [MS05, YBY^{+}13]. TileGAN [FAW19]. tiles [KCODL06, LD06, CSHD03]. tiling [vW09]. Tilings [NI24]. tilted [XCW^{+}20]. Time [ARMCO23, And83, AIH^{+}08, BWC^{+}23, BYG96, BJ05, BKCO16, CHTK24, CWTW17, DNZ^{+}17b, DLK18, DWS^{+}23, DLP^{+}23, GTR^{+}06, GXY^{+}17a, GZS^{+}22, GNHM15, GVNB18, HXZ^{+}19, HMI23, KZSR16, KKLD23, KKN^{+}22, KJGP23, KIM^{+}19, LBK17a, LZY^{+}21, MBGS15, MOR^{+}18, Mey91, MU22, MNV^{+}21, TZS^{+}18, TSLP14, TCS^{+}23, VTSSH15, WLF^{+}20, WS85, XLC^{+}23, ZXTZ15, ZZT^{+}21, ABW^{+}17, ASA^{+}09, ADM^{+}08, BWDL21, BHR13, BP08, BZ11, BMSR20, BAOR06, BM07, BK04, BBG21, CHWH17, CWLZ13, CHZ14, CBZB15, CWW^{+}16, CKH18, CCWL18, CAD^{+}21, CH02, CPD07, CBI13, CM11, CT05, CHP07, DNZ^{+}17a, DRvdP15, DLL^{+}18, DYN03, DHOO05, DKD^{+}16, DDF^{+}17, DCB^{+}22, EMU15, FKY08, FYK10, GO12, GCB^{+}17, GSKJ03, GRGC15, GXY^{+}17b, HLX^{+}21, HV04, HAK^{+}22, HED05, HFF18, HRE^{+}08, HHHW15, HDHN16, HSW^{+}17, HKA^{+}18, Hub96, HESL11, JBPS11, JSRV22, JP02, JTL^{+}12, JKT^{+}15, KWB^{+}13, KNS^{+}09, KUJH21]. time [KCODL06, KRF^{+}18, KAMJ05, LEN09, LH16, LES10, LZC11, LMLL21, LTK09, LLKP11, LHdG^{+}14, LGL^{+}19, LLJ22, LLX^{+}01, LCH^{+}21, LFTC13, LHLK10, LXC^{+}15, LBK17b, LZH^{+}20, LCX^{+}21, LB06, LCC21, MMCK14, MHM^{+}17, MBPY^{+}18, MP04, MP08, MSS^{+}17, MDB^{+}19, MCK13, MRNK21, NSX^{+}18, NMD^{+}17, NOP^{+}18, NZV^{+}11, NZIS13, PZ08, PO08, PVG19, POC05, PYA^{+}24, RSM^{+}10a, RSV^{+}23, RWS^{+}06, RTK^{+}15, RJ07, RHHL02, SAL^{+}08, SZT^{+}08, SGXT20, SHHW16, SCT^{+}15, SL17, SSII18b, SKS02, SXH^{+}21, SRNN05, SMPR07, TDSG15, TDL^{+}18, TWH^{+}22, TZN^{+}15, TZT^{+}18, TPT16, TLP06, TS12, VBG^{+}13, VRBC18, VSJ21, WKF^{+}21, WAO^{+}09, WWD^{+}05, WTL^{+}06a, WPP07, WP09b, WJBK15, WYM^{+}16, WSJP17, WJ19, WMB^{+}20, WXLY17, WGT^{+}05, WOG06, WZN^{+}14, WCRZ21, XUC^{+}14, XZY^{+}17, YZX21, YZH^{+}23, ZIT^{+}18, ZZI^{+}17, ZBYX19, ZHHZ20, ZZZ^{+}23, ZHWG08, ZRL^{+}08, ZNI^{+}14, dASTH10]. time-critical [Hub96, LMLL21]. time-domain [WJ19]. time-gated [PVG19, WCRZ21]. time-image [BMSR20]. time-independent [BBG21]. Time-lapse [MBGS15, BM07, HAK^{+}22, LEN09, SMPR07, TDSG15]. time-multiplexed [WGT^{+}05]. Time-of-Flight [BWC^{+}23, GNHM15, GVNB18, HMI23, KZSR16, KJGP23, ABW^{+}17, CHWH17, HHHW15, MHM^{+}17, NZV^{+}11, SHHW16]. Time-resolved [AIH^{+}08]. Time-travel [LZY^{+}21]. time-variant [WTL^{+}06a]. Time-varying [BKCO16, GTR^{+}06, BHR13, DRvdP15, HED05, XZY^{+}17]. Time/Space [BYG96]. times [SPDF13]. Timestep [FSKP23]. tissue [BBO^{+}09, DFW20, KPMP^{+}17]. tissues [PRWH^{+}18]. TM [GWY^{+}21]. TM-NET [GWY^{+}21]. TOG [Ols88]. together [GSKJ03, RTB17]. toil [DBWG15]. token [Zit13]. tolerance [MCSA15, YRPF09]. tolerant [SLWF14]. tomographic [WLHR11]. Tomography [SSW^{+}23, GKHH12, IYYI14, RWL^{+}22, ZIT^{+}18, ZIT^{+}19]. ton [CXW^{+}05]. Tonal [FL11, LFUS06]. Tone [SW18, WC21a, ASC^{+}14, BPD06, EMU15, EKM17, FFLS08, KO11, LCTS05, MDK08, MAF^{+}09, RSSF02, RTS^{+}07, WYX11, YZWH12, ZF03]. tool [BBR^{+}21, BDM09, FH04a, JRT^{+}15, MZB^{+}17, WAC07, WZL^{+}20, XFAT12]. toolkit [FH04b, MGDB05]. Tools [BLA12, BD86, HA92, LFL^{+}23, SB93, SLF22, PLKD18, RMD12]. toon [ZLWH16]. tooning [WXSC04]. Toonsynth [DLKS18]. TopoCut [FDBH22]. Topological [LDW97, VW94, vOV96, GMP09, LDK^{+}18, NGH04, TR98, Xia21, ZCLJ20, VW95]. Topologically [SH23, PSH^{+}21, PKZ04]. Topologically-Stable [SH23]. Topology [ALX^{+}14, ABA02, CZXL23, DFL^{+}15, HZCJ17, LDS^{+}22, MHCT23, MB12, NHS^{+}13, PSF09, Sar00, WKMH^{+}23, YJY23, ZJL14, ZSCM17b, ZHCJ15, AXZ^{+}15, ABO16, BHK14, BW13, BHLW12, BBB10b, DRvdP15, JZH07, LHM09, LHZ^{+}18, MBF04, Mus13, NKJF09, QJ21, SLS^{+}07, Sta03, WTGT10, WHDS04, YHZ^{+}14, ZPBK17, ZSCM17a]. Topology- [PSF09]. Topology-adaptive [MB12]. topology-aware [SLS^{+}07]. Topology-based [DFL^{+}15]. Topology-constrained [ZJL14, ZHCJ15]. Topology-controlled [HZCJ17]. Topology-driven [NHS^{+}13]. topology-preserving [LHM09]. Topology-reducing [ABA02]. Topology-varying [ALX^{+}14, AXZ^{+}15]. Toric [GPSZ11, LC15, MGA^{+}17]. ToRoS [MHCT23]. torque [JWDL19]. Torso [LJL23]. Total [BBG24, MGDA^{+}15, PEL^{+}21, XYXJ12]. Total-variation [BBG24]. touch [PRWH^{+}18, RP09]. tourism [SSS06]. tourist [GASP08]. tower [DFL^{+}15]. toy [ZXS^{+}12]. toys [MS04, MI07, SWT^{+}17]. trace [MKZ^{+}21]. traced [EDR11, HR13, PFHA10]. Tracer [GIF^{+}18]. Tracing [BK85, BK87, CFS^{+}18, DLTW90, FHL^{+}18, GHCC88, GRS^{+}17a, HYS23, JRSS21, Kaj83, KIM^{+}19, Lev90, NID20, NKK^{+}14, PP94, RS14b, RLU95, SLM^{+}17a, TB87, VKJ^{+}17, WQF^{+}21, WHG84, vW84, BDT99, BSS^{+}13, CRS^{+}16, CXW^{+}05, CTE05, DHW^{+}11, FSP^{+}22, GRS^{+}17b, HJW^{+}08, HJ11a, HQL^{+}10, HZ11, KMA^{+}15, LAA^{+}05, LADL18, LWL^{+}20, MKD^{+}16, Mor11, MRNK21, MHC^{+}16, NPP^{+}11, PBD^{+}10, PCS^{+}20, PBMH02, RSH05b, SHHD17, SLM^{+}17b, SLWF14, SLW22, SWF^{+}21, TOG22, WIK^{+}06, WBS07, WWB^{+}14, WSS05]. TrackCam [LWCT14]. tracked [CB04, JBM^{+}17, PSK^{+}16]. trackers [NDMKJ22]. Tracking [BHLW12, WKHA18, AHSS04, BHW13, CSHH21, CHZ14, CCWL18, CMSA20, CMMK15, DBG14, FBH21, HLW^{+}18, HK10a, HMT^{+}15, JTST10, KRF^{+}18, KHLN17, LWCT14, MB12, NSJ14, SWW^{+}20, TBC^{+}16, TTT^{+}17, TAH^{+}04, TPT16, TTR^{+}17, VGB^{+}14, WP09b, WMB^{+}20, WXLY17, WZC^{+}22, WSS18, ZLWH16, ZBGB19]. TRACKS [BMWG07]. trade [LDS02, SWC^{+}18]. trade-offs [LDS02, SWC^{+}18]. Tradeoffs [BYG96]. traditional [CWZ^{+}21a]. Traffic [SQSL22, LWL17, SWL11, WSL13]. train [WPKL17]. Trainable [EGP02]. Training [HL14, ZK22, MCS15, PCPW20]. Trajectories [PFP^{+}22, TFD^{+}18, RH16]. Trajectory [GM84, HNH19, LH18]. Tran [Ros20]. Tran-similar [Ros20]. Transactions [Bea88]. transcripts [SBLD15]. Transfer [AONA22, BBC^{+}23, HLC^{+}19, IWHH20, JPL22, LFZ15, RCL21, SHS^{+}18, YJLL22, ZTD^{+}23, ZHM^{+}23, AWL^{+}20, AHLG^{+}13, ABW14, ACSM12, BVGP09, BRM^{+}18, BSBC12, CNR08, DYT05, FZL^{+}15, GYQ^{+}18, HPB06, HHP^{+}21, HSC^{+}22, JAM^{+}10, JBP06, JFA^{+}15, JAG18, JLWM22, KAGS20, KAMJ05, LEN09, LYY^{+}17, LKWS16, MCW^{+}21, ODAO15, PCS^{+}20, PTMD07, SDKN18, SED16, SPB^{+}14, SHS^{+}17, SKS02, SLSS03, SHHS03, SLS05, SSBD03, SLL^{+}21b, SP04, SJA^{+}20, TZN^{+}15, TLJP18, TS06, VBPP05, WSH^{+}16, WJ19, WTBS07b, XWCH15, XCLT14, YWS^{+}11, YM16, ZHRB13, ZRB14, LAM^{+}11]. Transferring [HLR^{+}17, WAM02]. Transfiguring [KS16]. Transflower [VPHB^{+}21]. Transform [GSC^{+}15, LLF^{+}20, LWS^{+}15, PP94, Pag98, WWWG22, AKZ^{+}17, BHY15, FMR20, GO11, HJ11b, PSG^{+}06, YHCOZ18]. Transformation [NN90, YYL^{+}19, ZMW^{+}23, APCO21, DYT05, WKR99, WGT^{+}05]. Transformations [BSB16, NN90, Pat85, Pat87, Tur82, Ale02, BSB17, CPS11, JBK^{+}12, LSS^{+}17, NRC21, Spr82, VMW15]. transformed [HDHN16]. Transformer [HZL22, LLB24, PQF^{+}23, FYK10]. Transformer-based [LLB24]. Transformers [QZZ22, Wan23, WSML23, LSH^{+}22]. Transforming [XZM^{+}18]. transforms [LMAH^{+}18]. Transfusive [YJHS12]. Transient [IH20, LRT^{+}14, BL15, HHGH13, JMM^{+}14, OHX^{+}14, PKHK15, YKC^{+}21]. Transition [SYSP14, TLZ^{+}24, TWH^{+}22]. transitions [BLA12, DDD^{+}14, WB08]. translating [CLD^{+}13]. Translation [LXZ^{+}23, CLY18, FTP03, HPP05, MYC^{+}22, WSS^{+}19, YZX21]. Translational [LW15]. translucency [BATU18]. Translucent [BAU15, IRN^{+}22, RT90, DI11, DJ05, GXZ^{+}13, GLL^{+}04, HV04, JB02, PRJ^{+}13, WTL05, WZT^{+}08a]. transmission [AAR05, KV05, MP04]. transmittance [KDPN21, VJK21]. transparent [LWL^{+}20, SOA11, WZQ^{+}18, YTBK11]. Transport [BRSMD22, BJNJ18, DKHS14, LR15, RLLG^{+}20, SGSS22, SHS^{+}18, XLY^{+}22a, ZFT^{+}21, BH21, BJ17, BvdPPH11, BPC16, BC19, DHS^{+}05, GKDS12, GLDZ15, HPJ12, HKD14, Hac18, HXC^{+}20, IZT^{+}07, JM12, KHD14, KGH^{+}14, LCCS18, Leh07, LZT^{+}08, LKL^{+}13, Lip18, MSRB07, MCSK^{+}17, MRK^{+}14, MGJ19, NG18, NSCL08, OK10, ORK12, OHX^{+}14, OHHD18, Pan17, PML^{+}09, QSH^{+}15, RHJD18, SNM^{+}13, SHS^{+}17, SOHK16, SV19, SY21a, VKS^{+}14, VK16, WDT^{+}09, WHY20, ZSGJ21, dGBOD12, LLR^{+}15]. transport-and-pack [HXC^{+}20]. transport-based [SV19]. transportation [SdGP^{+}15]. TransPose [YZX21]. Trap [PFP^{+}22]. travel [LZY^{+}21]. traversal [BAM14, NPP^{+}11, PBvdP15, SNCH08, WIK^{+}06]. treatment [BFA02, HVTG08, KK87]. Tree [LLB24, Shn92, WLX^{+}18, AMA^{+}19, BO04, CNX^{+}08, LGB^{+}21, LYO^{+}10, LPC^{+}11, MGT^{+}03, NFD07, PHL^{+}09, PNDN12, PSK^{+}12, PNH^{+}14, PJH^{+}17, PHBC21, TZW^{+}07, TFX^{+}08, WLLS22, XLJ^{+}09, ZHWG08, JP04]. Tree-Maps [Shn92]. tree-modeling [NFD07]. TreeJuxtaposer [MGT^{+}03]. treemaps [BSW02]. TreePartNet [LGB^{+}21]. Trees [HTS^{+}22, AGDL09, DVS03, DIP^{+}18, LBAD^{+}06, LDS^{+}11, LKK^{+}21, LMPB^{+}13, PSK^{+}12, PNH^{+}14, RMD04, XGC07]. triage [CYW^{+}16]. Triangle [LS00, SS10b, ULP^{+}15, AFSR03, CSN^{+}12, GLLR11, LZKW10, PPW18, QHY^{+}16, SNB07, SW05, SC20, SOA11, SS21, SSP08, SGC18, SP04, WZHB09]. triangle-oriented [QHY^{+}16]. triangle-quad [PPW18]. triangle/quad [SW05]. Triangular [Sar00, FKY^{+}10, JSW05, Lip12, MC21, PU06, YHB05, ZFO^{+}22]. Triangulated [RS14b, HR05]. Triangulating [FM84, WS85]. Triangulation [CI84, EPO91, KLN91, WWX^{+}22, dFP95, FAB^{+}18, HSG^{+}19, LPS^{+}13, RAM^{+}21]. Triangulations [Kal14, LFXH17, Pet01, SG01, dGMMD14, Ale19, Ale20, ILSS06, MMdGD11, SSC19a, Tak22]. trichromatic [RZK11]. trigonometric [PKHK15]. trilinear [Csé19]. Trim [VKW^{+}23]. trimmed [LCBK19, SFL^{+}08]. trimming [GBK05, SF09]. trimodal [YCL^{+}20]. Trip [Pra89]. Triple [NRH04, SR09]. triple-product [SR09]. TriWild [HSG^{+}19]. trouble [DBWG15]. True [RC22]. True2Form [XCS^{+}14]. truly [MMG06]. truss [SHOW02]. try [LVKS21]. try-on [LVKS21]. TryOnGAN [LVKS21]. tuner [CLD^{+}13]. Tuning [GAA^{+}23, RMBCO23, VKM^{+}23]. Tunnel [LBW^{+}23, DLSCS08, She13]. turbulence [BWDL21, CQD^{+}18, KTJG08, KTT13, MBT^{+}15, NSCL08, PTSG09, PTC^{+}10, SDKN18]. Turbulent [LWP^{+}23, LL23, LCD^{+}20a, LLDL21]. Turning [BLCD02, SSJ^{+}11, WX91]. tutorials [GAL^{+}09]. Tutte [AL15, AL16, AKL17]. TV [FMR20, MP04]. twice [YRPF09]. twilight [HMS05]. TwinTex [XZP^{+}23]. Twistable [JS11]. Twister [LKG^{+}03b]. twisty [SZ15]. Two [AWL15, BPD06, Gla90, GWBN24, JTMW20, Las90, LD13, LD23, QZZ22, RMSG^{+}08, SJ94, SG11, TFD^{+}18, THG99, WCL^{+}23, ZLW^{+}18, ZSCM17a, ZSCM17b, AMB^{+}21, BB12, FQL^{+}20, Gal99, GLT^{+}21, HP17, HTYW22, HFG^{+}18, IGLF06, LWS02, LCD^{+}20a, LK20, LMLD22, LKG^{+}03b, MDB^{+}19, NAI^{+}18, NGL10, NØ13, RRC^{+}16, TB20, WAH^{+}10, WGH21, XNY^{+}16]. two-continua [NØ13]. Two-Dimensional [Gla90]. two-handed [LKG^{+}03b]. Two-layer [LD13]. two-level [LWS02]. Two-Phase [LD23, BB12, LMLD22]. two-piece [AMB^{+}21, NAI^{+}18]. two-player [WAH^{+}10, WGH21]. Two-Point [TFD^{+}18]. Two-Scale [ZSCM17b, BPD06, SG11, ZSCM17a, HP17]. Two-shot [AWL15, XNY^{+}16]. Two-Stage [QZZ22, ZLW^{+}18, HTYW22]. two-stream [GLT^{+}21]. Two-Way [GWBN24, RMSG^{+}08, WCL^{+}23, FQL^{+}20, HFG^{+}18, LCD^{+}20a, NGL10, TB20]. Type [LDW97]. typefaces [Sha03]. Typography [IVH^{+}23].

UAVs [XDF^{+}19]. ubiquitous [LGK^{+}16]. UI [KS21]. UIMS [Hil86, SG91]. Ultra [GLT^{+}23, TTT^{+}17, VLD^{+}13]. ultra-fast [TTT^{+}17]. Ultra-High [GLT^{+}23]. ultra-thin [VLD^{+}13]. ultrasound [LSCS14]. Umbrella [RKP^{+}22]. unactuated [YL08]. Unbiased [BLD20, GIGM22, MBGJ22, NDMKJ22, QSH^{+}15, YIC^{+}10, DJBJ19, KDPN21]. Unbounded [RSV^{+}23]. uncertain [WFH10]. uncertainty [UMK17]. unconstrained [YSN^{+}18]. uncontrolled [VWB^{+}12]. Unconventional [MV21]. Understanding [GXZ^{+}13, PKH^{+}17a, PKH^{+}17b, SN17, XADR12, HOM15, LRT^{+}14, LT20, NXS12, SMZ^{+}14, YZL^{+}22]. Underwater [OKRC10, WP12, MDZ^{+}21]. Unfolding [SK16, MS04]. uniaxial [WW08]. UniColor [HZL22]. Unified [GJ22, HZL22, MWM23, MMCK14, MWC^{+}23, MKB^{+}10, MHU19, RXL21, SHU^{+}16, TLZ^{+}24, ZTD^{+}23, ZZC^{+}22, CLC^{+}20, CLL^{+}22, DM13, GD04, LBB17a, LSD^{+}22, MAC22, SXH^{+}21, VdFG99, WMW15, YCL^{+}17]. uniform [AVR^{+}22, CADS09, LFS16, WW11]. uniformity [PBC^{+}22]. uniformly [HRV^{+}18]. Unifying [KGH^{+}14]. unit [DFM13, HAM07, WSS05]. units [LHLK10]. UniTune [VKM^{+}23]. unity [OBA^{+}03]. universal [CLF^{+}18]. Unknown [CHTK24, DCP^{+}14b, XDPT16, XZY^{+}17, ZSD^{+}21]. unlabeled [XWCH15]. Unmixing [AAPS16, AASP17b, AASP17a, AAPS17]. Unmixing-Based [AASP17b, AASP17a]. UnMousePad [RP09]. unordered [SSS^{+}08]. unorganized [HLZ^{+}09]. Unoriented [WZX^{+}23, HWW^{+}22]. Unpaired [AWL^{+}20, CLY18, GYQ^{+}18]. unparameterized [gDGPR02]. unreinforced [PBSH13]. unseen [SMZ^{+}14]. unsharp [LCD06, RSI^{+}08]. Unsigned [HCW^{+}23]. unsmoothed [SHM22]. Unstructured [BBPP10, JDH^{+}22, JGMR23, ZDF^{+}23, GCD^{+}20, HJM^{+}22, NLGK18, PKC^{+}16, TKKT12, YAB^{+}22]. Unsupervised [CRB23, CPW21, HFW^{+}19, LYF^{+}20, SvKK^{+}11, WSH^{+}16, YC21, BME21, FYW^{+}18, HWH^{+}18]. Unsynchronized [MCT15, YLC^{+}20]. Untangling [BWK03, BRB^{+}19]. Unwrap [RAKRF08]. UofA* [SG91]. Updated [HLSO12, HGMRT20]. updates [HSH20, LLKC21]. upper [LST09]. Upright [FCODS08]. ups [LJGH11]. Upsampling [BLDL21, SZW^{+}23, CAWH16, Fat07, KGBS11, KCLU07, SLJT08, WGP^{+}10]. upscaling [FF11]. Urban [GDAB^{+}17a, NPA^{+}22, VLA15, YWVW13, AVB08, CMZP14, GDAB^{+}17b, KFWM17, KCYW13, LCX^{+}21, NSZ^{+}10, NGDA^{+}16, SHFH11, SMGH18, VABW09, VGDA^{+}12, ZYX^{+}21, ZSW^{+}10, ZXH^{+}20]. Use [HC86, Tur82, BSW02, YYL22]. User [BD86, BPD09, BBPD12, BPB13, Fol86a, Fol86b, Fol86c, HC86, Hud94, Jac86, Pel05, RvE93, RO94, SG91, GB08a, HRE^{+}08, JKH^{+}22, KKB^{+}11, LZC11, Ols84, PCLC16, PTG02, SH08, WPC^{+}14, YCYW20, ZZI^{+}17]. User-assisted [BPD09, BPB13]. user-centered [GB08a]. User-configurable [Pel05]. user-created [HRE^{+}08]. User-guided [BBPD12, ZZI^{+}17]. User-Interface [RvE93]. user-specified [WPC^{+}14]. users [KP09, KP10]. Using [BIW93, BBB^{+}93, BJNJ18, BN90, CBYJ23, CM21, CFP^{+}21, CZB23, CGM91, CSS96, CJM21, DNZ^{+}17b, DGH16, DLW^{+}22, Duf17a, DKD^{+}17a, EC93, Fat14, GF82, GXY^{+}17a, HCOB10, HGM14, Hud94, HWZ^{+}20, IH20, JCY23, JWI^{+}21, JGN16, KL17a, KLN91, LDD^{+}23, LLK^{+}19, LLN^{+}14, LCK22, LH17a, MHS^{+}19a, MHNT15, Mer23, MU22, NID20, PMHD19, QLH^{+}22, RLY^{+}14, RYPZ23, SMR^{+}22, SDN18, ST16, SG17, SHD^{+}14, SHS^{+}18, SBN15, Spr82, TSLP14, TB87, VMKK00, WMB21, WWWZ23, WK95, War92, WLL23, WLS^{+}23, XZZ18, XLY^{+}22a, XLCB15, XNZ^{+}22, YZW^{+}16, YLC^{+}20, YFFA21, YCP16, ZB94, ZWK14, ZZW^{+}22a, ZWHB22, ZCP^{+}23, AZMW21, Ada21, Aga07, ARNL05, ALK^{+}17, APCO21, AZB09, AYL^{+}12, ABA02, ACSM12, ASL^{+}17, AAM03, BCT15, BKGK17, BAS14, BWSS09, BCN08, BP08, BdSP09, BGAM12, BAM13, BKKL15, BBO91, BHB^{+}11, Bel18, BM05, BBGB16, BBG^{+}13, BBB^{+}14, BL15]. using [BDK^{+}16, BWKS11, BvdPPH11, BPC16, BNTS07, BFK^{+}16, BSEH18, CHWH17, CK14b, CB04, CI97, CH07, CKS^{+}17, CRG^{+}20, CNX^{+}08, CLW^{+}14, CBW^{+}18, CM11, CLSA20, CPWAP08, CLQW08, CWL12, CLS03, CS09, CJN^{+}17, CK11, DNZ^{+}17a, DSB^{+}12, DH96, DLF12, DZS08, DYN03, DIO^{+}12, DZPZ09, Duf17b, DDP99, DKD^{+}17b, EKD^{+}17, EB08, FXBH16, FBH^{+}10, Fat09b, Fat11, FLB17, FKY08, FSH11b, FSP^{+}22, FCJ07, FLSG14, FBH21, GJTP17, GGG^{+}13, GLA^{+}19, GFT^{+}11, GLDZ15, GWP^{+}19, GNS^{+}12, GF12, GKJ^{+}05, GBAM11, GJWW14, GXY^{+}17b, GSH^{+}20, HJ11a, HTC^{+}14, HET^{+}14, HRL15, HE07, HHGH13, HLR^{+}14, HDN^{+}16, HSS98, HAB20, HTS^{+}22, HSTP11, HLHR09, HSHF10, HMLL14, HMLL15, HXC^{+}20, HZZ11, HLBR12, HKAK14, IOOI05, IMF^{+}21, JKSH13, JL11a, JNSJ11, JTL^{+}12, JZW^{+}15, JWDL19, JCRA11, JMA06, JKZS10, JMAK10, JZvdP^{+}08, KL17b, KCW^{+}18, KT03, KGS^{+}18]. using [KSES14, Kim10, KLM^{+}12, KLM24, KLF^{+}19, KSE^{+}03, KLV20, LJS^{+}15, LLDD09, LSC^{+}22, LHKR10, LWH^{+}11, LCXS09, LRR04, LCTS05, LZF10, LDF14, LLW04, LGX^{+}13, LLZM10, LLX^{+}12, LHZ16, LVS^{+}16, LWL17, LDPT17, LTT^{+}20, LRFH13, LWO19, LZCV20, LXW^{+}11, LCK^{+}14, LH17b, LH18, LGC^{+}23, LSCS14, LB05, LH04, LEQ^{+}07, MJC^{+}08, MTP^{+}18, MLR^{+}14, MWBR13, MPN^{+}02, MZD05, MTPS04, MRA^{+}13, MSL^{+}11, MBGJ22, MB12, MS04, MM06, MWM08, MdLH10, MWTK13, MGT^{+}03, MAB^{+}15, MHR^{+}16, NYY04, NSX^{+}18, Nah20, NZV^{+}11, NNC^{+}20, NSCL08, NKGR06, NFD07, NRH03, NL13, NZIS13, OLAH14, PZM13, PBH15, PRJ^{+}13, Par17, PCSS06, PMS12, PTMD07, PL07, PBvdP15, PBvdP16, PBYV17, PPW18, PTSG09, PTC^{+}10, PGZ^{+}19, PEVBC21, QZG^{+}19, RTF^{+}04, RAT06, RNd^{+}07, RGB16, RGF^{+}20, RWS^{+}06, RDL^{+}15, RKB04, RKZ11, RMBB^{+}13, SHM^{+}18, SMH^{+}11, SW85, SNCH08, SMW06, ST14, SvTSH14, SED16, SAN23]. using [SBSS12, SAL^{+}08, SSW^{+}23, SWTC14, SHS^{+}17, SOA11, SHK^{+}14, SHM^{+}14, SGG^{+}06, SLWS07, SRL^{+}15, TMRL14, TK14, TZK^{+}11, TGB13, TZN19, TS06, TYY^{+}19, TT09, UBW99, VABW09, VSJ21, VPB^{+}09b, WIK^{+}06, WBS07, WHSG97, WZT^{+}08a, WHDK12, WYY^{+}14, WLL^{+}14, WSXC16, WZK^{+}17, WMB19, WJL^{+}20, WG09, WZC12, WLHR12, WMP^{+}06, WJV^{+}05, WM03, WGP^{+}10, WGH22, Xia21, XLJ^{+}09, XWW^{+}14, XSZB15, XAW^{+}23, YCR^{+}15, YL10, YL12, YJB^{+}14, YYW^{+}12a, YBY^{+}13, YT13, YCHK15, ZRLK07, ZLY^{+}21, ZJMB11, ZF03, ZHS^{+}05, ZRL^{+}08, ZTF^{+}18, ZAFW21, ZXS^{+}21, ZKU^{+}04, Zit13, ZNI^{+}14]. UV [HDC07, KPWG24, NKS^{+}23, PTH^{+}17, Tar16]. UV-maps [Tar16].

v [LJGH11, Mir98]. V-Clip [Mir98]. v-style [LJGH11]. VAEs [LZCV20, WGH22]. valid [FP03, UIM12, WMC11]. validated [FCGH08, GWM^{+}08]. validation [RLR^{+}21]. validity [SSM15]. valley [OBS04]. Value [MSCG23, SCJ^{+}23, HF06, JSW05, LJH13a, TMB18]. values [KABL14, LFUS06]. variability [KMYG12, OLGM11, ROA^{+}13]. Variable [DPD22, LK20, ZF03]. variable-coefficient [ZF03]. Variance [HZE^{+}19, MCSK^{+}17, PSC^{+}15, SK13]. Variance-minimizing [MCSK^{+}17]. variant [BSD09, WTL^{+}06a, ZZV^{+}03]. variants [LL19]. Variates [CJM21, NRN^{+}23, MRKN20, RJN16]. Variation [MGDA^{+}15, BBG24, LBJK09, MLH^{+}09, XYXJ12]. Variational [ACSYD05, BCWG09, BHSH^{+}22, CSAD04, DSSS23, FSK04, HCJ19, LBB17a, Sar00, SC18b, WGS23, ZZWC12, BBB07, DK09, GWAB19, KS98, LMH^{+}15, MMTD07, SHM^{+}18, WP10, XLLW20, YI17]. Variations [BS90, BSW13, BL15, DMIF15, GBLM16, HOM15, ZHG^{+}16]. varied [HRE^{+}08, SSJ^{+}14]. variety [MLD^{+}08]. varifocal [ALK^{+}17]. various [SHU^{+}16]. Varrier [SMG^{+}05]. Varying [Fol87, MXZ^{+}23, ALX^{+}14, AXZ^{+}15, BJ10a, BHR13, BB17, BKCO16, BATU18, DRvdP15, DWP^{+}10, DTPG12, DCP^{+}14b, GTR^{+}06, HED05, HMP^{+}08, LXR^{+}18, MGS^{+}21, MAF^{+}09, MAG^{+}09, PSH^{+}21, PFB^{+}20, SSJC22, TDMS16, TDG18, WRG^{+}09, XDPT16, XZY^{+}17]. VASCO [ZZL^{+}23]. vast [HQT^{+}21]. VAXstation [Lev84]. VDAC [MAYZ^{+}20]. VDB [Mus13]. VDP [MKRH11]. Vector [AOCBC15, BSEH18, CM83, DRvdP14, DRvdP15, LTDD16, LABS23, SSC19b, SWWW15, WZYG10, WSML23, ZMT06, vFTS06, AVR^{+}22, BKKL15, BBG12, EBJ^{+}06, EPD09, FSH11a, FSDH07, GLdFN14, Gol85b, LLGRK20, LMPB^{+}13, MSSG^{+}21, NH08, OBW^{+}08, TLHD03, TWZ22, WWT^{+}06, WYZG11, WL21, YLL^{+}22, ZJL14]. vectorial [BBG12]. Vectorization [BS19, ZMW^{+}23, ZWRY21, ZCX^{+}22, DKT^{+}23, FLB16, FLB17, HDS^{+}18, LHM09, NHS^{+}13, PNCB21, SLWS07, XLY09, XSTN14]. Vectorized [HZD^{+}23]. vectors [GI04, ST14]. vegetation [PMG^{+}22]. vehicles [KCD09, NOP^{+}18]. Veiling [TAHL07]. velocimetry [XIAP^{+}17]. Velocity [CPAB22, Erl07, HMI23, GNS^{+}12, SS11, XIAP^{+}17]. Velocity-based [Erl07]. velocity-vorticity [GNS^{+}12]. VEMPIC [TBBC^{+}22]. Venant [BJ05, KTY09]. venation [RFL^{+}05]. ventral [WKF^{+}21]. Verbal [CZL^{+}14]. vergence [TDM^{+}14]. verification [QJ21]. Versatile [AIA^{+}12, AAT13, RYPZ23, HNB^{+}06, LLDL21, TKTS11]. versus [LD06, LDS02, WQF^{+}21]. vertex [GKDS12, Man86, SNB07, TH19, YWH13]. Vertices [YCP16, BDD11, LZKW10]. vertices-based [BDD11]. Very [JGC^{+}15]. Via [POK23, Pra89, AMZ99, AW20, AAPS16, AAPS17, ALX^{+}14, ASK^{+}22, ARS14, BPK^{+}13, BR21a, BHR13, BZC^{+}23, BVS16, BS19, Bou18, CCWL18, CXW^{+}23a, CA24, CLJ^{+}20, CKT^{+}23, CSSL21, CSL^{+}22, CZXL23, CPW21, CYT^{+}18, CPS13, DGHM93, DLX^{+}21, DKT^{+}23, ED04, FYW^{+}18, FW22, Fat07, FPBCO20, FCW^{+}17, FMR20, GGY18, GPHSH19, GZC15, HNO^{+}23, HFW^{+}19, HS13, HCS13, HvKW^{+}16, HWV^{+}18, HSS^{+}13, HCW15, HWK15, HXM^{+}18, HPC21, IYYI14, JBM^{+}17, JJJ^{+}21, JW15, JKT^{+}15, KEE13, KAEE20, KYS^{+}15, KSS06, KJDL09, KTL^{+}04, KLPCP18, LLF^{+}20, LMLH07, LVKS21, LSQ^{+}15, LLM21, LZS^{+}21, LCL^{+}23, LGHL23, LVS18, LCORL07, LZH^{+}20, LZBCJ21, LSVT15, MDK^{+}16, MGA^{+}17, MSL^{+}24, MIB15, MNB23, NBLCO20, OBS04, PCLC16, PFX^{+}22, PO18, PNCB21, QZZ22, RBvB^{+}04, RPWO18, SGM12, SGSS22, SLK^{+}24, SAJ21, SJ22b, SHX^{+}22, She13, SLWL23, SBK^{+}18, SPSH^{+}17, SvKK^{+}11, SOHK16, SLMR14, SYM^{+}24, SJA^{+}20]. via [TLG17a, TLG17b, TEG18, TWBO03, THW^{+}14, TNWK22, WYL^{+}14, WLY^{+}16, WLT16, WYL^{+}20, WL21, WWWG22, WSS^{+}19, WPL18, WTBS07b, WYXJ21, XHWW22, XZZ^{+}14, XLXJ11, XYXJ12, XCS^{+}14, YC21, YNK^{+}22, YNL^{+}21, ZXJ^{+}13, ZYM^{+}20, ZSSJL20, ZLC^{+}22, ZMW^{+}23, ZTD^{+}23, ZJY^{+}22, ZWS^{+}24, ZYL^{+}17]. vibrating [BF12]. Vibration [HXK^{+}19, JBP06]. Vibration-minimizing [HXK^{+}19]. Vid2Player [ZSAF21]. Video [AČMS10, BDG15, BMBRD24, BJS^{+}08, BGSF10, Bea88, BM05, BNTS07, CWL12, CK20, CAC^{+}02, DSJ^{+}11, DLX^{+}21, FJA^{+}14, GZX^{+}22, GZC^{+}16, GF12, GXSD23, HXZ^{+}19, HLSH18, JSSH15, LLK^{+}19, LYC^{+}22, LSS05, LHM^{+}18, LXC^{+}15, PCSS06, RKS^{+}14, ST04, SBSH18, SAA^{+}21, SLL^{+}21a, SGdA^{+}10, SDA^{+}23, VSHJ12, WXSC04, WMZ^{+}13, XLS^{+}11, XCZ^{+}18, YJLL22, ZSAF21, ZZZ^{+}22, ZMW^{+}23, AWL^{+}20, AZP^{+}05, AXR09, AGB^{+}16, ASC^{+}14, BWSS09, BAAR12, BBPP10, BM07, BLA12, BSHK04, BZCC10, BSPP13, BST^{+}14, BTS^{+}15, CAD^{+}21, CTMS03, CCS^{+}15, CM10, CSRP10, CWTW17, DRW^{+}14, DCD15, FZL^{+}15, FL11, FAC11, FF11, FTZ^{+}19, GVWT13, GZW^{+}16, GO11, GCSS06, GWN^{+}03, GB08b, HKAK16, IBP15, JST^{+}19, JMK^{+}22, JLF^{+}09, JMA06, KSB^{+}13, KUWS03, KC19, KOWD21, KGT^{+}18, KWB^{+}15, KDMW17, Kop16, KLHG09, KPB^{+}12, KSE^{+}03, LDTA17, LDS^{+}11, LJH13b, LYGC15, LFH15, LGJA09, LGW^{+}11, LYTS13, LWCT14, LCD^{+}20b, MKMS04]. video [MEMS06, MDC^{+}21, MCE^{+}17, MMP^{+}05, MZRT16, MCW^{+}21, PCHF18, RAKRF08, RTS^{+}07, RSA08, SSRB^{+}17, SLJT08, SMPR07, TKTS11, TKKT12, Van06, WRDF13, WBC^{+}05, WFS^{+}09, WLSL10, WHSL11, WZK^{+}17, WMB^{+}20, WC10, WOG06, WRS^{+}12, XYJ13, YGL^{+}14, YPL21, ZWZ^{+}16, ZCT^{+}21, ZLY^{+}21, ZQPM12, ZYQ^{+}14, ZKU^{+}04, dAST^{+}08, vdHDT^{+}07, BWSS09]. Video-audio [LXC^{+}15]. Video-based [SGdA^{+}10, VSHJ12, WMZ^{+}13, XLS^{+}11, BBPP10]. Video-Driven [ZZZ^{+}22, MCW^{+}21]. Video-guided [PCSS06]. videoconferencing [EMT^{+}20]. VideoDoodles [YBMN^{+}23]. videography [XYH^{+}18, ZMN^{+}19]. VideoMocap [WC10]. videorealistic [EGP02]. Videos [LXZ^{+}19, MHU19, TWLT19, YBMN^{+}23, ZYM^{+}23, BDG15, BBPP10, CWW^{+}13a, HXFW20, JTST10, KCS14, LLZ18, LCL^{+}22, MTM16, MGC^{+}19, MNBN07, PKM^{+}18, SWTC14, SBLD15, TZT^{+}18, WLZ^{+}09, WSZ^{+}14]. Videoscapes [TKKT12]. VideoSnapping [WSZ^{+}14]. VideoTrace [vdHDT^{+}07]. Vidgets [XBZN19]. View [ASN^{+}20, Gla90, HSX^{+}22, HNH19, KL23, KLR^{+}22, LJZ^{+}23, PNTK23, PVY90, RSV^{+}23, TCS^{+}23, WBF^{+}17a, WWT^{+}03, YPA^{+}18, ZFT^{+}21, ZZT^{+}21, BMSR20, CWW^{+}12, DSAF^{+}13, DFL^{+}15, DDD^{+}14, DSC^{+}20, FZBR16, GAF^{+}10, HHC18, HMLL15, HWK15, KWR16, KQG^{+}23, Kou16, KYC^{+}17, LD21, LACS08, LAGP09, LTJ18, LHR^{+}21, MLR^{+}14, MDC^{+}21, MSOC^{+}19, NMD^{+}17, NOP^{+}18, NZV^{+}11, ODAO15, PZ17, PGZ^{+}19, PMGD21, SHL^{+}17, SHZ^{+}20, VBK05, VBMP08, VPB^{+}09b, WBF^{+}17b, WLH^{+}13, XLS^{+}11, XLX^{+}16, XBS^{+}19, ZCW^{+}17, ZTF^{+}18, ZKU^{+}04, dAST^{+}08]. view- [BMSR20]. View-Dependent [PNTK23, WWT^{+}03]. view-enhanced [DFL^{+}15]. View-Synthesis [PNTK23]. viewer [NYY04, WTS^{+}23, YLL^{+}22]. viewer-perceived [YLL^{+}22]. viewers [SLV^{+}13]. viewfinder [BPK^{+}13]. Viewing [CLJ^{+}20, FKN17, KUDC07, KNC^{+}08]. Viewpoint [HNH19, HSV^{+}22, AAC^{+}06, CTMS03, CCS^{+}15, GCD^{+}20, HPP^{+}18, PMGD21, SLF^{+}11, TFK^{+}03, YZL^{+}22, ZLY^{+}21]. Views [SYZ^{+}23, HMC11, WOQS05]. ViRheometry [HNO^{+}23]. Virtual [ANL^{+}23, ACP^{+}01, AS21, DFYL19, DCT^{+}22, FSRS22, HKWB09, HC86, JWD^{+}23, KAW20, LLL22, LBW^{+}23, MNV^{+}21, NNDJ12, RSM^{+}23, TZS^{+}18, WBF^{+}17a, WBF^{+}17b, YNK^{+}22, ALY08, AGB^{+}16, BM05, CGP^{+}21, DKH^{+}10, Did18, EVC^{+}15, EAPL06, HMO12, HRZ^{+}13, JWW^{+}20, KDMW17, KKW20, KKB^{+}11, KOOP11, LSL^{+}18, LCL06, LHLY21, LNWB03, MGK17, MBB12, MIWB02, MSSG^{+}21, MBF04, OEE^{+}18, PSK^{+}16, RRS19, SMG^{+}05, SSRB^{+}17, SMG^{+}20, SSC10, SBK11, SWK16, SPW^{+}18, TGD04, ZCB^{+}22]. VirtualStudio2Go [GB08b]. viscoelastic [BGFAO17, FLGJ19, GBO04, SXH^{+}21, WT08]. viscoplastic [BWHT07, TLZ^{+}24]. viscosity [GWAB19, LBB17a, NSS^{+}19, PICT15, TB20]. Viscous [PGG^{+}23, BUAG12, BAV^{+}10, LBB17a, VRBC18]. viseme [ELFS16]. Visemenet [ZXL^{+}18]. Visibility [ASL^{+}17, SS00, TD23, Wil92, ZWRY21, BGAM12, BMW^{+}09, DSDD07, DD02a, DDP99, DDP02, EPD09, GBAM11, HJ11a, KTB07, LSCO03, MKRH11, MGT^{+}03, RAMN12, WWZ^{+}06]. Visibility-consistent [ASL^{+}17]. Visible [KL23, SG82, VKW^{+}23, WGY^{+}18, WS85, HDC07, MDC^{+}21]. Visio [MPK09]. Visio-lization [MPK09]. vision [MTA^{+}20, OPOD10, SMHW16, SARW^{+}15, WM14]. vision-guided [MTA^{+}20]. VisionWand [CB04]. Visual [CXW^{+}05, DA18, DGVG^{+}23, FR22, GWBN24, JGC^{+}15, JGMR23, LYY^{+}17, MGDA^{+}15, NWYM19, PKD^{+}19, RFWB07, SBLD15, VVCOS23, VMKK00, WK95, XGZ^{+}23, YPG01, ZCS^{+}22, ARS14, BB15, DRW^{+}14, DK99, DMHG13, DDD^{+}14, EML^{+}18, GSCO12, HWBR14, KRF^{+}18, KSSI17, LWW08, MKRH11, MWH^{+}09, ODGK03, POAR12, PCS23a, PCLC16, SCS^{+}08, SMHW16, SMGE11, WWS^{+}05, YPB16, YCL^{+}17, ZLE14]. VisualIDs [LRFN04]. Visualization [DII23, FSRS22, Shn92, BDM09, CKPS17, CGG^{+}04, DKP11, GCSS06, GGT17, HTER04, HZG09, NHAH03, RFL^{+}05, WKR99, vW02, vW09]. Visualizing [HFK94, KK91, WF96, KGFF14, VWJ^{+}13]. Visuomotor [EHSN20, YLNP12]. ViT [TBTA^{+}24]. Vivace [FTP16]. VizGen [YPB16]. VNect [MSS^{+}17]. VoCo [JMD^{+}17]. voice [TFK^{+}03]. void [LVS18]. VolCCD [TMY^{+}11]. Volume [ASGS23, AMG^{+}19, AMB^{+}21, AFC^{+}10, BBC22, HZE^{+}19, ISF07, KL23, KLM24, Lev90, LCORL07, LEQ^{+}07, Mal93, Tar16, ZZL^{+}23, AAM03, BTFN^{+}08, BKR^{+}05, DWW^{+}18, GZB^{+}13, HJ11b, JTSW17, KLL^{+}07, MAYZ^{+}20, MCSA15, McC00, NDMKJ22, ODAO15, TMY^{+}11, WBS07, WFP12]. Volume-aware [AMG^{+}19]. Volume-encoded [Tar16]. Volumes [SVB17a, SLL^{+}21a, CPS15, KHLN17, LAA^{+}05, LSS^{+}19, Mus13, PRK^{+}17, PSF09, SAJ21, SOA11, SVB17b, WYZG11, ZHRB13]. Volumetric [ASGS23, AONA22, DPW15, FPSG22, GLZ^{+}21, HC23, MJJG18, NCB23, OKH^{+}16, ONOI04, PBS20, RMD04, RKB^{+}23, TSNI10, TWR^{+}23, US24, ABL^{+}21, ACA^{+}19, BCRK^{+}10, BJ17, CSK^{+}22, CBI13, DJBJ19, DDF^{+}17, FLP14, GKH^{+}13, GWB05, GHV^{+}18, HR13, JNSJ11, KGB^{+}09, KGH^{+}14, LYP^{+}18, LCH^{+}21, LSS^{+}21, LSCS14, MPH^{+}20, MCK13, NJS^{+}11, PSNB13, SHM^{+}18, VJK21, WLT22, XFCT18, ZJMB11, ZHS^{+}05, ZDI^{+}15]. VoroCrust [ABE^{+}20]. Voronoi [LL10, ABE^{+}20, BLdG^{+}16, GS85, LWL^{+}09, LXY^{+}16, LFXH17, MDL16, MHSL18, RSLL18, SGG^{+}06, XWX^{+}22]. Vortex [IWC22, DBWG15, PTG12, SRF05, WP10, XTZ^{+}21]. vortical [XWWZ22]. vortices [GGT17]. vorticity [GNS^{+}12, ZBG15a]. vorticle [Ang17]. voting [LF09]. Voxel [YLJ18, KSA13, NZIS13]. voxelization [SS10a]. voxelized [SKOA14]. voxels [GM05, LLMZ16]. VR [AKG^{+}23, BYLR20, DWX^{+}21, DLP^{+}23, MWHL21, OLSL16, RAR^{+}21, SBSH18, WSS^{+}19, ZLC^{+}22]. VR/AR [ZLC^{+}22]. vs [FLB16]. VToonify [YJLL22].

Walk [HZE^{+}19, MSCG23, SCJ^{+}23, SMGC23]. Walk-on-Boundary [SCJ^{+}23]. Walking [DFYL19, CBYvdP08, CBvdP10, DFZ^{+}17, JKH^{+}22, SPW^{+}18, WFH09, WFH10]. Walks [PM95, LT20]. wall [AHM^{+}15, BTFN^{+}08, SWL^{+}22]. Wallpaper [WSH19]. WallPlan [SWL^{+}22]. wand [CB04]. Wang [CSHD03, KCODL06, LD06, LEQ^{+}07]. Warp [GSZ^{+}18, ZIT^{+}19, LKG^{+}03b, WLSL10]. Warp-and-project [ZIT^{+}19]. Warp-guided [GSZ^{+}18]. WarpDriver [WLP16]. Warped [XBLZ23, BLD20]. Warped-Area [XBLZ23, BLD20]. Warping [KL23, LKE18, ATDP11, HCS13, KC21, LSC^{+}12, NFL12, VPB^{+}09a, VBBF16]. warps [CAA10, CDSHD13, LGJA09, MJBF02]. Wasserstein [BPC16, QCHC17a, QCHC17b, SdGP^{+}15]. Water [JW15, JW17, JSMF^{+}18, JW23, WMT05, XAW^{+}23, BNK10, CMT^{+}16, CM11, EB14, EMF02, GSLF05, HHP^{+}21, IGLF06, LZJ16, LGF04, NØ13, SB12, SHW19, SRF05, SSJ^{+}20, SSK05a, TGK^{+}17]. watercolorization [BNTS07]. Watertight [SFL^{+}08]. Wave [JW15, LWO19, MRA^{+}13, SSJ^{+}20, TB87, XWH^{+}23, YMR^{+}13, YXW^{+}23, YHK07, AR15, BWC^{+}23, CMT^{+}16, CRG^{+}20, CGP^{+}21, CQD^{+}18, GJZ21, JW17, LGX^{+}13, LGK^{+}16, RSM^{+}10a, RS14a, RTK^{+}15, SHW19, WQLJ18, WVJH17, XWM^{+}20, YHW^{+}18, ZHLB10]. Wave-based [LWO19, MRA^{+}13, WQLJ18, ZHLB10]. wave-optical [WVJH17]. Wave-ray [YMR^{+}13]. Wave-Tracing [TB87]. Wavefront [JW15, QHY^{+}16]. Wavelet [CJAMJ05, CD05, HHL^{+}24, JLF^{+}23, KTJG08, MU22, GHBCO21, NRH03, NRH04, ODR09, SM06, SR09]. Wavelet-Based [JLF^{+}23]. Wavelet-domain [HHL^{+}24]. wavelet-driven [GHBCO21]. Wavelets [CSS96, Fat09a, JSMF^{+}18, LF08]. Waves [CCL^{+}22, JW23, TB87, HQT^{+}21, NSB13, SHL^{+}17, SSJ^{+}20]. Way [GWBN24, FQL^{+}20, HFG^{+}18, LCD^{+}20a, LTJ18, NGL10, RMSG^{+}08, TB20, WCL^{+}23]. weak [SZB18, ZLB16a]. Weakly [CHY21, LBW^{+}23, SSK^{+}17]. Weakly-Compressible [LBW^{+}23]. Weakly-supervised [CHY21, SSK^{+}17]. weaknesses [CHM^{+}12]. wearable [ZSZ^{+}14]. Weather [GDAB^{+}17a, GDAB^{+}17b]. weathering [BKCO16, BLR^{+}11, CXW^{+}05]. Weatherscapes [HHP^{+}21]. Weavecraft [WZL^{+}20]. Weaving [VZF^{+}19, ACXG09, CK14b, RPC^{+}21, STP12, WZL^{+}20]. web [PCLC16]. Webcam [LEN09]. Weight [BL18, LD13, LSZ^{+}14]. Weighted [DSZ17, Fol87, MCY14, PBDSH13, dGMMD14, Ale20, BN21, WYL^{+}14]. Weighting [NID20]. weights [JBPS11, WS21]. Weingarten [PKPP21]. well [CSD^{+}09, VSK^{+}17]. wet [WFS22]. Wetbrush [CKIW15]. wheels [GPD^{+}18]. Where [CGL^{+}08]. Which [SZC^{+}22]. while [SLS^{+}16]. Whippletree [SKB^{+}14]. Whirlpools [OGN^{+}23]. White [HHX^{+}18, BBPD12, HMP^{+}08, LYC18]. White-Box [HHX^{+}18]. whitewater [WFS22]. whole [MTA^{+}20]. whole-body [MTA^{+}20]. wide [CAA09, MLR^{+}14, MDC^{+}21, NYY04, SHL^{+}17, SLL19, TAV^{+}10]. wide-angle [CAA09, SLL19, TAV^{+}10]. widgets [BL15, XBZN19]. wiggling [KySK09]. wiggly [KA08]. Wikipedia [RMBB^{+}13]. Wild [SSSH17, BBS14b, FFBB21, HZG^{+}18, RRFG17]. wildfires [HBP^{+}21]. Willmore [GA20, SCD^{+}21]. Wind [LBW^{+}23, AR15, SWT^{+}17, UPSW16]. wind-up [SWT^{+}17]. Winding [FGC23, XDW^{+}23, BDS^{+}18, JKSH13]. Winding-Number [XDW^{+}23]. Window [HC86, SG86, Wes88, BG84]. Window-Based [HC86, Wes88]. Windy [PNH^{+}14]. Wire [GSFD^{+}14, HHC18, ILB15, LCL^{+}17, XKCB18, YXFH21]. wired [Xu18]. wireframe [WPGM16]. wireless [ICG17, RBvB^{+}04]. WireRoom [YXFH21]. wires [LFZ18]. within [MCSA15, PKCH18, SSC10, WWOH08]. Without [ABE^{+}20, FKN17, LXSW23, MYWI15, PGML^{+}19, SLV^{+}13, SJWG20]. Wood [IWHH20, LIY^{+}22, LDHM16, MWAM05, PJH^{+}17]. Word [IVH^{+}23]. Word-As-Image [IVH^{+}23]. words [BBGO11, SQRH^{+}16]. work [MYY^{+}10]. workbench [Ano03]. workflows [DTP15]. workloads [SKB^{+}14]. workplans [ZHPY21]. works [KLY^{+}14]. works-like [KLY^{+}14]. Workspaces [HC86, ZHPY21]. World [SBSH18, SRX^{+}23, ALY08, DvGNK99, HZ82, RBvB^{+}04, SGSS08, WRS^{+}12]. WorldBrush [EVC^{+}15]. worlds [EVC^{+}15, GJ22, TJ07, YYW^{+}12a]. Worst [McK87, PRZ17, ZPZ13, SZB18]. Worst-case [McK87, PRZ17, ZPZ13, SZB18]. Woven [GHCG17, CLMMO14, IM12, MGZJ20, ZJMB12]. WRAPD [BN21]. WrapIt [ILB15]. wrapped [ILB15]. Wrapping [LXG^{+}22, IRHSH20, PRLH^{+}22]. Wrinkle [CKSV23, RK13, WHRO10]. Wrinkled [HSF07]. wrinkles [RPC^{+}10, Wan21]. Wrinkling [CCK^{+}21, RPC^{+}10]. writing [PGML^{+}19, PFX^{+}22]. WYSIWYG [BPK^{+}13, KMM^{+}02].

X [BMSR20, IYYI14, PKLI^{+}19, SG86]. X-Fields [BMSR20]. X-ray [IYYI14]. X-Shells [PKLI^{+}19]. x86 [SCS^{+}08].

Yarn [CLMMO14, KJM08, KJM10, LWS^{+}18, SNW20, SNW21, SSBL^{+}22, YKJM12, ZLB16b]. yarn-based [KJM10]. Yarn-level [CLMMO14, LWS^{+}18, SNW20, SSBL^{+}22, YKJM12]. Year [Ano90b]. yields [FV96]. YIQ [SCB87].

Z  [JLBM05].  Z-

buffer [JLBM05]. Zero [CWL22, HZE^{+}19, HCW^{+}23, HZP^{+}22]. Zero-Shot [CWL22, HZP^{+}22]. Zero-Variance [HZE^{+}19]. zip [FTD21]. zippables [SPSH18]. ZoeMatrope [MIWI16]. zonal [NSF12]. Zones [RV89]. Zoom [WLS^{+}23, BGKS17, BGSF10]. Zoomorphic [DYYT15].

References

Adamson:2006:PSC

[AA06] Anders Adamson and Marc Alexa. Point-sampled cell complexes. ACM Transactions on Graphics, 25(3): 671–680, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alexa:2009:IPS

[AA09] Marc Alexa and Anders Adamson. Interpolatory point set surfaces — convexity and Hermite data. ACM Transactions on Graphics, 28(2): 20:1–20:20, April 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Agarwala:2006:PLS

[AAC^{+}06] Aseem Agarwala, Maneesh Agrawala, Michael Cohen, David Salesin, and Richard Szeliski. Photographing long scenes with multi-viewpoint panoramas. ACM Transactions on Graphics, 25(3): 853–861, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aittala:2016:RMN

[AAL16] Miika Aittala, Timo Aila, and Jaakko Lehtinen. Reflectance modeling by neural texture synthesis. ACM Transactions on Graphics, 35 (4):65:1–65:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Assarsson:2003:GBS

[AAM03] Ulf Assarsson and Tomas Akenine-Möller. A geometrybased soft shadow volume algorithm using graphics hardware. ACM Transactions on Graphics, 22(3):511–520, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ansari:2020:MII

[AAMSB20] Navid Ansari, Omid Alizadeh-Mousavi, Hans-Peter Seidel, and Vahid Babaei. Mixed integer ink selection for spectral reproduction. ACM Transactions on Graphics, 39(6): 255:1–255:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417761.

Aksoy:2016:IHQ

[AAPS16] Yagiz Aksoy, Tunç Ozan Aydin, Marc Pollefeys, and Aljosa Smolić. Interactive high-quality green-screen keying via color unmixing. ACM Transactions on Graphics, 35 (5):152:1–152:??, September 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aksoy:2017:IHQ

[AAPS17] Yagiz Aksoy, Tunç Ozan Aydin, Marc Pollefeys, and Aljoa Smolić. Interactive highquality green-screen keying via color unmixing. ACM Transactions on Graphics, 36 (4):61:1–61:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alregib:2005:ERT

[AAR05] Ghassan Alregib, Yucel Altunbasak, and Jarek Rossignac. Error-resilient transmission of 3D models. ACM Transactions on Graphics, 24(2): 182–208, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aksoy:2017:UBSb

[AASP17a] Yagiz Aksoy, Tunç Ozan Aydin, Aljoa Smoli, and Marc Pollefeys. Unmixing-based soft color segmentation for image manipulation. ACM Transactions on Graphics, 36 (4):61:1–61:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aksoy:2017:UBSa

[AASP17b] Yagiz Aksoy, Tunç Ozan Aydin, Aljosa Smolić, and Marc Pollefeys. Unmixing-based soft color segmentation for image manipulation. ACM Transactions on Graphics, 36 (2):19:1–19:??, April 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Akinci:2013:VST

[AAT13] Nadir Akinci, Gizem Akinci, and Matthias Teschner. Versatile surface tension and adhesion for SPH fluids. ACM Transactions on Graphics, 32 (6):182:1–182:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Abhyankar:1989:APR

[AB89] Shreeram S. Abhyankar and Chanderjit J. Bajaj. Automatic parameterization of rational curves and surfaces IV: algebraic space curves. ACM Transactions on Graphics, 8(4):325–334, October 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        77273.html.

Alexa:2008:SS

[AB08] Marc Alexa and Tamy Boubekeur. Subdivision shading. ACM Transactions on Graphics, 27 (5):142:1–142:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ando:2020:POL

[AB20] Ryoichi Ando and Christopher Batty. A practical octree liquid simulator with adaptive surface resolution. ACM Transactions on Graphics, 39 (4):32:1–32:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392460.

Andujar:2002:TRS

[ABA02] Carlos Andújar, Pere Brunet, and Dolors Ayala. Topologyreducing surface simplification using a discrete solid representation. ACM Transactions on Graphics, 21(2):88–105, April 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Abdelkader:2020:VVM

[ABE^{+}20] Ahmed Abdelkader, Chandrajit L. Bajaj, Mohamed S. Ebeida, Ahmed H. Mahmoud, Scott A. Mitchell, John D. Owens, and Ahmad A. Rushdi. Voro-Crust: Voronoi meshing without clipping. ACM Transactions on Graphics, 39(3): 23:1–23:16, June 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3337680.

Alterman:2021:ILS

[ABGL21] Marina Alterman, Chen Bar, Ioannis Gkioulekas, and Anat Levin. Imaging with local speckle intensity correlations: Theory and practice. ACM Transactions on Graphics, 40(3):30:1–30:22, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3447392.

Ayala:1985:ORM

[ABJN85] D. Ayala, P. Brunet, R. Juan, and I. Navazo. Object representation by means of nonminimal division quadtrees and octrees. ACM Transactions on Graphics, 4(1): 41–59, January 1985. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        3975.html.

Abdrashitov:2021:IMV

[ABL^{+}21] Rinat Abdrashitov, Seungbae Bang, David Levin, Karan Singh, and Alec Jacobson. Interactive modelling of volumetric musculoskeletal anatomy. ACM Transactions on Graphics, 40(4): 122:1–122:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459769.

Azevedo:2016:PGT

[ABO16] Vinicius C. Azevedo, Christopher Batty, and Manuel M. Oliveira. Preserving geometry and topology for fluid flows with thin obstacles and narrow gaps. ACM Transactions on Graphics, 35(4): 97:1–97:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ament:2014:RRT

[ABW14] Marco Ament, Christoph Bergmann, and Daniel Weiskopf. Refractive radiative transfer equation. ACM Transactions on Graphics, 33(2):17:1– 17:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Achar:2017:ETF

[ABW^{+}17] Supreeth Achar, Joseph R.

Bartels, William L. ‘Red’ Whittaker, Kiriakos N. Kutulakos, and Srinivasa G. Narasimhan. Epipolar timeof-flight imaging. ACM Transactions on Graphics, 36(4): 37:1–37:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Acar:2007:LSD

[Aca07]  Rüyam Acar.  Level set

driven flows. ACM Transactions on Graphics, 26(4):15:1– 15:15, October 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Araujo:2019:SSS

[ACA^{+}19] Chrystiano Araújo, Daniela Cabiddu, Marco Attene, Marco Livesu, Nicholas Vining, and Alla Sheffer. Surface2Volume: surface segmentation conforming assemblable volumetric partition. ACM Transactions on Graphics, 38(4):80:1–80:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Azencot:2017:CFC

[ACBCO17] Omri Azencot, Etienne Corman, Mirela Ben-Chen, and Maks Ovsjanikov. Consistent functional cross field design for mesh quadrangulation. ACM Transactions on Graphics, 36(4):92:1–92:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Abhyankar:1990:IIA

[ACC90] Shreeram S. Abhyankar, Srinivasan Chandrasekar, and Vijaya Chandru. Improper intersection of algebraic curves. ACM Transactions on Graphics, 9(2):147–159, April 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
78957.html.

Assa:2005:ASP

[ACCO05] Jackie Assa, Yaron Caspi, and Daniel Cohen-Or. Action synopsis: pose selection and illustration. ACM Transactions on Graphics, 24(3):667–676, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aydin:2010:VQA

[AČMS10] Tunç Ozan Aydin, Martin Čadík, Karol Myszkowski, and Hans-Peter Seidel. Video quality assessment for computer graphics applications. ACM Transactions on Graphics, 29(6):161:1–161:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aristidou:2018:DMM

[ACOH^{+}18] Andreas Aristidou, Daniel Cohen-Or, Jessica K. Hodgins, Yiorgos Chrysanthou, and Ariel Shamir. Deep motifs and motion signatures. ACM Transactions on Graphics, 37 (6):187:1–187:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Assa:2008:MOH

[ACOYL08] Jackie Assa, Daniel Cohen-Or, I-Cheng Yeh, and Tong-Yee Lee. Motion overview of human actions. ACM Transactions on Graphics, 27(5): 115:1–115:??, December 2008.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alonso:2001:VMG

[ACP^{+}01] L. Alonso, F. Cuny, S. Petitjean, J.-C. Paul, S. Lazard, and E. Wies. The virtual mesh: a geometric abstraction for efficiently computing radiosity. ACM Transactions on Graphics, 20(3):169–201, July 2001. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Allen:2002:ABD

[ACP02] Brett Allen, Brian Curless, and Zoran Popović. Articulated body deformation from range scan data. ACM Transactions on Graphics, 21(3): 612–619, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Allen:2003:SHB

[ACP03] Brett Allen, Brian Curless, and Zoran Popović. The space of human body shapes: reconstruction and parameterization from range scans. ACM Transactions on Graphics, 22 (3):587–594, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alliez:2003:APR

[ACSD^{+}03] Pierre Alliez, David Cohen-Steiner, Olivier Devillers, Bruno Lévy, and Mathieu Desbrun. Anisotropic polygonal remeshing. ACM Transactions on Graphics, 22(3): 485–493, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Antani:2012:ISP

[ACSM12] Lakulish Antani, Anish Chandak, Lauri Savioja, and Dinesh Manocha. Interactive sound propagation using compact acoustic transfer operators. ACM Transactions on Graphics, 31(1): 7:1–7:12, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alliez:2005:VTM

[ACSYD05] Pierre Alliez, David Cohen-Steiner, Mariette Yvinec, and Mathieu Desbrun. Variational tetrahedral meshing. ACM Transactions on Graphics, 24 (3):617–625, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Akleman:2009:CPW

[ACXG09] Ergun Akleman, Jianer Chen, Qing Xing, and Jonathan L. Gross. Cyclic plain-weaving on polygonal mesh surfaces with graph rotation systems. ACM Transactions on Graphics, 28(3):78:1–78:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adams:2003:IBO

[AD03] Bart Adams and Philip Dutré. Interactive boolean operations on surfel-bounded solids. ACM Transactions on Graphics, 22(3):651–656, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Agarwala:2004:IDP

[ADA^{+}04] Aseem Agarwala, Mira Dontcheva, Maneesh Agrawala, Steven Drucker, Alex Colburn, Brian Curless, David Salesin, and Michael Cohen. Interactive digital photomontage. ACM Transactions on Graphics, 23 (3):294–302, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adams:2021:FMF

[Ada21] Andrew Adams. Fast median filters using separable sorting networks. ACM Transactions on Graphics, 40 (4):70:1–70:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459773.

Annen:2008:RTA

[ADM^{+}08] Thomas Annen, Zhao Dong, Tom Mertens, Philippe Bekaert, Hans-Peter Seidel, and Jan Kautz. Real-time, allfrequency shadows in dynamic scenes. ACM Transactions on Graphics, 27(3):

34:1–34:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Averbuch-Elor:2015:RRO

[AECO15] Hadar Averbuch-Elor and Daniel Cohen-Or. RingIt: Ring-ordering casual photos of a temporal event. ACM Transactions on Graphics, 34 (3):33:1–33:??, April 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Averbuch-Elor:2017:BPL

[AECOKC17] Hadar Averbuch-Elor, Daniel Cohen-Or, Johannes Kopf, and Michael F. Cohen. Bringing portraits to life. ACM Transactions on Graphics, 36 (6):196:1–196:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arikan:2002:IMG

[AF02] Okan Arikan and D. A.

Forsyth. Interactive motion generation from examples. ACM Transactions on Graphics, 21(3):483–490, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Allard:2010:VCC

[AFC^{+}10] Jérémie Allard, François Faure, Hadrien Courtecuisse, Florent Falipou, Christian Duriez, and Paul G. Kry. Volume contact constraints at arbitrary resolution. ACM Transactions on Graphics, 29 (4):82:1–82:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adikusuma:2020:FCD

[AFH20] Yohanes Yudhi Adikusuma, Zheng Fang, and Ying He. Fast construction of discrete geodesic graphs. ACM Transactions on Graphics, 39(2): 14:1–14:14, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3144567.

Avrahami:2023:BLD

[AFL23] Omri Avrahami, Ohad Fried, and Dani Lischinski. Blended latent diffusion. ACM Transactions on Graphics, 42(4): 149:1–149:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592450.

Arikan:2003:MSA

[AFO03] Okan Arikan, David A. Forsyth, and James F. O’Brien. Motion synthesis from annotations. ACM Transactions on Graphics, 22 (3):402–408, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arikan:2005:FDA

[AFO05] Okan Arikan, David A.

Forsyth, and James F. O’Brien. Fast and detailed approximate global illumination by irradiance decomposition. ACM Transactions on Graphics, 24(3):1108–1114, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Auslander:1995:FEC

[AFP^{+}95] Joel Auslander, Alex Fukunaga, Hadi Partovi, Jon Christensen, Lloyd Hsu, Peter Reiss, Andrew Shuman, Joe Marks, and J. Thomas Ngo. Further experience with controller-based automatic motion synthesis for articulated figures. ACM Transactions on Graphics, 14 (4):311–336, October 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
225295.html.

Akyuz:2007:DHD

[AFR^{+}07] Ahmet Oğuz Akyüz, Roland Fleming, Bernhard E. Riecke, Erik Reinhard, and Heinrich H. Bülthoff. Do HDR displays support LDR content?: a psychophysical evaluation. ACM Transactions on Graphics, 26(3):38:1–38:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Attene:2003:SRT

[AFSR03] Marco Attene, Bianca Falcidieno, Michela Spagnuolo, and Jarek Rossignac. Swing-Wrapper: Retiling triangle meshes for better edgebreaker compression. ACM Transactions on Graphics, 22(4): 982–996, October 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Au:2007:HAI

[AFTCO07] Oscar Kin-Chung Au, Hongbo Fu, Chiew-Lan Tai, and Daniel Cohen-Or. Handleaware isolines for scalable shape editing. ACM Transactions on Graphics, 26(3): 83:1–83:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Apitz:2005:CCB

[AG05] Georg Apitz and François Guimbretière. CrossY: a crossing-based drawing application. ACM Transactions on Graphics, 24(3):930, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Agarwala:2007:EGD

[Aga07] Aseem Agarwala. Efficient gradient-domain compositing using quadtrees. ACM Transactions on Graphics, 26(3): 94:1–94:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anderson:2016:JVR

[AGB^{+}16] Robert Anderson, David Gallup, Jonathan T. Barron, Janne Kontkanen, Noah Snavely, Carlos Hernández, Sameer Agarwal, and Steven M. Seitz. Jump: virtual reality video. ACM Transactions on Graphics, 35(6): 198:1–198:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adams:2009:GKT

[AGDL09] Andrew Adams, Natasha Gelfand, Jennifer Dolson, and Marc Levoy. Gaussian KDtrees for fast high-dimensional filtering. ACM Transactions on Graphics, 28(3):21:1– 21:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2022:NJF

[AGK^{+}22] Noam Aigerman, Kunal Gupta, Vladimir G. Kim, Siddhartha Chaudhuri, Jun Saito, and Thibault Groueix. Neural Jacobian fields: learning intrinsic mappings of arbitrary meshes. ACM Transactions on Graphics, 41 (4):109:1–109:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530141.

Aanjaneya:2017:PDS

[AGL^{+}17] Mridul Aanjaneya, Ming Gao, Haixiang Liu, Christopher Batty, and Eftychios Sifakis. Power diagrams and sparse paged grids for high resolution adaptive liquids. ACM Transactions on Graphics, 36(4): 140:1–140:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ao:2022:RGR

[AGL^{+}22] Tenglong Ao, Qingzhe Gao, Yuke Lou, Baoquan Chen, and Libin Liu. Rhythmic gesticulator: Rhythm-aware co-speech gesture synthesis with hierarchical neural embeddings. ACM Transactions on Graphics, 41(6): 209:1–209:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555435.

Argudo:2020:SMA

[AGP^{+}20] Oscar Argudo, Eric Galin, Adrien Peytavie, Axel Paris, and Eric Guérin. Simulation, modeling and authoring of glaciers. ACM Transactions on Graphics, 39(6): 177:1–177:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417855.

Ahn:2021:KSL

[AGS21] Byeongjoo Ahn, Ioannis Gkioulekas, and Aswin C. Sankaranarayanan. Kaleidoscopic structured light. ACM Transactions on Graphics, 40 (6):214:1–214:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480524.

Andersson:2015:MDC

[AHAM15] Magnus Andersson, Jon Hasselgren, and Tomas Akenine-Möller. Masked depth culling for graphics hardware. ACM Transactions on Graphics, 34 (6):188:1–188:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Auzinger:2018:CDN

[AHB18] Thomas Auzinger, Wolfgang Heidrich, and Bernd Bickel. Computational design of nanostructural color for additive manufacturing. ACM Transactions on Graphics, 37(4):159:1–159:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ahmed:2015:APP

[AHD15] Abdalla G. M. Ahmed, Hui Huang, and Oliver Deussen. AA patterns for point sets with controlled spectral properties. ACM Transactions on Graphics, 34(6):212:1– 212:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alexa:2017:ODSa

[AHL17a] Marc Alexa, Kristian Hildebrand, and Sylvain Lefebvre. Optimal discrete slicing. ACM Transactions on Graphics, 36 (1):12:1–12:??, February 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alexa:2017:ODSb

[AHL17b] Marc Alexa, Kristian Hildebrand, and Sylvain Lefebvre. Optimal discrete slicing. ACM Transactions on Graphics, 36 (4):64:1–64:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ali-Hamadi:2013:AT

[AHLG^{+}13] Dicko Ali-Hamadi, Tiantian Liu, Benjamin Gilles, Ladislav Kavan, François Faure, Olivier Palombi, and Marie-Paule Cani. Anatomy transfer. ACM Transactions on Graphics, 32(6):188:1–188:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adib:2015:CHF

[AHM^{+}15] Fadel Adib, Chen-Yu Hsu, Hongzi Mao, Dina Katabi, and Frédo Durand. Capturing the human figure through a wall. ACM Transactions on Graphics, 34(6):219:1– 219:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Agarwala:2004:KBT

[AHSS04] Aseem Agarwala, Aaron Hertzmann, David H. Salesin, and Steven M. Seitz. Keyframebased tracking for rotoscoping and animation. ACM Transactions on Graphics, 23(3): 584–591, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Akinci:2012:VRF

[AIA^{+}12] Nadir Akinci, Markus Ihmsen, Gizem Akinci, Barbara Solenthaler, and Matthias Teschner. Versatile rigidfluid coupling for incompressible SPH. ACM Transactions on Graphics, 31(4): 62:1–62:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Atcheson:2008:TRC

[AIH^{+}08] Bradley Atcheson, Ivo Ihrke, Wolfgang Heidrich, Art Tevs, Derek Bradley, Marcus Magnor, and Hans-Peter Seidel. Time-resolved 3D capture of non-stationary gas flows. ACM Transactions on Graphics, 27(5):132:1– 132:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adams:2010:FEP

[AJD^{+}10] Andrew Adams, David E. Jacobs, Jennifer Dolson, Marius Tico, Kari Pulli, Eino-Ville Talvala, Boris Ajdin, Daniel Vaquero, Hendrik P. A. Lensch, Mark Horowitz, Sung Hee Park, Natasha Gelfand, Jongmin Baek, Wojciech Matusik, and Marc Levoy. The Frankencamera: an experimental platform for computational photography. ACM Transactions on Graphics, 29(4):29:1–29:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

An:2012:MDC

[AJM12] Steven S. An, Doug L. James, and Steve Marschner. Motiondriven concatenative synthesis of cloth sounds. ACM Transactions on Graphics, 31(4): 102:1–102:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Abdrashitov:2020:SEP

[AJS20] Rinat Abdrashitov, Alec Jacobson, and Karan Singh. A system for efficient 3D printed stop-motion face animation. ACM Transactions on Graphics, 39(1):1:1– 1:11, February 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3360510.

Amenta:2004:DPS

[AK04] Nina Amenta and Yong Joo Kil. Defining point-set surfaces. ACM Transactions on Graphics, 23(3):264– 270, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aizenman:2023:SEM

[AKG^{+}23] Avi M. Aizenman, George A.

Koulieris, Agostino Gibaldi, Vibhor Sehgal, Dennis M. Levi, and Martin S. Banks. The statistics of eye movements and binocular disparities during VR gaming: Implications for headset design. ACM Transactions on Graphics, 42(1):7:1–7:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3549529.

An:2008:OCE

[AKJ08] Steven S. An, Theodore Kim, and Doug L. James. Optimizing cubature for efficient integration of subspace deformations. ACM Transactions on Graphics, 27(5):

165:1–165:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2017:SOT

[AKL17] Noam Aigerman, Shahar Z. Kovalsky, and Yaron Lipman. Spherical orbifold Tutte embeddings. ACM Transactions on Graphics, 36(4): 90:1–90:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aberman:2017:DTS

[AKZ^{+}17] Kfir Aberman, Oren Katzir, Qiang Zhou, Zegang Luo, Andrei Sharf, Chen Greif, Baoquan Chen, and Daniel Cohen-Or. Dip transform for 3D shape reconstruction. ACM Transactions on Graphics, 36(4):79:1–79:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2013:IBD

[AL13] Noam Aigerman and Yaron Lipman. Injective and bounded distortion mappings in 3D. ACM Transactions on Graphics, 32(4):106:1– 106:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2015:OTE

[AL15] Noam Aigerman and Yaron Lipman. Orbifold Tutte embeddings. ACM Transactions on Graphics, 34(6): 190:1–190:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2016:HOT

[AL16] Noam Aigerman and Yaron Lipman. Hyperbolic orbifold Tutte embeddings. ACM Transactions on Graphics, 35 (6):217:1–217:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alexa:2002:LCT

[Ale02] Marc Alexa. Linear combination of transformations. ACM Transactions on Graphics, 21 (3):380–387, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alexa:2019:HT

[Ale19] Marc Alexa. Harmonic triangulations. ACM Transactions on Graphics, 38(4): 54:1–54:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alexa:2020:CWD

[Ale20]  Marc Alexa.  Conform-

ing weighted Delaunay triangulations. ACM Transactions on Graphics, 39(6): 248:1–248:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417776.

Aksit:2017:NEV

[ALK^{+}17] Kaan Aksit, Ward Lopes, Jonghyun Kim, Peter Shirley, and David Luebke. Neareye varifocal augmented reality display using see-through screens. ACM Transactions on Graphics, 36(6):189:1– 189:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aberman:2020:SAN

[ALL^{+}20] Kfir Aberman, Peizh Uo Li, Dani Lischinski, Olga Sorkine-Hornung, Daniel Cohen-Or, and Baoquan Chen. Skeletonaware networks for deep motion retargeting. ACM Transactions on Graphics, 39(4): 62:1–62:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392462.

Anderson:2017:AED

[ALLD17] Luke Anderson, Tzu-Mao Li, Jaakko Lehtinen, and Frédo Durand. Aether: an embedded domain specific sampling language for Monte Carlo rendering. ACM Transactions on Graphics, 36(4): 99:1–99:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aberman:2018:NBB

[ALS^{+}18] Kfir Aberman, Jing Liao, Mingyi Shi, Dani Lischinski, Baoquan Chen, and Daniel Cohen-Or. Neural bestbuddies: sparse cross-domain correspondence. ACM Transactions on Graphics, 37(4): 69:1–69:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alhashim:2014:TVS

[ALX^{+}14] Ibraheem Alhashim, Honghua Li, Kai Xu, Junjie Cao, Rui Ma, and Hao Zhang. Topology-varying 3D shape creation via structural blending. ACM Transactions on Graphics, 33(4):158:1– 158:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aliaga:2008:VRS

[ALY08] Daniel G. Aliaga, Alvin J.

Law, and Yu Hong Yeung. A virtual restoration stage for real-world objects. ACM Transactions on Graphics, 27 (5):149:1–149:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Albahar:2021:PSD

[ALY^{+}21] Badour Albahar, Jingwan Lu, Jimei Yang, Zhixin Shu, Eli Shechtman, and Jia-Bin Huang. Pose with style: detail-preserving pose-guided image synthesis with conditional StyleGAN. ACM Transactions on Graphics, 40 (6):218:1–218:11, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480559.

Alexa:2010:RI

[AM10] Marc Alexa and Wojciech Matusik. Reliefs as images. ACM Transactions on Graphics, 29 (4):60:1–60:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adams:2019:LOH

[AMA^{+}19] Andrew Adams, Karima Ma, Luke Anderson, Riyadh Baghdadi, Tzu-Mao Li, Michaël Gharbi, Benoit Steiner, Steven Johnson, Kayvon Fatahalian, Frédo Durand, and Jonathan Ragan-Kelley. Learning to optimize halide with tree search and random programs. ACM Transactions on Graphics, 38 (4):121:1–121:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alderighi:2021:VDT

[AMB^{+}21] Thomas Alderighi, Luigi Malomo, Bernd Bickel, Paolo Cignoni, and Nico Pietroni. Volume decomposition for two-piece rigid casting. ACM Transactions on Graphics, 40 (6):272:1–272:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480555.

Aiger:2008:PCS

[AMCO08] Dror Aiger, Niloy J. Mitra, and Daniel Cohen-Or. 4-points congruent sets for robust pairwise surface registration. ACM Transactions on Graphics, 27(3):85:1– 85:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alliez:2002:IGR

[AMD02] Pierre Alliez, Mark Meyer, and Mathieu Desbrun. Interactive geometry remeshing. ACM Transactions on Graphics, 21(3):347–354, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alderighi:2018:MCD

[AMG^{+}18] Thomas Alderighi, Luigi Malomo, Daniela Giorgi, Nico Pietroni, Bernd Bickel, and Paolo Cignoni. Metamolds: computational design of silicone molds. ACM Transactions on Graphics, 37(4): 136:1–136:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alderighi:2019:VAD

[AMG^{+}19] Thomas Alderighi, Luigi Malomo, Daniela Giorgi, Bernd Bickel, Paolo Cignoni, and Nico Pietroni. Volume-aware design of composite molds. ACM Transactions on Graphics, 38(4):110:1–110:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Atzmon:2018:PCN

[AML18] Matan Atzmon, Haggai Maron, and Yaron Lipman. Point convolutional neural networks by extension operators. ACM Transactions on Graphics, 37 (4):71:1–71:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aydin:2008:DRI

[AMMS08] Tunç Ozan Aydin, Rafał Mantiuk, Karol Myszkowski, and Hans-Peter Seidel. Dynamic range independent image quality assessment. ACM Transactions on Graphics, 27 (3):69:1–69:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aila:2003:DSG

[AMN03] Timo Aila, Ville Miettinen, and Petri Nordlund. Delay streams for graphics hardware. ACM Transactions on Graphics, 22(3):792–800, July 2003. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Akenine-Moller:2003:GMH

[AMS03] Tomas Akenine-Möller and Jacob Ström. Graphics for the masses: a hardware rasterization architecture for mobile phones. ACM Transactions on Graphics, 22(3):801–808, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aguado:1999:MGC

[AMZ99] Alberto S. Aguado, Eugenia Montiel, and Ed Zaluska. Modeling generalized cylinders via Fourier morphing. ACM Transactions on Graphics, 18(4):293–315, October 1999. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
1999-18-4/p293-aguado/.

Alexanderson:2023:LDA

[ANBH23] Simon Alexanderson, Rajmund Nagy, Jonas Beskow, and Gustav Eje Henter. Listen, denoise, action! audiodriven motion synthesis with diffusion models. ACM Transactions on Graphics, 42 (4):44:1–44:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592458.

Anderson:1982:HLE

[And82] D. P. Anderson. Hidden line elimination in projected grid surfaces. ACM Transactions on Graphics, 1(4):274– 288, October 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anderson:1983:TRP

[And83]  D. P. Anderson.  Tech-

niques for reducing pen plotting time. ACM Transactions on Graphics, 2(3): 197–212, July 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Angelidis:2017:MSV

[Ang17] Alexis Angelidis. Multi-scale vorticle fluids. ACM Transactions on Graphics, 36(4): 104:1–104:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ahmed:2017:APS

[ANHD17] Abdalla G. M. Ahmed, Till Niese, Hui Huang, and Oliver Deussen. An adaptive point sampler on a regular lattice. ACM Transactions on Graphics, 36(4):138:1–138:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adkins:2023:HID

[ANL^{+}23] Alex Adkins, Aline Normoyle, Lorraine Lin, Yu Sun, Yuting Ye, Massimiliano Di Luca, and Sophie Jörg. How important are detailed hand motions for communication for a virtual character through the lens of charades? ACM Transactions on Graphics, 42 (3):27:1–27:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3578575.

Anonymous:1982:IA

[Ano82] Anonymous. Information for authors. ACM Transactions on Graphics, 1(1):82– 84, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1983:IA

[Ano83]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 2(3): 213–216, July 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1984:IA

[Ano84]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 3(3): 238–240, July 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1985:AI

[Ano85a] Anonymous. Author index.

ACM Transactions on Graphics, 4(1):71–72, January 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1985:CP

[Ano85b] Anonymous. Call for papers. ACM Transactions on Graphics, 4(3):244, July 1985. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1986:IA

[Ano86]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 5(3): 276–278, July 1986. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1987:IA

[Ano87]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 6(3): 240–242, July 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1988:IA

[Ano88]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 7(3): 222–224, July 1988. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1989:IA

[Ano89]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 8(3): 255–257, July 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1990:C

[Ano90a] Anonymous. Corrigendum.

ACM Transactions on Graphics, 9(2):244, April 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1990:FYC

[Ano90b] Anonymous. Five-year cumulative author index. ACM Transactions on Graphics, 9 (1):142–144, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1990:IA

[Ano90c]  Anonymous.  Information

for authors. ACM Transactions on Graphics, 9(3): 342–344, July 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1992:AI

[Ano92a] Anonymous. Author index.

ACM Transactions on Graphics, 11(1):100–101, January 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1992:CP

[Ano92b] Anonymous. Call for papers.

ACM Transactions on Graphics, 11(4):299, October 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1993:AI

[Ano93] Anonymous. Author index. ACM Transactions on Graphics, 12(1):109–110, January 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1994:AI

[Ano94] Anonymous. Author index. ACM Transactions on Graphics, 13(4):425–426, October 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1995:AI

[Ano95] Anonymous. 1995 author index. ACM Transactions on Graphics, 14(4):412–413, October 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:1996:AI

[Ano96] Anonymous. 1996 author index. ACM Transactions on Graphics, 15(4):377–378, October 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:2003:AWC

[Ano03]  Anonymous.  The actu-

ated workbench: computercontrolled actuation in tabletop tangible interfaces. ACM Transactions on Graphics, 22 (3):699, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Anonymous:2010:AAP

[Ano10]  Anonymous.  Acknowledg-

ment — AIM@SHAPE project attribution. ACM Transactions on Graphics, 29(2): 20:1, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Akbay:2018:EPM

[ANZS18] Muzaffer Akbay, Nicholas Nobles, Victor Zordan, and Tamar Shinar. An extended partitioned method for conservative solid–fluid coupling. ACM Transactions on Graphics, 37(4):86:1–86:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Azencot:2015:DDV

[AOCBC15] Omri Azencot, Maks Ovsjanikov, Frédéric Chazal, and Mirela Ben-Chen. Discrete derivatives of vector fields on surfaces — an operator approach. ACM Transactions on Graphics, 34(3): 29:1–29:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aurand:2022:ENS

[AONA22] Joshua Aurand, Raphael Ortiz, Silvia Nauer, and Vinicius C. Azevedo. Efficient neural style transfer for volumetric simulations. ACM Transactions on Graphics, 41(6): 257:1–257:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555517.

Aksoy:2018:SSS

[AOP^{+}18] Yagiz Aksoy, Tae-Hyun Oh, Sylvain Paris, Marc Pollefeys, and Wojciech Matusik. Semantic soft segmentation. ACM Transactions on Graphics, 37(4):72:1–72:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

An:2008:AAP

[AP08] Xiaobo An and Fabio Pellacini. AppProp: all-pairs appearance-space edit propagation. ACM Transactions on Graphics, 27(3):40:1– 40:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ahmed:2016:LDB

[APC^{+}16] Abdalla G. M. Ahmed, Hélène Perrier, David Coeurjolly, Victor Ostromoukhov, Jianwei Guo, Dong-Ming Yan, Hui Huang, and Oliver Deussen. Low-discrepancy blue noise sampling. ACM Transactions on Graphics, 35(6): 247:1–247:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alaluf:2021:OMS

[APCO21] Yuval Alaluf, Or Patashnik, and Daniel Cohen-Or. Only a matter of style: age transformation using a stylebased regression model. ACM Transactions on Graphics, 40 (4):45:1–45:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459805.

Agrawala:2003:DES

[APH^{+}03] Maneesh Agrawala, Doantam Phan, Julie Heiser, John Haymaker, Jeff Klingner, Pat Hanrahan, and Barbara Tversky. Designing effective stepby-step assembly instructions. ACM Transactions on Graphics, 22(3):828–837, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aubry:2014:FLL

[APH^{+}14] Mathieu Aubry, Sylvain Paris, Samuel W. Hasinoff, Jan Kautz, and Frédo Durand. Fast local Laplacian filters: Theory and applications. ACM Transactions on Graphics, 33(5):167:1– 167:??, August 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Adams:2007:ASP

[APKG07] Bart Adams, Mark Pauly, Richard Keiser, and Leonidas J. Guibas. Adaptively sampled particle fluids. ACM Transactions on Graphics, 26(3): 48:1–48:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2014:LBL

[APL14] Noam Aigerman, Roi Poranne, and Yaron Lipman. Lifted bijections for low distortion surface mappings. ACM Transactions on Graphics, 33(4):69:1–69:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aigerman:2015:SSM

[APL15] Noam Aigerman, Roi Poranne, and Yaron Lipman. Seamless surface mappings. ACM Transactions on Graphics, 34(4):72:1–72:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arev:2014:AEF

[APS^{+}14] Ido Arev, Hyun Soo Park, Yaser Sheikh, Jessica Hodgins, and Ariel Shamir. Automatic editing of footage from multiple social cameras. ACM Transactions on Graphics, 33 (4):81:1–81:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ahmed:2023:AOS

[APW23] Abdalla G. M. Ahmed, Matt Pharr, and Peter Wonka. ART–Owen scrambling. ACM Transactions on Graphics, 42 (6):258:1–258:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618307.

Allen:2015:AFI

[AR15] Andrew Allen and Nikunj Raghuvanshi. Aerophones in flatland: interactive wave simulation of wind instruments. ACM Transactions on Graphics, 34(4):134:1– 134:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Agarwal:2003:SIS

[ARBJ03] Sameer Agarwal, Ravi Ramamoorthi, Serge Belongie, and Henrik Wann Jensen. Structured importance sampling of environment maps. ACM Transactions on Graphics, 22(3):605–612, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arikan:2006:CMC

[Ari06] Okan Arikan. Compression of motion capture databases. ACM Transactions on Graphics, 25(3):890–897, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alaluf:2023:NST

[ARMCO23] Yuval Alaluf, Elad Richardson, Gal Metzer, and Daniel Cohen-Or. A neural spacetime representation for textto-image personalization. ACM Transactions on Graphics, 42 (6):243:1–243:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618322.

Agrawal:2005:RPA

[ARNL05] Amit Agrawal, Ramesh Raskar, Shree K. Nayar, and Yuanzhen Li. Removing photography artifacts using gradient projection and flash-exposure sampling. ACM Transactions on Graphics, 24(3):828–835, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aubry:2014:PMA

[ARS14] Mathieu Aubry, Bryan C. Russell, and Josef Sivic. Painting-to-3D model alignment via discriminative visual elements. ACM Transactions on Graphics, 33(2): 14:1–14:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ahmed:2022:GBN

[ARW22] Abdalla G. M. Ahmed, Jing Ren, and Peter Wonka. Gaussian blue noise. ACM Transactions on Graphics, 41(6): 260:1–260:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555519.

Ashikhmin:2002:SIT

[AS02] Michael Ashikhmin and Peter Shirley. Steerable illumination textures. ACM Transactions on Graphics, 21(1): 1–19, January 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Avidan:2007:SCC

[AS07] Shai Avidan and Ariel Shamir.

Seam carving for contentaware image resizing. ACM Transactions on Graphics, 26 (3):10:1–10:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arora:2021:MAD

[AS21] Rahul Arora and Karan Singh. Mid-air drawing of curves on 3D surfaces in virtual reality. ACM Transactions on Graphics, 40(3):33:1–33:17, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3459090.

Alcantara:2009:RTP

[ASA^{+}09] Dan A. Alcantara, Andrei Sharf, Fatemeh Abbasinejad, Shubhabrata Sengupta, Michael Mitzenmacher, John D. Owens, and Nina Amenta. Real-time parallel hashing on the GPU. ACM Transactions on Graphics, 28(5): 154:1–154:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ansari:2022:MIN

[ASB22] Navid Ansari, Hans-Peter Seidel, and Vahid Babaei. Mixed integer neural inverse design. ACM Transactions on Graphics, 41(4):151:1–151:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530083.

Aydin:2014:TCL

[ASC^{+}14] Tunç Ozan Aydin, Nikolce Stefanoski, Simone Croci, Markus Gross, and Aljoscha Smolic. Temporally coherent local tone mapping of HDR video. ACM Transactions on Graphics, 33(6): 196:1–196:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arikan:2013:SOB

[ASF^{+}13] Murat Arikan, Michael Schwärzler, Simon Flöry, Michael Wimmer, and Stefan Maierhofer. O-snap: Optimization-based snapping for modeling architecture. ACM Transactions on Graphics, 32(1):6:1– 6:15, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Avron:2010:SRS

[ASGCO10] Haim Avron, Andrei Sharf, Chen Greif, and Daniel Cohen-Or. ℓ_{1}-sparse reconstruction of sharp point set surfaces. ACM Transactions on Graphics, 29(5):135:1– 135:12, October 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Abulnaga:2023:SVM

[ASGS23] S. Mazdak Abulnaga, Oded Stein, Polina Golland, and Justin Solomon. Symmetric volume maps: Order-invariant volumetric mesh correspondence with free boundary. ACM Transactions on Graphics, 42(3):25:1–25:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3572897.

Ahmed:2023:ASD

[ASHW23] Abdalla G. M. Ahmed, Mikhail Skopenkov, Markus Hadwiger, and Peter Wonka. Analysis and synthesis of digital dyadic sequences. ACM Transactions on Graphics, 42 (6):218:1–218:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618308.

Anguelov:2005:SSC

[ASK^{+}05] Dragomir Anguelov, Praveen Srinivasan, Daphne Koller, Sebastian Thrun, Jim Rodgers, and James Davis. SCAPE: shape completion and animation of people. ACM Transactions on Graphics, 24(3): 408–416, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Akhter:2012:BSB

[ASK^{+}12] Ijaz Akhter, Tomas Simon, Sohaib Khan, Iain Matthews, and Yaser Sheikh. Bilinear spatiotemporal basis models. ACM Transactions on Graphics, 31(2):17:1–17:12, April 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ashtari:2022:RBS

[ASK^{+}22] Amirsaman Ashtari, Chang Wook Seo, Cholmin Kang, Sihun Cha, and Junyong Noh. Reference based sketch extraction via attention mechanism. ACM Transactions on Graphics, 41(6):207:1– 207:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555504.

Aroudj:2017:VCT

[ASL^{+}17] Samir Aroudj, Patrick Seemann, Fabian Langguth, Stefan Guthe, and Michael Goesele. Visibility-consistent thin surface reconstruction using multi-scale kernels. ACM Transactions on Graphics, 36 (6):187:1–187:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ashtari:2020:CSF

[ASN^{+}20] Amirsaman Ashtari, Stefan Stevsić, Tobias Nägeli, Jean-Charles Bazin, and Otmar Hilliges. Capturing subjective first-person view shots with drones for automated cinematography. ACM Transactions on Graphics, 39 (5):159:1–159:14, September 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3378673.

Asente:2007:DPM

[ASP07] Paul Asente, Mike Schuster, and Teri Pettit. Dynamic planar map illustration. ACM Transactions on Graphics, 26 (3):30:1–30:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Au:2008:SEM

[ATC^{+}08] Oscar Kin-Chung Au, Chiew-Lan Tai, Hung-Kuo Chu, Daniel Cohen-Or, and Tong-Yee Lee. Skeleton extraction by mesh contraction. ACM Transactions on Graphics, 27 (3):44:1–44:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

An:2011:ARM

[ATDP11] Xiaobo An, Xin Tong, Jonathan D. Denning, and Fabio Pellacini. AppWarp: retargeting measured materials by appearance-space warping. ACM Transactions on Graphics, 30(6):147:1– 147:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Arabadzhiyska:2017:SLP

[ATM^{+}17] Elena Arabadzhiyska, Okan Tarhan Tursun, Karol Myszkowski, Hans-Peter Seidel, and Piotr Didyk. Saccade landing position prediction for gazecontingent rendering. ACM Transactions on Graphics, 36 (4):50:1–50:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ando:2013:HAL

[ATW13] Ryoichi Ando, Nils Thürey, and Chris Wojtan. Highly adaptive liquid simulations on tetrahedral meshes. ACM Transactions on Graphics, 32 (4):103:1–103:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ando:2015:SFS

[ATW15] Ryoichi Ando, Nils Thuerey, and Chris Wojtan. A stream function solver for liquid simulations. ACM Transactions on Graphics, 34(4):53:1– 53:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Angles:2017:SBI

[ATW^{+}17] Baptiste Angles, Marco Tarini, Brian Wyvill, Loïc Barthe, and Andrea Tagliasacchi. Sketch-based implicit blending. ACM Transactions on Graphics, 36(6):181:1– 181:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aliaga:2008:IEB

[AVB08] Daniel G. Aliaga, Carlos A.

Vanegas, and Bedřich Beneš. Interactive example-based urban layout synthesis. ACM Transactions on Graphics, 27 (5):160:1–160:??, December 2008. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Araujo:2023:SPR

[AVB^{+}23] Chrystiano Araújo, Nicholas Vining, Silver Burla, Manuel Ruivo De Oliveira, Enrique Rosales, and Alla Sheffer. Slippagepreserving reshaping of humanmade 3D content. ACM Transactions on Graphics, 42 (6):272:1–272:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3618391.

Agrawal:2016:TBL

[AvdP16] Shailen Agrawal and Michiel van de Panne. Task-based locomotion. ACM Transactions on Graphics, 35(4): 82:1–82:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

AlBorno:2017:DAE

[AVF17] Mazen Al Borno, Michiel Van De Panne, and Eugene Fiume. Domain of attraction expansion for physics-based character control. ACM Transactions on Graphics, 36(4): 145:1–145:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ainsley:2012:SPA

[AVGT12] Samantha Ainsley, Etienne Vouga, Eitan Grinspun, and Rasmus Tamstorf. Speculative parallel asynchronous contact mechanics. ACM Transactions on Graphics, 31 (6):151:1–151:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Araujo:2022:LUP

[AVR^{+}22] Chrystiano Araújo, Nicholas Vining, Enrique Rosales, Giorgio Gori, and Alla Sheffer. As-locally-uniform-as-possible reshaping of vector clip-art. ACM Transactions on Graphics, 41 (4):160:1–160:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530098.

Alexa:2011:DLG

[AW11] Marc Alexa and Max Wardetzky. Discrete Laplacians on general polygonal meshes. ACM Transactions on Graphics, 30(4):102:1–102:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ahmed:2020:SSB

[AW20] Abdalla G. M. Ahmed and Peter Wonka. Screen-space bluenoise diffusion of Monte Carlo sampling error via hierarchical ordering of pixels. ACM Transactions on Graphics, 39 (6):244:1–244:15, November 2020. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417881.

Ahmed:2021:ODN

[AW21] Abdalla G. M. Ahmed and Peter Wonka. Optimizing dyadic nets. ACM Transactions on Graphics, 40(4): 141:1–141:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459880.

Akeley:2004:SDP

[AWGB04] Kurt Akeley, Simon J. Watt, Ahna Reza Girshick, and Martin S. Banks. A stereo display prototype with multiple focal distances. ACM Transactions on Graphics, 23(3): 804–813, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aittala:2013:PSC

[AWL13] Miika Aittala, Tim Weyrich, and Jaakko Lehtinen. Practical SVBRDF capture in the frequency domain. ACM Transactions on Graphics, 32 (4):110:1–110:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aittala:2015:TSS

[AWL15] Miika Aittala, Tim Weyrich, and Jaakko Lehtinen. Twoshot SVBRDF capture for stationary materials. ACM Transactions on Graphics, 34(4):110:1–110:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aberman:2019:LCA

[AWL^{+}19] Kfir Aberman, Rundi Wu, Dani Lischinski, Baoquan Chen, and Daniel Cohen-Or. Learning character-agnostic motion for motion retargeting in 2D. ACM Transactions on Graphics, 38(4): 75:1–75:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aberman:2020:UMS

[AWL^{+}20] Kfir Aberman, Yijia Weng, Dani Lischinski, Daniel Cohen-Or, and Baoquan Chen. Unpaired motion style transfer from video to animation. ACM Transactions on Graphics, 39(4):64:1– 64:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392469.

Agrawal:2009:IMB

[AXR09] Amit Agrawal, Yi Xu, and Ramesh Raskar. Invertible motion blur in video. ACM Transactions on Graphics, 28 (3):95:1–95:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Alhashim:2015:DDT

[AXZ^{+}15] Ibraheem Alhashim, Kai Xu, Yixin Zhuang, Junjie Cao, Patricio Simari, and Hao Zhang. Deformation-driven topology-varying 3D shape correspondence. ACM Transactions on Graphics, 34(6): 236:1–236:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aliaga:2012:FHR

[AYL^{+}12] Daniel G. Aliaga, Yu Hong Yeung, Alvin Law, Behzad Sajadi, and Aditi Majumder. Fast high-resolution appearance editing using superimposed projections. ACM Transactions on Graphics, 31 (2):13:1–13:13, April 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Aliaga:2009:FMS

[AZB09] Daniel G. Aliaga, Ji Zhang, and Mireille Boutin. A framework for modeling 3D scenes using pose-free equations. ACM Transactions on Graphics, 29(1):7:1–7:15, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ao:2023:GGD

[AZL23] Tenglong Ao, Zeyi Zhang, and Libin Liu. GestureDiffuCLIP:

Gesture diffusion model with CLIP latents. ACM Transactions on Graphics, 42(4): 42:1–42:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592097.

Abdal:2021:SAC

[AZMW21] Rameen Abdal, Peihao Zhu, Niloy J. Mitra, and Peter Wonka. StyleFlow: Attributeconditioned exploration of StyleGAN-generated images using conditional continuous normalizing flows. ACM Transactions on Graphics, 40(3):21:1–21:21, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3447648.

Agarwala:2005:PVT

[AZP^{+}05] Aseem Agarwala, Ke Colin Zheng, Chris Pal, Maneesh Agrawala, Michael Cohen, Brian Curless, David Salesin, and Richard Szeliski. Panoramic video textures. ACM Transactions on Graphics, 24(3):821–827, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Burt:1983:MSA

[BA83] P. J. Burt and E. H. Adelson. A multiresolution spline with application to image mosaics. ACM Transactions on Graphics, 2(4):217–236, October 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boubekeur:2008:PT

[BA08] Tamy Boubekeur and Marc Alexa. Phong Tessellation. ACM Transactions on Graphics, 27(5):141:1– 141:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bai:2012:SAV

[BAAR12] Jiamin Bai, Aseem Agarwala, Maneesh Agrawala, and Ravi Ramamoorthi. Selectively deanimating video. ACM Transactions on Graphics, 31(4): 66:1–66:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bertails:2006:SHP

[BAC^{+}06] Florence Bertails, Basile Audoly, Marie-Paule Cani, Bernard Querleux, Frédéric Leroy, and Jean-Luc Lévêque. Super-helices for predicting the dynamics of natural hair. ACM Transactions on Graphics, 25(3):1180–1187, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Burley:2018:DED

[BAC^{+}18] Brent Burley, David Adler, Matt Jen-Yuan Chiang, Hank Driskill, Ralf Habel, Patrick Kelly, Peter Kutz, Yining Karl Li, and Daniel Teece. The design and evolution of Disney’s Hyperion renderer. ACM Transactions on Graphics, 37(3):33:1– 33:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3182159.

Baykal:2023:CGS

[BAC^{+}23] Ahmet Canberk Baykal, Abdul Basit Anees, Duygu Ceylan, Erkut Erdem, Aykut Erdem, and Deniz Yuret. CLIP-guided StyleGAN inversion for text-driven real image editing. ACM Transactions on Graphics, 42(5): 172:1–172:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3610287.

Bae:2010:CR

[BAD10] Soonmin Bae, Aseem Agarwala, and Frédo Durand. Computational rephotography. ACM Transactions on Graphics, 29(3):24:1– 24:15, June 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baecker:1982:SPR

[Bae82] R. M. Baecker. Sizing and positioning rectangles. ACM Transactions on Graphics, 1 (2):184–185, April 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baecher:2018:SDF

[Bae18] Moritz Baecher. Session details: Fabulously computed fashion. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ben-Artzi:2008:PPR

[BAERD08] Aner Ben-Artzi, Kevin Egan, Ravi Ramamoorthi, and Frédo Durand. A precomputed polynomial representation for interactive BRDF editing with global illumination. ACM Transactions on Graphics, 27(2):13:1– 13:14, April 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bar:2019:MCF

[BAGL19] Chen Bar, Marina Alterman, Ioannis Gkioulekas, and Anat Levin. A Monte Carlo framework for rendering speckle statistics in scattering media. ACM Transactions on Graphics, 38(4):39:1–39:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baker:1994:CIA

[Bak94]  Henry G. Baker.  Corri-

genda: “Intersection Algorithms for Lines and Circles”. ACM Transactions on Graphics, 13(3):308–310, July 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

197874.html. See [MST89].

Barringer:2013:AAA

[BAM13] Rasmus Barringer and Tomas Akenine-Möller. A 4: asynchronous adaptive anti-aliasing using shared memory. ACM Transactions on Graphics, 32 (4):100:1–100:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barringer:2014:DRS

[BAM14] Rasmus Barringer and Tomas Akenine-Möller. Dynamic ray stream traversal. ACM Transactions on Graphics, 33(4): 151:1–151:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ben-Artzi:2006:RTB

[BAOR06] Aner Ben-Artzi, Ryan Overbeck, and Ravi Ramamoorthi. Real-time BRDF editing in complex lighting. ACM Transactions on Graphics, 25 (3):945–954, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barth:1986:OOA

[Bar86] Paul S. Barth. An objectoriented approach to graphical interfaces. ACM Transactions on Graphics, 5(2): 142–172, April 1986. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        22951.html.

Baerentzen:2014:ISM

[BAS14] J. Andreas Bærentzen, Rinat Abdrashitov, and Karan Singh. Interactive shape modeling using a skeletonmesh co-representation. ACM Transactions on Graphics, 33 (4):132:1–132:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brunton:2018:PSV

[BATU18] Alan Brunton, Can Ates Arikan, Tejas Madan Tanksale, and Philipp Urban. 3D printing spatially varying color and translucency. ACM Transactions on Graphics, 37(4): 157:1–157:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brunton:2015:PLC

[BAU15] Alan Brunton, Can Ates Arikan, and Philipp Urban. Pushing the limits of 3D color printing: Error diffusion with translucent materials. ACM Transactions on Graphics, 35 (1):4:1–4:??, December 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bergou:2010:DVT

[BAV^{+}10] Miklós Bergou, Basile Audoly, Etienne Vouga, Max Wardetzky, and Eitan Grinspun. Discrete viscous threads. ACM Transactions on Graphics, 29 (4):116:1–116:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barsky:1983:LCB

[BB83] Brian A. Barsky and John C.

Beatty. Local control of bias and tension in beta-splines. ACM Transactions on Graphics, 2(2):109–134, April 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boyd:2012:MET

[BB12] Landon Boyd and Robert Bridson. MultiFLIP for energetic two-phase fluid simulation. ACM Transactions on Graphics, 31(2):16:1– 16:12, April 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Beeler:2014:RSF

[BB14] Thabo Beeler and Derek Bradley. Rigid stabilization of facial expressions. ACM Transactions on Graphics, 33 (4):44:1–44:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bell:2015:LVS

[BB15] Sean Bell and Kavita Bala. Learning visual similarity for product design with convolutional neural networks. ACM Transactions on Graphics, 34 (4):98:1–98:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2017:PEM

[BB17] Laurent Belcour and Pascal Barla. A practical extension to microfacet theory for the modeling of varying iridescence. ACM Transactions on Graphics, 36(4): 65:1–65:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brodt:2022:SEC

[BB22] Kirill Brodt and Mikhail Bessmeltsev. Sketch2Pose: estimating a 3D character pose from a bitmap sketch. ACM Transactions on Graphics, 41(4):85:1–85:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530106.

Bickel:2007:MSC

[BBA^{+}07] Bernd Bickel, Mario Botsch, Roland Angst, Wojciech Matusik, Miguel Otaduy, Hanspeter Pfister, and Markus Gross. Multi-scale capture of facial geometry and motion. ACM Transactions on Graphics, 26(3):33:1–33:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bartels:1993:ECS

[BBB^{+}93] Richard H. Bartels, John C.

Beatty, Kellogg S. Booth, Eric G. Bosch, and Pierre Jolicoeur. Experimental comparison of splines using the shapematching paradigm. ACM Transactions on Graphics, 12(3):179–208, July 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
169709.html.

Batty:2007:FVF

[BBB07] Christopher Batty, Florence Bertails, and Robert Bridson. A fast variational framework for accurate solidfluid coupling. ACM Transactions on Graphics, 26(3): 100:1–100:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Beeler:2010:HQS

[BBB^{+}10a] Thabo Beeler, Bernd Bickel, Paul Beardsley, Bob Sumner, and Markus Gross. Highquality single-shot capture of facial geometry. ACM Transactions on Graphics, 29(4): 40:1–40:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brochu:2010:MFS

[BBB10b] Tyson Brochu, Christopher Batty, and Robert Bridson. Matching fluid simulation elements to surface geometry and topology. ACM Transactions on Graphics, 29(4): 47:1–47:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bermano:2014:FPE

[BBB^{+}14] Amit H. Bermano, Derek Bradley, Thabo Beeler, Fabio Zund, Derek Nowrouzezahrai, Ilya Baran, Olga Sorkine-Hornung, Hanspeter Pfister, Robert W. Sumner, Bernd Bickel, and Markus Gross. Facial performance enhancement using dynamic shape space analysis. ACM Transactions on Graphics, 33(2): 13:1–13:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bruckler:2022:VPQ

[BBC22] Hendrik Brückler, David Bommes, and Marcel Campen. Volume parametrization quantization for hexahedral meshing. ACM Transactions on Graphics, 41(4):60:1– 60:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530123.

Bati:2023:CCC

[BBC^{+}23] Mégane Bati, Stéphane Blanco, Christophe Coustet, Vincent Eymet, Vincent Forest, Richard Fournier, Jacques Gautrais, Nicolas Mellado, Mathias Paulin, and Benjamin Piaud. Coupling conduction, convection and radiative transfer in a single path-space: Application to infrared rendering. ACM Transactions on Graphics, 42 (4):79:1–79:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592121.

Boye:2012:VSF

[BBG12] Simon Boyé, Pascal Barla, and Gaël Guennebaud. A vectorial solver for free-form vector gradients. ACM Transactions on Graphics, 31(6): 173:1–173:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bermano:2013:APA

[BBG^{+}13] Amit Bermano, Philipp Brüschweiler, Anselm Grundhöfer, Daisuke Iwai, Bernd Bickel, and Markus Gross. Augmenting physical avatars using projector-based illumination. ACM Transactions on Graphics, 32(6):189:1– 189:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brunel:2021:TID

[BBG21] Camille Brunel, Pierre Bénard, and Gaël Guennebaud. A time-independent deformer for elastic contacts. ACM Transactions on Graphics, 40(4):159:1–159:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3450626.3459879.

Brokman:2024:STV

[BBG24] Jonathan Brokman, Martin Burger, and Guy Gilboa. Spectral total-variation processing of shapes-theory and applications. ACM Transactions on Graphics, 43(2): 22:1–22:??, April 2024. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3641845.

Berard:2016:LEC

[BBGB16] Pascal Bérard, Derek Bradley, Markus Gross, and Thabo Beeler. Lightweight eye capture using a parametric model. ACM Transactions on Graphics, 35(4): 117:1–117:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bronstein:2011:SGG

[BBGO11] Alexander M. Bronstein, Michael M. Bronstein, Leonidas J. Guibas, and Maks Ovsjanikov. Shape google: Geometric words and expressions for invariant shape retrieval. ACM Transactions on Graphics, 30(1):1:1–1:20, January 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bacher:2012:FAC

[BBJP12] Moritz Bächer, Bernd Bickel, Doug L. James, and Hanspeter Pfister. Fabricating articulated characters from skinned meshes. ACM Transactions on Graphics, 31(4): 47:1–47:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bermano:2015:DST

[BBK^{+}15] Amit Bermano, Thabo Beeler, Yeara Kozlov, Derek Bradley, Bernd Bickel, and Markus Gross. Detailed spatiotemporal reconstruction of eyelids. ACM Transactions on Graphics, 34(4):44:1– 44:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Beeler:2012:CRS

[BBN^{+}12] Thabo Beeler, Bernd Bickel, Gioacchino Noris, Paul Beardsley, Steve Marschner, Robert W. Sumner, and Markus Gross. Coupled 3D reconstruction of sparse facial hair and skin. ACM Transactions on Graphics, 31(4):117:1–117:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Berard:2014:HQC

[BBN^{+}14] Pascal Bérard, Derek Bradley, Maurizio Nitti, Thabo Beeler, and Markus Gross. Highquality capture of eyes. ACM Transactions on Graphics, 33 (6):223:1–223:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Becker:1991:IMT

[BBO91] Shawn C. Becker, William A. Barrett, and Dan R. Olsen, Jr. Interactive measurement of three-dimensional objects using a depth buffer and linear probe. ACM Transactions on Graphics, 10(2):201–207, April 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        108446.html.

Bickel:2009:CMN

[BBO^{+}09] Bernd Bickel, Moritz Bächer, Miguel A. Otaduy, Wojciech Matusik, Hanspeter Pfister, and Markus Gross. Capture and modeling of non-linear heterogeneous soft tissue. ACM Transactions on Graphics, 28(3):89:1–89:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bickel:2010:DFM

[BBO^{+}10] Bernd Bickel, Moritz Bächer, Miguel A. Otaduy, Hyunho Richard Lee, Hanspeter Pfister, Markus Gross, and Wojciech Matusik. Design and fabrication of materials with desired deformation behavior. ACM Transactions on Graphics, 29(4): 63:1–63:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bati:2021:IME

[BBP21] Mégane Bati, Pascal Barla, and Romain Pacanowski. An inverse method for the exploration of layered material appearance. ACM Transactions on Graphics, 40(4): 176:1–176:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459857.

Boyadzhiev:2015:BSD

[BBPA15] Ivaylo Boyadzhiev, Kavita Bala, Sylvain Paris, and Edward Adelson. Band-sifting decomposition for imagebased material editing. ACM Transactions on Graphics, 34(5):163:1–163:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boyadzhiev:2012:UGW

[BBPD12] Ivaylo Boyadzhiev, Kavita Bala, Sylvain Paris, and Frédo Durand. User-guided white balance for mixed lighting conditions. ACM Transactions on Graphics, 31(6): 200:1–200:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ballan:2010:UVB

[BBPP10] Luca Ballan, Gabriel J. Brostow, Jens Puwein, and Marc Pollefeys. Unstructured videobased rendering: interactive exploration of casually captured videos. ACM Transactions on Graphics, 29(4): 87:1–87:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barton:2021:GTM

[BBR^{+}21] Michael Barton, Michal Bizzarri, Florian Rist, Oleksii Sliusarenko, and Helmut Pottmann. Geometry and tool motion planning for curvature adapted CNC machining. ACM Transactions on Graphics, 40(4): 180:1–180:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459837.

Bassett:2013:AAP

[BBS^{+}13] Katie Bassett, Ilya Baran, Johannes Schmid, Markus Gross, and Robert W. Sumner. Authoring and animating painterly characters. ACM Transactions on Graphics, 32 (5):156:1–156:12, September 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2014:LFA

[BBS14a] Laurent Belcour, Kavita Bala, and Cyril Soler. A local frequency analysis of light scattering and absorption. ACM Transactions on Graphics, 33(5):163:1–163:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bell:2014:IIW

[BBS14b] Sean Bell, Kavita Bala, and Noah Snavely. Intrinsic images in the wild. ACM Transactions on Graphics, 33(4): 159:1–159:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barrett:2002:OBI

[BC02] William A. Barrett and Alan S. Cheney. Object-based image editing. ACM Transactions on Graphics, 21(3):

777–784, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bargteil:2014:ADB

[BC14] Adam W. Bargteil and Elaine Cohen. Animation of deformable bodies with quadratic Bézier finite elements. ACM Transactions on Graphics, 33(3): 27:1–27:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ben-Chen:2018:SDN

[BC18] Mirela Ben-Chen. Session details: Nets, cages and meshes. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bonneel:2019:SSP

[BC19] Nicolas Bonneel and David Coeurjolly. SPOT: sliced partial optimal transport. ACM Transactions on Graphics, 38 (4):89:1–89:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bruckler:2023:CEC

[BC23] Hendrik Brückler and Marcel Campen. Collapsing embedded cell complexes for safer hexahedral meshing. ACM Transactions on Graphics, 42 (6):180:1–180:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618384.

Bern:2017:IDA

[BCC17] James M. Bern, Kai-Hung Chang, and Stelian Coros. Interactive design of animated plushies. ACM Transactions on Graphics, 36(4): 80:1–80:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bommes:2013:IGM

[BCE^{+}13] David Bommes, Marcel Campen, Hans-Christian Ebke, Pierre Alliez, and Leif Kobbelt. Integer-grid maps for reliable quad meshing. ACM Transactions on Graphics, 32(4): 98:1–98:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ben-Chen:2005:OSC

[BCG05] Mirela Ben-Chen and Craig Gotsman. On the optimality of spectral compression of mesh data. ACM Transactions on Graphics, 24(1): 60–80, January 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Benard:2013:SAE

[BCK^{+}13] Pierre Bénard, Forrester Cole, Michael Kass, Igor Mordatch, James Hegarty, Martin Sebastian Senn, Kurt Fleischer, Davide Pesare, and Katherine Breeden. Stylizing animation by example. ACM Transactions on Graphics, 32(4): 119:1–119:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bernardin:2023:CBS

[BCK^{+}23] Antonin Bernardin, Eulalie Coevoet, Paul Kry, Sheldon Andrews, Christian Duriez, and Maud Marchal. Constraint-based simulation of passive suction cups. ACM Transactions on Graphics, 42 (1):13:1–13:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3551889.

Bando:2008:EDM

[BCN08] Yosuke Bando, Bing-Yu Chen, and Tomoyuki Nishita. Extracting depth and matte using a color-filtered aperture. ACM Transactions on Graphics, 27(5):134:1– 134:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baran:2010:HVS

[BCRK^{+}10] Ilya Baran, Jiawen Chen, Jonathan Ragan-Kelley, Frédo Durand, and Jaakko Lehtinen. A hierarchical volumetric shadow algorithm for single scattering. ACM Transactions on Graphics, 29(6): 178:1–178:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bacher:2015:LIL

[BCT15] Moritz Bächer, Stelian Coros, and Bernhard Thomaszewski. LinkEdit: interactive linkage editing using symbolic kinematics. ACM Transactions on Graphics, 34(4):99:1– 99:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bessmeltsev:2015:MCC

[BCV^{+}15] Mikhail Bessmeltsev, Will Chang, Nicholas Vining, Alla Sheffer, and Karan Singh. Modeling character canvases from cartoon drawings. ACM Transactions on Graphics, 34(5):162:1–162:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bright:2017:HGP

[BCW17] Alon Bright, Edward Chien, and Ofir Weber. Harmonic global parametrization with rational holonomy. ACM Transactions on Graphics, 36 (4):89:1–89:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ben-Chen:2009:VHM

[BCWG09] Mirela Ben-Chen, Ofir Weber, and Craig Gotsman. Variational harmonic maps for space deformation. ACM Transactions on Graphics, 28 (3):34:1–34:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bajaj:1995:MCP

[BCX95] Chanderjit L. Bajaj, Jindon Chen, and Guoliang Xu. Modeling with cubic A-patches. ACM Transactions on Graphics, 14(2):103–133, April 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        221662.html.

Borning:1986:CBT

[BD86] Alan Borning and Robert Duisberg. Constraint-based tools for building user interfaces. ACM Transactions on Graphics, 5(4):345–374, October 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        29354.html.

Benson:2002:OT

[BD02a] David Benson and Joel Davis. Octree textures. ACM Transactions on Graphics, 21(3): 785–790, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brooks:2002:SSB

[BD02b] Stephen Brooks and Neil Dodgson. Self-similarity based texture editing. ACM Transactions on Graphics, 21(3): 653–656, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bertails-Descoubes:2011:NNS

[BDCDA11] Florence Bertails-Descoubes, Florent Cadoux, Gilles Daviet, and Vincent Acary. A nonsmooth Newton solver for capturing exact Coulomb friction in fiber assemblies. ACM Transactions on Graphics, 30 (1):6:1–6:14, January 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barki:2011:CVB

[BDD11] Hichem Barki, Florence Denis, and Florent Dupont. Contributing vertices-based Minkowski sum of a nonconvex– convex pair of polyhedra. ACM Transactions on Graphics, 30(1):3:1–3:16, January 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Balakrishnan:2015:VDH

[BDG15] Guha Balakrishnan, Frédo Durand, and John Guttag. Video diff: highlighting differences between similar actions in videos. ACM Transactions on Graphics, 34(6): 194:1–194:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Blumberg:2002:ILI

[BDI^{+}02] Bruce Blumberg, Marc Downie, Yuri Ivanov, Matt Berlin, Michael Patrick Johnson, and Bill Tomlinson. Integrated learning for interactive synthetic characters. ACM Transactions on Graphics, 21 (3):417–426, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boechat:2016:RSP

[BDK^{+}16] Pedro Boechat, Mark Dokter, Michael Kenzel, Hans-Peter Seidel, Dieter Schmalstieg, and Markus Steinberger. Representing and scheduling procedural generation using operator graphs. ACM Transactions on Graphics, 35(6): 183:1–183:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bergner:2009:TCI

[BDM09] Steven Bergner, Mark S. Drew, and Torsten Möller. A tool to create illuminant and reflectance spectra for light-driven graphics and visualization. ACM Transactions on Graphics, 28(1):5:1– 5:11, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bhunia:2020:PCS

[BDM^{+}20] Ayan Kumar Bhunia, Ayan Das, Umar Riaz Muhammad, Yongxin Yang, Timothy M. Hospedales, Tao Xiang, Yulia Gryaditskaya, and Yi-Zhe Song. Pixelor: a competitive sketching AI agent. so you think you can sketch? ACM Transactions on Graphics, 39(6): 166:1–166:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417840.

Briedis:2021:NFI

[BDM^{+}21] Karlis Martins Briedis, Abdelaziz Djelouah, Mark Meyer, Ian McGonigal, Markus Gross, and Christopher Schroers. Neural frame interpolation for rendered content. ACM Transactions on Graphics, 40 (6):239:1–239:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480553.

Barill:2018:FWN

[BDS^{+}18] Gavin Barill, Neil G. Dickson, Ryan Schmidt, David I. W. Levin, and Alec Jacobson. Fast winding numbers for soups and clouds. ACM Transactions on Graphics, 37 (4):43:1–43:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barbic:2009:DOA

[BdSP09] Jernej Barbič, Marco da Silva, and Jovan Popović. Deformable object animation using reduced optimal control. ACM Transactions on Graphics, 28(3):53:1–53:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bala:1999:RIA

[BDT99] Kavita Bala, Julie Dorsey, and Seth Teller. Radiance interpolants for accelerated bounded-error ray tracing. ACM Transactions on Graphics, 18(3):213–256, July 1999. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tog/
                                        1999-18-3/p213-bala/.

Bonneel:2008:FMS

[BDT^{+}08] Nicolas Bonneel, George Drettakis, Nicolas Tsingos, Isabelle Viaud-Delmon, and Doug James. Fast modal sounds with scalable frequency-domain synthesis. ACM Transactions on Graphics, 27(3):24:1–24:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Busaryev:2013:AFS

[BDW13] Oleksiy Busaryev, Tamal K.

Dey, and Huamin Wang. Adaptive fracture simulation of multi-layered thin plates. ACM Transactions on Graphics, 32(4):52:1–52:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Busaryev:2012:ABI

[BDWR12] Oleksiy Busaryev, Tamal K.

Dey, Huamin Wang, and Zhong Ren. Animating bubble interactions in a liquid foam. ACM Transactions on Graphics, 31(4): 63:1–63:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Beatty:1988:VAT

[Bea88] John C. Beatty. A video adjunct to Transactions on Graphics. ACM Transactions on Graphics, 7(4):229– 230, October 1988. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Beatty:1991:ENE

[Bea91] John Beatty. Editorial: New Editor-in-Chief. ACM Transactions on Graphics, 10(3): 209–210, July 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brochu:2012:EGE

[BEB12] Tyson Brochu, Essex Edwards, and Robert Bridson. Efficient geometrically exact continuous collision detection. ACM Transactions on Graphics, 31(4):96:1–96:7, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brandt:2018:HRP

[BEH18] Christopher Brandt, Elmar Eisemann, and Klaus Hildebrandt. Hyper-reduced projective dynamics. ACM Transactions on Graphics, 37 (4):80:1–80:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2018:ERL

[Bel18] Laurent Belcour. Efficient rendering of layered materials using an atomic decomposition with statistical operators. ACM Transactions on Graphics, 37(4):73:1–73:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bergeron:1982:EIa

[Ber82a] R. D. Bergeron. Editor’s introduction. ACM Transactions on Graphics, 1(1): 1–4, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bergeron:1982:EIb

[Ber82b] R. Daniel Bergeron. Editor’s introduction. ACM Transactions on Graphics, 1 (3):189, July 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Buss:2001:SAA

[BF01] Samuel R. Buss and Jay P.

Fillmore. Spherical averages and applications to spherical splines and interpolation. ACM Transactions on Graphics, 20(2):95–126, April 2001. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Burns:2008:ACC

[BF08] Michael Burns and Adam Finkelstein. Adaptive cutaways for comprehensible rendering of polygonal scenes. ACM Transactions on Graphics, 27(5):154:1–154:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Berthouzoz:2012:REV

[BF12] Floraine Berthouzoz and Raanan Fattal. Resolution enhancement by vibrating displays. ACM Transactions on Graphics, 31(2): 15:1–15:14, April 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bridson:2002:RTC

[BFA02] Robert Bridson, Ronald Fedkiw, and John Anderson. Robust treatment of collisions, contact and friction for cloth animation. ACM Transactions on Graphics, 21(3): 594–603, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bolz:2003:SMS

[BFGS03] Jeff Bolz, Ian Farmer, Eitan Grinspun, and Peter Schröder. Sparse matrix solvers on the GPU: conjugate gradients and multigrid. ACM Transactions on Graphics, 22 (3):917–924, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Buck:2004:BGS

[BFH^{+}04] Ian Buck, Tim Foley, Daniel Horn, Jeremy Sugerman, Kayvon Fatahalian, Mike Houston, and Pat Hanrahan. Brook for GPUs: stream computing on graphics hardware. ACM Transactions on Graphics, 23(3):777–786, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brainerd:2016:EGR

[BFK^{+}16] Wade Brainerd, Tim Foley, Manuel Kraemer, Henry Moreton, and Matthias Nießner. Efficient GPU rendering of subdivision surfaces using adaptive quadtrees. ACM Transactions on Graphics, 35 (4):113:1–113:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Badler:1984:WC

[BG84] Norman I. Badler and Tamar E.

Granor. The window controller. ACM Transactions on Graphics, 3(4):312–315, October 1984. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bartels:1989:GEIb

[BG89a] Richard H. Bartels and Ronald N. Goldman. Guest Editors’ introduction. ACM Transactions on Graphics, 8 (4):261, October 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bartels:1989:GEIa

[BG89b] Richard H. Bartels and Ronald N. Goldman. Guest Editors’ introduction: Special issue on computer-aided geometric design. ACM Transactions on Graphics, 8(3): 145–146, July 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bartels:1990:GEI

[BG90] Richard H. Bartels and Ronald N. Goldman. Guest Editors’ introduction. ACM Transactions on Graphics, 9 (2):145–146, April 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barringer:2012:HQC

[BGAM12] Rasmus Barringer, Carl Johan Gribel, and Tomas Akenine-Möller. High-quality curve rendering using line sampled visibility. ACM Transactions on Graphics, 31(6): 162:1–162:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Borgeat:2005:GID

[BGB^{+}05] Louis Borgeat, Guy Godin, François Blais, Philippe Massicotte, and Christian Lahanier. GoLD: interactive display of huge colored and textured models. ACM Transactions on Graphics, 24(3): 869–877, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belhe:2023:DAN

[BGF^{+}23] Yash Belhe, Michaël Gharbi, Matthew Fisher, Iliyan Georgiev, Ravi Ramamoorthi, and Tzu-Mao Li. Discontinuity-aware 2D neural fields. ACM Transactions on Graphics, 42(6): 217:1–217:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618379.

Barreiro:2017:CCE

[BGFAO17] Héctor Barreiro, Ignacio García-Fernández, Iván Alduán, and Miguel A. Otaduy. Conformation constraints for efficient viscoelastic fluid simulation. ACM Transactions on Graphics, 36(6):221:1– 221:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Band:2018:PBI

[BGI^{+}18] Stefan Band, Christoph Gissler, Markus Ihmsen, Jens Cornelis, Andreas Peer, and Matthias Teschner. Pressure boundaries for implicit incompressible SPH. ACM Transactions on Graphics, 37(2): 14:1–14:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Berthouzoz:2013:PSP

[BGK^{+}13] Floraine Berthouzoz, Akash Garg, Danny M. Kaufman, Eitan Grinspun, and Maneesh Agrawala. Parsing sewing patterns into 3D garments. ACM Transactions on Graphics, 32 (4):85:1–85:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2016:BSI

[BGK16] Seung-Hwan Baek, Diego Gutierrez, and Min H. Kim. Birefractive stereo imaging for single-shot depth acquisition. ACM Transactions on Graphics, 35(6):194:1– 194:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Badki:2017:CZF

[BGKS17] Abhishek Badki, Orazio Gallo, Jan Kautz, and Pradeep Sen. Computational zoom: a framework for postcapture image composition. ACM Transactions on Graphics, 36(4):46:1–46:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bar:2020:RNF

[BGL20] Chen Bar, Ioannis Gkioulekas, and Anat Levin. Rendering near-field speckle statistics in scattering media. ACM Transactions on Graphics, 39 (6):187:1–187:18, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417813.

Bargteil:2006:SLC

[BGOS06] Adam W. Bargteil, Tolga G. Goktekin, James F. O’Brien, and John A. Strain. A semi-Lagrangian contouring method for fluid simulation. ACM Transactions on Graphics, 25(1):19–38, January 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barnes:2010:VTC

[BGSF10] Connelly Barnes, Dan B.

Goldman, Eli Shechtman, and Adam Finkelstein. Video tapestries with continuous temporal zoom. ACM Transactions on Graphics, 29(4): 89:1–89:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2021:PST

[BH21] Seung-Hwan Baek and Felix Heide. Polarimetric spatio-temporal light transport probing. ACM Transactions on Graphics, 40(6): 212:1–212:18, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480517.

Beeler:2011:HQP

[BHB^{+}11] Thabo Beeler, Fabian Hahn, Derek Bradley, Bernd Bickel, Paul Beardsley, Craig Gotsman, Robert W. Sumner, and Markus Gross. Highquality passive facial performance capture using anchor frames. ACM Transactions on Graphics, 30(4): 75:1–75:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Back:2020:DCI

[BHHM20] Jonghee Back, Binh-Son Hua, Toshiya Hachisuka, and Bochang Moon. Deep combiner for independent and correlated pixel estimates. ACM Transactions on Graphics, 39 (6):242:1–242:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417847.

Benard:2014:CSS

[BHK14] Pierre Bénard, Aaron Hertzmann, and Michael Kass. Computing smooth surface contours with accurate topology. ACM Transactions on Graphics, 33(2):19:1– 19:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bojsen-Hansen:2012:TSE

[BHLW12] Morten Bojsen-Hansen, Hao Li, and Chris Wojtan. Tracking surfaces with evolving topology. ACM Transactions on Graphics, 31(4):53:1– 53:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ben-Hamu:2018:MCG

[BHMK^{+}18] Heli Ben-Hamu, Haggai Maron, Itay Kezurer, Gal Avineri, and Yaron Lipman. Multichart generative surface modeling. ACM Transactions on Graphics, 37(6):215:1– 215:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bajaj:1998:RPN

[BHN98] Chandrajit L. Bajaj, Robert L.

Holt, and Arun N. Netravali. Rational parametrizations of nonsingular real cubic surfaces. ACM Transactions on Graphics, 17(1): 1–31, January 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1998-17-1/
p1-bajaj/.

Bridson:2007:CNP

[BHN07] Robert Bridson, Jim Houriham, and Marcus Nordenstam. Curl-noise for procedural fluid flow. ACM Transactions on Graphics, 26(3): 46:1–46:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bradley:2010:HRP

[BHPS10] Derek Bradley, Wolfgang Heidrich, Tiberiu Popa, and Alla Sheffer. High resolution passive facial performance capture. ACM Transactions on Graphics, 29(4): 41:1–41:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bando:2013:NIB

[BHR13] Yosuke Bando, Henry Holtzman, and Ramesh Raskar. Near-invariant blur for depth and 2D motion via timevarying light field analysis. ACM Transactions on Graphics, 32(2):13:1–13:15, April 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bunge:2022:VQS

[BHSH^{+}22] Astrid Bunge, Philipp Herholz, Olga Sorkine-Hornung, Mario Botsch, and Michael Kazhdan. Variational quadratic shape functions for polygons and polyhedra. ACM Transactions on Graphics, 41(4):54:1–54:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530137.

Bojsen-Hansen:2013:LST

[BHW13] Morten Bojsen-Hansen and Chris Wojtan. Liquid surface tracking with error compensation. ACM Transactions on Graphics, 32(4): 68:1–68:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bojsen-Hansen:2016:GNR

[BHW16] Morten Bojsen-Hansen and Chris Wojtan. Generalized non-reflecting boundaries for fluid re-simulation. ACM Transactions on Graphics, 35 (4):96:1–96:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bi:2015:ITE

[BHY15] Sai Bi, Xiaoguang Han, and Yizhou Yu. An L_{1} image transform for edge-preserving smoothing and scene-level intrinsic decomposition. ACM Transactions on Graphics, 34 (4):78:1–78:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bajaj:1992:ASD

[BI92] Chanderjit L. Bajaj and Insung Ihm. Algebraic surface design with Hermite interpolation. ACM Transactions on Graphics, 11(1): 61–91, January 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
120081.html.

Bimber:2008:SDR

[BI08] Oliver Bimber and Daisuke Iwai. Superimposing dynamic range. ACM Transactions on Graphics, 27(5): 150:1–150:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bajaj:2001:RIC

[BIP01] Chandrajit Bajaj, Insung Ihm, and Sanghun Park. 3D RGB image compression for interactive applications. ACM Transactions on Graphics, 20 (1):10–38, January 2001. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/2001-20-1/
                                        p10-bajaj/.

Bajaj:1993:HOI

[BIW93] Chanderjit Bajaj, Ihm Insung, and Joe Warren. Higherorder interpolation and leastsquares approximation using implicit algebraic surfaces. ACM Transactions on Graphics, 12(4):327–347, October 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        159734.html.

Barbic:2005:RTS

[BJ05] Jernej Barbič and Doug James. Real-time subspace integration for St. Venant– Kirchhoff deformable models. ACM Transactions on Graphics, 24(3):982–990, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2010:ASV

[BJ10a] Jongmin Baek and David E. Jacobs. Accelerating spatially varying Gaussian filters. ACM Transactions on Graphics, 29 (6):169:1–169:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barbic:2010:SSC

[BJ10b] Jernej Barbič and Doug L.

James. Subspace self-collision culling. ACM Transactions on Graphics, 29(4): 81:1–81:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bitterli:2017:BPB

[BJ17] Benedikt Bitterli and Wojciech Jarosz. Beyond points and beams: higherdimensional photon samples for volumetric light transport. ACM Transactions on Graphics, 36(4):112:1–112:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Borosan:2012:RAR

[BJD^{+}12] Péter Borosán, Ming Jin, Doug DeCarlo, Yotam Gingold, and Andrew Nealen. RigMesh: automatic rigging for part-based shape modeling and deformation. ACM Transactions on Graphics, 31 (6):198:1–198:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bitterli:2018:RJM

[BJNJ18] Benedikt Bitterli, Wenzel Jakob, Jan Novák, and Wojciech Jarosz. Reversible jump Metropolis light transport using inverse mappings. ACM Transactions on Graphics, 37 (1):1:1–1:??, January 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barnes:2008:VPP

[BJS^{+}08] Connelly Barnes, David E. Jacobs, Jason Sanders, Dan B. Goldman, Szymon Rusinkiewicz, Adam Finkelstein, and Maneesh Agrawala. Video puppetry: a performative interface for cutout animation. ACM Transactions on Graphics, 27(5):124:1– 124:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2018:SAP

[BJTK18] Seung-Hwan Baek, Daniel S. Jeon, Xin Tong, and Min H. Kim. Simultaneous acquisition of polarimetric SVBRDF and normals. ACM Transactions on Graphics, 37(6): 268:1–268:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bronsvoort:1985:RTG

[BK85] Willem F. Bronsvoort and Fopke Klok. Ray tracing generalized cylinders.

ACM Transactions on Graphics, 4(4):291–303, October 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

6118.html. See corrigendum [BK87].

Bronsvoort:1987:CRT

[BK87] Willem F. Bronsvoort and Fopke Klok. Corrigendum: “Ray Tracing Generalized Cylinders”. ACM Transactions on Graphics, 6(3): 238–239, July 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
214383.html. See [BK85].

Botsch:2004:IFR

[BK04] Mario Botsch and Leif Kobbelt.

An intuitive framework for real-time freeform modeling. ACM Transactions on Graphics, 23(3):630–634, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bernstein:2016:WNP

[BK16] Gilbert Louis Bernstein and Fredrik Kjolstad. Why new programming languages for simulation? ACM Transactions on Graphics, 35(2): 20:1–20:??, May 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bellini:2016:TVW

[BKCO16] Rachele Bellini, Yanir Kleiman, and Daniel Cohen-Or. Timevarying weathering in texture space. ACM Transactions on Graphics, 35(4): 141:1–141:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bitouk:2008:FSA

[BKD^{+}08] Dmitri Bitouk, Neeraj Kumar, Samreen Dhillon, Peter Belhumeur, and Shree K. Nayar. Face swapping: automatically replacing faces in photographs. ACM Transactions on Graphics, 27(3): 39:1–39:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2017:CSS

[BKGK17] Seung-Hwan Baek, Incheol Kim, Kaist Diego Gutierrez, and Min H. Kim. Compact single-shot hyperspectral imaging using a prism. ACM Transactions on Graphics, 36 (6):217:1–217:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Batra:2015:AVG

[BKKL15] Vineet Batra, Mark J. Kilgard, Harish Kumar, and Tristan Lorach. Accelerating vector graphics rendering using the graphics hardware pipeline. ACM Transactions on Graphics, 34(4): 146:1–146:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bai:2016:ADD

[BKLP16] Yunfei Bai, Danny M. Kaufman, C. Karen Liu, and Jovan Popović. Artist-directed dynamics for 2D animation. ACM Transactions on Graphics, 35(4):145:1–145:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Burns:2005:LDV

[BKR^{+}05] Michael Burns, Janek Klawe, Szymon Rusinkiewicz, Adam Finkelstein, and Doug De-Carlo. Line drawings from volume data. ACM Transactions on Graphics, 24(3): 512–518, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bi:2017:PBO

[BKR17] Sai Bi, Nima Khademi Kalantari, and Ravi Ramamoorthi. Patch-based optimization for image-based texture mapping. ACM Transactions on Graphics, 36(4):106:1–106:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bickel:2012:PFC

[BKS^{+}12] Bernd Bickel, Peter Kaufmann, Mélina Skouras, Bernhard Thomaszewski, Derek Bradley, Thabo Beeler, Phil Jackson, Steve Marschner, Wojciech Matusik, and Markus Gross. Physical face cloning. ACM Transactions on Graphics, 31(4):118:1–118:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bernstein:2015:LUT

[BL15] Gilbert Louis Bernstein and Wilmot Li. Lillicon: using transient widgets to create scale variations of icons. ACM Transactions on Graphics, 34(4):144:1–144:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bang:2018:SII

[BL18] Seungbae Bang and Sung-Hee Lee. Spline interface for intuitive skinning weight editing. ACM Transactions on Graphics, 37(5): 174:1–174:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3186565.

Bauchet:2020:KSR

[BL20] Jean-Philippe Bauchet and Florent Lafarge. Kinetic shape reconstruction. ACM Transactions on Graphics, 39 (5):156:1–156:14, September 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3376918.

Berthouzoz:2012:TPC

[BLA12] Floraine Berthouzoz, Wilmot Li, and Maneesh Agrawala. Tools for placing cuts and transitions in interview video. ACM Transactions on Graphics, 31(4):67:1–67:8, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Berio:2022:SSB

[BLAE22] Daniel Berio, Frederic Fol Leymarie, Paul Asente, and Jose Echevarria. StrokeStyles: Stroke-based segmentation and stylization of fonts. ACM Transactions on Graphics, 41 (3):28:1–28:21, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3505246.

Bao:2022:HFD

[BLC^{+}22] Linchao Bao, Xiangkai Lin, Yajing Chen, Haoxian Zhang, Sheng Wang, Xuefei Zhe, Di Kang, Haozhi Huang, Xinwei Jiang, Jue Wang, Dong Yu, and Zhengyou Zhang. High-fidelity 3D digital human head creation from RGB-D selfies. ACM Transactions on Graphics, 41 (1):3:1–3:21, February 2022.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3472954.

Bregler:2002:TMM

[BLCD02] Christoph Bregler, Lorie Loeb, Erika Chuang, and Hrishi Deshpande. Turning to the masters: motion capturing cartoons. ACM Transactions on Graphics, 21(3): 399–407, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bangaru:2020:UWA

[BLD20] Sai Praveen Bangaru, Tzu-Mao Li, and Frédo Durand. Unbiased warpedarea sampling for differentiable rendering. ACM Transactions on Graphics, 39(6): 245:1–245:18, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417833.

Berthouzoz:2011:FCA

[BLDA11] Floraine Berthouzoz, Wilmot Li, Mira Dontcheva, and Maneesh Agrawala. A framework for content-adaptive photo manipulation macros: Application to face, landscape, and global manipulations. ACM Transactions on Graphics, 30(5):120:1–120:14, October 2011. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Budninskiy:2016:OVT

[BLdG^{+}16] Max Budninskiy, Beibei Liu, Fernando de Goes, Yiying Tong, Pierre Alliez, and Mathieu Desbrun. Optimal Voronoi tessellations with Hessian-based anisotropy. ACM Transactions on Graphics, 35(6): 242:1–242:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bai:2021:DUS

[BLDL21] Kai Bai, Wei Li, Mathieu Desbrun, and Xiaopei Liu. Dynamic upsampling of smoke through dictionarybased learning. ACM Transactions on Graphics, 40 (1):4:1–4:19, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3412360.

Blinn:1982:GAS

[Bli82] James F. Blinn. A generalization of algebraic surface drawing. ACM Transactions on Graphics, 1(3):235–256, July 1982. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Berger:2013:BSR

[BLN^{+}13] Matthew Berger, Joshua A.

Levine, Luis Gustavo Nonato, Gabriel Taubin, and Claudio T. Silva. A benchmark for surface reconstruction. ACM Transactions on Graphics, 32 (2):20:1–20:17, April 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brady:2014:GDN

[BLPW14] Adam Brady, Jason Lawrence, Pieter Peers, and Westley Weimer. genBRDF: discovering new analytic BRDFs with genetic programming. ACM Transactions on Graphics, 33 (4):114:1–114:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bosch:2011:IGW

[BLR^{+}11] Carles Bosch, Pierre-Yves Laffont, Holly Rushmeier, Julie Dorsey, and George Drettakis. Image-guided weathering: a new approach applied to flow phenomena. ACM Transactions on Graphics, 30(3):20:1–20:13, May 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bi:2021:DRA

[BLS^{+}21] Sai Bi, Stephen Lombardi, Shunsuke Saito, Tomas Simon, Shih-En Wei, Kevyn Mcphail, Ravi Ramamoorthi, Yaser Sheikh, and Jason Saragih. Deep relightable appearance models for animatable faces. ACM Transactions on Graphics, 40 (4):89:1–89:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459829.

Bharaj:2015:CDM

[BLT^{+}15] Gaurav Bharaj, David I. W.

Levin, James Tompkin, Yun Fei, Hanspeter Pfister, Wojciech Matusik, and Changxi Zheng. Computational design of metallophone contact sounds. ACM Transactions on Graphics, 34(6): 223:1–223:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Budninskiy:2016:PCG

[BLTD16] Max Budninskiy, Beibei Liu, Yiying Tong, and Mathieu Desbrun. Power coordinates: a geometric construction of barycentric coordinates on convex polytopes. ACM Transactions on Graphics, 35 (6):241:1–241:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Blythe:2006:DS

[Bly06] David Blythe. The Direct3D 10 system. ACM Transactions on Graphics, 25(3): 724–734, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bennett:2005:VEU

[BM05] Eric P. Bennett and Leonard McMillan. Video enhancement using per-pixel virtual exposures. ACM Transactions on Graphics, 24(3): 845–852, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bennett:2007:CTL

[BM07] Eric P. Bennett and Leonard McMillan. Computational time-lapse video. ACM Transactions on Graphics, 26(3): 102:1–102:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Banterle:2024:SSH

[BMBRD24] Francesco Banterle, Demetris Marnerides, Thomas Bashford-Rogers, and Kurt Debattista. Self-supervised high dynamic range imaging: What can be learned from a single 8-bit video? ACM Transactions on Graphics, 43 (2):24:1–24:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3648570.

Biermann:2002:CPE

[BMBZ02] Henning Biermann, Ioana Martin, Fausto Bernardini, and Denis Zorin. Cut-and-paste editing of multiresolution surfaces. ACM Transactions on Graphics, 21(3): 312–321, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bertiche:2021:PPB

[BME21] Hugo Bertiche, Meysam Madadi, and Sergio Escalera. PBNS: physically based neural simulation for unsupervised garment pose space deformation. ACM Transactions on Graphics, 40(6): 198:1–198:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480479.

Bertiche:2022:NCS

[BME22] Hugo Bertiche, Meysam Madadi, and Sergio Escalera. Neural cloth simulation. ACM Transactions on Graphics, 41 (6):220:1–220:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555491.

Bouaziz:2014:PDF

[BML^{+}14] Sofien Bouaziz, Sebastian Martin, Tiantian Liu, Ladislav Kavan, and Mark Pauly. Projective dynamics: fusing constraint projections for fast simulation. ACM Transactions on Graphics, 33(4): 154:1–154:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bangaru:2021:SDP

[BMM^{+}21] Sai Praveen Bangaru, Jesse Michel, Kevin Mu, Gilbert Bernstein, Tzu-Mao Li, and Jonathan Ragan-Kelley. Systematically differentiating parametric discontinuities. ACM Transactions on Graphics, 40(4):107:1–107:18, August 2021. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459775.

Bailey:2009:SGD

[BMSG09] Reynold Bailey, Ann Mc-Namara, Nisha Sudarsanam, and Cindy Grimm. Subtle gaze direction. ACM Transactions on Graphics, 28(4): 100:1–100:14, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bemana:2020:XFI

[BMSR20] Mojtaba Bemana, Karol Myszkowski, Hans-Peter Seidel, and Tobias Ritschel. X-Fields: implicit neural view- , light- and time-image interpolation. ACM Transactions on Graphics, 39(6): 257:1–257:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417827.

Bittner:2009:AGV

[BMW^{+}09] Jiří Bittner, Oliver Mattausch, Peter Wonka, Vlastimil Havran, and Michael Wimmer. Adaptive global visibility sampling. ACM Transactions on Graphics, 28(3): 94:1–94:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bergou:2007:TTD

[BMWG07] Miklós Bergou, Saurabh Mathur, Max Wardetzky, and Eitan Grinspun. TRACKS: toward directable thin shells. ACM Transactions on Graphics, 26(3):50:1–50:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brunet:1990:SRO

[BN90] Pere Brunet and Isabel Navazo. Solid representation and operation using extended octrees. ACM Transactions on Graphics, 9(2): 170–197, April 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
78959.html.

Brown:2021:WWR

[BN21] George E. Brown and Rahul Narain. WRAPD: weighted rotation-aware ADMM for parameterization and deformation. ACM Transactions on Graphics, 40(4): 82:1–82:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459942.

Bradley:2013:IBR

[BNB13] Derek Bradley, Derek Nowrouzezahrai, and Paul Beardsley. Imagebased reconstruction and synthesis of dense foliage. ACM Transactions on Graphics, 32 (4):74:1–74:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barnum:2010:MLD

[BNK10] Peter C. Barnum, Srinivasa G. Narasimhan, and Takeo Kanade. A multilayered display with water drops. ACM Transactions on Graphics, 29(4): 76:1–76:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bousseau:2007:VWU

[BNTS07] Adrien Bousseau, Fabrice Neyret, Joëlle Thollot, and David Salesin. Video watercolorization using bidirectional texture advection. ACM Transactions on Graphics, 26(3):104:1–104:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bradshaw:2004:AMA

[BO04] Gareth Bradshaw and Carol O’Sullivan. Adaptive medialaxis approximation for spheretree construction. ACM Transactions on Graphics, 23 (1):1–26, January 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bousseau:2013:GPP

[BOD^{+}13] Adrien Bousseau, James P.

O’Shea, Frédo Durand, Ravi Ramamoorthi, and Maneesh Agrawala. Gloss perception in painterly and cartoon rendering. ACM Transactions on Graphics, 32(2): 18:1–18:13, April 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bailey:2018:FDD

[BODO18] Stephen W. Bailey, Dave Otte, Paul Dilorenzo, and James F. O’Brien. Fast and deep deformation approximations. ACM Transactions on Graphics, 37(4): 119:1–119:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brown:2018:ADF

[BOFN18] George E. Brown, Matthew Overby, Zahra Forootaninia, and Rahul Narain. Accurate dissipative forces in optimization integrators. ACM Transactions on Graphics, 37(6):

282:1–282:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boissonnat:1984:GST

[Boi84] Jean-Daniel Boissonnat. Geometric structures for threedimensional shape representation. ACM Transactions on Graphics, 3(4):266–286, October 1984. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boubekeur:2018:SDA

[Bou18] Tamy Boubekeur. Session details: Acquiring and editing geometry via RGB (D) images. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baran:2007:ARA

[BP07] Ilya Baran and Jovan Popović. Automatic rigging and animation of 3D characters. ACM Transactions on Graphics, 26 (3):72:1–72:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barbic:2008:RTC

[BP08] Jernej Barbič and Jovan Popović. Real-time control of physically based simulations using gentle forces. ACM Transactions on Graphics, 27 (5):163:1–163:??, December 2008. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Bau:2012:RTF

[BP12] Olivier Bau and Ivan Poupyrev.

REVEL: tactile feedback technology for augmented reality. ACM Transactions on Graphics, 31(4):89:1–89:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boyadzhiev:2013:UAI

[BPB13] Ivaylo Boyadzhiev, Sylvain Paris, and Kavita Bala. Userassisted image compositing for photographic lighting. ACM Transactions on Graphics, 32 (4):36:1–36:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bonneel:2016:WBC

[BPC16] Nicolas Bonneel, Gabriel Peyré, and Marco Cuturi. Wasserstein barycentric coordinates: histogram regression using optimal transport. ACM Transactions on Graphics, 35 (4):71:1–71:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bae:2006:TST

[BPD06] Soonmin Bae, Sylvain Paris, and Frédo Durand. Two-scale tone management for photographic look. ACM Transactions on Graphics, 25(3):

637–645, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bousseau:2009:UAI

[BPD09] Adrien Bousseau, Sylvain Paris, and Frédo Durand. User-assisted intrinsic images. ACM Transactions on Graphics, 28(5):130:1– 130:10, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bauszat:2017:GDP

[BPE17] Pablo Bauszat, Victor Petitjean, and Elmar Eisemann. Gradient-domain path reusing. ACM Transactions on Graphics, 36(6):229:1– 229:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bischoff:2005:ARP

[BPK05] Stephan Bischoff, Darko Pavic, and Leif Kobbelt. Automatic restoration of polygon models. ACM Transactions on Graphics, 24(4): 1332–1352, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bo:2011:CAS

[BPK^{+}11] Pengbo Bo, Helmut Pottmann, Martin Kilian, Wenping Wang, and Johannes Wallner. Circular arc structures. ACM Transactions on Graphics, 30 (4):101:1–101:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2013:WCP

[BPK^{+}13] Jongmin Baek, Dawid Pajak, Kihwan Kim, Kari Pulli, and Marc Levoy. WYSIWYG computational photography via viewfinder editing. ACM Transactions on Graphics, 32 (6):198:1–198:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bradley:2008:MGC

[BPS^{+}08] Derek Bradley, Tiberiu Popa, Alla Sheffer, Wolfgang Heidrich, and Tamy Boubekeur. Markerless garment capture. ACM Transactions on Graphics, 27(3):99:1–99:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Borrel:1994:SCD

[BR94] Paul Borrel and Ari Rappoport. Simple constrained deformations for geometric modeling and interactive design. ACM Transactions on Graphics, 13(2): 137–155, April 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
176581.html.

Brown:2007:GNR

[BR07] Benedict J. Brown and Szymon Rusinkiewicz. Global non-rigid alignment of 3-D scans. ACM Transactions on Graphics, 26(3): 21:1–21:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baerentzen:2021:SLS

[BR21a] Andreas Bærentzen and Eva Rotenberg. Skeletonization via local separators. ACM Transactions on Graphics, 40(5):187:1–187:18, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3459233.

Brunton:2021:DSD

[BR21b] Alan Brunton and Lubna Abu Rmaileh. Displaced signed distance fields for additive manufacturing. ACM Transactions on Graphics, 40(4): 179:1–179:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459827.

Buffet:2019:IUR

[BRB^{+}19] Thomas Buffet, Damien Rohmer, Loïc Barthe, Laurence Boissieux, and Marie-Paule Cani. Implicit untangling: a robust solution for modeling layered clothing.

ACM Transactions on Graphics, 38(4):120:1–120:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bitterli:2018:RTF

[BRM^{+}18] Benedikt Bitterli, Srinath Ravichandran, Thomas Müller, Magnus Wrenninge, Jan Novák, Steve Marschner, and Wojciech Jarosz. A radiative transfer framework for non-exponential media. ACM Transactions on Graphics, 37 (6):225:1–225:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bashford-Rogers:2022:EML

[BRSMD22] Thomas Bashford-Rogers, Luís Paulo Santos, Demetris Marnerides, and Kurt Debattista. Ensemble Metropolis light transport. ACM Transactions on Graphics, 41 (1):5:1–5:15, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3472294.

Ball:1988:CTP

[BS88] A. A. Ball and D. J. T.

Storry. Conditions for tangent plane continuity over recursively generated B-spline surfaces. ACM Transactions on Graphics, 7(2):83–102, April 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        42459.html.

Ball:1990:ICV

[BS90] A. A. Ball and D. J. T. Storry. An investigation of curvature variations over recursively generated B-spline surfaces. ACM Transactions on Graphics, 9(4):424–437, October 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        88580.html.

Bessmeltsev:2019:VLD

[BS19] Mikhail Bessmeltsev and Justin Solomon. Vectorization of line drawings via polyvector fields. ACM Transactions on Graphics, 38 (1):9:1–9:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3202661.

Birklbauer:2016:NSD

[BSB16] Clemens Birklbauer, David C. Schedl, and Oliver Bimber. Nonuniform spatial deformation of light fields by locally linear transformations. ACM Transactions on Graphics, 35 (5):156:1–156:??, September 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Birklbauer:2017:NSD

[BSB17] Clemens Birklbauer, David C.

Schedl, and Oliver Bimber. Nonuniform spatial deformation of light fields by locally linear transformations. ACM Transactions on Graphics, 36 (4):143:1–143:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brouet:2012:DPG

[BSBC12] Remi Brouet, Alla Sheffer, Laurence Boissieux, and Marie-Paule Cani. Design preserving garment transfer. ACM Transactions on Graphics, 31(4):36:1–36:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Balzer:2009:CCP

[BSD09] Michael Balzer, Thomas Schlömer, and Oliver Deussen. Capacity-constrained point distributions: a variant of Lloyd’s method. ACM Transactions on Graphics, 28(3): 86:1–86:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brandt:2018:MSV

[BSEH18] Christopher Brandt, Leonardo Scandolo, Elmar Eisemann, and Klaus Hildebrandt. Modeling n-symmetry vector fields using higher-order energies. ACM Transactions on Graphics, 37(2):18:1–18:??, July 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barnes:2009:PRC

[BSFG09] Connelly Barnes, Eli Shechtman, Adam Finkelstein, and Dan B. Goldman. Patch-Match: a randomized correspondence algorithm for structural image editing. ACM Transactions on Graphics, 28(3):24:1–24:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barbic:2012:IED

[BSG12] Jernej Barbic, Funshing Sin, and Eitan Grinspun. Interactive editing of deformable simulations. ACM Transactions on Graphics, 31(4): 70:1–70:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bhat:2004:FBV

[BSHK04] Kiran S. Bhat, Steven M. Seitz, Jessica K. Hodgins, and Pradeep K. Khosla. Flowbased video synthesis and editing. ACM Transactions on Graphics, 23(3):360– 363, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bartle:2016:PDP

[BSK^{+}16] Aric Bartle, Alla Sheffer, Vladimir G. Kim, Danny M.

Kaufman, Nicholas Vining, and Floraine Berthouzoz. Physics-driven pattern adjustment for direct 3D garment editing. ACM Transactions on Graphics, 35(4): 50:1–50:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bako:2023:DAP

[BSK23] Steve Bako, Pradeep Sen, and Anton Kaplanyan. Deep appearance prefiltering. ACM Transactions on Graphics, 42 (2):23:1–23:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3570327.

Bai:2012:SCO

[BSL12] Yunfei Bai, Kristin Siu, and C. Karen Liu. Synthesis of concurrent object manipulation tasks. ACM Transactions on Graphics, 31(6): 156:1–156:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bernstein:2016:EDP

[BSL^{+}16] Gilbert Louis Bernstein, Chinmayee Shah, Crystal Lemire, Zachary Devito, Matthew Fisher, Philip Levis, and Pat Hanrahan. Ebb: a DSL for physical simulation on CPUs and GPUs. ACM Transactions on Graphics, 35(2):21:1–21:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bleser:1988:CSR

[BSM88] Teresa W. Bleser, John L. Sibert, and J. Patrick McGee. Charcoal sketching: Returning control to the artist. ACM Transactions on Graphics, 7(1):76–81, January 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        42230.html.

Breslav:2007:DPS

[BSM^{+}07] Simon Breslav, Karol Szerszen, Lee Markosian, Pascal Barla, and Joëlle Thollot. Dynamic 2D patterns for shading 3D scenes. ACM Transactions on Graphics, 26(3): 20:1–20:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Berger:2013:SAP

[BSM^{+}13] Itamar Berger, Ariel Shamir, Moshe Mahler, Elizabeth Carter, and Jessica Hodgins. Style and abstraction in portrait sketching. ACM Transactions on Graphics, 32(4): 55:1–55:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bagher:2016:NPF

[BSN16] Mahdi M. Bagher, John Snyder, and Derek Nowrouzezahrai. A non-parametric factor microfacet model for isotropic BRDFs. ACM Transactions on Graphics, 35(5):159:1– 159:??, September 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bau:2019:SPM

[BSP^{+}19] David Bau, Hendrik Strobelt, William Peebles, Jonas Wulff, Bolei Zhou, Jun-Yan Zhu, and Antonio Torralba. Semantic photo manipulation with a generative image prior. ACM Transactions on Graphics, 38 (4):59:1–59:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bonneel:2013:EBV

[BSPP13] Nicolas Bonneel, Kalyan Sunkavalli, Sylvain Paris, and Hanspeter Pfister. Examplebased video color grading. ACM Transactions on Graphics, 32(4):39:1–39:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Becker:2023:CSD

[BSR^{+}23] Quentin Becker, Seiichi Suzuki, Yingying Ren, Davide Pellis, Julian Panetta, and Mark Pauly. C-shells: Deployable gridshells with curved beams. ACM Transactions on Graphics, 42(6): 173:1–173:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618366.

Baran:2011:MOC

[BSS^{+}11] Ilya Baran, Johannes Schmid, Thomas Siegrist, Markus Gross, and Robert W. Sumner. Mixed-order compositing for 3D paintings. ACM Transactions on Graphics, 30(6): 132:1–132:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2013:CTE

[BSS^{+}13] Laurent Belcour, Cyril Soler, Kartic Subr, Nicolas Holzschuch, and Fredo Durand. 5D covariance tracing for efficient defocus and motion blur. ACM Transactions on Graphics, 32 (3):31:1–31:18, June 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bang:2023:AFM

[BSSJ23] Seungbae Bang, Kirill Serkh, Oded Stein, and Alec Jacobson. An adaptive fastmultipole-accelerated hybrid boundary integral equation method for accurate diffusion curves. ACM Transactions on Graphics, 42(6): 215:1–215:??, December 2023.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618374.

Bratkova:2009:ARM

[BST09] Margarita Bratkova, Peter Shirley, and William B. Thompson. Artistic rendering of mountainous terrain. ACM Transactions on Graphics, 28(4):102:1–102:18, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bonneel:2014:IIV

[BST^{+}14] Nicolas Bonneel, Kalyan Sunkavalli, James Tompkin, Deqing Sun, Sylvain Paris, and Hanspeter Pfister. Interactive intrinsic video editing. ACM Transactions on Graphics, 33(6):197:1– 197:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boissonnat:2015:ADM

[BSTY15] Jean-Daniel Boissonnat, Kan-Le Shi, Jane Tournois, and Mariette Yvinec. Anisotropic Delaunay meshes of surfaces. ACM Transactions on Graphics, 34(2):14:1–14:??, February 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bederson:2002:OQT

[BSW02] Benjamin B. Bederson, Ben Shneiderman, and Martin Wattenberg. Ordered and quantum treemaps: Making effective use of 2D space to display hierarchies. ACM Transactions on Graphics, 21 (4):833–854, October 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bao:2013:PFV

[BSW13] Fan Bao, Michael Schwarz, and Peter Wonka. Procedural facade variations from a single layout. ACM Transactions on Graphics, 32(1): 8:1–8:13, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bansal:2019:AIL

[BT19] Sumukh Bansal and Aditya Tatu. Affine interpolation in a lie group framework. ACM Transactions on Graphics, 38 (4):71:1–71:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Brown:2008:SHV

[BTFN^{+}08] Benedict J. Brown, Corey Toler-Franklin, Diego Nehab, Michael Burns, David Dobkin, Andreas Vlachopoulos, Christos Doumas, Szymon Rusinkiewicz, and Tim Weyrich. A system for high-volume acquisition and matching of fresco fragments: reassembling Theran wall paintings. ACM Transactions on Graphics, 27(3):84:1– 84:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bonneel:2015:BVT

[BTS^{+}15] Nicolas Bonneel, James Tompkin, Kalyan Sunkavalli, Deqing Sun, Sylvain Paris, and Hanspeter Pfister. Blind video temporal consistency. ACM Transactions on Graphics, 34 (6):196:1–196:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barda:2023:GDS

[BTSB23] Amir Barda, Guy Tevet, Adriana Schulz, and Amit Haim Bermano. Generative design of sheet metal structures. ACM Transactions on Graphics, 42(4):116:1– 116:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592444.

Batty:2012:DVS

[BUAG12] Christopher Batty, Andres Uribe, Basile Audoly, and Eitan Grinspun. Discrete viscous sheets. ACM Transactions on Graphics, 31(4): 113:1–113:7, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bell:2013:ORA

[BUSB13] Sean Bell, Paul Upchurch, Noah Snavely, and Kavita Bala. OpenSurfaces: a richly annotated catalog of surface appearance. ACM Transactions on Graphics, 32(4): 111:1–111:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Boksebeld:2022:HOD

[BV22] Iwan Boksebeld and Amir Vaxman. High-order directional fields. ACM Transactions on Graphics, 41(6): 254:1–254:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555455.

Bonneel:2011:DIU

[BvdPPH11] Nicolas Bonneel, Michiel van de Panne, Sylvain Paris, and Wolfgang Heidrich. Displacement interpolation using Lagrangian mass transport. ACM Transactions on Graphics, 30(6):158:1– 158:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Babaei:2017:CCP

[BVF^{+}17a] Vahid Babaei, Kiril Vidimce, Michael Foshey, Alexandre Kaspar, Piotr Didyk, and Wojciech Matusik. Color contoning for 3D printing. ACM Transactions on Graphics, 36 (4):124:1–124:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Borno:2017:DAEa

[BVF17b] Mazen Al Borno, Michiel Van De Panne, and Eugene Fiume. Domain of attraction expansion for physics-based character control. ACM Transactions on Graphics, 36(2): 17:1–17:??, April 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bermano:2011:ORO

[BVG11] Amit Bermano, Amir Vaxman, and Craig Gotsman. Online reconstruction of 3D objects from arbitrary crosssections. ACM Transactions on Graphics, 30(5):113:1– 113:14, October 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baran:2009:SDT

[BVGP09] Ilya Baran, Daniel Vlasic, Eitan Grinspun, and Jovan Popović. Semantic deformation transfer. ACM Transactions on Graphics, 28(3): 36:1–36:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bako:2017:KPC

[BVM^{+}17] Steve Bako, Thijs Vogels, Brian Mcwilliams, Mark Meyer, Jan NováK, Alex Harvill, Pradeep Sen, Tony Derose, and Fabrice Rousselle. Kernel-predicting convolutional networks for denoising Monte Carlo renderings. ACM Transactions on Graphics, 36(4):97:1–97:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bessmeltsev:2016:GPC

[BVS16] Mikhail Bessmeltsev, Nicholas Vining, and Alla Sheffer. Gesture3D: posing 3D characters via gesture drawings. ACM Transactions on Graphics, 35(6):165:1– 165:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bernstein:2013:PHH

[BW13] Gilbert Louis Bernstein and Chris Wojtan. Putting holes in holey geometry: topology change for arbitrary surfaces. ACM Transactions on Graphics, 32(4):34:1–34:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Balbao:2022:BIH

[BW22] Arthur E. Balbão and Marcelo Walter. A biologically inspired hair aging model. ACM Transactions on Graphics, 41 (6):223:1–223:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555444.

Bacher:2014:SIO

[BWBSH14] Moritz Bächer, Emily Whiting, Bernd Bickel, and Olga Sorkine-Hornung. Spin-it: optimizing moment of inertia for spinnable objects. ACM Transactions on Graphics, 33 (4):96:1–96:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baek:2023:CWF

[BWC^{+}23] Seung-Hwan Baek, Noah Walsh, Ilya Chugunov, Zheng Shi, and Felix Heide. Centimeterwave free-space neural timeof-flight imaging. ACM Transactions on Graphics, 42 (1):3:1–3:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3522671.

Bai:2021:PHR

[BWDL21] Kai Bai, Chunhao Wang, Mathieu Desbrun, and Xiaopei Liu. Predicting highresolution turbulence details in space and time. ACM Transactions on Graphics, 40 (6):200:1–200:16, December 2021. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478513.3480492.

Bala:2003:CEP

[BWG03] Kavita Bala, Bruce Walter, and Donald P. Greenberg. Combining edges and points for interactive high-quality rendering. ACM Transactions on Graphics, 22(3): 631–640, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bargteil:2007:FEM

[BWHT07] Adam W. Bargteil, Chris Wojtán, Jessica K. Hodgins, and Greg Turk. A finite element method for animating large viscoplastic flow. ACM Transactions on Graphics, 26 (3):16:1–16:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baraff:2003:UC

[BWK03] David Baraff, Andrew Witkin, and Michael Kass. Untangling cloth. ACM Transactions on Graphics, 22(3):862–870, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bokeloh:2011:PAS

[BWKS11] Martin Bokeloh, Michael Wand, Vladlen Koltun, and Hans-Peter Seidel. Patternaware shape deformation using sliding dockers. ACM Transactions on Graphics, 30 (6):123:1–123:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bangaru:2023:SDF

[BWL^{+}23] Sai Praveen Bangaru, Lifan Wu, Tzu-Mao Li, Jacob Munkberg, Gilbert Bernstein, Jonathan Ragan-Kelley, Frédo Durand, Aaron Lefohn, and Yong He. SLANG.D: Fast, modular and differentiable shader programming. ACM Transactions on Graphics, 42(6):264:1– 264:??, December 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618353.

Bouaziz:2013:OMR

[BWP13] Sofien Bouaziz, Yangang Wang, and Mark Pauly. Online modeling for realtime facial animation. ACM Transactions on Graphics, 32(4): 40:1–40:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bergou:2008:DER

[BWR^{+}08] Miklós Bergou, Max Wardetzky, Stephen Robinson, Basile Audoly, and Eitan Grinspun. Discrete elastic rods. ACM Transactions on Graphics, 27 (3):63:1–63:??, August 2008.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Baranoski:2005:SDA

[BWRB05] Gladimir V. G. Baranoski, Justin Wan, Jon G. Rokne, and Ian Bell. Simulating the dynamics of auroral phenomena. ACM Transactions on Graphics, 24(1): 37–59, January 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bokeloh:2010:CBP

[BWS10] Martin Bokeloh, Michael Wand, and Hans-Peter Seidel. A connection between partial symmetry and inverse procedural modeling. ACM Transactions on Graphics, 29(4): 104:1–104:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bagautdinov:2021:DSA

[BWS^{+}21] Timur Bagautdinov, Chenglei Wu, Tomas Simon, Fabián Prada, Takaaki Shiratori, Shih-En Wei, Weipeng Xu, Yaser Sheikh, and Jason Saragih. Driving-signal aware full-body avatars. ACM Transactions on Graphics, 40(4):143:1–143:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3450626.3459850.

Bokeloh:2012:AMP

[BWSK12] Martin Bokeloh, Michael Wand, Hans-Peter Seidel, and Vladlen Koltun. An algebraic model for parameterized shape editing. ACM Transactions on Graphics, 31(4): 78:1–78:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bai:2009:VSR

[BWSS09] Xue Bai, Jue Wang, David Simons, and Guillermo Sapiro. Video SnapCut: robust video object cutout using localized classifiers. ACM Transactions on Graphics, 28(3):70:1– 70:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bessmeltsev:2012:DDQ

[BWSS12] Mikhail Bessmeltsev, Caoyu Wang, Alla Sheffer, and Karan Singh. Design-driven quadrangulation of closed 3D curves. ACM Transactions on Graphics, 31(6):178:1– 178:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bowers:2010:PPD

[BWWM10] John Bowers, Rui Wang, Li-Yi Wei, and David Maletz. Parallel Poisson disk sampling with spectrum analysis on surfaces. ACM Transactions on Graphics, 29(6):

166:1–166:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bajaj:2003:ADS

[BX03] Chandrajit L. Bajaj and Guoliang Xu. Anisotropic diffusion of surfaces and functions on surfaces. ACM Transactions on Graphics, 22(1): 4–32, January 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2018:ICS

[BXH^{+}18] Laurent Belcour, Guofu Xie, Christophe Hery, Mark Meyer, Wojciech Jarosz, and Derek Nowrouzezahrai. Integrating clipped spherical harmonics expansions. ACM Transactions on Graphics, 37(2): 19:1–19:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bar-Yehuda:1996:TST

[BYG96] Rueven Bar-Yehuda and Craig Gotsman. Time/space tradeoffs for polygon mesh rendering. ACM Transactions on Graphics, 15(2): 141–152, April 1996. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bertel:2020:OCV

[BYLR20] Tobias Bertel, Mingze Yuan, Reuben Lindroos, and Christian Richardt. OmniPhotos: casual 360^{∘} VR photography. ACM Transactions on Graphics, 39(6): 266:1–266:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417770.

Bao:2013:GEG

[BYMW13] Fan Bao, Dong-Ming Yan, Niloy J. Mitra, and Peter Wonka. Generating and exploring good building layouts. ACM Transactions on Graphics, 32(4):122:1–122:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2017:ACGa

[BYRN17a] Laurent Belcour, Ling-Qi Yan, Ravi Ramamoorthi, and Derek Nowrouzezahrai. Antialiasing complex global illumination effects in path-space. ACM Transactions on Graphics, 36(1):9:1–9:??, February 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Belcour:2017:ACGb

[BYRN17b] Laurent Belcour, Ling-Qi Yan, Ravi Ramamoorthi, and Derek Nowrouzezahrai. Antialiasing complex global illumination effects in path-space. ACM Transactions on Graphics, 36(4):75:1–75:??, July 2017. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Barbic:2011:RTL

[BZ11] Jernej Barbič and Yili Zhao. Real-time large-deformation substructuring. ACM Transactions on Graphics, 30(4): 91:1–91:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Benchekroun:2023:FCD

[BZC^{+}23] Otman Benchekroun, Jiayi Eris Zhang, Siddartha Chaudhuri, Eitan Grinspun, Yi Zhou, and Alec Jacobson. Fast complementary dynamics via skinning eigenmodes. ACM Transactions on Graphics, 42(4): 106:1–106:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592404.

Bhat:2010:GGD

[BZCC10] Pravin Bhat, C. Lawrence Zitnick, Michael Cohen, and Brian Curless. GradientShop: a gradient-domain optimization framework for image and video filtering. ACM Transactions on Graphics, 29(2): 10:1–10:14, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Bharadwaj:2023:FFL

[BZH^{+}23] Shrisha Bharadwaj, Yufeng Zheng, Otmar Hilliges, Michael J. Black, and Victoria Fernandez Abrevaya. FLARE: Fast learning of animatable and relightable mesh avatars. ACM Transactions on Graphics, 42 (6):204:1–204:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618401.

Bommes:2009:MIQ

[BZK09] David Bommes, Henrik Zimmer, and Leif Kobbelt. Mixed-integer quadrangulation. ACM Transactions on Graphics, 28(3):77:1– 77:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Barnes:2015:PEP

[BZL^{+}15] Connelly Barnes, Fang-Lue Zhang, Liming Lou, Xian Wu, and Shi-Min Hu. PatchTable: efficient patch queries for large datasets and applications. ACM Transactions on Graphics, 34(4):97:1–97:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

iRibera:2017:FRA

[BZL^{+}17] Roger Blanco i Ribera, Eduard Zell, J. P. Lewis, Junyong Noh, and Mario Botsch.

Facial retargeting with automatic range of motion alignment. ACM Transactions on Graphics, 36(4):154:1– 154:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2000:TAS

[CA00] Min Chen and James Arvo. Theory and application of specular path perturbation. ACM Transactions on Graphics, 19(4):246–278, October 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
2000-19-4/p246-chen/.

Careaga:2024:IID

[CA24] Chris Careaga and Yagiz Aksoy. Intrinsic image decomposition via ordinal shading. ACM Transactions on Graphics, 43(1):12:1–12:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3630750.

Carroll:2009:OCP

[CAA09] Robert Carroll, Maneesh Agrawal, and Aseem Agarwala. Optimizing contentpreserving projections for wide-angle images. ACM Transactions on Graphics, 28 (3):43:1–43:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Carroll:2010:IWA

[CAA10] Robert Carroll, Aseem Agarwala, and Maneesh Agrawala. Image warps for artistic perspective manipulation. ACM Transactions on Graphics, 29 (4):127:1–127:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chuang:2002:VMC

[CAC^{+}02] Yung-Yu Chuang, Aseem Agarwala, Brian Curless, David H. Salesin, and Richard Szeliski. Video matting of complex scenes. ACM Transactions on Graphics, 21(3): 243–248, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chapiro:2019:LAM

[CAD19] Alexandre Chapiro, Robin Atkins, and Scott Daly. A luminance-aware model of judder perception. ACM Transactions on Graphics, 38(5):142:1–142:??, October 2019. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3338696.

Cao:2021:RTN

[CAD^{+}21] Chen Cao, Vasu Agrawal, Fernando De La Torre, Lele Chen, Jason Saragih, Tomas Simon, and Yaser Sheikh. Real-time 3D neural facial animation from binocular video. ACM Transactions on Graphics, 40(4):87:1–87:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459806.

Cashman:2009:NEP

[CADS09] Thomas J. Cashman, Ursula H. Augsdörfer, Neil A. Dodgson, and Malcolm A. Sabin. NURBS with extraordinary points: high-degree, non-uniform, rational subdivision schemes. ACM Transactions on Graphics, 28(3): 46:1–46:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chadwick:2009:HSP

[CAJ09] Jeffrey N. Chadwick, Steven S. An, and Doug L. James. Harmonic shells: a practical nonlinear sound model for nearrigid thin shells. ACM Transactions on Graphics, 28(5): 119:1–119:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chang:2009:SAE

[CAO09] Jianghao Chang, Benoît Alain, and Victor Ostromoukhov. Structure-aware error diffusion. ACM Transactions on Graphics, 28(5): 162:1–162:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chentanez:2009:ISS

[CAR^{+}09] Nuttapong Chentanez, Ron Alterovitz, Daniel Ritchie, Lita Cho, Kris K. Hauser, Ken Goldberg, Jonathan R. Shewchuk, and James F. O’Brien. Interactive simulation of surgical needle insertion and steering. ACM Transactions on Graphics, 28 (3):88:1–88:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Casner:1991:TAA

[Cas91] Stephen M. Casner. A taskanalytic approach to the automated design of graphic presentations. ACM Transactions on Graphics, 10(2): 111–151, April 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
108361.html.

Chefer:2023:AEA

[CAV^{+}23] Hila Chefer, Yuval Alaluf, Yael Vinker, Lior Wolf, and Daniel Cohen-Or. Attendand-excite: Attention-based semantic guidance for text-to-image diffusion models. ACM Transactions on Graphics, 42(4):148:1–148:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592116.

Chen:2016:BGU

[CAWH16] Jiawen Chen, Andrew Adams, Neal Wadhwa, and Samuel W. Hasinoff. Bilateral guided upsampling. ACM Transactions on Graphics, 35(6): 203:1–203:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2004:VIT

[CB04] Xiang Cao and Ravin Balakrishnan. VisionWand: interaction techniques for large displays using a passive wand tracked in 3D. ACM Transactions on Graphics, 23(3): 729, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chuang:2005:MSE

[CB05] Erika Chuang and Christoph Bregler. Mood swings: expressive speech animation. ACM Transactions on Graphics, 24 (2):331–347, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Calderon:2014:PM

[CB14] Stéphane Calderon and Tamy Boubekeur. Point morphology. ACM Transactions on Graphics, 33(4): 45:1–45:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Calderon:2017:BPS

[CB17] Stéphane Calderon and Tamy Boubekeur. Bounding proxies for shape approximation. ACM Transactions on Graphics, 36(4):57:1–57:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2002:LFM

[CBCG02] Wei-Chao Chen, Jean-Yves Bouguet, Michael H. Chu, and Radek Grzeszczuk. Light field mapping: efficient representation and hardware rendering of surface light fields. ACM Transactions on Graphics, 21 (3):447–456, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Casati:2013:SSC

[CBD13] Romain Casati and Florence Bertails-Descoubes. Super space clothoids. ACM Transactions on Graphics, 32(4): 48:1–48:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2013:SRTa

[CBI13] Jiawen Chen, Dennis Bautembach, and Shahram Izadi.

Scalable real-time volumetric surface reconstruction. ACM Transactions on Graphics, 32 (4):113:1–113:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2012:DLM

[CBK12] Marcel Campen, David Bommes, and Leif Kobbelt. Dual loops meshing: quality quad layouts on manifolds. ACM Transactions on Graphics, 31(4): 110:1–110:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2015:QGP

[CBK15] Marcel Campen, David Bommes, and Leif Kobbelt. Quantized global parametrization. ACM Transactions on Graphics, 34 (6):192:1–192:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Coevoet:2020:AMR

[CBK20] Eulalie Coevoet, Otman Benchekroun, and Paul G. Kry. Adaptive merging for rigid body simulation. ACM Transactions on Graphics, 39 (4):35:1–35:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392417.

Chen:2015:HMS

[CBKM15] Tenn F. Chen, Gladimir V. G.

Baranoski, Bradley W. Kimmel, and Erik Miranda. Hyperspectral modeling of skin appearance. ACM Transactions on Graphics, 34(3): 31:1–31:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Choi:2016:SSB

[CBL^{+}16] Byungkuk Choi, Roger Blanco i Ribera, J. P. Lewis, Yeongho Seol, Seokpyo Hong, Haegwang Eom, Sunjin Jung, and Junyong Noh. SketchiMo: sketch-based motion editing for articulated characters. ACM Transactions on Graphics, 35(4):146:1–146:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chakravarthula:2022:PAH

[CBS^{+}22]  Praneeth  Chakravarthula,

Seung-Hwan Baek, Florian Schiffers, Ethan Tseng, Grace Kuo, Andrew Maimone, Nathan Matsuda, Oliver Cossairt, Douglas Lanman, and Felix Heide. Pupil-aware holography. ACM Transactions on Graphics, 41(6): 212:1–212:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555508.

Coros:2009:RTB

[CBvdP09] Stelian Coros, Philippe Beaudoin, and Michiel van de Panne. Robust task-based control policies for physicsbased characters. ACM Transactions on Graphics, 28(5):170:1–170:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Coros:2010:GBW

[CBvdP10] Stelian Coros, Philippe Beaudoin, and Michiel van de Panne. Generalized biped walking control. ACM Transactions on Graphics, 29(4): 130:1–130:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2018:NCU

[CBW^{+}18] Jiong Chen, Hujun Bao, Tianyu Wang, Mathieu Desbrun, and Jin Huang. Numerical coarsening using discontinuous shape functions. ACM Transactions on Graphics, 37(4):120:1–120:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chae:2023:EEH

[CBYJ23] Minseok Chae, Kiseung Bang, Dongheon Yoo, and Yoonchan Jeong. Étendue expansion in holographic near eye displays through sparse eyebox generation using lens array eyepiece. ACM Transactions on Graphics, 42(4): 58:1–58:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592441.

Coros:2008:SCW

[CBYvdP08] Stelian Coros, Philippe Beaudoin, Kang Kang Yin, and Michiel van de Pann. Synthesis of constrained walking skills. ACM Transactions on Graphics, 27(5): 113:1–113:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2015:RTH

[CBZB15] Chen Cao, Derek Bradley, Kun Zhou, and Thabo Beeler. Real-time high-fidelity facial performance capture. ACM Transactions on Graphics, 34 (4):46:1–46:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Corman:2019:SMF

[CC19] Etienne Corman and Keenan Crane. Symmetric moving frames. ACM Transactions on Graphics, 38(4): 87:1–87:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Coiffier:2023:MMF

[CC23] Guillaume Coiffier and Etienne Corman. The method of moving frames for surface global parametrization. ACM Transactions on Graphics, 42(5):166:1–166:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3604282.

Cali:2012:PNA

[CCA^{+}12] Jacques Calì, Dan A. Calian, Cristina Amati, Rebecca Kleinberger, Anthony Steed, Jan Kautz, and Tim Weyrich. 3D-printing of non-assembly, articulated models. ACM Transactions on Graphics, 31 (6):130:1–130:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chandran:2022:LAC

[CCGB22] Prashanth Chandran, Loïc Ciccone, Markus Gross, and Derek Bradley. Local anatomically-constrained facial performance retargeting. ACM Transactions on Graphics, 41(4):168:1–168:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530114.

Cameron:1992:RMG

[CCK92] Stephen Cameron and Yap Chee-Keng. Refinement methods for geometric bounds in constructive solid geometry. ACM Transactions on Graphics, 11(1):12–39, January 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
123764.html.

Chen:2021:FWC

[CCK^{+}21] Zhen Chen, Hsiao-Yu Chen, Danny M. Kaufman, Mélina Skouras, and Etienne Vouga. Fine wrinkling on coarsely meshed thin shells. ACM Transactions on Graphics, 40(5):190:1–190:32, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3462758.

Cao:2012:ASM

[CCL12] Ying Cao, Antoni B. Chan, and Rynson W. H. Lau. Automatic stylistic manga layout. ACM Transactions on Graphics, 31(6):141:1– 141:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2022:EBS

[CCL^{+}22] Yadi Cao, Yunuo Chen, Minchen Li, Yin Yang, Xinxin Zhang, Mridul Aanjaneya, and Chenfanfu Jiang. An efficient B-spline Lagrangian/ Eulerian method for compressible flow, shock waves, and fracturing solids. ACM Transactions on Graphics, 41(5):169:1–169:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3519595.

Chu:2013:HQC

[CCLM13] Hung-Kuo Chu, Chia-Sheng Chang, Ruen-Rone Lee, and Niloy J. Mitra. Halftone QR codes. ACM Transactions on Graphics, 32(6): 217:1–217:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2005:AAH

[CCOST05] Doron Chen, Daniel Cohen-Or, Olga Sorkine, and Sivan Toledo. Algebraic analysis of high-pass quantization. ACM Transactions on Graphics, 24 (4):1259–1282, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Collet:2015:HQS

[CCS^{+}15] Alvaro Collet, Ming Chuang, Pat Sweeney, Don Gillett, Dennis Evseev, David Calabrese, Hugues Hoppe, Adam Kirk, and Steve Sullivan. High-quality streamable freeviewpoint video. ACM Transactions on Graphics, 34(4): 69:1–69:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2021:ERD

[CCS^{+}21] Marcel Campen, Ryan Capouellez, Hanxiao Shen, Leyi Zhu, Daniele Panozzo, and Denis Zorin. Efficient and robust discrete conformal equivalence with boundary. ACM Transactions on Graphics, 40(6): 261:1–261:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480557.

Chen:2009:SII

[CCT^{+}09] Tao Chen, Ming-Ming Cheng, Ping Tan, Ariel Shamir, and Shi-Min Hu. Sketch2Photo: Internet image montage. ACM Transactions on Graphics, 28(5):124:1–124:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:1993:SIS

[CCW93] Lin-Lin Chen, Shuo-Yan Chou, and Tony C. Woo. Separating and intersecting spherical polygons: Computing machinability on three- , four-, and five-axis numerically controlled machines. ACM Transactions on Graphics, 12(4):305–326, October 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        159732.html.

Chien:2016:BDH

[CCW16] Edward Chien, Renjie Chen, and Ofir Weber. Bounded distortion harmonic shape interpolation. ACM Transactions on Graphics, 35(4): 105:1–105:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2018:SRT

[CCWL18] Chen Cao, Menglei Chai, Oliver Woodford, and Linjie Luo. Stabilized real-time face tracking via a learned dynamic rigidity prior. ACM Transactions on Graphics, 37 (6):233:1–233:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cook:2005:WN

[CD05] Robert L. Cook and Tony DeRose. Wavelet noise. ACM Transactions on Graphics, 24 (3):803–811, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cutler:2002:PAA

[CDM^{+}02] Barbara Cutler, Julie Dorsey, Leonard McMillan, Matthias Müller, and Robert Jagnow. A procedural approach to authoring solid models. ACM Transactions on Graphics, 21 (3):302–311, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2014:RSE

[CDP^{+}14] Guojun Chen, Yue Dong, Pieter Peers, Jiawan Zhang, and Xin Tong. Reflectance scanning: estimating shading frame and BRDF with generalized linear light sources. ACM Transactions on Graphics, 33(4):117:1–117:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chaurasia:2013:DSL

[CDSHD13] Gaurav Chaurasia, Sylvain Duchene, Olga Sorkine-Hornung, and George Drettakis. Depth synthesis and local warps for plausible image-based navigation. ACM Transactions on Graphics, 32(3): 30:1–30:12, June 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2023:SPB

[CDY23] He Chen, Elie Diaz, and Cem Yuksel. Shortest path to boundary for self-intersecting meshes. ACM Transactions on Graphics, 42(4): 146:1–146:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592136.

Chen:2008:IPS

[CEW^{+}08] Guoning Chen, Gregory Esch, Peter Wonka, Pascal Müller, and Eugene Zhang. Interactive procedural street modeling. ACM Transactions on Graphics, 27(3):103:1– 103:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chang:2015:PBP

[CFL^{+}15] Huiwen Chang, Ohad Fried, Yiming Liu, Stephen Di-Verdi, and Adam Finkelstein. Palette-based photo recoloring. ACM Transactions on Graphics, 34(4): 139:1–139:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:OAC

[CFP^{+}21] Shiqi Chen, Huajun Feng, Dexin Pan, Zhihai Xu, Qi Li, and Yueting Chen. Optical aberrations correction in postprocessing using imaging simulation. ACM Transactions on Graphics, 40(5): 192:1–192:15, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3474088.

Christensen:2018:RAP

[CFS^{+}18] Per Christensen, Julian Fong, Jonathan Shade, Wayne Wooten, Brenden Schubert, Andrew Kensler, Stephen Friedman, Charlie Kilpatrick, Cliff Ramshaw, Marc Bannister, Brenton Rayner, Jonathan Brouillat, and Max Liani. RenderMan: an advanced path-tracing architecture for movie rendering. ACM Transactions on Graphics, 37 (3):30:1–30:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3182162.

Chen:2013:MFA

[CFW13] Zhili Chen, Renguo Feng, and Huamin Wang. Modeling friction and air effects between cloth and deformable bodies. ACM Transactions on Graphics, 32(4):88:1–88:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cheng:1989:PBS

[CG89] Fuhua Cheng and Ardeshir Goshtasby. A parallel Bspline surface fitting algorithm. ACM Transactions on Graphics, 8(1):41–50, January 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
214377.html.

Capell:2002:ISD

[CGC^{+}02] Steve Capell, Seth Green, Brian Curless, Tom Duchamp, and Zoran Popović. Interactive skeleton-driven dynamic deformations. ACM Transactions on Graphics, 21(3): 586–593, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chuang:2003:SMC

[CGC^{+}03] Yung-Yu Chuang, Dan B. Goldman, Brian Curless, David H. Salesin, and Richard Szeliski. Shadow matting and compositing. ACM Transactions on Graphics, 22(3): 494–500, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2009:BMS

[CGF09] Xiaobai Chen, Aleksey Golovinskiy, and Thomas Funkhouser. A benchmark for 3D mesh segmentation. ACM Transactions on Graphics, 28(3): 73:1–73:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cignoni:2004:ATE

[CGG^{+}04] Paolo Cignoni, Fabio Ganovelli, Enrico Gobbetti, Fabio Marton, Federico Ponchio, and Roberto Scopigno. Adaptive tetrapuzzles: efficient out-of-core construction and visualization of gigantic multiresolution polygonal models. ACM Transactions on Graphics, 23(3):796–803, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cordonnier:2017:ALC

[CGG^{+}17] Guillaume Cordonnier, Eric Galin, James Gain, Bedrich Benes, Eric Guérin, Adrien Peytavie, and Marie-Paule Cani. Authoring landscapes by combining ecosystem and terrain erosion simulation. ACM Transactions on Graphics, 36(4):134:1–134:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cole:2008:WDP

[CGL^{+}08] Forrester Cole, Aleksey Golovinskiy, Alex Limpaecher, Heather Stoddart Barros, Adam Finkelstein, Thomas Funkhouser, and Szymon Rusinkiewicz. Where do people draw lines? ACM Transactions on Graphics, 27(3):88:1–88:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chionh:1991:UMR

[CGM91] Eng-Wee Chionh, Ronald N.

Goldman, and James R. Miller. Using multivariate resultants to find the intersection of three quadric surfaces. ACM Transactions on Graphics, 10(4):378–400, October 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/

                                        toc/Abstracts/0730-0301/
                                        116917.html.

Chang:2011:FOB

[CGM11] Chia-Tche Chang, Bastien Gorissen, and Samuel Melchior. Fast oriented bounding box optimization on the rotation group SO(3, R). ACM Transactions on Graphics, 30(5):122:1–122:16, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chizhov:2022:PEO

[CGMS22] Vassillen Chizhov, Iliyan Georgiev, Karol Myszkowski, and Gurprit Singh. Perceptual error optimization for Monte Carlo rendering. ACM Transactions on Graphics, 41 (3):26:1–26:17, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3504002.

Choi:2021:NHL

[CGP^{+}21] Suyeon Choi, Manu Gopakumar, Yifan Peng, Jonghyun Kim, and Gordon Wetzstein. Neural 3D holography: learning accurate wave propagation models for 3D holographic virtual and augmented reality displays. ACM Transactions on Graphics, 40(6): 240:1–240:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480542.

Chen:2013:BBN

[CGW^{+}13] Jiating Chen, Xiaoyin Ge, Li-Yi Wei, Bin Wang, Yusu Wang, Huamin Wang, Yun Fei, Kang-Lai Qian, Jun-Hai Yong, and Wenping Wang. Bilateral blue noise sampling. ACM Transactions on Graphics, 32(6):216:1– 216:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chuang:2005:APS

[CGZ^{+}05] Yung-Yu Chuang, Dan B.

Goldman, Ke Colin Zheng, Brian Curless, David H. Salesin, and Richard Szeliski. Animating pictures with stochastic motion textures. ACM Transactions on Graphics, 24(3):853–860, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chong:2008:PBC

[CGZ08] Hamilton Y. Chong, Steven J.

Gortler, and Todd Zickler. A perception-based color space for illumination-invariant image processing. ACM Transactions on Graphics, 27(3): 61:1–61:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chuang:1989:LIA

[CH89] J. H. Chuang and C. M. Hoffmann. On local implicit approximation and its applications. ACM Transactions on Graphics, 8(4):298–324, October 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        77272.html.

Carr:2002:MAR

[CH02] Nathan A. Carr and John C. Hart. Meshed atlases for real-time procedural solid texturing. ACM Transactions on Graphics, 21(2): 106–131, April 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Carr:2004:PD

[CH04] Nathan A. Carr and John C. Hart. Painting detail. ACM Transactions on Graphics, 23 (3):845–852, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2005:PAL

[CH05] Jinxiang Chai and Jessica K. Hodgins. Performance animation from low-dimensional control signals. ACM Transactions on Graphics, 24(3): 686–696, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2007:CBM

[CH07] Jinxiang Chai and Jessica K.

Hodgins. Constraint-based motion optimization using a statistical dynamic model. ACM Transactions on Graphics, 26(3):8:1–8:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chosson:2014:BSR

[CH14] Sylvain M. Chosson and Roger D. Hersch. Beating shapes relying on Moiré level lines. ACM Transactions on Graphics, 34(1):9:1– 9:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cuypers:2012:RMD

[CHB^{+}12] Tom Cuypers, Tom Haber, Philippe Bekaert, Se Baek Oh, and Ramesh Raskar. Reflectance model for diffraction. ACM Transactions on Graphics, 31(5):122:1– 122:11, August 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cheng:1992:ESD

[Che92] Fuhua Cheng. Estimating subdivision depths for rational curves and surfaces. ACM Transactions on Graphics, 11(2):140–151, April 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        130829.html.

Chen:2013:AGP

[Che13] Baoquan Chen. Analyzing growing plants from 4D point cloud data. ACM Transactions on Graphics, 32(6): 157:1–157:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2010:CI

[CHM^{+}10] Hung-Kuo Chu, Wei-Hsin Hsu, Niloy J. Mitra, Daniel Cohen-Or, Tien-Tsin Wong, and Tong-Yee Lee. Camouflage images. ACM Transactions on Graphics, 29(4): 51:1–51:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cadik:2012:NMR

[CHM^{+}12] Martin Cadík, Robert Herzog, Rafal Mantiuk, Karol Myszkowski, and Hans-Peter Seidel. New measurements reveal weaknesses of image quality metrics in evaluating graphics artifacts. ACM Transactions on Graphics, 31 (6):147:1–147:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cooper:2007:ALR

[CHP07] Seth Cooper, Aaron Hertzmann, and Zoran Popović.

Active learning for real-time motion controllers. ACM Transactions on Graphics, 26 (3):5:1–5:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cook:2007:SSA

[CHPR07] Robert L. Cook, John Halstead, Maxwell Planck, and David Ryu. Stochastic simplification of aggregate detail. ACM Transactions on Graphics, 26(3):79:1–79:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2024:RTR

[CHTK24] Kinfung Chu, Jiawei Huang, Hidemasa Takana, and Yoshifumi Kitamura. Real-time reconstruction of fluid flow under unknown disturbance. ACM Transactions on Graphics, 43(1):4:1–4:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3624011.

Callenberg:2017:SDI

[CHWH17] Clara Callenberg, Felix Heide, Gordon Wetzstein, and Matthias B. Hullin. Snapshot difference imaging using correlation time-of-flight sensors. ACM Transactions on Graphics, 36 (6):220:1–220:??, November 2017. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Cho:2021:WSC

[CHY21] In-Young Cho, Yuchi Huo, and Sung-Eui Yoon. Weaklysupervised contrastive learning in path manifold for Monte Carlo image reconstruction. ACM Transactions on Graphics, 40(4): 38:1–38:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459876.

Cao:2014:DDE

[CHZ14] Chen Cao, Qiming Hou, and Kun Zhou. Displaced dynamic expression regression for real-time facial tracking and animation. ACM Transactions on Graphics, 33(4): 43:1–43:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chazelle:1984:TSC

[CI84] B. Chazelle and J. Incerpi. Triangulation and shapecomplexity. ACM Transactions on Graphics, 3(2): 135–152, April 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Castillo:1997:SCF

[CI97] Enrique Castillo and Andrés Iglesias. Some characterizations of families of surfaces using functional equations. ACM Transactions on Graphics, 16(3):296–318, July 1997. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
1997-16-3/p296-castillo/.

Chadwick:2011:AFS

[CJ11] Jeffrey N. Chadwick and Doug L. James. Animating fire with sound. ACM Transactions on Graphics, 30(4): 84:1–84:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Clarberg:2005:WIS

[CJAMJ05] Petrik Clarberg, Wojciech Jarosz, Tomas Akenine-Möller, and Henrik Wann Jensen. Wavelet importance sampling: efficiently evaluating products of complex functions. ACM Transactions on Graphics, 24(3): 1166–1175, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Crespo:2021:PSA

[CJM21] Miguel Crespo, Adrian Jarabo, and Adolfo Muñoz. Primaryspace adaptive control variates using piecewise-polynomial approximations. ACM Transactions on Graphics, 40(3): 25:1–25:15, July 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450627.

Choi:2017:HQH

[CJN^{+}17] Inchang Choi, Daniel S. Jeon, Giljoo Nam, Diego Gutierrez, and Min H. Kim. High-quality hyperspectral reconstruction using a spectral prior. ACM Transactions on Graphics, 36 (6):218:1–218:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cordonnier:2023:FTG

[CJP^{+}23] Guillaume Cordonnier, Guillaume Jouvet, Adrien Peytavie, Jean Braun, Marie-Paule Cani, Bedrich Benes, Eric Galin, Eric Guérin, and James Gain. Forming terrains by glacial erosion. ACM Transactions on Graphics, 42 (4):61:1–61:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592422.

Choi:2002:SRC

[CK02] Kwang-Jin Choi and Hyeong-Seok Ko. Stable but responsive cloth. ACM Transactions on Graphics, 21(3): 604–611, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chaudhuri:2010:DDS

[CK10] Siddhartha Chaudhuri and Vladlen Koltun. Data-driven suggestions for creativity support in 3D modeling. ACM Transactions on Graphics, 29 (6):183:1–183:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chuang:2011:IAG

[CK11] Ming Chuang and Michael Kazhdan. Interactive and anisotropic geometry processing using the screened Poisson equation. ACM Transactions on Graphics, 30(4): 57:1–57:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campbell:2014:LMF

[CK14a] Neill D. F. Campbell and Jan Kautz. Learning a manifold of fonts. ACM Transactions on Graphics, 33(4): 91:1–91:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2014:DSW

[CK14b] Marcel Campen and Leif Kobbelt. Dual strip weaving: interactive design of quad layouts using elastica strips. ACM Transactions on Graphics, 33(6):183:1– 183:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Choi:2020:DIF

[CK20] Jinsoo Choi and In So Kweon. Deep iterative frame interpolation for full-frame video stabilization. ACM Transactions on Graphics, 39(1): 4:1–4:9, February 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3363550.

Chaudhuri:2011:PRA

[CKGK11] Siddhartha Chaudhuri, Evangelos Kalogerakis, Leonidas Guibas, and Vladlen Koltun. Probabilistic reasoning for assembly-based 3D modeling. ACM Transactions on Graphics, 30(4):35:1–35:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2018:RTH

[CKH18] Yan-Pei Cao, Leif Kobbelt, and Shi-Min Hu. Realtime high-accuracy threedimensional reconstruction with consumer RGB-D cameras. ACM Transactions on Graphics, 37(5):171:1– 171:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3182157.

Chen:2015:WGB

[CKIW15] Zhili Chen, Byungmoon Kim, Daichi Ito, and Huamin Wang. Wetbrush: GPU-based 3D painting simulation at the bristle level. ACM Transactions on Graphics, 34(6): 200:1–200:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Coros:2011:LSS

[CKJ^{+}11] Stelian Coros, Andrej Karpathy, Ben Jones, Lionel Reveret, and Michiel van de Panne. Locomotion skills for simulated quadrupeds. ACM Transactions on Graphics, 30 (4):59:1–59:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:MCI

[CKMR^{+}21] Jingyu Chen, Victoria Kala, Alan Marquez-Razon, Elias Gueidon, David A. B. Hyde, and Joseph Teran. A momentum-conserving implicit material point method for surface tension with contact angles and spatial gradients. ACM Transactions on Graphics, 40(4): 111:1–111:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459874.

Chern:2016:SS

[CKP^{+}16] Albert Chern, Felix Knöppel, Ulrich Pinkall, Peter Schröder, and Steffen Weißmann. Schrödinger’s smoke. ACM Transactions on Graphics, 35(4): 77:1–77:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cho:2021:MRO

[CKP^{+}21] Kyungmin Cho, Chaelin Kim, Jungjin Park, Joonkyu Park, and Junyong Noh. Motion recommendation for online character control. ACM Transactions on Graphics, 40(6): 196:1–196:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480512.

Chern:2017:IFC

[CKPS17] Albert Chern, Felix Knöppel, Ulrich Pinkall, and Peter Schröder. Inside fluids: Clebsch maps for visualization and processing. ACM Transactions on Graphics, 36(4): 142:1–142:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chern:2018:SM

[CKPS18] Albert Chern, Felix Knöppel, Ulrich Pinkall, and Peter Schröder. Shape from metric. ACM Transactions on Graphics, 37(4):63:1–63:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chaitanya:2017:IRM

[CKS^{+}17] Chakravarty R. Alla Chaitanya, Anton S. Kaplanyan, Christoph Schied, Marco Salvi, Aaron Lefohn, Derek Nowrouzezahrai, and Timo Aila. Interactive reconstruction of Monte Carlo image sequences using a recurrent denoising autoencoder. ACM Transactions on Graphics, 36 (4):98:1–98:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chang:2018:TMD

[CKS18] Jen-Hao Rick Chang, B. V.

K. Vijaya Kumar, and Aswin C. Sankaranarayanan. Towards multifocal displays with dense focal stacks. ACM Transactions on Graphics, 37 (6):198:1–198:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2023:CWF

[CKSV23] Zhen Chen, Danny Kaufman, Mélina Skouras, and Etienne Vouga. Complex wrinkle field evolution. ACM Transactions on Graphics, 42 (4):72:1–72:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592397.

Chao:2023:CPA

[CKT^{+}23] Cheng-Kang Ted Chao, Jason Klein, Jianchao Tan, Jose Echevarria, and Yotam Gingold. ColorfulCurves: Palette-aware lightness control and color editing via sparse optimization. ACM Transactions on Graphics, 42 (4):98:1–98:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592405.

Chen:2008:SRR

[CKX^{+}08] Xuejin Chen, Sing Bing Kang, Ying-Qing Xu, Julie Dorsey, and Heung-Yeung Shum. Sketching reality: Realistic interpretation of architectural designs. ACM Transactions on Graphics, 27(2): 11:1–11:15, April 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cho:2009:FMD

[CL09] Sunghyun Cho and Seungyong Lee. Fast motion deblurring. ACM Transactions on Graphics, 28(5):145:1– 145:8, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chang:1996:IST

[CLC96] Meng-Chou Chang, Feipei Lai, and Wei-Chao Chen. Image shaping taking into account relativistic effects. ACM Transactions on Graphics, 15(4):265–300, October 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
234537.html.

Cao:2014:LHA

[CLC14] Ying Cao, Rynson W. H. Lau, and Antoni B. Chan. Look over here: attention-directing composition of manga elements. ACM Transactions on Graphics, 33(4): 94:1–94:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2020:MLS

[CLC^{+}20] Xiao-Song Chen, Chen-Feng Li, Geng-Chen Cao, Yun-Tao Jiang, and Shi-Min Hu. A moving least square reproducing kernel particle method for unified multiphase continuum simulation. ACM Transactions on Graphics, 39(6): 176:1–176:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417809.

Chen:2013:SRTb

[CLD^{+}13] Desai Chen, David I. W. Levin, Piotr Didyk, Pitchaya Sitthi-Amorn, and Wojciech Matusik. Spec2Fab: a reducer-tuner model for translating specifications to 3D prints. ACM Transactions on Graphics, 32(4):135:1– 135:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2018:FUB

[CLF^{+}18] Xuelin Chen, Honghua Li, Chi-Wing Fu, Hao Zhang, Daniel Cohen-Or, and Baoquan Chen. 3D fabrication with universal building blocks and pyramidal shells. ACM Transactions on Graphics, 37 (6):189:1–189:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cirio:2016:CSS

[CLG^{+}16] Gabriel Cirio, Dingzeyu Li, Eitan Grinspun, Miguel A. Otaduy, and Changxi Zheng. Crumpling sound synthesis. ACM Transactions on Graphics, 35(6):181:1–181:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cha:2020:EIV

[CLJ^{+}20] Seunghoon Cha, Jungjin Lee, Seunghwa Jeong, Younghui Kim, and Junyong Noh. Enhanced interactive 360^{∘} viewing via automatic guidance. ACM Transactions on Graphics, 39(5):154:1– 154:15, September 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3183794.

Chen:2020:CDS

[CLJL20] Honglin Chen, Hsueh-TI Derek Liu, Alec Jacobson, and David I. W. Levin. Chordal decomposition for spectral coarsening. ACM Transactions on Graphics, 39(6): 265:1–265:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417789.

Cho:2014:BTF

[CLKL14] Hojin Cho, Hyunjoon Lee, Henry Kang, and Seungyong Lee. Bilateral texture filtering. ACM Transactions on Graphics, 33(4): 128:1–128:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:DDF

[CLL^{+}21] Shu-Yu Chen, Feng-Lin Liu, Yu-Kun Lai, Paul L. Rosin, Chunpeng Li, Hongbo Fu, and Lin Gao. DeepFaceEditing: deep face generation and editing with disentangled geometry and appearance control. ACM Transactions on Graphics, 40(4): 90:1–90:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459760.

Chen:2022:UNB

[CLL^{+}22] Yunuo Chen, Minchen Li, Lei Lan, Hao Su, Yin Yang, and Chenfanfu Jiang. A unified Newton barrier method for multibody dynamics. ACM Transactions on Graphics, 41(4):66:1–66:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530076.

Ceylan:2013:DFM

[CLM^{+}13] Duygu Ceylan, Wilmot Li, Niloy J. Mitra, Maneesh Agrawala, and Mark Pauly. Designing and fabricating mechanical automata from mocap sequences. ACM Transactions on Graphics, 32(6): 186:1–186:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2017:DAN

[CLMK17] Desai Chen, David I. W. Levin, Wojciech Matusik, and Danny M. Kaufman. Dynamics-aware numerical coarsening for fabrication design. ACM Transactions on Graphics, 36(4): 84:1–84:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cirio:2014:YLS

[CLMMO14] Gabriel Cirio, Jorge Lopez-Moreno, David Miraut, and Miguel A. Otaduy. Yarnlevel simulation of woven cloth. ACM Transactions on Graphics, 33(6):207:1– 207:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chi:2008:SAI

[CLQW08] Ming-Te Chi, Tong-Yee Lee, Yingge Qu, and Tien-Tsin Wong. Self-animating images: illusory motion using repeated asymmetric patterns. ACM Transactions on Graphics, 27 (3):62:1–62:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cohen:1985:ADR

[CLS85] Elaine Cohen, Tom Lyche, and Larry L. Schumaker. Algorithms for degree-raising of splines. ACM Transactions on Graphics, 4(3):171–181, July 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Choi:2003:PBL

[CLS03] Min Gyu Choi, Jehee Lee, and Sung Yong Shin. Planning biped locomotion using motion capture data and probabilistic roadmaps. ACM Transactions on Graphics, 22 (2):182–203, April 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2015:HQH

[CLS^{+}15] Menglei Chai, Linjie Luo, Kalyan Sunkavalli, Nathan Carr, Sunil Hadap, and Kun Zhou. High-quality hair modeling from a single portrait photo. ACM Transactions on Graphics, 34(6):204:1– 204:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cholewiak:2017:CRC

[CLS^{+}17] Steven A. Cholewiak, Gordon D. Love, Pratul P. Srinivasan, Ren Ng, and Martin S. Banks. ChromaBlur: rendering chromatic eye aberration improves accommodation and realism. ACM Transactions on Graphics, 36(6): 210:1–210:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cherchi:2020:FRM

[CLSA20] Gianmarco Cherchi, Marco Livesu, Riccardo Scateni, and Marco Attene. Fast and robust mesh arrangements using floating-point arithmetic. ACM Transactions on Graphics, 39(6): 250:1–250:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417818.

Cui:2021:SSF

[CLSK21] Qiaodong Cui, Timothy Langlois, Pradeep Sen, and Theodore Kim. Spiralspectral fluid simulation. ACM Transactions on Graphics, 40(6):202:1–202:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480536.

Chen:2015:DDF

[CLSM15] Desai Chen, David I. W.

Levin, Shinjiro Sueda, and Wojciech Matusik. Datadriven finite elements for geometry and material design. ACM Transactions on Graphics, 34(4):74:1–74:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Christensen:1997:CGG

[CLSS97] Per H. Christensen, Dani Lischinski, Eric J. Stollnitz, and David H. Salesin. Clustering for glossy global illumination. ACM Transactions on Graphics, 16(1): 3–33, January 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1997-16-1/
                                        p3-christensen/.

Chen:2022:PMR

[CLT^{+}22] Qiang Chen, Tingsong Lu, Yang Tong, Guoliang Luo, Xiaogang Jin, and Zhigang Deng. A practical model for realistic butterfly flight simulation. ACM Transactions on Graphics, 41(3): 31:1–31:12, June 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3510459.

Chen:2014:ASM

[CLW^{+}14] Kang Chen, Yu-Kun Lai, Yu-Xin Wu, Ralph Martin, and Shi-Min Hu. Automatic semantic modeling of indoor scenes from low-quality RGB-D data using contextual information. ACM Transactions on Graphics, 33(6): 208:1–208:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chien:2016:BDP

[CLW16] Edward Chien, Zohar Levi, and Ofir Weber. Bounded distortion parametrization in the space of metrics. ACM Transactions on Graphics, 35(6):

215:1–215:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2022:SPI

[CLX^{+}22] Anpei Chen, Ruiyang Liu, Ling Xie, Zhang Chen, Hao Su, and Jingyi Yu. SofGAN: a portrait image generator with dynamic styling. ACM Transactions on Graphics, 41 (1):1:1–1:26, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3470848.

Cao:2018:CUP

[CLY18] Kaidi Cao, Jing Liao, and Lu Yuan. CariGANs: unpaired photo-to-caricature translation. ACM Transactions on Graphics, 37(6): 244:1–244:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2022:PIN

[CLZ^{+}22] Mengyu Chu, Lingjie Liu, Quan Zheng, Erik Franz, Hans-Peter Seidel, Christian Theobalt, and Rhaleb Zayer. Physics informed neural fields for smoke reconstruction with sparse data. ACM Transactions on Graphics, 41 (4):119:1–119:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530169.

Carlbom:1983:QAV

[CM83] I. Carlbom and J. Michener. Quantitative analysis of vector graphics system performance. ACM Transactions on Graphics, 2(1):57–88, January 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Correa:2010:DVN

[CM10] Carlos D. Correa and Kwan-Liu Ma. Dynamic video narratives. ACM Transactions on Graphics, 29(4): 88:1–88:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chentanez:2011:RTE

[CM11] Nuttapong Chentanez and Matthias Müller. Real-time Eulerian water simulation using a restricted tall cell grid. ACM Transactions on Graphics, 30(4):82:1–82:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Clarberg:2014:DSB

[CM14] Petrik Clarberg and Jacob Munkberg. Deep shading buffers on commodity GPUs. ACM Transactions on Graphics, 33(6):227:1– 227:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:DSS

[CM21] Weixuan Chen and Daniel McDuff. DeepMag: Sourcespecific change magnification using gradient ascent. ACM Transactions on Graphics, 40 (1):2:1–2:14, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3408865.

Chen:2017:FTD

[CML^{+}17] Weikai Chen, Yuexin Ma, Sylvain Lefebvre, Shiqing Xin, Jonàs Martínez, and Wenping Wang. Fabricable tile decors. ACM Transactions on Graphics, 36(6):175:1– 175:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chentanez:2015:FGF

[CMMK15] Nuttapong Chentanez, Matthias Müller, Miles Macklin, and Tae-Yong Kim. Fast grid-free surface tracking. ACM Transactions on Graphics, 34(4): 148:1–148:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Christensen:1995:ESA

[CMS95] Jon Christensen, Joe Marks, and Stuart Shieber. An empirical study of algorithms for point-feature label placement. ACM Transactions on Graphics, 14(3):203–232, July 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
212334.html.

Chen:2020:ECC

[CMSA20] Yi-Lu Chen, Jonathan Meier, Barbara Solenthaler, and Vinicius C. Azevedo. An extended cut-cell method for sub-grid liquids tracking with surface tension. ACM Transactions on Graphics, 39(6): 169:1–169:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417859.

Carlson:2004:RFA

[CMT04] Mark Carlson, Peter J. Mucha, and Greg Turk. Rigid fluid: animating the interplay between rigid bodies and fluid. ACM Transactions on Graphics, 23(3):377–384, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Coros:2012:DOA

[CMT^{+}12] Stelian Coros, Sebastian Martin, Bernhard Thomaszewski, Christian Schumacher, Robert Sumner, and Markus Gross. Deformable objects alive! ACM Transactions on Graphics, 31(4):69:1–69:9, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Canabal:2016:DKW

[CMT^{+}16] José A. Canabal, David Miraut, Nils Thuerey, Theodore Kim, Javier Portilla, and Miguel A. Otaduy. Dispersion kernels for water wave simulation. ACM Transactions on Graphics, 35(6): 202:1–202:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ceylan:2014:CSM

[CMZP14] Duygu Ceylan, Niloy J. Mitra, Youyi Zheng, and Mark Pauly. Coupled structurefrom-motion and 3D symmetry detection for urban facades. ACM Transactions on Graphics, 33(1):2:1– 2:15, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cossairt:2008:LFT

[CNR08] Oliver Cossairt, Shree Nayar, and Ravi Ramamoorthi. Light field transfer: global illumination between real and synthetic objects. ACM Transactions on Graphics, 27(3): 57:1–57:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2008:SBT

[CNX^{+}08] Xuejin Chen, Boris Neubert, Ying-Qing Xu, Oliver Deussen, and Sing Bing Kang. Sketch-based tree modeling using Markov random field. ACM Transactions on Graphics, 27(5): 109:1–109:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2022:SOM

[CNZ^{+}22] Xuwen Chen, Xingyu Ni, Bo Zhu, Bin Wang, and Baoquan Chen. Simulation and optimization of magnetoelastic thin shells. ACM Transactions on Graphics, 41(4):61:1–61:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530142.

Corman:2019:FCD

[CO19] Etienne Corman and Maks Ovsjanikov. Functional characterization of deformation fields. ACM Transactions on Graphics, 38(1):8:1–8:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3292480.

Cohen:1987:NLB

[Coh87] Elaine Cohen. A new local basis for designing with tensioned splines. ACM Transactions on Graphics, 6(2): 81–122, April 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
31337.html.

Cook:1986:SSC

[Coo86] Robert L. Cook. Stochastic sampling in computer graphics. ACM Transactions on Graphics, 5(1):51–72, January 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

8927.html. See remarks [Pav90, WP90]. Also in Tutorial: Computer Graphics: Image Synthesis, Computer Society Press, Washington, 1988, pp. 283–304.

Coros:2018:SDC

[Cor18] Stelian Coros. Session details:

Character animation. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ciccone:2019:TSO

[CÖS19] Loïc Ciccone, Cengiz Öztireli, and Robert W. Sumner. Tangent-space optimization for interactive animation control. ACM Transactions on Graphics, 38(4):101:1– 101:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cohen-Or:2006:CH

[COSG^{+}06] Daniel Cohen-Or, Olga Sorkine, Ran Gal, Tommer Leyvand, and Ying-Qing Xu. Color harmonization. ACM Transactions on Graphics, 25(3): 624–630, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cohen-Or:1998:TDD

[COSL98] Daniel Cohen-Or, Amira Solomovic, and David Levin. Three-dimensional distance field metamorphosis. ACM Transactions on Graphics, 17 (2):116–141, April 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/tog/1998-17-2/
                                        p116-cohen-or/.

Chang:2022:CFB

[CPAB22] Jumyung Chang, Ruben Partono, Vinicius C. Azevedo, and Christopher Batty. Curlflow: Boundary-respecting pointwise incompressible velocity interpolation for gridbased fluids. ACM Transactions on Graphics, 41(6): 243:1–243:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555498.

Cherchi:2022:IRM

[CPAL22] Gianmarco Cherchi, Fabio Pellacini, Marco Attene, and Marco Livesu. Interactive and robust mesh booleans. ACM Transactions on Graphics, 41 (6):248:1–248:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555460.

Chen:2007:RTE

[CPD07] Jiawen Chen, Sylvain Paris, and Frédo Durand. Realtime edge-aware image processing with the bilateral grid. ACM Transactions on Graphics, 26(3):103:1–103:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:CDD

[CPMK21] He Chen, Hyojoon Park, Kutay Macit, and Ladislav Kavan. Capturing detailed deformations of moving human bodies. ACM Transactions on Graphics, 40(4): 85:1–85:18, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459792.

Cignoni:2014:FAM

[CPMS14] Paolo Cignoni, Nico Pietroni, Luigi Malomo, and Roberto Scopigno. Field-aligned mesh joinery. ACM Transactions on Graphics, 33(1):11:1– 11:12, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cleary:2007:BFL

[CPPK07] Paul W. Cleary, Soon Hyoung Pyo, Mahesh Prakash, and Bon Ki Koo. Bubbling and frothing liquids. ACM Transactions on Graphics, 26 (3):97:1–97:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Crane:2011:STD

[CPS11] Keenan Crane, Ulrich Pinkall, and Peter Schröder. Spin transformations of discrete surfaces. ACM Transactions on Graphics, 30(4): 104:1–104:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Crane:2013:RFC

[CPS13] Keenan Crane, Ulrich Pinkall, and Peter Schröder. Robust fairing via conformal curvature flow. ACM Transactions on Graphics, 32(4): 61:1–61:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chern:2015:CCD

[CPS15] Albert Chern, Ulrich Pinkall, and Peter Schröder. Closeto-conformal deformations of volumes. ACM Transactions on Graphics, 34(4):56:1– 56:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:BAS

[CPSP21] Tian Chen, Julian Panetta, Max Schnaubelt, and Mark Pauly. Bistable auxetic surface structures. ACM Transactions on Graphics, 40 (4):39:1–39:9, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459940.

Chao:2010:SGM

[CPSS10] Isaac Chao, Ulrich Pinkall, Patrick Sanan, and Peter Schröder. A simple geometric model for elastic deformations. ACM Transactions on Graphics, 29(4): 38:1–38:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Charalambous:2023:GCC

[CPV^{+}23] Panayiotis Charalambous, Julien Pettre, Vassilis Vassiliades, Yiorgos Chrysanthou, and Nuria Pelechano. GREIL-Crowds: Crowd simulation with deep reinforcement learning and examples. ACM Transactions on Graphics, 42(4):137:1– 137:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592459.

Chu:2021:USC

[CPW21] Lei Chu, Hao Pan, and Wenping Wang. Unsupervised shape completion via deep prior in the neural tangent kernel perspective. ACM Transactions on Graphics, 40(3):32:1–32:17, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3459234.

Chen:2023:RLP

[CPW^{+}23] Zhen Chen, Zherong Pan, Kui Wu, Etienne Vouga, and Xifeng Gao. Robust low-poly meshing for general 3D models. ACM Transactions on Graphics, 42(4): 119:1–119:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592396.

Cheslack-Postava:2008:FRL

[CPWAP08] Ewen Cheslack-Postava, Rui Wang, Oskar Akerlund, and Fabio Pellacini. Fast, realistic lighting and material design using nonlinear cut approximation. ACM Transactions on Graphics, 27(5): 128:1–128:??, December 2008.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2022:THS

[CPY^{+}22] Xin Chen, Anqi Pang, Wei Yang, Peihao Wang, Lan Xu, and Jingyi Yu. TightCap: 3D human shape capture with clothing tightness field. ACM Transactions on Graphics, 41 (1):9:1–9:17, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478518.

Cirio:2018:MSS

[CQD^{+}18] Gabriel Cirio, Ante Qu, George Drettakis, Eitan Grinspun, and Changxi Zheng. Multi-scale simulation of nonlinear thin-shell sound with wave turbulence. ACM Transactions on Graphics, 37(4): 110:1–110:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2023:MSS

[CQS^{+}23] Rulin Chen, Pengyun Qiu, Peng Song, Bailin Deng, Ziqi Wang, and Ying He. Masonry Shell structures with discrete equivalence classes. ACM Transactions on Graphics, 42(4):115:1–115:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592095.

Carroll:2011:IDM

[CRA11] Robert Carroll, Ravi Ramamoorthi, and Maneesh Agrawala. Illumination decomposition for material recoloring with consistent interreflections. ACM Transactions on Graphics, 30(4): 43:1–43:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2023:ULR

[CRB23] Dongliang Cao, Paul Roetzer, and Florian Bernard. Unsupervised learning of robust spectral shape matching. ACM Transactions on Graphics, 42(4):132:1– 132:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592107.

Cui:2023:DIF

[CRCM23] Qiaodong Cui, Victor Rong, Desai Chen, and Wojciech Matusik. Dense, interlockingfree and scalable spectral packing of generic 3D objects. ACM Transactions on Graphics, 42(4):141:1– 141:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592126.

Chaitanya:2020:DSL

[CRG^{+}20] Chakravarty R. Alla Chaitanya, Nikunj Raghuvanshi, Keith W. Godin, Zechen Zhang, Derek Nowrouzezahrai, and John M. Snyder. Directional sources and listeners in interactive sound propagation using reciprocal wave field coding. ACM Transactions on Graphics, 39(4): 44:1–44:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392459.

Cao:2016:ISP

[CRS^{+}16] Chunxiao Cao, Zhong Ren, Carl Schissler, Dinesh Manocha, and Kun Zhou. Interactive sound propagation with bidirectional path tracing. ACM Transactions on Graphics, 35 (6):180:1–180:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cant:2000:TPM

[CS00] R. J. Cant and P. A. Shrubsole. Texture potential MIP mapping, a new highquality texture antialiasing algorithm. ACM Transactions on Graphics, 19(3):164–184, 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tog/2000-
19-3/p164-cant/p164-cant.

                                        pdf;   http://www.acm.org/
                                        pubs/citations/journals/
                                        tog/2000-19-3/p164-cant/.

Choi:2009:FSM

[CS09] Jaeil Choi and Andrzej Szymczak. Fitting solid meshes to animated surfaces using linear elasticity. ACM Transactions on Graphics, 28(1): 6:1–6:10, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cohen-Steiner:2004:VSA

[CSAD04] David Cohen-Steiner, Pierre Alliez, and Mathieu Desbrun. Variational shape approximation. ACM Transactions on Graphics, 23(3): 905–914, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:SHS

[CSAP21] Yu Ju (Edwin) Chen, Seung Heon Sheen, Uri M. Ascher, and Dinesh K. Pai. SIERE: a hybrid semi-implicit exponential integrator for efficiently simulating stiff deformable objects. ACM Transactions on Graphics, 40 (1):3:1–3:12, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3410527.

Corman:2017:FCIa

[CSBC^{+}17a] Etienne Corman, Justin Solomon, Mirela Ben-Chen, Leonidas Guibas, and Maks Ovsjanikov. Functional characterization of intrinsic and extrinsic geometry. ACM Transactions on Graphics, 36 (2):14:1–14:??, April 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Corman:2017:FCIb

[CSBC^{+}17b] Etienne Corman, Justin Solomon, Mirela Ben-Chen, Leonidas Guibas, and Maks Ovsjanikov. Functional characterization of intrinsic and extrinsic geometry. ACM Transactions on Graphics, 36 (4):59:1–59:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cole:2009:HWD

[CSD^{+}09] Forrester Cole, Kevin Sanik, Doug DeCarlo, Adam Finkelstein, Thomas Funkhouser, Szymon Rusinkiewicz, and Manish Singh. How well do line drawings depict shape? ACM Transactions on Graphics, 28(3):28:1–28:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Csebfalvi:2019:BTI

[Csé19] Balázs Csébfalvi. Beyond trilinear interpolation: higher quality for free. ACM Transactions on Graphics, 38(4): 56:1–56:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cohen:2003:WTI

[CSHD03] Michael F. Cohen, Jonathan Shade, Stefan Hiller, and Oliver Deussen. Wang Tiles for image and texture generation. ACM Transactions on Graphics, 22(3):287–294, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:MCP

[CSHD21] Jiong Chen, Florian Schäfer, Jin Huang, and Mathieu Desbrun. Multiscale Cholesky preconditioning for ill-conditioned problems. ACM Transactions on Graphics, 40 (4):81:1–81:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459851.

Callenberg:2021:LCS

[CSHH21] Clara Callenberg, Zheng Shi, Felix Heide, and Matthias B. Hullin. Low-cost SPAD sensing for non-line-of-sight tracking, material classification and depth imaging. ACM Transactions on Graphics, 40 (4):61:1–61:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459824.

Cui:2018:SLE

[CSK18] Qiaodong Cui, Pradeep Sen, and Theodore Kim. Scalable Laplacian eigenfluids. ACM Transactions on Graphics, 37 (4):87:1–87:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2022:AVA

[CSK^{+}22] Chen Cao, Tomas Simon, Jin Kyu Kim, Gabe Schwartz, Michael Zollhoefer, Shun-Suke Saito, Stephen Lombardi, Shih-En Wei, Danielle Belko, Shoou-I Yu, Yaser Sheikh, and Jason Saragih. Authentic volumetric avatars from a phone scan. ACM Transactions on Graphics, 41 (4):163:1–163:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530143.

Cheng:2022:EPG

[CSL^{+}22] Yingjie Cheng, Peng Song, Yukun Lu, Wen Jie Jeremy Chew, and Ligang Liu. Exact 3D path generation via 3D cam-linkage mechanisms. ACM Transactions on Graphics, 41(6): 225:1–225:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555431.

Chakravarthula:2023:TSN

[CSL^{+}23]  Praneeth  Chakravarthula,

Jipeng Sun, Xiao Li, Chenyang Lei, Gene Chou, Mario Bijelic, Johannes Froesch, Arka Majumdar, and Felix Heide. Thin on-sensor nanophotonic array cameras. ACM Transactions on Graphics, 42(6): 249:1–249:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618398.

Chen:2012:DPT

[CSN^{+}12] Ge Chen, Pedro V. Sander, Diego Nehab, Lei Yang, and Liang Hu. Depth-presorted triangle lists. ACM Transactions on Graphics, 31(6): 160:1–160:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2012:SPS

[CSPF12] Xiaobai Chen, Abulhair Saparov, Bill Pang, and Thomas Funkhouser. Schelling points on 3D surface meshes. ACM Transactions on Graphics, 31(4):29:1–29:12, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Criminisi:2010:GIV

[CSRP10] Antonio Criminisi, Toby Sharp, Carsten Rother, and Patrick Pérez. Geodesic image and video editing. ACM Transactions on Graphics, 29(5):134:1–134:15, October 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Christensen:1996:GIG

[CSS96] Per H. Christensen, Eric J.

Stollnitz, and David H. Salesin. Global illumination of glossy environments using wavelets and importance. ACM Transactions on Graphics, 15(1):37–71, January 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
225888.html;   http://www.
acm.org/pubs/toc/Abstracts/
0730-0301/226153.html.

Cheng:2021:STM

[CSSL21] Yingjie Cheng, Yucheng Sun, Peng Song, and Ligang Liu. Spatial-temporal motion control via composite camfollower mechanisms. ACM Transactions on Graphics, 40 (6):270:1–270:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480477.

Calabrese:2016:CSC

[CSTP16] Claudio Calabrese, Gabriele Salvati, Marco Tarini, and Fabio Pellacini. cSculpt: a system for collaborative sculpting. ACM Transactions on Graphics, 35(4): 91:1–91:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2018:PSE

[CSvRV18] Hsiao-Yu Chen, Arnav Sastry, Wim M. van Rees, and Etienne Vouga. Physical simulation of environmentally induced thin shell deformation. ACM Transactions on Graphics, 37(4):146:1– 146:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2016:AFA

[CSW^{+}16] Menglei Chai, Tianjia Shao, Hongzhi Wu, Yanlin Weng, and Kun Zhou. AutoHair: fully automatic hair modeling from a single image. ACM Transactions on Graphics, 35 (4):116:1–116:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2016:BMS

[CSZ16] Marcel Campen, Cláudio T. Silva, and Denis Zorin. Bijective maps from simplicial foliations. ACM Transactions on Graphics, 35(4):

74:1–74:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2020:SPA

[CSZZ20] Marcel Campen, Hanxiao Shen, Jiaran Zhou, and Denis Zorin. Seamless parametrization with arbitrary cones for arbitrary genus. ACM Transactions on Graphics, 39(1): 2:1–2:19, February 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3360511.

Cook:1982:RMC

[CT82] R. L. Cook and K. E. Torrance. A reflectance model for computer graphics. ACM Transactions on Graphics, 1 (1):7–24, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2005:MRT

[CT05] Nelson S.-H. Chu and Chiew-Lan Tai. MoXi: real-time ink dispersion in absorbent paper. ACM Transactions on Graphics, 24(3):504–511, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2017:DDS

[CT17] Mengyu Chu and Nils Thuerey.

Data-driven synthesis of smoke flows with CNN-based feature descriptors. ACM Transactions on Graphics, 36 (4):69:1–69:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cline:2005:ERP

[CTE05] David Cline, Justin Talbot, and Parris Egbert. Energy redistribution path tracing. ACM Transactions on Graphics, 24(3):1186–1195, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chakravarthula:2022:HFH

[CTFH22]  Praneeth  Chakravarthula,

Ethan Tseng, Henry Fuchs, and Felix Heide. Hogel-free holography. ACM Transactions on Graphics, 41(5): 178:1–178:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3516428.

Cao:2005:ESD

[CTFP05] Yong Cao, Wen C. Tien, Petros Faloutsos, and Frédéric Pighin. Expressive speechdriven facial animation. ACM Transactions on Graphics, 24 (4):1283–1302, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2022:NDC

[CTFZ22] Zhiqin Chen, Andrea Tagliasacchi, Thomas Funkhouser, and Hao Zhang. Neural dual contouring. ACM Transactions on Graphics, 41 (4):104:1–104:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530108.

Clarberg:2014:AAM

[CTH^{+}14] Petrik Clarberg, Robert Toth, Jon Hasselgren, Jim Nilsson, and Tomas Akenine-Möller. AMFS: adaptive multi-frequency shading for future graphics processors. ACM Transactions on Graphics, 33(4):141:1–141:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:CCO

[CTL^{+}21] Kang Chen, Zhipeng Tan, Jin Lei, Song-Hai Zhang, Yuan-Chen Guo, Weidong Zhang, and Shi-Min Hu. ChoreoMaster: choreographyoriented music-driven dance synthesis. ACM Transactions on Graphics, 40(4): 145:1–145:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459932.

Clarberg:2013:SBD

[CTM13] Petrik Clarberg, Robert Toth, and Jacob Munkberg. A sortbased deferred shading architecture for decoupled sampling. ACM Transactions on Graphics, 32(4):141:1– 141:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Carranza:2003:FVV

[CTMS03] Joel Carranza, Christian Theobalt, Marcus A. Magnor, and Hans-Peter Seidel. Freeviewpoint video of human actors. ACM Transactions on Graphics, 22(3):569–577, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Coros:2013:CDM

[CTN^{+}13] Stelian Coros, Bernhard Thomaszewski, Gioacchino Noris, Shinjiro Sueda, Moira Forberg, Robert W. Sumner, Wojciech Matusik, and Bernd Bickel. Computational design of mechanical characters. ACM Transactions on Graphics, 32(4):83:1–83:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chakravarthula:2020:LHL

[CTS^{+}20]  Praneeth  Chakravarthula,

Ethan Tseng, Tarun Srivastava, Henry Fuchs, and Felix Heide. Learned hardwarein-the-loop phase retrieval for holographic near-eye displays. ACM Transactions on Graphics, 39(6):186:1– 186:18, November 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417846.

Chu:2021:LMC

[CTS^{+}21] Mengyu Chu, Nils Thuerey, Hans-Peter Seidel, Christian Theobalt, and Rhaleb Zayer. Learning meaningful controls for fluids. ACM Transactions on Graphics, 40(4): 100:1–100:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459845.

Clegg:2015:AHD

[CTTL15] Alexander Clegg, Jie Tan, Greg Turk, and C. Karen Liu. Animating human dressing. ACM Transactions on Graphics, 34(4):116:1– 116:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2004:STF

[CTW^{+}04] Yanyun Chen, Xin Tong, Jiaping Wang, Stephen Lin, Baining Guo, and Heung-Yeung Shum. Shell texture functions. ACM Transactions on Graphics, 23(3): 343–353, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2009:NBI

[CTW09] Jia Chen, Chi-Keung Tang, and Jue Wang. Noise brush: interactive high quality image-noise separation. ACM Transactions on Graphics, 28(5):146:1–146:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Custers:2020:SDF

[CV20] Bram Custers and Amir Vaxman. Subdivision directional fields. ACM Transactions on Graphics, 39(2): 11:1–11:20, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3375659.

Chen:2015:BDH

[CW15] Renjie Chen and Ofir Weber. Bounded distortion harmonic mappings in the plane. ACM Transactions on Graphics, 34 (4):73:1–73:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2017:GAL

[CW17] Renjie Chen and Ofir Weber. GPU-accelerated locally injective shape deformation. ACM Transactions on Graphics, 36 (6):214:1–214:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2011:NRC

[CWC11] Hsiang-Ting Chen, Li-Yi Wei, and Chun-Fa Chang. Nonlinear revision control for images. ACM Transactions on Graphics, 30(4):105:1–105:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2020:LFE

[CWK^{+}20] Wenzheng Chen, Fangyin Wei, Kiriakos N. Kutulakos, Szymon Rusinkiewicz, and Felix Heide. Learned feature embeddings for non-line-of-sight imaging and recognition. ACM Transactions on Graphics, 39(6):230:1– 230:18, November 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417825.

Chen:2013:PSI

[CWKBC13] Renjie Chen, Ofir Weber, Daniel Keren, and Mirela Ben-Chen. Planar shape interpolation with bounded distortion. ACM Transactions on Graphics, 32(4): 108:1–108:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cho:2012:VDH

[CWL12] Sunghyun Cho, Jue Wang, and Seungyong Lee. Video deblurring for hand-held cameras using patch-based synthesis. ACM Transactions on Graphics, 31(4): 64:1–64:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2022:TZS

[CWL22] Zhaoxi Chen, Guangcong Wang, and Ziwei Liu. Text2Light: Zero-shot text-driven HDR panorama generation. ACM Transactions on Graphics, 41 (6):195:1–195:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555447.

Cao:2013:SRR

[CWLZ13] Chen Cao, Yanlin Weng, Stephen Lin, and Kun Zhou. 3D shape regression for realtime facial animation. ACM Transactions on Graphics, 32 (4):41:1–41:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2003:RMB

[cWP03] Jia chi Wu and Zoran Popović. Realistic modeling of bird flight animations. ACM Transactions on Graphics, 22 (3):888–895, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2010:TAB

[cWP10] Jia chi Wu and Zoran Popović. Terrain-adaptive bipedal locomotion control. ACM Transactions on Graphics, 29(4):72:1–72:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2022:CDH

[CWSB22] Rulin Chen, Ziqi Wang, Peng Song, and Bernd Bickel. Computational design of high-level interlocking puzzles. ACM Transactions on Graphics, 41 (4):150:1–150:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530071.

Clausen:2013:SLS

[CWSO13] Pascal Clausen, Martin Wicke, Jonathan R. Shewchuk, and James F. O’Brien. Simulating liquids and solidliquid interactions with Lagrangian meshes. ACM Transactions on Graphics, 32(2): 17:1–17:15, April 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cui:2017:TSV

[CWTW17] Zhaopeng Cui, Oliver Wang, Ping Tan, and Jue Wang. Time slice video synthesis by robust video alignment. ACM Transactions on Graphics, 36 (4):131:1–131:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2012:SVH

[CWW^{+}12] Menglei Chai, Lvdi Wang, Yanlin Weng, Yizhou Yu, Baining Guo, and Kun Zhou. Single-view hair modeling for portrait manipulation. ACM Transactions on Graphics, 31 (4):116:1–116:8, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2013:DHM

[CWW^{+}13a] Menglei Chai, Lvdi Wang, Yanlin Weng, Xiaogang Jin, and Kun Zhou. Dynamic hair manipulation in images and videos. ACM Transactions on Graphics, 32(4): 75:1–75:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Crane:2013:GHN

[CWW13b] Keenan Crane, Clarisse Weischedel, and Max Wardetzky. Geodesics in heat: a new approach to computing distance based on heat flow. ACM Transactions on Graphics, 32(5):152:1– 152:11, September 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2016:RTF

[CWW^{+}16] Chen Cao, Hongzhi Wu, Yanlin Weng, Tianjia Shao, and Kun Zhou. Real-time facial animation with image-based dynamic avatars. ACM Transactions on Graphics, 35(4): 126:1–126:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chandran:2021:RSC

[CWZ^{+}21a] Prashanth Chandran, Sebastian Winberg, Gaspard Zoss, Jérémy Riviere, Markus Gross, Paulo Gotardo, and Derek Bradley. Rendering with style: combining traditional and neural approaches for high-quality face rendering. ACM Transactions on Graphics, 40(6): 223:1–223:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480509.

Chen:2021:MSN

[CWZ^{+}21b] Kang Chen, Yupan Wang, Song-Hai Zhang, Sen-Zhe Xu, Weidong Zhang, and Shi-Min Hu. MoCap-solver: a neural solver for optical motion capture data. ACM Transactions on Graphics, 40 (4):84:1–84:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459681.

Chen:2002:MRR

[CXGS02] Yanyun Chen, Yingqing Xu, Baining Guo, and Heung-Yeung Shum. Modeling and rendering of realistic feathers.

ACM Transactions on Graphics, 21(3):630–636, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2005:VSW

[CXW^{+}05] Yanyun Chen, Lin Xia, Tien-Tsin Wong, Xin Tong, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Visual simulation of weathering by γ-ton tracing. ACM Transactions on Graphics, 24(3): 1127–1133, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cao:2023:SAR

[CXW^{+}23a] Hezhi Cao, Xi Xia, Guan Wu, Ruizhen Hu, and Ligang Liu. ScanBot: Autonomous reconstruction via deep reinforcement learning. ACM Transactions on Graphics, 42(4): 157:1–157:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592113.

Chen:2023:DFL

[CXW^{+}23b] Anpei Chen, Zexiang Xu, Xinyue Wei, Siyu Tang, Hao Su, and Andreas Geiger. Dictionary fields: Learning a neural basis decomposition. ACM Transactions on Graphics, 42(4): 156:1–156:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592135.

Chen:2015:MDA

[CXY^{+}15] Kang Chen, Kun Xu, Yizhou Yu, Tian-Yi Wang, and Shi-Min Hu. Magic decorator: automatic material suggestion for indoor digital scenes. ACM Transactions on Graphics, 34 (6):232:1–232:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2014:PIA

[CYFW14] Zhili Chen, Miaojun Yao, Renguo Feng, and Huamin Wang. Physics-inspired adaptive fracture refinement. ACM Transactions on Graphics, 33 (4):113:1–113:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Clegg:2018:LDS

[CYT^{+}18] Alexander Clegg, Wenhao Yu, Jie Tan, C. Karen Liu, and Greg Turk. Learning to dress: synthesizing human dressing motion via deep reinforcement learning. ACM Transactions on Graphics, 37(6): 179:1–179:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chang:2016:ATP

[CYW^{+}16] Huiwen Chang, Fisher Yu, Jue Wang, Douglas Ashley, and Adam Finkelstein. Automatic triage for a photo series. ACM Transactions on Graphics, 35 (4):148:1–148:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chang:2011:GRD

[CZ11] Will Chang and Matthias Zwicker. Global registration of dynamic range scans for articulated model reconstruction. ACM Transactions on Graphics, 30(3):26:1– 26:15, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Campen:2017:SMF

[CZ17] Marcel Campen and Denis Zorin. Similarity maps and field-guided T-splines: a perfect couple. ACM Transactions on Graphics, 36(4): 91:1–91:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2021:NMC

[CZ21] Zhiqin Chen and Hao Zhang. Neural marching cubes. ACM Transactions on Graphics, 40 (6):251:1–251:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480518.

Capouellez:2023:MOP

[CZ23] Ryan Capouellez and Denis Zorin. Metric optimization in Penner coordinates. ACM Transactions on Graphics, 42(6): 234:1–234:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618394.

Chen:2023:CAR

[CZB23] Huanyu Chen, Danyong Zhao, and Jernej Barbic. Capturing animation-ready isotropic materials using systematic poking. ACM Transactions on Graphics, 42(6): 223:1–223:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618406.

Chia:2011:SCI

[CZG^{+}11] Alex Yong-Sang Chia, Shaojie Zhuo, Raj Kumar Gupta, Yu-Wing Tai, Siu-Yeung Cho, Ping Tan, and Stephen Lin. Semantic colorization with Internet images. ACM Transactions on Graphics, 30(6): 156:1–156:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chadwick:2012:PAN

[CZJ12] Jeffrey N. Chadwick, Changxi Zheng, and Doug L. James.

Precomputed acceleration noise for improved rigidbody sound. ACM Transactions on Graphics, 31(4): 103:1–103:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cheng:2014:IVG

[CZL^{+}14] Ming-Ming Cheng, Shuai Zheng, Wen-Yan Lin, Vibhav Vineet, Paul Sturgess, Nigel Crook, Niloy J. Mitra, and Philip Torr. ImageSpirit: Verbal guided image parsing. ACM Transactions on Graphics, 34(1):3:1–3:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2015:GMD

[CZL^{+}15a] Xiaowu Chen, Bin Zhou, Feixiang Lu, Lin Wang, Lang Bi, and Ping Tan. Garment modeling with a depth camera. ACM Transactions on Graphics, 34(6):203:1– 203:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2015:DDP

[CZL^{+}15b] Xuelin Chen, Hao Zhang, Jinjie Lin, Ruizhen Hu, Lin Lu, Qixing Huang, Bedrich Benes, Daniel Cohen-Or, and Baoquan Chen. Dapper: decompose-and-pack for 3D printing. ACM Transactions on Graphics, 34(6):213:1– 213:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cheng:2010:RFA

[CZM^{+}10] Ming-Ming Cheng, Fang-Lue Zhang, Niloy J. Mitra, Xiaolei Huang, and Shi-Min Hu. RepFinder: finding approximately repeated scene elements for image editing. ACM Transactions on Graphics, 29 (4):83:1–83:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chermain:2023:ODC

[CZM^{+}23] Xavier Chermain, Cédric Zanni, Jonàs Martínez, Pierre-Alexandre Hugron, and Sylvain Lefebvre. Orientable dense cyclic infill for anisotropic appearance fabrication. ACM Transactions on Graphics, 42 (4):68:1–68:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592412.

Cossairt:2010:DCP

[CZN10] Oliver Cossairt, Changyin Zhou, and Shree Nayar. Diffusion coded photography for extended depth of field. ACM Transactions on Graphics, 29 (4):31:1–31:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2013:SEE

[CZS^{+}13] Tao Chen, Zhe Zhu, Ariel Shamir, Shi-Min Hu, and Daniel Cohen-Or. 3-Sweep: extracting editable objects from a single photo. ACM Transactions on Graphics, 32 (6):195:1–195:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2016:SFD

[CZX^{+}16] Weikai Chen, Xiaolong Zhang, Shiqing Xin, Yang Xia, Sylvain Lefebvre, and Wenping Wang. Synthesis of filigrees for digital fabrication. ACM Transactions on Graphics, 35 (4):98:1–98:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Cheng:2023:TDA

[CZXL23] Jin-San Cheng, Bingwei Zhang, Yikun Xiao, and Ming Li. Topology driven approximation to rational surface-surface intersection via interval algebraic topology analysis. ACM Transactions on Graphics, 42(4): 38:1–38:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592452.

Chen:2014:ANM

[CZXZ14] Xiang Chen, Changxi Zheng, Weiwei Xu, and Kun Zhou.

An asymptotic numerical method for inverse elastic shape design. ACM Transactions on Graphics, 33(4): 95:1–95:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2017:SCPa

[CZY17a] Jieyu Chu, Nafees Bin Zafar, and Xubo Yang. A Schur complement preconditioner for scalable parallel fluid simulation. ACM Transactions on Graphics, 36(4): 139:1–139:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chu:2017:SCPb

[CZY17b] Jieyu Chu, Nafees Bin Zafar, and Xubo Yang. A Schur complement preconditioner for scalable parallel fluid simulation. ACM Transactions on Graphics, 36(5): 163:1–163:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chai:2014:RMI

[CZZ14] Menglei Chai, Changxi Zheng, and Kun Zhou. A reduced model for interactive hairs. ACM Transactions on Graphics, 33(4):124:1–124:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Chen:2012:MPE

[CZZT12] Xiaowu Chen, Dongqing Zou, Qinping Zhao, and Ping Tan. Manifold preserving edit propagation. ACM Transactions on Graphics, 31(6):132:1– 132:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Davis:2018:VRB

[DA18] Abe Davis and Maneesh Agrawala. Visual rhythm and beat. ACM Transactions on Graphics, 37(4): 122:1–122:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Diazzi:2021:CPM

[DA21] Lorenzo Diazzi and Marco Attene. Convex polyhedral meshing for robust solid modeling. ACM Transactions on Graphics, 40(6): 259:1–259:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480564.

Demir:2015:CSS

[DAB15] Ilke Demir, Daniel G. Aliaga, and Bedrich Benes. Coupled segmentation and similarity detection for architectural models. ACM Transactions on Graphics, 34(4): 104:1–104:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deschaintre:2018:SIS

[DAD^{+}18] Valentin Deschaintre, Miika Aittala, Fredo Durand, George Drettakis, and Adrien Bousseau. Single-image SVBRDF capture with a rendering-aware deep network. ACM Transactions on Graphics, 37(4):128:1– 128:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deAguiar:2008:PCS

[dAST^{+}08] Edilson de Aguiar, Carsten Stoll, Christian Theobalt, Naveed Ahmed, Hans-Peter Seidel, and Sebastian Thrun. Performance capture from sparse multi-view video. ACM Transactions on Graphics, 27 (3):98:1–98:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deAguiar:2010:SSR

[dASTH10] Edilson de Aguiar, Leonid Sigal, Adrien Treuille, and Jessica K. Hodgins. Stable spaces for real-time clothing. ACM Transactions on Graphics, 29 (4):106:1–106:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Daviet:2020:SSF

[Dav20] Gilles Daviet. Simple and scalable frictional contacts for thin nodal objects. ACM Transactions on Graphics, 39 (4):61:1–61:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392439.

Day:1990:IAF

[Day90] A. M. Day. The implementation of an algorithm to find the convex hull of a set of three-dimensional points. ACM Transactions on Graphics, 9(1):105–132, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77640.html.

DeRose:1988:GCS

[DB88] Tony D. DeRose and Brian A. Barsky. Geometric continuity, shape parameters, and geometric constructions for Catmull-Rom splines. ACM Transactions on Graphics, 7 (1):1–41, January 1988. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
42265.html.

Dvoroznak:2017:EBE

[DBB^{+}17] Marek Dvoroznák, Pierre Bénard, Pascal Barla, Oliver Wang, and Daniel Sýkora. Example-based expressive animation of 2D rigid bodies. ACM Transactions on Graphics, 36(4):127:1–127:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Daviet:2016:SIM

[DBD16] Gilles Daviet and Florence Bertails-Descoubes. A semiimplicit material point method for the continuum simulation of granular materials. ACM Transactions on Graphics, 35 (4):102:1–102:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Daviet:2011:HIS

[DBDB11] Gilles Daviet, Florence Bertails-Descoubes, and Laurence Boissieux. A hybrid iterative solver for robustly capturing Coulomb friction in hair dynamics. ACM Transactions on Graphics, 30(6): 139:1–139:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2006:SSQ

[DBG^{+}06] Shen Dong, Peer-Timo Bremer, Michael Garland, Valerio Pascucci, and John C. Hart. Spectral surface quadrangulation. ACM Transactions on Graphics, 25(3): 1057–1066, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Da:2014:MMB

[DBG14] Fang Da, Christopher Batty, and Eitan Grinspun. Multimaterial mesh-based surface tracking. ACM Transactions on Graphics, 33(4): 112:1–112:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Diamanti:2015:SCI

[DBP^{+}15] Olga Diamanti, Connelly Barnes, Sylvain Paris, Eli Shechtman, and Olga Sorkine-Hornung. Synthesis of complex image appearance from limited exemplars. ACM Transactions on Graphics, 34 (2):22:1–22:??, February 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Da:2015:DBS

[DBWG15] Fang Da, Christopher Batty, Chris Wojtan, and Eitan Grinspun. Double bubbles sans toil and trouble: discrete circulation-preserving vortex sheets for soap films and foams. ACM Transactions on Graphics, 34(4): 149:1–149:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duinkharjav:2022:IFI

[DCB^{+}22] Budmonde Duinkharjav, Praneeth Chakravarthula, Rachel Brown, Anjul Patney, and Qi Sun. Image features influence reaction time: a learned probabilistic perceptual model for saccade latency. ACM Transactions on Graphics, 41(4):144:1– 144:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530055.

Davis:2015:ISM

[DCD15] Abe Davis, Justin G. Chen, and Frédo Durand. Imagespace modal bases for plausible manipulation of objects in video. ACM Transactions on Graphics, 34(6): 239:1–239:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Drori:2003:FBI

[DCOY03] Iddo Drori, Daniel Cohen-Or, and Hezy Yeshurun. Fragment-based image completion. ACM Transactions on Graphics, 22(3):303–312, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DiRenzo:2014:ALS

[DCP14a] Francesco Di Renzo, Claudio Calabrese, and Fabio Pellacini. AppIm: linear spaces for image-based appearance editing. ACM Transactions on Graphics, 33(6):194:1– 194:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2014:AMR

[DCP^{+}14b] Yue Dong, Guojun Chen, Pieter Peers, Jiawan Zhang, and Xin Tong. Appearancefrom-motion: recovering spatially varying surface reflectance under unknown lighting. ACM Transactions on Graphics, 33(6):193:1– 193:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duinkharjav:2022:CPG

[DCT^{+}22] Budmonde Duinkharjav, Kenneth Chen, Abhishek Tyagi, Jiayi He, Yuhao Zhu, and Qi Sun. Color-perception-guided display power reduction for virtual reality. ACM Transactions on Graphics, 41 (6):210:1–210:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555473.

Duguet:2002:REV

[DD02a] Florent Duguet and George Drettakis. Robust epsilon visibility. ACM Transactions on Graphics, 21(3):567–575, July 2002. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Durand:2002:FBF

[DD02b] Frédo Durand and Julie Dorsey. Fast bilateral filtering for the display of highdynamic-range images. ACM Transactions on Graphics, 21 (3):257–266, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Du:2014:IVQ

[DDD^{+}14] Song-Pei Du, Piotr Didyk, Frédo Durand, Shi-Min Hu, and Wojciech Matusik. Improving visual quality of view transitions in automultiscopic displays. ACM Transactions on Graphics, 33(6):192:1– 192:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dou:2017:MRT

[DDF^{+}17] Mingsong Dou, Philip Davidson, Sean Ryan Fanello, Sameh Khamis, Adarsh Kowdle, Christoph Rhemann, Vladimir Tankovich, and Shahram Izadi. Motion2fusion: real-time volumetric performance capture. ACM Transactions on Graphics, 36(6): 246:1–246:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Durand:1999:FAH

[DDP99] Frédo Durand, George Drettakis, and Claude Puech. Fast and accurate hierarchical radiosity using global visibility. ACM Transactions on Graphics, 18(2): 128–170, April 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/tog/1999-18-2/
                                        p128-durand/.

Durand:2002:VC

[DDP02] Frédo Durand, George Drettakis, and Claude Puech. The 3D visibility complex. ACM Transactions on Graphics, 21 (2):176–206, April 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Decoret:2003:BCE

[DDSD03] Xavier Décoret, Frédo Durand, François X. Sillion, and Julie Dorsey. Billboard clouds for extreme model simplification. ACM Transactions on Graphics, 22(3):689–696, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2015:MBE

[DDTP15] Bo Dong, Yue Dong, Xin Tong, and Pieter Peers. Measurement-based editing of diffuse albedo with consistent interreflections. ACM Transactions on Graphics, 34(4):

112:1–112:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dinerstein:2005:FML

[DE05] Jonathan Dinerstein and Parris K. Egbert. Fast multi-level adaptation for interactive autonomous characters. ACM Transactions on Graphics, 24 (2):262–288, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deering:2005:PAM

[Dee05] Michael F. Deering. A photon accurate model of the human eye. ACM Transactions on Graphics, 24(3):649–658, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dobkin:1996:CDA

[DEM96] David P. Dobkin, David Eppstein, and Don P. Mitchell. Computing the discrepancy with applications to supersampling patterns. ACM Transactions on Graphics, 15 (4):354–376, October 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
234536.html.

DeRose:1988:CBS

[DeR88] Tony D. DeRose. Composing Bézier simplexes. ACM Transactions on Graphics, 7 (3):198–221, July 1988. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
44482.html.

Didyk:2010:ADR

[DER^{+}10] Piotr Didyk, Elmar Eisemann, Tobias Ritschel, Karol Myszkowski, and Hans-Peter Seidel. Apparent display resolution enhancement for moving images. ACM Transactions on Graphics, 29(4): 113:1–113:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeFloriani:1988:HBM

[DF88] Leila De Floriani and Bianca Falcidieno. A hierarchical boundary model for solid object representation. ACM Transactions on Graphics, 7(1):42–60, January 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
46164.html.

Doraiswamy:2015:TBC

[DFL^{+}15] Harish Doraiswamy, Nivan Ferreira, Marcos Lage, Huy Vo, Luc Wilson, Heidi Werner, Muchan Park, and Cláudio Silva. Topology-based catalogue exploration framework for identifying viewenhanced tower designs. ACM Transactions on Graphics, 34 (6):230:1–230:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duce:1988:FSS

[DFM88] D. A. Duce, E. V. C.

Fielding, and L. S. Marshall. Formal specification of a small example based on GKS. ACM Transactions on Graphics, 7(3):180–197, July 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
44481.html.

Doyle:2013:HUF

[DFM13] Michael J. Doyle, Colin Fowler, and Michael Manzke. A hardware unit for fast SAHoptimised BVH construction. ACM Transactions on Graphics, 32(4):139:1–139:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deFloriani:1995:HTM

[dFP95] Leila de Floriani and Enrico Puppo. Hierarchical triangulation for multiresolution surface description. ACM Transactions on Graphics, 14(4):363–411, October 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/

                                        toc/Abstracts/0730-0301/
                                        225297.html.

DeCarlo:2003:SCC

[DFRS03] Doug DeCarlo, Adam Finkelstein, Szymon Rusinkiewicz, and Anthony Santella. Suggestive contours for conveying shape. ACM Transactions on Graphics, 22(3):848–855, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeGomensoroMalheiros:2020:LNC

[DFW20] Marcelo De Gomensoro Malheiros, Henrique Fensterseifer, and Marcelo Walter. The leopard never changes its spots: realistic pigmentation pattern formation by coupling tissue growth with reaction– diffusion. ACM Transactions on Graphics, 39(4): 63:1–63:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392478.

Dong:2019:RSM

[DFYL19] Zhi-Chao Dong, Xiao-Ming Fu, Zeshi Yang, and Ligang Liu. Redirected smooth mappings for multiuser real walking in virtual reality. ACM Transactions on Graphics, 38(5):149:1–149:??, October 2019. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.

acm.org/ft_gateway.cfm?
id=3345554.

Dong:2017:SAM

[DFZ^{+}17] Zhi-Chao Dong, Xiao-Ming Fu, Chi Zhang, Kang Wu, and Ligang Liu. Smooth assembled mappings for large-scale real walking. ACM Transactions on Graphics, 36(6): 211:1–211:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deGoes:2013:ESM

[dGAOD13] Fernando de Goes, Pierre Alliez, Houman Owhadi, and Mathieu Desbrun. On the equilibrium of simplicial masonry structures. ACM Transactions on Graphics, 32(4): 93:1–93:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deGoes:2012:BNT

[dGBOD12] Fernando de Goes, Katherine Breeden, Victor Ostromoukhov, and Mathieu Desbrun. Blue noise through optimal transport. ACM Transactions on Graphics, 31(6): 171:1–171:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deGoes:2016:SEC

[dGDMD16] Fernando de Goes, Mathieu Desbrun, Mark Meyer, and Tony DeRose. Subdivision exterior calculus for geometry processing. ACM Transactions on Graphics, 35(4): 133:1–133:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Damberg:2016:HBH

[DGH16] Gerwin Damberg, James Gregson, and Wolfgang Heidrich. High brightness HDR projection using dynamic freeform lensing. ACM Transactions on Graphics, 35(3): 24:1–24:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeRose:1993:FCA

[DGHM93] Tony D. DeRose, Ronald N. Goldman, Hans Hagen, and Stephen Mann. Functional composition algorithms via blossoming. ACM Transactions on Graphics, 12(2): 113–135, April 1993. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
151290.html.

deGoes:2014:WTG

[dGMMD14] Fernando de Goes, Pooran Memari, Patrick Mullen, and Mathieu Desbrun. Weighted triangulations for geometry processing. ACM Transactions on Graphics, 33(3): 28:1–28:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deschaintre:2023:VLF

[DGVG^{+}23] Valentin Deschaintre, Julia Guerrero-Viu, Diego Gutierrez, Tamy Boubekeur, and Belen Masia. The visual language of fabrics. ACM Transactions on Graphics, 42 (4):50:1–50:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592391.

deGoes:2015:PPI

[dGWH^{+}15] Fernando de Goes, Corentin Wallez, Jin Huang, Dmitry Pavlov, and Mathieu Desbrun. Power particles: an incompressible fluid solver based on power diagrams. ACM Transactions on Graphics, 34(4):50:1–50:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Davidson:1996:DGN

[DH96] Ron Davidson and David Harel. Drawing graphs nicely using simulated annealing. ACM Transactions on Graphics, 15(4):301–331, October 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
234538.html.

Dunbar:2006:SDS

[DH06] Daniel Dunbar and Greg Humphreys. A spatial data structure for fast Poissondisk sample generation. ACM Transactions on Graphics, 25 (3):503–508, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Da:2016:SOL

[DHB^{+}16] Fang Da, David Hahn, Christopher Batty, Chris Wojtan, and Eitan Grinspun. Surface-only liquids. ACM Transactions on Graphics, 35 (4):78:1–78:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dupuy:2017:SCP

[DHB17] Jonathan Dupuy, Eric Heitz, and Laurent Belcour. A spherical cap preserving parameterization for spherical distributions. ACM Transactions on Graphics, 36(4): 139:1–139:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deng:2021:PGI

[DHC^{+}21] Xi Deng, Milos Hasan, Nathan Carr, Zexiang Xu, and Steve Marschner. Path graphs: iterative path space filtering. ACM Transactions on Graphics, 40(6): 276:1–276:15, December 2021.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480547.

Dupuy:2013:LEA

[DHI^{+}13] Jonathan Dupuy, Eric Heitz, Jean-Claude Iehl, Pierre Poulin, Fabrice Neyret, and Victor Ostromoukhov. Linear efficient antialiased displacement and reflectance mapping. ACM Transactions on Graphics, 32(6):211:1– 211:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dumas:2014:BGA

[DHL14] Jérémie Dumas, Jean Hergel, and Sylvain Lefebvre. Bridging the gap: automated steady scaffoldings for 3D printing. ACM Transactions on Graphics, 33(4): 98:1–98:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dobbyn:2005:GRT

[DHOO05] Simon Dobbyn, John Hamill, Keith O’Conor, and Carol O’Sullivan. Geopostors: a real-time geometry/impostor crowd rendering system. ACM Transactions on Graphics, 24 (3):933, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Durand:2005:FAL

[DHS^{+}05] Frédo Durand, Nicolas Holzschuch, Cyril Soler, Eric Chan, and François X. Sillion. A frequency analysis of light transport. ACM Transactions on Graphics, 24(3): 1115–1126, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Djeu:2011:RAD

[DHW^{+}11] Peter Djeu, Warren Hunt, Rui Wang, Ikrima Elhassan, Gordon Stoll, and William R. Mark. Razor: an architecture for dynamic multiresolution ray tracing. ACM Transactions on Graphics, 30(5): 115:1–115:26, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DEon:2011:QDM

[DI11] Eugene D’Eon and Geoffrey Irving. A quantized-diffusion model for rendering translucent materials. ACM Transactions on Graphics, 30(4): 56:1–56:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Didyk:2018:SDA

[Did18] Piotr Didyk. Session details: Acquisition, rendering and display for virtual reality. ACM Transactions on Graphics, 37(6), November 2018.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dobashi:2023:EVL

[DII23] Yoshinori Dobashi, Naoto Ishikawa, and Kei Iwasaki. Efficient visualization of light pollution for the night sky. ACM Transactions on Graphics, 42(6):219:1–219:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618337.

Dobashi:2012:IPA

[DIO^{+}12] Yoshinori Dobashi, Wataru Iwasaki, Ayumi Ono, Tsuyoshi Yamamoto, Yonghao Yue, and Tomoyuki Nishita. An inverse problem approach for automatically adjusting the parameters for rendering clouds using photographs. ACM Transactions on Graphics, 31 (6):145:1–145:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Du:2018:IAC

[DIP^{+}18] Tao Du, Jeevana Priya Inala, Yewen Pu, Andrew Spielberg, Adriana Schulz, Daniela Rus, Armando Solar-Lezama, and Wojciech Matusik. InverseCSG: automatic conversion of 3D models to CSG trees. ACM Transactions on Graphics, 37(6):213:1– 213:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Donner:2005:LDM

[DJ05] Craig Donner and Henrik Wann Jensen. Light diffusion in multi-layered translucent materials. ACM Transactions on Graphics, 24(3): 1032–1039, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeGoes:2017:RKS

[DJ17] Fernando De Goes and Doug L. James. Regularized kelvinlets: sculpting brushes based on fundamental solutions of elasticity. ACM Transactions on Graphics, 36 (4):40:1–40:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goes:2018:DKS

[DJ18a] Fernando De Goes and Doug L. James. Dynamic kelvinlets: secondary motions based on fundamental solutions of elastodynamics. ACM Transactions on Graphics, 37 (4):81:1–81:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dupuy:2018:APE

[DJ18b] Jonathan Dupuy and Wenzel Jakob. An adaptive parameterization for efficient material acquisition and rendering. ACM Transactions on Graphics, 37(6):274:1– 274:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Derouet-Jourdan:2013:IDH

[DJBDDT13] Alexandre Derouet-Jourdan, Florence Bertails-Descoubes, Gilles Daviet, and Joëlle Thollot. Inverse dynamic hair modeling with frictional contact. ACM Transactions on Graphics, 32(6):159:1– 159:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Derouet-Jourdan:2010:SID

[DJBDT10] Alexandre Derouet-Jourdan, Florence Bertails-Descoubes, and Joëlle Thollot. Stable inverse dynamic curves. ACM Transactions on Graphics, 29 (6):137:1–137:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deng:2019:PSR

[DJBJ19] Xi Deng, Shaojie Jiao, Benedikt Bitterli, and Wojciech Jarosz. Photon surfaces for robust, unbiased volumetric density estimation. ACM Transactions on Graphics, 38 (4):46:1–46:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Douglas:1999:MRE

[DK99] Sarah A. Douglas and Arthur E. Kirkpatrick. Model and representation: the effect of visual feedback on human performance in a color picker interface. ACM Transactions on Graphics, 18(2): 96–127, April 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1999-18-2/
p96-douglas/.

Degener:2009:VAA

[DK09] Patrick Degener and Reinhard Klein. A variational approach for automatic generation of panoramic maps. ACM Transactions on Graphics, 28(1): 2:1–2:14, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Djuren:2023:KSB

[DKA23] Tobias Djuren, Maximilian Kohlbrenner, and Marc Alexa. K-surfaces: Béziersplines interpolating at Gaussian curvature extrema. ACM Transactions on Graphics, 42 (6):210:1–210:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618383.

Dou:2016:FRT

[DKD^{+}16] Mingsong Dou, Sameh Khamis, Yury Degtyarev, Philip Davidson, Sean Ryan Fanello, Adarsh Kowdle, Sergio Orts Escolano, Christoph Rhemann, David Kim, Jonathan Taylor, Pushmeet Kohli, Vladimir Tankovich, and Shahram Izadi. Fusion4D: real-time performance capture of challenging scenes. ACM Transactions on Graphics, 35 (4):114:1–114:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Durupinar:2017:PPAa

[DKD^{+}17a] Funda Durupinar, Mubbasir Kapadia, Susan Deutsch, Michael Neff, and Norman I. Badler. PERFORM: Perceptual approach for adding OCEAN personality to human motion using laban movement analysis. ACM Transactions on Graphics, 36(1): 6:1–6:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Durupinar:2017:PPAb

[DKD^{+}17b] Funda Durupinar, Mubbasir Kapadia, Susan Deutsch, Michael Neff, and Norman I. Badler. Perform: perceptual approach for adding OCEAN personality to human motion using laban movement analysis. ACM Transactions on Graphics, 36(4):

48:1–48:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Davidovic:2010:CGL

[DKH^{+}10] Tomáš Davidovič, Jaroslav Křivánek, Miloš Hašan, Philipp Slusallek, and Kavita Bala. Combining global and local virtual lights for detailed glossy illumination. ACM Transactions on Graphics, 29 (6):143:1–143:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Davidovic:2014:PLT

[DKHS14] Tomás Davidovic, Jaroslav Krivánek, Milos Hasan, and Philipp Slusallek. Progressive light transport simulation on the GPU: Survey and improvements. ACM Transactions on Graphics, 33(3): 29:1–29:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dobashi:2008:FCC

[DKNY08] Yoshinori Dobashi, Katsutoshi Kusumoto, Tomoyuki Nishita, and Tsuyoshi Yamamoto. Feedback control of cumuliform cloud formation based on computational fluid dynamics. ACM Transactions on Graphics, 27(3): 94:1–94:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Denning:2011:MIV

[DKP11] Jonathan D. Denning, William B.

Kerr, and Fabio Pellacini. MeshFlow: interactive visualization of mesh construction sequences. ACM Transactions on Graphics, 30(4): 66:1–66:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Du:2023:IVE

[DKT^{+}23] Zheng-Jun Du, Liang-Fu Kang, Jianchao Tan, Yotam Gingold, and Kun Xu. Image vectorization and editing via linear gradient layer decomposition. ACM Transactions on Graphics, 42(4): 97:1–97:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592128.

Du:2021:OGI

[DKZ^{+}21] Xingyi Du, Danny M. Kaufman, Qingnan Zhou, Shahar Z. Kovalsky, Yajie Yan, Noam Aigerman, and Tao Ju. Optimizing global injectivity for constrained parameterization. ACM Transactions on Graphics, 40(6): 260:1–260:18, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480556.

Dang:2015:IDP

[DLC^{+}15] Minh Dang, Stefan Lienhard, Duygu Ceylan, Boris Neubert, Peter Wonka, and Mark Pauly. Interactive design of probability density functions for shape grammars. ACM Transactions on Graphics, 34 (6):206:1–206:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeWitt:2012:FSU

[DLF12] Tyler De Witt, Christian Lessig, and Eugene Fiume. Fluid simulation using Laplacian eigenfunctions. ACM Transactions on Graphics, 31 (1):10:1–10:11, January 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dyn:1990:BSS

[DLG90] Nira Dyn, David Levin, and John A. Gregory. A butterfly subdivision scheme for surface interpolation with tension control. ACM Transactions on Graphics, 9(2): 160–169, April 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        78958.html.

Dinev:2018:SIR

[DLK18] Dimitar Dinev, Tiantian Liu, and Ladislav Kavan. Stabilizing integrators for realtime physics. ACM Transactions on Graphics, 37(1): 9:1–9:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dvoroznak:2018:TEB

[DLKS18] Marek Dvoroznák, Wilmot Li, Vladimir G. Kim, and Daniel Sýkora. Toonsynth: example-based synthesis of hand-colored cartoon animations. ACM Transactions on Graphics, 37(4):167:1– 167:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dumas:2015:ESS

[DLL^{+}15] Jérémie Dumas, An Lu, Sylvain Lefebvre, Jun Wu, T. U. München, Christian Dick, and T. U. München. By-example synthesis of structurally sound patterns. ACM Transactions on Graphics, 34(4):137:1– 137:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dinev:2018:FFE

[DLL^{+}18] Dimitar Dinev, Tiantian Liu, Jing Li, Bernhard Thomaszewski, and Ladislav Kavan. FEPR: fast energy projection for real-time simulation of deformable objects. ACM Transactions on Graphics, 37(4):79:1–79:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

deLasa:2010:FBL

[dLMH10] Martin de Lasa, Igor Mordatch, and Aaron Hertzmann. Feature-based locomotion controllers. ACM Transactions on Graphics, 29(4): 131:1–131:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duinkharjav:2023:SRA

[DLP^{+}23] Budmonde Duinkharjav, Benjamin Liang, Anjul Patney, Rachel Brown, and Qi Sun. The shortest route is not always the fastest: Probability-modeled stereoscopic eye movement completion time in VR. ACM Transactions on Graphics, 42(6): 220:1–220:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618334.

Donner:2009:EBM

[DLR^{+}09] Craig Donner, Jason Lawrence, Ravi Ramamoorthi, Toshiya Hachisuka, Henrik Wann Jensen, and Shree Nayar. An empirical BSSRDF model. ACM Transactions on Graphics, 28(3):30:1–30:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dey:2008:CGA

[DLSCS08] Tamal K. Dey, Kuiyu Li, Jian Sun, and David Cohen-Steiner. Computing geometryaware handle and tunnel loops in 3D models. ACM Transactions on Graphics, 27(3): 45:1–45:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dobkin:1990:CTP

[DLTW90] David P. Dobkin, Silvio V. F. Levy, William P. Thurston, and Allan R. Wilks. Contour tracing by piecewise linear approximations. ACM Transactions on Graphics, 9(4):389–423, October 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
88575.html.

Deng:2022:CCS

[DLW^{+}22] Hong Deng, Yang Liu, Beibei Wang, Jian Yang, Lei Ma, Nicolas Holzschuch, and Ling-Qi Yan. Constant-cost spatioangular prefiltering of glinty appearance using tensor decomposition. ACM Transactions on Graphics, 41 (2):22:1–22:17, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3507915.

Du:2021:VRS

[DLX^{+}21] Zheng-Jun Du, Kai-Xiang Lei, Kun Xu, Jianchao Tan, and Yotam Gingold. Video recoloring via spatial-temporal geometric palettes. ACM Transactions on Graphics, 40(4): 150:1–150:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459675.

Deng:2013:UIS

[DM13] Chongyang Deng and Weiyin Ma. A unified interpolatory subdivision scheme for quadrilateral meshes. ACM Transactions on Graphics, 32(3): 23:1–23:11, June 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Delbracio:2014:BMC

[DMB^{+}14] Mauricio Delbracio, Pablo Musé, Antoni Buades, Julien Chauvier, Nicholas Phelps, and Jean-Michel Morel. Boosting Monte Carlo rendering by ray histogram fusion. ACM Transactions on Graphics, 33 (1):8:1–8:15, January 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Du:2013:MVC

[DMHG13] Song-Pei Du, Belen Masia, Shi-Min Hu, and Diego Gutierrez. A metric of visual comfort for stereoscopic motion. ACM Transactions on Graphics, 32(6):222:1– 222:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dekel:2015:RMN

[DMIF15] Tali Dekel, Tomer Michaeli, Michal Irani, and William T. Freeman. Revealing and modifying non-local variations in a single image. ACM Transactions on Graphics, 34(6): 227:1–227:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dym:2017:DFS

[DML17] Nadav Dym, Haggai Maron, and Yaron Lipman. DS++: a flexible, scalable and provably tight relaxation for matching problems. ACM Transactions on Graphics, 36(6): 184:1–184:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Devito:2017:ODS

[DMZ^{+}17] Zachary Devito, Michael Mara, Michael Zollhöfer, Gilbert Bernstein, Jonathan Ragan-Kelley, Christian Theobalt, Pat Hanrahan, Matthew Fisher, and Matthias Niessner. Opt: a domain specific language for non-linear least squares optimization in graphics and imaging. ACM Transactions on Graphics, 36(5): 171:1–171:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deering:2002:SGA

[DN02] Michael Deering and David Naegle. The SAGE graphics architecture. ACM Transactions on Graphics, 21(3): 683–692, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duca:2005:RDE

[DNB^{+}05] Nathaniel Duca, Krzysztof Niski, Jonathan Bilodeau, Matthew Bolitho, Yuan Chen, and Jonathan Cohen. A relational debugging engine for the graphics pipeline. ACM Transactions on Graphics, 24 (3):453–463, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dai:2017:BRTb

[DNZ^{+}17a] Angela Dai, Matthias Nießner, Michael Zollhöfer, Shahram Izadi, and Christian Theobalt. BundleFusion: real-time globally consistent 3D reconstruction using on-the-fly surface re-integration. ACM Transactions on Graphics, 36(4):

76:1–76:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dai:2017:BRTa

[DNZ^{+}17b] Angela Dai, Matthias Nießner, Michael Zollhöfer, Shahram Izadi, and Christian Theobalt. BundleFusion: Real-time globally consistent 3D reconstruction using on-the-fly surface reintegration. ACM Transactions on Graphics, 36 (3):24:1–24:??, June 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Denning:2013:MDM

[DP13] Jonathan D. Denning and Fabio Pellacini. MeshGit: diffing and merging meshes for polygonal modeling. ACM Transactions on Graphics, 32 (4):35:1–35:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Diolatzis:2022:AEN

[DPD22] Stavros Diolatzis, Julien Philip, and George Drettakis. Active exploration for neural global illumination of variable scenes. ACM Transactions on Graphics, 41(5): 171:1–171:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3522735.

Dumont:2003:PDD

[DPF03] Reynald Dumont, Fabio Pellacini, and James A. Ferwerda. Perceptually-driven decision theory for interactive realistic rendering. ACM Transactions on Graphics, 22 (2):152–181, April 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Diazzi:2023:CDT

[DPVA23] Lorenzo Diazzi, Daniele Panozzo, Amir Vaxman, and Marco Attene. Constrained Delaunay tetrahedrization: a robust and practical approach. ACM Transactions on Graphics, 42 (6):181:1–181:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3618352.

Deuss:2014:ASS

[DPW^{+}14] Mario Deuss, Daniele Panozzo, Emily Whiting, Yang Liu, Philippe Block, Olga Sorkine-Hornung, and Mark Pauly. Assembling self-supporting structures. ACM Transactions on Graphics, 33(6): 214:1–214:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dansereau:2015:LVF

[DPW15] Donald G. Dansereau, Oscar Pizarro, and Stefan B.

Williams. Linear volumetric focus for light field cameras. ACM Transactions on Graphics, 34(2):15:1–15:??, February 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duchene:2015:MII

[DRC^{+}15] Sylvain Duchêne, Clement Riant, Gaurav Chaurasia, Jorge Lopez Moreno, Pierre-Yves Laffont, Stefan Popov, Adrien Bousseau, and George Drettakis. Multiview intrinsic images of outdoors scenes with an application to relighting. ACM Transactions on Graphics, 34(5): 164:1–164:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Didyk:2011:PMD

[DRE^{+}11] Piotr Didyk, Tobias Ritschel, Elmar Eisemann, Karol Myszkowski, and Hans-Peter Seidel. A perceptual model for disparity. ACM Transactions on Graphics, 30(4):96:1–96:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Didyk:2012:LCA

[DRE^{+}12] Piotr Didyk, Tobias Ritschel, Elmar Eisemann, Karol Myszkowski, Hans-Peter Seidel, and Wojciech Matusik. A luminancecontrast-aware disparity model and applications. ACM Transactions on Graphics, 31(6):

184:1–184:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dalstein:2014:VGC

[DRvdP14] Boris Dalstein, Rémi Ronfard, and Michiel van de Panne. Vector graphics complexes. ACM Transactions on Graphics, 33(4):133:1–133:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dalstein:2015:VGA

[DRvdP15] Boris Dalstein, Rémi Ronfard, and Michiel van de Panne. Vector graphics animation with time-varying topology. ACM Transactions on Graphics, 34(4):145:1– 145:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Davis:2014:VMP

[DRW^{+}14] Abe Davis, Michael Rubinstein, Neal Wadhwa, Gautham J. Mysore, Frédo Durand, and William T. Freeman. The visual microphone: passive recovery of sound from video. ACM Transactions on Graphics, 33 (4):79:1–79:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Desaulniers:1992:EMB

[DS92] H. Desaulniers and N. F.

Stewart. An extension of manifold boundary representations to the r-sets. ACM Transactions on Graphics, 11(1):40–60, January 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
111777.html.

DeCarlo:2002:SAP

[DS02] Doug DeCarlo and Anthony Santella. Stylization and abstraction of photographs. ACM Transactions on Graphics, 21(3):769–776, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DePaoli:2015:SSB

[DS15] Chris De Paoli and Karan Singh. SecondSkin: sketchbased construction of layered 3D models. ACM Transactions on Graphics, 34(4): 126:1–126:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Didyk:2013:JVE

[DSAF^{+}13] Piotr Didyk, Pitchaya Sitthi-Amorn, William Freeman, Frédo Durand, and Wojciech Matusik. Joint view expansion and filtering for automultiscopic 3D displays. ACM Transactions on Graphics, 32 (6):221:1–221:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

daSilva:2008:ISS

[dSAP08] Marco da Silva, Yeuhi Abe, and Jovan Popović. Interactive simulation of stylized human locomotion. ACM Transactions on Graphics, 27(3): 82:1–82:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Darabi:2012:IMC

[DSB^{+}12] Soheil Darabi, Eli Shechtman, Connelly Barnes, Dan B. Goldman, and Pradeep Sen. Image melding: combining inconsistent images using patchbased synthesis. ACM Transactions on Graphics, 31(4): 82:1–82:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dvoroznak:2020:MMS

[DSC^{+}20] Marek Dvorožňák, Daniel Sýkora, Cassidy Curtis, Brian Curless, Olga Sorkine-Hornung, and David Salesin. Monster mash: a single-view approach to casual 3D modeling and animation. ACM Transactions on Graphics, 39(6): 214:1–214:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417805.

Dachsbacher:2007:IVA

[DSDD07] Carsten Dachsbacher, Marc Stamminger, George Drettakis, and Frédo Durand. Implicit visibility and antiradiance for interactive global illumination. ACM Transactions on Graphics, 26(3): 61:1–61:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

daSilva:2009:LBC

[dSDP09] Marco da Silva, Frédo Durand, and Jovan Popović. Linear Bellman combination for control of character animation. ACM Transactions on Graphics, 28(3):82:1– 82:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeGoes:2022:CAT

[DSF22] Fernando De Goes, William Sheffler, and Kurt Fleischer. Character articulation through profile curves. ACM Transactions on Graphics, 41 (4):139:1–139:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530060.

Doersch:2012:WMP

[DSG^{+}12] Carl Doersch, Saurabh Singh, Abhinav Gupta, Josef Sivic, and Alexei A. Efros. What makes Paris look like Paris? ACM Transactions on Graphics, 31(4):101:1–101:9, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dale:2011:VFR

[DSJ^{+}11] Kevin Dale, Kalyan Sunkavalli, Micah K. Johnson, Daniel Vlasic, Wojciech Matusik, and Hanspeter Pfister. Video face replacement. ACM Transactions on Graphics, 30(6): 130:1–130:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Diamond:2021:DPT

[DSJA^{+}21] Steven Diamond, Vincent Sitzmann, Frank Julca-Aguilar, Stephen Boyd, Gordon Wetzstein, and Felix Heide. Dirty pixels: Towards end-to-end image processing and perception. ACM Transactions on Graphics, 40(3): 23:1–23:15, July 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3446918.

Der:2006:IKR

[DSP06] Kevin G. Der, Robert W. Sumner, and Jovan Popović. Inverse kinematics for reduced deformable models. ACM Transactions on Graphics, 25 (3):1174–1179, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Daniels:2008:QMS

[DSSC08] Joel Daniels, Cláudio T. Silva, Jason Shepherd, and Elaine Cohen. Quadrilateral mesh simplification. ACM Transactions on Graphics, 27(5): 148:1–148:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dodik:2023:VBC

[DSSS23] Ana Dodik, Oded Stein, Vincent Sitzmann, and Justin Solomon. Variational barycentric coordinates. ACM Transactions on Graphics, 42(6): 255:1–255:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618403.

Du:2016:CMD

[DSZ^{+}16] Tao Du, Adriana Schulz, Bo Zhu, Bernd Bickel, and Wojciech Matusik. Computational multicopter design. ACM Transactions on Graphics, 35(6):227:1– 227:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deussen:2017:WLB

[DSZ17] Oliver Deussen, Marc Spicker, and Qian Zheng. Weighted Linde–Buzo–gray stippling. ACM Transactions on Graphics, 36(6):233:1–233:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Diebel:2006:BMP

[DTB06] James R. Diebel, Sebastian Thrun, and Michael Brünig. A Bayesian method for probable surface reconstruction and decimation. ACM Transactions on Graphics, 25(1): 39–59, January 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Denning:2015:FCS

[DTP15] Jonathan D. Denning, Valentina Tibaldo, and Fabio Pellacini. 3DFlow: continuous summarization of mesh editing workflows. ACM Transactions on Graphics, 34(4):140:1– 140:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duenser:2023:NCM

[DTPC23] Simon Duenser, Bernhard Thomaszewski, Roi Poranne, and Stelian Coros. Nonlinear compliant modes for large-deformation analysis of flexible structures. ACM Transactions on Graphics, 42 (2):21:1–21:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3568952.

Dong:2011:AIM

[DTPG11] Yue Dong, Xin Tong, Fabio Pellacini, and Baining Guo. AppGen: interactive material modeling from a single image. ACM Transactions on Graphics, 30(6): 146:1–146:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2012:PSV

[DTPG12] Yue Dong, Xin Tong, Fabio Pellacini, and Baining Guo. Printing spatially-varying reflectance for reproducing HDR images. ACM Transactions on Graphics, 31(4): 40:1–40:7, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duff:2017:DCUa

[Duf17a] Tom Duff. Deep compositing using Lie algebras. ACM Transactions on Graphics, 36 (3):26:1–26:??, June 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duff:2017:DCUb

[Duf17b] Tom Duff. Deep compositing using lie algebras. ACM Transactions on Graphics, 36 (4):120:1–120:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dunlavey:1983:EPF

[Dun83] M. R. Dunlavey. Efficient polygon-filling algorithms for raster displays. ACM Transactions on Graphics, 2(4): 264–273, October 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Damera-Venkata:2009:DS

[DVC09] Niranjan Damera-Venkata and Nelson L. Chang. Display supersampling. ACM Transactions on Graphics, 28(1): 9:1–9:19, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dana:1999:RTR

[DvGNK99] Kristin J. Dana, Bram van Ginneken, Shree K. Nayar, and Jan J. Koenderink. Reflectance and texture of real-world surfaces. ACM Transactions on Graphics, 18(1):1–34, January 1999.

                                        URL http://www.acm.org:
                                        80/pubs/citations/journals/
                                        tog/1999-18-1/p1-dana/.

Diamanti:2015:IPF

[DVPSH15] Olga Diamanti, Amir Vaxman, Daniele Panozzo, and Olga Sorkine-Hornung. Integrable PolyVector fields. ACM Transactions on Graphics, 34 (4):38:1–38:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dachsbacher:2003:SPT

[DVS03] Carsten Dachsbacher, Christian Vogelgsang, and Marc Stamminger. Sequential point trees. ACM Transactions on Graphics, 22(3):657–662, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Donner:2008:LHR

[DWd^{+}08] Craig Donner, Tim Weyrich, Eugene d’Eon, Ravi Ramamoorthi, and Szymon Rusinkiewicz. A layered, heterogeneous reflectance model for acquiring and rendering human skin. ACM Transactions on Graphics, 27(5): 140:1–140:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deng:2022:MEL

[DWK^{+}22] Yitong Deng, Mengdi Wang, Xiangxin Kong, Shiying Xiong, Zangyueyang Xian, and Bo Zhu. A moving Eulerian–Lagrangian particle method for thin film and foam simulation. ACM Transactions on Graphics, 41 (4):154:1–154:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530174.

Du:2022:DDP

[DWM^{+}22] Tao Du, Kui Wu, Pingchuan Ma, Sebastien Wah, Andrew Spielberg, Daniela Rus, and Wojciech Matusik. DiffPD: Differentiable projective dynamics. ACM Transactions on Graphics, 41(2):13:1– 13:21, April 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3490168.

Dong:2015:PAM

[DWMG15] Zhao Dong, Bruce Walter, Steve Marschner, and Donald P. Greenberg. Predicting appearance from measured microgeometry of metal surfaces. ACM Transactions on Graphics, 35(1):9:1– 9:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2010:FSV

[DWP^{+}10] Yue Dong, Jiaping Wang, Fabio Pellacini, Xin Tong, and Baining Guo. Fabricating spatially-varying subsurface scattering. ACM Transactions on Graphics, 29(4): 62:1–62:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Du:2020:FOF

[DWS^{+}20] Tao Du, Kui Wu, Andrew Spielberg, Wojciech Matusik, Bo Zhu, and Eftychios Sifakis. Functional optimization of fluidic devices with differentiable Stokes flow. ACM Transactions on Graphics, 39(6): 197:1–197:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417795.

Duan:2023:BBN

[DWS^{+}23] Hao-Bin Duan, Miao Wang, Jin-Chuan Shi, Xu-Chuan Chen, and Yan-Pei Cao. BakedAvatar: Baking neural fields for real-time head avatar synthesis. ACM Transactions on Graphics, 42(6): 225:1–225:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618399.

Debevec:2002:LRA

[DWT^{+}02] Paul Debevec, Andreas Wenger, Chris Tchou, Andrew Gardner, Jamie Waese, and Tim Hawkins. A lighting reproduction approach to live-action compositing. ACM Transactions on Graphics, 21(3): 547–556, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2010:MBS

[DWT^{+}10] Yue Dong, Jiaping Wang, Xin Tong, John Snyder, Yanxiang Lan, Moshe Ben-Ezra, and Baining Guo. Manifold bootstrapping for SVBRDF capture. ACM Transactions on Graphics, 29(4): 98:1–98:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dai:2018:SFV

[DWW^{+}18] Chengkai Dai, Charlie C. L. Wang, Chenming Wu, Sylvain Lefebvre, Guoxin Fang, and Yong-Jin Liu. Supportfree volume printing by multiaxis motion. ACM Transactions on Graphics, 37(4): 134:1–134:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dong:2021:TRP

[DWX^{+}21] Zhi-Chao Dong, Wenming Wu, Zenghao Xu, Qi Sun, Guanjie Yuan, Ligang Liu, and Xiao-Ming Fu. Tailored reality: Perceptionaware scene restructuring for adaptive VR navigation. ACM Transactions on Graphics, 40(5):193:1–193:15, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3470847.

Dong:2023:SSR

[DXG^{+}23] Zheng Dong, Ke Xu, Yaoan Gao, Qilin Sun, Hujun Bao, Weiwei Xu, and Rynson W. H. Lau. SAILOR: Synergizing radiance and occupancy fields for live human performance capture. ACM Transactions on Graphics, 42(6): 205:1–205:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618370.

Dong:2019:MRC

[DXZ^{+}19] Siyan Dong, Kai Xu, Qiang Zhou, Andrea Tagliasacchi, Shiqing Xin, Matthias Nießner, and Baoquan Chen. Multi-robot collaborative dense scene reconstruction. ACM Transactions on Graphics, 38 (4):84:1–84:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dobashi:2003:RTR

[DYN03] Yoshinori Dobashi, Tsuyoshi Yamamoto, and Tomoyuki Nishita. Real-time rendering of aerodynamic sound using sound textures based on computational fluid dynamics. ACM Transactions on Graphics, 22(3):732–740, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dontcheva:2003:LAC

[DYP03] Mira Dontcheva, Gary Yngve, and Zoran Popović. Layered acting for character animation. ACM Transactions on Graphics, 22(3):409–416, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Dinh:2005:TTD

[DYT05] Huong Quynh Dinh, Anthony Yezzi, and Greg Turk. Texture transfer during shape transformation. ACM Transactions on Graphics, 24(2): 289–310, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duncan:2016:ICH

[DYY16] Noah Duncan, Lap-Fai Yu, and Sai-Kit Yeung. Interchangeable components for hands-on assembly based modelling. ACM Transactions on Graphics, 35(6): 234:1–234:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duncan:2015:ZD

[DYYT15] Noah Duncan, Lap-Fai Yu, Sai-Kit Yeung, and Demetri Terzopoulos. Zoomorphic design. ACM Transactions on Graphics, 34(4):95:1– 95:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Duncan:2017:AD

[DYYT17] Noah Duncan, Lap-Fai Yu, Sai-Kit Yeung, and Demetri Terzopoulos. Approximate dissections. ACM Transactions on Graphics, 36(6): 182:1–182:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Deng:2023:FSN

[DYZ^{+}23] Yitong Deng, Hong-Xing Yu, Diyang Zhang, Jiajun Wu, and Bo Zhu. Fluid simulation on neural flow maps. ACM Transactions on Graphics, 42 (6):248:1–248:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618392.

Du:2021:BSH

[DZCJ21] Xingyi Du, Qingnan Zhou, Nathan Carr, and Tao Ju. Boundary-sampled halfspaces: a new representation for constructive solid modeling. ACM Transactions on Graphics, 40(4): 53:1–53:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459870.

Du:2022:RCI

[DZCJ22] Xingyi Du, Qingnan Zhou, Nathan Carr, and Tao Ju. Robust computation of implicit surface networks for piecewise linear functions. ACM Transactions on Graphics, 41(4):41:1–41:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530176.

Dong:2009:OIR

[DZPZ09] Weiming Dong, Ning Zhou, Jean-Claude Paul, and Xiaopeng Zhang. Optimized image resizing using seam carving and scaling. ACM Transactions on Graphics, 28(5): 125:1–125:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DiLorenzo:2008:LLC

[DZS08] Paul C. DiLorenzo, Victor B. Zordan, and Benjamin L. Sanders. Laughing out loud: control for modeling anatomically inspired laughter using audio. ACM Transactions on Graphics, 27(5): 125:1–125:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Esteves:2006:APV

[EAPL06] Claudia Esteves, Gustavo Arechavaleta, Julien Pettré, and Jean-Paul Laumond. Animation planning for virtual characters cooperation. ACM Transactions on Graphics, 25 (2):319–339, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

English:2008:ADS

[EB08] Elliot English and Robert Bridson. Animating developable surfaces using nonconforming elements. ACM Transactions on Graphics, 27 (3):66:1–66:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Edwards:2014:DWC

[EB14] Essex Edwards and Robert Bridson. Detailed water with coarse grids: combining surface meshes and adaptive discontinuous Galerkin. ACM Transactions on Graphics, 33 (4):136:1–136:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ebke:2013:QRQ

[EBCK13] Hans-Christian Ebke, David Bommes, Marcel Campen, and Leif Kobbelt. QEx: robust quad mesh extraction. ACM Transactions on Graphics, 32(6):168:1– 168:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Echevarria:2014:CSH

[EBGB14] Jose I. Echevarria, Derek Bradley, Diego Gutierrez, and Thabo Beeler. Capturing and stylizing hair for 3D fabrication. ACM Transactions on Graphics, 33(4): 125:1–125:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Edwards:2006:HVD

[EBJ^{+}06] Dave Edwards, Solomon Boulos, Jared Johnson, Peter Shirley, Michael Ashikhmin, Michael Stark, and Chris Wyman. The halfway vector disk for BRDF modeling. ACM Transactions on Graphics, 25(1):1–18, January 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Elber:1993:SOS

[EC93] Gershon Elber and Elaine Cohen. Second-order surface analysis using hybrid symbolic and numeric operators. ACM Transactions on Graphics, 12(2):160–178, April 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        151283.html.

Elber:1996:AIB

[EC96] Gershon Elber and Elaine Cohen. Adaptive isocurve-based rendering for freeform surfaces. ACM Transactions on Graphics, 15(3):249–263, July 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        230537.html.

Ebke:2014:LDQ

[ECBK14] Hans-Christian Ebke, Marcel Campen, David Bommes, and Leif Kobbelt. Level-of-detail quad meshing. ACM Transactions on Graphics, 33(6): 184:1–184:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Eisemann:2004:FPE

[ED04] Elmar Eisemann and Frédo Durand. Flash photography enhancement via intrinsic relighting. ACM Transactions on Graphics, 23(3): 673–678, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Efrat:2016:CLS

[EDF^{+}16] Netalee Efrat, Piotr Didyk, Mike Foshey, Wojciech Matusik, and Anat Levin. Cinema 3D: large scale automultiscopic display. ACM Transactions on Graphics, 35(4): 59:1–59:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ebeida:2011:EMP

[EDP^{+}11] Mohamed S. Ebeida, Andrew A. Davidson, Anjul Patney, Patrick M. Knupp, Scott A. Mitchell, and John D. Owens. Efficient maximal Poisson-disk sampling. ACM Transactions on Graphics, 30 (4):49:1–49:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Egan:2011:PFE

[EDR11] Kevin Egan, Frédo Durand, and Ravi Ramamoorthi. Practical filtering for efficient ray-traced directional occlusion. ACM Transactions on Graphics, 30(6):180:1– 180:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ezzat:2002:TVS

[EGP02] Tony Ezzat, Gadi Geiger, and Tomaso Poggio. Trainable videorealistic speech animation. ACM Transactions on Graphics, 21(3):388–398, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Eitz:2012:HDH

[EHA12] Mathias Eitz, James Hays, and Marc Alexa. How do humans sketch objects? ACM Transactions on Graphics, 31(4):44:1–44:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Egan:2011:FAS

[EHDR11] Kevin Egan, Florian Hecht, Frédo Durand, and Ravi Ramamoorthi. Frequency analysis and sheared filtering for shadow light fields of complex occluders. ACM Transactions on Graphics, 30(2): 9:1–9:13, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Eom:2020:MPC

[EHSN20] Haegwang Eom, Daseong Han, Joseph S. Shin (formerly Sung Yong Shin), and Junyong Noh. Model predictive control with a visuomotor system for physics-based character animation. ACM Transactions on Graphics, 39(1): 3:1–3:11, February 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3360905.

Elber:1998:BSR

[EK98] Gershon Elber and Myung-Soo Kim. The bisector surface of rational space curves. ACM Transactions on Graphics, 17 (1):32–49, January 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1998-17-1/
p32-elber/.

Edahiro:1984:NPL

[EKA84] M. Edahiro, I. Kokubo, and Ta. Asano. A new pointlocation algorithm and its practical efficiency: comparison with existing algorithms. ACM Transactions on Graphics, 3(2):86–109, April 1984. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Eilertsen:2017:HIR

[EKD^{+}17] Gabriel Eilertsen, Joel Kronander, Gyorgy Denes, Rafal K. Mantiuk, and Jonas Unger. HDR image reconstruction from a single exposure using deep CNNs. ACM Transactions on Graphics, 36(6): 178:1–178:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Endo:2017:DRT

[EKM17] Yuki Endo, Yoshihiro Kanamori, and Jun Mitani. Deep reverse tone mapping. ACM Transactions on Graphics, 36(6): 177:1–177:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Eigensatz:2010:PAF

[EKS^{+}10] Michael Eigensatz, Martin Kilian, Alexander Schiftner, Niloy J. Mitra, Helmut Pottmann, and Mark Pauly. Paneling architectural freeform surfaces. ACM Transactions on Graphics, 29 (4):45:1–45:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Edwards:2016:JAC

[ELFS16] Pif Edwards, Chris Landreth, Eugene Fiume, and Karan Singh. JALI: an animatorcentric viseme model for expressive lip synchronization.

ACM Transactions on Graphics, 35(4):127:1–127:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Edelsbrunner:1990:SST

[EM90] Herbert Edelsbrunner and Ernst Peter Mucke. Simulation of simplicity: a technique to cope with degenerate cases in geometric algorithms. ACM Transactions on Graphics, 9(1):66–104, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77639.html.

Edelsbrunner:1994:TDA

[EM94] Herbert Edelsbrunner and Ernst P. Mücke. Threedimensional alpha shapes. ACM Transactions on Graphics, 13(1):43–72, January 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
156635.html.

Ezquerra:1996:APD

[EM96] Norberto Ezquerra and Rakesh Mullick. An approach to 3D pose determination. ACM Transactions on Graphics, 15 (2):99–120, April 1996. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Enright:2002:ARC

[EMF02] Douglas Enright, Stephen Marschner, and Ronald Fedkiw. Animation and rendering of complex water surfaces. ACM Transactions on Graphics, 21(3):736–744, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ephrat:2018:LLC

[EML^{+}18] Ariel Ephrat, Inbar Mosseri, Oran Lang, Tali Dekel, Kevin Wilson, Avinatan Hassidim, William T. Freeman, and Michael Rubinstein. Looking to listen at the cocktail party: a speaker-independent audio-visual model for speech separation. ACM Transactions on Graphics, 37(4): 112:1–112:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ennis:2010:SBB

[EMO10] Cathy Ennis, Rachel McDonnell, and Carol O’Sullivan. Seeing is believing: body motion dominates in multisensory conversations. ACM Transactions on Graphics, 29 (4):91:1–91:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Elgharib:2020:EV

[EMT^{+}20] Mohamed Elgharib, Mohit Mendiratta, Justus Thies, Matthias Niessner, Hans-Peter Seidel, Ayush Tewari, Vladislav Golyanik, and Christian Theobalt. Egocentric videoconferencing. ACM Transactions on Graphics, 39 (6):268:1–268:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417808.

Eilertsen:2015:RTN

[EMU15] Gabriel Eilertsen, Rafał K.

Mantiuk, and Jonas Unger. Real-time noise-aware tone mapping. ACM Transactions on Graphics, 34(6):198:1– 198:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ecormier-Nocca:2021:ACL

[ENCC^{+}21] Pierre Ecormier-Nocca, Guillaume Cordonnier, Philippe Carrez, Anne-Marie Moigne, Pooran Memari, Bedrich Benes, and Marie-Paule Cani. Authoring consistent landscapes with flora and fauna. ACM Transactions on Graphics, 40(4):105:1–105:13, August 2021. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459952.

Eisemann:2009:VAC

[EPD09] Elmar Eisemann, Sylvain Paris, and Frédo Durand. A visibility algorithm for converting 3D meshes into editable 2D vector graphics. ACM Transactions on Graphics, 28(3):83:1–83:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ebeida:2014:KDD

[EPM^{+}14] Mohamed S. Ebeida, Anjul Patney, Scott A. Mitchell, Keith R. Dalbey, Andrew A. Davidson, and John D. Owens. k–d Darts: Sampling by k-dimensional flat searches. ACM Transactions on Graphics, 33(1):3:1–3:16, January 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ekoule:1991:TAA

[EPO91] A. B. Ekoule, F. C. Peyrin, and C. L. Odet. A triangulation algorithm from arbitrary shaped multiple planar contours. ACM Transactions on Graphics, 10(2): 182–199, April 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        108363.html.

Eitz:2012:SBS

[ERB^{+}12] Mathias Eitz, Ronald Richter, Tamy Boubekeur, Kristian Hildebrand, and Marc Alexa. Sketch-based shape retrieval. ACM Transactions on Graphics, 31(4):31:1–31:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Erleben:2007:VBS

[Erl07] Kenny Erleben. Velocitybased shock propagation for multibody dynamics animation. ACM Transactions on Graphics, 26(2): 12:1–12:??, June 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Erleben:2018:MAM

[Erl18] Kenny Erleben. Methodology for assessing mesh-based contact point methods. ACM Transactions on Graphics, 37 (3):39:1–39:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3096239.

Etienne:2019:CSC

[ERP^{+}19] Jimmy Etienne, Nicolas Ray, Daniele Panozzo, Samuel Hornus, Charlie C. L. Wang, Jonàs Martínez, Sara Mc-Mains, Marc Alexa, Brian Wyvill, and Sylvain Lefebvre. CurviSlicer: slightly curved slicing for 3-axis printers. ACM Transactions on Graphics, 38(4):81:1–81:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Esturo:2014:SQE

[ERT14] Janick Martinez Esturo, Christian Rössl, and Holger Theisel. Smoothed quadratic energies on meshes. ACM Transactions on Graphics, 34 (1):2:1–2:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ezuz:2019:RHM

[ESBC19] Danielle Ezuz, Justin Solomon, and Mirela Ben-Chen. Reversible harmonic maps between discrete surfaces. ACM Transactions on Graphics, 38 (2):15:1–15:??, April 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3202660.

Ebke:2016:ICQ

[ESCK16] Hans-Christian Ebke, Patrick Schmidt, Marcel Campen, and Leif Kobbelt. Interactively controlled quad remeshing of high resolution 3D models. ACM Transactions on Graphics, 35(6):218:1– 218:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Egger:2020:MFM

[EST^{+}20] Bernhard Egger, William A. P. Smith, Ayush Tewari, Stefanie Wuhrer, Michael Zollhoefer, Thabo Beeler, Florian Bernard, Timo Bolkart, Adam Kortylewski, Sami Romdhani, Christian Theobalt, Volker Blanz, and Thomas Vetter. 3D morphable face models-past, present, and future. ACM Transactions on Graphics, 39(5):157:1– 157:38, September 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3395208.

Elek:2017:SAT

[ESZ^{+}17] Oskar Elek, Denis Sumin, Ran Zhang, Tim Weyrich, Karol Myszkowski, Bernd Bickel, Alexander Wilkie, and Jaroslav Krivánek. Scatteringaware texture reproduction for 3D printing. ACM Transactions on Graphics, 36(6): 241:1–241:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Egan:2009:FAS

[ETH^{+}09] Kevin Egan, Yu-Ting Tseng, Nicolas Holzschuch, Frédo Durand, and Ravi Ramamoorthi. Frequency analysis and sheared reconstruction for rendering motion blur. ACM Transactions on Graphics, 28(3):93:1–93:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Elcott:2007:SCP

[ETK^{+}07] Sharif Elcott, Yiying Tong, Eva Kanso, Peter Schröder, and Mathieu Desbrun. Stable, circulation-preserving, simplicial fluids. ACM Transactions on Graphics, 26(1):4:1– 4:12, January 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Emilien:2015:WIE

[EVC^{+}15] Arnaud Emilien, Ulysse Vimont, Marie-Paule Cani, Pierre Poulin, and Bedrich Benes. WorldBrush: interactive example-based synthesis of procedural virtual worlds. ACM Transactions on Graphics, 34(4):106:1– 106:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2018:COD

[FAB^{+}18] Leman Feng, Pierre Alliez, Laurent Busé, Hervé Delingette, and Mathieu Desbrun. Curved optimal Delaunay triangulation. ACM Transactions on Graphics, 37 (4):61:1–61:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fiss:2011:CPS

[FAC11] Juliet Fiss, Aseem Agarwala, and Brian Curless. Candid portrait selection from video. ACM Transactions on Graphics, 30(6):128:1– 128:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Francu:2021:LPT

[FAER21] Mihai Frâncu, Arni Asgeirsson, Kenny Erleben, and Mads J. L. Rønnow. Lockingproof tetrahedra. ACM Transactions on Graphics, 40 (2):12:1–12:17, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3444949.

Farin:1989:CCO

[Far89] Gerald Farin. Curvature continuity and offsets for piecewise conics. ACM Transactions on Graphics, 8(2): 89–99, April 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
62056.html.

Fattal:2007:MSD

[FAR07] Raanan Fattal, Maneesh Agrawala, and Szymon Rusinkiewicz. Multiscale shape and detail enhancement from multilight image collections. ACM Transactions on Graphics, 26 (3):51:1–51:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2007:IUI

[Fat07] Raanan Fattal. Image upsampling via imposed edge statistics. ACM Transactions on Graphics, 26(3): 95:1–95:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2008:SID

[Fat08] Raanan Fattal. Single image dehazing. ACM Transactions on Graphics, 27(3): 72:1–72:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2009:EAW

[Fat09a] Raanan Fattal. Edge-avoiding wavelets and their applications. ACM Transactions on Graphics, 28(3):22:1– 22:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2009:PMI

[Fat09b] Raanan Fattal. Participating media illumination using light propagation maps. ACM Transactions on Graphics, 28(1):7:1–7:11, January 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2011:BNP

[Fat11] Raanan Fattal. Blue-noise point sampling using kernel density model. ACM Transactions on Graphics, 30(4): 48:1–48:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2014:DUC

[Fat14] Raanan Fattal. Dehazing using color-lines. ACM Transactions on Graphics, 34(1):13:1– 13:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fruhstuck:2019:TSL

[FAW19] Anna Frühstück, Ibraheem Alhashim, and Peter Wonka. TileGAN: synthesis of largescale non-homogeneous textures. ACM Transactions on Graphics, 38(4):58:1– 58:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Forsey:1995:SFH

[FB95] David Forsey and Richard H.

Bartels. Surface fitting with hierarchical splines. ACM Transactions on Graphics, 14(2):134–161, April 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
221665.html.

Farchi:2018:IOC

[FBC18] Nahum Farchi and Mirela Ben-Chen. Integer-only cross field computation. ACM Transactions on Graphics, 37 (4):91:1–91:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Finnendahl:2023:EEE

[FBCA23] Ugo Finnendahl, Dimitrios Bogiokas, Pablo Robles Cervantes, and Marc Alexa. Efficient embeddings in exact arithmetic. ACM Transactions on Graphics, 42(4): 71:1–71:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592445.

Fei:2018:MSM

[FBGZ18] Yun (Raymond) Fei, Christopher Batty, Eitan Grinspun, and Changxi Zheng. A multi-scale model for simulating liquid-fabric interactions. ACM Transactions on Graphics, 37(4):51:1–51:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fatahalian:2010:RSG

[FBH^{+}10] Kayvon Fatahalian, Solomon Boulos, James Hegarty, Kurt Akeley, William R. Mark, Henry Moreton, and Pat Hanrahan. Reducing shading on GPUs using quad-fragment merging. ACM Transactions on Graphics, 29(4): 67:1–67:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fussell:2021:SMT

[FBH21] Levi Fussell, Kevin Bergamin, and Daniel Holden. SuperTrack: motion tracking for physically simulated characters using supervised learning. ACM Transactions on Graphics, 40(6): 197:1–197:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480527.

Fuchs:2007:ASR

[FBLS07] Martin Fuchs, Volker Blanz, Hendrik P. A. Lensch, and Hans-Peter Seidel. Adaptive sampling of reflectance fields. ACM Transactions on Graphics, 26(2):10:1–10:??, June 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Freire:2023:PLY

[FBS^{+}23] Marco Freire, Manas Bhargava, Camille Schreck, Pierre-Alexandre Hugron, Bernd Bickel, and Sylvain Lefebvre. PCBend: Light up your 3D shapes with foldable circuit boards. ACM Transactions on Graphics, 42(4): 142:1–142:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592411.

Fang:2018:QTM

[FBT^{+}18] Xianzhong Fang, Hujun Bao, Yiying Tong, Mathieu Desbrun, and Jin Huang. Quadrangulation through morseparameterization hybridization. ACM Transactions on Graphics, 37(4):92:1– 92:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2009:EBI

[FCA09] Raanan Fattal, Robert Carroll, and Maneesh Agrawala. Edge-based image coarsening. ACM Transactions on Graphics, 29(1):6:1–6:11, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Filip:2008:PVM

[FCGH08] Jiří Filip, Michael J. Chantler, Patrick R. Green, and Michal Haindl. A psychophysically validated metric for bidirectional texture data reduction. ACM Transactions on Graphics, 27(5):138:1– 138:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Frisvad:2007:CSP

[FCJ07] Jeppe Revall Frisvad, Niels Jørgen Christensen, and Henrik Wann Jensen. Computing the scattering properties of participating media using Lorenz-Mie theory. ACM Transactions on Graphics, 26(3): 60:1–60:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2022:SBF

[FCK22] Linxu Fan, Floyd M. Chitalu, and Taku Komura. Simulating brittle fracture with material points. ACM Transactions on Graphics, 41(5): 177:1–177:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3522573.

Fleishman:2003:PPS

[FCOAS03] Shachar Fleishman, Daniel Cohen-Or, Marc Alexa, and Cláudio T. Silva. Progressive point set surfaces. ACM Transactions on Graphics, 22 (4):997–1011, October 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fu:2008:UOM

[FCODS08] Hongbo Fu, Daniel Cohen-Or, Gideon Dror, and Alla Sheffer. Upright orientation of manmade objects. ACM Transactions on Graphics, 27(3): 42:1–42:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fleishman:2005:RML

[FCOS05] Shachar Fleishman, Daniel Cohen-Or, and Cláudio T. Silva. Robust moving leastsquares fitting with sharp features. ACM Transactions on Graphics, 24(3):544–552, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fu:2017:ASI

[FCW^{+}17] Qiang Fu, Xiaowu Chen, Xiaotian Wang, Sijia Wen, Bin Zhou, and Hongbo Fu. Adaptive synthesis of indoor scenes via activity-associated object relation graphs. ACM Transactions on Graphics, 36(6): 201:1–201:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Frederickx:2017:FSD

[FD17] Roald Frederickx and Philip Dutré. A forward scattering dipole model from a functional integral approximation. ACM Transactions on Graphics, 36(4):109:1–109:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fang:2022:TFR

[FDBH22] Xianzhong Fang, Mathieu Desbrun, Hujun Bao, and Jin Huang. TopoCut: fast and robust planar cutting of arbitrary domains. ACM Transactions on Graphics, 41(4):40:1–40:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530149.

Fleishman:2003:BMD

[FDCO03] Shachar Fleishman, Iddo Drori, and Daniel Cohen-Or. Bilateral mesh denoising. ACM Transactions on Graphics, 22(3):950–953, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fournier:1988:PFB

[FF88] Alain Fournier and Donald Fussell. On the power of the frame buffer. ACM Transactions on Graphics, 7(2):103–128, April 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
42460.html.

Freedman:2011:IVU

[FF11] Gilad Freedman and Raanan Fattal. Image and video upscaling from local selfexamples. ACM Transactions on Graphics, 30(2):12:1– 12:11, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fisher:2009:DPC

[FFB^{+}09] Matthew Fisher, Kayvon Fatahalian, Solomon Boulos, Kurt Akeley, William R. Mark, and Pat Hanrahan. DiagSplit: parallel, crack-free, adaptive tessellation for micropolygon rendering. ACM Transactions on Graphics, 28 (5):150:1–150:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2021:LAD

[FFBB21] Yao Feng, Haiwen Feng, Michael J. Black, and Timo Bolkart. Learning an animatable detailed 3D face model from in-the-wild images. ACM Transactions on Graphics, 40 (4):88:1–88:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459936.

Farbman:2010:DME

[FFL10] Zeev Farbman, Raanan Fattal, and Dani Lischinski. Diffusion maps for edge-aware image editing. ACM Transactions on Graphics, 29(6): 145:1–145:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Farbman:2011:CP

[FFL11] Zeev Farbman, Raanan Fattal, and Dani Lischinski. Convolution pyramids. ACM Transactions on Graphics, 30 (6):175:1–175:??, December 2011. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Farbman:2008:EPD

[FFLS08] Zeev Farbman, Raanan Fattal, Dani Lischinski, and Richard Szeliski. Edgepreserving decompositions for multi-scale tone and detail manipulation. ACM Transactions on Graphics, 27(3): 67:1–67:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

TOG-147190061

[FFWL^{+}22] José Antonio Fernández-Fernández, Lukas Westhofen, Fabian Löschner, Stefan Rhys Jeske, Andreas Longva, and Jan Bender. Fast octree neighborhood search for SPH simulations. ACM Transactions on Graphics, 41(6): 242:1–242:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555523.

Ferguson:1990:CSI

[FG90] David R. Ferguson and Thomas A. Grandine. On the construction of surface interpolating curves: I. A method for handling nonconstant parameter curves. ACM Transactions on Graphics, 9(2):212–225, April 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        78961.html.

Fuhrmann:2011:FDM

[FG11] Simon Fuhrmann and Michael Goesele. Fusion of depth maps with multiple scales. ACM Transactions on Graphics, 30 (6):148:1–148:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fuhrmann:2014:FSS

[FG14] Simon Fuhrmann and Michael Goesele. Floating scale surface reconstruction. ACM Transactions on Graphics, 33(4): 46:1–46:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Faure:2011:SMM

[FGBP11] François Faure, Benjamin Gilles, Guillaume Bousquet, and Dinesh K. Pai. Sparse meshless models of complex deformable solids. ACM Transactions on Graphics, 30 (4):73:1–73:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2023:WND

[FGC23] Nicole Feng, Mark Gillespie, and Keenan Crane. Winding numbers on discrete surfaces. ACM Transactions on Graphics, 42(4):36:1–36:??, August 2023. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592401.

Fu:2017:PPC

[FGG^{+}17] Chuyuan Fu, Qi Guo, Theodore Gast, Chenfanfu Jiang, and Joseph Teran. A polynomial particle-in-cell method. ACM Transactions on Graphics, 36 (6):222:1–222:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Forrest:1984:GEI

[FGN84] Robin Forrest, Leo Guibas, and Jurg Nievergelt. Guest Editor’s introduction to special issue on computational geometry. ACM Transactions on Graphics, 3(4):241– 243, October 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fei:2021:RIM

[FGW^{+}21] Yun (Raymond) Fei, Qi Guo, Rundong Wu, Li Huang, and Ming Gao. Revisiting integration in the material point method: a scheme for easier separation and less dissipation. ACM Transactions on Graphics, 40(4): 109:1–109:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459678.

Fellner:1993:RRG

[FH93] Dieter W. Fellner and Christoph Helmberg. Robust rendering of general ellipses and elliptical arcs. ACM Transactions on Graphics, 12(3): 251–276, July 1993. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        169704.html.

Fudos:1997:GCA

[FH97] Ioannis Fudos and Christoph M. Hoffmann. A graph-constructive approach to solving systems of geometric constraints. ACM Transactions on Graphics, 16 (2):179–216, April 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1997-16-2/
                                        p179-fudos/.

Fang:2004:TTS

[FH04a] Hui Fang and John C. Hart. Textureshop: texture synthesis as a photograph editing tool. ACM Transactions on Graphics, 23(3):354– 359, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fogarty:2004:GTO

[FH04b] James Fogarty and Scott E. Hudson. GADGET: a toolkit for optimization-based approaches to interface and display generation. ACM Transactions on Graphics, 23(3): 730, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fang:2007:DPS

[FH07] Hui Fang and John C. Hart.

Detail preserving shape deformation in image editing. ACM Transactions on Graphics, 26 (3):12:1–12:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fisher:2010:CBS

[FH10] Matthew Fisher and Pat Hanrahan. Context-based search for 3D models. ACM Transactions on Graphics, 29(6): 182:1–182:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:2011:SMC

[FH11] Tim Foley and Pat Hanrahan.

Spark: modular, composable shaders for graphics hardware. ACM Transactions on Graphics, 30(4):107:1–107:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2023:MPG

[FHG^{+}23] Zhimin Fan, Pengpei Hong, Jie Guo, Changqing Zou, Yanwen Guo, and Ling-Qi Yan.

Manifold path guiding for importance sampling specular chains. ACM Transactions on Graphics, 42(6): 257:1–257:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618360.

Frisvad:2014:DDM

[FHK14] Jeppe Revall Frisvad, Toshiya Hachisuka, and Thomas Kim Kjeldsen. Directional dipole model for subsurface scattering. ACM Transactions on Graphics, 34(1):5:1–5:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Farbman:2009:CII

[FHL^{+}09] Zeev Farbman, Gil Hoffer, Yaron Lipman, Daniel Cohen-Or, and Dani Lischinski. Coordinates for instant image cloning. ACM Transactions on Graphics, 28(3): 67:1–67:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fascione:2018:MBS

[FHL^{+}18] Luca Fascione, Johannes Hanika, Mark Leone, Marc Droske, Jorge Schwarzhaupt, Tomás Davidovic, Andrea Weidlich, and Johannes Meng. Manuka: a batch-shading architecture for spectral path tracing in movie production. ACM Transactions on Graphics, 37(3):31:1– 31:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3182161.

Flynn:2021:GFC

[FHM^{+}21] Sean Flynn, David Hart, Bryan Morse, Seth Holladay, and Parris Egbert. Generalized fluid carving with fast lattice-guided seam computation. ACM Transactions on Graphics, 40(6): 255:1–255:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480544.

Feng:2022:LBB

[FHXW22] Xudong Feng, Wenchao Huang, Weiwei Xu, and Huamin Wang. Learning-based bending stiffness parameter estimation by a drape tester. ACM Transactions on Graphics, 41(6):221:1–221:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555464.

Field:1985:ILI

[Fie85] Dan Field. Incremental linear interpolation. ACM Transactions on Graphics, 4(1):

1–11, January 1985. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        3976.html.

Filip:1989:BPS

[Fil89] Daniel J. Filip. Blending parametric surfaces. ACM Transactions on Graphics, 8 (3):164–173, July 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        77057.html.

Fiume:2000:AFA

[Fiu00]  Eugene Fiume.  Alain

Fournier: 1943–2000: An appreciation. ACM Transactions on Graphics, 19(4):243– 245, October 2000. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/2000-19-4/
                                        p243-fiume/.

Fyffe:2014:DHR

[FJA^{+}14] Graham Fyffe, Andrew Jones, Oleg Alexander, Ryosuke Ichikari, and Paul Debevec. Driving high-resolution facial scans with video performance capture. ACM Transactions on Graphics, 34(1): 8:1–8:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fiser:2016:SIG

[FJL^{+}16] Jakub Fiser, Ondrej Jamriska, Michal Lukác, Eli Shechtman, Paul Asente, Jingwan Lu, and Daniel Sýkora. StyLit: illumination-guided example-based stylization of 3D renderings. ACM Transactions on Graphics, 35(4): 92:1–92:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fiser:2017:EBS

[FJS^{+}17] Jakub Fiser, Ondrej Jamriska, David Simons, Eli Shechtman, Jingwan Lu, Paul Asente, Michal Lukác, and Daniel Sýkora. Example-based synthesis of stylized facial animations. ACM Transactions on Graphics, 36(4): 155:1–155:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fanello:2014:LDC

[FKI^{+}14] Sean Ryan Fanello, Cem Keskin, Shahram Izadi, Pushmeet Kohli, David Kim, David Sweeney, Antonio Criminisi, Jamie Shotton, Sing Bing Kang, and Tim Paek. Learning to be a depth camera for close-range human capture and interaction. ACM Transactions on Graphics, 33 (4):86:1–86:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fukiage:2017:HPB

[FKN17] Taiki Fukiage, Takahiro Kawabe, and Shin’ya Nishida. Hiding of phase-based stereo disparity for ghost-free viewing without glasses. ACM Transactions on Graphics, 36(4): 147:1–147:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Funkhouser:2004:ME

[FKS^{+}04] Thomas Funkhouser, Michael Kazhdan, Philip Shilane, Patrick Min, William Kiefer, Ayellet Tal, Szymon Rusinkiewicz, and David Dobkin. Modeling by example. ACM Transactions on Graphics, 23(3): 652–663, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2008:RTD

[FKY08] Wei-Wen Feng, Byung-Uck Kim, and Yizhou Yu. Realtime data driven deformation using kernel canonical correlation analysis. ACM Transactions on Graphics, 27(3): 91:1–91:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2010:FPT

[FKY^{+}10] Wei-Wen Feng, Byung-Uck Kim, Yizhou Yu, Liang Peng, and John Hart. Featurepreserving triangular geometry images for level-of-detail representation of static and skinned meshes. ACM Transactions on Graphics, 29(2): 11:1–11:13, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2004:TDS

[FL04] Raanan Fattal and Dani Lischinski. Target-driven smoke animation. ACM Transactions on Graphics, 23 (3):441–448, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Farbman:2011:TSV

[FL11] Zeev Farbman and Dani Lischinski. Tonal stabilization of video. ACM Transactions on Graphics, 30(4): 89:1–89:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fu:2016:CIF

[FL16] Xiao-Ming Fu and Yang Liu. Computing inversionfree mappings by simplex assembly. ACM Transactions on Graphics, 35(6):216:1– 216:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Favreau:2016:FVS

[FLB16] Jean-Dominique Favreau, Florent Lafarge, and Adrien Bousseau. Fidelity vs. simplicity: a global approach to line drawing vectorization. ACM Transactions on Graphics, 35 (4):120:1–120:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Favreau:2017:PIA

[FLB17] Jean-Dominique Favreau, Florent Lafarge, and Adrien Bousseau. Photo2clipart: image abstraction and vectorization using layered linear gradients. ACM Transactions on Graphics, 36(6): 180:1–180:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fu:2015:CLI

[FLG15] Xiao-Ming Fu, Yang Liu, and Baining Guo. Computing locally injective mappings by advanced MIPS. ACM Transactions on Graphics, 34(4): 71:1–71:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fang:2019:SRI

[FLGJ19] Yu Fang, Minchen Li, Ming Gao, and Chenfanfu Jiang. Silly rubber: an implicit material point method for simulating non-equilibrated viscoelastic and elastoplastic solids. ACM Transactions on Graphics, 38(4): 118:1–118:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fu:2010:STS

[FLHCO10] Chi-Wing Fu, Chi-Fu Lai, Ying He, and Daniel Cohen-Or. K-set tilable surfaces. ACM Transactions on Graphics, 29(4):44:1–44:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fang:2021:GGI

[FLJK21] Yu Fang, Minchen Li, Chenfanfu Jiang, and Danny M. Kaufman. Guaranteed globally injective 3D deformation processing. ACM Transactions on Graphics, 40(4): 75:1–75:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459757.

Fan:2013:ELS

[FLLP13] Ye Fan, Joshua Litven, David I. W. Levin, and Dinesh K. Pai. Eulerian-on-Lagrangian simulation. ACM Transactions on Graphics, 32(3): 22:1–22:9, June 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2014:AVM

[FLP14] Ye Fan, Joshua Litven, and Dinesh K. Pai. Active volumetric musculoskeletal systems. ACM Transactions on Graphics, 33(4):152:1– 152:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ferguson:2021:IFR

[FLS^{+}21] Zachary Ferguson, Minchen Li, Teseo Schneider, Francisca Gil-Ureta, Timothy Langlois, Chenfanfu Jiang, Denis Zorin, Danny M. Kaufman, and Daniele Panozzo. Intersection-free rigid body dynamics. ACM Transactions on Graphics, 40(4): 183:1–183:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459802.

Fu:2014:ASM

[FLSG14] Xiao-Ming Fu, Yang Liu, John Snyder, and Baining Guo. Anisotropic simplicial meshing using local convex functions. ACM Transactions on Graphics, 33(6): 182:1–182:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fattal:2002:GDH

[FLW02] Raanan Fattal, Dani Lischinski, and Michael Werman.

Gradient domain high dynamic range compression. ACM Transactions on Graphics, 21(3):249–256, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fournier:1984:TSP

[FM84] A. Fournier and D. Y. Montuno. Triangulating simple polygons and equivalent problems. ACM Transactions on Graphics, 3(2):153–174, April 1984. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fei:2017:MSM

[FMB^{+}17] Yun (Raymond) Fei, Henrique Teles Maia, Christopher Batty, Changxi Zheng, and Eitan Grinspun. A multi-scale model for simulating liquidhair interactions. ACM Transactions on Graphics, 36(4): 56:1–56:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Funkhouser:2003:SEM

[FMK^{+}03] Thomas Funkhouser, Patrick Min, Michael Kazhdan, Joyce Chen, Alex Halderman, David Dobkin, and David Jacobs. A search engine for 3D models. ACM Transactions on Graphics, 22(1):83–105, January 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2014:SCF

[FMLW14] Lubin Fan, Przemyslaw Musialski, Ligang Liu, and Peter Wonka. Structure completion for facade layouts. ACM Transactions on Graphics, 33 (6):210:1–210:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fumero:2020:NSG

[FMR20] Marco Fumero, Michael Möller, and Emanuele Rodolà. Nonlinear spectral geometry processing via the TV transform. ACM Transactions on Graphics, 39(6):199:1– 199:16, November 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417849.

Forootaninia:2020:FDS

[FN20] Zahra Forootaninia and Rahul Narain. Frequency-domain smoke guiding. ACM Transactions on Graphics, 39(6): 172:1–172:10, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417842.

Farouki:1989:APD

[FNO89] R. T. Farouki, C. A. Neff, and M. A. O’Connor. Automatic parsing of degenerate quadric-surface intersections. ACM Transactions on Graphics, 8(3):174–203, July 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77058.html.

Feiner:1982:ESC

[FNvD82] Steven Feiner, Sandor Nagy, and Andries van Dam. An experimental system for creating and presenting interactive graphical documents. ACM Transactions on Graphics, 1 (1):59–77, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feldman:2003:ASP

[FOA03] Bryan E. Feldman, James F.

O’Brien, and Okan Arikan. Animating suspended particle explosions. ACM Transactions on Graphics, 22(3): 708–715, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feldman:2005:AGH

[FOK05] Bryan E. Feldman, James F.

O’Brien, and Bryan M. Klingner. Animating gases with hybrid meshes. ACM Transactions on Graphics, 24 (3):904–909, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1986:GEIa

[Fol86a] James Foley. Guest Editor’s introduction: Special issue on user interface software. ACM Transactions on Graphics, 5 (2):75–78, April 1986. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1986:GEIb

[Fol86b] James Foley. Guest Editor’s introduction: Special issue on user interface software. ACM Transactions on Graphics, 5 (3):175–178, July 1986. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1986:GEIc

[Fol86c] James Foley. Guest Editor’s introduction: Special issue on user interface software. ACM Transactions on Graphics, 5(4):279–282, October 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1987:WBS

[Fol87] Thomas A. Foley. Weighted bicubic spline interpolation to rapidly varying data. ACM Transactions on Graphics, 6(1):1–18, January 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        27626.html.

Foley:1991:ELB

[Fol91] Jim Foley. Editorial: Looking back, looking ahead. ACM Transactions on Graphics, 10 (4):321–322, October 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1992:E

[Fol92] Jim Foley. Editorial. ACM Transactions on Graphics, 11 (4):297–298, October 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1994:SC

[Fol94] Jim Foley. Scope and charter.

ACM Transactions on Graphics, 13(1):1, January 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1995:E

[Fol95a] Jim Foley. Editorial. ACM Transactions on Graphics, 14 (3):201, July 1995. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Foley:1995:SC

[Fol95b] Jim Foley. Scope and charter. ACM Transactions on Graphics, 14(1):1–2, January 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fang:2021:CSC

[FOL^{+}21] Qing Fang, Wenqing Ouyang, Mo Li, Ligang Liu, and Xiao-Ming Fu. Computing sparse cones with bounded distortion for conformal parameterizations. ACM Transactions on Graphics, 40(6): 262:1–262:9, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480526.

Fang:2003:ESP

[FP03] Anthony C. Fang and Nancy S. Pollard. Efficient synthesis of physically valid human motion. ACM Transactions on Graphics, 22(3):417–426, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fish:2020:SSS

[FPBCO20] Noa Fish, Lilach Perry, Amit Bermano, and Daniel Cohen-Or. SketchPatch: sketch stylization via seamless patchlevel synthesis. ACM Transactions on Graphics, 39(6): 227:1–227:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417816.

Fanni:2022:PDP

[FPSG22] Filippo Andrea Fanni, Fabio Pellacini, Riccardo Scateni, and Andrea Giachetti. PAVEL:

Decorative patterns with packed volumetric elements. ACM Transactions on Graphics, 41(2):19:1–19:15, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3502802.

Fang:2020:IMI

[FQL^{+}20] Yu Fang, Ziyin Qu, Minchen Li, Xinxin Zhang, Yixin Zhu, Mridul Aanjaneya, and Chenfanfu Jiang. IQ-MPM: an interface quadrature material point method for nonsticky strongly two-way coupled nonlinear solids and fluids. ACM Transactions on Graphics, 39(4):51:1– 51:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392438.

Fournier:1987:GEI

[FR87] A. Fournier and W. T. Reeves.

Guest Editors’ introduction. ACM Transactions on Graphics, 6(3):165–166, July 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fischer:2022:MML

[FR22] Michael Fischer and Tobias Ritschel. Metappearance: Meta-learning for visual appearance reproduction. ACM Transactions on Graphics, 41 (6):245:1–245:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555458.

Freeman:2016:DAF

[Fre16] Bill Freeman. The diffractive achromat full spectrum computational imaging with diffractive optics. ACM Transactions on Graphics, 35 (4):31:1–31:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fisher:2012:EBS

[FRS^{+}12] Matthew Fisher, Daniel Ritchie, Manolis Savva, Thomas Funkhouser, and Pat Hanrahan. Examplebased synthesis of 3D object arrangements. ACM Transactions on Graphics, 31(6): 135:1–135:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Friston:2019:PRH

[FRS19] Sebastian Friston, Tobias Ritschel, and Anthony Steed. Perceptual rasterization for head-mounted display image synthesis. ACM Transactions on Graphics, 38(4): 97:1–97:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fuchs:2008:TPR

[FRSL08] Martin Fuchs, Ramesh Raskar, Hans-Peter Seidel, and Hendrik P. A. Lensch. Towards passive 6D reflectance field displays. ACM Transactions on Graphics, 27(3): 58:1–58:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fisher:2007:DTV

[FSDH07] Matthew Fisher, Peter Schröder, Mathieu Desbrun, and Hugues Hoppe. Design of tangent vector fields. ACM Transactions on Graphics, 26(3): 56:1–56:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fried:2016:PAM

[FSGF16] Ohad Fried, Eli Shechtman, Dan B. Goldman, and Adam Finkelstein. Perspectiveaware manipulation of portrait photos. ACM Transactions on Graphics, 35(4): 128:1–128:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fergus:2006:RCS

[FSH^{+}06] Rob Fergus, Barun Singh, Aaron Hertzmann, Sam T. Roweis, and William T. Freeman. Removing camera shake from a single photograph. ACM Transactions on Graphics, 25(3):787–794, July 2006.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Finch:2011:FVG

[FSH11a] Mark Finch, John Snyder, and Hugues Hoppe. Freeform vector graphics with controlled thin-plate splines. ACM Transactions on Graphics, 30 (6):166:1–166:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fisher:2011:CSR

[FSH11b] Matthew Fisher, Manolis Savva, and Pat Hanrahan. Characterizing structural relationships in scenes using graph kernels. ACM Transactions on Graphics, 30(4): 34:1–34:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Friedel:2004:VNM

[FSK04] Ilja Friedel, Peter Schröder, and Andrei Khodakovsky. Variational normal meshes. ACM Transactions on Graphics, 23(4):1061–1073, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ferguson:2023:TRC

[FSKP23] Zachary Ferguson, Teseo Schneider, Danny Kaufman, and Daniele Panozzo. Intimestep remeshing for contacting elastodynamics. ACM Transactions on Graphics, 42(4):145:1–145:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592428.

Fisher:2015:ACS

[FSL^{+}15] Matthew Fisher, Manolis Savva, Yangyan Li, Pat Hanrahan, and Matthias Nießner. Activity-centric scene synthesis for functional 3D scene modeling. ACM Transactions on Graphics, 34(6): 179:1–179:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fouladi:2022:RLL

[FSP^{+}22] Sadjad Fouladi, Brennan Shacklett, Fait Poms, Arjun Arora, Alex Ozdemir, Deepti Raghavan, Pat Hanrahan, Kayvon Fatahalian, and Keith Winstein. R2E2: low-latency path tracing of terabyte-scale scenes using thousands of cloud CPUs. ACM Transactions on Graphics, 41(4): 76:1–76:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530171.

Freiwald:2022:CLR

[FSRS22] Jann Philipp Freiwald, Susanne Schmidt, Bernhard E. Riecke, and Frank Steinicke.

The continuity of locomotion: Rethinking conventions for locomotion and its visualization in shared virtual reality spaces. ACM Transactions on Graphics, 41(6): 211:1–211:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555522.

Fu:2015:CIF

[FSY^{+}15] Chi-Wing Fu, Peng Song, Xiaoqi Yan, Lee Wei Yang, Pradeep Kumar Jayaraman, and Daniel Cohen-Or. Computational interlocking furniture assembly. ACM Transactions on Graphics, 34(4): 91:1–91:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2021:QZS

[FTD21] Leman Feng, Yiying Tong, and Mathieu Desbrun. Q-zip: singularity editing primitive for quad meshes. ACM Transactions on Graphics, 40(6): 258:1–258:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480523.

Freeman:2003:LST

[FTP03] William T. Freeman, Joshua B. Tenenbaum, and Egon C. Pasztor. Learning style translation for the lines of a drawing. ACM Transactions on Graphics, 22(1): 33–46, January 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fratarcangeli:2016:VPG

[FTP16] Marco Fratarcangeli, Valentina Tibaldo, and Fabio Pellacini. Vivace: a practical Gauss– Seidel method for stable soft body dynamics. ACM Transactions on Graphics, 35(6): 214:1–214:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fried:2019:TBE

[FTZ^{+}19] Ohad Fried, Ayush Tewari, Michael Zollhöfer, Adam Finkelstein, Eli Shechtman, Dan B. Goldman, Kyle Genova, Zeyu Jin, Christian Theobalt, and Maneesh Agrawala. Text-based editing of talking-head video. ACM Transactions on Graphics, 38 (4):68:1–68:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fuchs:1982:GEI

[Fuc82] Henry Fuchs. Guest Editor’s introduction. ACM Transactions on Graphics, 1 (1):5–6, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fortune:1996:SAY

[FV96] Steven Fortune and Christopher J. Van Wyk. Static analysis yields efficient exact integer arithmetic for computational geometry. ACM Transactions on Graphics, 15(3):223–248, July 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        230533.html.

Fish:2016:SON

[FvKBCO16] Noa Fish, Oliver van Kaick, Amit Bermano, and Daniel Cohen-Or. Structure-oriented networks of shape collections. ACM Transactions on Graphics, 35(6):171:1– 171:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2012:DBL

[FW12] Powei Feng and Joe Warren. Discrete bi-Laplacians and biharmonic B-splines. ACM Transactions on Graphics, 31 (4):115:1–115:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2016:PME

[FW16] Lubin Fan and Peter Wonka. A probabilistic model for exteriors of residential buildings. ACM Transactions on Graphics, 35(5):155:1– 155:??, September 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fargion:2022:GIF

[FW22] Guy Fargion and Ofir Weber. Globally injective flattening via a reduced harmonic subspace. ACM Transactions on Graphics, 41(6): 253:1–253:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555449.

Fang:2016:AHM

[FXBH16] Xianzhong Fang, Weiwei Xu, Hujun Bao, and Jin Huang. All-hex meshing using closedform induced polycube. ACM Transactions on Graphics, 35 (4):124:1–124:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2010:DTR

[FYK10] Wei-Wen Feng, Yizhou Yu, and Byung-Uck Kim. A deformation transformer for realtime cloth animation. ACM Transactions on Graphics, 29 (4):108:1–108:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2018:ISU

[FYW^{+}18] Qingnan Fan, Jiaolong Yang, David Wipf, Baoquan Chen, and Xin Tong. Image smoothing via unsupervised learning. ACM Transactions on Graphics, 37(6):259:1– 259:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Feng:2016:CDM

[FYY^{+}16] Tian Feng, Lap-Fai Yu, Sai-Kit Yeung, KangKang Yin, and Kun Zhou. Crowd-driven mid-scale layout design. ACM Transactions on Graphics, 35 (4):132:1–132:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2016:AVP

[FZBR16] Xinyi Fan, Linguang Zhang, Benedict Brown, and Szymon Rusinkiewicz. Automated view and path planning for scalable multi-object 3D scanning. ACM Transactions on Graphics, 35(6): 239:1–239:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fan:2015:JNS

[FZL^{+}15] Qingnan Fan, Fan Zhong, Dani Lischinski, Daniel Cohen-Or, and Baoquan Chen. JumpCut: non-successive mask transfer and interpolation for video cutout. ACM Transactions on Graphics, 34 (6):195:1–195:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fu:2011:ACL

[FZLM11] Hongbo Fu, Shizhe Zhou, Ligang Liu, and Niloy J. Mitra. Animated construction of line drawings. ACM Transactions on Graphics, 30(6): 133:1–133:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Fang:2020:RFM

[FZZ^{+}20] Guoxin Fang, Tianyu Zhang, Sikai Zhong, Xiangjia Chen, Zichun Zhong, and Charlie C. L. Wang. Reinforced FDM: multi-axis filament alignment with controlled anisotropic strength. ACM Transactions on Graphics, 39(6): 204:1–204:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417834.

Gruber:2020:CWF

[GA20] Anthony Gruber and Eugenio Aulisa. Computational p-Willmore flow with conformal penalty. ACM Transactions on Graphics, 39(5):161:1– 161:16, September 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3369387.

Gal:2023:EBD

[GAA^{+}23] Rinon Gal, Moab Arar, Yuval Atzmon, Amit H. Bermano, Gal Chechik, and Daniel Cohen-Or. Encoder-based domain tuning for fast personalization of text-to-image models. ACM Transactions on Graphics, 42(4): 150:1–150:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592133.

Goldade:2020:CBA

[GAB20] Ryan Goldade, Mridul Aanjaneya, and Christopher Batty. Constraint bubbles and affine regions: reduced fluid models for efficient immersed bubbles and flexible spatial coarsening. ACM Transactions on Graphics, 39 (4):43:1–43:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392455.

Goesele:2010:APC

[GAF^{+}10] Michael Goesele, Jens Ackermann, Simon Fuhrmann, Carsten Haubold, and Ronny Klowsky. Ambient point clouds for view interpolation. ACM Transactions on Graphics, 29(4):95:1–95:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garces:2014:SMI

[GAGH14] Elena Garces, Aseem Agarwala, Diego Gutierrez, and Aaron Hertzmann. A similarity measure for illustration style. ACM Transactions on Graphics, 33(4): 93:1–93:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gallier:1999:SMD

[Gal99]  Jean Gallier.  A simple

method for drawing a rational curve as two Bézier segments. ACM Transactions on Graphics, 18(4):316–328, October 1999. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
1999-18-4/p316-gallier/.

Grabler:2009:GPM

[GAL^{+}09] Floraine Grabler, Maneesh Agrawala, Wilmot Li, Mira Dontcheva, and Takeo Igarashi. Generating photo manipulation tutorials by demonstration. ACM Transactions on Graphics, 28(3):66:1– 66:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garces:2023:TMD

[GARP^{+}23] Elena Garces, Victor Arellano, Carlos Rodriguez-Pardo, David Pascual-Hernandez, Sergio Suja, and Jorge Lopez-Moreno. Towards material digitization with a dual-scale optical system. ACM Transactions on Graphics, 42(4): 152:1–152:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592147.

Grabler:2008:AGT

[GASP08] Floraine Grabler, Maneesh Agrawala, Robert W. Sumner, and Mark Pauly. Automatic generation of tourist maps. ACM Transactions on Graphics, 27(3):100:1– 100:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gain:2008:SSD

[GB08a] James Gain and Dominique Bechmann. A survey of spatial deformation from a usercentered perspective. ACM Transactions on Graphics, 27(4):107:1–107:32, October 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Grundhofer:2008:VDV

[GB08b] Anselm Grundhöfer and Oliver Bimber. VirtualStudio2Go: digital video composition for real environments. ACM Transactions on Graphics, 27 (5):151:1–151:??, December 2008. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Gerszewski:2013:PBA

[GB13] Dan Gerszewski and Adam W.

Bargteil. Physics-based animation of large-scale splashing liquids. ACM Transactions on Graphics, 32(6): 185:1–185:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gribel:2011:HQS

[GBAM11] Carl Johan Gribel, Rasmus Barringer, and Tomas Akenine-Möller. High-quality spatio-temporal rendering using semi-analytical visibility. ACM Transactions on Graphics, 30(4):54:1–54:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gourmel:2013:GBI

[GBC^{+}13] Olivier Gourmel, Loic Barthe, Marie-Paule Cani, Brian Wyvill, Adrien Bernhardt, Mathias Paulin, and Herbert Grasberger. A gradient-based implicit blend. ACM Transactions on Graphics, 32(2): 12:1–12:12, April 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guendelman:2003:NRB

[GBF03] Eran Guendelman, Robert Bridson, and Ronald Fedkiw. Nonconvex rigid bodies with stacking. ACM Transactions on Graphics, 22(3): 871–878, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gilles:2011:FBE

[GBFP11] Benjamin Gilles, Guillaume Bousquet, François Faure, and Dinesh K. Pai. Frame-based elastic models. ACM Transactions on Graphics, 30(2): 15:1–15:12, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guthe:2005:GBT

[GBK05] Michael Guthe, Aákos Balázs, and Reinhard Klein. GPUbased trimming and tessellation of NURBS and T-Spline surfaces. ACM Transactions on Graphics, 24(3): 1016–1023, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guerrero:2016:PEP

[GBLM16] Paul Guerrero, Gilbert Bernstein, Wilmot Li, and Niloy J. Mitra. PATEX: exploring pattern variations. ACM Transactions on Graphics, 35(4): 48:1–48:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goktekin:2004:MAV

[GBO04] Tolga G. Goktekin, Adam W.

Bargteil, and James F. O’Brien. A method for animating viscoelastic fluids. ACM Transactions on Graphics, 23(3):463–468, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gharbi:2017:DBL

[GCB^{+}17] Michaël Gharbi, Jiawen Chen, Jonathan T. Barron, Samuel W. Hasinoff, and Frédo Durand. Deep bilateral learning for real-time image enhancement. ACM Transactions on Graphics, 36(4):118:1–118:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2020:DNL

[GCD^{+}20] Duan Gao, Guojun Chen, Yue Dong, Pieter Peers, Kun Xu, and Xin Tong. Deferred neural lighting: free-viewpoint relighting from unstructured photographs. ACM Transactions on Graphics, 39(6): 258:1–258:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417767.

Guo:2019:FGF

[GCH^{+}19] Jie Guo, Yanjun Chen, Bingyang Hu, Ling-Qi Yan, Yanwen Guo, and Yuntao Liu. Fractional Gaussian fields for modeling and rendering of spatially-correlated media. ACM Transactions on Graphics, 38(4):45:1–45:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gal:2006:SGF

[GCO06] Ran Gal and Daniel Cohen-Or. Salient geometric features for partial shape matching and similarity. ACM Transactions on Graphics, 25(1): 130–150, January 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ghosh:2010:CPS

[GCP^{+}10] Abhijeet Ghosh, Tongbo Chen, Pieter Peers, Cyrus A. Wilson, and Paul Debevec. Circularly polarized spherical illumination reflectometry. ACM Transactions on Graphics, 29(6):162:1– 162:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gharbi:2016:DJD

[GCPD16] Michaël Gharbi, Gaurav Chaurasia, Sylvain Paris, and Frédo Durand. Deep joint demosaicking and denoising. ACM Transactions on Graphics, 35(6):191:1– 191:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guay:2013:LAI

[GCR13] Martin Guay, Marie-Paule Cani, and Rémi Ronfard. The line of action: an intuitive interface for expressive character posing. ACM Transactions on Graphics, 32(6): 205:1–205:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldman:2006:SSV

[GCSS06] Dan B. Goldman, Brian Curless, David Salesin, and Steven M. Seitz. Schematic storyboarding for video visualization and editing. ACM Transactions on Graphics, 25 (3):862–871, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gandoin:2002:PLC

[GD02] Pierre-Marie Gandoin and Olivier Devillers. Progressive lossless compression of arbitrary simplicial complexes. ACM Transactions on Graphics, 21(3):372–379, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Granier:2004:FRA

[GD04] Xavier Granier and George Drettakis. A final reconstruction approach for a unified global illumination algorithm. ACM Transactions on Graphics, 23(2):163–189, April 2004.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garcia-Dorado:2017:FWSa

[GDAB^{+}17a] Ignacio Garcia-Dorado, Daniel G. Aliaga, Saiprasanth Bhalachandran, Paul Schmid, and Dev Niyogi. Fast weather simulation for inverse procedural design of 3D urban models. ACM Transactions on Graphics, 36(2):21:1–21:??, April 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garcia-Dorado:2017:FWSb

[GDAB^{+}17b] Ignacio Garcia-Dorado, Daniel G. Aliaga, Saiprasanth Bhalachandran, Paul Schmid, and Dev Niyogi. Fast weather simulation for inverse procedural design of 3D urban models. ACM Transactions on Graphics, 36(4):133:1–133:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2015:HMR

[GDC15] Xifeng Gao, Zhigang Deng, and Guoning Chen. Hexahedral mesh re-parameterization from aligned base-complex. ACM Transactions on Graphics, 34(4):142:1–142:??, August 2015. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guerin:2017:IEB

[GDG^{+}17] Éric Guérin, Julie Digne, Éric Galin, Adrien Peytavie, Christian Wolf, Bedrich Benes, and Benoît Martinez. Interactive example-based terrain authoring with conditional generative adversarial networks. ACM Transactions on Graphics, 36(6): 228:1–228:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

DeBry:2002:PRT

[gDGPR02] David (grue) DeBry, Jonathan Gibbs, Devorah DeLeon Petty, and Nate Robins. Painting and rendering textures on unparameterized models. ACM Transactions on Graphics, 21(3): 763–768, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garrett:1982:GPU

[GF82] M. T. Garrett and J. D. Foley.

Graphics programming using a database system with dependency declarations. ACM Transactions on Graphics, 1 (2):109–128, April 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Golovinskiy:2008:RCM

[GF08] Aleksey Golovinskiy and Thomas Funkhouser. Randomized cuts for 3D mesh analysis. ACM Transactions on Graphics, 27(5):145:1– 145:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldstein:2012:VSU

[GF12] Amit Goldstein and Raanan Fattal. Video stabilization using epipolar geometry. ACM Transactions on Graphics, 31(5):126:1–126:10, August 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guenter:2012:FG

[GFD^{+}12] Brian Guenter, Mark Finch, Steven Drucker, Desney Tan, and John Snyder. Foveated 3D graphics. ACM Transactions on Graphics, 31(6): 164:1–164:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Grandia:2023:DDO

[GFK^{+}23] Ruben Grandia, Farbod Farshidian, Espen Knoop, Christian Schumacher, Marco Hutter, and Moritz Bächer. DOC: Differentiable optimal control for retargeting motions onto legged robots. ACM Transactions on Graphics, 42 (4):96:1–96:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592454.

Gauthier:2022:MNN

[GFL^{+}22] Alban Gauthier, Robin Faury, Jérémy Levallois, Théo Thonat, Jean-Marc Thiery, and Tamy Boubekeur. MIPNet: Neural normal-to-anisotropic-roughness MIP mapping. ACM Transactions on Graphics, 41(6):246:1–246:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555487.

Glassner:1995:DDR

[GFMS95] A. S. Glassner, K. P. Fishkin, D. H. Marimont, and M. C. Stone. Device-directed rendering. ACM Transactions on Graphics, 14(1):58– 76, January 1995. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
200977.html.

Ghosh:2011:MFC

[GFT^{+}11] Abhijeet Ghosh, Graham Fyffe, Borom Tunwattanapong, Jay Busch, Xueming Yu, and Paul Debevec. Multiview face capture using polarized spherical gradient illumination. ACM Transactions on Graphics, 30(6): 129:1–129:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guennebaud:2007:APS

[GG07] Gaël Guennebaud and Markus Gross. Algebraic point set surfaces. ACM Transactions on Graphics, 26(3): 23:1–23:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Genevaux:2013:TGU

[GGG^{+}13] Jean-David Génevaux, Éric Galin, Eric Guérin, Adrien Peytavie, and Bedrich Benes. Terrain generation using procedural models based on hydrology. ACM Transactions on Graphics, 32(4):143:1– 143:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gu:2002:GI

[GGH02] Xianfeng Gu, Steven J. Gortler, and Hugues Hoppe. Geometry images. ACM Transactions on Graphics, 21 (3):355–361, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goesele:2003:ALS

[GGHS03] Michael Goesele, Xavier Granier, Wolfgang Heidrich, and Hans-Peter Seidel. Accurate light source acquisition and rendering. ACM Transactions on Graphics, 22 (3):621–630, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gamboa:2018:SAF

[GGN18] Luis E. Gamboa, Jean-Philippe Guertin, and Derek Nowrouzezahrai. Scalable appearance filtering for complex lighting effects. ACM Transactions on Graphics, 37(6): 277:1–277:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gavriil:2020:CDC

[GGP^{+}20] Konstantinos Gavriil, Ruslan Guseinov, Jesús Pérez, Davide Pellis, Paul Henderson, Florian Rist, Helmut Pottmann, and Bernd Bickel. Computational design of cold bent glass façades. ACM Transactions on Graphics, 39(6): 208:1–208:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417843.

Gotsman:2003:FSP

[GGS03] Craig Gotsman, Xianfeng Gu, and Alla Sheffer. Fundamentals of spherical parameterization for 3D meshes. ACM Transactions on Graphics, 22 (3):358–363, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gunther:2017:GOV

[GGT17] Tobias Günther, Markus Gross, and Holger Theisel. Generic objective vortices for flow visualization. ACM Transactions on Graphics, 36 (4):141:1–141:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ge:2018:ISR

[GGY18] Weifeng Ge, Bingchen Gong, and Yizhou Yu. Image superresolution via deterministicstochastic synthesis and local statistical rectification. ACM Transactions on Graphics, 37 (6):260:1–260:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Greiner:1998:ECA

[GH98] Günther Greiner and Kai Hormann. Efficient clipping of arbitrary polygons. ACM Transactions on Graphics, 17 (2):71–83, April 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/tog/1998-17-2/
                                        p71-greiner/.

Gissler:2020:ICS

[GHB^{+}20] Christoph Gissler, Andreas Henne, Stefan Band, Andreas Peer, and Matthias Teschner. An implicit compressible SPH solver for snow simulation. ACM Transactions on Graphics, 39(4): 36:1–36:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392431.

Gal:2021:SSB

[GHBCO21] Rinon Gal, Dana Cohen Hochberg, Amit Bermano, and Daniel Cohen-Or. SWA-GAN: a style-based waveletdriven generative model. ACM Transactions on Graphics, 40(4):134:1–134:11, August 2021. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459836.

Gaudet:1988:MEH

[GHCC88] Severin Gaudet, Richard Hobson, Pradeep Chilka, and Thomas Calvert. Multiprocessor experiments for high speed ray tracing. ACM Transactions on Graphics, 7 (3):151–179, July 1988. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
44480.html.

Guarnera:2017:WFM

[GHCG17] Giuseppe Claudio Guarnera, Peter Hall, Alain Chesnais, and Mashhuda Glencross. Woven fabric model creation from a single image. ACM Transactions on Graphics, 36(5):165:1–165:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldenthal:2007:ESI

[GHF^{+}07] Rony Goldenthal, David Harmon, Raanan Fattal, Michel Bercovier, and Eitan Grinspun. Efficient simulation of inextensible cloth. ACM Transactions on Graphics, 26 (3):49:1–49:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2018:MPM

[GHF^{+}18] Qi Guo, Xuchen Han, Chuyuan Fu, Theodore Gast, Rasmus Tamstorf, and Joseph Teran. A material point method for thin shells with frictional contact. ACM Transactions on Graphics, 37(4):147:1– 147:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gryaditskaya:2020:LFC

[GHL^{+}20] Yulia Gryaditskaya, Felix Hähnlein, Chenxi Liu, Alla Sheffer, and Adrien Bousseau. Lifting freehand concept sketches into 3D. ACM Transactions on Graphics, 39(6): 167:1–167:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417851.

Ghosh:2008:PMA

[GHP^{+}08] Abhijeet Ghosh, Tim Hawkins, Pieter Peers, Sune Frederiksen, and Paul Debevec. Practical modeling and acquisition of layered facial reflectance. ACM Transactions on Graphics, 27(5):139:1– 139:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guerrero:2022:MGM

[GHS^{+}22] Paul Guerrero, Miloš Hašan, Kalyan Sunkavalli, Radomír Měch, Tamy Boubekeur, and Niloy J. Mitra. MatFormer: a generative model for procedural materials. ACM Transactions on Graphics, 41(4):46:1–46:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530173.

Gruson:2018:GDV

[GHV^{+}18] Adrien Gruson, Binh-Son Hua, Nicolas Vibert, Derek Nowrouzezahrai, and Toshiya Hachisuka. Gradient-domain volumetric photon density estimation. ACM Transactions on Graphics, 37(4):82:1– 82:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2018:PFM

[GHZ18] Yu Guo, Milos Hasan, and Shuang Zhao. Position-free Monte Carlo simulation for arbitrary layered BSDFs. ACM Transactions on Graphics, 37 (6):279:1–279:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Geilinger:2020:AAD

[GHZ^{+}20] Moritz Geilinger, David Hahn, Jonas Zehnder, Moritz Bächer, Bernhard Thomaszewski, and Stelian Coros. ADD: analytically differentiable dynamics for multi-body systems with frictional contact. ACM Transactions on Graphics, 39 (6):190:1–190:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417766.

Goldfeather:2004:NCO

[GI04] Jack Goldfeather and Victoria Interrante. A novel cubic-order algorithm for approximating principal direction vectors. ACM Transactions on Graphics, 23(1): 45–63, January 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Georgiev:2018:ABF

[GIF^{+}18] Iliyan Georgiev, Thiago Ize, Mike Farnsworth, Ramón Montoya-Vozmediano, Alan King, Brecht Van Lommel, Angel Jimenez, Oscar Anson, Shinji Ogaki, Eric Johnston, Adrien Herubel, Declan Russell, Frédéric Servant, and Marcos Fajardo. Arnold: a brute-force production path tracer. ACM Transactions on Graphics, 37 (3):32:1–32:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3182160.

Gu:2022:NJS

[GIGM22] Jeongmin Gu, Jose A. Iglesias-Guitian, and Bochang Moon. Neural James–Stein combiner for unbiased and biased renderings. ACM Transactions on Graphics, 41(6): 262:1–262:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555496.

Gregson:2014:CSC

[GITH14] James Gregson, Ivo Ihrke, Nils Thuerey, and Wolfgang Heidrich. From capture to simulation: connecting forward and inverse problems in fluids. ACM Transactions on Graphics, 33(4): 139:1–139:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gingold:2009:SAM

[GIZ09] Yotam Gingold, Takeo Igarashi, and Denis Zorin. Structured annotations for 2D-to-3D modeling. ACM Transactions on Graphics, 28(5): 148:1–148:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goel:2022:UMW

[GJ22] Purvi Goel and Doug L. James. Unified many-worlds browsing of arbitrary physicsbased animations. ACM Transactions on Graphics, 41 (4):156:1–156:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530082.

Guo:2020:IPM

[GJB^{+}20] Jianwei Guo, Haiyong Jiang, Bedrich Benes, Oliver Deussen, Xiaopeng Zhang, Dani Lischinski, and Hui Huang. Inverse procedural modeling of branching structures by inferring L-systems. ACM Transactions on Graphics, 39 (5):155:1–155:13, September 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3394105.

Garg:2016:CDR

[GJG16] Akash Garg, Alec Jacobson, and Eitan Grinspun. Computational design of reconfigurables. ACM Transactions on Graphics, 35(4): 90:1–90:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2017:RHD

[GJTP17] Xifeng Gao, Wenzel Jakob, Marco Tarini, and Daniele Panozzo. Robust hexdominant mesh generation using field-guided polyhedral agglomeration. ACM Transactions on Graphics, 36(4): 114:1–114:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guerrero:2014:EPU

[GJWW14] Paul Guerrero, Stefan Jeschke, Michael Wimmer, and Peter Wonka. Edit propagation using geometric relationship functions. ACM Transactions on Graphics, 33(2): 15:1–15:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guerrero:2015:LSP

[GJWW15] Paul Guerrero, Stefan Jeschke, Michael Wimmer, and Peter Wonka. Learning shape placements by example. ACM Transactions on Graphics, 34(4):108:1–108:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2021:BMT

[GJZ21] Yu Guo, Adrian Jarabo, and Shuang Zhao. Beyond Mie theory: systematic computation of bulk scattering parameters based on microphysical wave optics. ACM Transactions on Graphics, 40(6): 285:1–285:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480543.

Georgiev:2012:LTS

[GKDS12] Iliyan Georgiev, Jaroslav Krivánek, Tomás Davidovic, and Philipp Slusallek. Light transport simulation with vertex connection and merging. ACM Transactions on Graphics, 31(6):192:1– 192:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Georgiev:2013:JIS

[GKH^{+}13] Iliyan Georgiev, Jaroslav Krivánek, Toshiya Hachisuka, Derek Nowrouzezahrai, and Wojciech Jarosz. Joint importance sampling of low-order volumetric scattering. ACM Transactions on Graphics, 32 (6):164:1–164:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gregson:2012:STA

[GKHH12] James Gregson, Michael Krimerman, Matthias B. Hullin, and Wolfgang Heidrich. Stochastic tomography and its applications in 3D imaging of mixing fluids. ACM Transactions on Graphics, 31(4):52:1–52:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Govindaraju:2005:ICD

[GKJ^{+}05] Naga K. Govindaraju, David Knott, Nitin Jain, Ilknur Kabul, Rasmus Tamstorf, Russell Gayle, Ming C. Lin, and Dinesh Manocha. Interactive collision detection between deformable models using chromatic decomposition. ACM Transactions on Graphics, 24(3):991–999, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garanzha:2021:FFM

[GKK^{+}21] Vladimir Garanzha, Igor Kaporin, Liudmila Kudryavtseva, François Protais, Nicolas Ray, and Dmitry Sokolov. Foldover-free maps in 50 lines of code. ACM Transactions on Graphics, 40(4): 102:1–102:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459847.

Garanzha:2024:QSO

[GKK^{+}24] Vladimir Garanzha, Igor Kaporin, Liudmila Kudryavtseva, Francois Protais, and Dmitry Sokolov. In the quest for scale-optimal mappings. ACM Transactions on Graphics, 43(1):8:1–8:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3627102.

Grinspun:2002:CSF

[GKS02] Eitan Grinspun, Petr Krysl, and Peter Schröder. CHARMS: a simple framework for adaptive simulation. ACM Transactions on Graphics, 21(3): 281–290, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Granados:2013:ANM

[GKTT13] Miguel Granados, Kwang In Kim, James Tompkin, and Christian Theobalt. Automatic noise modeling for ghost-free HDR reconstruction. ACM Transactions on Graphics, 32(6):201:1– 201:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Glassner:1990:TDV

[Gla90] Andrew S. Glassner. A twodimensional view controller. ACM Transactions on Graphics, 9(1):138–141, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77642.html.

Glassner:1995:E

[Gla95] Andrew S. Glassner. Editorial. ACM Transactions on Graphics, 14(4):309– 310, October 1995. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Glassner:1997:E

[Gla97] Andrew Glassner. Editorial.

ACM Transactions on Graphics, 16(1):1–2, January 1997. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gharbi:2019:SBM

[GLA^{+}19] Michaël Gharbi, Tzu-Mao Li, Miika Aittala, Jaakko Lehtinen, and Frédo Durand. Sample-based Monte Carlo denoising using a kernelsplatting network. ACM Transactions on Graphics, 38 (4):125:1–125:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Galvane:2018:DCD

[GLC^{+}18] Quentin Galvane, Christophe Lino, Marc Christie, Julien Fleureau, Fabien Servant, Frano̧is-Louis Tariolle, and Philippe Guillotel. Directing cinematographic drones.

ACM Transactions on Graphics, 37(3):34:1–34:??, August 2018. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3181975.

Gao:2023:SSB

[GLC^{+}23] Lin Gao, Feng-Lin Liu, Shu-Yu Chen, Kaiwen Jiang, Chun-Peng Li, Yu-Kun Lai, and Hongbo Fu. Sketch-FaceNeRF: Sketch-based facial generation and editing in neural radiance fields. ACM Transactions on Graphics, 42(4):159:1–159:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592100.

Gao:2019:DIR

[GLD^{+}19] Duan Gao, Xiao Li, Yue Dong, Pieter Peers, Kun Xu, and Xin Tong. Deep inverse rendering for high-resolution SVBRDF estimation from an arbitrary number of images. ACM Transactions on Graphics, 38(4):134:1–134:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ganacim:2014:MPV

[GLdFN14] Francisco Ganacim, Rodolfo S. Lima, Luiz Henrique de Figueiredo, and Diego Nehab. Massivelyparallel vector graphics. ACM Transactions on Graphics, 33 (6):229:1–229:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gkioulekas:2015:MSL

[GLDZ15] Ioannis Gkioulekas, Anat Levin, Frédo Durand, and Todd Zickler. Micron-scale light transport decomposition using interferometry. ACM Transactions on Graphics, 34 (4):37:1–37:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2023:MML

[GLH^{+}23] Jie Guo, Zeru Li, Xueyan He, Beibei Wang, Wenbin Li, Yanwen Guo, and Ling-Qi Yan. MetaLayer: a metalearned BSDF model for layered materials. ACM Transactions on Graphics, 42(6): 222:1–222:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618365.

Garcia:2011:CPH

[GLHL11] Ismael García, Sylvain Lefebvre, Samuel Hornus, and Anass Lasram. Coherent parallel hashing. ACM Transactions on Graphics, 30(6): 161:1–161:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goesele:2004:DAT

[GLL^{+}04] Michael Goesele, Hendrik P. A. Lensch, Jochen Lang, Christian Fuchs, and Hans-Peter Seidel. DISCO: acquisition of translucent objects. ACM Transactions on Graphics, 23(3):835–844, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2016:EFD

[GLL^{+}16] Lin Gao, Yu-Kun Lai, Dun Liang, Shu-Yu Chen, and Shihong Xia. Efficient and flexible deformation representation for data-driven surface modeling. ACM Transactions on Graphics, 35(5):158:1– 158:??, September 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Galerne:2012:GNE

[GLLD12] Bruno Galerne, Ares Lagae, Sylvain Lefebvre, and George Drettakis. Gabor noise by example. ACM Transactions on Graphics, 31(4): 73:1–73:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gurung:2011:LCC

[GLLR11] Topraj Gurung, Mark Luffel, Peter Lindstrom, and Jarek Rossignac. LR: compact connectivity representation for triangle meshes. ACM Transactions on Graphics, 30(4):

67:1–67:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2022:CCR

[GLP^{+}22] Haoxiang Guo, Shilin Liu, Hao Pan, Yang Liu, Xin Tong, and Baining Guo. ComplexGen: CAD reconstruction by B-rep chain complex generation. ACM Transactions on Graphics, 41 (4):129:1–129:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530078.

Guo:2021:HAT

[GLT^{+}21] Jie Guo, Shuichang Lai, Chengzhi Tao, Yuelong Cai, Lei Wang, Yanwen Guo, and Ling-Qi Yan. Highlightaware two-stream network for single-image SVBRDF acquisition. ACM Transactions on Graphics, 40(4): 123:1–123:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459854.

Guo:2023:UHR

[GLT^{+}23] Jie Guo, Shuichang Lai, Qinghao Tu, Chengzhi Tao, Changqing Zou, and Yanwen Guo. Ultra-high resolution SVBRDF recovery from a single image. ACM Transactions on Graphics, 42 (3):33:1–33:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3593798.

Guo:2022:ACA

[GLX^{+}22] Junfu Guo, Changhao Li, Xi Xia, Ruizhen Hu, and Ligang Liu. Asynchronous collaborative autoscanning with mode switching for multi-robot scene reconstruction. ACM Transactions on Graphics, 41(6):198:1– 198:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555483.

Govindaraju:2003:ISG

[GLY^{+}03] Naga K. Govindaraju, Brandon Lloyd, Sung-Eui Yoon, Avneesh Sud, and Dinesh Manocha. Interactive shadow generation in complex environments. ACM Transactions on Graphics, 22(3): 501–510, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2021:VAS

[GLZ^{+}21] Jie Guo, Mengtian Li, Zijing Zong, Yuntao Liu, Jingwu He, Yanwen Guo, and Ling-Qi Yan. Volumetric appearance stylization with stylizing kernel prediction network. ACM Transactions on Graphics, 40(4):162:1– 162:15, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459799.

Ghosh:1984:BTA

[GM84] P. K. Ghosh and S. P. Mudur.

The bush-trajectory approach to figure specification: Some algebraic solutions. ACM Transactions on Graphics, 3 (2):110–134, April 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gobbetti:2005:FVM

[GM05] Enrico Gobbetti and Fabio Marton. Far voxels: a multiresolution framework for interactive rendering of huge complex 3D models on commodity graphics platforms. ACM Transactions on Graphics, 24(3):878–885, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gamito:2009:AMP

[GM09] Manuel N. Gamito and Steve C. Maddock. Accurate multidimensional Poissondisk sampling. ACM Transactions on Graphics, 29(1):8:1– 8:19, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guseinov:2017:CSO

[GMB17] Ruslan Guseinov, Eder Miguel, and Bernd Bickel. Curve-Ups: shaping objects from flat plates with tension-actuated curvature. ACM Transactions on Graphics, 36(4): 64:1–64:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guillen:2020:GFP

[GMG^{+}20] Ibón Guillén, Julio Marco, Diego Gutierrez, Wenzel Jakob, and Adrian Jarabo. A general framework for pearlescent materials. ACM Transactions on Graphics, 39(6): 253:1–253:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417782.

Grochow:2004:SBI

[GMHP04] Keith Grochow, Steven L. Martin, Aaron Hertzmann, and Zoran Popović. Stylebased inverse kinematics. ACM Transactions on Graphics, 23(3):522–531, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Golovinskiy:2006:SMS

[GMP^{+}06] Aleksey Golovinskiy, Wojciech Matusik, Hanspeter Pfister, Szymon Rusinkiewicz, and Thomas Funkhouser. A statistical model for synthesis of detailed facial geometry. ACM Transactions on Graphics, 25(3):1025–1034, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gomes:2009:BBA

[GMP09] Abel J. P. Gomes, José F. M. Morgado, and Edgar S. Pereira. A BSP-based algorithm for dimensionally nonhomogeneous planar implicit curves with topological guarantees. ACM Transactions on Graphics, 28(2): 17:1–17:24, April 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Glauser:2016:RAT

[GMP^{+}16] Oliver Glauser, Wan-Chun Ma, Daniele Panozzo, Alec Jacobson, Otmar Hilliges, and Olga Sorkine-Hornung. Rig animation with a tangible and modular input device. ACM Transactions on Graphics, 35 (4):144:1–144:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guerrero:2016:RRA

[GMW16] Paul Guerrero, Niloy J. Mitra, and Peter Wonka. RAID: a relation-augmented image descriptor. ACM Transactions on Graphics, 35(4): 46:1–46:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garg:2006:PRR

[GN06] Kshitiz Garg and Shree K. Nayar. Photorealistic rendering of rain streaks. ACM Transactions on Graphics, 25 (3):996–1002, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gupta:2015:PIG

[GNHM15] Mohit Gupta, Shree K. Nayar, Matthias B. Hullin, and Jaime Martin. Phasor imaging: a generalization of correlationbased time-of-flight imaging. ACM Transactions on Graphics, 34(5):156:1–156:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Golas:2012:LSF

[GNS^{+}12] Abhinav Golas, Rahul Narain, Jason Sewall, Pavel Krajcevski, Pradeep Dubey, and Ming Lin. Large-scale fluid simulation using velocityvorticity domain decomposition. ACM Transactions on Graphics, 31(6):148:1– 148:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gastal:2011:DTE

[GO11] Eduardo S. L. Gastal and Manuel M. Oliveira. Domain transform for edge-aware image and video processing. ACM Transactions on Graphics, 30(4):69:1–69:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gastal:2012:AMR

[GO12] Eduardo S. L. Gastal and Manuel M. Oliveira. Adaptive manifolds for real-time highdimensional filtering. ACM Transactions on Graphics, 31(4):33:1–33:13, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gastal:2017:SRI

[GO17] Eduardo S. L. Gastal and Manuel M. Oliveira. Spectral remapping for image downscaling. ACM Transactions on Graphics, 36(4): 145:1–145:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldman:1984:MCC

[Gol84] Ronald N. Goldman. Markov chains and computer-aided geometric design: Part I — problems and constraints. ACM Transactions on Graphics, 3(3):204–222, July 1984. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldman:1985:MCC

[Gol85a] R. N. Goldman. Markov chains and computer aided geometric design II — examples and subdivision matrices. ACM Transactions on Graphics, 4(1):12–40, January 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
3974.html.

Goldman:1985:IEV

[Gol85b] Ronald N. Goldman. Illicit expressions in vector algebra. ACM Transactions on Graphics, 4(3):223–243, July 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldman:2002:AGF

[Gol02] Ron Goldman. On the algebraic and geometric foundations of computer graphics. ACM Transactions on Graphics, 21(1):52–86, January 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gonzalez-Ochoa:1998:CMO

[GOMP98] Carlos Gonzalez-Ochoa, Scott McCammon, and Jörg Peters. Computing moments of objects enclosed by piecewise polynomial surfaces. ACM Transactions on Graphics, 17 (3):143–157, July 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1998-17-3/
p143-gonzalez-ochoa/.

Goshtasby:2000:GPI

[Gos00] A. Ardeshir Goshtasby. Grouping and parameterizing irregularly spaced points for curve fitting. ACM Transactions on Graphics, 19(3):185–203, 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tog/2000-
19-3/p185-goshtasby/p185-
goshtasby.pdf;   http://
www.acm.org/pubs/citations/
journals/tog/2000-19-3/
p185-goshtasby/.

Gooch:2005:CSP

[GOTG05] Amy A. Gooch, Sven C.

Olsen, Jack Tumblin, and Bruce Gooch. Color2Gray: salience-preserving color removal. ACM Transactions on Graphics, 24(3):634–639, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garcia:2008:IIG

[GP08] Ismael García and Gustavo Patow. IGT: inverse geometric textures. ACM Transactions on Graphics, 27(5): 137:1–137:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gonzalez:2009:CMM

[GP09] Francisco González and Gustavo Patow. Continuity mapping for multi-chart textures. ACM Transactions on Graphics, 28(5):109:1– 109:8, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gissler:2019:ISP

[GPB^{+}19] Christoph Gissler, Andreas Peer, Stefan Band, Jan Bender, and Matthias Teschner. Interlinked SPH pressure solvers for strong fluid-rigid coupling. ACM Transactions on Graphics, 38(1):5:1– 5:??, February 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3284980.

Garcia:2013:CMM

[GPCP13] Francisco González García, Teresa Paradinas, Narcís Coll, and Gustavo Patow. *Cages: a multilevel, multi-cage-based system for mesh deformation. ACM Transactions on Graphics, 32(3):24:1–24:13, June 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Geilinger:2018:SOB

[GPD^{+}18] Moritz Geilinger, Roi Poranne, Ruta Desai, Bernhard Thomaszewski, and Stelian Coros. Skaterbots: optimization-based design and motion synthesis for robotic creatures with legs and wheels. ACM Transactions on Graphics, 37(4): 160:1–160:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2018:AFS

[GPH^{+}18] Ming Gao, Andre Pradhana, Xuchen Han, Qi Guo, Grant Kot, Eftychios Sifakis, and Chenfanfu Jiang. Animating fluid sediment mixture in particle-laden flows. ACM Transactions on Graphics, 37(4):149:1–149:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Glauser:2019:DCS

[GPHSH19] Oliver Glauser, Daniele Panozzo, Otmar Hilliges, and Olga Sorkine-Hornung. Deformation capture via soft and stretchable sensor arrays. ACM Transactions on Graphics, 38(2):16:1– 16:??, April 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3311972.

Gal:2022:SNC

[GPM^{+}22] Rinon Gal, Or Patashnik, Haggai Maron, Amit H. Bermano, Gal Chechik, and Daniel Cohen-Or. StyleGAN-NADA: CLIP-guided domain adaptation of image generators. ACM Transactions on Graphics, 41(4):141:1– 141:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530164.

Garcia-Puente:2011:TDB

[GPSZ11] Luis David García-Puente, Frank Sottile, and Chungang Zhu. Toric degenerations of Bézier patches. ACM Transactions on Graphics, 30(5): 110:1–110:10, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2017:RSS

[GPW^{+}17] Xifeng Gao, Daniele Panozzo, Wenping Wang, Zhigang Deng, and Guoning Chen. Robust structure simplification for hex re-meshing. ACM Transactions on Graphics, 36 (6):185:1–185:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gotardo:2018:PDF

[GRB^{+}18] Paulo Gotardo, Jérémy Riviere, Derek Bradley, Abhijeet Ghosh, and Thabo Beeler. Practical dynamic facial appearance modeling and acquisition. ACM Transactions on Graphics, 37(6):232:1– 232:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gu:2009:RIA

[GRBN09] Jinwei Gu, Ravi Ramamoorthi, Peter Belhumeur, and Shree Nayar. Removing image artifacts due to dirty camera lenses and thin occluders. ACM Transactions on Graphics, 28(5):144:1– 144:10, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Green:1986:STD

[Gre86]  Mark Green.  A sur-

vey of three dialogue models. ACM Transactions on Graphics, 5(3):244–275, July 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
24057.html.

Gooch:2004:HFI

[GRG04] Bruce Gooch, Erik Reinhard, and Amy Gooch. Human facial illustrations: Creation and psychophysical evaluation. ACM Transactions on Graphics, 23(1):27– 44, January 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guay:2015:STS

[GRGC15] Martin Guay, Rémi Ronfard, Michael Gleicher, and Marie-Paule Cani. Spacetime sketching of character animation. ACM Transactions on Graphics, 34(4): 118:1–118:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guan:2012:DDP

[GRH^{+}12] Peng Guan, Loretta Reiss, David A. Hirshberg, Alexander Weiss, and Michael J. Black. DRAPE: DRessing Any PErson. ACM Transactions on Graphics, 31(4): 35:1–35:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Geist:1993:MFD

[GRS93] Robert Geist, Robert Reynolds, and Darrell Suggs. A Markovian framework for digital halftoning. ACM Transactions on Graphics, 12(2): 136–159, April 1993. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        151281.html.

Gruson:2017:STFa

[GRS^{+}17a] Adrien Gruson, Mickaël Ribardière, Martin Sik, Jirí Vorba, Rémi Cozot, Kadi Bouatouch, and Jaroslav Krivánek. A spatial target function for Metropolis photon tracing. ACM Transactions on Graphics, 36(1):4:1– 4:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gruson:2017:STFb

[GRS^{+}17b] Adrien Gruson, Mickaël Ribardière, Martin Sik, Jirí Vorba, Rémi Cozot, Kadi Bouatouch, and Jaroslav Krivánek. A spatial target function for Metropolis photon tracing. ACM Transactions on Graphics, 36(4): 75:1–75:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gunther:2013:OOL

[GRT13] Tobias Günther, Christian Rössl, and Holger Theisel. Opacity optimization for 3D line fields. ACM Transactions on Graphics, 32(4): 120:1–120:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guibas:1982:LBM

[GS82] L. J. Guibas and J. Stolfi. A language for bitmap manipulation. ACM Transactions on Graphics, 1(3):191–214, July 1982. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guibas:1985:PMG

[GS85] Leonidas Guibas and Jorge Stolfi. Primitives for the manipulation of general subdivisions and computation of Voronoi diagrams. ACM Transactions on Graphics, 4 (2):74–123, April 1985. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guy:2004:GGR

[GS04] Stephane Guy and Cyril Soler. Graphics gems revisited: fast and physically-based rendering of gemstones. ACM Transactions on Graphics, 23(3): 231–238, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gharbi:2015:TRE

[GSC^{+}15] Michaël Gharbi, YiChang Shih, Gaurav Chaurasia, Jonathan Ragan-Kelley, Sylvain Paris, and Frédo Durand. Transform recipes for efficient cloud photo enhancement. ACM Transactions on Graphics, 34(6): 228:1–228:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gillespie:2021:ICI

[GSC21a] Mark Gillespie, Nicholas Sharp, and Keenan Crane. Integer coordinates for intrinsic geometry processing. ACM Transactions on Graphics, 40 (6):252:1–252:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480522.

Gillespie:2021:DCE

[GSC21b] Mark Gillespie, Boris Springborn, and Keenan Crane. Discrete conformal equivalence of polyhedral surfaces. ACM Transactions on Graphics, 40(4):103:1–103:20, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459763.

Gingold:2012:MPH

[GSCO12] Yotam Gingold, Ariel Shamir, and Daniel Cohen-Or. Micro perceptual human computation for visual tasks. ACM Transactions on Graphics, 31(5):119:1–119:12, August 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garg:2014:WMD

[GSFD^{+}14] Akash Garg, Andrew O.

Sageman-Furnas, Bailin Deng, Yonghao Yue, Eitan Grinspun, Mark Pauly, and Max Wardetzky. Wire mesh design. ACM Transactions on Graphics, 33(4):66:1–66:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gebhardt:2018:OAP

[GSH18] Christoph Gebhardt, Stefan Stevsić, and Otmar Hilliges. Optimizing for aesthetically pleasing quadrotor camera motion. ACM Transactions on Graphics, 37(4):90:1– 90:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2020:MRC

[GSH^{+}20] Yu Guo, Cameron Smith, Milos Hasan, Kalyan Sunkavalli, and Shuang Zhao. MaterialGAN: reflectance capture using a generative SVBRDF model. ACM Transactions on Graphics, 39(6): 254:1–254:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417779.

Gleicher:2003:STM

[GSKJ03] Michael Gleicher, Hyun Joon Shin, Lucas Kovar, and Andrew Jepsen. Snap-together motion: assembling run-time animations. ACM Transactions on Graphics, 22(3): 702, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guendelman:2005:CWS

[GSLF05] Eran Guendelman, Andrew Selle, Frank Losasso, and Ronald Fedkiw. Coupling water and smoke to thin deformable and rigid shells. ACM Transactions on Graphics, 24(3):973–981, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gutierrez:2008:DPC

[GSLM^{+}08] Diego Gutierrez, Francisco J.

Seron, Jorge Lopez-Moreno, Maria P. Sanchez, Jorge Fandos, and Erik Reinhard. Depicting procedural caustics in single images. ACM Transactions on Graphics, 27(5): 120:1–120:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gal:2009:IAE

[GSMCO09] Ran Gal, Olga Sorkine, Niloy J. Mitra, and Daniel Cohen-Or. iWIRES: an analyze-and-edit approach to shape manipulation. ACM Transactions on Graphics, 28 (3):33:1–33:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Green:2007:MAP

[GSMD07] Paul Green, Wenyang Sun, Wojciech Matusik, and Frédo Durand. Multi-aperture photography. ACM Transactions on Graphics, 26(3): 68:1–68:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gross:2023:MSC

[GSP^{+}23] Oliver Gross, Yousuf Soliman, Marcel Padilla, Felix Knöppel, Ulrich Pinkall, and Peter Schröder. Motion from shape change. ACM Transactions on Graphics, 42(4): 107:1–107:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592417.

Granskog:2021:NSG

[GSRN21] Jonathan Granskog, Till N. Schnabel, Fabrice Rousselle, and Jan Novák. Neural scene graph rendering. ACM Transactions on Graphics, 40(4): 164:1–164:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459848.

Gilet:2014:LRP

[GSV^{+}14] Guillaume Gilet, Basile Sauvage, Kenneth Vanhoey, Jean-Michel Dischler, and Djamchid Ghazanfarpour. Local random-phase noise for procedural texturing. ACM Transactions on Graphics, 33(6): 195:1–195:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gori:2017:FDC

[GSV^{+}17] Giorgio Gori, Alla Sheffer, Nicholas Vining, Enrique Rosales, Nathan Carr, and Tao Ju. FlowRep: descriptive curve networks for free-form design shapes. ACM Transactions on Graphics, 36(4): 59:1–59:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gardner:2017:LPI

[GSY^{+}17] Marc-André Gardner, Kalyan Sunkavalli, Ersin Yumer, Xiaohui Shen, Emiliano Gambaretto, Christian Gagné, and Jean-François Lalonde. Learning to predict indoor illumination from a single image. ACM Transactions on Graphics, 36 (6):176:1–176:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Geng:2018:WGG

[GSZ^{+}18] Jiahao Geng, Tianjia Shao, Youyi Zheng, Yanlin Weng, and Kun Zhou. Warp-guided GANs for single-photo facial animation. ACM Transactions on Graphics, 37(6): 231:1–231:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guenter:1996:QPH

[GT96] Brian Guenter and Jack Tumblin. Quadrature prefiltering for high quality antialiasing. ACM Transactions on Graphics, 15(4):332–353, October 1996. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        234540.html.

Gryka:2015:LRS

[GTB15] Maciej Gryka, Michael Terry, and Gabriel J. Brostow. Learning to remove soft shadows. ACM Transactions on Graphics, 34(5):153:1– 153:??, October 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Grabli:2010:PRL

[GTDS10] Stéphane Grabli, Emmanuel Turquin, Frédo Durand, and François X. Sillion. Programmable rendering of line drawing from 3D scenes. ACM Transactions on Graphics, 29 (2):18:1–18:20, March 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gardner:2003:LLS

[GTHD03] Andrew Gardner, Chris Tchou, Tim Hawkins, and Paul Debevec. Linear light source reflectometry. ACM Transactions on Graphics, 22(3): 749–758, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2017:AGI

[GTJS17] Ming Gao, Andre Pradhana Tampubolon, Chenfanfu Jiang, and Eftychios Sifakis. An adaptive generalized interpolation material point method for simulating elastoplastic materials. ACM Transactions on Graphics, 36(6): 223:1–223:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gu:2006:TVS

[GTR^{+}06] Jinwei Gu, Chien-I Tu, Ravi Ramamoorthi, Peter Belhumeur, Wojciech Matusik, and Shree Nayar. Timevarying surface appearance: acquisition, modeling and rendering. ACM Transactions on Graphics, 25(3):762–771, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guenter:2007:ESD

[Gue07] Brian Guenter. Efficient symbolic differentiation for graphics applications. ACM Transactions on Graphics, 26(3): 108:1–108:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gupta:2018:SDR

[Gup18] Mohit Gupta. Session details: Rendering & reflectance. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gil-Ureta:2020:RGS

[GUPZ20] Francisca Gil-Ureta, Nico Pietroni, and Denis Zorin. Reinforcement of general shell structures. ACM Transactions on Graphics, 39 (5):153:1–153:19, September 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3375677.

Guy:2012:SSM

[GvdBL^{+}12] Stephen J. Guy, Jur van den Berg, Wenxi Liu, Rynson Lau, Ming C. Lin, and Dinesh Manocha. A statistical similarity measure for aggregate crowd dynamics. ACM Transactions on Graphics, 31(6): 190:1–190:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Geijtenbeek:2013:FMB

[GvdPvdS13] Thomas Geijtenbeek, Michiel van de Panne, and A. Frank van der Stappen. Flexible muscle-based locomotion for bipedal creatures. ACM Transactions on Graphics, 32 (6):206:1–206:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gupta:2018:WOC

[GVNB18] Mohit Gupta, Andreas Velten, Shree K. Nayar, and Eric Breitbach. What are optimal coding functions for time-of-flight imaging? ACM Transactions on Graphics, 37(2): 13:1–13:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garrido:2013:RDD

[GVWT13] Pablo Garrido, Levi Valgaert, Chenglei Wu, and Christian Theobalt. Reconstructing detailed dynamic face geometry from monocular video. ACM Transactions on Graphics, 32 (6):158:1–158:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guitard:1990:CSE

[GW90] Richard Guitard and Colin Ware. A color sequence editor. ACM Transactions on Graphics, 9(3):338–341, July 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldade:2019:AVF

[GWAB19] Ryan Goldade, Yipeng Wang, Mridul Aanjaneya, and Christopher Batty. An adaptive variational finite difference framework for efficient symmetric octree viscosity. ACM Transactions on Graphics, 38(4): 94:1–94:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Grossman:2005:MFG

[GWB05] Tovi Grossman, Daniel Wigdor, and Ravin Balakrishnan. Multi-finger gestural interaction with 3D volumetric displays. ACM Transactions on Graphics, 24(3):931, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gupta:2024:LCF

[GWBN24] Mohit Gupta, Jian Wang, Karl Bayer, and Shree K. Nayar. Light codes for fast two-way human-centric visual communication. ACM Transactions on Graphics, 43 (1):1:1–1:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3617682.

Grosse:2010:CAP

[GWGB10] Max Grosse, Gordon Wetzstein, Anselm Grundhöfer, and Oliver Bimber. Coded aperture projection. ACM Transactions on Graphics, 29 (3):22:1–22:12, June 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gan:2023:HHI

[GWLG23] Ji Gan, Weiqiang Wang, Jiaxu Leng, and Xinbo Gao. HiGAN+: Handwriting imitation GAN with disentangled representations. ACM Transactions on Graphics, 42 (1):11:1–11:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550070.

Glencross:2008:PVM

[GWM^{+}08] Mashhuda Glencross, Gregory J. Ward, Francho Melendez, Caroline Jay, Jun Liu, and Roger Hubbold. A perceptually validated model for surface depth hallucination. ACM Transactions on Graphics, 27(3):59:1–59:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gross:2003:BCS

[GWN^{+}03] Markus Gross, Stephan Würmlin, Martin Naef, Edouard Lamboray, Christian Spagno, Andreas Kunz, Esther Koller-Meier, Tomas Svoboda, Luc Van Gool, Silke Lang, Kai Strehlke, Andrew Vande Moere, and Oliver Staadt. bluec: a spatially immersive display and 3D video portal for telepresence. ACM Transactions on Graphics, 22(3): 819–827, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Glauser:2019:IHP

[GWP^{+}19] Oliver Glauser, Shihao Wu, Daniele Panozzo, Otmar Hilliges, and Olga Sorkine-Hornung. Interactive hand pose estimation using a stretch-sensing soft glove. ACM Transactions on Graphics, 38(4):41:1–41:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2018:GOM

[GWW^{+}18] Ming Gao, Xinlei Wang, Kui Wu, Andre Pradhana, Eftychios Sifakis, Cem Yuksel, and Chenfanfu Jiang. GPU optimization of material point methods. ACM Transactions on Graphics, 37(6):254:1– 254:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2021:TND

[GWY^{+}21] Lin Gao, Tong Wu, Yu-Jie Yuan, Ming-Xian Lin, Yu-Kun Lai, and Hao Zhang. TM-NET: deep generative networks for textured meshes. ACM Transactions on Graphics, 40(6):263:1–263:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480503.

Gu:2023:FRV

[GXSD23] Zeqi Gu, Wenqi Xian, Noah Snavely, and Abe Davis. FactorMatte: Redefining video matting for re-composition tasks. ACM Transactions on Graphics, 42(4):47:1– 47:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592423.

Guo:2017:RTGa

[GXY^{+}17a] Kaiwen Guo, Feng Xu, Tao Yu, Xiaoyang Liu, Qionghai Dai, and Yebin Liu. Real-time geometry, albedo, and motion reconstruction using a single RGB-D camera. ACM Transactions on Graphics, 36(3): 32:1–32:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2017:RTGb

[GXY^{+}17b] Kaiwen Guo, Feng Xu, Tao Yu, Xiaoyang Liu, Qionghai Dai, and Yebin Liu. Real-time geometry, albedo and motion reconstruction using a single RGBD camera. ACM Transactions on Graphics, 36(4): 44:1–44:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gkioulekas:2013:URP

[GXZ^{+}13] Ioannis Gkioulekas, Bei Xiao, Shuang Zhao, Edward H. Adelson, Todd Zickler, and Kavita Bala. Understanding the role of phase function in translucent appearance. ACM Transactions on Graphics, 32 (5):147:1–147:19, September 2013. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Grittmann:2022:EAM

[GYGS22] Pascal Grittmann, Ömercan Yazici, Iliyan Georgiev, and Philipp Slusallek. Efficiencyaware multiple importance sampling for bidirectional rendering algorithms. ACM Transactions on Graphics, 41(4):80:1–80:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530126.

Gao:2018:AUS

[GYQ^{+}18] Lin Gao, Jie Yang, Yi-Ling Qiao, Yu-Kun Lai, Paul L. Rosin, Weiwei Xu, and Shihong Xia. Automatic unpaired shape deformation transfer. ACM Transactions on Graphics, 37(6):237:1– 237:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garland:2005:QBS

[GZ05] Michael Garland and Yuan Zhou. Quadric-based simplification in any dimension. ACM Transactions on Graphics, 24 (2):209–239, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gingold:2008:SBS

[GZ08] Yotam Gingold and Denis Zorin. Shading-based surface editing. ACM Transactions on Graphics, 27(3): 95:1–95:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gkioulekas:2013:IVR

[GZB^{+}13] Ioannis Gkioulekas, Shuang Zhao, Kavita Bala, Todd Zickler, and Anat Levin. Inverse volume rendering with material dictionaries. ACM Transactions on Graphics, 32(6): 162:1–162:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2015:MLD

[GZC15] Kan Guo, Dongqing Zou, and Xiaowu Chen. 3D mesh labeling via deep convolutional neural networks. ACM Transactions on Graphics, 35(1): 3:1–3:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Garrido:2016:RPF

[GZC^{+}16] Pablo Garrido, Michael Zollhöfer, Dan Casas, Levi Valgaerts, Kiran Varanasi, Patrick Pérez, and Christian Theobalt. Reconstruction of personalized 3D face rigs from monocular video. ACM Transactions on Graphics, 35(3):

28:1–28:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Goldberg:2008:AN

[GZD08] Alexander Goldberg, Matthias Zwicker, and Frédo Durand. Anisotropic noise. ACM Transactions on Graphics, 27 (3):54:1–54:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Glasner:2014:RD

[GZL14] Daniel Glasner, Todd Zickler, and Anat Levin. A reflectance display. ACM Transactions on Graphics, 33(4): 61:1–61:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Guo:2022:ELP

[GZS^{+}22] Jie Guo, Zijing Zong, Yadong Song, Xihao Fu, Chengzhi Tao, Yanwen Guo, and Ling-Qi Yan. Efficient light probes for real-time global illumination. ACM Transactions on Graphics, 41(6): 202:1–202:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555452.

Garrido:2016:CRL

[GZW^{+}16] Pablo Garrido, Michael Zollhöfer, Chenglei Wu, Derek Bradley, Patrick Pérez, Thabo Beeler, and Christian Theobalt. Corrective 3D reconstruction of lips from monocular video. ACM Transactions on Graphics, 35(6):219:1– 219:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Gao:2022:RPS

[GZX^{+}22] Xuan Gao, Chenglai Zhong, Jun Xiang, Yang Hong, Yudong Guo, and Juyong Zhang. Reconstructing personalized semantic facial NeRF models from monocular video. ACM Transactions on Graphics, 41(6): 200:1–200:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555501.

Hansen:1992:AGN

[HA92] Allan Hansen and Farhad Arbab. An algorithm for generating NC tools paths for arbitrarily shaped pockets with islands. ACM Transactions on Graphics, 11(2):152–182, April 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
130832.html.

Herholz:2018:FOR

[HA18] Philipp Herholz and Marc Alexa. Factor once: reusing Cholesky factorizations on sub-meshes. ACM Transactions on Graphics, 37(6): 230:1–230:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Haines:2016:MTY

[HAB16] Tom S. F. Haines, Oisin Mac Aodha, and Gabriel J. Brostow. My text in your handwriting. ACM Transactions on Graphics, 35(3): 26:1–26:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Henter:2020:MPC

[HAB20] Gustav Eje Henter, Simon Alexanderson, and Jonas Beskow. MoGlow: probabilistic and controllable motion synthesis using normalising flows. ACM Transactions on Graphics, 39(6): 236:1–236:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417836.

Hachisuka:2018:SDB

[Hac18] Toshiya Hachisuka. Session details: Beyond light transport. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harkonen:2022:DRC

[HAK^{+}22] Erik Härkönen, Miika Aittala, Tuomas Kynkäänniemi, Samuli Laine, Timo Aila, and Jaakko Lehtinen. Disentangling random and cyclic effects in time-lapse sequences. ACM Transactions on Graphics, 41(4):72:1–72:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530170.

Hasselgren:2007:PPC

[HAM07] Jon Hasselgren and Thomas Akenine-Möller. PCU: the programmable culling unit. ACM Transactions on Graphics, 26(3):92:1–92:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hart:2003:Ea

[Har03a] John C. Hart. Editorial. ACM Transactions on Graphics, 22 (1):2, January 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hart:2003:Eb

[Har03b] John C. Hart. Editorial. ACM Transactions on Graphics, 22 (4):981, October 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hart:2004:E

[Har04] John C. Hart. Editorial. ACM Transactions on Graphics, 23 (4):929, October 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hart:2005:E

[Har05] John C. Hart. Editorial. ACM Transactions on Graphics, 24 (2):181, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hedman:2017:CP

[HASK17] Peter Hedman, Suhib Alsisan, Richard Szeliski, and Johannes Kopf. Casual 3D photography. ACM Transactions on Graphics, 36(6): 234:1–234:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hohmeyer:1989:RCP

[HB89] M. E. Hohmeyer and B. A. Barsky. Rational continuity: parametric, geometric, and Frenet frame continuity of rational curves. ACM Transactions on Graphics, 8(4):335–359, October 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77274.html.

Hafner:2021:DSP

[HB21] Christian Hafner and Bernd Bickel. The design space of plane elastic curves. ACM Transactions on Graphics, 40(4):126:1–126:20, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459800.

Hafner:2023:DSK

[HB23] Christian Hafner and Bernd Bickel. The design space of Kirchhoff rods. ACM Transactions on Graphics, 42(5): 171:1–171:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3606033.

Hegarty:2014:DCH

[HBD^{+}14] James Hegarty, John Brunhaver, Zachary DeVito, Jonathan Ragan-Kelley, Noy Cohen, Steven Bell, Artem Vasilyev, Mark Horowitz, and Pat Hanrahan. Darkroom: compiling high-level image processing code into hardware pipelines. ACM Transactions on Graphics, 33(4):144:1–144:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holroyd:2011:CFM

[HBLM11] Michael Holroyd, Ilya Baran, Jason Lawrence, and Wojciech Matusik. Computing and fabricating multilayer models. ACM Transactions on Graphics, 30(6): 187:1–187:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hadrich:2021:FPM

[HBP^{+}21] Torsten Hädrich, Daniel T. Banuti, Wojtek Pałubicki, Sören Pirk, and Dominik L. Michels. Fire in paradise: mesoscale simulation of wildfires. ACM Transactions on Graphics, 40(4): 163:1–163:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459954.

Henderson:1986:RUM

[HC86] D. Austin Henderson, Jr. and Stuart K. Card. Rooms: the use of multiple virtual workspaces to reduce space contention in a windowbased graphical user interface. ACM Transactions on Graphics, 5(3):211–243, July 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
24056.html.

Hersch:2004:BMI

[HC04] Roger David Hersch and Sylvain Chosson. Band moiré images. ACM Transactions on Graphics, 23(3):239– 247, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hinderink:2023:GML

[HC23] Steffen Hinderink and Marcel Campen. Galaxy maps: Localized foliations for bijective volumetric mapping. ACM Transactions on Graphics, 42(4):129:1–129:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592410.

Hersch:2003:RCI

[HCE03] Roger D. Hersch, Fabien Collaud, and Patrick Emmel. Reproducing color images with embedded metallic patterns. ACM Transactions on Graphics, 22(3):427–434, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2022:NGS

[HCH22] Jiahui Huang, Hao-Xiang Chen, and Shi-Min Hu. A neural Galerkin solver for accurate surface reconstruction. ACM Transactions on Graphics, 41(6):229:1– 229:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555457.

Huang:2019:VIP

[HCJ19] Zhiyang Huang, Nathan Carr, and Tao Ju. Variational implicit point set surfaces. ACM Transactions on Graphics, 38 (4):124:1–124:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2018:DEB

[HCL^{+}18] Mingming He, Dongdong Chen, Jing Liao, Pedro V. Sander, and Lu Yuan. Deep exemplar-based colorization. ACM Transactions on Graphics, 37(4):47:1–47:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2024:GFS

[HCLK24] Kemeng Huang, Floyd M. Chitalu, Huancheng Lin, and Taku Komura. GIPC: Fast and stable Gauss–Newton optimization of IPC barrier energy. ACM Transactions on Graphics, 43(2): 23:1–23:??, April 2024. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3643028.

Held:2010:UBA

[HCOB10] Robert T. Held, Emily A. Cooper, James F. O’Brien, and Martin S. Banks. Using blur to affect perceived distance and size. ACM Transactions on Graphics, 29(2):

19:1–19:16, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2013:RPI

[HCS13] Kaiming He, Huiwen Chang, and Jian Sun. Rectangling panoramic images via warping. ACM Transactions on Graphics, 32(4): 79:1–79:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2011:LMC

[HCTW11] Haoda Huang, Jinxiang Chai, Xin Tong, and Hsiang-Tao Wu. Leveraging motion capture and 3D scanning for highfidelity facial performance acquisition. ACM Transactions on Graphics, 30(4): 74:1–74:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2015:LFS

[HCW15] Fu-Chung Huang, Kevin Chen, and Gordon Wetzstein. The light field stereoscope: immersive computer graphics via factored near-eye light field displays with focus cues. ACM Transactions on Graphics, 34(4):60:1–60:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hou:2023:RZL

[HCW^{+}23] Fei Hou, Xuhui Chen, Wencheng Wang, Hong Qin, and Ying He. Robust zero level-set extraction from unsigned distance fields based on double covering. ACM Transactions on Graphics, 42(6): 245:1–245:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618314.

Hadadan:2021:NR

[HCZ21] Saeed Hadadan, Shuhong Chen, and Matthias Zwicker. Neural radiosity. ACM Transactions on Graphics, 40(6): 236:1–236:11, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480569.

Herholz:2017:LSS

[HDA17] Philipp Herholz, Timothy A. Davis, and Marc Alexa. Localized solutions of sparse linear systems for geometry processing. ACM Transactions on Graphics, 36(6): 183:1–183:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hersch:2007:CIV

[HDC07] Roger D. Hersch, Philipp Donzé, and Sylvain Chosson. Color images visible under UV light. ACM Transactions on Graphics, 26(3): 75:1–75:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heitz:2015:SMD

[HDCD15] Eric Heitz, Jonathan Dupuy, Cyril Crassin, and Carsten Dachsbacher. The SGGX microflake distribution. ACM Transactions on Graphics, 34 (4):48:1–48:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hegarty:2016:RFM

[HDD^{+}16] James Hegarty, Ross Daly, Zachary DeVito, Jonathan Ragan-Kelley, Mark Horowitz, and Pat Hanrahan. Rigel: flexible multi-rate image processing hardware. ACM Transactions on Graphics, 35 (4):85:1–85:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2017:LTC

[HDGN17] Jingwei Huang, Angela Dai, Leonidas Guibas, and Matthias Niessner. 3Dlite: towards commodity 3D scanning for content creation. ACM Transactions on Graphics, 36(6): 203:1–203:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heitz:2016:RTP

[HDHN16] Eric Heitz, Jonathan Dupuy, Stephen Hill, and David Neubelt. Real-time polygonallight shading with linearly transformed cosines. ACM Transactions on Graphics, 35 (4):41:1–41:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hornung:2007:CAP

[HDK07] Alexander Hornung, Ellen Dekkers, and Leif Kobbelt. Character animation from 2D pictures and 3D motion data. ACM Transactions on Graphics, 26(1):1:1–1:9, January 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Henzler:2021:GMB

[HDMR21] Philipp Henzler, Valentin Deschaintre, Niloy J. Mitra, and Tobias Ritschel. Generative modelling of BRDF textures from flash images. ACM Transactions on Graphics, 40 (6):284:1–284:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478513.3480507.

Heide:2016:PEI

[HDN^{+}16] Felix Heide, Steven Diamond, Matthias Nießner, Jonathan Ragan-Kelley, Wolfgang Heidrich, and Gordon Wetzstein.

ProxImaL: efficient image optimization using proximal algorithms. ACM Transactions on Graphics, 35(4): 84:1–84:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hoshyari:2018:PDS

[HDS^{+}18] Shayan Hoshyari, Edoardo Alberto Dominici, Alla Sheffer, Nathan Carr, Zhaowen Wang, Duygu Ceylan, and I-Chao Shen. Perception-driven semistructured boundary vectorization. ACM Transactions on Graphics, 37(4):118:1– 118:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hays:2007:SCU

[HE07] James Hays and Alexei A.

Efros. Scene completion using millions of photographs. ACM Transactions on Graphics, 26 (3):4:1–4:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hawkins:2005:ATV

[HED05] Tim Hawkins, Per Einarsson, and Paul Debevec. Acquisition of time-varying participating media. ACM Transactions on Graphics, 24(3): 812–815, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hoiem:2005:APP

[HEH05] Derek Hoiem, Alexei A. Efros, and Martial Hebert. Automatic photo pop-up. ACM Transactions on Graphics, 24 (3):577–584, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hullin:2011:PBR

[HESL11] Matthias Hullin, Elmar Eisemann, Hans-Peter Seidel, and Sungkil Lee. Physically-based real-time lens flare rendering. ACM Transactions on Graphics, 30(4):108:1–108:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hamalainen:2014:OMS

[HET^{+}14] Perttu Hämäläinen, Sebastian Eriksson, Esa Tanskanen, Ville Kyrki, and Jaakko Lehtinen. Online motion synthesis using sequential Monte Carlo. ACM Transactions on Graphics, 33(4):51:1–51:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hormann:2006:MVC

[HF06] Kai Hormann and Michael S. Floater. Mean value coordinates for arbitrary planar polygons. ACM Transactions on Graphics, 25(4): 1424–1441, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2016:SRE

[HFF16] Yong He, Tim Foley, and Kayvon Fatahalian. A system for rapid exploration of shader optimization choices. ACM Transactions on Graphics, 35(4):112:1–112:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2018:SLM

[HFF18] Yong He, Kayvon Fatahalian, and Tim Foley. Slang: language mechanisms for extensible real-time shading systems. ACM Transactions on Graphics, 37(4):141:1– 141:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2006:RFO

[HFG^{+}06] Qi-Xing Huang, Simon Flöry, Natasha Gelfand, Michael Hofer, and Helmut Pottmann. Reassembling fractured objects by geometric matching. ACM Transactions on Graphics, 25(3):569–578, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2018:MLS

[HFG^{+}18] Yuanming Hu, Yu Fang, Ziheng Ge, Ziyin Qu, Yixin Zhu, Andre Pradhana, and Chenfanfu Jiang. A moving least squares material point method with displacement discontinuity and two-way rigid body coupling. ACM Transactions on Graphics, 37(4):150:1–150:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2017:SCM

[HFH^{+}17] Yong He, Tim Foley, Teguh Hofstee, Haomin Long, and Kayvon Fatahalian. Shader components: modular and high performance shader development. ACM Transactions on Graphics, 36(4): 100:1–100:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hullin:2008:FIR

[HFI^{+}08] Matthias B. Hullin, Martin Fuchs, Ivo Ihrke, Hans-Peter Seidel, and Hendrik P. A. Lensch. Fluorescent immersion range scanning. ACM Transactions on Graphics, 27 (3):87:1–87:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hart:1994:VQR

[HFK94] John C. Hart, George K. Francis, and Louis H. Kauffman. Visualizing quaternion rotation. ACM Transactions on Graphics, 13(3):256–276, July 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/

toc/Abstracts/0730-0301/
197480.html.

Huang:2014:DDS

[HFL14] Zhe Huang, Hongbo Fu, and Rynson W. H. Lau. Data-driven segmentation and labeling of freehand sketches. ACM Transactions on Graphics, 33(6):175:1– 175:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hasan:2010:PRM

[HFM^{+}10] Miloš Hašan, Martin Fuchs, Wojciech Matusik, Hanspeter Pfister, and Szymon Rusinkiewicz. Physical reproduction of materials with specified subsurface scattering. ACM Transactions on Graphics, 29(4): 61:1–61:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2015:SRA

[HFTF15] Yong He, Tim Foley, Natalya Tatarchuk, and Kayvon Fatahalian. A system for rapid, automatic shader levelof-detail. ACM Transactions on Graphics, 34(6):187:1– 187:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hanocka:2019:APS

[HFW^{+}19] Rana Hanocka, Noa Fish, Zhenhua Wang, Raja Giryes, Shachar Fleishman, and Daniel Cohen-Or. ALIGNet: Partial-shape agnostic alignment via unsupervised learning. ACM Transactions on Graphics, 38(1):1:1–1:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3267347.

Horvath:2009:DHR

[HG09] Christopher Horvath and Willi Geiger. Directable, highresolution simulation of fire on the GPU. ACM Transactions on Graphics, 28(3): 41:1–41:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2012:FGR

[HGCO^{+}12] Hui Huang, Minglun Gong, Daniel Cohen-Or, Yaobin Ouyang, Fuwen Tan, and Hao Zhang. Field-guided registration for feature-conforming shape composition. ACM Transactions on Graphics, 31 (6):179:1–179:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2014:EGP

[HGF14] Yong He, Yan Gu, and Kayvon Fatahalian. Extending the graphics pipeline with adaptive, multi-rate shading. ACM Transactions on Graphics, 33(4):142:1–142:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heinzle:2011:CSC

[HGG^{+}11] Simon Heinzle, Pierre Greisen, David Gallup, Christine Chen, Daniel Saner, Aljoscha Smolic, Andreas Burg, Wojciech Matusik, and Markus Gross. Computational stereo camera system with programmable control loop. ACM Transactions on Graphics, 30 (4):94:1–94:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2014:NRS

[HGM14] Qixing Huang, Leonidas J.

Guibas, and Niloy J. Mitra. Near-regular structure discovery using linear programming. ACM Transactions on Graphics, 33(3): 23:1–23:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hyde:2020:IUL

[HGMRT20] David A. B. Hyde, Steven W.

Gagniere, Alan Marquez-Razon, and Joseph Teran. An implicit updated Lagrangian formulation for liquids with large surface energy. ACM Transactions on Graphics, 39 (6):183:1–183:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417845.

Hachet:2004:CEI

[HGRT04] Martin Hachet, Pascal Guitton, Patrick Reuter, and Florence Tyndiuk. The CAT for efficient 2D and 3D interaction as an alternative to mouse adaptations. ACM Transactions on Graphics, 23 (3):731, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hua:2023:RCM

[HGS23] Qingqin Hua, Pascal Grittmann, and Philipp Slusallek. Revisiting controlled mixture sampling for rendering applications. ACM Transactions on Graphics, 42(4): 64:1–64:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592435.

Han:2017:DDL

[HGY17] Xiaoguang Han, Chang Gao, and Yizhou Yu. DeepSketch2Face: a deep learning based sketching system for 3D face and caricature modeling. ACM Transactions on Graphics, 36(4):126:1–126:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Henry:1990:MI

[HH90] Tyson R. Henry and Scott E.

Hudson. Multidimensional icons. ACM Transactions on Graphics, 9(1):133– 137, January 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77641.html.

Han:2010:OCM

[HH10] Charles Han and Hugues Hoppe. Optimizing continuity in multiscale imagery. ACM Transactions on Graphics, 29 (6):171:1–171:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hill:2016:EFS

[HH16] David J. Hill and Ronald D.

Henderson. Efficient fluid simulation on the surface of a sphere. ACM Transactions on Graphics, 35(2): 16:1–16:??, May 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hullin:2010:AAB

[HHA^{+}10] Matthias B. Hullin, Johannes Hanika, Boris Ajdin, Hans-Peter Seidel, Jan Kautz, and Hendrik P. A. Lensch. Acquisition and analysis of bispectral bidirectional reflectance and reradiation distribution functions. ACM Transactions on Graphics, 29(4): 97:1–97:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsiao:2018:MVW

[HHC18] Kai-Wen Hsiao, Jia-Bin Huang, and Hung-Kuo Chu. Multiview wire art. ACM Transactions on Graphics, 37(6): 242:1–242:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hong:2019:PBF

[HHC^{+}19] Seokpyo Hong, Daseong Han, Kyungmin Cho, Joseph S. Shin, and Junyong Noh. Physics-based full-body soccer motion control for dribbling and shooting. ACM Transactions on Graphics, 38 (4):74:1–74:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2022:IPM

[HHD^{+}22] Yiwei Hu, Chengan He, Valentin Deschaintre, Julie Dorsey, and Holly Rushmeier. An inverse procedural modeling pipeline for SVBRDF maps. ACM Transactions on Graphics, 41 (2):18:1–18:17, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502431.

Heitz:2016:MSM

[HHdD16] Eric Heitz, Johannes Hanika, Eugene d’Eon, and Carsten Dachsbacher. Multiplescattering microfacet BSDFs with the Smith model. ACM Transactions on Graphics, 35 (4):58:1–58:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hanocka:2019:MNE

[HHF^{+}19] Rana Hanocka, Amir Hertz, Noa Fish, Raja Giryes, Shachar Fleishman, and Daniel Cohen-Or. MeshCNN: a network with an edge. ACM Transactions on Graphics, 38 (4):90:1–90:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heide:2013:LBT

[HHGH13] Felix Heide, Matthias B.

Hullin, James Gregson, and Wolfgang Heidrich. Lowbudget transient imaging using photonic mixer devices. ACM Transactions on Graphics, 32(4):45:1–45:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heide:2015:DTF

[HHHW15] Felix Heide, Wolfgang Heidrich, Matthias Hullin, and Gordon Wetzstein. Doppler time-of-flight imaging. ACM Transactions on Graphics, 34 (4):36:1–36:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2024:NWD

[HHL^{+}24] Jingyu Hu, Ka-Hei Hui, Zhengzhe Liu, Ruihui Li, and Chi-Wing Fu. Neural wavelet-domain diffusion for 3D shape generation, inversion, and manipulation. ACM Transactions on Graphics, 43 (2):16:1–16:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3635304.

Huang:2019:ALS

[HHM19] Libo Huang, Torsten Hädrich, and Dominik L. Michels. On the accurate large-scale simulation of ferrofluids. ACM Transactions on Graphics, 38(4):93:1–93:15, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Humphreys:2002:CSP

[HHN^{+}02] Greg Humphreys, Mike Houston, Ren Ng, Randall Frank, Sean Ahern, Peter D. Kirchner, and James T. Klosowski. Chromium: a stream-processing framework for interactive rendering on clusters. ACM Transactions on Graphics, 21 (3):693–702, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Herrera:2021:WNH

[HHP^{+}21] Jorge Alejandro Amador Herrera, Torsten Hädrich, Wojtek Pałubicki, Daniel T. Banuti, Sören Pirk, and Dominik L. Michels. Weatherscapes: nowcasting heat transfer and water continuity. ACM Transactions on Graphics, 40(6): 204:1–204:19, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480532.

Halperin:2021:ELD

[HHV^{+}21] Tavi Halperin, Hanit Hakim, Orestis Vantzos, Gershon Hochman, Netai Benaim, Lior Sassy, Michael Kupchik, Ofir Bibi, and Ohad Fried. Endless loops: detecting and animating periodic patterns in still images. ACM Transactions on Graphics, 40(4): 142:1–142:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459935.

Hu:2018:EWB

[HHX^{+}18] Yuanming Hu, Hao He, Chenxi Xu, Baoyuan Wang, and Stephen Lin. Exposure: a white-box photo postprocessing framework. ACM Transactions on Graphics, 37 (2):26:1–26:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2020:CSS

[HIK^{+}20] Weizhen Huang, Julian Iseringhausen, Tom Kneiphof, Ziyin Qu, Chenfanfu Jiang, and Matthias B. Hullin. Chemomechanical simulation of soap film flow on spherical bubbles. ACM Transactions on Graphics, 39(4): 41:1–41:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392094.

Hill:1986:SCC

[Hil86] Ralph D. Hill. Supporting concurrency, communication, and synchronization in human-computer interaction – the Sassafras UIMS. ACM Transactions on Graphics, 5(3):179–210, July 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        24055.html.

Hill:1987:ADR

[Hil87] Ralph D. Hill. Adaptive 2-D rotation control. ACM Transactions on Graphics, 6 (2):159–161, April 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hachisuka:2009:SPP

[HJ09] Toshiya Hachisuka and Henrik Wann Jensen. Stochastic progressive photon mapping. ACM Transactions on Graphics, 28(5):141:1– 141:8, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hachisuka:2011:RAP

[HJ11a] Toshiya Hachisuka and Henrik Wann Jensen. Robust adaptive photon tracing using photon path visibility. ACM Transactions on Graphics, 30(5):114:1–114:11, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hunt:2011:APT

[HJ11b] Warren A. Hunt and Gregory S. Johnson. The area perspective transform: a homogeneous transform for efficient in-volume queries. ACM Transactions on Graphics, 30 (2):8:1–8:6, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hachisuka:2010:PEE

[HJJ10] Toshiya Hachisuka, Wojciech Jarosz, and Henrik Wann Jensen. A progressive error estimation framework for photon density estimation. ACM Transactions on Graphics, 29 (6):144:1–144:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hwang:2022:SEP

[HJM^{+}22] Inseung Hwang, Daniel S. Jeon, Adolfo Muñoz, Diego Gutierrez, Xin Tong, and Min H. Kim. Sparse ellipsometry: portable acquisition of polarimetric SVBRDF and shape with unstructured flash photography. ACM Transactions on Graphics, 41 (4):133:1–133:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530075.

Huang:2014:BCP

[HJS^{+}14] Jin Huang, Tengfei Jiang, Zeyun Shi, Yiying Tong, Hujun Bao, and Mathieu Desbrun. l_{1}-based construction of polycube maps from complex shapes. ACM Transactions on Graphics, 33(3): 25:1–25:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hachisuka:2008:MAS

[HJW^{+}08] Toshiya Hachisuka, Wojciech Jarosz, Richard Peter Weistroffer, Kevin Dale, Greg Humphreys, Matthias Zwicker, and Henrik Wann Jensen. Multidimensional adaptive sampling and reconstruction for ray tracing.

ACM Transactions on Graphics, 27(3):33:1–33:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hong:2005:DF

[HK05] Jeong-Mo Hong and Chang-Hun Kim. Discontinuous fluids. ACM Transactions on Graphics, 24(3):915–920, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heo:2010:DPF

[HK10a] Nambin Heo and Hyeong-Seok Ko. Detail-preserving fully-Eulerian interface tracking framework. ACM Transactions on Graphics, 29(6): 176:1–176:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2010:PO

[HK10b] Shu-Wei Hsu and John Keyser. Piles of objects. ACM Transactions on Graphics, 29 (6):155:1–155:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2012:ACP

[HK12] Shu-Wei Hsu and John Keyser. Automated constraint placement to maintain pile shape. ACM Transactions on Graphics, 31(6): 150:1–150:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hedman:2018:IP

[HK18a] Peter Hedman and Johannes Kopf. Instant 3D photography. ACM Transactions on Graphics, 37(4):101:1– 101:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holynski:2018:FDD

[HK18b] Aleksander Holynski and Johannes Kopf. Fast depth densification for occlusion-aware augmented reality. ACM Transactions on Graphics, 37 (6):194:1–194:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2018:DIP

[HKA^{+}18] Yinghao Huang, Manuel Kaufmann, Emre Aksan, Michael J. Black, Otmar Hilliges, and Gerard Pons-Moll. Deep inertial poser: learning to reconstruct human pose from sparse inertial measurements in real time. ACM Transactions on Graphics, 37 (6):185:1–185:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2014:ICU

[HKAK14] Jia-Bin Huang, Sing Bing Kang, Narendra Ahuja, and Johannes Kopf. Image completion using planar structure guidance. ACM Transactions on Graphics, 33(4): 129:1–129:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2016:TCC

[HKAK16] Jia-Bin Huang, Sing Bing Kang, Narendra Ahuja, and Johannes Kopf. Temporally coherent completion of dynamic video. ACM Transactions on Graphics, 35(6): 196:1–196:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2018:LLS

[HKC^{+}18] Haibin Huang, Evangelos Kalogerakis, Siddhartha Chaudhuri, Duygu Ceylan, Vladimir G. Kim, and Ersin Yumer. Learning local shape descriptors from part correspondences with multiview convolutional networks. ACM Transactions on Graphics, 37(1):6:1– 6:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hachisuka:2014:MML

[HKD14] Toshiya Hachisuka, Anton S.

Kaplanyan, and Carsten Dachsbacher. Multiplexed Metropolis light transport. ACM Transactions on Graphics, 33(4):100:1–100:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2011:JSS

[HKG11] Qixing Huang, Vladlen Koltun, and Leonidas Guibas. Joint shape segmentation with linear programming. ACM Transactions on Graphics, 30 (6):125:1–125:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holden:2020:LMM

[HKPP20] Daniel Holden, Oussama Kanoun, Maksym Perepichka, and Tiberiu Popa. Learned motion matching. ACM Transactions on Graphics, 39 (4):53:1–53:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392440.

Kim:2003:RMS

[hKPS03] Tae hoon Kim, Sang Il Park, and Sung Yong Shin. Rhythmic-motion synthesis based on motion-beat analysis. ACM Transactions on Graphics, 22(3):392–401, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holden:2017:PFN

[HKS17] Daniel Holden, Taku Komura, and Jun Saito. Phasefunctioned neural networks for character control. ACM Transactions on Graphics, 36 (4):42:1–42:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ho:2010:SRP

[HKT10] Edmond S. L. Ho, Taku Komura, and Chiew-Lan Tai. Spatial relationship preserving character motion adaptation. ACM Transactions on Graphics, 29(4): 33:1–33:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hasan:2009:VSL

[HKWB09] Miloš Hašan, Jaroslav Křivánek, Bruce Walter, and Kavita Bala. Virtual spherical lights for many-light rendering of glossy scenes. ACM Transactions on Graphics, 28(5): 143:1–143:6, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ha:2014:ITD

[HL14] Sehoon Ha and C. Karen Liu. Iterative training of dynamic skills inspired by human coaching techniques. ACM Transactions on Graphics, 34(1):1:1–1:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2012:COA

[HLBR12] Fu-Chung Huang, Douglas Lanman, Brian A. Barsky, and Ramesh Raskar. Correcting for optical aberrations using multilayer displays. ACM Transactions on Graphics, 31 (6):185:1–185:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2019:PCT

[HLC^{+}19] Mingming He, Jing Liao, Dongdong Chen, Lu Yuan, and Pedro V. Sander. Progressive color transfer with dense semantic correspondences. ACM Transactions on Graphics, 38(2): 13:1–13:??, April 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3292482.

Hu:2022:SBM

[HLG^{+}22] Shi-Min Hu, Zheng-Ning Liu, Meng-Hao Guo, Jun-Xiong Cai, Jiahui Huang, Tai-Jiang Mu, and Ralph R. Martin. Subdivision-based mesh convolution networks. ACM Transactions on Graphics, 41 (3):25:1–25:16, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3506694.

Hirsch:2009:BST

[HLHR09] Matthew Hirsch, Douglas Lanman, Henry Holtzman, and Ramesh Raskar. BiDi screen: a thin, depth-sensing LCD for 3D interaction using light fields. ACM Transactions on Graphics, 28(5): 159:1–159:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holroyd:2008:PAE

[HLHZ08] Michael Holroyd, Jason Lawrence, Greg Humphreys, and Todd Zickler. A photometric approach for estimating normals and tangents. ACM Transactions on Graphics, 27(5): 133:1–133:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Himeur:2022:PLN

[HLP^{+}22] Chems-Eddine Himeur, Thibault Lejemble, Thomas Pellegrini, Mathias Paulin, Loic Barthe, and Nicolas Mellado. PCED-Net: a lightweight neural network for fast and interactive edge detection in 3D point clouds. ACM Transactions on Graphics, 41(1): 10:1–10:21, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3481804.

Heide:2014:CDS

[HLR^{+}14] Felix Heide, Douglas Lanman, Dikpal Reddy, Jan Kautz, Kari Pulli, and David Luebke. Cascaded displays: spatiotemporal superresolution using offset pixel layers. ACM Transactions on Graphics, 33 (4):60:1–60:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hennessey:2017:TIB

[HLR^{+}17] James W. Hennessey, Wilmot Li, Bryan Russell, Eli Shechtman, and Niloy J. Mitra. Transferring image-based edits for multi-channel compositing. ACM Transactions on Graphics, 36(6):179:1– 179:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2018:GPV

[HLSH18] Mingming He, Jing Liao, Pedro V. Sander, and Hugues Hoppe. Gigapixel panorama video loops. ACM Transactions on Graphics, 37(1): 3:1–3:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hecht:2012:USC

[HLSO12] Florian Hecht, Yeon Jin Lee, Jonathan R. Shewchuk, and James F. O’Brien. Updated sparse Cholesky factors for corotational elastodynamics. ACM Transactions on Graphics, 31(5): 123:1–123:13, August 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2017:CLSa

[HLV^{+}17a] Ruizhen Hu, Wenchao Li, Oliver Van Kaick, Hui Huang, Melinos Averkiou, Daniel Cohen-Or, and Hao Zhang. Co-locating style-defining elements on 3D shapes. ACM Transactions on Graphics, 36 (3):33:1–33:??, June 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2017:CLSb

[HLV^{+}17b] Ruizhen Hu, Wenchao Li, Oliver Van Kaick, Hui Huang, Melinos Averkiou, Daniel Cohen-Or, and Hao Zhang. Co-locating style-defining elements on 3D shapes. ACM Transactions on Graphics, 36 (4):50:1–50:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2017:LPP

[HLV^{+}17c] Ruizhen Hu, Wenchao Li, Oliver Van Kaick, Ariel Shamir, Hao Zhang, and Hui Huang. Learning to predict part mobility from a single static snapshot. ACM Transactions on Graphics, 36(6): 227:1–227:??, November 2017.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2012:SMS

[HLW^{+}12] Xiaowei He, Ning Liu, Guoping Wang, Fengjun Zhang, Sheng Li, Songdong Shao, and Hongan Wang. Staggered meshless solid-fluid coupling. ACM Transactions on Graphics, 31(6):149:1– 149:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Han:2018:OOM

[HLW^{+}18] Shangchen Han, Beibei Liu, Robert Wang, Yuting Ye, Christopher D. Twigg, and Kenrick Kin. Online optical marker-based hand tracking with deep labels. ACM Transactions on Graphics, 37(4): 166:1–166:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2019:TMS

[HLW^{+}19] Qixing Huang, Zhenxiao Liang, Haoyun Wang, Simiao Zuo, and Chandrajit Bajaj. Tensor maps for synchronizing heterogeneous shape collections. ACM Transactions on Graphics, 38(4): 106:1–106:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Habermann:2021:RTD

[HLX^{+}21] Marc Habermann, Lingjie Liu, Weipeng Xu, Michael Zollhoefer, Gerard Pons-Moll, and Christian Theobalt. Realtime deep dynamic characters. ACM Transactions on Graphics, 40(4):94:1–94:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459749.

Hu:2021:QCQ

[HLY^{+}21] Yuanming Hu, Jiafeng Liu, Xuanda Yang, Mingkuan Xu, Ye Kuang, Weiwei Xu, Qiang Dai, William T. Freeman, and Frédo Durand. QuanTaichi: a compiler for quantized simulations. ACM Transactions on Graphics, 40(4): 182:1–182:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459671.

Hong:2008:BA

[HLYK08] Jeong-Mo Hong, Ho-Young Lee, Jong-Chul Yoon, and Chang-Hun Kim. Bubbles alive. ACM Transactions on Graphics, 27(3):48:1– 48:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2009:CUP

[HLZ^{+}09] Hui Huang, Dan Li, Hao Zhang, Uri Ascher, and Daniel Cohen-Or. Consolidation of unorganized point clouds for surface reconstruction. ACM Transactions on Graphics, 28(5):176:1–176:7, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holroyd:2010:COS

[HLZ10] Michael Holroyd, Jason Lawrence, and Todd Zickler. A coaxial optical scanner for synchronous acquisition of 3D geometry and surface reflectance. ACM Transactions on Graphics, 29(4): 99:1–99:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2014:APS

[HLZCO14] Ruizhen Hu, Honghua Li, Hao Zhang, and Daniel Cohen-Or. Approximate pyramidal shape decomposition. ACM Transactions on Graphics, 33(6): 213:1–213:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Haase:1992:MPM

[HM92] Chet S. Haase and Gary W. Meyer. Modeling pigmented materials for realistic image synthesis. ACM Transactions on Graphics, 11(4):305– 335, October 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
146452.html.

Huang:2020:SOF

[HM20] Libo Huang and Dominik L.

Michels. Surface-only ferrofluids. ACM Transactions on Graphics, 39(6): 174:1–174:17, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417799.

Hasselgren:2009:APT

[HMAM09] Jon Hasselgren, Jacob Munkberg, and Tomas Akenine-Möller. Automatic pre-tessellation culling. ACM Transactions on Graphics, 28(2):19:1– 19:10, April 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2011:RFM

[HMC11] Wei-Hsien Hsu, Kwan-Liu Ma, and Carlos Correa. A rendering framework for multiscale views of 3D models. ACM Transactions on Graphics, 30(6):131:1–131:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hillesland:2003:NOF

[HMG03] Karl E. Hillesland, Sergey Molinov, and Radek Grzeszczuk.

Nonlinear optimization framework for image-based modeling on programmable graphics hardware. ACM Transactions on Graphics, 22(3): 925–934, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2023:DFH

[HMI23] Yunpu Hu, Leo Miyashita, and Masatoshi Ishikawa. Differential frequency heterodyne time-of-flight imaging for instantaneous depth and velocity estimation. ACM Transactions on Graphics, 42 (1):9:1–9:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3546939.

Hu:2014:CBH

[HML^{+}14] Liwen Hu, Chongyang Ma, Linjie Luo, Li-Yi Wei, and Hao Li. Capturing braided hairstyles. ACM Transactions on Graphics, 33(6): 225:1–225:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harish:2016:PIK

[HMLB16] Pawan Harish, Mentar Mahmudi, Benoît Le Callennec, and Ronan Boulic. Parallel inverse kinematics for multithreaded architectures. ACM Transactions on Graphics, 35(2):19:1–19:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2014:RHC

[HMLL14] Liwen Hu, Chongyang Ma, Linjie Luo, and Hao Li. Robust hair capture using simulated examples. ACM Transactions on Graphics, 33(4): 126:1–126:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2015:SVH

[HMLL15] Liwen Hu, Chongyang Ma, Linjie Luo, and Hao Li. Single-view hair modeling using a hairstyle database. ACM Transactions on Graphics, 34(4):125:1–125:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2021:SCO

[HMM^{+}21] Shi-Sheng Huang, Ze-Yu Ma, Tai-Jiang Mu, Hongbo Fu, and Shi-Min Hu. Supervoxel convolution for online 3D semantic segmentation. ACM Transactions on Graphics, 40(3):34:1–34:15, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hoyet:2012:PIR

[HMO12] Ludovic Hoyet, Rachel Mc-Donnell, and Carol O’Sullivan. Push it real: perceiving causality in virtual interactions. ACM Transactions on Graphics, 31(4): 90:1–90:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2008:LME

[HMP^{+}08] Eugene Hsu, Tom Mertens, Sylvain Paris, Shai Avidan, and Frédo Durand. Light mixture estimation for spatially varying white balance. ACM Transactions on Graphics, 27 (3):70:1–70:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hadrich:2020:SSC

[HMP^{+}20] Torsten Hädrich, Miłosz Makowski, Wojtek Pałubicki, Daniel T. Banuti, Sören Pirk, and Dominik L. Michels. Stormscapes: simulating cloud dynamics in the now. ACM Transactions on Graphics, 39(6):175:1–175:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417801.

Haber:2005:PBS

[HMS05] Jörg Haber, Marcus Magnor, and Hans-Peter Seidel. Physically-based simulation of twilight phenomena. ACM Transactions on Graphics, 24 (4):1353–1373, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hahn:2012:RSP

[HMT^{+}12] Fabian Hahn, Sebastian Martin, Bernhard Thomaszewski, Robert Sumner, Stelian Coros, and Markus Gross. Rigspace physics. ACM Transactions on Graphics, 31(4): 72:1–72:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2015:IMT

[HMT^{+}15] Jiawei Huang, Tsuyoshi Mori, Kazuki Takashima, Shuichiro Hashi, and Yoshifumi Kitamura. IM6D: magnetic tracking system with 6-DOF passive markers for dexterous 3D interaction and motion. ACM Transactions on Graphics, 34 (6):217:1–217:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Houston:2006:HRL

[HNB^{+}06] Ben Houston, Michael B.

Nielsen, Christopher Batty, Ola Nilsson, and Ken Museth. Hierarchical RLE level set: a compact and versatile deformable surface representation. ACM Transactions on Graphics, 25(1):151– 175, January 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hepp:2019:PVT

[HNH19] Benjamin Hepp, Matthias Nießner, and Otmar Hilliges. Plan3D: Viewpoint and trajectory optimization for aerial multi-view stereo reconstruction. ACM Transactions on Graphics, 38(1):4:1–4:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3233794.

Hamamichi:2023:NNV

[HNO^{+}23] Mitsuki Hamamichi, Kentaro Nagasawa, Masato Okada, Ryohei Seto, and Yonghao Yue. Non-Newtonian ViRheometry via similarity analysis. ACM Transactions on Graphics, 42(6): 193:1–193:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618310.

Hobby:1990:RNC

[Hob90]  John D. Hobby.  Ras-

terization of nonparametric curves. ACM Transactions on Graphics, 9(3):262–277, July 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        78966.html.

Hobby:1991:NSI

[Hob91] John D. Hobby. Numerically stable implicitization of cubic curves. ACM Transactions on Graphics, 10(3):255–296, July 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
108546.html.

Hodgins:2000:E

[Hod00]  Jessica Hodgins.  Edito-

rial. ACM Transactions on Graphics, 19(3):163, 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tog/2000-
19-3/p163-hodgins/p163-
hodgins.pdf;   http://www.
acm.org/pubs/citations/
journals/tog/2000-19-3/
p163-hodgins/.

Hodgins:2002:A

[Hod02a] Jessica Hodgins. Acknowledgments. ACM Transactions on Graphics, 21(2):230, April 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hodgins:2002:E

[Hod02b] Jessica Hodgins. Editorial.

ACM Transactions on Graphics, 21(2):87, April 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hodgins:2003:E

[Hod03] Jessica Hodgins. Editorial. ACM Transactions on Graphics, 22(1):1, January 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hachisuka:2008:PPM

[HOJ08] Toshiya Hachisuka, Shinji Ogaki, and Henrik Wann Jensen. Progressive photon mapping. ACM Transactions on Graphics, 27(5): 130:1–130:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hoyet:2016:PES

[HOKP16] Ludovic Hoyet, Anne-Helene Olivier, Richard Kulpa, and Julien Pettré. Perceptual effect of shoulder motions on crowd animations. ACM Transactions on Graphics, 35 (4):53:1–53:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holden:2018:RSO

[Hol18] Daniel Holden. Robust solving of optical motion capture data by denoising. ACM Transactions on Graphics, 37(4): 165:1–165:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hueting:2015:CJU

[HOM15] Moos Hueting, Maks Ovsjanikov, and Niloy J. Mitra. CrossLink: joint understanding of image and 3D model collections through shape and camera pose variations. ACM Transactions on Graphics, 34 (6):233:1–233:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heide:2019:NLS

[HOZ^{+}19] Felix Heide, Matthew O’Toole, Kai Zang, David B. Lindell, Steven Diamond, and Gordon Wetzstein. Non-line-of-sight imaging with partial occluders and surface normals. ACM Transactions on Graphics, 38 (3):22:1–22:??, June 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3269977.

Han:2003:MBT

[HP03] Jefferson Y. Han and Ken Perlin. Measuring bidirectional texture reflectance with a kaleidoscope. ACM Transactions on Graphics, 22(3): 741–748, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hofer:2004:EMS

[HP04] Michael Hofer and Helmut Pottmann. Energyminimizing splines in manifolds. ACM Transactions on Graphics, 23(3):284–293, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holzschuch:2017:TSM

[HP17] Nicolas Holzschuch and Romain Pacanowski. A twoscale microfacet reflectance model combining reflection and diffraction. ACM Transactions on Graphics, 36(4): 66:1–66:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hasan:2006:DIT

[HPB06] Miloš Hašan, Fabio Pellacini, and Kavita Bala. Directto-indirect transfer for cinematic relighting. ACM Transactions on Graphics, 25(3): 1089–1097, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hasan:2007:MRC

[HPB07] Miloš Hašan, Fabio Pellacini, and Kavita Bala. Matrix row-column sampling for the many-light problem. ACM Transactions on Graphics, 26 (3):26:1–26:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huber:2021:DAS

[HPC21] Simon Huber, Roi Poranne, and Stelian Coros. Designing actuation systems for animatronic figures via globally optimal discrete search. ACM Transactions on Graphics, 40(4):174:1–174:10, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459867.

Hertz:2022:SEI

[HPG^{+}22] Amir Hertz, Or Perel, Raja Giryes, Olga Sorkine-Hornung, and Daniel Cohen-Or. SPAGHETTI: editing implicit shapes through part aware generation. ACM Transactions on Graphics, 41 (4):106:1–106:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530084.

Hachisuka:2012:PSE

[HPJ12] Toshiya Hachisuka, Jacopo Pantaleoni, and Henrik Wann Jensen. A path space extension for robust light transport simulation. ACM Transactions on Graphics, 31(6): 191:1–191:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2017:MPF

[HPK^{+}17] Fu-Chung Huang, Dawid Pajak, Jonghyun Kim, Jan Kautz, and David Luebke. Mixed-primary factorization for dual-frame computational displays. ACM Transactions on Graphics, 36(4): 149:1–149:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2005:STH

[HPP05] Eugene Hsu, Kari Pulli, and Jovan Popović. Style translation for human motion. ACM Transactions on Graphics, 24 (3):1082–1089, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hedman:2018:DBF

[HPP^{+}18] Peter Hedman, Julien Philip, True Price, Jan-Michael Frahm, George Drettakis, and Gabriel Brostow. Deep blending for free-viewpoint imagebased rendering. ACM Transactions on Graphics, 37(6): 257:1–257:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2022:NFL

[HPP^{+}22] Zhaoyang Huang, Xiaokun Pan, Weihong Pan, Weikang Bian, Yan Xu, Ka Chun Cheung, Guofeng Zhang, and Hongsheng Li. NeuralMarker: a framework for learning general marker correspondence. ACM Transactions on Graphics, 41(6):

271:1–271:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555468.

Harmon:2011:IAG

[HPSZ11] David Harmon, Daniele Panozzo, Olga Sorkine, and Denis Zorin. Interference-aware geometric modeling. ACM Transactions on Graphics, 30(6): 137:1–137:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hou:2010:MRT

[HQL^{+}10] Qiming Hou, Hao Qin, Wenyao Li, Baining Guo, and Kun Zhou. Micropolygon ray tracing with defocus and motion blur. ACM Transactions on Graphics, 29(4): 64:1–64:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2021:SSW

[HQT^{+}21] Libo Huang, Ziyin Qu, Xun Tan, Xinxin Zhang, Dominik L. Michels, and Chenfanfu Jiang. Ships, splashes, and waves on a vast ocean. ACM Transactions on Graphics, 40(6): 203:1–203:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480495.

Hable:2005:BGB

[HR05] John Hable and Jarek Rossignac. Blister: GPU-based rendering of Boolean combinations of free-form triangulated shapes. ACM Transactions on Graphics, 24(3):1024–1031, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hasan:2013:IAE

[HR13] Milovs Hasan and Ravi Ramamoorthi. Interactive albedo editing in path-traced volumetric materials. ACM Transactions on Graphics, 32 (2):11:1–11:11, April 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hedman:2016:SII

[HRDB16] Peter Hedman, Tobias Ritschel, George Drettakis, and Gabriel Brostow. Scalable inside-out image-based rendering. ACM Transactions on Graphics, 35 (6):231:1–231:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hecker:2008:RTM

[HRE^{+}08] Chris Hecker, Bernd Raabe, Ryan W. Enslow, John De-Weese, Jordan Maynard, and Kees van Prooijen. Realtime motion retargeting to highly varied user-created morphologies. ACM Transactions on Graphics, 27(3): 27:1–27:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heide:2013:HQC

[HRH^{+}13] Felix Heide, Mushfiqur Rouf, Matthias B. Hullin, Bjorn Labitzke, Wolfgang Heidrich, and Andreas Kolb. Highquality computational imaging through simple lenses. ACM Transactions on Graphics, 32(5):149:1–149:14, September 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hamalainen:2015:OCS

[HRL15] Perttu Hämäläinen, Joose Rajamäki, and C. Karen Liu. Online control of simulated humanoids using particle belief propagation. ACM Transactions on Graphics, 34(4): 81:1–81:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Han:2008:MTS

[HRRG08] Charles Han, Eric Risser, Ravi Ramamoorthi, and Eitan Grinspun. Multiscale texture synthesis. ACM Transactions on Graphics, 27(3): 51:1–51:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2023:PEP

[HRS^{+}23] Xingchang Huang, Tobias Ritschel, Hans-Peter Seidel, Pooran Memari, and Gurprit Singh. Patternshop: Editing point patterns by image manipulation. ACM Transactions on Graphics, 42 (4):53:1–53:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592418.

Hill:1997:CAQ

[HRV97] Bernhard Hill, Thomas Roger, and Friedrich Wilhelm Vorhagen. Comparative analysis of the quantization of color spaces on the basis of the CIELAB color-difference formula. ACM Transactions on Graphics, 16(2): 109–154, April 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1997-16-2/
p109-hill/.

Hermosilla:2018:MCC

[HRV^{+}18] Pedro Hermosilla, Tobias Ritschel, Pere-Pau Vázquez, Àlvar Vinacua, and Timo Ropinski. Monte Carlo convolution for learning on non-uniformly sampled point clouds. ACM Transactions on Graphics, 37(6):235:1– 235:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harrison:2004:OLC

[HRvdP04] Jason Harrison, Ronald A.

Rensink, and Michiel van de Panne. Obscuring length changes during animated motion. ACM Transactions on Graphics, 23(3):569– 573, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hoyet:2013:EDA

[HRZ^{+}13] Ludovic Hoyet, Kenneth Ryall, Katja Zibrek, Hwangpil Park, Jehee Lee, Jessica Hodgins, and Carol O’Sullivan. Evaluating the distinctiveness and attractiveness of human motions on realistic virtual bodies. ACM Transactions on Graphics, 32(6): 204:1–204:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2013:MDM

[HS13] Lei He and Scott Schaefer. Mesh denoising via L_{0} minimization. ACM Transactions on Graphics, 32(4): 64:1–64:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hadwiger:2012:SPM

[HSB^{+}12] Markus Hadwiger, Ronell Sicat, Johanna Beyer, Jens Krüger, and Torsten Möller. Sparse PDF maps for nonlinear multi-resolution image operations. ACM Transactions on Graphics, 31(6): 133:1–133:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2022:PSM

[HSC^{+}22] Ruizhen Hu, Xiangyu Su, Xiangkai Chen, Oliver Van Kaick, and Hui Huang. Phototo-shape material transfer for diverse structures. ACM Transactions on Graphics, 41 (4):131:1–131:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530088.

Heck:2013:BNS

[HSD13] Daniel Heck, Thomas Schlömer, and Oliver Deussen. Blue noise sampling with controlled aliasing. ACM Transactions on Graphics, 32(3): 25:1–25:12, June 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hong:2007:WFC

[HSF07] Jeong-Mo Hong, Tamar Shinar, and Ronald Fedkiw. Wrinkled flames and cellular patterns. ACM Transactions on Graphics, 26(3): 47:1–47:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2013:FGS

[HSG13] Qi-Xing Huang, Hao Su, and Leonidas Guibas. Finegrained semi-supervised labeling of large shape collections. ACM Transactions on Graphics, 32(6):190:1– 190:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hasinoff:2016:BPH

[HSG^{+}16] Samuel W. Hasinoff, Dillon Sharlet, Ryan Geiss, Andrew Adams, Jonathan T. Barron, Florian Kainz, Jiawen Chen, and Marc Levoy. Burst photography for high dynamic range and low-light imaging on mobile cameras. ACM Transactions on Graphics, 35 (6):192:1–192:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2019:TRT

[HSG^{+}19] Yixin Hu, Teseo Schneider, Xifeng Gao, Qingnan Zhou, Alec Jacobson, Denis Zorin, and Daniele Panozzo. Tri-Wild: robust triangulation with curve constraints. ACM Transactions on Graphics, 38 (4):52:1–52:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

HaCohen:2011:NRD

[HSGL11] Yoav HaCohen, Eli Shechtman, Dan B. Goldman, and Dani Lischinski. Non-rigid dense correspondence with applications for image enhancement. ACM Transactions on Graphics, 30(4): 70:1–70:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

HaCohen:2013:OCC

[HSGL13] Yoav HaCohen, Eli Shechtman, Dan B. Goldman, and Dani Lischinski. Optimizing color consistency in photo collections. ACM Transactions on Graphics, 32(4): 38:1–38:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

He:2010:PSS

[HSH10] Lei He, Scott Schaefer, and Kai Hormann. Parameterizing subdivision surfaces. ACM Transactions on Graphics, 29 (4):120:1–120:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Herholz:2020:SCU

[HSH20] Philipp Herholz and Olga Sorkine-Hornung. Sparse Cholesky updates for interactive mesh parameterization. ACM Transactions on Graphics, 39(6):

202:1–202:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417828.

Hoskinson:2010:LRH

[HSHF10] Reynald Hoskinson, Boris Stoeber, Wolfgang Heidrich, and Sidney Fels. Light reallocation for high contrast projection using an analog micromirror array. ACM Transactions on Graphics, 29(6): 165:1–165:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Holden:2016:DLF

[HSK16] Daniel Holden, Jun Saito, and Taku Komura. A deep learning framework for character motion synthesis and editing. ACM Transactions on Graphics, 35(4):138:1–138:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2006:SGD

[HSL^{+}06] Jin Huang, Xiaohan Shi, Xinguo Liu, Kun Zhou, Li-Yi Wei, Shang-Hua Teng, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Subspace gradient domain mesh deformation. ACM Transactions on Graphics, 25(3): 1126–1134, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Han:2007:FDN

[HSRG07] Charles Han, Bo Sun, Ravi Ramamoorthi, and Eitan Grinspun. Frequency domain normal map filtering. ACM Transactions on Graphics, 26 (3):28:1–28:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heidrich:1998:SPS

[HSS98] Wolfgang Heidrich, Philip Slusallek, and Hans-Peter Seidel. Sampling procedural shaders using affine arithmetic. ACM Transactions on Graphics, 17(3):158–176, July 1998. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tog/
                                        1998-17-3/p158-heidrich/.

Huang:2013:QOC

[HSS^{+}13] Shi-Sheng Huang, Ariel Shamir, Chao-Hui Shen, Hao Zhang, Alla Sheffer, Shi-Min Hu, and Daniel Cohen-Or. Qualitative organization of collections of shapes via quartet analysis. ACM Transactions on Graphics, 32(4):71:1–71:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Halli:2010:ERM

[HSST10] Akram Halli, Abderrahim Saaidi, Khalid Satori, and Hamid Tairi. Extrusion and revolution mapping. ACM Transactions on Graphics, 29(5):132:1–132:14, October 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heide:2014:FFC

[HST^{+}14] Felix Heide, Markus Steinberger, Yun-Ta Tsai, Mushfiqur Rouf, Dawid Pajak, Dikpal Reddy, Orazio Gallo, Jing Liu, Wolfgang Heidrich, Karen Egiazarian, Jan Kautz, and Kari Pulli. FlexISP: a flexible camera image processing framework. ACM Transactions on Graphics, 33(6): 231:1–231:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hildebrandt:2011:ISM

[HSTP11] Klaus Hildebrandt, Christian Schulz, Christoph Von Tycowicz, and Konrad Polthier. Interactive surface modeling using modal analysis. ACM Transactions on Graphics, 30(5):119:1–119:11, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hladky:2022:QQB

[HSV^{+}22] Jozef Hladky, Michael Stengel, Nicholas Vining, Bernhard Kerbl, Hans-Peter Seidel, and Markus Steinberger. QuadStream: a quad-based scene streaming architecture for novel viewpoint reconstruction. ACM Transactions on Graphics, 41(6): 233:1–233:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555524.

Hildebrandt:2012:ISC

[HSvTP12] Klaus Hildebrandt, Christian Schulz, Christoph von Tycowicz, and Konrad Polthier. Interactive spacetime control of deformable objects. ACM Transactions on Graphics, 31 (4):71:1–71:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2017:ADS

[HSW^{+}17] Liwen Hu, Shunsuke Saito, Lingyu Wei, Koki Nagano, Jaewoo Seo, Jens Fursund, Iman Sadeghi, Carrie Sun, Yen-Chun Chen, and Hao Li. Avatar digitization from a single image for real-time rendering. ACM Transactions on Graphics, 36(6):195:1– 195:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Halimi:2022:PBC

[HSX^{+}22] Oshri Halimi, Tuur Stuyck, Donglai Xiang, Timur Bagautdinov, He Wen, Ron Kimmel, Takaaki Shiratori, Chenglei Wu, Yaser Sheikh, and Fabian Prada. Pattern-based cloth registration and sparseview animation. ACM Transactions on Graphics, 41(6): 196:1–196:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555448.

Hahn:2014:SCS

[HTC^{+}14] Fabian Hahn, Bernhard Thomaszewski, Stelian Coros, Robert W. Sumner, Forrester Cole, Mark Meyer, Tony DeRose, and Markus Gross. Subspace clothing simulation using adaptive bases. ACM Transactions on Graphics, 33(4): 105:1–105:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2015:HSS

[HTCH15] Peng Huang, Margara Tejera, John Collomosse, and Adrian Hilton. Hybrid skeletalsurface motion graphs for character animation from 4D performance capture. ACM Transactions on Graphics, 34 (2):17:1–17:??, February 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Healey:2004:PBB

[HTER04] Christopher G. Healey, Laura Tateosian, James T. Enns, and Mark Remple. Perceptually based brush strokes for nonphotorealistic visualization. ACM Transactions on Graphics, 23(1):64– 96, January 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harary:2014:CBC

[HTG14] Gur Harary, Ayellet Tal, and Eitan Grinspun. Contextbased coherent surface completion. ACM Transactions on Graphics, 33(1):5:1– 5:12, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Herholz:2022:SSC

[HTS^{+}22] Philipp Herholz, Xuan Tang, Teseo Schneider, Shoaib Kamil, Daniele Panozzo, and Olga Sorkine-Hornung. Sparsity-specific code optimization using expression trees. ACM Transactions on Graphics, 41(5):175:1– 175:??, October 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3520484.

Huang:2011:BAS

[HTWB11] Jin Huang, Yiying Tong, Hongyu Wei, and Hujun Bao. Boundary aligned smooth 3D cross-frame field. ACM Transactions on Graphics, 30(6): 143:1–143:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2022:GTS

[HTYW22] Jerry Hsu, Nghia Truong, Cem Yuksel, and Kui Wu. A general two-stage initialization for sag-free deformable simulations. ACM Transactions on Graphics, 41(4): 64:1–64:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530165.

Hubbard:1996:APS

[Hub96] Philip M. Hubbard. Approximating polyhedra with spheres for time-critical collision detection. ACM Transactions on Graphics, 15(3): 179–210, July 1996. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
230518.html. MPEG an-

imations showing the algorithm’s performance are available on the World-Wide

Web at http://www.acm.
org/tog/hubbard96/index.
html.

Hudson:1992:ASC

[Hud92] Scott E. Hudson. Adding shadows to a 3D cursor. ACM Transactions on Graphics, 11 (2):193–199, April 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hudson:1994:UIS

[Hud94]  Scott E. Hudson.  User

interface specification using an enhanced spreadsheet model. ACM Transactions on Graphics, 13(3):209–239, July 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        195787.html.

Hao:2004:RTR

[HV04] Xuejun Hao and Amitabh Varshney. Real-time rendering of translucent meshes. ACM Transactions on Graphics, 23(2):120–142, April 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2016:LHO

[HvKW^{+}16] Ruizhen Hu, Oliver van Kaick, Bojian Wu, Hui Huang, Ariel Shamir, and Hao Zhang. Learning how objects function via co-analysis of interactions. ACM Transactions on Graphics, 35(4):47:1–47:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harmon:2009:ACM

[HVS^{+}09] David Harmon, Etienne Vouga, Breannan Smith, Rasmus Tamstorf, and Eitan Grinspun. Asynchronous contact mechanics. ACM Transactions on Graphics, 28(3):

87:1–87:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harmon:2008:RTS

[HVTG08] David Harmon, Etienne Vouga, Rasmus Tamstorf, and Eitan Grinspun. Robust treatment of simultaneous collisions. ACM Transactions on Graphics, 27(3): 23:1–23:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hosek:2012:AMF

[HW12] Lukas Hosek and Alexander Wilkie. An analytic model for full spectral skydome radiance. ACM Transactions on Graphics, 31(4): 95:1–95:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hahn:2015:HRB

[HW15] David Hahn and Chris Wojtan. High-resolution brittle fracture simulation with boundary elements. ACM Transactions on Graphics, 34(4):151:1–151:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hahn:2016:FAB

[HW16] David Hahn and Chris Wojtan. Fast approximations for boundary element based brittle fracture simulation. ACM Transactions on Graphics, 35 (4):104:1–104:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heistermann:2023:MDF

[HWB23] Martin Heistermann, Jethro Warnett, and David Bommes. Min-deviation-flow in bidirected graphs for T-mesh quantization. ACM Transactions on Graphics, 42(4): 70:1–70:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592437.

Huang:2014:EFD

[HWBR14] Fu-Chung Huang, Gordon Wetzstein, Brian A. Barsky, and Ramesh Raskar. Eyeglassesfree display: towards correcting visual aberrations with computational light field displays. ACM Transactions on Graphics, 33(4): 59:1–59:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2013:MSP

[HWCO^{+}13] Hui Huang, Shihao Wu, Daniel Cohen-Or, Minglun Gong, Hao Zhang, Guiqing Li, and Baoquan Chen. L_{1}- medial skeleton of point cloud. ACM Transactions on Graphics, 32(4):65:1–65:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2013:EAP

[HWG^{+}13] Hui Huang, Shihao Wu, Minglun Gong, Daniel Cohen-Or, Uri Ascher, and Hao (Richard) Zhang. Edge-aware point set resampling. ACM Transactions on Graphics, 32(1):9:1– 9:12, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2014:FMN

[HWG14] Qixing Huang, Fan Wang, and Leonidas Guibas. Functional map networks for analyzing and exploring large shape collections. ACM Transactions on Graphics, 33(4): 36:1–36:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huo:2016:AMC

[HWH^{+}16] Yuchi Huo, Rui Wang, Tianlei Hu, Wei Hua, and Hujun Bao. Adaptive matrix column sampling and completion for rendering participating media. ACM Transactions on Graphics, 35(6): 167:1–167:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Han:2018:DUP

[HWH^{+}18] Chu Han, Qiang Wen, Shengfeng He, Qianshu Zhu, Yinjie Tan, Guoqiang Han, and Tien-Tsin Wong. Deep unsupervised pixelization. ACM Transactions on Graphics, 37(6): 243:1–243:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huo:2015:MSR

[HWJ^{+}15] Yuchi Huo, Rui Wang, Shihao Jin, Xinguo Liu, and Hujun Bao. A matrix sampling-and-recovery approach for manylights rendering. ACM Transactions on Graphics, 34(6): 210:1–210:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2015:SVR

[HWK15] Qixing Huang, Hai Wang, and Vladlen Koltun. Singleview reconstruction via joint analysis of image and shape collections. ACM Transactions on Graphics, 34(4):87:1– 87:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hsu:2023:SFI

[HWP^{+}23] Jerry Hsu, Tongtong Wang, Zherong Pan, Xifeng Gao, Cem Yuksel, and Kui Wu. Sag-free initialization for strand-based hybrid hair simulation. ACM Transactions on Graphics, 42(4): 74:1–74:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592143.

Hirsch:2014:CLF

[HWR14] Matthew Hirsch, Gordon Wetzstein, and Ramesh Raskar. A compressive light field projection system. ACM Transactions on Graphics, 33(4): 58:1–58:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Heide:2013:AIS

[HWRH13] Felix Heide, Gordon Wetzstein, Ramesh Raskar, and Wolfgang Heidrich. Adaptive image synthesis for compressive displays. ACM Transactions on Graphics, 32(4): 132:1–132:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2018:SOR

[HWV^{+}18] Ruizhen Hu, Cheng Wen, Oliver Van Kaick, Luanmin Chen, Di Lin, Daniel Cohen-Or, and Hui Huang. Semantic object reconstruction via casual handheld scanning. ACM Transactions on Graphics, 37 (6):219:1–219:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hou:2022:IPS

[HWW^{+}22] Fei Hou, Chiyu Wang, Wencheng Wang, Hong Qin, Chen Qian, and Ying He. Iterative Poisson surface reconstruction (iPSR) for unoriented points. ACM Transactions on Graphics, 41 (4):128:1–128:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530096.

He:2014:RSS

[HWZ^{+}14] Xiaowei He, Huamin Wang, Fengjun Zhang, Hongan Wang, Guoping Wang, and Kun Zhou. Robust simulation of sparsely sampled thin features in SPH-based free surface flows. ACM Transactions on Graphics, 34(1):7:1– 7:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huo:2020:AIR

[HWZ^{+}20] Yuchi Huo, Rui Wang, Ruzahng Zheng, Hualin Xu, Hujun Bao, and Sung-Eui Yoon. Adaptive incident radiance field sampling and reconstruction using deep reinforcement learning. ACM Transactions on Graphics, 39(1): 6:1–6:17, February 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.

acm.org/doi/abs/10.1145/
3368313.

Hu:2020:TNT

[HXC^{+}20] Ruizhen Hu, Juzhan Xu, Bin Chen, Minglun Gong, Hao Zhang, and Hui Huang. TAP-Net: transport-and-pack using reinforcement learning. ACM Transactions on Graphics, 39(6):232:1– 232:15, November 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417796.

Hu:2020:MBV

[HXFW20] Wenbo Hu, Menghan Xia, Chi-Wing Fu, and Tien-Tsin Wong. Mononizing binocular videos. ACM Transactions on Graphics, 39(6): 228:1–228:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417764.

Hoshyari:2019:VMM

[HXK^{+}19] Shayan Hoshyari, Hongyi Xu, Espen Knoop, Stelian Coros, and Moritz Bächer. Vibration-minimizing motion retargeting for robotic characters. ACM Transactions on Graphics, 38(4):102:1– 102:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2013:IIE

[HXM^{+}13] Shi-Min Hu, Kun Xu, Li-Qian Ma, Bin Liu, Bi-Ye Jiang, and Jue Wang. Inverse image editing: recovering a semantic editing history from a before-and-after image pair. ACM Transactions on Graphics, 32(6): 194:1–194:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2018:AMP

[HXM^{+}18] Hui Huang, Ke Xie, Lin Ma, Dani Lischinski, Minglun Gong, Xin Tong, and Daniel Cohen-Or. Appearance modeling via proxy-to-image alignment. ACM Transactions on Graphics, 37(1):10:1– 10:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Habermann:2019:LRT

[HXZ^{+}19] Marc Habermann, Weipeng Xu, Michael Zollhöfer, Gerard Pons-Moll, and Christian Theobalt. LiveCap: Real-time human performance capture from monocular video. ACM Transactions on Graphics, 38 (2):14:1–14:??, April 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3311970.

He:2020:ISD

[HXZW20] Feixiang He, Yuanhang Xiang, Xi Zhao, and He Wang. Informative scene decomposition for crowd analysis, comparison and simulation guidance. ACM Transactions on Graphics, 39(4): 50:1–50:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392407.

Huang:2013:MGT

[HYG^{+}13] Hui Huang, Kangxue Yin, Minglun Gong, Dani Lischinski, Daniel Cohen-Or, Uri Ascher, and Baoquan Chen. “mind the gap”: teleregistration for structuredriven image completion. ACM Transactions on Graphics, 32(6):174:1–174:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ha:2012:FLM

[HYL12] Sehoon Ha, Yuting Ye, and C. Karen Liu. Falling and landing motion control for character animation. ACM Transactions on Graphics, 31 (6):155:1–155:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harvey:2020:RMB

[HYNP20] Félix G. Harvey, Mike Yurick, Derek Nowrouzezahrai, and Christopher Pal. Robust motion in-betweening. ACM Transactions on Graphics, 39 (4):60:1–60:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392480.

Haydel:2023:LAL

[HYS23] Jacob Haydel, Cem Yuksel, and Larry Seiler. Locallyadaptive level-of-detail for hardware-accelerated ray tracing. ACM Transactions on Graphics, 42(6):196:1– 196:??, December 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618359.

Hu:2018:PGN

[HYZ^{+}18] Ruizhen Hu, Zihao Yan, Jingwen Zhang, Oliver Van Kaick, Ariel Shamir, Hao Zhang, and Hui Huang. Predictive and generative neural networks for object functionality. ACM Transactions on Graphics, 37(4):151:1–151:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hubschman:1982:FFC

[HZ82] H. Hubschman and S. W. Zucker. Frame-to-frame coherence and the hidden surface computation: constraints for a convex world. ACM Transactions on Graphics, 1 (2):129–162, April 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hou:2011:SRM

[HZ11] Qiming Hou and Kun Zhou.

A shading reuse method for efficient micropolygon ray tracing. ACM Transactions on Graphics, 30(6):151:1– 151:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Harmon:2013:SIL

[HZ13] David Harmon and Denis Zorin. Subspace integration with local deformations. ACM Transactions on Graphics, 32 (4):107:1–107:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2022:PCG

[HZC^{+}22] Tianxin Huang, Jiangning Zhang, Jun Chen, Zhonggan Ding, Ying Tai, Zhenyu Zhang, Chengjie Wang, and Yong Liu. 3QNet: 3D point cloud geometry quantization compression network. ACM Transactions on Graphics, 41 (6):187:1–187:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555481.

Huang:2017:TCR

[HZCJ17] Zhiyang Huang, Ming Zou, Nathan Carr, and Tao Ju. Topology-controlled reconstruction of multi-labelled domains from cross-sections. ACM Transactions on Graphics, 36(4):76:1–76:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2023:ALS

[HZD^{+}23] Jingwei Huang, Shanshan Zhang, Bo Duan, Yanfeng Zhang, Xiaoyang Guo, Mingwei Sun, and Li Yi. ArrangementNet: Learning scene arrangements for vectorized indoor scene modeling. ACM Transactions on Graphics, 42 (4):51:1–51:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592122.

Herholz:2019:VPG

[HZE^{+}19] Sebastian Herholz, Yangyang Zhao, Oskar Elek, Derek Nowrouzezahrai, Hendrik P. A. Lensch, and Jaroslav Krivánek. Volume path guiding based on zero-variance random walk theory. ACM Transactions on Graphics, 38(3): 25:1–25:??, June 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3230635.

Hou:2008:BBS

[HZG08] Qiming Hou, Kun Zhou, and Baining Guo. BSGP: bulk-synchronous GPU programming. ACM Transactions on Graphics, 27(3):19:1– 19:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hou:2009:DGS

[HZG09] Qiming Hou, Kun Zhou, and Baining Guo. Debugging GPU stream programs through automatic dataflow recording and visualization. ACM Transactions on Graphics, 28(5): 153:1–153:11, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2012:OAE

[HZG^{+}12] Qi-Xing Huang, Guo-Xin Zhang, Lin Gao, Shi-Min Hu, Adrian Butscher, and Leonidas Guibas. An optimization approach for extracting and encoding consistent maps in a shape collection. ACM Transactions on Graphics, 31(6):167:1– 167:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2018:TMW

[HZG^{+}18] Yixin Hu, Qingnan Zhou, Xifeng Gao, Alec Jacobson, Denis Zorin, and Daniele Panozzo. Tetrahedral meshing in the wild. ACM Transactions on Graphics, 37(4): 60:1–60:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2016:FRF

[HZH^{+}16] Yijiang Huang, Juyong Zhang, Xin Hu, Guoxian Song, Zhongyuan Liu, Lei Yu, and Ligang Liu. Frame-Fab: robotic fabrication of frame shapes. ACM Transactions on Graphics, 35(6): 224:1–224:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2022:UUF

[HZL22] Zhitong Huang, Nanxuan Zhao, and Jing Liao. Uni-Color: a unified framework for multi-modal colorization with transformer. ACM Transactions on Graphics, 41(6): 205:1–205:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555471.

Huang:2008:SQO

[HZM^{+}08] Jin Huang, Muyang Zhang, Jin Ma, Xinguo Liu, Leif Kobbelt, and Hujun Bao. Spectral quadrangulation with orientation and alignment control. ACM Transactions on Graphics, 27(5):147:1– 147:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hong:2022:AZS

[HZP^{+}22] Fangzhou Hong, Mingyuan Zhang, Liang Pan, Zhongang Cai, Lei Yang, and Ziwei Liu. AvatarCLIP: zero-shot textdriven generation and animation of 3D avatars. ACM Transactions on Graphics, 41 (4):161:1–161:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530094.

Hu:2015:ICI

[HZvK^{+}15] Ruizhen Hu, Chenyang Zhu, Oliver van Kaick, Ligang Liu, Ariel Shamir, and Hao Zhang. Interaction context (ICON): towards a geometric functionality descriptor. ACM Transactions on Graphics, 34(4): 83:1–83:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Herrera:2012:LHI

[HZW12] Tomas Lay Herrera, Arno Zinke, and Andreas Weber. Lighting hair from the inside: a thermal approach to hair reconstruction. ACM Transactions on Graphics, 31(6): 146:1–146:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Hu:2013:PPB

[HZW^{+}13] Shi-Min Hu, Fang-Lue Zhang, Miao Wang, Ralph R. Martin, and Jue Wang. PatchNet: a patch-based image representation for interactive librarydriven image editing. ACM Transactions on Graphics, 32 (6):196:1–196:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Huang:2011:ALC

[HZZ11] Hua Huang, Lei Zhang, and Hong-Chao Zhang. Arcimboldolike collage using Internet images. ACM Transactions on Graphics, 30(6):155:1– 155:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ikemoto:2009:GME

[IAF09] Leslie Ikemoto, Okan Arikan, and David Forsyth. Generalizing motion edits with Gaussian processes. ACM Transactions on Graphics, 28(1): 1:1–1:12, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iarussi:2015:BRC

[IBB15] Emmanuel Iarussi, David Bommes, and Adrien Bousseau. BendFields: Regularized curvature fields from rough concept sketches. ACM Transactions on Graphics, 34(3): 24:1–24:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ichim:2015:DAC

[IBP15] Alexandru Eugen Ichim, Sofien Bouaziz, and Mark Pauly. Dynamic 3D avatar creation from hand-held video input. ACM Transactions on Graphics, 34(4):45:1– 45:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iyer:2017:PWC

[ICG17] Vikram Iyer, Justin Chan, and Shyamnath Gollakota. 3D printing wireless connected objects. ACM Transactions on Graphics, 36(6):242:1– 242:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iwasaki:2012:IBS

[IDN12] Kei Iwasaki, Yoshinori Dobashi, and Tomoyuki Nishita. Interactive bi-scale editing of highly glossy materials. ACM Transactions on Graphics, 31 (6):144:1–144:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Isenburg:2003:CCG

[IG03] Martin Isenburg and Stefan Gumhold. Out-of-core compression for gigantic polygon meshes. ACM Transactions on Graphics, 22(3): 935–942, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Irving:2006:ESL

[IGLF06] Geoffrey Irving, Eran Guendelman, Frank Losasso, and Ronald Fedkiw. Efficient simulation of large bodies of water by coupling two and three dimensional techniques. ACM Transactions on Graphics, 25 (3):805–811, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iseringhausen:2017:ITS

[IGP^{+}17] Julian Iseringhausen, Bastian Goldlücke, Nina Pesheva, Stanimir Iliev, Alexander Wender, Martin Fuchs, and Matthias B. Hullin. 4D imaging through spray-on optics. ACM Transactions on Graphics, 36(4):35:1–35:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Igarashi:2003:CM

[IH03] Takeo Igarashi and John F. Hughes. Clothing manipulation. ACM Transactions on Graphics, 22(3):697, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iseringhausen:2020:NLS

[IH20] Julian Iseringhausen and Matthias B. Hullin. Non-line-of-sight reconstruction using efficient transient rendering. ACM Transactions on Graphics, 39(1):8:1– 8:14, February 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3368314.

Igarashi:2012:BIB

[IIM12] Yuki Igarashi, Takeo Igarashi, and Jun Mitani. Beady: interactive beadwork design and construction. ACM Transactions on Graphics, 31(4): 49:1–49:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ilbery:2013:BDC

[IKCM13] Peter Ilbery, Luke Kendall, Cyril Concolato, and Michael McCosker. Biharmonic diffusion curve images from boundary elements. ACM Transactions on Graphics, 32(6): 219:1–219:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ichim:2017:PPB

[IKKP17] Alexandru-Eugen Ichim, Petr Kadlecek, Ladislav Kavan, and Mark Pauly. Phace: physics-based face modeling and animation. ACM Transactions on Graphics, 36(4): 153:1–153:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iarussi:2015:WCA

[ILB15] Emmanuel Iarussi, Wilmot Li, and Adrien Bousseau. WrapIt: computer-assisted crafting of wire wrapped jewelry. ACM Transactions on Graphics, 34 (6):221:1–221:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Isenburg:2006:SCD

[ILSS06] Martin Isenburg, Yuanxin Liu, Jonathan Shewchuk, and Jack Snoeyink. Streaming computation of Delaunay triangulations. ACM Transactions on Graphics, 25(3): 1049–1056, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Igarashi:2010:ALO

[IM10] Takeo Igarashi and Jun Mitani. Apparent layer operations for the manipulation of deformable objects. ACM Transactions on Graphics, 29 (4):110:1–110:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Irawan:2012:SRW

[IM12] Piti Irawan and Steve Marschner.

Specular reflection from woven cloth. ACM Transactions on Graphics, 31(1):11:1– 11:20, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Isik:2021:IMC

[IMF^{+}21] Mustafa Isik, Krishna Mullia, Matthew Fisher, Jonathan Eisenmann, and Michaël Gharbi. Interactive Monte Carlo denoising using affinity of neural features. ACM Transactions on Graphics, 40 (4):37:1–37:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459793.

Igarashi:2005:RPS

[IMH05] Takeo Igarashi, Tomer Moscovich, and John F. Hughes. Asrigid-as-possible shape manipulation. ACM Transactions on Graphics, 24(3): 1134–1141, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ijiri:2005:FDI

[IOOI05] Takashi Ijiri, Shigeru Owada, Makoto Okabe, and Takeo Igarashi. Floral diagrams and inflorescences: interactive flower modeling using botanical structural constraints.

ACM Transactions on Graphics, 24(3):720–726, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Isik:2023:HHF

[IRG^{+}23] Mustafa Işık, Martin Rünz, Markos Georgopoulos, Taras Khakhulin, Jonathan Starck, Lourdes Agapito, and Matthias Nießner. HumanRF: Highfidelity neural radiance fields for humans in motion. ACM Transactions on Graphics, 42(4):160:1–160:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592415.

Ion:2020:SAD

[IRHSH20] Alexandra Ion, Michael Rabinovich, Philipp Herholz, and Olga Sorkine-Hornung. Shape approximation by developable wrapping. ACM Transactions on Graphics, 39(6): 200:1–200:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417835.

Iser:2022:ASM

[IRN^{+}22] Tomáš Iser, Tobias Rittig, Emilie Nogué, Thomas Klaus Nindel, and Alexander Wilkie. Affordable spectral measurements of translucent materials. ACM Transactions on Graphics, 41(6):

199:1–199:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555499.

Inza:2023:DQM

[IRWP23] Victor Ceballos Inza, Florian Rist, Johannes Wallner, and Helmut Pottmann. Developable quad meshes and contact element nets. ACM Transactions on Graphics, 42 (6):183:1–183:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618355.

Ivrissimtzis:2004:SRS

[ISD04] Ioannis P. Ivrissimtzis, Malcolm A. Sabin, and Neil A. Dodgson. On the support of recursive subdivision. ACM Transactions on Graphics, 23 (4):1043–1060, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Irving:2007:VCF

[ISF07] Geoffrey Irving, Craig Schroeder, and Ronald Fedkiw. Volume conserving finite element simulations of deformable models. ACM Transactions on Graphics, 26(3):13:1–13:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ishida:2020:MSF

[ISN^{+}20] Sadashige Ishida, Peter Synak, Fumiya Narita, Toshiya Hachisuka, and Chris Wojtan. A model for soap film dynamics with evolving thickness. ACM Transactions on Graphics, 39 (4):31:1–31:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392405.

Iizuka:2016:LTC

[ISSI16] Satoshi Iizuka, Edgar Simo-Serra, and Hiroshi Ishikawa. Let there be color!: joint endto-end learning of global and local image priors for automatic image colorization with simultaneous classification. ACM Transactions on Graphics, 35(4):110:1–110:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iizuka:2017:GLC

[ISSI17] Satoshi Iizuka, Edgar Simo-Serra, and Hiroshi Ishikawa. Globally and locally consistent image completion. ACM Transactions on Graphics, 36 (4):107:1–107:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ito:2014:CEP

[ITM^{+}14] Atsushi Ito, Salil Tambe, Kaushik Mitra, Aswin C. Sankaranarayanan, and Ashok Veeraraghavan. Compressive epsilon photography for post-capture control in digital imaging. ACM Transactions on Graphics, 33(4): 88:1–88:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Iluz:2023:WIS

[IVH^{+}23] Shir Iluz, Yael Vinker, Amir Hertz, Daniel Berio, Daniel Cohen-Or, and Ariel Shamir. Word-as-image for semantic typography. ACM Transactions on Graphics, 42(4): 151:1–151:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592123.

Ishida:2022:HDF

[IWC22] Sadashige Ishida, Chris Wojtan, and Albert Chern. Hidden degrees of freedom in implicit vortex filaments. ACM Transactions on Graphics, 41 (6):241:1–241:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555459.

Iseringhausen:2020:CPF

[IWHH20] Julian Iseringhausen, Michael Weinmann, Weizhen Huang, and Matthias B. Hullin. Computational parquetry: Fabricated style transfer with wood pixels. ACM Transactions on Graphics, 39(2): 12:1–12:14, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3378541.

Ishigaki:2009:PBC

[IWZL09] Satoru Ishigaki, Timothy White, Victor B. Zordan, and C. Karen Liu. Performancebased control interface for character animation. ACM Transactions on Graphics, 28 (3):61:1–61:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ishida:2017:HGF

[IYAH17] Sadashige Ishida, Masafumi Yamamoto, Ryoichi Ando, and Toshiya Hachisuka. A hyperbolic geometric flow for evolving films and foams. ACM Transactions on Graphics, 36(6):199:1–199:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ijiri:2014:FMX

[IYYI14] Takashi Ijiri, Shin Yoshizawa, Hideo Yokota, and Takeo Igarashi. Flower modeling via X-ray computed tomography. ACM Transactions on Graphics, 33(4):48:1–48:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Izadi:2018:SDM

[Iza18] Shahram Izadi. Session details: Modeling things on (and in) your head. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ishiwaka:2021:FBI

[IZE^{+}21] Yuko Ishiwaka, Xiao S.

Zeng, Michael Lee Eastman, Sho Kakazu, Sarah Gross, Ryosuke Mizutani, and Masaki Nakada. Foids: bioinspired fish simulation for generating synthetic datasets. ACM Transactions on Graphics, 40(6):207:1–207:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480520.

Ihrke:2007:ERE

[IZT^{+}07] Ivo Ihrke, Gernot Ziegler, Art Tevs, Christian Theobalt, Marcus Magnor, and Hans-Peter Seidel. Eikonal rendering: efficient light transport in refractive objects. ACM Transactions on Graphics, 26 (3):59:1–59:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacob:1986:SLD

[Jac86] Robert J. K. Jacob. A specification language for directmanipulation user interfaces. ACM Transactions on Graphics, 5(4):283–317, October 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
27624.html.

Jarabo:2018:RTF

[JAG18] Adrian Jarabo, Carlos Aliaga, and Diego Gutierrez. A radiative transfer framework for spatially-correlated materials. ACM Transactions on Graphics, 37(4):83:1–83:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jakob:2010:RTF

[JAM^{+}10] Wenzel Jakob, Adam Arbree, Jonathan T. Moon, Kavita Bala, and Steve Marschner. A radiative transfer framework for rendering materials with anisotropic structure. ACM Transactions on Graphics, 29 (4):53:1–53:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

James:2020:PDB

[Jam20] Doug L. James. Phong deformation: a better C_{0} interpolant for embedded deformation. ACM Transactions on Graphics, 39(4): 56:1–56:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392371.

Jansen:1991:DOP

[Jan91] Frederik W. Jansen. Depthorder point classification techniques for CSG display algorithms. ACM Transactions on Graphics, 10(1):40– 70, January 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
99904.html.

Joan-Arinyo:1999:CCE

[JASR99] R. Joan-Arinyo and A. Soto-Riera. Combining constructive and equational geometric constraint-solving techniques. ACM Transactions on Graphics, 18 (1):35–55, January 1999.

URL http://www.acm.org:
80/pubs/citations/journals/
tog/1999-18-1/p35-joan-
arinyo/.

Jensen:2002:RHR

[JB02] Henrik Wann Jensen and Juan Buhler. A rapid hierarchical rendering technique for translucent materials. ACM Transactions on Graphics, 21 (3):576–581, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacobson:2012:FAS

[JBK^{+}12] Alec Jacobson, Ilya Baran, Ladislav Kavan, Jovan Popović, and Olga Sorkine. Fast automatic skinning transformations. ACM Transactions on Graphics, 31(4): 77:1–77:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jang:2018:HNE

[JBLL18] Changwon Jang, Kiseung Bang, Gang Li, and Byoungho Lee. Holographic near-eye display with expanded eye-box. ACM Transactions on Graphics, 37(6): 195:1–195:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jang:2017:RAR

[JBM^{+}17] Changwon Jang, Kiseung Bang, Seokil Moon, Jonghyun Kim, Seungjae Lee, and Byoungho Lee. Retinal 3D: augmented reality neareye display via pupil-tracked light field projection on retina. ACM Transactions on Graphics, 36(6):190:1– 190:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

James:2006:PAT

[JBP06] Doug L. James, Jernej Barbič, and Dinesh K. Pai. Precomputed acoustic transfer: output-sensitive, accurate sound generation for geometrically complex vibration sources. ACM Transactions on Graphics, 25(3): 987–995, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacobson:2011:BBW

[JBPS11] Alec Jacobson, Ilya Baran, Jovan Popović, and Olga Sorkine. Bounded biharmonic weights for real-time deformation. ACM Transactions on Graphics, 30(4): 78:1–78:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2020:SLG

[JBX^{+}20] R. Kenny Jones, Theresa Barton, Xianghao Xu, Kai Wang, Ellen Jiang, Paul Guerrero, Niloy J. Mitra, and Daniel Ritchie. ShapeAssembly: learning to generate programs for 3D shape structure synthesis. ACM Transactions on Graphics, 39(6): 234:1–234:20, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417812.

Jeon:2019:CSH

[JBY^{+}19] Daniel S. Jeon, Seung-Hwan Baek, Shinyoung Yi, Qiang Fu, Xiong Dun, Wolfgang Heidrich, and Min H. Kim. Compact snapshot hyperspectral imaging with diffracted rotation. ACM Transactions on Graphics, 38(4): 117:1–117:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2023:NID

[JCFG23] Kaiwen Jiang, Shu-Yu Chen, Hongbo Fu, and Lin Gao. NeRFFaceLighting: Implicit and disentangled face lighting representation leveraging generative prior in neural radiance fields. ACM Transactions on Graphics, 42 (3):35:1–35:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3597300.

Jones:2021:SMO

[JCG^{+}21] R. Kenny Jones, David Charatan, Paul Guerrero, Niloy J. Mitra, and Daniel Ritchie. ShapeMOD: macro operation discovery for 3D shape programs. ACM Transactions on Graphics, 40(4): 153:1–153:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459821.

Ju:2010:MC

[JCP^{+}10] Eunjung Ju, Myung Geol Choi, Minji Park, Jehee Lee, Kang Hoon Lee, and Shigeo Takahashi. Morphable crowds. ACM Transactions on Graphics, 29(6):140:1– 140:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Johnson:2011:MCU

[JCRA11] Micah K. Johnson, Forrester Cole, Alvin Raj, and Edward H. Adelson. Microgeometry capture using an elastomeric sensor. ACM Transactions on Graphics, 30(4): 46:1–46:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeschke:2009:GLS

[JCW09a] Stefan Jeschke, David Cline, and Peter Wonka. A GPU Laplacian solver for diffusion curves and Poisson image editing. ACM Transactions on Graphics, 28(5): 116:1–116:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeschke:2009:RSD

[JCW09b] Stefan Jeschke, David Cline, and Peter Wonka. Rendering surface details with diffusion curves. ACM Transactions on Graphics, 28(5): 117:1–117:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2021:CKS

[JCW^{+}21] Hongda Jiang, Marc Christie, Xi Wang, Libin Liu, Bin Wang, and Baoquan Chen. Camera keyframing with style and control. ACM Transactions on Graphics, 40(6): 209:1–209:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480533.

Jia:2023:SCR

[JCY23] Xiaohong Jia, Falai Chen, and Shanshan Yao. Singularity computation for rational parametric surfaces using moving planes. ACM Transactions on Graphics, 42 (1):12:1–12:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3551387.

Judd:2007:ARL

[JDA07] Tilke Judd, Frédo Durand, and Edward Adelson. Apparent ridges for line drawing. ACM Transactions on Graphics, 26(3):19:1–19:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2003:NIF

[JDD03] Thouis R. Jones, Frédo Durand, and Mathieu Desbrun. Non-iterative, featurepreserving mesh smoothing. ACM Transactions on Graphics, 22(3):943–949, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2022:DSU

[JDH^{+}22] Zhongshi Jiang, Jiacheng Dai, Yixin Hu, Yunfan Zhou, Jeremie Dumas, Qingnan Zhou, Gurkirat Singh Bajwa, Denis Zorin, Daniele Panozzo, and Teseo Schneider. Declarative specification for unstructured mesh editing algorithms. ACM Transactions on Graphics, 41(6): 251:1–251:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555513.

Jakob:2014:CFR

[JdJM14] Wenzel Jakob, Eugene d’Eon, Otto Jakob, and Steve Marschner. A comprehensive framework for rendering layered materials. ACM Transactions on Graphics, 33(4): 118:1–118:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jagnow:2004:STS

[JDR04] Robert Jagnow, Julie Dorsey, and Holly Rushmeier. Stereological techniques for solid textures. ACM Transactions on Graphics, 23(3):329– 335, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jarosz:2008:RCP

[JDZJ08] Wojciech Jarosz, Craig Donner, Matthias Zwicker, and Henrik Wann Jensen. Radiance caching for participating media. ACM Transactions on Graphics, 27(1): 7:1–7:11, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

James:2003:PID

[JF03] Doug L. James and Kayvon Fatahalian. Precomputing interactive dynamic deformable scenes. ACM Transactions on Graphics, 22(3):879–887, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jamriska:2015:LAT

[JFA^{+}15] Ondrej Jamriska, Jakub Fiser, Paul Asente, Jingwan Lu, Eli Shechtman, and Daniel Sýkora. LazyFluids: appearance transfer for fluid animations. ACM Transactions on Graphics, 34(4):92:1– 92:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2015:FFG

[JFH^{+}15] Tengfei Jiang, Xianzhong Fang, Jin Huang, Hujun Bao, Yiying Tong, and Mathieu Desbrun. Frame field generation through metric customization. ACM Transactions on Graphics, 34(4):40:1– 40:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacobs:2015:SVE

[JGC^{+}15] David E. Jacobs, Orazio Gallo, Emily A. Cooper, Kari Pulli, and Marc Levoy. Simulating the visual experience of very bright and very dark scenes. ACM Transactions on Graphics, 34(3): 25:1–25:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jin:2015:AIA

[JGGN15] Ming Jin, Dan Gopstein, Yotam Gingold, and Andrew Nealen. AniMesh: interleaved animation, modeling, and editing. ACM Transactions on Graphics, 34(6): 207:1–207:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2023:SDA

[JGMR23] R. Kenny Jones, Paul Guerrero, Niloy J. Mitra, and Daniel Ritchie. ShapeCoder:

Discovering abstractions for visual programs from unstructured primitives. ACM Transactions on Graphics, 42 (4):49:1–49:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592416.

Jo:2016:DDC

[JGN16] Kensei Jo, Mohit Gupta, and Shree K. Nayar. DisCo: Display-camera communication using rolling shutter sensors. ACM Transactions on Graphics, 35(5):150:1– 150:??, September 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2017:AEC

[JGT17] Chenfanfu Jiang, Theodore Gast, and Joseph Teran. Anisotropic elastoplasticity for cloth, knit and hair frictional contact. ACM Transactions on Graphics, 36(4): 152:1–152:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2021:ADL

[JHC^{+}21] Benjamin Jones, Dalton Hildreth, Duowen Chen, Ilya Baran, Vladimir G. Kim, and Adriana Schulz. AutoMate: a dataset and learning approach for automatic mating of CAD assemblies. ACM Transactions on Graphics, 40(6):

227:1–227:18, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480562.

Jung:2015:SFD

[JHR^{+}15] Amaury Jung, Stefanie Hahmann, Damien Rohmer, Antoine Begault, Laurence Boissieux, and Marie-Paule Cani. Sketching folds: Developable surfaces from nonplanar silhouettes. ACM Transactions on Graphics, 34(5):155:1–155:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2022:SSR

[JHR22] R. Kenny Jones, Aalia Habib, and Daniel Ritchie. SHRED: 3D shape region decomposition with learned local operations. ACM Transactions on Graphics, 41(6): 186:1–186:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555440.

Jorg:2012:DDF

[JHS12] Sophie Jörg, Jessica Hodgins, and Alla Safonova. Datadriven finger motion synthesis for gesturing characters. ACM Transactions on Graphics, 31 (6):189:1–189:??, November 2012. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Jourdan:2023:SMP

[JHS^{+}23] David Jourdan, Pierre-Alexandre Hugron, Camille Schreck, Jonàs Martínez, and Sylvain Lefebvre. Shrink & morph: 3D-printed self-shaping shells actuated by a shape memory effect. ACM Transactions on Graphics, 42(6): 187:1–187:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618386.

Jakob:2014:DSM

[JHY^{+}14] Wenzel Jakob, Milos Hasan, Ling-Qi Yan, Jason Lawrence, Ravi Ramamoorthi, and Steve Marschner. Discrete stochastic microfacet models. ACM Transactions on Graphics, 33 (4):115:1–115:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jia:2021:SSA

[Jia21] Kai Jia. SANM: a symbolic asymptotic numerical solver with applications in mesh deformation. ACM Transactions on Graphics, 40 (4):79:1–79:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459755.

Jang:2021:SCG

[JJJ^{+}21] Wonjong Jang, Gwangjin Ju, Yucheol Jung, Jiaolong Yang, Xin Tong, and Seungyong Lee. StyleCari-GAN: caricature generation via StyleGAN feature map modulation. ACM Transactions on Graphics, 40(4): 116:1–116:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459860.

Jazar:2023:TSI

[JK23] Kavosh Jazar and Paul G.

Kry. Temporal set inversion for animated implicits. ACM Transactions on Graphics, 42(4):134:1–134:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592448.

Jeon:2022:DOS

[JKH^{+}22] Sang-Bin Jeon, Soon-Uk Kwon, June-Young Hwang, Yong-Hun Cho, Hayeon Kim, Jinhyung Park, and In-Kwon Lee. Dynamic optimal space partitioning for redirected walking in multiuser environment. ACM Transactions on Graphics, 41(4):90:1–90:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530113.

Jacobson:2013:RIO

[JKSH13] Alec Jacobson, Ladislav Kavan, and Olga Sorkine-Hornung. Robust insideoutside segmentation using generalized winding numbers. ACM Transactions on Graphics, 32(4):33:1–33:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Joshi:2015:RTH

[JKT^{+}15] Neel Joshi, Wolf Kienzle, Mike Toelle, Matt Uyttendaele, and Michael F. Cohen. Real-time hyperlapse creation via optimal frame selection. ACM Transactions on Graphics, 34 (4):63:1–63:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Joshi:2010:IDU

[JKZS10] Neel Joshi, Sing Bing Kang, C. Lawrence Zitnick, and Richard Szeliski. Image deblurring using inertial measurement sensors. ACM Transactions on Graphics, 29 (4):30:1–30:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jain:2011:CPB

[JL11a] Sumit Jain and C. Karen Liu. Controlling physicsbased characters using soft contacts. ACM Transactions on Graphics, 30(6):163:1– 163:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jain:2011:MSC

[JL11b] Sumit Jain and C. Karen Liu.

Modal-space control for articulated characters. ACM Transactions on Graphics, 30(5):118:1–118:12, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Johnson:2005:IZB

[JLBM05] Gregory S. Johnson, Juhyun Lee, Christopher A. Burns, and William R. Mark. The irregular Z-buffer: Hardware acceleration for irregular data structures. ACM Transactions on Graphics, 24 (4):1462–1482, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2009:AEC

[JLF^{+}09] Andrew Jones, Magnus Lang, Graham Fyffe, Xueming Yu, Jay Busch, Ian McDowall, Mark Bolas, and Paul Debevec. Achieving eye contact in a one-to-many 3D video teleconferencing system. ACM Transactions on Graphics, 28 (3):64:1–64:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2023:LLI

[JLF^{+}23] Hai Jiang, Ao Luo, Haoqiang Fan, Songchen Han, and Shuaicheng Liu. Lowlight image enhancement with wavelet-based diffusion models. ACM Transactions on Graphics, 42(6): 238:1–238:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618373.

Jacobs:2003:AGB

[JLS^{+}03] Charles Jacobs, Wilmot Li, Evan Schrier, David Bargeron, and David Salesin. Adaptive grid-based document layout. ACM Transactions on Graphics, 22(3): 838–847, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ju:2002:DCH

[JLSW02] Tao Ju, Frank Losasso, Scott Schaefer, and Joe Warren. Dual contouring of Hermite data. ACM Transactions on Graphics, 21(3):339–346, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jin:2022:NLB

[JLWM22] Xutong Jin, Sheng Li, Guoping Wang, and Dinesh Manocha. NeuralSound: learning-based modal sound synthesis with acoustic transfer. ACM Transactions on Graphics, 41 (4):121:1–121:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530184.

Jakob:2012:MEM

[JM12] Wenzel Jakob and Steve Marschner. Manifold exploration: a Markov Chain Monte Carlo technique for rendering scenes with difficult specular transport. ACM Transactions on Graphics, 31(4):58:1–58:13, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Joshi:2006:NVM

[JMA06] Neel Joshi, Wojciech Matusik, and Shai Avidan. Natural video matting using camera arrays. ACM Transactions on Graphics, 25(3):779–786, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Joshi:2010:PPE

[JMAK10] Neel Joshi, Wojciech Matusik, Edward H. Adelson, and David J. Kriegman. Personal photo enhancement using example images. ACM Transactions on Graphics, 29 (2):12:1–12:15, March 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jarabo:2014:HDP

[JMB^{+}14] Adrian Jarabo, Belen Masia, Adrien Bousseau, Fabio Pellacini, and Diego Gutierrez. How do people edit light fields? ACM Transactions on Graphics, 33(4): 146:1–146:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jang:2020:DFF

[JMB^{+}20] Changwon Jang, Olivier Mercier, Kiseung Bang, Gang Li, Yang Zhao, and Douglas Lanman. Design and fabrication of freeform holographic optical elements. ACM Transactions on Graphics, 39(6): 184:1–184:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417762.

Joshi:2007:HCC

[JMD^{+}07] Pushkar Joshi, Mark Meyer, Tony DeRose, Brian Green, and Tom Sanocki. Harmonic coordinates for character articulation. ACM Transactions on Graphics, 26(3): 71:1–71:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jin:2017:VTB

[JMD^{+}17] Zeyu Jin, Gautham J. Mysore, Stephen Diverdi, Jingwan Lu, and Adam Finkelstein. VoCo: text-based insertion and replacement in audio narration. ACM Transactions on Graphics, 36(4):96:1–96:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jang:2022:ESR

[JMK^{+}22] Hyeonjoong Jang, Andréas Meuleman, Dahyun Kang, Donggun Kim, Christian Richardt, and Min H. Kim. Egocentric scene reconstruction from an omnidirectional video. ACM Transactions on Graphics, 41(4):100:1– 100:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530074.

Jakob:2009:CHA

[JMM09] Wenzel Jakob, Jonathan T.

Moon, and Steve Marschner. Capturing hair assemblies fiber by fiber. ACM Transactions on Graphics, 28(5): 164:1–164:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jarabo:2014:FTR

[JMM^{+}14] Adrian Jarabo, Julio Marco, Adolfo Muñoz, Raul Buisan, Wojciech Jarosz, and Diego Gutierrez. A framework for transient rendering. ACM Transactions on Graphics, 33 (6):177:1–177:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2007:RIL

[JMY^{+}07] Andrew Jones, Ian McDowall, Hideshi Yamada, Mark Bolas, and Paul Debevec. Rendering for an interactive 360^{∘} light field display. ACM Transactions on Graphics, 26(3): 40:1–40:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2022:CDK

[JMZ^{+}22] Benjamin Jones, Yuxuan Mei, Haisen Zhao, Taylor Gotfrid, Jennifer Mankoff, and Adriana Schulz. Computational design of knit templates. ACM Transactions on Graphics, 41 (2):16:1–16:16, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3488006.

Jones:2023:BRM

[JNK^{+}23] Benjamin Jones, James Noeckel, Milin Kodnongbua, Ilya Baran, and Adriana Schulz. B-rep matching for collaborating across CAD systems. ACM Transactions on Graphics, 42(4):104:1– 104:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592125.

Jarosz:2011:CTV

[JNSJ11] Wojciech Jarosz, Derek Nowrouzezahrai, Iman Sadeghi, and Henrik Wann Jensen. A comprehensive theory of volumetric radiance estimation using photon points and beams. ACM Transactions on Graphics, 30(1):5:1–5:19, January 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jarosz:2011:PPB

[JNT^{+}11] Wojciech Jarosz, Derek Nowrouzezahrai, Robert Thomas, Peter-Pike Sloan, and Matthias Zwicker. Progressive photon beams. ACM Transactions on Graphics, 30(6):181:1–181:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Joe:1989:MKR

[Joe89] Barry Joe. Multiple-knot and rational cubic betasplines. ACM Transactions on Graphics, 8(2):100–120, April 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
62055.html.

Joe:1990:KIB

[Joe90a] Barry Joe. Knot insertion for beta-spline curves and surfaces. ACM Transactions on Graphics, 9(1):41–65, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77638.html.

Joe:1990:QBS

[Joe90b] Barry Joe. Quartic betasplines. ACM Transactions on Graphics, 9(3):301–337, July 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
78968.html.

James:2002:DDR

[JP02] Doug L. James and Dinesh K. Pai. DyRT: dynamic response textures for real time deformation simulation with graphics hardware. ACM Transactions on Graphics, 21(3): 582–585, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

James:2003:MGF

[JP03] Doug L. James and Dinesh K. Pai. Multiresolution Green’s function methods for interactive simulation of large-scale elastostatic objects. ACM Transactions on Graphics, 22 (1):47–82, January 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

James:2004:BTO

[JP04] Doug L. James and Dinesh K. Pai. BD-tree: outputsensitive collision detection for reduced deformable models. ACM Transactions on Graphics, 23(3):393–398, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacobson:2014:TMI

[JPG^{+}14] Alec Jacobson, Daniele Panozzo, Oliver Glauser, Cédric Pradalier, Otmar Hilliges, and Olga Sorkine-Hornung. Tangible and modular input device for character articulation. ACM Transactions on Graphics, 33 (4):82:1–82:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jang:2022:MPA

[JPL22] Deok-Kyeong Jang, Soomin Park, and Sung-Hee Lee. Motion puzzle: Arbitrary motion style transfer by body part. ACM Transactions on Graphics, 41(3):33:1–33:16, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3516429.

Jiang:2020:FQB

[JRPW20] Caigui Jiang, Florian Rist, Helmut Pottmann, and Johannes Wallner. Freeform quad-based kirigami. ACM Transactions on Graphics, 39 (6):209:1–209:11, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417844.

Jaros:2021:GAP

[JRSS21] Milan Jaros, Lubomír Ríha, Petr Strakos, and Matej Spetko. GPU accelerated path tracing of massive scenes. ACM Transactions on Graphics, 40(2):16:1–16:17, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3447807.

Joubert:2015:ITD

[JRT^{+}15] Niels Joubert, Mike Roberts, Anh Truong, Floraine Berthouzoz, and Pat Hanrahan. An interactive tool for designing quadrotor camera shots. ACM Transactions on Graphics, 34 (6):238:1–238:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jacobson:2011:STB

[JS11] Alec Jacobson and Olga Sorkine. Stretchable and twistable bones for skeletal shape deformation. ACM Transactions on Graphics, 30 (6):165:1–165:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jimenez:2010:PAM

[JSB^{+}10] Jorge Jimenez, Timothy Scully, Nuno Barbosa, Craig Donner, Xenxo Alvarez, Teresa Vieira, Paul Matts, Verónica Orvalho, Diego Gutierrez, and Tim Weyrich. A practical appearance model for dynamic facial color. ACM Transactions on Graphics, 29 (6):141:1–141:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jarosz:2012:TAA

[JSKJ12] Wojciech Jarosz, Volker Schönefeld, Leif Kobbelt, and Henrik Wann Jensen. Theory, analysis and applications of 2D global illumination. ACM Transactions on Graphics, 31(5):125:1–125:21, August 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeschke:2018:WSW

[JSMF^{+}18] Stefan Jeschke, Tomás Skrivan, Matthias Müller-Fischer, Nuttapong Chentanez, Miles Macklin, and Chris Wojtan. Water surface wavelets. ACM Transactions on Graphics, 37 (4):94:1–94:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jain:2012:TDP

[JSMH12] Eakta Jain, Yaser Sheikh, Moshe Mahler, and Jessica Hodgins. Three-dimensional proxies for hand-drawn characters. ACM Transactions on Graphics, 31(1):8:1– 8:16, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2017:SCA

[JSP17] Zhongshi Jiang, Scott Schaefer, and Daniele Panozzo. Simplicial complex augmentation framework for bijective maps. ACM Transactions on Graphics, 36(6): 186:1–186:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jakob:2022:DJJ

[JSRV22] Wenzel Jakob, Sébastien Speierer, Nicolas Roussel, and Delio Vicini. DR.JIT: a justin-time compiler for differentiable rendering. ACM Transactions on Graphics, 41 (4):124:1–124:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530099.

Jiang:2015:APC

[JSS^{+}15] Chenfanfu Jiang, Craig Schroeder, Andrew Selle, Joseph Teran, and Alexey Stomakhin. The affine particle-in-cell method. ACM Transactions on Graphics, 34(4):51:1–51:??, August 2015. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Jain:2015:GDV

[JSSH15] Eakta Jain, Yaser Sheikh, Ariel Shamir, and Jessica Hodgins. Gaze-driven video re-editing. ACM Transactions on Graphics, 34(2):21:1– 21:??, February 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jamriska:2019:SVE

[JST^{+}19] Ondrej Jamriska, Sárka Sochorová, Ondrej Texler, Michal Lukác, Jakub Fiser, Jingwan Lu, Eli Shechtman, and Daniel Sýkora. Stylizing video by example. ACM Transactions on Graphics, 38(4): 107:1–107:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jia:2006:DDP

[JSTS06] Jiaya Jia, Jian Sun, Chi-Keung Tang, and Heung-Yeung Shum. Drag-and-drop pasting. ACM Transactions on Graphics, 25(3): 631–637, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ju:2005:MVC

[JSW05] Tao Ju, Scott Schaefer, and Joe Warren. Mean value coordinates for closed triangular meshes. ACM Transactions on Graphics, 24(3): 561–566, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2020:BPS

[JSZP20] Zhongshi Jiang, Teseo Schneider, Denis Zorin, and Daniele Panozzo. Bijective projection in a shell. ACM Transactions on Graphics, 39(6): 247:1–247:18, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417769.

James:2005:SMA

[JT05] Doug L. James and Christopher D. Twigg. Skinning mesh animations. ACM Transactions on Graphics, 24(3): 399–407, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2009:SAM

[JTC09] Nianjuan Jiang, Ping Tan, and Loong-Fah Cheong. Symmetric architecture modeling with a single image. ACM Transactions on Graphics, 28(5):113:1–113:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

James:2007:MEM

[JTCW07] Doug L. James, Christopher D. Twigg, Andrew Cove, and Robert Y. Wang. Mesh Ensemble Motion Graphs: Data-driven mesh animation with constraints. ACM Transactions on Graphics, 26(4): 17:1–17:16, October 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Je:2012:PRT

[JTL^{+}12] Changsoo Je, Min Tang, Youngeun Lee, Minkyoung Lee, and Young J. Kim. Poly-Depth: Real-time penetration depth computation using iterative contact-space projection. ACM Transactions on Graphics, 31(1):5:1– 5:14, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jia:2020:CCE

[JTMW20] Xiaohong Jia, Changhe Tu, Bernard Mourrain, and Wenping Wang. Complete classification and efficient determination of arrangements formed by two ellipsoids. ACM Transactions on Graphics, 39(3): 27:1–27:12, June 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3388540.

Jakob:2015:IFA

[JTPSH15] Wenzel Jakob, Marco Tarini, Daniele Panozzo, and Olga Sorkine-Hornung. Instant field-aligned meshes. ACM Transactions on Graphics, 34 (6):189:1–189:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jain:2012:MMA

[JTRS12] Arjun Jain, Thorsten Thormählen, Tobias Ritschel, and Hans-Peter Seidel. Material memex: automatic material suggestions for 3D objects. ACM Transactions on Graphics, 31 (6):143:1–143:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jones:2016:EBP

[JTSB16] Ben Jones, Nils Thuerey, Tamar Shinar, and Adam W. Bargteil. Example-based plastic deformation of rigid bodies. ACM Transactions on Graphics, 35(4): 34:1–34:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jain:2010:MTR

[JTST10] Arjun Jain, Thorsten Thormählen, Hans-Peter Seidel, and Christian Theobalt. MovieReshape: tracking and reshaping of humans in videos. ACM Transactions on Graphics, 29(6): 148:1–148:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2017:DVO

[JTSW17] Caigui Jiang, Chengcheng Tang, Hans-Peter Seidel, and Peter Wonka. Design and volume optimization of space structures. ACM Transactions on Graphics, 36(4): 159:1–159:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2015:PP

[JTV^{+}15] Caigui Jiang, Chengcheng Tang, Amir Vaxman, Peter Wonka, and Helmut Pottmann. Polyhedral patterns. ACM Transactions on Graphics, 34(6):172:1– 172:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ju:2004:RRP

[Ju04] Tao Ju. Robust repair of polygonal models. ACM Transactions on Graphics, 23 (3):888–895, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeschke:2015:WWA

[JW15] Stefan Jeschke and Chris Wojtan. Water wave animation via wavefront parameter interpolation. ACM Transactions on Graphics, 34(3): 27:1–27:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeschke:2017:WWP

[JW17] Stefan Jeschke and Chris Wojtan. Water wave packets. ACM Transactions on Graphics, 36(4):103:1–103:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeschke:2023:GSW

[JW23] Stefan Jeschke and Chris Wojtan. Generalizing shallow water simulations with dispersive surface waves. ACM Transactions on Graphics, 42 (4):83:1–83:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592098.

Jiang:2023:CKD

[JWD^{+}23] Haiyan Jiang, Dongdong Weng, Xiaonuo Dongye, Le Luo, and Zhenliang Zhang. Commonsense knowledgedriven joint reasoning approach for object retrieval in virtual reality. ACM Transactions on Graphics, 42(6): 198:1–198:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618320.

Jiang:2019:SBR

[JWDL19] Yifeng Jiang, Tom Van Wouwe, Friedl De Groote, and C. Karen Liu. Synthesis of biologically realistic human motion using joint torque actuation. ACM Transactions on Graphics, 38(4): 72:1–72:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2021:UIC

[JWI^{+}21] Caigui Jiang, Hui Wang, Victor Ceballos Inza, Felix Dellinger, Florian Rist, Johannes Wallner, and Helmut Pottmann. Using isometries for computational design and fabrication. ACM Transactions on Graphics, 40 (4):42:1–42:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459839.

Jones:2014:DEP

[JWJ^{+}14] Ben Jones, Stephen Ward, Ashok Jallepalli, Joseph Perenia, and Adam W. Bargteil. Deformation embedding for point-based elastoplastic simulation. ACM Transactions on Graphics, 33(2):21:1– 21:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ju:2013:DDC

[JWL^{+}13] Eunjung Ju, Jungdam Won, Jehee Lee, Byungkuk Choi, Junyong Noh, and Min Gyu Choi. Data-driven control of flapping flight. ACM Transactions on Graphics, 32 (5):151:1–151:12, September 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jeske:2024:IST

[JWL^{+}24] Stefan Rhys Jeske, Lukas Westhofen, Fabian Löschner, José Antonio Fernándezfernández, and Jan Bender. Implicit surface tension for SPH fluid simulation. ACM Transactions on Graphics, 43 (1):13:1–13:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3631936.

Jiang:2023:PPP

[JWT^{+}23] Caigui Jiang, Cheng Wang, Xavier Tellier, Johannes Wallner, and Helmut Pottmann. Planar panels and planar supporting beams in architectural structures. ACM Transactions on Graphics, 42 (2):19:1–19:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3561050.

Jiang:2020:EDV

[JWW^{+}20] Hongda Jiang, Bin Wang, Xi Wang, Marc Christie, and Baoquan Chen. Exampledriven virtual cinematography by learning camera behaviors. ACM Transactions on Graphics, 39(4):

45:1–45:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392427.

Jeng:1996:MCP

[JX96] Elvis Ko-Yung Jeng and Zhigang Xiang. Moving cursor plane for interactive sculpting. ACM Transactions on Graphics, 15(3):211–222, July 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
230523.html.

Jain:2009:OBI

[JYL09] Sumit Jain, Yuting Ye, and C. Karen Liu. Optimizationbased interactive motion synthesis. ACM Transactions on Graphics, 28(1):10:1–10:12, January 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2022:TTD

[JYQ^{+}22] Yuming Jiang, Shuai Yang, Haonan Qiu, Wayne Wu, Chen Change Loy, and Ziwei Liu. Text2Human: textdriven controllable human image generation. ACM Transactions on Graphics, 41 (4):162:1–162:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530104.

Jiang:2023:NII

[JYW^{+}23] Weiwei Jiang, Difeng Yu, Chaofan Wang, Zhanna Sarsenbayeva, Niels van Berkel, Jorge Goncalves, and Vassilis Kostakos. Near-infrared imaging for information embedding and extraction with layered structures. ACM Transactions on Graphics, 42 (1):4:1–4:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3533426.

Ju:2007:ETM

[JZH07] Tao Ju, Qian-Yi Zhou, and Shi-Min Hu. Editing the topology of 3D models by sketching. ACM Transactions on Graphics, 26(3): 42:1–42:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2021:BCH

[JZH^{+}21] Zhongshi Jiang, Ziyi Zhang, Yixin Hu, Teseo Schneider, Denis Zorin, and Daniele Panozzo. Bijective and coarse high-order tetrahedral meshes. ACM Transactions on Graphics, 40(4): 157:1–157:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459840.

Ju:2008:RST

[JZvdP^{+}08] Tao Ju, Qian-Yi Zhou, Michiel van de Panne, Daniel Cohen-Or, and Ulrich Neumann. Reusable skinning templates using cage-based deformations. ACM Transactions on Graphics, 27(5): 122:1–122:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Jiang:2015:BNS

[JZW^{+}15] Min Jiang, Yahan Zhou, Rui Wang, Richard Southern, and Jian Jun Zhang. Blue noise sampling using an SPH-based method. ACM Transactions on Graphics, 34(6):211:1– 211:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kass:2008:AOM

[KA08] Michael Kass and John Anderson. Animating oscillatory motion with overlap: wiggly splines. ACM Transactions on Graphics, 27(3):28:1– 28:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2010:AGD

[KAB^{+}10] Johannes Kopf, Maneesh Agrawala, David Bargeron, David Salesin, and Michael Cohen. Automatic generation of destination maps. ACM Transactions on Graphics, 29 (6):158:1–158:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kovalsky:2014:CSV

[KABL14] Shahar Z. Kovalsky, Noam Aigerman, Ronen Basri, and Yaron Lipman. Controlling singular values with semidefinite programming. ACM Transactions on Graphics, 33 (4):68:1–68:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kovalsky:2015:LSB

[KABL15] Shahar Z. Kovalsky, Noam Aigerman, Ronen Basri, and Yaron Lipman. Largescale bounded distortion mappings. ACM Transactions on Graphics, 34(6):191:1– 191:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karacan:2020:MAN

[KAEE20] Levent Karacan, Zeynep Akata, Aykut Erdem, and Erkut Erdem. Manipulating attributes of natural scenes via hallucination. ACM Transactions on Graphics, 39 (1):7:1–7:17, February 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3368312.

Kim:2020:LNS

[KAGS20] Byungsoo Kim, Vinicius C.

Azevedo, Markus Gross, and Barbara Solenthaler. Lagrangian neural style transfer for fluids. ACM Transactions on Graphics, 39(4): 52:1–52:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392473.

Kajiya:1983:NTR

[Kaj83] James T. Kajiya. New techniques for ray tracing procedurally defined objects. ACM Transactions on Graphics, 2 (3):161–181, July 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). Also appeared in SIG-GRAPH ’83 Proceedings, and in Tutorial: Computer Graphics: Image Synthesis, Computer Society Press, Washington, 1988, pp. 168–188.

Kallmann:2014:DRL

[Kal14] Marcelo Kallmann. Dynamic and robust local clearance triangulations. ACM Transactions on Graphics, 33(5): 161:1–161:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karras:2017:ADF

[KAL^{+}17] Tero Karras, Timo Aila, Samuli Laine, Antti Herva, and Jaakko Lehtinen. Audiodriven facial animation by joint end-to-end learning of pose and emotion. ACM Transactions on Graphics, 36 (4):94:1–94:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalogerakis:2018:SDL

[Kal18] Evangelos Kalogerakis. Session details: Learning geometry. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kristensen:2005:PLR

[KAMJ05] Anders Wang Kristensen, Tomas Akenine-Möller, and Henrik Wann Jensen. Precomputed local radiance transfer for real-time lighting design. ACM Transactions on Graphics, 24(3):1208–1215, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kang:2015:HCE

[Kan15] Sing Bing Kang. Homogeneous codes for energyefficient illumination and imaging. ACM Transactions on Graphics, 34(4):35:1– 35:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaufman:2018:SDS

[Kau18] Danny Kaufman. Session details: Structured simulation. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Konrad:2020:GCO

[KAW20] Robert Konrad, Anastasios Angelopoulos, and Gordon Wetzstein. Gaze-contingent ocular parallax rendering for virtual reality. ACM Transactions on Graphics, 39(2): 10:1–10:12, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3361330.

Kilgard:2012:GAP

[KB12] Mark J. Kilgard and Jeff Bolz. GPU-accelerated path rendering. ACM Transactions on Graphics, 31(6): 172:1–172:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koulieris:2017:ACH

[KBBD17] George-Alex Koulieris, Bee Bui, Martin S. Banks, and George Drettakis. Accommodation and comfort in headmounted displays. ACM Transactions on Graphics, 36 (4):87:1–87:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kimmel:2013:SAC

[KBC^{+}13] Bradley W. Kimmel, Gladimir V. G. Baranoski, T. F. Chen, Daniel Yim, and Erik Miranda. Spectral appearance changes induced by light exposure. ACM Transactions on Graphics, 32(1):10:1– 10:13, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kautz:2007:IEM

[KBD07] Jan Kautz, Solomon Boulos, and Frédo Durand. Interactive editing and modeling of bidirectional texture functions. ACM Transactions on Graphics, 26(3):53:1–53:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalantari:2015:MLA

[KBS15] Nima Khademi Kalantari, Steve Bako, and Pradeep Sen. A machine learning approach for filtering Monte Carlo noise. ACM Transactions on Graphics, 34(4): 122:1–122:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koschier:2017:REF

[KBT17] Dan Koschier, Jan Bender, and Nils Thuerey. Robust eXtended finite elements for complex cutting of deformables. ACM Transactions on Graphics, 36(4): 55:1–55:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kovacs:2015:DMS

[KBZ15] Denis Kovacs, Justin Bisceglio, and Denis Zorin. Dyadic T-mesh subdivision. ACM Transactions on Graphics, 34(4):143:1–143:??, August 2015. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kang:2019:IAN

[KC19] Kyoungkook Kang and Sunghyun Cho. Interactive and automatic navigation for 360^{∘} video playback. ACM Transactions on Graphics, 38(4): 108:1–108:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2021:DCC

[KC21] Suzi Kim and Sunghee Choi.

Dynamic closest color warping to sort and compare palettes. ACM Transactions on Graphics, 40(4): 95:1–95:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459776.

Khanteimouri:2023:BGB

[KC23] Payam Khanteimouri and Marcel Campen. 3D Bézier guarding: Boundary-conforming curved tetrahedral meshing. ACM Transactions on Graphics, 42(6):182:1– 182:??, December 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618332.

Kobilarov:2009:LGI

[KCD09] Marin Kobilarov, Keenan Crane, and Mathieu Desbrun. Lie group integrators for animation and control of vehicles. ACM Transactions on Graphics, 28(2):16:1–16:14, April 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Konakovic:2016:BDC

[KCD^{+}16] Mina Konaković, Keenan Crane, Bailin Deng, Sofien Bouaziz, Daniel Piker, and Mark Pauly. Beyond developable: computational design and fabrication with auxetic materials. ACM Transactions on Graphics, 35(4): 89:1–89:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2014:SHC

[KCGF14] Vladimir G. Kim, Siddhartha Chaudhuri, Leonidas Guibas, and Thomas Funkhouser. Shape2Pose: human-centric shape analysis. ACM Transactions on Graphics, 33(4): 120:1–120:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalogerakis:2012:PMC

[KCKK12] Evangelos Kalogerakis, Siddhartha Chaudhuri, Daphne Koller, and Vladlen Koltun. A probabilistic model for component-based shape synthesis. ACM Transactions on Graphics, 31(4): 55:1–55:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2007:JBU

[KCLU07] Johannes Kopf, Michael F.

Cohen, Dani Lischinski, and Matt Uyttendaele. Joint bilateral upsampling. ACM Transactions on Graphics, 26(3): 96:1–96:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kilian:2023:MSF

[KCMP23] Martin Kilian, Anthony S Ramos Cisneros, Christian Müller, and Helmut Pottmann. Meshes with spherical faces. ACM Transactions on Graphics, 42(6):

184:1–184:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618345.

Kopf:2006:RWT

[KCODL06] Johannes Kopf, Daniel Cohen-Or, Oliver Deussen, and Dani Lischinski. Recursive Wang tiles for real-time blue noise. ACM Transactions on Graphics, 25(3):509–518, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knoppel:2013:GOD

[KCPS13] Felix Knöppel, Keenan Crane, Ulrich Pinkall, and Peter Schröder. Globally optimal direction fields. ACM Transactions on Graphics, 32(4): 59:1–59:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knoppel:2015:SPS

[KCPS15] Felix Knöppel, Keenan Crane, Ulrich Pinkall, and Peter Schröder. Stripe patterns on surfaces. ACM Transactions on Graphics, 34(4): 39:1–39:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2014:FPH

[KCS14] Johannes Kopf, Michael F. Cohen, and Richard Szeliski.

First-person hyper-lapse videos. ACM Transactions on Graphics, 33(4):78:1–78:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2010:SSB

[KCSC10] Johannes Kopf, Billy Chen, Richard Szeliski, and Michael Cohen. Street slide: browsing street level imagery. ACM Transactions on Graphics, 29 (4):96:1–96:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kulla:2018:SPI

[KCSG18] Christopher Kulla, Alejandro Conty, Clifford Stein, and Larry Gritz. Sony Pictures Imageworks Arnold. ACM Transactions on Graphics, 37 (3):29:1–29:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3180495.

Kang:2018:ERC

[KCW^{+}18] Kaizhang Kang, Zimin Chen, Jiaping Wang, Kun Zhou, and Hongzhi Wu. Efficient reflectance capture using an autoencoder. ACM Transactions on Graphics, 37(4): 127:1–127:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kuang:2013:CRA

[KCYW13] Zhengzheng Kuang, Bin Chan, Yizhou Yu, and Wenping Wang. A compact random-access representation for urban modeling and rendering. ACM Transactions on Graphics, 32(6):172:1– 172:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kavan:2008:GSA

[KCŽO08] Ladislav Kavan, Steven Collins, Jiří Žára, and Carol O’Sullivan. Geometric skinning with approximate dual quaternion blending. ACM Transactions on Graphics, 27(4):105:1– 105:23, October 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaplanyan:2013:APP

[KD13a] Anton S. Kaplanyan and Carsten Dachsbacher. Adaptive progressive photon mapping. ACM Transactions on Graphics, 32(2):16:1– 16:13, April 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2013:SFR

[KD13b] Theodore Kim and John Delaney. Subspace fluid resimulation. ACM Transactions on Graphics, 32(4): 62:1–62:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2022:MRN

[KDH22] Seung-Wook Kim, Jaehyung Doh, and Junghyun Han. Modeling and rendering noneuclidean spaces approximated with concatenated polytopes. ACM Transactions on Graphics, 41(4): 78:1–78:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530186.

Kim:2019:AEI

[KDI19] Theodore Kim, Fernando De Goes, and Hayley Iben. Anisotropic elasticity for inversion-safety and element rehabilitation. ACM Transactions on Graphics, 38(4): 69:1–69:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kellnhofer:2016:GSD

[KDM^{+}16] Petr Kellnhofer, Piotr Didyk, Karol Myszkowski, Mohamed M. Hefeeda, Hans-Peter Seidel, and Wojciech Matusik. GazeStereo3D: seamless disparity manipulations. ACM Transactions on Graphics, 35 (4):68:1–68:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalnins:2003:CSS

[KDMF03] Robert D. Kalnins, Philip L. Davidson, Lee Markosian, and Adam Finkelstein. Coherent stylized silhouettes. ACM Transactions on Graphics, 22 (3):856–861, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Konrad:2017:STL

[KDMW17] Robert Konrad, Donald G. Dansereau, Aniq Masood, and Gordon Wetzstein. SpinVR: towards live-streaming 3D virtual reality video. ACM Transactions on Graphics, 36 (6):209:1–209:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kettunen:2021:URM

[KDPN21] Markus Kettunen, Eugene D’Eon, Jacopo Pantaleoni, and Jan Novák. An unbiased ray-marching transmittance estimator. ACM Transactions on Graphics, 40(4): 137:1–137:20, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459937.

Kellnhofer:2016:MPS

[KDR^{+}16] Petr Kellnhofer, Piotr Didyk, Tobias Ritschel, Belen Masia, Karol Myszkowski, and Hans-Peter Seidel. Motion parallax in stereo 3D: model and applications. ACM Transactions on Graphics, 35(6): 176:1–176:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kellnhofer:2017:THE

[KDW^{+}17] Petr Kellnhofer, Piotr Didyk, Szu-Po Wang, Pitchaya Sitthi-Amorn, William Freeman, Fredo Durand, and Wojciech Matusik. 3DTV at home: Eulerian–Lagrangian stereo-to-multiview conversion. ACM Transactions on Graphics, 36(4):146:1– 146:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kanamori:2018:RHO

[KE18] Yoshihiro Kanamori and Yuki Endo. Relighting humans: occlusion-aware inverse rendering for full-body human images. ACM Transactions on Graphics, 37(6):270:1– 270:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwatra:2005:TOE

[KEBK05] Vivek Kwatra, Irfan Essa, Aaron Bobick, and Nipun Kwatra. Texture optimization for example-based synthesis. ACM Transactions on Graphics, 24(3):795–802, July 2005. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Karacan:2013:SPI

[KEE13] Levent Karacan, Erkut Erdem, and Aykut Erdem. Structure-preserving image smoothing via region covariances. ACM Transactions on Graphics, 32(6):176:1– 176:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaufman:2005:FFD

[KEP05] Danny M. Kaufman, Timothy Edmunds, and Dinesh K. Pai. Fast frictional dynamics for rigid bodies. ACM Transactions on Graphics, 24 (3):946–956, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kurlander:1993:ICM

[KF93] David Kurlander and Steven Feiner. Inferring constraints from multiple snapshots. ACM Transactions on Graphics, 12(4):277–304, October 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        159731.html.

Krishnan:2009:DFP

[KF09] Dilip Krishnan and Rob Fergus. Dark flash photography. ACM Transactions on Graphics, 28(3):96:1–96:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krivanek:2010:EGI

[KFB10] Jaroslav Křivánek, James A.

Ferwerda, and Kavita Bala. Effects of global illumination approximations on material appearance. ACM Transactions on Graphics, 29(4): 112:1–112:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kilian:2008:CF

[KFC^{+}08] Martin Kilian, Simon Flöry, Zhonggui Chen, Niloy J. Mitra, Alla Sheffer, and Helmut Pottmann. Curved folding. ACM Transactions on Graphics, 27(3):75:1–75:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Klingner:2006:FAD

[KFCO06] Bryan M. Klingner, Bryan E.

Feldman, Nuttapong Chentanez, and James F. O’Brien. Fluid animation with dynamic meshes. ACM Transactions on Graphics, 25(3): 820–825, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2007:STS

[KFCO^{+}07] Johannes Kopf, Chi-Wing Fu, Daniel Cohen-Or, Oliver Deussen, Dani Lischinski, and Tien-Tsin Wong. Solid texture synthesis from 2D exemplars. ACM Transactions on Graphics, 26(3):2:1–2:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kazhdan:2004:SMA

[KFR04] Michael Kazhdan, Thomas Funkhouser, and Szymon Rusinkiewicz. Shape matching and anisotropy. ACM Transactions on Graphics, 23 (3):623–629, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krishnan:2013:EPL

[KFS13] Dilip Krishnan, Raanan Fattal, and Richard Szeliski. Efficient preconditioning of Laplacian matrices for computer graphics. ACM Transactions on Graphics, 32(4): 142:1–142:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kelly:2017:BLS

[KFWM17] Tom Kelly, John Femiani, Peter Wonka, and Niloy J. Mitra. BigSUR: large-scale structured urban reconstruction. ACM Transactions on Graphics, 36(6):204:1– 204:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kovar:2004:AEP

[KG04] Lucas Kovar and Michael Gleicher. Automated extraction and parameterization of motions in large data sets. ACM Transactions on Graphics, 23(3):559–568, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kho:2005:SMD

[KG05] Youngihn Kho and Michael Garland. Sketching mesh deformations. ACM Transactions on Graphics, 24(3): 934, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kircher:2006:EAD

[KG06] Scott Kircher and Michael Garland. Editing arbitrarily deforming surface animations. ACM Transactions on Graphics, 25(3):1098–1107, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kircher:2008:FFM

[KG08] Scott Kircher and Michael Garland. Free-form motion processing. ACM Transactions on Graphics, 27(2):12:1– 12:13, April 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kainz:2009:RCM

[KGB^{+}09] Bernhard Kainz, Markus Grabner, Alexander Bornik, Stefan Hauswiesner, Judith Muehl, and Dieter Schmalstieg. Ray casting of multiple volumetric datasets with polyhedral boundaries on manycore GPUs. ACM Transactions on Graphics, 28(5): 152:1–152:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kavan:2011:PIU

[KGBS11] Ladislav Kavan, Dan Gerszewski, Adam W. Bargteil, and Peter-Pike Sloan. Physicsinspired upsampling for cloth simulation in games. ACM Transactions on Graphics, 30 (4):93:1–93:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karsch:2014:CAV

[KGFF14] Kevin Karsch, Mani Golparvar-Fard, and David Forsyth. ConstructAide: analyzing and visualizing construction sites through photographs and building models. ACM Transactions on Graphics, 33(6): 176:1–176:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kapp:2020:DDA

[KGG^{+}20] Konrad Kapp, James Gain, Eric Guérin, Eric Galin, and Adrien Peytavie. Datadriven authoring of large-scale ecosystems. ACM Transactions on Graphics, 39(6): 217:1–217:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417848.

Krivanek:2014:UPB

[KGH^{+}14] Jaroslav Krivánek, Iliyan Georgiev, Toshiya Hachisuka, Petr Vévoda, Martin Sik, Derek Nowrouzezahrai, and Wojciech Jarosz. Unifying points, beams, and paths in volumetric light transport simulation. ACM Transactions on Graphics, 33(4): 103:1–103:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kovalsky:2016:AQP

[KGL16] Shahar Z. Kovalsky, Meirav Galun, and Yaron Lipman. Accelerated quadratic proxy for geometric optimization. ACM Transactions on Graphics, 35(4):134:1–134:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kodnongbua:2022:CDP

[KGL^{+}22] Milin Kodnongbua, Ian Good, Yu Lou, Jeffrey Lipton, and Adriana Schulz. Computational design of passive grippers. ACM Transactions on Graphics, 41(4):149:1– 149:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530162.

Kovar:2002:MG

[KGP02] Lucas Kovar, Michael Gleicher, and Frédéric Pighin. Motion graphs. ACM Transactions on Graphics, 21(3): 473–482, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Klar:2016:DPE

[KGP^{+}16] Gergely Klár, Theodore Gast, Andre Pradhana, Chuyuan Fu, Craig Schroeder, Chenfanfu Jiang, and Joseph Teran. Drucker–Prager elastoplasticity for sand animation. ACM Transactions on Graphics, 35(4):103:1–103:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kelly:2018:FGD

[KGS^{+}18] Tom Kelly, Paul Guerrero, Anthony Steed, Peter Wonka, and Niloy J. Mitra. Franken-GAN: guided detail synthesis for building mass models using style-synchonized GANs. ACM Transactions on Graphics, 37(6):216:1– 216:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2018:DVP

[KGT^{+}18] Hyeongwoo Kim, Pablo Garrido, Ayush Tewari, Weipeng Xu, Justus Thies, Matthias Niessner, Patrick Pérez, Christian Richardt, Michael Zollhöfer, and Christian Theobalt. Deep video portraits. ACM Transactions on Graphics, 37(4):163:1– 163:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karpenko:2006:SFF

[KH06] Olga A. Karpenko and John F. Hughes. SmoothSketch: 3D free-form shapes from complex sketches. ACM Transactions on Graphics, 25 (3):589–598, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kazhdan:2008:SMG

[KH08] Michael Kazhdan and Hugues Hoppe. Streaming multigrid for gradient-domain operations on large images. ACM Transactions on Graphics, 27 (3):21:1–21:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kazhdan:2010:MAP

[KH10] Michael Kazhdan and Hugues Hoppe. Metric-aware processing of spherical imagery. ACM Transactions on Graphics, 29 (6):149:1–149:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kazhdan:2013:SPS

[KH13] Michael Kazhdan and Hugues Hoppe. Screened Poisson surface reconstruction. ACM Transactions on Graphics, 32 (3):29:1–29:13, June 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwon:2017:MMIa

[KH17a] Taesoo Kwon and Jessica K. Hodgins. Momentum-mapped inverted pendulum models for controlling dynamic human motions. ACM Transactions on Graphics, 36(1):10:1– 10:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwon:2017:MMIb

[KH17b] Taesoo Kwon and Jessica K. Hodgins. Momentum-mapped inverted pendulum models for controlling dynamic human motions. ACM Transactions on Graphics, 36(4): 145:1–145:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaplanyan:2014:NCR

[KHD14] Anton S. Kaplanyan, Johannes Hanika, and Carsten Dachsbacher. The naturalconstraint representation of the path space for efficient light transport simulation. ACM Transactions on Graphics, 33(4):102:1–102:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karsch:2011:RSO

[KHFH11] Kevin Karsch, Varsha Hedau, David Forsyth, and Derek Hoiem. Rendering synthetic objects into legacy photographs. ACM Transactions on Graphics, 30(6): 157:1–157:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2011:MPS

[KHH^{+}11] Changil Kim, Alexander Hornung, Simon Heinzle, Wojciech Matusik, and Markus Gross. Multi-perspective stereoscopy from light fields. ACM Transactions on Graphics, 30(6):190:1–190:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2009:SMC

[KHKL09] Manmyung Kim, Kyunglyul Hyun, Jongmin Kim, and Jehee Lee. Synchronized multicharacter motion editing.

ACM Transactions on Graphics, 28(3):79:1–79:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2011:HDF

[KHKR11] Jaewon Kim, Roarke Horstmeyer, Ig-Jae Kim, and Ramesh Raskar. Highlighted depthof-field photography: Shining light on focus. ACM Transactions on Graphics, 30(3): 24:1–24:9, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kettunen:2019:DCR

[KHL19] Markus Kettunen, Erik Härkönen, and Jaakko Lehtinen. Deep convolutional reconstruction for gradient-domain rendering. ACM Transactions on Graphics, 38(4):126:1– 126:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kutz:2017:SDT

[KHLN17] Peter Kutz, Ralf Habel, Yining Karl Li, and Jan Novák. Spectral and decomposition tracking for rendering heterogeneous volumes. ACM Transactions on Graphics, 36 (4):111:1–111:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kahler:2003:RDR

[KHS03] Kolja Kähler, Jörg Haber, and Hans-Peter Seidel. Reanimating the dead: reconstruction of expressive faces from skull data. ACM Transactions on Graphics, 22(3):554–561, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalogerakis:2010:LMS

[KHS10] Evangelos Kalogerakis, Aaron Hertzmann, and Karan Singh. Learning 3D mesh segmentation and labeling. ACM Transactions on Graphics, 29 (4):102:1–102:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kadlecek:2016:RPA

[KIL^{+}16] Petr Kadlecek, Alexandru-Eugen Ichim, Tiantian Liu, Jaroslav Krivánek, and Ladislav Kavan. Reconstructing personalized anatomical models for physics-based body animation. ACM Transactions on Graphics, 35(6):213:1– 213:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2010:MPF

[Kim10]  Byungmoon Kim.  Multi-

phase fluid simulations using regional level sets. ACM Transactions on Graphics, 29 (6):175:1–175:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2018:SDI

[Kim18] Min H. Kim. Session details: IM-material. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koskela:2019:BMO

[KIM^{+}19] Matias Koskela, Kalle Immonen, Markku Mäkitalo, Alessandro Foi, Timo Viitanen, Pekka Jääskeläinen, Heikki Kultala, and Jarmo Takala. Blockwise multiorder feature regression for real-time path-tracing reconstruction. ACM Transactions on Graphics, 38(5): 138:1–138:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3269978.

Keinert:2015:SFM

[KISS15] Benjamin Keinert, Matthias Innmann, Michael Sänger, and Marc Stamminger. Spherical Fibonacci mapping. ACM Transactions on Graphics, 34 (6):193:1–193:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2009:SSD

[KJ09] Theodore Kim and Doug L.

James. Skipping steps in deformable simulation with online model reduction. ACM Transactions on Graphics, 28(5):123:1–123:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2009:RCG

[KJDL09] Yongjin Kim, Cheolhun Jang, Julien Demouth, and Seungyong Lee. Robust color-to-gray via nonlinear global mapping. ACM Transactions on Graphics, 28(5):161:1– 161:4, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2023:DTF

[KJGP23] Juhyeon Kim, Wojciech Jarosz, Ioannis Gkioulekas, and Adithya Pediredla. Doppler time-of-flight rendering. ACM Transactions on Graphics, 42 (6):271:1–271:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618335.

Kaldor:2008:SKC

[KJM08] Jonathan M. Kaldor, Doug L.

James, and Steve Marschner. Simulating knitted cloth at the yarn level. ACM Transactions on Graphics, 27(3):

65:1–65:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaldor:2010:EYB

[KJM10] Jonathan M. Kaldor, Doug L. James, and Steve Marschner. Efficient yarn-based cloth with adaptive contact linearization. ACM Transactions on Graphics, 29(4): 105:1–105:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2019:FAD

[KJS^{+}19] Jonghyun Kim, Youngmo Jeong, Michael Stengel, Kaan Aksit, Rachel Albert, Ben Boudaoud, Trey Greer, Joohwan Kim, Ward Lopes, Zander Majercik, Peter Shirley, Josef Spjut, Morgan McGuire, and David Luebke. Foveated AR: dynamically-foveated augmented reality display. ACM Transactions on Graphics, 38 (4):99:1–99:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kamada:1987:ETH

[KK87] Tomihisa Kamada and Satoru Kawai. An enhanced treatment of hidden lines. ACM Transactions on Graphics, 6 (4):308–323, October 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
35042.html.

Kamada:1991:GFV

[KK91] Tomihisa Kamada and Satoru Kawai. A general framework for visualizing abstract objects and relations. ACM Transactions on Graphics, 10(1):1–39, January 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
99903.html.

Kulik:2011:CSS

[KKB^{+}11] Alexander Kulik, André Kunert, Stephan Beck, Roman Reichel, Roland Blach, Armin Zink, and Bernd Froehlich. C1x6: a stereoscopic six-user display for co-located collaboration in shared virtual environments. ACM Transactions on Graphics, 30(6):188:1–188:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2012:QPI

[KKDK12] Johannes Kopf, Wolf Kienzle, Steven Drucker, and Sing Bing Kang. Quality prediction for image completion. ACM Transactions on Graphics, 31(6):131:1– 131:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kerbl:2023:GSR

[KKLD23] Bernhard Kerbl, Georgios Kopanas, Thomas Leimkuehler, and George Drettakis. 3D Gaussian splatting for realtime radiance field rendering. ACM Transactions on Graphics, 42(4):139:1– 139:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592433.

Kim:2013:NEP

[KKN^{+}13] Doyub Kim, Woojong Koh, Rahul Narain, Kayvon Fatahalian, Adrien Treuille, and James F. O’Brien. Nearexhaustive precomputation of secondary cloth effects. ACM Transactions on Graphics, 32 (4):87:1–87:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kider:2014:FEC

[KKN^{+}14] Joseph T. Kider, Jr., Daniel Knowlton, Jeremy Newlin, Yining Karl Li, and Donald P. Greenberg. A framework for the experimental comparison of solar and skydome illumination. ACM Transactions on Graphics, 33(6): 180:1–180:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2022:THQ

[KKN^{+}22] Jungeon Kim, Hyomin Kim, Hyeonseo Nam, Jaesik Park, and Seungyong Lee. TextureMe: High-quality textured scene reconstruction in real time. ACM Transactions on Graphics, 41(3): 24:1–24:18, June 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3503926.

Kjolstad:2016:SLP

[KKRK^{+}16] Fredrik Kjolstad, Shoaib Kamil, Jonathan Ragan-Kelley, David I. W. Levin, Shinjiro Sueda, Desai Chen, Etienne Vouga, Danny M. Kaufman, Gurtej Kanwar, Wojciech Matusik, and Saman Amarasinghe. Simit: a language for physical simulation. ACM Transactions on Graphics, 35(2):20:1–20:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kenzel:2018:HPS

[KKSS18] Michael Kenzel, Bernhard Kerbl, Dieter Schmalstieg, and Markus Steinberger. A high-performance software graphics pipeline architecture for the GPU. ACM Transactions on Graphics, 37(4): 140:1–140:??, August 2018.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krajancich:2020:ODP

[KKW20] Brooke Krajancich, Petr Kellnhofer, and Gordon Wetzstein. Optimizing depth perception in virtual and augmented reality through gaze-contingent stereo rendering. ACM Transactions on Graphics, 39(6): 269:1–269:10, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417820.

Krajancich:2021:PME

[KKW21] Brooke Krajancich, Petr Kellnhofer, and Gordon Wetzstein. A perceptual model for eccentricity-dependent spatiotemporal flicker fusion and its applications to foveated graphics. ACM Transactions on Graphics, 40(4): 47:1–47:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459784.

Krajancich:2023:TAA

[KKW23] Brooke Krajancich, Petr Kellnhofer, and Gordon Wetzstein. Towards attentionaware foveated rendering. ACM Transactions on Graphics, 42(4):77:1–77:??, August 2023. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592406.

Kopf:2011:DPA

[KL11] Johannes Kopf and Dani Lischinski. Depixelizing pixel art. ACM Transactions on Graphics, 30(4): 99:1–99:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2012:DRH

[KL12] Johannes Kopf and Dani Lischinski. Digital reconstruction of halftoned color comics. ACM Transactions on Graphics, 31(6):140:1– 140:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kang:2017:MCLa

[KL17a] Changgu Kang and Sung-Hee Lee. Multi-contact locomotion using a contact graph with feasibility predictors. ACM Transactions on Graphics, 36 (2):22:1–22:??, April 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kang:2017:MCLb

[KL17b] Changgu Kang and Sung-Hee Lee. Multi-contact locomotion using a contact graph with feasibility predictors. ACM Transactions on Graphics, 36 (4):145:1–145:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Korosteleva:2022:NRS

[KL22] Maria Korosteleva and Sung-Hee Lee. NeuralTailor: reconstructing sewing pattern structures from 3D point clouds of garments. ACM Transactions on Graphics, 41 (4):158:1–158:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530179.

Kim:2023:PVH

[KL23] Janghun Kim and Sungkil Lee. Potentially visible hidden-volume rendering for multi-view warping. ACM Transactions on Graphics, 42 (4):86:1–86:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592108.

Klassen:1987:MEA

[Kla87] R. Victor Klassen. Modeling the effect of the atmosphere on light. ACM Transactions on Graphics, 6(3): 215–237, July 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
35071.html.

Klassen:1991:DAC

[Kla91a] R. Victor Klassen. Drawing antialiased cubic spline curves. ACM Transactions on Graphics, 10(1):92– 108, January 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
99906.html.

Klassen:1991:IFD

[Kla91b] R. Victor Klassen. Integer forward differencing of cubic polynomials: Analysis and algorithms. ACM Transactions on Graphics, 10(2):152–181, April 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
108364.html.

Klassen:1994:EIH

[Kla94] R. Victor Klassen. Exact integer hybrid subdivision and forward differencing of cubics. ACM Transactions on Graphics, 13(3):240–255, July 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
197476.html.

Kim:2011:BIM

[KLF11] Vladimir G. Kim, Yaron Lipman, and Thomas Funkhouser.

Blended intrinsic maps. ACM Transactions on Graphics, 30 (4):79:1–79:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2012:SGT

[KLF12] Vladimir G. Kim, Yaron Lipman, and Thomas Funkhouser. Symmetry-guided texture synthesis and manipulation. ACM Transactions on Graphics, 31(3):22:1–22:14, May 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kuo:2019:CIC

[KLF^{+}19] Calvin Kuo, Ziheng Liang, Ye Fan, Jean-Sébastien Blouin, and Dinesh K. Pai. Creating impactful characters: correcting human impact accelerations using high rate IMUs in dynamic activities. ACM Transactions on Graphics, 38 (4):47:1–47:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krahenbuhl:2009:SRS

[KLHG09] Philipp Krähenbühl, Manuel Lang, Alexander Hornung, and Markus Gross. A system for retargeting of streaming video. ACM Transactions on Graphics, 28(5): 126:1–126:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2013:SLD

[KLKL13] Yongjin Kim, Yunjin Lee, Henry Kang, and Seungyong Lee. Stereoscopic 3D line drawing. ACM Transactions on Graphics, 32(4): 57:1–57:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2007:SBF

[KLL^{+}07] Byungmoon Kim, Yingjie Liu, Ignacio Llamas, Xiangmin Jiao, and Jarek Rossignac. Simulation of bubbles in foam with the volume control method. ACM Transactions on Graphics, 26(3): 98:1–98:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwon:2008:GME

[KLLT08] Taesoo Kwon, Kang Hoon Lee, Jehee Lee, and Shigeo Takahashi. Group motion editing. ACM Transactions on Graphics, 27(3): 80:1–80:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2012:ECM

[KLM^{+}12] Vladimir G. Kim, Wilmot Li, Niloy J. Mitra, Stephen Di-Verdi, and Thomas Funkhouser. Exploring collections of 3D models using fuzzy correspondences. ACM Transactions on Graphics, 31(4):

54:1–54:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2013:LPB

[KLM^{+}13] Vladimir G. Kim, Wilmot Li, Niloy J. Mitra, Siddhartha Chaudhuri, Stephen DiVerdi, and Thomas Funkhouser. Learning part-based templates from large collections of 3D shapes. ACM Transactions on Graphics, 32(4): 70:1–70:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2024:NHR

[KLM24] Doyub Kim, Minjae Lee, and Ken Museth. NeuralVDB: High-resolution sparse volume representation using hierarchical neural networks. ACM Transactions on Graphics, 43 (2):20:1–20:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3641817.

Karasick:1991:EDT

[KLN91] Michael Karasick, Derek Lieber, and Lee R. Nackman. Efficient Delaunay triangulation using rational arithmetic. ACM Transactions on Graphics, 10(1):71–91, January 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
99905.html.

Konakovic-Lukovic:2018:RDC

[KLPCP18] Mina Konaković-Luković, Julian Panetta, Keenan Crane, and Mark Pauly. Rapid deployment of curved surfaces via programmable auxetics. ACM Transactions on Graphics, 37(4):106:1– 106:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopanas:2022:NPC

[KLR^{+}22] Georgios Kopanas, Thomas Leimkühler, Gilles Rainer, Clément Jambon, and George Drettakis. Neural point catacaustics for novel-view synthesis of reflections. ACM Transactions on Graphics, 41(6): 201:1–201:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555497.

Khodakovsky:2003:GSP

[KLS03] Andrei Khodakovsky, Nathan Litke, and Peter Schröder. Globally smooth parameterizations with low distortion. ACM Transactions on Graphics, 22(3):350–357, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2013:IBR

[KLS^{+}13] Johannes Kopf, Fabian Langguth, Daniel Scharstein, Richard Szeliski, and Michael Goesele. Image-based rendering in the gradient domain. ACM Transactions on Graphics, 32(6): 199:1–199:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwon:2020:FFM

[KLV20] Taesoo Kwon, Yoonsang Lee, and Michiel Van De Panne. Fast and flexible multilegged locomotion using learned centroidal dynamics. ACM Transactions on Graphics, 39 (4):46:1–46:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392432.

Koo:2014:CWL

[KLY^{+}14] Bongjin Koo, Wilmot Li, JiaXian Yao, Maneesh Agrawala, and Niloy J. Mitra. Creating works-like prototypes of mechanical objects. ACM Transactions on Graphics, 33(6): 217:1–217:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krishnan:1997:ESI

[KM97] Shankar Krishnan and Dinesh Manocha. An efficient surface intersection algorithm based on lower-dimensional formulation. ACM Transactions on Graphics, 16(1): 74–106, January 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1997-16-1/
p74-krishnan/.

Khungurn:2017:ASE

[KM17] Pramook Khungurn and Steve Marschner. Azimuthal scattering from elliptical hair fibers. ACM Transactions on Graphics, 36(2):13:1– 13:??, April 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kettunen:2015:GDP

[KMA^{+}15] Markus Kettunen, Marco Manzi, Miika Aittala, Jaakko Lehtinen, Frédo Durand, and Matthias Zwicker. Gradientdomain path tracing. ACM Transactions on Graphics, 34(4):123:1–123:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaufmann:2009:ETD

[KMB^{+}09] Peter Kaufmann, Sebastian Martin, Mario Botsch, Eitan Grinspun, and Markus Gross. Enrichment textures for detailed cutting of shells. ACM Transactions on Graphics, 28 (3):50:1–50:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalnins:2002:WND

[KMM^{+}02] Robert D. Kalnins, Lee Markosian, Barbara J. Meier, Michael A. Kowalski, Joseph C. Lee, Philip L. Davidson, Matthew Webb, John F. Hughes, and Adam Finkelstein. WYSIWYG NPR: drawing strokes directly on 3D models. ACM Transactions on Graphics, 21(3): 755–762, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kallweit:2017:DSR

[KMM^{+}17a] Simon Kallweit, Thomas Müller, Brian Mcwilliams, Markus Gross, and Jan Novák. Deep scattering: rendering atmospheric clouds with radiance-predicting neural networks. ACM Transactions on Graphics, 36(6): 231:1–231:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kilian:2017:SACa

[KMM17b] Martin Kilian, Aron Monszpart, and Niloy J. Mitra. String actuated curved folded surfaces. ACM Transactions on Graphics, 36(3): 25:1–25:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kilian:2017:SACb

[KMM17c] Martin Kilian, Aron Monszpart, and Niloy J. Mitra. String actuated curved folded surfaces. ACM Transactions on Graphics, 36(4): 64:1–64:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kharevych:2009:NCI

[KMOD09] Lily Kharevych, Patrick Mullen, Houman Owhadi, and Mathieu Desbrun. Numerical coarsening of inhomogeneous elastic materials. ACM Transactions on Graphics, 28(3): 51:1–51:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kilian:2007:GMS

[KMP07] Martin Kilian, Niloy J. Mitra, and Helmut Pottmann. Geometric modeling in shape space. ACM Transactions on Graphics, 26(3):64:1– 64:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kuznetsov:2021:NMR

[KMX^{+}21] Alexandr Kuznetsov, Krishna Mullia, Zexiang Xu, Milos Hasan, and Ravi Ramamoorthi. NeuMIP: multi-resolution neural materials. ACM Transactions on Graphics, 40(4): 175:1–175:13, August 2021.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459795.

Kim:2012:AIE

[KMYG12] Young Min Kim, Niloy J. Mitra, Dong-Ming Yan, and Leonidas Guibas. Acquiring 3D indoor environments with variability and repetition. ACM Transactions on Graphics, 31(6):138:1– 138:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2002:IMH

[KN02] Tae-Yong Kim and Ulrich Neumann. Interactive multiresolution hair modeling and editing. ACM Transactions on Graphics, 21(3):620–629, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kuthirummal:2006:MRC

[KN06] Sujit Kuthirummal and Shree K. Nayar. Multiview radial catadioptric imaging for scene capture. ACM Transactions on Graphics, 25(3):916–923, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalogerakis:2012:LHP

[KNBH12] Evangelos Kalogerakis, Derek Nowrouzezahrai, Simon Breslav, and Aaron Hertzmann. Learning hatching for pen-and-ink illustration of surfaces. ACM Transactions on Graphics, 31 (1):1:1–1:17, January 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2008:DPM

[KNC^{+}08] Johannes Kopf, Boris Neubert, Billy Chen, Michael Cohen, Daniel Cohen-Or, Oliver Deussen, Matt Uyttendaele, and Dani Lischinski. Deep photo: modelbased photograph enhancement and viewing. ACM Transactions on Graphics, 27 (5):116:1–116:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2022:LDC

[KNK^{+}22] Hyomin Kim, Hyeonseo Nam, Jungeon Kim, Jaesik Park, and Seungyong Lee. LaplacianFusion: Detailed 3D clothed-human body reconstruction. ACM Transactions on Graphics, 41(6): 216:1–216:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555511.

Kim:2022:AHI

[KNL^{+}22] Dongyeon Kim, Seung-Woo Nam, Byounghyo Lee, Jong-Mo Seo, and Byoungho Lee. Accommodative holography: improving accommodation response for perceptually realistic holographic displays. ACM Transactions on Graphics, 41 (4):111:1–111:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530147.

Kalogerakis:2009:DDC

[KNS^{+}09] Evangelos Kalogerakis, Derek Nowrouzezahrai, Patricio Simari, James Mccrae, Aaron Hertzmann, and Karan Singh. Data-driven curvature for real-time line drawing of dynamic scenes. ACM Transactions on Graphics, 28(1):11:1– 11:13, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knuth:1987:DHD

[Knu87] Donald E. Knuth. Digital halftones by dot diffusion. ACM Transactions on Graphics, 6(4):245–273, October 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
35040.html.

Kirk:2011:PBT

[KO11] Adam G. Kirk and James F. O’Brien. Perceptually based tone mapping for low-light conditions. ACM Transactions on Graphics, 30(4): 42:1–42:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kuang:2022:NNR

[KOC^{+}22] Zhengfei Kuang, Kyle Olszewski, Menglei Chai, Zeng Huang, Panos Achlioptas, and Sergey Tulyakov. NeROIC: neural rendering of objects from online image collections. ACM Transactions on Graphics, 41(4):56:1–56:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530177.

Kee:2013:EPM

[KOF13] Eric Kee, James O’Brien, and Hany Farid. Exposing photo manipulation with inconsistent shadows. ACM Transactions on Graphics, 32(3): 28:1–28:12, June 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kee:2014:EPM

[KOF14] Eric Kee, James F. O’Brien, and Hany Farid. Exposing photo manipulation from shading and shadows. ACM Transactions on Graphics, 33(5):165:1–165:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kulpa:2011:IRC

[KOOP11] Richard Kulpa, Anne-Hélène Olivier, Jan Ondřej, and Julien Pettré. Imperceptible relaxation of collision avoidance constraints in virtual crowds. ACM Transactions on Graphics, 30(6): 138:1–138:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2016:VS

[Kop16] Johannes Kopf. 360^{∘} video stabilization. ACM Transactions on Graphics, 35(6): 195:1–195:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koutaki:2016:BCI

[Kou16] Gou Koutaki. Binary continuous image decomposition for multi-view display. ACM Transactions on Graphics, 35 (4):69:1–69:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kasten:2021:LNA

[KOWD21] Yoni Kasten, Dolev Ofri, Oliver Wang, and Tali Dekel. Layered neural atlases for consistent video editing. ACM Transactions on Graphics, 40 (6):210:1–210:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480546.

Kim:2011:CBF

[KOY^{+}11] Yong-Joon Kim, Young-Taek Oh, Seung-Hyun Yoon, Myung-Soo Kim, and Gershon Elber. Coons BVH for freeform geometric models. ACM Transactions on Graphics, 30(6):169:1–169:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kasson:1992:ASC

[KP92] James M. Kasson and Wil Plouffe. An analysis of selected computer interchange color spaces. ACM Transactions on Graphics, 11 (4):373–405, October 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
146479.html.

Kim:2002:JIM

[KP02] Junhwan Kim and Fabio Pellacini. Jigsaw image mosaics. ACM Transactions on Graphics, 21(3):657–664, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kry:2003:CCS

[KP03] Paul G. Kry and Dinesh K.

Pai. Continuous contact simulation for smooth surfaces. ACM Transactions on Graphics, 22(1):106–129, January 2003. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Kry:2006:ICS

[KP06] Paul G. Kry and Dinesh K. Pai. Interaction capture and synthesis. ACM Transactions on Graphics, 25(3): 872–880, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karciauskas:2007:BPS

[KP07] Kȩstutis Karčiauskas and Jörg Peters. Bicubic polar subdivision. ACM Transactions on Graphics, 26(4):14:1– 14:6, October 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kerr:2009:TEL

[KP09] William B. Kerr and Fabio Pellacini. Toward evaluating lighting design interface paradigms for novice users. ACM Transactions on Graphics, 28(3):26:1–26:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kerr:2010:TEM

[KP10] William B. Kerr and Fabio Pellacini. Toward evaluating material design interface paradigms for novice users. ACM Transactions on Graphics, 29(4):35:1–35:??, July 2010. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Kass:2011:CNN

[KP11a] Michael Kass and Davide Pesare. Coherent noise for non-photorealistic rendering. ACM Transactions on Graphics, 30(4):30:1–30:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2011:FSS

[KP11b] Junggon Kim and Nancy S.

Pollard. Fast simulation of skeleton-driven deformable body characters. ACM Transactions on Graphics, 30(5): 121:1–121:19, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2018:DCA

[KP18] Tae-Hoon Kim and Sang Il Park. Deep context-aware descreening and rescreening of halftone images. ACM Transactions on Graphics, 37(4): 48:1–48:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kafri:2022:SDS

[KPACO22] Omer Kafri, Or Patashnik, Yuval Alaluf, and Daniel Cohen-Or. StyleFusion: Disentangling spatial segments in StyleGAN-generated images. ACM Transactions on Graphics, 41(5):179:1–179:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3527168.

Kuster:2012:GCH

[KPB^{+}12] Claudia Kuster, Tiberiu Popa, Jean-Charles Bazin, Craig Gotsman, and Markus Gross. Gaze correction for home video conferencing. ACM Transactions on Graphics, 31 (6):174:1–174:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2018:MDM

[KPH18] Seung-Wook Kim, Sun Young Park, and Junghyun Han. Magnetization dynamics for magnetic object interactions. ACM Transactions on Graphics, 37(4):121:1–121:??, August 2018. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krajcevski:2016:GGD

[KPM16] Pavel Krajcevski, Srihari Pratapa, and Dinesh Manocha. GST: GPU-decodable supercompressed textures. ACM Transactions on Graphics, 35 (6):230:1–230:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Konrad:2017:AIC

[KPM^{+}17] Robert Konrad, Nitish Padmanaban, Keenan Molner, Emily A. Cooper, and Gordon Wetzstein. Accommodationinvariant computational neareye displays. ACM Transactions on Graphics, 36(4): 88:1–88:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2017:DDP

[KPMP^{+}17] Meekyoung Kim, Gerard Pons-Moll, Sergi Pujades, Seungbae Bang, Jinwook Kim, Michael J. Black, and Sung-Hee Lee. Data-driven physics for human soft tissue animation. ACM Transactions on Graphics, 36(4): 54:1–54:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karciauskas:2017:JSS

[KPP17] Kestutis Karciauskas, Daniele Panozzo, and Jörg Peters. Tjunctions in spline surfaces. ACM Transactions on Graphics, 36(5):170:1–170:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knodt:2024:JUO

[KPWG24] Julian Knodt, Zherong Pan, Kui Wu, and Xifeng Gao. Joint UV optimization and texture baking. ACM Transactions on Graphics, 43 (1):2:1–2:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3617683.

Kilian:2017:MMF

[KPWP17] Martin Kilian, Davide Pellis, Johannes Wallner, and Helmut Pottmann. Materialminimizing forms and structures. ACM Transactions on Graphics, 36(6):173:1– 173:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knapitsch:2017:TTB

[KPZK17] Arno Knapitsch, Jaesik Park, Qian-Yi Zhou, and Vladlen Koltun. Tanks and temples: benchmarking largescale scene reconstruction. ACM Transactions on Graphics, 36(4):78:1–78:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kirschstein:2023:NMV

[KQG^{+}23] Tobias Kirschstein, Shenhan Qian, Simon Giebenhain, Tim Walter, and Matthias Nießner. NeRSemble: Multiview radiance field reconstruction of human heads. ACM Transactions on Graphics, 42(4):161:1–161:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592455.

Kalantari:2017:DHD

[KR17] Nima Khademi Kalantari and Ravi Ramamoorthi. Deep high dynamic range imaging of dynamic scenes. ACM Transactions on Graphics, 36 (4):144:1–144:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2012:ISM

[KRD^{+}12] Min H. Kim, Holly Rushmeier, Julie Dorsey, Todd Alan Harvey, Richard O. Prum, David S. Kittle, and David J. Brady. 3D imaging spectroscopy for measuring hyperspectral patterns on solid objects. ACM Transactions on Graphics, 31(4): 38:1–38:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kowdle:2018:NSR

[KRF^{+}18] Adarsh Kowdle, Christoph Rhemann, Sean Fanello, Andrea Tagliasacchi, Jonathan Taylor, Philip Davidson, Mingsong Dou, Kaiwen Guo, Cem Keskin, Sameh Khamis, David Kim, Danhang Tang, Vladimir Tankovich, Julien Valentin, and Shahram Izadi. The need 4 speed in real-time dense visual tracking. ACM Transactions on Graphics, 37 (6):220:1–220:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Khan:2006:IBM

[KRFB06] Erum Arif Khan, Erik Reinhard, Roland W. Fleming, and Heinrich H. Bülthoff. Image-based material editing. ACM Transactions on Graphics, 25(3):654–663, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2011:EAC

[KRK11] Min H. Kim, Tobias Ritschel, and Jan Kautz. Edge-aware color appearance. ACM Transactions on Graphics, 30 (2):13:1–13:9, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krogh:1982:AAP

[Kro82] F. T. Krogh. ACM algorithms policy. ACM Transactions on Graphics, 1(1):78–81, January 1982. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karasick:1995:ISM

[KS95] Michael Karasick and David Strip. Intersecting solids on a massively parallel processor. ACM Transactions on Graphics, 14(1):21–57, January 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
200976.html.

Kobbelt:1998:MFV

[KS98] Leif Kobbelt and Peter Schröder. A multiresolution framework for variational subdivision. ACM Transactions on Graphics, 17(4): 209–237, October 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1998-17-4/
p209-kobbelt/.

Kaplan:2004:ISP

[KS04a] Craig S. Kaplan and David H.

Salesin. Islamic star patterns in absolute geometry. ACM Transactions on Graphics, 23 (2):97–119, April 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kraevoy:2004:CPC

[KS04b] Vladislav Kraevoy and Alla Sheffer. Cross-parameterization and compatible remeshing of 3D models. ACM Transactions on Graphics, 23(3): 861–869, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kass:2010:SLH

[KS10] Michael Kass and Justin Solomon. Smoothed local histogram filters. ACM Transactions on Graphics, 29(4):

100:1–100:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Krishnan:2011:MMP

[KS11] Dilip Krishnan and Richard Szeliski. Multigrid and multilevel preconditioners for computational photography. ACM Transactions on Graphics, 30 (6):177:1–177:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kavan:2012:EID

[KS12] Ladislav Kavan and Olga Sorkine. Elasticity-inspired deformers for character articulation. ACM Transactions on Graphics, 31(6): 196:1–196:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kemelmacher-Shlizerman:2016:TP

[KS16] Ira Kemelmacher-Shlizerman. Transfiguring portraits. ACM Transactions on Graphics, 35 (4):94:1–94:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2021:OUL

[KS21] Joonho Kim and Karan Singh. Optimizing UI layouts for deformable face-rig manipulation. ACM Transactions on Graphics, 40(4): 172:1–172:12, August 2021.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459842.

Kampe:2013:HRS

[KSA13] Viktor Kämpe, Erik Sintorn, and Ulf Assarsson. High resolution sparse voxel DAGs. ACM Transactions on Graphics, 32(4):101:1–101:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalantari:2013:PBH

[KSB^{+}13] Nima Khademi Kalantari, Eli Shechtman, Connelly Barnes, Soheil Darabi, Dan B. Goldman, and Pradeep Sen. Patchbased high dynamic range video. ACM Transactions on Graphics, 32(6):202:1– 202:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwatra:2003:GTI

[KSE^{+}03] Vivek Kwatra, Arno Schödl, Irfan Essa, Greg Turk, and Aaron Bobick. Graphcut textures: image and video synthesis using graph cuts. ACM Transactions on Graphics, 22 (3):277–286, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kholgade:2014:OMS

[KSES14] Natasha Kholgade, Tomas Simon, Alexei Efros, and Yaser Sheikh. 3D object manipulation in a single photograph using stock 3D models. ACM Transactions on Graphics, 33 (4):127:1–127:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kraevoy:2003:MCC

[KSG03] Vladislav Kraevoy, Alla Sheffer, and Craig Gotsman. Matchmaker: constructing constrained texture maps. ACM Transactions on Graphics, 22(3):326–333, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kazhdan:2010:DGD

[KSH10] Michael Kazhdan, Dinoj Surendran, and Hugues Hoppe. Distributed gradient-domain processing of planar and spherical images. ACM Transactions on Graphics, 29(2): 14:1–14:11, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Karsch:2014:ASI

[KSH^{+}14] Kevin Karsch, Kalyan Sunkavalli, Sunil Hadap, Nathan Carr, Hailin Jin, Rafael Fonte, Michael Sittig, and David Forsyth. Automatic scene inference for 3D object compositing. ACM Transactions on Graphics, 33(3): 32:1–32:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kwan:2016:PAD

[KSH^{+}16] Kin Chung Kwan, Lok Tsun Sinn, Chu Han, Tien-Tsin Wong, and Chi-Wing Fu. Pyramid of arclength descriptor for generating collage of shapes. ACM Transactions on Graphics, 35(6): 229:1–229:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Korosteleva:2023:GPP

[KSH23] Maria Korosteleva and Olga Sorkine-Hornung. Garment-Code: Programming parametric sewing patterns. ACM Transactions on Graphics, 42 (6):199:1–199:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618351.

Karamouzas:2018:CSP

[KSHG18] Ioannis Karamouzas, Nick Sohre, Ran Hu, and Stephen J. Guy. Crowd space: a predictive crowd analysis technique. ACM Transactions on Graphics, 37(6):186:1– 186:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaufman:2008:SPF

[KSJP08] Danny M. Kaufman, Shinjiro Sueda, Doug L. James, and Dinesh K. Pai. Staggered projections for frictional contact in multibody systems. ACM Transactions on Graphics, 27 (5):164:1–164:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2014:IML

[KSKL14] Jongmin Kim, Yeongho Seol, Taesoo Kwon, and Jehee Lee. Interactive manipulation of large-scale crowd animation. ACM Transactions on Graphics, 33(4):83:1–83:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kuo:2023:MH

[KSL^{+}23] Grace Kuo, Florian Schiffers, Douglas Lanman, Oliver Cossairt, and Nathan Matsuda. Multisource holography. ACM Transactions on Graphics, 42 (6):203:1–203:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3618380.

Karamouzas:2017:ICO

[KSNG17] Ioannis Karamouzas, Nick Sohre, Rahul Narain, and Stephen J. Guy. Implicit crowds: optimization integrator for robust crowd simulation. ACM Transactions on Graphics, 36(4):136:1– 136:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2013:CAI

[KSP13] Johannes Kopf, Ariel Shamir, and Pieter Peers. Contentadaptive image downscaling. ACM Transactions on Graphics, 32(6):173:1– 173:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kharevych:2006:DCM

[KSS06] Liliya Kharevych, Boris Springborn, and Peter Schröder. Discrete conformal mappings via circle patterns. ACM Transactions on Graphics, 25(2): 412–438, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koyama:2015:ACD

[KSS^{+}15] Yuki Koyama, Shinjiro Sueda, Emma Steinhardt, Takeo Igarashi, Ariel Shamir, and Wojciech Matusik. Auto-Connect: computational design of 3D-printable connectors. ACM Transactions on Graphics, 34(6):231:1– 231:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kraevoy:2008:NHR

[KSSCO08] Vladislav Kraevoy, Alla Sheffer, Ariel Shamir, and Daniel Cohen-Or. Non-homogeneous resizing of complex models. ACM Transactions on Graphics, 27(5):111:1–111:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kemelmacher-Shlizerman:2011:EP

[KSSGS11] Ira Kemelmacher-Shlizerman, Eli Shechtman, Rahul Garg, and Steven M. Seitz. Exploring photobios. ACM Transactions on Graphics, 30(4): 61:1–61:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koyama:2017:SLS

[KSSI17] Yuki Koyama, Issei Sato, Daisuke Sakamoto, and Takeo Igarashi. Sequential line search for efficient visual design optimization by crowds. ACM Transactions on Graphics, 36(4):48:1–48:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kolomenkin:2008:DCS

[KST08] Michael Kolomenkin, Ilan Shimshoni, and Ayellet Tal. Demarcating curves for shape illustration. ACM Transactions on Graphics, 27(5): 157:1–157:??, December 2008.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Khungurn:2015:MRF

[KSZ^{+}15] Pramook Khungurn, Daniel Schroeder, Shuang Zhao, Kavita Bala, and Steve Marschner. Matching real fabrics with micro-appearance models. ACM Transactions on Graphics, 35(1):1:1– 1:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Katz:2003:HMD

[KT03] Sagi Katz and Ayellet Tal.

Hierarchical mesh decomposition using fuzzy clustering and cuts. ACM Transactions on Graphics, 22(3):954–961, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Katz:2007:DVP

[KTB07] Sagi Katz, Ayellet Tal, and Ronen Basri. Direct visibility of point sets. ACM Transactions on Graphics, 26(3): 24:1–24:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kailkhura:2016:SBN

[KTBV16] Bhavya Kailkhura, Jayaraman J. Thiagarajan, Peer-Timo Bremer, and Pramod K. Varshney. Stair blue noise sampling. ACM Transactions on Graphics, 35(6): 248:1–248:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2008:WTF

[KTJG08] Theodore Kim, Nils Thürey, Doug James, and Markus Gross. Wavelet turbulence for fluid simulation. ACM Transactions on Graphics, 27(3): 50:1–50:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Koller:2004:PIG

[KTL^{+}04] David Koller, Michael Turitzin, Marc Levoy, Marco Tarini, Giuseppe Croccia, Paolo Cignoni, and Roberto Scopigno. Protected interactive 3D graphics via remote rendering. ACM Transactions on Graphics, 23(3): 695–703, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaufman:2014:ANC

[KTS^{+}14] Danny M. Kaufman, Rasmus Tamstorf, Breannan Smith, Jean-Marie Aubry, and Eitan Grinspun. Adaptive nonlinearity for collisions in complex rod assemblies. ACM Transactions on Graphics, 33(4): 123:1–123:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2013:CPT

[KTT13] Theodore Kim, Jerry Tessendorf, and Nils Thürey. Closest point turbulence for liquid surfaces. ACM Transactions on Graphics, 32(2):15:1–15:13, April 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kikuuwe:2009:EBC

[KTY09] Ryo Kikuuwe, Hiroaki Tabuchi, and Motoji Yamamoto. An edge-based computationally efficient formulation of Saint Venant–Kirchhoff tetrahedral finite elements. ACM Transactions on Graphics, 28(1): 8:1–8:13, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kopf:2007:CVG

[KUDC07] Johannes Kopf, Matt Uyttendaele, Oliver Deussen, and Michael F. Cohen. Capturing and viewing gigapixel images. ACM Transactions on Graphics, 26(3):93:1–93:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kee:2021:CPD

[KUJH21] Min Hyung Kee, Kiwon Um, Wooseok Jeong, and Junghyun Han. Constrained projective dynamics: realtime simulation of deformable objects with energy-momentum conservation. ACM Transactions on Graphics, 40(4): 160:1–160:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459878.

Kang:2003:HDR

[KUWS03] Sing Bing Kang, Matthew Uyttendaele, Simon Winder, and Richard Szeliski. High dynamic range video. ACM Transactions on Graphics, 22 (3):319–325, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalaiah:2005:SGR

[KV05] Aravind Kalaiah and Amitabh Varshney. Statistical geometry representation for efficient transmission and rendering. ACM Transactions on Graphics, 24(2):348–373, April 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kondapaneni:2019:OMI

[KVG^{+}19] Ivo Kondapaneni, Petr Vevoda, Pascal Grittmann, Tomás Skrivan, Philipp Slusallek, and Jaroslav Krivánek. Optimal multiple importance sampling. ACM Transactions on Graphics, 38(4): 37:1–37:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kleiman:2015:SSE

[KvKSHCO15] Yanir Kleiman, Oliver van Kaick, Olga Sorkine-Hornung, and Daniel Cohen-Or. SHED: shape edit distance for finegrained shape similarity. ACM Transactions on Graphics, 34 (6):235:1–235:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kruger:2003:LAO

[KW03] Jens Krüger and Rüdiger Westermann. Linear algebra operators for GPU implementation of numerical algorithms. ACM Transactions on Graphics, 22(3):908–916, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kelly:2011:IAM

[KW11] Tom Kelly and Peter Wonka.

Interactive architectural modeling with procedural extrusions. ACM Transactions on Graphics, 30(2):14:1– 14:15, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kadambi:2013:CTF

[KWB^{+}13] Achuta Kadambi, Refael Whyte, Ayush Bhandari, Lee Streeter, Christopher Barsi, Adrian Dorrington, and Ramesh Raskar. Coded time of flight cameras: sparse deconvolution to address multipath interference and recover time profiles. ACM Transactions on Graphics, 32(6): 167:1–167:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Klose:2015:SBS

[KWB^{+}15] Felix Klose, Oliver Wang, Jean-Charles Bazin, Marcus Magnor, and Alexander Sorkine-Hornung. Sampling based scene-space video processing. ACM Transactions on Graphics, 34(4):67:1– 67:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2009:MHC

[KWK09] Min H. Kim, Tim Weyrich, and Jan Kautz. Modeling human color perception under extended luminance levels. ACM Transactions on Graphics, 28(3):27:1–27:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kaspar:2021:KSC

[KWL^{+}21] Alexandre Kaspar, Kui Wu, Yiyue Luo, Liane Makatura, and Wojciech Matusik. Knit sketching: from cut & sew patterns to machine-knit garments. ACM Transactions on Graphics, 40(4): 63:1–63:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459752.

Khungurn:2017:FRF

[KWN^{+}17] Pramook Khungurn, Rundong Wu, James Noeckel, Steve Marschner, and Kavita Bala. Fast rendering of fabric micro-appearance models under directional and spherical Gaussian lights. ACM Transactions on Graphics, 36(6): 232:1–232:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kalantari:2016:LBV

[KWR16] Nima Khademi Kalantari, Ting-Chun Wang, and Ravi Ramamoorthi. Learningbased view synthesis for light field cameras. ACM Transactions on Graphics, 35(6): 193:1–193:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Keller:2023:SST

[KWS^{+}23] Marilyn Keller, Keenon Werling, Soyong Shin, Scott Delp, Sergi Pujades, C. Karen Liu, and Michael J. Black. From skin to skeleton: Towards biomechanically accurate 3D digital humans. ACM Transactions on Graphics, 42(6): 253:1–253:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618381.

Krs:2017:SSV

[KYC^{+}17] Vojtech Krs, Ersin Yumer, Nathan Carr, Bedrich Benes, and Radomír Mech. Skippy: single view 3D curve interactive modeling. ACM Transactions on Graphics, 36(4): 128:1–128:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kauvar:2015:ACD

[KYS^{+}15] Isaac Kauvar, Samuel J. Yang, Liang Shi, Ian McDowall, and Gordon Wetzstein. Adaptive color display via perceptuallydriven factored spectral projection. ACM Transactions on Graphics, 34(6):165:1– 165:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2009:SWL

[KySK09] Doyub Kim, Oh young Song, and Hyeong-Seok Ko. Stretching and wiggling liquids. ACM Transactions on Graphics, 28(5):120:1–120:7, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2010:PSD

[KySK10] Doyub Kim, Oh young Song, and Hyeong-Seok Ko. A practical simulation of dispersed bubble flow. ACM Transactions on Graphics, 29(4): 70:1–70:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2008:LAI

[KYYL08] Yongjin Kim, Jingyi Yu, Xuan Yu, and Seungyong Lee. Lineart illustration of dynamic and specular surfaces. ACM Transactions on Graphics, 27 (5):156:1–156:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Knaus:2011:PPM

[KZ11] Claude Knaus and Matthias Zwicker. Progressive photon mapping: a probabilistic approach. ACM Transactions on Graphics, 30(3): 25:1–25:13, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kim:2013:SRH

[KZP^{+}13] Changil Kim, Henning Zimmer, Yael Pritch, Alexander Sorkine-Hornung, and Markus Gross. Scene reconstruction from high spatio-angular resolution light fields. ACM Transactions on Graphics, 32 (4):73:1–73:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Kadambi:2016:OIT

[KZSR16] Achuta Kadambi, Hang Zhao, Boxin Shi, and Ramesh Raskar. Occluded imaging with time-of-flight sensors. ACM Transactions on Graphics, 35(2):15:1–15:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Laine:2005:SSV

[LAA^{+}05] Samuli Laine, Timo Aila, Ulf Assarsson, Jaakko Lehtinen, and Tomas Akenine-Möller. Soft shadow volumes for ray tracing. ACM Transactions on Graphics, 24(3): 1156–1165, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2023:SPD

[LABS23] Chenxi Liu, Toshiki Aoki, Mikhail Bessmeltsev, and Alla Sheffer. StripMaker: Perception-driven learned vector sketch consolidation. ACM Transactions on Graphics, 42 (4):55:1–55:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592130.

Lehtinen:2011:TLF

[LAC^{+}11] Jaakko Lehtinen, Timo Aila, Jiawen Chen, Samuli Laine, and Frédo Durand. Temporal light field reconstruction for rendering distribution effects. ACM Transactions on Graphics, 30(4): 55:1–55:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2008:AGI

[LACS08] Wilmot Li, Maneesh Agrawala, Brian Curless, and David Salesin. Automated generation of interactive 3D exploded view diagrams. ACM Transactions on Graphics, 27(3):101:1–101:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lenaerts:2008:PFP

[LAD08] Toon Lenaerts, Bart Adams, and Philip Dutré. Porous flow in particle-based fluid simulations. ACM Transactions on Graphics, 27(3):49:1– 49:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2018:DMC

[LADL18] Tzu-Mao Li, Miika Aittala, Frédo Durand, and Jaakko Lehtinen. Differentiable Monte Carlo ray tracing through edge sampling. ACM Transactions on Graphics, 37 (6):222:1–222:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2009:RSV

[LAGP09] Hao Li, Bart Adams, Leonidas J. Guibas, and Mark Pauly. Robust single-view geometry and motion reconstruction. ACM Transactions on Graphics, 28 (5):175:1–175:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2021:LSA

[LAH^{+}21] Peizhuo Li, Kfir Aberman, Rana Hanocka, Libin Liu, Olga Sorkine-Hornung, and Baoquan Chen. Learning skeletal articulations with neural blend shapes. ACM Transactions on Graphics, 40(4):130:1–130:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3450626.3459852.

Langlois:2014:ECM

[LAJJ14] Timothy R. Langlois, Steven S. An, Kelvin K. Jin, and Doug L. James. Eigenmode compression for modal sound models. ACM Transactions on Graphics, 33(4): 40:1–40:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Laine:2011:CDS

[LAKL11] Samuli Laine, Timo Aila, Tero Karras, and Jaakko Lehtinen. Clipless dual-space bounds for faster stochastic rasterization. ACM Transactions on Graphics, 30(4):106:1–106:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lehtinen:2012:RIL

[LALD12] Jaakko Lehtinen, Timo Aila, Samuli Laine, and Frédo Durand. Reconstructing the indirect light field for global illumination. ACM Transactions on Graphics, 31(4): 51:1–51:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loos:2011:MRT

[LAM^{+}11] Bradford J. Loos, Lakulish Antani, Kenny Mitchell, Derek Nowrouzezahrai, Wojciech Jarosz, and Peter-Pike Sloan. Modular Radiance Transfer. ACM Transactions on Graphics, 30(6): 178:1–178:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lasser:1990:TRT

[Las90]  Dieter Lasser.  Two re-

marks on tau-splines. ACM Transactions on Graphics, 9(2):198–211, April 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
78960.html.

Lau:2018:SDH

[Lau18] Manfred Lau. Session details: How people look and move. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2012:S

[LAZ^{+}12] Honghua Li, Ibraheem Alhashim, Hao Zhang, Ariel Shamir, and Daniel Cohen-Or. Stackabilization. ACM Transactions on Graphics, 31 (6):158:1–158:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2022:GNM

[LAZ^{+}22] Peizhuo Li, Kfir Aberman, Zihan Zhang, Rana Hanocka, and Olga Sorkine-Hornung. GANimator: neural motion synthesis from a single sequence. ACM Transactions on Graphics, 41(4):138:1– 138:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530157.

Liang:1984:NCM

[LB84] Y.-D. Liang and B. A. Barsky. A new concept and method for line clipping. ACM Transactions on Graphics, 3(1): 1–22, January 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loop:2005:RIC

[LB05] Charles Loop and Jim Blinn.

Resolution independent curve rendering using programmable graphics hardware. ACM Transactions on Graphics, 24 (3):1000–1009, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loop:2006:RTG

[LB06] Charles Loop and Jim Blinn.

Real-time GPU rendering of piecewise algebraic surfaces. ACM Transactions on Graphics, 25(3):664–670, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2018:ISI

[LB18] Yijing Li and Jernej Barbic.

Immersion of self-intersecting solids and surfaces. ACM Transactions on Graphics, 37 (4):45:1–45:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2023:LMF

[LB23] Heng Liu and David Bommes.

Locally meshable frame fields. ACM Transactions on Graphics, 42(4):112:1–112:??, August 2023. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592457.

Lawrence:2006:IST

[LBAD^{+}06] Jason Lawrence, Aner Ben-Artzi, Christopher DeCoro, Wojciech Matusik, Hanspeter Pfister, Ravi Ramamoorthi, and Szymon Rusinkiewicz. Inverse shade trees for nonparametric material representation and editing. ACM Transactions on Graphics, 25 (3):735–745, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2002:EA

[LBB02] Sooha Park Lee, Jeremy B. Badler, and Norman I. Badler. Eyes alive. ACM Transactions on Graphics, 21(3): 637–644, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Larionov:2017:VSU

[LBB17a] Egor Larionov, Christopher Batty, and Robert Bridson. Variational Stokes: a unified pressure–viscosity solver for accurate viscous liquids. ACM Transactions on Graphics, 36(4):101:1–101:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2017:LMF

[LBB^{+}17b] Tianye Li, Timo Bolkart, Michael J. Black, Hao Li, and Javier Romero. Learning a model of facial shape and expression from 4D scans. ACM Transactions on Graphics, 36 (6):194:1–194:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lamb:2022:DLJ

[LBB22] Nikolas Lamb, Sean Banerjee, and Natasha Kholgade Banerjee. DeepJoin: Learning a joint occupancy, signed distance, and normal field function for shape repair. ACM Transactions on Graphics, 41 (6):230:1–230:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555470.

Leake:2021:MFF

[LBDA21] Mackenzie Leake, Gilbert Bernstein, Abe Davis, and Maneesh Agrawala. A mathematical foundation for foundation paper pieceable quilts. ACM Transactions on Graphics, 40(4):65:1–65:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459853.

Lu:2013:RPE

[LBDF13] Jingwan Lu, Connelly Barnes, Stephen DiVerdi, and Adam Finkelstein. RealBrush: painting with examples of physical media. ACM Transactions on Graphics, 32(4): 117:1–117:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2023:CAO

[LBHH23] Chenxi Liu, Pierre Bénard, Aaron Hertzmann, and Shayan Hoshyari. ConTesse: Accurate occluding contours for subdivision surfaces. ACM Transactions on Graphics, 42 (1):5:1–5:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3544778.

Lau:2009:MST

[LBJK09] Manfred Lau, Ziv Bar-Joseph, and James Kuffner. Modeling spatial and temporal variation in motion data. ACM Transactions on Graphics, 28(5): 171:1–171:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lyon:2016:HRH

[LBK16] Max Lyon, David Bommes, and Leif Kobbelt. HexEx: robust hexahedral mesh extraction. ACM Transactions on Graphics, 35(4): 123:1–123:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:QNMa

[LBK17a] Tiantian Liu, Sofien Bouaziz, and Ladislav Kavan. Quasi-Newton methods for realtime simulation of hyperelastic materials. ACM Transactions on Graphics, 36(3): 23:1–23:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:QNMb

[LBK17b] Tiantian Liu, Sofien Bouaziz, and Ladislav Kavan. Quasinewton methods for realtime simulation of hyperelastic materials. ACM Transactions on Graphics, 36(4): 116:1–116:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2013:FSM

[LBOK13] Tiantian Liu, Adam W.

Bargteil, James F. O’Brien, and Ladislav Kavan. Fast simulation of mass-spring systems. ACM Transactions on Graphics, 32(6):214:1– 214:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Laffont:2012:CII

[LBP^{+}12] Pierre-Yves Laffont, Adrien Bousseau, Sylvain Paris, Frédo Durand, and George Drettakis. Coherent intrinsic images from photo collections. ACM Transactions on Graphics, 31(6):202:1– 202:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Luo:2012:CPM

[LBRM12] Linjie Luo, Ilya Baran, Szymon Rusinkiewicz, and Wojciech Matusik. Chopper: partitioning models into 3Dprintable parts. ACM Transactions on Graphics, 31(6): 129:1–129:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lu:2014:DDS

[LBW^{+}14] Jingwan Lu, Connelly Barnes, Connie Wan, Paul Asente, Radomir Mech, and Adam Finkelstein. DecoBrush: drawing structured decorative patterns by example. ACM Transactions on Graphics, 33 (4):90:1–90:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lyu:2023:BVW

[LBW^{+}23] Chaoyang Lyu, Kai Bai, Yiheng Wu, Mathieu Desbrun, Changxi Zheng, and Xiaopei Liu. Building a virtual weakly-compressible wind tunnel testing facility. ACM Transactions on Graphics, 42(4):125:1– 125:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592394.

Luken:1996:CSD

[LC96] William L. Luken and Fuhua (Frank)

Cheng. Comparison of surface and derivative evaluation methods for the rendering of NURB surfaces. ACM Transactions on Graphics, 15(2): 153–178, April 1996. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lino:2015:IEC

[LC15] Christophe Lino and Marc Christie. Intuitive and efficient camera control with the toric space. ACM Transactions on Graphics, 34(4):82:1– 82:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ly:2018:IES

[LCBD^{+}18] Mickaël Ly, Romain Casati, Florence Bertails-Descoubes, Mélina Skouras, and Laurence Boissieux. Inverse elastic shell design with contact and friction. ACM Transactions on Graphics, 37(6): 201:1–201:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lyon:2019:PQF

[LCBK19] Max Lyon, Marcel Campen, David Bommes, and Leif Kobbelt. Parametrization quantization with free boundaries for trimmed quad meshing. ACM Transactions on Graphics, 38(4):51:1– 51:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:CRT

[LCC^{+}18] Lingjie Liu, Nenglun Chen, Duygu Ceylan, Christian Theobalt, Wenping Wang, and Niloy J. Mitra. Curve-Fusion: reconstructing thin structures from RGBD sequences. ACM Transactions on Graphics, 37(6):218:1– 218:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lu:2021:LSP

[LCC21] Yuanxun Lu, Jinxiang Chai, and Xun Cao. Live speech portraits: real-time photorealistic talking-head animation. ACM Transactions on Graphics, 40(6): 220:1–220:17, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480484.

Liu:2022:RFA

[LCC^{+}22] Shichen Liu, Yunxuan Cai, Haiwei Chen, Yichao Zhou, and Yajie Zhao. Rapid face asset acquisition with recurrent feature alignment. ACM Transactions on Graphics, 41 (6):214:1–214:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555509.

Lavenant:2018:DOT

[LCCS18] Hugo Lavenant, Sebastian Claici, Edward Chien, and Justin Solomon. Dynamical optimal transport on discrete surfaces. ACM Transactions on Graphics, 37(6): 250:1–250:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Luft:2006:IEU

[LCD06] Thomas Luft, Carsten Colditz, and Oliver Deussen. Image enhancement by unsharp masking the depth buffer. ACM Transactions on Graphics, 25 (3):1206–1213, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2019:CSO

[LCD^{+}19] Jongho Lee, Jenu Varghese Chacko, Bing Dai, Syed Azer Reza, Abdul Kader Sagar, Kevin W. Eliceiri, Andreas Velten, and Mohit Gupta. Coding scheme optimization for fast fluorescence lifetime imaging. ACM Transactions on Graphics, 38(3): 26:1–26:??, June 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3325136.

Li:2020:FST

[LCD^{+}20a] Wei Li, Yixin Chen, Mathieu Desbrun, Changxi Zheng, and Xiaopei Liu. Fast and scalable turbulent flow simulation with two-way coupling. ACM Transactions on Graphics, 39 (4):47:1–47:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392400.

Lu:2020:LNR

[LCD^{+}20b] Erika Lu, Forrester Cole, Tali Dekel, Weidi Xie, Andrew Zisserman, David Salesin, William T. Freeman, and Michael Rubinstein. Layered neural rendering for retiming people in video. ACM Transactions on Graphics, 39(6): 256:1–256:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417760.

Lipman:2010:SFE

[LCDF10] Yaron Lipman, Xiaobai Chen, Ingrid Daubechies, and Thomas Funkhouser. Symmetry factored embedding and distance. ACM Transactions on Graphics, 29(4):103:1– 103:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liao:2021:RTL

[LCH^{+}21] Wentao Liao, Renjie Chen, Yuchen Hua, Ligang Liu, and Ofir Weber. Realtime locally injective volumetric deformation. ACM Transactions on Graphics, 40 (4):74:1–74:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459794.

Liu:2014:CCS

[LCK^{+}14] Tianqiang Liu, Siddhartha Chaudhuri, Vladimir G. Kim, Qixing Huang, Niloy J. Mitra, and Thomas Funkhouser. Creating consistent scene graphs using a probabilistic grammar. ACM Transactions on Graphics, 33(6):211:1– 211:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2022:IAE

[LCK22] Huancheng Lin, Floyd M.

Chitalu, and Taku Komura. Isotropic ARAP energy using Cauchy–Green invariants. ACM Transactions on Graphics, 41(6):275:1– 275:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555507.

Lee:2006:MPB

[LCL06] Kang Hoon Lee, Myung Geol Choi, and Jehee Lee. Motion patches: building blocks for virtual environments annotated with motion data. ACM Transactions on Graphics, 25 (3):898–906, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:IBR

[LCL^{+}17] Lingjie Liu, Duygu Ceylan, Cheng Lin, Wenping Wang, and Niloy J. Mitra. Image-based reconstruction of wire art. ACM Transactions on Graphics, 36(4): 63:1–63:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2022:DSB

[LCL^{+}22] Feng-Lin Liu, Shu-Yu Chen, Yu-Kun Lai, Chunpeng Li, Yue-Ren Jiang, Hongbo Fu, and Lin Gao. DeepFace-VideoEditing: sketch-based deep editing of face videos. ACM Transactions on Graphics, 41(4):167:1–167:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530056.

Li:2023:EBM

[LCL^{+}23] Weiyu Li, Xuelin Chen, Peizhuo Li, Olga Sorkine-Hornung, and Baoquan Chen. Example-based motion synthesis via generative motion matching. ACM Transactions on Graphics, 42(4): 94:1–94:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592395.

Leyvand:2008:DDE

[LCODL08] Tommer Leyvand, Daniel Cohen-Or, Gideon Dror, and Dani Lischinski. Data-driven enhancement of facial attractiveness. ACM Transactions on Graphics, 27(3): 38:1–38:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2007:PFP

[LCOLTE07] Yaron Lipman, Daniel Cohen-Or, David Levin, and Hillel Tal-Ezer. Parameterizationfree projection for geometry reconstruction. ACM Transactions on Graphics, 26(3): 22:1–22:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2007:VSP

[LCORL07] Yaron Lipman, Daniel Cohen-Or, Gal Ran, and David Levin. Volume and shape preservation via moving frame manipulation. ACM Transactions on Graphics, 26(1): 5:1–5:14, January 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2011:SPR

[LCOZ^{+}11] Jinjie Lin, Daniel Cohen-Or, Hao Zhang, Cheng Liang, Andrei Sharf, Oliver Deussen, and Baoquan Chen. Structure-preserving retargeting of irregular 3D architecture. ACM Transactions on Graphics, 30(6):183:1– 183:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2002:ICA

[LCR^{+}02] Jehee Lee, Jinxiang Chai, Paul S. A. Reitsma, Jessica K. Hodgins, and Nancy S. Pollard. Interactive control of avatars animated with human motion data. ACM Transactions on Graphics, 21(3): 491–500, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2019:PBS

[LCT19] Shiguang Liu, Haonan Cheng, and Yiying Tong. Physicallybased statistical simulation of rain sound. ACM Transactions on Graphics, 38(4): 123:1–123:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:NMN

[LCT23] Yue Li, Stelian Coros, and Bernhard Thomaszewski. Neural metamaterial networks for nonlinear material design. ACM Transactions on Graphics, 42(6):186:1– 186:??, December 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618325.

Ledda:2005:ETM

[LCTS05] Patrick Ledda, Alan Chalmers, Tom Troscianko, and Helge Seetzen. Evaluation of tone mapping operators using a High Dynamic Range display. ACM Transactions on Graphics, 24(3):640–648, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levoy:2004:SAC

[LCV^{+}04] Marc Levoy, Billy Chen, Vaibhav Vaish, Mark Horowitz, Ian McDowall, and Mark Bolas. Synthetic aperture confocal imaging. ACM Transactions on Graphics, 23(3): 825–834, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lv:2016:DDI

[LCX16] Xiaolei Lv, Jinxiang Chai, and Shihong Xia. Datadriven inverse dynamics for human motion. ACM Transactions on Graphics, 35(6): 163:1–163:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2021:APP

[LCX^{+}21] Yilin Liu, Ruiqi Cui, Ke Xie, Minglun Gong, and Hui Huang. Aerial path planning for online real-time exploration and offline high-quality reconstruction of large-scale urban scenes. ACM Transactions on Graphics, 40(6): 226:1–226:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480491.

Lau:2009:FPI

[LCXS09] Manfred Lau, Jinxiang Chai, Ying-Qing Xu, and Heung-Yeung Shum. Face poser: Interactive modeling of 3D facial expressions using facial priors. ACM Transactions on Graphics, 29(1):3:1–3:17, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loop:1989:MGB

[LD89] Charles T. Loop and Tony D. DeRose. A multisided generalization of Bézier surfaces. ACM Transactions on Graphics, 8(3):204–234, July 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77059.html.

Lagae:2005:POD

[LD05] Ares Lagae and Philip Dutré.

A procedural object distribution function. ACM Transactions on Graphics, 24(4): 1442–1461, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lagae:2006:AWT

[LD06] Ares Lagae and Philip Dutré.

An alternative for Wang tiles: colored edges versus colored corners. ACM Transactions on Graphics, 25(4):1442– 1459, October 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lagae:2011:FSG

[LD11] Ares Lagae and George Drettakis. Filtering solid Gabor noise. ACM Transactions on Graphics, 30(4): 51:1–51:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Le:2012:SSD

[LD12] Binh Huy Le and Zhigang Deng. Smooth skinning decomposition with rigid bones. ACM Transactions on Graphics, 31(6):199:1– 199:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Le:2013:TLS

[LD13] Binh Huy Le and Zhigang Deng. Two-layer sparse compression of dense-weight blend skinning. ACM Transactions on Graphics, 32(4): 124:1–124:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Le:2014:RAS

[LD14] Binh Huy Le and Zhigang Deng. Robust and accurate skeletal rigging from mesh sequences. ACM Transactions on Graphics, 33(4): 84:1–84:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Leimkuhler:2021:FFV

[LD21] Thomas Leimkühler and George Drettakis. FreeStyle-GAN: free-view editable portrait rendering with the camera manifold. ACM Transactions on Graphics, 40(6): 224:1–224:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480538.

Li:2023:FSC

[LD23] Wei Li and Mathieu Desbrun. Fluid-solid coupling in kinetic two-phase flow simulation. ACM Transactions on Graphics, 42(4): 123:1–123:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592138.

Le:2023:CGC

[LDD^{+}23] Nhat Le, Tuong Do, Khoa Do, Hien Nguyen, Erman Tjiputra, Quang D. Tran, and Anh Nguyen. Controllable group choreography using contrastive diffusion. ACM Transactions on Graphics, 42 (6):224:1–224:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618356.

Lessig:2014:CTS

[LDF14] Christian Lessig, Mathieu Desbrun, and Eugene Fiume. A constructive theory of sampling for image synthesis using reproducing kernel bases. ACM Transactions on Graphics, 33(4):55:1–55:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2016:SST

[LDHM16] Albert Julius Liu, Zhao Dong, Milos Hasan, and Steve Marschner. Simulating the structure and texture of solid wood. ACM Transactions on Graphics, 35(6):170:1– 170:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lazar:2018:ROT

[LDK^{+}18] Roee Lazar, Nadav Dym, Yam Kushinsky, Zhiyang Huang, Tao Ju, and Yaron Lipman. Robust optimization for topological surface reconstruction. ACM Transactions on Graphics, 37(4):46:1–46:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2018:IFC

[LDN^{+}18] Jie Li, Gilles Daviet, Rahul Narain, Florence Bertails-Descoubes, Matthew Overby, George E. Brown, and Laurence Boissieux. An implicit frictional contact solver for adaptive cloth simulation. ACM Transactions on Graphics, 37(4):52:1–52:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:1984:AFE

[LdPS84] Y. T. Lee, A. de Pennington, and N. K. Shaw. Automatic finite-element mesh generation from geometric models — A point-based approach. ACM Transactions on Graphics, 3(4):287–311, October 1984. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lan:2013:BSA

[LDPT13] Yanxiang Lan, Yue Dong, Fabio Pellacini, and Xin Tong. Bi-scale appearance fabrication. ACM Transactions on Graphics, 32(4): 145:1–145:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2017:MSA

[LDPT17] Xiao Li, Yue Dong, Pieter Peers, and Xin Tong. Modeling surface appearance from a single photograph using selfaugmented convolutional neural networks. ACM Transactions on Graphics, 36(4): 45:1–45:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Littlewood:2002:POF

[LDS02] D. J. Littlewood, P. A.

Drakopoulos, and G. Subbarayan. Pareto-optimal formulations for cost versus colorimetric accuracy trade-offs in printer color management. ACM Transactions on Graphics, 21(2):132–175, April 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2003:PEC

[LDS03] Haeyoung Lee, Mathieu Desbrun, and Peter Schröder. Progressive encoding of complex isosurfaces. ACM Transactions on Graphics, 22(3): 471–476, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2011:MGM

[LDS^{+}11] Chuan Li, Oliver Deussen, Yi-Zhe Song, Phil Willis, and Peter Hall. Modeling and generating moving trees from video. ACM Transactions on Graphics, 30(6): 127:1–127:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lau:2016:TMS

[LDS^{+}16] Manfred Lau, Kapil Dev, Weiqi Shi, Julie Dorsey, and Holly Rushmeier. Tactile mesh saliency. ACM Transactions on Graphics, 35(4): 52:1–52:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2022:FTO

[LDS^{+}22] Yifei Li, Tao Du, Sangeetha Grama Srinivasan, Kui Wu, Bo Zhu, Eftychios Sifakis, and Wojciech Matusik. Fluidic topology optimization with an anisotropic mixture model. ACM Transactions on Graphics, 41(6):239:1–239:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555429.

Leake:2017:CVE

[LDTA17] Mackenzie Leake, Abe Davis, Anh Truong, and Maneesh Agrawala. Computational video editing for dialoguedriven scenes. ACM Transactions on Graphics, 36(4): 130:1–130:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lounsbery:1997:MAS

[LDW97] Michael Lounsbery, Tony D.

DeRose, and Joe Warren. Multiresolution analysis for surfaces of arbitrary topological type. ACM Transactions on Graphics, 16(1): 34–73, January 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1997-16-1/
p34-lounsbery/.

Li:2023:DDC

[LDW^{+}23] Yifei Li, Tao Du, Kui Wu, Jie Xu, and Wojciech Matusik. DiffCloth: Differentiable cloth simulation with dry frictional contact. ACM Transactions on Graphics, 42 (1):2:1–2:??, February 2023.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3527660.

Lee:2005:FBM

[Lee05] Sang Hun Lee. Featurebased multiresolution modeling of solids. ACM Transactions on Graphics, 24(4): 1417–1441, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2018:SDL

[Lee18]  Seungyong Lee.  Session

details: Low-level imaging. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lehtinen:2007:FPC

[Leh07] Jaakko Lehtinen. A framework for precomputed and captured light transport. ACM Transactions on Graphics, 26(4):13:1–13:22, October 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lalonde:2009:WCA

[LEN09] Jean-François Lalonde, Alexei A. Efros, and Srinivasa G. Narasimhan. Webcam clip art: appearance and illuminant transfer from timelapse sequences. ACM Transactions on Graphics, 28(5):

131:1–131:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lecouat:2022:HDR

[LEPM22] Bruno Lecouat, Thomas Eboli, Jean Ponce, and Julien Mairal. High dynamic range and super-resolution from raw image bursts. ACM Transactions on Graphics, 41(4):38:1–38:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530180.

Lu:2007:VIU

[LEQ^{+}07] Aidong Lu, David S. Ebert, Wei Qiao, Martin Kraus, and Benjamin Mora. Volume illustration using Wang Cubes. ACM Transactions on Graphics, 26(2):11:1–11:??, June 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2009:DFR

[LES09] Sungkil Lee, Elmar Eisemann, and Hans-Peter Seidel. Depthof-field rendering with multiview synthesis. ACM Transactions on Graphics, 28(5): 134:1–134:6, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2010:RTL

[LES10] Sungkil Lee, Elmar Eisemann, and Hans-Peter Seidel. Realtime lens blur effects and focus control. ACM Transactions on Graphics, 29(4): 65:1–65:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lessig:2020:LFS

[Les20]  Christian Lessig.  Local

Fourier slice photography. ACM Transactions on Graphics, 39(3):24:1–24:16, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3339307.

Levy:1984:VSG

[Lev84] Henry M. Levy. VAXstation: a general-purpose raster graphics architecture. ACM Transactions on Graphics, 3 (1):70–83, January 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levoy:1990:ERT

[Lev90] Marc Levoy. Efficient ray tracing of volume data. ACM Transactions on Graphics, 9 (3):245–261, July 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
78965.html.

Levy:2003:DDE

[Lév03] Bruno Lévy. Dual domain extrapolation. ACM Transactions on Graphics, 22(3): 364–369, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levin:2006:MSS

[Lev06] Adi Levin. Modified subdivision surfaces with continuous curvature. ACM Transactions on Graphics, 25(3): 1035–1040, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levi:2021:DSP

[Lev21] Zohar Levi. Direct seamless parametrization. ACM Transactions on Graphics, 40 (1):6:1–6:14, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3439828.

Levi:2023:SPC

[Lev23]  Zohar Levi.  Seamless

parametrization with cone and partial loop control. ACM Transactions on Graphics, 42(5):164:1–164:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3600087.

Lewis:1987:GSS

[Lew87]  J. P. Lewis.  General-

ized stochastic subdivision. ACM Transactions on Graphics, 6(3):167–190, July 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
35069.html.

Lamorlette:2002:SMF

[LF02] Arnauld Lamorlette and Nick Foster. Structural modeling of flames for a production environment. ACM Transactions on Graphics, 21(3): 729–735, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lessig:2008:SOS

[LF08] Christian Lessig and Eugene Fiume. SOHO: Orthogonal and symmetric Haar wavelets on the sphere. ACM Transactions on Graphics, 27(1): 4:1–4:11, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2009:MVS

[LF09] Yaron Lipman and Thomas Funkhouser. Möbius voting for surface correspondence. ACM Transactions on Graphics, 28(3):72:1–72:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lukac:2013:PFT

[LFB^{+}13] Michal Lukác, Jakub Fiser, Jean-Charles Bazin, Ondrej Jamriska, Alexander Sorkine-Hornung, and Daniel Sýkora. Painting by feature: texture boundaries for example-based image creation. ACM Transactions on Graphics, 32(4): 116:1–116:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levin:2007:IDC

[LFDF07] Anat Levin, Rob Fergus, Frédo Durand, and William T. Freeman. Image and depth from a conventional camera with a coded aperture. ACM Transactions on Graphics, 26 (3):70:1–70:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liao:2015:FCS

[LFH15] Jing Liao, Mark Finch, and Hugues Hoppe. Fast computation of seamless video loops. ACM Transactions on Graphics, 34(6):197:1– 197:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:SSE

[LFJG17] Songrun Liu, Zachary Ferguson, Alec Jacobson, and Yotam Gingold. Seamless: seam erasure and seam-aware decoupling of shape from mesh resolution. ACM Transactions on Graphics, 36(6): 216:1–216:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lo:2009:PP

[LFL09] Kui-Yip Lo, Chi-Wing Fu, and Hongwei Li. 3D polyomino puzzle. ACM Transactions on Graphics, 28(5): 157:1–157:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lakshmipathy:2023:CEA

[LFL^{+}23] Arjun Sriram Lakshmipathy, Nicole Feng, Yu Xi Lee, Moshe Mahler, and Nancy Pollard. Contact Edit: Artist tools for intuitive modeling of handobject interactions. ACM Transactions on Graphics, 42 (4):45:1–45:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592117.

Li:2022:CSI

[LFO^{+}22] Mo Li, Qing Fang, Wenqing Ouyang, Ligang Liu, and Xiao-Ming Fu. Computing sparse integer-constrained cones for conformal parameterizations. ACM Transactions on Graphics, 41(4): 58:1–58:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530118.

Larionov:2021:FCS

[LFP21] Egor Larionov, Ye Fan, and Dinesh K. Pai. Frictional contact on smooth elastic solids. ACM Transactions on Graphics, 40(2):15:1–15:17, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3446663.

Li:2016:GNU

[LFS16] Xin Li, G. Thomas Finnigan, and Thomas W. Sederberg. G^{1} non-uniform Catmull– Clark surfaces. ACM Transactions on Graphics, 35(4): 135:1–135:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2020:IPC

[LFS^{+}20] Minchen Li, Z Ac Hary Ferguson, Teseo Schneider, Timothy Langlois, Denis Zorin, Daniele Panozzo, Chenfanfu Jiang, and Danny M. Kaufman. Incremental potential contact: intersectionand inversion-free, largedeformation dynamics. ACM Transactions on Graphics, 39 (4):49:1–49:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.

                                        acm.org/doi/abs/10.1145/
                                        3386569.3392425.

Limpaecher:2013:RTD

[LFTC13] Alex Limpaecher, Nicolas Feltman, Adrien Treuille, and Michael Cohen. Real-time drawing assistance through crowdsourcing. ACM Transactions on Graphics, 32(4): 54:1–54:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lischinski:2006:ILA

[LFUS06] Dani Lischinski, Zeev Farbman, Matt Uyttendaele, and Richard Szeliski. Interactive local adjustment of tonal values. ACM Transactions on Graphics, 25(3):646–653, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:CID

[LFXH17] Yong-Jin Liu, Dian Fan, Chun-Xu Xu, and Ying He. Constructing intrinsic Delaunay triangulations from the dual of geodesic Voronoi diagrams. ACM Transactions on Graphics, 36(2): 15:1–15:??, April 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2019:ARB

[LFY^{+}19] Hao-Yu Liu, Xiao-Ming Fu, Chunyang Ye, Shuangming Chai, and Ligang Liu. Atlas refinement with bounded packing efficiency. ACM Transactions on Graphics, 38 (4):33:1–33:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:IAT

[LFZ15] Dingzeyu Li, Yun Fei, and Changxi Zheng. Interactive acoustic transfer approximation for modal sound. ACM Transactions on Graphics, 35 (1):2:1–2:??, December 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lira:2018:FEW

[LFZ18] Wallace Lira, Chi-Wing Fu, and Hao Zhang. Fabricable Eulerian wires for 3D shape abstraction. ACM Transactions on Graphics, 37(6): 240:1–240:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:ECS

[LFZ^{+}23] Mo Li, Qing Fang, Zheng Zhang, Ligang Liu, and Xiao-Ming Fu. Efficient cone singularity construction for conformal parameterizations. ACM Transactions on Graphics, 42(6): 235:1–235:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618407.

Li:2018:DPI

[LGA^{+}18] Tzu-Mao Li, Michaël Gharbi, Andrew Adams, Frédo Durand, and Jonathan Ragan-Kelley. Differentiable programming for image processing and deep learning in halide. ACM Transactions on Graphics, 37(4):139:1– 139:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lawrence:2021:PSH

[LGA^{+}21] Jason Lawrence, Danb Goldman, Supreeth Achar, Gregory Major Blascovich, Joseph G. Desloge, Tommy Fortes, Eric M. Gomez, Sascha Häberling, Hugues Hoppe, Andy Huibers, Claude Knaus, Brian Kuschak, Ricardo Martin-Brualla, Harris Nover, Andrew Ian Russell, Steven M. Seitz, and Kevin Tong. Project starline: a highfidelity telepresence system. ACM Transactions on Graphics, 40(6):242:1–242:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480490.

Liu:2021:TND

[LGB^{+}21] Yanchao Liu, Jianwei Guo, Bedrich Benes, Oliver Deussen, Xiaopeng Zhang, and Hui Huang. TreePartNet: neural decomposition of point clouds for 3D tree reconstruction. ACM Transactions on Graphics, 40(6): 232:1–232:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480486.

Liu:2023:SSU

[LGC^{+}23] Hsueh-Ti Derek Liu, Mark Gillespie, Benjamin Chislett, Nicholas Sharp, Alec Jacobson, and Keenan Crane. Surface simplification using intrinsic error metrics. ACM Transactions on Graphics, 42(4):118:1–118:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592403.

Losasso:2004:SWS

[LGF04] Frank Losasso, Frédéric Gibou, and Ron Fedkiw. Simulating water and smoke with an octree data structure. ACM Transactions on Graphics, 23(3):457–462, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lu:2007:CAT

[LGG^{+}07] Jianye Lu, Athinodoros S.

Georghiades, Andreas Glaser, Hongzhi Wu, Li-Yi Wei, Baining Guo, Julie Dorsey, and Holly Rushmeier. Contextaware textures. ACM Transactions on Graphics, 26(1): 3:1–3:22, January 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:OSC

[LGHL23] Changhao Li, Junfu Guo, Ruizhen Hu, and Ligang Liu. Online scene CAD recomposition via autonomous scanning. ACM Transactions on Graphics, 42(6): 251:1–251:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618339.

Liu:2009:CPW

[LGJA09] Feng Liu, Michael Gleicher, Hailin Jin, and Aseem Agarwala. Content-preserving warps for 3D video stabilization. ACM Transactions on Graphics, 28(3):44:1– 44:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lien:2016:SUG

[LGK^{+}16] Jaime Lien, Nicholas Gillian, M. Emre Karagozler, Patrick Amihood, Carsten Schwesig, Erik Olson, Hakim Raja, and Ivan Poupyrev. Soli: ubiquitous gesture sensing with millimeter wave radar. ACM Transactions on Graphics, 35 (4):142:1–142:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2019:DOT

[LGL^{+}19] Minchen Li, Ming Gao, Timothy Langlois, Chenfanfu Jiang, and Danny M. Kaufman. Decomposed optimization time integrator for largestep elastodynamics. ACM Transactions on Graphics, 38 (4):70:1–70:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lloyd:2008:LPS

[LGQ^{+}08] D. Brandon Lloyd, Naga K.

Govindaraju, Cory Quammen, Steven E. Molnar, and Dinesh Manocha. Logarithmic perspective shadow maps. ACM Transactions on Graphics, 27(4):106:1–106:32, October 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2011:SVS

[LGW^{+}11] Feng Liu, Michael Gleicher, Jue Wang, Hailin Jin, and Aseem Agarwala. Subspace video stabilization. ACM Transactions on Graphics, 30 (1):4:1–4:10, January 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levin:2013:FBH

[LGX^{+}13] Anat Levin, Daniel Glasner, Ying Xiong, Frédo Durand, William Freeman, Wojciech Matusik, and Todd Zickler. Fabricating BRDFs at high spatial resolution using wave optics. ACM Transactions on Graphics, 32(4): 144:1–144:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2013:SDR

[LGZ^{+}13] Hui Lin, Jizhou Gao, Yu Zhou, Guiliang Lu, Mao Ye, Chenxi Zhang, Ligang Liu, and Ruigang Yang. Semantic decomposition and reconstruction of residential scenes from LiDAR data. ACM Transactions on Graphics, 32(4): 66:1–66:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Losasso:2004:GCT

[LH04] Frank Losasso and Hugues Hoppe. Geometry clipmaps: terrain rendering using nested regular grids. ACM Transactions on Graphics, 23(3): 769–776, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lefebvre:2005:PCT

[LH05] Sylvain Lefebvre and Hugues Hoppe. Parallel controllable texture synthesis. ACM Transactions on Graphics, 24 (3):777–786, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lefebvre:2006:AST

[LH06a] Sylvain Lefebvre and Hugues Hoppe. Appearance-space texture synthesis. ACM Transactions on Graphics, 25 (3):541–548, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lefebvre:2006:PSH

[LH06b] Sylvain Lefebvre and Hugues Hoppe. Perfect spatial hashing. ACM Transactions on Graphics, 25(3):579–588, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Le:2016:RTS

[LH16] Binh Huy Le and Jessica K.

Hodgins. Real-time skeletal skinning with optimized centers of rotation. ACM Transactions on Graphics, 35(4): 37:1–37:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:LSCa

[LH17a] Libin Liu and Jessica Hodgins. Learning to schedule control fragments for physicsbased characters using deep Q-learning. ACM Transactions on Graphics, 36(3):

29:1–29:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2017:LSCb

[LH17b] Libin Liu and Jessica Hodgins. Learning to schedule control fragments for physicsbased characters using deep Q-learning. ACM Transactions on Graphics, 36(4): 42:1–42:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:LBD

[LH18] Libin Liu and Jessica Hodgins. Learning basketball dribbling skills using trajectory optimization and deep reinforcement learning. ACM Transactions on Graphics, 37(4):142:1–142:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:FF

[LHAZ15] Honghua Li, Ruizhen Hu, Ibraheem Alhashim, and Hao Zhang. Foldabilizing furniture. ACM Transactions on Graphics, 34(4):90:1– 90:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2014:STE

[LHdG^{+}14] Siwang Li, Jin Huang, Fernando de Goes, Xiaogang Jin, Hujun Bao, and Mathieu Desbrun. Space-time editing of elastic motion through material optimization and reduction. ACM Transactions on Graphics, 33(4): 108:1–108:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lalonde:2007:PCA

[LHE^{+}07] Jean-François Lalonde, Derek Hoiem, Alexei A. Efros, Carsten Rother, John Winn, and Antonio Criminisi. Photo clip art. ACM Transactions on Graphics, 26(3): 3:1–3:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levin:2009:FAC

[LHG^{+}09] Anat Levin, Samuel W. Hasinoff, Paul Green, Frédo Durand, and William T. Freeman. 4D frequency analysis of computational cameras for depth of field extension. ACM Transactions on Graphics, 28(3):97:1–97:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2023:EHE

[LHH^{+}23] Zhengzhe Liu, Jingyu Hu, Ka-Hei Hui, Xiaojuan Qi, Daniel Cohen-Or, and Chi-Wing Fu. EXIM: a hybrid explicit-implicit representation for text-guided 3D shape generation. ACM Transactions on Graphics, 42(6): 228:1–228:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618312.

Ling:2014:SQF

[LHJ^{+}14] Ruotian Ling, Jin Huang, Bert Jüttler, Feng Sun, Hujun Bao, and Wenping Wang. Spectral quadrangulation with feature curve alignment and element size control. ACM Transactions on Graphics, 34(1):11:1–11:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Laine:2020:MPH

[LHK^{+}20] Samuli Laine, Janne Hellsten, Tero Karras, Yeongho Seol, Jaakko Lehtinen, and Timo Aila. Modular primitives for high-performance differentiable rendering. ACM Transactions on Graphics, 39(6): 194:1–194:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417861.

Lanman:2010:CAP

[LHKR10] Douglas Lanman, Matthew Hirsch, Yunhee Kim, and Ramesh Raskar. Contentadaptive parallax barriers: optimizing dual-layer 3D displays using low-rank light field factorization. ACM Transactions on Graphics, 29(6): 163:1–163:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lefebvre:2010:ESA

[LHL10] Sylvain Lefebvre, Samuel Hornus, and Anass Lasram. Byexample synthesis of architectural textures. ACM Transactions on Graphics, 29(4): 84:1–84:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2015:SCF

[LHLF15] Tianqiang Liu, Aaron Hertzmann, Wilmot Li, and Thomas Funkhouser. Style compatibility for 3D furniture models. ACM Transactions on Graphics, 34(4): 85:1–85:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2010:RTC

[LHLK10] Fuchang Liu, Takahiro Harada, Youngeun Lee, and Young J. Kim. Real-time collision culling of a million bodies on graphics processing units. ACM Transactions on Graphics, 29(6):154:1– 154:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2021:SSA

[LHLY21] Changyang Li, Haikun Huang, Jyh-Ming Lien, and Lap-Fai Yu. Synthesizing sceneaware virtual reality teleport graphs. ACM Transactions on Graphics, 40(6): 229:1–229:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480478.

Lai:2009:ATP

[LHM09] Yu-Kun Lai, Shi-Min Hu, and Ralph R. Martin. Automatic and topology-preserving gradient mesh generation for image vectorization. ACM Transactions on Graphics, 28 (3):85:1–85:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liang:2018:VFA

[LHM^{+}18] Shu Liang, Xiufeng Huang, Xianyu Meng, Kunyao Chen, Linda G. Shapiro, and Ira Kemelmacher-Shlizerman. Video to fully automatic 3D hair model. ACM Transactions on Graphics, 37(6): 206:1–206:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2005:LPB

[LHP05] C. Karen Liu, Aaron Hertzmann, and Zoran Popović.

Learning physics-based motion style with nonlinear inverse optimization. ACM Transactions on Graphics, 24 (3):1071–1081, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2021:NAN

[LHR^{+}21] Lingjie Liu, Marc Habermann, Viktor Rudnev, Kripasindhu Sarkar, Jiatao Gu, and Christian Theobalt. Neural actor: neural free-view synthesis of human actors with pose control. ACM Transactions on Graphics, 40 (6):219:1–219:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480528.

Loi:2017:PAEa

[LHVT17a] Hugo Loi, Thomas Hurtut, Romain Vergne, and Joelle Thollot. Programmable 2D arrangements for element texture design. ACM Transactions on Graphics, 36(3): 27:1–27:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loi:2017:PAEb

[LHVT17b] Hugo Loi, Thomas Hurtut, Romain Vergne, and Joëlle Thollot. Programmable 2D arrangements for element texture design. ACM Transactions on Graphics, 36(4): 105:1–105:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lang:2010:NDM

[LHW^{+}10] Manuel Lang, Alexander Hornung, Oliver Wang, Steven Poulakos, Aljoscha Smolic, and Markus Gross. Nonlinear disparity mapping for stereoscopic 3D. ACM Transactions on Graphics, 29(4): 75:1–75:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2016:EGP

[LHZ16] Rui Li, Qiming Hou, and Kun Zhou. Efficient GPU path rendering using scanline rasterization. ACM Transactions on Graphics, 35(6): 228:1–228:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:NBT

[LHZ^{+}18] Haixiang Liu, Yuanming Hu, Bo Zhu, Wojciech Matusik, and Eftychios Sifakis. Narrow-band topology optimization on a sparsely populated grid. ACM Transactions on Graphics, 37(6): 251:1–251:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2021:KGE

[LHZ^{+}21] Zishun Liu, Xingjian Han, Yuchen Zhang, Xiangjia Chen, Yu-Kun Lai, Eugeni L. Doubrovski, Emily Whiting, and Charlie C. L. Wang. Knitting 4D garments with elasticity controlled for body motion. ACM Transactions on Graphics, 40 (4):62:1–62:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459868.

Li:2018:SDF

[Li18]  Hao Li.  Session details:

Faces, faces, faces. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2012:BDM

[Lip12] Yaron Lipman. Bounded distortion mapping spaces for triangular meshes. ACM Transactions on Graphics, 31(4): 108:1–108:13, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2018:SDM

[Lip18] Yaron Lipman. Session details: Mapping + transport. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2009:DMG

[Liu09] C. Karen Liu. Dextrous manipulation from a grasping pose. ACM Transactions on Graphics, 28(3):59:1– 59:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:SDA

[Liu18] Karen Liu. Session details: Aerial propagation. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Larsson:2022:PTS

[LIY^{+}22] Maria Larsson, Takashi Ijiri, Hironori Yoshida, Johannes A. J. Huber, Magnus Fredriksson, Olof Broman, and Takeo Igarashi. Procedural texturing of solid wood with knots. ACM Transactions on Graphics, 41(4):45:1–45:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530081.

Langlois:2014:IFA

[LJ14] Timothy R. Langlois and Doug L. James. Inverse-Foley animation: synchronizing rigid-body motions to sound. ACM Transactions on Graphics, 33(4): 41:1–41:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ly:2020:PDD

[LJBBD20] MickaÃLy, Jean Jouve, Laurence Boissieux, and Florence Bertails-Descoubes. Projective dynamics with dry frictional contact. ACM Transactions on Graphics, 39(4): 57:1–57:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392396.

Liu:2014:SCB

[LJG14] Songrun Liu, Alec Jacobson, and Yotam Gingold. Skinning cubic Bézier splines and Catmull–Clark subdivision surfaces. ACM Transactions on Graphics, 33(6): 190:1–190:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2011:GSV

[LJGH11] Xian-Ying Li, Tao Ju, Yan Gu, and Shi-Min Hu. A geometric study of v-style popups: theories and algorithms. ACM Transactions on Graphics, 30(4):98:1–98:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2013:CMV

[LJH13a] Xian-Ying Li, Tao Ju, and Shi-Min Hu. Cubic mean value coordinates. ACM Transactions on Graphics, 32 (4):126:1–126:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liao:2013:AVL

[LJH13b] Zicheng Liao, Neel Joshi, and Hugues Hoppe. Automated video looping with progressive dynamism. ACM Transactions on Graphics, 32(4): 77:1–77:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2018:PMS

[LJJ^{+}18] Joo Ho Lee, Adrian Jarabo, Daniel S. Jeon, Diego Gutierrez, and Min H. Kim. Practical multiple scattering for rough surfaces. ACM Transactions on Graphics, 37(6): 275:1–275:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2023:ADS

[LJL23] Seunghwan Lee, Yifeng Jiang, and C. Karen Liu. Anatomically detailed simulation of human torso. ACM Transactions on Graphics, 42(4): 40:1–40:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592425.

Lee:2016:ALF

[LJM^{+}16] Seungjae Lee, Changwon Jang, Seokil Moon, Jaebum Cho, and Byoungho Lee. Additive light field displays: realization of augmented reality with holographic optical elements. ACM Transactions on Graphics, 35(4): 60:1–60:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2019:SCG

[LJO19] Hsueh-Ti Derek Liu, Alec Jacobson, and Maks Ovsjanikov. Spectral coarsening of geometric operators. ACM Transactions on Graphics, 38 (4):105:1–105:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ladicky:2015:DDF

[LJS^{+}15] Lúbor Ladický, SoHyeon Jeong, Barbara Solenthaler, Marc Pollefeys, and Markus Gross. Data-driven fluid simulations using regression forests. ACM Transactions on Graphics, 34(6):199:1– 199:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:EEM

[LJZ^{+}23] Chenghong Li, Leyang Jin, Yujian Zheng, Yizhou Yu, and Xiaoguang Han. EMS: 3D eyebrow modeling from singleview images. ACM Transactions on Graphics, 42(6): 269:1–269:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618323.

Latta:2002:HFB

[LK02] Lutz Latta and Andreas Kolb. Homomorphic factorization of BRDF-based lighting computation. ACM Transactions on Graphics, 21(3):509–516, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2020:SLF

[LK20] Qinbo Li and Nima Khademi Kalantari. Synthesizing light field from a single image with variable MPI and two network fusion. ACM Transactions on Graphics, 39(6): 229:1–229:10, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417785.

Lee:2022:RDA

[LKB22a] Joon Hyub Lee, Hanbit Kim, and Seok-Hyung Bae. Rapid design of articulated objects.

ACM Transactions on Graphics, 41(4):89:1–89:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530092.

Lin:2022:GRI

[LKB^{+}22b] Daqi Lin, Markus Kettunen, Benedikt Bitterli, Jacopo Pantaleoni, Cem Yuksel, and Chris Wyman. Generalized resampled importance sampling: foundations of ReSTIR. ACM Transactions on Graphics, 41(4):75:1–75:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530158.

Lee:2018:IDW

[LKE18] Sungkil Lee, Younguk Kim, and Elmar Eisemann. Iterative depth warping. ACM Transactions on Graphics, 37(5):177:1–177:??, November 2018. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3190859.

Lipman:2012:SFQ

[LKF12] Yaron Lipman, Vladimir G.

Kim, and Thomas A. Funkhouser. Simple formulas for quasiconformal plane deformations. ACM Transactions on Graphics, 31(5):124:1– 124:13, August 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lensch:2003:IBR

[LKG^{+}03a] Hendrik P. A. Lensch, Jan Kautz, Michael Goesele, Wolfgang Heidrich, and Hans-Peter Seidel. Image-based reconstruction of spatial appearance and geometric detail. ACM Transactions on Graphics, 22(2):234–257, April 2003. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Llamas:2003:TSW

[LKG^{+}03b] Ignacio Llamas, Byungmoon Kim, Joshua Gargus, Jarek Rossignac, and Chris D. Shaw. Twister: a space-warp operator for the two-handed editing of 3D shapes. ACM Transactions on Graphics, 22 (3):663–668, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2021:CIP

[LKJ21] Minchen Li, Danny M. Kaufman, and Chenfanfu Jiang. Codimensional incremental potential contact. ACM Transactions on Graphics, 40(4):170:1–170:24, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3450626.3459767.

Li:2021:ICM

[LKJC21] Yong Li, Shoaib Kamil, Alec Jacobson, and Yotam Cingold. I[HEART]LA: compilable markdown for linear algebra. ACM Transactions on Graphics, 40(6): 264:1–264:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480506.

Lee:2016:ROS

[LKK^{+}16] Jungjin Lee, Bumki Kim, Kyehyun Kim, Younghui Kim, and Junyong Noh. Rich360: optimized spherical representation from structured panoramic camera arrays. ACM Transactions on Graphics, 35(4):63:1–63:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2018:OJO

[LKK^{+}18] Minchen Li, Danny M. Kaufman, Vladimir G. Kim, Justin Solomon, and Alla Sheffer. OptCuts: joint optimization of surface cuts and parameterization. ACM Transactions on Graphics, 37(6): 247:1–247:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2021:LRB

[LKK^{+}21] Bosheng Li, Jacek Kałuzny, Jonathan Klein, Dominik L.

Michels, Wojtek Pałubicki, Bedrich Benes, and Sören Pirk. Learning to reconstruct botanical trees from single images. ACM Transactions on Graphics, 40(6): 231:1–231:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480525.

Lee:2010:DDB

[LKL10] Yoonsang Lee, Sungeun Kim, and Jehee Lee. Data-driven biped control. ACM Transactions on Graphics, 29(4): 129:1–129:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lehtinen:2013:GDM

[LKL^{+}13] Jaakko Lehtinen, Tero Karras, Samuli Laine, Miika Aittala, Frédo Durand, and Timo Aila. Gradient-domain Metropolis light transport. ACM Transactions on Graphics, 32(4): 95:1–95:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lan:2022:ABD

[LKL^{+}22] Lei Lan, Danny M. Kaufman, Minchen Li, Chenfanfu Jiang, and Yin Yang. Affine body dynamics: fast, stable and intersection-free simulation of stiff materials. ACM Transactions on Graphics, 41(4):67:1–67:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530064.

Lee:2023:CMI

[LKL23] Dohae Lee, Hyun Kang, and In-Kwon Lee. ClothCombo: Modeling inter-cloth interaction for draping multilayered clothes. ACM Transactions on Graphics, 42(6): 247:1–247:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618376.

Li:2023:RCF

[LKM^{+}23] Bosheng Li, Jonathan Klein, Dominik L. Michels, Bedrich Benes, Sören Pirk, and Wojtek Pałubicki. Rhizomorph: The coordinated function of shoots and roots. ACM Transactions on Graphics, 42 (4):59:1–59:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592145.

Lun:2015:ESL

[LKS15] Zhaoliang Lun, Evangelos Kalogerakis, and Alla Sheffer. Elements of style: learning perceptual shape style similarity. ACM Transactions on Graphics, 34(4):84:1– 84:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levine:2010:GC

[LKTK10] Sergey Levine, Philipp Krähenbühl, Sebastian Thrun, and Vladlen Koltun. Gesture controllers. ACM Transactions on Graphics, 29(4):124:1–124:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loffler:2014:CDF

[LKvK^{+}14] Maarten Löffler, Mira Kaiser, Tim van Kapel, Gerwin Klappe, Marc van Kreveld, and Frank Staals. The Connect-The-Dots family of puzzles: design and automatic generation. ACM Transactions on Graphics, 33(4): 72:1–72:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lun:2016:FPS

[LKWS16] Zhaoliang Lun, Evangelos Kalogerakis, Rui Wang, and Alla Sheffer. Functionality preserving shape style transfer. ACM Transactions on Graphics, 35(6):209:1– 209:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Low:2012:BMA

[LKYU12] Joakim Löw, Joel Kronander, Anders Ynnerman, and Jonas Unger. BRDF models for accurate and efficient rendering of glossy surfaces. ACM Transactions on Graphics, 31(1):9:1–9:14, January 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2020:DFA

[LKZ^{+}20] Jiaman Li, Zhengfei Kuang, Yajie Zhao, Mingming He, Karl Bladin, and Hao Li. Dynamic facial asset and rig generation from a single scan. ACM Transactions on Graphics, 39(6): 215:1–215:18, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417817.

Levy:2010:CVT

[LL10] Bruno Lévy and Yang Liu.

L_{p} Centroidal Voronoi Tessellation and its applications. ACM Transactions on Graphics, 29(4):119:1–119:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lepage:2011:MM

[LL11] Daniel Lepage and Jason Lawrence. Material matting. ACM Transactions on Graphics, 30(6):144:1– 144:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lanman:2013:NEL

[LL13] Douglas Lanman and David Luebke. Near-eye light field displays. ACM Transactions on Graphics, 32(6):220:1– 220:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Le:2019:DDM

[LL19] Binh Huy Le and J. P. Lewis. Direct delta mush skinning and variants. ACM Transactions on Graphics, 38(4): 113:1–113:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2023:PKS

[LL23] Mengyun Liu and Xiaopei Liu. A parametric kinetic solver for simulating boundary-dominated turbulent flow phenomena. ACM Transactions on Graphics, 42 (6):189:1–189:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3618313.

Lee:2024:LST

[LLB24] Jae Joong Lee, Bosheng Li, and Bedrich Benes. Latent L-systems: Transformerbased tree generator. ACM Transactions on Graphics, 43 (1):7:1–7:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3627101.

Lipman:2008:GC

[LLCO08] Yaron Lipman, David Levin, and Daniel Cohen-Or. Green Coordinates. ACM Transactions on Graphics, 27(3):78:1– 78:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lagae:2009:PNU

[LLDD09] Ares Lagae, Sylvain Lefebvre, George Drettakis, and Philip Dutré. Procedural noise using sparse Gabor convolution. ACM Transactions on Graphics, 28(3):54:1–54:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lyu:2021:FVF

[LLDL21] Chaoyang Lyu, Wei Li, Mathieu Desbrun, and Xiaopei Liu. Fast and versatile fluidsolid coupling for turbulent flow simulation. ACM Transactions on Graphics, 40(6): 201:1–201:18, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480493.

Lan:2020:MEE

[LLF^{+}20] Lei Lan, Ran Luo, Marco Fratarcangeli, Weiwei Xu, Huamin Wang, Xiaohu Guo, Junfeng Yao, and Yin Yang. Medial elastics: Efficient and collision-ready deformation via medial axis transform. ACM Transactions on Graphics, 39(3):20:1–20:17, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3384515.

Li:2020:DVG

[LLGRK20] Tzu-Mao Li, Michal Lukác, Michaël Gharbi, and Jonathan Ragan-Kelley. Differentiable vector graphics rasterization for editing and learning. ACM Transactions on Graphics, 39 (6):193:1–193:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417871.

Liu:2004:NRT

[LLH04] Yanxi Liu, Wen-Chieh Lin, and James Hays. Nearregular texture analysis and manipulation. ACM Transactions on Graphics, 23(3): 368–376, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2022:LRD

[LLH^{+}22] Yilin Liu, Liqiang Lin, Yue Hu, Ke Xie, Chi-Wing Fu, Hao Zhang, and Hui Huang. Learning reconstructability for drone aerial path planning. ACM Transactions on Graphics, 41(6):197:1– 197:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555433.

Li:2021:SGS

[LLHF21] Ruihui Li, Xianzhi Li, Ka-Hei Hui, and Chi-Wing Fu. SP-GAN: sphere-guided 3D shape generation and manipulation. ACM Transactions on Graphics, 40(4): 151:1–151:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459766.

Li:2022:IAR

[LLHY22] Changyang Li, Wanwan Li, Haikun Huang, and Lap-Fai Yu. Interactive augmented reality storytelling guided by scene semantics. ACM Transactions on Graphics, 41(4):91:1–91:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530061.

Levin:2011:ESS

[LLJ^{+}11] David I. W. Levin, Joshua Litven, Garrett L. Jones, Shinjiro Sueda, and Dinesh K. Pai. Eulerian solid simulation with contact. ACM Transactions on Graphics, 30(4): 36:1–36:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2022:ECI

[LLJ22] Xuan Li, Minchen Li, and Chenfanfu Jiang. Energetically consistent inelasticity for optimization time integration. ACM Transactions on Graphics, 41(4):52:1–52:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530072.

Lan:2023:SOS

[LLJ^{+}23] Lei Lan, Minchen Li, Chenfanfu Jiang, Huamin Wang, and Yin Yang. Second-order stencil descent for interiorpoint hyperelasticity. ACM Transactions on Graphics, 42(4):108:1–108:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592104.

Lee:2015:PRS

[LLK^{+}15] Yoonsang Lee, Kyungho Lee, Soon-Sun Kwon, Jiwon Jeong, Carol O’Sullivan, Moon Seok Park, and Jehee Lee. Pushrecovery stability of biped locomotion. ACM Transactions on Graphics, 34(6): 180:1–180:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2019:VEU

[LLK^{+}19] Sangwoo Lee, Jungjin Lee, Bumki Kim, Kyehyun Kim, and Junyong Noh. Video extrapolation using neighboring frames. ACM Transactions on Graphics, 38(3): 20:1–20:??, June 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3196492.

Longva:2020:HOF

[LLK^{+}20] Andreas Longva, Fabian Löschner, Tassilo Kugelstadt, José Antonio Fernández-Fernández, and Jan Bender. Higher-order finite elements for embedded simulation. ACM Transactions on Graphics, 39(6): 181:1–181:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417853.

Li:2021:ICT

[LLKC21] Jing Li, Tiantian Liu, Ladislav Kavan, and Baoquan Chen. Interactive cutting and tearing in projective dynamics with progressive Cholesky updates. ACM Transactions on Graphics, 40(6): 254:1–254:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480505.

Levine:2011:STP

[LLKP11] Sergey Levine, Yongjoon Lee, Vladlen Koltun, and Zoran Popović. Space-time planning with parameterized locomotion controllers. ACM Transactions on Graphics, 30(3): 23:1–23:11, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2018:ICA

[LLL18] Kyungho Lee, Seyoung Lee, and Jehee Lee. Interactive character animation by learning multi-objective control. ACM Transactions on Graphics, 37(6):180:1– 180:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2022:LVC

[LLL22] Seyoung Lee, Jiye Lee, and Jehee Lee. Learning virtual chimeras by dynamic motion reassembly. ACM Transactions on Graphics, 41(6): 182:1–182:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555489.

Lee:2021:LFM

[LLLL21] Seyoung Lee, Sunmin Lee, Yongwoo Lee, and Jehee Lee.

Learning a family of motor skills from a single motion clip. ACM Transactions on Graphics, 40(4): 93:1–93:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459774.

Li:2021:BAG

[LLM21] Sijia Li, Shiguang Liu, and Dinesh Manocha. Binaural audio generation via multitask learning. ACM Transactions on Graphics, 40(6): 243:1–243:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480560.

Li:2016:AVC

[LLMZ16] Dingzeyu Li, David I. W.

Levin, Wojciech Matusik, and Changxi Zheng. Acoustic voxels: computational optimization of modular acoustic filters. ACM Transactions on Graphics, 35(4): 88:1–88:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liao:2014:AIM

[LLN^{+}14] Jing Liao, Rodolfo S. Lima, Diego Nehab, Hugues Hoppe, Pedro V. Sander, and Jinhui Yu. Automating image morphing using structural similarity on a halfway domain. ACM Transactions on Graphics, 33(5):168:1–168:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2009:CCC

[LLP09] Yongjoon Lee, Seong Jae Lee, and Zoran Popović. Compact character controllers. ACM Transactions on Graphics, 28(5):169:1–169:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Luo:2013:SAH

[LLR13] Linjie Luo, Hao Li, and Szymon Rusinkiewicz. Structureaware hair capture. ACM Transactions on Graphics, 32 (4):76:1–76:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:AGM

[LLR^{+}15] Tzu-Mao Li, Jaakko Lehtinen, Ravi Ramamoorthi, Wenzel Jakob, and Frédo Durand. Anisotropic Gaussian mutations for Metropolis Light Transport through Hessian– Hamiltonian dynamics. ACM Transactions on Graphics, 34 (6):209:1–209:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2012:TCC

[LLV^{+}12] Hao Li, Linjie Luo, Daniel Vlasic, Pieter Peers, Jovan Popović, Mark Pauly, and Szymon Rusinkiewicz. Temporally coherent completion of dynamic shapes. ACM Transactions on Graphics, 31(1): 2:1–2:11, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levin:2004:CUO

[LLW04] Anat Levin, Dani Lischinski, and Yair Weiss. Colorization using optimization. ACM Transactions on Graphics, 23(3):689–694, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liang:2008:PAP

[LLW^{+}08] Chia-Kai Liang, Tai-Hsu Lin, Bing-Yi Wong, Chi Liu, and Homer H. Chen. Programmable aperture photography: multiplexed light field acquisition. ACM Transactions on Graphics, 27(3):55:1– 55:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2017:DEM

[LLW17] Chengze Li, Xueting Liu, and Tien-Tsin Wong. Deep extraction of manga structural lines. ACM Transactions on Graphics, 36(4):117:1–117:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2014:SSE

[LLWD14] Xing Lin, Yebin Liu, Jiamin Wu, and Qionghai Dai. Spatial-spectral encoded compressive hyperspectral imaging. ACM Transactions on Graphics, 33(6):233:1– 233:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liang:2001:RTT

[LLX^{+}01] Lin Liang, Ce Liu, Ying-Qing Xu, Baining Guo, and Heung-Yeung Shum. Real-time texture synthesis by patchbased sampling. ACM Transactions on Graphics, 20(3): 127–150, July 2001. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2012:AHM

[LLX^{+}12] Yufei Li, Yang Liu, Weiwei Xu, Wenping Wang, and Baining Guo. All-hex meshing using singularityrestricted field. ACM Transactions on Graphics, 31(6): 177:1–177:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2018:SAA

[LLZ18] Dingzeyu Li, Timothy R. Langlois, and Changxi Zheng. Scene-aware audio for 360^{∘} videos. ACM Transactions on Graphics, 37(4):111:1– 111:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2020:CCS

[LLZ^{+}20] Zehui Lin, Sheng Li, Xinlu Zeng, Congyi Zhang, Jinzhu Jia, Guoping Wang, and Dinesh Manocha. CPPM: chi-squared progressive photon mapping. ACM Transactions on Graphics, 39(6): 240:1–240:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417822.

Li:2010:ARM

[LLZM10] Guo Li, Ligang Liu, Hanlin Zheng, and Niloy J. Mitra. Analysis, reconstruction and manipulation using arterial snakes. ACM Transactions on Graphics, 29(6): 152:1–152:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:1991:CSP

[LM91] S. L. Lee and A. A. Majid. Closed smooth piecewise bicubic surfaces. ACM Transactions on Graphics, 10 (4):342–365, October 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
116915.html.

Liu:1997:OAE

[LM97] Wayne Liu and Stephen Mann. An optimal algorithm for expanding the composition of polynomials. ACM Transactions on Graphics, 16 (2):155–178, April 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1997-16-2/
                                        p155-liu/.

Li:2018:CFR

[LMAH^{+}18] Shuhua Li, Ali Mahdavi-Amiri, Ruizhen Hu, Han Liu, Changqing Zou, Oliver Van Kaick, Xiuping Liu, Hui Huang, and Hao Zhang. Construction and fabrication of reversible shape transforms. ACM Transactions on Graphics, 37(6):190:1– 190:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2016:SSC

[LMAS16] Haixiang Liu, Nathan Mitchell, Mridul Aanjaneya, and Eftychios Sifakis. A scalable Schur-complement fluids solver for heterogeneous compute platforms. ACM Transactions on Graphics, 35(6): 201:1–201:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loper:2014:MMS

[LMB14] Matthew Loper, Naureen Mahmood, and Michael J. Black. MoSh: motion and shape capture from sparse markers. ACM Transactions on Graphics, 33(6):220:1– 220:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2015:MRV

[LMH^{+}15] Beibei Liu, Gemma Mason, Julian Hodgson, Yiying Tong, and Mathieu Desbrun. Modelreduced variational fluid simulation. ACM Transactions on Graphics, 34(6):244:1– 244:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2022:EKS

[LMLD22] Wei Li, Yihui Ma, Xiaopei Liu, and Mathieu Desbrun. Efficient kinetic simulation of two-phase flows. ACM Transactions on Graphics, 41 (4):114:1–114:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530132.

Lee:2007:LDA

[LMLH07] Yunjin Lee, Lee Markosian, Seungyong Lee, and John F. Hughes. Line drawings via abstracted shading. ACM Transactions on Graphics, 26(3):

18:1–18:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2021:LTC

[LMLL21] Kyungho Lee, Sehee Min, Sunmin Lee, and Jehee Lee. Learning time-critical responses for interactive character control. ACM Transactions on Graphics, 40(4): 147:1–147:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459826.

Li:2022:EHR

[LMM^{+}22] Gengyan Li, Abhimitra Meka, Franziska Mueller, Marcel C. Buehler, Otmar Hilliges, and Thabo Beeler. EyeNeRF: a hybrid representation for photorealistic synthesis, animation and relighting of human eyes. ACM Transactions on Graphics, 41 (4):166:1–166:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530130.

Lopez-Moreno:2013:DSM

[LMPB^{+}13] Jorge Lopez-Moreno, Stefan Popov, Adrien Bousseau, Maneesh Agrawala, and George Drettakis. Depicting stylized materials with vector shade trees. ACM Transactions on Graphics, 32(4):

118:1–118:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lane:1983:AFR

[LMR83] J. M. Lane, R. Magedson, and M. Rarick. An algorithm for filling regions on graphics display devices. ACM Transactions on Graphics, 2 (3):192–196, July 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loper:2015:SSM

[LMR^{+}15] Matthew Loper, Naureen Mahmood, Javier Romero, Gerard Pons-Moll, and Michael J. Black. SMPL: a skinned multi-person linear model. ACM Transactions on Graphics, 34(6):248:1–248:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Laga:2013:GCS

[LMS13] Hamid Laga, Michela Mortara, and Michela Spagnuolo. Geometry and context for semantic correspondences and functionality recognition in man-made 3D shapes. ACM Transactions on Graphics, 32 (5):150:1–150:16, September 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lagunas:2019:SMM

[LMS^{+}19] Manuel Lagunas, Sandra Malpica, Ana Serrano, Elena Garces, Diego Gutierrez, and Belen Masia. A similarity measure for material appearance. ACM Transactions on Graphics, 38(4): 135:1–135:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2013:SCA

[LMY^{+}13] Xueting Liu, Xiangyu Mao, Xuan Yang, Linling Zhang, and Tien-Tsin Wong. Stereoscopizing cel animations. ACM Transactions on Graphics, 32(6):223:1–223:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lan:2022:PFP

[LMY^{+}22] Lei Lan, Guanqun Ma, Yin Yang, Changxi Zheng, Minchen Li, and Chenfanfu Jiang. Penetration-free projective dynamics on the GPU. ACM Transactions on Graphics, 41(4):69:1–69:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530069.

Lantz:1984:SGD

[LN84] K. A. Lantz and W. I. Nowicki. Structured graphics for distributed systems. ACM Transactions on Graphics, 3 (1):23–51, January 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lang:2022:PLG

[LN22] Johannes Lang and Miguel A.

Nacenta. Perception of letter glyph parameters for InfoTypography. ACM Transactions on Graphics, 41 (4):147:1–147:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530111.

Levoy:2006:LFM

[LNA^{+}06] Marc Levoy, Ren Ng, Andrew Adams, Matthew Footer, and Mark Horowitz. Light field microscopy. ACM Transactions on Graphics, 25(3): 924–934, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2023:SSM

[LNI^{+}23] Jenny Lin, Vidya Narayanan, Yuka Ikarashi, Jonathan Ragan-Kelley, Gilbert Bernstein, and James McCann. Semantics and scheduling for machine knitting compilers. ACM Transactions on Graphics, 42(4):143:1– 143:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592449.

Lou:2016:IPA

[LNLB16] Liming Lou, Paul Nguyen, Jason Lawrence, and Connelly Barnes. Image perforation: Automatically accelerating image pipelines by intelligently skipping samples. ACM Transactions on Graphics, 35 (5):153:1–153:??, September 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lok:2003:IDR

[LNWB03] Benjamin Lok, Samir Naik, Mary Whitton, and Frederick P. Brooks. Incorporating dynamic real objects into immersive virtual environments. ACM Transactions on Graphics, 22(3):701, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:GLG

[LNZ^{+}23] Xingqiao Li, Xingyu Ni, Bo Zhu, Bin Wang, and Baoquan Chen. GARM-LS: a gradient-augmented reference-map method for level-set fluid simulation. ACM Transactions on Graphics, 42(6):192:1–192:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618377.

Lau:2011:CFM

[LOMI11] Manfred Lau, Akira Ohgawara, Jun Mitani, and Takeo Igarashi. Converting 3D furniture models to fabricatable parts and connectors. ACM Transactions on Graphics, 30 (4):85:1–85:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lindell:2018:SPI

[LOW18] David B. Lindell, Matthew O’Toole, and Gordon Wetzstein. Single-photon 3D imaging with deep sensor fusion. ACM Transactions on Graphics, 37(4):113:1– 113:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2002:SCD

[LP02] C. Karen Liu and Zoran Popović. Synthesis of complex dynamic character motion from simple animations. ACM Transactions on Graphics, 21(3):408–416, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2010:LBS

[LP10] Seong Jae Lee and Zoran Popović. Learning behavior styles with inverse reinforcement learning. ACM Transactions on Graphics, 29(4): 122:1–122:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2020:SSC

[LPBM20] Changjian Li, Hao Pan, Adrien Bousseau, and Niloy J. Mitra. Sketch2CAD: sequential CAD modeling by sketching in context. ACM Transactions on Graphics, 39(6): 164:1–164:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417807.

Li:2022:FPF

[LPBM22] Changjian Li, Hao Pan, Adrien Bousseau, and Niloy J. Mitra. Free2CAD: parsing freehand drawings into CAD commands. ACM Transactions on Graphics, 41(4):93:1–93:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530133.

Livny:2011:TLT

[LPC^{+}11] Yotam Livny, Soeren Pirk, Zhanglin Cheng, Feilong Yan, Oliver Deussen, Daniel Cohen-Or, and Baoquan Chen. Texture-lobes for tree modelling. ACM Transactions on Graphics, 30(4): 53:1–53:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Livesu:2022:ODS

[LPC22] Marco Livesu, Luca Pitzalis, and Gianmarco Cherchi. Optimal dual schemes for adaptive grid based hexmeshing. ACM Transactions on Graphics, 41(2):15:1–15:14, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3494456.

Liu:2023:DSS

[LPC^{+}23] Daoming Liu, Davide Pellis, Yu-Chou Chiang, Florian Rist, Johannes Wallner, and Helmut Pottmann. Deployable strip structures. ACM Transactions on Graphics, 42(4):103:1–103:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592393.

Lee:2014:LCM

[LPKL14] Yoonsang Lee, Moon Seok Park, Taesoo Kwon, and Jehee Lee. Locomotion control for many-muscle humanoids. ACM Transactions on Graphics, 33(6):218:1– 218:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2017:BMF

[LPL^{+}17] Changjian Li, Hao Pan, Yang Liu, Xin Tong, Alla Sheffer, and Wenping Wang. BendSketch: modeling freeform surfaces through 2D sketching. ACM Transactions on Graphics, 36(4):125:1–125:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2018:RFG

[LPL^{+}18] Changjian Li, Hao Pan, Yang Liu, Xin Tong, Alla Sheffer, and Wenping Wang. Robust flow-guided neural prediction for sketch-based freeform surface modeling. ACM Transactions on Graphics, 37(6): 238:1–238:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2019:SMA

[LPLL19] Seunghwan Lee, Moonseok Park, Kyoungmin Lee, and Jehee Lee. Scalable muscleactuated human simulation and control. ACM Transactions on Graphics, 38(4): 73:1–73:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levy:2002:LSC

[LPRM02] Bruno Lévy, Sylvain Petitjean, Nicolas Ray, and Jérome Maillot. Least squares conformal maps for automatic texture atlas generation. ACM Transactions on Graphics, 21 (3):362–371, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2013:CSS

[LPS^{+}13] Yang Liu, Hao Pan, John Snyder, Wenping Wang, and Baining Guo. Computing selfsupporting surfaces by regular triangulation. ACM Transactions on Graphics, 32(4): 92:1–92:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2006:GMC

[LPW^{+}06] Yang Liu, Helmut Pottmann, Johannes Wallner, Yong-Liang Yang, and Wenping Wang. Geometric modeling with conical meshes and developable surfaces. ACM Transactions on Graphics, 25 (3):681–689, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2019:GGR

[LPX^{+}19] Manyi Li, Akshay Gadi Patil, Kai Xu, Siddhartha Chaudhuri, Owais Khan, Ariel Shamir, Changhe Tu, Baoquan Chen, Daniel Cohen-Or, and Hao Zhang. GRAINS: Generative recursive autoencoders for INdoor scenes. ACM Transactions on Graphics, 38(2):12:1–12:??, April 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/ft_gateway.
                                        cfm?id=3303766.

Li:2024:LAS

[LQGY24] Pu Li, Weize Quan, Jianwei Guo, and Dong-Ming Yan. Layout-aware single-image document flattening. ACM Transactions on Graphics, 43 (1):9:1–9:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3627818.

Lamming:1990:SMI

[LR90] Michael G. Lamming and Warren L. Rhodes. A simple method for improved color printing of monitor images. ACM Transactions on Graphics, 9(4):345–375, October 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/

88567.html. See corrigenda

[LR91].

Lamming:1991:CSM

[LR91] Michael G. Lamming and Warren L. Rhodes. Corrigenda: “A Simple Method for Improved Color Printing of Monitor Images”. ACM Transactions on Graphics, 10 (1):109, January 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [LR90].

Liang:2015:LTF

[LR15] Chia-Kai Liang and Ravi Ramamoorthi. A light transport framework for lenslet light field cameras. ACM Transactions on Graphics, 34(2):16:1– 16:??, February 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2007:ICI

[LRA^{+}07] Wilmot Li, Lincoln Ritter, Maneesh Agrawala, Brian Curless, and David Salesin. Interactive cutaway illustrations of complex 3D models. ACM Transactions on Graphics, 26(3):31:1–31:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lanman:2008:SFM

[LRAT08] Douglas Lanman, Ramesh Raskar, Amit Agrawal, and Gabriel Taubin. Shield fields: modeling and capturing 3D occluders. ACM Transactions on Graphics, 27(5): 131:1–131:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2010:BD

[LRF10] Yaron Lipman, Raif M.

Rustamov, and Thomas A. Funkhouser. Biharmonic distance. ACM Transactions on Graphics, 29(3): 27:1–27:11, June 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2013:PCN

[LRFH13] Sharon Lin, Daniel Ritchie, Matthew Fisher, and Pat Hanrahan. Probabilistic colorby-numbers: suggesting pattern colorizations using factor graphs. ACM Transactions on Graphics, 32(4): 37:1–37:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lewis:2004:VAD

[LRFN04] J. P. Lewis, Ruth Rosenholtz, Nickson Fong, and Ulrich Neumann. VisualIDs: automatic distinctive icons for desktop interfaces. ACM Transactions on Graphics, 23 (3):416–423, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lawrence:2004:EBI

[LRR04] Jason Lawrence, Szymon Rusinkiewicz, and Ravi Ramamoorthi. Efficient BRDF importance sampling using a factored representation. ACM Transactions on Graphics, 23 (3):496–505, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:SCR

[LRS18] Chenxi Liu, Enrique Rosales, and Alla Sheffer. StrokeAggregator: consolidating raw sketches into artist-intended curve drawings. ACM Transactions on Graphics, 37(4): 97:1–97:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Laffont:2014:TAH

[LRT^{+}14] Pierre-Yves Laffont, Zhile Ren, Xiaofeng Tao, Chao Qian, and James Hays. Transient attributes for high-level understanding and editing of outdoor scenes. ACM Transactions on Graphics, 33(4): 149:1–149:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2000:NTT

[LS00] Michael Lee and Hanan Samet. Navigating through triangle meshes implemented as linear quadtrees. ACM Transactions on Graphics, 19(2):79–121, April 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tog/2000-
19-2/p79-lee/p79-lee.pdf;
 http://www.acm.org/pubs/
citations/journals/tog/
2000-19-2/p79-lee/.

Lalonde:2002:SDC

[LS02] Paul Lalonde and Eric Schenk.

Shader-driven compilation of rendering assets. ACM Transactions on Graphics, 21(3): 713–720, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Labelle:2007:ISF

[LS07] François Labelle and Jonathan Richard Shewchuk. Isosurface stuffing: fast tetrahedral meshes with good dihedral angles. ACM Transactions on Graphics, 26 (3):57:1–57:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Loop:2008:ACC

[LS08] Charles Loop and Scott Schaefer. Approximating Catmull–Clark subdivision surfaces with bicubic patches. ACM Transactions on Graphics, 27(1):8:1–8:11, March 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2005:CCH

[LSA05] Yuanzhen Li, Lavanya Sharan, and Edward H. Adelson. Compressing and companding high dynamic range images with subband architectures. ACM Transactions on Graphics, 24(3):836–844, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lockerman:2016:MSL

[LSA^{+}16] Yitzchak David Lockerman, Basile Sauvage, Rémi Allègre, Jean-Michel Dischler, Julie Dorsey, and Holly Rushmeier. Multi-scale label-map extraction for texture synthesis. ACM Transactions on Graphics, 35(4):140:1–140:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levin:2008:MIP

[LSC^{+}08] Anat Levin, Peter Sand, Taeg Sang Cho, Frédo Durand, and William T. Freeman. Motion-invariant photography. ACM Transactions on Graphics, 27(3):71:1– 71:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Luo:2012:PAW

[LSC^{+}12] Sheng-Jie Luo, I-Chao Shen, Bing-Yu Chen, Wen-Huang Cheng, and Yung-Yu Chuang. Perspective-aware warping for seamless stereoscopic image cloning. ACM Transactions on Graphics, 31(6):182:1– 182:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lai:2022:FDU

[LSC^{+}22] Wei-Sheng Lai, Yichang Shih, Lun-Cheng Chu, Xiaotong Wu, Sung-Fang Tsai, Michael Krainin, Deqing Sun, and Chia-Kai Liang. Face deblurring using dual camera fusion on mobile phones. ACM Transactions on Graphics, 41 (4):148:1–148:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530131.

Luo:2020:CCA

[LSCC20] Ying-Sheng Luo, Jonathan Hans Soeseno, Trista Pei-Chun Chen, and Wei-Chao Chen. CARL: controllable agent with reinforcement learning for quadruped locomotion. ACM Transactions on Graphics, 39(4): 38:1–38:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392433.

Leyvand:2003:RSF

[LSCO03] Tommer Leyvand, Olga Sorkine, and Daniel Cohen-Or. Ray space factorization for fromregion visibility. ACM Transactions on Graphics, 22(3): 595–604, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Long:2014:RVH

[LSCS14] Benjamin Long, Sue Ann Seah, Tom Carter, and Sriram Subramanian. Rendering volumetric haptic shapes in midair using ultrasound. ACM Transactions on Graphics, 33 (6):181:1–181:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Langlois:2016:SSA

[LSD^{+}16] Timothy Langlois, Ariel Shamir, Daniel Dror, Wojciech Matusik, and David I. W. Levin. Stochastic structural analysis for contextaware design and fabrication. ACM Transactions on Graphics, 35(6):226:1– 226:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lesser:2022:LUM

[LSD^{+}22] Steve Lesser, Alexey Stomakhin, Gilles Daviet, Joel Wretborn, John Edholm, Noh-Hoon Lee, Eston Schweickart, Xiao Zhai, Sean Flynn, and Andrew Moffat. Loki: a unified multiphysics simulation framework for production. ACM Transactions on Graphics, 41(4): 50:1–50:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530058.

Li:2018:FEG

[LSGV18] Minchen Li, Alla Sheffer, Eitan Grinspun, and Nicholas Vining. Foldsketch: enriching garments with physically reproducible folds. ACM Transactions on Graphics, 37(4): 133:1–133:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2010:PAP

[LSH^{+}10] Xian-Ying Li, Chao-Hui Shen, Shi-Sheng Huang, Tao Ju, and Shi-Min Hu. Popup: automatic paper architectures from 3D models. ACM Transactions on Graphics, 29(4): 111:1–111:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2022:AAS

[LSH^{+}22] Difan Liu, Sandesh Shetty, Tobias Hinz, Matthew Fisher, Richard Zhang, Taesung Park, and Evangelos Kalogerakis. ASSET: autoregressive semantic scene editing with transformers at high resolutions. ACM Transactions on Graphics, 41(4): 74:1–74:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530172.

Lefohn:2006:GGE

[LSK^{+}06] Aaron E. Lefohn, Shubhabrata Sengupta, Joe Kniss, Robert Strzodka, and John D. Owens. Glift: Generic, efficient, random-access GPU data structures. ACM Transactions on Graphics, 25(1): 60–99, January 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Langbehn:2018:BEL

[LSL^{+}18] Eike Langbehn, Frank Steinicke, Markus Lappe, Gregory F. Welch, and Gerd Bruder. In the blink of an eye: leveraging blink-induced suppression for imperceptible position and orientation redirection in virtual reality. ACM Transactions on Graphics, 37(4): 66:1–66:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipman:2005:LRI

[LSLCO05] Yaron Lipman, Olga Sorkine, David Levin, and Daniel Cohen-Or. Linear rotationinvariant coordinates for meshes. ACM Transactions on Graphics, 24(3): 479–487, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:EEP

[LSM23] Beichen Li, Liang Shi, and Wojciech Matusik. End-to-end procedural material capture with proxy-free mixedinteger optimization. ACM Transactions on Graphics, 42(4):153:1–153:??, August 2023. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592132.

Loop:2009:ASS

[LSNC09] Charles Loop, Scott Schaefer, Tianyun Ni, and Ignacio Castaño. Approximating subdivision surfaces with Gregory patches for hardware tessellation. ACM Transactions on Graphics, 28(5): 151:1–151:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2013:TSE

[LSNP13] Duo Li, Shinjiro Sueda, Debanga R. Neog, and Dinesh K. Pai. Thin skin elastodynamics. ACM Transactions on Graphics, 32(4): 49:1–49:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lefohn:2007:RMS

[LSO07] Aaron E. Lefohn, Shubhabrata Sengupta, and John D. Owens. Resolution-matched shadow maps. ACM Transactions on Graphics, 26(4):20:1– 20:17, October 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:JES

[LSQ^{+}15] Yangyan Li, Hao Su, Charles Ruizhongtai Qi, Noa Fish, Daniel Cohen-Or, and Leonidas J. Guibas. Joint embeddings of shapes and images via CNN image purification. ACM Transactions on Graphics, 34 (6):234:1–234:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Leimkuhler:2018:LKS

[LSR18] Thomas Leimkühler, Hans-Peter Seidel, and Tobias Ritschel. Laplacian kernel splatting for efficient depthof-field and motion blur synthesis or reconstruction. ACM Transactions on Graphics, 37 (4):55:1–55:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2005:VOC

[LSS05] Yin Li, Jian Sun, and Heung-Yeung Shum. Video object cut and paste. ACM Transactions on Graphics, 24(3): 595–600, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2009:PS

[LSS09] Jiangyu Liu, Jian Sun, and Heung-Yeung Shum. Paint selection. ACM Transactions on Graphics, 28(3):69:1– 69:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lukac:2017:NRR

[LSS^{+}17] Michal Lukác, Daniel Sýkora, Kalyan Sunkavalli, Eli Shechtman, Ondrej Jamriska, Nathan Carr, and Tomás Pajdla. Nautilus: recovering regional symmetry transformations for image editing. ACM Transactions on Graphics, 36(4): 108:1–108:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lombardi:2019:NVL

[LSS^{+}19] Stephen Lombardi, Tomas Simon, Jason Saragih, Gabriel Schwartz, Andreas Lehrmann, and Yaser Sheikh. Neural volumes: learning dynamic renderable volumes from images. ACM Transactions on Graphics, 38(4):65:1–65:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lombardi:2021:MVP

[LSS^{+}21] Stephen Lombardi, Tomas Simon, Gabriel Schwartz, Michael Zollhoefer, Yaser Sheikh, and Jason Saragih. Mixture of volumetric primitives for efficient neural rendering. ACM Transactions on Graphics, 40(4): 59:1–59:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459863.

Losasso:2006:MIL

[LSSF06] Frank Losasso, Tamar Shinar, Andrew Selle, and Ronald Fedkiw. Multiple interacting liquids. ACM Transactions on Graphics, 25(3):812–819, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lombardi:2018:DAM

[LSSS18] Stephen Lombardi, Jason Saragih, Tomas Simon, and Yaser Sheikh. Deep appearance models for face rendering. ACM Transactions on Graphics, 37(4):68:1– 68:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lu:2019:SRB

[LSSW19] Wenjia Lu, Zuoqiang Shi, Jian Sun, and Bin Wang. Surface reconstruction based on the modified Gauss formula. ACM Transactions on Graphics, 38 (1):2:1–2:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3233984.

Lee:2009:CBM

[LST09] Sung-Hee Lee, Eftychios Sifakis, and Demetri Terzopoulos. Comprehensive biomechanical modeling and simulation of the upper body. ACM Transactions on Graphics, 28(4):99:1–99:17, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2004:LS

[LSTS04] Yin Li, Jian Sun, Chi-Keung Tang, and Heung-Yeung Shum. Lazy snapping. ACM Transactions on Graphics, 23(3):303–308, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Livesu:2015:PHM

[LSVT15] Marco Livesu, Alla Sheffer, Nicholas Vining, and Marco Tarini. Practical hex-mesh optimization via edge-cone rectification. ACM Transactions on Graphics, 34(4): 141:1–141:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lu:2014:BLS

[LSZ^{+}14] Lin Lu, Andrei Sharf, Haisen Zhao, Yuan Wei, Qingnan Fan, Xuelin Chen, Yann Savoye, Changhe Tu, Daniel Cohen-Or, and Baoquan Chen. Build-to-last: strength to weight 3D printed objects. ACM Transactions on Graphics, 33(4):97:1–97:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2022:AQP

[LSZ^{+}22] Jiafeng Liu, Haoyang Shi, Siyuan Zhang, Yin Yang, Chongyang Ma, and Weiwei Xu. Automatic quantization for physics-based simulation. ACM Transactions on Graphics, 41(4):51:1–51:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530154.

Lindstrom:2000:IDS

[LT00] Peter Lindstrom and Greg Turk. Image-driven simplification. ACM Transactions on Graphics, 19(3):204–241, 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tog/2000-
19-3/p204-lindstrom/p204-
lindstrom.pdf;   http://
www.acm.org/pubs/citations/
journals/tog/2000-19-3/
p204-lindstrom/.

Lee:2006:HBM

[LT06] Sung-Hee Lee and Demetri Terzopoulos. Heads up!: biomechanical modeling and neuromuscular control of the neck. ACM Transactions on Graphics, 25(3): 1188–1198, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2008:SJM

[LT08] Sung-Hee Lee and Demetri Terzopoulos. Spline joints for multibody dynamics. ACM Transactions on Graphics, 27 (3):22:1–22:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2009:PAS

[LT09] Qingde Li and Jie Tian. 2D piecewise algebraic splines for implicit modeling. ACM Transactions on Graphics, 28 (2):13:1–13:19, April 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lahav:2020:MDM

[LT20] Alon Lahav and Ayellet Tal. MeshWalker: deep mesh understanding by random walks. ACM Transactions on Graphics, 39(6): 263:1–263:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417806.

Liu:2016:DCC

[LTDD16] Beibei Liu, Yiying Tong, Fernando De Goes, and Mathieu Desbrun. Discrete connection and covariant derivative for vector field analysis and design. ACM Transactions on Graphics, 35(3): 23:1–23:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2005:MM

[LTF^{+}05] Ce Liu, Antonio Torralba, William T. Freeman, Frédo Durand, and Edward H. Adelson. Motion magnification. ACM Transactions on Graphics, 24(3):519–526, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lyu:2023:DPI

[LTH^{+}23] Linjie Lyu, Ayush Tewari, Marc Habermann, Shunsuke Saito, Michael Zollhöfer, Thomas Leimkühler, and Christian Theobalt. Diffusion posterior illumination for ambiguity-aware inverse rendering. ACM Transactions on Graphics, 42(6): 233:1–233:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618357.

Liu:2018:PSE

[LTJ18] Hsueh-Ti Derek Liu, Michael Tao, and Alec Jacobson. Paparazzi: surface editing by way of multi-view image processing. ACM Transactions on Graphics, 37(6):221:1– 221:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levine:2009:RTP

[LTK09] Sergey Levine, Christian Theobalt, and Vladlen Koltun.

Real-time prosody-driven synthesis of body language. ACM Transactions on Graphics, 28 (5):172:1–172:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:FPS

[LTO^{+}15] Hao Li, Laura Trutoiu, Kyle Olszewski, Lingyu Wei, Tristan Trutna, Pei-Lun Hsieh, Aaron Nicholls, and Chongyang Ma. Facial performance sensing head-mounted display. ACM Transactions on Graphics, 34(4):47:1– 47:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2020:PCI

[LTT^{+}20] Cheng Li, Min Tang, Ruofeng Tong, Ming Cai, Jieyi Zhao, and Dinesh Manocha. P-cloth: interactive complex cloth simulation on multi-GPU systems using dynamic matrix assembly and pipelined implicit integrators. ACM Transactions on Graphics, 39(6): 180:1–180:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417763.

Lo:2010:SCP

[LvBK^{+}10] Wan-Yen Lo, Jeroen van Baar, Claude Knaus, Matthias Zwicker, and Markus Gross.

Stereoscopic 3D copy & paste. ACM Transactions on Graphics, 29(6):147:1– 147:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2013:SP

[LVG^{+}13] Hao Li, Etienne Vouga, Anton Gudym, Linjie Luo, Jonathan T. Barron, and Gleb Gusev. 3D selfportraits. ACM Transactions on Graphics, 32(6): 187:1–187:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Le:2021:DDM

[LVGO21] Binh Huy Le, Keven Villeneuve, and Carlos Gonzalez-Ochoa. Direct delta mush skinning compression with continuous examples. ACM Transactions on Graphics, 40 (4):72:1–72:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459779.

Lee:2005:MS

[LVJ05] Chang Ha Lee, Amitabh Varshney, and David W. Jacobs. Mesh saliency. ACM Transactions on Graphics, 24 (3):659–666, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lewis:2021:TBA

[LVKS21] Kathleen M. Lewis, Srivatsan Varadharajan, and Ira Kemelmacher-Shlizerman. TryOnGAN: body-aware tryon via layered interpolation. ACM Transactions on Graphics, 40(4):115:1– 115:10, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459884.

Livesu:2013:PMG

[LVS^{+}13] Marco Livesu, Nicholas Vining, Alla Sheffer, James Gregson, and Riccardo Scateni. PolyCut: monotone graphcuts for PolyCube basecomplex construction. ACM Transactions on Graphics, 32 (6):171:1–171:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2016:RAP

[LVS^{+}16] Wenbin Li, Fabio Viola, Jonathan Starck, Gabriel J. Brostow, and Neill D. F. Campbell. Roto++: accelerating professional rotoscoping using shape manifolds. ACM Transactions on Graphics, 35 (4):62:1–62:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Limper:2018:BCA

[LVS18] Max Limper, Nicholas Vining, and Alla Sheffer. Box cutter: atlas refinement for efficient packing via void elimination. ACM Transactions on Graphics, 37(4):153:1– 153:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2016:GLC

[LVY16] Libin Liu, Michiel Van De Panne, and Kangkang Yin. Guided learning of control graphs for physics-based characters. ACM Transactions on Graphics, 35(3): 29:1–29:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:ATB

[LW15] Chuan Li and Michael Wand.

Approximate translational building blocks for image decomposition and synthesis. ACM Transactions on Graphics, 34(5):158:1–158:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levi:2016:CFB

[LW16] Zohar Levi and Ofir Weber.

On the convexity and feasibility of the bounded distortion harmonic mapping problem. ACM Transactions on Graphics, 35(4):106:1–106:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lang:2012:PTC

[LWA^{+}12] Manuel Lang, Oliver Wang, Tunc Aydin, Aljoscha Smolic, and Markus Gross. Practical temporal consistency for image-based graphics applications. ACM Transactions on Graphics, 31(4): 34:1–34:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2010:MFI

[LWB^{+}10] Yongjoon Lee, Kevin Wampler, Gilbert Bernstein, Jovan Popović, and Zoran Popović. Motion fields for interactive character locomotion. ACM Transactions on Graphics, 29 (6):138:1–138:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2011:GCF

[LWC^{+}11] Yangyan Li, Xiaokun Wu, Yiorgos Chrysathou, Andrei Sharf, Daniel Cohen-Or, and Niloy J. Mitra. GlobFit: consistently fitting primitives by discovering global relations. ACM Transactions on Graphics, 30(4):52:1–52:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2012:SBO

[LWC12] Tzu-Mao Li, Yu-Ting Wu, and Yung-Yu Chuang. SUREbased optimization for adaptive sampling and reconstruction. ACM Transactions on Graphics, 31(6):194:1– 194:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2013:NRM

[LWC^{+}13] Yiming Liu, Jue Wang, Sunghyun Cho, Adam Finkelstein, and Szymon Rusinkiewicz. A no-reference metric for evaluating the quality of motion deblurring. ACM Transactions on Graphics, 32(6): 175:1–175:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2014:TAT

[LWCT14] Shuaicheng Liu, Jue Wang, Sunghyun Cho, and Ping Tan. TrackCam: 3D-aware tracking shots from consumer video. ACM Transactions on Graphics, 33(6):198:1– 198:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2022:HHS

[LWF^{+}22] Jinyuan Liu, Mengdi Wang, Fan Feng, Annie Tang, Qiqin Le, and Bo Zhu. Hydrophobic and hydrophilic solid-fluid interaction. ACM Transactions on Graphics, 41(6): 256:1–256:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555478.

Lai:2023:PDP

[LWF^{+}23] Zeqiang Lai, Kaixuan Wei, Ying Fu, Philipp Härtel, and Felix Heide. Δ-prox: Differentiable proximal algorithm modeling for large-scale optimization. ACM Transactions on Graphics, 42(4): 105:1–105:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592144.

Lanman:2011:PFD

[LWH^{+}11] Douglas Lanman, Gordon Wetzstein, Matthew Hirsch, Wolfgang Heidrich, and Ramesh Raskar. Polarization fields: dynamic light field display using multi-layer LCDs. ACM Transactions on Graphics, 30 (6):186:1–186:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levine:2012:CCC

[LWH^{+}12] Sergey Levine, Jack M. Wang, Alexis Haraux, Zoran Popović, and Vladlen Koltun. Continuous character control with low-dimensional embeddings. ACM Transactions on Graphics, 31(4): 28:1–28:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2015:CAS

[LWH15] Xueting Liu, Tien-Tsin Wong, and Pheng-Ann Heng. Closureaware sketch simplification. ACM Transactions on Graphics, 34(6):168:1–168:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2009:CVT

[LWL^{+}09] Yang Liu, Wenping Wang, Bruno Lévy, Feng Sun, Dong-Ming Yan, Lin Lu, and Chenglei Yang. On centroidal Voronoi tessellation — energy smoothness and fast computation. ACM Transactions on Graphics, 28(4): 101:1–101:17, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2017:CST

[LWL17] Weizi Li, David Wolinski, and Ming C. Lin. City-scale traffic animation using statistical learning and metamodelbased optimization. ACM Transactions on Graphics, 36 (6):200:1–200:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lyu:2020:DRT

[LWL^{+}20] Jiahui Lyu, Bojian Wu, Dani Lischinski, Daniel Cohen-Or, and Hui Huang. Differentiable refraction-tracing for mesh reconstruction of transparent objects. ACM Transactions on Graphics, 39(6): 195:1–195:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417815.

Li:2023:OMG

[LWL23a] Jiaman Li, Jiajun Wu, and C. Karen Liu. Object motion guided human motion synthesis. ACM Transactions on Graphics, 42(6): 197:1–197:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618333.

Liu:2023:NNG

[LWL^{+}23b] Yuan Liu, Peng Wang, Cheng Lin, Xiaoxiao Long, Jiepeng Wang, Lingjie Liu, Taku Komura, and Wenping Wang. NeRO: Neural geometry and BRDF reconstruction of reflective objects from multiview images. ACM Transactions on Graphics, 42(4): 114:1–114:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592134.

Lipp:2014:P

[LWM14] Markus Lipp, Peter Wonka, and Pascal Müller. Push-Pull++. ACM Transactions on Graphics, 33(4):

130:1–130:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lindell:2019:WBN

[LWO19] David B. Lindell, Gordon Wetzstein, and Matthew O’Toole. Wave-based nonline-of-sight imaging using fast f-k migration. ACM Transactions on Graphics, 38 (4):116:1–116:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2010:EBF

[LWP10] Hao Li, Thibaut Weise, and Mark Pauly. Example-based facial rigging. ACM Transactions on Graphics, 29(4): 32:1–32:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:HOM

[LWP^{+}23] Wei Li, Tongtong Wang, Zherong Pan, Xifeng Gao, Kui Wu, and Mathieu Desbrun. High-order momentencoded kinetic simulation of turbulent flows. ACM Transactions on Graphics, 42(6): 190:1–190:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618341.

Liu:2008:IC

[LWQ^{+}08] Xiaopei Liu, Liang Wan, Yingge Qu, Tien-Tsin Wong, Stephen Lin, Chi-Sing Leung, and Pheng-Ann Heng. Intrinsic colorization. ACM Transactions on Graphics, 27(5): 152:1–152:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2002:MTT

[LWS02] Yan Li, Tianshu Wang, and Heung-Yeung Shum. Motion texture: a two-level statistical model for character motion synthesis. ACM Transactions on Graphics, 21(3): 465–472, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2015:QMC

[LWS^{+}15] Pan Li, Bin Wang, Feng Sun, Xiaohu Guo, Caiming Zhang, and Wenping Wang. Q-MAT: Computing medial axis transform by quadratic error minimization. ACM Transactions on Graphics, 35(1):8:1– 8:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Leaf:2018:IDP

[LWS^{+}18] Jonathan Leaf, Rundong Wu, Eston Schweickart, Doug L. James, and Steve Marschner. Interactive design of periodic yarn-level cloth patterns. ACM Transactions on Graphics, 37(6):202:1– 202:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2010:ABN

[LWSF10] Hongwei Li, Li-Yi Wei, Pedro V. Sander, and Chi-Wing Fu. Anisotropic blue noise sampling. ACM Transactions on Graphics, 29(6): 167:1–167:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lipp:2008:IVE

[LWW08] Markus Lipp, Peter Wonka, and Michael Wimmer. Interactive visual editing of grammars for procedural architecture. ACM Transactions on Graphics, 27(3): 102:1–102:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2015:VAD

[LXC^{+}15] Yilong Liu, Feng Xu, Jinxiang Chai, Xin Tong, Lijuan Wang, and Qiang Huo. Video-audio driven real-time facial animation. ACM Transactions on Graphics, 34(6): 182:1–182:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2017:GGR

[LXC^{+}17] Jun Li, Kai Xu, Siddhartha Chaudhuri, Ersin Yumer, Hao Zhang, and Leonidas Guibas. GRASS: generative recursive autoencoders for shape structures. ACM Transactions on Graphics, 36(4): 52:1–52:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2015:ECS

[LXFH15] Yong-Jin Liu, Chun-Xu Xu, Dian Fan, and Ying He. Efficient construction and simplification of Delaunay meshes. ACM Transactions on Graphics, 34(6):174:1– 174:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2022:COW

[LXG^{+}22] Jian Liu, Shiqing Xin, Xifeng Gao, Kaihang Gao, Kai Xu, Baoquan Chen, and Changhe Tu. Computational objectwrapping rope nets. ACM Transactions on Graphics, 41 (1):6:1–6:16, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3476829.

Luo:2022:AAN

[LXJ^{+}22] Haimin Luo, Teng Xu, Yuheng Jiang, Chenglin Zhou, Qiwei Qiu, Yingliang Zhang, Wei Yang, Lan Xu, and Jingyi Yu. Artemis: articulated neural pets with appearance and motion synthesis. ACM Transactions on Graphics, 41 (4):164:1–164:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530086.

Liu:2023:TGS

[LXL^{+}23] Lijuan Liu, Xiangyu Xu, Zhijie Lin, Jiabin Liang, and Shuicheng Yan. Towards garment sewing pattern reconstruction from a single image. ACM Transactions on Graphics, 42(6): 200:1–200:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618319.

Li:2018:LRS

[LXR^{+}18] Zhengqin Li, Zexiang Xu, Ravi Ramamoorthi, Kalyan Sunkavalli, and Manmohan Chandraker. Learning to reconstruct shape and spatiallyvarying reflectance from a single image. ACM Transactions on Graphics, 37(6): 269:1–269:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:OAG

[LXS^{+}18] Ligang Liu, Xi Xia, Han Sun, Qi Shen, Juzhan Xu, Bin Chen, Hui Huang, and Kai Xu. Object-aware guidance for autonomous scene reconstruction. ACM Transactions on Graphics, 37(4): 104:1–104:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2023:SRP

[LXSW23] Siyou Lin, Dong Xiao, Zuoqiang Shi, and Bin Wang. Surface reconstruction from point clouds without normals by parametrizing the Gauss formula. ACM Transactions on Graphics, 42(2): 14:1–14:??, April 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3554730.

Liu:2011:GPQ

[LXW^{+}11] Yang Liu, Weiwei Xu, Jun Wang, Lifeng Zhu, Baining Guo, Falai Chen, and Guoping Wang. General planar quadrilateral mesh design using conjugate direction field. ACM Transactions on Graphics, 30(6): 140:1–140:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2016:MDE

[LXY^{+}16] Yong-Jin Liu, Chun-Xu Xu, Ran Yi, Dian Fan, and Ying He. Manifold differential evolution (MDE): a global optimization method for geodesic centroidal Voronoi tessellations on meshes. ACM Transactions on Graphics, 35(6): 243:1–243:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:DDS

[LXY^{+}23] Zhehao Li, Qingyu Xu, Xiaohan Ye, Bo Ren, and Ligang Liu. DiffFR: Differentiable SPH-based fluidrigid coupling for rigid body control. ACM Transactions on Graphics, 42(6): 179:1–179:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618318.

Liu:2019:NRR

[LXZ^{+}19] Lingjie Liu, Weipeng Xu, Michael Zollhöfer, Hyeongwoo Kim, Florian Bernard, Marc Habermann, Wenping Wang, and Christian Theobalt. Neural rendering and reenactment of human actor videos. ACM Transactions on Graphics, 38 (5):139:1–139:14, November 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3333002.

Li:2023:PCA

[LXZ^{+}23] Zhansheng Li, Yangyang Xu, Nanxuan Zhao, Yang Zhou, Yongtuo Liu, Dahua Lin, and Shengfeng He. Parsingconditioned anime translation: a new dataset and method. ACM Transactions on Graphics, 42(3): 30:1–30:??, June 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3585002.

Li:2023:GAS

[LY23] Changyang Li and Lap-Fai Yu. Generating activity snippets by learning human-scene interactions. ACM Transactions on Graphics, 42(4): 136:1–136:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592096.

Lin:2018:SAB

[LYC18] You-En Lin, Yong-Liang Yang, and Hung-Kuo Chu. Scale-aware black-and-white abstraction of 3D shapes. ACM Transactions on Graphics, 37(4):117:1–117:??, August 2018. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2022:PGS

[LYC^{+}22] Dawon Lee, Jung Eun Yoo, Kyungmin Cho, Bumki Kim, Gyeonghun Im, and Junyong Noh. PopStage: The generation of stage crossediting video based on spatiotemporal matching. ACM Transactions on Graphics, 41 (6):194:1–194:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555467.

Li:2020:UDD

[LYF^{+}20] Xianzhi Li, Lequan Yu, Chi-Wing Fu, Daniel Cohen-Or, and Pheng-Ann Heng. Unsupervised detection of distinctive regions on 3D shapes. ACM Transactions on Graphics, 39(5):158:1– 158:14, September 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3366785.

Lu:2012:HEB

[LYFD12] Jingwan Lu, Fisher Yu, Adam Finkelstein, and Stephen DiVerdi. HelpingHand: example-based stroke stylization. ACM Transactions on Graphics, 31(4): 46:1–46:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liao:2015:AMD

[LYGC15] Zicheng Liao, Yizhou Yu, Bingchen Gong, and Lechao Cheng. audeosynth: musicdriven video montage. ACM Transactions on Graphics, 34 (4):68:1–68:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Luo:2015:LOL

[LYH^{+}15] Sheng-Jie Luo, Yonghao Yue, Chun-Kai Huang, Yu-Huan Chung, Sei Imai, Tomoyuki Nishita, and Bing-Yu Chen. Legolization: optimizing LEGO designs. ACM Transactions on Graphics, 34 (6):222:1–222:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lan:2021:MIA

[LYK^{+}21] Lei Lan, Yin Yang, Danny Kaufman, Junfeng Yao, Minchen Li, and Chenfanfu Jiang. Medial IPC: accelerated incremental potential contact with medial elastics. ACM Transactions on Graphics, 40(4): 158:1–158:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459753.

LeGendre:2016:PML

[LYL^{+}16] Chloe LeGendre, Xueming Yu, Dai Liu, Jay Busch, Andrew Jones, Sumanta Pattanaik, and Paul Debevec. Practical multispectral lighting reproduction. ACM Transactions on Graphics, 35 (4):32:1–32:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lin:2008:DIS

[LYLL08] Shujin Lin, Fang You, Xiaonan Luo, and Zheng Li. Deducing interpolating subdivision schemes from approximating subdivision schemes. ACM Transactions on Graphics, 27(5):146:1–146:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:PP

[LYNF18] Ligang Liu, Chunyang Ye, Ruiqi Ni, and Xiao-Ming Fu. Progressive parameterizations. ACM Transactions on Graphics, 37(4):41:1– 41:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Livny:2010:ART

[LYO^{+}10] Yotam Livny, Feilong Yan, Matt Olson, Baoquan Chen, Hao Zhang, and Jihad El-Sana. Automatic reconstruction of tree skeletal structures from point clouds. ACM Transactions on Graphics, 29 (6):151:1–151:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2023:SCH

[LYO^{+}23] Zhengqin Li, Li Yu, Mikhail Okunev, Manmohan Chandraker, and Zhao Dong. Spatiotemporally consistent HDR indoor lighting estimation. ACM Transactions on Graphics, 42(3):34:1–34:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3595921.

Lipman:2014:FMB

[LYP^{+}14] Yaron Lipman, Stav Yagev, Roi Poranne, David W. Jacobs, and Ronen Basri. Feature matching with bounded distortion. ACM Transactions on Graphics, 33(3): 26:1–26:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lee:2018:DMC

[LYP^{+}18] Seunghwan Lee, Ri Yu, Jungnam Park, Mridul Aanjaneya, Eftychios Sifakis, and Jehee Lee. Dexterous manipulation and control with volumetric muscles. ACM Transactions on Graphics, 37(4): 57:1–57:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2014:FBI

[LYT^{+}14] Ziwei Liu, Lu Yuan, Xiaoou Tang, Matt Uyttendaele, and Jian Sun. Fast burst images denoising. ACM Transactions on Graphics, 33(6): 232:1–232:??, November 2014.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2013:BCP

[LYTS13] Shuaicheng Liu, Lu Yuan, Ping Tan, and Jian Sun. Bundled camera paths for video stabilization. ACM Transactions on Graphics, 32(4): 78:1–78:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2010:SBC

[LYvdP^{+}10] Libin Liu, KangKang Yin, Michiel van de Panne, Tianjia Shao, and Weiwei Xu. Sampling-based contact-rich motion control. ACM Transactions on Graphics, 29(4): 128:1–128:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2012:TRC

[LYvdPG12] Libin Liu, KangKang Yin, Michiel van de Panne, and Baining Guo. Terrain runner: control, parameterization, composition, and planning for highly dynamic motions. ACM Transactions on Graphics, 31(6):154:1– 154:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2013:SCS

[LYWG13] Libin Liu, KangKang Yin, Bin Wang, and Baining Guo. Simulation and control of skeleton-driven soft body characters. ACM Transactions on Graphics, 32(6): 215:1–215:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liao:2017:VAT

[LYY^{+}17] Jing Liao, Yuan Yao, Lu Yuan, Gang Hua, and Sing Bing Kang. Visual attribute transfer through deep image analogy. ACM Transactions on Graphics, 36(4): 120:1–120:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2013:RFA

[LYYB13] Hao Li, Jihun Yu, Yuting Ye, and Chris Bregler. Realtime facial animation with on-the-fly correctives. ACM Transactions on Graphics, 32(4): 42:1–42:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

LaViola:2004:MSC

[LZ04] Joseph J. LaViola, Jr. and Robert C. Zeleznik. MathPad^{2}: a system for the creation and exploration of mathematical sketches. ACM Transactions on Graphics, 23(3):

432–440, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Levi:2014:SMG

[LZ14] Zohar Levi and Denis Zorin.

Strict minimizers for geometric optimization. ACM Transactions on Graphics, 33(6): 185:1–185:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2021:SMI

[LZBCJ21] Hsueh-Ti Derek Liu, Jiayi Eris Zhang, Mirela Ben-Chen, and Alec Jacobson. Surface multigrid via intrinsic prolongation. ACM Transactions on Graphics, 40 (4):80:1–80:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459768.

Lee:2011:SRT

[LZC11] Yong Jae Lee, C. Lawrence Zitnick, and Michael F. Cohen. ShadowDraw: realtime user guidance for freehand drawing. ACM Transactions on Graphics, 30(4): 27:1–27:9, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2018:SCO

[LZC^{+}18] Heng Liu, Paul Zhang, Edward Chien, Justin Solomon, and David Bommes. Singularityconstrained octahedral fields for hexahedral meshing. ACM Transactions on Graphics, 37 (4):93:1–93:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ling:2020:CCU

[LZCV20] Hung Yu Ling, Fabio Zinno, George Cheng, and Michiel Van De Panne. Character controllers using motion VAEs. ACM Transactions on Graphics, 39(4): 40:1–40:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392422.

Lian:2019:ESL

[LZCX19] Zhouhui Lian, Bo Zhao, Xudong Chen, and Jianguo Xiao. EasyFont: a style learning-based system to easily build your largescale handwriting fonts. ACM Transactions on Graphics, 38 (1):6:1–6:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3213767.

Lentine:2010:NAI

[LZF10] Michael Lentine, Wen Zheng, and Ronald Fedkiw. A novel algorithm for incompressible flow using only a coarse grid projection. ACM Transactions on Graphics, 29(4): 114:1–114:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2019:CPA

[LZF^{+}19] Hao Liu, Xiao-Teng Zhang, Xiao-Ming Fu, Zhi-Chao Dong, and Ligang Liu. Computational peeling art design. ACM Transactions on Graphics, 38(4):64:1–64:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Lun:2017:LGD

[LZH^{+}17] Zhaoliang Lun, Changqing Zou, Haibin Huang, Evangelos Kalogerakis, Ping Tan, Marie-Paule Cani, and Hao Zhang. Learning to group discrete graphical patterns. ACM Transactions on Graphics, 36 (6):225:1–225:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2020:RTI

[LZH^{+}20] Wei Liu, Pingping Zhang, Xiaolin Huang, Jie Yang, Chunhua Shen, and Ian Reid. Realtime image smoothing via iterative least squares. ACM Transactions on Graphics, 39 (3):28:1–28:24, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3388887.

Loubet:2020:SSI

[LZHJ20] Guillaume Loubet, Tizian Zeltner, Nicolas Holzschuch, and Wenzel Jakob. Slopespace integrals for specular next event estimation. ACM Transactions on Graphics, 39 (6):239:1–239:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417811.

Langlois:2016:TAW

[LZJ16] Timothy R. Langlois, Changxi Zheng, and Doug L. James. Toward animating water with complex acoustic bubbles. ACM Transactions on Graphics, 35(4):95:1–95:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2010:EOI

[LZKW10] Yuanyuan Li, Eugene Zhang, Yoshihiro Kobayashi, and Peter Wonka. Editing operations for irregular vertices in triangle meshes. ACM Transactions on Graphics, 29(6): 153:1–153:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Li:2022:NNR

[LZQ^{+}22] Yuwei Li, Longwen Zhang, Zesong Qiu, Yingwenqi Jiang, Nianyi Li, Yuexin Ma, Yuyao Zhang, Lan Xu, and Jingyi Yu. NIMBLE: a non-rigid hand model with bones and muscles. ACM Transactions on Graphics, 41(4): 120:1–120:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530079.

Li:2021:IAH

[LZS^{+}21] Lingxiao Li, Paul Zhang, Dmitriy Smirnov, S. Mazdak Abulnaga, and Justin Solomon. Interactive allhex meshing via cuboid decomposition. ACM Transactions on Graphics, 40(6): 256:1–256:17, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480568.

Lehtinen:2008:MHR

[LZT^{+}08] Jaakko Lehtinen, Matthias Zwicker, Emmanuel Turquin, Janne Kontkanen, Frédo Durand, François X. Sillion, and Timo Aila. A meshless hierarchical representation for light transport. ACM Transactions on Graphics, 27(3): 37:1–37:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Liu:2019:NAS

[LZT^{+}19] Lijuan Liu, Youyi Zheng, Di Tang, Yi Yuan, Changjie Fan, and Kun Zhou. NeuroSkinning: automatic skin binding for production characters with deep graph networks. ACM Transactions on Graphics, 38(4):114:1– 114:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Luo:2021:TTR

[LZY^{+}21] Xuan Luo, Xuaner (Cecilia) Zhang, Paul Yoo, Ricardo Martin-Brualla, Jason Lawrence, and Steven M. Seitz. Time-travel rephotography. ACM Transactions on Graphics, 40(6): 213:1–213:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480485.

Liu:2021:MFS

[LZZ^{+}21] Zhong-Yuan Liu, Zhan Zhang, Di Zhang, Chunyang Ye, Ligang Liu, and Xiao-Ming Fu. Modeling and fabrication with specified discrete equivalence classes. ACM Transactions on Graphics, 40 (4):41:1–41:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459843.

Meyer:2006:SAA

[MA06] Mark Meyer and John Anderson. Statistical acceleration for animated global illumination. ACM Transactions on Graphics, 25(3): 1075–1080, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meyer:2007:KPS

[MA07] Mark Meyer and John Anderson. Key Point Subspace Acceleration and soft caching. ACM Transactions on Graphics, 26(3):74:1–74:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Musialski:2015:ROS

[MAB^{+}15] Przemyslaw Musialski, Thomas Auzinger, Michael Birsak, Michael Wimmer, and Leif Kobbelt. Reduced-order shape optimization using offset surfaces. ACM Transactions on Graphics, 34(4): 102:1–102:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mackinlay:1986:ADG

[Mac86] Jock Mackinlay. Automating the design of graphical presentations of relational information. ACM Transactions on Graphics, 5(2):110–141, April 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/

toc/Abstracts/0730-0301/
22950.html.

Mantiuk:2022:SUM

[MAC22] Rafał K. Mantiuk, Maliha Ashraf, and Alexandre Chapiro. stelaCSF: a unified model of contrast sensitivity as the function of spatiotemporal frequency, eccentricity, luminance and area. ACM Transactions on Graphics, 41 (4):145:1–145:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530115.

Masia:2009:ERT

[MAF^{+}09] Belen Masia, Sandra Agustin, Roland W. Fleming, Olga Sorkine, and Diego Gutierrez. Evaluation of reverse tone mapping through varying exposure conditions. ACM Transactions on Graphics, 28(5):160:1–160:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matusik:2009:PSV

[MAG^{+}09] Wojciech Matusik, Boris Ajdin, Jinwei Gu, Jason Lawrence, Hendrik P. A. Lensch, Fabio Pellacini, and Szymon Rusinkiewicz. Printing spatially-varying reflectance. ACM Transactions on Graphics, 28(5):128:1–128:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maillot:1992:NFM

[Mai92]  Patrick-Gilles Maillot.  A

new, fast method for 2-D polygon clipping: Analysis and software implementation. ACM Transactions on Graphics, 11(3):276–290, July 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
130894.html.

Mercier-Aubin:2022:ARE

[MAKWL22] Alexandre Mercier-Aubin, Paul G. Kry, Alexandre Winter, and David I. W. Levin. Adaptive rigidification of elastic solids. ACM Transactions on Graphics, 41(4):71:1–71:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530124.

Mallet:1989:DSI

[Mal89] Jean-Laurent Mallet. Discrete smooth interpolation. ACM Transactions on Graphics, 8(2):121–144, April 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
62057.html.

Malzbender:1993:FVR

[Mal93] Tom Malzbender. Fourier volume rendering. ACM Transactions on Graphics, 12(3):233–250, July 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        169705.html.

Mantyla:1986:BOM

[Man86] Martti Mantyla. Boolean operations of 2-manifolds through vertex neighborhood classification. ACM Transactions on Graphics, 5(1): 1–29, January 1986. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        7530.html.

McCann:2016:CMK

[MAN^{+}16] James McCann, Lea Albaugh, Vidya Narayanan, April Grow, Wojciech Matusik, Jennifer Mankoff, and Jessica Hodgins. A compiler for 3D machine knitting. ACM Transactions on Graphics, 35 (4):49:1–49:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mullapudi:2016:ASH

[MAS^{+}16] Ravi Teja Mullapudi, Andrew Adams, Dillon Sharlet, Jonathan Ragan-Kelley, and Kayvon Fatahalian. Automatically scheduling halide image processing pipelines. ACM Transactions on Graphics, 35 (4):83:1–83:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitchell:2015:NML

[MASS15] Nathan Mitchell, Mridul Aanjaneya, Rajsekhar Setaluri, and Eftychios Sifakis. Nonmanifold level sets: a multivalued implicit surface representation with applications to self-collision processing. ACM Transactions on Graphics, 34 (6):247:1–247:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mahdavi-Amiri:2020:VVD

[MAYZ^{+}20] Ali Mahdavi-Amiri, Fenggen Yu, Haisen Zhao, Adriana Schulz, and Hao Zhang. VDAC: volume decomposeand-carve for subtractive manufacturing. ACM Transactions on Graphics, 39(6): 203:1–203:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417772.

Misztal:2012:TAI

[MB12] Marek Krzysztof Misztal and Jakob Andreas Bærentzen. Topology-adaptive interface tracking using the deformable simplicial complex. ACM Transactions on Graphics, 31(3):24:1–24:12, May 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Michel:2021:DAI

[MB21] Élie Michel and Tamy Boubekeur. DAG amendment for inverse control of parametric shapes. ACM Transactions on Graphics, 40(4): 173:1–173:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459823.

McDonnell:2012:RMR

[MBB12] Rachel McDonnell, Martin Breidt, and Heinrich H. Bülthoff. Render me real?: investigating the effect of render style on the perception of animated virtual humans. ACM Transactions on Graphics, 31(4):91:1–91:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Molino:2004:VNA

[MBF04] Neil Molino, Zhaosheng Bao, and Ron Fedkiw. A virtual node algorithm for changing mesh topology during simulation. ACM Transactions on Graphics, 23(3):385– 392, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Misso:2022:UCR

[MBGJ22] Zackary Misso, Benedikt Bitterli, Iliyan Georgiev, and Wojciech Jarosz. Unbiased and consistent rendering using biased estimators. ACM Transactions on Graphics, 41(4):48:1–48:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530160.

Martin-Brualla:2015:TLM

[MBGS15] Ricardo Martin-Brualla, David Gallup, and Steven M. Seitz. Time-lapse mining from Internet photos. ACM Transactions on Graphics, 34(4): 62:1–62:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moon:2010:COR

[MBK^{+}10] Bochang Moon, Yongyoung Byun, Tae-Joon Kim, Pio Claudio, Hye-Sun Kim, Yun-Ji Ban, Seung Woo Nam, and Sung-Eui Yoon. Cacheoblivious ray reordering. ACM Transactions on Graphics, 29 (3):28:1–28:10, June 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martin-Brualla:2018:LEP

[MBPY^{+}18] Ricardo Martin-Brualla, Rohit Pandey, Shuoran Yang, Pavel Pidlypenskyi, Jonathan Taylor, Julien Valentin, Sameh Khamis, Philip Davidson, Anastasia Tkach, Peter Lincoln, Adarsh Kowdle, Christoph Rhemann, Dan B. Goldman, Cem Keskin, Steve Seitz, Shahram Izadi, and Sean Fanello. LookinGood: enhancing performance capture with real-time neural re-rendering. ACM Transactions on Graphics, 37(6): 255:1–255:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maharik:2011:DM

[MBS^{+}11] Ron Maharik, Mikhail Bessmeltsev, Alla Sheffer, Ariel Shamir, and Nathan Carr. Digital micrography. ACM Transactions on Graphics, 30 (4):100:1–100:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mercier:2015:STP

[MBT^{+}15] Olivier Mercier, Cynthia Beauchemin, Nils Thuerey, Theodore Kim, and Derek Nowrouzezahrai. Surface turbulence for particle-based liquid simulations. ACM Transactions on Graphics, 34(6): 202:1–202:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

TOG-126230047

[MBU22] Mostafa Morsy Abdelkader Morsy, Alan Brunton, and Philipp Urban. Shape dithering for 3D printing. ACM Transactions on Graphics, 41(4):82:1–82:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530129.

Museth:2002:LSS

[MBWB02] Ken Museth, David E. Breen, Ross T. Whitaker, and Alan H. Barr. Level set surface editing operators. ACM Transactions on Graphics, 21 (3):330–338, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

MacIntyre:1992:PAC

[MC92] Blair MacIntyre and William B.

Cowan. A practical approach to calculating luminance contrast on a CRT. ACM Transactions on Graphics, 11(4):336–347, October 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
146467.html.

Muller:2011:SSO

[MC11] Matthias Müller and Nuttapong Chentanez. Solid simulation with oriented particles. ACM Transactions on Graphics, 30(4):92:1– 92:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Min:2012:MGC

[MC12] Jianyuan Min and Jinxiang Chai. Motion graphs++: a compact generative model for semantic motion analysis and synthesis. ACM Transactions on Graphics, 31(6): 153:1–153:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mandad:2021:GQH

[MC21] Manish Mandad and Marcel Campen. Guaranteed-quality higher-order triangular meshing of 2D domains. ACM Transactions on Graphics, 40(4):154:1–154:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459673.

McCool:1999:ADM

[McC99] Michael D. McCool. Anisotropic diffusion for Monte Carlo noise reduction. ACM Transactions on Graphics, 18(2): 171–194, April 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1999-18-2/
p171-mccool/.

McCool:2000:SVR

[McC00] Michael D. McCool. Shadow volume reconstruction from depth maps. ACM Transactions on Graphics, 19(1): 1–26, January 2000. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/2000-19-1/
p1-mccool/.

Min:2009:IGH

[MCC09] Jianyuan Min, Yen-Lin Chen, and Jinxiang Chai. Interactive generation of human animation with deformable motion models. ACM Transactions on Graphics, 29(1):9:1– 9:12, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matzen:2017:LCS

[MCE^{+}17] Kevin Matzen, Michael F. Cohen, Bryce Evans, Johannes Kopf, and Richard Szeliski. Low-cost 360 stereo photography and video capture. ACM Transactions on Graphics, 36 (4):148:1–148:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Munkberg:2006:HDR

[MCHAM06] Jacob Munkberg, Petrik Clarberg, Jon Hasselgren, and Tomas Akenine-Möller. High dynamic range texture compression for graphics hardware. ACM Transactions on Graphics, 25(3):698–706, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McIlroy:1983:BAC

[McI83] M. McIlroy. Best approximate circles on integer grids. ACM Transactions on Graphics, 2 (4):237–263, October 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McIlroy:1992:GRE

[McI92] M. Douglas McIlroy. Getting raster ellipses right. ACM Transactions on Graphics, 11(3):259–275, July 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
130892.html.

McKenna:1987:WCO

[McK87] Michael McKenna. Worstcase optimal hidden-surface removal. ACM Transactions on Graphics, 6(1):19– 28, January 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
27627.html.

Muller:2013:RTD

[MCK13] Matthias Müller, Nuttapong Chentanez, and Tae-Yong Kim. Real time dynamic fracture with volumetric approximate convex decompositions. ACM Transactions on Graphics, 32(4):115:1–115:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2015:AMR

[MCKM15] Matthias Müller, Nuttapong Chentanez, Tae-Yong Kim, and Miles Macklin. Air meshes for robust collision handling. ACM Transactions on Graphics, 34(4):133:1– 133:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitra:2009:EI

[MCL^{+}09] Niloy J. Mitra, Hung-Kuo Chu, Tong-Yee Lee, Lior Wolf, Hezy Yeshurun, and Daniel Cohen-Or. Emerging images. ACM Transactions on Graphics, 28(5): 163:1–163:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mullen:2009:EPI

[MCP^{+}09] Patrick Mullen, Keenan Crane, Dmitry Pavlov, Yiying Tong, and Mathieu Desbrun. Energypreserving integrators for fluid animation. ACM Transactions on Graphics, 28(3):38:1– 38:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitchell:2015:GIA

[MCS15] Nathan Mitchell, Court Cutting, and Eftychios Sifakis. GRIDiron: an interactive authoring and cognitive training foundation for reconstructive plastic surgery procedures. ACM Transactions on Graphics, 34(4):43:1–43:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mandad:2015:IAW

[MCSA15] Manish Mandad, David Cohen-Steiner, and Pierre Alliez. Isotopic approximation within a tolerance volume. ACM Transactions on Graphics, 34 (4):64:1–64:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mandad:2017:VMT

[MCSK^{+}17] Manish Mandad, David Cohen-Steiner, Leif Kobbelt, Pierre Alliez, and Mathieu Desbrun. Variance-minimizing transport plans for intersurface mapping. ACM Transactions on Graphics, 36(4): 39:1–39:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moreno:2015:USL

[MCT15] Daniel Moreno, Fatih Calakli, and Gabriel Taubin. Unsynchronized structured light. ACM Transactions on Graphics, 34(6):178:1–178:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moser:2021:SSV

[MCW^{+}21] Lucio Moser, Chinyu Chien, Mark Williams, Jose Serra, Darren Hendler, and Doug Roble. Semi-supervised videodriven facial animation transfer for production. ACM Transactions on Graphics, 40 (6):222:1–222:18, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480515.

Moon:2014:ARB

[MCY14] Bochang Moon, Nathan Carr, and Sung-Eui Yoon. Adaptive rendering based on weighted local regression. ACM Transactions on Graphics, 33(5): 170:1–170:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Manocha:1994:AIP

[MD94] Dinesh Manocha and James Demmel. Algorithms for intersecting parametric and algebraic curves I: Simple intersections. ACM Transactions on Graphics, 13(1):73– 100, January 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
174617.html.

Mueller:2019:RTP

[MDB^{+}19] Franziska Mueller, Micah Davis, Florian Bernard, Oleksandr Sotnychenko, Mickeal Verschoor, Miguel A. Otaduy, Dan Casas, and Christian Theobalt. Real-time pose and shape reconstruction of two interacting hands with a single depth camera. ACM Transactions on Graphics, 38(4): 49:1–49:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mantiuk:2021:FVD

[MDC^{+}21] Rafał K. Mantiuk, Gyorgy Denes, Alexandre Chapiro, Anton Kaplanyan, Gizem Rufo, Romain Bachy, Trisha Lian, and Anjul Patney. FovVideoVDP: a visible difference predictor for wide field-of-view video. ACM Transactions on Graphics, 40 (4):49:1–49:19, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459831.

Maestre:2023:DSP

[MDH^{+}23] Juan Sebastian Montes Maestre, Yinwei Du, Ronan Hinchet, Stelian Coros, and Bernhard Thomaszewski. Differentiable stripe patterns for inverse design of structured surfaces. ACM Transactions on Graphics, 42(4): 102:1–102:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592114.

Mantiuk:2008:DAT

[MDK08] Rafał Mantiuk, Scott Daly, and Louis Kerofsky. Display adaptive tone mapping. ACM Transactions on Graphics, 27 (3):68:1–68:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maron:2016:PRE

[MDK^{+}16] Haggai Maron, Nadav Dym, Itay Kezurer, Shahar Kovalsky, and Yaron Lipman. Point registration via efficient convex relaxation. ACM Transactions on Graphics, 35(4): 73:1–73:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Murmann:2016:CBF

[MDKD16] Lukas Murmann, Abe Davis, Jan Kautz, and Frédo Durand. Computational bounce flash for indoor portraits. ACM Transactions on Graphics, 35(6):190:1–190:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martinez:2016:PVF

[MDL16] Jonàs Martínez, Jérémie Dumas, and Sylvain Lefebvre. Procedural Voronoi foams for additive manufacturing. ACM Transactions on Graphics, 35 (4):44:1–44:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mordatch:2010:RPB

[MdLH10] Igor Mordatch, Martin de Lasa, and Aaron Hertzmann. Robust physics-based locomotion using low-dimensional planning. ACM Transactions on Graphics, 29(4): 71:1–71:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martinez:2015:SAO

[MDLW15] Jonàs Martínez, Jérémie Dumas, Sylvain Lefebvre, and Li-Yi Wei. Structure and appearance optimization for controllable shape design. ACM Transactions on Graphics, 34 (6):229:1–229:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McCool:2004:SA

[MDP^{+}04] Michael McCool, Stefanus Du Toit, Tiberiu Popa, Bryan Chan, and Kevin Moule. Shader algebra. ACM Transactions on Graphics, 23(3): 787–795, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2021:DDC

[MDZ^{+}21] Pingchuan Ma, Tao Du, John Z. Zhang, Kui Wu, Andrew Spielberg, Robert K. Katzschmann, and Wojciech Matusik. DiffAqua: a differentiable computational design pipeline for soft underwater swimmers with shape interpolation. ACM Transactions on Graphics, 40(4): 132:1–132:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459832.

Mora:2005:LCM

[ME05] Benjamin Mora and David S.

Ebert. Low-complexity maximum intensity projection. ACM Transactions on Graphics, 24(4):1392–1416, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitchell:2018:SDH

[MEA^{+}18] Scott A. Mitchell, Mohamed S. Ebeida, Muhammad A. Awad, Chonhyon Park, Anjul Patney, Ahmad A. Rushdi, Laura P. Swiler, Dinesh Manocha, and Li-Yi Wei. Spoke-darts for high-dimensional blue-noise sampling. ACM Transactions on Graphics, 37(2):

22:1–22:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Macklin:2019:NSN

[MEM^{+}19] Miles Macklin, Kenny Erleben, Matthias Müller, Nuttapong Chentanez, Stefan Jeschke, and Viktor Makoviychuk. Non-smooth Newton methods for deformable multibody dynamics. ACM Transactions on Graphics, 38(5): 140:1–140:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mantiuk:2006:BCH

[MEMS06] Rafał Mantiuk, Alexander Efremov, Karol Myszkowski, and Hans-Peter Seidel. Backward compatible high dynamic range MPEG video compression. ACM Transactions on Graphics, 25(3): 713–723, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Merrell:2023:EBP

[Mer23]  Paul Merrell.  Example-

based procedural modeling using graph grammars. ACM Transactions on Graphics, 42 (4):60:1–60:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592119.

Muller:2022:ING

[MESK22] Thomas Müller, Alex Evans, Christoph Schied, and Alexander Keller. Instant neural graphics primitives with a multiresolution hash encoding. ACM Transactions on Graphics, 41(4):102:1– 102:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530127.

Meyer:1991:LTO

[Mey91] Alan Meyer. A linear time Oslo algorithm. ACM Transactions on Graphics, 10(3): 312–318, July 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
108552.html.

Matsuda:2017:FSD

[MFL17] Nathan Matsuda, Alexander Fix, and Douglas Lanman. Focal surface displays. ACM Transactions on Graphics, 36 (4):86:1–86:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2010:DDI

[MFR^{+}10] Kerstin Müller, Christoph Fünfzig, Lars Reusche, Dianne Hansford, Gerald Farin, and Hans Hagen. DINUS:

Double Insertion, Nonuniform, Stationary subdivision surfaces. ACM Transactions on Graphics, 29(3): 25:1–25:21, June 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mohr:2003:BEA

[MG03] Alex Mohr and Michael Gleicher. Building efficient, accurate character skins from examples. ACM Transactions on Graphics, 22(3):562–568, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maron:2017:CNN

[MGA^{+}17] Haggai Maron, Meirav Galun, Noam Aigerman, Miri Trope, Nadav Dym, Ersin Yumer, Vladimir G. Kim, and Yaron Lipman. Convolutional neural networks on surfaces via seamless toric covers. ACM Transactions on Graphics, 36 (4):71:1–71:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2022:SFD

[MGA^{+}22] Karima Ma, Michael Gharbi, Andrew Adams, Shoaib Kamil, Tzu-Mao Li, Connelly Barnes, and Jonathan Ragan-Kelley. Searching for fast demosaicking algorithms. ACM Transactions on Graphics, 41(5): 172:1–172:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3508461.

Mark:2003:CSP

[MGAK03] William R. Mark, R. Steven Glanville, Kurt Akeley, and Mark J. Kilgard. Cg: a system for programming graphics hardware in a C-like language. ACM Transactions on Graphics, 22(3):896–907, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Monszpart:2019:IIG

[MGC^{+}19] Aron Monszpart, Paul Guerrero, Duygu Ceylan, Ersin Yumer, and Niloy J. Mitra. iMapper: interaction-guided scene mapping from monocular videos. ACM Transactions on Graphics, 38(4): 92:1–92:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Montalto:2015:TVA

[MGDA^{+}15] Carlos Montalto, Ignacio Garcia-Dorado, Daniel Aliaga, Manuel M. Oliveira, and Feng Meng. A total variation approach for customizing imagery to improve visual acuity. ACM Transactions on Graphics, 34(3):28:1–28:??, April 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

MacIntyre:2005:DTR

[MGDB05] Blair MacIntyre, Maribeth Gandy, Steven Dow, and Jay David Bolter. DART: a toolkit for rapid design exploration of augmented reality experiences. ACM Transactions on Graphics, 24 (3):932, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miller:2019:NSP

[MGJ19] Bailey Miller, Iliyan Georgiev, and Wojciech Jarosz. A nullscattering path integral formulation of light transport. ACM Transactions on Graphics, 38(4):44:1–44:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maimone:2017:HNE

[MGK17] Andrew Maimone, Andreas Georgiou, and Joel S. Kollin. Holographic near-eye displays for virtual and augmented reality. ACM Transactions on Graphics, 36(4): 85:1–85:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitra:2006:PAS

[MGP06] Niloy J. Mitra, Leonidas J. Guibas, and Mark Pauly. Partial and approximate symmetry detection for 3D geometry. ACM Transactions on Graphics, 25(3):560–568, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitra:2007:S

[MGP07] Niloy J. Mitra, Leonidas J.

Guibas, and Mark Pauly. Symmetrization. ACM Transactions on Graphics, 26(3): 63:1–63:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miklos:2010:DSA

[MGP10] Balint Miklos, Joachim Giesen, and Mark Pauly. Discrete scale axis representations for 3D geometry. ACM Transactions on Graphics, 29(4): 101:1–101:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Makatura:2021:PGE

[MGS^{+}21] Liane Makatura, Minghao Guo, Adriana Schulz, Justin Solomon, and Wojciech Matusik. Pareto gamuts: exploring optimal designs across varying contexts. ACM Transactions on Graphics, 40(4): 171:1–171:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459750.

Munzner:2003:TST

[MGT^{+}03] Tamara Munzner, François Guimbretière, Serdar Tasiran, Li Zhang, and Yunhong Zhou. TreeJuxtaposer: scalable tree comparison using Focus+Context with guaranteed visibility. ACM Transactions on Graphics, 22(3): 453–462, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Montazeri:2020:PPB

[MGZJ20] Zahra Montazeri, Søren B. Gammelmark, Shuang Zhao, and Henrik Wann Jensen. A practical ply-based appearance model of woven fabrics. ACM Transactions on Graphics, 39(6): 251:1–251:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417777.

Munkberg:2016:TSC

[MHC^{+}16] Jacob Munkberg, Jon Hasselgren, Petrik Clarberg, Magnus Andersson, and Tomas Akenine-Möller. Texture space caching and reconstruction for ray tracing. ACM Transactions on Graphics, 35 (6):249:1–249:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maestre:2023:TTO

[MHCT23] Juan Montes Maestre, Ronan Hinchet, Stelian Coros, and Bernhard Thomaszewski. ToRoS: a topology optimization approach for designing robotic skins. ACM Transactions on Graphics, 42(6): 194:1–194:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618382.

Metzer:2021:SSN

[MHGCO21] Gal Metzer, Rana Hanocka, Raja Giryes, and Daniel Cohen-Or. Self-sampling for neural point cloud consolidation. ACM Transactions on Graphics, 40(5): 191:1–191:14, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3470645.

Mahajan:2009:MGP

[MHM^{+}09] Dhruv Mahajan, Fu-Chung Huang, Wojciech Matusik, Ravi Ramamoorthi, and Peter Belhumeur. Moving gradients: a path-based method for plausible image interpolation. ACM Transactions on Graphics, 28(3):42:1–42:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marco:2017:DSR

[MHM^{+}17] Julio Marco, Quercus Hernandez, Adolfo Muñoz, Yue Dong, Adrian Jarabo, Min H. Kim, Xin Tong, and Diego Gutierrez. DeepToF: offthe-shelf real-time correction of multipath interference in time-of-flight imaging. ACM Transactions on Graphics, 36 (6):219:1–219:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mazhar:2015:UNM

[MHNT15] Hammad Mazhar, Toby Heyn, Dan Negrut, and Alessandro Tasora. Using Nesterov’s method to accelerate multibody dynamics with friction and contact. ACM Transactions on Graphics, 34(3): 32:1–32:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meka:2019:DRF

[MHP^{+}19] Abhimitra Meka, Christian Häne, Rohit Pandey, Michael Zollhöfer, Sean Fanello, Graham Fyffe, Adarsh Kowdle, Xueming Yu, Jay Busch, Jason Dourgarian, Peter Denny, Sofien Bouaziz, Peter Lincoln, Matt Whalen, Geoff Harvey, Jonathan Taylor, Shahram Izadi, Andrea Tagliasacchi, Paul Debevec, Christian Theobalt, Julien Valentin, and Christoph Rhemann. Deep reflectance fields: high-quality facial reflectance field inference from color gradient illumination. ACM Transactions on Graphics, 38 (4):77:1–77:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Musialski:2016:NLS

[MHR^{+}16] Przemyslaw Musialski, Christian Hafner, Florian Rist, Michael Birsak, Michael Wimmer, and Leif Kobbelt. Nonlinear shape optimization using local subspace projections. ACM Transactions on Graphics, 35(4):87:1–87:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2019:CSS

[MHS^{+}19a] Long Ma, Ying He, Qian Sun, Yuanfeng Zhou, Caiming Zhang, and Wenping Wang. Constructing 3D self-supporting surfaces with isotropic stress using 4D minimal hypersurfaces of revolution. ACM Transactions on Graphics, 38(5):144:1– 144:??, October 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Makowski:2019:SSM

[MHS^{+}19b] Miłosz Makowski, Torsten Hädrich, Jan Scheffczyk, Dominik L. Michels, Sören Pirk, and Wojtek Pałubicki. Synthetic silviculture: multi-scale modeling of plant ecosystems. ACM Transactions on Graphics, 38(4):131:1–131:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martinez:2018:PVD

[MHSL18] Jonàs Martínez, Samuel Hornus, Haichuan Song, and Sylvain Lefebvre. Polyhedral Voronoi diagrams for additive manufacturing. ACM Transactions on Graphics, 37(4): 129:1–129:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2005:MDB

[MHTG05] Matthias Müller, Bruno Heidelberger, Matthias Teschner, and Markus Gross. Meshless deformations based on shape matching. ACM Transactions on Graphics, 24(3): 471–478, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miandji:2019:UFC

[MHU19] Ehsan Miandji, Saghi Hajisharif, and Jonas Unger. A unified framework for compression and compressed sensing of light fields and light field videos. ACM Transactions on Graphics, 38(3): 23:1–23:??, June 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.

acm.org/ft_gateway.cfm?
id=3269980.

Mara:2021:TSH

[MHZ^{+}21a] Michael Mara, Felix Heide, Michael Zollhöfer, Matthias Nießner, and Pat Hanrahan. Thallo — scheduling for high-performance largescale non-linear least-squares solvers. ACM Transactions on Graphics, 40(5): 184:1–184:14, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3453986.

Metzer:2021:OPC

[MHZ^{+}21b] Gal Metzer, Rana Hanocka, Denis Zorin, Raja Giryes, Daniele Panozzo, and Daniel Cohen-Or. Orienting point clouds with dipole propagation. ACM Transactions on Graphics, 40(4): 165:1–165:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459835.

Mori:2007:PID

[MI07] Yuki Mori and Takeo Igarashi.

Plushie: an interactive design system for plush toys. ACM Transactions on Graphics, 26 (3):45:1–45:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miki:2015:PSS

[MIB15] Masaaki Miki, Takeo Igarashi, and Philippe Block. Parametric self-supporting surfaces via direct computation of Airy stress functions. ACM Transactions on Graphics, 34(4): 89:1–89:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moon:2015:ARL

[MIGYM15] Bochang Moon, Jose A. Iglesias-Guitian, Sung-Eui Yoon, and Kenny Mitchell. Adaptive rendering with linear predictions. ACM Transactions on Graphics, 34(4): 121:1–121:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miller:1987:GAN

[Mil87] James R. Miller. Geometric approaches to nonplanar quadric surface intersection curves. ACM Transactions on Graphics, 6(4):274–307, October 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        35041.html.

Mirtich:1998:VCF

[Mir98] Brian Mirtich. V-Clip: fast and robust polyhedral collision detection. ACM Transactions on Graphics, 17(3):

177–208, July 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1998-17-3/
p177-mirtich/.

Mitra:2018:SDF

[Mit18] Niloy Mitra. Session details: Fun in geometry & fabrication. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meehan:2002:PMP

[MIWB02] Michael Meehan, Brent Insko, Mary Whitton, and Frederick P. Brooks, Jr. Physiological measures of presence in stressful virtual environments. ACM Transactions on Graphics, 21(3):645–652, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miyashita:2016:ZSP

[MIWI16] Leo Miyashita, Kota Ishihara, Yoshihiro Watanabe, and Masatoshi Ishikawa. Zoe-Matrope: a system for physical material design. ACM Transactions on Graphics, 35 (4):66:1–66:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Milliron:2002:FGW

[MJBF02] Tim Milliron, Robert J. Jensen, Ronen Barzel, and Adam Finkelstein. A framework for geometric warps and deformations. ACM Transactions on Graphics, 21(1): 20–51, January 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marschner:2003:LSH

[MJC^{+}03] Stephen R. Marschner, Henrik Wann Jensen, Mike Cammarano, Steve Worley, and Pat Hanrahan. Light scattering from human hair fibers. ACM Transactions on Graphics, 22(3):780–791, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2008:FPS

[MJC^{+}08] Wan-Chun Ma, Andrew Jones, Jen-Yuan Chiang, Tim Hawkins, Sune Frederiksen, Pieter Peers, Marko Vukovic, Ming Ouhyoung, and Paul Debevec. Facial performance synthesis using deformationdriven polynomial displacement maps. ACM Transactions on Graphics, 27(5): 121:1–121:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marco:2018:SOO

[MJJG18] Julio Marco, Adrian Jarabo, Wojciech Jarosz, and Diego Gutierrez. Second-order occlusion-aware volumetric radiance caching. ACM Transactions on Graphics, 37 (2):20:1–20:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mukai:2005:GMI

[MK05] Tomohiko Mukai and Shigeru Kuriyama. Geostatistical motion interpolation. ACM Transactions on Graphics, 24 (3):1062–1070, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mukai:2016:EDS

[MK16] Tomohiko Mukai and Shigeru Kuriyama. Efficient dynamic skinning with low-rank helper bone controllers. ACM Transactions on Graphics, 35(4): 36:1–36:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martin:2010:USE

[MKB^{+}10] Sebastian Martin, Peter Kaufmann, Mario Botsch, Eitan Grinspun, and Markus Gross. Unified simulation of elastic rods, shells, and solids. ACM Transactions on Graphics, 29 (4):39:1–39:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Manzi:2016:TGD

[MKD^{+}16] Marco Manzi, Markus Kettunen, Frédo Durand, Matthias Zwicker, and Jaakko Lehtinen. Temporal gradientdomain path tracing. ACM Transactions on Graphics, 35 (6):246:1–246:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mantiuk:2004:PMH

[MKMS04] Rafal Mantiuk, Grzegorz Krawczyk, Karol Myszkowski, and Hans-Peter Seidel. Perceptionmotivated high dynamic range video encoding. ACM Transactions on Graphics, 23(3): 733–741, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mantiuk:2011:HVC

[MKRH11] Rafat Mantiuk, Kil Joong Kim, Allan G. Rempel, and Wolfgang Heidrich. HDR-VDP-2: a calibrated visual metric for visibility and quality predictions in all luminance conditions. ACM Transactions on Graphics, 30 (4):40:1–40:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2021:FFS

[MKZ^{+}21] Xiaohe Ma, Kaizhang Kang, Ruisheng Zhu, Hongzhi Wu, and Kun Zhou. Free-form scanning of non-planar appearance with neural trace photography. ACM Transactions on Graphics, 40(4): 124:1–124:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459679.

Madan:2022:FES

[ML22] Abhishek Madan and David I. W. Levin. Fast evaluation of smooth distance constraints on co-dimensional geometry. ACM Transactions on Graphics, 41(4): 68:1–68:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530093.

Miguel:2016:CDS

[MLB16] Eder Miguel, Mathias Lepoutre, and Bernd Bickel. Computational design of stable planar-rod structures. ACM Transactions on Graphics, 35(4):86:1–86:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McDonnell:2008:CAP

[MLD^{+}08] Rachel McDonnell, Michéal Larkin, Simon Dobbyn, Steven Collins, and Carol O’Sullivan. Clone attack! Perception of crowd variety. ACM Transactions on Graphics, 27(3):

26:1–26:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McDonnell:2009:ECC

[MLH^{+}09] Rachel McDonnell, Michéal Larkin, Benjamín Hernández, Isaac Rudomin, and Carol O’Sullivan. Eye-catching crowds: saliency based selective variation. ACM Transactions on Graphics, 28(3): 55:1–55:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martel:2021:AAC

[MLL^{+}21] Julien N. P. Martel, David B. Lindell, Connor Z. Lin, Eric R. Chan, Marco Monteiro, and Gordon Wetzstein. Acorn: adaptive coordinate networks for neural scene representation. ACM Transactions on Graphics, 40(4): 58:1–58:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459785.

Ma:2022:NPD

[MLL^{+}22] Li Ma, Xiaoyu Li, Jing Liao, Xuan Wang, Qi Zhang, Jue Wang, and Pedro V. Sander. Neural parameterization for dynamic human head editing. ACM Transactions on Graphics, 41(6): 236:1–236:??, December 2022.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555494.

Muico:2009:CAN

[MLPP09] Uldarico Muico, Yongjoon Lee, Jovan Popović, and Zoran Popović. Contact-aware nonlinear control of dynamic characters. ACM Transactions on Graphics, 28(3):81:1– 81:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maimone:2014:PDW

[MLR^{+}14] Andrew Maimone, Douglas Lanman, Kishore Rathinavel, Kurtis Keller, David Luebke, and Henry Fuchs. Pinlight displays: wide field of view augmented reality eyeglasses using defocused point light sources. ACM Transactions on Graphics, 33(4): 89:1–89:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mercier:2022:MLD

[MLR^{+}22] Corentin Mercier, Thibault Lescoat, Pierre Roussillon, Tamy Boubekeur, and Jean-Marc Thiery. Moving levelof-detail surfaces. ACM Transactions on Graphics, 41 (4):130:1–130:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530151.

Muntoni:2018:AAH

[MLS^{+}18] Alessandro Muntoni, Marco Livesu, Riccardo Scateni, Alla Sheffer, and Daniele Panozzo. Axis-aligned heightfield block decomposition of 3D shapes. ACM Transactions on Graphics, 37(5): 169:1–169:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3204458.

Michels:2017:SAI

[MLT17] Dominik L. Michels, Vu Thai Luan, and Mayya Tokman. A stiffly accurate integrator for elastodynamic problems. ACM Transactions on Graphics, 36(4):116:1–116:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maia:2019:LOB

[MLYZ19] Henrique Teles Maia, Dingzeyu Li, Yuan Yang, and Changxi Zheng. LayerCode: optical barcodes for 3D printed shapes. ACM Transactions on Graphics, 38(4):112:1– 112:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2016:ADI

[MLZ^{+}16] Rui Ma, Honghua Li, Changqing Zou, Zicheng Liao, Xin Tong, and Hao Zhang. Actiondriven 3D indoor scene evolution. ACM Transactions on Graphics, 35(6):173:1– 173:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moon:2006:SMS

[MM06] Jonathan T. Moon and Stephen R. Marschner. Simulating multiple scattering in hair using a photon mapping approach. ACM Transactions on Graphics, 25(3): 1067–1074, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Merrell:2008:CMS

[MM08] Paul Merrell and Dinesh Manocha. Continuous model synthesis. ACM Transactions on Graphics, 27(5): 158:1–158:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Macklin:2013:PBF

[MM13] Miles Macklin and Matthias Müller. Position based fluids. ACM Transactions on Graphics, 32(4):104:1–104:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miki:2022:IET

[MM22] Masaaki Miki and Toby Mitchell. Interactive exploration of tension-compression mixed shells. ACM Transactions on Graphics, 41(6): 263:1–263:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555438.

Monszpart:2015:RRM

[MMBM15] Aron Monszpart, Nicolas Mellado, Gabriel J. Brostow, and Niloy J. Mitra. RAPter: rebuilding man-made scenes with regular arrangements of planes. ACM Transactions on Graphics, 34(4):103:1– 103:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Macklin:2014:UPP

[MMCK14] Miles Macklin, Matthias Müller, Nuttapong Chentanez, and Tae-Yong Kim. Unified particle physics for real-time applications. ACM Transactions on Graphics, 33 (4):153:1–153:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mullen:2011:HHO

[MMdGD11] Patrick Mullen, Pooran Memari, Fernando de Goes, and Mathieu Desbrun. HOT:

Hodge-optimized triangulations. ACM Transactions on Graphics, 30(4):103:1– 103:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Merry:2006:AST

[MMG06] Bruce Merry, Patrick Marais, and James Gain. Animation space: a truly linear framework for character animation. ACM Transactions on Graphics, 25(4):1400–1423, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Montano-Murillo:2023:OLL

[MMHP23] Roberto Montano-Murillo, Ryuji Hirayama, and Diego Martinez Plasencia. Open-MPD: a low-level presentation engine for multimodal particle-based displays. ACM Transactions on Graphics, 42 (2):24:1–24:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3572896.

Moon:2016:APR

[MMMG16] Bochang Moon, Steven Mc-Donagh, Kenny Mitchell, and Markus Gross. Adaptive polynomial rendering. ACM Transactions on Graphics, 35 (4):40:1–40:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McGuire:2005:DVM

[MMP^{+}05] Morgan McGuire, Wojciech Matusik, Hanspeter Pfister, John F. Hughes, and Frédo Durand. Defocus video matting. ACM Transactions on Graphics, 24(3):567–576, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2019:NIS

[MMR^{+}19] Thomas Müller, Brian Mcwilliams, Fabrice Rousselle, Markus Gross, and Jan Novák. Neural importance sampling. ACM Transactions on Graphics, 38(5):145:1–145:??, October 2019. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3341156.

Meyron:2018:LPG

[MMT18] Jocelyn Meyron, Quentin Mérigot, and Boris Thibert. Light in power: a general and parameter-free algorithm for caustic design. ACM Transactions on Graphics, 37(6): 224:1–224:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maggiordomo:2023:MMC

[MMT23] Andrea Maggiordomo, Henry Moreton, and Marco Tarini. Micro-mesh construction. ACM Transactions on Graphics, 42(4):121:1–121:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592440.

Mullen:2007:VAE

[MMTD07] Patrick Mullen, Alexander McKenzie, Yiying Tong, and Mathieu Desbrun. A variational approach to Eulerian geometry processing. ACM Transactions on Graphics, 26 (3):66:1–66:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Myronova:2023:DOS

[MNB23] Mariia Myronova, William Neveu, and Mikhail Bessmeltsev. Differential operators on sketches via alpha contours. ACM Transactions on Graphics, 42(4): 69:1–69:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592420.

Moreno-Noguer:2007:ARI

[MNBN07] Francesc Moreno-Noguer, Peter N. Belhumeur, and Shree K. Nayar. Active refocusing of images and videos. ACM Transactions on Graphics, 26(3):67:1–67:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mueller:2021:TAS

[MNV^{+}21] Joerg H. Mueller, Thomas Neff, Philip Voglreiter, Markus Steinberger, and Dieter Schmalstieg. Temporally adaptive shading reuse for real-time rendering and virtual reality. ACM Transactions on Graphics, 40(2):11:1–11:14, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3446790.

Mora:2011:NRT

[Mor11] Benjamin Mora. Naive raytracing: a divide-and-conquer approach. ACM Transactions on Graphics, 30(5): 117:1–117:12, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Melzi:2018:DTE

[MOR^{+}18] Simone Melzi, Maks Ovsjanikov, Giorgio Roffo, Marco Cristani, and Umberto Castellani. Discrete time evolution process descriptor for shape analysis and matching. ACM Transactions on Graphics, 37 (1):4:1–4:??, January 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matusik:2004:TSS

[MP04] Wojciech Matusik and Hanspeter Pfister. 3D TV: a scalable system for real-time acquisition, transmission, and autostereoscopic display of dynamic scenes. ACM Transactions on Graphics, 23(3): 814–824, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McCann:2007:RCM

[MP07] James McCann and Nancy Pollard. Responsive characters from motion fragments. ACM Transactions on Graphics, 26(3):6:1–6:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McCann:2008:RTG

[MP08] James McCann and Nancy S.

Pollard. Real-time gradientdomain painting. ACM Transactions on Graphics, 27(3): 93:1–93:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McCann:2009:LL

[MP09a] James McCann and Nancy Pollard. Local layering. ACM Transactions on Graphics, 28 (3):84:1–84:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitra:2009:SA

[MP09b] Niloy J. Mitra and Mark Pauly. Shadow art. ACM Transactions on Graphics, 28(5):156:1–156:7, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Myles:2009:BPS

[MP09c] Ashish Myles and Jörg Peters. Bi-3 C^{2} polar subdivision. ACM Transactions on Graphics, 28(3):48:1–48:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Montanari:2017:IGAa

[MPB17a] Mattia Montanari, Nik Petrinic, and Ettore Barbieri. Improving the GJK algorithm for faster and more reliable distance queries between convex objects. ACM Transactions on Graphics, 36(3): 30:1–30:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Montanari:2017:IGAb

[MPB17b] Mattia Montanari, Nik Petrinic, and Ettore Barbieri. Improving the GJK algorithm for faster and more reliable distance queries between convex objects. ACM Transactions on Graphics, 36(4): 151:1–151:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Malomo:2016:FAD

[MPBC16] Luigi Malomo, Nico Pietroni, Bernd Bickel, and Paolo Cignoni. FlexMolds: automatic design of flexible shells for molding. ACM Transactions on Graphics, 35(6): 223:1–223:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matusik:2003:DDR

[MPBM03] Wojciech Matusik, Hanspeter Pfister, Matt Brand, and Leonard McMillan. A datadriven reflectance model. ACM Transactions on Graphics, 22(3):759–769, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Masselus:2003:RIL

[MPDW03] Vincent Masselus, Pieter Peers, Philip Dutré, and Yves D. Willems. Relighting with 4D incident light fields. ACM Transactions on Graphics, 22(3):613–620, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mendiratta:2023:ATD

[MPE^{+}23] Mohit Mendiratta, Xingang Pan, Mohamed Elgharib, Kartik Teotia, Mallikarjun B R, Ayush Tewari, Vladislav Golyanik, Adam Kortylewski, and Christian Theobalt. AvatarStudio: Text-driven editing of 3D dynamic human head avatars. ACM Transactions on Graphics, 42(6):226:1–226:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618368.

Ma:2018:LDS

[MPF^{+}18] Rui Ma, Akshay Gadi Patil, Matthew Fisher, Manyi Li, Sören Pirk, Binh-Son Hua, Sai-Kit Yeung, Xin Tong, Leonidas Guibas, and Hao Zhang. Language-driven synthesis of 3D scenes from scene databases. ACM Transactions on Graphics, 37(6): 212:1–212:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2016:ERH

[MPG^{+}16] Thomas Müller, Marios Papas, Markus Gross, Wojciech Jarosz, and Jan Novák. Efficient rendering of heterogeneous polydisperse granular media. ACM Transactions on Graphics, 35(6): 168:1–168:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meng:2015:MSM

[MPH^{+}15] Johannes Meng, Marios Papas, Ralf Habel, Carsten Dachsbacher, Steve Marschner, Markus Gross, and Wojciech Jarosz. Multi-scale modeling and rendering of granular materials. ACM Transactions on Graphics, 34(4): 49:1–49:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meka:2020:DRT

[MPH^{+}20] Abhimitra Meka, Rohit Pandey, Christian Häne, Sergio Orts-Escolano, Peter Barnum, Philip David-Son, Daniel Erickson, Yinda Zhang, Jonathan Taylor, Sofien Bouaziz, Chloe Legendre, Wan-Chun Ma, Ryan Overbeck, Thabo Beeler, Paul Debevec, Shahram Izadi, Christian Theobalt, Christoph Rhemann, and Sean Fanello. Deep relightable textures: volumetric performance capture with neural rendering. ACM Transactions on Graphics, 39(6): 259:1–259:21, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417814.

Malomo:2018:FCD

[MPI^{+}18] Luigi Malomo, Jesús Pérez, Emmanuel Iarussi, Nico Pietroni, Eder Miguel, Paolo Cignoni, and Bernd Bickel. FlexMaps: computational design of flat flexible shells for shaping 3D objects. ACM Transactions on Graphics, 37 (6):241:1–241:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mohammed:2009:VLG

[MPK09] Umar Mohammed, Simon J. D. Prince, and Jan Kautz. Visio-lization: generating novel facial images. ACM Transactions on Graphics, 28 (3):57:1–57:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Myles:2010:FAM

[MPKZ10] Ashish Myles, Nico Pietroni, Denis Kovacs, and Denis Zorin. Feature-aligned T - meshes. ACM Transactions on Graphics, 29(4):117:1– 117:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matusik:2002:IBP

[MPN^{+}02] Wojciech Matusik, Hanspeter Pfister, Addy Ngan, Paul Beardsley, Remo Ziegler, and Leonard McMillan. Imagebased 3D photography using opacity hulls. ACM Transactions on Graphics, 21(3): 427–437, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mahmoud:2021:RGM

[MPO21] Ahmed H. Mahmoud, Serban D. Porumbescu, and John D. Owens. RXMesh: a GPU mesh data structure. ACM Transactions on Graphics, 40(4):104:1– 104:16, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459748.

Muico:2011:CCP

[MPP11] Uldarico Muico, Jovan Popović, and Zoran Popović. Composite control of physically simulated characters. ACM Transactions on Graphics, 30(3): 16:1–16:11, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Myles:2014:RFA

[MPZ14] Ashish Myles, Nico Pietroni, and Denis Zorin. Robust fieldaligned global parametrization. ACM Transactions on Graphics, 33(4):135:1– 135:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mehra:2013:WBS

[MRA^{+}13] Ravish Mehra, Nikunj Raghuvanshi, Lakulish Antani, Anish Chandak, Sean Curtis, and Dinesh Manocha. Wavebased sound propagation in large open scenes using an equivalent source formulation. ACM Transactions on Graphics, 32(2):19:1–19:13, April 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matveev:2022:DDE

[MRA^{+}22] Albert Matveev, Ruslan Rakhimov, Alexey Artemov, Gleb Bobrovskikh, Vage Egiazarian, Emil Bogomolov, Daniele Panozzo, Denis Zorin, and Evgeny Burnaev. DEF: deep estimation of sharp geometric features in 3D shapes. ACM Transactions on Graphics, 41(4):108:1–108:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530140.

Meyer:1986:EEC

[MRC^{+}86] Gary W. Meyer, Holly E. Rushmeier, Michael F. Cohen, Donald P. Greenberg, and Kenneth E. Torrance. An experimental evaluation of computer graphics imagery. ACM Transactions on Graphics, 5(1):30–50, January 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
7920.html.

Muller:2005:ECB

[MRC05] Meinard Müller, Tido Röder, and Michael Clausen. Efficient content-based retrieval of motion capture data. ACM Transactions on Graphics, 24 (3):677–685, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2006:ESS

[MRF06] Kerstin Müller, Lars Reusche, and Dieter Fellner. Extended subdivision surfaces: Building a bridge between NURBS and Catmull–Clark surfaces. ACM Transactions on Graphics, 25 (2):268–292, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Manakov:2013:RCA

[MRK^{+}13] Alkhazur Manakov, John F.

Restrepo, Oliver Klehm, Ramon Hegedüs, Elmar Eisemann, Hans-Peter Seidel, and Ivo Ihrke. A reconfigurable camera add-on for high dynamic range, multispectral, polarization, and light-field imaging. ACM Transactions on Graphics, 32(4): 47:1–47:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Manzi:2014:ISG

[MRK^{+}14] Marco Manzi, Fabrice Rousselle, Markus Kettunen, Jaakko Lehtinen, and Matthias Zwicker. Improved sampling for gradient-domain Metropolis light transport. ACM Transactions on Graphics, 33 (6):178:1–178:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2020:NCV

[MRKN20] Thomas Müller, Fabrice Rousselle, Alexander Keller, and Jan Novák. Neural control variates. ACM Transactions on Graphics, 39(6): 243:1–243:19, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417804.

Muller:2021:RTN

[MRNK21] Thomas Müller, Fabrice Rousselle, Jan Novák, and Alexander Keller. Real-time neural radiance caching for path tracing. ACM Transactions on Graphics, 40(4): 36:1–36:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459812.

Mitani:2004:MPT

[MS04] Jun Mitani and Hiromasa Suzuki. Making papercraft toys from meshes using stripbased approximate unfolding. ACM Transactions on Graphics, 23(3):259–263, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Majumder:2005:PPS

[MS05] Aditi Majumder and Rick Stevens. Perceptual photometric seamlessness in projection-based tiled displays. ACM Transactions on Graphics, 24(1):118–139, January 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Manson:2013:CCT

[MS13] Josiah Manson and Scott Schaefer. Cardinality-constrained texture filtering. ACM Transactions on Graphics, 32(4): 140:1–140:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miller:2023:BVC

[MSCG23] Bailey Miller, Rohan Sawhney, Keenan Crane, and Ioannis Gkioulekas. Boundary value caching for walk on spheres. ACM Transactions on Graphics, 42(4): 82:1–82:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592400.

Martinez:2017:ONF

[MSDL17] Jonàs Martínez, Haichuan Song, Jérémie Dumas, and Sylvain Lefebvre. Orthotropic k-nearest foams for additive manufacturing. ACM Transactions on Graphics, 36(4): 121:1–121:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martinet:2006:ADS

[MSHS06] Aurélien Martinet, Cyril Soler, Nicolas Holzschuch, and François X. Sillion. Accurate detection of symmetries in 3D shapes. ACM Transactions on Graphics, 25(2): 439–464, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Merrell:2010:CGR

[MSK10] Paul Merrell, Eric Schkufza, and Vladlen Koltun. Computergenerated residential building layouts. ACM Transactions on Graphics, 29(6): 181:1–181:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maloisel:2023:ODR

[MSK^{+}23] Guirec Maloisel, Christian Schumacher, Espen Knoop, Ruben Grandia, and Moritz Bächer. Optimal design of robotic character kinematics. ACM Transactions on Graphics, 42(6): 195:1–195:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618404.

Merrell:2011:IFL

[MSL^{+}11] Paul Merrell, Eric Schkufza, Zeyang Li, Maneesh Agrawala, and Vladlen Koltun. Interactive furniture layout using interior design guidelines. ACM Transactions on Graphics, 30 (4):87:1–87:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mahapatra:2023:TGS

[MSL^{+}23] Aniruddha Mahapatra, Aliaksandr Siarohin, Hsin-Ying Lee, Sergey Tulyakov, and Jun-Yan Zhu. Text-guided synthesis of Eulerian cinemagraphs. ACM Transactions on Graphics, 42(6): 259:1–259:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618326.

Menapace:2024:PGM

[MSL^{+}24] Willi Menapace, Aliaksandr Siarohin, Stéphane Lathuilière, Panos Achlioptas, Vladislav Golyanik, Sergey Tulyakov, and Elisa Ricci. Promptable game models: Text-guided game simulation via masked diffusion models. ACM Transactions on Graphics, 43(2):17:1–17:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3635705.

McCrae:2011:SSP

[MSM11] James McCrae, Karan Singh, and Niloy J. Mitra. Slices: a shape-proxy based on planar sections. ACM Transactions on Graphics, 30(6): 168:1–168:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mercier:2017:FGC

[MSM^{+}17] Olivier Mercier, Yusufu Sulai, Kevin Mackenzie, Marina Zannoli, James Hillis, Derek Nowrouzezahrai, and Douglas Lanman. Fast gaze-contingent optimal decompositions for multifocal displays. ACM Transactions on Graphics, 36 (6):237:1–237:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2023:SPC

[MSM^{+}23] Sizhuo Ma, Varun Sundar, Paul Mos, Claudio Bruschini, Edoardo Charbon, and Mohit Gupta. Seeing photons in color. ACM Transactions on Graphics, 42(4): 99:1–99:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592438.

Mildenhall:2019:LLF

[MSOC^{+}19] Ben Mildenhall, Pratul P. Srinivasan, Rodrigo Ortiz-Cayon, Nima Khademi Kalantari, Ravi Ramamoorthi, Ren Ng, and Abhishek Kar. Local light field fusion: practical view synthesis with prescriptive sampling guidelines. ACM Transactions on Graphics, 38(4):29:1–29:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mashayekhi:2018:ADE

[MSQ^{+}18] Omid Mashayekhi, Chinmayee Shah, Hang Qu, Andrew Lim, and Philip Levis. Automatically distributing Eulerian and hybrid fluid simulations in the cloud. ACM Transactions on Graphics, 37 (2):24:1–24:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mahajan:2007:TLL

[MSRB07] Dhruv Mahajan, Ira Kemelmacher Shlizerman, Ravi Ramamoorthi, and Peter Belhumeur. A theory of locally low dimensional light transport. ACM Transactions on Graphics, 26(3): 62:1–62:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meyers:1992:SC

[MSS92] David Meyers, Shelley Skinner, and Kenneth Sloan. Surfaces from contours. ACM Transactions on Graphics, 11(3):228–258, July 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/

                                        toc/Abstracts/0730-0301/
                                        131213.html.

Malzbender:2012:PRF

[MSS^{+}12] Tom Malzbender, Ramin Samadani, Steven Scher, Adam Crume, Douglas Dunn, and James Davis. Printing reflectance functions. ACM Transactions on Graphics, 31(3):20:1–20:11, May 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mehta:2017:VRT

[MSS^{+}17] Dushyant Mehta, Srinath Sridhar, Oleksandr Sotnychenko, Helge Rhodin, Mohammad Shafiei, Hans-Peter Seidel, Weipeng Xu, Dan Casas, and Christian Theobalt. VNect: real-time 3D human pose estimation with a single RGB camera. ACM Transactions on Graphics, 36(4): 44:1–44:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martinez:2019:SSM

[MSS^{+}19] Jonàs Martínez, Mélina Skouras, Christian Schumacher, Samuel Hornus, Sylvain Lefebvre, and Bernhard Thomaszewski. Star-shaped metrics for mechanical metamaterial design. ACM Transactions on Graphics, 38(4):82:1–82:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mo:2021:GVS

[MSSG^{+}21] Haoran Mo, Edgar Simo-Serra, Chengying Gao, Changqing Zou, and Ruomei Wang. General virtual sketching framework for vector line art. ACM Transactions on Graphics, 40 (4):51:1–51:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459833.

Middleditch:1989:IAL

[MST89] A. E. Middleditch, T. W.

Stacey, and S. B. Tor. Intersection algorithms for lines and circles. ACM Transactions on Graphics, 8(1): 25–40, January 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

49157.html. See corrigenda [Bak94].

McAdams:2009:DPC

[MSW^{+}09] Aleka McAdams, Andrew Selle, Kelly Ward, Eftychios Sifakis, and Joseph Teran. Detail preserving continuum simulation of straight hair. ACM Transactions on Graphics, 28 (3):62:1–62:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Michels:2014:EIS

[MSW14] Dominik L. Michels, Gerrit A.

Sobottka, and Andreas G. Weber. Exponential integrators for stiff elastodynamic problems. ACM Transactions on Graphics, 33(1):7:1– 7:20, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Merel:2020:CCR

[MTA^{+}20] Josh Merel, Saran Tunyasuvunakool, Arun Ahuja, Yuval Tassa, Leonard Hasenclever, Vu Pham, Tom Erez, Greg Wayne, and Nicolas Heess. Catch & carry: reusable neural controllers for vision-guided whole-body tasks. ACM Transactions on Graphics, 39(4):39:1– 39:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392474.

Miguel:2013:MEI

[MTB^{+}13] Eder Miguel, Rasmus Tamstorf, Derek Bradley, Sara C. Schvartzman, Bernhard Thomaszewski, Bernd Bickel, Wojciech Matusik, Steve Marschner, and Miguel A. Otaduy. Modeling and estimation of internal friction in cloth. ACM Transactions on Graphics, 32(6): 212:1–212:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Martin:2011:EBE

[MTGG11] Sebastian Martin, Bernhard Thomaszewski, Eitan Grinspun, and Markus Gross. Example-based elastic materials. ACM Transactions on Graphics, 30(4): 72:1–72:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Monszpart:2016:SPG

[MTM16] Aron Monszpart, Nils Thuerey, and Niloy J. Mitra. SMASH: physics-guided reconstruction of collisions from videos. ACM Transactions on Graphics, 35 (6):199:1–199:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Megaro:2015:IDP

[MTN^{+}15] Vittorio Megaro, Bernhard Thomaszewski, Maurizio Nitti, Otmar Hilliges, Markus Gross, and Stelian Coros. Interactive design of 3D-printable robotic creatures. ACM Transactions on Graphics, 34(6): 216:1–216:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mordatch:2012:DCB

[MTP12] Igor Mordatch, Emanuel Todorov, and Zoran Popović. Discovery of complex behaviors through contact-invariant optimization. ACM Transactions on Graphics, 31(4):

43:1–43:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marcias:2015:DDI

[MTP^{+}15] Giorgio Marcias, Kenshi Takayama, Nico Pietroni, Daniele Panozzo, Olga Sorkine-Hornung, Enrico Puppo, and Paolo Cignoni. Data-driven interactive quadrangulation. ACM Transactions on Graphics, 34(4):65:1–65:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2018:FDR

[MTP^{+}18] Pingchuan Ma, Yunsheng Tian, Zherong Pan, Bo Ren, and Dinesh Manocha. Fluid directed rigid body control using deep reinforcement learning. ACM Transactions on Graphics, 37(4):96:1– 96:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McNamara:2004:FCU

[MTPS04] Antoine McNamara, Adrien Treuille, Zoran Popović, and Jos Stam. Fluid control using the adjoint method. ACM Transactions on Graphics, 23(3):449–456, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moroto:2022:CTM

[MU22] Yuji Moroto and Nobuyuki Umetani. Constant time median filter using 2D wavelet matrix. ACM Transactions on Graphics, 41(6): 267:1–267:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555512.

Martin:2015:ODD

[MUB15] Tobias Martin, Nobuyuki Umetani, and Bernd Bickel. OmniAD: data-driven omnidirectional aerodynamics. ACM Transactions on Graphics, 34(4):113:1–113:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Museth:2013:VHR

[Mus13] Ken Museth. VDB: Highresolution sparse volumes with dynamic topology. ACM Transactions on Graphics, 32 (3):27:1–27:22, June 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Meekes:2021:UPS

[MV21] Merel Meekes and Amir Vaxman. Unconventional patterns on surfaces. ACM Transactions on Graphics, 40(4): 101:1–101:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459933.

Mueller:2018:SAS

[MVD^{+}18] Joerg H. Mueller, Philip Voglreiter, Mark Dokter, Thomas Neff, Mina Makar, Markus Steinberger, and Dieter Schmalstieg. Shading atlas streaming. ACM Transactions on Graphics, 37(6): 199:1–199:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mellado:2017:CPS

[MVH^{+}17] Nicolas Mellado, David Vanderhaeghe, Charlotte Hoarau, Sidonie Christophe, Mathieu Brédif, and Loic Barthe. Constrained palette-space exploration. ACM Transactions on Graphics, 36(4): 60:1–60:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marschner:2005:MMA

[MWAM05] Stephen R. Marschner, Stephen H. Westin, Adam Arbree, and Jonathan T. Moon. Measuring and modeling the appearance of finished wood. ACM Transactions on Graphics, 24 (3):727–734, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marwah:2013:CLF

[MWBR13] Kshitij Marwah, Gordon Wetzstein, Yosuke Bando, and Ramesh Raskar. Compressive light field photography using overcomplete dictionaries and optimized projections. ACM Transactions on Graphics, 32 (4):46:1–46:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Makatura:2023:PMU

[MWC^{+}23] Liane Makatura, Bohan Wang, Yi-Lu Chen, Bolei Deng, Chris Wojtan, Bernd Bickel, and Wojciech Matusik. Procedural metamaterials: a unified procedural graph for metamaterial design. ACM Transactions on Graphics, 42(5): 168:1–168:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3605389.

Ma:2009:MFT

[MWGZ09] Chongyang Ma, Li-Yi Wei, Baining Guo, and Kun Zhou. Motion field texture synthesis. ACM Transactions on Graphics, 28(5):110:1– 110:8, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2006:PMB

[MWH^{+}06] Pascal Müller, Peter Wonka, Simon Haegler, Andreas Ulmer, and Luc Van Gool. Procedural modeling of buildings. ACM Transactions on Graphics, 25(3):614–623, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mohan:2009:BIV

[MWH^{+}09] Ankit Mohan, Grace Woo, Shinsaku Hiura, Quinn Smithwick, and Ramesh Raskar. Bokode: imperceptible visual tags for camera based interaction from a distance. ACM Transactions on Graphics, 28 (3):98:1–98:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Maimone:2013:FCA

[MWH^{+}13] Andrew Maimone, Gordon Wetzstein, Matthew Hirsch, Douglas Lanman, Ramesh Raskar, and Henry Fuchs. Focus 3D: Compressive accommodation display. ACM Transactions on Graphics, 32 (5):153:1–153:13, September 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matsuda:2021:VSC

[MWHL21] Nathan Matsuda, Brian Wheelwright, Joel Hegland, and Douglas Lanman. VR social copresence with light field displays. ACM Transactions on Graphics, 40(6): 244:1–244:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480481.

Miyashita:2018:MPM

[MWI18] Leo Miyashita, Yoshihiro Watanabe, and Masatoshi Ishikawa. MIDAS projection: markerless and modelless dynamic projection mapping for material representation. ACM Transactions on Graphics, 37 (6):196:1–196:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2013:DET

[MWLT13] Chongyang Ma, Li-Yi Wei, Sylvain Lefebvre, and Xin Tong. Dynamic element textures. ACM Transactions on Graphics, 32(4): 90:1–90:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Moon:2008:EMS

[MWM08] Jonathan T. Moon, Bruce Walter, and Steve Marschner. Efficient multiple scattering in hair using spherical harmonics. ACM Transactions on Graphics, 27(3):31:1– 31:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2023:USS

[MWM23] Weiyin Ma, Xu Wang, and Yue Ma. An unified λsubdivision scheme for quadrilateral meshes with optimal curvature performance in extraordinary regions. ACM Transactions on Graphics, 42 (6):209:1–209:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618400.

Mehta:2012:AAF

[MWR12] Soham Uday Mehta, Brandon Wang, and Ravi Ramamoorthi. Axis-aligned filtering for interactive sampled soft shadows. ACM Transactions on Graphics, 31(6): 163:1–163:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mehta:2013:AAF

[MWRD13] Soham Uday Mehta, Brandon Wang, Ravi Ramamoorthi, and Fredo Durand. Axisaligned filtering for interactive physically-based diffuse indirect lighting. ACM Transactions on Graphics, 32(4): 96:1–96:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2011:DET

[MWT11] Chongyang Ma, Li-Yi Wei, and Xin Tong. Discrete element textures. ACM Transactions on Graphics, 30(4): 62:1–62:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mordatch:2013:AHL

[MWTK13] Igor Mordatch, Jack M.

Wang, Emanuel Todorov, and Vladlen Koltun. Animating human lower limbs using contact-invariant optimization. ACM Transactions on Graphics, 32(6):203:1– 203:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2023:ODM

[MXZ^{+}23] Xiaohe Ma, Xianmin Xu, Leyao Zhang, Kun Zhou, and Hongzhi Wu. OpenSVBRDF: a database of measured spatially-varying reflectance. ACM Transactions on Graphics, 42(6):254:1–254:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618358.

Men:2022:DND

[MYC^{+}22] Yifang Men, Yuan Yao, Miaomiao Cui, Zhouhui Lian, and Xuansong Xie. DCT-net: domain-calibrated translation for portrait stylization. ACM Transactions on Graphics, 41 (4):140:1–140:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530159.

Moss:2010:SLA

[MYH^{+}10] William Moss, Hengchin Yeh, Jeong-Mo Hong, Ming C. Lin, and Dinesh Manocha. Sounding liquids: Automatic sound synthesis from fluid simulation. ACM Transactions on Graphics, 29(3): 21:1–21:13, June 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mehta:2014:FAA

[MYRD14] Soham Uday Mehta, JiaXian Yao, Ravi Ramamoorthi, and Fredo Durand. Factored axis-aligned filtering for rendering multiple distribution effects. ACM Transactions on Graphics, 33(4): 57:1–57:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Miyashita:2015:MSO

[MYWI15] Leo Miyashita, Ryota Yonezawa, Yoshihiro Watanabe, and Masatoshi Ishikawa. 3D motion sensing of any object without prior knowledge. ACM Transactions on Graphics, 34(6):218:1– 218:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mitra:2010:IHM

[MYY^{+}10] Niloy J. Mitra, Yong-Liang Yang, Dong-Ming Yan, Wilmot Li, and Maneesh Agrawala. Illustrating how mechanical assemblies work. ACM Transactions on Graphics, 29(4): 58:1–58:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Myles:2012:GPI

[MZ12] Ashish Myles and Denis Zorin.

Global parametrization by incremental flattening. ACM Transactions on Graphics, 31 (4):109:1–109:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Myles:2013:CDC

[MZ13] Ashish Myles and Denis Zorin.

Controlled-distortion constrained global parametrization. ACM Transactions on Graphics, 32(4):105:1– 105:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Megaro:2017:CDT

[MZB^{+}17] Vittorio Megaro, Jonas Zehnder, Moritz Bächer, Stelian Coros, Markus Gross, and Bernhard Thomaszewski. A computational design tool for compliant mechanisms. ACM Transactions on Graphics, 36 (4):82:1–82:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Matusik:2005:TDU

[MZD05] Wojciech Matusik, Matthias Zwicker, and Frédo Durand. Texture design using a simplicial complex of morphable textures. ACM Transactions on Graphics, 24(3): 787–794, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mehra:2009:AMM

[MZL^{+}09] Ravish Mehra, Qingnan Zhou, Jeremy Long, Alla Sheffer, Amy Gooch, and Niloy J. Mitra. Abstraction of manmade shapes. ACM Transactions on Graphics, 28(5): 137:1–137:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ma:2017:CDF

[MZL^{+}17] Li-Ke Ma, Yizhonc Zhang, Yang Liu, Kun Zhou, and Xin Tong. Computational design and fabrication of soft pneumatic objects with desired deformations. ACM Transactions on Graphics, 36(6): 239:1–239:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Marschner:2021:SSG

[MZPS21] Zoë Marschner, Paul Zhang, David Palmer, and Justin Solomon. Sum-of-squares geometry processing. ACM Transactions on Graphics, 40 (6):253:1–253:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480551.

Meka:2016:LIV

[MZRT16] Abhimitra Meka, Michael Zollhöfer, Christian Richardt, and Christian Theobalt. Live intrinsic video. ACM Transactions on Graphics, 35(4): 109:1–109:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Macchietto:2009:MCB

[MZS09] Adriano Macchietto, Victor Zordan, and Christian R. Shelton. Momentum control for balance. ACM Transactions on Graphics, 28(3):80:1– 80:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

McAdams:2011:EEC

[MZS^{+}11] Aleka McAdams, Yongning Zhu, Andrew Selle, Mark Empey, Rasmus Tamstorf, Joseph Teran, and Eftychios Sifakis. Efficient elasticity for character skinning with contact and collisions. ACM Transactions on Graphics, 30(4): 37:1–37:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Muller:2007:IBP

[MZWV07] Pascal Müller, Gang Zeng, Peter Wonka, and Luc Van Gool. Image-based procedural modeling of facades. ACM Transactions on Graphics, 26(3): 85:1–85:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narain:2015:OPI

[NAB^{+}15] Rahul Narain, Rachel A. Albert, Abdullah Bulbul, Gregory J. Ward, Martin S. Banks, and James F. O’Brien. Optimal presentation of imagery with focus cues on multi-plane displays. ACM Transactions on Graphics, 34 (4):59:1–59:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narayanan:2018:AMK

[NAH^{+}18] Vidya Narayanan, Lea Albaugh, Jessica Hodgins, Stelian Coros, and James Mccann. Automatic machine knitting of 3D meshes. ACM Transactions on Graphics, 37 (3):35:1–35:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3186265.

Nah:2020:QFE

[Nah20] Jae-Ho Nah. QuickETC2:

Fast ETC2 texture compression using Luma differences. ACM Transactions on Graphics, 39(6): 270:1–270:10, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417787.

Nitzan:2022:MPG

[NAH^{+}22] Yotam Nitzan, Kfir Aberman, Qiurui He, Orly Liba, Michal Yarom, Yossi Gandelsman, Inbar Mosseri, Yael Pritch, and Daniel Cohen-Or. MyStyle: a personalized generative prior. ACM Transactions on Graphics, 41(6): 206:1–206:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555436.

Naiman:1998:JEW

[Nai98] Avi C. Naiman. Jagged edges:

when is filtering needed? ACM Transactions on Graphics, 17(4):238–258, October 1998. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
1998-17-4/p238-naiman/.

Nakashima:2018:CIS

[NAI^{+}18] Kazutaka Nakashima, Thomas Auzinger, Emmanuel Iarussi, Ran Zhang, Takeo Igarashi, and Bernd Bickel. CoreCavity: interactive shell decomposition for fabrication with two-piece rigid molds. ACM Transactions on Graphics, 37(4):135:1–135:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nasri:1987:PSM

[Nas87] Ahmad H. Nasri. Polyhedral subdivision methods for free-form surfaces. ACM Transactions on Graphics, 6(1):29–73, January 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        27628.html.

Nielsen:2011:GSH

[NB11] Michael B. Nielsen and Robert Bridson. Guide shapes for high resolution naturalistic liquid simulation. ACM Transactions on Graphics, 30 (4):83:1–83:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nayar:2004:LSD

[NBB04] Shree K. Nayar, Peter N. Belhumeur, and Terry E. Boult. Lighting sensitive display.

ACM Transactions on Graphics, 23(4):963–979, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nielsen:2022:PBC

[NBHSB22] Michael B. Nielsen, Morten Bojsen-Hansen, Konstantinos Stamatelos, and Robert Bridson. Physics-based combustion simulation. ACM Transactions on Graphics, 41(5): 176:1–176:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3526213.

Nitzan:2020:FID

[NBLCO20] Yotam Nitzan, Amit Bermano, Yangyan Li, and Daniel Cohen-Or. Face identity disentanglement via latent space mapping. ACM Transactions on Graphics, 39(6): 225:1–225:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417826.

Nigolian:2023:ECP

[NCB23] Valentin Zénon Nigolian, Marcel Campen, and David Bommes. Expansion cones: a progressive volumetric mapping framework. ACM Transactions on Graphics, 42(4): 131:1–131:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592421.

Nonato:2005:BCG

[NCVMO05] Luis Gustavo Nonato, Alex Jesus Cuadros-Vargas, Rosane Minghim, and Maria Cristina F. De Oliveira. Betaconnection: Generating a family of models from planar cross sections. ACM Transactions on Graphics, 24 (4):1239–1258, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nehme:2023:TMQ

[NDD^{+}23] Yana Nehmé, Johanna Delanoy, Florent Dupont, Jean-Philippe Farrugia, Patrick Le Callet, and Guillaume Lavoué. Textured mesh quality assessment: Large-scale dataset and deep learningbased quality metric. ACM Transactions on Graphics, 42 (3):31:1–31:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592786.

Nimier-David:2022:UIV

[NDMKJ22] Merlin Nimier-David, Thomas Müller, Alexander Keller, and Wenzel Jakob. Unbiased inverse volume rendering with differential trackers. ACM Transactions on Graphics, 41(4):44:1–44:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530073.

Nsampi:2023:NFC

[NDS^{+}23] Ntumba Elie Nsampi, Adarsh Djeacoumar, Hans-Peter Seidel, Tobias Ritschel, and Thomas Leimkühler. Neural field convolutions by repeated differentiation. ACM Transactions on Graphics, 42(6): 206:1–206:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618340.

Nagano:2015:SMD

[NFA^{+}15] Koki Nagano, Graham Fyffe, Oleg Alexander, Jernej Barbiç, Hao Li, Abhijeet Ghosh, and Paul Debevec. Skin microstructure deformation with displacement map convolution. ACM Transactions on Graphics, 34(4):109:1– 109:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Neubert:2007:AIB

[NFD07] Boris Neubert, Thomas Franken, and Oliver Deussen. Approximate image-based treemodeling using particle flows. ACM Transactions on Graphics, 26(3):88:1–88:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nguyen:2002:PBM

[NFJ02] Duc Quang Nguyen, Ronald Fedkiw, and Henrik Wann Jensen. Physically based modeling and animation of fire. ACM Transactions on Graphics, 21(3):721–728, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Niu:2012:EWS

[NFL12] Yuzhen Niu, Wu-Chi Feng, and Feng Liu. Enabling warping on stereoscopic images. ACM Transactions on Graphics, 31(6):183:1– 183:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ng:2005:FSP

[Ng05] Ren Ng. Fourier slice photography. ACM Transactions on Graphics, 24(3):735–744, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nader:2018:ITM

[NG18] Georges Nader and Gael Guennebaud. Instant transport maps on 2D grids. ACM Transactions on Graphics, 37 (6):249:1–249:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narain:2009:ADD

[NGCL09] Rahul Narain, Abhinav Golas, Sean Curtis, and Ming C.

Lin. Aggregate dynamics for dense crowd simulation. ACM Transactions on Graphics, 28(5):122:1–122:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narasimhan:2006:ASP

[NGD^{+}06] Srinivasa G. Narasimhan, Mohit Gupta, Craig Donner, Ravi Ramamoorthi, Shree K. Nayar, and Henrik Wann Jensen. Acquiring scattering properties of participating media by dilution. ACM Transactions on Graphics, 25 (3):1003–1012, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nishida:2016:ISU

[NGDA^{+}16] Gen Nishida, Ignacio Garcia-Dorado, Daniel G. Aliaga, Bedrich Benes, and Adrien Bousseau. Interactive sketching of urban procedural models. ACM Transactions on Graphics, 35(4):130:1– 130:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ni:2004:FMF

[NGH04] Xinlai Ni, Michael Garland, and John C. Hart. Fair Morse functions for extracting the topological structure of a surface mesh. ACM Transactions on Graphics, 23(3): 613–622, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narain:2010:FFG

[NGL10] Rahul Narain, Abhinav Golas, and Ming C. Lin. Free-flowing granular materials with twoway solid coupling. ACM Transactions on Graphics, 29 (6):173:1–173:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nehab:2008:RAR

[NH08] Diego Nehab and Hugues Hoppe. Random-access rendering of general vector graphics. ACM Transactions on Graphics, 27(5):135:1–135:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nasikun:2022:HSI

[NH22] Ahmad Nasikun and Klaus Hildebrandt. The hierarchical subspace iteration method for Laplace–Beltrami eigenproblems. ACM Transactions on Graphics, 41(2):17:1– 17:14, April 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3495208.

Niederauer:2003:NII

[NHAH03] Christopher Niederauer, Mike Houston, Maneesh Agrawala, and Greg Humphreys. Noninvasive interactive visualization of dynamic architectural environments. ACM Transactions on Graphics, 22 (3):700, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Noris:2013:TDV

[NHS^{+}13] Gioacchino Noris, Alexander Hornung, Robert W. Sumner, Maryann Simmons, and Markus Gross. Topologydriven vectorization of clean line drawings. ACM Transactions on Graphics, 32(1): 4:1–4:11, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nagata:2022:ELD

[NI22] Yuichi Nagata and Shinji Imahori. Escherization with large deformations based on as-rigid-as-possible shape modeling. ACM Transactions on Graphics, 41(2):11:1– 11:16, April 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3487017.

Nagata:2024:CDE

[NI24] Yuichi Nagata and Shinji Imahori. Creation of dihedral Escher-like tilings based on asrigid-as-possible deformation. ACM Transactions on Graphics, 43(2):18:1–18:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3638048.

Nabata:2020:RAW

[NID20] Kosuke Nabata, Kei Iwasaki, and Yoshinori Dobashi. Resamplingaware weighting functions for bidirectional path tracing using multiple light sub-paths. ACM Transactions on Graphics, 39(2):15:1–15:11, April 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3338994.

Nindel:2021:GBF

[NIR^{+}21] Thomas Klaus Nindel, Tomás Iser, Tobias Rittig, Alexander Wilkie, and Jaroslav Krivánek. A gradient-based framework for 3D print appearance optimization. ACM Transactions on Graphics, 40(4):178:1–178:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3450626.3459844.

Nealen:2007:FDF

[NISA07] Andrew Nealen, Takeo Igarashi, Olga Sorkine, and Marc Alexa. FiberMesh: designing freeform surfaces with 3D curves. ACM Transactions on Graphics, 26(3): 41:1–41:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nicolet:2021:LSI

[NJJ21] Baptiste Nicolet, Alec Jacobson, and Wenzel Jakob. Large steps in inverse rendering of geometry. ACM Transactions on Graphics, 40(6): 248:1–248:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480501.

Nielsen:2015:OMB

[NJR15] Jannik Boll Nielsen, Henrik Wann Jensen, and Ravi Ramamoorthi. On optimal, minimal BRDF sampling for reflectance acquisition. ACM Transactions on Graphics, 34 (6):186:1–186:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nowrouzezahrai:2011:PSA

[NJS^{+}11] Derek Nowrouzezahrai, Jared Johnson, Andrew Selle, Dylan Lacewell, Michael Kaschalk, and Wojciech Jarosz. A programmable system for artistic volumetric lighting. ACM Transactions on Graphics, 30 (4):29:1–29:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Neff:2008:GMA

[NKAS08] Michael Neff, Michael Kipp, Irene Albrecht, and Hans-Peter Seidel. Gesture modeling and animation based on a probabilistic re-creation of speaker style. ACM Transactions on Graphics, 27(1): 5:1–5:24, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nayar:2006:FSD

[NKGR06] Shree K. Nayar, Gurunandan Krishnan, Michael D. Grossberg, and Ramesh Raskar. Fast separation of direct and global components of a scene using high frequency illumination. ACM Transactions on Graphics, 25(3):935–944, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nesme:2009:PTE

[NKJF09] Matthieu Nesme, Paul G. Kry, Lenka Jeřábková, and François Faure. Preserving topology and elasticity for embedded deformable models. ACM Transactions on Graphics, 28(3):52:1–52:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nah:2014:RRT

[NKK^{+}14] Jae-Ho Nah, Hyuck-Joo Kwon, Dong-Seok Kim, Cheol-Ho Jeong, Jinhong Park, Tack-Don Han, Dinesh Manocha, and Woo-Chan Park. Ray-Core: A ray-tracing hardware architecture for mobile devices. ACM Transactions on Graphics, 33(5): 162:1–162:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nam:2023:DHP

[NKKJ23] Seung-Woo Nam, Youngjin Kim, Dongyeon Kim, and Yoonchan Jeong. Depolarized holography with polarizationmultiplexing metasurface. ACM Transactions on Graphics, 42(6):202:1–202:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618395.

Narumi:2023:IPS

[NKS^{+}23] Koya Narumi, Kazuki Koyama, Kai Suto, Yuta Noma, Hiroki Sato, Tomohiro Tachi, Masaaki Sugimoto, Takeo Igarashi, and Yoshihiro Kawahara. Inkjet 4D print: Selffolding tessellated origami objects by inkjet UV printing. ACM Transactions on Graphics, 42(4):117:1– 117:??, August 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592409.

Niessner:2013:ADM

[NL13] Matthias Nießner and Charles Loop. Analytic displacement mapping using hardware tessellation. ACM Transactions on Graphics, 32(3): 26:1–26:9, June 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nam:2018:PSA

[NLGK18] Giljoo Nam, Joo Ho Lee, Diego Gutierrez, and Min H. Kim. Practical SVBRDF acquisition of 3D objects with unstructured flash photography. ACM Transactions on Graphics, 37(6):267:1– 267:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Niessner:2012:FAG

[NLMD12] Matthias Nießner, Charles Loop, Mark Meyer, and Tony Derose. Feature-adaptive GPU rendering of Catmull– Clark subdivision surfaces. ACM Transactions on Graphics, 31(1):6:1–6:11, January 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nam:2016:SAM

[NLW^{+}16] Giljoo Nam, Joo Ho Lee, Hongzhi Wu, Diego Gutierrez, and Min H. Kim. Simultaneous acquisition of microscale reflectance and normals. ACM Transactions on Graphics, 35 (6):185:1–185:??, November 2016. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Nehab:2016:PRF

[NM16] Diego Nehab and André Maximo. Parallel recursive filtering of infinite input extensions. ACM Transactions on Graphics, 35(6):204:1– 204:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nageli:2017:RTP

[NMD^{+}17] Tobias Nägeli, Lukas Meier, Alexander Domahidi, Javier Alonso-Mora, and Otmar Hilliges. Real-time planning for automated multi-view drone cinematography. ACM Transactions on Graphics, 36 (4):132:1–132:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nehab:2011:GER

[NMLH11] Diego Nehab, André Maximo, Rodolfo S. Lima, and Hugues Hoppe. GPU-efficient recursive filtering and summedarea tables. ACM Transactions on Graphics, 30(6): 176:1–176:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See errata [NMLH14].

Nehab:2014:EGE

[NMLH14] Diego Nehab, André Maximo, Rodolfo S. Lima, and Hugues Hoppe. Errata for GPUefficient recursive filtering and summed-area tables. ACM Transactions on Graphics, 33(3):33:1–33:??, May 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [NMLH11].

Nicholl:1990:PGT

[NN90] Robin A. Nicholl and Tina M. Nicholl. Performing geometric transformations by program transformation. ACM Transactions on Graphics, 9(1):28–40, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77637.html.

Neumann:1995:RHM

[NN95] László Neumann and Attila Neumann. Radiosity and hybrid methods. ACM Transactions on Graphics, 14(3):233–265, July 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
212347.html.

Nishino:2004:ER

[NN04] Ko Nishino and Shree K. Nayar. Eyes for relighting. ACM Transactions on Graphics, 23(3):704–711, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nakanishi:2020:RLA

[NNC^{+}20] Rafael Nakanishi, Filipe Nascimento, Rafael Campos, Paulo Pagliosa, and Afonso Paiva. RBF liquids: an adaptive PIC solver using RBF-FD. ACM Transactions on Graphics, 39(6): 170:1–170:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417794.

Novak:2012:VRL

[NNDJ12] Jan Novák, Derek Nowrouzezahrai, Carsten Dachsbacher, and Wojciech Jarosz. Virtual ray lights for rendering scenes with participating media. ACM Transactions on Graphics, 31(4):60:1–60:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nielsen:2007:CCL

[NNSM07] Michael B. Nielsen, Ola Nilsson, Andreas Söderström, and Ken Museth. Out-of-core and compressed level set methods. ACM Transactions on Graphics, 26(4):16:1–16:26, October 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nielsen:2013:TCA

[NØ13] Michael B. Nielsen and Ole Østerby. A two-continua approach to Eulerian simulation of water spray. ACM Transactions on Graphics, 32(4): 67:1–67:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nishita:1985:SMP

[NON85] T. Nishita, I. Okamura, and E. Nakamae. Shading models for point and linear sources. ACM Transactions on Graphics, 4(2):124–146, April 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nageli:2018:FRT

[NOP^{+}18] Tobias Nägeli, Samuel Oberholzer, Silvan Plüss, Javier Alonso-Mora, and Otmar Hilliges. Flycon: real-time environment-independent multiview human pose estimation with aerial vehicles. ACM Transactions on Graphics, 37 (6):182:1–182:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Niese:2022:PUF

[NPA^{+}22] Till Niese, Sören Pirk, Matthias Albrecht, Bedrich Benes, and Oliver Deussen. Procedural urban forestry. ACM Transactions on Graphics, 41 (2):20:1–20:18, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3502220.

Nuvoli:2022:SBA

[NPC^{+}22] Stefano Nuvoli, Nico Pietroni, Paolo Cignoni, Riccardo Scateni, and Marco Tarini. SkinMixer: Blending 3D animated models. ACM Transactions on Graphics, 41(6): 250:1–250:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555503.

Nguyen-Phuoc:2022:SSN

[NPLX22] Thu Nguyen-Phuoc, Feng Liu, and Lei Xiao. SNeRF: stylized neural implicit representations for 3D scenes. ACM Transactions on Graphics, 41 (4):142:1–142:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530107.

Narain:2013:FCA

[NPO13] Rahul Narain, Tobias Pfaff, and James F. O’Brien. Folding and crumpling adaptive sheets. ACM Transactions on Graphics, 32(4): 51:1–51:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nah:2011:TET

[NPP^{+}11] Jae-Ho Nah, Jeong-Soo Park, Chanmin Park, Jin-Woo Kim, Yun-Hye Jung, Woo-Chan Park, and Tack-Don Han.

T&I engine: traversal and intersection engine for hardware accelerated ray tracing. ACM Transactions on Graphics, 30 (6):160:1–160:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nazzaro:2022:GID

[NPP22] Giacomo Nazzaro, Enrico Puppo, and Fabio Pellacini. geoTangle: Interactive design of geodesic tangle patterns on surfaces. ACM Transactions on Graphics, 41 (2):12:1–12:17, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3487909.

Nader:2021:KFS

[NQC^{+}21] Georges Nader, Yu Han Quek, Pei Zhi Chia, Oliver Weeger, and Sai-Kit Yeung. KnitKit: a flexible system for machine knitting of customizable textiles. ACM Transactions on Graphics, 40 (4):64:1–64:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459790.

Nabizadeh:2021:KTS

[NRC21] Mohammad Sina Nabizadeh, Ravi Ramamoorthi, and Albert Chern. Kelvin transformations for simulations on infinite domains. ACM Transactions on Graphics, 40 (4):97:1–97:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459809.

Nehab:2005:ECP

[NRDR05] Diego Nehab, Szymon Rusinkiewicz, James Davis, and Ravi Ramamoorthi. Efficiently combining positions and normals for precise 3D geometry. ACM Transactions on Graphics, 24 (3):536–543, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ng:2003:AFS

[NRH03] Ren Ng, Ravi Ramamoorthi, and Pat Hanrahan. Allfrequency shadows using nonlinear wavelet lighting approximation. ACM Transactions on Graphics, 22(3): 376–381, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ng:2004:TPW

[NRH04] Ren Ng, Ravi Ramamoorthi, and Pat Hanrahan. Triple product wavelet integrals for all-frequency relighting. ACM Transactions on Graphics, 23 (3):477–487, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Naderi:2017:DSH

[NRH17] Kourosh Naderi, Joose Rajamäki, and Perttu Hämäläinen. Discovering and synthesizing humanoid climbing movements. ACM Transactions on Graphics, 36(4): 43:1–43:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nicolet:2023:RCV

[NRN^{+}23] Baptiste Nicolet, Fabrice Rousselle, Jan Novak, Alexander Keller, Wenzel Jakob, and Thomas Müller. Recursive control variates for inverse rendering. ACM Transactions on Graphics, 42 (4):62:1–62:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592139.

Nguyen:2015:DDC

[NRS15] Chuong H. Nguyen, Tobias Ritschel, and Hans-Peter Seidel. Data-driven color manifolds. ACM Transactions on Graphics, 34(2):20:1– 20:??, February 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nealen:2005:SBI

[NSACO05] Andrew Nealen, Olga Sorkine, Marc Alexa, and Daniel Cohen-Or. A sketch-based interface for detail-preserving mesh editing. ACM Transactions on Graphics, 24(3): 1142–1147, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nielsen:2013:SWA

[NSB13] Michael B. Nielsen, Andreas Söderström, and Robert Bridson. Synthesizing waves from animated height fields. ACM Transactions on Graphics, 32(1):2:1–2:9, January 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narain:2008:FAT

[NSCL08] Rahul Narain, Jason Sewall, Mark Carlson, and Ming C. Lin. Fast animation of turbulence using energy transport and procedural synthesis. ACM Transactions on Graphics, 27(5):166:1– 166:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nowrouzezahrai:2012:SZH

[NSF12] Derek Nowrouzezahrai, Patricio Simari, and Eugene Fiume. Sparse zonal harmonic factorization for efficient SH rotation. ACM Transactions on Graphics, 31 (3):23:1–23:9, May 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Novak:2014:RRT

[NSJ14] Jan Novák, Andrew Selle, and Wojciech Jarosz. Residual ratio tracking for estimating attenuation in participating media. ACM Transactions on Graphics, 33(6): 179:1–179:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Narain:2012:AAR

[NSO12] Rahul Narain, Armin Samii, and James F. O’Brien. Adaptive anisotropic remeshing for cloth simulation. ACM Transactions on Graphics, 31(6): 152:1–152:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nagasawa:2019:MSV

[NSS^{+}19] Kentaro Nagasawa, Takayuki Suzuki, Ryohei Seto, Masato Okada, and Yonghao Yue. Mixing sauces: a viscosity blending model for shear thinning fluids. ACM Transactions on Graphics, 38(4): 95:1–95:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nan:2011:CGR

[NSX^{+}11] Liangliang Nan, Andrei Sharf, Ke Xie, Tien-Tsin Wong, Oliver Deussen, Daniel Cohen-Or, and Baoquan Chen. Conjoining gestalt rules for abstraction of architectural drawings. ACM Transactions on Graphics, 30(6): 185:1–185:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nagano:2018:PRT

[NSX^{+}18] Koki Nagano, Jaewoo Seo, Jun Xing, Lingyu Wei, Zimo Li, Shunsuke Saito, Aviral Agarwal, Jens Fursund, and Hao Li. paGAN: real-time avatars using dynamic textures. ACM Transactions on Graphics, 37(6):258:1– 258:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nan:2010:SIU

[NSZ^{+}10] Liangliang Nan, Andrei Sharf, Hao Zhang, Daniel Cohen-Or, and Baoquan Chen. Smart-Boxes for interactive urban reconstruction. ACM Transactions on Graphics, 29(4): 93:1–93:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Neumann:2013:SLD

[NVW^{+}13] Thomas Neumann, Kiran Varanasi, Stephan Wenger, Markus Wacker, Marcus Magnor, and Christian Theobalt. Sparse localized deformation components. ACM Transactions on Graphics, 32(6): 179:1–179:??, November 2013.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nabizadeh:2022:CF

[NWRC22] Mohammad Sina Nabizadeh, Stephanie Wang, Ravi Ramamoorthi, and Albert Chern. Covector fluids. ACM Transactions on Graphics, 41 (4):113:1–113:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530120.

Narayanan:2019:VKM

[NWYM19] Vidya Narayanan, Kui Wu, Cem Yuksel, and James Mc-Cann. Visual knitting machine programming. ACM Transactions on Graphics, 38 (4):63:1–63:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nan:2012:SCA

[NXS12] Liangliang Nan, Ke Xie, and Andrei Sharf. A searchclassify approach for cluttered indoor scene understanding. ACM Transactions on Graphics, 31(6):137:1– 137:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Niizeki:1994:PII

[NY94] Masatoshi Niizeki and Fujio Yamaguchi. Projectively invariant intersection detections for solid modeling. ACM Transactions on Graphics, 13(3):277–299, July 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
197485.html.

Nagahara:2004:SWV

[NYY04] Hajime Nagahara, Yasushi Yagi, and Masahiko Yachida. Super wide viewer using catadioptrical optics. ACM Transactions on Graphics, 23(3): 732, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Nakada:2018:DLB

[NZC^{+}18] Masaki Nakada, Tao Zhou, Honglin Chen, Tomer Weiss, and Demetri Terzopoulos. Deep learning of biomimetic sensorimotor control for biomechanical human animation. ACM Transactions on Graphics, 37(4):56:1–56:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Niessner:2013:RTR

[NZIS13] Matthias Nießner, Michael Zollhöfer, Shahram Izadi, and Marc Stamminger. Real-time 3D reconstruction at scale using voxel hashing. ACM Transactions on Graphics, 32 (6):169:1–169:??, November 2013. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Naik:2011:SVR

[NZV^{+}11] Nikhil Naik, Shuang Zhao, Andreas Velten, Ramesh Raskar, and Kavita Bala. Single view reflectance capture using multiplexed scattering and time-of-flight imaging. ACM Transactions on Graphics, 30(6):171:1– 171:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ni:2020:LSM

[NZWC20] Xingyu Ni, Bo Zhu, Bin Wang, and Baoquan Chen. A level-set method for magnetic substance simulation. ACM Transactions on Graphics, 39 (4):29:1–29:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392445.

Oztireli:2010:SSM

[ÖAG10] A. Cengiz Öztireli, Marc Alexa, and Markus Gross. Spectral sampling of manifolds. ACM Transactions on Graphics, 29(6):168:1– 168:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

ODonovan:2011:CCL

[OAH11] Peter O’Donovan, Aseem Agarwala, and Aaron Hertzmann. Color compatibility from large datasets. ACM Transactions on Graphics, 30 (4):63:1–63:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ohtake:2003:MLP

[OBA^{+}03] Yutaka Ohtake, Alexander Belyaev, Marc Alexa, Greg Turk, and Hans-Peter Seidel. Multi-level partition of unity implicits. ACM Transactions on Graphics, 22(3): 463–470, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ovsjanikov:2012:FMF

[OBCS^{+}12] Maks Ovsjanikov, Mirela Ben-Chen, Justin Solomon, Adrian Butscher, and Leonidas Guibas. Functional maps: a flexible representation of maps between shapes. ACM Transactions on Graphics, 31(4): 30:1–30:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

OBrien:2002:GMA

[OBH02] James F. O’Brien, Adam W.

Bargteil, and Jessica K. Hodgins. Graphical modeling and animation of ductile fracture. ACM Transactions on Graphics, 21(3):291–294, July 2002.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ohtake:2004:RVL

[OBS04] Yutaka Ohtake, Alexander Belyaev, and Hans-Peter Seidel. Ridge-valley lines on meshes via implicit surface fitting. ACM Transactions on Graphics, 23(3): 609–612, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Orzan:2008:DCV

[OBW^{+}08] Alexandrina Orzan, Adrien Bousseau, Holger Winnemöller, Pascal Barla, Joëlle Thollot, and David Salesin. Diffusion curves: a vector representation for smooth-shaded images. ACM Transactions on Graphics, 27(3):92:1– 92:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Owens:2016:MDI

[OCH^{+}16] Andrew Owens, Mikolaj Cieslak, Jeremy Hart, Regine Classen-Bockhoff, and Przemyslaw Prusinkiewicz. Modeling dense inflorescences. ACM Transactions on Graphics, 35 (4):136:1–136:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ouyang:2021:SSF

[OCNG21] Peichang Ouyang, Kwok Wai Chung, Alain Nicolas, and Krzysztof Gdawiec. Selfsimilar fractal drawings inspired by M. C. Escher’s print Square Limit. ACM Transactions on Graphics, 40(3):31:1–31:34, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3456298.

OSullivan:2001:CP

[OD01] Carol O’Sullivan and John Dingliana. Collisions and perception. ACM Transactions on Graphics, 20(3): 151–168, July 2001. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Okabe:2015:FVM

[ODAO15] Makoto Okabe, Yoshinori Dobashi, Ken Anjyo, and Rikio Onai. Fluid volume modeling from sparse multiview images by appearance transfer. ACM Transactions on Graphics, 34(4):93:1– 93:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

OSullivan:2003:EVF

[ODGK03] Carol O’Sullivan, John Dingliana, Thanh Giang, and Mary K. Kaiser. Evaluating the visual fidelity of physically based animations. ACM Transactions on Graphics, 22(3): 527–536, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ostromoukhov:2004:FHI

[ODJ04] Victor Ostromoukhov, Charles Donohue, and Pierre-Marc Jodoin. Fast hierarchical importance sampling with blue noise properties. ACM Transactions on Graphics, 23(3): 488–495, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Overbeck:2009:AWR

[ODR09] Ryan S. Overbeck, Craig Donner, and Ravi Ramamoorthi. Adaptive wavelet rendering. ACM Transactions on Graphics, 28(5):140:1– 140:12, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Overbeck:2018:SAP

[OEE^{+}18] Ryan S. Overbeck, Daniel Erickson, Daniel Evangelakos, Matt Pharr, and Paul Debevec. A system for acquiring, processing, and rendering panoramic light field stills for virtual reality. ACM Transactions on Graphics, 37(6): 197:1–197:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ouellette:2001:NSO

[OF01] Marc J. Ouellette and Eugene Fiume. On numerical solutions to one-dimensional integration problems with applications to linear light sources. ACM Transactions on Graphics, 20(4):232–279, October 2001. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

OBrien:2012:EPM

[OF12] James F. O’Brien and Hany Farid. Exposing photo manipulation with inconsistent reflections. ACM Transactions on Graphics, 31(1):4:1– 4:11, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Osada:2002:SD

[OFCD02] Robert Osada, Thomas Funkhouser, Bernard Chazelle, and David Dobkin. Shape distributions. ACM Transactions on Graphics, 21(4):807–832, October 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Oztireli:2012:ASP

[ÖG12] A. Cengiz Öztireli and Markus Gross. Analysis and synthesis of point distributions based on pair correlation. ACM Transactions on Graphics, 31(6): 170:1–170:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Oztireli:2015:PBD

[ÖG15] A. Cengiz Öztireli and Markus Gross. Perceptually based downscaling of images. ACM Transactions on Graphics, 34 (4):77:1–77:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ouyang:2023:ISD

[OGN^{+}23] Peichang Ouyang, Krzysztof Gdawiec, Alain Nicolas, David Bailey, and Kwok Wai Chung. Interlocking spiral drawings inspired by M. C. Escher’s print whirlpools. ACM Transactions on Graphics, 42(2):18:1–18:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3560711.

Oskam:2011:OOS

[OHB^{+}11] Thomas Oskam, Alexander Hornung, Huw Bowles, Kenny Mitchell, and Markus Gross. OSCAM — optimized stereoscopic camera control for interactive 3D. ACM Transactions on Graphics, 30(6): 189:1–189:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Otsu:2018:GAM

[OHHD18] Hisanari Otsu, Johannes Hanika, Toshiya Hachisuka, and Carsten Dachsbacher. Geometry-aware Metropolis light transport. ACM Transactions on Graphics, 37(6): 278:1–278:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ochiai:2014:PDG

[OHR14] Yoichi Ochiai, Takayuki Hoshi, and Jun Rekimoto. Pixie dust: graphics generated by levitated and animated objects in computational acoustic-potential field. ACM Transactions on Graphics, 33(4):85:1–85:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

OToole:2014:TFP

[OHX^{+}14] Matthew O’Toole, Felix Heide, Lei Xiao, Matthias B. Hullin, Wolfgang Heidrich, and Kiriakos N. Kutulakos. Temporal frequency probing for 5D transient analysis of global light transport. ACM Transactions on Graphics, 33(4): 87:1–87:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

OToole:2010:OCF

[OK10] Matthew O’Toole and Kiriakos N. Kutulakos. Optical computing for fast light transport analysis. ACM Transactions on Graphics, 29(6): 164:1–164:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ochiai:2016:FLF

[OKH^{+}16] Yoichi Ochiai, Kota Kumagai, Takayuki Hoshi, Jun Rekimoto, Satoshi Hasegawa, and Yoshio Hayasaki. Fairy lights in femtoseconds: Aerial and volumetric graphics rendered by focused femtosecond laser combined with computational holographic fields. ACM Transactions on Graphics, 35(2):17:1–17:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Otsu:2017:FSS

[OKH^{+}17] Hisanari Otsu, Anton S. Kaplanyan, Johannes Hanika, Carsten Dachsbacher, and Toshiya Hachisuka. Fusing state spaces for Markov chain Monte Carlo rendering. ACM Transactions on Graphics, 36 (4):74:1–74:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ozgen:2010:UCS

[OKRC10] Oktar Ozgen, Marcelo Kallmann, Lynnette Es Ramirez, and Carlos Fm Coimbra. Underwater cloth simulation with fractional derivatives.

ACM Transactions on Graphics, 29(3):23:1–23:9, June 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Otaduy:2003:SPS

[OL03] Miguel A. Otaduy and Ming C.

Lin. Sensation preserving simplification for haptic rendering. ACM Transactions on Graphics, 22(3):543–553, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

ODonovan:2014:EFS

[OLAH14] Peter O’Donovan, Janis Libeks, Aseem Agarwala, and Aaron Hertzmann. Exploratory font selection using crowdsourced attributes. ACM Transactions on Graphics, 33(4): 92:1–92:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ovsjanikov:2011:ECV

[OLGM11] Maks Ovsjanikov, Wilmot Li, Leonidas Guibas, and Niloy J. Mitra. Exploration of continuous variability in collections of 3D shapes. ACM Transactions on Graphics, 30 (4):33:1–33:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Olsen:1984:PAU

[Ols84] Dan R. Olsen, Jr. Pushdown automata for user interface management. ACM Transactions on Graphics, 3 (3):177–203, July 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Olsen:1986:MMI

[Ols86] Dan R. Olsen, Jr. MIKE: The menu interaction kontrol environment. ACM Transactions on Graphics, 5 (4):318–344, October 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
28868.html.

Olsen:1988:CST

[Ols88] Dan Olsen. Call for submissions to the TOG interactive techniques notebook. ACM Transactions on Graphics, 7 (4):227–228, October 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Olsen:1992:BES

[Ols92] Dan R. Olsen. Bookmarks: An enhanced scroll bar. ACM Transactions on Graphics, 11 (3):291–295, July 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Olszewski:2016:HFF

[OLSL16] Kyle Olszewski, Joseph J. Lim, Shunsuke Saito, and Hao Li. High-fidelity facial and speech animation for VR HMDs. ACM Transactions on Graphics, 35(6): 221:1–221:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Owada:2004:VID

[ONOI04] Shigeru Owada, Frank Nielsen, Makoto Okabe, and Takeo Igarashi. Volumetric illustration: designing 3D models with internal textures. ACM Transactions on Graphics, 23 (3):322–328, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ou:2011:LMS

[OP11] Jiawei Ou and Fabio Pellacini. LightSlice: matrix slice sampling for the manylights problem. ACM Transactions on Graphics, 30(6): 179:1–179:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ondrej:2010:SVB

[OPOD10] Jan Ondřej, Julien Pettré, Anne-Hélène Olivier, and Stéphane Donikian. A synthetic-vision based steering approach for crowd simulation. ACM Transactions on Graphics, 29(4):123:1– 123:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

OToole:2012:PDC

[ORK12] Matthew O’Toole, Ramesh Raskar, and Kiriakos N. Kutulakos. Primal-dual coding to probe light transport. ACM Transactions on Graphics, 31(4):39:1–39:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ostromoukhov:2007:SP

[Ost07] Victor Ostromoukhov. Sampling with polyominoes. ACM Transactions on Graphics, 26 (3):78:1–78:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Oliva:2006:HI

[OTS06] Aude Oliva, Antonio Torralba, and Philippe G. Schyns. Hybrid images. ACM Transactions on Graphics, 25(3): 527–532, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Oztireli:2016:ISP

[Özt16] A. Cengiz Öztireli. Integration with stochastic point processes. ACM Transactions on Graphics, 35(5):160:1– 160:??, September 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pintore:2021:DLR

[PAAG21] Giovanni Pintore, Eva Almansa, Marco Agus, and Enrico Gobbetti. Deep3DLayout: 3D reconstruction of an indoor layout from a spherical panoramic image. ACM Transactions on Graphics, 40 (6):250:1–250:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480480.

Perel:2021:LMA

[PABE^{+}21] Or Perel, Oron Anschel, Omri Ben-Eliezer, Shai Mazor, and Hadar Averbuch-Elor. Learning multimodal affinities for textual editing in images. ACM Transactions on Graphics, 40(3):26:1–26:16, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3451340.

Paglieroni:1998:DPP

[Pag98] David W. Paglieroni. The directional parameter plane transform of a height field. ACM Transactions on Graphics, 17(1):50–70, January 1998. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL http:

//www.acm.org:80/pubs/citations/
journals/tog/1998-17-1/
p50-paglieroni/.

Peiret:2019:SCB

[PAK^{+}19] Albert Peiret, Sheldon Andrews, József Kövecses, Paul G. Kry, and Marek Teichmann. Schur complement-based substructuring of stiff multibody systems with contact. ACM Transactions on Graphics, 38(5):150:1–150:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2018:DEG

[PALvdP18] Xue Bin Peng, Pieter Abbeel, Sergey Levine, and Michiel van de Panne. DeepMimic: example-guided deep reinforcement learning of physicsbased character skills. ACM Transactions on Graphics, 37(4):143:1–143:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pantaleoni:2017:CML

[Pan17] Jacopo Pantaleoni. Charted Metropolis light transport. ACM Transactions on Graphics, 36(4):75:1–75:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panozzo:2018:SDM

[Pan18] Daniele Panozzo. Session details: Meshing. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paris:2017:CMO

[Par17] Sylvain Paris. CoLux: multiobject 3D micro-motion analysis using speckle imaging. ACM Transactions on Graphics, 36(4):34:1–34:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Petikam:2021:SRD

[PAR21] Lohit Petikam, Ken Anjyo, and Taehyun Rhee. Shading rig: Dynamic artdirectable stylised shading for 3D characters. ACM Transactions on Graphics, 40(5): 189:1–189:14, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3461696.

Patterson:1985:PTP

[Pat85] Richard R. Patterson. Projective transformations of the parameter of a Bernstein-Bézier curve. ACM Transactions on Graphics, 4(4):276–290, October 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

6119.html. See corrigendum [Pat87].

Patterson:1987:CPT

[Pat87] Richard R. Patterson. Corrigendum: “Projective Transformations of the Parameter of a Bernstein-Bézier Curve”.

ACM Transactions on Graphics, 6(1):79, January 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [Pat85].

Pavlidis:1983:CFC

[Pav83] Theodosios Pavlidis. Curve fitting with conic splines. ACM Transactions on Graphics, 2(1):1–31, January 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pavlidis:1990:RCS

[Pav90] Theo Pavlidis. Re: Comments on “Stochastic Sampling in Computer Graphics”. ACM Transactions on Graphics, 9 (2):233–236, April 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [Coo86, WP90].

Pullen:2002:MCA

[PB02] Katherine Pullen and Christoph Bregler. Motion capture assisted animation: texturing and synthesis. ACM Transactions on Graphics, 21(3): 501–508, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paulin:2022:MMS

[PBC^{+}22] Loïs Paulin, Nicolas Bonneel, David Coeurjolly, Jean-Claude Iehl, Alexander Keller, and Victor Ostromoukhov. MatBuilder: mastering sampling uniformity over projections. ACM Transactions on Graphics, 41(4): 84:1–84:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530063.

Paoluzzi:1993:DIM

[PBCF93] A. Paoluzzi, F. Bernardini, C. Cattani, and V. Ferrucci. Dimension-independent modeling with simplicial complexes. ACM Transactions on Graphics, 12(1):56– 102, January 1993. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
169719.html.

Parker:2010:OGP

[PBD^{+}10] Steven G. Parker, James Bigler, Andreas Dietrich, Heiko Friedrich, Jared Hoberock, David Luebke, David McAllister, Morgan McGuire, Keith Morley, Austin Robison, and Martin Stich. OptiX: a general purpose ray tracing engine. ACM Transactions on Graphics, 29(4): 66:1–66:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panozzo:2013:WAS

[PBDSH13] Daniele Panozzo, Ilya Baran, Olga Diamanti, and Olga Sorkine-Hornung. Weighted averages on surfaces. ACM Transactions on Graphics, 32 (4):60:1–60:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Porumbescu:2005:SM

[PBFJ05] Serban D. Porumbescu, Brian Budge, Louis Feng, and Kenneth I. Joy. Shell maps. ACM Transactions on Graphics, 24 (3):626–633, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2015:SDS

[PBH15] Zherong Pan, Hujun Bao, and Jin Huang. Subspace dynamic simulation using rotationstrain coordinates. ACM Transactions on Graphics, 34 (6):242:1–242:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2014:EQ

[PBJW14] Chi-Han Peng, Michael Barton, Caigui Jiang, and Peter Wonka. Exploring quadrangulations. ACM Transactions on Graphics, 33(1):12:1– 12:13, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pidhorskyi:2022:DFA

[PBM^{+}22] Stanislav Pidhorskyi, Timur Bagautdinov, Shugao Ma, Jason Saragih, Gabriel Schwartz, Yaser Sheikh, and Tomas Simon. Depth of field aware differentiable rendering. ACM Transactions on Graphics, 41 (6):190:1–190:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555521.

Pellacini:2007:LP

[PBMF07] Fabio Pellacini, Frank Battaglia, R. Keith Morley, and Adam Finkelstein. Lighting with paint. ACM Transactions on Graphics, 26(2):9:1–9:??, June 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Purcell:2002:RTP

[PBMH02] Timothy J. Purcell, Ian Buck, William R. Mark, and Pat Hanrahan. Ray tracing on programmable graphics hardware. ACM Transactions on Graphics, 21(3):703–712, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paris:2004:CHG

[PBS04] Sylvain Paris, Hector M.

Briceño, and François X. Sillion. Capture of hair geometry from multiple images. ACM Transactions on Graphics, 23(3):712–719, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Palmer:2020:ARV

[PBS20] David Palmer, David Bommes, and Justin Solomon. Algebraic representations for volumetric frame fields. ACM Transactions on Graphics, 39 (2):16:1–16:17, April 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/doi/abs/10.
                                        1145/3366786.

Panozzo:2013:DUM

[PBSH13] Daniele Panozzo, Philippe Block, and Olga Sorkine-Hornung. Designing unreinforced masonry models. ACM Transactions on Graphics, 32 (4):91:1–91:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2015:DTT

[PBvdP15] Xue Bin Peng, Glen Berseth, and Michiel van de Panne. Dynamic terrain traversal skills using reinforcement learning. ACM Transactions on Graphics, 34(4):80:1– 80:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2016:TAL

[PBvdP16] Xue Bin Peng, Glen Berseth, and Michiel van de Panne. Terrain-adaptive locomotion skills using deep reinforcement learning. ACM Transactions on Graphics, 35(4):

81:1–81:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Preiner:2019:GPS

[PBW19] Reinhold Preiner, Tamy Boubekeur, and Michael Wimmer. Gaussian-product subdivision surfaces. ACM Transactions on Graphics, 38 (4):35:1–35:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2017:DDL

[PBYV17] Xue Bin Peng, Glen Berseth, Kangkang Yin, and Michiel Van De Panne. DeepLoco: dynamic locomotion skills using hierarchical deep reinforcement learning. ACM Transactions on Graphics, 36 (4):41:1–41:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Potmesil:1982:SIG

[PC82] M. Potmesil and I. Chakravarty.

Synthetic image generation with a lens and aperture camera model. ACM Transactions on Graphics, 1(2): 85–108, April 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Piovarci:2023:SSC

[PCB23] Michal Piovarci, Alexandre Chapiro, and Bernd Bickel.

Skin-screen: a computational fabrication framework for color tattoos. ACM Transactions on Graphics, 42 (4):67:1–67:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592432.

Poms:2018:SEV

[PCHF18] Alex Poms, Will Crichton, Pat Hanrahan, and Kayvon Fatahalian. Scanner: efficient video analysis at scale. ACM Transactions on Graphics, 37(4):138:1–138:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paulin:2021:CSS

[PCI^{+}21] Loïs Paulin, David Coeurjolly, Jean-Claude Iehl, Nicolas Bonneel, Alexander Keller, and Victor Ostromoukhov. Cascaded Sobol’ sampling. ACM Transactions on Graphics, 40(6):275:1–275:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480482.

Paris:2008:HPG

[PCK^{+}08] Sylvain Paris, Will Chang, Oleg I. Kozhushnyan, Wojciech Jarosz, Wojciech Matusik, Matthias Zwicker, and Frédo Durand. Hair photobooth: geometric and photometric acquisition of real hairstyles. ACM Transactions on Graphics, 27(3):30:1– 30:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Padilla:2019:BRI

[PCK^{+}19] Marcel Padilla, Albert Chern, Felix Knöppel, Ulrich Pinkall, and Peter Schröder. On bubble rings and ink chandeliers. ACM Transactions on Graphics, 38(4):129:1–129:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2012:RMC

[PCL^{+}12] Hao Pan, Yi-King Choi, Yang Liu, Wenchao Hu, Qiang Du, Konrad Polthier, Caiming Zhang, and Wenping Wang. Robust modeling of constant mean curvature surfaces. ACM Transactions on Graphics, 31(4):85:1–85:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pang:2016:DUA

[PCLC16] Xufang Pang, Ying Cao, Rynson W. H. Lau, and Antoni B. Chan. Directing user attention via visual flow on web designs. ACM Transactions on Graphics, 35(6): 240:1–240:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2020:NHC

[PCPW20] Yifan Peng, Suyeon Choi, Nitish Padmanaban, and Gordon Wetzstein. Neural holography with camera-in-the-loop training. ACM Transactions on Graphics, 39(6): 185:1–185:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417802.

Pediredla:2020:PTE

[PCS^{+}20] Adithya Pediredla, Yasin Karimi Chalmiani, Matteo Giuseppe Scopelliti, Maysamreza Chamanzar, Srinivasa Narasimhan, and Ioannis Gkioulekas. Path tracing estimators for refractive radiative transfer. ACM Transactions on Graphics, 39 (6):241:1–241:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417793.

Pandey:2023:JIV

[PCS23a] Karran Pandey, Fanny Chevalier, and Karan Singh. Juxtaform: interactive visual summarization for exploratory shape design. ACM Transactions on Graphics, 42(4): 52:1–52:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592436.

Pietroni:2023:HMG

[PCS^{+}23b] Nico Pietroni, Marcel Campen, Alla Sheffer, Gianmarco Cherchi, David Bommes, Xifeng Gao, Riccardo Scateni, Franck Ledoux, Jean Remacle, and Marco Livesu. Hex-mesh generation and processing: a survey. ACM Transactions on Graphics, 42(2): 16:1–16:??, April 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3554920.

Park:2006:VGM

[PCSS06] Min Je Park, Min Gyu Choi, Yoshihisa Shinagawa, and Sung Yong Shin. Videoguided motion synthesis using example motions. ACM Transactions on Graphics, 25 (4):1327–1359, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pietroni:2022:CPM

[PDF^{+}22] Nico Pietroni, Corentin Dumery, Raphael Falque, Mark Liu, Teresa Vidal-Calleja, and Olga Sorkine-Hornung. Computational pattern making from 3D garment models. ACM Transactions on Graphics, 41(4):157:1–157:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530145.

Peng:2017:MMH

[PDSH17] Yifan Peng, Xiong Dun, Qilin Sun, and Wolfgang Heidrich. Mix-and-match holography. ACM Transactions on Graphics, 36(6):191:1– 191:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2018:AAG

[PDZ^{+}18] Yue Peng, Bailin Deng, Juyong Zhang, Fanyu Geng, Wenjie Qin, and Ligang Liu. Anderson acceleration for geometry optimization and physics simulation. ACM Transactions on Graphics, 37(4):42:1– 42:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellacini:2005:UCA

[Pel05]  Fabio Pellacini.  User-

configurable automatic shader simplification. ACM Transactions on Graphics, 24(3): 445–452, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellacini:2010:EIE

[Pel10] Fabio Pellacini. envyLight: an interface for editing natural illumination. ACM Transactions on Graphics, 29(4): 34:1–34:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pandey:2021:TRL

[PEL^{+}21] Rohit Pandey, Sergio Orts Escolano, Chloe Legendre, Christian HÃne, Sofien Bouaziz, Christoph Rhemann, Paul Debevec, and Sean Fanello. Total relighting: learning to relight portraits for background replacement. ACM Transactions on Graphics, 40 (4):43:1–43:21, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459872.

Perlin:2002:IN

[Per02] Ken Perlin. Improving noise.

ACM Transactions on Graphics, 21(3):681–682, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peters:1989:LGH

[Pet89]  Jörg Peters.  Local gen-

eralized Hermite interpolation by quartic C^{2} space curves. ACM Transactions on Graphics, 8(3):235–242, July 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77060.html.

Peters:1995:SPM

[Pet95] Jörg Peters. Smoothing polyhedra made easy. ACM Transactions on Graphics, 14(2):162–170, April 1995.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
221670.html.

Peters:2001:SPR

[Pet01] Jörg Peters. Smooth patching of refined triangulations. ACM Transactions on Graphics, 20(1):1–9, January 2001. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
2001-20-1/p1-peters/.

Peters:2021:BIS

[Pet21] Christoph Peters. BRDF importance sampling for polygonal lights. ACM Transactions on Graphics, 40(4): 140:1–140:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459672.

Pluta:2021:PCP

[PEVBC21] Kacper Pluta, Michal Edelstein, Amir Vaxman, and Mirela Ben-Chen. PH-CPF: planar hexagonal meshing using coordinate power fields. ACM Transactions on Graphics, 40(4):156:1– 156:19, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459770.

Posch:1989:CBA

[PF89] K. C. Posch and W. D. Fellner. The circle-brush algorithm. ACM Transactions on Graphics, 8(1):1–24, January 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
49156.html.

Piovarci:2020:TSV

[PFB^{+}20] Michal Piovarči, Michael Foshey, Vahid Babaei, Szymon Rusinkiewicz, Wojciech Matusik, and Piotr Didyk. Towards spatially varying gloss reproduction for 3D printing. ACM Transactions on Graphics, 39(6): 206:1–206:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417850.

Pantaleoni:2010:PFR

[PFHA10] Jacopo Pantaleoni, Luca Fascione, Martin Hill, and Timo Aila. PantaRay: fast raytraced occlusion caching of massive scenes. ACM Transactions on Graphics, 29(4): 37:1–37:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paneva:2022:OOT

[PFP^{+}22] Viktorija Paneva, Arthur Fleig, Diego MartíNez Plasencia, Timm Faulwasser, and Jörg Müller. OptiTrap: Optimal trap trajectories for acoustic levitation displays. ACM Transactions on Graphics, 41(5):173:1–173:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3517746.

Piovarci:2022:CLC

[PFX^{+}22] Michal Piovarči, Michael Foshey, Jie Xu, Timmothy Erps, Vahid Babaei, Piotr Didyk, Szymon Rusinkiewicz, Wojciech Matusik, and Bernd Bickel. Closed-loop control of direct ink writing via reinforcement learning. ACM Transactions on Graphics, 41 (4):112:1–112:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530144.

Perez:2003:PIE

[PGB03] Patrick Pérez, Michel Gangnet, and Andrew Blake. Poisson image editing. ACM Transactions on Graphics, 22(3): 313–318, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paris:2023:ASM

[PGCG23] Axel Paris, Eric Guérin, Pauline Collon, and Eric Galin. Authoring and simulating meandering rivers. ACM Transactions on Graphics, 42 (6):239:1–239:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618350.

Panuelos:2023:PPM

[PGG^{+}23] Jonathan Panuelos, Ryan Goldade, Eitan Grinspun, David Levin, and Christopher Batty. PolyStokes: a polynomial model reduction method for viscous fluid simulation. ACM Transactions on Graphics, 42(4): 124:1–124:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592146.

Peytavie:2024:DID

[PGG^{+}24] Adrien Peytavie, James Gain, Eric Guérin, Oscar Argudo, and Eric Galin. Dead-Wood: Including disturbance and decay in the depiction of digital nature. ACM Transactions on Graphics, 43 (2):21:1–21:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3641816.

Peng:2022:ALS

[PGH^{+}22] Xue Bin Peng, Yunrong Guo, Lina Halper, Sergey Levine, and Sanja Fidler. ASE: largescale reusable adversarial skill embeddings for physically simulated characters. ACM Transactions on Graphics, 41(4):94:1–94:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530110.

Padilla:2022:FBP

[PGK^{+}22] Marcel Padilla, Oliver Gross, Felix Knöppel, Albert Chern, Ulrich Pinkall, and Peter Schröder. Filament based plasma. ACM Transactions on Graphics, 41(4):153:1– 153:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530102.

Perard-Gayot:2019:RGR

[PGML^{+}19] Arsène Pérard-Gayot, Richard Membarth, Roland Leißa, Sebastian Hack, and Philipp Slusallek. Rodent: generating renderers without writing a generator. ACM Transactions on Graphics, 38(4): 40:1–40:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paris:2019:TAI

[PGP^{+}19] Axel Paris, Eric Galin, Adrien Peytavie, Eric Guérin, and James Gain. Terrain amplification with implicit 3D features. ACM Transactions on Graphics, 38(5): 147:1–147:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3342765.

Philip:2019:MVR

[PGZ^{+}19] Julien Philip, Michaël Gharbi, Tinghui Zhou, Alexei A. Efros, and George Drettakis. Multi-view relighting using a geometry-aware network. ACM Transactions on Graphics, 38(4):78:1–78:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Praun:2003:SPR

[PH03] Emil Praun and Hugues Hoppe. Spherical parametrization and remeshing. ACM Transactions on Graphics, 22 (3):340–349, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:2006:CAS

[PH06] Sang Il Park and Jessica K.

Hodgins. Capturing and animating skin deformation in human motion. ACM Transactions on Graphics, 25(3):

881–889, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:2008:DDM

[PH08] Sang Il Park and Jessica K. Hodgins. Data-driven modeling of skin and muscle deformation. ACM Transactions on Graphics, 27(3):96:1– 96:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pjanic:2015:CCE

[PH15a] Petar Pjanic and Roger D. Hersch. Color changing effects with anisotropic halftone prints on metal. ACM Transactions on Graphics, 34(6): 167:1–167:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pjanic:2015:CIP

[PH15b] Petar Pjanic and Roger D. Hersch. Color imaging and pattern hiding on a metallic substrate. ACM Transactions on Graphics, 34(4): 130:1–130:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pharr:2018:GEI

[Pha18] Matt Pharr. Guest Editor’s introduction: Special issue on production rendering. ACM Transactions on Graphics, 37 (3):28:1–28:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3212511.

Polasek:2021:IAP

[PHBC21] Tomas Polasek, David Hrusa, Bedrich Benes, and Martin Cadík. ICTree: automatic perceptual metrics for tree models. ACM Transactions on Graphics, 40(6): 230:1–230:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480519.

Pottmann:2010:GP

[PHD^{+}10] Helmut Pottmann, Qixing Huang, Bailin Deng, Alexander Schiftner, Martin Kilian, Leonidas Guibas, and Johannes Wallner. Geodesic patterns. ACM Transactions on Graphics, 29(4): 43:1–43:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paris:2011:LLF

[PHK11] Sylvain Paris, Samuel W.

Hasinoff, and Jan Kautz. Local Laplacian filters: edgeaware image processing with a Laplacian pyramid. ACM Transactions on Graphics, 30 (4):68:1–68:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Palubicki:2009:SOT

[PHL^{+}09] Wojciech Palubicki, Kipp Horel, Steven Longay, Adam Runions, Brendan Lane, Radomír Měch, and Przemyslaw Prusinkiewicz. Selforganizing tree models for image synthesis. ACM Transactions on Graphics, 28(3): 58:1–58:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:2023:CCP

[PHM^{+}23] Keunhong Park, Philipp Henzler, Ben Mildenhall, Jonathan T. Barron, and Ricardo Martin-Brualla. CamP: Camera preconditioning for neural radiance fields. ACM Transactions on Graphics, 42 (6):208:1–208:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3618321.

Papas:2012:MLR

[PHN^{+}12] Marios Papas, Thomas Houit, Derek Nowrouzezahrai, Markus Gross, and Wojciech Jarosz. The magic lens: refractive steganography. ACM Transactions on Graphics, 31(6): 186:1–186:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Portenier:2018:FDS

[PHS^{+}18] Tiziano Portenier, Qiyang Hu, Attila Szabó, Siavash Arjomand Bigdeli, Paolo Favaro, and Matthias Zwicker. Faceshop: deep sketch-based face image editing. ACM Transactions on Graphics, 37(4): 99:1–99:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2013:ILL

[PHT^{+}13] Zherong Pan, Jin Huang, Yiying Tong, Changxi Zheng, and Hujun Bao. Interactive localized liquid motion editing. ACM Transactions on Graphics, 32(6): 184:1–184:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panetta:2021:CID

[PIC^{+}21] Julian Panetta, Florin Isvoranu, Tian Chen, Emmanuel Siéfert, Benoît Roman, and Mark Pauly. Computational inverse design of surfacebased inflatables. ACM Transactions on Graphics, 40 (4):40:1–40:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459789.

Peer:2015:IVF

[PICT15] Andreas Peer, Markus Ihmsen, Jens Cornelis, and Matthias Teschner. An implicit viscosity formulation for SPH fluids. ACM Transactions on Graphics, 34(4): 114:1–114:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pike:1983:GOB

[Pik83] R. Pike. Graphics in overlapping bitmap layers. ACM Transactions on Graphics, 2 (2):135–160, April 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pirk:2017:IWC

[PJH^{+}17] Sören Pirk, Michal Jarzabek, Torsten Hädrich, Dominik L. Michels, and Wojciech Palubicki. Interactive wood combustion for botanical tree models. ACM Transactions on Graphics, 36(6): 197:1–197:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Prosser:1983:IMG

[PK83] Colin J. Prosser and Alistair C. Kilgour. An integer method for the graphical output of conic sections. ACM Transactions on Graphics, 2 (3):182–191, July 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2005:GGP

[PK05] Jingliang Peng and C.-C. Jay Kuo. Geometry-guided progressive lossless 3D mesh coding with octree (OT) decomposition. ACM Transactions on Graphics, 24(3): 609–616, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Philbrick:2022:PMH

[PK22] Greg Philbrick and Craig S.

Kaplan. A primitive for manual hatching. ACM Transactions on Graphics, 41 (2):21:1–21:17, April 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3503460.

Pauly:2005:MAF

[PKA^{+}05] Mark Pauly, Richard Keiser, Bart Adams, Philip Dutré, Markus Gross, and Leonidas J. Guibas. Meshless animation of fracturing solids. ACM Transactions on Graphics, 24 (3):957–964, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Prada:2016:MGU

[PKC^{+}16] Fabián Prada, Misha Kazhdan, Ming Chuang, Alvaro Collet, and Hugues Hoppe. Motion graphs for unstructured textured meshes. ACM Transactions on Graphics, 35 (4):108:1–108:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Prada:2017:SAP

[PKC^{+}17] Fabián Prada, Misha Kazhdan, Ming Chuang, Alvaro Collet, and Hugues Hoppe. Spatiotemporal atlas parameterization for evolving meshes. ACM Transactions on Graphics, 36(4):58:1–58:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Prada:2018:GDP

[PKCH18] Fabián Prada, Misha Kazhdan, Ming Chuang, and Hugues Hoppe. Gradientdomain processing within a texture atlas. ACM Transactions on Graphics, 37(4): 154:1–154:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellis:2019:VSP

[PKD^{+}19] Davide Pellis, Martin Kilian, Felix Dellinger, Johannes Wallner, and Helmut Pottmann. Visual smoothness of polyhedral surfaces. ACM Transactions on Graphics, 38 (4):31:1–31:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pauly:2006:PBM

[PKG06] Mark Pauly, Leif P. Kobbelt, and Markus Gross. Pointbased multiscale surface representation. ACM Transactions on Graphics, 25(2): 177–193, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pirk:2017:UEOa

[PKH^{+}17a] Sören Pirk, Vojtech Krs, Kaimo Hu, Suren Deepak Rajasekaran, Hao Kang, Yusuke Yoshiyasu, Bedrich Benes, and Leonidas J. Guibas. Understanding and exploiting object interaction landscapes. ACM Transactions on Graphics, 36(3):31:1–31:??, June 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pirk:2017:UEOb

[PKH^{+}17b] Sören Pirk, Vojtech Krs, Kaimo Hu, Suren Deepak Rajasekaran, Hao Kang, Yusuke Yoshiyasu, Bedrich Benes, and Leonidas J. Guibas. Understanding and exploiting object interaction landscapes. ACM Transactions on Graphics, 36(4):52:1–52:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peters:2015:STM

[PKHK15] Christoph Peters, Jonathan Klein, Matthias B. Hullin, and Reinhard Klein. Solving trigonometric moment problems for fast transient imaging. ACM Transactions on Graphics, 34(6):220:1– 220:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pauly:2003:SMP

[PKKG03] Mark Pauly, Richard Keiser, Leif P. Kobbelt, and Markus Gross. Shape modeling with point-sampled geometry. ACM Transactions on Graphics, 22(3):641–650, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panetta:2019:XSN

[PKLI^{+}19] J. Panetta, M. Konaković-Luković, F. Isvoranu, E. Bouleau, and M. Pauly. X-Shells: a new class of deployable beam structures. ACM Transactions on Graphics, 38(4): 83:1–83:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paczkowski:2011:ISA

[PKM^{+}11] Patrick Paczkowski, Min H. Kim, Yann Morvan, Julie Dorsey, Holly Rushmeier, and Carol O’Sullivan. Insitu: sketching architectural designs in context. ACM Transactions on Graphics, 30(6): 182:1–182:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2018:SRL

[PKM^{+}18] Xue Bin Peng, Angjoo Kanazawa, Jitendra Malik, Pieter Abbeel, and Sergey Levine. SFV: reinforcement learning of physical skills from videos. ACM Transactions on Graphics, 37(6): 178:1–178:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellis:2021:CDW

[PKPP21] Davide Pellis, Martin Kilian, Helmut Pottmann, and Mark Pauly. Computational design of Weingarten surfaces. ACM Transactions on Graphics, 40(4): 114:1–114:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459939.

Peng:2004:IMT

[PKZ04] Jianbo Peng, Daniel Kristjansson, and Denis Zorin. Interactive modeling of topologically complex geometric detail. ACM Transactions on Graphics, 23(3):635– 643, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellacini:2007:AEM

[PL07] Fabio Pellacini and Jason Lawrence. AppWand: editing measured materials using appearance-driven optimization. ACM Transactions on Graphics, 26(3): 54:1–54:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Poranne:2014:PGP

[PL14] Roi Poranne and Yaron Lipman. Provably good planar mappings. ACM Transactions on Graphics, 33(4): 76:1–76:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pitzalis:2021:GAR

[PLC^{+}21] Luca Pitzalis, Marco Livesu, Gianmarco Cherchi, Enrico Gobbetti, and Riccardo Scateni. Generalized adaptive refinement for grid-based hexahedral meshing. ACM Transactions on Graphics, 40(6): 257:1–257:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480508.

Piovarci:2018:PAM

[PLKD18] Michal Piovarci, David I. W. Levin, Danny M. Kaufman, and Piotr Didyk. Perceptionaware modeling and fabrication of digital drawing tools. ACM Transactions on Graphics, 37(4):123:1– 123:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pereira:2017:PAA

[PLMR17] Thiago Pereira, Carolina L. A. Paes Leme, Steve Marschner, and Szymon Rusinkiewicz. Printing anisotropic appearance with magnetic flakes. ACM Transactions on Graphics, 36(4): 123:1–123:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panozzo:2012:FSS

[PLPZ12] Daniele Panozzo, Yaron Lipman, Enrico Puppo, and Denis Zorin. Fields on symmetric surfaces. ACM Transactions on Graphics, 31(4): 111:1–111:12, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Piovarci:2016:IAP

[PLR^{+}16] Michal Piovarci, David I. W.

Levin, Jason Rebello, Desai Chen, Roman Durikovic, Hanspeter Pfister, Wojciech Matusik, and Piotr Didyk. An interaction-aware, perceptual model for non-linear elastic objects. ACM Transactions on Graphics, 35(4): 55:1–55:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2015:FAS

[PLS^{+}15] Hao Pan, Yang Liu, Alla Sheffer, Nicholas Vining, Chang-Jian Li, and Wenping Wang. Flow aligned surfacing of curve networks. ACM Transactions on Graphics, 34(4): 127:1–127:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pottmann:2007:GML

[PLW^{+}07] Helmut Pottmann, Yang Liu, Johannes Wallner, Alexander Bobenko, and Wenping Wang. Geometry of multilayer freeform structures for architecture. ACM Transactions on Graphics, 26(3): 65:1–65:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pattanaik:1995:AER

[PM95] S. N. Pattanaik and S. P. Mudur. Adjoint equations and random walks for illumination computation. ACM Transactions on Graphics, 14(1):77–102, January 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        200985.html.

Peyre:2005:SCG

[PM05] Gabriel Peyré and Stéphane Mallat. Surface compression with geometric bandelets. ACM Transactions on Graphics, 24(3):601–608, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2017:ESSa

[PM17a] Zherong Pan and Dinesh Manocha. Efficient solver for spacetime control of smoke. ACM Transactions on Graphics, 36(4):68:1–68:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2017:ESSb

[PM17b] Zherong Pan and Dinesh Manocha. Efficient solver for spacetime control of smoke. ACM Transactions on Graphics, 36(5):162:1–162:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2018:AAR

[PM18] Zherong Pan and Dinesh Manocha. Active animations of reduced deformable models with environment interactions. ACM Transactions on Graphics, 37(3):36:1– 36:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3197565.

Pillwein:2021:GDE

[PM21] Stefan Pillwein and Przemyslaw Musialski. Generalized deployable elastic geodesic grids. ACM Transactions on Graphics, 40(6): 271:1–271:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480516.

Preiner:2014:CPF

[PMA^{+}14] Reinhold Preiner, Oliver Mattausch, Murat Arikan, Renato Pajarola, and Michael Wimmer. Continuous projection for fast L_{1} reconstruction. ACM Transactions on Graphics, 33(4):47:1–47:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2021:AAM

[PMA^{+}21] Xue Bin Peng, Ze Ma, Pieter Abbeel, Sergey Levine, and Angjoo Kanazawa. AMP: adversarial motion priors for stylized physics-based character control. ACM Transactions on Graphics, 40(4): 144:1–144:20, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459670.

Palubicki:2022:ECR

[PMG^{+}22] Wojtek Pałubicki, Miłosz Makowski, Weronika Gajda, Torsten Hädrich, Dominik L. Michels, and Sören Pirk. Ecoclimates: climate-response modeling of vegetation. ACM Transactions on Graphics, 41 (4):155:1–155:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530146.

Philip:2021:FVI

[PMGD21] Julien Philip, Sébastien Morgenthaler, Michaël Gharbi, and George Drettakis. Freeviewpoint indoor neural relighting from multi-view stereo. ACM Transactions on Graphics, 40(5):194:1– 194:18, October 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3469842.

Peters:2019:UMR

[PMHD19] Christoph Peters, Sebastian Merzbach, Johannes Hanika, and Carsten Dachsbacher. Using moments to represent bounded signals for spectral rendering. ACM Transactions on Graphics, 38(4): 136:1–136:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Puhachov:2023:RMM

[PMKB23] Ivan Puhachov, Cedric Martens, Paul G. Kry, and Mikhail Bessmeltsev. Reconstruction of machine-made shapes from bitmap sketches. ACM Transactions on Graphics, 42(6): 268:1–268:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618361.

Peers:2009:CLT

[PML^{+}09] Pieter Peers, Dhruv K. Mahajan, Bruce Lamond, Abhijeet Ghosh, Wojciech Matusik, Ravi Ramamoorthi, and Paul Debevec. Compressive light transport sensing. ACM Transactions on Graphics, 28(1):3:1–3:18, January 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panetta:2022:SRI

[PMLB22] Julian Panetta, Haleh Mohammadian, Emiliano Luci, and Vahid Babaei. Shape from release: Inverse design and fabrication of controlled release structures. ACM Transactions on Graphics, 41(6): 274:1–274:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555518.

Pamplona:2010:NID

[PMOR10] Vitor F. Pamplona, Ankit Mohan, Manuel M. Oliveira, and Ramesh Raskar. NETRA: interactive display for estimating refractive errors and focal range. ACM Transactions on Graphics, 29(4): 77:1–77:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pons-Moll:2017:CSC

[PMPHB17] Gerard Pons-Moll, Sergi Pujades, Sonny Hu, and Michael J. Black. ClothCap: seamless 4D clothing capture and retargeting. ACM Transactions on Graphics, 36(4): 73:1–73:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pons-Moll:2015:DMD

[PMRMB15] Gerard Pons-Moll, Javier Romero, Naureen Mahmood, and Michael J. Black. Dyna: a model of dynamic human shape in motion. ACM Transactions on Graphics, 34(4): 120:1–120:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Patterson:2012:SCN

[PMS12] Taylor Patterson, Nathan Mitchell, and Eftychios Sifakis. Simulation of complex nonlinear elastic bodies using lattice deformers. ACM Transactions on Graphics, 31(6): 197:1–197:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pauly:2008:DSR

[PMW^{+}08] Mark Pauly, Niloy J. Mitra, Johannes Wallner, Helmut Pottmann, and Leonidas J. Guibas. Discovering structural regularity in 3D geometry. ACM Transactions on Graphics, 27(3):43:1– 43:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pietroni:2021:RFL

[PNA^{+}21] Nico Pietroni, Stefano Nuvoli, Thomas Alderighi, Paolo Cignoni, and Marco Tarini. Reliable feature-line driven quad-remeshing. ACM Transactions on Graphics, 40(4): 155:1–155:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459941.

Puhachov:2021:KDL

[PNCB21] Ivan Puhachov, William Neveu, Edward Chien, and Mikhail Bessmeltsev. Keypointdriven line drawing vectorization via PolyVector flow. ACM Transactions on Graphics, 40(6):266:1–266:17, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480529.

Pfaff:2014:ATC

[PNdJO14] Tobias Pfaff, Rahul Narain, Juan Miguel de Joya, and James F. O’Brien. Adaptive tearing and cracking of thin sheets. ACM Transactions on Graphics, 33(4): 110:1–110:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pirk:2012:CAM

[PNDN12] Sören Pirk, Till Niese, Oliver Deussen, and Boris Neubert. Capturing and animating the morphogenesis of polygonal tree models. ACM Transactions on Graphics, 31(6): 169:1–169:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pirk:2014:WTC

[PNH^{+}14] Sören Pirk, Till Niese, Torsten Hädrich, Bedrich Benes, and Oliver Deussen. Windy trees: computing stress response for developmental tree models. ACM Transactions on Graphics, 33(6):204:1– 204:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paliwal:2023:RVD

[PNTK23] Avinash Paliwal, Brandon G.

Nguyen, Andrii Tsarov, and Nima Khademi Kalantari. ReShader: View-dependent highlights for single image view-synthesis. ACM Transactions on Graphics, 42(6): 216:1–216:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618393.

Patney:2008:RTR

[PO08] Anjul Patney and John D. Owens. Real-time Reyesstyle adaptive surface subdivision. ACM Transactions on Graphics, 27(5):143:1– 143:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Poulenard:2018:MDG

[PO18] Adrien Poulenard and Maks Ovsjanikov. Multi-directional geodesic neural networks via equivariant convolution. ACM Transactions on Graphics, 37 (6):236:1–236:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pamplona:2012:TDC

[POAR12] Vitor F. Pamplona, Manuel M. Oliveira, Daniel G. Aliaga, and Ramesh Raskar. Tailored displays to compensate for visual aberrations. ACM Transactions on Graphics, 31(4): 81:1–81:12, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pamplona:2009:PMP

[POB09] Vitor F. Pamplona, Manuel M.

Oliveira, and Gladimir V. G. Baranoski. Photorealistic models for pupil light reflex and iridal pattern deformation. ACM Transactions on Graphics, 28(4): 106:1–106:12, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Policarpo:2005:RTR

[POC05] Fábio Policarpo, Manuel M.

Oliveira, and João L. D. Comba. Real-time relief mapping on arbitrary polygonal surfaces. ACM Transactions on Graphics, 24(3):935, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Poya:2023:GOS

[POK23] Roman Poya, Rogelio Ortigosa, and Theodore Kim. Geometric optimisation via spectral shifting. ACM Transactions on Graphics, 42 (3):29:1–29:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3585003.

Pottmann:1991:LCC

[Pot91]  Helmut Pottmann.  Lo-

cally controllable conic splines with curvature continuity. ACM Transactions on Graphics, 10(4):366–377, October 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
116916.html.

Perez:2017:CDA

[POT17] Jesús Pérez, Miguel A. Otaduy, and Bernhard Thomaszewski. Computational design and automated fabrication of Kirchhoff-plateau surfaces. ACM Transactions on Graphics, 36(4):62:1–62:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paluszny:1993:FTC

[PP93] Marco Paluszny and Richard R. Patterson. A family of tangent continuous cubic algebraic splines. ACM Transactions on Graphics, 12(3):209–232, July 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
169707.html.

Paglieroni:1994:HDD

[PP94] David W. Paglieroni and Sidney M. Petersen. Height distributional distance transform methods for height field ray tracing. ACM Transactions on Graphics, 13 (4):376–399, October 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
197312.html.

Panozzo:2014:FFA

[PPTSH14] Daniele Panozzo, Enrico Puppo, Marco Tarini, and Olga Sorkine-Hornung. Frame fields: anisotropic and nonorthogonal cross fields. ACM Transactions on Graphics, 33 (4):134:1–134:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paoluzzi:1995:GPP

[PPV95] Alberto Paoluzzi, Valerio Pascucci, and Michele Vicentino. Geometric programming: a programming approach to geometric design. ACM Transactions on Graphics, 14(3):266–306, July 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
212349.html.

Peng:2018:DPU

[PPW18] Chi-Han Peng, Helmut Pottmann, and Peter Wonka. Designing patterns using triangle-quad hybrid meshes. ACM Transactions on Graphics, 37(4): 107:1–107:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pamplona:2011:CIM

[PPZ^{+}11] Vitor F. Pamplona, Erick B. Passos, Jan Zizka, Manuel M. Oliveira, Everett Lawson, Esteban Clua, and Ramesh Raskar. CATRA: interactive measuring and modeling of cataracts. ACM Transactions on Graphics, 30(4): 47:1–47:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pang:2023:BSG

[PQF^{+}23] Kunkun Pang, Dafei Qin, Yingruo Fan, Julian Habekost, Takaaki Shiratori, Junichi Yamagishi, and Taku Komura. BodyFormer: Semanticsguided 3D body gesture synthesis with transformer. ACM Transactions on Graphics, 42 (4):43:1–43:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592456.

Pang:2008:SAH

[PQW^{+}08] Wai-Man Pang, Yingge Qu, Tien-Tsin Wong, Daniel Cohen-Or, and Pheng-Ann Heng. Structure-aware halftoning. ACM Transactions on Graphics, 27(3):89:1– 89:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:1997:SII

[PR97a] F. C. Park and Bahram Ravani. Smooth invariant interpolation of rotations. ACM Transactions on Graphics, 16 (3):277–295, July 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1997-16-3/
p277-park/.

Peters:1997:SSS

[PR97b] Jörg Peters and Ulrich Reif.

The simplest subdivision scheme for smoothing polyhedra. ACM Transactions on Graphics, 16(4):420–431, October 1997. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1997-16-4/
p420-peters/.

Popescu:2006:FR

[PR06] Voicu Popescu and Paul Rosen. Forward rasterization. ACM Transactions on Graphics, 25(2):375–411, April 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Prautzsch:1989:RTB

[Pra89] Hartmut Prautzsch. A round trip to B-splines via De Casteljau. ACM Transactions on Graphics, 8(3): 243–254, July 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77061.html.

Popescu:2009:GC

[PRAV09] Voicu Popescu, Paul Rosen, and Nicoletta Adamo-Villani. The graph camera. ACM Transactions on Graphics, 28(5):158:1–158:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:2018:PPM

[PRFS18] Keunhong Park, Konstantinos Rematas, Ali Farhadi, and Steven M. Seitz. PhotoShape: photorealistic materials for large-scale shape collections. ACM Transactions on Graphics, 37(6): 192:1–192:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Papas:2013:FTM

[PRJ^{+}13] Marios Papas, Christian Regg, Wojciech Jarosz, Bernd Bickel, Philip Jackson, Wojciech Matusik, Steve Marschner, and Markus Gross. Fabricating translucent materials using continuous pigment mixtures. ACM Transactions on Graphics, 32(4): 146:1–146:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Palacios:2017:TFD

[PRK^{+}17] Jonathan Palacios, Lawrence Roy, Prashant Kumar, Chen-Yuan Hsu, Weikai Chen, Chongyang Ma, Li-Yi Wei, and Eugene Zhang. Tensor field design in volumes. ACM Transactions on Graphics, 36 (6):188:1–188:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Portaneri:2022:AWO

[PRLH^{+}22] Cédric Portaneri, Mael Rouxel-Labbé, Michael Hemmer, David Cohen-Steiner, and Pierre Alliez. Alpha wrapping with an offset. ACM Transactions on Graphics, 41 (4):127:1–127:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530152.

Pereira:2014:CLR

[PRM14] Thiago Pereira, Szymon Rusinkiewicz, and Wojciech Matusik. Computational light routing: 3D printed optical fibers for sensing and display. ACM Transactions on Graphics, 33(3):24:1–24:??, May 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pejsa:2016:ADG

[PRMG16] Tomislav Pejsa, Daniel Rakita, Bilge Mutlu, and Michael Gleicher. Authoring directed gaze for full-body motion capture. ACM Transactions on Graphics, 35(6):161:1– 161:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Paille:2015:DAB

[PRP^{+}15] Gilles-Philippe Paillé, Nicolas Ray, Pierre Poulin, Alla Sheffer, and Bruno Lévy. Dihedral angle-based maps of tetrahedral meshes. ACM Transactions on Graphics, 34(4): 54:1–54:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pai:2018:HTM

[PRWH^{+}18] Dinesh K. Pai, Austin Rothwell, Pearson Wyder-Hodge, Alistair Wick, Ye Fan, Egor Larionov, Darcy Harrison, Debanga Raj Neog, and Cole Shing. The human touch: measuring contact with real human soft tissues. ACM Transactions on Graphics, 37 (4):58:1–58:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panetta:2017:WCS

[PRZ17] Julian Panetta, Abtin Rahimian, and Denis Zorin. Worstcase stress relief for microstructures. ACM Transactions on Graphics, 36(4): 122:1–122:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peters:2004:CDS

[PS04] Jörg Peters and Le-Jeng Shiue. Combining 4- and 3direction subdivision. ACM Transactions on Graphics, 23 (4):980–1003, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Petschnigg:2004:DPF

[PSA^{+}04] Georg Petschnigg, Richard Szeliski, Maneesh Agrawala, Michael Cohen, Hugues Hoppe, and Kentaro Toyama. Digital photography with flash and no-flash image pairs. ACM Transactions on Graphics, 23 (3):664–672, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pottmann:2008:FSS

[PSB^{+}08] Helmut Pottmann, Alexander Schiftner, Pengbo Bo, Heinz Schmiedhofer, Wenping Wang, Niccolo Baldassini, and Johannes Wallner. Freeform surfaces from single curved panels. ACM Transactions on Graphics, 27 (3):76:1–76:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pascucci:2007:RLC

[PSBM07] Valerio Pascucci, Giorgio Scorzelli, Peer-Timo Bremer, and Ajith Mascarenhas. Robust on-line computation of Reeb graphs: simplicity and speed. ACM Transactions on Graphics, 26(3): 58:1–58:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pilleboue:2015:VAM

[PSC^{+}15] Adrien Pilleboue, Gurprit Singh, David Coeurjolly, Michael Kazhdan, and Victor Ostromoukhov. Variance analysis for Monte Carlo integration. ACM Transactions on Graphics, 34(4): 124:1–124:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Popovic:2003:MSC

[PSE03] Jovan Popović, Steven M. Seitz, and Michael Erdmann. Motion sketching for control of rigid-body simulations. ACM Transactions on Graphics, 22(4):1034–1054, October 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Patane:2009:TED

[PSF09] Giuseppe Patanè, Michela Spagnuolo, and Bianca Falcidieno. Topology- and errordriven extension of scalar functions from surfaces to volumes. ACM Transactions on Graphics, 29(1):4:1– 4:20, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Podolak:2006:PRS

[PSG^{+}06] Joshua Podolak, Philip Shilane, Aleksey Golovinskiy, Szymon Rusinkiewicz, and Thomas Funkhouser. A planar-reflective symmetry transform for 3D shapes. ACM Transactions on Graphics, 25(3):549–559, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:2021:HHD

[PSH^{+}21] Keunhong Park, Utkarsh Sinha, Peter Hedman, Jonathan T. Barron, Sofien Bouaziz, Dan B. Goldman, Ricardo Martin-Brualla, and Steven M. Seitz. HyperNeRF: a higherdimensional representation for topologically varying neural radiance fields. ACM Transactions on Graphics, 40(6): 238:1–238:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480487.

Pirk:2012:PTI

[PSK^{+}12] Sören Pirk, Ondrej Stava, Julian Kratt, Michel Abdul Massih Said, Boris Neubert, Radomír Mech, Bedrich Benes, and Oliver Deussen. Plastic trees: interactive self-adapting botanical tree models. ACM Transactions on Graphics, 31(4): 50:1–50:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Patney:2016:TFR

[PSK^{+}16] Anjul Patney, Marco Salvi, Joohwan Kim, Anton Kaplanyan, Chris Wyman, Nir Benty, David Luebke, and Aaron Lefohn. Towards foveated rendering for gazetracked virtual reality. ACM Transactions on Graphics, 35 (6):179:1–179:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Park:2020:NCN

[PSN20] Sanghun Park, Kwanggyoon Seo, and Junyong Noh. Neural crossbreed: neural based image metamorphosis. ACM Transactions on Graphics, 39 (6):224:1–224:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417797.

Patel:2013:ICS

[PSNB13] Daniel Patel, Veronika Soltészová, Jan Martin Nordbotten, and Stefan Bruckner. Instant convolution shadows for volumetric detail mapping. ACM Transactions on Graphics, 32 (5):154:1–154:18, September 2013. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Pfaff:2010:SFS

[PTC^{+}10] Tobias Pfaff, Nils Thuerey, Jonathan Cohen, Sarah Tariq, and Markus Gross. Scalable fluid simulation using anisotropic turbulence particles. ACM Transactions on Graphics, 29(6):174:1– 174:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Perez:2015:DFF

[PTC^{+}15] Jesús Pérez, Bernhard Thomaszewski, Stelian Coros, Bernd Bickel, José A. Canabal, Robert Sumner, and Miguel A. Otaduy. Design and fabrication of flexible rod meshes. ACM Transactions on Graphics, 34(4): 138:1–138:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellacini:2002:UII

[PTG02] Fabio Pellacini, Parag Tole, and Donald P. Greenberg. A user interface for interactive cinematic shadow design. ACM Transactions on Graphics, 21(3):563–566, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pfaff:2012:LVS

[PTG12] Tobias Pfaff, Nils Thuerey, and Markus Gross. Lagrangian vortex sheets for animating fluids. ACM Transactions on Graphics, 31(4): 112:1–112:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Poranne:2017:ASD

[PTH^{+}17] Roi Poranne, Marco Tarini, Sandro Huber, Daniele Panozzo, and Olga Sorkine-Hornung. Autocuts: simultaneous distortion and cut optimization for UV mapping. ACM Transactions on Graphics, 36(6): 215:1–215:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peers:2007:PPF

[PTMD07] Pieter Peers, Naoki Tamura, Wojciech Matusik, and Paul Debevec. Post-production facial performance relighting using reflectance transfer. ACM Transactions on Graphics, 26 (3):52:1–52:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pfaff:2009:STU

[PTSG09] Tobias Pfaff, Nils Thuerey, Andrew Selle, and Markus Gross. Synthetic turbulence using artificial boundary layers. ACM Transactions on Graphics, 28(5): 121:1–121:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Patney:2015:PFA

[PTSO15] Anjul Patney, Stanley Tzeng, Kerry A. Seitz, Jr., and John D. Owens. Piko: a framework for authoring programmable graphics pipelines. ACM Transactions on Graphics, 34(4):147:1– 147:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pietroni:2011:GPR

[PTSZ11] Nico Pietroni, Marco Tarini, Olga Sorkine, and Denis Zorin. Global parametrization of range image sets. ACM Transactions on Graphics, 30 (6):149:1–149:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pietroni:2017:PBT

[PTV^{+}17] Nico Pietroni, Marco Tarini, Amir Vaxman, Daniele Panozzo, and Paolo Cignoni. Positionbased tensegrity design. ACM Transactions on Graphics, 36 (6):172:1–172:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Prautzsch:2006:PTS

[PU06] Hartmut Prautzsch and Georg Umlauf. Parametrizations for triangular G^{k} spline surfaces of low degree. ACM Transactions on Graphics, 25 (4):1281–1293, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Plantinga:2006:CCG

[PV06] Simon Plantinga and Gert Vegter. Computing contour generators of evolving implicit surfaces. ACM Transactions on Graphics, 25(4): 1243–1280, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peers:2006:CFR

[PvBM^{+}06] Pieter Peers, Karl vom Berge, Wojciech Matusik, Ravi Ramamoorthi, Jason Lawrence, Szymon Rusinkiewicz, and Philip Dutré. A compact factored representation of heterogeneous subsurface scattering. ACM Transactions on Graphics, 25(3):746–753, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pediredla:2019:EPC

[PVG19] Adithya Pediredla, Ashok Veeraraghavan, and Ioannis Gkioulekas. Ellipsoidal path connections for timegated rendering. ACM Transactions on Graphics, 38(4): 38:1–38:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellacini:2005:LHH

[PVL^{+}05] Fabio Pellacini, Kiril Vidimče, Aaron Lefohn, Alex Mohr, Mark Leone, and John Warren. Lpics: a hybrid hardwareaccelerated relighting engine for computer cinematography. ACM Transactions on Graphics, 24(3):464–470, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pellerin:2018:TSH

[PVR18] Jeanne Pellerin, Kilian Verhetsel, and Jean-FranÇois Remacle. There are 174 subdivisions of the hexahedron into tetrahedra. ACM Transactions on Graphics, 37(6): 266:1–266:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Preparata:1990:CAV

[PVY90] Franco P. Preparata, Jeffrey Scott Vitter, and Mariette Yvinec. Computation of the axial view of a set of isothetic parallelepipeds. ACM Transactions on Graphics, 9(3):278–300, July 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
78967.html.

Prevost:2013:MIS

[PWLSH13] Romain Prévost, Emily Whiting, Sylvain Lefebvre, and Olga Sorkine-Hornung. Make it stand: balancing shapes for 3D fabrication. ACM Transactions on Graphics, 32(4): 81:1–81:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2018:AS

[PXW18] Mengqi Peng, Jun Xing, and Li-Yi Wei. Autocomplete 3D sculpting. ACM Transactions on Graphics, 37(4): 132:1–132:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ponton:2024:SRT

[PYA^{+}24] Jose Luis Ponton, Haoran Yun, Andreas Aristidou, Carlos Andujar, and Nuria Pelechano. Sparse-Poser: Real-time full-body motion reconstruction from sparse data. ACM Transactions on Graphics, 43 (1):5:1–5:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3625264.

Peng:2016:CND

[PYB^{+}16] Chi-Han Peng, Yong-Liang Yang, Fan Bao, Daniel Fink, Dong-Ming Yan, Peter Wonka, and Niloy J. Mitra. Computational network design from functional specifications. ACM Transactions on Graphics, 35(4): 131:1–131:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2014:CLD

[PYW14] Chi-Han Peng, Yong-Liang Yang, and Peter Wonka. Computing layouts with deformable templates. ACM Transactions on Graphics, 33 (4):99:1–99:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Palacios:2007:RSF

[PZ07] Jonathan Palacios and Eugene Zhang. Rotational symmetry field design on surfaces. ACM Transactions on Graphics, 26(3):55:1–55:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Parilov:2008:RTR

[PZ08] Evgueni Parilov and Denis Zorin. Real-time rendering of textures with feature curves. ACM Transactions on Graphics, 27(1): 3:1–3:15, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Penner:2017:SRV

[PZ17] Eric Penner and Li Zhang.

Soft 3D reconstruction for view synthesis. ACM Transactions on Graphics, 36(6):

235:1–235:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Peng:2011:CEQ

[PZKW11] Chi-Han Peng, Eugene Zhang, Yoshihiro Kobayashi, and Peter Wonka. Connectivity editing for quadrilateral meshes. ACM Transactions on Graphics, 30(6):141:1– 141:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pan:2013:EPD

[PZM13] Jia Pan, Xinyu Zhang, and Dinesh Manocha. Efficient penetration depth approximation using active learning. ACM Transactions on Graphics, 32(6):191:1– 191:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Panetta:2015:ETA

[PZM^{+}15] Julian Panetta, Qingnan Zhou, Luigi Malomo, Nico Pietroni, Paolo Cignoni, and Denis Zorin. Elastic textures for additive fabrication. ACM Transactions on Graphics, 34(4):135:1–135:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Pang:2023:LGE

[PZWW23] Bo Pang, Zhongtian Zheng, Guoping Wang, and Peng-Shuai Wang. Learning the geodesic embedding with graph neural networks. ACM Transactions on Graphics, 42 (6):236:1–236:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618317.

Qin:2017:WBNa

[QCHC17a] Hongxing Qin, Yi Chen, Jinlong He, and Baoquan Chen. Wasserstein blue noise sampling. ACM Transactions on Graphics, 36(4):137:1– 137:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qin:2017:WBNb

[QCHC17b] Hongxing Qin, Yi Chen, Jinlong He, and Baoquan Chen. Wasserstein blue noise sampling. ACM Transactions on Graphics, 36(5):168:1– 168:??, October 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qin:2023:SLM

[QCOS23] Yingsi Qin, Wei-Yu Chen, Matthew O’Toole, and Aswin C. Sankaranarayanan. Split-Lohmann multifocal displays. ACM Transactions on Graphics, 42(4):57:1–57:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592110.

Qin:2016:FED

[QHY^{+}16] Yipeng Qin, Xiaoguang Han, Hongchuan Yu, Yizhou Yu, and Jianjun Zhang. Fast and exact discrete geodesic computation based on triangleoriented wavefront propagation. ACM Transactions on Graphics, 35(4):125:1– 125:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qu:2021:FLN

[QJ21] Ante Qu and Doug L. James. Fast linking numbers for topology verification of loopy structures. ACM Transactions on Graphics, 40(4): 106:1–106:19, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459778.

Qu:2022:PPC

[QLDJ22] Ziyin Qu, Minchen Li, Fernando De Goes, and Chenfanfu Jiang. The power particle-in-cell method. ACM Transactions on Graphics, 41 (4):118:1–118:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530066.

Qiu:2022:SSC

[QLH^{+}22] Zesong Qiu, Yuwei Li, Dongming He, Qixuan Zhang, Longwen Zhang, Yinghao Zhang, Jingya Wang, Lan Xu, Xudong Wang, Yuyao Zhang, and Jingyi Yu. SCULPTOR: Skeleton-consistent face creation using a learned parametric generator. ACM Transactions on Graphics, 41(6): 213:1–213:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555462.

Qu:2023:PPH

[QLY^{+}23] Ziyin Qu, Minchen Li, Yin Yang, Chenfanfu Jiang, and Fernando De Goes. Power plastics: a hybrid Lagrangian/Eulerian solver for mesoscale inelastic flows. ACM Transactions on Graphics, 42(6):240:1–240:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618344.

Qu:2008:RPM

[QPWH08] Yingge Qu, Wai-Man Pang, Tien-Tsin Wong, and Pheng-Ann Heng. Richness-preserving manga screening. ACM Transactions on Graphics, 27 (5):155:1–155:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qiu:2023:SDG

[QRL^{+}23] Yuxing Qiu, Samuel Temple Reeve, Minchen Li, Yin Yang, Stuart Ryan Slattery, and Chenfanfu Jiang. A sparse distributed gigascale resolution material point method. ACM Transactions on Graphics, 42 (2):22:1–22:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3570160.

Qin:2015:UPG

[QSH^{+}15] Hao Qin, Xin Sun, Qiming Hou, Baining Guo, and Kun Zhou. Unbiased photon gathering for light transport simulation. ACM Transactions on Graphics, 34(6): 208:1–208:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Quan:2006:IBP

[QTZ^{+}06] Long Quan, Ping Tan, Gang Zeng, Lu Yuan, Jingdong Wang, and Sing Bing Kang. Image-based plant modeling. ACM Transactions on Graphics, 25(3):599–604, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qu:2006:MC

[QWH06] Yingge Qu, Tien-Tsin Wong, and Pheng-Ann Heng. Manga colorization. ACM Transactions on Graphics, 25(3): 1214–1220, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qu:2019:ECF

[QZG^{+}19] Ziyin Qu, Xinxin Zhang, Ming Gao, Chenfanfu Jiang, and Baoquan Chen. Efficient and conservative fluids using bidirectional mapping. ACM Transactions on Graphics, 38 (4):128:1–128:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Qin:2022:MBT

[QZZ22] Jia Qin, Youyi Zheng, and Kun Zhou. Motion inbetweening via two-stage transformers. ACM Transactions on Graphics, 41(6): 184:1–184:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555454.

Rivers:2012:SN

[RAD12] Alec Rivers, Andrew Adams, and Frédo Durand. Sculpting by numbers. ACM Transactions on Graphics, 31(6): 157:1–157:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Roimela:2006:HDR

[RAI06] Kimmo Roimela, Tomi Aarnio, and Joonas Itäranta. High dynamic range texture compression. ACM Transactions on Graphics, 25(3):707–712, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rav-Acha:2008:UMN

[RAKRF08] Alex Rav-Acha, Pushmeet Kohli, Carsten Rother, and Andrew Fitzgibbon. Unwrap mosaics: a new representation for video editing. ACM Transactions on Graphics, 27(3): 17:1–17:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rakotosaona:2021:DST

[RAM^{+}21] Marie-Julie Rakotosaona, Noam Aigerman, Niloy J. Mitra, Maks Ovsjanikov, and Paul Guerrero. Differentiable surface triangulation. ACM Transactions on Graphics, 40 (6):267:1–267:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478513.3480554.

Ramamoorthi:2012:TMC

[RAMN12] Ravi Ramamoorthi, John Anderson, Mark Meyer, and Derek Nowrouzezahrai. A theory of Monte Carlo visibility sampling. ACM Transactions on Graphics, 31(5): 121:1–121:16, August 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rappoport:1991:RCS

[Rap91]  Ari Rappoport.  Render-

ing curves and surfaces with hybrid subdivision and forward differencing. ACM Transactions on Graphics, 10 (4):323–341, October 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
116914.html.

Rosales:2021:AAG

[RAR^{+}21] Enrique Rosales, Chrystiano Araújo, Jafet Rodriguez, Nicholas Vining, Dongwook Yoon, and Alla Sheffer. AdaptiBrush: adaptive general and predictable VR ribbon brush. ACM Transactions on Graphics, 40(6): 247:1–247:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480511.

Raskar:2006:CEP

[RAT06] Ramesh Raskar, Amit Agrawal, and Jack Tumblin. Coded exposure photography: motion deblurring using fluttered shutter. ACM Transactions on Graphics, 25(3): 795–804, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raskar:2008:GAP

[RAWV08] Ramesh Raskar, Amit Agrawal, Cyrus A. Wilson, and Ashok Veeraraghavan. Glare aware photography: 4D ray sampling for reducing glare effects of camera lenses. ACM Transactions on Graphics, 27(3): 56:1–56:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rmaileh:2023:MFG

[RB23] Lubna Abu Rmaileh and Alan Brunton. Meso-facets for goniochromatic 3D printing. ACM Transactions on Graphics, 42(4):66:1–66:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592137.

Rusinkiewicz:2006:ESD

[RBD06] Szymon Rusinkiewicz, Michael Burns, and Doug DeCarlo. Exaggerated shading for depicting shape and detail. ACM Transactions on Graphics, 25(3):1199–1205, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ramanarayanan:2008:PCA

[RBF08] Ganesh Ramanarayanan, Kavita Bala, and James A. Ferwerda. Perception of complex aggregates. ACM Transactions on Graphics, 27(3):60:1– 60:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rother:2006:A

[RBHB06] Carsten Rother, Lucas Bordeaux, Youssef Hamadi, and Andrew Blake. AutoCollage. ACM Transactions on Graphics, 25(3):847–852, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ribardiere:2019:MBG

[RBSM19] Mickaël Ribardière, Benjamin Bringier, Lionel Simonot, and Daniel Meneveaux. Microfacet BSDFs generated from NDFs and explicit microgeometry. ACM Transactions on Graphics, 38(5): 143:1–143:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3338697.

Raskar:2004:RLI

[RBvB^{+}04] Ramesh Raskar, Paul Beardsley, Jeroen van Baar, Yao Wang, Paul Dietz, Johnny Lee, Darren Leigh, and Thomas Willwacher. RFIG lamps: interacting with a self-describing world via photosensing wireless tags and projectors. ACM Transactions on Graphics, 23(3):406– 415, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rodriguez:2022:TSM

[RC22] Alejandro Rodríguez and Gabriel Cirio. True seams: modeling seams in digital garments. ACM Transactions on Graphics, 41(4): 62:1–62:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530128.

Romero:2022:CCD

[RCCO22] Cristian Romero, Dan Casas, Maurizio M. Chiaramonte, and Miguel A. Otaduy. Contact-centric deformation learning. ACM Transactions on Graphics, 41(4): 70:1–70:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530182.

Randrianandrasana:2021:TMB

[RCL21] Joël Randrianandrasana, Patrick Callet, and Laurent Lucas. Transfer matrix based layered materials rendering. ACM Transactions on Graphics, 40(4):177:1–177:16, August 2021. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459859.

Roussel:2019:DCR

[RCLM19] Robin Roussel, Marie-Paule Cani, Jean-Claude Léon, and Niloy J. Mitra. Designing chain reaction contraptions from causal graphs. ACM Transactions on Graphics, 38 (4):43:1–43:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rosenberger:2009:LSS

[RCOL09] Amir Rosenberger, Daniel Cohen-Or, and Dani Lischinski. Layered shape synthesis: automatic generation of control maps for non-stationary textures. ACM Transactions on Graphics, 28(5):107:1– 107:5, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Romero:2021:LCC

[RCPO21] Cristian Romero, Dan Casas, Jesús Pérez, and Miguel Otaduy. Learning contact corrections for handle-based subspace dynamics. ACM Transactions on Graphics, 40(4): 131:1–131:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459875.

Rivers:2010:MS

[RDI10] Alec Rivers, Frédo Durand, and Takeo Igarashi. 3D modeling with silhouettes. ACM Transactions on Graphics, 29 (4):109:1–109:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2015:IBR

[RDL^{+}15] Peiran Ren, Yue Dong, Stephen Lin, Xin Tong, and Baining Guo. Image based relighting using neural networks. ACM Transactions on Graphics, 34(4):111:1– 111:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reeves:1983:PST

[Ree83] W. T. Reeves. Particle systems – a technique for modeling a class of fuzzy objects. ACM Transactions on Graphics, 2(2):91–108, April 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ritschel:2009:MRS

[REG^{+}09] T. Ritschel, T. Engelhardt, T. Grosch, H.-P. Seidel, J. Kautz, and C. Dachsbacher. Micro-rendering for scalable, parallel final gathering. ACM Transactions on Graphics, 28(5):132:1–132:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ressler:1987:IGT

[Res87] Sanford Ressler. The incrementor: a graphical technique for manipulating parameters. ACM Transactions on Graphics, 6(1):74–78, January 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
214381.html.

Runions:2005:MVL

[RFL^{+}05] Adam Runions, Martin Fuhrer, Brendan Lane, Pavol Federl, Anne-Gaëlle Rolland-Lagan, and Przemyslaw Prusinkiewicz. Modeling and visualization of leaf venation patterns. ACM Transactions on Graphics, 24 (3):702–711, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ruckert:2022:AAD

[RFS22] Darius Rückert, Linus Franke, and Marc Stamminger. ADOP: approximate differentiable one-pixel point rendering. ACM Transactions on Graphics, 41(4):99:1–99:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530122.

Ren:2023:DDD

[RFW^{+}23] Haocheng Ren, Hangming Fan, Rui Wang, Yuchi Huo, Rui Tang, Lei Wang, and Hujun Bao. Data-driven digital lighting design for residential indoor spaces. ACM Transactions on Graphics, 42 (3):28:1–28:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3582001.

Ramanarayanan:2007:VET

[RFWB07] Ganesh Ramanarayanan, James Ferwerda, Bruce Walter, and Kavita Bala. Visual equivalence: towards a new standard for image fidelity. ACM Transactions on Graphics, 26 (3):76:1–76:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raymond:2016:MSR

[RGB16] Boris Raymond, Gaël Guennebaud, and Pascal Barla. Multi-scale rendering of scratched materials using a structured SV-BRDF model. ACM Transactions on Graphics, 35 (4):57:1–57:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reddy:2020:DPS

[RGF^{+}20] Pradyumna Reddy, Paul Guerrero, Matt Fisher, Wilmot Li, and Niloy J. Mitra. Discovering pattern structure using differentiable compositing. ACM Transactions on Graphics, 39(6):

262:1–262:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417830.

Rath:2022:EEA

[RGH^{+}22] Alexander Rath, Pascal Grittmann, Sebastian Herholz, Philippe Weier, and Philipp Slusallek. EARS: efficiency-aware Russian roulette and splitting. ACM Transactions on Graphics, 41(4):81:1–81:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530168.

Ritschel:2008:ISM

[RGK^{+}08] T. Ritschel, T. Grosch, M. H.

Kim, H.-P. Seidel, C. Dachsbacher, and J. Kautz. Imperfect shadow maps for efficient computation of indirect illumination. ACM Transactions on Graphics, 27(5): 129:1–129:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Redon:2005:ADA

[RGL05] Stephane Redon, Nico Galoppo, and Ming C. Lin. Adaptive dynamics of articulated bodies. ACM Transactions on Graphics, 24(3): 936–945, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rubinstein:2010:CSI

[RGSS10] Michael Rubinstein, Diego Gutierrez, Olga Sorkine, and Ariel Shamir. A comparative study of image retargeting. ACM Transactions on Graphics, 29(6):160:1– 160:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ramamoorthi:2002:FSE

[RH02] Ravi Ramamoorthi and Pat Hanrahan. Frequency space environment map rendering. ACM Transactions on Graphics, 21(3):517–526, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ramamoorthi:2004:SPF

[RH04] Ravi Ramamoorthi and Pat Hanrahan. A signal-processing framework for reflection. ACM Transactions on Graphics, 23(4):1004–1042, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Roberts:2016:GDF

[RH16] Mike Roberts and Pat Hanrahan. Generating dynamically feasible trajectories for quadrotor cameras. ACM Transactions on Graphics, 35 (4):61:1–61:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Risser:2010:SSI

[RHDG10] Eric Risser, Charles Han, Rozenn Dahyot, and Eitan Grinspun. Synthesizing structured image hybrids. ACM Transactions on Graphics, 29 (4):85:1–85:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rusinkiewicz:2002:RTM

[RHHL02] Szymon Rusinkiewicz, Olaf Hall-Holt, and Marc Levoy. Real-time 3D model acquisition. ACM Transactions on Graphics, 21(3):438–446, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reibold:2018:SGS

[RHJD18] Florian Reibold, Johannes Hanika, Alisa Jung, and Carsten Dachsbacher. Selective guided sampling with complete light transport paths. ACM Transactions on Graphics, 37(6):223:1– 223:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rabinovich:2018:DGN

[RHSH18a] Michael Rabinovich, Tim Hoffmann, and Olga Sorkine-Hornung. Discrete geodesic nets for modeling developable surfaces. ACM Transactions on Graphics, 37(2): 16:1–16:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rabinovich:2018:SSD

[RHSH18b] Michael Rabinovich, Tim Hoffmann, and Olga Sorkine-Hornung. The shape space of discrete orthogonal geodesic nets. ACM Transactions on Graphics, 37(6):228:1– 228:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rappoport:1994:IDS

[RHW94] Ari Rappoport, Yaacov Hel-Or, and Michael Werman. Interactive design of smooth objects with probabilistic point constraints. ACM Transactions on Graphics, 13(2): 156–176, April 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
176582.html.

Rivers:2010:CM

[RID10] Alec Rivers, Takeo Igarashi, and Frédo Durand. 2.5D cartoon models. ACM Transactions on Graphics, 29(4): 59:1–59:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ritchie:2018:SDL

[Rit18] Daniel Ritchie. Session details: Learning to compose & decompose. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rivers:2007:FFL

[RJ07] Alec R. Rivers and Doug L.

James. FastLSM: fast lattice shape matching for robust real-time deformation. ACM Transactions on Graphics, 26 (3):82:1–82:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rousselle:2016:ISC

[RJN16] Fabrice Rousselle, Wojciech Jarosz, and Jan Novák. Image-space control variates for rendering. ACM Transactions on Graphics, 35(6): 169:1–169:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Remillard:2013:ETS

[RK13] Olivier Rémillard and Paul G.

Kry. Embedded thin shells for wrinkle simulation. ACM Transactions on Graphics, 32 (4):50:1–50:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ragan-Kelley:2012:DAS

[RKAP^{+}12] Jonathan Ragan-Kelley, Andrew Adams, Sylvain Paris, Marc Levoy, Saman Amarasinghe, and Frédo Durand. Decoupling algorithms from schedules for easy optimization of image processing pipelines. ACM Transactions on Graphics, 31(4): 32:1–32:12, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rother:2004:GIF

[RKB04] Carsten Rother, Vladimir Kolmogorov, and Andrew Blake. “GrabCut”: interactive foreground extraction using iterated graph cuts. ACM Transactions on Graphics, 23 (3):309–314, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reed:2023:NVR

[RKB^{+}23] Albert Reed, Juhyeon Kim, Thomas Blanford, Adithya Pediredla, Daniel Brown, and Suren Jayasuriya. Neural volumetric reconstruction for coherent synthetic aperture sonar. ACM Transactions on Graphics, 42(4): 113:1–113:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592141.

Ragan-Kelley:2007:LAI

[RKKS^{+}07] Jonathan Ragan-Kelley, Charlie Kilpatrick, Brian W.

Smith, Doug Epps, Paul Green, Christophe Hery, and Frédo Durand. The lightspeed automatic interactive lighting preview system. ACM Transactions on Graphics, 26(3): 25:1–25:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ragan-Kelley:2011:DSG

[RKLC^{+}11] Jonathan Ragan-Kelley, Jaakko Lehtinen, Jiawen Chen, Michael Doggett, and Frédo Durand. Decoupled sampling for graphics pipelines. ACM Transactions on Graphics, 30(3):17:1–17:17, May 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2022:UME

[RKP^{+}22] Yingying Ren, Uday Kusupati, Julian Panetta, Florin Isvoranu, Davide Pellis, Tian Chen, and Mark Pauly. Umbrella meshes: elastic mechanisms for freeform shape deployment. ACM Transactions on Graphics, 41 (4):152:1–152:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530089.

Rogge:2014:GRM

[RKS^{+}14] Lorenz Rogge, Felix Klose, Michael Stengel, Martin Eisemann, and Marcus Magnor. Garment replacement in monocular video sequences. ACM Transactions on Graphics, 34(1):6:1–6:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rousselle:2011:ASR

[RKZ11] Fabrice Rousselle, Claude Knaus, and Matthias Zwicker. Adaptive sampling and reconstruction using greedy error minimization. ACM Transactions on Graphics, 30(6): 159:1–159:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rousselle:2012:ARN

[RKZ12] Fabrice Rousselle, Claude Knaus, and Matthias Zwicker. Adaptive rendering with nonlocal means filtering. ACM Transactions on Graphics, 31 (6):195:1–195:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ray:2006:PGP

[RLL^{+}06] Nicolas Ray, Wan Chiu Li, Bruno Lévy, Alla Sheffer, and Pierre Alliez. Periodic global parameterization. ACM Transactions on Graphics, 25(4):1460–1485, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rioux-Lavoie:2020:DRM

[RLLG^{+}20] Damien Rioux-Lavoie, Joey Litalien, Adrien Gruson, Toshiya Hachisuka, and Derek Nowrouzezahrai. Delayed rejection Metropolis light transport. ACM Transactions on Graphics, 39(3): 26:1–26:14, June 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3388538.

Rodriguez:2020:GPR

[RLP^{+}20] Simon Rodriguez, Thomas Leimkühler, Siddhant Prakash, Chris Wyman, Peter Shirley, and George Drettakis. Glossy probe reprojection for interactive global illumination. ACM Transactions on Graphics, 39 (6):237:1–237:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417823.

Romero:2021:PVS

[RLR^{+}21] Victor Romero, MickaÃLy, Abdullah Haroon Rasheed, RaphaÃCharrondière, Arnaud Lazarus, Sébastien Neukirch, and Florence Bertails-Descoubes. Physical validation of simulators in computer graphics: a new framework dedicated to slender elastic structures and frictional contact. ACM Transactions on Graphics, 40(4):

66:1–66:19, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459931.

Rioux-Lavoie:2022:MCM

[RLSÖ^{+}22] Damien Rioux-Lavoie, Ryusuke Sugimoto, Tümay Özdemir, Naoharu H. Shimada, Christopher Batty, Derek Nowrouzezahrai, and Toshiya Hachisuka. A Monte Carlo method for fluid simulation. ACM Transactions on Graphics, 41(6): 240:1–240:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555450.

Redner:1995:SBS

[RLU95] Richard A. Redner, Mark E. Lee, and Samuel P. Uselton. Smooth B-spline illumination maps for bidirectional ray tracing. ACM Transactions on Graphics, 14 (4):337–362, October 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/

225296.html. Corrections to Figures 4–9 are available on the World-Wide Web at http:

                                        //www.acm.org/tog/AandE.
                                        html.

Ren:2014:MFS

[RLY^{+}14] Bo Ren, Chenfeng Li, Xiao Yan, Ming C. Lin, Javier Bonet, and Shi-Min Hu. Multiple-fluid SPH simulation using a mixture model. ACM Transactions on Graphics, 33(5):171:1–171:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ruan:2021:SFI

[RLZ^{+}21] Liangwang Ruan, Jinyuan Liu, Bo Zhu, Shinjiro Sueda, Bin Wang, and Baoquan Chen. Solid-fluid interaction with surface-tension-dominant contact. ACM Transactions on Graphics, 40(4):120:1–120:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459862.

Ramamoorthi:2007:FOA

[RMB07] Ravi Ramamoorthi, Dhruv Mahajan, and Peter Belhumeur. A first-order analysis of lighting, shading, and shadows. ACM Transactions on Graphics, 26(1):2:1– 2:21, January 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Russell:2013:WUO

[RMBB^{+}13] Bryan C. Russell, Ricardo Martin-Brualla, Daniel J. Butler, Steven M. Seitz, and Luke Zettlemoyer. 3D Wikipedia: using online text to automatically label and navigate reconstructed geometry. ACM Transactions on Graphics, 32(6):193:1– 193:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Roich:2023:PTL

[RMBCO23] Daniel Roich, Ron Mokady, Amit H. Bermano, and Daniel Cohen-Or. Pivotal tuning for latent-based editing of real images. ACM Transactions on Graphics, 42 (1):6:1–6:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3544777.

Reche:2004:VRI

[RMD04] Alex Reche, Ignacio Martin, and George Drettakis. Volumetric reconstruction and interactive rendering of trees from photographs. ACM Transactions on Graphics, 23 (3):720–727, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rivers:2012:PCT

[RMD12] Alec Rivers, Ilan E. Moyer, and Frédo Durand. Positioncorrecting tools for 2D digital fabrication. ACM Transactions on Graphics, 31(4): 88:1–88:7, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ritchie:2015:CPM

[RMGH15] Daniel Ritchie, Ben Mildenhall, Noah D. Goodman, and Pat Hanrahan. Controlling procedural modeling programs with stochasticallyordered sequential Monte Carlo. ACM Transactions on Graphics, 34(4):105:1– 105:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2020:MRM

[RMOW20] Jing Ren, Simone Melzi, Maks Ovsjanikov, and Peter Wonka. MapTree: recovering multiple solutions in the space of maps. ACM Transactions on Graphics, 39(6): 264:1–264:17, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417800.

Roessle:2023:GLD

[RMP^{+}23] Barbara Roessle, Norman Müller, Lorenzo Porzi, Samuel Rota Bulò, Peter Kontschieder, and Matthias Niessner. GANeRF: Leveraging discriminators to optimize neural radiance fields. ACM Transactions on Graphics, 42(6): 207:1–207:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618402.

Robinson-Mosher:2008:TWC

[RMSG^{+}08] Avi Robinson-Mosher, Tamar Shinar, Jon Gretarsson, Jonathan Su, and Ronald Fedkiw. Twoway coupling of fluids to rigid and deformable solids and shells. ACM Transactions on Graphics, 27(3):46:1– 46:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raskar:2007:PLA

[RNd^{+}07] Ramesh Raskar, Hideaki Nii, Bert deDecker, Yuki Hashimoto, Jay Summet, Dylan Moore, Yong Zhao, Jonathan Westhues, Paul Dietz, John Barnwell, Shree Nayar, Masahiko Inami, Philippe Bekaert, Michael Noland, Vlad Branzoi, and Erich Bruns. Prakash: lighting aware motion capture using photosensing markers and multiplexed illuminators. ACM Transactions on Graphics, 26(3):36:1–36:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rasmussen:2003:SSL

[RNGF03] Nick Rasmussen, Duc Quang Nguyen, Willi Geiger, and Ronald Fedkiw. Smoke simulation for large scale phenomena. ACM Transactions on Graphics, 22(3):703–707, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Riso:2022:BBO

[RNP^{+}22] Marzia Riso, Giacomo Nazzaro, Enrico Puppo, Alec Jacobson, Qingnan Zhou, and Fabio Pellacini. Bool-Surf: Boolean operations on surfaces. ACM Transactions on Graphics, 41(6): 247:1–247:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555466.

Robertson:1985:ASS

[RO85] Philip K. Robertson and John F. O’Callaghan. The application of scene synthesis techniques to the display of multidimensional image data. ACM Transactions on Graphics, 4(4):247–274, October 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

6117.html. See corrigenda [RO87].

Robertson:1987:CAS

[RO87] Philip K. Robertson and John F. O’Callaghan. Corrigenda: “The Application of Scene Synthesis Techniques to the Display of Multidimensional Image Data”. ACM Transactions on Graphics, 6 (2):162, April 1987. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [RO85].

Rodham:1994:STM

[RO94] Kenneth J. Rodham and Dan R. Olsen, Jr. Smart telepointers: Maintaining telepointer consistency in the presence of user interface customization. ACM Transactions on Graphics, 13(3): 300–307, July 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
197492.html.

Rustamov:2013:MBE

[ROA^{+}13] Raif M. Rustamov, Maks Ovsjanikov, Omri Azencot, Mirela Ben-Chen, Frédéric Chazal, and Leonidas Guibas. Map-based exploration of intrinsic shape differences and variability. ACM Transactions on Graphics, 32(4): 72:1–72:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rockwood:1989:DMI

[Roc89] A. P. Rockwood. The displacement method for implicit blending surfaces in solid models. ACM Transactions on Graphics, 8(4): 279–297, October 1989. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77271.html.

Ringham:2021:MFP

[ROC^{+}21] Lee Ringham, Andrew Owens, Mikolaj Cieslak, Lawrence D. Harder, and Przemyslaw Prusinkiewicz. Modeling flower pigmentation patterns. ACM Transactions on Graphics, 40(6):233:1–233:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480548.

Rossignac:1994:ISI

[Ros94] Jarek Rossignac. Introduction to the special issue on interactive sculpting. ACM Transactions on Graphics, 13 (2):101–102, April 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rossignac:2020:COT

[Ros20] Jarek Rossignac. Corneroperated tran-similar (COTS) maps, patterns, and lattices. ACM Transactions on Graphics, 39(1):5:1–5:14, February 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3267346.

Ritschel:2009:IRE

[ROTS09] Tobias Ritschel, Makoto Okabe, Thorsten Thormählen, and Hans-Peter Seidel. Interactive reflection editing. ACM Transactions on Graphics, 28(5):129:1–129:7, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reitsma:2003:PMC

[RP03] Paul S. A. Reitsma and Nancy S. Pollard. Perceptual metrics for character animation: sensitivity to errors in ballistic motion. ACM Transactions on Graphics, 22 (3):537–542, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reitsma:2007:EMG

[RP07] Paul S. A. Reitsma and Nancy S. Pollard. Evaluating motion graphs for character animation. ACM Transactions on Graphics, 26(4):18:1– 18:24, October 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rosenberg:2009:UIM

[RP09] Ilya Rosenberg and Ken Perlin. The UnMousePad: an interpolating multi-touch forcesensing input pad. ACM Transactions on Graphics, 28 (3):65:1–65:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rohmer:2010:AWA

[RPC^{+}10] Damien Rohmer, Tiberiu Popa, Marie-Paule Cani, Stefanie Hahmann, and Alla Sheffer. Animation wrinkling: augmenting coarse cloth simulations with realistic-looking wrinkles. ACM Transactions on Graphics, 29(6):157:1– 157:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2021:WCR

[RPC^{+}21] Yingying Ren, Julian Panetta, Tian Chen, Florin Isvoranu, Samuel Poincloux, Christopher Brandt, Alison Martin, and Mark Pauly. 3D weaving with curved ribbons. ACM Transactions on Graphics, 40(4):127:1–127:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459788.

Ren:2005:DDA

[RPE^{+}05] Liu Ren, Alton Patrick, Alexei A. Efros, Jessica K. Hodgins, and James M. Rehg. A data-driven approach to quantifying natural human motion. ACM Transactions on Graphics, 24(3): 1090–1097, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reinhard:2012:CIA

[RPK^{+}12] Erik Reinhard, Tania Pouli, Timo Kunkel, Ben Long, Anders Ballestad, and Gerwin Damberg. Calibrated image appearance reproduction. ACM Transactions on Graphics, 31(6):201:1– 201:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rabinovich:2017:SLIa

[RPPSH17a] Michael Rabinovich, Roi Poranne, Daniele Panozzo, and Olga Sorkine-Hornung. Scalable locally injective mappings. ACM Transactions on Graphics, 36(2):16:1– 16:??, April 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rabinovich:2017:SLIb

[RPPSH17b] Michael Rabinovich, Roi Poranne, Daniele Panozzo, and Olga Sorkine-Hornung. Scalable locally injective mappings. ACM Transactions on Graphics, 36(4):37:1– 37:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2018:COP

[RPWO18] Jing Ren, Adrien Poulenard, Peter Wonka, and Maks Ovsjanikov. Continuous and orientation-preserving correspondences via functional maps. ACM Transactions on Graphics, 37(6):248:1– 248:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rhodin:2016:EEM

[RRC^{+}16] Helge Rhodin, Christian Richardt, Dan Casas, Eldar Insafutdinov, Mohammad Shafiei, Hans-Peter Seidel, Bernt Schiele, and Christian Theobalt. EgoCap: egocentric marker-less motion capture with two fisheye cameras. ACM Transactions on Graphics, 35(6): 162:1–162:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Riviere:2017:PIR

[RRFG17] Jérémy Riviere, Ilya Reshetouski, Luka Filipi, and Abhijeet Ghosh. Polarization imaging reflectometry in the wild. ACM Transactions on Graphics, 36(6):206:1– 206:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Regg:2010:CHH

[RRMG10] Christian Regg, Szymon Rusinkiewicz, Wojciech Matusik, and Markus Gross. Computational highlight holography. ACM Transactions on Graphics, 29(6):170:1– 170:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reinert:2013:IED

[RRS13] Bernhard Reinert, Tobias Ritschel, and Hans-Peter Seidel. Interactive by-example design of artistic packing layouts. ACM Transactions on Graphics, 32(6):218:1– 218:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rosales:2019:SVR

[RRS19] Enrique Rosales, Jafet Rodriguez, and Alla Sheffer. SurfaceBrush: from virtual reality drawings to manifold surfaces. ACM Transactions on Graphics, 38(4): 96:1–96:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raghothama:1998:BRD

[RS98] Srinivas Raghothama and Vadim Shapiro. Boundary representation deformation in parametric solid modeling. ACM Transactions on Graphics, 17(4):259–286, October 1998. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/

journals/tog/1998-17-4/
p259-raghothama/.

Raghuvanshi:2014:PWF

[RS14a] Nikunj Raghuvanshi and John Snyder. Parametric wave field coding for precomputed sound propagation. ACM Transactions on Graphics, 33(4): 38:1–38:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ray:2014:RPT

[RS14b] Nicolas Ray and Dmitry Sokolov. Robust polylines tracing for N-symmetry direction field on triangulated surfaces. ACM Transactions on Graphics, 33(3): 30:1–30:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raghuvanshi:2018:PDC

[RS18] Nikunj Raghuvanshi and John Snyder. Parametric directional coding for precomputed sound propagation. ACM Transactions on Graphics, 37(4):108:1–108:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rubinstein:2008:ISC

[RSA08] Michael Rubinstein, Ariel Shamir, and Shai Avidan. Improved seam carving for video retargeting. ACM Transactions on Graphics, 27(3):16:1– 16:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rubinstein:2009:MOM

[RSA09] Michael Rubinstein, Ariel Shamir, and Shai Avidan. Multi-operator media retargeting. ACM Transactions on Graphics, 28(3):23:1– 23:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2005:LSF

[RSH^{+}05a] Liu Ren, Gregory Shakhnarovich, Jessica K. Hodgins, Hanspeter Pfister, and Paul Viola. Learning silhouette features for control of human motion. ACM Transactions on Graphics, 24(4):1303–1331, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reshetov:2005:MLR

[RSH05b] Alexander Reshetov, Alexei Soupikov, and Jim Hurley. Multi-level ray tracing algorithm. ACM Transactions on Graphics, 24(3): 1176–1185, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ritschel:2008:UMS

[RSI^{+}08] Tobias Ritschel, Kaleigh Smith, Matthias Ihrke, Thorsten Grosch, Karol Myszkowski, and Hans-Peter Seidel. 3D unsharp masking for scene coherent enhancement. ACM Transactions on Graphics, 27 (3):90:1–90:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ray:2016:PFF

[RSL16] Nicolas Ray, Dmitry Sokolov, and Bruno Lévy. Practical 3D frame field generation. ACM Transactions on Graphics, 35 (6):233:1–233:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ray:2018:MVG

[RSLL18] Nicolas Ray, Dmitry Sokolov, Sylvain Lefebvre, and Bruno Lévy. Meshless Voronoi on the GPU. ACM Transactions on Graphics, 37(6): 265:1–265:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raghuvanshi:2010:PWS

[RSM^{+}10a] Nikunj Raghuvanshi, John Snyder, Ravish Mehra, Ming Lin, and Naga Govindaraju. Precomputed wave simulation for real-time sound propagation of dynamic sources in complex scenes. ACM Transactions on Graphics, 29(4): 68:1–68:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Reshetov:2010:CNI

[RSM10b] Alexander Reshetov, Alexei Soupikov, and William R. Mark. Consistent normal interpolation. ACM Transactions on Graphics, 29(6): 142:1–142:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Royo:2023:VMN

[RSM^{+}23] Diego Royo, Talha Sultan, Adolfo Muñoz, Khadijeh Masumnia-Bisheh, Eric Brandt, Diego Gutierrez, Andreas Velten, and Julio Marco. Virtual mirrors: Non-line-of-sight imaging beyond the third bounce. ACM Transactions on Graphics, 42(4): 140:1–140:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592429.

Rinaldi:2023:NDM

[RSP23] Eduardo Rinaldi, Davide Sforza, and Fabio Pellacini. NodeGit: Diffing and merging node graphs. ACM Transactions on Graphics, 42(6): 265:1–265:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618343.

Reinhard:2002:PTR

[RSSF02] Erik Reinhard, Michael Stark, Peter Shirley, and James Ferwerda. Photographic tone reproduction for digital images. ACM Transactions on Graphics, 21(3):267–276, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2024:DTD

[RSSH24] Jing Ren, Aviv Segall, and Olga Sorkine-Hornung. Digital three-dimensional smocking design. ACM Transactions on Graphics, 43(2): 14:1–14:??, April 2024. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3631945.

Reiser:2023:MME

[RSV^{+}23] Christian Reiser, Rick Szeliski, Dor Verbin, Pratul Srinivasan, Ben Mildenhall, Andreas Geiger, Jon Barron, and Peter Hedman. MERF: Memory-efficient radiance fields for real-time view synthesis in unbounded scenes. ACM Transactions on Graphics, 42(4):89:1–89:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592426.

Rushmeier:1990:ERM

[RT90] Holly E. Rushmeier and Kenneth E. Torrance. Extending the radiosity method to include specularly reflecting and translucent materials. ACM Transactions on Graphics, 9(1):1–27, January 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        77636.html.

Romero:2017:EHM

[RTB17] Javier Romero, Dimitrios Tzionas, and Michael J. Black. Embodied hands: modeling and capturing hands and bodies together. ACM Transactions on Graphics, 36 (6):245:1–245:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ritschel:2010:ISS

[RTD^{+}10] Tobias Ritschel, Thorsten Thormählen, Carsten Dachsbacher, Jan Kautz, and Hans-Peter Seidel. Interactive onsurface signal deformation. ACM Transactions on Graphics, 29(4):36:1–36:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

R:2021:PPA

[RTD^{+}21] Mallikarjun B R, Ayush Tewari, Abdallah Dib, Tim Weyrich, Bernd Bickel, Hans-Peter Seidel, Hanspeter Pfister, Wojciech Matusik, Louis Chevallier, Mohamed Elgharib, and Christian Theobalt.

PhotoApp: photorealistic appearance editing of head portraits. ACM Transactions on Graphics, 40(4): 44:1–44:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459765.

Raskar:2004:NPC

[RTF^{+}04] Ramesh Raskar, Kar-Han Tan, Rogerio Feris, Jingyi Yu, and Matthew Turk. Nonphotorealistic camera: depth edge detection and stylized rendering using multi-flash imaging. ACM Transactions on Graphics, 23(3):679– 688, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rhodin:2015:GWG

[RTK^{+}15] Helge Rhodin, James Tompkin, Kwang In Kim, Edilson de Aguiar, Hanspeter Pfister, Hans-Peter Seidel, and Christian Theobalt. Generalizing wave gestures from sparse examples for real-time character control. ACM Transactions on Graphics, 34(6): 181:1–181:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rempel:2007:LFR

[RTS^{+}07] Allan G. Rempel, Matthew Trentacoste, Helge Seetzen, H. David Young, Wolfgang Heidrich, Lorne Whitehead, and Greg Ward. Ldr2Hdr: on-the-fly reverse tone mapping of legacy video and photographs. ACM Transactions on Graphics, 26(3): 39:1–39:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rusinkiewicz:2019:SOF

[Rus19]  Szymon Rusinkiewicz.  A

symmetric objective function for ICP. ACM Transactions on Graphics, 38(4): 85:1–85:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rossignac:1989:AZC

[RV89] Jaroslaw R. Rossignac and Herbert B. Voelcker. Active zones in CSG for accelerating boundary evaluation, redundancy elimination, interference detection, and shading algorithms. ACM Transactions on Graphics, 8(1):51–87, January 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        51123.html.

Rossignac:2011:SAM

[RV11] Jarek Rossignac and Álvar Vinacua. Steady affine motions and morphs. ACM Transactions on Graphics, 30(5):116:1–116:16, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ray:2009:GAD

[RVAL09] Nicolas Ray, Bruno Vallet, Laurent Alonso, and Bruno Levy. Geometry-aware direction field processing. ACM Transactions on Graphics, 29 (1):1:1–1:11, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raskar:2003:IGA

[RvBB^{+}03] Ramesh Raskar, Jeroen van Baar, Paul Beardsley, Thomas Willwacher, Srinivas Rao, and Clifton Forlines. iLamps: geometrically aware and selfconfiguring projectors. ACM Transactions on Graphics, 22 (3):809–818, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rappoport:1993:UID

[RvE93] Ari Rappoport and Maarten van Emmerik. User-interface devices for rapid and exact number specification. ACM Transactions on Graphics, 12(4):348–354, October 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
214380.html.

Ray:2008:SDF

[RVLL08] Nicolas Ray, Bruno Vallet, Wan Chiu Li, and Bruno Lévy. N-symmetry direction field design. ACM Transactions on Graphics, 27(2): 10:1–10:13, April 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2013:GIR

[RWG^{+}13] Peiran Ren, Jiaping Wang, Minmin Gong, Stephen Lin, Xin Tong, and Baining Guo. Global illumination with radiance regression functions. ACM Transactions on Graphics, 32(4):130:1–130:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ruckert:2022:NNA

[RWL^{+}22] Darius Rückert, Yuanhao Wang, Rui Li, Ramzi Idoughi, and Wolfgang Heidrich. NeAT: neural adaptive tomography. ACM Transactions on Graphics, 41(4):55:1–55:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530121.

Ren:2006:RTS

[RWS^{+}06] Zhong Ren, Rui Wang, John Snyder, Kun Zhou, Xinguo Liu, Bo Sun, Peter-Pike Sloan, Hujun Bao, Qunsheng Peng, and Baining Guo. Realtime soft shadows in dynamic scenes using spherical harmonic exponentiation. ACM Transactions on Graphics, 25 (3):977–986, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2011:PR

[RWS^{+}11] Peiran Ren, Jiaping Wang, John Snyder, Xin Tong, and Baining Guo. Pocket reflectometry. ACM Transactions on Graphics, 30(4): 45:1–45:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Raveendran:2014:BL

[RWTT14] Karthik Raveendran, Chris Wojtan, Nils Thuerey, and Greg Turk. Blending liquids. ACM Transactions on Graphics, 33(4):137:1–137:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Rokne:1990:FLS

[RWW90] J. G. Rokne, Brian Wyvill, and Xiaolin Wu. Fast line scan-conversion. ACM Transactions on Graphics, 9 (4):376–388, October 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
88572.html.

Ren:2021:UPS

[RXL21] Bo Ren, Ben Xu, and Chenfeng Li. Unified particle system for multiple-fluid flow and porous material. ACM Transactions on Graphics, 40(4): 118:1–118:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459764.

Rokne:1992:DSI

[RY92] J. Rokne and Y. Yao. Doublestep incremental linear interpolation. ACM Transactions on Graphics, 11(2): 183–192, April 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        130833.html. See [RY93].

Rokne:1993:C

[RY93] J. Rokne and Y. Yao. Corrigendum. ACM Transactions on Graphics, 12(1): 108, January 1993. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [RY92].

Ren:2013:EGP

[RYL13] Zhimin Ren, Hengchin Yeh, and Ming C. Lin. Exampleguided physically based modal sound synthesis. ACM Transactions on Graphics, 32(1): 1:1–1:16, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2023:VCF

[RYPZ23] Bo Ren, Xiaohan Ye, Zherong Pan, and Taiyuan Zhang. Versatile control of fluid-directed solid objects using multi-task reinforcement learning. ACM Transactions on Graphics, 42 (2):15:1–15:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3554731.

Rao:2022:ISA

[RYW^{+}22] Chaolin Rao, Huangjie Yu, Haochuan Wan, Jindong Zhou, Yueyang Zheng, Minye Wu, Yu Ma, Anpei Chen, Binzhe Yuan, Pingqiang Zhou, Xin Lou, and Jingyi Yu. ICARUS: a specialized architecture for neural radiance fields rendering. ACM Transactions on Graphics, 41(6): 234:1–234:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555505.

Rump:2011:PSC

[RZK11] Martin Rump, Arno Zinke, and Reinhard Klein. Practical spectral characterization of trichromatic cameras. ACM Transactions on Graphics, 30 (6):170:1–170:??, December 2011. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2010:IHR

[RZL^{+}10] Zhong Ren, Kun Zhou, Tengfei Li, Wei Hua, and Baining Guo. Interactive hair rendering under environment lighting. ACM Transactions on Graphics, 29(4): 55:1–55:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ren:2021:IER

[RZW^{+}21] Jing Ren, Biao Zhang, Bojian Wu, Jianqiang Huang, Lubin Fan, Maks Ovsjanikov, and Peter Wonka. Intuitive and efficient roof modeling for reconstruction and synthesis. ACM Transactions on Graphics, 40(6): 249:1–249:17, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480494.

Shi:2021:MHM

[SAA^{+}21] Mingyi Shi, Kfir Aberman, Andreas Aristidou, Taku Komura, Dani Lischinski, Daniel Cohen-Or, and Baoquan Chen. MotioNet: 3D human motion reconstruction from monocular video with skeleton consistency. ACM Transactions on Graphics, 40 (1):1:1–1:15, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3407659.

Setaluri:2014:SSP

[SABS14] Rajsekhar Setaluri, Mridul Aanjaneya, Sean Bauer, and Eftychios Sifakis. SPGrid: a sparse paged grid structure applied to adaptive smoke simulation. ACM Transactions on Graphics, 33(6): 205:1–205:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharf:2004:CBS

[SACO04] Andrei Sharf, Marc Alexa, and Daniel Cohen-Or. Contextbased surface completion. ACM Transactions on Graphics, 23(3):878–887, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharp:2022:DDA

[SACO22] Nicholas Sharp, Souhaib Attaiki, Keenan Crane, and Maks Ovsjanikov. Diffusion-Net: Discretization agnostic learning on surfaces. ACM Transactions on Graphics, 41 (3):27:1–27:16, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3507905.

Sahillioglu:2018:GIS

[Sah18] Yusuf Sahillioglu. A genetic isometric shape correspondence algorithm with adaptive sampling. ACM Transactions on Graphics, 37(5): 175:1–175:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3243593.

Said:1989:GBC

[Sai89] H. B. Said. A generalized ball curve and its recursive algorithm. ACM Transactions on Graphics, 8(4):360–371, October 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        77275.html.

Sellan:2021:SVS

[SAJ21] Silvia Sellán, Noam Aigerman, and Alec Jacobson. Swept volumes via spacetime numerical continuation. ACM Transactions on Graphics, 40 (4):55:1–55:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459780.

Sharf:2008:STS

[SAL^{+}08] Andrei Sharf, Dan A. Alcantara, Thomas Lewiner, Chen Greif, Alla Sheffer, Nina Amenta, and Daniel Cohen-Or. Space-time surface reconstruction using incompressible flow. ACM Transactions on Graphics, 27(5): 110:1–110:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sitthi-amorn:2008:ARB

[SaLY^{+}08] Pitchaya Sitthi-amorn, Jason Lawrence, Lei Yang, Pedro V. Sander, Diego Nehab, and Jiahe Xi. Automated reprojection-based pixel shader optimization. ACM Transactions on Graphics, 27(5):127:1–127:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sitthi-Amorn:2011:GPS

[SAMWL11] Pitchaya Sitthi-Amorn, Nicholas Modly, Westley Weimer, and Jason Lawrence. Genetic programming for shader simplification. ACM Transactions on Graphics, 30(6):152:1– 152:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Seo:2023:SSR

[SAN23] Chang Wook Seo, Amirsaman Ashtari, and Junyong Noh. Semi-supervised reference-based sketch extraction using a contrastive learning framework. ACM Transactions on Graphics, 42 (4):56:1–56:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592392.

Schreiner:2004:ISM

[SAPH04] John Schreiner, Arul Asirvatham, Emil Praun, and Hugues Hoppe. Inter-surface mapping. ACM Transactions on Graphics, 23(3):870– 877, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sarraga:2000:VMM

[Sar00] Ramon F. Sarraga. A variational method to model G^{1} surfaces over triangular meshes of arbitrary topology in R^{3}. ACM Transactions on Graphics, 19(4): 279–301, October 2000. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/2000-19-4/
p279-sarraga/.

Sitthi-Amorn:2015:MMV

[SARW^{+}15] Pitchaya Sitthi-Amorn, Javier E. Ramos, Yuwang Wangy, Joyce Kwan, Justin Lan, Wenshou Wang, and Wojciech Matusik. MultiFab: a machine vision assisted platform for multimaterial 3D printing. ACM Transactions on Graphics, 34(4):129:1–129:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sitharam:2006:SSN

[SAZK06] Meera Sitharam, Adam Arbree, Yong Zhou, and Naganandhini Kohareswaran. Solution space navigation for geometric constraint systems. ACM Transactions on Graphics, 25(2):194–213, April 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Salesin:1993:ATO

[SB93] David Salesin and Ronen Barzel. Adjustable tools: An object-oriented interaction metaphor. ACM Transactions on Graphics, 12 (1):103–107, January 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
214378.html.

Sapidis:1995:DCP

[SB95] Nickolas S. Sapidis and Paul J. Best. Direct construction of polynomial surfaces from dense range images through region growing. ACM Transactions on Graphics, 14(2):171–200, April 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
221672.html.

Schechter:2012:GSA

[SB12] Hagit Schechter and Robert Bridson. Ghost SPH for animating water. ACM Transactions on Graphics, 31(4): 61:1–61:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2022:STO

[SBB^{+}22] Zheng Shi, Yuval Bahat, Seung-Hwan Baek, Qiang Fu, Hadi Amata, Xiao Li, Praneeth Chakravarthula, Wolfgang Heidrich, and Felix Heide. Seeing through obstructions with diffractive cloaking. ACM Transactions on Graphics, 41(4): 37:1–37:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530185.

Sadeghi:2013:PMA

[SBdDJ13] Iman Sadeghi, Oleg Bisker, Joachim de Deken, and Henrik Wann Jensen. A practical microcylinder appearance model for cloth rendering. ACM Transactions on Graphics, 32(2):14:1–14:12, April 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sangkloy:2016:SDL

[SBHH16] Patsorn Sangkloy, Nathan Burnell, Cusuh Ham, and James Hays. The sketchy database: learning to retrieve badly drawn bunnies. ACM Transactions on Graphics, 35 (4):119:1–119:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Steinicke:2011:RPP

[SBK11] Frank Steinicke, Gerd Bruder, and Scott Kuhl. Realistic perspective projections for virtual objects and environments. ACM Transactions on Graphics, 30(5): 112:1–112:10, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2018:DMP

[SBK^{+}18] Liang Shi, Vahid Babaei, Changil Kim, Michael Foshey, Yuanming Hu, Pitchaya Sitthi-Amorn, Szymon Rusinkiewicz, and Wojciech Matusik. Deep multispectral painting reproduction via multi-layer, custom-ink printing. ACM Transactions on Graphics, 37 (6):271:1–271:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shin:2015:VTL

[SBLD15] Hijung Valentina Shin, Floraine Berthouzoz, Wilmot Li, and Frédo Durand. Visual transcripts: lecture notes from blackboard-style lecture videos. ACM Transactions on Graphics, 34(6):240:1– 240:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Soler:2015:EAS

[SBN15] Cyril Soler, Mahdi M. Bagher, and Derek Nowrouzezahrai. Efficient and accurate spherical kernel integrals using isotropic decomposition. ACM Transactions on Graphics, 34(5):161:1–161:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schumacher:2015:MCE

[SBR^{+}15] Christian Schumacher, Bernd Bickel, Jan Rys, Steve Marschner, Chiara Daraio, and Markus Gross. Microstructures to control elasticity in 3D printing. ACM Transactions on Graphics, 34(4):136:1–136:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sanchez-Banderas:2020:REL

[SBRBO20] Rosa M. Sánchez-Banderas, Alejandro Rodríguez, Héctor Barreiro, and Miguel A. Otaduy. Robust Eulerian-on-Lagrangian rods. ACM Transactions on Graphics, 39(4): 59:1–59:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392489.

Schroers:2018:OVP

[SBSH18] Christopher Schroers, Jean-Charles Bazin, and Alexander Sorkine-Hornung. An omnistereoscopic video pipeline for capture and display of real-world VR. ACM Transactions on Graphics, 37(3): 37:1–37:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3225150.

Shao:2012:CSC

[SBSS12] Cloud Shao, Adrien Bousseau, Alla Sheffer, and Karan Singh. CrossShade: shading concept sketches using crosssection curves. ACM Transactions on Graphics, 31(4): 45:1–45:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2019:SIP

[SBT^{+}19] Tiancheng Sun, Jonathan T.

Barron, Yun-Ta Tsai, Zexiang Xu, Xueming Yu, Graham Fyffe, Christoph Rhemann, Jay Busch, Paul Debevec, and Ravi Ramamoorthi. Single image portrait relighting. ACM Transactions on Graphics, 38(4):79:1–79:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2009:CMS

[SBZ09] Xiaohan Shi, Hujun Bao, and Kun Zhou. Out-of-core multigrid solver for streaming meshes. ACM Transactions on Graphics, 28(5): 173:1–173:7, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sawhney:2018:BFF

[SC18a] Rohan Sawhney and Keenan Crane. Boundary first flattening. ACM Transactions on Graphics, 37(1):5:1– 5:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharp:2018:VSC

[SC18b] Nicholas Sharp and Keenan Crane. Variational surface cutting. ACM Transactions on Graphics, 37(4):156:1– 156:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharp:2020:YCF

[SC20] Nicholas Sharp and Keenan Crane. You can find geodesic paths in triangle meshes by just flipping edges. ACM Transactions on Graphics, 39 (6):249:1–249:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417839.

Soler:2002:HPM

[SCA02] Cyril Soler, Marie-Paule Cani, and Alexis Angelidis. Hierarchical pattern mapping. ACM Transactions on Graphics, 21 (3):673–680, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schwarz:1987:ECR

[SCB87] Michael W. Schwarz, William B.

Cowan, and John C. Beatty. An experimental comparison of RGB, YIQ, LAB, HSV, and opponent color models. ACM Transactions on Graphics, 6(2):123–158, April 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
31338.html.

Stone:1988:CGM

[SCB88] Maureen C. Stone, William B.

Cowan, and John C. Beatty. Color gamut mapping and the printing of digital color images. ACM Transactions on Graphics, 7(4):249–292, October 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
48045.html.

Sayed:2022:LIC

[SCCB22] Mohamed Sayed, Robert Cinca, Enrico Costanza, and Gabriel Brostow. Look-Out! interactive camera gimbal controller for filming long takes. ACM Transactions on Graphics, 41(3): 30:1–30:16, June 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3506693.

Soliman:2021:CWS

[SCD^{+}21] Yousuf Soliman, Albert Chern, Olga Diamanti, Felix Knöppel, Ulrich Pinkall, and Peter Schröder. Constrained Willmore surfaces. ACM Transactions on Graphics, 40(4): 112:1–112:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459759.

Sederberg:2004:SSL

[SCF^{+}04] Thomas W. Sederberg, David L. Cardon, G. Thomas Finnigan, Nicholas S. North, Jianmin Zheng, and Tom Lyche. Tspline simplification and local refinement. ACM Transactions on Graphics, 23(3): 276–283, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sen:2005:DP

[SCG^{+}05] Pradeep Sen, Billy Chen, Gaurav Garg, Stephen R. Marschner, Mark Horowitz, Marc Levoy, and Hendrik P. A. Lensch. Dual photography. ACM Transactions on Graphics, 24(3):745–755, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Skouras:2015:ISD

[SCGT15] Mélina Skouras, Stelian Coros, Eitan Grinspun, and Bernhard Thomaszewski. Interactive surface design with interlocking elements. ACM Transactions on Graphics, 34(6): 224:1–224:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sen:2003:SSM

[SCH03] Pradeep Sen, Mike Cammarano, and Pat Hanrahan. Shadow silhouette maps. ACM Transactions on Graphics, 22(3):521–526, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Savva:2014:SIA

[SCH^{+}14] Manolis Savva, Angel X.

Chang, Pat Hanrahan, Matthew Fisher, and Matthias Nießner. SceneGrok: inferring action maps in 3D environments. ACM Transactions on Graphics, 33(6):212:1– 212:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Savva:2016:PLI

[SCH^{+}16] Manolis Savva, Angel X. Chang, Pat Hanrahan, Matthew Fisher, and Matthias Nießner. PiGraphs: learning interaction snapshots from observations. ACM Transactions on Graphics, 35(4): 139:1–139:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sugimoto:2023:PWB

[SCJ^{+}23] Ryusuke Sugimoto, Terry Chen, Yiti Jiang, Christopher Batty, and Toshiya Hachisuka. A practical walk-on-boundary method for boundary value problems. ACM Transactions on Graphics, 42(4): 81:1–81:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592109.

Sendik:2017:DCTa

[SCO17a] Omry Sendik and Daniel Cohen-Or. Deep correlations for texture synthesis. ACM Transactions on Graphics, 36 (4):105:1–105:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sendik:2017:DCTb

[SCO17b] Omry Sendik and Daniel Cohen-Or. Deep correlations for texture synthesis.

ACM Transactions on Graphics, 36(5):161:1–161:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Seiler:2008:LMC

[SCS^{+}08] Larry Seiler, Doug Carmean, Eric Sprangle, Tom Forsyth, Michael Abrash, Pradeep Dubey, Stephen Junkins, Adam Lake, Jeremy Sugerman, Robert Cavin, Roger Espasa, Ed Grochowski, Toni Juan, and Pat Hanrahan. Larrabee: a many-core x86 architecture for visual computing. ACM Transactions on Graphics, 27(3):18:1– 18:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Siegl:2015:RTP

[SCT^{+}15] Christian Siegl, Matteo Colaianni, Lucas Thies, Justus Thies, Michael Zollhöfer, Shahram Izadi, Marc Stamminger, and Frank Bauer. Real-time pixel luminance optimization for dynamic multiprojection mapping. ACM Transactions on Graphics, 34 (6):237:1–237:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Serrano:2021:ESI

[SCW^{+}21] Ana Serrano, Bin Chen, Chao Wang, Michal Piovarci, Hans-Peter Seidel, Piotr Didyk, and Karol Myszkowski. The effect of shape and illumination on material perception: model and applications. ACM Transactions on Graphics, 40(4): 125:1–125:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459813.

Stone:1989:GCP

[SD89] Maureen C. Stone and Tony D. DeRose. A geometric characterization of parametric cubic curves. ACM Transactions on Graphics, 8(3):147–163, July 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77056.html.

Stamminger:2002:PSM

[SD02] Marc Stamminger and George Drettakis. Perspective shadow maps. ACM Transactions on Graphics, 21(3):557–562, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sen:2012:FNR

[SD12] Pradeep Sen and Soheil Darabi. On filtering the noise from the random parameters in Monte Carlo rendering. ACM Transactions on Graphics, 31(3):18:1–18:15, May 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2023:EIV

[SDA^{+}23] Jiatian Sun, Longxiulin Deng, Triantafyllos Afouras, Andrew Owens, and Abe Davis. Eventfulness for interactive video alignment. ACM Transactions on Graphics, 42 (4):46:1–46:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592118.

Schneider:2019:PSF

[SDG^{+}19] Teseo Schneider, Jérémie Dumas, Xifeng Gao, Mario Botsch, Daniele Panozzo, and Denis Zorin. Poly-spline finite-element method. ACM Transactions on Graphics, 38 (3):19:1–19:??, June 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3313797.

Solomon:2015:CWD

[SdGP^{+}15] Justin Solomon, Fernando de Goes, Gabriel Peyré, Marco Cuturi, Adrian Butscher, Andy Nguyen, Tao Du, and Leonidas Guibas. Convolutional Wasserstein distances: efficient optimal transportation on geometric domains. ACM Transactions on Graphics, 34(4):66:1–66:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sakurai:2018:FRD

[SDIN18] Kaisei Sakurai, Yoshinori Dobashi, Kei Iwasaki, and Tomoyuki Nishita. Fabricating reflectors for displaying multiple images. ACM Transactions on Graphics, 37(4): 158:1–158:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2018:SNH

[SDK18] Breannan Smith, Fernando De Goes, and Theodore Kim. Stable neo-Hookean flesh simulation. ACM Transactions on Graphics, 37(2): 12:1–12:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2019:AEI

[SDK19] Breannan Smith, Fernando De Goes, and Theodore Kim. Analytic eigensystems for isotropic distortion energies. ACM Transactions on Graphics, 38(1):3:1– 3:??, February 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3241041.

Sato:2021:SGS

[SDK21] Syuhei Sato, Yoshinori Dobashi, and Theodore Kim. Streamguided smoke simulations. ACM Transactions on Graphics, 40(4):161:1–161:7, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459846.

Sato:2018:EBT

[SDKN18] Syuhei Sato, Yoshinori Dobashi, Theodore Kim, and Tomoyuki Nishita. Example-based turbulence style transfer. ACM Transactions on Graphics, 37 (4):84:1–84:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sato:2018:EFA

[SDN18] Syuhei Sato, Yoshinori Dobashi, and Tomoyuki Nishita. Editing fluid animation using flow interpolation. ACM Transactions on Graphics, 37(5): 173:1–173:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3213771.

Stone:2004:SHC

[SDO^{+}04] Matthew Stone, Doug De-Carlo, Insuk Oh, Christian Rodriguez, Adrian Stere, Alyssa Lees, and Chris Bregler. Speaking with hands: creating animated conversational characters from recordings of human performance. ACM Transactions on Graphics, 23(3):506–513, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sitzmann:2018:EEO

[SDP^{+}18] Vincent Sitzmann, Steven Diamond, Yifan Peng, Xiong Dun, Stephen Boyd, Wolfgang Heidrich, Felix Heide, and Gordon Wetzstein. Endto-end optimization of optics and image processing for achromatic extended depth of field and super-resolution imaging. ACM Transactions on Graphics, 37(4):114:1– 114:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sheffer:2002:SOG

[SdS02] Alla Sheffer and Eric de Sturler. Smoothing an overlay grid to minimize linear distortion in texture mapping. ACM Transactions on Graphics, 21 (4):874–890, October 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2016:CCF

[SDW^{+}16] Peng Song, Bailin Deng, Ziqi Wang, Zhichao Dong, Wei Li, Chi-Wing Fu, and Ligang Liu. CofiFab: coarseto-fine fabrication of large 3D objects. ACM Transactions on Graphics, 35(4): 45:1–45:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Selim:2016:PST

[SED16] Ahmed Selim, Mohamed Elgharib, and Linda Doyle. Painting style transfer for head portraits using convolutional neural networks. ACM Transactions on Graphics, 35 (4):129:1–129:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Seidel:1993:PFG

[Sei93]  Hans-Peter Seidel.  Polar

forms for geometrically continuous spline curves of arbitrary degree. ACM Transactions on Graphics, 12(1): 1–34, January 1993. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
169726.html.

Shilane:2007:DRS

[SF07] Philip Shilane and Thomas Funkhouser. Distinctive regions of 3D surfaces. ACM Transactions on Graphics, 26 (2):7:1–7:??, June 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schollmeyer:2009:DTN

[SF09] Andre Schollmeyer and Bernd Fröhlich. Direct trimming of NURBS surfaces on the GPU. ACM Transactions on Graphics, 28(3):47:1–47:??, August 2009. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Stokes:1992:PRD

[SFB92] Mike Stokes, Mark D. Fairchild, and Roy S. Berns. Precision requirements for digital color reproduction. ACM Transactions on Graphics, 11 (4):406–422, October 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
146482.html.

Sugerman:2009:GPM

[SFB^{+}09] Jeremy Sugerman, Kayvon Fatahalian, Solomon Boulos, Kurt Akeley, and Pat Hanrahan. GRAMPS: a programming model for graphics pipelines. ACM Transactions on Graphics, 28(1): 4:1–4:11, January 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Su:2023:SAA

[SFC^{+}23] Zejia Su, Qingnan Fan, Xuelin Chen, Oliver Van Kaick, Hui Huang, and Ruizhen Hu. Scene-aware activity program generation with language guidance. ACM Transactions on Graphics, 42(6): 252:1–252:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618338.

Shen:2012:SRP

[SFCH12] Chao-Hui Shen, Hongbo Fu, Kang Chen, and Shi-Min Hu. Structure recovery by part assembly. ACM Transactions on Graphics, 31(6): 180:1–180:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2012:RIP

[SFCO12] Peng Song, Chi-Wing Fu, and Daniel Cohen-Or. Recursive interlocking puzzles. ACM Transactions on Graphics, 31 (6):128:1–128:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2022:GDM

[SFD^{+}22] Yuefan Shen, Hongbo Fu, Zhongshuo Du, Xiang Chen, Evgeny Burnaev, Denis Zorin, Kun Zhou, and Youyi Zheng. GCN-Denoiser: Mesh denoising with graph convolutional networks. ACM Transactions on Graphics, 41 (1):8:1–8:14, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3480168.

Song:2013:RFS

[SFG^{+}13] Peng Song, Chi-Wing Fu, Prashant Goswami, Jianmin Zheng, Niloy J. Mitra, and Daniel Cohen-Or. Reciprocal frame structures made easy. ACM Transactions on Graphics, 32(4):94:1–94:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2017:RIF

[SFJ^{+}17] Peng Song, Chi-Wing Fu, Yueming Jin, Hongfei Xu, Ligang Liu, Pheng-Ann Heng, and Daniel Cohen-Or. Reconfigurable interlocking furniture. ACM Transactions on Graphics, 36(6):174:1– 174:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sederberg:2008:WTN

[SFL^{+}08] Thomas W. Sederberg, G. Thomas Finnigan, Xin Li, Hongwei Lin, and Heather Ipson. Watertight trimmed NURBS. ACM Transactions on Graphics, 27(3):79:1–79:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Saund:2004:PSI

[SFLM04] Eric Saund, David Fleet, Daniel Larner, and James Mahoney. Perceptuallysupported image editing of text and graphics. ACM Transactions on Graphics, 23 (3):728, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stokes:2004:PIC

[SFWG04] William A. Stokes, James A.

Ferwerda, Bruce Walter, and Donald P. Greenberg. Perceptual illumination components: a new approach to efficient, high quality global illumination rendering. ACM Transactions on Graphics, 23(3): 742–749, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sechrest:1982:VPR

[SG82] S. Sechrest and D. P. Greenberg. A visible polygon reconstruction algorithm. ACM Transactions on Graphics, 1 (1):25–42, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Scheifler:1986:XWS

[SG86] Robert W. Scheifler and Jim Gettys. The X window system. ACM Transactions on Graphics, 5(2):79–109, April 1986. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
24053.html.

Singh:1991:ALS

[SG91] Gurminder Singh and Mark Green. Automating the lexical and syntactic design of graphical user interfaces: The UofA* UIMS. ACM Transactions on Graphics, 10(3):213–254, July 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        108543.html.

Surazhsky:2001:CMC

[SG01] Vitaly Surazhsky and Craig Gotsman. Controllable morphing of compatible planar triangulations. ACM Transactions on Graphics, 20(4): 203–231, October 2001. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Solenthaler:2011:TSP

[SG11] Barbara Solenthaler and Markus Gross. Two-scale particle simulation. ACM Transactions on Graphics, 30(4): 81:1–81:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2017:IRT

[SG17] Li-Yong Shen and Ron Goldman. Implicitizing rational tensor product surfaces using the resultant of three moving planes. ACM Transactions on Graphics, 36(5): 167:1–167:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stein:2018:DTM

[SGC18] Oded Stein, Eitan Grinspun, and Keenan Crane. Developability of triangle meshes. ACM Transactions on Graphics, 37(4):77:1–77:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sonlu:2021:CAF

[SGD21] Sinan Sonlu, Ugur Güdükbay, and Funda Durupinar. A conversational agent framework with multi-modal personality expression. ACM Transactions on Graphics, 40 (1):7:1–7:16, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3439795.

Stoll:2010:VBR

[SGdA^{+}10] Carsten Stoll, Juergen Gall, Edilson de Aguiar, Sebastian Thrun, and Christian Theobalt. Video-based reconstruction of animatable human characters. ACM Transactions on Graphics, 29(6): 139:1–139:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sud:2006:FPC

[SGG^{+}06] Avneesh Sud, Naga Govindaraju, Russell Gayle, Ilknur Kabul, and Dinesh Manocha. Fast proximity computation among deformable models using discrete Voronoi diagrams. ACM Transactions on Graphics, 25(3):1144–1153, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Salaun:2022:RBM

[SGH^{+}22] Corentin Salaün, Adrien Gruson, Binh-Son Hua, Toshiya Hachisuka, and Gurprit Singh. Regression-based Monte Carlo integration. ACM Transactions on Graphics, 41(4): 79:1–79:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530095.

Sajadi:2012:EGR

[SGM12] Behzad Sajadi, M. Gopi, and Aditi Majumder. Edgeguided resolution enhancement in projectors via optical pixel sharing. ACM Transactions on Graphics, 31(4): 79:1–79:12, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Serrano:2016:ICS

[SGM^{+}16] Ana Serrano, Diego Gutierrez, Karol Myszkowski, Hans-Peter Seidel, and Belen Masia. An intuitive control space for material appearance. ACM Transactions on Graphics, 35 (6):186:1–186:??, November 2016. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Shimada:2023:DMD

[SGPT23] Soshi Shimada, Vladislav Golyanik, Patrick Pérez, and Christian Theobalt. Decaf: Monocular deformation capture for face and hand interactions. ACM Transactions on Graphics, 42(6): 262:1–262:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618329.

Snavely:2008:FPT

[SGSS08] Noah Snavely, Rahul Garg, Steven M. Seitz, and Richard Szeliski. Finding paths through the world’s photos. ACM Transactions on Graphics, 27(3):15:1–15:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Salaun:2022:SMC

[SGSS22] Corentin Salaün, Iliyan Georgiev, Hans-Peter Seidel, and Gurprit Singh. Scalable multiclass sampling via filtered sliced optimal transport. ACM Transactions on Graphics, 41(6):261:1–261:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555484.

Schmidt:2006:IDC

[SGW06] Ryan Schmidt, Cindy Grimm, and Brian Wyvill. Interactive decal compositing with discrete exponential maps. ACM Transactions on Graphics, 25 (3):605–613, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stein:2018:NBC

[SGWJ18] Oded Stein, Eitan Grinspun, Max Wardetzky, and Alec Jacobson. Natural boundary conditions for smoothing in geometry processing. ACM Transactions on Graphics, 37 (2):23:1–23:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shimada:2021:NMH

[SGX^{+}21] Soshi Shimada, Vladislav Golyanik, Weipeng Xu, Patrick Pérez, and Christian Theobalt. Neural monocular 3D human motion capture with physical awareness. ACM Transactions on Graphics, 40 (4):83:1–83:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459825.

Shimada:2020:PPP

[SGXT20] Soshi Shimada, Vladislav Golyanik, Weipeng Xu, and Christian Theobalt. PhysCap: physically plausible monocular 3D motion capture in real time. ACM Transactions on Graphics, 39(6): 235:1–235:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417877.

Safonova:2007:COS

[SH07] Alla Safonova and Jessica K.

Hodgins. Construction and optimal search of interpolated motion graphs. ACM Transactions on Graphics, 26(3): 106:1–106:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shiratori:2008:ABU

[SH08] Takaaki Shiratori and Jessica K. Hodgins. Accelerometerbased user interfaces for the control of a physically simulated character. ACM Transactions on Graphics, 27(5): 123:1–123:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sahillioglu:2023:APR

[SH23] Yusuf Sahillioglu and Devin Horsman. Augmented paths and reodesics for topologicallystable matching. ACM Transactions on Graphics, 42 (2):17:1–17:??, April 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3554978.

Shamir:2003:CBA

[Sha03] Ariel Shamir. Constraintbased approach for automatic hinting of digital typefaces. ACM Transactions on Graphics, 22(2):131–151, April 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2014:LFR

[SHD^{+}14] Lixin Shi, Haitham Hassanieh, Abe Davis, Dina Katabi, and Fredo Durand. Light field reconstruction using sparsity in the continuous Fourier domain. ACM Transactions on Graphics, 34(1): 12:1–12:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schneider:2018:DSA

[SHD^{+}18] Teseo Schneider, Yixin Hu, Jérémie Dumas, Xifeng Gao, Daniele Panozzo, and Denis Zorin. Decoupling simulation accuracy from mesh quality. ACM Transactions on Graphics, 37(6):280:1– 280:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sheffer:2013:ECH

[She13] Alla Sheffer. An efficient computation of handle and tunnel loops via Reeb graphs. ACM Transactions on Graphics, 32 (4):32:1–32:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2011:APU

[SHFH11] Chao-Hui Shen, Shi-Sheng Huang, Hongbo Fu, and Shi-Min Hu. Adaptive partitioning of urban facades. ACM Transactions on Graphics, 30 (6):184:1–184:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schneider:2022:LSC

[SHG^{+}22] Teseo Schneider, Yixin Hu, Xifeng Gao, Jérémie Dumas, Denis Zorin, and Daniele Panozzo. A large-scale comparison of tetrahedral and hexahedral elements for solving elliptic PDEs with the finite element method. ACM Transactions on Graphics, 41 (3):23:1–23:14, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3508372.

Suri:1999:ABB

[SHH99] Subhash Suri, Philip M. Hubbard, and John F. Hughes. Analyzing bounding boxes for object intersection. ACM Transactions on Graphics, 18 (3):257–277, July 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1999-18-3/
                                        p257-suri/.

Schussler:2017:MBN

[SHHD17] Vincent Schüssler, Eric Heitz, Johannes Hanika, and Carsten Dachsbacher. Microfacetbased normal mapping for robust Monte Carlo path tracing. ACM Transactions on Graphics, 36(6):205:1– 205:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sloan:2003:CPC

[SHHS03] Peter-Pike Sloan, Jesse Hall, John Hart, and John Snyder. Clustered principal components for precomputed radiance transfer. ACM Transactions on Graphics, 22(3): 382–391, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shrestha:2016:CIM

[SHHW16] Shikhar Shrestha, Felix Heide, Wolfgang Heidrich, and Gordon Wetzstein. Computational imaging with multicamera time-of-flight systems. ACM Transactions on Graphics, 35(4):33:1–33:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stanton:2014:SRG

[SHK^{+}14] Matt Stanton, Ben Humberston, Brandon Kase, James F. O’Brien, Kayvon Fatahalian, and Adrien Treuille. Selfrefining games using player analytics. ACM Transactions on Graphics, 33(4): 73:1–73:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2017:PGF

[SHK^{+}17] Qi Sun, Fu-Chung Huang, Joohwan Kim, Li-Yi Wei, David Luebke, and Arie Kaufman. Perceptually-guided foveation for light field displays. ACM Transactions on Graphics, 36(6):192:1– 192:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2017:NEL

[SHL^{+}17] Liang Shi, Fu-Chung Huang, Ward Lopes, Wojciech Matusik, and David Luebke. Near-eye light field holographic rendering with spherical waves for wide field of view interactive 3D computer graphics. ACM Transactions on Graphics, 36(6): 236:1–236:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Su:2014:EID

[SHM^{+}14] Hao Su, Qixing Huang, Niloy J. Mitra, Yangyan Li, and Leonidas Guibas. Estimating image depth using shape collections. ACM Transactions on Graphics, 33 (4):37:1–37:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Saito:2018:HSU

[SHM^{+}18] Shunsuke Saito, Liwen Hu, Chongyang Ma, Hikaru Ibayashi, Linjie Luo, and Hao Li. 3D hair synthesis using volumetric variational autoencoders. ACM Transactions on Graphics, 37(6):208:1– 208:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shao:2022:FUA

[SHM22] Han Shao, Libo Huang, and Dominik L. Michels. A fast unsmoothed aggregation algebraic multigrid framework for the large-scale simulation of incompressible flow. ACM Transactions on Graphics, 41(4):49:1–49:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530109.

Shneiderman:1992:TVT

[Shn92] Ben Shneiderman. Tree visualization with tree-maps: a 2-D space-filling approach. ACM Transactions on Graphics, 11(1):92–99, January 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
115768.html.

Smith:2002:CMT

[SHOW02] Jeffrey Smith, Jessica Hodgins, Irving Oppenheim, and Andrew Witkin. Creating models of truss structures with optimization. ACM Transactions on Graphics, 21 (3):295–301, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Safonova:2004:SPR

[SHP04] Alla Safonova, Jessica K.

Hodgins, and Nancy S. Pollard. Synthesizing physically realistic human motion in low-dimensional, behaviorspecific spaces. ACM Transactions on Graphics, 23(3): 514–521, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Seetzen:2004:HDR

[SHS^{+}04] Helge Seetzen, Wolfgang Heidrich, Wolfgang Stuerzlinger, Greg Ward, Lorne Whitehead, Matthew Trentacoste, Abhijeet Ghosh, and Andrejs Vorozcovs. High dynamic range display systems.

ACM Transactions on Graphics, 23(3):760–768, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shu:2017:PLT

[SHS^{+}17] Zhixin Shu, Sunil Hadap, Eli Shechtman, Kalyan Sunkavalli, Sylvain Paris, and Dimitris Samaras. Portrait lighting transfer using a mass transport approach. ACM Transactions on Graphics, 36(4): 145:1–145:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shu:2018:PLT

[SHS^{+}18] Zhixin Shu, Sunil Hadap, Eli Shechtman, Kalyan Sunkavalli, Sylvain Paris, and Dimitris Samaras. Portrait lighting transfer using a mass transport approach. ACM Transactions on Graphics, 37(1): 2:1–2:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Satoi:2016:UMP

[SHU^{+}16] Daiki Satoi, Mikihiro Hagiwara, Akira Uemoto, Hisanao Nakadai, and Junichi Hoshino. Unified motion planner for fishes with various swimming styles. ACM Transactions on Graphics, 35(4): 80:1–80:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schreck:2019:FSW

[SHW19] Camille Schreck, Christian Hafner, and Chris Wojtan. Fundamental solutions for water wave animation. ACM Transactions on Graphics, 38 (4):130:1–130:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schiftner:2009:PCS

[SHWP09] Alexander Schiftner, Mathias Höbinger, Johannes Wallner, and Helmut Pottmann. Packing circles and spheres on surfaces. ACM Transactions on Graphics, 28(5): 139:1–139:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

She:2022:LHD

[SHX^{+}22] Qijin She, Ruizhen Hu, Juzhan Xu, Min Liu, Kai Xu, and Hui Huang. Learning high-DOF reaching-and-grasping via dynamic representation of gripper-object interaction. ACM Transactions on Graphics, 41(4): 97:1–97:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530091.

Shi:2020:SLP

[SHZ^{+}20] Yifei Shi, Junwen Huang, Hongjia Zhang, Xin Xu, Szymon Rusinkiewicz, and Kai Xu. SymmetryNet: learning to predict reflectional and rotational symmetries of 3D shapes from single-view RGB-D images. ACM Transactions on Graphics, 39(6): 213:1–213:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417775.

Seo:2011:CDM

[SILN11] Jaewoo Seo, Geoffrey Irving, J. P. Lewis, and Junyong Noh. Compression and direct manipulation of complex blendshape models. ACM Transactions on Graphics, 30(6): 164:1–164:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shene:1994:LDI

[SJ94] Ching-Kuang Shene and John K. Johnstone. On the lower degree intersections of two natural quadrics. ACM Transactions on Graphics, 13(4):400–424, October 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
197316.html.

Spencer:2013:PPR

[SJ13] Ben Spencer and Mark W.

Jones. Progressive photon relaxation. ACM Transactions on Graphics, 32(1):7:1– 7:11, January 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Singh:2017:CAA

[SJ17] Gurprit Singh and Wojciech Jarosz. Convergence analysis for anisotropic Monte Carlo sampling spectra. ACM Transactions on Graphics, 36 (4):137:1–137:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sochorova:2021:PPM

[SJ21] Sárka Sochorová and Ondrej Jamriska. Practical pigment mixing for digital painting. ACM Transactions on Graphics, 40(6): 234:1–234:11, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480549.

Sellan:2022:SPS

[SJ22a] Silvia Sellán and Alec Jacobson. Stochastic Poisson surface reconstruction. ACM Transactions on Graphics, 41 (6):227:1–227:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555441.

Sharp:2022:SDG

[SJ22b] Nicholas Sharp and Alec Jacobson. Spelunking the deep: guaranteed queries on general neural implicit surfaces via range analysis. ACM Transactions on Graphics, 41 (4):107:1–107:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530155.

Shan:2008:HQM

[SJA08] Qi Shan, Jiaya Jia, and Aseem Agarwala. High-quality motion deblurring from a single image. ACM Transactions on Graphics, 27(3):73:1– 73:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sung:2020:DDT

[SJA^{+}20] Minhyuk Sung, Zhenyu Jiang, Panos Achlioptas, Niloy J. Mitra, and Leonidas J. Guibas. DeformSyncNet: Deformation transfer via synchronized shape deformation spaces. ACM Transactions on Graphics, 39(6): 261:1–261:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417783.

Schwarzhaupt:2012:PHB

[SJJ12] Jorge Schwarzhaupt, Henrik Wann Jensen, and Wojciech Jarosz. Practical Hessian-based error control for irradiance caching. ACM Transactions on Graphics, 31 (6):193:1–193:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sueda:2011:LSD

[SJLP11] Shinjiro Sueda, Garrett L.

Jones, David I. W. Levin, and Dinesh K. Pai. Large-scale dynamic simulation of highly constrained strands. ACM Transactions on Graphics, 30 (4):39:1–39:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schweickart:2017:AER

[SJM17] Eston Schweickart, Doug L.

James, and Steve Marschner. Animating elastic rods with sound. ACM Transactions on Graphics, 36(4):115:1– 115:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sunkavalli:2010:MSI

[SJMP10] Kalyan Sunkavalli, Micah K.

Johnson, Wojciech Matusik, and Hanspeter Pfister. Multiscale image harmonization. ACM Transactions on Graphics, 29(4):125:1–125:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shiue:2005:RGS

[SJP05] Le-Jeng Shiue, Ian Jones, and Jörg Peters. A realtime GPU subdivision kernel. ACM Transactions on Graphics, 24 (3):1010–1015, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2018:CMB

[SJR18] Tiancheng Sun, Henrik Wann Jensen, and Ravi Ramamoorthi. Connecting measured BRDFs to analytic BRDFs by data-driven diffuse-specular separation. ACM Transactions on Graphics, 37(6): 273:1–273:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2004:PM

[SJTS04] Jian Sun, Jiaya Jia, Chi-Keung Tang, and Heung-Yeung Shum. Poisson matting. ACM Transactions on Graphics, 23(3):315–321, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stein:2020:SEB

[SJWG20] Oded Stein, Alec Jacobson, Max Wardetzky, and Eitan Grinspun. A smoothness energy without boundary distortion for curved surfaces. ACM Transactions on Graphics, 39 (3):18:1–18:17, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3377406.

Shen:2019:PE

[SJZP19] Hanxiao Shen, Zhongshi Jiang, Denis Zorin, and Daniele Panozzo. Progressive embedding. ACM Transactions on Graphics, 38(4): 32:1–32:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Subr:2013:FAS

[SK13] Kartic Subr and Jan Kautz.

Fourier analysis of stochastic sampling strategies for assessing bias and variance in integration. ACM Transactions on Graphics, 32(4): 128:1–128:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sahillioglu:2016:DPM

[SK16] Yusuf Sahillioglu and Ladislav Kavan. Detail-preserving mesh unfolding for nonrigid shape retrieval. ACM Transactions on Graphics, 35(3): 27:1–27:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sung:2015:DDS

[SKAG15] Minhyuk Sung, Vladimir G. Kim, Roland Angst, and Leonidas Guibas. Data-driven structural priors for shape completion. ACM Transactions on Graphics, 34(6): 175:1–175:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Steinberger:2014:WTB

[SKB^{+}14] Markus Steinberger, Michael Kenzel, Pedro Boechat, Bernhard Kerbl, Mark Dokter, and Dieter Schmalstieg. Whippletree: task-based scheduling of dynamic workloads on the GPU. ACM Transactions on Graphics, 33(6): 228:1–228:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schulz:2021:MCI

[SKB^{+}21] Christoph Schulz, Kin Chung Kwan, Michael Becher, Daniel Baumgartner, Guido Reina, Oliver Deussen, and Daniel Weiskopf. Multi-class inverted stippling. ACM Transactions on Graphics, 40(6): 245:1–245:12, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480534.

Sykora:2014:IRB

[SKC^{+}14] Daniel Sýkora, Ladislav Kavan, Martin Cadík, Ondrej Jamriska, Alec Jacobson, Brian Whited, Maryann Simmons, and Olga Sorkine-Hornung. Ink-and-ray: Basrelief meshes for adding global illumination effects to handdrawn characters. ACM Transactions on Graphics, 33 (2):16:1–16:??, March 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sinha:2012:IBR

[SKG^{+}12] Sudipta N. Sinha, Johannes Kopf, Michael Goesele, Daniel Scharstein, and Richard Szeliski. Image-based rendering for scenes with reflections. ACM Transactions on Graphics, 31(4):100:1–100:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Steinberger:2012:SDS

[SKK^{+}12] Markus Steinberger, Bernhard Kainz, Bernhard Kerbl, Stefan Hauswiesner, Michael Kenzel, and Dieter Schmalstieg. Softshell: dynamic scheduling on GPUs. ACM Transactions on Graphics, 31 (6):161:1–161:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sok:2007:SBB

[SKL07] Kwang Won Sok, Manmyung Kim, and Jehee Lee. Simulating biped behaviors from human motion data. ACM Transactions on Graphics, 26 (3):107:1–107:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Soderstrom:2010:PBN

[SKM10] Andreas Söderström, Matts Karlsson, and Ken Museth. A PML-based nonreflective boundary for free surface fluid animation. ACM Transactions on Graphics, 29(5): 136:1–136:17, October 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sintorn:2014:CPV

[SKOA14] Erik Sintorn, Viktor Kämpe, Ola Olsson, and Ulf Assarsson. Compact precomputed voxelized shadows. ACM Transactions on Graphics, 33 (4):150:1–150:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sueda:2008:MSH

[SKP08] Shinjiro Sueda, Andrew Kaufman, and Dinesh K. Pai. Musculotendon simulation for hand animation. ACM Transactions on Graphics, 27(3): 83:1–83:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sloan:2002:PRT

[SKS02] Peter-Pike Sloan, Jan Kautz, and John Snyder. Precomputed radiance transfer for real-time rendering in dynamic, low-frequency lighting environments. ACM Transactions on Graphics, 21(3): 527–536, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sellan:2020:OCS

[SKSJ20] Silvia Sellán, Jacob Kesten, Ang Yan Sheng, and Alec Jacobson. Opening and closing surfaces. ACM Transactions on Graphics, 39(6): 198:1–198:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417778.

Schmidt:2009:ADS

[SKSK09] Ryan Schmidt, Azam Khan, Karan Singh, and Gord Kurtenbach. Analytic drawing of 3D scaffolds. ACM Transactions on Graphics, 28 (5):149:1–149:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shum:2008:IPM

[SKSY08] Hubert P. H. Shum, Taku Komura, Masashi Shiraishi, and Shuntaro Yamazaki. Interaction patches for multicharacter animation. ACM Transactions on Graphics, 27 (5):114:1–114:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2012:RSI

[SKV^{+}12] Breannan Smith, Danny M. Kaufman, Etienne Vouga, Rasmus Tamstorf, and Eitan Grinspun. Reflections on simultaneous impact. ACM Transactions on Graphics, 31 (4):106:1–106:12, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sen:2012:RPB

[SKY^{+}12] Pradeep Sen, Nima Khademi Kalantari, Maziar Yaesoubi, Soheil Darabi, Dan B. Goldman, and Eli Shechtman. Robust patch-based HDR reconstruction of dynamic scenes. ACM Transactions on Graphics, 31(6):203:1– 203:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Silvennoinen:2017:RTG

[SL17] Ari Silvennoinen and Jaakko Lehtinen. Real-time global illumination by precomputed local reconstruction from sparse radiance probes. ACM Transactions on Graphics, 36 (6):230:1–230:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shugrina:2017:PPI

[SLD17] Maria Shugrina, Jingwan Lu, and Stephen Diverdi. Playful palette: an interactive parametric color mixer for artists. ACM Transactions on Graphics, 36(4):61:1–61:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Selle:2008:MSM

[SLF08] Andrew Selle, Michael Lentine, and Ronald Fedkiw. A mass spring model for hair simulation. ACM Transactions on Graphics, 27(3):64:1– 64:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Secord:2011:PMV

[SLF^{+}11] Adrian Secord, Jingwan Lu, Adam Finkelstein, Manish Singh, and Andrew Nealen. Perceptual models of viewpoint preference. ACM Transactions on Graphics, 30(5): 109:1–109:12, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shugrina:2022:NBE

[SLF22] Maria Shugrina, Chin-Ying Li, and Sanja Fidler. Neural brushstroke engine: Learning a latent style space of interactive drawing tools. ACM Transactions on Graphics, 41 (6):269:1–269:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555472.

Shin:2001:CPI

[SLGS01] Hyun Joon Shin, Jehee Lee, Michael Gleicher, and Sung Yong Shin. Computer puppetry: An importancebased approach. ACM Transactions on Graphics, 20(2): 67–94, April 2001. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2020:MDM

[SLH^{+}20] Liang Shi, Beichen Li, Miloš Hašan, Kalyan Sunkavalli, Tamy Boubekeur, Radomir Mech, and Wojciech Matusik. MATch: differentiable material graphs for procedural material capture. ACM Transactions on Graphics, 39(6): 196:1–196:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417781.

Shan:2008:FIV

[SLJT08] Qi Shan, Zhaorong Li, Jiaya Jia, and Chi-Keung Tang. Fast image/video upsampling. ACM Transactions on Graphics, 27(5):

153:1–153:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sawhney:2024:DRS

[SLK^{+}24] Rohan Sawhney, Daqi Lin, Markus Kettunen, Benedikt Bitterli, Ravi Ramamoorthi, Chris Wyman, and Matt Pharr. Decorrelating Re-STIR samplers via MCMC mutations. ACM Transactions on Graphics, 43(1): 10:1–10:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3629166.

Sun:2006:FM

[SLKS06] Jian Sun, Yin Li, Sing Bing Kang, and Heung-Yeung Shum. Flash matting. ACM Transactions on Graphics, 25 (3):772–778, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shih:2019:DFW

[SLL19] YiChang Shih, Wei-Sheng Lai, and Chia-Kai Liang. Distortion-free wide-angle portraits on camera phones. ACM Transactions on Graphics, 38(4):61:1–61:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Son:2021:RVD

[SLL^{+}21a] Hyeongseok Son, Junyong Lee, Jonghyeop Lee, Sunghyun Cho, and Seungyong Lee. Recurrent video deblurring with blur-invariant motion estimation and pixel volumes. ACM Transactions on Graphics, 40(5):185:1–185:18, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3453720.

Song:2021:ASP

[SLL^{+}21b] Guoxian Song, Linjie Luo, Jing Liu, Wan-Chun Ma, Chunpong Lai, Chuanxia Zheng, and Tat-Jen Cham. AgileGAN: stylizing portraits by inversion-consistent transfer learning. ACM Transactions on Graphics, 40(4): 117:1–117:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459771.

Selgrad:2017:CRRa

[SLM^{+}17a] Kai Selgrad, Alexander Lier, Magdalena Martinek, Christoph Buchenau, Michael Guthe, Franziska Kranz, Henry Schäfer, and Marc Stamminger. A compressed representation for ray tracing parametric surfaces. ACM Transactions on Graphics, 36(1):5:1–5:??, February 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Selgrad:2017:CRRb

[SLM^{+}17b] Kai Selgrad, Alexander Lier, Magdalena Martinek, Christoph Buchenau, Michael Guthe, Franziska Kranz, Henry Schäfer, and Marc Stamminger. A compressed representation for ray tracing parametric surfaces. ACM Transactions on Graphics, 36(4):100:1– 100:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sellan:2023:BGF

[SLM^{+}23] Silvia Sellán, Jack Luong, Leticia Mattos Da Silva, Aravind Ramakrishnan, Yuchuan Yang, and Alec Jacobson. Breaking good: Fracture modes for realtime destruction. ACM Transactions on Graphics, 42(1): 10:1–10:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3549540.

Sheffer:2005:AFR

[SLMB05] Alla Sheffer, Bruno Lévy, Maxim Mogilnitsky, and Alexander Bogomyakov. ABF++: fast and robust angle based flattening. ACM Transactions on Graphics, 24(2): 311–330, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2014:MSS

[SLMR14] Ran Song, Yonghuai Liu, Ralph R. Martin, and Paul L. Rosin. Mesh saliency via spectral processing. ACM Transactions on Graphics, 33(1): 6:1–6:17, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shao:2016:DFM

[SLR^{+}16] Tianjia Shao, Dongping Li, Yuliang Rong, Changxi Zheng, and Kun Zhou. Dynamic furniture modeling through assembly instructions. ACM Transactions on Graphics, 35 (6):172:1–172:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sloan:2005:LDP

[SLS05] Peter-Pike Sloan, Ben Luna, and John Snyder. Local, deformable precomputed radiance transfer. ACM Transactions on Graphics, 24(3): 1216–1224, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharf:2007:ITA

[SLS^{+}07] Andrei Sharf, Thomas Lewiner, Gil Shklarski, Sivan Toledo, and Daniel Cohen-Or. Interactive topology-aware surface reconstruction. ACM Transactions on Graphics, 26(3): 43:1–43:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Seol:2012:SEC

[SLS^{+}12] Yeongho Seol, J. P. Lewis, Jaewoo Seo, Byungkuk Choi, Ken Anjyo, and Junyong Noh. Spacetime expression cloning for blendshapes. ACM Transactions on Graphics, 31(2): 14:1–14:12, April 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shih:2016:BHB

[SLS^{+}16] Kuang-Tsu Shih, Jen-Shuo Liu, Frank Shyu, Su-Ling Yeh, and Homer H. Chen. Blocking harmful blue light while preserving image color appearance. ACM Transactions on Graphics, 35(6): 175:1–175:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sloan:2003:BSR

[SLSS03] Peter-Pike Sloan, Xinguo Liu, Heung-Yeung Shum, and John Snyder. Bi-scale radiance transfer. ACM Transactions on Graphics, 22(3): 370–375, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Si:2014:RBS

[SLST14] Weiguang Si, Sung-Hee Lee, Eftychios Sifakis, and Demetri Terzopoulos. Realistic biomechanical simulation and control of human swimming. ACM Transactions on Graphics, 34(1):10:1–10:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Scher:2013:TDN

[SLV^{+}13] Steven Scher, Jing Liu, Rajan Vaish, Prabath Gunawardane, and James Davis. 3D+2DTV: 3D displays with no ghosting for viewers without glasses. ACM Transactions on Graphics, 32(3):21:1–21:10, June 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Su:2022:SSB

[SLW22] Fujia Su, Sheng Li, and Guoping Wang. SPCBPT: subspace-based probabilistic connections for bidirectional path tracing. ACM Transactions on Graphics, 41(4):77:1–77:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530183.

Su:2014:EST

[SLWF14] Qingkun Su, Wing Ho Andy Li, Jue Wang, and Hongbo Fu. EZ-sketching: three-level optimization for error-tolerant image tracing. ACM Transactions on Graphics, 33(4): 54:1–54:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2023:SBR

[SLWL23] Pengfei Shen, Ruizeng Li, Beibei Wang, and Ligang Liu. Scratch-based reflection art via differentiable rendering. ACM Transactions on Graphics, 42(4): 65:1–65:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592142.

Sun:2007:IVU

[SLWS07] Jian Sun, Lin Liang, Fang Wen, and Heung-Yeung Shum. Image vectorization using optimized gradient meshes. ACM Transactions on Graphics, 26(3):11:1–11:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shao:2013:ICS

[SLZ^{+}13] Tianjia Shao, Wilmot Li, Kun Zhou, Weiwei Xu, Baining Guo, and Niloy J. Mitra. Interpreting concept sketches. ACM Transactions on Graphics, 32(4):56:1–56:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2006:GWP

[SM06] Weifeng Sun and Amar Mukherjee. Generalized wavelet product integral for rendering dynamic glossy objects. ACM Transactions on Graphics, 25(3):955–966, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schwarz:2015:APM

[SM15] Michael Schwarz and Pascal Müller. Advanced procedural modeling of architecture. ACM Transactions on Graphics, 34(4):107:1– 107:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schissler:2017:ISPa

[SM17a] Carl Schissler and Dinesh Manocha. Interactive sound propagation and rendering for large multi-source scenes. ACM Transactions on Graphics, 36(1):2:1–2:??, February 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schissler:2017:ISPb

[SM17b] Carl Schissler and Dinesh Manocha. Interactive sound propagation and rendering for large multi-source scenes. ACM Transactions on Graphics, 36(4):114:1–114:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2019:CFF

[SMB^{+}19] Haichuan Song, Jonàs Martínez, Pierre Bedell, Noémie Vennin, and Sylvain Lefebvre. Colored fused filament fabrication. ACM Transactions on Graphics, 38(5): 141:1–141:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3183793.

Schissler:2021:FDP

[SMC21] Carl Schissler, Gregor Mückl, and Paul Calamia. Fast diffraction pathfinding for dynamic sound propagation. ACM Transactions on Graphics, 40(4):138:1– 138:13, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459751.

Schumacher:2018:MCS

[SMCT18] Christian Schumacher, Steve Marschner, Markus Cross, and Bernhard Thomaszewski. Mechanical characterization of structured sheet materials. ACM Transactions on Graphics, 37(4):148:1– 148:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sigal:2015:PCS

[SMD^{+}15] Leonid Sigal, Moshe Mahler, Spencer Diaz, Kyna McIntosh, Elizabeth Carter, Timothy Richards, and Jessica Hodgins. A perceptual control space for garment simulation. ACM Transactions on Graphics, 34(4):117:1– 117:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sandin:2005:VAV

[SMG^{+}05] Daniel J. Sandin, Todd Margolis, Jinghua Ge, Javier Girado, Tom Peterka, and Thomas A. DeFanti. The Varrier^{TM} autostereoscopic virtual reality display. ACM Transactions on Graphics, 24 (3):894–903, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Serrano:2020:IML

[SMG^{+}20] Ana Serrano, Daniel Martin, Diego Gutierrez, Karol Myszkowski, and Belen Masia. Imperceptible manipulation of lateral camera motion for improved virtual reality applications. ACM Transactions on Graphics, 39(6): 267:1–267:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417773.

Sawhney:2023:WSG

[SMGC23] Rohan Sawhney, Bailey Miller, Ioannis Gkioulekas, and Keenan Crane. Walk on Stars: a grid-free Monte Carlo method for PDEs with Neumann boundary conditions. ACM Transactions on Graphics, 42(4):80:1–80:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592398.

Shrivastava:2011:DDV

[SMGE11] Abhinav Shrivastava, Tomasz Malisiewicz, Abhinav Gupta, and Alexei A. Efros. Datadriven visual similarity for cross-domain image matching. ACM Transactions on Graphics, 30(6):154:1– 154:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2018:APP

[SMGH18] Neil Smith, Nils Moehrle, Michael Goesele, and Wolfgang Heidrich. Aerial path planning for urban scene reconstruction: a continuous optimization method and benchmark. ACM Transactions on Graphics, 37(6): 183:1–183:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sajadi:2011:SPU

[SMH^{+}11] Behzad Sajadi, Aditi Majumder, Kazuhiro Hiwada, Atsuto Maki, and Ramesh Raskar. Switchable primaries using shiftable layers of color filter arrays. ACM Transactions on Graphics, 30(4): 65:1–65:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2023:FIE

[SMH^{+}23] Tianchang Shen, Jacob Munkberg, Jon Hasselgren, Kangxue Yin, Zian Wang, Wenzheng Chen, Zan Gojcic, Sanja Fidler, Nicholas Sharp, and Jun Gao. Flexible isosurface extraction for gradient-based mesh optimization. ACM Transactions on Graphics, 42 (4):37:1–37:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592430.

Shen:2016:SVS

[SMHW16] Wuyao Shen, Xiangyu Mao, Xinghong Hu, and Tien-Tsin Wong. Seamless visual sharing with color vision deficiencies. ACM Transactions on Graphics, 35(4):70:1–70:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Starke:2022:DPA

[SMK22] Sebastian Starke, Ian Mason, and Taku Komura. DeepPhase: periodic autoencoders for learning motion phase manifolds. ACM Transactions on Graphics, 41 (4):136:1–136:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530178.

Sadeghi:2012:PBS

[SML^{+}12] Iman Sadeghi, Adolfo Munoz, Philip Laven, Wojciech Jarosz, Francisco Seron, Diego Gutierrez, and Henrik Wann Jensen. Physically-based simulation of rainbows. ACM Transactions on Graphics, 31(1): 3:1–3:12, January 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schissler:2014:HOD

[SMM14] Carl Schissler, Ravish Mehra, and Dinesh Manocha. Highorder diffraction and diffuse reflections for interactive sound propagation in large environments. ACM Transactions on Graphics, 33(4): 39:1–39:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sand:2003:CCS

[SMP03] Peter Sand, Leonard McMillan, and Jovan Popović. Continuous capture of skin deformation. ACM Transactions on Graphics, 22(3): 578–586, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sunkavalli:2007:FTL

[SMPR07] Kalyan Sunkavalli, Wojciech Matusik, Hanspeter Pfister, and Szymon Rusinkiewicz. Factored time-lapse video. ACM Transactions on Graphics, 26(3):101:1–101:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shilkrot:2015:AAC

[SMPZ15] Roy Shilkrot, Pattie Maes, Joseph A. Paradiso, and Amit Zoran. Augmented airbrush for computer aided painting (CAP). ACM Transactions on Graphics, 34(2):19:1– 19:??, February 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Salehi:2022:DAS

[SMR^{+}22] Farnood Salehi, Marco Manzi, Gerhard Roethlin, Romann Weber, Christopher Schroers, and Marios Papas. Deep adaptive sampling and reconstruction using analytic distributions. ACM Transactions on Graphics, 41(6): 259:1–259:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555515.

Schaefer:2006:IDU

[SMW06] Scott Schaefer, Travis McPhail, and Joe Warren. Image deformation using moving least squares. ACM Transactions on Graphics, 25(3): 533–540, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shao:2014:IUS

[SMZ^{+}14] Tianjia Shao, Aron Monszpart, Youyi Zheng, Bongjin Koo, Weiwei Xu, Kun Zhou, and Niloy J. Mitra. Imagining the unseen: stability-based cuboid arrangements for scene understanding. ACM Transactions on Graphics, 33(6): 209:1–209:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2017:UIA

[SN17] Harrison Jesse Smith and Michael Neff. Understanding the impact of animated gesture performance on personality perceptions. ACM Transactions on Graphics, 36(4): 49:1–49:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sander:2007:FTR

[SNB07] Pedro V. Sander, Diego Nehab, and Joshua Barczak. Fast triangle reordering for vertex locality and reduced overdraw. ACM Transactions on Graphics, 26(3): 89:1–89:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sander:2008:ETM

[SNCH08] Pedro V. Sander, Diego Nehab, Eden Chlamtac, and Hugues Hoppe. Efficient traversal of mesh edges using adjacency primitives. ACM Transactions on Graphics, 27 (5):144:1–144:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sifakis:2005:ADF

[SNF05] Eftychios Sifakis, Igor Neverov, and Ronald Fedkiw. Automatic determination of facial muscle activations from sparse motion capture marker data. ACM Transactions on Graphics, 24(3):417–425, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schmidt:2013:PSM

[SNM^{+}13] Thorsten-Walther Schmidt, Jan Novák, Johannes Meng, Anton S. Kaplanyan, Tim Reiner, Derek Nowrouzezahrai, and Carsten Dachsbacher. Path-space manipulation of physically-based light transport. ACM Transactions on Graphics, 32(4):129:1– 129:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sperl:2020:HYL

[SNW20] Georg Sperl, Rahul Narain, and Chris Wojtan. Homogenized yarn-level cloth. ACM Transactions on Graphics, 39 (4):48:1–48:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392412.

Sperl:2021:MAD

[SNW21] Georg Sperl, Rahul Narain, and Chris Wojtan. Mechanicsaware deformation of yarn pattern geometry. ACM Transactions on Graphics, 40(4):168:1–168:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459816.

Sun:2021:MPM

[SNZ^{+}21] Yuchen Sun, Xingyu Ni, Bo Zhu, Bin Wang, and Baoquan Chen. A material point method for nonlinearly magnetized materials. ACM Transactions on Graphics, 40 (6):205:1–205:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480541.

Sharir:1992:SOS

[SO92] Micha Sharir and Mark H.

Overmars. A simple outputsensitive algorithm for hidden surface removal. ACM Transactions on Graphics, 11(1):1–11, January 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
112141.html.

Sintorn:2011:EAF

[SOA11] Erik Sintorn, Ola Olsson, and Ulf Assarsson. An efficient alias-free shadow algorithm for opaque and transparent objects using per-triangle shadow volumes. ACM Transactions on Graphics, 30(6): 153:1–153:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stutz:2022:SFF

[SOG^{+}22] Florian Cyril Stutz, Tim Felle Olsen, Jeroen Peter Groen, Tuan Nguyen Trung, Niels Aage, Ole Sigmund, Justin Solomon, and Jakob Andreas Bærentzen. Synthesis of frame field-aligned multi-laminar structures. ACM Transactions on Graphics, 41(5): 170:1–170:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3516522.

Sik:2016:RLT

[SOHK16] Martin Sik, Hisanari Otsu, Toshiya Hachisuka, and Jaroslav Krivánek. Robust light transport simulation via metropolised bidirectional estimators. ACM Transactions on Graphics, 35(6):245:1– 245:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2004:IAI

[SOS04] Chen Shen, James F. O’Brien, and Jonathan R. Shewchuk. Interpolating and approximating implicit surfaces from polygon soup. ACM Transactions on Graphics, 23(3): 896–904, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sumner:2004:DTT

[SP04] Robert W. Sumner and Jovan Popović. Deformation transfer for triangle meshes. ACM Transactions on Graphics, 23(3):399–405, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sulejmanpasic:2005:APB

[SP05] Adnan Sulejmanpašić and Jovan Popović. Adaptation of performed ballistic motion. ACM Transactions on Graphics, 24(1):165–179, January 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Solenthaler:2009:PCI

[SP09] B. Solenthaler and R. Pajarola. Predictive-corrective incompressible SPH. ACM Transactions on Graphics, 28 (3):40:1–40:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Santoni:2016:GGP

[SP16] Christian Santoni and Fabio Pellacini. gTangle: a grammar for the procedural generation of tangle patterns. ACM Transactions on Graphics, 35 (6):182:1–182:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shih:2014:STH

[SPB^{+}14] YiChang Shih, Sylvain Paris, Connelly Barnes, William T. Freeman, and Frédo Durand. Style transfer for headshot portraits. ACM Transactions on Graphics, 33(4): 148:1–148:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shih:2013:DDH

[SPDF13] Yichang Shih, Sylvain Paris, Frédo Durand, and William T. Freeman. Data-driven hallucination of different times of day from a single outdoor photo. ACM Transactions on Graphics, 32(6): 200:1–200:??, November 2013.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Spencer:2003:EAS

[Spe03] Stephen N. Spencer. Errata: ACM SIGGRAPH 2002 Papers. ACM Transactions on Graphics, 22(2):258, April 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schott:2023:LST

[SPF^{+}23] Hugo Schott, Axel Paris, Lucie Fournier, Eric Guérin, and Eric Galin. Large-scale terrain authoring through interactive erosion simulation. ACM Transactions on Graphics, 42(5):162:1–162:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592787.

Schuller:2016:CT

[SPG^{+}16] Christian Schüller, Daniele Panozzo, Anselm Grundhöfer, Henning Zimmer, Evgeni Sorkine, and Olga Sorkine-Hornung. Computational thermoforming. ACM Transactions on Graphics, 35(4): 43:1–43:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharma:2023:MSS

[SPG^{+}23] Prafull Sharma, Julien Philip, Michaël Gharbi, Bill Freeman, Fredo Durand, and Valentin Deschaintre. Materialistic: Selecting similar materials in images. ACM Transactions on Graphics, 42(4): 154:1–154:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592390.

Sodhi:2013:AIT

[SPGI13] Rajinder Sodhi, Ivan Poupyrev, Matthew Glisson, and Ali Israr. AIREAL: interactive tactile experiences in free air. ACM Transactions on Graphics, 32(4):134:1–134:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schertler:2018:GMG

[SPGT18] Nico Schertler, Daniele Panozzo, Stefan Gumhold, and Marco Tarini. Generalized motorcycle graphs for imperfect quad-dominant meshes. ACM Transactions on Graphics, 37(4):155:1–155:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sadeghi:2010:AFH

[SPJT10] Iman Sadeghi, Heather Pritchett, Henrik Wann Jensen, and Rasmus Tamstorf. An artist friendly hair shading system. ACM Transactions on Graphics, 29(4):56:1–56:??, July 2010. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Solomon:2016:EMA

[SPKS16] Justin Solomon, Gabriel Peyré, Vladimir G. Kim, and Suvrit Sra. Entropic metric alignment for correspondence problems. ACM Transactions on Graphics, 35(4): 72:1–72:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schvartzman:2010:SCE

[SPO10] Sara C. Schvartzman, Álvaro G.

Pérez, and Miguel A. Otaduy. Star-contours for efficient hierarchical self-collision detection. ACM Transactions on Graphics, 29(4): 80:1–80:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sproull:1982:UPT

[Spr82] R. F. Sproull. Using program transformations to derive linedrawing algorithms. ACM Transactions on Graphics, 1 (4):259–273, October 1982. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shiratori:2011:MCB

[SPS^{+}11] Takaaki Shiratori, Hyun Soo Park, Leonid Sigal, Yaser Sheikh, and Jessica K. Hodgins. Motion capture from body-mounted cameras. ACM Transactions on Graphics, 30 (4):31:1–31:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schuller:2014:AMS

[SPSH14] Christian Schüller, Daniele Panozzo, and Olga Sorkine-Hornung. Appearancemimicking surfaces. ACM Transactions on Graphics, 33 (6):216:1–216:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shtengel:2017:GOC

[SPSH^{+}17] Anna Shtengel, Roi Poranne, Olga Sorkine-Hornung, Shahar Z. Kovalsky, and Yaron Lipman. Geometric optimization via composite majorization. ACM Transactions on Graphics, 36(4): 38:1–38:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schuller:2018:SRZ

[SPSH18] Christian Schüller, Roi Poranne, and Olga Sorkine-Hornung. Shape representation by zippables. ACM Transactions on Graphics, 37 (4):78:1–78:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shin:2016:REE

[SPV^{+}16] Hijung V. Shin, Christopher F. Porst, Etienne Vouga, John Ochsendorf, and Frédo Durand. Reconciling elastic and equilibrium methods for static analysis. ACM Transactions on Graphics, 35(2): 13:1–13:??, May 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2018:TVR

[SPW^{+}18] Qi Sun, Anjul Patney, Li-Yi Wei, Omer Shapira, Jingwan Lu, Paul Asente, Suwen Zhu, Morgan Mcguire, David Luebke, and Arie Kaufman. Towards virtual reality infinite walking: dynamic saccadic redirection. ACM Transactions on Graphics, 37(4):67:1– 67:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Streuber:2016:BTC

[SQRH^{+}16] Stephan Streuber, M. Alejandra Quiros-Ramirez, Matthew Q. Hill, Carina A. Hahn, Silvia Zuffi, Alice O’Toole, and Michael J. Black. Body talk: crowdshaping realistic 3D avatars with words. ACM Transactions on Graphics, 35 (4):54:1–54:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Son:2022:DHT

[SQSL22] Sanghyun Son, Yi-Ling Qiao, Jason Sewall, and Ming C. Lin. Differentiable hybrid traffic simulation. ACM Transactions on Graphics, 41 (6):258:1–258:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555492.

Sanchez-Reyes:1997:SAP

[SR97] J. Sánchez-Reyes. The symmetric analogue of the polynomial power basis. ACM Transactions on Graphics, 16 (3):319–357, July 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1997-16-3/
                                        p319-sanchez-reyes/.

Sanchez-Reyes:2000:APP

[SR00] Javier Sánchez-Reyes. Applications of the polynomial s-power basis in geometry processing. ACM Transactions on Graphics, 19(1): 27–55, January 2000. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/2000-19-1/
                                        p27-sachez-reyes/.

Sun:2009:ADT

[SR09] Bo Sun and Ravi Ramamoorthi. Affine double- and tripleproduct wavelet integrals for rendering. ACM Transactions on Graphics, 28(2):14:1– 14:17, April 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sumin:2019:GAS

[SRB^{+}19] Denis Sumin, Tobias Rittig, Vahid Babaei, Thomas Nindel, Alexander Wilkie, Piotr Didyk, Bernd Bickel, Jaroslav Krivánek, Karol Myszkowski, and Tim Weyrich. Geometryaware scattering compensation for 3D printing. ACM Transactions on Graphics, 38 (4):111:1–111:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Selle:2005:VPM

[SRF05] Andrew Selle, Nick Rasmussen, and Ronald Fedkiw. A vortex particle method for smoke, water and explosions. ACM Transactions on Graphics, 24(3):910–914, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Solomon:2014:EMD

[SRGB14] Justin Solomon, Raif Rustamov, Leonidas Guibas, and Adrian Butscher. Earth mover’s distances on discrete surfaces. ACM Transactions on Graphics, 33(4): 67:1–67:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schreck:2015:NDG

[SRH^{+}15] Camille Schreck, Damien Rohmer, Stefanie Hahmann, Marie-Paule Cani, Shuo Jin, Charlie C. L. Wang, and Jean-Francis Bloch. Nonsmooth developable geometry for interactively animating paper crumpling. ACM Transactions on Graphics, 35(1):10:1– 10:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Swedish:2015:ESD

[SRL^{+}15] Tristan Swedish, Karin Roesch, Ik-Hyun Lee, Krishna Rastogi, Shoshana Bernstein, and Ramesh Raskar. eyeSelfie: self directed eye alignment using reciprocal eye box imaging. ACM Transactions on Graphics, 34(4):58:1–58:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2005:PAS

[SRNN05] Bo Sun, Ravi Ramamoorthi, Srinivasa G. Narasimhan, and Shree K. Nayar. A practical analytic single scattering model for real time rendering. ACM Transactions on Graphics, 24(3):1040–1049, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sokolov:2016:HDM

[SRUL16] Dmitry Sokolov, Nicolas Ray, Lionel Untereiner, and Bruno Lévy. Hexahedral-dominant meshing. ACM Transactions on Graphics, 35(5):157:1– 157:??, September 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sokolov:2017:HDM

[SRUL17] Dmitry Sokolov, Nicolas Ray, Lionel Untereiner, and Bruno Lévy. Hexahedral-dominant meshing. ACM Transactions on Graphics, 36(4): 114:1–114:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shacklett:2023:EDO

[SRX^{+}23] Brennan Shacklett, Luc Guy Rosenzweig, Zhiqiang Xie, Bidipta Sarkar, Andrew Szot, Erik Wijmans, Vladlen Koltun, Dhruv Batra, and Kayvon Fatahalian. An extensible, data-oriented architecture for high-performance, many-world simulation. ACM Transactions on Graphics, 42 (4):90:1–90:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592427.

Soler:2000:TBV

[SS00] Cyril Soler and F. X. Sillion.

Texture-based visibility for efficient lighting simulation. ACM Transactions on Graphics, 19(4):302–342, October 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tog/
                                        2000-19-4/p302-soler/.

Schwarz:2010:FPS

[SS10a] Michael Schwarz and Hans-Peter Seidel. Fast parallel surface and solid voxelization on GPUs. ACM Transactions on Graphics, 29(6): 179:1–179:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Singh:2010:TSD

[SS10b] Mayank Singh and Scott Schaefer. Triangle surfaces with discrete equivalence classes. ACM Transactions on Graphics, 29(4): 46:1–46:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stam:2011:VIS

[SS11] Jos Stam and Ryan Schmidt. On the velocity of an implicit surface. ACM Transactions on Graphics, 30(3): 21:1–21:7, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sadri:2014:FCB

[SS14] Bardia Sadri and Karan Singh. Flow-complex-based shape reconstruction from 3D curves. ACM Transactions on Graphics, 33(2):20:1– 20:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2015:BPF

[SS15] Jason Smith and Scott Schaefer. Bijective parameterization with free boundaries. ACM Transactions on Graphics, 34(4):70:1–70:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stomakhin:2017:FAB

[SS17] Alexey Stomakhin and Andrew Selle. Fluxed animated boundary method. ACM Transactions on Graphics, 36 (4):68:1–68:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Saragadam:2019:KKS

[SS19] Vishwanath Saragadam and Aswin C. Sankaranarayanan. KRISM–Krylov subspacebased optical computing of hyperspectral images. ACM Transactions on Graphics, 38(5):148:1–148:??, October 2019. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3345553.

Smirnov:2021:HLS

[SS21] Dmitriy Smirnov and Justin Solomon. HodgeNet: learning spectral geometry on triangle meshes. ACM Transactions on Graphics, 40(4): 166:1–166:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459797.

Sachdeva:2015:BSC

[SSB^{+}15] Prashant Sachdeva, Shinjiro Sueda, Susanne Bradley, Mikhail Fain, and Dinesh K. Pai. Biomechanical simulation and control of hands and tendinous systems. ACM Transactions on Graphics, 34 (4):42:1–42:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schulz:2017:RPSa

[SSB^{+}17a] Adriana Schulz, Ariel Shamir, Ilya Baran, David I. W. Levin, Pitchaya Sitthi-Amorn, and Wojciech Matusik. Retrieval on parametric shape collections. ACM Transactions on Graphics, 36(1):11:1– 11:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schulz:2017:RPSb

[SSB^{+}17b] Adriana Schulz, Ariel Shamir, Ilya Baran, David I. W. Levin, Pitchaya Sitthi-Amorn, and Wojciech Matusik. Retrieval on parametric shape collections. ACM Transactions on Graphics, 36(4): 52:1–52:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Soler:2003:EIA

[SSBD03] Cyril Soler, François X. Sillion, Frédéric Blaise, and Philippe Dereffye. An efficient instantiation algorithm for simulating radiant energy transfer in plant models. ACM Transactions on Graphics, 22 (2):204–233, April 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schmid:2010:PME

[SSBG10] Johannes Schmid, Robert W.

Sumner, Huw Bowles, and Markus Gross. Programmable motion effects. ACM Transactions on Graphics, 29(4): 57:1–57:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sperl:2022:EYL

[SSBL^{+}22] Georg Sperl, Rosa M. Sánchez-Banderas, Manwen Li, Chris Wojtan, and Miguel A. Otaduy. Estimation of yarnlevel simulation models for production fabrics. ACM Transactions on Graphics, 41(4):65:1–65:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530167.

Slater:2010:SVE

[SSC10] Mel Slater, Bernhard Spanlang, and David Corominas. Simulating virtual environments within virtual environments as the basis for a psychophysics of presence. ACM Transactions on Graphics, 29 (4):92:1–92:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stomakhin:2013:MPM

[SSC^{+}13] Alexey Stomakhin, Craig Schroeder, Lawrence Chai, Joseph Teran, and Andrew Selle. A material point method for snow simulation. ACM Transactions on Graphics, 32(4):102:1–102:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Soliman:2018:OCS

[SSC18] Yousuf Soliman, Dejan Slepcev, and Keenan Crane. Optimal cone singularities for conformal flattening. ACM Transactions on Graphics, 37(4): 105:1–105:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharp:2019:NIT

[SSC19a] Nicholas Sharp, Yousuf Soliman, and Keenan Crane. Navigating intrinsic triangulations. ACM Transactions on Graphics, 38(4): 55:1–55:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sharp:2019:VHM

[SSC19b] Nicholas Sharp, Yousuf Soliman, and Keenan Crane. The vector heat method. ACM Transactions on Graphics, 38 (3):24:1–24:??, June 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3243651.

Soler:2009:FDF

[SSD^{+}09a] Cyril Soler, Kartic Subr, Frédo Durand, Nicolas Holzschuch, and François Sillion. Fourier depth of field. ACM Transactions on Graphics, 28(2): 18:1–18:12, April 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Subr:2009:EPM

[SSD09b] Kartic Subr, Cyril Soler, and Frédo Durand. Edgepreserving multiscale image decomposition based on local extrema. ACM Transactions on Graphics, 28(5): 147:1–147:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schmid:2011:OIC

[SSGS11] Johannes Schmid, Martin Sebastian Senn, Markus Gross, and Robert W. Sumner. Over-Coat: an implicit canvas for 3D painting. ACM Transactions on Graphics, 30(4): 28:1–28:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Simo-Serra:2018:MSA

[SSII18a] Edgar Simo-Serra, Satoshi Iizuka, and Hiroshi Ishikawa. Mastering sketching: Adversarial augmentation for structured prediction. ACM Transactions on Graphics, 37(1): 11:1–11:??, January 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Simo-Serra:2018:RTD

[SSII18b] Edgar Simo-Serra, Satoshi Iizuka, and Hiroshi Ishikawa. Real-time data-driven interactive rough sketch inking. ACM Transactions on Graphics, 37 (4):98:1–98:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Simo-Serra:2016:LSF

[SSISI16] Edgar Simo-Serra, Satoshi Iizuka, Kazuma Sasaki, and Hiroshi Ishikawa. Learning to simplify: fully convolutional networks for rough sketch cleanup. ACM Transactions on Graphics, 35(4): 121:1–121:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Summa:2011:IEM

[SSJ^{+}11] Brian Summa, Giorgio Scorzelli, Ming Jiang, Peer-Timo Bremer, and Valerio Pascucci. Interactive editing of massive imagery made simple: Turning Atlanta into Atlantis. ACM Transactions on Graphics, 30(2):7:1–7:13, April 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stomakhin:2014:AMP

[SSJ^{+}14] Alexey Stomakhin, Craig Schroeder, Chenfanfu Jiang, Lawrence Chai, Joseph Teran, and Andrew Selle. Augmented MPM for phase-change and varied materials. ACM Transactions on Graphics, 33(4): 138:1–138:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Skrivan:2020:WCS

[SSJ^{+}20] Tomas Skrivan, Andreas Soderstrom, John Johansson, Christoph Sprenger, Ken Museth, and Chris Wojtan. Wave curves: simulating Lagrangian water waves on dynamically deforming surfaces. ACM Transactions on Graphics, 39(4): 65:1–65:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392466.

Sawhney:2022:GFM

[SSJC22] Rohan Sawhney, Dario Seyb, Wojciech Jarosz, and Keenan Crane. Grid-free Monte Carlo for PDEs with spatially varying coefficients. ACM Transactions on Graphics, 41(4):53:1–53:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530134.

Song:2005:SNW

[SSK05a] Oh-Young Song, Hyuncheol Shin, and Hyeong-Seok Ko. Stable but nondissipative water. ACM Transactions on Graphics, 24(1):81–97, January 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Surazhsky:2005:FEA

[SSK^{+}05b] Vitaly Surazhsky, Tatiana Surazhsky, Danil Kirsanov, Steven J. Gortler, and Hugues Hoppe. Fast exact and approximate geodesics on meshes. ACM Transactions on Graphics, 24(3): 553–560, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Seol:2011:AFF

[SSK^{+}11] Yeongho Seol, Jaewoo Seo, Paul Hyunjin Kim, J. P. Lewis, and Junyong Noh. Artist friendly facial animation retargeting. ACM Transactions on Graphics, 30(6): 162:1–162:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sung:2017:CWS

[SSK^{+}17] Minhyuk Sung, Hao Su, Vladimir G. Kim, Siddhartha Chaudhuri, and Leonidas Guibas. Complementme: weakly-supervised component suggestions for 3D modeling. ACM Transactions on Graphics, 36(6):226:1– 226:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Suwajanakorn:2017:SOL

[SSKS17] Supasorn Suwajanakorn, Steven M.

Seitz, and Ira Kemelmacher-Shlizerman. Synthesizing Obama: learning lip sync from audio. ACM Transactions on Graphics, 36(4): 95:1–95:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schulz:2014:DFE

[SSL^{+}14] Adriana Schulz, Ariel Shamir, David I. W. Levin, Pitchaya Sitthi-amorn, and Wojciech Matusik. Design and fabrication by example. ACM Transactions on Graphics, 33(4): 62:1–62:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shugrina:2015:FFC

[SSM15] Maria Shugrina, Ariel Shamir, and Wojciech Matusik. Fab forms: customizable objects for fabrication with validity and geometry caching. ACM Transactions on Graphics, 34(4):100:1–100:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sumner:2007:EDS

[SSP07] Robert W. Sumner, Johannes Schmid, and Mark Pauly. Embedded deformation for shape manipulation. ACM Transactions on Graphics, 26(3): 80:1–80:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Springborn:2008:CET

[SSP08] Boris Springborn, Peter Schröder, and Ulrich Pinkall. Conformal equivalence of triangle meshes. ACM Transactions on Graphics, 27(3):77:1– 77:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2020:AFR

[SSR20] Steven L. Song, Weiqi Shi, and Michael Reed. Accurate face rig approximation with deep differential subspace reconstruction. ACM Transactions on Graphics, 39(4): 34:1–34:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392491.

Serrano:2017:MEC

[SSRB^{+}17] Ana Serrano, Vincent Sitzmann, Jaime Ruiz-Borau, Gordon Wetzstein, Diego Gutierrez, and Belen Masia. Movie editing and cognitive event segmentation in virtual reality video. ACM Transactions on Graphics, 36(4): 47:1–47:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Snavely:2006:PTE

[SSS06] Noah Snavely, Steven M.

Seitz, and Richard Szeliski. Photo tourism: exploring photo collections in 3D. ACM Transactions on Graphics, 25 (3):835–846, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sinha:2008:IAM

[SSS^{+}08] Sudipta N. Sinha, Drew Steedly, Richard Szeliski, Maneesh Agrawala, and Marc Pollefeys. Interactive 3D architectural modeling from unordered photo collections. ACM Transactions on Graphics, 27(5):159:1–159:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shu:2017:EEE

[SSSH17] Zhixin Shu, Eli Shechtman, Dimitris Samaras, and Sunil Hadap. EyeOpener: Editing eyes in the wild. ACM Transactions on Graphics, 36(1): 1:1–1:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sproull:1983:D

[SST^{+}83] R. F. Sproull, I. E. Sutherland, A. Thomson, S. Gupta, and C. Minter. The 8 by 8 display. ACM Transactions on Graphics, 2(1):32–56, January 1983. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Salvati:2015:MCM

[SSTP15] Gabriele Salvati, Christian Santoni, Valentina Tibaldo, and Fabio Pellacini. Mesh-Histo: collaborative modeling by sharing and retargeting editing histories. ACM Transactions on Graphics, 34(6):

205:1–205:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stanton:2013:NPG

[SSW^{+}13] Matt Stanton, Yu Sheng, Martin Wicke, Federico Perazzi, Amos Yuen, Srinivasa Narasimhan, and Adrien Treuille. Non-polynomial Galerkin projection on deforming meshes. ACM Transactions on Graphics, 32(4): 86:1–86:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2023:CHF

[SSW^{+}23] Yuefan Shen, Shunsuke Saito, Ziyan Wang, Olivier Maury, Chenglei Wu, Jessica Hodgins, Youyi Zheng, and Giljoo Nam. CT2Hair: High-fidelity 3D hair modeling using computed tomography. ACM Transactions on Graphics, 42 (4):75:1–75:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592106.

Shum:2004:PLF

[SSY^{+}04] Heung-Yeung Shum, Jian Sun, Shuntaro Yamazaki, Yin Li, and Chi-Keung Tang. Popup light field: an interactive image-based modeling and rendering system. ACM Transactions on Graphics, 23 (2):143–162, April 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Steinberg:2022:TPP

[SSY22] Shlomi Steinberg, Pradeep Sen, and Ling-Qi Yan. Towards practical physicaloptics rendering. ACM Transactions on Graphics, 41 (4):132:1–132:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530119.

Shalom:2010:CCS

[SSZCO10] Shy Shalom, Ariel Shamir, Hao Zhang, and Daniel Cohen-Or. Cone carving for surface reconstruction. ACM Transactions on Graphics, 29 (6):150:1–150:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sand:2004:VM

[ST04] Peter Sand and Seth Teller. Video matching. ACM Transactions on Graphics, 23(3): 592–599, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schneider:2014:SCC

[ST14] Rosália G. Schneider and Tinne Tuytelaars. Sketch classification and classificationdriven analysis using Fisher vectors. ACM Transactions on Graphics, 33(6):174:1– 174:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schneider:2016:EBS

[ST16] Rosália G. Schneider and Tinne Tuytelaars. Examplebased sketch segmentation and labeling using CRFs. ACM Transactions on Graphics, 35(5):151:1–151:??, September 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stam:2003:FSA

[Sta03] Jos Stam. Flows on surfaces of arbitrary topology. ACM Transactions on Graphics, 22 (3):724–731, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Skouras:2013:CDA

[STC^{+}13] Mélina Skouras, Bernhard Thomaszewski, Stelian Coros, Bernd Bickel, and Markus Gross. Computational design of actuated deformable characters. ACM Transactions on Graphics, 32(4): 82:1–82:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Steinberg:2020:ARL

[Ste20] Shlomi Steinberg. Accurate rendering of liquid-crystals and inhomogeneous optically anisotropic media. ACM Transactions on Graphics, 39 (3):22:1–22:23, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3381748.

Schertler:2017:FAO

[STJ^{+}17] Nico Schertler, Marco Tarini, Wenzel Jakob, Misha Kazhdan, Stefan Gumhold, and Daniele Panozzo. Fieldaligned online surface reconstruction. ACM Transactions on Graphics, 36(4): 77:1–77:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Skouras:2014:DIS

[STK^{+}14] Mélina Skouras, Bernhard Thomaszewski, Peter Kaufmann, Akash Garg, Bernd Bickel, Eitan Grinspun, and Markus Gross. Designing inflatable structures. ACM Transactions on Graphics, 33 (4):63:1–63:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stone:1992:SIC

[Sto92] Maureen C. Stone. Special issue on color. ACM Transactions on Graphics, 11(4): 300–304, October 1992. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Summa:2012:PWF

[STP12] Brian Summa, Julien Tierny, and Valerio Pascucci. Panorama weaving: fast and flexible seam processing. ACM Transactions on Graphics, 31(4): 83:1–83:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2009:SRE

[STPP09] Ying Song, Xin Tong, Fabio Pellacini, and Pieter Peers. SubEdit: a representation for editing measured heterogeneous subsurface scattering. ACM Transactions on Graphics, 28(3):31:1–31:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schwartzburg:2014:HCC

[STTP14] Yuliy Schwartzburg, Romain Testuz, Andrea Tagliasacchi, and Mark Pauly. Highcontrast computational caustic design. ACM Transactions on Graphics, 33(4): 74:1–74:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2015:BAR

[STXJ15] Jianping Shi, Xin Tao, Li Xu, and Jiaya Jia. Break Ames room illusion: depth from general single images. ACM Transactions on Graphics, 34 (6):225:1–225:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2014:FMR

[STZ14] Timothy Sun, Papoj Thamjaroenporn, and Changxi Zheng. Fast multipole representation of diffusion curves and points. ACM Transactions on Graphics, 33(4): 53:1–53:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2016:RFM

[STZ^{+}16] Xiaoyong Shen, Xin Tao, Chao Zhou, Hongyun Gao, and Jiaya Jia. Regional foremost matching for Internet scene images. ACM Transactions on Graphics, 35(6): 178:1–178:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2006:RBI

[Sun06] Yinlong Sun. Rendering biological iridescences with RGBbased renderers. ACM Transactions on Graphics, 25(1): 100–129, January 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shapiro:1993:SBC

[SV93] Vadim Shapiro and Donald L. Vossler. Separation for boundary to CSG conversion. ACM Transactions on Graphics, 12(1):35–55, January 1993. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
169723.html.

Solomon:2019:OTB

[SV19] Justin Solomon and Amir Vaxman. Optimal transportbased polar interpolation of directional fields. ACM Transactions on Graphics, 38(4): 88:1–88:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Stava:2012:SRI

[SVB^{+}12] Ondrej Stava, Juraj Vanek, Bedrich Benes, Nathan Carr, and Radomír Mech. Stress relief: improving structural strength of 3D printable objects. ACM Transactions on Graphics, 31(4): 48:1–48:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Solomon:2017:BEOa

[SVB17a] Justin Solomon, Amir Vaxman, and David Bommes. Boundary element octahedral fields in volumes. ACM Transactions on Graphics, 36(3): 28:1–28:??, June 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Solomon:2017:BEOb

[SVB17b] Justin Solomon, Amir Vaxman, and David Bommes. Boundary element octahedral fields in volumes. ACM Transactions on Graphics, 36(4): 114:1–114:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sacht:2015:NC

[SVJ15] Leonardo Sacht, Etienne Vouga, and Alec Jacobson. Nested cages. ACM Transactions on Graphics, 34(6): 170:1–170:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sidi:2011:UCS

[SvKK^{+}11] Oana Sidi, Oliver van Kaick, Yanir Kleiman, Hao Zhang, and Daniel Cohen-Or. Unsupervised co-segmentation of a set of shapes via descriptor-space spectral clustering. ACM Transactions on Graphics, 30(6):126:1– 126:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schulz:2014:ADO

[SvTSH14] Christian Schulz, Christoph von Tycowicz, Hans-Peter Seidel, and Klaus Hildebrandt. Animating deformable objects using sparse spacetime constraints. ACM Transactions on Graphics, 33(4):

109:1–109:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Samet:1985:SCP

[SW85] Hanan Samet and Robert E.

Webber. Storing a collection of polygons using quadtrees. ACM Transactions on Graphics, 4(3):182–222, July 1985. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schaefer:2005:TQS

[SW05] Scott Schaefer and Joe Warren. On C^{2} triangle/quad subdivision. ACM Transactions on Graphics, 24(1): 28–36, January 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schwarz:2014:PDE

[SW14] Michael Schwarz and Peter Wonka. Procedural design of exterior lighting for buildings with complex constraints. ACM Transactions on Graphics, 33(5):166:1– 166:??, August 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shu:2018:LAR

[SW18] Xiao Shu and Xiaolin Wu. Locally adaptive rank-constrained optimal tone mapping. ACM Transactions on Graphics, 37 (3):38:1–38:??, August 2018.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3225219.

Schulz:2018:IED

[SWC^{+}18] Adriana Schulz, Harrison Wang, Eitan Crinspun, Justin Solomon, and Wojciech Matusik. Interactive exploration of design trade-offs. ACM Transactions on Graphics, 37(4):131:1–131:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2021:EEC

[SWF^{+}21] Qilin Sun, Congli Wang, Qiang Fu, Xiong Dun, and Wolfgang Heidrich. End-to-end complex lens design with differentiate ray tracing. ACM Transactions on Graphics, 40 (4):71:1–71:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459674.

Sun:2016:MVP

[SWK16] Qi Sun, Li-Yi Wei, and Arie Kaufman. Mapping virtual and physical reality. ACM Transactions on Graphics, 35 (4):64:1–64:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sewall:2011:IHS

[SWL11] Jason Sewall, David Wilkie, and Ming C. Lin. Interactive hybrid simulation of largescale traffic. ACM Transactions on Graphics, 30(6): 135:1–135:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2022:WSF

[SWL^{+}22] Jiahui Sun, Wenming Wu, Ligang Liu, Wenjie Min, Gaofeng Zhang, and Liping Zheng. WallPlan: synthesizing floorplans by learning to generate wall graphs. ACM Transactions on Graphics, 41(4):92:1–92:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530135.

Srinivasan:2021:LAQ

[SWR^{+}21] Sangeetha Grama Srinivasan, Qisi Wang, Junior Rojas, Gergely Klár, Ladislav Kavan, and Eftychios Sifakis. Learning active quasistatic physics-based models from data. ACM Transactions on Graphics, 40(4):129:1– 129:14, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459883.

Sun:2022:IID

[SWS^{+}22] Jingxiang Sun, Xuan Wang, Yichun Shi, Lizhen Wang, Jue Wang, and Yebin Liu. IDE-3D: Interactive disentangled editing for highresolution 3D-aware portrait synthesis. ACM Transactions on Graphics, 41(6): 270:1–270:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555506.

Song:2017:CDW

[SWT^{+}17] Peng Song, Xiaofei Wang, Xiao Tang, Chi-Wing Fu, Hongfei Xu, Ligang Liu, and Niloy J. Mitra. Computational design of wind-up toys. ACM Transactions on Graphics, 36(6):238:1– 238:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2014:AAH

[SWTC14] Fuhao Shi, Hsiang-Tao Wu, Xin Tong, and Jinxiang Chai. Automatic acquisition of highfidelity facial performances using monocular videos. ACM Transactions on Graphics, 33 (6):222:1–222:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Smith:2020:CDH

[SWW^{+}20] Breannan Smith, Chenglei Wu, He Wen, Patrick Peluse, Yaser Sheikh, Jessica K. Hodgins, and Takaaki Shiratori. Constraining dense hand surface tracking with elasticity. ACM Transactions on Graphics, 39(6): 219:1–219:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417768.

Song:2015:VRF

[SWWW15] Ying Song, Jiaping Wang, Li-Yi Wei, and Wencheng Wang. Vector regression functions for texture compression. ACM Transactions on Graphics, 35 (1):5:1–5:??, December 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shirley:1996:MCT

[SWZ96] Peter Shirley, Changyaw Wang, and Kurt Zimmerman. Monte Carlo techniques for direct lighting calculations. ACM Transactions on Graphics, 15(1):1–36, January 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
225887.html;   http://www.
acm.org/pubs/toc/Abstracts/
0730-0301/226151.html.

Sun:2012:DCT

[SXD^{+}12] Xin Sun, Guofu Xie, Yue Dong, Stephen Lin, Weiwei Xu, Wencheng Wang, Xin Tong, and Baining Guo. Diffusion curve textures for resolution independent texture mapping. ACM Transactions on Graphics, 31(4): 74:1–74:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Su:2021:USO

[SXH^{+}21] Haozhe Su, Tao Xue, Chengguizi Han, Chenfanfu Jiang, and Mridul Aanjaneya. A unified second-order accurate in time MPM formulation for simulating viscoelastic liquids with phase change. ACM Transactions on Graphics, 40(4): 119:1–119:18, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459820.

Shao:2012:IAS

[SXZ^{+}12] Tianjia Shao, Weiwei Xu, Kun Zhou, Jingdong Wang, Dongping Li, and Baining Guo. An interactive approach to semantic modeling of indoor scenes with an RGBD camera. ACM Transactions on Graphics, 31(6): 136:1–136:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schulz:2017:IDS

[SXZ^{+}17] Adriana Schulz, Jie Xu, Bo Zhu, Changxi Zheng, Eitan Grinspun, and Wojciech Matusik. Interactive design space exploration and optimization for CAD models. ACM Transactions on Graphics, 36(4):157:1–157:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2020:LSS

[SXZ^{+}20] Tiancheng Sun, Zexiang Xu, Xiuming Zhang, Sean Fanello, Christoph Rhemann, Paul Debevec, Yun-Ta Tsai, Jonathan T. Barron, and Ravi Ramamoorthi. Light stage super-resolution: continuous high-frequency relighting. ACM Transactions on Graphics, 39(6): 260:1–260:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417821.

Shi:2005:CSA

[SY05] Lin Shi and Yizhou Yu. Controllable smoke animation with guiding objects. ACM Transactions on Graphics, 24 (1):140–164, January 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Steinberg:2021:GFP

[SY21a] Shlomi Steinberg and Ling-Qi Yan. A generic framework for physical light transport. ACM Transactions on Graphics, 40(4):139:1– 139:20, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459791.

Steinberg:2021:PLM

[SY21b] Shlomi Steinberg and Ling-Qi Yan. Physical light-matter interaction in Hermite–Gauss space. ACM Transactions on Graphics, 40(6):283:1– 283:17, December 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480530.

Steinberg:2022:RSS

[SY22] Shlomi Steinberg and Ling-Qi Yan. Rendering of subjective speckle formed by rough statistical surfaces. ACM Transactions on Graphics, 41 (1):2:1–2:23, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3472293.

Shi:2006:FMA

[SYBF06] Lin Shi, Yizhou Yu, Nathan Bell, and Wei-Wen Feng. A fast multigrid algorithm for mesh deformation. ACM Transactions on Graphics, 25 (3):1108–1117, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2005:ICS

[SYJS05] Jian Sun, Lu Yuan, Jiaya Jia, and Heung-Yeung Shum. Image completion with structure propagation. ACM Transactions on Graphics, 24(3): 861–868, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2024:HHA

[SYM^{+}24] Jia-Mu Sun, Jie Yang, Kaichun Mo, Yu-Kun Lai, Leonidas Guibas, and Lin Gao. Haisor: Human-aware indoor scene optimization via deep reinforcement learning. ACM Transactions on Graphics, 43(2):15:1–15:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3632947.

Shen:2021:HOD

[SYS^{+}21] Siyuan Shen, Yin Yang, Tianjia Shao, He Wang, Chenfanfu Jiang, Lei Lan, and Kun Zhou. High-order differentiable autoencoder for nonlinear model reduction. ACM Transactions on Graphics, 40 (4):68:1–68:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459754.

Shi:2014:CTS

[SYSP14] Kan-Le Shi, Jun-Hai Yong, Jia-Guang Sun, and Jean-Claude Paul. Continuity transition with a single regular curved-knot spline surface. ACM Transactions on Graphics, 33(5):164:1– 164:??, August 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shuai:2023:RCH

[SYZ^{+}23] Qing Shuai, Zhiyuan Yu, Zhize Zhou, Lixin Fan, Haijun Yang, Can Yang, and Xiaowei Zhou. Reconstructing close human interactions from multiple views. ACM Transactions on Graphics, 42(6): 273:1–273:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618336.

Sun:2015:CDT

[SZ15] Timothy Sun and Changxi Zheng. Computational design of twisty joints and puzzles. ACM Transactions on Graphics, 34(4):101:1– 101:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Schumacher:2018:SSW

[SZB18] Christian Schumacher, Jonas Zehnder, and Moritz Bächer. Set-in-stone: worst-case optimization of structures weak in tension. ACM Transactions on Graphics, 37(6): 252:1–252:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sederberg:2003:SN

[SZBN03] Thomas W. Sederberg, Jianmin Zheng, Almaz Bakenov, and Ahmad Nasri. T-splines and T-NURCCs. ACM Transactions on Graphics, 22(3): 477–484, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2007:IRD

[SZC^{+}07] Xin Sun, Kun Zhou, Yanyun Chen, Stephen Lin, Jiaoying Shi, and Baining Guo. Interactive relighting with dynamic BRDFs. ACM Transactions on Graphics, 26(3): 27:1–27:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shen:2022:WCF

[SZC^{+}22] Hanxiao Shen, Leyi Zhu, Ryan Capouellez, Daniele Panozzo, Marcel Campen, and Denis Zorin. Which cross fields can be quadrangulated?: global parameterization from prescribed holonomy signatures. ACM Transactions on Graphics, 41(4):59:1–59:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530187.

Sun:2020:EEL

[SZD^{+}20] Qilin Sun, Jian Zhang, Xiong Dun, Bernard Ghanem, Yifan Peng, and Wolfgang Heidrich. End-to-end learned, optically coded super-resolution SPAD camera. ACM Transactions on Graphics, 39(2): 9:1–9:14, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3372261.

Szeliski:2006:LAH

[Sze06]  Richard Szeliski.  Locally

adapted hierarchical basis preconditioning. ACM Transactions on Graphics, 25(3): 1135–1143, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2013:LSS

[SZG^{+}13] Xin Sun, Kun Zhou, Jie Guo, Guofu Xie, Jingui Pan, Wencheng Wang, and Baining Guo. Line segment sampling with blue-noise properties. ACM Transactions on Graphics, 32(4):127:1–127:??, July 2013. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Sumner:2005:MBI

[SZGP05] Robert W. Sumner, Matthias Zwicker, Craig Gotsman, and Jovan Popović. Mesh-based inverse kinematics. ACM Transactions on Graphics, 24 (3):488–495, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Saito:2015:CBA

[SZK15] Shunsuke Saito, Zi-Ye Zhou, and Ladislav Kavan. Computational bodybuilding: anatomicallybased modeling of human bodies. ACM Transactions on Graphics, 34(4):41:1– 41:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Starke:2020:LMP

[SZKZ20] Sebastian Starke, Yiwei Zhao, Taku Komura, and Kazi Zaman. Local motion phases for learning multi-contact character movements. ACM Transactions on Graphics, 39(4): 54:1–54:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392450.

Smith:2023:MAC

[SZL^{+}23] Harrison Jesse Smith, Qingyuan Zheng, Yifei Li, Somya Jain, and Jessica K. Hodgins. A method for animating children’s drawings of the human figure. ACM Transactions on Graphics, 42(3): 32:1–32:??, June 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592788.

Sun:2010:LSG

[SZLG10] Xin Sun, Kun Zhou, Stephen Lin, and Baining Guo. Line space gathering for single scattering in large scenes. ACM Transactions on Graphics, 29 (4):54:1–54:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Sun:2008:IRD

[SZS^{+}08] Xin Sun, Kun Zhou, Eric Stollnitz, Jiaoying Shi, and Baining Guo. Interactive relighting of dynamic refractive objects. ACM Transactions on Graphics, 27(3): 35:1–35:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2007:MPC

[SZT^{+}07] Xiaohan Shi, Kun Zhou, Yiying Tong, Mathieu Desbrun, Hujun Bao, and Baining Guo. Mesh puppetry: cascading optimization of mesh deformation with inverse kinematics. ACM Transactions on Graphics, 26(3):81:1–81:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Shi:2008:EBD

[SZT^{+}08] Xiaohan Shi, Kun Zhou, Yiying Tong, Mathieu Desbrun, Hujun Bao, and Baining Guo. Example-based dynamic skinning in real time. ACM Transactions on Graphics, 27(3): 29:1–29:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Song:2023:GLU

[SZW^{+}23] Shuangbing Song, Fan Zhong, Tianju Wang, Xueying Qin, and Changhe Tu. Guided linear upsampling. ACM Transactions on Graphics, 42(4): 100:1–100:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592453.

Starke:2021:NAL

[SZZK21] Sebastian Starke, Yiwei Zhao, Fabio Zinno, and Taku Komura. Neural animation layering for synthesizing martial arts movements. ACM Transactions on Graphics, 40 (4):92:1–92:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459881.

Theobalt:2004:PBT

[TAH^{+}04] Christian Theobalt, Irene Albrecht, Jörg Haber, Marcus Magnor, and Hans-Peter Seidel. Pitching a baseball: tracking high-speed motion with multi-exposure images. ACM Transactions on Graphics, 23(3):540–547, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Talvala:2007:VGH

[TAHL07] Eino-Ville Talvala, Andrew Adams, Mark Horowitz, and Marc Levoy. Veiling glare in high dynamic range imaging. ACM Transactions on Graphics, 26(3):37:1–37:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Takayama:2022:CIT

[Tak22] Kenshi Takayama. Compatible intrinsic triangulations. ACM Transactions on Graphics, 41(4):57:1–57:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530175.

Tursun:2019:LCA

[TAKW^{+}19] Okan Tarhan Tursun, Elena Arabadzhiyska-Koleva, Marek Wernikowski, Radosław Mantiuk, Hans-Peter Seidel, Karol Myszkowski, and Piotr Didyk. Luminance-contrast-aware foveated rendering. ACM Transactions on Graphics, 38(4): 98:1–98:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tanner:1983:GEI

[Tan83]  Peter P. Tanner.  Guest

editor introduction. ACM Transactions on Graphics, 2 (2):89, April 1983. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tov:2021:DES

[TAN^{+}21] Omer Tov, Yuval Alaluf, Yotam Nitzan, Or Patashnik, and Daniel Cohen-Or. Designing an encoder for StyleGAN image manipulation. ACM Transactions on Graphics, 40(4): 133:1–133:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459838.

Tarini:2016:VEU

[Tar16]  Marco Tarini.  Volume-

encoded UV-maps. ACM Transactions on Graphics, 35 (4):107:1–107:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Taubin:1994:DAR

[Tau94]  Gabriel Taubin.  Discrete

approximations for rasterizing implicit curves. ACM Transactions on Graphics, 13(1):3–42, January 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
174531.html.

Taguchi:2010:ACM

[TAV^{+}10] Yuichi Taguchi, Amit Agrawal, Ashok Veeraraghavan, Srikumar Ramalingam, and Ramesh Raskar. Axial-cones: modeling spherical catadioptric cameras for wide-angle light field rendering. ACM Transactions on Graphics, 29(6): 172:1–172:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tso:1987:MRW

[TB87] Pauline Y. Ts’o and Brian A. Barsky. Modeling and rendering waves: Wave-tracing using beta-splines and reflective and refractive texture mapping. ACM Transactions on Graphics, 6(3):191–214, July 1987. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
35070.html.

Takahashi:2020:MMP

[TB20] Tetsuya Takahashi and Christopher Batty. Monolith: a monolithic pressure-viscosity-contact solver for strong two-way rigid-rigid rigid-fluid coupling. ACM Transactions on Graphics, 39(6): 182:1–182:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417798.

Takahashi:2021:FMO

[TB21] Tetsuya Takahashi and Christopher Batty. Frictional-Monolith: a monolithic optimization-based approach for granular flow with contactaware rigid-body coupling. ACM Transactions on Graphics, 40(6):206:1–206:20, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480539.

Takahashi:2022:EMO

[TB22] Tetsuya Takahashi and Christopher Batty. ElastoMonolith: a monolithic optimizationbased liquid solver for contactaware elastic-solid coupling. ACM Transactions on Graphics, 41(6):255:1–255:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555474.

Tao:2022:VPP

[TBBC^{+}22] Michael Tao, Christopher Batty, Mirela Ben-Chen, Eugene Fiume, and David I. W. Levin. VEMPIC: particle-in-polyhedron fluid simulation for intricate solid boundaries. ACM Transactions on Graphics, 41(4):115:1–115:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530138.

Taylor:2016:EPI

[TBC^{+}16] Jonathan Taylor, Lucas Bordeaux, Thomas Cashman, Bob Corish, Cem Keskin, Toby Sharp, Eduardo Soto, David Sweeney, Julien Valentin, Benjamin Luff, Arran Topalian, Erroll Wood, Sameh Khamis, Pushmeet Kohli, Shahram Izadi, Richard Banks, Andrew Fitzgibbon, and Jamie Shotton. Efficient and precise interactive hand tracking through joint, continuous optimization of pose and correspondences. ACM Transactions on Graphics, 35(4): 143:1–143:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tumanyan:2024:DSA

[TBTA^{+}24] Narek Tumanyan, Omer Bar-Tal, Shir Amir, Shai Bagon, and Tali Dekel. Disentangling structure and appearance in ViT feature space. ACM Transactions on Graphics, 43 (1):11:1–11:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3630096.

Tai:2008:TAR

[TBTS08] Yu-Wing Tai, Michael S.

Brown, Chi-Keung Tang, and Heung-Yeung Shum. Texture amendment: reducing texture distortion in constrained parameterization. ACM Transactions on Graphics, 27(5): 136:1–136:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tonge:2012:MSJ

[TBV12] Richard Tonge, Feodor Benevolenski, and Andrey Voroshilov. Mass splitting for jitter-free parallel rigid body simulation. ACM Transactions on Graphics, 31(4):105:1–105:8, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thorne:2004:MDI

[TBvdP04] Matthew Thorne, David Burke, and Michiel van de Panne. Motion doodles: an interface for sketching character motion. ACM Transactions on Graphics, 23(3): 424–431, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tevs:2012:ACI

[TBW^{+}12] Art Tevs, Alexander Berner, Michael Wand, Ivo Ihrke, Martin Bokeloh, Jens Kerber, and Hans-Peter Seidel. Animation cartography-intrinsic reconstruction of shape and motion. ACM Transactions on Graphics, 31(2):12:1– 12:15, April 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2016:IDD

[TBWP16] Chengcheng Tang, Pengbo Bo, Johannes Wallner, and Helmut Pottmann. Interactive design of developable surfaces. ACM Transactions on Graphics, 35(2):12:1–12:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thomaszewski:2014:CDL

[TCG^{+}14] Bernhard Thomaszewski, Stelian Coros, Damien Gauge, Vittorio Megaro, Eitan Grinspun, and Markus Gross. Computational design of linkage-based characters. ACM Transactions on Graphics, 33(4): 64:1–64:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Trusty:2021:SME

[TCL21] Ty Trusty, Honglin Chen, and David I. W. Levin. The shape matching element method: direct animation of curved surface models. ACM Transactions on Graphics, 40 (4):69:1–69:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459772.

Treuille:2006:CC

[TCP06] Adrien Treuille, Seth Cooper, and Zoran Popović. Continuum crowds. ACM Transactions on Graphics, 25(3): 1160–1168, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Trevithick:2023:RTR

[TCS^{+}23] Alex Trevithick, Matthew Chan, Michael Stengel, Eric Chan, Chao Liu, Zhiding Yu, Sameh Khamis, Manmohan Chandraker, Ravi Ramamoorthi, and Koki Nagano. Real-time radiance fields for single-image portrait view synthesis. ACM Transactions on Graphics, 42(4): 135:1–135:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592460.

Tang:2023:BCC

[TCT23] Pengbin Tang, Stelian Coros, and Bernhard Thomaszewski. Beyond Chainmail: Computational modeling of discrete interlocking materials. ACM Transactions on Graphics, 42 (4):84:1–84:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592112.

Tasse:2016:SSB

[TD16] Flora Ponjou Tasse and Neil Dodgson. Shape2Vec: semantic-based descriptors for 3D shapes, sketches and images. ACM Transactions on Graphics, 35(6):208:1– 208:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tursun:2023:PVM

[TD23] Cara Tursun and Piotr Didyk. Perceptual visibility model for temporal contrast changes in periphery. ACM Transactions on Graphics, 42(2): 20:1–20:??, April 2023. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3564241.

Toisoul:2018:ASV

[TDG18] Antoine Toisoul, Daljit Singh Dhillon, and Abhijeet Ghosh. Acquiring spatially varying appearance of printed holographic surfaces. ACM Transactions on Graphics, 37(6): 272:1–272:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2018:RTC

[TDL^{+}18] Danhang Tang, Mingsong Dou, Peter Lincoln, Philip Davidson, Kaiwen Guo, Jonathan Taylor, Sean Fanello, Cem Keskin, Adarsh Kowdle, Sofien Bouaziz, Shahram Izadi, and Andrea Tagliasacchi. Realtime compression and streaming of 4D performances. ACM Transactions on Graphics, 37 (6):256:1–256:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tena:2011:IRB

[TDM11] J. Rafael Tena, Fernando De la Torre, and Iain Matthews. Interactive region-based linear 3D face models. ACM Transactions on Graphics, 30(4): 76:1–76:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Templin:2014:MOE

[TDM^{+}14] Krzysztof Templin, Piotr Didyk, Karol Myszkowski, Mohamed M. Hefeeda, Hans-Peter Seidel, and Wojciech Matusik. Modeling and optimizing eye vergence response to stereoscopic cuts. ACM Transactions on Graphics, 33 (4):145:1–145:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Templin:2016:EDC

[TDMS16] Krzysztof Templin, Piotr Didyk, Karol Myszkowski, and Hans-Peter Seidel. Emulating displays with continuously varying frame rates. ACM Transactions on Graphics, 35(4):67:1–67:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Templin:2012:HMI

[TDR^{+}12] Krzysztof Templin, Piotr Didyk, Tobias Ritschel, Karol Myszkowski, and Hans-Peter Seidel. Highlight microdisparity for improved gloss depiction. ACM Transactions on Graphics, 31(4): 92:1–92:5, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2015:DTL

[TDSG15] Jianchao Tan, Marek Dvoroznák, Daniel Sýkora, and Yotam Gingold. Decomposing timelapse paintings into layers. ACM Transactions on Graphics, 34(4):61:1–61:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tanner:1982:R

[TE82] Peter P. Tanner and Kenneth B. Evans. The rack. ACM Transactions on Graphics, 1(2):186–188, April 1982. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2018:EPB

[TEG18] Jianchao Tan, Jose Echevarria, and Yotam Gingold. Efficient palette-based decomposition and recoloring of images via RGBXY-space geometry. ACM Transactions on Graphics, 37(6):262:1– 262:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

TenBosch:2020:DRB

[Ten20]  Marc Ten Bosch.  N-

dimensional rigid body dynamics. ACM Transactions on Graphics, 39(4): 55:1–55:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392483.

Terran:2018:SDO

[Ter18] Joseph Terran. Session details: Optimizing structures & materials. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tewari:2020:PPI

[TER^{+}20] Ayush Tewari, Mohamed Elgharib, Mallikarjun B R, Florian Bernard, Hans-Peter Seidel, Patrick Pérez, Michael Zollhöfer, and Christian Theobalt. PIE: portrait image embedding for semantic control. ACM Transactions on Graphics, 39(6): 223:1–223:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417803.

Tricard:2019:PPN

[TEZ^{+}19] Thibault Tricard, Semyon Efremov, Cédric Zanni, Fabrice Neyret, Jonàs Martínez, and Sylvain Lefebvre. Procedural phasor noise. ACM Transactions on Graphics, 38 (4):57:1–57:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Toler-Franklin:2010:MFM

[TFBW^{+}10] Corey Toler-Franklin, Benedict Brown, Tim Weyrich, Thomas Funkhouser, and Szymon Rusinkiewicz. Multifeature matching of fresco fragments. ACM Transactions on Graphics, 29(6): 185:1–185:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tonneau:2018:TPA

[TFD^{+}18] Steve Tonneau, Pierre Fernbach, Andrea Del Prete, Julien Pettré, and Nicolas Mansard. 2PAC: Two-point attractors for center of mass trajectories in multi-contact scenarios. ACM Transactions on Graphics, 37(5): 176:1–176:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3213773.

Tunwattanapong:2013:ARS

[TFG^{+}13] Borom Tunwattanapong, Graham Fyffe, Paul Graham, Jay Busch, Xueming Yu, Abhijeet Ghosh, and Paul Debevec. Acquiring reflectance and shape from continuous spherical harmonic illumination. ACM Transactions on Graphics, 32 (4):109:1–109:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsang:2003:BCS

[TFK^{+}03] Michael Tsang, George W.

Fitzmzurice, Gordon Kurtenbach, Azam Khan, and Bill Buxton. Boom chameleon: simultaneous capture of 3D viewpoint, voice and gesture annotations on a spatiallyaware display. ACM Transactions on Graphics, 22(3): 698, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2008:SIT

[TFX^{+}08] Ping Tan, Tian Fang, Jianxiong Xiao, Peng Zhao, and Long Quan. Single image tree modeling. ACM Transactions on Graphics, 27(5): 108:1–108:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Toisoul:2017:PARa

[TG17a] Antoine Toisoul and Abhijeet Ghosh. Practical acquisition and rendering of diffraction effects in surface reflectance. ACM Transactions on Graphics, 36(4): 64:1–64:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Toisoul:2017:PARb

[TG17b] Antoine Toisoul and Abhijeet Ghosh. Practical acquisition and rendering of diffraction effects in surface reflectance. ACM Transactions on Graphics, 36(5):166:1–166:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thiery:2013:SMS

[TGB13] Jean-Marc Thiery, Émilie Guy, and Tamy Boubekeur. Sphere-Meshes: shape approximation using spherical quadric error metrics. ACM Transactions on Graphics, 32 (6):178:1–178:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thiery:2016:AMA

[TGBE16] Jean-Marc Thiery, Émilie Guy, Tamy Boubekeur, and Elmar Eisemann. Animated mesh approximation with sphere-meshes. ACM Transactions on Graphics, 35(3): 30:1–30:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsingos:2004:PAR

[TGD04] Nicolas Tsingos, Emmanuel Gallo, and George Drettakis. Perceptual audio rendering of complex virtual environments. ACM Transactions on Graphics, 23(3):249–258, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tampubolon:2017:MSS

[TGK^{+}17] Andre Pradhana Tampubolon, Theodore Gast, Gergely Klár, Chuyuan Fu, Joseph Teran, Chenfanfu Jiang, and Ken Museth. Multi-species simulation of porous sand and water mixtures. ACM Transactions on Graphics, 36(4): 105:1–105:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2014:LBS

[TGLT14] Jie Tan, Yuting Gu, C. Karen Liu, and Greg Turk. Learning bicycle stunts. ACM Transactions on Graphics, 33(4):

50:1–50:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thomaszewski:2008:MM

[TGPS08] Bernhard Thomaszewski, Andreas Gumann, Simon Pabst, and Wolfgang Straßer. Magnets in motion. ACM Transactions on Graphics, 27(5): 162:1–162:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2011:ASC

[TGTL11] Jie Tan, Yuting Gu, Greg Turk, and C. Karen Liu. Articulated swimming creatures. ACM Transactions on Graphics, 30(4):58:1–58:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Talton:2009:EMC

[TGY^{+}09] Jerry O. Talton, Daniel Gibson, Lingfeng Yang, Pat Hanrahan, and Vladlen Koltun. Exploratory modeling with collaborative design spaces. ACM Transactions on Graphics, 28(5):167:1–167:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tymms:2018:QPM

[TGZ18] Chelsea Tymms, Esther P. Gardner, and Denis Zorin. A quantitative perceptual model for tactile roughness. ACM Transactions on Graphics, 37(5):168:1– 168:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3186267.

Tokuyoshi:2019:HRR

[TH19] Yusuke Tokuyoshi and Takahiro Harada. Hierarchical Russian roulette for vertex connections. ACM Transactions on Graphics, 38(4): 36:1–36:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tarini:2004:PM

[THCM04] Marco Tarini, Kai Hormann, Paolo Cignoni, and Claudio Montani. PolyCube-maps. ACM Transactions on Graphics, 23(3):853–860, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tumblin:1999:TMD

[THG99] Jack Tumblin, Jessica K. Hodgins, and Brian K. Guenter. Two methods for display of high contrast images. ACM Transactions on Graphics, 18(1):56–94, January 1999.

URL http://www.acm.org:
80/pubs/citations/journals/
tog/1999-18-1/p56-tumblin/

.

Tompkin:2013:CAL

[THKM13] James Tompkin, Simon Heinzle, Jan Kautz, and Wojciech Matusik. Content-adaptive lenticular prints. ACM Transactions on Graphics, 32(4): 133:1–133:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thuerey:2017:ISLa

[Thu17a]  Nils Thuerey.  Interpola-

tions of smoke and liquid simulations. ACM Transactions on Graphics, 36(1):3:1– 3:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thuerey:2017:ISLb

[Thu17b]  Nils Thuerey.  Interpola-

tions of smoke and liquid simulations. ACM Transactions on Graphics, 36(4): 68:1–68:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tevs:2014:RSG

[THW^{+}14] Art Tevs, Qixing Huang, Michael Wand, Hans-Peter Seidel, and Leonidas Guibas. Relating shapes via geometric symmetries and regularities. ACM Transactions on Graphics, 33(4):119:1–119:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Todo:2007:LCS

[TiABI07] Hideki Todo, Ken ichi Anjyo, William Baxter, and Takeo Igarashi. Locally controllable stylized shading. ACM Transactions on Graphics, 26(3): 17:1–17:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Takezawa:2016:FFO

[TISM16] Masahito Takezawa, Takuma Imai, Kentaro Shida, and Takashi Maekawa. Fabrication of freeform objects by principal strips. ACM Transactions on Graphics, 35(6): 225:1–225:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Twigg:2007:MWB

[TJ07] Christopher D. Twigg and Doug L. James. Many-worlds browsing for control of multibody dynamics. ACM Transactions on Graphics, 26(3): 14:1–14:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Twigg:2008:BSR

[TJ08] Christopher D. Twigg and Doug L. James. Backward steps in rigid body simulation. ACM Transactions on Graphics, 27(3):25:1–25:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tamstorf:2015:SAM

[TJM15] Rasmus Tamstorf, Toby Jones, and Stephen F. Mc-Cormick. Smoothed aggregation multigrid for cloth simulation. ACM Transactions on Graphics, 34(6): 245:1–245:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tak:2005:PBM

[TK05] Seyoon Tak and Hyeong-Seok Ko. A physically-based motion retargeting filter. ACM Transactions on Graphics, 24(1):98–117, January 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2014:IGP

[TK14] Min Tang and Young J. Kim. Interactive generalized penetration depth computation for rigid and articulated models using object norm. ACM Transactions on Graphics, 33 (1):1:1–1:15, January 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tabellion:2023:CLE

[TKG^{+}23] Eric Tabellion, Nikhil Karnad, Noa Glaser, Ben Weiss, David E. Jacobs, and Yael Pritch. Computational long exposure mobile photography. ACM Transactions on Graphics, 42(4):48:1–48:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592124.

Tompkin:2012:VES

[TKKT12] James Tompkin, Kwang In Kim, Jan Kautz, and Christian Theobalt. Videoscapes: exploring sparse, unstructured video collections. ACM Transactions on Graphics, 31(4):68:1–68:12, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tocci:2011:VHV

[TKTS11] Michael D. Tocci, Chris Kiser, Nora Tocci, and Pradeep Sen. A versatile HDR video production system. ACM Transactions on Graphics, 30(4): 41:1–41:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Taylor:2017:DLA

[TKY^{+}17] Sarah Taylor, Taehwan Kim, Yisong Yue, Moshe Mahler, James Krahe, Anastasio Garcia Rodriguez, Jessica Hodgins, and Iain Matthews. A deep learning approach for generalized speech animation. ACM Transactions on Graphics, 36(4):93:1–93:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tabellion:2004:AGI

[TL04] Eric Tabellion and Arnauld Lamorlette. An approximate global illumination system for computer generated films. ACM Transactions on Graphics, 23(3):469–476, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2017:DILa

[TLG17a] Jianchao Tan, Jyh-Ming Lien, and Yotam Gingold. Decomposing images into layers via RGB-space geometry. ACM Transactions on Graphics, 36 (1):7:1–7:??, February 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2017:DILb

[TLG17b] Jianchao Tan, Jyh-Ming Lien, and Yotam Gingold. Decomposing images into layers via RGB-space geometry. ACM Transactions on Graphics, 36 (4):61:1–61:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tong:2003:DMV

[TLHD03] Yiying Tong, Santiago Lombeyda, Anil N. Hirani, and Mathieu Desbrun. Discrete multiscale vector field decomposition. ACM Transactions on Graphics, 22(3):445–452, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thul:2018:ACD

[TLJP18] Daniel Thul, L’ubor Ladický, Sohyeon Jeong, and Marc Pollefeys. Approximate convex decomposition and transfer for animated meshes. ACM Transactions on Graphics, 37 (6):226:1–226:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2009:IHD

[TLK09] Min Tang, Minkyoung Lee, and Young J. Kim. Interactive Hausdorff distance computation for general polygonal models. ACM Transactions on Graphics, 28(3):74:1– 74:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Teng:2016:ESF

[TLK16] Yun Teng, David I. W. Levin, and Theodore Kim. Eulerian solid-fluid coupling. ACM Transactions on Graphics, 35 (6):200:1–200:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Talton:2011:MPM

[TLL^{+}11] Jerry O. Talton, Yu Lou, Steve Lesser, Jared Duke, Radomír Měch, and Vladlen Koltun. Metropolis procedural modeling. ACM Transactions on Graphics, 30(2): 11:1–11:14, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Treuille:2006:MRR

[TLP06] Adrien Treuille, Andrew Lewis, and Zoran Popović. Model reduction for real-time fluids. ACM Transactions on Graphics, 25(3):826–834, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Treuille:2007:NOC

[TLP07] Adrien Treuille, Yongjoon Lee, and Zoran Popović. Near-optimal character animation with continuous control. ACM Transactions on Graphics, 26(3):7:1–7:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tu:2024:UMF

[TLZ^{+}24] Zaili Tu, Chen Li, Zipeng Zhao, Long Liu, Chenhui Wang, Changbo Wang, and Hong Qin. A unified MPM framework supporting phasefield models and elasticviscoplastic phase transition. ACM Transactions on Graphics, 43(2):19:1–19:??, April 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3638047.

Tor:1984:CDS

[TM84] S. B. Tor and A. E. Middleditch. Convex decomposition of simple polygons. ACM Transactions on Graphics, 3(4): 244–265, October 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsoli:2014:BLS

[TMB14] Aggeliki Tsoli, Naureen Mahmood, and Michael J. Black. Breathing life into shape: capturing, modeling and animating 3D human breathing. ACM Transactions on Graphics, 33(4):52:1–52:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thiery:2018:MVC

[TMB18] Jean-Marc Thiery, Pooran Memari, and Tamy Boubekeur. Mean value coordinates for quad cages in 3D. ACM Transactions on Graphics, 37 (6):229:1–229:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Teng:2015:SCF

[TMDK15] Yun Teng, Mark Meyer, Tony DeRose, and Theodore Kim. Subspace condensation: full space adaptivity for subspace deformations. ACM Transactions on Graphics, 34(4): 76:1–76:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tseng:2021:DCO

[TMM^{+}21] Ethan Tseng, Ali Mosleh, Fahim Mannan, Karl St-Arnaud, Avinash Sharma, Yifan Peng, Alexander Braun, Derek Nowrouzezahrai, Jean-François Lalonde, and Felix Heide. Differentiable compound optics and processing pipeline optimization for endto-end camera design. ACM Transactions on Graphics, 40 (2):18:1–18:19, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3446791.

Tang:2012:CPF

[TMOT12] Min Tang, Dinesh Manocha, Miguel A. Otaduy, and Ruofeng Tong. Continuous penalty forces. ACM Transactions on Graphics, 31(4): 107:1–107:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Treuille:2003:KCS

[TMPS03] Adrien Treuille, Antoine Mc-Namara, Zoran Popović, and Jos Stam. Keyframe control of smoke simulations. ACM Transactions on Graphics, 22 (3):716–723, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tam:2014:DPR

[TMRL14] Gary K. L. Tam, Ralph R.

Martin, Paul L. Rosin, and Yu-Kun Lai. Diffusion pruning for rapidly and robustly selecting global correspondences using local isometry. ACM Transactions on Graphics, 33(1):4:1–4:17, January 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2011:VFC

[TMY^{+}11] Min Tang, Dinesh Manocha, Sung-Eui Yoon, Peng Du, Jae-Pil Heo, and Ruo-Feng Tong. VolCCD: Fast continuous collision culling between deforming volume meshes. ACM Transactions on Graphics, 30(5):111:1–111:15, October 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tournier:2015:SCD

[TNGF15] Maxime Tournier, Matthieu Nesme, Benjamin Gilles, and François Faure. Stable constrained dynamics. ACM Transactions on Graphics, 34(4):132:1–132:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Trettner:2022:EEM

[TNWK22] Philip Trettner, Julius Nehring-Wirxel, and Leif Kobbelt. EMBER: exact mesh booleans via efficient & robust local arrangements. ACM Transactions on Graphics, 41(4):39:1–39:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530181.

Turk:2002:MIS

[TO02] Greg Turk and James F. O’Brien. Modelling with implicit surfaces that interpolate. ACM Transactions on Graphics, 21(4):855– 873, October 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Teh:2022:ANR

[TOG22] Arjun Teh, Matthew O’Toole, and Ioannis Gkioulekas. Adjoint nonlinear ray tracing. ACM Transactions on Graphics, 41(4):126:1–126:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530077.

Takayama:2008:LST

[TOII08] Kenshi Takayama, Makoto Okabe, Takashi Ijiri, and Takeo Igarashi. Lapped solid textures: filling a model with anisotropic textures. ACM Transactions on Graphics, 27 (3):53:1–53:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Teng:2014:SAS

[TOK14] Yun Teng, Miguel A. Otaduy, and Theodore Kim. Simulating articulated subspace self-contact. ACM Transactions on Graphics, 33(4): 106:1–106:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Staff:2003:LR

[TOP03] TOPLAS Staff. List of reviewers. ACM Transactions on Graphics, 22(1):3, January 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsumura:2003:IBS

[TOS^{+}03] Norimichi Tsumura, Nobutoshi Ojima, Kayoko Sato, Mitsuhiro Shiraishi, Hideto Shimizu, Hirohide Nabeshima, Syuuichi Akazaki, Kimihiko Hori, and Yoichi Miyake. Image-based skin color and texture analysis/synthesis by extracting hemoglobin and melanin information in the skin. ACM Transactions on Graphics, 22(3):770–779, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tarini:2011:SQD

[TPP^{+}11] Marco Tarini, Enrico Puppo, Daniele Panozzo, Nico Pietroni, and Paolo Cignoni. Simple quad domains for field aligned mesh parametrization. ACM Transactions on Graphics, 30 (6):142:1–142:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Takayama:2013:SBG

[TPSHSH13] Kenshi Takayama, Daniele Panozzo, Alexander Sorkine-Hornung, and Olga Sorkine-Hornung. Sketch-based generation and editing of quad meshes. ACM Transactions on Graphics, 32(4): 97:1–97:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tkach:2016:SMR

[TPT16] Anastasia Tkach, Mark Pauly, and Andrea Tagliasacchi. Sphere-meshes for real-time hand modeling and tracking. ACM Transactions on Graphics, 35(6):222:1– 222:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tole:2002:IGI

[TPWG02] Parag Tole, Fabio Pellacini, Bruce Walter, and Donald P. Greenberg. Interactive global illumination in dynamic scenes. ACM Transactions on Graphics, 21(3): 537–546, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Terzopoulos:1994:DNG

[TQ94] Demetri Terzopoulos and Hong Qin. Dynamic NURBS with geometric constraints to interactive sculpting. ACM Transactions on Graphics, 13(2):103–136, April 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
176580.html.

Taubin:1998:GCT

[TR98] Gabriel Taubin and Jarek Rossignac. Geometric compression through topological surgery. ACM Transactions on Graphics, 17(2): 84–115, April 1998. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1998-17-2/
p84-taubin/.

Torres:2016:HRI

[TREO16] Rosell Torres, Alejandro Rodríiguez, José M. Espadero, and Miguel A. Otaduy. Highresolution interaction with corotational coarsening models. ACM Transactions on Graphics, 35(6):211:1–211:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsai:2006:AFP

[TS06] Yu-Ting Tsai and Zen-Chung Shih. All-frequency precomputed radiance transfer using spherical radial basis functions and clustered tensor approximation. ACM Transactions on Graphics, 25(3): 967–976, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thormahlen:2008:MOI

[TS08] Thorsten Thormählen and Hans-Peter Seidel. 3Dmodeling by ortho-image generation from image sequences. ACM Transactions on Graphics, 27(3):86:1–86:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsai:2012:CTA

[TS12] Yu-Ting Tsai and Zen-Chung Shih. K-clustered tensor approximation: a sparse multilinear model for real-time rendering. ACM Transactions on Graphics, 31(3):19:1– 19:17, May 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsai:2015:MCT

[Tsa15]  Yu-Ting Tsai.  Multi-

way K-clustered tensor approximation: Toward highperformance photorealistic data-driven rendering. ACM Transactions on Graphics, 34(5):157:1–157:??, October 2015. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2014:FFP

[TSG^{+}14] Chengcheng Tang, Xiang Sun, Alexandra Gomes, Johannes Wallner, and Helmut Pottmann. Form-finding with polyhedral meshes made simple. ACM Transactions on Graphics, 33(4): 70:1–70:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tsai:2016:SLS

[TSL^{+}16] Yi-Hsuan Tsai, Xiaohui Shen, Zhe Lin, Kalyan Sunkavalli, and Ming-Hsuan Yang. Sky is not the limit: semantic-aware sky replacement. ACM Transactions on Graphics, 35(4): 149:1–149:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tompson:2014:RTC

[TSLP14] Jonathan Tompson, Murphy Stein, Yann Lecun, and Ken Perlin. Real-time continuous pose recovery of human hands using convolutional networks. ACM Transactions on Graphics, 33(5): 169:1–169:??, August 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Takayama:2010:VMD

[TSNI10] Kenshi Takayama, Olga Sorkine, Andrew Nealen, and Takeo Igarashi. Volumetric modeling with diffusion surfaces. ACM Transactions on Graphics, 29(6):180:1– 180:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tzur:2009:FPT

[TT09] Yochay Tzur and Ayellet Tal. FlexiStickers: photogrammetric texture mapping using casual images. ACM Transactions on Graphics, 28(3): 45:1–45:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tariq:2022:NBE

[TTD22] Taimoor Tariq, Cara Tursun, and Piotr Didyk. Noise-based enhancement for foveated rendering. ACM Transactions on Graphics, 41(4):143:1– 143:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530101.

Tan:2012:SBL

[TTL12] Jie Tan, Greg Turk, and C. Karen Liu. Soft body locomotion. ACM Transactions on Graphics, 31(4): 26:1–26:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tkach:2017:OGM

[TTR^{+}17] Anastasia Tkach, Andrea Tagliasacchi, Edoardo Remelli, Mark Pauly, and Andrew Fitzgibbon. Online generative model personalization for hand tracking. ACM Transactions on Graphics, 36(6): 243:1–243:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Taylor:2017:ADF

[TTT^{+}17] Jonathan Taylor, Vladimir Tankovich, Danhang Tang, Cem Keskin, David Kim, Philip Davidson, Adarsh Kowdle, and Shahram Izadi. Articulated distance fields for ultra-fast tracking of hands interacting. ACM Transactions on Graphics, 36(6): 244:1–244:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2014:FEC

[TTWM14] Min Tang, Ruofeng Tong, Zhendong Wang, and Dinesh Manocha. Fast and exact continuous collision detection with Bernstein sign classification. ACM Transactions on Graphics, 33(6): 186:1–186:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tricard:2020:FOM

[TTZ^{+}20] Thibault Tricard, Vincent Tavernier, Cédric Zanni, Jonàs Martínez, Pierre-Alexandre Hugron, Fabrice Neyret, and Sylvain Lefebvre. Freely orientable microstructures for designing deformable 3D prints. ACM Transactions on Graphics, 39(6):211:1– 211:16, November 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417790.

Tongbuasirilai:2022:SNP

[TUGM22] Tanaboon Tongbuasirilai, Jonas Unger, Christine Guillemot, and Ehsan Miandji. A sparse non-parametric BRDF model. ACM Transactions on Graphics, 41(5):181:1–181:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3533427.

Turkowski:1982:AAT

[Tur82] K. Turkowski. Anti-aliasing through the use of coordinate transformations. ACM Transactions on Graphics, 1 (3):215–234, July 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thomas:2020:RPN

[TVLF20] Manu Mathew Thomas, Karthik Vaidyanathan, Gabor Liktor, and Angus G. Forbes. A reduced-precision network for image reconstruction. ACM Transactions on Graphics, 39 (6):231:1–231:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417786.

Tournois:2009:IDR

[TWAD09] Jane Tournois, Camille Wormser, Pierre Alliez, and Mathieu Desbrun. Interleaving Delaunay refinement and optimization for practical isotropic tetrahedron mesh generation. ACM Transactions on Graphics, 28(3):75:1–75:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tasdizen:2003:GSP

[TWBO03] Tolga Tasdizen, Ross Whitaker, Paul Burchard, and Stanley Osher. Geometric surface processing via normal maps. ACM Transactions on Graphics, 22(4):1012–1033, October 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thurey:2010:MAM

[TWGT10] Nils Thürey, Chris Wojtan, Markus Gross, and Greg Turk. A multiscale approach to mesh-based surface tension flows. ACM Transactions on Graphics, 29(4): 48:1–48:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2022:RTC

[TWH^{+}22] Xiangjun Tang, He Wang, Bo Hu, Xu Gong, Ruifan Yi, Qilong Kou, and Xiaogang Jin. Real-time controllable motion transition for characters. ACM Transactions on Graphics, 41 (4):137:1–137:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530090.

Tong:2005:MRQ

[TWL^{+}05] Xin Tong, Jiaping Wang, Stephen Lin, Baining Guo, and Heung-Yeung Shum. Modeling and rendering of quasi-homogeneous materials. ACM Transactions on Graphics, 24(3):1054–1061, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2018:CIC

[TWL^{+}18] Min Tang, Tongtong Wang, Zhongyuan Liu, Ruofeng Tong, and Dinesh Manocha. I-cloth: incremental collision handling for GPU-based interactive cloth simulation. ACM Transactions on Graphics, 37 (6):204:1–204:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2019:JSD

[TWLT19] Chengzhou Tang, Oliver Wang, Feng Liu, and Ping Tan. Joint stabilization and direction of 360^{∘} videos. ACM Transactions on Graphics, 38 (2):18:1–18:??, April 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3211889.

Tang:2023:HRV

[TWR^{+}23] Sicong Tang, Guangyuan Wang, Qing Ran, Lingzhi Li, Li Shen, and Ping Tan. High-resolution volumetric reconstruction for clothed humans. ACM Transactions on Graphics, 42(5): 170:1–170:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3606032.

Tu:2020:CCT

[TWY^{+}20] Peihan Tu, Li-Yi Wei, Koji Yatani, Takeo Igarashi, and Matthias Zwicker. Continuous curve textures. ACM Transactions on Graphics, 39(6): 168:1–168:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417780.

Tymms:2020:APT

[TWZ20] Chelsea Tymms, Siqi Wang, and Denis Zorin. Appearancepreserving tactile optimization. ACM Transactions on Graphics, 39(6):212:1– 212:16, November 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417857.

Tu:2022:CVT

[TWZ22] Peihan Tu, Li-Yi Wei, and Matthias Zwicker. Clustered vector textures. ACM Transactions on Graphics, 41 (4):159:1–159:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530062.

Tao:2009:SAB

[TYS09] Litian Tao, Lu Yuan, and Jian Sun. SkyFinder: attributebased sky image search. ACM Transactions on Graphics, 28 (3):68:1–68:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tseng:2019:HOB

[TYY^{+}19] Ethan Tseng, Felix Yu, Yuting Yang, Fahim Mannan, Karl St. Arnaud, Derek Nowrouzezahrai, Jean-François Lalonde, and Felix Heide. Hyperparameter optimization in black-box image processing using differentiable proxies. ACM Transactions on Graphics, 38(4):27:1–27:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tagliasacchi:2009:CSE

[TZCO09] Andrea Tagliasacchi, Hao Zhang, and Daniel Cohen-Or. Curve skeleton extraction from incomplete point cloud. ACM Transactions on Graphics, 28(3):71:1–71:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2020:HBA

[TZCT20] Pengbin Tang, Jonas Zehnder, Stelian Coros, and Bernhard Thomaszewski. A harmonic balance approach for designing compliant mechanical systems with nonlinear periodic motions. ACM Transactions on Graphics, 39(6): 191:1–191:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417765.

Tseng:2022:NPF

[TZJ^{+}22] Ethan Tseng, Yuxuan Zhang, Lars Jebe, Xuaner Zhang, Zhihao Xia, Yifei Fan, Felix Heide, and Jiawen Chen. Neural photo-finishing. ACM Transactions on Graphics, 41 (6):238:1–238:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555526.

Tautges:2011:MRU

[TZK^{+}11] Jochen Tautges, Arno Zinke, Björn Krüger, Jan Baumann, Andreas Weber, Thomas Helten, Meinard Müller, Hans-Peter Seidel, and Bernd Eberhardt. Motion reconstruction using sparse accelerometer data. ACM Transactions on Graphics, 30(3): 18:1–18:12, May 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tong:2002:SBT

[TZL^{+}02] Xin Tong, Jingdan Zhang, Ligang Liu, Xi Wang, Baining Guo, and Heung-Yeung Shum. Synthesis of bidirectional texture functions on arbitrary surfaces. ACM Transactions on Graphics, 21(3): 665–672, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thies:2015:RTE

[TZN^{+}15] Justus Thies, Michael Zollhöfer, Matthias Nießner, Levi Valgaerts, Marc Stamminger, and Christian Theobalt. Real-time expression transfer for facial reenactment. ACM Transactions on Graphics, 34(6): 183:1–183:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thies:2019:DNR

[TZN19] Justus Thies, Michael Zollhöfer, and Matthias Nießner. Deferred neural rendering: image synthesis using neural textures. ACM Transactions on Graphics, 38(4): 66:1–66:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thies:2018:FRT

[TZS^{+}18] Justus Thies, Michael Zollhöfer, Marc Stamminger, Christian Theobalt, and Matthias Nießner. FaceVR: Real-time gaze-aware facial reenactment in virtual reality. ACM Transactions on Graphics, 37(2): 25:1–25:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Thies:2018:HRT

[TZT^{+}18] Justus Thies, Michael Zollhöfer, Christian Theobalt, Marc Stamminger, and Matthias Niessner. Headon: real-time reenactment of human portrait videos. ACM Transactions on Graphics, 37(4): 164:1–164:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tan:2007:IBT

[TZW^{+}07] Ping Tan, Gang Zeng, Jingdong Wang, Sing Bing Kang, and Long Quan. Image-based tree modeling. ACM Transactions on Graphics, 26(3): 87:1–87:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Tang:2023:MFM

[TZY^{+}23] Yijie Tang, Jiazhao Zhang, Zhinan Yu, He Wang, and Kai Xu. MIPS-Fusion: Multiimplicit-submaps for scalable and robust online neural RGB-D reconstruction. ACM Transactions on Graphics, 42 (6):246:1–246:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618363.

Tozoni:2021:OCB

[TZZ21] Davi Colli Tozoni, Yunfan Zhou, and Denis Zorin. Optimizing contact-based assemblies. ACM Transactions on Graphics, 40(6): 269:1–269:19, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480552.

Umetani:2018:LTD

[UB18] Nobuyuki Umetani and Bernd Bickel. Learning threedimensional flow for interactive aerodynamic design. ACM Transactions on Graphics, 37(4):89:1–89:??, August 2018. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Ugail:1999:TID

[UBW99] Hassan Ugail, Malcolm I. G.

Bloor, and Michael J. Wilson. Techniques for interactive design using the PDE method. ACM Transactions on Graphics, 18(2): 195–212, April 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org:80/pubs/citations/
journals/tog/1999-18-2/
p195-ugail/.

Um:2017:PEL

[UHT17] Kiwon Um, Xiangyu Hu, and Nils Thuerey. Perceptual evaluation of liquid simulation methods. ACM Transactions on Graphics, 36(4): 143:1–143:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Umetani:2012:GEP

[UIM12] Nobuyuki Umetani, Takeo Igarashi, and Niloy J. Mitra. Guided exploration of physically valid shapes for furniture design. ACM Transactions on Graphics, 31(4): 86:1–86:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Umetani:2011:SCI

[UKIG11] Nobuyuki Umetani, Danny M. Kaufman, Takeo Igarashi, and Eitan Grinspun. Sensitive couture for interactive garment modeling and editing. ACM Transactions on Graphics, 30(4):90:1–90:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Umetani:2014:PID

[UKSI14] Nobuyuki Umetani, Yuki Koyama, Ryan Schmidt, and Takeo Igarashi. Pteromys: interactive design and optimization of free-formed free-flight model airplanes. ACM Transactions on Graphics, 33(4): 65:1–65:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Usai:2015:EQL

[ULP^{+}15] Francesco Usai, Marco Livesu, Enrico Puppo, Marco Tarini, and Riccardo Scateni. Extraction of the quad layout of a triangle mesh guided by its curve skeleton. ACM Transactions on Graphics, 35(1):6:1– 6:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ulu:2017:LSD

[UMK17] Erva Ulu, James Mccann, and Levent Burak Kara. Lightweight structure design under force location uncertainty. ACM Transactions on Graphics, 36(4):158:1– 158:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Umetani:2016:PIR

[UPSW16] Nobuyuki Umetani, Athina Panotopoulou, Ryan Schmidt, and Emily Whiting. Printone: interactive resonance simulation for free-form printwind instrument design. ACM Transactions on Graphics, 35 (6):184:1–184:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Uchytil:2024:FBA

[US24] Christopher Uchytil and Duane Storti. A functionbased approach to interactive high-precision volumetric design and fabrication. ACM Transactions on Graphics, 43 (1):3:1–3:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3622934.

Urban:2019:RRT

[UTB^{+}19] Philipp Urban, Tejas Madan Tanksale, Alan Brunton, Bui Minh Vu, and Shigeki Nakauchi. Redefining A in RGBA: Towards a standard for graphical 3D printing. ACM Transactions on Graphics, 38(3):21:1– 21:??, June 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3319910.

Upadhyay:2023:EDM

[UZB^{+}23] Rishi Upadhyay, Howard Zhang, Yunhao Ba, Ethan Yang, Blake Gella, Sicheng Jiang, Alex Wong, and Achuta Kadambi. Enhancing diffusion models with 3D perspective geometry constraints. ACM Transactions on Graphics, 42(6): 237:1–237:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618389.

Veenstra:1988:LDO

[VA88] Jack Veenstra and Narendra Ahuja. Line drawings of octree-represented objects. ACM Transactions on Graphics, 7(1):61–75, January 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
42189.html.

Vanegas:2009:IDU

[VABW09] Carlos A. Vanegas, Daniel G. Aliaga, Bedřich Beneš, and Paul A. Waddell. Interactive design of urban spaces using geometrical and behavioral modeling. ACM Transactions on Graphics, 28(5): 111:1–111:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Velazquez-Armendariz:2015:CLI

[VADWG15] Edgar Velázquez-Armendáriz, Zhao Dong, Bruce Walter, and Donald P. Greenberg. Complex luminaires: Illumination and appearance rendering. ACM Transactions on Graphics, 34(3): 26:1–26:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

VanWyk:1982:HLL

[Van82] Christopher J. Van Wyk. A high-level language for specifying pictures. ACM Transactions on Graphics, 1(2): 163–182, April 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

VanHateren:2006:EHD

[Van06] J. H. Van Hateren. Encoding of high dynamic range video with a model of human cones. ACM Transactions on Graphics, 25(4):1380–1399, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vlasic:2007:PMC

[VAV^{+}07] Daniel Vlasic, Rolf Adelsberger, Giovanni Vannucci, John Barnwell, Markus Gross, Wojciech Matusik, and Jovan Popović. Practical motion capture in everyday surroundings. ACM Transactions on Graphics, 26(3): 35:1–35:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Velazquez-Armendariz:2009:ABP

[VAZH^{+}09] Edgar Velázquez-Armendáriz, Shuang Zhao, Miloš Hašan, Bruce Walter, and Kavita Bala. Automatic bounding of programmable shaders for efficient global illumination. ACM Transactions on Graphics, 28(5):142:1– 142:9, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vanraes:2006:TSS

[VB06] Evelyne Vanraes and Adhemar Bultheel. A tangent subdivision scheme. ACM Transactions on Graphics, 25(2): 340–355, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vergne:2016:FGW

[VBBF16] Romain Vergne, Pascal Barla, Georges-Pierre Bonneau, and Roland W. Fleming. Flowguided warping for imagebased shape manipulation. ACM Transactions on Graphics, 35(4):93:1–93:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaxman:2010:MRA

[VBCG10] Amir Vaxman, Mirela Ben-Chen, and Craig Gotsman. A multi-resolution approach to heat kernels on discrete surfaces. ACM Transactions on Graphics, 29(4): 121:1–121:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vergne:2012:SFI

[VBFG12] Romain Vergne, Pascal Barla, Roland W. Fleming, and Xavier Granier. Surface flows for image-based shading design. ACM Transactions on Graphics, 31(4): 94:1–94:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaillant:2013:ISR

[VBG^{+}13] Rodolphe Vaillant, Loïc Barthe, Gaël Guennebaud, Marie-Paule Cani, Damien Rohmer, Brian Wyvill, Olivier Gourmel, and Mathias Paulin. Implicit skinning: real-time skin deformation with contact modeling. ACM Transactions on Graphics, 32(4): 125:1–125:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vedula:2005:IBS

[VBK05] Sundar Vedula, Simon Baker, and Takeo Kanade. Imagebased spatio-temporal modeling and view interpolation of dynamic events. ACM Transactions on Graphics, 24(2): 240–261, April 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vlasic:2008:AMA

[VBMP08] Daniel Vlasic, Ilya Baran, Wojciech Matusik, and Jovan Popović. Articulated mesh animation from multiview silhouettes. ACM Transactions on Graphics, 27(3): 97:1–97:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vlasic:2005:FTM

[VBPP05] Daniel Vlasic, Matthew Brand, Hanspeter Pfister, and Jovan Popović. Face transfer with multilinear models. ACM Transactions on Graphics, 24 (3):426–433, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vilesov:2022:BCG

[VCA^{+}22] Alexander Vilesov, Pradyumna Chari, Adnan Armouti, Anirudh Bindiganavale Harish, Kimaya Kulkarni, Ananya Deoghare, Laleh Jalilian, and Achuta Kadambi. Blending camera and 77 GHz radar sensing for equitable, robust plethysmography. ACM Transactions on Graphics, 41(4): 36:1–36:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530161.

Velho:1999:UAH

[VdFG99] Luiz Velho, Luiz Henrique de Figueiredo, and Jonas Gomes. A unified approach for hierarchical adaptive tesselation of surfaces. ACM Transactions on Graphics, 18(4): 329–360, October 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tog/1999-18-4/
p329-velho/.

vandenHengel:2007:VRI

[vdHDT^{+}07] Anton van den Hengel, Anthony Dick, Thorsten Thormählen, Ben Ward, and Philip H. S. Torr. Video-Trace: rapid interactive scene modelling from video. ACM Transactions on Graphics, 26 (3):86:1–86:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

VanKaick:2014:SSA

[VFK^{+}14] Oliver Van Kaick, Noa Fish, Yanir Kleiman, Shmuel Asafi, and Daniel Cohen-Or. Shape segmentation by approximate convexity analysis. ACM Transactions on Graphics, 34 (1):4:1–4:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

vonFunck:2006:VFB

[vFTS06] Wolfram von Funck, Holger Theisel, and Hans-Peter Seidel. Vector field based shape deformations. ACM Transactions on Graphics, 25(3): 1118–1125, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaillant:2014:RIS

[VGB^{+}14] Rodolphe Vaillant, Gäel Guennebaud, Loïc Barthe, Brian Wyvill, and Marie-Paule Cani. Robust isosurface tracking for interactive character skinning. ACM Transactions on Graphics, 33 (6):189:1–189:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vanegas:2012:IDU

[VGDA^{+}12] Carlos A. Vanegas, Ignacio Garcia-Dorado, Daniel G. Aliaga, Bedrich Benes, and Paul Waddell. Inverse design of urban procedural models. ACM Transactions on Graphics, 31(6):168:1– 168:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vouga:2012:DSS

[VHWP12] Etienne Vouga, Mathias Höbinger, Johannes Wallner, and Helmut Pottmann. Design of self-supporting surfaces. ACM Transactions on Graphics, 31(4):87:1– 87:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Verschoor:2019:EAC

[VJ19] Mickeal Verschoor and Andrei C. Jalba. Efficient and accurate collision response for elastically deformable models. ACM Transactions on Graphics, 38 (2):17:1–17:??, April 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3209887.

Vicini:2021:NET

[VJK21] Delio Vicini, Wenzel Jakob, and Anton Kaplanyan. A nonexponential transmittance model for volumetric scene representations. ACM Transactions on Graphics, 40(4): 136:1–136:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459815.

Vorba:2016:ADR

[VK16] Jirí Vorba and Jaroslav Krivánek. Adjoint-driven Russian roulette and splitting in light transport simulation. ACM Transactions on Graphics, 35(4):42:1–42:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Valentin:2018:DMS

[VKB^{+}18] Julien Valentin, Adarsh Kowdle, Jonathan T. Barron, Neal Wadhwa, Max Dzitsiuk, Michael Schoenberg, Vivek Verma, Ambrus Csaszar, Eric Turner, Ivan Dryanovski, Joao Afonso, Jose Pascoal, Konstantine Tsotsos, Mira Leung, Mirko Schmidt, Onur Guleryuz, Sameh Khamis, Vladimir Tankovitch, Sean Fanello, Shahram Izadi, and Christoph Rhemann. Depth from motion for smartphone AR. ACM Transactions on Graphics, 37(6):193:1– 193:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Viitanen:2017:MFH

[VKJ^{+}17] Timo Viitanen, Matias Koskela, Pekka Jääskeläinen, Heikki Kultala, and Jarmo Takala. MergeTree: a fast hardware HLBVH constructor for animated ray tracing. ACM Transactions on Graphics, 36(5):169:1–169:??, October 2017. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Vicini:2019:LSA

[VKJ19] Delio Vicini, Vladlen Koltun, and Wenzel Jakob. A learned shape-adaptive subsurface scattering model. ACM Transactions on Graphics, 38 (4):127:1–127:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vevoda:2018:BOR

[VKK18] Petr Vévoda, Ivo Kondapaneni, and Jaroslav Krivánek. Bayesian online regression for adaptive direct illumination sampling. ACM Transactions on Graphics, 37(4): 125:1–125:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Valevski:2023:UTD

[VKM^{+}23] Dani Valevski, Matan Kalman, Eyal Molad, Eyal Segalis, Yossi Matias, and Yaniv Leviathan. UniTune: Textdriven image editing by fine tuning a diffusion model on a single image. ACM Transactions on Graphics, 42(4): 128:1–128:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592451.

Vorba:2014:LLP

[VKS^{+}14] Jirí Vorba, Ondrej Karlík, Martin Sik, Tobias Ritschel, and Jaroslav Krivánek. Online learning of parametric mixture models for light transport simulation. ACM Transactions on Graphics, 33 (4):101:1–101:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Voglreiter:2023:TRO

[VKW^{+}23] Philip Voglreiter, Bernhard Kerbl, Alexander Weinrauch, Joerg Hermann Mueller, Thomas Neff, Markus Steinberger, and Dieter Schmalstieg. Trim regions for online computation of from-region potentially visible sets. ACM Transactions on Graphics, 42 (4):85:1–85:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592434.

vanKaick:2013:CHA

[vKXZ^{+}13] Oliver van Kaick, Kai Xu, Hao Zhang, Yanzhen Wang, Shuyang Sun, Ariel Shamir, and Daniel Cohen-Or. Cohierarchical analysis of shape structures. ACM Transactions on Graphics, 32(4): 69:1–69:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Verdie:2015:LGU

[VLA15] Yannick Verdie, Florent Lafarge, and Pierre Alliez. LOD generation for urban scenes. ACM Transactions on Graphics, 34(3):30:1–30:??, April 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vangorp:2007:ISP

[VLD07] Peter Vangorp, Jurgen Laurijssen, and Philip Dutré. The influence of shape on the perception of material reflectance. ACM Transactions on Graphics, 26(3): 77:1–77:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Venkataraman:2013:PUT

[VLD^{+}13] Kartik Venkataraman, Dan Lelescu, Jacques Duparré, Andrew McMahon, Gabriel Molina, Priyam Chatterjee, Robert Mullis, and Shree Nayar. PiCam: an ultra-thin high performance monolithic camera array. ACM Transactions on Graphics, 32(6): 166:1–166:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Mossel:2021:SJP

[VLV^{+}21] Dave Pagurek Van Mossel, Chenxi Liu, Nicholas Vining, Mikhail Bessmeltsev, and Alla Sheffer. StrokeStrip: joint parameterization and fitting of stroke clusters. ACM Transactions on Graphics, 40 (4):50:1–50:18, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459777.

Vaidyanathan:2015:LLF

[VMCS15] Karthik Vaidyanathan, Jacob Munkberg, Petrik Clarberg, and Marco Salvi. Layered light field reconstruction for defocus blur. ACM Transactions on Graphics, 34(2):23:1– 23:??, February 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vangorp:2015:MLA

[VMGM15] Peter Vangorp, Karol Myszkowski, Erich W. Graf, and Rafał K. Mantiuk. A model of local adaptation. ACM Transactions on Graphics, 34(6): 166:1–166:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Volevich:2000:UVD

[VMKK00] Valdimir Volevich, Karol Myszkowski, Andrei Khodulev, and Edward A. Kopylov. Using the visual differences predictor to improve performance of progressive global illumination computation. ACM Transactions on Graphics, 19(2):122–161, April 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
articles/journals/tog/2000-
19-2/p122-volevich/p122-
volevich.pdf;   http://
www.acm.org/pubs/citations/
journals/tog/2000-19-2/
p122-volevich/.

Volino:2006:RSC

[VMT06] Pascal Volino and Nadia Magnenat-Thalmann. Resolving surface collisions through intersection contour minimization. ACM Transactions on Graphics, 25(3): 1154–1159, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Volino:2009:SAN

[VMTF09] Pascal Volino, Nadia Magnenat-Thalmann, and François Faure. A simple approach to nonlinear tensile stiffness for accurate cloth simulation. ACM Transactions on Graphics, 28(4):105:1– 105:16, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaxman:2015:CMD

[VMW15] Amir Vaxman, Christian Müller, and Ofir Weber. Conformal mesh deformations with Möbius transformations. ACM Transactions on Graphics, 34(4):55:1–55:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaxman:2017:RMP

[VMW17] Amir Vaxman, Christian Müller, and Ofir Weber. Regular meshes from polygonal patterns. ACM Transactions on Graphics, 36(4): 113:1–113:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaxman:2018:CMS

[VMW18] Amir Vaxman, Christian Müller, and Ofir Weber. Canonical Möbius subdivision. ACM Transactions on Graphics, 37(6):227:1– 227:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

VanAken:1985:CDA

[VN85] J. Van Aken and M. Novak. Curve-drawing algorithms for raster displays. ACM Transactions on Graphics, 4(2): 147–169, April 1985. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

vanOverveld:1996:SSD

[vOV96] C. W. A. M. van Overveld and Marie Luce Viaud. Sticky splines: Definition and manipulation of spline structures with maintained topological relations. ACM Transactions on Graphics, 15(1): 72–98, January 1996. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
225889.html;   http://www.
acm.org/pubs/toc/Abstracts/
0730-0301/226154.html.

Vergne:2009:LWE

[VPB^{+}09a] Romain Vergne, Romain Pacanowski, Pascal Barla, Xavier Granier, and Christophe Schlick. Light warping for enhanced surface depiction. ACM Transactions on Graphics, 28(3):25:1–25:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vlasic:2009:DSC

[VPB^{+}09b] Daniel Vlasic, Pieter Peers, Ilya Baran, Paul Debevec, Jovan Popović, Szymon Rusinkiewicz, and Wojciech Matusik. Dynamic shape capture using multi-view photometric stereo. ACM Transactions on Graphics, 28(5):174:1– 174:11, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Velinov:2018:ACM

[VPB^{+}18] Zdravko Velinov, Marios Papas, Derek Bradley, Paulo Gotardo, Parsa Mirdehghan, Steve Marschner, Jan Novák, and Thabo Beeler. Appearance capture and modeling of human teeth. ACM Transactions on Graphics, 37(6): 207:1–207:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vinker:2022:CSA

[VPB^{+}22] Yael Vinker, Ehsan Pajouheshgar, Jessica Y. Bo, Roman Christian Bachmann, Amit Haim Bermano, Daniel Cohen-Or, Amir Zamir, and Ariel Shamir. CLIPasso: semantically-aware object sketching. ACM Transactions on Graphics, 41(4): 86:1–86:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530068.

Valle-Perez:2021:TPA

[VPHB^{+}21] Guillermo Valle-Pérez, Gustav Eje Henter, Jonas Beskow, Andre Holzapfel, Pierre-Yves Oudeyer, and Simon Alexanderson. Transflower: probabilistic autoregressive dance generation with multimodal attention. ACM Transactions on Graphics, 40(6): 195:1–195:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480570.

Verhetsel:2019:FHS

[VPR19] Kilian Verhetsel, Jeanne Pellerin, and Jean-François Remacle. Finding hexahedrizations for small quadrangulations of the sphere. ACM Transactions on Graphics, 38(4):53:1–53:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vemuri:1994:MSH

[VR94] B. C. Vemuri and A. Radisavljevic. Multiresolution stochastic hybrid shape models with fractal priors. ACM Transactions on Graphics, 13(2):177–207, April 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
176583.html.

Veeraraghavan:2007:DPM

[VRA^{+}07] Ashok Veeraraghavan, Ramesh Raskar, Amit Agrawal, Ankit Mohan, and Jack Tumblin. Dappled photography: mask enhanced cameras for heterodyned light fields and coded aperture refocusing. ACM Transactions on Graphics, 26 (3):69:1–69:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vantzos:2018:RTV

[VRBC18] Orestis Vantzos, Saar Raz, and Mirela Ben-Chen. Realtime viscous thin films. ACM Transactions on Graphics, 37 (6):281:1–281:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vangorp:2013:PPD

[VRC^{+}13] Peter Vangorp, Christian Richardt, Emily A. Cooper, Gaurav Chaurasia, Martin S. Banks, and George Drettakis. Perception of perspective distortions in imagebased rendering. ACM Transactions on Graphics, 32(4): 58:1–58:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vogels:2018:DKP

[VRM^{+}18] Thijs Vogels, Fabrice Rousselle, Brian Mcwilliams, Gerhard Röthlin, Alex Harvill, David Adler, Mark Meyer, and Jan Novák. Denoising with kernel prediction and asymmetric loss functions. ACM Transactions on Graphics, 37(4):124:1– 124:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vidulis:2023:CEM

[VRP^{+}23] Michele Vidulis, Yingying Ren, Julian Panetta, Eitan Grinspun, and Mark Pauly. Computational exploration of multistable elastic knots. ACM Transactions on Graphics, 42(4): 73:1–73:??, August 2023.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592399.

Vondrak:2012:VBM

[VSHJ12] Marek Vondrak, Leonid Sigal, Jessica Hodgins, and Odest Jenkins. Video-based 3D motion capture through biped control. ACM Transactions on Graphics, 31(4): 27:1–27:12, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vicini:2021:PRB

[VSJ21] Delio Vicini, Sébastien Speierer, and Wenzel Jakob. Path replay backpropagation: differentiating light paths using constant memory and linear time. ACM Transactions on Graphics, 40(4): 108:1–108:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459804.

Vicini:2022:DSD

[VSJ22] Delio Vicini, Sébastien Speierer, and Wenzel Jakob. Differentiable signed distance function rendering. ACM Transactions on Graphics, 41 (4):125:1–125:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530139.

Vouga:2017:AWE

[VSK^{+}17] Etienne Vouga, Breannan Smith, Danny M. Kaufman, Rasmus Tamstorf, and Eitan Grinspun. All’s well that ends well: guaranteed resolution of simultaneous rigid body impact. ACM Transactions on Graphics, 36(4): 151:1–151:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vanhoey:2013:FMS

[VSLD13] Kenneth Vanhoey, Basile Sauvage, Frédéric Larue, and Jean-Michel Dischler. Onthe-fly multi-scale infinite texturing from example. ACM Transactions on Graphics, 32 (6):208:1–208:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vaidyanathan:2023:RAN

[VSW^{+}23] Karthik Vaidyanathan, Marco Salvi, Bartlomiej Wronski, Tomas Akenine-Moller, Pontus Ebelin, and Aaron Lefohn. Random-access neural compression of material textures. ACM Transactions on Graphics, 42(4):88:1–88:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592407.

Vasilescu:2004:TMI

[VT04] M. Alex O. Vasilescu and Demetri Terzopoulos. Tensor-Textures: multilinear imagebased rendering. ACM Transactions on Graphics, 23(3): 336–342, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

vonTycowicz:2013:ECR

[vTSSH13] Christoph von Tycowicz, Christian Schulz, Hans-Peter Seidel, and Klaus Hildebrandt. An efficient construction of reduced deformable objects. ACM Transactions on Graphics, 32(6):213:1– 213:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Von-Tycowicz:2015:RTN

[VTSSH15] Christoph Von-Tycowicz, Christian Schulz, Hans-Peter Seidel, and Klaus Hildebrandt. Real-time nonlinear shape interpolation. ACM Transactions on Graphics, 34(3): 34:1–34:??, April 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Valentin:2015:SIL

[VVC^{+}15] Julien Valentin, Vibhav Vineet, Ming-Ming Cheng, David Kim, Jamie Shotton, Pushmeet Kohli, Matthias Nießner, Antonio Criminisi, Shahram Izadi, and Philip Torr. SemanticPaint: Interactive 3D labeling and learning at your fingertips. ACM Transactions on Graphics, 34(5):154:1–154:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vinker:2023:CDV

[VVCOS23] Yael Vinker, Andrey Voynov, Daniel Cohen-Or, and Ariel Shamir. Concept decomposition for visual exploration and inspiration. ACM Transactions on Graphics, 42(6): 241:1–241:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618315.

Verhoeven:2022:DPQ

[VVHSH22] Floor Verhoeven, Amir Vaxman, Tim Hoffmann, and Olga Sorkine-Hornung. Dev2PQ: Planar quadrilateral strip remeshing of developable surfaces. ACM Transactions on Graphics, 41(3): 29:1–29:18, June 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3510002.

vanWijk:1984:RTO

[vW84] Jarke J. van Wijk. Ray tracing objects defined by sweeping planar cubic splines. ACM Transactions on Graphics, 3 (3):223–237, July 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

VanGelder:1994:TCI

[VW94] Allen Van Gelder and Jane Wilhelms. Topological considerations in isosurface generation. ACM Transactions on Graphics, 13(4):337– 375, October 1994. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/

195828.html. See corrigendum: [VW95].

VanGelder:1995:CTC

[VW95] Allen Van Gelder and Jane Wilhelms. Corrigendum: “Topological Considerations in Isosurface Generation”. ACM Transactions on Graphics, 14(3):307–308, July 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
215264.html. See [VW94].

VanOverveld:1997:PNI

[VW97] C. W. A. M. Van Overveld and B. Wyvill. Phong normal interpolation revisited. ACM Transactions on Graphics, 16(4):397–419, October 1997. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/tog/1997-16-4/
                                        p397-van_overveld/.

vanWijk:2002:IBF

[vW02] Jarke J. van Wijk. Image based flow visualization. ACM Transactions on Graphics, 21 (3):745–754, July 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

vanWijk:2009:STC

[vW09] Jarke J. van Wijk. Symmetric tiling of closed surfaces: visualization of regular maps. ACM Transactions on Graphics, 28(3):49:1–49:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Valgaerts:2012:LBF

[VWB^{+}12] Levi Valgaerts, Chenglei Wu, Andrés Bruhn, Hans-Peter Seidel, and Christian Theobalt. Lightweight binocular facial performance capture under uncontrolled lighting. ACM Transactions on Graphics, 31(6):187:1– 187:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Velten:2013:FPC

[VWJ^{+}13] Andreas Velten, Di Wu, Adrian Jarabo, Belen Masia, Christopher Barsi, Chinmaya Joshi, Everett Lawson, Moungi Bawendi, Diego Gutierrez, and Ramesh Raskar. Femto-photography: capturing and visualizing the propagation of light. ACM Transactions on Graphics, 32(4): 44:1–44:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vidimce:2013:OPP

[VWRKM13] Kiril Vidimce, Szu-Po Wang, Jonathan Ragan-Kelley, and Wojciech Matusik. Open-Fab: a programmable pipeline for multi-material fabrication. ACM Transactions on Graphics, 32(4):136:1–136:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Vekhter:2019:WGF

[VZF^{+}19] Josh Vekhter, Jiacheng Zhuo, Luisa F. Gil Fandino, Qixing Huang, and Etienne Vouga. Weaving geodesic foliations. ACM Transactions on Graphics, 38(4):34:1–34:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Worchel:2023:DRP

[WA23] Markus Worchel and Marc Alexa. Differentiable rendering of parametric geometry. ACM Transactions on Graphics, 42(6): 232:1–232:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618387.

Walter:2006:ML

[WABG06] Bruce Walter, Adam Arbree, Kavita Bala, and Donald P. Greenberg. Multidimensional lightcuts. ACM Transactions on Graphics, 25 (3):1081–1088, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2007:SSI

[WAC07] Jue Wang, Maneesh Agrawala, and Michael F. Cohen. Soft scissors: an interactive tool for realtime high quality matting. ACM Transactions on Graphics, 26(3):9:1–9:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wampler:2010:CAT

[WAH^{+}10] Kevin Wampler, Erik Andersen, Evan Herbst, Yongjoon Lee, and Zoran Popović. Character animation in twoplayer adversarial games. ACM Transactions on Graphics, 29(3):26:1–26:13, June 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Winchenbach:2020:SAB

[WAK20] Rene Winchenbach, Rustam Akhunov, and Andreas Kolb. Semi-analytic boundary handling below particle resolution for smoothed particle hydrodynamics. ACM Transactions on Graphics, 39(6): 173:1–173:17, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417829.

Wills:2009:TPS

[WAKB09] Josh Wills, Sameer Agarwal, David Kriegman, and Serge Belongie. Toward a perceptual space for gloss. ACM Transactions on Graphics, 28(4): 103:1–103:15, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Welsh:2002:TCG

[WAM02] Tomihisa Welsh, Michael Ashikhmin, and Klaus Mueller. Transferring color to greyscale images. ACM Transactions on Graphics, 21(3):277–280, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wampler:2016:FRE

[Wam16] Kevin Wampler. Fast and reliable example-based mesh IK for stylized deformations. ACM Transactions on Graphics, 35(6):235:1– 235:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2014:DCC

[Wan14] Huamin Wang. Defending continuous collision detection against errors. ACM Transactions on Graphics, 33(4): 122:1–122:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2015:CSI

[Wan15] Huamin Wang. A Chebyshev semi-iterative approach for accelerating projective and position-based dynamics. ACM Transactions on Graphics, 34(6):246:1–246:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2018:RFS

[Wan18a]  Huamin Wang.  Rule-free

sewing pattern adjustment with precision and efficiency. ACM Transactions on Graphics, 37(4):53:1–53:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2018:SDM

[Wan18b] Oliver Wang. Session details: Mixed reality. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:GBS

[Wan21] Huamin Wang. GPU-based simulation of cloth wrinkles at submillimeter levels. ACM Transactions on Graphics, 40(4):169:1–169:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3450626.3459787.

Wang:2023:OOB

[Wan23]  Peng-Shuai Wang.  Oct-

Former: Octree-based transformers for 3D point clouds. ACM Transactions on Graphics, 42(4):155:1–155:??, August 2023. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592131.

Wand:2009:ERN

[WAO^{+}09] Michael Wand, Bart Adams, Maksim Ovsjanikov, Alexander Berner, Martin Bokeloh, Philipp Jenke, Leonidas Guibas, Hans-Peter Seidel, and Andreas Schilling. Efficient reconstruction of nonrigid shape and motion from real-time 3D scanner data. ACM Transactions on Graphics, 28(2):15:1–15:15, April 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Warren:1989:BAS

[War89]  J. Warren.  Blending al-

gebraic surfaces. ACM Transactions on Graphics, 8 (4):263–278, October 1989. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
77270.html.

Warren:1992:CMR

[War92] Joe Warren. Creating multisided rational Bézier surfaces using base points. ACM Transactions on Graphics, 11(2):127–139, April 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
130828.html.

Wang:2012:ACA

[WAvK^{+}12] Yunhai Wang, Shmulik Asafi, Oliver van Kaick, Hao Zhang, Daniel Cohen-Or, and Baoquan Chen. Active co-analysis of a set of shapes. ACM Transactions on Graphics, 31 (6):165:1–165:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2008:SEL

[WB08] Jing Wang and Bobby Bodenheimer. Synthesis and evaluation of linear motion transitions. ACM Transactions on Graphics, 27(1): 1:1–1:22, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wen:2023:KLS

[WB23] Jiahao Wen and Jernej Barbic. Kirchhoff–Love shells with arbitrary hyperelastic materials. ACM Transactions on Graphics, 42(6): 174:1–174:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618405.

Wang:2005:IVC

[WBC^{+}05] Jue Wang, Pravin Bhat, R. Alex Colburn, Maneesh Agrawala, and Michael F. Cohen. Interactive video cutout. ACM Transactions on Graphics, 24(3):585–594, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2019:SSG

[WBCPS19] Yu Wang, Mirela Ben-Chen, Iosif Polterovich, and Justin Solomon. Steklov spectral geometry for extrinsic shape analysis. ACM Transactions on Graphics, 38(1):7:1– 7:??, February 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3152156.

Waechter:2017:VRNa

[WBF^{+}17a] Michael Waechter, Mate Beljan, Simon Fuhrmann, Nils Moehrle, Johannes Kopf, and Michael Goesele. Virtual rephotography: Novel view prediction error for 3D reconstruction. ACM Transactions on Graphics, 36(1):8:1– 8:??, February 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Waechter:2017:VRNb

[WBF^{+}17b] Michael Waechter, Mate Beljan, Simon Fuhrmann, Nils Moehrle, Johannes Kopf, and Michael Goesele. Virtual rephotography: novel view prediction error for 3D reconstruction. ACM Transactions on Graphics, 36(4): 45:1–45:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2016:MBT

[WBG^{+}16] Chenglei Wu, Derek Bradley, Pablo Garrido, Michael Zollhöfer, Christian Theobalt, Markus Gross, and Thabo Beeler. Model-based teeth reconstruction. ACM Transactions on Graphics, 35(6):220:1– 220:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2016:ACL

[WBGB16] Chenglei Wu, Derek Bradley, Markus Gross, and Thabo Beeler. An anatomicallyconstrained local deformation model for monocular face capture. ACM Transactions on Graphics, 35(4): 115:1–115:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weise:2011:RPB

[WBLP11] Thibaut Weise, Sofien Bouaziz, Hao Li, and Mark Pauly. Realtime performance-based facial animation. ACM Transactions on Graphics, 30(4): 77:1–77:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wald:2007:RTD

[WBS07] Ingo Wald, Solomon Boulos, and Peter Shirley. Ray tracing deformable scenes using dynamic bounding volume hierarchies. ACM Transactions on Graphics, 26(1):6:1– 6:18, January 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2022:RGR

[WBZ22] Sheng-Yu Wang, David Bau, and Jun-Yan Zhu. Rewriting geometric rules of a GAN. ACM Transactions on Graphics, 41(4):73:1–73:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530065.

Ware:1990:RCG

[WC90] Colin Ware and William Cowan. The RGYB color geometry. ACM Transactions on Graphics, 9(2):226–232, April 1990. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/

78962.html. See corrigenda [WC91].

Ware:1991:CRC

[WC91] Colin Ware and William Cowan. Corrigenda: “The RGYB Color Geometry”. ACM Transactions on Graphics, 10(3):319, July 1991. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [WC90].

Wei:2010:VMP

[WC10] Xiaolin Wei and Jinxiang Chai. VideoMocap: modeling physically realistic human motion from monocular video sequences. ACM Transactions on Graphics, 29(4): 42:1–42:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:RED

[WC21a] Minqi Wang and Emily A. Cooper. A re-examination of dichoptic tone mapping. ACM Transactions on Graphics, 40 (2):13:1–13:15, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3443702.

Wang:2021:CMS

[WC21b] Stephanie Wang and Albert Chern. Computing minimal surfaces with differential forms. ACM Transactions on Graphics, 40(4): 113:1–113:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459781.

White:2007:CAO

[WCF07] Ryan White, Keenan Crane, and D. A. Forsyth. Capturing and animating occluded cloth. ACM Transactions on Graphics, 26(3):34:1–34:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2022:CMC

[WCFL22] Kang Wu, Renjie Chen, Xiao-Ming Fu, and Ligang Liu. Computational mirror cup and saucer art. ACM Transactions on Graphics, 41(5): 174:1–174:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3517120.

Wolper:2020:AAA

[WCL^{+}20] Joshuah Wolper, Yunuo Chen, Minchen Li, Yu Fang, Ziyin Qu, Jiecong Lu, Meggie Cheng, and Chenfanfu Jiang. AnisoMPM: animating anisotropic damage mechanics. ACM Transactions on Graphics, 39(4): 37:1–37:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392428.

Wang:2023:FGB

[WCL^{+}23] Tianyu Wang, Jiong Chen, Dongping Li, Xiaowei Liu, Huamin Wang, and Kun Zhou. Fast GPU-based two-way continuous collision handling. ACM Transactions on Graphics, 42(5): 167:1–167:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3604551.

Wang:2018:LSS

[WCPM18] Tuanfeng Y. Wang, Duygu Ceylan, Jovan Popović, and Niloy J. Mitra. Learning a shared shape space for multimodal garment design. ACM Transactions on Graphics, 37 (6):203:1–203:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2021:DTG

[WCRZ21] Lifan Wu, Guangyan Cai, Ravi Ramamoorthi, and Shuang Zhao. Differentiable time-gated rendering. ACM Transactions on Graphics, 40 (6):287:1–287:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480489.

Wang:2022:JNP

[WCSC22] Yujie Wang, Praneeth Chakravarthula, Qi Sun, and Baoquan Chen. Joint neural phase retrieval and compression for energyand computation-efficient holography on the edge. ACM Transactions on Graphics, 41 (4):110:1–110:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530070.

Wu:2022:MYO

[WCZ^{+}22] Zongwei Wu, Liangyu Chai, Nanxuan Zhao, Bailin Deng, Yongtuo Liu, Qiang Wen, Junle Wang, and Shengfeng He. Make your own sprites: Aliasing-aware and cell-controllable pixelization. ACM Transactions on Graphics, 41(6):193:1–193:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555482.

Wang:2006:CAF

[WDAC06] Jue Wang, Steven M. Drucker, Maneesh Agrawala, and Michael F. Cohen. The cartoon animation filter. ACM Transactions on Graphics, 25 (3):1169–1173, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weyrich:2007:DBR

[WDB^{+}07] Tim Weyrich, Jia Deng, Connelly Barnes, Szymon Rusinkiewicz, and Adam Finkelstein. Digital bas-relief from 3D scenes. ACM Transactions on Graphics, 26(3): 32:1–32:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weber:2008:PAA

[WDB^{+}08] Ofir Weber, Yohai S. Devir, Alexander M. Bronstein, Michael M. Bronstein, and Ron Kimmel. Parallel algorithms for approximation of distance maps on parametric surfaces. ACM Transactions on Graphics, 27(4): 104:1–104:16, October 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:TFS

[WDK^{+}21] Mengdi Wang, Yitong Deng, Xiangxin Kong, Aditya H. Prasad, Shiying Xiong, and Bo Zhu. Thin-film smoothed particle hydrodynamics fluid. ACM Transactions on Graphics, 40(4):110:1–110:16, August 2021. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459864.

Wu:2011:PBI

[WDR11] Hongzhi Wu, Julie Dorsey, and Holly Rushmeier. Physicallybased interactive bi-scale material design. ACM Transactions on Graphics, 30(6): 145:1–145:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2013:IBS

[WDR13] Hongzhi Wu, Julie Dorsey, and Holly Rushmeier. Inverse bi-scale material design. ACM Transactions on Graphics, 32 (6):163:1–163:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:KNM

[WDT^{+}09] Jiaping Wang, Yue Dong, Xin Tong, Zhouchen Lin, and Baining Guo. Kernel Nyström method for light transport. ACM Transactions on Graphics, 28(3):29:1–29:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wadhwa:2015:DMR

[WDW^{+}15] Neal Wadhwa, Tali Dekel, Donglai Wei, Frédo Durand, and William T. Freeman. Deviation magnification: revealing departures from ideal geometries. ACM Transactions on Graphics, 34(6): 226:1–226:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weiss:2006:FMB

[Wei06] Ben Weiss. Fast median and bilateral filtering. ACM Transactions on Graphics, 25 (3):519–526, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wei:2008:PPD

[Wei08] Li-Yi Wei. Parallel Poisson disk sampling. ACM Transactions on Graphics, 27(3): 20:1–20:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wei:2010:MCB

[Wei10] Li-Yi Wei. Multi-class blue noise sampling. ACM Transactions on Graphics, 29(4): 79:1–79:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Westmore:1988:WBG

[Wes88] Richard J. Westmore. A window-based graphics frame store architecture. ACM Transactions on Graphics, 7 (4):233–248, October 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
46166.html.

West:2021:PBF

[Wes21] Rex West. Physically-based feature line rendering. ACM Transactions on Graphics, 40 (6):246:1–246:11, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480550.

Ware:1996:ESM

[WF96] Colin Ware and Glenn Franck.

Evaluating stereo and motion cues for visualizing information nets in three dimensions. ACM Transactions on Graphics, 15(2):121–140, April 1996. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Walter:2005:LSA

[WFA^{+}05] Bruce Walter, Sebastian Fernandez, Adam Arbree, Kavita Bala, Michael Donikian, and Donald P. Greenberg. Lightcuts: a scalable approach to illumination. ACM Transactions on Graphics, 24(3): 1098–1107, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2018:MAO

[WFDH18] Congli Wang, Qiang Fu, Xiong Dun, and Wolfgang Heidrich. Megapixel adaptive optics: towards correcting large-scale distortions in computational cameras. ACM Transactions on Graphics, 37(4):115:1–115:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weyrich:2007:HAS

[WFH^{+}07] Tim Weyrich, Cyril Flaig, Simon Heinzle, Simon Mall, Timo Aila, Kaspar Rohrer, Daniel B. Fasnacht, Norbert Felber, Stephan Oetiker, Hubert Kaeslin, Mario Botsch, and Markus Gross. A hardware architecture for surface splatting. ACM Transactions on Graphics, 26(3): 90:1–90:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:OWC

[WFH09] Jack M. Wang, David J. Fleet, and Aaron Hertzmann. Optimizing walking controllers. ACM Transactions on Graphics, 28(5): 168:1–168:8, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2010:OWC

[WFH10] Jack M. Wang, David J. Fleet, and Aaron Hertzmann. Optimizing walking controllers for uncertain inputs and environments. ACM Transactions on Graphics, 29(4): 73:1–73:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2015:RGN

[WFL^{+}15] Peng-Shuai Wang, Xiao-Ming Fu, Yang Liu, Xin Tong, Shi-Lin Liu, and Baining Guo. Rolling guidance normal filter for geometric processing. ACM Transactions on Graphics, 34(6):173:1– 173:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wolper:2019:CMC

[WFL^{+}19] Joshuah Wolper, Yu Fang, Minchen Li, Jiecong Lu, Ming Gao, and Chenfanfu Jiang. CD-MPM: continuum damage material point methods for dynamic fracture animation. ACM Transactions on Graphics, 38(4):119:1–119:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2012:AIB

[WFP12] Bin Wang, François Faure, and Dinesh K. Pai. Adaptive image-based intersection volume. ACM Transactions on Graphics, 31(4): 97:1–97:9, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:MAT

[WFS^{+}09] Yu-Shuen Wang, Hongbo Fu, Olga Sorkine, Tong-Yee Lee, and Hans-Peter Seidel. Motion-aware temporal coherence for video resizing. ACM Transactions on Graphics, 28 (5):127:1–127:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:LSB

[WFS^{+}21] Bolun Wang, Zachary Ferguson, Teseo Schneider, Xin Jiang, Marco Attene, and Daniele Panozzo. A largescale benchmark and an inclusion-based algorithm for continuous collision detection. ACM Transactions on Graphics, 40(5):188:1–188:16, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3460775.

Wretborn:2022:GBW

[WFS22] Joel Wretborn, Sean Flynn, and Alexey Stomakhin. Guided bubbles and wet foam for realistic whitewater simulation. ACM Transactions on Graphics, 41(4):117:1–117:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530059.

Wu:2010:MRI

[WFY^{+}10] Tai-Pang Wu, Chi-Wing Fu, Sai-Kit Yeung, Jiaya Jia, and Chi-Keung Tang. Modeling and rendering of impossible figures. ACM Transactions on Graphics, 29(2): 13:1–13:15, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weber:2009:CFC

[WG09] Andrew J. Weber and Galen Gornowicz. Collision-free construction of animated feathers using implicit constraint surfaces. ACM Transactions on Graphics, 28(2): 12:1–12:7, April 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weber:2010:CCM

[WG10] Ofir Weber and Craig Gotsman. Controllable conformal maps for shape deformation and interpolation. ACM Transactions on Graphics, 29 (4):78:1–78:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wronski:2019:HMF

[WGDE^{+}19] Bartlomiej Wronski, Ignacio Garcia-Dorado, Manfred Ernst, Damien Kelly, Michael Krainin, Chia-Kai Liang, Marc Levoy, and Peyman Milanfar. Handheld multi-frame super-resolution. ACM Transactions on Graphics, 38(4): 28:1–28:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2018:SM

[WGF^{+}18] Kui Wu, Xifeng Gao, Zachary Ferguson, Daniele Panozzo, and Cem Yuksel. Stitch meshing. ACM Transactions on Graphics, 37(4):130:1– 130:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Won:2020:SAC

[WGH20] Jungdam Won, Deepak Gopinath, and Jessica Hodgins. A scalable approach to control diverse behaviors for physically simulated characters. ACM Transactions on Graphics, 39 (4):33:1–33:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392381.

Won:2021:CSP

[WGH21] Jungdam Won, Deepak Gopinath, and Jessica Hodgins. Control strategies for physically simulated characters performing two-player competitive sports. ACM Transactions on Graphics, 40(4): 146:1–146:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459761.

Won:2022:PBC

[WGH22] Jungdam Won, Deepak Gopinath, and Jessica Hodgins. Physicsbased character controllers using conditional VAEs. ACM Transactions on Graphics, 41(4):96:1–96:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530067.

Wadhwa:2018:SDF

[WGJ^{+}18] Neal Wadhwa, Rahul Garg, David E. Jacobs, Bryan E. Feldman, Nori Kanazawa, Robert Carroll, Yair Movshovitz-Attias, Jonathan T. Barron, Yael Pritch, and Marc Levoy. Synthetic depth-of-field with a single-camera mobile phone. ACM Transactions on Graphics, 37(4):64:1–64:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Whelan:2018:RSM

[WGL^{+}18] Thomas Whelan, Michael Goesele, Steven J. Lovegrove, Julian Straub, Simon Green, Richard Szeliski, Steven Butterfield, Shobhit Verma, and Richard Newcombe. Reconstructing scenes with mirror and glass surfaces. ACM Transactions on Graphics, 37(4):102:1–102:??, August 2018. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Wilson:2010:TUP

[WGP^{+}10] Cyrus A. Wilson, Abhijeet Ghosh, Pieter Peers, Jen-Yuan Chiang, Jay Busch, and Paul Debevec. Temporal upsampling of performance geometry using photometric alignment. ACM Transactions on Graphics, 29(2):17:1– 17:11, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2023:VQH

[WGS23] Yu Wang, Minghao Guo, and Justin Solomon. Variational quasi-harmonic maps for computing diffeomorphisms. ACM Transactions on Graphics, 42(4):130:1–130:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592105.

Wenger:2005:PRR

[WGT^{+}05] Andreas Wenger, Andrew Gardner, Chris Tchou, Jonas Unger, Tim Hawkins, and Paul Debevec. Performance relighting and reflectance transformation with time-multiplexed illumination. ACM Transactions on Graphics, 24(3):756–764, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2013:PAS

[WGW^{+}13] Yunhai Wang, Minglun Gong, Tianhua Wang, Daniel Cohen-Or, Hao Zhang, and Baoquan Chen. Projective analysis for 3D shape segmentation. ACM Transactions on Graphics, 32 (6):192:1–192:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wolski:2018:DMP

[WGY^{+}18] Krzysztof Wolski, Daniele Giunchi, Nanyang Ye, Piotr Didyk, Karol Myszkowski, Radoslaw Mantiuk, Hans-Peter Seidel, Anthony Steed, and Rafal K. Mantiuk. Dataset and metrics for predicting local visible differences. ACM Transactions on Graphics, 37(5):172:1– 172:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2012:HQI

[WHB^{+}12] Sen Wang, Tingbo Hou, John Border, Hong Qin, and Rodney Miller. High-quality image deblurring with panchromatic pixels. ACM Transactions on Graphics, 31(5): 120:1–120:11, August 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2012:OLC

[WHDK12] Jack M. Wang, Samuel R.

Hamner, Scott L. Delp, and Vladlen Koltun. Optimizing locomotion controllers using biologically-based actuators and objectives. ACM Transactions on Graphics, 31(4): 25:1–25:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wood:2004:RET

[WHDS04] Zoë Wood, Hugues Hoppe, Mathieu Desbrun, and Peter Schröder. Removing excess topology from isosurfaces. ACM Transactions on Graphics, 23(2):190–208, April 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weghorst:1984:ICM

[WHG84] Hank Weghorst, Gary Hooper, and Donald P. Greenberg. Improved computational methods for ray tracing. ACM Transactions on Graphics, 3 (1):52–69, January 1984. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2015:DPC

[WHG^{+}15] Shihao Wu, Hui Huang, Minglun Gong, Matthias Zwicker, and Daniel Cohen-Or. Deep points consolidation. ACM Transactions on Graphics, 34(6):176:1– 176:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2020:EBM

[WHHY20] Beibei Wang, Milos Hasan, Nicolas Holzschuch, and Ling-Qi Yan. Example-based microstructure rendering with constant storage. ACM Transactions on Graphics, 39 (5):162:1–162:12, September 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3406836.

Winchenbach:2017:ICA

[WHK17] Rene Winchenbach, Hendrik Hochstetter, and Andreas Kolb. Infinite continuous adaptivity for incompressible SPH. ACM Transactions on Graphics, 36(4): 102:1–102:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2010:EBW

[WHRO10] Huamin Wang, Florian Hecht, Ravi Ramamoorthi, and James O’Brien. Examplebased wrinkle synthesis for clothing animation. ACM Transactions on Graphics, 29 (4):107:1–107:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Walter:1997:GIU

[WHSG97] Bruce Walter, Philip M. Hubbard, Peter Shirley, and Donald P. Greenberg. Global illumination using local linear density estimation. ACM Transactions on Graphics, 16 (3):217–259, July 1997. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1997-16-3/
                                        p217-walter/.

Wang:2011:SCV

[WHSL11] Yu-Shuen Wang, Jen-Hung Hsiao, Olga Sorkine, and Tong-Yee Lee. Scalable and coherent video resizing with per-frame optimization. ACM Transactions on Graphics, 30 (4):88:1–88:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2013:GBC

[WHY^{+}13] Rui Wang, Yuchi Huo, Yazhen Yuan, Kun Zhou, Wei Hua, and Hujun Bao. GPU-based out-of-core many-lights rendering. ACM Transactions on Graphics, 32(6):210:1– 210:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2020:PCE

[WHY20] Beibei Wang, Milos Hasan, and Ling-Qi Yan. Path cuts: efficient rendering of pure specular light transport. ACM Transactions on Graphics, 39 (6):238:1–238:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417792.

Wei:2008:ITS

[WHZ^{+}08] Li-Yi Wei, Jianwei Han, Kun Zhou, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Inverse texture synthesis. ACM Transactions on Graphics, 27(3):52:1–52:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wald:2006:RTA

[WIK^{+}06] Ingo Wald, Thiago Ize, Andrew Kensler, Aaron Knoll, and Steven G. Parker. Ray tracing animated scenes using coherent grid traversal. ACM Transactions on Graphics, 25 (3):485–493, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Williams:1992:VOM

[Wil92] Peter L. Williams. Visibility ordering meshed polyhedra. ACM Transactions on Graphics, 11(2):103–126, April 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/

                                        toc/Abstracts/0730-0301/
                                        130899.html.

Wimmer:2014:MRS

[Wim14]  Michael Wimmer.  Meta-

representation of shape families. ACM Transactions on Graphics, 33(4):34:1– 34:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2019:KOM

[WJ19] Jui-Hsien Wang and Doug L. James. KleinPAT: optimal mode conflation for timedomain precomputation of acoustic transfer. ACM Transactions on Graphics, 38 (4):122:1–122:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2015:LSD

[WJBK15] Yu Wang, Alec Jacobson, Jernej Barbic, and Ladislav Kavan. Linear subspace design for real-time shape deformation. ACM Transactions on Graphics, 34(4):57:1– 57:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2022:PFM

[WJF^{+}22] Beibei Wang, Wenhua Jin, Jiahui Fan, Jian Yang, Nicolas Holzschuch, and Ling-Qi Yan. Position-free multiple-bounce computations for Smith microfacet BSDFs. ACM Transactions on Graphics, 41 (4):134:1–134:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530112.

Wang:2023:SLM

[WJHY23] Beibei Wang, Wenhua Jin, Milos Hasan, and Ling-Qi Yan. SpongeCake: a layered microflake surface appearance model. ACM Transactions on Graphics, 42 (1):8:1–8:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3546940.

Wang:2020:CST

[WJL^{+}20] Hui Wang, Yongxu Jin, Anqi Luo, Xubo Yang, and Bo Zhu. Codimensional surface tension flow using moving-least-squares particles. ACM Transactions on Graphics, 39 (4):42:1–42:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3386569.3392487.

Wilburn:2005:HPI

[WJV^{+}05] Bennett Wilburn, Neel Joshi, Vaibhav Vaish, Eino-Ville Talvala, Emilio Antunez, Adam Barth, Andrew Adams, Mark Horowitz, and Marc Levoy.

High performance imaging using large camera arrays. ACM Transactions on Graphics, 24 (3):765–776, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2008:CRM

[WJZL08] Wenping Wang, Bert Jüttler, Dayue Zheng, and Yang Liu. Computation of rotation minimizing frames. ACM Transactions on Graphics, 27(1): 2:1–2:19, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ware:1995:UVT

[WK95] Colin Ware and William Knight. Using visual texture for information display. ACM Transactions on Graphics, 14(1):3–20, January 1995. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        200974.html.

Winchenbach:2021:ORS

[WK21] Rene Winchenbach and Andreas Kolb. Optimized refinement for spatially adaptive SPH. ACM Transactions on Graphics, 40(1):8:1– 8:15, January 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3363555.

Walter:2012:BL

[WKB12] Bruce Walter, Pramook Khungurn, and Kavita Bala. Bidirectional lightcuts. ACM Transactions on Graphics, 31(4):59:1–59:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Walton:2021:BBR

[WKF^{+}21] David R. Walton, Rafael Kuffner Dos Anjos, Sebastian Friston, David Swapp, Kaan Akşit, Anthony Steed, and Tobias Ritschel. Beyond blur: real-time ventral metamers for foveated rendering. ACM Transactions on Graphics, 40 (4):48:1–48:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459943.

Wang:2018:TGO

[WKHA18] Xi Wang, Sebastian Koch, Kenneth Holmqvist, and Marc Alexa. Tracking the gaze on objects in 3D: how do people really look at the bunny? ACM Transactions on Graphics, 37(6): 188:1–188:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2023:TCT

[WKMH^{+}23] Ziqi Wang, Florian Kennel-Maushart, Yijiang Huang, Bernhard Thomaszewski, and Stelian Coros. A temporal coherent topology optimization approach for assembly planning of bespoke frame structures. ACM Transactions on Graphics, 42(4): 144:1–144:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592102.

Weiskopf:1999:SDE

[WKR99] Daniel Weiskopf, Ute Kraus, and Hanns Ruder. Searchlight and Doppler effects in the visualization of special relativity: a corrected derivation of the transformation of radiance. ACM Transactions on Graphics, 18(3):278–292, July 1999. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
1999-18-3/p278-weiskopf/.

Won:2016:STD

[WL16] Jungdam Won and Jehee Lee. Shadow theatre: discovering human motion from a sequence of silhouettes. ACM Transactions on Graphics, 35 (4):147:1–147:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:DSH

[WL21] Yizhi Wang and Zhouhui Lian. DeepVecFont: synthesizing high-quality vector fonts via dual-modality learning. ACM Transactions on Graphics, 40(6): 265:1–265:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480488.

Wang:2020:HOT

[WLF^{+}20] Xinlei Wang, Minchen Li, Yu Fang, Xinxin Zhang, Ming Gao, Min Tang, Danny M. Kaufman, and Chenfanfu Jiang. Hierarchical optimization time integration for CFLrate MPM stepping. ACM Transactions on Graphics, 39 (3):21:1–21:16, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3386760.

Wang:2017:COB

[WLG^{+}17] Peng-Shuai Wang, Yang Liu, Yu-Xiao Guo, Chun-Yu Sun, and Xin Tong. O-CNN: octree-based convolutional neural networks for 3D shape analysis. ACM Transactions on Graphics, 36(4): 72:1–72:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wong:2013:RVB

[WLH^{+}13] Sai-Keung Wong, Wen-Chieh Lin, Chun-Hung Hung, Yi-Jheng Huang, and Shing-Yeu Lii. Radial view based culling for continuous self-collision detection of skeletal models. ACM Transactions on Graphics, 32(4):114:1–114:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wetzstein:2011:LTI

[WLHR11] Gordon Wetzstein, Douglas Lanman, Wolfgang Heidrich, and Ramesh Raskar. Layered 3D: tomographic image synthesis for attenuation-based light field and high dynamic range displays. ACM Transactions on Graphics, 30(4): 95:1–95:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wetzstein:2012:TDC

[WLHR12] Gordon Wetzstein, Douglas Lanman, Matthew Hirsch, and Ramesh Raskar. Tensor displays: compressive light field synthesis using multilayer displays with directional backlighting. ACM Transactions on Graphics, 31(4): 80:1–80:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2022:NGC

[WLJ^{+}22] Yusen Wang, Zongcheng Li, Yu Jiang, Kaixuan Zhou, Tuo Cao, Yanping Fu, and Chunxia Xiao. NeuralRoom: Geometry-constrained neural implicit surfaces for indoor scene reconstruction. ACM Transactions on Graphics, 41 (6):226:1–226:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555514.

Wang:2014:BDD

[WLL^{+}14] Miao Wang, Yu-Kun Lai, Yuan Liang, Ralph R. Martin, and Shi-Min Hu. BiggerPicture: data-driven image extrapolation using graph matching. ACM Transactions on Graphics, 33(6): 173:1–173:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2023:DOG

[WLL23] Jiaxian Wu, Yue Lin, and Dehui Lu. DR-Occluder: Generating occluders using differentiable rendering. ACM Transactions on Graphics, 42(6): 231:1–231:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618346.

Wei:2022:ACD

[WLLS22] Xinyue Wei, Minghua Liu, Zhan Ling, and Hao Su. Approximate convex decomposition for 3D meshes with collision-aware concavity and tree search. ACM Transactions on Graphics, 41(4):

42:1–42:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530103.

Wei:2015:ILF

[WLM^{+}15] Li-Yi Wei, Chia-Kai Liang, Graham Myhre, Colvin Pitts, and Kurt Akeley. Improving light field camera sample design with irregularity and aberration. ACM Transactions on Graphics, 34(4): 152:1–152:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Won:2014:GRD

[WLO^{+}14] Jungdam Won, Kyungho Lee, Carol O’Sullivan, Jessica K. Hodgins, and Jehee Lee. Generating and ranking diverse multi-character interactions. ACM Transactions on Graphics, 33(6):219:1– 219:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wolinski:2016:WCA

[WLP16] David Wolinski, Ming C. Lin, and Julien Pettré. Warp-Driver: context-aware probabilistic motion prediction for crowd simulation. ACM Transactions on Graphics, 35 (6):164:1–164:??, November 2016. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2023:EHZ

[WLS^{+}23] Xiaotong Wu, Wei-Sheng Lai, Yichang Shih, Charles Herrmann, Michael Krainin, Deqing Sun, and Chia-Kai Liang. Efficient hybrid zoom using camera fusion on mobile phones. ACM Transactions on Graphics, 42(6): 263:1–263:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618362.

Wang:2010:MBV

[WLSL10] Yu-Shuen Wang, Hui-Chih Lin, Olga Sorkine, and Tong-Yee Lee. Motion-based video retargeting with optimized crop-and-warp. ACM Transactions on Graphics, 29(4): 90:1–90:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2016:MDC

[WLT16] Peng-Shuai Wang, Yang Liu, and Xin Tong. Mesh denoising via cascaded normal regression. ACM Transactions on Graphics, 35(6): 232:1–232:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2022:DOG

[WLT22] Peng-Shuai Wang, Yang Liu, and Xin Tong. Dual octree graph networks for learning adaptive volumetric shape representations. ACM Transactions on Graphics, 41 (4):103:1–103:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530087.

Wang:2019:PPI

[WLW^{+}19] Kai Wang, Yu-An Lin, Ben Weissmann, Manolis Savva, Angel X. Chang, and Daniel Ritchie. PlanIT: planning and instantiating indoor scenes with relation graph and spatial prior networks. ACM Transactions on Graphics, 38 (4):132:1–132:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2018:SSB

[WLX^{+}18] Guan Wang, Hamid Laga, Ning Xie, Jinyuan Jia, and Hedi Tabia. The shape space of 3D botanical tree models. ACM Transactions on Graphics, 37(1):7:1–7:??, January 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2016:CMC

[WLY^{+}16] Ruimin Wang, Ligang Liu, Zhouwang Yang, Kang Wang, Wen Shan, Jiansong Deng, and Falai Chen. Construction of manifolds via compatible sparse representations. ACM Transactions on Graphics, 35(2):14:1–14:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2020:SMA

[WLY20] Hanqing Wang, Wei Liang, and Lap-Fai Yu. Scene mover: automatic move planning for scene arrangement by deep reinforcement learning. ACM Transactions on Graphics, 39 (6):233:1–233:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417788.

Wang:2009:PGL

[WLZ^{+}09] Huamin Wang, Miao Liao, Qing Zhang, Ruigang Yang, and Greg Turk. Physically guided liquid surface modeling from videos. ACM Transactions on Graphics, 28(3): 90:1–90:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:PSL

[WLZ^{+}21] Yupan Wang, Guiqing Li, Huiqian Zhang, Xinyi Zou, Yuxin Liu, and Yongwei Nie. PanoMan: Sparse localized components-based model for full human motions. ACM Transactions on Graphics, 40 (2):19:1–19:17, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3447244.

Wilson:2003:SCE

[WM03] Andrew Wilson and Dinesh Manocha. Simplifying complex environments using incremental textured depth meshes. ACM Transactions on Graphics, 22(3): 678–688, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wanat:2014:SCC

[WM14] Robert Wanat and Rafal K. Mantiuk. Simulating and compensating changes in appearance between day and night vision. ACM Transactions on Graphics, 33(4): 147:1–147:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2019:HMS

[WMB19] Bohan Wang, George Matcuk, and Jernej Barbic. Hand modeling and simulation using stabilized magnetic resonance imaging. ACM Transactions on Graphics, 38(4): 115:1–115:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2020:RRT

[WMB^{+}20] Jiayi Wang, Franziska Mueller, Florian Bernard, Suzanne Sorli, Oleksandr Sotnychenko, Neng Qian, Miguel A. Otaduy, Dan Casas, and Christian Theobalt. RGB2Hands: realtime tracking of 3D hand interactions from monocular RGB video. ACM Transactions on Graphics, 39(6): 218:1–218:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417852.

Wang:2021:MPA

[WMB21] Bohan Wang, George Matcuk, and Jernej Barbic. Modeling of personalized anatomy using plastic strains. ACM Transactions on Graphics, 40 (2):14:1–14:21, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3443703.

Wei:2011:PVS

[WMC11] Xiaolin Wei, Jianyuan Min, and Jinxiang Chai. Physically valid statistical models for human motion generation. ACM Transactions on Graphics, 30(3):19:1–19:10, May 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weyrich:2006:AHF

[WMP^{+}06] Tim Weyrich, Wojciech Matusik, Hanspeter Pfister, Bernd Bickel, Craig Donner, Chien Tu, Janet McAndless, Jinho Lee, Addy Ngan, Henrik Wann Jensen, and Markus Gross. Analysis of human faces using a measurementbased skin reflectance model. ACM Transactions on Graphics, 25(3):1013–1024, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2005:WDS

[WMT05] Huamin Wang, Peter J. Mucha, and Greg Turk. Water drops on surfaces. ACM Transactions on Graphics, 24 (3):921–929, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2015:UAS

[WMW15] Xiaofeng Wu, Rajaditya Mukherjee, and Huamin Wang. A unified approach for subspace simulation of deformable bodies in multiple domains. ACM Transactions on Graphics, 34(6): 241:1–241:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2013:VBH

[WMZ^{+}13] Yangang Wang, Jianyuan Min, Jianjie Zhang, Yebin Liu, Feng Xu, Qionghai Dai, and Jinxiang Chai. Videobased hand manipulation capture through composite motion control. ACM Transactions on Graphics, 32(4): 43:1–43:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wiersma:2022:DAO

[WNEH22] Ruben Wiersma, Ahmad Nasikun, Elmar Eisemann, and Klaus Hildebrandt. Delta-Conv: anisotropic operators for geometric deep learning on point clouds. ACM Transactions on Graphics, 41 (4):105:1–105:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530166.

Whiting:2009:PMS

[WOD09] Emily Whiting, John Ochsendorf, and Frédo Durand. Procedural modeling of structurallysound masonry buildings. ACM Transactions on Graphics, 28(5):112:1–112:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Winnemoller:2006:RTV

[WOG06] Holger Winnemöller, Sven C.

Olsen, and Bruce Gooch. Real-time video abstraction. ACM Transactions on Graphics, 25(3):1221–1226, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wei:2005:MHM

[WOQS05] Yichen Wei, Eyal Ofek, Long Quan, and Heung-Yeung Shum. Modeling hair from multiple views. ACM Transactions on Graphics, 24(3): 816–820, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2010:MRI

[WOR10] Huamin Wang, James O’Brien, and Ravi Ramamoorthi. Multi-resolution isotropic strain limiting. ACM Transactions on Graphics, 29(6):156:1– 156:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2011:DDE

[WOR11] Huamin Wang, James F. O’Brien, and Ravi Ramamoorthi. Data-driven elastic models for cloth: modeling and measurement. ACM Transactions on Graphics, 30 (4):71:1–71:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wold:1990:RCS

[WP90] Erling Wold and Kim Pepard. Re: Comments on “Stochastic Sampling in Computer Graphics”. ACM Transactions on Graphics, 9(2): 237–243, April 1990. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). See [Coo86, Pav90].

Wallner:2006:ISS

[WP06] Johannes Wallner and Helmut Pottmann. Intrinsic subdivision with smooth limits for graphics and animation. ACM Transactions on Graphics, 25 (2):356–374, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wampler:2009:OGF

[WP09a] Kevin Wampler and Zoran Popović. Optimal gait and form for animal locomotion. ACM Transactions on Graphics, 28(3):60:1–60:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:RTH

[WP09b] Robert Y. Wang and Jovan Popović. Real-time handtracking with a color glove. ACM Transactions on Graphics, 28(3):63:1–63:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weissmann:2010:FBS

[WP10] Steffen Weißmann and Ulrich Pinkall. Filament-based smoke with vortex shedding and variational reconnection.

ACM Transactions on Graphics, 29(4):115:1–115:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weissmann:2012:URB

[WP12] Steffen Weißmann and Ulrich Pinkall. Underwater rigid body dynamics. ACM Transactions on Graphics, 31(4): 104:1–104:7, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wachtel:2014:FTB

[WPC^{+}14] Florent Wachtel, Adrien Pilleboue, David Coeurjolly, Katherine Breeden, Gurprit Singh, Gaël Cathelin, Fernando de Goes, Mathieu Desbrun, and Victor Ostromoukhov. Fast tilebased adaptive sampling with user-specified Fourier spectra. ACM Transactions on Graphics, 33(4):56:1–56:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2016:PAM

[WPGM16] Rundong Wu, Huaishu Peng, François Guimbretière, and Steve Marschner. Printing arbitrary meshes with a 5DOF wireframe printer. ACM Transactions on Graphics, 35 (4):101:1–101:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Won:2017:HTY

[WPKL17] Jungdam Won, Jongho Park, Kwanyu Kim, and Jehee Lee. How to train your dragon: example-guided control of flapping flight. ACM Transactions on Graphics, 36(6): 198:1–198:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2006:FBS

[WPL06] Wenping Wang, Helmut Pottmann, and Yang Liu. Fitting B-spline curves to point clouds by curvaturebased squared distance minimization. ACM Transactions on Graphics, 25(2): 214–238, April 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Won:2018:ACF

[WPL18] Jungdam Won, Jungnam Park, and Jehee Lee. Aerobatics control of flying creatures via self-regulated learning. ACM Transactions on Graphics, 37(6):181:1– 181:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Willis:2021:FGD

[WPL^{+}21] Karl D. D. Willis, Yewen Pu, Jieliang Luo, Hang Chu, Tao Du, Joseph G. Lambourne, Armando Solar-Lezama, and Wojciech Matusik. Fusion 360 gallery: a dataset and environment for programmatic CAD construction from human design sequences. ACM Transactions on Graphics, 40 (4):54:1–54:24, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459818.

Weidner:2018:ELC

[WPLS18] Nicholas J. Weidner, Kyle Piddington, David I. W. Levin, and Shinjiro Sueda. Eulerian-on-Lagrangian cloth simulation. ACM Transactions on Graphics, 37(4):50:1– 50:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weyrich:2009:FMC

[WPMR09] Tim Weyrich, Pieter Peers, Wojciech Matusik, and Szymon Rusinkiewicz. Fabricating microgeometry for custom surface reflectance. ACM Transactions on Graphics, 28 (3):32:1–32:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2007:RTE

[WPP07] Robert Y. Wang, Kari Pulli, and Jovan Popović. Realtime enveloping with rotational regression. ACM Transactions on Graphics, 26(3):

73:1–73:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wampler:2014:GLS

[WPP14] Kevin Wampler, Zoran Popović, and Jovan Popović. Generalizing locomotion style to new animals with inverse optimal regression. ACM Transactions on Graphics, 33(4): 49:1–49:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weissmann:2014:SRS

[WPS14] Steffen Weißmann, Ulrich Pinkall, and Peter Schröder. Smoke rings from smoke. ACM Transactions on Graphics, 33(4):140:1–140:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:TVF

[WQF^{+}21] Zeyu Wang, Sherry Qiu, Nicole Feng, Holly Rushmeier, Leonard McMillan, and Julie Dorsey. Tracing versus freehand for evaluating computergenerated drawings. ACM Transactions on Graphics, 40 (4):52:1–52:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459819.

Wang:2018:TWB

[WQLJ18] Jui-Hsien Wang, Ante Qu, Timothy R. Langlois, and Doug L. James. Toward wave-based sound synthesis for computer animation. ACM Transactions on Graphics, 37(4):109:1–109:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2020:MPS

[WQS^{+}20] Xinlei Wang, Yuxing Qiu, Stuart R. Slattery, Yu Fang, Minchen Li, Song-Chun Zhu, Yixin Zhu, Min Tang, Dinesh Manocha, and Chenfanfu Jiang. A massively parallel and scalable multi-CPU material point method. ACM Transactions on Graphics, 39 (4):30:1–30:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392442.

Wang:2018:ASH

[WR18] Jingwen Wang and Ravi Ramamoorthi. Analytic spherical harmonic coefficients for polygonal area lights. ACM Transactions on Graphics, 37 (4):54:1–54:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wadhwa:2013:PBV

[WRDF13] Neal Wadhwa, Michael Rubinstein, Frédo Durand, and William T. Freeman. Phasebased video motion processing. ACM Transactions on Graphics, 32(4): 80:1–80:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:AFR

[WRG^{+}09] Jiaping Wang, Peiran Ren, Minmin Gong, John Snyder, and Baining Guo. Allfrequency rendering of dynamic, spatially-varying reflectance. ACM Transactions on Graphics, 28(5):133:1– 133:10, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wicke:2010:DLR

[WRK^{+}10] Martin Wicke, Daniel Ritchie, Bryan M. Klingner, Sebastian Burke, Jonathan R. Shewchuk, and James F. O’Brien. Dynamic local remeshing for elastoplastic simulation. ACM Transactions on Graphics, 29(4): 49:1–49:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2012:EVM

[WRS^{+}12] Hao-Yu Wu, Michael Rubinstein, Eugene Shih, John Guttag, Frédo Durand, and William Freeman. Eulerian video magnification for revealing subtle changes in the world. ACM Transactions on Graphics, 31(4): 65:1–65:8, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Woo:1985:LTA

[WS85] T. C. Woo and S. Y. Shin. A linear time algorithm for triangulating a point-visible polygon. ACM Transactions on Graphics, 4(1):60– 69, January 1985. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ward:1999:HRC

[WS99] Gregory Ward and Maryann Simmons. The holodeck ray cache: an interactive rendering system for global illumination in nondiffuse environments. ACM Transactions on Graphics, 18(4): 361–398, October 1999. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tog/1999-18-4/
                                        p361-ward/.

Wu:2017:IRSa

[WS17a] Jung-Hsuan Wu and Suguru Saito. Interactive relighting in single low-dynamic range images. ACM Transactions on Graphics, 36(2): 18:1–18:??, April 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2017:IRSb

[WS17b] Jung-Hsuan Wu and Suguru Saito. Interactive relighting in single low-dynamic range images. ACM Transactions on Graphics, 36(4): 143:1–143:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:FQH

[WS21] Yu Wang and Justin Solomon.

Fast quasi-harmonic weights for geometric data interpolation. ACM Transactions on Graphics, 40(4): 73:1–73:15, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459801.

Wang:2018:DCP

[WSCR18] Kai Wang, Manolis Savva, Angel X. Chang, and Daniel Ritchie. Deep convolutional priors for indoor scene synthesis. ACM Transactions on Graphics, 37(4):70:1– 70:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2016:UTT

[WSH^{+}16] Tuanfeng Y. Wang, Hao Su, Qixing Huang, Jingwei Huang, Leonidas Guibas, and Niloy J. Mitra. Unsupervised texture transfer from images to model collections. ACM Transactions on Graphics, 35 (6):177:1–177:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2018:GLG

[WSH^{+}18] Hao Wang, Nadav Schor, Ruizhen Hu, Haibin Huang, Daniel Cohen-Or, and Hui Huang. Global-to-local generative model for 3D shapes. ACM Transactions on Graphics, 37(6):214:1– 214:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wolff:2019:WPA

[WSH19] Katja Wolff and Olga Sorkine-Hornung. Wallpaper pattern alignment along garment seams. ACM Transactions on Graphics, 38(4): 62:1–62:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2017:BMI

[WSJP17] Jui-Hsien Wang, Rajsekhar Setaluri, Doug L. James, and Dinesh K. Pai. Bounce maps: an improved restitution model for real-time rigidbody impact. ACM Transactions on Graphics, 36(4): 150:1–150:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wilkie:2013:FRD

[WSL13] David Wilkie, Jason Sewall, and Ming Lin. Flow reconstruction for data-driven traffic animation. ACM Transactions on Graphics, 32(4): 89:1–89:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2014:QDP

[WSL^{+}14] Shihao Wu, Wei Sun, Pinxin Long, Hui Huang, Daniel Cohen-Or, Minglun Gong, Oliver Deussen, and Baoquan Chen. Quality-driven Poissonguided autoscanning. ACM Transactions on Graphics, 33 (6):203:1–203:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2019:DGC

[WSL^{+}19] Yue Wang, Yongbin Sun, Ziwei Liu, Sanjay E. Sarma, Michael M. Bronstein, and Justin M. Solomon. Dynamic graph CNN for learning on point clouds. ACM Transactions on Graphics, 38(5): 146:1–146:??, October 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326362.

Wang:2018:ACP

[WSLT18] Peng-Shuai Wang, Chun-Yu Sun, Yang Liu, and Xin Tong. Adaptive O-CNN: a patchbased deep representation of 3D shapes. ACM Transactions on Graphics, 37(6): 217:1–217:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2011:EDS

[WSM11] Chun-Po Wang, Noah Snavely, and Steve Marschner. Estimating dual-scale properties of glossy surfaces from stepedge lighting. ACM Transactions on Graphics, 30(6): 172:1–172:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2023:ITG

[WSML23] Ronghuan Wu, Wanchao Su, Kede Ma, and Jing Liao. Icon-Shop: Text-guided vector icon synthesis with autoregressive transformers. ACM Transactions on Graphics, 42(6): 230:1–230:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618364.

Wang:2023:ASE

[WSND^{+}23] Zian Wang, Tianchang Shen, Merlin Nimier-David, Nicholas Sharp, Jun Gao, Alexander Keller, Sanja Fidler, Thomas Müller, and Zan Gojcic. Adaptive shells for efficient neural radiance field rendering. ACM Transactions on Graphics, 42(6): 260:1–260:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618390.

Wang:2018:DGF

[WSP18] Ziqi Wang, Peng Song, and Mark Pauly. DESIA: a general framework for designing interlocking assemblies. ACM Transactions on Graphics, 37 (6):191:1–191:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:MMO

[WSP21] Ziqi Wang, Peng Song, and Mark Pauly. MOCCA: modeling and optimizing conejoints for complex assemblies. ACM Transactions on Graphics, 40(4):181:1– 181:14, August 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459680.

Wang:2023:INR

[WSP^{+}23] Chao Wang, Ana Serrano, Xingang Pan, Krzysztof Wolski, Bin Chen, Karol Myszkowski, Hans-Peter Seidel, Christian Theobalt, and Thomas Leimkühler. An implicit neural representation for the image stack: Depth, all in focus, and high dynamic range. ACM Transactions on Graphics, 42(6): 221:1–221:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618367.

Woop:2005:RPR

[WSS05] Sven Woop, Jörg Schmittler, and Philipp Slusallek. RPU: a programmable ray processing unit for realtime ray tracing. ACM Transactions on Graphics, 24(3):434–444, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2018:DIL

[WSS18] Chenglei Wu, Takaaki Shiratori, and Yaser Sheikh. Deep incremental learning for efficient high-fidelity face tracking. ACM Transactions on Graphics, 37(6):234:1– 234:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wei:2019:VFA

[WSS^{+}19] Shih-En Wei, Jason Saragih, Tomas Simon, Adam W. Harley, Stephen Lombardi, Michal Perdoch, Alexander Hypes, Dawei Wang, Hernan Badino, and Yaser Sheikh.

VR facial animation via multiview image translation. ACM Transactions on Graphics, 38 (4):67:1–67:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2013:HPE

[WSSK13] He Wang, Kirill A. Sidorov, Peter Sandilands, and Taku Komura. Harmonic parameterization by electrostatics. ACM Transactions on Graphics, 32(5):155:1– 155:12, September 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wicke:2009:MBF

[WST09] Martin Wicke, Matt Stanton, and Adrien Treuille. Modular bases for fluid dynamics. ACM Transactions on Graphics, 28(3):39:1–39:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2008:INR

[WSTS08] Tai-Pang Wu, Jian Sun, Chi-Keung Tang, and Heung-Yeung Shum. Interactive normal reconstruction from a single image. ACM Transactions on Graphics, 27(5): 119:1–119:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2013:SPC

[WSVT13] Chenglei Wu, Carsten Stoll, Levi Valgaerts, and Christian Theobalt. On-set performance capture of multiple actors with a stereo camera. ACM Transactions on Graphics, 32 (6):161:1–161:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Whiting:2012:SOM

[WSW^{+}12] Emily Whiting, Hijung Shin, Robert Wang, John Ochsendorf, and Frédo Durand. Structural optimization of 3D masonry buildings. ACM Transactions on Graphics, 31(6): 159:1–159:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2016:REG

[WSXC16] Congyi Wang, Fuhao Shi, Shihong Xia, and Jinxiang Chai. Realtime 3D eye gaze animation using a single RGB camera. ACM Transactions on Graphics, 35(4): 118:1–118:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2019:KSM

[WSY19] Kui Wu, Hannah Swan, and Cem Yuksel. Knittable stitch meshes. ACM Transactions on Graphics, 38(1): 10:1–10:??, February 2019.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3292481.

Wang:2014:VIS

[WSZ^{+}14] Oliver Wang, Christopher Schroers, Henning Zimmer, Markus Gross, and Alexander Sorkine-Hornung. VideoSnapping: interactive synchronization of multiple videos. ACM Transactions on Graphics, 33 (4):77:1–77:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2018:DSD

[WSZ^{+}18] Lijun Wang, Xiaohui Shen, Jianming Zhang, Oliver Wang, Zhe Lin, Chih-Yao Hsieh, Sarah Kong, and Huchuan Lu. DeepLens: shallow depth of field from a single image. ACM Transactions on Graphics, 37(6): 245:1–245:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wojtan:2008:FVB

[WT08] Chris Wojtan and Greg Turk.

Fast viscoelastic behavior with thin features. ACM Transactions on Graphics, 27 (3):47:1–47:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2007:NSM

[WTBS07a] Tai-Pang Wu, Chi-Keung Tang, Michael S. Brown, and Heung-Yeung Shum. Natural shadow matting. ACM Transactions on Graphics, 26 (2):8:1–8:??, June 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2007:SIN

[WTBS07b] Tai-Pang Wu, Chi-Keung Tang, Michael S. Brown, and Heung-Yeung Shum. ShapePalettes: interactive normal transfer via sketching. ACM Transactions on Graphics, 26(3):44:1–44:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wolski:2022:GMP

[WTD^{+}22] Krzysztof Wolski, Laura Trutoiu, Zhao Dong, Zhengyang Shen, Kevin Mackenzie, and Alexandre Chapiro. Geometric: a perceptual dataset of distortions on faces. ACM Transactions on Graphics, 41 (6):215:1–215:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555475.

Wojtan:2009:DMS

[WTGT09] Chris Wojtan, Nils Thürey, Markus Gross, and Greg Turk. Deforming meshes that split and merge. ACM Transactions on Graphics, 28(3): 76:1–76:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wojtan:2010:PIT

[WTGT10] Chris Wojtan, Nils Thürey, Markus Gross, and Greg Turk. Physics-inspired topology changes for thin fluid features. ACM Transactions on Graphics, 29(4): 50:1–50:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2005:AFI

[WTL05] Rui Wang, John Tran, and David Luebke. Allfrequency interactive relighting of translucent objects with single and multiple scattering. ACM Transactions on Graphics, 24(3):1202–1207, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2006:AMM

[WTL^{+}06a] Jiaping Wang, Xin Tong, Stephen Lin, Minghao Pan, Chao Wang, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Appearance manifolds for modeling timevariant appearance of materials. ACM Transactions on Graphics, 25(3):754–761, July 2006. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2006:AFR

[WTL06b] Rui Wang, John Tran, and David Luebke. All-frequency relighting of glossy objects. ACM Transactions on Graphics, 25(2):293–318, April 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weinrauch:2023:EBM

[WTS^{+}23] Alexander Weinrauch, Wolfgang Tatzgern, Pascal Stadlbauer, Alexis Crickx, Jozef Hladky, Arno Coomans, Martin Winter, Joerg H. Mueller, and Markus Steinberger. Effect-based multiviewer caching for cloudnative rendering. ACM Transactions on Graphics, 42 (4):87:1–87:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592431.

Wang:2008:OSS

[WTSL08] Yu-Shuen Wang, Chiew-Lan Tai, Olga Sorkine, and Tong-Yee Lee. Optimized scaleand-stretch for image resizing. ACM Transactions on Graphics, 27(5):118:1– 118:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:1992:CQD

[Wu92] Xialin Wu. Color quantization by dynamic programming and principal analysis. ACM Transactions on Graphics, 11(4):348–372, October 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
146475.html.

Wilhelms:1992:OFI

[WV92] Jane Wilhelms and Allen Van Gelder. Octrees for faster isosurface generation. ACM Transactions on Graphics, 11(3):201–227, July 1992. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
130882.html.

Wilkie:2021:FRA

[WVBR^{+}21] Alexander Wilkie, Petr Vevoda, Thomas Bashford-Rogers, Lukáš Hošek, Tomáš Iser, Monika Kolářová, Tobias Rittig, and Jaroslav Křivánek. A fitted radiance and attenuation model for realistic atmospheres. ACM Transactions on Graphics, 40(4): 135:1–135:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459758.

Werner:2017:SIW

[WVJH17] Sebastian Werner, Zdravko Velinov, Wenzel Jakob, and Matthias B. Hullin. Scratch iridescence: wave-optical rendering of diffractive surface structure. ACM Transactions on Graphics, 36(6): 207:1–207:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2022:DSI

[WVY^{+}22] Ying Wang, Jasper Verheul, Sang-Hoon Yeo, Nima Khademi Kalantari, and Shinjiro Sueda. Differentiable simulation of inertial musculotendons. ACM Transactions on Graphics, 41 (6):272:1–272:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555490.

Whitted:1982:STD

[WW82] T. Whitted and D. M. Weimer. A software testbed for the development of 3D raster graphics systems. ACM Transactions on Graphics, 1 (1):43–57, January 1982. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weidlich:2008:RRB

[WW08] Andrea Weidlich and Alexander Wilkie. Realistic rendering of birefringency in uniaxial crystals. ACM Transactions on Graphics, 27(1): 6:1–6:12, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wei:2011:DDA

[WW11] Li-Yi Wei and Rui Wang. Differential domain analysis for non-uniform sampling. ACM Transactions on Graphics, 30 (4):50:1–50:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Willis:2013:IFI

[WW13] Karl D. D. Willis and Andrew D. Wilson. InfraStructs: fabricating information inside physical objects for imaging in the terahertz region. ACM Transactions on Graphics, 32 (4):138:1–138:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weber:2016:RDP

[WWA^{+}16] Nicolas Weber, Michael Waechter, Sandra C. Amend, Stefan Guthe, and Michael Goesele. Rapid, detail-preserving image downscaling. ACM Transactions on Graphics, 35(6): 205:1–205:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wald:2014:EKF

[WWB^{+}14] Ingo Wald, Sven Woop, Carsten Benthin, Gregory S. Johnson, and Manfred Ernst. Embree: a kernel framework for efficient CPU ray tracing. ACM Transactions on Graphics, 33(4):143:1–143:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2019:REA

[WWB^{+}19] Ying Wang, Nicholas J. Weidner, Margaret A. Baxter, Yura Hwang, Danny M. Kaufman, and Shinjiro Sueda. RedMax: efficient & flexible approach for articulated dynamics. ACM Transactions on Graphics, 38(4): 104:1–104:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2005:RTR

[WWD^{+}05] Lifeng Wang, Wenle Wang, Julie Dorsey, Xu Yang, Baining Guo, and Heung-Yeung Shum. Real-time rendering of plant leaves. ACM Transactions on Graphics, 24(3): 712–719, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2010:RSS

[WWF^{+}10] Huisi Wu, Yu-Shuen Wang, Kun-Chuan Feng, Tien-Tsin Wong, Tong-Yee Lee, and Pheng-Ann Heng. Resizing by symmetry-summarization. ACM Transactions on Graphics, 29(6):159:1–159:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Watson:2004:STC

[WWH04] Benjamin Watson, Neff Walker, and Larry F. Hodges. Suprathreshold control of peripheral LOD. ACM Transactions on Graphics, 23(3): 750–759, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2006:DCI

[WWH06] Guangyu Wang, Tien-Tsin Wong, and Pheng-Ann Heng. Deringing cartoons by image analogies. ACM Transactions on Graphics, 25(4): 1360–1379, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2019:SSA

[WWL^{+}19] Zhijie Wu, Xiang Wang, Di Lin, Dani Lischinski, Daniel Cohen-Or, and Hui Huang. SAGNet: structureaware generative network for 3D-shape modeling. ACM Transactions on Graphics, 38 (4):91:1–91:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2021:LCR

[WWLC21] Yu-Chen Wang, Yu-Ting Wu, Tzu-Mao Li, and Yung-Yu Chuang. Learning to cluster for rendering with many lights. ACM Transactions on Graphics, 40(6): 277:1–277:10, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480561.

Wang:2008:FRC

[WWOH08] Huamin Wang, Yonatan Wexler, Eyal Ofek, and Hugues Hoppe. Factoring repeated content within and among images. ACM Transactions on Graphics, 27(3): 14:1–14:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2005:CTA

[WWS^{+}05] Hongcheng Wang, Qing Wu, Lin Shi, Yizhou Yu, and Narendra Ahuja. Out-of-core tensor approximation of multi-dimensional matrices of visual data. ACM Transactions on Graphics, 24(3): 527–535, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2023:RSP

[WWSP23] Bolun Wang, Hui Wang, Eike Schling, and Helmut Pottmann. Rectifying strip patterns. ACM Transactions on Graphics, 42(6): 256:1–256:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618378.

Wonka:2003:IA

[WWSR03] Peter Wonka, Michael Wimmer, François Sillion, and William Ribarsky. Instant architecture. ACM Transactions on Graphics, 22(3): 669–677, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2003:VDD

[WWT^{+}03] Lifeng Wang, Xi Wang, Xin Tong, Stephen Lin, Shimin Hu, Baining Guo, and Heung-Yeung Shum. View-dependent displacement mapping. ACM Transactions on Graphics, 22 (3):334–339, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2006:ESS

[WWT^{+}06] Ke Wang, Weiwei, Yiying Tong, Mathieu Desbrun, and Peter Schröder. Edge subdivision schemes and the construction of smooth vector fields. ACM Transactions on Graphics, 25(3): 1041–1048, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2022:GBM

[WWW22] Botao Wu, Zhendong Wang, and Huamin Wang. A GPU-based multilevel additive Schwarz preconditioner for cloth and deformable body simulation. ACM Transactions on Graphics, 41(4):63:1–63:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530085.

Wang:2022:CMA

[WWWG22] Ningna Wang, Bin Wang, Wenping Wang, and Xiaohu Guo. Computing medial axis transform with feature preservation via restricted power diagram. ACM Transactions on Graphics, 41(6): 188:1–188:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555465.

Wang:2023:ASP

[WWWZ23] Yu-Chen Wang, Chris Wyman, Lifan Wu, and Shuang Zhao. Amortizing samples in physics-based inverse rendering using ReSTIR. ACM Transactions on Graphics, 42 (6):214:1–214:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618331.

Wang:2022:RDT

[WWX^{+}22] Pengfei Wang, Zixiong Wang, Shiqing Xin, Xifeng Gao, Wenping Wang, and Changhe Tu. Restricted Delaunay triangulation for explicit surface reconstruction. ACM Transactions on Graphics, 41(5): 180:1–180:??, October 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3533768.

Wang:2013:CEP

[WWY^{+}13] Weiming Wang, Tuanfeng Y.

Wang, Zhouwang Yang, Ligang Liu, Xin Tong, Weihua Tong, Jiansong Deng, Falai Chen, and Xiuping Liu. Costeffective printing of 3D objects with skin-frame structures. ACM Transactions on Graphics, 32(6):177:1– 177:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2015:DCM

[WWY^{+}15] Bin Wang, Longhua Wu, KangKang Yin, Uri Ascher, Libin Liu, and Hui Huang. Deformation capture and modeling of soft objects. ACM Transactions on Graphics, 34(4):94:1–94:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2021:SFR

[WWYW21] Longhua Wu, Botao Wu, Yin Yang, and Huamin Wang. A safe and fast repulsion method for GPU-based cloth self collisions. ACM Transactions on Graphics, 40 (1):5:1–5:18, January 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3430025.

Wonka:2006:GVS

[WWZ^{+}06] Peter Wonka, Michael Wimmer, Kaichi Zhou, Stefan Maierhofer, Gerd Hesina, and Alexander Reshetov. Guided visibility sampling. ACM Transactions on Graphics, 25 (3):494–502, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:EGB

[WWZ^{+}09] Rui Wang, Rui Wang, Kun Zhou, Minghao Pan, and Hujun Bao. An efficient GPUbased approach for interactive global illumination. ACM Transactions on Graphics, 28 (3):91:1–91:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Walton:1991:TPP

[WX91] D. J. Walton and R. Xu. Turning point preserving planar interpolation. ACM Transactions on Graphics, 10(3):297–311, July 1991. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
108548.html.

Wen:2017:RTE

[WXLY17] Quan Wen, Feng Xu, Ming Lu, and Jun-Hai Yong. Realtime 3D eyelids tracking from semantic edges. ACM Transactions on Graphics, 36(6): 193:1–193:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2004:VT

[WXSC04] Jue Wang, Yingqing Xu, Heung-Yeung Shum, and Michael F. Cohen. Video tooning. ACM Transactions on Graphics, 23(3):574– 583, August 2004. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2022:SNI

[WXZ^{+}22] Xiuchao Wu, Jiamin Xu, Zihan Zhu, Hujun Bao, Qixing Huang, James Tompkin, and Weiwei Xu. Scalable neural indoor scene rendering. ACM Transactions on Graphics, 41(4):98:1–98:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530153.

Wu:2023:SSB

[WXZ^{+}23] Xiuchao Wu, Jiamin Xu, Xin Zhang, Hujun Bao, Qixing Huang, Yujun Shen, James Tompkin, and Weiwei Xu. ScaNeRF: Scalable bundle-adjusting neural radiance fields for large-scale scene rendering. ACM Transactions on Graphics, 42(6): 261:1–261:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618369.

Wu:2004:FMD

[WY04] Qing Wu and Yizhou Yu. Feature matching and deformation for texture synthesis. ACM Transactions on Graphics, 23(3):364–367, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2016:DME

[WY16] Huamin Wang and Yin Yang. Descent methods for elastic body simulation on the GPU. ACM Transactions on Graphics, 35(6):212:1– 212:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2014:IPM

[WYD^{+}14] Fuzhang Wu, Dong-Ming Yan, Weiming Dong, Xiaopeng Zhang, and Peter Wonka. Inverse procedural modeling of facade layouts. ACM Transactions on Graphics, 33(4): 121:1–121:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2014:DNF

[WYL^{+}14] Ruimin Wang, Zhouwang Yang, Ligang Liu, Jiansong Deng, and Falai Chen. Decoupling noise and features via weighted l_{1}-analysis compressed sensing. ACM Transactions on Graphics, 33(2): 18:1–18:??, March 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2020:SIP

[WYL^{+}20] Zhibo Wang, Xin Yu, Ming Lu, Quan Wang, Chen Qian, and Feng Xu. Single image portrait relighting via explicit multiple reflectance channel modeling. ACM Transactions on Graphics, 39(6): 220:1–220:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417824.

Wyman:2005:AIS

[Wym05]  Chris Wyman.  An ap-

proximate image-space approach for interactive refraction. ACM Transactions on Graphics, 24(3): 1050–1053, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2016:RTR

[WYM^{+}16] Rui Wang, Bowen Yu, Julio Marco, Tianlei Hu, Diego Gutierrez, and Hujun Bao. Real-time rendering on a power budget. ACM Transactions on Graphics, 35(4): 111:1–111:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2010:DDI

[WYW^{+}10] Baoyuan Wang, Yizhou Yu, Tien-Tsin Wong, Chun Chen, and Ying-Qing Xu. Datadriven image color theme enhancement. ACM Transactions on Graphics, 29(6): 146:1–146:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2023:SDB

[WYW23] Zhendong Wang, Yin Yang, and Huamin Wang. Stable discrete bending by analytic eigensystem and adaptive orthotropic geometric stiffness. ACM Transactions on Graphics, 42(6): 175:1–175:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618372.

Wang:2011:EBI

[WYX11] Baoyuan Wang, Yizhou Yu, and Ying-Qing Xu. Examplebased image color and tone style enhancement. ACM Transactions on Graphics, 30 (4):64:1–64:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2021:CFF

[WYXJ21] Yiqian Wu, Yong-Liang Yang, Qinjie Xiao, and Xiaogang Jin. Coarse-to-fine: facial structure editing of portrait images via latent space classifications. ACM Transactions on Graphics, 40(4): 46:1–46:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459814.

Wang:2014:ASS

[WYY^{+}14] Rui Wang, Xianjin Yang, Yazhen Yuan, Wei Chen, Kavita Bala, and Hujun Bao. Automatic shader simplification using surface signal approximation. ACM Transactions on Graphics, 33(6): 226:1–226:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2009:EBH

[WYZG09] Lvdi Wang, Yizhou Yu, Kun Zhou, and Baining Guo.

Example-based hair geometry synthesis. ACM Transactions on Graphics, 28(3): 56:1–56:??, August 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2011:MVV

[WYZG11] Lvdi Wang, Yizhou Yu, Kun Zhou, and Baining Guo. Multiscale vector volumes. ACM Transactions on Graphics, 30 (6):167:1–167:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weber:2014:LIP

[WZ14] Ofir Weber and Denis Zorin. Locally injective parametrization with arbitrary fixed boundaries. ACM Transactions on Graphics, 33(4): 75:1–75:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2022:LGS

[WZ22] Rundi Wu and Changxi Zheng. Learning to generate 3D shapes from a single example. ACM Transactions on Graphics, 41(6): 224:1–224:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555480.

Wang:2017:BMB

[WZB17] Bohan Wang, Yili Zhao, and Jernej Barbic. Botanical materials based on biomechanics. ACM Transactions on Graphics, 36(4):135:1–135:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wei:2012:ARF

[WZC12] Xiaolin Wei, Peizhao Zhang, and Jinxiang Chai. Accurate realtime full-body motion capture using a single depth camera. ACM Transactions on Graphics, 31(6): 188:1–188:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wisessing:2020:EMI

[WZC^{+}20] Pisut Wisessing, Katja Zibrek, Douglas W. Cunningham, John Dingliana, and Rachel McDonnell. Enlighten me: Importance of brightness and shadow for character emotion and appeal. ACM Transactions on Graphics, 39(3): 19:1–19:12, June 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3383195.

Winberg:2022:FHT

[WZC^{+}22] Sebastian Winberg, Gaspard Zoss, Prashanth Chandran, Paulo Gotardo, and Derek Bradley. Facial hair tracking for high fidelity performance capture. ACM Transactions on Graphics, 41 (4):165:1–165:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530116.

Wang:2018:LGL

[WZF^{+}18] Xiaogang Wang, Bin Zhou, Haiyue Fang, Xiaowu Chen, Qinping Zhao, and Kai Xu. Learning to group and label fine-grained shape components. ACM Transactions on Graphics, 37(6):210:1– 210:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Walter:2009:SSR

[WZHB09] Bruce Walter, Shuang Zhao, Nicolas Holzschuch, and Kavita Bala. Single scattering in refractive media with triangle mesh boundaries. ACM Transactions on Graphics, 28 (3):92:1–92:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2023:LBP

[WZHL23] Zhenwei Wang, Nanxuan Zhao, Gerhard Hancke, and Rynson W. H. Lau. Languagebased photo color adjustment for graphic designs. ACM Transactions on Graphics, 42(4):101:1–101:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592111.

Wang:2017:LFV

[WZK^{+}17] Ting-Chun Wang, Jun-Yan Zhu, Nima Khademi Kalantari, Alexei A. Efros, and Ravi Ramamoorthi. Light field video capture using a learning-based hybrid imaging system. ACM Transactions on Graphics, 36(4): 133:1–133:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Weier:2023:NPC

[WZK^{+}23] Philippe Weier, Tobias Zirr, Anton Kaplanyan, Ling-Qi Yan, and Philipp Slusallek. Neural prefiltering for correlationaware levels of detail. ACM Transactions on Graphics, 42 (4):78:1–78:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592443.

Wu:2020:WID

[WZL^{+}20] Rundong Wu, Joy Xiaoji Zhang, Jonathan Leaf, Xinru Hua, Ante Qu, Claire Harvey, Emily Holtzman, Joy Ko, Brooks Hagan, Doug James, François Guimbretière, and Steve Marschner. Weavecraft: an interactive design and simulation tool for 3D weaving. ACM Transactions on Graphics, 39(6): 210:1–210:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417865.

Wolski:2022:DSI

[WZMM22] Krzysztof Wolski, Fangcheng Zhong, Karol Myszkowski, and Rafał K. Mantiuk. Dark stereo: improving depth perception under low luminance. ACM Transactions on Graphics, 41(4):146:1–146:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3528223.3530136.

Wu:2014:RTS

[WZN^{+}14] Chenglei Wu, Michael Zollhöfer, Matthias Nießner, Marc Stamminger, Shahram Izadi, and Christian Theobalt. Realtime shading-based refinement for consumer depth cameras. ACM Transactions on Graphics, 33(6):200:1– 200:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2018:FRT

[WZQ^{+}18] Bojian Wu, Yang Zhou, Yiming Qian, Minglun Cong, and Hui Huang. Full 3D reconstruction of transparent objects. ACM Transactions on Graphics, 37(4): 103:1–103:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2008:MRH

[WZT^{+}08a] Jiaping Wang, Shuang Zhao, Xin Tong, Stephen Lin, Zhouchen Lin, Yue Dong, Baining Guo, and Heung-Yeung Shum. Modeling and rendering of heterogeneous translucent materials using the diffusion equation. ACM Transactions on Graphics, 27 (1):9:1–9:19, March 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2008:MAS

[WZT^{+}08b] Jiaping Wang, Shuang Zhao, Xin Tong, John Snyder, and Baining Guo. Modeling anisotropic surface reflectance with example-based microfacet synthesis. ACM Transactions on Graphics, 27(3): 41:1–41:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wang:2023:NSH

[WZX^{+}23] Zixiong Wang, Yunxiao Zhang, Rui Xu, Fan Zhang, Peng-Shuai Wang, Shuangmin Chen, Shiqing Xin, Wenping Wang, and Changhe Tu. Neural-singular-Hessian:

Implicit neural representation of unoriented point clouds by enforcing singular Hessian. ACM Transactions on Graphics, 42(6): 274:1–274:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3618311.

Wang:2010:VST

[WZYG10] Lvdi Wang, Kun Zhou, Yizhou Yu, and Baining Guo. Vector solid textures. ACM Transactions on Graphics, 29 (4):86:1–86:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Wu:2019:AAP

[WZYR19] Lifan Wu, Shuang Zhao, Ling-Qi Yan, and Ravi Ramamoorthi. Accurate appearance preserving prefiltering for rendering displacement-mapped surfaces. ACM Transactions on Graphics, 38(4): 137:1–137:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xue:2012:UIR

[XADR12] Su Xue, Aseem Agarwala, Julie Dorsey, and Holly Rushmeier. Understanding and improving the realism of image composites. ACM Transactions on Graphics, 31(4): 84:1–84:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xue:2023:IWS

[XAW^{+}23] Kangrui Xue, Ryan M. Aronson, Jui-Hsien Wang, Timothy R. Langlois, and Doug L. James. Improved water sound synthesis using coupled bubbles. ACM Transactions on Graphics, 42(4): 127:1–127:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592424.

Xu:2016:PSS

[XB16] Hongyi Xu and Jernej Barbic. Pose-space subspace dynamics. ACM Transactions on Graphics, 35(4): 35:1–35:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2017:EBD

[XB17] Hongyi Xu and Jernej Barbic. Example-based damping design. ACM Transactions on Graphics, 36(4): 53:1–53:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2023:WAR

[XBLZ23] Peiyu Xu, Sai Bangaru, Tzu-Mao Li, and Shuang Zhao.

Warped-area reparameterization of differential path integrals. ACM Transactions on Graphics, 42(6): 213:1–213:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618330.

Xu:2019:DVS

[XBS^{+}19] Zexiang Xu, Sai Bi, Kalyan Sunkavalli, Sunil Hadap, Hao Su, and Ravi Ramamoorthi. Deep view synthesis from sparse photometric images. ACM Transactions on Graphics, 38(4): 76:1–76:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiang:2022:DAD

[XBS^{+}22] Donglai Xiang, Timur Bagautdinov, Tuur Stuyck, Fabian Prada, Javier Romero, Weipeng Xu, Shunsuke Saito, Jingfan Guo, Breannan Smith, Takaaki Shiratori, Yaser Sheikh, Jessica Hodgins, and Chenglei Wu. Dressing avatars: Deep photorealistic appearance for physically simulated clothing. ACM Transactions on Graphics, 41(6): 222:1–222:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555456.

Xiao:2019:VMM

[XBZN19] Chang Xiao, Karl Bayer, Changxi Zheng, and Shree K. Nayar. Vidgets: modular mechanical widgets for mobile devices. ACM Transactions on Graphics, 38(4): 100:1–100:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2013:SSB

[XCF^{+}13] Kun Xu, Kang Chen, Hongbo Fu, Wei-Lun Sun, and Shi-Min Hu. Sketch2scene: sketch-based co-retrieval and co-placement of 3D models. ACM Transactions on Graphics, 32(4):123:1–123:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2014:CHF

[XCLT14] Feng Xu, Jinxiang Chai, Yilong Liu, and Xin Tong. Controllable high-fidelity facial performance transfer. ACM Transactions on Graphics, 33 (4):42:1–42:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2014:PAR

[XCM^{+}14] Kun Xu, Yan-Pei Cao, Li-Qian Ma, Zhao Dong, Rui Wang, and Shi-Min Hu. A practical algorithm for rendering interreflections with allfrequency BRDFs. ACM Transactions on Graphics, 33 (1):10:1–10:16, January 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2009:FAS

[XCOJ^{+}09] Kai Xu, Daniel Cohen-Or, Tao Ju, Ligang Liu, Hao Zhang, Shizhe Zhou, and Yueshan Xiong. Feature-aligned shape texturing. ACM Transactions on Graphics, 28(5): 108:1–108:7, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2014:TCN

[XCS^{+}14] Baoxuan Xu, William Chang, Alla Sheffer, Adrien Bousseau, James McCrae, and Karan Singh. True2Form: 3D curve networks from 2D sketches via selective regularization. ACM Transactions on Graphics, 33 (4):131:1–131:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xing:2014:APR

[XCW14] Jun Xing, Hsiang-Ting Chen, and Li-Yi Wei. Autocomplete painting repetitions. ACM Transactions on Graphics, 33 (6):172:1–172:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2020:AST

[XCW^{+}20] Yuwei Xiao, Szeyu Chan, Siqi Wang, Bo Zhu, and Xubo Yang. An adaptive staggeredtilted grid for incompressible flow simulation. ACM Transactions on Graphics, 39(6): 171:1–171:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417837.

Xu:2018:MHP

[XCZ^{+}18] Weipeng Xu, Avishek Chatterjee, Michael Zollhöfer, Helge Rhodin, Dushyant Mehta, Hans-Peter Seidel, and Christian Theobalt. MonoPerfCap: Human performance capture from monocular video. ACM Transactions on Graphics, 37(2): 27:1–27:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2019:LFC

[XDF^{+}19] Jie Xu, Tao Du, Michael Foshey, Beichen Li, Bo Zhu, Adriana Schulz, and Wojciech Matusik. Learning to fly: computational controller design for hybrid UAVs with reinforcement learning. ACM Transactions on Graphics, 38(4): 42:1–42:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xia:2016:RSS

[XDPT16] Rui Xia, Yue Dong, Pieter Peers, and Xin Tong. Recovering shape and spatiallyvarying surface reflectance under unknown illumination. ACM Transactions on Graphics, 35(6):187:1– 187:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2023:GCN

[XDW^{+}23] Rui Xu, Zhiyang Dou, Ningna Wang, Shiqing Xin, Shuangmin Chen, Mingyan Jiang, Xiaohu Guo, Wenping Wang, and Changhe Tu. Globally consistent normal orientation for point clouds by regularizing the windingnumber field. ACM Transactions on Graphics, 42(4): 111:1–111:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592129.

Xu:2012:LSS

[XFAT12] Pengfei Xu, Hongbo Fu, Oscar Kin-Chung Au, and Chiew-Lan Tai. Lazy selection: a scribble-based tool for smart shape elements selection. ACM Transactions on Graphics, 31(6):142:1– 142:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xie:2018:TTC

[XFCT18] You Xie, Erik Franz, Mengyu Chu, and Nils Thuerey. tempoGAN: a temporally coherent, volumetric GAN for super-resolution fluid flow. ACM Transactions on Graphics, 37(4):95:1–95:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2008:IBF

[XFT^{+}08] Jianxiong Xiao, Tian Fang, Ping Tan, Peng Zhao, Eyal Ofek, and Long Quan. Imagebased façade modeling. ACM Transactions on Graphics, 27 (5):161:1–161:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2009:IBS

[XFZ^{+}09] Jianxiong Xiao, Tian Fang, Peng Zhao, Maxime Lhuillier, and Long Quan. Imagebased street-side city modeling. ACM Transactions on Graphics, 28(5):114:1– 114:12, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2007:KHB

[XGC07] Hui Xu, Nathan Gossett, and Baoquan Chen. Knowledge and heuristic-based modeling of laser-scanned trees. ACM Transactions on Graphics, 26 (4):19:1–19:13, October 2007.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2023:NPV

[XGZ^{+}23] Juzhan Xu, Minglun Gong, Hao Zhang, Hui Huang, and Ruizhen Hu. Neural packing: from visual sensing to reinforcement learning. ACM Transactions on Graphics, 42 (6):267:1–267:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618354.

Xie:2018:MSD

[XH18] Feng Xie and Pat Hanrahan. Multiple scattering from distributions of specular v-grooves. ACM Transactions on Graphics, 37(6): 276:1–276:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2015:ACS

[XHS^{+}15] Kai Xu, Hui Huang, Yifei Shi, Hao Li, Pinxin Long, Jianong Caichen, Wei Sun, and Baoquan Chen. Autoscanning for coupled scene reconstruction and proactive object analysis. ACM Transactions on Graphics, 34(6): 177:1–177:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xia:2022:DIC

[XHWW22] Menghan Xia, Wenbo Hu, Tien-Tsin Wong, and Jue Wang. Disentangled image colorization via global anchors. ACM Transactions on Graphics, 41(6): 204:1–204:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555432.

Xiang:1997:CIQ

[Xia97] Zhigang Xiang. Color image quantization by minimizing the maximum intercluster distance. ACM Transactions on Graphics, 16(3):260–276, July 1997. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
1997-16-3/p260-xiang/.

Xiang:2021:ETE

[Xia21] Sitao Xiang. Eliminating topological errors in neural network rotation estimation using self-selecting ensembles. ACM Transactions on Graphics, 40(4): 167:1–167:21, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459882.

Xiong:2017:RPI

[XIAP^{+}17] Jinhui Xiong, Ramzi Idoughi, Andres A. Aguirre-Pablo, Abdulrahman B. Aljedaani, Xiong Dun, Qiang Fu, Sigurdur T. Thoroddsen, and Wolfgang Heidrich. Rainbow particle imaging velocimetry for dense 3D fluid velocity imaging. ACM Transactions on Graphics, 36(4): 36:1–36:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xie:2018:PPS

[XIM18] Haoran Xie, Takeo Igarashi, and Kazunori Miyata. Precomputed panel solver for aerodynamics simulation. ACM Transactions on Graphics, 37 (2):17:1–17:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2007:IGM

[XK07] Jie Xu and Craig S. Kaplan. Image-guided maze construction. ACM Transactions on Graphics, 26(3): 29:1–29:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2018:BID

[XKCB18] Hongyi Xu, Espen Knoop, Stelian Coros, and Moritz Bächer. Bend-it: design and fabrication of kinetic wire characters. ACM Transactions on Graphics, 37(6): 239:1–239:??, November 2018.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2018:DLI

[XKF^{+}18] Lei Xiao, Anton Kaplanyan, Alexander Fix, Matthew Chapman, and Douglas Lanman. DeepFocus: learned image synthesis for computational displays. ACM Transactions on Graphics, 37(6): 200:1–200:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xin:2016:CPD

[XLC^{+}16] Shi-Qing Xin, Bruno Lévy, Zhonggui Chen, Lei Chu, Yaohui Yu, Changhe Tu, and Wenping Wang. Centroidal power diagrams with capacity constraints: computation, applications, and extension. ACM Transactions on Graphics, 35(6):244:1– 244:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiong:2023:EET

[XLC^{+}23] Ruicheng Xiong, Yang Lu, Cong Chen, Jiaming Zhu, Yajun Zeng, and Ligang Liu. ETER: Elastic tessellation for real-time pixel-accurate rendering of large-scale NURBS models. ACM Transactions on Graphics, 42(4): 133:1–133:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592419.

Xu:2015:IMD

[XLCB15] Hongyi Xu, Yijing Li, Yong Chen, and Jernej Barbivc. Interactive material design using model reduction. ACM Transactions on Graphics, 34 (2):18:1–18:??, February 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xin:2011:MBP

[XLF^{+}11] Shiqing Xin, Chi-Fu Lai, Chi-Wing Fu, Tien-Tsin Wong, Ying He, and Daniel Cohen-Or. Making burr puzzles from 3D models. ACM Transactions on Graphics, 30(4): 97:1–97:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2009:EAB

[XLJ^{+}09] Kun Xu, Yong Li, Tao Ju, Shi-Min Hu, and Tian-Qiang Liu. Efficient affinity-based edit propagation using K-D tree. ACM Transactions on Graphics, 28(5):118:1– 118:6, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xie:2020:MFS

[XLLW20] Minshan Xie, Chengze Li, Xueting Liu, and Tien-Tsin Wong. Manga filling style conversion with screentone variational autoencoder. ACM Transactions on Graphics, 39 (6):226:1–226:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417873.

Xu:2011:VBC

[XLS^{+}11] Feng Xu, Yebin Liu, Carsten Stoll, James Tompkin, Gaurav Bharaj, Qionghai Dai, Hans-Peter Seidel, Jan Kautz, and Christian Theobalt. Videobased characters: creating new human performances from a multi-view video database. ACM Transactions on Graphics, 30(4): 32:1–32:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xia:2018:IG

[XLW18] Menghan Xia, Xueting Liu, and Tien-Tsin Wong. Invertible grayscale. ACM Transactions on Graphics, 37(6): 246:1–246:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2016:IMM

[XLX^{+}16] Mingliang Xu, Mingyuan Li, Weiwei Xu, Zhigang Deng, Yin Yang, and Kun Zhou. Interactive mechanism modeling from multi-view images. ACM Transactions on Graphics, 35 (6):236:1–236:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2011:ISG

[XLXJ11] Li Xu, Cewu Lu, Yi Xu, and Jiaya Jia. Image smoothing via L_{0} gradient minimization. ACM Transactions on Graphics, 30(6):174:1– 174:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xia:2009:PBI

[XLY09] Tian Xia, Binbin Liao, and Yizhou Yu. Patch-based image vectorization with automatic curvilinear feature alignment. ACM Transactions on Graphics, 28(5): 115:1–115:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xing:2022:DRU

[XLY^{+}22a] Jiankai Xing, Fujun Luan, Ling-Qi Yan, Xuejun Hu, Houde Qian, and Kun Xu. Differentiable rendering using RGBXY derivatives and optimal transport. ACM Transactions on Graphics, 41(6): 189:1–189:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555479.

Xu:2022:HLB

[XLY^{+}22b] Pengfei Xu, Yifan Li, Zhijin Yang, Weiran Shi, Hongbo Fu, and Hui Huang. Hierarchical layout blending with recursive optimal correspondence. ACM Transactions on Graphics, 41(6): 249:1–249:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555446.

Xie:2023:CPS

[XLYJ23] Tianyi Xie, Minchen Li, Yin Yang, and Chenfanfu Jiang. A contact proxy splitting method for Lagrangian solidfluid coupling. ACM Transactions on Graphics, 42(4): 122:1–122:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592115.

Xu:2010:SCS

[XLZ^{+}10] Kai Xu, Honghua Li, Hao Zhang, Daniel Cohen-Or, Yueshan Xiong, and Zhi-Quan Cheng. Style-content separation by anisotropic part scales. ACM Transactions on Graphics, 29(6):184:1– 184:??, December 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2011:IHR

[XMR^{+}11] Kun Xu, Li-Qian Ma, Bo Ren, Rui Wang, and Shi-Min Hu. Interactive hair rendering and appearance editing under environment lighting. ACM Transactions on Graphics, 30 (6):173:1–173:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2014:OHS

[XMZ^{+}14] Kai Xu, Rui Ma, Hao Zhang, Chenyang Zhu, Ariel Shamir, Daniel Cohen-Or, and Hui Huang. Organizing heterogeneous scene collections through contextual focal points. ACM Transactions on Graphics, 33(4): 35:1–35:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2016:MBS

[XNY^{+}16] Zexiang Xu, Jannik Boll Nielsen, Jiyang Yu, Henrik Wann Jensen, and Ravi Ramamoorthi. Minimal BRDF sampling for two-shot near-field reflectance acquisition. ACM Transactions on Graphics, 35(6):188:1– 188:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2022:HFA

[XNZ^{+}22] Yabin Xu, Liangliang Nan, Laishui Zhou, Jun Wang, and Charlie C. L. Wang. HRBF-Fusion: Accurate 3D reconstruction from RGB-D data using on-the-fly implicits. ACM Transactions on Graphics, 41(3):35:1–35:19, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3516521.

Xiang:2021:MCS

[XPB^{+}21] Donglai Xiang, Fabian Prada, Timur Bagautdinov, Weipeng Xu, Yuan Dong, He Wen, Jessica Hodgins, and Chenglei Wu. Modeling clothing as a separate layer for an animatable human avatar. ACM Transactions on Graphics, 40 (6):199:1–199:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3478513.3480545.

Xue:2015:CAO

[XRLF15] Tianfan Xue, Michael Rubinstein, Ce Liu, and William T. Freeman. A computational approach for obstruction-free photography. ACM Transactions on Graphics, 34(4): 79:1–79:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xing:2022:PBS

[XRW^{+}22] Jingrui Xing, Liangwang Ruan, Bin Wang, Bo Zhu, and Baoquan Chen. Positionbased surface tension flow. ACM Transactions on Graphics, 41(6):244:1–244:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555476.

Xu:2013:ASG

[XSD^{+}13] Kun Xu, Wei-Lun Sun, Zhao Dong, Dan-Yong Zhao, Run-Dong Wu, and Shi-Min Hu. Anisotropic spherical Gaussians. ACM Transactions on Graphics, 32(6):209:1– 209:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xue:2020:NDN

[XSH^{+}20] Tao Xue, Haozhe Su, Chengguizi Han, Chenfanfu Jiang, and Mridul Aanjaneya. A novel discretization and numerical solver for non–Fourier diffusion. ACM Transactions on Graphics, 39(6): 178:1–178:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417863.

Xu:2018:DIB

[XSHR18] Zexiang Xu, Kalyan Sunkavalli, Sunil Hadap, and Ravi Ramamoorthi. Deep imagebased relighting from optimal sparse samples. ACM Transactions on Graphics, 37(4): 126:1–126:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2022:DHD

[XSL^{+}22] Chufeng Xiao, Wanchao Su, Jing Liao, Zhouhui Lian, Yi-Zhe Song, and Hongbo Fu. DifferSketching: How differently do people sketch 3D objects? ACM Transactions on Graphics, 41(6): 264:1–264:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555493.

Xie:2014:HDC

[XSTN14] Guofu Xie, Xin Sun, Xin Tong, and Derek Nowrouzezahrai. Hierarchical diffusion curves for accurate automatic image vectorization. ACM Transactions on Graphics, 33(6): 230:1–230:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2016:ADD

[XSZ^{+}16] Kai Xu, Yifei Shi, Lintao Zheng, Junyu Zhang, Min Liu, Hui Huang, Hao Su, Daniel Cohen-Or, and Baoquan Chen. 3D attentiondriven depth acquisition for object identification. ACM Transactions on Graphics, 35 (6):238:1–238:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2015:NMD

[XSZB15] Hongyi Xu, Funshing Sin, Yufeng Zhu, and Jernej Barbic. Nonlinear material design using principal stretches. ACM Transactions on Graphics, 34(4):75:1–75:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2023:CML

[XSZK23] Pei Xu, Xiumin Shang, Victor Zordan, and Ioannis Karamouzas. Composite motion learning with task control. ACM Transactions on Graphics, 42(4): 93:1–93:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592447.

Xiong:2021:IFS

[XTZ^{+}21] Shiying Xiong, Rui Tao, Yaorui Zhang, Fan Feng, and Bo Zhu. Incompressible flow simulation on vortex segment clouds. ACM Transactions on Graphics, 40 (4):98:1–98:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459865.

Xu:2018:SDG

[Xu18] Kai Xu. Session details: Get wired. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2014:SOR

[XUC^{+}14] Weiwei Xu, Nobuyuki Umentani, Qianwen Chao, Jie Mao, Xiaogang Jin, and Xin Tong. Sensitivity-optimized rigging for example-based real-time clothing synthesis. ACM Transactions on Graphics, 33 (4):107:1–107:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xin:2009:ICH

[XW09] Shi-Qing Xin and Guo-Jin Wang. Improving Chen and Han’s algorithm on the discrete geodesic problem. ACM Transactions on Graphics, 28 (4):104:1–104:8, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xin:2016:IGF

[XWC^{+}16] Shi-Qing Xin, Wenping Wang, Shuangmin Chen, Jieyu Zhao, and Zhenyu Shu. Intrinsic girth function for shape processing. ACM Transactions on Graphics, 35(3): 25:1–25:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xia:2015:RST

[XWCH15] Shihong Xia, Congyi Wang, Jinxiang Chai, and Jessica Hodgins. Realtime style transfer for unlabeled heterogeneous human motion. ACM Transactions on Graphics, 34(4):119:1–119:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2022:RRF

[XWD^{+}22] Rui Xu, Zixiong Wang, Zhiyang Dou, Chen Zong, Shiqing Xin, Mingyan Jiang, Tao Ju, and Changhe Tu. RFEPS: Reconstructing featureline equipped polygonal surface. ACM Transactions on Graphics, 41(6):228:1– 228:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555443.

Xia:2023:PWO

[XWH^{+}23] Mengqi Xia, Bruce Walter, Christophe Hery, Olivier Maury, Eric Michielssen, and Steve Marschner. A practical wave optics reflection model for hair and fur. ACM Transactions on Graphics, 42 (4):39:1–39:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592446.

Xu:2008:AAM

[XWL^{+}08] Xuemiao Xu, Liang Wan, Xiaopei Liu, Tien-Tsin Wong, Liansheng Wang, and Chi-Sing Leung. Animating animal motion from still. ACM Transactions on Graphics, 27 (5):117:1–117:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xia:2020:WOB

[XWM^{+}20] Mengqi (Mandy) Xia, Bruce Walter, Eric Michielssen, David Bindel, and Steve Marschner. A wave optics based fiber scattering model. ACM Transactions on Graphics, 39(6): 252:1–252:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417841.

Xing:2015:AHD

[XWSY15] Jun Xing, Li-Yi Wei, Takaaki Shiratori, and Koji Yatani. Autocomplete hand-drawn animations. ACM Transactions on Graphics, 34(6): 169:1–169:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2014:DHC

[XWW^{+}14] Zexiang Xu, Hsiang-Tao Wu, Lvdi Wang, Changxi Zheng, Xin Tong, and Yue Qi. Dynamic hair capture using spacetime optimization. ACM Transactions on Graphics, 33 (6):224:1–224:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiong:2022:CMF

[XWWZ22] Shiying Xiong, Zhecheng Wang, Mengdi Wang, and Bo Zhu. A Clebsch method for free-surface vortical flow simulation. ACM Transactions on Graphics, 41 (4):116:1–116:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530150.

Xin:2022:SEC

[XWX^{+}22] Shiqing Xin, Pengfei Wang, Rui Xu, Dongming Yan, Shuangmin Chen, Wenping Wang, Caiming Zhang, and Changhe Tu. SurfaceVoronoi: Efficiently computing Voronoi diagrams over mesh surfaces with arbitrary distance solvers. ACM Transactions on Graphics, 41(6): 185:1–185:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555453.

Xu:2009:JAM

[XWY^{+}09] Weiwei Xu, Jun Wang, KangKang Yin, Kun Zhou, Michiel van de Panne, Falai Chen, and Baining Guo.

Joint-aware manipulation of deformable models. ACM Transactions on Graphics, 28 (3):35:1–35:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2021:SIB

[XWZ^{+}21] Jiamin Xu, Xiuchao Wu, Zihan Zhu, Qixing Huang, Yin Yang, Hujun Bao, and Weiwei Xu. Scalable imagebased indoor scene rendering with reflections. ACM Transactions on Graphics, 40 (4):60:1–60:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459849.

Xu:2023:APA

[XXA^{+}23] Pei Xu, Kaixiang Xie, Sheldon Andrews, Paul G. Kry, Michael Neff, Morgan Mcguire, Ioannis Karamouzas, and Victor Zordan. AdaptNet: Policy adaptation for physicsbased character control. ACM Transactions on Graphics, 42 (6):177:1–177:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618375.

Xu:2006:ACP

[XXK^{+}06] Songhua Xu, Yingqing Xu, Sing Bing Kang, David H. Salesin, Yunhe Pan, and Heung-Yeung Shum. Animating Chinese paintings through stroke-based decomposition. ACM Transactions on Graphics, 25(2):239–267, April 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xie:2021:SMI

[XXL^{+}21] Minshan Xie, Menghan Xia, Xueting Liu, Chengze Li, and Tien-Tsin Wong. Seamless manga inpainting with semantics awareness. ACM Transactions on Graphics, 40 (4):96:1–96:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459822.

Xie:2018:CCC

[XYH^{+}18] Ke Xie, Hao Yang, Shengqiu Huang, Dani Lischinski, Marc Christie, Kai Xu, Minglun Gong, Daniel Cohen-Or, and Hui Huang. Creating and chaining camera moves for quadrotor videography. ACM Transactions on Graphics, 37 (4):88:1–88:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2021:SDS

[XYH^{+}21] Chufeng Xiao, Deng Yu, Xiaoguang Han, Youyi Zheng, and Hongbo Fu. SketchHair-Salon: deep sketch-based hair image synthesis. ACM Transactions on Graphics, 40(6): 216:1–216:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480502.

Xu:2013:SCM

[XYJ13] Li Xu, Qiong Yan, and Jiaya Jia. A sparse control model for image and video editing. ACM Transactions on Graphics, 32(6): 197:1–197:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2012:SET

[XYXJ12] Li Xu, Qiong Yan, Yang Xia, and Jiaya Jia. Structure extraction from texture via relative total variation. ACM Transactions on Graphics, 31 (6):139:1–139:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2012:FDS

[XZCOC12] Kai Xu, Hao Zhang, Daniel Cohen-Or, and Baoquan Chen. Fit and diverse: set evolution for inspiring 3D shape galleries. ACM Transactions on Graphics, 31(4): 57:1–57:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2012:MSP

[XZJ^{+}12] Kai Xu, Hao Zhang, Wei Jiang, Ramsay Dyer, Zhiquan Cheng, Ligang Liu, and Baoquan Chen. Multiscale partial intrinsic symmetry detection. ACM Transactions on Graphics, 31(6): 181:1–181:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2020:RNR

[XZK^{+}20] Zhan Xu, Yang Zhou, Evangelos Kalogerakis, Chris Landreth, and Karan Singh. RigNet: neural rigging for articulated characters. ACM Transactions on Graphics, 39 (4):58:1–58:??, July 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3386569.3392379.

Xiao:2018:CDT

[XZM^{+}18] Nan Xiao, Zhe Zhu, Ralph R. Martin, Kun Xu, Jia-Ming Lu, and Shi-Min Hu. Computational design of transforming pop-up books. ACM Transactions on Graphics, 37(1): 8:1–8:??, January 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiong:2023:TGA

[XZP^{+}23] Weidan Xiong, Hongqian Zhang, Botao Peng, Ziyu Hu, Yongli Wu, Jianwei Guo, and Hui Huang. TwinTex: Geometry-aware texture generation for abstracted 3D architectural models. ACM Transactions on Graphics, 42 (6):227:1–227:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618328.

Xu:2009:PIR

[XZT^{+}09] Kai Xu, Hao Zhang, Andrea Tagliasacchi, Ligang Liu, Guo Li, Min Meng, and Yueshan Xiong. Partial intrinsic reflectional symmetry of 3D shapes. ACM Transactions on Graphics, 28(5): 138:1–138:10, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2010:SBA

[XZW10] Xuemiao Xu, Linling Zhang, and Tien-Tsin Wong. Structurebased ASCII art. ACM Transactions on Graphics, 29(4): 52:1–52:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2007:GDE

[XZY^{+}07] Weiwei Xu, Kun Zhou, Yizhou Yu, Qifeng Tan, Qunsheng Peng, and Baining Guo. Gradient domain editing of deforming mesh sequences. ACM Transactions on Graphics, 26(3):84:1–84:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2017:ARU

[XZY^{+}17] Kai Xu, Lintao Zheng, Zihao Yan, Guohang Yan, Eugene Zhang, Matthias Niessner, Oliver Deussen, Daniel Cohen-Or, and Hui Huang. Autonomous reconstruction of unknown indoor scenes guided by time-varying tensor fields. ACM Transactions on Graphics, 36(6): 202:1–202:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xu:2011:PIM

[XZZ^{+}11] Kai Xu, Hanlin Zheng, Hao Zhang, Daniel Cohen-Or, Ligang Liu, and Yueshan Xiong. Photo-inspired model-driven 3D object modeling. ACM Transactions on Graphics, 30 (4):80:1–80:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiong:2014:RSR

[XZZ^{+}14] Shiyao Xiong, Juyong Zhangy, Jianmin Zheng, Jianfei Cai, and Ligang Liu. Robust surface reconstruction via dictionary learning. ACM Transactions on Graphics, 33(6): 201:1–201:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2018:FEI

[XZZ18] Chang Xiao, Cheng Zhang, and Changxi Zheng. Font-Code: Embedding information in text documents using glyph perturbation. ACM Transactions on Graphics, 37 (2):15:1–15:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Xiao:2021:EDB

[XZZ^{+}21] Qinjie Xiao, Hanyuan Zhang, Zhaorui Zhang, Yiqian Wu, Luyuan Wang, Xiaogang Jin, Xinwei Jiang, Yong-Liang Yang, Tianjia Shao, and Kun Zhou. EyelashNet: a dataset and a baseline method for eyelash matting. ACM Transactions on Graphics, 40(6): 217:1–217:17, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480540.

Yu:2022:PSS

[YAB^{+}22] Emilie Yu, Rahul Arora, J. Andreas Bærentzen, Karan Singh, and Adrien Bousseau. Piecewise-smooth surface fitting onto unstructured 3D sketches. ACM Transactions on Graphics, 41(4): 88:1–88:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530100.

Yang:2020:DDH

[YAV^{+}20] Jinfan Yang, Chrystiano Araujo, Nicholas Vining, Zachary Ferguson, Enrique Rosales, Daniele Panozzo, Sylvain Lefevbre, Paolo Cignoni, and Alla Sheffer. DHFSlicer: double height-field slicing for milling fixed-height materials. ACM Transactions on Graphics, 39(6): 205:1–205:17, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417810.

Yang:2022:ADP

[YBAF22] Yuting Yang, Connelly Barnes, Andrew Adams, and Adam Finkelstein. A δ:autodiff for discontinuous programs — applied to shaders. ACM Transactions on Graphics, 41 (4):135:1–135:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530125.

Yu:2023:VHD

[YBMN^{+}23] Emilie Yu, Kevin Blackburn-Matzen, Cuong Nguyen, Oliver Wang, Rubaiat Habib Kazi, and Adrien Bousseau. VideoDoodles: Hand-drawn animations on videos with scene-aware canvases. ACM Transactions on Graphics, 42 (4):54:1–54:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592413.

Yu:2021:RS

[YBSC21] Chris Yu, Caleb Brakensiek, Henrik Schumacher, and Keenan Crane. Repulsive surfaces. ACM Transactions on Graphics, 40(6): 268:1–268:19, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480521.

Yeh:2013:STP

[YBY^{+}13] Yi-Ting Yeh, Katherine Breeden, Lingfeng Yang, Matthew Fisher, and Pat Hanrahan. Synthesis of tiled patterns using factor graphs. ACM Transactions on Graphics, 32 (1):3:1–3:13, January 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2021:ULC

[YC21] Kaizhi Yang and Xuejin Chen. Unsupervised learning for cuboid shape abstraction via joint segmentation from point clouds. ACM Transactions on Graphics, 40(4): 152:1–152:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459873.

Yin:2008:CMA

[YCBvdP08] KangKang Yin, Stelian Coros, Philippe Beaudoin, and Michiel van de Panne. Continuation methods for adapting simulated skills. ACM Transactions on Graphics, 27(3):81:1– 81:??, August 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2017:CDT

[YCC17] Christopher Yu, Keenan Crane, and Stelian Coros. Computational design of telescoping structures. ACM Transactions on Graphics, 36 (4):83:1–83:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yumer:2015:SSE

[YCHK15] Mehmet Ersin Yumer, Siddhartha Chaudhuri, Jessica K. Hodgins, and Levent Burak Kara. Semantic shape editing using deformation handles. ACM Transactions on Graphics, 34(4): 86:1–86:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yao:2015:LSB

[YCL^{+}15] Miaojun Yao, Zhili Chen, Linjie Luo, Rui Wang, and Huamin Wang. Level-set-based partitioning and packing optimization of a printable model. ACM Transactions on Graphics, 34(6): 214:1–214:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2017:UPS

[YCL^{+}17] Tao Yang, Jian Chang, Ming C. Lin, Ralph R. Martin, Jian J. Zhang, and Shi-Min Hu. A unified particle system framework for multiphase, multi-material visual simulations. ACM Transactions on Graphics, 36(6): 224:1–224:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yoon:2020:SGG

[YCL^{+}20] Youngwoo Yoon, Bok Cha, Joo-Haeng Lee, Minsu Jang, Jaeyeon Lee, Jaehong Kim, and Geehyuk Lee. Speech gesture generation from the trimodal context of text, audio, and speaker identity. ACM Transactions on Graphics, 39 (6):222:1–222:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417838.

Yeung:2016:ICC

[YCP16] Yu-Hong Yeung, Jessica Crouch, and Alex Pothen. Interactively cutting and constraining vertices in meshes using augmented matrices. ACM Transactions on Graphics, 35(2):18:1–18:??, May 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2015:FMF

[YCR^{+}15] Tao Yang, Jian Chang, Bo Ren, Ming C. Lin, Jian Jun Zhang, and Shi-Min Hu. Fast multiple-fluid simulation using Helmholtz free energy. ACM Transactions on Graphics, 34(6): 201:1–201:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2020:ILS

[YCYW20] Guowei Yan, Zhili Chen, Jimei Yang, and Huamin Wang. Interactive liquid splash modeling by user sketches. ACM Transactions on Graphics, 39(6): 165:1–165:13, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417832.

Yuan:2011:PGS

[YCZ11] Zhi Yuan, Fan Chen, and Ye Zhao. Pattern-guided smoke animation with Lagrangian Coherent Structure. ACM Transactions on Graphics, 30(6):136:1– 136:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yao:2021:ITB

[YFFA21] Xinwei Yao, Ohad Fried, Kayvon Fatahalian, and Maneesh Agrawala. Iterative text-based editing of talking-heads using neural retargeting. ACM Transactions on Graphics, 40(3): 20:1–20:14, June 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yi:2017:LHS

[YGH^{+}17] Li Yi, Leonidas Guibas, Aaron Hertzmann, Vladimir G. Kim, Hao Su, and Ersin Yumer. Learning hierarchical shape segmentation and labeling from online repositories. ACM Transactions on Graphics, 36 (4):70:1–70:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ye:2014:IVA

[YGL^{+}14] Genzhi Ye, Elena Garces, Yebin Liu, Qionghai Dai, and Diego Gutierrez. Intrinsic video and applications. ACM Transactions on Graphics, 33 (4):80:1–80:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yun:1997:LCC

[YGM97] Hee Cheol Yun, Brian K. Guenter, and Russell M. Mersereau. Lossless compression of computer generated animation frames. ACM Transactions on Graphics, 16(4):359–396, October 1997. CODEN AT-GRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL http://www.
                                        acm.org:80/pubs/citations/
                                        journals/tog/1997-16-4/
                                        p359-yun/.

Yvart:2005:HTS

[YHB05] Alex Yvart, Stefanie Hahmann, and Georges-Pierre Bonneau. Hierarchical triangular splines. ACM Transactions on Graphics, 24(4): 1374–1391, October 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yin:2018:PNB

[YHCOZ18] Kangxue Yin, Hui Huang, Daniel Cohen-Or, and Hao Zhang. P2P-NET: bidirectional point displacement net for shape transform. ACM Transactions on Graphics, 37(4):152:1–152:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2014:RGH

[YHJ^{+}14] Ling-Qi Yan, Milos Hasan, Wenzel Jakob, Jason Lawrence, Steve Marschner, and Ravi Ramamoorthi. Rendering glints on high-resolution normal-mapped specular surfaces. ACM Transactions on Graphics, 33(4):116:1– 116:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuksel:2007:WP

[YHK07] Cem Yuksel, Donald H.

House, and John Keyser. Wave particles. ACM Transactions on Graphics, 26(3): 99:1–99:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yi:2018:DPI

[YHL^{+}18] Li Yi, Haibin Huang, Difan Liu, Evangelos Kalogerakis, Hao Su, and Leonidas Guibas. Deep part induction from articulated object pairs. ACM Transactions on Graphics, 37 (6):209:1–209:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2016:PND

[YHMR16] Ling-Qi Yan, Milos Hasan, Steve Marschner, and Ravi Ramamoorthi. Positionnormal distributions for efficient rendering of specular microstructure. ACM Transactions on Graphics, 35(4): 56:1–56:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2018:RSM

[YHW^{+}18] Ling-Qi Yan, Milos Hasan, Bruce Walter, Steve Marschner, and Ravi Ramamoorthi. Rendering specular microgeometry with wave optics. ACM Transactions on Graphics, 37 (4):75:1–75:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yin:2014:MIS

[YHZ^{+}14] Kangxue Yin, Hui Huang, Hao Zhang, Minglun Gong, Daniel Cohen-Or, and Baoquan Chen. Morfit: interactive surface reconstruction from incomplete point clouds with curve-driven topology and geometry control. ACM Transactions on Graphics, 33 (6):202:1–202:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yamamoto:2017:FPB

[YI17] Kazuhiko Yamamoto and Takeo Igarashi. Fully perceptual-based 3D spatial sound individualization with an adaptive variational autoencoder. ACM Transactions on Graphics, 36(6): 212:1–212:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yue:2010:UAS

[YIC^{+}10] Yonghao Yue, Kei Iwasaki, Bing-Yu Chen, Yoshinori Dobashi, and Tomoyuki Nishita. Unbiased, adaptive stochastic sampling for rendering inhomogeneous participating media. ACM Transactions on Graphics, 29(6): 177:1–177:??, December 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yue:2014:PBC

[YIC^{+}14] Yonghao Yue, Kei Iwasaki, Bing-Yu Chen, Yoshinori Dobashi, and Tomoyuki Nishita. Poisson-based continuous surface generation for goal-based caustics. ACM Transactions on Graphics, 33(3):31:1–31:??, May 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yoshida:2015:ASH

[YIO^{+}15] Hironori Yoshida, Takeo Igarashi, Yusuke Obuchi, Yosuke Takami, Jun Sato, Mika Araki, Masaaki Miki, Kosuke Nagata, Kazuhide Sakai, and Syunsuke Igarashi. Architecture-scale humanassisted additive manufacturing. ACM Transactions on Graphics, 34(4):88:1– 88:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ye:2014:TBD

[YJB^{+}14] Genzhi Ye, Sundeep Jolly, V. Michael Bove, Jr., Qionghai Dai, Ramesh Raskar, and Gordon Wetzstein. Toward BxDF display using multilayer diffraction. ACM Transactions on Graphics, 33(6): 191:1–191:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yucer:2012:TIM

[YJHS12] Kaan Yücer, Alec Jacobson, Alexander Hornung, and Olga Sorkine. Transfusive image manipulation. ACM Transactions on Graphics, 31(6): 176:1–176:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2016:MSS

[YJL^{+}16] Xiao Yan, Yun-Tao Jiang, Chen-Feng Li, Ralph R. Martin, and Shi-Min Hu. Multiphase SPH simulation for interactive fluids and solids. ACM Transactions on Graphics, 35(4):79:1–79:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2022:VCH

[YJLL22] Shuai Yang, Liming Jiang, Ziwei Liu, and Chen Change Loy. VToonify: Controllable high-resolution portrait video style transfer. ACM Transactions on Graphics, 41(6):

203:1–203:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555437.

Yan:2017:EPN

[YJR17] Ling-Qi Yan, Henrik Wann Jensen, and Ravi Ramamoorthi. An efficient and practical near and far field fur reflectance model. ACM Transactions on Graphics, 36(4): 67:1–67:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2023:TGB

[YJY23] Jieyin Yang, Xiaohong Jia, and Dong-Ming Yan. Topology guaranteed B-spline surface/surface intersection. ACM Transactions on Graphics, 42 (6):211:1–211:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618349.

Yumer:2012:CAS

[YK12] Mehmet Ersin Yumer and Levent Burak Kara. Coabstraction of shape collections. ACM Transactions on Graphics, 31(6):166:1– 166:??, November 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yumer:2014:CCH

[YK14] Mehmet Ersin Yumer and Levent Burak Kara. Coconstrained handles for deformation in shape collections. ACM Transactions on Graphics, 33(6):187:1– 187:??, November 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yi:2016:SAF

[YKC^{+}16] Li Yi, Vladimir G. Kim, Duygu Ceylan, I-Chao Shen, Mengyan Yan, Hao Su, Arcewu Lu, Qixing Huang, Alla Sheffer, and Leonidas Guibas. A scalable active framework for region annotation in 3D shape collections. ACM Transactions on Graphics, 35(6):210:1– 210:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yi:2021:DTR

[YKC^{+}21] Shinyoung Yi, Donggun Kim, Kiseok Choi, Adrian Jarabo, Diego Gutierrez, and Min H. Kim. Differentiable transient rendering. ACM Transactions on Graphics, 40(6): 286:1–286:11, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480498.

Yao:2017:IDSa

[YKGA17a] Jiaxian Yao, Danny M. Kaufman, Yotam Gingold, and Maneesh Agrawala. Interactive design and stability analysis of decorative joinery for furniture. ACM Transactions on Graphics, 36(2): 20:1–20:??, April 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yao:2017:IDSb

[YKGA17b] Jiaxian Yao, Danny M. Kaufman, Yotam Gingold, and Maneesh Agrawala. Interactive design and stability analysis of decorative joinery for furniture. ACM Transactions on Graphics, 36(4): 157:1–157:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yamane:2004:SAH

[YKH04] Katsu Yamane, James J.

Kuffner, and Jessica K. Hodgins. Synthesizing animations of human manipulation tasks. ACM Transactions on Graphics, 23(3):532–539, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuksel:2010:MC

[YKH10] Cem Yuksel, John Keyser, and Donald H. House. Mesh colors. ACM Transactions on Graphics, 29(2):15:1– 15:11, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuksel:2012:SMM

[YKJM12] Cem Yuksel, Jonathan M. Kaldor, Doug L. James, and Steve Marschner. Stitch meshes for modeling knitted clothing with yarn-level detail. ACM Transactions on Graphics, 31(4): 37:1–37:12, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2022:INR

[YKZ^{+}22] Lingchen Yang, Byungsoo Kim, Gaspard Zoss, Baran Gözcü, Markus Gross, and Barbara Solenthaler. Implicit neural representation for physics-driven actuated soft bodies. ACM Transactions on Graphics, 41 (4):122:1–122:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530156.

Ye:2008:ARC

[YL08] Yuting Ye and C. Karen Liu. Animating responsive characters with dynamic constraints in near-unactuated coordinates. ACM Transactions on Graphics, 27(5): 112:1–112:??, December 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ye:2010:OFC

[YL10] Yuting Ye and C. Karen Liu.

Optimal feedback control for character animation using an abstract model. ACM Transactions on Graphics, 29(4): 74:1–74:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ye:2012:SDH

[YL12] Yuting Ye and C. Karen Liu.

Synthesis of detailed hand manipulations using contact sampling. ACM Transactions on Graphics, 31(4):41:1– 41:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2022:EEB

[YLB^{+}22] Kai Yan, Christoph Lassner, Brian Budge, Zhao Dong, and Shuang Zhao. Efficient estimation of boundary integrals for path-space differentiable rendering. ACM Transactions on Graphics, 41 (4):123:1–123:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530080.

Yang:2020:NRR

[YLC^{+}20] Sheng Yang, Beichen Li, Yan-Pei Cao, Hongbo Fu, Yu-Kun Lai, Leif Kobbelt, and Shi-Min Hu. Noiseresilient reconstruction of panoramas and 3D scenes using robot-mounted unsynchronized commodity RGB-D cameras. ACM Transactions on Graphics, 39(5):152:1– 152:15, September 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3389412.

Yi:2018:DMS

[YLH18] Ran Yi, Yong-Jin Liu, and Ying He. Delaunay mesh simplification with differential evolution. ACM Transactions on Graphics, 37(6): 263:1–263:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2018:VCE

[YLJ18] Yajie Yan, David Letscher, and Tao Ju. Voxel cores: efficient, robust, and provably good approximation of 3D medial axes. ACM Transactions on Graphics, 37(4):44:1– 44:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yin:2022:DVP

[YLL^{+}22] Jerry Yin, Chenxi Liu, Rebecca Lin, Nicholas Vining, Helge Rhodin, and Alla Sheffer. Detecting viewerperceived intended vector sketch connectivity. ACM Transactions on Graphics, 41(4):87:1–87:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530097.

Yeo:2012:ESV

[YLNP12] Sang Hoon Yeo, Martin Lesmana, Debanga R. Neog, and Dinesh K. Pai. Eyecatch: simulating visuomotor coordination for object interception. ACM Transactions on Graphics, 31(4): 42:1–42:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yoon:2005:COM

[YLPM05] Sung-Eui Yoon, Peter Lindstrom, Valerio Pascucci, and Dinesh Manocha. Cacheoblivious mesh layouts. ACM Transactions on Graphics, 24 (3):886–893, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yin:2007:SSB

[YLvdP07] KangKang Yin, Kevin Loken, and Michiel van de Panne. SIMBICON: simple biped locomotion control. ACM Transactions on Graphics, 26(3):105:1–105:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2015:EPR

[YLX^{+}15] Yin Yang, Dingzeyu Li, Weiwei Xu, Yuan Tian, and Changxi Zheng. Expediting precomputation for reduced deformable simulation. ACM Transactions on Graphics, 34 (6):243:1–243:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2018:IDM

[YLYW18] Guowei Yan, Wei Li, Ruigang Yang, and Huamin Wang. Inexact descent methods for elastic parameter optimization. ACM Transactions on Graphics, 37(6):253:1– 253:??, November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yumer:2016:SST

[YM16] M. Ersin Yumer and Niloy J. Mitra. Spectral style transfer for human motion between independent actions. ACM Transactions on Graphics, 35 (4):137:1–137:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yaldiz:2021:DEE

[YMJ^{+}21] Mustafa B. Yaldiz, Andreas Meuleman, Hyeonjoong Jang, Hyunho Ha, and Min H. Kim. DeepFormableTag: endto-end generation and recognition of deformable fiducial markers. ACM Transactions on Graphics, 40(4): 67:1–67:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459762.

Yang:2023:DNL

[YML^{+}23] Jie Yang, Kaichun Mo, Yu-Kun Lai, Leonidas J. Guibas, and Lin Gao. DSG-Net: Learning disentangled structure and geometry for 3D shape generation. ACM Transactions on Graphics, 42 (1):1:1–1:??, February 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3526212.

Yeh:2013:WRC

[YMR^{+}13] Hengchin Yeh, Ravish Mehra, Zhimin Ren, Lakulish Antani, Dinesh Manocha, and Ming Lin. Wave-ray coupling for interactive sound propagation in large complex scenes. ACM Transactions on Graphics, 32(6): 165:1–165:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2015:FSF

[YMRD15] Ling-Qi Yan, Soham Uday Mehta, Ravi Ramamoorthi, and Fredo Durand. Fast 4D sheared filtering for interactive rendering of distribution effects. ACM Transactions on Graphics, 35(1):7:1– 7:??, December 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yeh:2022:LRP

[YNK^{+}22] Yu-Ying Yeh, Koki Nagano, Sameh Khamis, Jan Kautz, Ming-Yu Liu, and Ting-Chun Wang. Learning to relight portrait images via a virtual light stage and synthetic-to-real adaptation. ACM Transactions on Graphics, 41(6): 231:1–231:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555442.

Yu:2021:MCD

[YNL^{+}21] Jiaqi Yu, Yongwei Nie, Chengjiang Long, Wenju Xu, Qing Zhang, and Guiqing Li. Monte Carlo denoising via auxiliary feature guided self-attention. ACM Transactions on Graphics, 40(6): 273:1–273:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480565.

Yang:2009:AS

[YNS^{+}09] Lei Yang, Diego Nehab, Pedro V. Sander, Pitchaya Sitthi-amorn, Jason Lawrence, and Hugues Hoppe. Amortized supersampling. ACM Transactions on Graphics, 28 (5):135:1–135:12, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yaniv:2019:FAL

[YNS19] Jordan Yaniv, Yael Newman, and Ariel Shamir. The face of art: landmark detection and geometric style in portraits. ACM Transactions on Graphics, 38(4): 60:1–60:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2016:BAG

[YNW16] Feilong Yan, Liangliang Nan, and Peter Wonka. Block assembly for global registration of building scans. ACM Transactions on Graphics, 35 (6):237:1–237:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yin:2023:FC

[YNW^{+}23] Hang Yin, Mohammad Sina Nabizadeh, Baichuan Wu, Stephanie Wang, and Albert Chern. Fluid cohomology. ACM Transactions on Graphics, 42(4): 126:1–126:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592402.

Yang:2018:PIG

[YPA^{+}18] Shan Yang, Zherong Pan, Tanya Amert, Ke Wang, Licheng Yu, Tamara Berg, and Ming C. Lin. Physicsinspired garment recovery from a single-view image. ACM Transactions on Graphics, 37(5):170:1–170:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2016:VAV

[YPB16] Yuting Yang, Sam Prestwood, and Connelly Barnes. VizGen: accelerating visual computing prototypes in dynamic languages. ACM Transactions on Graphics, 35(6): 206:1–206:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yee:2001:SSV

[YPG01] H. Yee, S. Pattanaik, and D. P. Greenberg. Spatiotemporal sensitivity and visual attention for efficient rendering of dynamic environments. ACM Transactions on Graphics, 20(1):39–65, January 2001. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL http:

                                        //www.acm.org/tog/yee01/
                                        index.html.

Yu:2021:HDM

[YPL21] Ri Yu, Hwangpil Park, and Jehee Lee. Human dynamics from monocular video with dynamic camera movements. ACM Transactions on Graphics, 40(6): 208:1–208:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480504.

Yang:2023:LBI

[YPL^{+}23] Zeshi Yang, Zherong Pan, Manyi Li, Kui Wu, and Xifeng Gao. Learning based 2D irregular shape packing. ACM Transactions on Graphics, 42 (6):266:1–266:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618348.

Yan:2023:HDR

[YR23] Han Yan and Bo Ren.

High density ratio multifluid simulation with peridynamics. ACM Transactions on Graphics, 42(6): 191:1–191:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618347.

Yeh:2009:FMT

[YRPF09] Thomas Y. Yeh, Glenn Reinman, Sanjay J. Patel, and Petros Faloutsos. Fool me twice: Exploring and exploiting error tolerance in physics-based animation. ACM Transactions on Graphics, 29(1):5:1– 5:11, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yue:2015:CFM

[YSB^{+}15] Yonghao Yue, Breannan Smith, Christopher Batty, Changxi Zheng, and Eitan Grinspun. Continuum foam: a material point method for shear-dependent flows. ACM Transactions on Graphics, 34(5):160:1–160:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2016:ETM

[YSC^{+}16] Yajie Yan, Kyle Sykes, Erin Chambers, David Letscher, and Tao Ju. Erosion thickness on medial axes of 3D shapes. ACM Transactions on Graphics, 35(4):38:1–38:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yue:2018:HGA

[YSC^{+}18] Yonghao Yue, Breannan Smith, Peter Yichen Chen, Maytee Chantharayukhonthorn, Ken Kamrin, and Eitan Grinspun. Hybrid grains: adaptive coupling of discrete and continuum simulations of granular media. ACM Transactions on Graphics, 37(6): 283:1–283:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2021:RC

[YSC21] Chris Yu, Henrik Schumacher, and Keenan Crane. Repulsive curves. ACM Transactions on Graphics, 40(2): 10:1–10:21, June 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3439429.

Yang:2023:GGA

[YSC^{+}23] Haitao Yang, Bo Sun, Liyan Chen, Amy Pavel, and Qixing Huang. GeoLatent: a geometric approach to latent space design for deformable shape generators. ACM Transactions on Graphics, 42(6): 242:1–242:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618371.

Yao:2022:CMB

[YSCL22] Heyuan Yao, Zhenhua Song, Baoquan Chen, and Libin Liu. ControlVAE: Modelbased learning of generative controllers for physics-based characters. ACM Transactions on Graphics, 41(6): 183:1–183:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555434.

Yucer:2016:EOS

[YSHWSH16] Kaan Yücer, Alexander Sorkine-Hornung, Oliver Wang, and Olga Sorkine-Hornung. Efficient 3D object segmentation from densely sampled light fields with applications to 3D reconstruction. ACM Transactions on Graphics, 35(3): 22:1–22:??, June 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2017:BME

[YSJR17] Ling-Qi Yan, Weilun Sun, Henrik Wann Jensen, and Ravi Ramamoorthi. A BSS-RDF model for efficient rendering of fur with global illumination. ACM Transactions on Graphics, 36(6): 208:1–208:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuksel:2009:HM

[YSK09] Cem Yuksel, Scott Schaefer, and John Keyser. Hair meshes. ACM Transactions on Graphics, 28(5):166:1– 166:7, December 2009. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2014:PSI

[YSL^{+}14] Feilong Yan, Andrei Sharf, Wenzhen Lin, Hui Huang, and Baoquan Chen. Proactive 3D scanning of inaccessible parts.

ACM Transactions on Graphics, 33(4):157:1–157:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2011:AR

[YSLH11] Lei Yang, Pedro V. Sander, Jason Lawrence, and Hugues Hoppe. Antialiasing recovery. ACM Transactions on Graphics, 30(3):22:1–22:9, May 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yamaguchi:2018:HFF

[YSN^{+}18] Shuco Yamaguchi, Shunsuke Saito, Koki Nagano, Yajie Zhao, Weikai Chen, Kyle Olszewski, Shigeo Morishima, and Hao Li. High-fidelity facial reflectance and geometry inference from an unconstrained image. ACM Transactions on Graphics, 37(4): 162:1–162:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuan:2007:IDB

[YSQS07] Lu Yuan, Jian Sun, Long Quan, and Heung-Yeung Shum. Image deblurring with blurred/noisy image pairs. ACM Transactions on Graphics, 26(3):1:1–1:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuan:2008:PIS

[YSQS08] Lu Yuan, Jian Sun, Long Quan, and Heung-Yeung Shum. Progressive inter-scale and intra-scale non-blind image deconvolution. ACM Transactions on Graphics, 27 (3):74:1–74:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2017:CIL

[YSW^{+}17] Zhipei Yan, Stephen Schiller, Gregg Wilensky, Nathan Carr, and Scott Schaefer. kcurves: interpolation at local maximum curvature. ACM Transactions on Graphics, 36 (4):129:1–129:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2020:IMG

[YSW^{+}20] Lingchen Yang, Zefeng Shi, Yiqian Wu, Xiang Li, Kun Zhou, Hongbo Fu, and Youyi Zheng. iOrthoPredictor: model-guided deep prediction of teeth alignment. ACM Transactions on Graphics, 39 (6):216:1–216:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3414685.3417771.

Yan:2023:IDA

[YSW^{+}23] Zihao Yan, Fubao Su, Mingyang Wang, Ruizhen Hu, Hao Zhang, and Hui Huang. Interaction-driven active 3D reconstruction with object interiors. ACM Transactions on Graphics, 42(6): 250:1–250:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618327.

Yu:2013:RSP

[YT13] Jihun Yu and Greg Turk.

Reconstructing surfaces of particle-based fluids using anisotropic kernels. ACM Transactions on Graphics, 32 (1):5:1–5:12, January 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yeung:2011:MCT

[YTBK11] Sai-Kit Yeung, Chi-Keung Tang, Michael S. Brown, and Sing Bing Kang. Matting and compositing of transparent and refractive objects. ACM Transactions on Graphics, 30(1):2:1–2:13, January 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yan:2015:PAR

[YTJR15] Ling-Qi Yan, Chi-Wei Tseng, Henrik Wann Jensen, and Ravi Ramamoorthi. Physicallyaccurate fur reflectance: modeling, measurement and rendering. ACM Transactions on Graphics, 34(6):185:1– 185:??, November 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2018:LSL

[YTL18] Wenhao Yu, Greg Turk, and C. Karen Liu. Learning symmetric and low-energy locomotion. ACM Transactions on Graphics, 37(4):144:1– 144:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2011:IBB

[YTS^{+}11] Lei Yang, Yu-Chiu Tse, Pedro V. Sander, Jason Lawrence, Diego Nehab, Hugues Hoppe, and Clara L. Wilkins. Image-based bidirectional scene reprojection. ACM Transactions on Graphics, 30(6):150:1–150:??, December 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuksel:2020:CCI

[Yuk20] Cem Yuksel. A class of C^{2} interpolating splines. ACM Transactions on Graphics, 39 (5):160:1–160:14, September 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3400301.

Yan:2020:BRS

[YVG20] Chuan Yan, David Vanderhaeghe, and Yotam Gingold. A benchmark for rough sketch cleanup. ACM Transactions on Graphics, 39(6): 163:1–163:14, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417784.

Yan:2013:GPA

[YW13] Dong-Ming Yan and Peter Wonka. Gap processing for adaptive maximal Poisson-disk sampling. ACM Transactions on Graphics, 32 (5):148:1–148:15, September 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ying:2013:SVG

[YWH13] Xiang Ying, Xiaoning Wang, and Ying He. Saddle vertex graph (SVG): a novel solution to the discrete geodesic problem. ACM Transactions on Graphics, 32(6):170:1– 170:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2011:EFA

[YWS^{+}11] Fei Yang, Jue Wang, Eli Shechtman, Lubomir Bourdev, and Dimitri Metaxas. Expression flow for 3D-aware face component transfer.

ACM Transactions on Graphics, 30(4):60:1–60:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2013:UPL

[YWVW13] Yong-Liang Yang, Jun Wang, Etienne Vouga, and Peter Wonka. Urban pattern: layout design by hierarchical domain splitting. ACM Transactions on Graphics, 32(6): 181:1–181:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2021:WMG

[YXFH21] Zhijin Yang, Pengfei Xu, Hongbo Fu, and Hui Huang. WireRoom: model-guided explorative design of abstract wire art. ACM Transactions on Graphics, 40(4): 128:1–128:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459796.

Ying:2014:PCH

[YXH14] Xiang Ying, Shi-Qing Xin, and Ying He. Parallel Chen– Han (PCH) algorithm for discrete geodesics. ACM Transactions on Graphics, 33(1): 9:1–9:11, January 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2022:MCE

[YXK^{+}22] Chang Yu, Yi Xu, Ye Kuang, Yuanming Hu, and Tiantian Liu. MeshTaichi: a compiler for efficient mesh-based operations. ACM Transactions on Graphics, 41(6): 252:1–252:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555430.

Yu:2023:FWR

[YXW^{+}23] Yunchen Yu, Mengqi Xia, Bruce Walter, Eric Michielssen, and Steve Marschner. A full-wave reference simulator for computing surface reflectance. ACM Transactions on Graphics, 42(4): 109:1–109:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592414.

Yang:2021:CGF

[YXZ^{+}21] Shuqi Yang, Shiying Xiong, Yaorui Zhang, Fan Feng, Jinyuan Liu, and Bo Zhu. Clebsch gauge fluid. ACM Transactions on Graphics, 40 (4):99:1–99:11, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459866.

Yuksel:2017:LGH

[YY17] Can Yuksel and Cem Yuksel. Lighting grid hierarchy for self-illuminating explosions. ACM Transactions on Graphics, 36(4):110:1–110:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2019:LCC

[YYL^{+}19] Minjing Yu, Zipeng Ye, Yong-Jin Liu, Ying He, and Charlie C. L. Wang. LineUp: Computing chain-based physical transformation. ACM Transactions on Graphics, 38 (1):11:1–11:??, February 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3269979.

Yang:2022:LUC

[YYL22] Zeshi Yang, Kangkang Yin, and Libin Liu. Learning to use chopsticks in diverse gripping styles. ACM Transactions on Graphics, 41(4):95:1–95:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530057.

Yang:2011:SSE

[YYPM11] Yong-Liang Yang, Yi-Jun Yang, Helmut Pottmann, and Niloy J. Mitra. Shape space exploration of constrained meshes. ACM Transactions on Graphics, 30(6):124:1– 124:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2011:MIH

[YYT^{+}11] Lap-Fai Yu, Sai-Kit Yeung, Chi-Keung Tang, Demetri Terzopoulos, Tony F. Chan, and Stanley J. Osher. Make it home: automatic optimization of furniture arrangement. ACM Transactions on Graphics, 30(4):86:1–86:??, July 2011. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2012:DOS

[YYTC12] Lap-Fai Yu, Sai-Kit Yeung, Demetri Terzopoulos, and Tony F. Chan. DressUp!: outfit synthesis through automatic optimization. ACM Transactions on Graphics, 31 (6):134:1–134:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yin:2021:DDA

[YYVY21] Zhiqi Yin, Zeshi Yang, Michiel Van De Panne, and Kangkang Yin. Discovering diverse athletic jumping strategies. ACM Transactions on Graphics, 40 (4):91:1–91:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459817.

Yeh:2012:SOW

[YYW^{+}12a] Yi-Ting Yeh, Lingfeng Yang, Matthew Watson, Noah D. Goodman, and Pat Hanrahan. Synthesizing open worlds with constraints using locally annealed reversible jump MCMC. ACM Transactions on Graphics, 31(4): 56:1–56:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yuan:2012:OSM

[YYW12b] Zhan Yuan, Yizhou Yu, and Wenping Wang. Object-space multiphase implicit functions. ACM Transactions on Graphics, 31(4):114:1–114:10, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Ying:2004:SMB

[YZ04] Lexing Ying and Denis Zorin. A simple manifold-based construction of surfaces of arbitrary smoothness. ACM Transactions on Graphics, 23 (3):271–275, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yi:2023:ERT

[YZH^{+}23] Xinyu Yi, Yuxiao Zhou, Marc Habermann, Vladislav Golyanik, Shaohua Pan, Christian Theobalt, and Feng Xu. EgoLocate: Real-time motion capture, localization, and mapping with sparse body-mounted sensors. ACM Transactions on Graphics, 42 (4):76:1–76:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592099.

Yang:2022:NRR

[YZL^{+}22] Bangbang Yang, Yinda Zhang, Yijin Li, Zhaopeng Cui, Sean Fanello, Hujun Bao, and Guofeng Zhang. Neural rendering in a room: amodal 3D understanding and freeviewpoint rendering for the closed scene composed of pre-captured objects. ACM Transactions on Graphics, 41 (4):101:1–101:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530163.

Yu:2022:EDP

[YZN^{+}22] Zihan Yu, Cheng Zhang, Derek Nowrouzezahrai, Zhao Dong, and Shuang Zhao. Efficient differentiation of pixel reconstruction filters for path-space differentiable rendering. ACM Transactions on Graphics, 41(6): 191:1–191:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555500.

Yan:2016:APA

[YZW^{+}16] Zhicheng Yan, Hao Zhang, Baoyuan Wang, Sylvain Paris, and Yizhou Yu. Automatic photo adjustment using deep neural networks. ACM Transactions on Graphics, 35(2): 11:1–11:??, May 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yang:2012:BTM

[YZWH12] Xuan Yang, Linling Zhang, Tien-Tsin Wong, and Pheng-Ann Heng. Binocular tone mapping. ACM Transactions on Graphics, 31(4):93:1– 93:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2004:MEP

[YZX^{+}04] Yizhou Yu, Kun Zhou, Dong Xu, Xiaohan Shi, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Mesh editing with Poisson-based gradient field manipulation. ACM Transactions on Graphics, 23 (3):644–651, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yu:2018:SSC

[YZX^{+}18] Fenggen Yu, Yan Zhang*, Kai Xu, Ali Mahdavi-Amiri, and Hao Zhang. Semi-supervised co-analysis of 3D shape styles from projected lines. ACM Transactions on Graphics, 37 (2):21:1–21:??, July 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Yi:2021:TRT

[YZX21] Xinyu Yi, Yuxiao Zhou, and Feng Xu. TransPose: realtime 3D human translation and pose estimation with six inertial sensors. ACM Transactions on Graphics, 40 (4):86:1–86:13, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459786.

Zhang:2021:CDP

[ZAB21] Ran Zhang, Thomas Auzinger, and Bernd Bickel. Computational design of planar multistable compliant structures. ACM Transactions on Graphics, 40(5): 186:1–186:16, October 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3453477.

Zhang:2017:FAR

[ZAC^{+}17] Ran Zhang, Thomas Auzinger, Duygu Ceylan, Wilmot Li, and Bernd Bickel. Functionalityaware retargeting of mechanisms to 3D shapes. ACM Transactions on Graphics, 36 (4):81:1–81:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2014:MMC

[ZAE^{+}14] Jun-Yan Zhu, Aseem Agarwala, Alexei A. Efros, Eli Shechtman, and Jue Wang. Mirror mirror: crowdsourcing better portraits. ACM Transactions on Graphics, 33(6): 234:1–234:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2021:BGB

[ZAFW21] Peihao Zhu, Rameen Abdal, John Femiani, and Peter Wonka. Barbershop: GANbased image compositing using segmentation masks. ACM Transactions on Graphics, 40 (6):215:1–215:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3478513.3480537.

Zell:2015:SSE

[ZAJ^{+}15] Eduard Zell, Carlos Aliaga, Adrian Jarabo, Katja Zibrek, Diego Gutierrez, Rachel Mc-Donnell, and Mario Botsch. To stylize or not to stylize?: the effect of shape and material stylization on the perception of computergenerated faces. ACM Transactions on Graphics, 34(6): 184:1–184:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:1994:IKP

[ZB94] Jianmin Zhao and Norman I. Badler. Inverse kinematics positioning using nonlinear programming for highly articulated figures. ACM Transactions on Graphics, 13(4):313–336, October 1994. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
toc/Abstracts/0730-0301/
195827.html.

Zhu:2005:ASF

[ZB05] Yongning Zhu and Robert Bridson. Animating sand as a fluid. ACM Transactions on Graphics, 24(3):965–972, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2013:IAS

[ZB13] Yili Zhao and Jernej Barbic. Interactive authoring of simulation-ready plants. ACM Transactions on Graphics, 32 (4):84:1–84:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2014:PFS

[ZB14] Xinxin Zhang and Robert Bridson. A PPPM fast summation method for fluids and beyond. ACM Transactions on Graphics, 33(6):

206:1–206:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zoss:2018:ERJ

[ZBBB18] Gaspard Zoss, Derek Bradley, Pascal Bérard, and Thabo Beeler. An empirical rig for jaw animation. ACM Transactions on Graphics, 37(4): 59:1–59:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2015:RMV

[ZBG15a] Xinxin Zhang, Robert Bridson, and Chen Greif. Restoring the missing vorticity in advection-projection fluid solvers. ACM Transactions on Graphics, 34(4):52:1– 52:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2015:SRB

[ZBG15b] Yufeng Zhu, Robert Bridson, and Chen Greif. Simulating rigid body fracture with surface meshes. ACM Transactions on Graphics, 34(4): 150:1–150:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zoss:2019:AMJ

[ZBGB19] Gaspard Zoss, Thabo Beeler, Markus Gross, and Derek Bradley. Accurate markerless jaw tracking for facial performance capture. ACM Transactions on Graphics, 38(4): 50:1–50:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:CDF

[ZBJ^{+}23] Zhan Zhang, Christopher Brandt, Jean Jouve, Yue Wang, Tian Chen, Mark Pauly, and Julian Panetta. Computational design of flexible planar microstructures. ACM Transactions on Graphics, 42(6):185:1–185:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618396.

Zhu:2018:BCQ

[ZBK18] Yufeng Zhu, Robert Bridson, and Danny M. Kaufman. Blended cured quasi-Newton for distortion optimization. ACM Transactions on Graphics, 37(4):40:1– 40:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2020:CD

[ZBLJ20] Jiayi Eris Zhang, Seungbae Bang, David I. W. Levin, and Alec Jacobson. Complementary dynamics. ACM Transactions on Graphics, 39(6): 179:1–179:11, November 2020.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417819.

Zheng:2020:NLF

[ZBW^{+}20] Quan Zheng, Vahid Babaei, Gordon Wetzstein, Hans-Peter Seidel, Matthias Zwicker, and Gurprit Singh. Neural light field 3D printing. ACM Transactions on Graphics, 39 (6):207:1–207:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417879.

Zhu:2021:NCL

[ZBX^{+}21] Junqiu Zhu, Yaoyi Bai, Zilin Xu, Steve Bako, Edgar Velázquez-Armendáriz, Lu Wang, Pradeep Sen, Miloš Hašan, and Ling-Qi Yan. Neural complex luminaires: representation and rendering. ACM Transactions on Graphics, 40 (4):57:1–57:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459798.

Zhang:2019:IRT

[ZBYX19] Hao Zhang, Zi-Hao Bo, Jun-Hai Yong, and Feng Xu. InteractionFusion: real-time reconstruction of hand poses and deformable objects in handobject interactions. ACM Transactions on Graphics, 38 (4):48:1–48:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhi:2022:SSA

[ZCB^{+}22] Tiancheng Zhi, Bowei Chen, Ivaylo Boyadzhiev, Sing Bing Kang, Martial Hebert, and Srinivasa G. Narasimhan. Semantically supervised appearance decomposition for virtual staging from a single panorama. ACM Transactions on Graphics, 41(4): 83:1–83:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3528223.3530148.

Zheng:2012:IIC

[ZCC^{+}12] Youyi Zheng, Xiang Chen, Ming-Ming Cheng, Kun Zhou, Shi-Min Hu, and Niloy J. Mitra. Interactive images: cuboid proxies for smart image manipulation. ACM Transactions on Graphics, 31(4):99:1–99:11, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2016:ERR

[ZCC16] Edward Zhang, Michael F.

Cohen, and Brian Curless. Emptying, refurnishing, and relighting indoor spaces. ACM Transactions on Graphics, 35 (6):174:1–174:??, November 2016. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2016:SPS

[ZCD^{+}16] Zhiming Zhou, Guojun Chen, Yue Dong, David Wipf, Yong Yu, John Snyder, and Xin Tong. Sparse-as-possible SVBRDF acquisition. ACM Transactions on Graphics, 35 (6):189:1–189:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2018:WCP

[ZCL18] Nanxuan Zhao, Ying Cao, and Rynson W. H. Lau. What characterizes personalities of graphic designs? ACM Transactions on Graphics, 37(4): 116:1–116:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zeng:2020:CFG

[ZCLJ20] Dan Zeng, Erin Chambers, David Letscher, and Tao Ju. To cut or to fill: a global optimization approach to topological simplification. ACM Transactions on Graphics, 39(6): 201:1–201:18, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417854.

Zhang:2022:MGD

[ZCM22] Meng Zhang, Duygu Ceylan, and Niloy J. Mitra. Motion guided deep dynamic 3D garments. ACM Transactions on Graphics, 41(6): 219:1–219:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555485.

Zhou:2023:GHQ

[ZCP^{+}23] Yuxiao Zhou, Menglei Chai, Alessandro Pepe, Markus Gross, and Thabo Beeler. GroomGen: a high-quality generative hair model using hierarchical latent representations. ACM Transactions on Graphics, 42(6): 270:1–270:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618309.

Zou:2016:LCC

[ZCR^{+}16] Changqing Zou, Junjie Cao, Warunika Ranaweera, Ibraheem Alhashim, Ping Tan, Alla Sheffer, and Hao Zhang. Legible compact calligrams. ACM Transactions on Graphics, 35(4):122:1–122:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zoss:2022:PRF

[ZCS^{+}22] Gaspard Zoss, Prashanth Chandran, Eftychios Sifakis, Markus Gross, Paulo Gotardo, and Derek Bradley. Production-ready face reaging for visual effects. ACM Transactions on Graphics, 41 (6):237:1–237:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3550454.3555520.

Zehnder:2016:DSS

[ZCT16] Jonas Zehnder, Stelian Coros, and Bernhard Thomaszewski. Designing structurally-sound ornamental curve networks. ACM Transactions on Graphics, 35(4):99:1–99:??, July 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:CDM

[ZCT^{+}21] Zhoutong Zhang, Forrester Cole, Richard Tucker, William T. Freeman, and Tali Dekel. Consistent depth of moving objects in video. ACM Transactions on Graphics, 40(4): 148:1–148:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459871.

Zehnder:2022:SSG

[ZCT22] Jonas Zehnder, Stelian Coros, and Bernhard Thomaszewski. SGN: Sparse Gauss–Newton for accelerated sensitivity analysis. ACM Transactions on Graphics, 41(1):4:1– 4:10, February 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3470005.

Zhang:2017:DDA

[ZCW^{+}17] Meng Zhang, Menglei Chai, Hongzhi Wu, Hao Yang, and Kun Zhou. A data-driven approach to four-view imagebased hair modeling. ACM Transactions on Graphics, 36 (4):156:1–156:??, July 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2022:TSB

[ZCX^{+}22] Haikuan Zhu, Juan Cao, Yanyang Xiao, Zhonggui Chen, Zichun Zhong, and Yongjie Jessica Zhang. TCBspline-based image vectorization. ACM Transactions on Graphics, 41(3): 34:1–34:17, June 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3513132.

Zirr:2020:PDI

[ZD20] Tobias Zirr and Carsten Dachsbacher. Path differentialinformed stratified MCMC and adaptive forward path sampling. ACM Transactions on Graphics, 39(6): 246:1–246:19, November 2020.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3414685.3417856.

Zhang:2021:ASM

[ZDDZ21] Cheng Zhang, Zhao Dong, Michael Doggett, and Shuang Zhao. Antithetic sampling for Monte Carlo differentiable rendering. ACM Transactions on Graphics, 40(4): 77:1–77:12, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459783.

Zhang:2022:PSC

[ZDF^{+}22] Jiayi Eris Zhang, Jérémie Dumas, Yun (Raymond) Fei, Alec Jacobson, Doug L. James, and Danny M. Kaufman. Progressive simulation for cloth quasistatics. ACM Transactions on Graphics, 41 (6):218:1–218:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555510.

Zhang:2023:PSQ

[ZDF^{+}23] Jiayi Eris Zhang, Jérémie Dumas, Yun (Raymond) Fei, Alec Jacobson, Doug L. James, and Danny M. Kaufman. Progressive shell qasistatics for unstructured meshes. ACM Transactions on Graphics, 42(6):

176:1–176:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618388.

Zollhofer:2015:SBR

[ZDI^{+}15] Michael Zollhöfer, Angela Dai, Matthias Innmann, Chenglei Wu, Marc Stamminger, Christian Theobalt, and Matthias Nießner. Shadingbased refinement on volumetric signed distance functions. ACM Transactions on Graphics, 34(4):96:1–96:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2014:LBC

[ZDL^{+}14] Juyong Zhang, Bailin Deng, Zishun Liu, Giuseppe Patanè, Sofien Bouaziz, Kai Hormann, and Ligang Liu. Local barycentric coordinates. ACM Transactions on Graphics, 33 (6):188:1–188:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:PPS

[ZDT^{+}23] Yuxin Zhang, Weiming Dong, Fan Tang, Nisha Huang, Haibin Huang, Chongyang Ma, Tong-Yee Lee, Oliver Deussen, and Changsheng Xu. ProSpect: Prompt spectrum for attribute-aware personalization of diffusion models. ACM Transactions on Graphics, 42(6): 244:1–244:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618342.

Zhang:2022:IPM

[ZEF^{+}22] Congyi Zhang, Mohamed Elgharib, Gereon Fox, Min Gu, Christian Theobalt, and Wenping Wang. An implicit parametric morphable dental model. ACM Transactions on Graphics, 41(6): 217:1–217:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555469.

Zhou:2003:IMT

[ZF03] Bingfeng Zhou and Xifeng Fang. Improving mid-tone quality of variable-coefficient error diffusion using threshold modulation. ACM Transactions on Graphics, 22(3): 437–444, July 2003. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2010:PRH

[ZFL^{+}10] Shizhe Zhou, Hongbo Fu, Ligang Liu, Daniel Cohen-Or, and Xiaoguang Han. Parametric reshaping of human bodies in images. ACM Transactions on Graphics, 29(4): 126:1–126:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2022:DDP

[ZFO^{+}22] Zheng-Yu Zhao, Qing Fang, Wenqing Ouyang, Zheng Zhang, Ligang Liu, and Xiao-Ming Fu. Developabilitydriven piecewise approximations for triangular meshes. ACM Transactions on Graphics, 41(4):43:1–43:??, July 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3528223.3530117.

Zhang:2019:CDF

[ZFS^{+}19] Xiaoting Zhang, Guoxin Fang, Melina Skouras, Gwenda Gieseler, Charlie C. L. Wang, and Emily Whiting. Computational design of fabric formwork. ACM Transactions on Graphics, 38(4): 109:1–109:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:NLT

[ZFT^{+}21] Xiuming Zhang, Sean Fanello, Yun-Ta Tsai, Tiancheng Sun, Tianfan Xue, Rohit Pandey, Sergio Orts-Escolano, Philip Davidson, Christoph Rhemann, Paul Debevec, Jonathan T. Barron, Ravi Ramamoorthi, and William T. Freeman. Neural light transport for relighting and view synthesis. ACM Transactions on Graphics, 40(1):9:1– 9:17, January 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3446328.

Zsolnai-Feher:2018:GMS

[ZFWW18] Károly Zsolnai-Fehér, Peter Wonka, and Michael Wimmer. Gaussian material synthesis. ACM Transactions on Graphics, 37(4):76:1–76:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zelinka:2002:PGP

[ZG02] Steve Zelinka and Michael Garland. Permission grids: practical, error-bounded simplification. ACM Transactions on Graphics, 21(2): 207–229, April 2002. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zelinka:2004:JMB

[ZG04] Steve Zelinka and Michael Garland. Jump map-based interactive texture synthesis. ACM Transactions on Graphics, 23(4):930–962, October 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2016:CFS

[ZGH^{+}16] Haisen Zhao, Fanglin Gu, Qi-Xing Huang, Jorge Garcia, Yong Chen, Changhe Tu, Bedrich Benes, Hao Zhang, Daniel Cohen-Or, and Baoquan Chen. Connected Fermat spirals for layered fabrication. ACM Transactions on Graphics, 35(4): 100:1–100:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhong:2013:PBA

[ZGW^{+}13] Zichun Zhong, Xiaohu Guo, Wenping Wang, Bruno Lévy, Feng Sun, Yang Liu, and Weihua Mao. Particle-based anisotropic surface meshing. ACM Transactions on Graphics, 32(4):99:1–99:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2023:CDL

[ZGXF23] Mingjun Zhou, Jiahao Ge, Hao Xu, and Chi-Wing Fu. Computational design of LEGO(R) sketch art. ACM Transactions on Graphics, 42 (6):201:1–201:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3618306.

Zhou:2016:MAS

[ZGZJ16] Qingnan Zhou, Eitan Grinspun, Denis Zorin, and Alec Jacobson. Mesh arrangements for solid geometry. ACM Transactions on Graphics, 35 (4):39:1–39:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2018:SDA

[Zha18] Shuang Zhao. Session details: Advanced SVBRDF. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zou:2015:TCS

[ZHCJ15] Ming Zou, Michelle Holloway, Nathan Carr, and Tao Ju. Topology-constrained surface reconstruction from crosssections. ACM Transactions on Graphics, 34(4):128:1– 128:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2016:RTC

[ZHG^{+}16] Xi Zhao, Ruizhen Hu, Paul Guerrero, Niloy Mitra, and Taku Komura. Relationship templates for creating scene variations. ACM Transactions on Graphics, 35(6): 207:1–207:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2020:RTR

[ZHHZ20] Jinta Zheng, Shih-Hsuan Hung, Kyle Hiebel, and Yue Zhang. Real-time rendering of decorative sound textures for soundscapes. ACM Transactions on Graphics, 39(6): 271:1–271:12, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417875.

Zhou:2005:PSF

[ZHL^{+}05] Kun Zhou, Yaohua Hu, Stephen Lin, Baining Guo, and Heung-Yeung Shum. Precomputed shadow fields for dynamic scenes. ACM Transactions on Graphics, 24(3): 1196–1201, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2010:WBA

[ZHLB10] Muyang Zhang, Jin Huang, Xinguo Liu, and Hujun Bao. A wave-based anisotropic quadrangulation method. ACM Transactions on Graphics, 29 (4):118:1–118:??, July 2010. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2023:NOO

[ZHM^{+}23] Chuankun Zheng, Yuchi Huo, Shaohua Mo, Zhihua Zhong, Zhizhen Wu, Wei Hua, Rui Wang, and Hujun Bao. NeLT: Object-oriented neural light transfer. ACM Transactions on Graphics, 42(5): 163:1–163:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3596491.

Zhou:2018:SDC

[Zho18] Kun Zhou. Session details: Capturing 4D performances. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:JCD

[ZHPY21] Yongqi Zhang, Haikun Huang, Erion Plaku, and Lap-Fai Yu. Joint computational design of workspaces and workplans. ACM Transactions on Graphics, 40(6): 228:1–228:16, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480500.

Zhou:2009:RIR

[ZHR^{+}09] Kun Zhou, Qiming Hou, Zhong Ren, Minmin Gong, Xin Sun, and Baining Guo. RenderAnts: interactive Reyes rendering on GPUs. ACM Transactions on Graphics, 28 (5):155:1–155:11, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2013:MFT

[ZHRB13] Shuang Zhao, Milos Hasan, Ravi Ramamoorthi, and Kavita Bala. Modular flux transfer: efficient rendering of high-resolution volumes with repeated structures. ACM Transactions on Graphics, 32 (4):131:1–131:??, July 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2005:LMD

[ZHS^{+}05] Kun Zhou, Jin Huang, John Snyder, Xinguo Liu, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Large mesh deformation using the volumetric graph Laplacian. ACM Transactions on Graphics, 24(3):496–503, July 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2020:MSA

[ZHS^{+}20] Yang Zhou, Xintong Han, Eli Shechtman, Jose Echevarria, Evangelos Kalogerakis, and Dingzeyu Li. MakeltTalk: speaker-aware talking-head animation. ACM Transactions on Graphics, 39(6): 221:1–221:15, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417774.

Zhu:2018:SDI

[Zhu18a] Jun-Yan Zhu. Session details: Image processing. ACM Transactions on Graphics, 37 (6), November 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhuang:2018:SDG

[Zhu18b] Richard (Hao) Zhuang. Session details: Geometry generation. ACM Transactions on Graphics, 37(6), November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2006:MQG

[ZHW^{+}06] Kun Zhou, Xin Huang, Xi Wang, Yiying Tong, Mathieu Desbrun, Baining Guo, and Heung-Yeung Shum. Mesh quilting for geometric texture synthesis. ACM Transactions on Graphics, 25 (3):690–697, July 2006. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2008:RTK

[ZHWG08] Kun Zhou, Qiming Hou, Rui Wang, and Baining Guo. Real-time KD-tree construction on graphics hardware. ACM Transactions on Graphics, 27(5):126:1– 126:??, December 2008. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2012:PSG

[ZHWW12] Yahan Zhou, Haibin Huang, Li-Yi Wei, and Rui Wang. Point sampling with general noise spectrum. ACM Transactions on Graphics, 31(4):

76:1–76:11, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2007:DMS

[ZHX^{+}07] Kun Zhou, Xin Huang, Weiwei Xu, Baining Guo, and Heung-Yeung Shum. Direct manipulation of subdivision surfaces on GPUs. ACM Transactions on Graphics, 26 (3):91:1–91:??, July 2007. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2011:SBD

[ZIH^{+}11] Bo Zhu, Michiaki Iwata, Ryo Haraguchi, Takashi Ashihara, Nobuyuki Umetani, Takeo Igarashi, and Kazuo Nakazawa. Sketch-based dynamic illustration of fluid systems. ACM Transactions on Graphics, 30(6):134:1– 134:??, December 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zitnick:2013:HBU

[Zit13] C. Lawrence Zitnick. Handwriting beautification using token means. ACM Transactions on Graphics, 32(4): 53:1–53:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zang:2018:STT

[ZIT^{+}18] Guangming Zang, Ramzi Idoughi, Ran Tao, Gilles Lubineau, Peter Wonka, and Wolfgang Heidrich. Spacetime tomography for continuously deforming objects. ACM Transactions on Graphics, 37(4):100:1–100:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zang:2019:WPT

[ZIT^{+}19] Guangming Zang, Ramzi Idoughi, Ran Tao, Gilles Lubineau, Peter Wonka, and Wolfgang Heidrich. Warpand-project tomography for rapidly deforming objects. ACM Transactions on Graphics, 38(4):86:1–86:??, July 2019. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2009:HF

[ZJ09] Changxi Zheng and Doug L. James. Harmonic fluids. ACM Transactions on Graphics, 28 (3):37:1–37:??, August 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2010:RBF

[ZJ10] Changxi Zheng and Doug L. James. Rigid-body fracture sound with precomputed soundbanks. ACM Transactions on Graphics, 29(4): 69:1–69:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2011:THQ

[ZJ11] Changxi Zheng and Doug L.

James. Toward high-quality modal contact sound. ACM Transactions on Graphics, 30 (4):38:1–38:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2012:EBS

[ZJ12] Changxi Zheng and Doug L.

James. Energy-based selfcollision culling for arbitrary mesh deformations. ACM Transactions on Graphics, 31(4):98:1–98:12, July 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zeltner:2018:LLC

[ZJ18] Tizian Zeltner and Wenzel Jakob. The layer laboratory: a calculus for additive and subtractive composition of anisotropic surface reflectance. ACM Transactions on Graphics, 37(4):74:1– 74:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2014:TCS

[ZJL14] Shizhe Zhou, Changyun Jiang, and Sylvain Lefebvre. Topology-constrained synthesis of vector patterns. ACM Transactions on Graphics, 33 (6):215:1–215:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2011:BVA

[ZJMB11] Shuang Zhao, Wenzel Jakob, Steve Marschner, and Kavita Bala. Building volumetric appearance models of fabric using micro CT imaging. ACM Transactions on Graphics, 30 (4):44:1–44:??, July 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2012:SAS

[ZJMB12] Shuang Zhao, Wenzel Jakob, Steve Marschner, and Kavita Bala. Structure-aware synthesis for predictive woven fabric appearance. ACM Transactions on Graphics, 31(4): 75:1–75:10, July 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:PTF

[ZJNZ23] Qing Zhang, Hao Jiang, Yongwei Nie, and Wei-Shi Zheng. Pyramid texture filtering. ACM Transactions on Graphics, 42(4): 110:1–110:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592120.

Zhong:2021:RRH

[ZJY^{+}21] Fangcheng Zhong, Akshay Jindal, Ali Özgür Yöntem, Param Hanji, Simon J. Watt, and Rafał K. Mantiuk. Reproducing reality with a highdynamic-range multi-focal stereo display. ACM Transactions on Graphics, 40(6): 241:1–241:14, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480513.

Zhao:2022:HPM

[ZJY^{+}22] Fuqiang Zhao, Yuheng Jiang, Kaixin Yao, Jiakai Zhang, Liao Wang, Haizhao Dai, Yuhui Zhong, Yingliang Zhang, Minye Wu, Lan Xu, and Jingyi Yu. Human performance modeling and rendering via neural animated mesh. ACM Transactions on Graphics, 41(6): 235:1–235:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555451.

Zhou:2013:DSR

[ZK13] Qian-Yi Zhou and Vladlen Koltun. Dense scene reconstruction with points of interest. ACM Transactions on Graphics, 32(4):112:1– 112:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2014:CMO

[ZK14] Qian-Yi Zhou and Vladlen Koltun. Color map optimization for 3D reconstruction with consumer depth cameras. ACM Transactions on Graphics, 33(4):155:1–155:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2022:LAT

[ZK22] Xilong Zhou and Nima Khademi Kalantari. Look-ahead training with learned reflectance loss for single-image SVBRDF estimation. ACM Transactions on Graphics, 41(6): 266:1–266:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555495.

Zehnder:2017:MDF

[ZKBT17] Jonas Zehnder, Espen Knoop, Moritz Bächer, and Bernhard Thomaszewski. MetaSilicone: design and fabrication of composite silicone with desired mechanical properties. ACM Transactions on Graphics, 36 (6):240:1–240:??, November 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zitnick:2004:HQV

[ZKU^{+}04] C. Lawrence Zitnick, Sing Bing Kang, Matthew Uyttendaele, Simon Winder, and Richard Szeliski. High-quality video view interpolation using a layered representation. ACM Transactions on Graphics, 23 (3):600–608, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2016:RFB

[ZLB16a] Xinxin Zhang, Minchen Li, and Robert Bridson. Resolving fluid boundary layers with particle strength exchange and weak adaptivity. ACM Transactions on Graphics, 35 (4):76:1–76:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2016:FPY

[ZLB16b] Shuang Zhao, Fujun Luan, and Kavita Bala. Fitting procedural yarn models for realistic cloth rendering. ACM Transactions on Graphics, 35 (4):51:1–51:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2013:NGS

[ZLC^{+}13] Bo Zhu, Wenlong Lu, Matthew Cong, Byungmoon Kim, and Ronald Fedkiw. A new grid structure for domain extension. ACM Transactions on Graphics, 32(4): 63:1–63:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2022:FAI

[ZLC^{+}22] Yunxiang Zhang, Benjamin Liang, Boyuan Chen, Paul M. Torrens, S. Farokh Atashzar, Dahua Lin, and Qi Sun. Force-aware interface via electromyography for natural VR/AR interaction. ACM Transactions on Graphics, 41 (6):268:1–268:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3550454.3555461.

Zhu:2014:AIE

[ZLE14] Jun-Yan Zhu, Yong Jae Lee, and Alexei A. Efros. Average-Explorer: interactive exploration and alignment of visual data collections. ACM Transactions on Graphics, 33(4): 160:1–160:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhong:2021:AGO

[ZLH^{+}21] Lei Zhong, Feng-Heng Li, Hao-Zhi Huang, Yong Zhang, Shao-Ping Lu, and Jue Wang. Aesthetic-guided outward image cropping. ACM Transactions on Graphics, 40(6): 211:1–211:13, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480566.

Zhang:2015:PMP

[ZLP^{+}15] Xiaoting Zhang, Xinyi Le, Athina Panotopoulou, Emily Whiting, and Charlie C. L. Wang. Perceptual models of preference in 3D printing direction. ACM Transactions on Graphics, 34(6): 215:1–215:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2015:CNN

[ZLQF15] Bo Zhu, Minjae Lee, Ed Quigley, and Ronald Fedkiw. Codimensional non-Newtonian fluids. ACM Transactions on Graphics, 34(4):115:1– 115:??, August 2015. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2016:PPL

[ZLW^{+}16] Haisen Zhao, Lin Lu, Yuan Wei, Dani Lischinski, Andrei Sharf, Daniel Cohen-Or, and Baoquan Chen. Printed perforated lampshades for continuous projective images. ACM Transactions on Graphics, 35 (5):154:1–154:??, September 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2018:TSS

[ZLW^{+}18] Lvmin Zhang, Chengze Li, Tien-Tsin Wong, Yi Ji, and Chunping Liu. Two-stage sketch colorization. ACM Transactions on Graphics, 37 (6):261:1–261:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2016:GOT

[ZLWH16] Haichao Zhu, Xueting Liu, Tien-Tsin Wong, and Pheng-Ann Heng. Globally optimal toon tracking. ACM Transactions on Graphics, 35(4): 75:1–75:??, July 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:EFV

[ZLY^{+}21] Jiakai Zhang, Xinhang Liu, Xinyi Ye, Fuqiang Zhao, Yanshun Zhang, Minye Wu, Yingliang Zhang, Lan Xu, and Jingyi Yu. Editable free-viewpoint video using a layered neural representation. ACM Transactions on Graphics, 40(4): 149:1–149:18, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459756.

Zhao:2023:EPD

[ZLZ^{+}23] Zheng-Yu Zhao, Mo Li, Zheng Zhang, Qing Fang, Ligang Liu, and Xiao-Ming Fu. Evolutionary piecewise developable approximations. ACM Transactions on Graphics, 42(4):120:1–120:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3592140.

Zhang:2011:ESC

[ZM11] Ren-Jiang Zhang and Weiyin Ma. An efficient scheme for curve and surface construction based on a set of interpolatory basis functions. ACM Transactions on Graphics, 30(2): 10:1–10:11, April 2011. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2013:STE

[ZM13] Yubo Zhang and Kwan-Liu Ma. Spatio-temporal extrapolation for fluid animation. ACM Transactions on Graphics, 32(6):183:1– 183:??, November 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zordan:2005:DRM

[ZMCF05] Victor Brian Zordan, Anna Majkowska, Bill Chiu, and Matthew Fast. Dynamic response for motion capture animation. ACM Transactions on Graphics, 24(3): 697–701, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2019:SDL

[ZMN^{+}19] Xuaner Zhang, Kevin Matzen, Vivien Nguyen, Dillon Yao, You Zhang, and Ren Ng. Synthetic defocus and lookahead autofocus for casual videography. ACM Transactions on Graphics, 38(4): 30:1–30:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zayer:2018:LFN

[ZMSS18] Rhaleb Zayer, Daniel Mlakar, Markus Steinberger, and Hans-Peter Seidel. Layered fields for natural tessellations on surfaces. ACM Transactions on Graphics, 37(6): 264:1–264:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2005:FBS

[ZMT05] Eugene Zhang, Konstantin Mischaikow, and Greg Turk. Feature-based surface parameterization and texture mapping. ACM Transactions on Graphics, 24(1):1–27, January 2005. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2006:VFD

[ZMT06] Eugene Zhang, Konstantin Mischaikow, and Greg Turk. Vector field design on surfaces. ACM Transactions on Graphics, 25(4):1294–1326, October 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:EMG

[ZMW^{+}23] Sharon Zhang, Jiaju Ma, Jiajun Wu, Daniel Ritchie, and Maneesh Agrawala. Editing motion graphics video via motion vectorization and transformation. ACM Transactions on Graphics, 42(6): 229:1–229:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618316.

Zhang:2006:PDA

[ZN06] Li Zhang and Shree Nayar.

Projection defocus analysis for scene capture and image display. ACM Transactions on Graphics, 25(3):907–915, July 2006. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zollhofer:2014:RTN

[ZNI^{+}14] Michael Zollhöfer, Matthias Nießner, Shahram Izadi, Christoph Rehmann, Christopher Zach, Matthew Fisher, Chenglei Wu, Andrew Fitzgibbon, Charles Loop, Christian Theobalt, and Marc Stamminger. Real-time non-rigid reconstruction using an RGB-D camera. ACM Transactions on Graphics, 33(4): 156:1–156:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zehnder:2018:ARS

[ZNT18] Jonas Zehnder, Rahul Narain, and Bernhard Thomaszewski. An advection-reflection solver for detail-preserving fluid simulation. ACM Transactions on Graphics, 37(4):85:1– 85:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zimmermann:2019:PRA

[ZPBC19] Simon Zimmermann, Roi Poranne, James M. Bern, and Stelian Coros. PuppetMaster: robotic animation of marionettes. ACM Transactions on Graphics, 38(4): 103:1–103:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2017:PIE

[ZPBK17] Yufeng Zhu, Jovan Popović, Robert Bridson, and Danny M. Kaufman. Planar interpolation with extreme deformation, topology change and dynamics. ACM Transactions on Graphics, 36(6):213:1– 213:??, November 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zwicker:2002:PIS

[ZPKG02] Matthias Zwicker, Mark Pauly, Oliver Knoll, and Markus Gross. Pointshop 3D: an interactive system for point-based surface editing. ACM Transactions on Graphics, 21(3):322–329, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2023:LAS

[ZPW^{+}23] Xin-Yang Zheng, Hao Pan, Peng-Shuai Wang, Xin Tong, Yang Liu, and Heung-Yeung Shum. Locally attentional SDF diffusion for controllable 3D shape generation. ACM Transactions on Graphics, 42 (4):91:1–91:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592103.

Zhao:2023:LPR

[ZPYX23] Hang Zhao, Zherong Pan, Yang Yu, and Kai Xu. Learning physically realizable skills for online packing of general 3D shapes. ACM Transactions on Graphics, 42(5): 165:1–165:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3603544.

Zhou:2013:WCS

[ZPZ13] Qingnan Zhou, Julian Panetta, and Denis Zorin. Worst-case structural analysis. ACM Transactions on Graphics, 32 (4):137:1–137:??, July 2013.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2014:CST

[ZQC^{+}14] Bo Zhu, Ed Quigley, Matthew Cong, Justin Solomon, and Ronald Fedkiw. Codimensional surface tension flow on simplicial complexes. ACM Transactions on Graphics, 33 (4):111:1–111:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2019:CAG

[ZQCL19] Xinru Zheng, Xiaotian Qiao, Ying Cao, and Rynson W. H. Lau. Content-aware generative modeling of graphic design layouts. ACM Transactions on Graphics, 38(4): 133:1–133:??, July 2019. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:DPG

[ZQL^{+}23] Longwen Zhang, Qiwei Qiu, Hongyang Lin, Qixuan Zhang, Cheng Shi, Wei Yang, Ye Shi, Sibei Yang, Lan Xu, and Jingyi Yu. DreamFace: Progressive generation of animatable 3D faces under text guidance. ACM Transactions on Graphics, 42(4): 138:1–138:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592094.

Zhong:2012:DAV

[ZQPM12] Fan Zhong, Xueying Qin, Qunsheng Peng, and Xiangxu Meng. Discontinuity-aware video object cutout. ACM Transactions on Graphics, 31 (6):175:1–175:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2014:HOS

[ZRB14] Shuang Zhao, Ravi Ramamoorthi, and Kavita Bala. High-order similarity relations in radiative transfer. ACM Transactions on Graphics, 33 (4):104:1–104:??, July 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:PSD

[ZRJ23] Ziyi Zhang, Nicolas Roussel, and Wenzel Jakob. Projective sampling for differentiable rendering of geometry. ACM Transactions on Graphics, 42(6): 212:1–212:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618385.

Zhou:2008:RTS

[ZRL^{+}08] Kun Zhou, Zhong Ren, Stephen Lin, Hujun Bao, Baining Guo, and Heung-Yeung Shum. Real-time smoke rendering using compensated ray marching. ACM Transactions on Graphics, 27 (3):36:1–36:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zinke:2009:PAP

[ZRL^{+}09] Arno Zinke, Martin Rump, Tomás Lay, Andreas Weber, Anton Andriyenko, and Reinhard Klein. A practical approach for photometric acquisition of hair color. ACM Transactions on Graphics, 28(5):165:1–165:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2007:CCD

[ZRLK07] Xinyu Zhang, Stephane Redon, Minkyoung Lee, and Young J. Kim. Continuous collision detection for articulated models using Taylor models and temporal culling. ACM Transactions on Graphics, 26(3):15:1–15:??, July 2007. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2018:ASP

[ZRSM18] Zechen Zhang, Nikunj Raghuvanshi, John Snyder, and Steve Marschner. Ambient sound propagation. ACM Transactions on Graphics, 37 (6):184:1–184:??, November 2018. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2000:ETP

[ZS00] Jianmin Zheng and Thomas W.

Sederberg. Estimating tessellation parameter intervals for rational curves and surfaces. ACM Transactions on Graphics, 19(1):56–77, January 2000. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

http://www.acm.org/pubs/
citations/journals/tog/
2000-19-1/p56-zheng/.

Zhang:2021:VCV

[ZSAF21] Haotian Zhang, Cristobal Sciutto, Maneesh Agrawala, and Kayvon Fatahalian. Vid2Player: Controllable video sprites that behave and appear like professional tennis players. ACM Transactions on Graphics, 40(3):24:1–24:16, July 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3448978.

Zhu:2017:TSTa

[ZSCM17a] Bo Zhu, Melina Skouras, Desai Chen, and Wojciech Matusik. Two-scale topology optimization with microstructures. ACM Transactions on Graphics, 36(4): 120:1–120:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2017:TSTb

[ZSCM17b] Bo Zhu, Mélina Skouras, Desai Chen, and Wojciech Matusik. Two-scale topology optimization with microstructures. ACM Transactions on Graphics, 36(5): 164:1–164:??, October 2017. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2004:SFH

[ZSCS04] Li Zhang, Noah Snavely, Brian Curless, and Steven M. Seitz. Spacetime faces: high resolution capture for modeling and animation. ACM Transactions on Graphics, 23 (3):548–558, August 2004. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:NNF

[ZSD^{+}21] Xiuming Zhang, Pratul P. Srinivasan, Boyang Deng, Paul Debevec, William T. Freeman, and Jonathan T. Barron. NeRFactor: neural factorization of shape and reflectance under an unknown illumination. ACM Transactions on Graphics, 40(6): 237:1–237:18, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480496.

Zeltner:2021:MCE

[ZSGJ21] Tizian Zeltner, Sébastien Speierer, Iliyan Georgiev, and Wenzel Jakob. Monte Carlo estimators for differential light transport. ACM Transactions on Graphics, 40 (4):78:1–78:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459807.

Zhang:2018:MAN

[ZSKS18] He Zhang, Sebastian Starke, Taku Komura, and Jun Saito. Mode-adaptive neural networks for quadruped motion control. ACM Transactions on Graphics, 37(4): 145:1–145:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2014:BFO

[ZSMS14] Yahan Zhou, Shinjiro Sueda, Wojciech Matusik, and Ariel Shamir. Boxelization: folding 3D objects into boxes. ACM Transactions on Graphics, 33 (4):71:1–71:??, July 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2020:GDP

[ZSSJL20] Lvmin Zhang, Edgar Simo-Serra, Yi Ji, and Chunping Liu. Generating digital painting lighting effects via RGBspace geometry. ACM Transactions on Graphics, 39(2): 13:1–13:13, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3372176.

Zhu:2010:EMM

[ZSTB10] Yongning Zhu, Eftychios Sifakis, Joseph Teran, and Achi Brandt. An efficient multigrid method for the simulation of high-resolution elastic solids. ACM Transactions on Graphics, 29(2):16:1– 16:18, March 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2010:NLS

[ZSW^{+}10] Qian Zheng, Andrei Sharf, Guowei Wan, Yangyan Li, Niloy J. Mitra, Daniel Cohen-Or, and Baoquan Chen. Nonlocal scan consolidation for 3D urban scenes. ACM Transactions on Graphics, 29(4): 94:1–94:??, July 2010. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2014:LDC

[ZSZ^{+}14] Peizhao Zhang, Kristin Siu, Jianjie Zhang, C. Karen Liu, and Jinxiang Chai. Leveraging depth cameras and wearable pressure sensors for fullbody kinematics and dynamics capture. ACM Transactions on Graphics, 33(6):

221:1–221:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:UAS

[ZTD^{+}23] Yuxin Zhang, Fan Tang, Weiming Dong, Haibin Huang, Chongyang Ma, Tong-Yee Lee, and Changsheng Xu. A unified arbitrary style transfer framework via adaptive contrastive learning. ACM Transactions on Graphics, 42(5): 169:1–169:??, October 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3605548.

Zhou:2018:SML

[ZTF^{+}18] Tinghui Zhou, Richard Tucker, John Flynn, Graham Fyffe, and Noah Snavely. Stereo magnification: learning view synthesis using multiplane images. ACM Transactions on Graphics, 37(4):65:1– 65:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:SSR

[ZTNW23] Biao Zhang, Jiapeng Tang, Matthias Nießner, and Peter Wonka. 3DShape2VecSet: a 3D shape representation for neural fields and generative diffusion models. ACM Transactions on Graphics, 42 (4):92:1–92:??, August 2023.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592442.

Zatzarinni:2009:RAE

[ZTS09] Rony Zatzarinni, Ayellet Tal, and Ariel Shamir. Relief analysis and extraction. ACM Transactions on Graphics, 28(5):136:1–136:9, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2020:OFF

[ZVC^{+}20] Paul Zhang, Josh Vekhter, Edward Chien, David Bommes, Etienne Vouga, and Justin Solomon. Octahedral frames for feature-aligned cross fields. ACM Transactions on Graphics, 39(3):25:1–25:13, June 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        abs/10.1145/3374209.

Zhang:2021:DNG

[ZWCM21] Meng Zhang, Tuanfeng Y. Wang, Duygu Ceylan, and Niloy J. Mitra. Dynamic neural garments. ACM Transactions on Graphics, 40(6): 235:1–235:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3478513.3480497.

Zhao:2016:DSP

[ZWDR16] Shuang Zhao, Lifan Wu, Frédo Durand, and Ravi Ramamoorthi. Downsampling scattering parameters for rendering anisotropic media. ACM Transactions on Graphics, 35(6):166:1– 166:??, November 2016. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2002:FBL

[ZWGS02] Zhunping Zhang, Lifeng Wang, Baining Guo, and Heung-Yeung Shum. Featurebased light field morphing. ACM Transactions on Graphics, 21(3):457–464, July 2002. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2022:SHA

[ZWHB22] Mianlun Zheng, Bohan Wang, Jingtao Huang, and Jernej Barbič. Simulation of hand anatomy using medical imaging. ACM Transactions on Graphics, 41(6):273:1– 273:??, December 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555486.

Zhao:2014:ISU

[ZWK14] Xi Zhao, He Wang, and Taku Komura. Indexing 3D scenes using the interaction bisector surface. ACM Transactions on Graphics, 33(3): 22:1–22:??, May 2014. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhong:2018:CHD

[ZWL^{+}18] Zichun Zhong, Wenping Wang, Bruno Lévy, Jing Hua, and Xiaohu Guo. Computing a high-dimensional Euclidean embedding from an arbitrary smooth Riemannian metric. ACM Transactions on Graphics, 37(4):62:1–62:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2022:SSC

[ZWL22] Lvmin Zhang, Tien-Tsin Wong, and Yuxin Liu. Spritefrom-sprite: Cartoon animation decomposition with self-supervised Sprite estimation. ACM Transactions on Graphics, 41(6): 192:1–192:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555439.

Zhou:2021:VFA

[ZWRY21] Yang Zhou, Lifan Wu, Ravi Ramamoorthi, and Ling-Qi Yan. Vectorization for fast, analytic, and differentiable visibility. ACM Transactions on Graphics, 40(3): 27:1–27:21, July 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3452097.

Zhao:2024:HHF

[ZWS^{+}24] Xiaochen Zhao, Lizhen Wang, Jingxiang Sun, Hongwen Zhang, Jinli Suo, and Yebin Liu. HAvatar: Highfidelity head avatar via facial model conditioned neural radiance field. ACM Transactions on Graphics, 43 (1):6:1–6:??, February 2024. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3626316.

Zhou:2005:T

[ZWT^{+}05] Kun Zhou, Xi Wang, Yiying Tong, Mathieu Desbrun, Baining Guo, and Heung-Yeung Shum. Texture-Montage. ACM Transactions on Graphics, 24(3): 1148–1155, July 2005. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2023:FGR

[ZWTP23] Kaixuan Zhang, Jingxian Wang, Daizong Tian, and Thrasyvoulos N. Pappas. Film grain rendering and parameter estimation. ACM Transactions on Graphics, 42(4): 63:1–63:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592127.

Zhang:2018:MHR

[ZWW^{+}18] Meng Zhang, Pan Wu, Hongzhi Wu, Yanlin Weng, Youyi Zheng, and Kun Zhou. Modeling hair from an RGB-D camera. ACM Transactions on Graphics, 37(6): 205:1–205:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2016:RBI

[ZWZ^{+}16] Fang-Lue Zhang, Xian Wu, Hao-Tian Zhang, Jue Wang, and Shi-Min Hu. Robust background identification for dynamic video editing. ACM Transactions on Graphics, 35 (6):197:1–197:??, November 2016. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2022:COD

[ZWZ^{+}22] Haisen Zhao, Max Willsey, Amy Zhu, Chandrakana Nandi, Zachary Tatlock, Justin Solomon, and Adriana Schulz. Co-optimization of design and fabrication plans for carpentry. ACM Transactions on Graphics, 41 (3):32:1–32:13, June 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3508499.

Zhu:2018:SSC

[ZXC^{+}18] Chenyang Zhu, Kai Xu, Siddhartha Chaudhuri, Renjiao Yi, and Hao Zhang. SCORES: shape composition with recursive substructure priors. ACM Transactions on Graphics, 37 (6):211:1–211:??, November 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2020:OAP

[ZXH^{+}20] Xiaohui Zhou, Ke Xie, Kai Huang, Yilin Liu, Yang Zhou, Minglun Gong, and Hui Huang. Offsite aerial path planning for efficient urban scene reconstruction. ACM Transactions on Graphics, 39 (6):192:1–192:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417791.

Zhang:2013:LAI

[ZXJ^{+}13] Hao Zhang, Kai Xu, Wei Jiang, Jinjie Lin, Daniel Cohen-Or, and Baoquan Chen. Layered analysis of irregular facades via symmetry maximization. ACM Transactions on Graphics, 32(4): 121:1–121:??, July 2013. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2020:RGG

[ZXKL^{+}20] Allan Zhao, Jie Xu, Mina Konaković-Luković, Josephine Hughes, Andrew Spielberg, Daniela Rus, and Wojciech Matusik. RoboGrammar: graph grammar for terrainoptimized robot design. ACM Transactions on Graphics, 39 (6):188:1–188:16, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

https://dl.acm.org/doi/
10.1145/3414685.3417831.

Zhou:2018:VAD

[ZXL^{+}18] Yang Zhou, Zhan Xu, Chris Landreth, Evangelos Kalogerakis, Subhransu Maji, and Karan Singh. Visemenet: audio-driven animator-centric speech animation. ACM Transactions on Graphics, 37(4):161:1–161:??, August 2018. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2012:MGM

[ZXS^{+}12] Lifeng Zhu, Weiwei Xu, John Snyder, Yang Liu, Guoping Wang, and Baining Guo. Motion-guided mechanical toy modeling. ACM Transactions on Graphics, 31(6): 127:1–127:??, November 2012. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2021:HNR

[ZXS^{+}21] Shilin Zhu, Zexiang Xu, Tiancheng Sun, Alexandr Kuznetsov, Mark Meyer, Henrik Wann Jensen, Hao Su, and Ravi Ramamoorthi. Hierarchical neural reconstruction for path guiding using hybrid path and photon samples. ACM Transactions on Graphics, 40(4): 35:1–35:16, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459810.

Zhu:2022:PDN

[ZXS^{+}22] Shilin Zhu, Zexiang Xu, Tiancheng Sun, Alexandr Kuznetsov, Mark Meyer, Henrik Wann Jensen, Hao Su, and Ravi Ramamoorthi. Photondriven neural reconstruction for path guiding. ACM Transactions on Graphics, 41 (1):7:1–7:15, February 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3476828.

Zong:2023:PFS

[ZXS^{+}23] Chen Zong, Jiacheng Xu, Jiantao Song, Shuangmin Chen, Shiqing Xin, Wenping Wang, and Changhe Tu. P2M: a fast solver for querying distance from point to mesh surface. ACM Transactions on Graphics, 42(4): 147:1–147:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592439.

Zhang:2015:OSA

[ZXTZ15] Yizhong Zhang, Weiwei Xu, Yiying Tong, and Kun Zhou. Online structure analysis for real-time indoor scene reconstruction. ACM Transactions on Graphics, 34(5): 159:1–159:??, October 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhong:2023:CPE

[ZXZL23] Fanchao Zhong, Yonglai Xu, Haisen Zhao, and Lin Lu. As-continuous-as-possible extrusion-based fabrication of surface models. ACM Transactions on Graphics, 42 (3):26:1–26:??, June 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3575859.

Zyda:1988:DAC

[Zyd88] Michael J. Zyda. A decomposable algorithm for contour surface display generation. ACM Transactions on Graphics, 7(2):129–148, April 1988. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        http://www.acm.org/pubs/
                                        toc/Abstracts/0730-0301/
                                        42461.html.

Zhou:2015:GCD

[ZYH^{+}15] Yang Zhou, Kangxue Yin, Hui Huang, Hao Zhang, Minglun Gong, and Daniel Cohen-Or. Generalized cylinder decomposition. ACM Transactions on Graphics, 34(6): 171:1–171:??, November 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhu:2017:DDS

[ZYL^{+}17] Chenyang Zhu, Renjiao Yi, Wallace Lira, Ibraheem Alhashim, Kai Xu, and Hao Zhang. Deformation-driven shape correspondence via shape recognition. ACM Transactions on Graphics, 36 (4):51:1–51:??, July 2017. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2020:LMA

[ZYL^{+}20] Yunbo Zhang, Wenhao Yu, C. Karen Liu, Charlie Kemp, and Greg Turk. Learning to manipulate amorphous materials. ACM Transactions on Graphics, 39(6): 189:1–189:11, November 2020. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3414685.3417868.

Zhang:2020:DGM

[ZYM^{+}20] Zaiwei Zhang, Zhenpei Yang, Chongyang Ma, Linjie Luo, Alexander Huth, Etienne Vouga, and Qixing Huang. Deep generative modeling for scene synthesis via hybrid representations. ACM Transactions on Graphics, 39(2): 17:1–17:21, April 2020. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3381866.

Zhang:2023:LPS

[ZYM^{+}23] Haotian Zhang, Ye Yuan, Viktor Makoviychuk, Yunrong Guo, Sanja Fidler, Xue Bin Peng, and Kayvon Fatahalian. Learning physically simulated tennis skills from broadcast videos. ACM Transactions on Graphics, 42 (4):95:1–95:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3592408.

Zhong:2014:SFB

[ZYQ^{+}14] Fan Zhong, Song Yang, Xueying Qin, Dani Lischinski, Daniel Cohen-Or, and Baoquan Chen. Slippage-free background replacement for hand-held video. ACM Transactions on Graphics, 33(6): 199:1–199:??, November 2014. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:MNM

[ZYSK21] He Zhang, Yuting Ye, Takaaki Shiratori, and Taku Komura. ManipNet: neural manipulation synthesis with a hand-object spatial representation. ACM Transactions on Graphics, 40(4): 121:1–121:14, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3450626.3459830.

Zinke:2008:DSA

[ZYWK08] Arno Zinke, Cem Yuksel, Andreas Weber, and John Keyser. Dual scattering approximation for fast multiple scattering in hair. ACM Transactions on Graphics, 27 (3):32:1–32:??, August 2008. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:CAP

[ZYX^{+}21] Han Zhang, Yucong Yao, Ke Xie, Chi-Wing Fu, Hao Zhang, and Hui Huang. Continuous aerial path planning for 3D urban scene reconstruction. ACM Transactions on Graphics, 40(6): 225:1–225:15, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480483.

Zhang:2021:PSD

[ZYZ21] Cheng Zhang, Zihan Yu, and Shuang Zhao. Pathspace differentiable rendering of participating media. ACM Transactions on Graphics, 40 (4):76:1–76:15, August 2021.

CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459782.

Zhang:2015:CHP

[ZYZZ15] Yizhong Zhang, Chunji Yin, Changxi Zheng, and Kun Zhou. Computational hydrographic printing. ACM Transactions on Graphics, 34(4): 131:1–131:??, August 2015. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhou:2018:NST

[ZZB^{+}18] Yang Zhou, Zhen Zhu, Xiang Bai, Dani Lischinski, Daniel Cohen-Or, and Hui Huang. Non-stationary texture synthesis by adversarial expansion. ACM Transactions on Graphics, 37(4):49:1– 49:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2022:LHP

[ZZC^{+}22] Shaokun Zheng, Zhiqian Zhou, Xin Chen, Difei Yan, Chuyan Zhang, Yuefeng Geng, Yan Gu, and Kun Xu. LuisaRender: a highperformance rendering framework with layered and unified interfaces on stream architectures. ACM Transactions on Graphics, 41(6): 232:1–232:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3550454.3555463.

Zhang:2023:HLP

[ZZC^{+}23] Longwen Zhang, Zijun Zhao, Xinzhou Cong, Qixuan Zhang, Shuqi Gu, Yuchong Gao, Rui Zheng, Wei Yang, Lan Xu, and Jingyi Yu. HACK: Learning a parametric head and neck model for highfidelity animation. ACM Transactions on Graphics, 42 (4):41:1–41:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3592093.

Zhuang:2013:GEM

[ZZCJ13] Yixin Zhuang, Ming Zou, Nathan Carr, and Tao Ju. A general and efficient method for finding cycles in 3D curve networks. ACM Transactions on Graphics, 32(6): 180:1–180:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2017:RTU

[ZZI^{+}17] Richard Zhang, Jun-Yan Zhu, Phillip Isola, Xinyang Geng, Angela S. Lin, Tianhe Yu, and Alexei A. Efros. Real-time user-guided image colorization with learned deep priors. ACM Transactions on Graphics, 36(4):119:1–119:??, July 2017. CODEN ATGRDF.

ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2021:SRB

[ZZL^{+}21] Nanxuan Zhao, Quanlong Zheng, Jing Liao, Ying Cao, Hanspeter Pfister, and Rynson W. H. Lau. Selective region-based photo color adjustment for graphic designs. ACM Transactions on Graphics, 40(2):17:1–17:16, June 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3447647.

Zhong:2023:VVS

[ZZL^{+}23] Fanchao Zhong, Haisen Zhao, Haochen Li, Xin Yan, Jikai Liu, Baoquan Chen, and Lin Lu. VASCO: Volume and surface co-decomposition for hybrid manufacturing. ACM Transactions on Graphics, 42 (6):188:1–188:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3618324.

Zhu:2023:NCP

[ZZLH23] Qingxu Zhu, He Zhang, Mengting Lan, and Lei Han. Neural categorical priors for physics-based character control. ACM Transactions on Graphics, 42(6): 178:1–178:??, December 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3618397.

Zhao:2013:RRP

[ZZMC13] Wenping Zhao, Jianjie Zhang, Jianyuan Min, and Jinxiang Chai. Robust realtime physics-based motion control for human grasping. ACM Transactions on Graphics, 32 (6):207:1–207:??, November 2013. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2021:SDV

[ZZT^{+}21] Hao Zhang, Yuxiao Zhou, Yifei Tian, Jun-Hai Yong, and Feng Xu. Single depth view based real-time reconstruction of hand-object interactions. ACM Transactions on Graphics, 40(3): 29:1–29:12, July 2021. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3451341.

Zhang:2003:SPV

[ZZV^{+}03] Jingdan Zhang, Kun Zhou, Luiz Velho, Baining Guo, and Heung-Yeung Shum. Synthesis of progressively-variant textures on arbitrary surfaces. ACM Transactions on Graphics, 22(3):295–302, July 2003. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2022:CRM

[ZZW^{+}22a] Chuankun Zheng, Ruzhang Zheng, Rui Wang, Shuang Zhao, and Hujun Bao. A compact representation of measured BRDFs using neural processes. ACM Transactions on Graphics, 41(2):14:1– 14:15, April 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3490385.

Zhu:2022:PLD

[ZZW^{+}22b] Junqiu Zhu, Sizhe Zhao, Lu Wang, Yanning Xu, and Ling-Qi Yan. Practical levelof-detail aggregation of fur appearance. ACM Transactions on Graphics, 41(4): 47:1–47:??, July 2022. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3528223.3530105.

Zhang:2012:VMD

[ZZWC12] Juyong Zhang, Jianmin Zheng, Chunlin Wu, and Jianfei Cai. Variational mesh decomposition. ACM Transactions on Graphics, 31(3): 21:1–21:14, May 2012. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhao:2018:DDS

[ZZX^{+}18] Haisen Zhao, Hao Zhang, Shiqing Xin, Yuanmin Deng, Changhe Tu, Wenping Wang, Daniel Cohen-Or, and Baoquan Chen. DSCarver: decompose-and-spiral-carve for subtractive manufacturing. ACM Transactions on Graphics, 37(4):137:1– 137:??, August 2018. CO-DEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zheng:2021:EDM

[ZZXY21] Shaokun Zheng, Fengshi Zheng, Kun Xu, and Ling-Qi Yan. Ensemble denoising for Monte Carlo renderings. ACM Transactions on Graphics, 40(6): 274:1–274:17, December 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3478513.3480510.

Zeng:2009:IPP

[ZZXZ09] Kun Zeng, Mingtian Zhao, Caiming Xiong, and Song-Chun Zhu. From image parsing to painterly rendering. ACM Transactions on Graphics, 29(1):2:1–2:11, December 2009. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic).

Zhang:2022:VDN

[ZZZ^{+}22] Longwen Zhang, Chuxiao Zeng, Qixuan Zhang, Hongyang Lin, Ruixiang Cao, Wei Yang, Lan Xu, and Jingyi Yu. Video-driven neural physically-based facial asset for production. ACM Transactions on Graphics, 41(6): 208:1–208:??, December 2022. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3550454.3555445.

Zheng:2023:ART

[ZZZ^{+}23] Zerong Zheng, Xiaochen Zhao, Hongwen Zhang, Boning Liu, and Yebin Liu. AvatarReX: Real-time expressive full-body avatars. ACM Transactions on Graphics, 42(4):158:1–158:??, August 2023. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3592101.

Zhang:2021:RRO

[ZZZX21] Jiazhao Zhang, Chenyang Zhu, Lintao Zheng, and Kai Xu. ROSEFusion: random optimization for online dense reconstruction under fast camera motion. ACM Transactions on Graphics, 40 (4):56:1–56:17, August 2021. CODEN ATGRDF. ISSN 0730-0301 (print), 1557-7368 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3450626.3459676.